A Complete Bibliography of Publications in Progress in Oceanography

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

25 June 2024 Version 1.19

Title word cross-reference

$19.00 [Hof81]. $300.00 [Ang79a]. $52.00 [Hof81]. $84.00 [Ang80]. > [DGMM85]. ^{1} [Tho77]. ^{13} [CGC^{+}20, OMK^{+}22, PRL^{+}18, SKH^{+}23, WRS^{+}92]. 134 [IHT^{+}21]. ^{137} [AFH^{+}11, KKS^{+}03, PAF^{+}11, SCMAR^{+}99, SCLG^{+}11, TAH^{+}11]. ^{14} [OE65, PRL^{+}18, SBK^{+}95, WRS^{+}92]. ^{15} [CGC^{+}20, CSC^{+}12, OMK^{+}22, SKH^{+}23, WRS^{+}92]. ^{15}N [GSVB23]. ^{18} [BSF95, LCJ^{+}17]. ^{210} [KKS^{+}03, MSL^{+}07, SCMAR^{+}99]. ^{222} [Gam14]. ^{227} [LvBS^{+}24]. ^{228} [IHT^{+}21]. ^{230} [DTKvH15, ST65]. ^{231} [DTKvH15]. ^{234} [AYK^{+}05, HPZC21, ST65]. ^{238} [AYK^{+}05]. ^{3} [SBK^{+}95]. ^{39} [SBK^{+}95]. _{2} [ARELAK24, AEPW93, BSF95, BK19, BF01, CKP^{+}20, CPG^{+}18, CKM^{+}21, CVBG21, CFG07, CF12, CMF15, DCD^{+}23, EFC^{+}23, EHSI12, FGS^{+}15, FC07, GCB^{+}22, GDSCU09, GSF^{+}15, GLLB22, HSS^{+}12, LL97, LM00, LGZ^{+}20, lLdZQ^{+}22, LGG18, MKOLA20, OYKK^{+}23, OKdA^{+}19, PPKR14, RCGC^{+}16, RVS^{+}21, SPK^{+}19, TŠT^{+}17, TAM^{+}15, VSA^{+}21, WD94, WNNI21, WBA^{+}22, WWL^{+}24, WST^{+}16, YMI88, ZCD08, ZDG^{+}21, ZKK^{+}16]. _{37} [STHM02]. _{6} [MRO^{+}08]. _{86} [SBH^{+}14]. a [RBR^{+}23]. α [KKKS14, SCCJ^{+}18, WPB^{+}08]. ≈ [BM07, CPNL07, CFG07, GMAB07, MSV^{+}14, ZHSMM14]. ≈ 150 [SLG^{+}12]. β [KKKS14, WPB^{+}08]. Δ [PRL^{+}18, CGC^{+}20, CSC^{+}12, GSVB23, LCJ^{+}17, OMK^{+}22, SKH^{+}23, WPB^{+}08]. f [EFC^{+}23, GDSCU09]. γ [WPB^{+}08]. K_{1} [AK97, CGW^{+}22, DJ92]. M_{2} [CGW^{+}22, DJ92]. n [SOH21]. N_{2} [DJ92]. O_{1} [DJ92]. ω [SCCJ^{+}18]. p [LGG18, ZDG^{+}21]. S_{2} [DJ92]. × [Ang79a, Hof81]. x [RCGC^{+}16].

* [WNNI21].

-alkanes [SOH21]. -dicarboxylic [SCCJ^{+}18]. -oxoacids [SCCJ^{+}18]. -unsaturated [WPB^{+}08].

/U [ST65].

0 [Ang79a, Ang88, SW81]. 0-08-020919-X [Hof81]. 0-08-021953-5 [Ang79a]. 0-08-022960-3 [SW81]. 0-08-026248-1 [Bak83]. 0-08-036649-X [Ang88]. 0-group [SEG^{+}22b]. 00039-3 [SDS22a]. 000m [DGMM85]. 02 [NBR^{+}08]. 08^{o} [MM80].

1 [CPB^{+}15, CNSHT15, GAM98a, MFS^{+}07, MR03, SFMT12, SJ02c, VOJD02a]. 1-D [CNSHT15]. 1.4 [SKWWGV18]. 1.4-REcoM2 [SKWWGV18]. 10 [AdAK^{+}18]. 10-year [HFO^{+}22]. 1000m [Lev88, OT19]. 123 [ÁBMÁS15]. 130 [GFB^{+}15b]. 141pp [Ang88]. 143 [MFS^{+}16a]. 150^{o} [AMY^{+}23, YKH^{+}24]. 155^{o} [YMA^{+}17]. 158^{o} [HBH^{+}17]. 159 [Kru19]. 16th [FJhT^{+}14]. 170^{o} [MRRC73]. 173 [FDH20]. 175 [RBS^{+}20]. 176 [BJMP20]. 18-year [LGZ^{+}20]. 180 [KMF^{+}20a, MHS^{+}20a]. 182 [JLP^{+}20a]. 1920s [Dri06]. 1926/1927 [WMWR08]. 1930s [Dri06]. 1958 [McG64]. 1965 [Ano65i]. 1970s [SKH00]. 1976/77 [WXH07]. 1979 [FHP83]. 1980s [FMW91, GDM^{+}15, MBB^{+}96, SBK^{+}95, SKSK06, TKWI08]. 1982/83 [LBH^{+}87]. 1984 [MN88]. 1986 [YJW88]. 1987 [KDB95]. 1988/89 [WXH07]. 1990 [STB^{+}92]. 1990s [INI^{+}17, JSKM02, RKM^{+}07, SiSI^{+}02, Smi05]. 1991 [PGC^{+}96, RAP95]. 1992/1995 [KF11]. 1994 [HFW^{+}98, KTB^{+}99]. 1995 [GCZ^{+}00, GSPMAI99, KTB^{+}99]. 1998 [Ano98c, Ano98b, Ano98a, Ano98d, Ano98e, JRW01]. 1998/1999 [IIS^{+}17]. 1998/99 [Min02]. 1999 [Ano99c, Ano99e, Ano99b, Ano99d].

2 [Ano65i, Ano09h, GAM98b, MSF^{+}07, SFMT14, SJ02b, VOJD02b]. 20-year [CMG15]. 2000 [Ano00e, Ano00g, Ano00d, Ano00a, Ano00b, Ano00c, Ano00h]. 2000s [GDM^{+}15]. 2002 [Ano02a, Ano02b, Ano02c, Ano22-33]. 2003 [Ano03f, Ano03e, Ano03c, Ano03a, Ano03d, Ano03b]. 2003/2004 [KF11].

2004 [Ano04b, Ano04f, Ano04c, Ano04d, Ano04g, Ano04e, HBV^{+}10]. 2005 [Ano05a, Ano05d, Ano05c, Ano05b]. 2006 [Ano06c, Ano06a, Ano06d, Ano06b]. 2007 [Ano07n, Ano07q, Ano07k, Ano07m, Ano07o, Ano07r, Ano07l, Ano07p, KMU^{+}12, LM10]. 2008 [Ano08p, Ano08r, Ano08s, Ano08o, Ano08n, Ano08u, Ano08w, Ano08q, Ano08t, Ano08v, MHA^{+}11]. 2009 [Ano09m, Ano09i, Ano09j, Ano09k, Ano09l, Ano22-27, Rud15]. 2010 [Ano10k, Ano10l, Ano10m]. 2011 [Ano11i, Ano11h, Ano11g, Ano11j, Ano22-28]. 2012 [Ano12m, Ano12n, Ano12p, Ano12o, Ano12l, Ano12k]. 2013 [Ano13n, Ano13l, Ano13p, Ano13m, Ano13o, Ano22-31]. 2014 [Ano14m, Ano14p, Ano14q, Ano14n, Ano14r, Ano14o, Ano14l]. 2015 [Ano15n, Ano15o, Ano15m, Ano15r, Ano15p, Ano15q, GMDD^{+}22b, MPD^{+}22]. 2016 [Ano16t, Ano16o, Ano16n, Ano16p, Ano16m, Ano16q, Ano16k, Ano16l, Ano16s, Ano16r]. 2017 [Ano17r, Ano17n, Ano17o, Ano17m, Ano17q, Ano17s, Ano17l, Ano17p]. 2018 [Ano18j, Ano18k, Ano18m, Ano18n, Ano18l, Ano22-29, Ano22-32, CBB^{+}19, CBB^{+}22b]. 2018/2019 [MST^{+}23a, SSN23]. 2019 [Ano19a, Ano19j, Ano19n, Ano19m, Ano19k, Ano19l, Ano19p, Ano22-30, FBB^{+}21, MPTMK22]. 2020 [Ano20a, Ano20b, Ano20m, Ano20n, Ano20p, Ano20o, Ano20r, Ano20s, Ano20t]. 2021 [Ano21b, Ano21a, Ano21c, Ano21n, Ano21o, Ano21q, Ano21p, Ano21s, Ano21t]. 2022 [Ano22b, Ano22a, Ano22c, Ano22u, Ano22v, Ano22x, Ano22w, Ano22z, Ano22-34]. 2023 [Ano23b, Ano23a, Ano23c, Ano23n, Ano23o, Ano23q, Ano23p, Ano23r, Ano23s, Ano23t]. 2024 [Ano24a, Ano24h, Ano24i, Ano24k, Ano24j, Ano24l]. 20^{o} [JG07]. 20th [HMH07]. 212-layer [MY92]. 21^{o} [BHC^{+}18]. 21st [DML^{+}16, MWS^{+}10, TLP^{+}16]. 22-year [HGBG20]. 224 [ZLRVB24a]. 22^{o} [FMSBW13, dSPF^{+}23]. 234/uranium [KG65]. 23^{o} [BM76, CBB^{+}02, VSGC21, dSPF^{+}23]. 24.5N [HGPFN^{+}14]. 24^{o} [BHC^{+}18]. 25-year [MAH^{+}15]. 25^{o} [BDBJ01]. 26^{o} [MSJ^{+}15]. 27^{o} [MM80]. 29 [CBB^{+}22b]. 2nd [Bak83].

3 [BHH^{+}16, MMR^{+}09, SW81, SW01, XRC^{+}15]. 3-D [BHH^{+}16, MMR^{+}09, SW01, XRC^{+}15]. 30-year [AUE^{+}14]. 3000m [YNMY23]. 30^{o} [Ang79b, BM76, HGT16, HGTP^{+}19, RKFD07]. 30m [AMEV07]. 31 [Ano94c]. 31^{o} [AAM^{+}14]. 32^{o} [APHGC^{+}22, MBKS08]. 33^{o} [FMSBW13, MCGS^{+}16]. 36.5^{o} [FC07]. 36^{o} [BM07, CPNL07, CFG07, GMAB07, HYM^{+}12, MSL^{+}07, MDL^{+}12, SLG^{+}12]. 38^{o} [SJM^{+}19]. 3D [ASR^{+}20, AUE^{+}14, MRM^{+}14, SNS^{+}22]. 3rd [PBH^{+}10].

4-dimensional [MAB^{+}11c, MAB^{+}11a, MAB^{+}11b]. 40^{o} [CF12, EM12]. 41 [BLAM00]. 41CP [KSK21]. 42^{o} [MSV^{+}14, ZHSMM14]. 44.7^{o} [PSP^{+}21]. 44^{o} [GMAGH^{+}17]. 46^{o} [MCGS^{+}16, SJM^{+}19]. 47^{o} [QOS^{+}22]. 4822m [RLT^{+}22]. 48^{o} [HM98, PMA^{+}14, PTPY^{+}23, RBL^{+}19, WBH15]. 49 [HHW22].

4Concepción [SSL07]. 4D [GAM98a, GAM98b, GA00, MZGA^{+}20, WSO01]. 4D-Var [MZGA^{+}20]. 4D-variational [GAM98a, GAM98b, GA00].

5 [Ang79a]. 50-year [CLdPHL23]. 50^{o} [BPP^{+}98, GMAGH^{+}17, GBC^{+}00, GCD^{+}13, HM98, Hen85, JTD^{+}14, vWHdS^{+}98, CBB^{+}02, RKFD07]. 51^{o} [PTPY^{+}23]. 52^{o} [STPHM^{+}23]. 53 [SDS22a]. 54^{o} [QOS^{+}22]. 56 [RG03a]. 56^{o} [PMA^{+}14, TSRF14]. 58 [Fei04]. 59 [VHK04].

’60 [Fug63]. 60^{o} [SAA^{+}15]. 648pp [Bak83]. 66^{o} [CMJPH^{+}18]. 69 [WF07].

7.5N [HGPFN^{+}14]. 70 [RLSF07]. 71 [SHC^{+}07]. 73 [Yas07b]. 76 [Ano94k]. 79 [SE09].

80 [VH09a]. 87 [KN11].

9-year [AT07]. 91 [Rea00, dMGS^{+}11b].

= [AHRT90].

A10 [KMWF11]. A2 [LSS^{+}10, SYB^{+}15]. A204 [Ano94j]. A209 [Ano94j]. A219 [Ano94j]. aanderaa [Dah69]. abdominalis [HTG15]. abiotic [GBT^{+}19]. ablation [LDHW20]. absolute [DPCS87, Emi65, Nee85]. absorption [HOY^{+}21a, KM08, SPWH21]. Abundance [BM07, EKB06, FBT^{+}22, KOT^{+}21, PSL87, STW^{+}15, SPB93, Ver91, ADV^{+}18, BMO12, BLI^{+}99, BM01, BCGN^{+}18, BPP^{+}98, BMG^{+}21a, BGWP^{+}17, CMF11, Cra09, CCW^{+}18, DDE^{+}95, DBR03, DAU22, DPM^{+}09, DBM17, DBR20, ECGP01, EBS^{+}18, EHG^{+}12, FELJ16, FGGDF^{+}04, FBM^{+}08, GFGGD^{+}23, GGAA^{+}23, HL05, HCAFD^{+}20, HLSX22, IIS^{+}17, IAFD02, JTQ^{+}18, JMZ23, LCGH07, MVN^{+}15, MCG^{+}02, MPSD15, MSA^{+}22, MWFH02, MMPG07, Nag01, NMLBCM^{+}01, PL87, PLHLF05, PCH08a, Reb02, SGL^{+}18, SCC^{+}19, SDH^{+}14, SAB^{+}21, TB15, THP21, VDDA^{+}08, VR03, WMB^{+}21, Whe06, XWL^{+}18, XYL^{+}22, YMA^{+}17, YNMY23, YPGE^{+}10, ZLS^{+}04]. abundances [CLX^{+}20, KYS^{+}17, MA12]. Abundant [ZBY^{+}22]. Abyssal [BBR^{+}01, BKD^{+}20, BGS^{+}04, DBJ^{+}15, HBD^{+}21, HCV^{+}20, IBW^{+}01, NRA^{+}21, RBL90, SBG16, Thu90, VSC01, VDP^{+}01, WDK^{+}01, dJSL^{+}20, AP20, BBRM20, BC01, CES^{+}19, CBL^{+}19, DBR20, GSV^{+}01, GA01, GHSC19, Hau18, HGTP^{+}19, HTV^{+}20, JZ19, JSLA^{+}21, JPBB20, KO19, KRHS14, LMPB^{+}16, LBK^{+}01, MDR20, QLY^{+}22, RMG90, RMB^{+}01, SSKA19, SLBH^{+}19, TDL^{+}17, VCSG^{+}01, VPW01, WHBW03, XWL^{+}22]. Ac [LvBS^{+}24]. Acartia [BD20, SJJ^{+}03]. ACC [LI10]. acceleration [HHWW20]. acceptors [MBCB88]. access [dlHRA^{+}18]. Accomplishments [WHG^{+}16]. according [MBCB88]. accounting [ARDP14, MCH^{+}12]. accretionary [JFG^{+}90]. accumulated [SL13]. Accumulation [CJRÁ^{+}13, CFC^{+}18, CBD^{+}24, LCJ^{+}07, MDB^{+}20, PAF^{+}11, SCMAR^{+}99, TRLA^{+}13, Tur65, Whe93, dFHR^{+}24, vWHdS^{+}98, vWdSBdH02]. accuracy [SSM^{+}18, TSG^{+}04]. accurate [McD81a]. achaeta [GS19]. achievements [PV18, VBL^{+}21]. acid [CSG^{+}15, CDP14, GVKD^{+}13, HMPZ11, KSG^{+}17, LSV14, PAG^{+}18, PPHM18, RBPGJ^{+}20, TSRF14, TRP^{+}23, VKJ^{+}23, WOW^{+}14, YGL^{+}10]. acidification [AFBT^{+}22, BHK^{+}16, CKP^{+}20, JCM^{+}21, JCF^{+}23, KBHML17, MCL^{+}15, OHH^{+}22, SPK^{+}19]. acidity [GCV^{+}24]. acids [BPP^{+}98, GSVB23, GGA^{+}05, KO19, MRBS^{+}24, MFDH22, SBC^{+}16, SCCJ^{+}18, WPB^{+}08]. acknowledgement [Ano08y]. Acoustic [BLR^{+}23b, BCLD^{+}17, CS16, GIC20, ON22, BBLD^{+}11, BPSN^{+}21, CBB^{+}15, DHD^{+}23, DEW^{+}97, FRV^{+}19, MIN^{+}20, PBBH^{+}22, PO15, UPPS^{+}21]. acoustically [ZBLF23]. acoustically-tracked [ZBLF23]. acoustics [ATC^{+}19, BOG20]. acquisition [BK19]. Across [HFW^{+}98, ABSDC07, BAM^{+}09, BGB^{+}08, BWB^{+}09, BBL^{+}18, BGWP^{+}17, CLV^{+}19, CMJPH^{+}18, CLY22, CS06, CP07, FWH^{+}17, FLdST98, FMCG15, FJH10, FWL^{+}15, GDM^{+}20, GIC20, GHC^{+}17, GLAHH^{+}22, GGAA^{+}23, GHG^{+}24, GRdSS^{+}22, HTG15, HFPS^{+}06, HLTB^{+}17, JTD^{+}14, JLRB20, JSKM02, JMZ23, KHM^{+}88, LSV14, LGR^{+}02, LLGS21, LBC^{+}15, LB14, LHEB98, LvIKB07, MCD^{+}14, MRBS^{+}24, MPB^{+}23, MTL05, MVBC^{+}21, NRA^{+}21, OHC^{+}17, OCH^{+}18, ORMR^{+}19, PCSMC12, PLEF^{+}23, PO15, RBL^{+}19, SW01, SSV^{+}11, TOKLC08, VVV21, VDDA^{+}08, VKGP^{+}13, XWW^{+}21, YGMR^{+}23, ZYN^{+}24, ZMW^{+}23, dLLdAWL^{+}23]. across- [PCSMC12]. Across-slope [HFW^{+}98]. Acta [Ang86]. Actiniaria [vPRT90]. Actinopterygii [SM21]. action [IST^{+}88]. actions [CBB^{+}22b]. Active [SLBVRR^{+}22, GSA^{+}20, Kit03, PCC^{+}19, SWH^{+}24, VWDF14, YYK88]. activities [BTG^{+}03, GGQ07, NCC^{+}15, OvdSN94, RMB^{+}01, XHW^{+}20]. activity [BMNW01, CPB^{+}15, CM14a, CED09, DLM^{+}96, ECGP01, FPS^{+}09, GASV^{+}09, GGPG^{+}19, IPG^{+}16, LKDL14, LvBS^{+}24, LS12, MKHO96, MCGS^{+}16, PS98, PRA^{+}18, PBD^{+}88, RBS^{+}09, SAM^{+}04, SS03, VDDA^{+}08, ZHSMM14]. actuelle [Ber65c]. acuminata [DRVMC^{+}22, PMMN^{+}22]. acutorostrata [HBK^{+}24]. adaptation [ÁLC22]. Adaptations [Sma10a, Sma10b, KV18, THBA19, VJJ^{+}22]. Adaptive [PVC^{+}08, Sma10a]. ADCP [ADS^{+}22, BPC^{+}05]. addition [BPA^{+}21, TRP^{+}23]. Additional [MTL05]. Aden [BF12]. Adjacent [ALV^{+}21, AP20, ACL^{+}18, BEH19, BCGN^{+}18, BBRM20, CN22, CES^{+}19, CFML22, CTR^{+}19, DSB^{+}24, Dri11, ESTM13, FBR^{+}13, GM19, HDM19, KFC^{+}13, LSV14, PCD^{+}18, RCSVGP^{+}16, SSB19, SSKA19, SCB^{+}16, TCDPP^{+}22, TAF^{+}22, WGZZ19, WL16, YLL19, Yao88]. adjustment [MZGA^{+}20]. Adriatic [KK20, PL09, BBB^{+}14, BPC^{+}05, CDDF11, CD65, CFML22, KS15, MVC^{+}11, PVG^{+}20, ŠPM^{+}22, ŠGM^{+}18, SK21, SBC^{+}24, TAF^{+}22, UCB^{+}18]. adult [CCHV^{+}21, CCS^{+}21, DBM17, LMM03]. adults [dSSDS^{+}20]. Advancements [ALV^{+}21]. Advances [PHKS17, Ang88, BTK^{+}99, CDL19, Don87, FJhT^{+}14, IPF23, KA85, MSI17, SGPdM18, SFK^{+}99]. Advantages [BMG13, KSY^{+}19]. Advection [HDA^{+}16, SWZS^{+}21, ARDP14, BRE^{+}24, CBB^{+}22a, CNSHT15, FRK^{+}09, IAN13, SST^{+}17, TPPG10, WKS^{+}15]. Advective [Men21, KFC^{+}23]. advice [Ric01]. Aegean [TBK^{+}99, TPP^{+}00, BTK^{+}99, BTG^{+}03, LT06, PST^{+}15, RKM^{+}07, SKP99, VKT15]. Aeolian [FAAV^{+}15]. Aequiyoldia [GWGR^{+}19]. aeration [Tit20]. Aerobic [CWS^{+}21]. Aerosol [WH20]. aerosols [TPN^{+}18]. affect [GLV12, LBP^{+}21, MPM^{+}18, NGLSSG14, XRC^{+}15]. affected [HLPL05]. affecting [AH10, BSF^{+}21, FPIJ85, KSK^{+}15, Li14, NMC^{+}09, QCdS^{+}07]. affects [RFS10]. affinities [BS90]. Africa [AAM^{+}14, BAM^{+}09, GASV^{+}09, VFCC^{+}22, Ang79b, ABT^{+}04, BBPHG^{+}11, BWB^{+}09, Cra09, GMAMB04, MM80, Mit83, OAD22, SJP10, Tom81a, VNMS91]. African [Pai20, ARD^{+}03, BN03, LPARF^{+}20, Med87, Mit91, RMHL09, SAM^{+}04]. After [CBM^{+}21, IL20, LBH^{+}87, LCZ^{+}24, LWBD^{+}17, LGG18, TSFA22, YFY05, dJSL^{+}20]. again [Mun97, PPPdS20]. against [SOA^{+}23]. age [Emi65, Koc65a, OYKK^{+}23]. Aggregate [KSKN21, JB15, KVNT20]. aggregate-associated [KVNT20]. Aggregate-colonizing [KSKN21]. aggregates [RPG^{+}18]. aggregating [ICB^{+}19]. Aggregation [Gir15, KNSN^{+}09, MPSD15]. aggregations [RTN90, VBJ^{+}20]. AGI [CWS^{+}21]. Agulhas [OKdA^{+}19]. aid [SVIA14]. Air [FLUC08, LC22, WWL^{+}24, XCH^{+}16, dlPHF^{+}15, BBB^{+}14, CFG07, CF12, CEF^{+}13, EHSI12, GLLB22, LGZ^{+}20, OKdA^{+}19, Rou65, WBA^{+}22, WST^{+}16, YN20, ZPY^{+}20]. Air-sea [WWL^{+}24, dlPHF^{+}15, YN20]. Airborne [Ric94, HKL^{+}15]. akashiwo [MPMFL^{+}23]. alalunga [GA10, LOBG^{+}10, MRAP22, NXT^{+}17, XNT^{+}17]. Alaska [CHC^{+}12, CD07, CCH^{+}12, CBT07, CM18b, DPM^{+}09, FMM^{+}20, HHDS02, Kli10, LHC^{+}21, LHW^{+}20, MDAW^{+}19, MCL^{+}15, Mil93a, PCH08a, PCR^{+}22, STGR^{+}23, WM13, WW02, ZJZ^{+}21]. Alaskan [CQC15, Emi65, GAF15, PSM^{+}22, WO15]. albacares [MRAP22]. albacore [DAIS10, GA10, LOBG^{+}10, NXT^{+}17, PYKF15, XNT^{+}17]. albatross [CLB^{+}13]. albatrosses [KST^{+}10, Ric15, Ric22, Sac16]. Alboran [CLG^{+}22, FGR^{+}06, GR85, OMR^{+}22, SGLF^{+}13, SBPGP^{+}23, BPGC^{+}20, BPSGP^{+}23, MCGR07, Pre86, YHM^{+}18, VBVYT05]. alcalinité [Rot65]. Aldabra [WLKM10]. aldehydes [WPB^{+}08, WL16]. Aleutian [RLT^{+}22]. Alfonso [AAML22]. algae [Blu88, BLMR^{+}20, CML^{+}16, KSG^{+}17]. algae-produced [KSG^{+}17]. algal [IPF23, KGB^{+}23, KSC10, KHJ^{+}10, LMA^{+}15, MPMFL^{+}23, MP04, NNFL21, PP10, PFHM10, TPRS10, WDC^{+}11, ZBY^{+}22]. Algeria [OAD22]. Algerian [MTL05, TG05]. algorithm [DHL^{+}21, XYWY23]. algorithms [CGG08]. aliphatic [FTG^{+}18]. Alkaline [IPG^{+}16, SAM^{+}04]. alkalinity [Rot65]. alkanes [SOH21]. alkenones [STHM02]. allochthonous [LSV14, YTB^{+}21]. allometric [KSKN21]. ALOHA [KLB^{+}21, KBC^{+}22]. along [ACK^{+}13, Ang79b, ABT^{+}04, BJ17, BSMC15, BECR^{+}22, BBFS19, BHC^{+}18, BGWP^{+}17, BTV^{+}17, CGM^{+}02, CBB^{+}02, CdTH^{+}16, CCHM02, CLB^{+}14, CSG^{+}15, CB17, CBL^{+}19, CJG88, DN07, DHDM22, EBW^{+}23, EBR^{+}14, FAAV^{+}15, FGGDF^{+}04, GMAMB04, HCAFD^{+}20, HBH^{+}17, JJA^{+}17, JG07, KY23, KLC^{+}15, KMWF11, KGJ^{+}10, KRL^{+}22, LRAE23, LKDL14, LM00, LHEB98, LPARF^{+}20, MG02, Mar20, MTC14, MAAS^{+}00, MIN^{+}20, MB07, Mit91, MMPG07, NHN^{+}21, PGT^{+}13, PV07, PHFK14, PS98, PRA^{+}18, PCC^{+}19, PHC^{+}19, RFSCF19, RN02, SSB19, SSB20a, SH09, Sol00, SAd^{+}17, SST^{+}17, SJ02a, SJM^{+}19, STR01, SGR^{+}22, TTMM^{+}17, TCDPP^{+}22, TSRF14, WHT86, YMA^{+}17, dWDB^{+}98]. along-track [LKDL14]. alongshore [LMM03, PCSMC12, TMPM^{+}16a]. alpha [RTBR^{+}22]. alpha-diversities [RTBR^{+}22]. alter [GMR^{+}23]. alteration [MHVS19]. altered [Cra09]. alternative [OMS^{+}15]. alters [KZD^{+}19]. Altimeter [SJ02c, SJ02b, CGG08, TM13]. Altimeter-derived [SJ02c, SJ02b]. altimetric [CPSM20, KDL^{+}01, SJ02a]. altimetry [CED09, LS20, LKDL14, RM97, RBS^{+}09, WLM07]. Aluminium [DTKvH15]. Alvinella [JP90]. alvinellid [JG90]. Amazon [GdRGC^{+}14]. Amerasian [GCFS06]. America [CAH^{+}22, CBB^{+}02, CCHM02, JJS03, STR01]. American [JAS^{+}20]. amino [AB90, BPP^{+}98, GSVB23, GGA^{+}05, HMPZ11, VKJ^{+}23]. Ammodytes [RHBS13]. ammonium [FUOG^{+}16, MMF^{+}12]. AMOC [HMH^{+}15, SF15]. among [BL02, CMM^{+}04, DDK^{+}18, GSV^{+}01, HGB^{+}21, JJL^{+}19, SBL^{+}23, SPMVP05]. amongst [BGA^{+}21]. Amphan [PTZ^{+}23]. Ampharetidae [AP20]. amphipod [AE09, FJ19, HTV^{+}20, JZ19]. Amphipoda [Thu90, gWjNfLyD20]. amphipods [BS90, BTV^{+}17, HS22, HCV^{+}20, JM19, RJT84, Thu90]. amplification [LCZ^{+}24]. amplitude [ITO^{+}14]. amplitudes [DLM91]. AMT [AB00, BJ17, PHCA17, SWT^{+}17, ABD^{+}17]. AMT-19 [SWT^{+}17]. Amundsen [DHD^{+}23, FTG^{+}11, JYK^{+}14, SDL^{+}19, YLL19]. Amur [FMWW14, SMN^{+}14, YAI^{+}14]. Anadyr [NHN^{+}21, NAH^{+}21]. analogous [MHS^{+}09]. analyses [FTG^{+}11, KSG^{+}17, LSM^{+}22, SBM91, SM16, SDO^{+}14, YGL^{+}10]. Analysing [SSB14]. Analysis [ANMP15, ÁBMÁS15, BHHR15, BPSGP^{+}23, FBD18, FRCH15, FACM^{+}23, FHL^{+}24, MMGL^{+}07, MEMC05, TMKJ^{+}09, WHT86, YPM^{+}10, YYhT^{+}17, ÁBMÁS14, ASR^{+}20, BFPS06, BGB^{+}08, BD19, BPGD^{+}14, BBF^{+}22, CDDF11, CPG^{+}18, CHC^{+}12, CSK^{+}12, CP83, CHB02, CR20, DM13, DGVGR24, Dom84, Fuk91, GRDS10, GJ00, GTNK21, HKL^{+}15, HPS^{+}01, HSGJ23, Hob10, HSC^{+}16, IBW^{+}01, JE92, Mac98, MERB12, MNT14, MHR^{+}10, McK08, MGC^{+}18, MLHM09, MC88, MDR20, MCT03, Ola65b, PPVG12, PM13, PZA^{+}15, PTG95, Reb02, RCGC^{+}16, RDL^{+}91, SBMB18, SMN^{+}13, SF02, ŠPM^{+}22, SBFP21, SDJ14, SHS^{+}05, TSS^{+}12, Tom81a, Tom81b, VCM04, WD94, WDC^{+}11, YNMY23, ZNR^{+}24, MAH^{+}15]. analysis/forecast [CP83]. Analytical [PTF10, LPA^{+}11]. Analyzing [DYO^{+}10, LTJ^{+}15]. anammox [GMBU12]. Anatomical [YKWF21].

Anatomy [PKA19]. Anaximenes [GPP22]. anchoita [DMC^{+}18, dFKdLZTT17]. anchor [AVK91, BC91, CB91, CJMI^{+}91, MIW91, PWMIM91, Ver91, WP91]. anchoveta [BGB^{+}08, CGC^{+}20, FBM^{+}08, GRB^{+}08, PCSMC12, SMPC^{+}12]. anchovies [APC^{+}12]. Anchovy [IFC^{+}07, PVA24, YPGE^{+}10, AIA^{+}15, BCT^{+}09, BFB^{+}20, BDL08, BLT^{+}08, CMS^{+}13, EB08, FRCH15, FVLC^{+}23, GCD^{+}18, HLS^{+}14a, HPW10, JBB^{+}14, KYS^{+}17, MPB^{+}23, OACB^{+}15, PST^{+}15, RFC^{+}15, SGWF^{+}19, SAY^{+}16, SBG^{+}08, TOKLC08, TIOM16, VOG^{+}08, XRC^{+}15, YFY^{+}22]. and/or [HSG^{+}15]. Andaman [JFUR20]. Andvord [LWT^{+}20, ZCLS20]. anemones [vPRT90]. Angel [WR03]. Angeles [aHFS92]. Angola [GDSCU09]. anguillid [KMF^{+}20a, KMF^{+}20b]. anguilliform [FMC^{+}15]. angular [CR97]. animal [GDN^{+}18]. animals [BC01, Zez90]. Annaba [OAD22]. Annelida [AP20]. annelids [JG90]. anniversary [Ano85a]. annotation [GPAB^{+}16]. Announcement [Ano80a, Ano82a]. Annual [AMG^{+}16, CMHM18, Mol04, PMC16, RWD01, TSFA22, AMFY20, BCOL^{+}19, BM07, CLB^{+}13, CLL^{+}18, CRHM12, DZ04, DBM17, EHG^{+}07, FCN^{+}19, GLLB22, HDZY15, Her97, JSdSS^{+}21, KFKO03, LMPB^{+}16, LC12, LW13, MTC12, MCGR07, NRA17, OACB^{+}15, RM93, RGC^{+}01, RG09, SGS^{+}23, SNZ^{+}20, SGR^{+}22, SBC^{+}24, TM13, WSO01]. Annular [ZHD^{+}20]. Anobothrus [AP20]. Anomalies [BLAM00, ATS01, HZCZ16, HHSR07, Kat18, SMdG02, SJ02b, BLAM98]. Anomalous [ABP15, BASS^{+}20, GFGGD^{+}23, VDGGD^{+}22]. Anomalously [AAML22, TSFA22]. Anomaly [Leg91, KWI20, LBNBM13, SCHBC^{+}22, SPV^{+}15, SD07, UKM^{+}14, DMML88]. Anomura [Mar20]. answer [BBLD^{+}11]. Antarctic [FDH20, HWPLvW20, HGH^{+}19, KY23, LWT^{+}20, MD07, AMY^{+}23, BHA^{+}14, BH85, CdD^{+}15, CP07, DCD^{+}23, DSR21, DHDM22, dCFK17, GNH19, Hay65, HSH^{+}19, IIM^{+}23, IGG^{+}19, IG19, LS12, LS13, MWS^{+}10, NPO^{+}19, NHG19, OYKK^{+}23, OJB99, PVM^{+}20, SSTD^{+}95, TNGP22, TFZS14, VPM^{+}19, VKDS^{+}18, WCC^{+}20, YKH^{+}24, ZSI^{+}05, Zen08, ZCV^{+}19, ZHD^{+}20]. Antarctica [JYK^{+}14, SDL^{+}19, TSFA22, ZCLS20, AYH^{+}23, DCD^{+}23, GBB96, LSF^{+}17, SAT^{+}22, SDL^{+}19]. antarcticus [NPO^{+}19]. antennatus [CPG08, CHSB^{+}21, SCB^{+}09]. Anthropogenic [FGS^{+}15, aHFS92, JJ08, lLdZQ^{+}22, PPKR14, CPG^{+}18, DNNNN16, EMK^{+}17, GSF^{+}15, Har82, JJA^{+}08, KFC^{+}13, LPA^{+}11, OYKK^{+}23, PKV18, RJO^{+}19, RLDC^{+}13, RLR^{+}18, RAB^{+}11, dMGS^{+}11b, dMGS^{+}11a]. anthropogenically [MGC^{+}18, TCDPP^{+}23]. anti [XHC^{+}20]. anti-cyclonic [XHC^{+}20]. anticyclonic [CBB^{+}22c, MSS^{+}02, QYF^{+}24, RAE^{+}05, WOW^{+}14]. Antikythira [KHC^{+}99]. AO [NBR^{+}08]. AO-02 [NBR^{+}08]. Aotearoa [JRR^{+}24]. AOU [WMC^{+}89]. APECOSM [Mau10]. Aperture [VOT^{+}99, Ric94]. Apex [Mau10]. Aplacophora [BHB^{+}19]. apparent [GJ00, NMK^{+}03, XYGJ23]. applicable [Ano17i, Ano17j, Ano17k]. Application [CN22, GLS08, iIYO^{+}10, SSL08, Suk88, TGJT09, VBL^{+}09, WCB20a, WPH^{+}10, XD95, XYL^{+}22, AC85, HKK12, KM22, KAH^{+}16, Kvi69, LFG10, MCKS17, MAB^{+}11a, RWD01, TGR05, vRGW10, BPA^{+}21, GMDS20]. Applications [MLHM09, BMM97, BDE03, Dev87, KSY^{+}19, MZGA^{+}20, RLSF06, RLSF07, SHG12, WD94, ZD17, vFB82]. applied [BIL03, BBF^{+}22, CPSM20, GAS^{+}22, Man04]. Applying [BMG13]. appreciation [Ano65e]. approach [ANMP15, BMG13, CSBL^{+}15, CLX^{+}20, CRF^{+}10, DYO^{+}10, DLL^{+}23, DVL^{+}99, DFM^{+}15, DFH^{+}16, ESTM^{+}12, FVLC^{+}23, FGS^{+}23, GWM^{+}22, GNH19, GWS^{+}23, GD91, HMRB^{+}03, HHMB^{+}09, HNL14, HAP^{+}16, hHRW^{+}05, iIYO^{+}10, JPM^{+}08, KKK04a, KSE^{+}09, KSB^{+}22, LM18, LM14, LL21, LSD^{+}15, MPM^{+}18, MPB^{+}23, OWH14, ORVES17, PMMN^{+}22, PBO10, PVA24, RGB^{+}17, RSMIS03, SJP10, STM10, SEW11, SPN98, TMR^{+}21, TAM^{+}13, TS10, VSGC21, Val99b, VAEP24, WFS^{+}15, YPGE^{+}10]. Approaches [Ano09h, MHA^{+}11, CP10, FBA09, HAA^{+}14, KSK^{+}15, Man04, WMB^{+}18, dlGFM^{+}23]. appropriate [FCEZ10, GKC^{+}14]. April [Ano00e, Ano02a, Ano04b, Ano05a, Ano06c, Ano07n, Ano08p, Ano12m, Ano13n, Ano14m, Ano16t, Ano17r, Ano19n, Ano20a, Ano21a, Ano22a, Ano23a, Ano24a, FHP83, GMDD^{+}22b, MPD^{+}22]. Aptenodytes [STC10, STEB16]. Aqua [WM13]. Aquarium [GMD^{+}22]. aquatic [Pas22]. aquuatic [BVB88]. Arabian [PLN^{+}23, AJA^{+}22, ABS^{+}20, ADS^{+}22, BNC05, CBM^{+}21, Cow05, GCB^{+}22, LEDR^{+}22, LO85, MGG22, MB05, MKHO96, Men21, NBG^{+}05, PJS^{+}22, SDP^{+}22, SRAV19, SGS^{+}23, Smi05, SM05, SGR^{+}22, VJJ^{+}22, VKJ^{+}23, WHBK05, WGG^{+}08]. Arauco [CTF07]. Arc [KTB^{+}99, KSPK99, TBK^{+}99, VOT^{+}99, GSPMAI99, SKP99]. Arcachon [LDD^{+}22]. Arcane [MS17]. archaeal [CTR^{+}19, TVT^{+}23]. Archeological [IHY^{+}01]. archipelago [TSRF14, MIH06, PMG15, PCH^{+}08b]. Archipelagos [Ché14]. architecture [CARBML^{+}22]. archival [BMC^{+}10, FFT^{+}18]. Arctic [FTG^{+}11, FMCG15, JCM^{+}21, KPM^{+}23, MIH06, SHC^{+}07, YGMR^{+}23, BF11, GvOSW11, WXH07, AMFY20, AvD15, BBE^{+}15, BSC^{+}19, BHM^{+}15, BSF95, BvdLA^{+}11, BLP^{+}20, BRD^{+}15, BOG20, BKC15, BS95, BD18, BMG^{+}21b, CKB^{+}17, CML^{+}16, CKP^{+}20, CW06, CM11, CGZ^{+}16, CKT^{+}13, CDP14, CRPS^{+}15, CCW^{+}18, DWH^{+}14, DPB06, DWFP^{+}19, DRD^{+}07, DHD^{+}23, DLD15, DS65, DSV^{+}24, DWC06, EBW^{+}23, FSVL10, FWL^{+}15, GML^{+}23, GWK17, GSSWK20, GTNK21, GCFS06, GBC^{+}15, GSC^{+}20, HBG^{+}21, HMO^{+}13, HKGH^{+}06, HKE^{+}10, HHSR07, HK65, IPHW^{+}23, JAC^{+}12, JBH20, JPIP22, JLS^{+}22, Kiv97, KHBA^{+}24, KH09, KB65, LSH^{+}11, LMA^{+}15, LLH^{+}21, LRJ^{+}15, MGWZ20, MVN^{+}15, MRO^{+}08, MOS^{+}13, MHA^{+}11, MMKS^{+}21, MBH^{+}01, MKSvA^{+}22, MHH^{+}15, MS15, MHVS19, NYH^{+}22, OWR^{+}07, OV24, PTM^{+}22, PAPL15, PFW15, PNF^{+}21, RJO^{+}19, RCB^{+}20, RSB^{+}15, RvBD^{+}22, Rud15, RKS^{+}15, RN06, SSB19]. Arctic [SBMB18, SBK^{+}95, SKWWGV18, SW92, SON^{+}20, SEW11, SHC^{+}06, SWZS^{+}21, SLY^{+}15, SPW22, Tit20, TAM^{+}15, VMB^{+}22b, VMH^{+}21, WMC^{+}89, WDC^{+}11, WO15, WHS17, Was06, Was11, WKS^{+}15, Was15, WCB^{+}20b, WZBK^{+}21, WC15, WHS^{+}23, WBD^{+}15, Woo18, YSD15, YWUK15, ZJZ^{+}21, ZBRJ23, ZWP23]. Arctic-Atlantic [PNF^{+}21]. Arctic-FVCOM [CGZ^{+}16]. Arcto [LNB13]. Arcto-boreal [LNB13]. area [AS20, AMEV07, BATNP04, BC19, BM07, CN22, CGD^{+}18, DMC^{+}18, DOP87, EAB^{+}23, FTSF21, Fra69, GCCY^{+}14, GKR20, GMAB07, HYM^{+}12, His22, IVT^{+}12, JSLA^{+}21, KKS^{+}19, KTW^{+}22, LWBD^{+}17, MA20, Mit83, MMF^{+}12, MDC^{+}07, MA12, MSL^{+}07, NH83, PPSVC^{+}13, PFE10, SVL^{+}23, STS^{+}12, VSGC21]. areas [AR18, BJMP19, BJMP20, BFV^{+}17, CDTM^{+}21, CMM^{+}04, CQO^{+}15, GGA^{+}05, GAS^{+}22, GRB^{+}08, HMP^{+}13, HLP^{+}16, HvDL^{+}17, KLC^{+}15, KFH^{+}15, KMF^{+}20a, KMF^{+}20b, LT06, LBH^{+}21, LLS01, ROBRB^{+}22, SFS^{+}12, SC65, Sie69, SAY^{+}16, SPN98, YYK88, dFKdLZTT17]. Argentina [ILA21]. Argentine [FGS^{+}23, GGE^{+}65]. Argo [RBS^{+}20, CM18b, FC05, MMF^{+}17, PL18, RBS^{+}22, RG09, WCX^{+}21]. argonauts [AG22]. arising [BAP^{+}22]. Aristeidae [CHSB^{+}21]. Aristeus [CPG08, CHSB^{+}21, SCB^{+}09]. Armorican [PL89]. aromatic [FTG^{+}18, SGL^{+}17]. Array [Ben85, BBL^{+}09, GW89, LPF^{+}20, SCY^{+}23, YHZ^{+}22]. arrays [SSM^{+}18]. arriving [SL13]. Arsenate [HSK^{+}19]. arsenic [TCDPP^{+}23]. art [CAA^{+}07, MPC12, SRM^{+}10, SBB^{+}14]. Arthropoda [PKA19]. article [SDS22a]. Artificial [BVJE19]. Asellota [BBFS19, GM19, MB20]. asellote [SBS90]. Asia [SSVP00]. Asian [LMW^{+}12, FJhT^{+}14, KJZ^{+}12, PO00, Qiu15, SOB^{+}08]. Ask [MKB00]. Aspects [TFZS14, BHS^{+}15, Bre06, MP04, Ola65a]. assay [AIA^{+}15]. assemblage [BCM^{+}02, Dol09, GSA^{+}20, HSL96, LMPB^{+}16, MDAW^{+}19, RSW^{+}23, RMHL09, VMB^{+}22a]. Assemblages [DN07, ACL^{+}18, AMEV07, ALG^{+}21, ATC^{+}19, BECR^{+}22, BM07, BJ90, CSV^{+}07, CLD22, CTR^{+}19, CCB^{+}20, DTOD00, DWS^{+}24, DSC^{+}19, DMBB02, EMBS13, ERT^{+}22, FTC^{+}16, FSVL10, Gal17, GIHJ23, GBB96, GRD^{+}23, GGA^{+}16, GGG^{+}18, GGSM^{+}20, GDM^{+}15, GLV12, HDM19, HTV^{+}20, KSVT00, LSW^{+}21, MST^{+}23a, MDGC^{+}12, MTH^{+}10, NFMCS^{+}22, PGGG17, ROBRB^{+}22, RLGC10, RTBR^{+}22, SPF^{+}23, SBG16, VDP^{+}01]. assembly [CGS23, VGM^{+}23]. assess [SAB^{+}22]. assessed [MERB12, RSM^{+}23, SAT^{+}22]. Assessing [JF13, LM14, LDHW20, SGL^{+}18, SMPC^{+}12, dlGFM^{+}23, EBS^{+}18, KSE^{+}09]. Assessment [ÁBMÁS14, ÁBMÁS15, AKAL20, JST^{+}24, SBH^{+}14, SOH21, CSBL^{+}15, CAO^{+}20, CLB^{+}14, EMU21, FC05, GBC^{+}00, HFS^{+}20, HHMB^{+}09, Hof10, JPBB20, KMB01, KKO10, LCBN14, Li14, MCL^{+}15, ORPRGIS22, PSP^{+}21, SNS^{+}22, WPH^{+}10, ZHSMM14, ZL01]. assessments [GZCL23]. assimilating [REG^{+}15]. Assimilation [CTMV^{+}14, KDL^{+}01, GBM^{+}01, GAM98a, GAM98b, GA00, JRW01, Kiv97, KNI^{+}05, MAB^{+}11c, MAB^{+}11a, MAB^{+}11b, MZGA^{+}20, SO91, WSO01, Whe93]. assimilative [WLM^{+}22]. assist [CN22]. associated [BM76, BJMP19, BJMP20, CGM^{+}02, CCA^{+}02, DLD15, EMBS13, FBR^{+}13, FRV^{+}19, FKH^{+}13, GGT^{+}15, GS19, GBC^{+}15, GPC^{+}03, Gri22, GLV12, Jón07, KAK^{+}22a, KGB^{+}23, KVNT20, MZH^{+}23, MS17, MPM^{+}18, NAH^{+}21, PKP14, PVA24, Sch83, ŠGM^{+}18, SKH00, TMN^{+}12, VDP^{+}01, ZAC^{+}23]. associations [DHD^{+}23, HFO90, MDR20, WMB^{+}21]. Assumptions [KSY^{+}19, BMG13]. Asteroidea [MJD^{+}21]. Aswan [Ore69]. asymmetry [GXX^{+}22, HZCZ16]. Atacama [FAAV^{+}15]. Atlantic [ABD^{+}17, ÁSÁB^{+}14, ÁBMÁS14, ÁBMÁS15, AKAL20, ALT10, BGMP03, BLHB07, BMdMS^{+}21, BLAM00, BBR^{+}01, BS02, CSV^{+}07, CKM^{+}21, CSC^{+}12, DML^{+}16, DMC^{+}18, EvdZSH02, FPD^{+}01, FGS^{+}15, FMH02, GMDD^{+}22b, GTB07, GSF^{+}15, HØ00, HBV^{+}10, HBR11, HCV^{+}20, IBW^{+}01, JLP^{+}20a, JG07, KSR^{+}01, KAH^{+}16, LS20, LMT^{+}19, MHS^{+}20a, MMF^{+}17, MRH^{+}14, NJCD01, OKdA^{+}19, ORMR^{+}19, PMG15, PS91, RWD01, RSB^{+}01, RFFL21, Rei89, Rei94, RHM^{+}19, STPHM^{+}23, SLOP^{+}22, Tom81a, TRMV15, UB10, VDP^{+}01, WWL^{+}22, WHS17, WLM^{+}22, WG82, vAB96, AHP19, AQVB^{+}10, AS96, AdAK^{+}18, Ang79a, Ang79b, Ang84, ABSDC07, AGL^{+}15, ASR^{+}20, APP21, BM76, BDTC15, BGMP03, BNCC15, BEH19, BLAM98, BSC^{+}07, BGM^{+}99, BMNW01, BKD^{+}20, BAOC^{+}07, BHS^{+}15, BHC^{+}18, BCGN^{+}18, BRG^{+}23, BHPC06, BDBJ01, BSH^{+}20, BMG^{+}21a, BTJ^{+}17, BMG^{+}21b, BGS^{+}04, BHMS09, BGWP^{+}17, BTV^{+}17, CRGA17]. Atlantic [CSM^{+}15, CPG^{+}18, CGV13a, CB06, CHG^{+}18, CMJPH^{+}18, CPC^{+}15, CPHR98, CBB^{+}22c, Cia14, CRF^{+}10, DHC^{+}20, DN07, DBC^{+}18, DL69, DGMM85, DMML88, DLM^{+}96, Dom84, Dri06, DP13, DFH^{+}16, DK07, ECGP01, EBM^{+}20, EBM^{+}21, Eme65, ED82, Emi65, EMU21, EMK^{+}17, FGSA97, FAAF88, FHG03, FCMCÁS19, dCFK17, FBT^{+}22, FWH^{+}17, FKZ^{+}15, FMP19, FJH10, FKH^{+}13, FMSBW13, FW91, FMW91, GSV^{+}01, GIPC^{+}15, GMAGH^{+}17, GSM^{+}17, GBC^{+}00, GdRGC^{+}14, GHC^{+}17, GLAHH^{+}22, GGAA^{+}23, GA01, GD85, GTNK21, GPC^{+}03, GAM98a, GAM98b, GA00, GEPC15, HLR17, HØH^{+}03, HMRA^{+}03, HBL^{+}13, HBD^{+}21, HPB^{+}09, HLM^{+}16, HDZY15, HMP^{+}13, HGPFN^{+}14, HGTP^{+}19, HMPZ11, HAA^{+}14, HHSR07, HMH07, HDB13, ICB^{+}19, JSdSS^{+}21, JLB^{+}08, JLP^{+}20b, JST^{+}24, JCM^{+}21, JPIP22, KSV08, KGL22, KKNT23, KY15, KDL^{+}01, KBHML17, KT04, KJH^{+}22]. Atlantic [Kos93, KHP^{+}18, KFM15, KVLA06, LBK^{+}01, LM00, LFC^{+}15, LLH^{+}21, LNB13, LAGM^{+}23, LB14, LvIKB07, LMP22, LOC95, LBF^{+}22, MMGL^{+}07, MHS^{+}20b, MHR^{+}10, MSJ^{+}15, McC92, MFS^{+}16a, MFS^{+}16b, McG64, MBH^{+}23, MAAS^{+}00, MIN^{+}20, MBF^{+}14, MRW^{+}14, MHCS^{+}23, MFM85, MGK^{+}86, Mol04, Mol22, MST^{+}23b, NCH^{+}07, Nof96, Nof00, NGNV12, Ola65b, OHC^{+}17, OCH^{+}18, OMS^{+}15, OÁSG^{+}16, Owe91, PTM^{+}22, PMM^{+}23, PVC^{+}08, PFHM16, PL01, PPHM18, PMDR06, Peñ24, PJH^{+}15, PMH17, PGGG17, PP85, PGC^{+}96, PHCA17, PNF^{+}21, Pug84, PGG^{+}22, QPR03, RGC^{+}01, Rea00, RBR^{+}23, RDL^{+}91, RCD^{+}94, RKFD07, RTN90, Ric85, Ric93, RAP95, RAB^{+}84, Roe84a, Roe84b, RB84, RJT84, RFKC16, RFPG15, RFS10, RR01, RKS^{+}15, RKC^{+}10, RDP^{+}21, SCLG^{+}11, SHK^{+}14, SF85, SIR^{+}07, SFMT12, SFMT14, dSSDS^{+}20, SWT^{+}17, Sme93, SQJ^{+}17, SNdSR^{+}24, SBD01, SJD10, SWZS^{+}21]. Atlantic [SBE^{+}20, SPW22, SPK^{+}22, SD07, THBA19, Thu90, TŠT^{+}17, TLM^{+}17, THM^{+}14, TBW00, Tsu86, VBL^{+}21, VMB^{+}22a, VHV^{+}12, VSC01, VCSG^{+}01, VPW01, VDDA^{+}08, VFS^{+}15, VBJ^{+}20, WMB^{+}21, WLD^{+}15, WSO^{+}13, WHBW03, WBB^{+}01, WDK^{+}01, WWSJ07, Yas07a, YSD15, YS15, YSN20, ZAC^{+}23, ZLZ^{+}17, ZBLF23, dPGSHL23, dlPHF^{+}15, ARH^{+}00, AB00, Ano17a, BJ17, PHCA17, RNP^{+}17, SWT^{+}17, ZSBL00]. atlantica [BLCL14]. atlanticum [OBD^{+}20]. atlantid [WPBG^{+}18]. Atlantification [WHS^{+}23, MGWZ20]. ATLANTIS [LFG10, KHL12]. Atlas [Ang79a, CBB^{+}02, TTL^{+}04, FMW91]. atmosphere [Bla63, LSW02, MLL^{+}22, NBG^{+}05, Pie01]. atmosphere-ocean-ecosystem [MLL^{+}22]. Atmospheric [AALM06, BJ17, DL69, DVB^{+}18, LGG18, PSM^{+}22, SBPGP^{+}23, ZPY^{+}20, BDTC15, DCL^{+}13a, GCD^{+}99, HDZY15, Kaz17, LZL^{+}22, MM99, OAB^{+}16, OAD22, RMK^{+}21, RDD^{+}18, RHM^{+}19, SMdG02, TPN^{+}18, VBA^{+}18, XY21]. Atmospheric-Driven [SBPGP^{+}23]. atmospherically [CPSM20]. atmospherically-forced [CPSM20]. Atoll [HWF^{+}21, PS23, RHB23]. atolls [PCH^{+}08b]. atoms [Ros65]. ATP [GGQ07]. ATP-P [GGQ07]. attenuation [KM08]. August [Ano13g, Ano22-29, Ano22-30, Ano98c, Ano99c, Ano00g, Ano08r, Ano09m, Ano12n, Ano16o, Ano17n, Ano20b, Ano21b, Ano22b, Ano23b, HFW^{+}98, JRW01]. Auks [KGJ^{+}10]. aurita [KLP^{+}17, TCF^{+}18]. austral [AMY^{+}23, MST^{+}23a, SSN23, YLL19]. Australia [DPF^{+}20, DWS^{+}24, Hob10, LHF^{+}16, MB07, TTF^{+}22, WOW^{+}14]. Australian [Her97, HT97, ORCH^{+}19, RD03, KY23, MD07, NC80, SOB^{+}08]. australis [FCN^{+}19]. Author [Ano65a, Ano65b, Ano69a, Ano73a, Ano85b]. auto [MSMH19]. auto- [MSMH19]. autocorrelation [BAOM^{+}12]. Automated [DHB^{+}21, MERB12, SPH^{+}15b]. automatic [iIYO^{+}10]. autotrophic [BLP93]. autumn [BC88, CWB^{+}22, EMU^{+}23, HBG^{+}21, KFC^{+}23, NST^{+}23, PFHM16, SSV^{+}11, WSH^{+}22]. Auxis [KTIT22]. availability [CSV^{+}07, FVLC^{+}23, JJL^{+}19, KZD^{+}19, LPF^{+}18, OMR^{+}22, OIC^{+}23, VMC^{+}19]. available [MBCB88, MFDH22]. average [Tur65, Wun24]. Avilés [RCSVGP^{+}16]. avoid [LPHL^{+}05b]. Avoiding [AF10]. avoids [McK08]. awakening [LMA^{+}15]. AXBT [WLM07]. axial [CSR90]. Azores [Ang89, CGMP14, FPIJ85, Gou85, NJCD01, PP85, SGMP15]. Azov [FBS^{+}18, KAG^{+}19].

B [Ang80]. back [PPPdS20]. background [BCF^{+}03, HMH^{+}15, Pie01]. backscatter [ADS^{+}22, BPSN^{+}21, MIN^{+}20, PBBH^{+}22, UPPS^{+}21]. backscattering [BLR^{+}23b, SBBV04]. backtracking [TMR^{+}21]. bacteria [FCG88, GMBU12, JP90, MY23, Sie88, SHS^{+}05]. Bacterial [ASC92, DMD^{+}00, ECGP01, BFJ18, CC88, CTR^{+}19, DDE^{+}95, DGP^{+}13, GMAGH^{+}17, GMDD^{+}22b, GMDD^{+}22a, HLR17, Her88, RMB^{+}01, SST^{+}17, SWH^{+}24, TAW^{+}15, TVT^{+}23, VCB^{+}00, ZKK^{+}16]. Bacterioplankton [BSMC15, BGM^{+}01, DLM^{+}12, DDP^{+}00, EGPM^{+}15, MDC^{+}07]. Baffin [MFM15, TRY^{+}04]. Bahamas [Ché14]. bairdi [RKCH15]. Baja [BGM^{+}10, DB02, Dur09, JOGM^{+}10, LAA12, LABD^{+}24, PELAA18]. Baker [RBL^{+}19]. Balaenoptera [HBK^{+}24]. balance [BSF95, HPZC21, LHP^{+}05, RGC^{+}01, Tur65, YTNK00]. balanced [Mau10]. balances [AHW99, BS95]. Balancing [SCB^{+}07, Fly03]. Balearic [BZD^{+}21, CHSB^{+}21, AQVB^{+}10, CPG08, OMR^{+}22, PTG95, PLJR22, dPAJ07]. Baleen [MHVS19, GVBV^{+}21]. Bali [SNR^{+}10]. ballast [SL13]. Ballenas [SRFHDH22]. BALTEX [OELP04]. Baltic [APC^{+}12, BMC17, BLC23, BBF^{+}22, CS18, HVRR15, HKPV12, HLP^{+}16, HvDL^{+}17, HCGK11, MHTG10, OELP04, OEL^{+}14, PBB^{+}12a, PDV12, PBB^{+}12b, PHFK14, RBF^{+}09, SPB^{+}12, Seg69, Sei63, TMKJ^{+}09, VH09a, VH09b, VPH^{+}12, VHK03, VHK04]. bamboo [PRA^{+}18]. bamboo-coral [PRA^{+}18]. band [SMFM^{+}21]. Bank [CSV^{+}07, HØ07, JJA^{+}17, RCSA01, CHC^{+}12, ESA^{+}13, LTSG13, SSI13, TSP^{+}13, BBL^{+}09, DGP^{+}13, EMBS13, GGJ^{+}10, LTSG13, MEST13, PIS13, SCB^{+}07, SEO13]. Banner [Ang80]. Barbados [JFG^{+}90, OSH^{+}96]. Barbara [AHW99, SC23, SPB^{+}02]. barbatus [MSC^{+}15]. Barcelona [SCMAR^{+}99]. barcoding [JM19]. Barents [EBW^{+}23, AEP^{+}23, AAMB^{+}24, AOMZ^{+}23, BSF^{+}21, BRG^{+}23, CIL^{+}23, CGV13a, CGV13b, DM13, DAvD^{+}20, Dol09, Dri11, DCL^{+}13a, EFC^{+}23, ESGP17, EGP^{+}18, EBD^{+}20, ESD^{+}21, HBL^{+}13, HBK^{+}24, JCF^{+}23, JMSB^{+}23, KFC^{+}23, KMS^{+}24, KSB^{+}22, KGB^{+}23, LJM^{+}16, LNB13, MGA^{+}23, RCS^{+}11, SHP^{+}23, SDH^{+}14, SAB^{+}21, SEG^{+}22b, SEG22a, SDO^{+}14, TVT^{+}23, Tit20, WRH^{+}06, WHS^{+}23, YS15, dFHR^{+}24, dlGFM^{+}23]. Barents/Norwegian/Greenland [HBL^{+}13]. Bari [CFML22, SCC^{+}19]. Barkley [CLSD18, CMHM18, DOS^{+}18]. Baroclinic [RZW^{+}23, CSLJ03, Hog85, HNSP^{+}19, KT97]. barotropic [DEW^{+}97, Hut87, LL97, RZW^{+}23, Sak86]. Barrier [BECR^{+}22, GHC^{+}17]. barriers [JLRB20, TKC^{+}22]. Barrow [HFO^{+}22]. bartramii [IIS^{+}17]. Base [CLdPHL23, KH09, SJP10]. Based [YN20, BEP02, BSH^{+}20, BMN19, BGV^{+}23, BLT^{+}08, CGMP14, CMS^{+}13, CMG15, CHC^{+}12, CSK^{+}12, CKT^{+}13, DPR^{+}18, DSBP15, DHL^{+}21, DPF^{+}20, ERBV21, FTSF21, FFA06, Fuk91, GSFP^{+}09, GRMB18, GNH19, GBH^{+}20, HSS^{+}12, HFS^{+}20, HMX^{+}23, JPM^{+}08, JAC^{+}12, JHDT12, KGL22, KSS^{+}23, KPSB22, LLH^{+}21, LMT^{+}19, LB20, NGLL^{+}22, PCSMC12, PMMN^{+}22, PGS^{+}22, Pra91, SGMVF14, SPSV^{+}20, SOA^{+}23, SGR^{+}22, TAM^{+}13, VVV21, VSPP14, WZFW16, WCX^{+}21, WSS15, WFJ^{+}15, WLM07, WPB05, WJPHB15, YWUK15, ZL01, ZZM^{+}24]. Baseline [JLS^{+}22, EBD^{+}20, MRAP22, OPH^{+}24, SKH^{+}23]. basic [Ken88]. Basin [AAML22, GGAA^{+}23, Hic92, HGBG20, IMM^{+}22, KKS^{+}19, LXC^{+}22, MPCNC^{+}19, UB10, AVS23, AIA^{+}18, BMK12, DSC^{+}19, GBC^{+}00, GPC^{+}03, HMTL05, HS07, IHT^{+}21, JLP^{+}20a, JLP^{+}20b, LBH^{+}21, LH08, MDB^{+}20, MZGA^{+}20, MJA^{+}07, NGLSSG14, PBB^{+}20, PKV18, SSL08, SiSI^{+}02, SPB^{+}02, SSM^{+}18, SSW^{+}09, TCDPP^{+}22, dlPHF^{+}15, SBB^{+}14, ABM^{+}05, BSW86, BS95, BPTT19, CMPNC^{+}22, CJ92, EFC^{+}23, FGS^{+}23, GGE^{+}65, Hau18, HHW01, HHW22, HKE^{+}10, JMSB^{+}23, KZSH85, KY23, LTG85, LSV14, LPA92, MHGP06, MD07, NHN^{+}21, Ore69, ÖHÜ89, Rot65, RKS^{+}15, SBK^{+}95, SPB^{+}12, SE92, TG05, VK92, WWN^{+}99, WRS^{+}92]. Basin-scale [GGAA^{+}23, HGBG20, IMM^{+}22, UB10, BMK12, GPC^{+}03, IHT^{+}21, JLP^{+}20a, JLP^{+}20b, MJA^{+}07, NGLSSG14, SSL08]. basin-wide [PKV18, SSM^{+}18, CMPNC^{+}22]. basinal [YNTS22]. basins [BHB^{+}19, BKC15, FWL^{+}15, aHFS92, JJL^{+}19, YSD15, Gor92, WJE^{+}92]. basis [vdS94a]. bassanus [SWP^{+}13a]. bassin [Rot65]. batch [TGR05]. batch-dissolution [TGR05]. Bathyal [SW21, Car98, DMD^{+}00, HFO90, HWBT03, SS03, TPM^{+}00]. Bathymetric [CGM^{+}02, GSSWK20, KKKS14, ACB^{+}13, BHS^{+}15, CMM^{+}04, FBD18, MMPG07, VMB^{+}22b]. bathymetries [Pra04]. bathymetry [LW13, STEB16]. bathypelagic [CRC^{+}19, PMFNGQ21, SLOP^{+}22, ZPC^{+}16]. Bathysquillidae [MKD90]. bathythermograph [Mol22]. Bay [CÁM06, CLMR23, FB05, HL05, HJLLN07, HPHL^{+}05, KKS^{+}03, KNS^{+}03, LZCZ05, LWT^{+}20, MGS90, NMK^{+}03, NKK03, OAD22, PHLL05, PLHLF05, RVS^{+}21, TWMY08, TSFA22, iUMY86, VOG^{+}08, VGM^{+}23, VK90, ZCLS20, BLT^{+}15, KOhL^{+}10, LDD^{+}22, USH15b, XYK^{+}22, AIA^{+}15, ADV^{+}18, ALT10, BCGN^{+}18, BHLU^{+}07, Bre06, BhTW10, CDS90, CSC^{+}12, DBC^{+}18, DPH^{+}18, DPR^{+}18, DHHP18, ERT^{+}22, EHG^{+}12, GCD^{+}18, GHL15, GA10, GRD^{+}23, HBV^{+}10, HCC02, HLD^{+}21, HPW10, IFC^{+}07, JJA^{+}13, JX18, JFUR20, KFKO03, KSP^{+}23, LAD^{+}18, LLL^{+}11, LCBN14, LML^{+}23, LOBG^{+}10, LSIC12, LSIB23, MMR^{+}12, MCG^{+}02, MLD^{+}03, MFM15, MJA^{+}07, MFH86, NP00, NMN08, OMS^{+}09, PMFNGQ21, PTZ^{+}23, PGRP^{+}18, PVV23, RCSVGP^{+}16, SMN^{+}13, SBL^{+}23, SS03, Soh03, SRT^{+}18, TMN^{+}12, TRY^{+}04, TFM03, USH15b, VLUC^{+}07, VDB^{+}20, VJJ^{+}22, VBM21, VKJ^{+}23, WCB20a, WPH^{+}10, XWW^{+}21, ZHF^{+}24]. Bay [Ang80]. Bayesian [GL23, OWH14, PVA24]. bays [HGH^{+}19]. BBMO [VGM^{+}23]. BCB [CQO^{+}15]. be [QSC^{+}15]. Beach [Let87, SCS87]. Beagle [CAH^{+}22, ILA21]. BEAGLE2003} [AFH^{+}11, KMWF11]. beaked [SGL^{+}18]. beam [ON22]. bearded [CQC15, MSC^{+}15]. Beaufort [CQO^{+}15, CBB^{+}15, LPF^{+}21, BD18, BPM^{+}14, CDP14, DLD^{+}19, DWC06, FMCG15, GDL^{+}15, HSG^{+}15, KFH^{+}15, LPW^{+}23, LBC^{+}15, MSC^{+}15, NNFL21, OACA20, PSM^{+}22, WLCG23]. bed [DXH^{+}02, HHK^{+}02, WAH^{+}20]. been [MKB00]. beer [GAF15]. before [LBH^{+}87]. behavior [CdTH^{+}16, HHP10, JBB^{+}14, MPM^{+}18, PO15, SBLA10, SK17, TNS^{+}05]. behavioral [BVJE19, BAP^{+}22, KSK^{+}15]. behavioral-hydrodynamic [KSK^{+}15]. behaviors [CZG^{+}21]. behaviour [ASC07, BLP^{+}20, BGB^{+}08, DCM16, DMBHG10, ESA^{+}13, FDB^{+}21, LSF^{+}17, MKOLA20, NRS^{+}19, RHB23, RAE^{+}05, STC10, STEB16, Sma10b, ST10]. behavioural [QCdS^{+}07]. behind [BCD^{+}20, DLL^{+}23, DMT15, LMA^{+}15, OMS^{+}15, SD07]. belly [GAF15]. belonging [CSH^{+}23]. belongs [RK20]. below [BHK^{+}19, OT19]. Beneath [BH07, MSL^{+}07]. benefits [BPA^{+}21]. BENGAL [BR01, RWD01, JFUR20, KSP^{+}23, MMR^{+}12, MJA^{+}07, MFH86, PTZ^{+}23, PVV23, RVS^{+}21, VJJ^{+}22, VBM21, VKJ^{+}23, XWW^{+}21]. Benguela [AH80, AVK91, AE09, BC91, BHAJ12, CB91, CJMI^{+}91, CS04, DBR03, FUOG^{+}16, GDSCU09, HMRB^{+}03, HSC09, HVEF09, HvdLS^{+}09, MIW91, NH83, PWMIM91, RBS^{+}09, SE16, SKRM^{+}95, TFM03, TS10, VPS09, Ver91, WP91, ZHBW01]. Benthic [BHE^{+}98, BRC^{+}18, Car98, CBL^{+}19, DDDT99, DTW^{+}00, FJA^{+}21, GBB96, HFO90, HG04, JPIP22, LTSG13, LGR^{+}02, RBL90, TDH^{+}95, ZHSMM14, vWM02a, Ano94k, BBMR19, BTG^{+}03, BD18, CLSP17, CMHM18, CSG^{+}15, CDP14, CTR^{+}19, DLL^{+}23, DGVGR24, DL17, DBJ^{+}15, DBR20, FLdST98, GRMB18, GPP22, GvOSW11, GD85, GWK17, GLV12, GEP^{+}08, HLS^{+}14b, IBW^{+}01, JPBB20, KGdS^{+}08, KLC^{+}15, KRHS14, LJM^{+}16, LRJ^{+}15, MGS90, MBH^{+}23, MKSvA^{+}22, MDR20, MRW^{+}14, NRA^{+}21, OB98, PPHM18, PRC^{+}20, PS98, QSC^{+}15, QOS^{+}22, RGC^{+}01, RCC^{+}18, SS03, SBG16, TAW^{+}15, TSG^{+}04, TvW98, TvG02, VKDS^{+}18, WLP^{+}21, ZCV^{+}19, ZWM^{+}15, ZBRJ23, vWMH98]. Benthic/midwater [RBL90]. benthivory [GBC^{+}15]. benthonic [Phl65, Sai65]. benthopelagic [BC01, GD85]. Benthos [CSR90, GSSWK20, JSHB90, RSB^{+}15, RSD^{+}90, VMB^{+}22b]. bentincki [CCM^{+}14, SYB^{+}15]. bergii [Cra09]. Bering [MLPN06, ANH21, AIHB^{+}07, AT07, BE99, BDC^{+}08, CQO^{+}15, CBB^{+}15, CHB02, CHS^{+}24, CP02, CQC15, DWH^{+}14, DBC^{+}23, FJA^{+}21, FMCG15, GCFS06, GTS^{+}21, HKN^{+}14, HOY^{+}21a, HMH^{+}15, HS02, ISM^{+}02, IAFD02, LDAM^{+}07, LCJ^{+}17, LSW02, MSC^{+}15, MRSS02, MOSN^{+}13, Min02, MSS^{+}02, MWFH02, NSE^{+}24, NHN^{+}21, PDAM^{+}15, PST^{+}02, RKS01, RKCH15, iSIS02, SOH21, STHM02, SiSI^{+}02, SNMW10, SYN^{+}21, TFY02, WFH^{+}22, WMC^{+}89, WD94, WDC^{+}11, WHI^{+}02, Woo18, YNM^{+}02, ZK06, ZWP23]. Bering/Chukchi [WMC^{+}89]. Bermuda [Ang79b, SS69]. Bernard [CRiI^{+}15b]. better [CRiI^{+}15a]. between [ALT10, BLP^{+}20, CCS^{+}21, CRF^{+}10, CBL^{+}19, CQC15, CTI^{+}19, DDDT99, DP18, DLD^{+}19, DCL^{+}13a, DL17, FMC^{+}20, GVBV^{+}21, GBC^{+}00, GA01, GDM^{+}15, HL05, HBK^{+}24, HJLLN07, HFW^{+}98, hHCK01, HBH^{+}17, HLS^{+}14b, HM06, IIS^{+}17, KF11, KKKS14, KTIT22, KAK^{+}22b, LSF^{+}17, LDB^{+}02, LHP^{+}05, LOBG^{+}10, LLL^{+}24, LB20, LHC^{+}19, LMP22, MSC^{+}15, MPV12, MNT14, MMF^{+}07, MSd^{+}16, MCG^{+}14, MM90, NYH^{+}22, PAM^{+}88, PL09, RFFL21, RGPB^{+}23, RLR^{+}18, SMFM^{+}21, SCB^{+}09, SRF^{+}19, Val99a, VCM04, VDB^{+}20, VHK03, VHK04, WNNI21, WSH15, XYGJ23, XC14, YFK21, Yu23, ZL01, ZSBL00]. between-region [MPV12]. Beyond [MBH^{+}01]. BGC [WCX^{+}21]. BGC-Argo [WCX^{+}21]. Bi [LDD^{+}22, BCLD^{+}17, OMS^{+}15, RNBP^{+}19]. Bi-decadal [LDD^{+}22, OMS^{+}15]. bi-frequency [BCLD^{+}17]. bi-phasic [RNBP^{+}19]. bias [CDB^{+}24, KSK21, MRH^{+}18]. bibliography [Ano65f, SMB88]. Bifurcation [Sak86, CF07, DCD^{+}23]. big [MVV^{+}19]. bigeye [HLTB^{+}17, HHP10, LSS^{+}10]. Bight [AG22, ASC92, BHPC06, BTS22, CB06, DIM09, Epp92, Ham87, Her97, HT97, KMMC09, KC15, KVLA06, LLS01, PG10, WWL^{+}22, WLM^{+}22, ZAC^{+}23, dFKdLZTT17]. Bill [SSB^{+}20b]. billfish [McI10]. Bio [GNH19, TII^{+}14, BDB^{+}04, BMB^{+}16, HPB^{+}09, KTH^{+}21, PTP^{+}22, PKP14, SWP^{+}13a]. bio-geographical [HPB^{+}09]. Bio-optical [GNH19, TII^{+}14, BDB^{+}04, BMB^{+}16, KTH^{+}21, PTP^{+}22]. bio-physical [PKP14, SWP^{+}13a]. Bioaccumulation [ORB^{+}18, FDM^{+}13]. Bioavailable [LFBP^{+}13]. Biochemical [DDP^{+}00, DDD^{+}00, MPC^{+}17]. Biodegradation [RPG^{+}18]. Biodiversity [BBRM20, GCLD19, MA20, MFA^{+}15, SSB20a, BD19, BHC^{+}18, EBM^{+}20, FAB^{+}09, MDAW^{+}19, MVV^{+}19, NCC^{+}15, RSB^{+}15, SMR^{+}20, SPH^{+}15b, STGR^{+}23, TKC^{+}22, WPA^{+}24, dLLdAWL^{+}23]. Bioen [MSB^{+}23]. Bioen-OSMOSE [MSB^{+}23]. Bioenergetic [GCD^{+}18, MPB^{+}23, MSB^{+}23]. Bioenergetics [JSdSS^{+}21, LLS01, MLHE23, YWUK15]. Biofilm [WST^{+}16]. Biofilm-like [WST^{+}16]. Biogenic [FTG^{+}11, GTR01, NEI^{+}22, ÁSÁB^{+}14, ÁBMÁS14, ÁBMÁS15, BT07, CWZ^{+}20, CE84, GLY23, IU14, LBNBM13, ORW^{+}01, PBP^{+}99, RGC^{+}01, TRP^{+}23, THM^{+}06, TGR05, WGZZ19, WSC^{+}21]. Biogeochemical [CLV^{+}19, DFM^{+}21, HBV^{+}10, HBW17, HWF^{+}21, NBG^{+}05, NMY^{+}14, NYH^{+}22, NST^{+}23, RGB^{+}17, VPW01, ÁSFP^{+}03, AH15, BFPS06, BHM^{+}15, BTS22, DVB^{+}18, DCS^{+}22, FMP19, FCEZ10, FYYC05, GLF^{+}17, GL23, HHR^{+}19, JB15, JLB^{+}08, KKO10, LK13, LRW^{+}15, LSH^{+}22, MMGL^{+}07, MCG^{+}14, PFHM16, PST^{+}15, RBL^{+}19, RRLS22, Smi05, TR99, TWBC^{+}13, TAO05, TDL^{+}17, UBB^{+}23, WSO^{+}13, WHBK05, WL16, XC14, ZHF^{+}24, ZDM^{+}20]. biogeochemically [DBRK17]. Biogeochemistry [CPNL07, KHC^{+}99, NCH^{+}07, CTMV^{+}14, Cow05, GCD^{+}99, HM15, KHS^{+}14, PBB^{+}20, Pai20, PLP99, SKWWGV18, SMP^{+}22b]. biogeographic [MB20, SW21, XLL^{+}20]. Biogeographical [FPY^{+}16, OHC^{+}17]. Biogeography [BC16, MKD90, WPBG^{+}18, Bol94, BBRM20, DSR21, ERBV21, GdRGC^{+}14, GC09, JAC^{+}12, KA94, OH94, OvdSN94, Sou94a, VWDF14, VFS^{+}15, WGCS13, Whi94, vdS94a, vdS94b]. Bioindicator [WWW^{+}23]. Biological [CPC^{+}02, HKE^{+}10, JLP^{+}20a, JLP^{+}20b, Seg69, SW22, BBE^{+}15, BHM^{+}15, BP02, BKC15, CW06, CMC^{+}16, CPG08, CLCBB19, CMF15, CEF^{+}13, DLL^{+}23, DRVMC^{+}22, FDE^{+}22, GdRGL^{+}01, GDI^{+}09, HVRR15, HFW^{+}98, HKPV12, HMKF08, HFPS^{+}06, KYT^{+}16, KCPM09, Law04, Leh01, LBSP01, LSMG01, LSS^{+}09, LHF^{+}16, LH89, LFBP^{+}13, MMGL^{+}07, MHS^{+}20a, MHS^{+}20b, MKHO96, MMN^{+}24, MJA^{+}07, Ore69, PMC16, RLT^{+}22, RDD^{+}18, RGI05, RBL90, RGM01, SOS^{+}07, SCD^{+}07, SMP^{+}22a, SMB88, SG91, SDJ14, TDGY22, TCL20, Tur15, WB03, YNM^{+}02, SHK^{+}14]. biologically [BCB^{+}05, KFH^{+}15, MCMT^{+}17]. biologist [Bil01]. Biology [Har82, BM76, CH07b, FARRL^{+}13, HWBT03, Nie07]. bioluminescence [MPMA13, MNT14, MSMH19]. biomarker [BSC^{+}19, WPW^{+}14]. biomarkers [LSV14, PPCWJ18]. Biomass [AMY^{+}23, KH09, AGL^{+}15, BMO12, BBMR19, BM07, CLdPHL23, CCW^{+}18, DDE^{+}95, DDP^{+}00, DAU22, DBJ^{+}15, FGGDF^{+}04, GWK17, GSSWK20, GBC^{+}15, GAPM16, HVEF09, HG04, ILA21, INO^{+}24, Igu04, IVT^{+}12, JSdSS^{+}21, KSVT00, KBC^{+}22, KGB^{+}23, KDB95, LLL^{+}11, LAP10, LMC^{+}20, LWBD^{+}17, MIW91, MA12, MMPG07, PD15, PS98, QPR03, SVL^{+}23, SiSI^{+}02, SSN23, SEG22a, SMM^{+}90, SOA^{+}23, VAEP24, VDGGD^{+}22, WSS15, WCS^{+}23, YMA^{+}17, YNMY23, KVNT20]. biomass-size [QPR03]. Biome [ERBV21]. Biophysical [LAHI10, PTP^{+}22, PDAM^{+}15, CKL^{+}14, KLP^{+}17, LBC^{+}23, MPM^{+}18, SNV^{+}18, VMH^{+}21]. bioprobes [NBLI20]. bioregionalisation [LML^{+}23]. bioregions [PGY^{+}22, PYKF15]. BiOS [CGB^{+}23]. biota [RCSA01]. Biotic [KHBA^{+}24, GBT^{+}19]. Biotransformation [CDS90]. bioturbation [HG04, MSL^{+}07, QSC^{+}15, Whe06]. biovolume [GWM^{+}22]. bird [DLD15]. birds [SPS^{+}99, SHT^{+}01]. Bisagno [CLD22, DSC^{+}19]. Biscay [CÁM06, CSC^{+}12, HBV^{+}10, AIA^{+}15, ADV^{+}18, ALT10, BCGN^{+}18, BHLU^{+}07, CDS90, DBC^{+}18, DPH^{+}18, DPR^{+}18, DHHP18, ERT^{+}22, GCD^{+}18, GA10, GRD^{+}23, HLD^{+}21, HPW10, IFC^{+}07, LAD^{+}18, LLL^{+}11, LCBN14, LML^{+}23, LOBG^{+}10, LSIB23, MGS90, NP00, PMFNGQ21, PGRP^{+}18, RCSVGP^{+}16, SBL^{+}23, SRT^{+}18, VLUC^{+}07, VK90, VDB^{+}20, WPH^{+}10]. Biscay [?]IrishLeJoncour:2023:WBE. BIT [SMN^{+}14]. Bivalve [Kam19, RvBD^{+}22]. bivalves [AS96, BLES16, LDH90]. Bivalvia [RVC^{+}13]. Black [WFD^{+}07, MNFY21, FBS^{+}18, GRS08, KBSB18, KAG^{+}19, LDMH09, MCKS17, MSGGM18, ÖÜT93, SÖÜ94b, SI97, Tol85a, Tol85b]. Black-legged [WFD^{+}07]. Blanc [FRK^{+}09, NIF^{+}15]. Blanes [ASFB^{+}13, ACL^{+}18, DCL^{+}13b, FBR^{+}13, IVR^{+}13, JFEC13, KFC^{+}13, LFBP^{+}13, LFCSV^{+}13, PPSV^{+}18, PPSVC^{+}13, RCF^{+}13, SCB^{+}09, VGM^{+}23, ZFSV^{+}09]. blending [BBM^{+}14]. Block [SCB^{+}16]. bloom [AW13, CKB^{+}17, CMC^{+}16, FHP83, GMDD^{+}22b, GHL15, GKS^{+}13, HBV^{+}10, INO^{+}24, IMM^{+}22, KHBA^{+}24, KFM15, LGL^{+}18, MCKS17, MMD^{+}16, NEI^{+}22, NSE^{+}24, OTNI20, PMMN^{+}22, PMS^{+}15, RSM^{+}23, RHM^{+}19, STB^{+}92, STF^{+}13, SMP^{+}22a, STHM02, ST10, SPH83, SDL^{+}19, TSAM^{+}22, TPRS10, VEM^{+}21, VBM21, WFH^{+}22, WLM^{+}13, WSH^{+}22, ZLG17a, ZLG17b, ZWM^{+}15]. Blooms [GBH^{+}20, BTS^{+}15a, BPGD^{+}14, DLC^{+}08, Fro05, FB05, HLPL05, HPHL^{+}05, ISM^{+}02, IPF23, KMMC09, KSC10, KHJ^{+}10, KTW^{+}22, LPHL^{+}05b, LPHL^{+}05a, LMA^{+}15, MPMFL^{+}23, OMS^{+}09, PHLL05, PP10, PFHM10, RPSC22, iSIS02, ST10, VHV^{+}12, VSPP14, WDC^{+}11, WQ08, ZBY^{+}22]. blows [HLM^{+}16]. blue [BPSN^{+}21, HPB^{+}09, MAFS^{+}22]. bluefin [AQVB^{+}10, BMC^{+}10, DFH^{+}16, DAIS10, FFT^{+}18, GA10, KKKY10, KTIT22, MMIB10, PMM^{+}23, RDP^{+}21]. Board [Ano17e, Ano17f, Ano17g, Ano17h, Ano63a, Ano64b, Ano73c, Ano76a, Ano79a, Ano80b, Ano81a, Ano82b, Ano83a, Ano84a, Ano85d, Ano85e, Ano86b, Ano86c, Ano87b, Ano87c, Ano88a, Ano88b, Ano89b, Ano89c, Ano90a, Ano90b, Ano91a, Ano91b, Ano91c, Ano92b, Ano92c, Ano93a, Ano93b, Ano94a, Ano94b, Ano95a, Ano95b, Ano96a, Ano96b, Ano97a, Ano97b, Ano07a, Ano07b, Ano07c, Ano07d, Ano07e, Ano07f, Ano07g, Ano07h, Ano07i, Ano07j, Ano08a, Ano08b, Ano08c, Ano08d, Ano08e, Ano08f, Ano08g, Ano08h, Ano08i, Ano08j, Ano08k, Ano08l, Ano08m, Ano09a, Ano09b, Ano09c, Ano09d, Ano09e, Ano09f, Ano09g, Ano10c, Ano10d, Ano10e, Ano10f, Ano10g, Ano10h, Ano10i, Ano11a, Ano11b, Ano11c, Ano11d, Ano11e, Ano11f, Ano12a, Ano12b, Ano12c, Ano12d]. Board [Ano12e, Ano12f, Ano12g, Ano12h, Ano12i, Ano12j, Ano13a, Ano13b, Ano13c, Ano13d, Ano13e, Ano13f, Ano17i, Ano17j, Ano17k, Ano18b, Ano18c, Ano18d, Ano18e, Ano18f, Ano18g, Ano18h, Ano18i, Ano19b, Ano19c, Ano19d, Ano19e, Ano19f, Ano19g, Ano19h, Ano19i, Ano20c, Ano20d, Ano20e, Ano20f, Ano20g, Ano20h, Ano20i, Ano20j, Ano20k, Ano20l, Ano21d, Ano21e, Ano21f, Ano21g, Ano21h, Ano21i, Ano21j, Ano21k, Ano21l, Ano21m, Ano22d, Ano22e, Ano22f, Ano22g, Ano22h, Ano22i, Ano22j, Ano22k, Ano22l, Ano22m, Ano22n, Ano22o, Ano22p, Ano22q, Ano22r, Ano22s, Ano22t, Ano23d, Ano23e, Ano23f, Ano23g, Ano23h, Ano23i, Ano23j, Ano23k, Ano23l, Ano23m, Ano24b, Ano24c, Ano24d, Ano24e, Ano24f, Ano24g]. bodies [IST^{+}88]. Body [RSW^{+}23, BFV^{+}17, GDL^{+}15, HSG^{+}15, HBG^{+}21, IAFD02, MMPG07, SE92, VDB^{+}20, YKNO23]. BOFS [STB^{+}92]. Bohai [LSXT01, LW13, SW12, YSY^{+}19]. bomb [KMWF11]. bomb-produced [KMWF11]. Bonelliida [GS19]. Bongo [JF13]. Book [Ang79a, Ang80, Ang88, Bak83, Hof81, SW81]. bord [Ber65c]. Borderland [Gor92]. boreal [LNB13, PTP^{+}22]. borealization [OV24]. Boreogadus [BF11, KSG^{+}17, VMH^{+}21]. bores [HNSP^{+}19]. boring [RVC^{+}13]. borne [SGR^{+}22]. Bornholm [BLC23, SPB^{+}12]. both [Kit03, WWL^{+}22]. bottlenose [LPF^{+}18]. bottles [MK86]. Bottom [DJW^{+}18, dCFK17, HS22, MLPN06, OYKK^{+}23, OJB99, PSA^{+}19, SPN98, SIB^{+}06, YFY^{+}22, Zen08, AH10, AR18, AF10, ASB^{+}08, BVJE19, BSF95, CML^{+}16, CB06, CHB02, DGVGR24, ESTM^{+}12, ESTM13, FZY^{+}23, FC07, FFA06, FWL^{+}15, Gam14, Hop64, HWB^{+}18, HM06, KKS^{+}19, LTG85, LL97, LYS^{+}22, MPB^{+}23, MM80, PPSV^{+}18, PAB^{+}21, PdMS^{+}13, RGMPR23, RKK^{+}21, SCHS^{+}24, VK90, VSC01, VOJD02b, VB14, XYK^{+}22, XWL^{+}22, YMK^{+}04, ZCV^{+}19, ZZWL06, dPCS23]. bottom- [XYK^{+}22]. bottom-ice [CML^{+}16]. bottom-living [MM80]. Bottom-up [HS22, MLPN06, PSA^{+}19, SIB^{+}06, YFY^{+}22, AH10, AF10, ASB^{+}08, BVJE19, ESTM^{+}12, ESTM13, FFA06, HM06, MPB^{+}23, PAB^{+}21, RGMPR23, SCHS^{+}24, ZCV^{+}19]. bottoms [QOS^{+}22]. boundaries [Ban65, BW65, CMJPH^{+}18, SCLS10, SJ02a, vdS94a]. Boundary [Ano09h, Car98, CM09, EBR^{+}14, FBA09, FAB^{+}09, GBC^{+}16, SPN98, SÖÜ94b, WO85, BTG^{+}03, BPM^{+}14, CMHM18, CDP14, DWS^{+}24, DSC^{+}21, DL17, EBvdL^{+}09, FBS22, HHMB^{+}09, HBW17, HVEF09, Hut92, KKK04a, KHD22, LG22, LG23, McC92, MBP65, MAAS^{+}00, MLK^{+}09, MC15, Mit91, PP10, Rog00, Sel65, TWAL^{+}11, TvW98, TvG02, TPRS10, VNMS91, VFS^{+}15, VOJD02b, VB14, WZFW16, WWN^{+}99, XY21]. Bowhead [DHD^{+}23, GDL^{+}15, CQO^{+}15, CBB^{+}15]. box [APHGC^{+}22, JPBB20, KKK04a, LB20]. box-corer [LB20]. brachiatus [HE07]. brackish [PBD^{+}88, Wen88]. branch [JC88, MR06, RKS^{+}15]. branches [PLB^{+}23]. Branching [KL86, Sek86]. Branica [SW81].

Bransfield [VKDS^{+}18, VAGMDRS22]. brash [NST^{+}23]. brasiliensis [DPGC14]. Brazil [AG22, BRB^{+}01, BTS22, PG10, DSC^{+}21, GRDS10, Has06, RPPM^{+}23, WR00, dSPF^{+}23]. Brazilian [dFKdLZTT17, BKB85, DPGC14]. breadth [LVGH^{+}15]. break [CMC^{+}16, DLD^{+}19, LMM03, OC06]. break-up [DLD^{+}19]. breaking [HLFL23, ITO^{+}14, PVG^{+}20, vHVAT22]. breaks [ZMW^{+}23]. breakup [NNFL21]. breakwater [Iwa23]. bred [HKK12]. breeders [SM16]. breeding [DWNN04, FJH10]. brevis [BK19]. bridged [LLX^{+}21]. Bridging [DP18, LMS10, MCG^{+}14]. Brief [CBGC^{+}08, Ano10a, Kaw98, SSM^{+}90b]. Britain [HPHW21]. British [CMHM18, FDM^{+}13, GKS^{+}13, GDM^{+}15, LMH^{+}13, LC10, PMC16, RHBS13, STF^{+}13, SL13, TSC03, CM14b, HHWW20, MB01]. Brittany [RSM^{+}23]. Broad [GAS^{+}22, MBH^{+}23, BB14, CMF11, OOTA15]. Broad-scale [GAS^{+}22, MBH^{+}23]. browser [VR03]. Bryozoa [Gri22]. bryozoan [GHSC19]. bryozoans [Gri22]. bubbles [Bla63]. budget [BDTC15, CBC^{+}06, FMP19, KVNT20, MMR^{+}09, RSB^{+}01, RKS^{+}15, VK92]. budgets [GSFP^{+}09, JVJ^{+}17, JWD^{+}02, RDD^{+}18, SGMP15, SCB^{+}07, ZLR^{+}07]. Building [AHC^{+}13, CSH^{+}23, Ber65a, TSL10]. Bulk [ZZWL06, CDP14, GASV^{+}09]. Bullet [KTIT22]. bungii [MFB^{+}84, THP21]. buoy [CPB^{+}15, UCB^{+}18, WZC20]. Buoyancy [SFMT12, PHFK14, SFMT14]. Buoyancy-forced [SFMT12]. Buoyant [MSd^{+}16, ILI^{+}12]. buoys [Kvi69]. burial [EvdZSH02, HM98, vWHdS^{+}98]. Burst [WLCG23]. burying [RHBS13]. Bussol’ [YYT^{+}14, TYO^{+}14]. bycatch [RSK^{+}23].

C [CGC^{+}20, CRC^{+}19, DVB^{+}18, JJA^{+}08, OMK^{+}22, OE65, PHK^{+}17, PRL^{+}18, SKH^{+}23, SBK^{+}95, STHM02, WRS^{+}92]. C. [DK07, WSH15]. cabbeling [Har05a]. Cabled [VGJ^{+}19, DOS^{+}18]. Cabliers [CLG^{+}22]. Cabo [CCRS20, VDGGD^{+}22]. Cadiz [SBPGP^{+}23, GPE^{+}17, LdCSB^{+}20, SRM^{+}10, LR07, RGPB^{+}23]. Cadmium [Sim81, FCG88]. Caicos [Ché14]. Calanid [MN88]. Calanidae [Mil88]. calanoid [AHSS22, BHS^{+}15, CÁM06, IS19, MSMR93, Nie07, SSTD^{+}95, THBA19, WSH15, WGG^{+}08]. Calanoida [FMT15, GBG05, HLPL05, Mil88, NMN08, PMH17]. Calanoides [AVK91, CÁM06, Ver91]. Calanus [AEP^{+}23, BMK12, BTNK13, BRE^{+}24, BRH^{+}05, BMG^{+}21a, CÁM06, DWFP^{+}19, DBM17, GHF^{+}21, GPC^{+}03, HHY03, HMP^{+}13, HRA^{+}08, HBR11, HE07, JAC^{+}12, JC04, KHP^{+}18, LBC^{+}23, LPHL^{+}05b, MMG^{+}13, MAH^{+}15, MRH^{+}14, NMN08, PD15, PHLL05, PLHLF05, SDH^{+}14, SAB^{+}21, SBE^{+}20, THP21, UB10, WPB^{+}08, WBC^{+}22, WSH15]. calcareous [Gal17]. calcification [KLIRK17, Kru19, ZCD08]. calcified [BHK^{+}19]. calcifiers [AOMZ^{+}23]. calcifying [GBB^{+}19]. Calcium [Wis65]. CalCOFI [Reb02]. Calculating [GGPG^{+}19]. Calculation [YJW88, Kaw86, YJ88, ZPY^{+}20]. calculations [Tur65]. Caledonia [CKL^{+}14]. calibrate [iIYO^{+}10, ORVES17]. calibration [HM00b, SBH^{+}14]. California [ASC92, Ban65, BGM^{+}10, DIM09, DLJ^{+}21, Dur09, Epp92, FFA06, KMMC09, KC15, LLS01, LABD^{+}24, MPC^{+}17, OBD^{+}20, PBBH^{+}22, PO15, RBE^{+}12, SRFHDH22, AAML22, AH10, AASJ23, ARG11, ABP15, BB14, BWB^{+}09, BL02, BM86, CCW^{+}02, CPC^{+}02, CB09, CCA^{+}02, CCD^{+}13, DFM^{+}21, DW02, DSBP15, DB02, ESTM^{+}12, FELMGM^{+}22, FSAO22, FWO15, FRCH15, FMC^{+}20, FWBC02, GFGGD^{+}23, GMD^{+}22, Gor92, HSMLDC^{+}22, Hau84, Hic79, HW02, hHRW^{+}05, Huy83, HSF02, JE92, JSA^{+}08, JOGM^{+}10, JC04, KHL12, KSD84, Kos02, KC02, LOG^{+}09, LJPGC02, LO07, Lav09, LAA12, LO21, LBP15, LCANAS^{+}07, LCPSMR^{+}10, LABD^{+}24, LB02, MS02, MAB^{+}11a, MAB^{+}11b, MJC^{+}17, MCT03, NFMCS^{+}22, NMLBCM^{+}01, PELAA18, PK02, PBS22, REG^{+}15, ROBRB^{+}22, RMK^{+}21, RB20, RFC^{+}15, RZTD17, RN02, SCHBC^{+}22, San15, SGMVF14, SSS^{+}11]. California [SFS^{+}12, SLBVRR^{+}22, SC23, SDK84, Sim84, SKHD84, SBM^{+}23, SHT^{+}01, STGR^{+}14, TBW09, VK92, Ven12, WFBN^{+}13, WDMC02, WRS^{+}92]. californicus [STS^{+}12]. call [Jac10]. Calvi [GHL15]. calycophoran [LSIC12]. Calyptogena [AB90]. camera [TSAM^{+}22]. campaign [JST^{+}24]. campechanus [ZLRVB24a, ZLRVB24b]. Can [CRS04, LPHL^{+}05a, Mau17, QSC^{+}15, RCD^{+}94, SvN04, TDL^{+}17, HMX^{+}23, RSG06, WFJ^{+}15]. Canada [CMHM18, GDM^{+}15, BBSN04, DLM91, EHG^{+}12, FDM^{+}13, GKS^{+}13, HGD22, HKE^{+}10, LC10, Man69, MPN09, PM13, RHBS13, STF^{+}13, TSC03, VSGD21]. Canadian [DHD^{+}23, CML^{+}16, EHSI12, MIH06, WSL20, WTH12, YGMR^{+}23]. Canal [Ore69]. CANALES [PLJR22]. Canaria [Ano09h]. Canaries [GMAMB04, RMHL09, RBHLA04]. Canary [BAM^{+}09, GASV^{+}09, AAM^{+}14, ABÁS^{+}09, AGL^{+}15, BAT^{+}98, BA04, BATNP04, BFR13, CLdPHL23, HLGA07, KZSH85, MHGP06, PVB23, SGWF^{+}19, SFMA20, SHL13, SAd^{+}17]. Cantabrian [CSV^{+}07, CÁM06, RCSVGP^{+}16]. Canyon [AHA^{+}16, CLSD18, CMHM18, DJW^{+}18, DOS^{+}18, EvdZSH02, LPF^{+}21, LPBM17, LRGV^{+}18, AHD18, CHG^{+}18, CFM^{+}18, CJRÁ^{+}13, CQZ^{+}18, CHSB^{+}21, CVHM^{+}18, CRC^{+}19, DCRR^{+}22, DCL^{+}13b, FVA^{+}19, FBR^{+}13, GCF^{+}19, GBB^{+}20, GIPG17, HDM19, IVR^{+}13, JOBT05, KCL^{+}12, KFC^{+}13, LFCSV^{+}13, MRH^{+}18, PGLG^{+}05, PRC^{+}20, RCC^{+}18, RCF^{+}13, RCSVGP^{+}16, SCB^{+}09, ZFSV^{+}09, ASFB^{+}13, ACL^{+}18, CFML22, EVM^{+}15, JFEC13, KGdS^{+}08, LPF^{+}20, LFBP^{+}13, PPSV^{+}18, PPSVC^{+}13, PGGG17, PRA^{+}18, SCC^{+}19]. canyon-slope [IVR^{+}13]. Canyons [MRH^{+}18, BFP^{+}18, BRC^{+}18, BD19, Bou65, CCM^{+}13, CDL19, CLD22, CTR^{+}19, DP18, DSB^{+}24, DSC^{+}19, DAU22, FBD18, GvOS^{+}08, HSN^{+}18, IHR18, KKS^{+}18, PPdM^{+}12, PCD^{+}18, PGT^{+}13, PCC^{+}19, RVC^{+}13, TCDPP^{+}22, TCDPP^{+}23, TAF^{+}22, TPPG10, TTF^{+}22, TCL^{+}15, GCLD19]. Cap [DCRR^{+}22, LRGV^{+}18]. capability [Dah69, PRC^{+}20]. Capacity [BK08, PS08]. cape [PK02, AAM^{+}14, ABT^{+}04, FRK^{+}09, GL06, MB07, NIF^{+}15, SVL^{+}23, VFCC^{+}22]. capelin [BSF^{+}21, CGV13a, CGV13b, DM13, MMD^{+}16, SEG22a]. capture [PTF12]. captured [HPW10]. Capturing [BTNK13, DLJ^{+}21, MVV^{+}19]. Carbon [DCKB13, GDI^{+}09, HOY^{+}21b, KAAK^{+}16, PPCWJ18, RN06, SHK^{+}14, SLOP^{+}22, SC65, VKGP^{+}13, WMC^{+}89, AOMZ^{+}23, AYK^{+}05, ABT^{+}04, AGL^{+}15, BEI^{+}20, BSMC15, BB24, BK19, BS02, CWB^{+}22, CBC^{+}06, CJ92, CCW^{+}18, DOS^{+}18, DHDM22, DGH^{+}20, DWC06, EvdZSH02, EMK^{+}17, FZY^{+}23, FTG^{+}11, FWBC02, FLUC08, FAH^{+}13, GCV^{+}24, GCD^{+}13, GSSWK20, HM98, HPS^{+}01, HW02, HKGH^{+}06, HPZC21, HMKF08, HAH^{+}22, IVT^{+}12, JCF^{+}23, JJ08, JJA^{+}08, KBHML17, KBF^{+}08, Kli10, KSS^{+}23, KFC^{+}23, KSG^{+}17, KGdS^{+}08, KVNT20, LSV14, LRW^{+}15, LBP^{+}21, MHS^{+}20a, MHS^{+}20b, MPM^{+}17, MRAP22, MMKS^{+}21, MLHE23, MMG^{+}11, MDC^{+}07, MNFY21, OB98, PG13, PHK^{+}17, PPSVC^{+}13, PDD^{+}22, RJO^{+}19, RCS^{+}11, RGE22, RSD^{+}90, SLBR18, SFMA20, SHd13, SJA^{+}23, SPG^{+}06, SMP^{+}22b, SW01, SE92, SIS^{+}14, SDJ14, TvG02, TMKJ^{+}09, TFM03, THM^{+}06, TDK^{+}16, VK92, WCC^{+}20]. carbon [WRH^{+}06, WXH07, XYGJ23, XCH^{+}16, YHM^{+}18, YSN20, ZCD08, ZMW^{+}23, ZKK^{+}16, ZZWL06, dFHR^{+}24, dJSL^{+}20, vWHdS^{+}98]. Carbon-14 [SC65]. Carbonate [PST^{+}02, BRG^{+}15, GLLB22, HHK^{+}22, KMSTK23, MLB^{+}20, Ola65a, RSB^{+}01, RDC^{+}21, Wis65, YGC^{+}21, BHK^{+}19]. carbonic [RAP95]. carboxylating [SPH83]. carcasses [YHRT22]. Carcinus [YFK21]. Caretta [MHR^{+}10]. Cariaco [BPTT19]. Caribbean [CMJPH^{+}18, MKMF^{+}89, Nof00, PGG^{+}22]. carinatus [AVK91, CÁM06, Ver91]. carnivorous [GSC^{+}20, MGS90, SE92]. carpenteri [RTN90, VBJ^{+}20]. Carrying [BK08, PS08]. carryover [WBF^{+}21]. Cascades [ISH^{+}04, dZTG05, LPHL^{+}05a, WDC^{+}11]. Cascadian [GLV12]. cascading [PPdM^{+}12, PdMS^{+}13, RCC^{+}18, SGL^{+}17, TPPG10]. cascading-dominated [TPPG10]. Case [MZK^{+}23, AHA^{+}16, Ang80, ALT10, BCLD^{+}17, CCRS20, CSC^{+}12, ESTM13, GKC^{+}14, HVS10, HPNDC15, HPW10, JBB^{+}14, KV18, KFC^{+}13, KGdS^{+}08, KS15, LJM^{+}16, LDD^{+}22, LB20, MDGC^{+}12, NHG19, NMC^{+}09, PKA19, PPdS21, RZW^{+}23, RSM^{+}23, SHd13, TIOM16, Val99b, YFY^{+}22]. Cassidaigne [FVA^{+}19]. Castle [NW87]. Catalan [DDDT99, ACE^{+}07, CGM^{+}02, RAE^{+}05, SCAA07, SVHM^{+}13]. Catalonia [CMF^{+}09]. catch [ARD^{+}03, JOGM^{+}10]. catch-and-release [JOGM^{+}10]. catches [FELMGM^{+}22, SNR^{+}10]. cause [AAMB^{+}24, CMF11, GJ00, SGLF^{+}13, dSSDS^{+}20, SA97, ZBY^{+}22]. caused [GW89, PPSV^{+}18, SMFM^{+}21, TSP^{+}13]. Causes [HMH07, BN03, Bea04, RTN90]. cautionary [Wai21]. caves [SCS87]. CCCC [BK08]. cell [AV23, BPA^{+}21, GASV^{+}09]. cells [WLL^{+}23]. cellular [LCB18]. cellular- [LCB18]. Celtic [CHG^{+}18, DGP^{+}13, DJ92, EMBS13, ERT^{+}22, FHP83, GWM^{+}22, LTSG13, LDB^{+}02, MZF^{+}08, MEST13, MPD^{+}22, PIS13, PM85, PL89, PDD^{+}22, SMP^{+}22a, VHV^{+}12]. census [APC13, Ric93, RBZ00]. centennial [KYS^{+}17]. center [LLL^{+}24, MDL^{+}12, SLG^{+}12]. Central [CCA^{+}02, HDM19, HLTB^{+}17, LBSP01, PGC^{+}96, Tom81a, VDDA^{+}08, AMEV07, AMG^{+}16, Ang79b, AT07, BB14, BD18, BM07, BCP09, CCW^{+}02, CTF07, CGC^{+}20, CPC^{+}02, CBM^{+}21, CFML22, CFG07, CF12, ES07, EHG^{+}07, EM12, FC07, FP03, FGGDF^{+}04, FWBC02, GMBU12, GSPP^{+}20, GRS08, GEP^{+}08, HSMLDC^{+}22, HE07, HEF^{+}12, HKY^{+}11, HHP10, IAFD02, KFKO03, KWI20, KP03, KNS^{+}03, KC02, LC12, LQU07, LDHW20, MERB12, MMES16, MMF^{+}12, MWFH02, MGH^{+}07, MTH^{+}10, MHCR^{+}12, MA12, MSL^{+}07, OWR^{+}07, OAWAN18, PCSMC12, PAM^{+}88, PKF02, PCC^{+}19, RM93, RFSCF19, RMG90, RBNJ^{+}12, RBPGJ^{+}20, RCSVGP^{+}16, RN02, SLM^{+}16, SFS^{+}12, SDGVE17, SPB^{+}12, SS03, SGO^{+}08, SBD^{+}07, SSL07, Soh03, SMPC^{+}12, STGR^{+}14, TFY02, TCF^{+}18, VAEP24, gWjNfLyD20, WJPHB15, YNM^{+}02, YCP^{+}12, YFY^{+}22, BM86, BASS^{+}20]. Central [DLM^{+}12, SCC^{+}19]. central-south [SDGVE17]. central-southern [EHG^{+}07, EM12, HE07, HEF^{+}12, MERB12, MMF^{+}12, MTH^{+}10, MHCR^{+}12, MA12, MSL^{+}07, RFSCF19, RBNJ^{+}12, SLM^{+}16, SMPC^{+}12, VAEP24]. central/southern [ES07, MMES16]. Centropages [ASC07, BLHB07, BHLU^{+}07, CCG07, CH07b, CBHL07, DK07, GTB07, HE07, IMHL07, MCD^{+}07]. Century [ALV^{+}21, BDLW14, DML^{+}16, HMH07, LSW02, MWS^{+}10, RD11, TLP^{+}16, WBH15]. Cephalaspidea [CES^{+}19]. cephalopod [KQP^{+}17, OÁSG^{+}16, RAG^{+}19]. certain [BM76]. Cetacean [MWFH02, BCM^{+}02, MCD^{+}14]. cetaceans [BPF06, BGA^{+}21, SRT^{+}18]. CFCs [JJA^{+}08, MBB^{+}96]. chaetognath [FB05, MMN12]. chaetognaths [GSC^{+}20]. chain [SOH21]. chalcogramma [MLPN06, YNM^{+}02]. chalcogrammus [GTS^{+}21]. Challenges [HHAR23, HAA^{+}14, CCM^{+}13, LCANAS^{+}07, MPC12, RBD^{+}07, VBL^{+}21, WMB^{+}18, Yas07a]. chamber [TDH^{+}95, TSG^{+}04]. chambers [TSG^{+}04]. Change [BK08, HSG^{+}15, WNNI21, AT07, ADV^{+}18, BM01, BBR^{+}01, BDE03, BHMS09, CDTM^{+}21, CM11, CALS^{+}23, CLMR23, Con87, DPCS87, Dri11, DBR20, FDH20, HRSM08, HGB^{+}21, HSH^{+}19, HWLT10, HSC^{+}16, HS02, KKK04a, KA94, LAHI10, LHW^{+}20, MPMFL^{+}23, McI10, MNM06, MCB^{+}10, NDEG22, NKK03, NPO^{+}19, OVG16, ORPRGIS22, PYKF15, PG10, Reb02, RK03a, RAB^{+}11, SRFHDH22, SYB^{+}15, SAY^{+}16, SON^{+}20, SHF01, SBM^{+}23, ŠGM^{+}18, SCS87, SHT^{+}01, STGR^{+}14, VMV^{+}23, WFD^{+}07, Was11, WHBW03, WBD^{+}15, WJPHB15, dlHRA^{+}18]. changed [LBH^{+}21]. Changes [ABSDC07, BCM^{+}02, CMJPH^{+}18, CCA^{+}02, HPHW21, JLRB20, KKB00, KF11, LF12, LSW02, MWS^{+}10, PBB^{+}20, PO15, RBE^{+}12, TWMY08, Ven12, WHK23, AMFY20, ABE^{+}15, AAMB^{+}24, AVG^{+}19, BAM^{+}09, BDLW14, BMG^{+}21a, Bro82, CGV13a, CJMO87, CGM^{+}02, CMF^{+}09, CGC^{+}20, CBOP15, CBM^{+}21, CS04, DMD^{+}00, DTOD00, DDD^{+}00, DMT15, DBC^{+}18, DLM^{+}96, DYL^{+}15, DMF^{+}09, ERT^{+}22, ESGP17, FGS^{+}15, FC07, FHG03, FFS^{+}20, FMSBW13, GSV^{+}01, GMDD^{+}22a, GMD^{+}22, GGA^{+}05, GW89, HVRR15, IVR^{+}13, Iwa23, JTQ^{+}18, JSdSS^{+}21, JKBH87, KBE^{+}22, KOhL^{+}10, KRHS14, LRNK99, LDD^{+}22, LXC^{+}22, LWY07, LDMH09, MMG^{+}13, MKB00, MMES16, Med87, MCGR07, MSGGM18, Min02, MIW91, Nag01, Ola65b, PVG^{+}20, PRTC13, PBO10, Pir87, PVA24, PBN13, RMC^{+}15, SMR^{+}20, SPC^{+}23, SKSK06, SC65, SJH^{+}90, SMKK21, SKH00, TKW06, THP21, VDB^{+}20, Whe93, WHBW03]. changes [Wil65, XYK^{+}22, YNMY23, YNM^{+}02, Yas07b, Yas07c, YFY05, ZWM^{+}15, dLLdAWL^{+}23]. Changing [LCPSMR^{+}10, MGE^{+}12, Tol85a, Tol85b, Yas07a, BRG^{+}15, BHM^{+}15, GCV^{+}24, IGG^{+}19, IPF23, JPIP22, JLS^{+}22, KLIRK17, Kru19, LRW^{+}15, LSH^{+}11, LRJ^{+}15, MTC14, MHH^{+}15, NGPH10, TAM^{+}15, Tur99, Val99a, Val99b, Ven12]. Changjiang [Ken88, ZLR^{+}07, ZCH^{+}17]. Channel [ALT10, CAH^{+}22, FWO15, ILA21, SPSR^{+}14, SRFHDH22, SC23, SAW^{+}15, CFM^{+}18, CTI^{+}19, MKM86, MWO^{+}12, ZWP23, AHW99, BTS^{+}15a, BTS^{+}15b, BLCL14, BHLU^{+}07, GCD^{+}18, GGA^{+}16, GGG^{+}18, GGSM^{+}20, HØ07, KM10, LSD^{+}15, MAH^{+}15, OAB^{+}16, PMS^{+}15, RMC^{+}15, SGA^{+}19, STW^{+}15, TAW^{+}15, TB15]. channels [ABC^{+}99, CCM^{+}14, IPF23, PMA^{+}14, PTPY^{+}23, SPF^{+}23, PLJR22]. Chaos [OAM00]. chapters [Ano10a]. character [TJ90]. Characterisation [JMG^{+}13, ASR^{+}20, CPHR98, CLG^{+}22]. characterises [MB01]. Characteristics [AS88, BH07, LW85, LEDR^{+}22, MNS^{+}24, WGM^{+}24, ABC^{+}99, CDB^{+}22, CQO^{+}15, GBB96, GBC^{+}15, HFPS^{+}06, Igu04, MLM09, Mol04, NHG19, NBR^{+}08, NST^{+}23, PTP^{+}22, Par86, PO15, SWP^{+}13a, SSTL16, SRG^{+}19, STR01, WBB^{+}01, YPM^{+}10, ZKT88, ZSH^{+}24, ZLX^{+}20]. Characterization [AQVB^{+}10, BZD^{+}21, BSC^{+}07, CFM^{+}18, KTH^{+}21, LKDL14, NIC^{+}19, QLY^{+}22, BGR^{+}15, GMBU12, MCGS^{+}16, YT06]. characterize [GNH19, LGL^{+}18, MSMH19, PYKF15, SW22]. Characterizing [AH15, LLH^{+}20, SJLW23, Fla02]. charts [PSK96]. chaude [Ber65b]. Checklist [OT19]. Chemical [BC91, BD85, CPHR98, Sim84, CPC^{+}02, FW91, GTR01, Hey78, ILA21, IPHW^{+}23, KSR^{+}01, KKS^{+}19, WNNI21, YT06]. chemicals [Bon88]. chemistry [BRG^{+}15, BSW86, Bro82, GLLB22, Har82, MLB^{+}20, PST^{+}02]. Chemolithoautotrophic [PLN^{+}23]. chemosynthetic [MSV^{+}14, MCGS^{+}16, ZHSMM14]. Chemotaxonomy [LMT^{+}19]. Chemotaxonomy-based [LMT^{+}19]. Chesapeake [JX18, WCB20a, ZHF^{+}24]. Chile [AMG^{+}16, CRHM12, GGQ07, SPF^{+}23, AMEV07, ACN01, BRR^{+}22, BWB^{+}09, BM07, CTF07, CGC^{+}20, CPNL07, CFG07, CF12, DLM^{+}12, ES07, EHG^{+}07, EM12, EHFD12, FC07, FAAV^{+}15, GMBU12, GKC^{+}14, GMAB07, GDI^{+}09, HE07, HEF^{+}12, IVT^{+}12, LC12, LQU07, MERB12, MMES16, MS02, MMF^{+}12, MGH^{+}07, MTH^{+}10, MHCR^{+}12, MA12, MSL^{+}07, MDL^{+}12, MCGS^{+}16, OAWAN18, PCSMC12, PMMN^{+}22, PAM^{+}88, PTPY^{+}23, PVA24, RFSCF19, RBL^{+}19, RBNJ^{+}12, RBPGJ^{+}20, SLM^{+}16, SLG^{+}12, SDGVE17, SYB^{+}15, SAY^{+}16, SBD^{+}07, SSL07, SMPC^{+}12, SJM^{+}19, VOG^{+}08, VSGC21, VMV^{+}23, VAEP24, YBS^{+}01, YPGE^{+}10, YCP^{+}12, ZHSMM14]. Chilean [BPSN^{+}21, LSV14, MSV^{+}14, SPSR^{+}14, CCM^{+}14, CTI^{+}19, FCN^{+}19, GCD^{+}13, IPD14, JTD^{+}14, LC12, MVBC^{+}21, QOS^{+}22, SV14, VLCCP14]. chilensis [HE07]. chimney [TJ90]. China [CGL^{+}20, CWZ^{+}20, CSW96, CLX^{+}20, Cho86, DFC^{+}21, DL17, DGH^{+}20, DHL^{+}21, FZ88, GCCY^{+}14, GC14, HZD^{+}23, HOY^{+}21b, HPZC21, HSH97, HLSX22, HHZ^{+}22, HWF^{+}21, JZZY24, KHM^{+}88, Li14, LYZ16, LGZ^{+}20, LBH^{+}21, LZL^{+}22, LYS^{+}22, LC16, LCJ^{+}07, LLH^{+}20, LGH^{+}21, LGD^{+}20, MXC^{+}21, MLL^{+}22, MY23, MZZ^{+}23, MH14, MKS^{+}22, NXY15, QLW10, QLY^{+}22, RLX^{+}24, RI86, RDC^{+}21, SSTL16, SW12, SWH^{+}24, TM13, WGZZ19, gWjNfLyD20, WST^{+}21, WL16, WZC20, XWL^{+}18, XLX^{+}20, XHC^{+}20, YKS^{+}12, YMI88, YGC^{+}21, Yao88, YJS86, YJ88, Yux88, ZLR^{+}07, ZGZ19, ZZPL18, ZDG^{+}21, ZMW^{+}23, ZDM^{+}20, ZSY^{+}22, ZZWL06, ZSW^{+}22, ITM86, LZCZ05, LW13, YSY^{+}19, YHZ^{+}22, ZKT88]. Chinese [SPWH21]. Chinook [BL02, CCS^{+}21, DIQJ21]. Chionoecetes [MGKW19, RKCH15]. chlorofluorocarbon [OYKK^{+}23]. chlorofluorocarbons [Sme93]. chlorophorum [RSM^{+}23]. Chlorophyll [Epp92, PHKS17, ABS^{+}20, BMK12, Ban96, BMGN15, BBL^{+}18, BDLW14, CLY22, CRHM12, CW02, EALF08, EBR^{+}14, FPIJ85, FYYC05, HSMLDC^{+}22, HZCZ16, HZD^{+}23, HJLLN07, ILA21, LHE^{+}13, LLL^{+}24, LFI^{+}13, LW13, LHC^{+}19, MHGGS19, MVC^{+}11, MGH^{+}07, NM17, PLHLF05, PHKS01, PVV23, RBR^{+}23, SNR^{+}10, SC23, SMGL01, STW^{+}15, SLH^{+}19, TBW09, VDS^{+}18, WM13, WDMC02, WQ08, XHW^{+}20, XWW^{+}21, YIY^{+}04, YKS^{+}12, ZL24]. chlorophyll-a [BMGN15, CLY22, CRHM12, EBR^{+}14, HSMLDC^{+}22, HJLLN07, LHE^{+}13, LW13, LHC^{+}19, MVC^{+}11, MGH^{+}07, NM17, PLHLF05, SNR^{+}10, STW^{+}15, SLH^{+}19, WM13, YIY^{+}04, YKS^{+}12]. choice [VBL04]. chromatic [ÁLC22]. chromophoric [CMPNC^{+}22, GF19, HOY^{+}21a, RFSCF19]. Chronology [CFC^{+}18, SCMAR^{+}99]. chronometers [Coo65]. chub [ABE^{+}15, GiIKX22, TMÁGC^{+}21, YKNO23]. Chukchi [CQO^{+}15, CBB^{+}15, OACA20, APC^{+}21, BSC^{+}19, BDG^{+}17, CP17, CQC15, DWH^{+}14, FJA^{+}21, FMCG15, GMR^{+}23, GCFS06, HOY^{+}21a, HKE^{+}10, KPM^{+}23, KFH^{+}15, LCJ^{+}17, LPS^{+}19, LPBM17, LBC^{+}15, LDHW20, MSC^{+}15, OPL^{+}21, PLB^{+}23, PST^{+}02, SYN^{+}21, WMC^{+}89, WD94]. Chukchi-East [KPM^{+}23]. chum [SKSK06, YWUK15]. ciliate [DRVMC^{+}22, JYK^{+}14, WWW^{+}23]. ciliates [SPB93]. Ciliophora [WWW^{+}23]. CINCS [GCZ^{+}00, TP00]. circuits [BLP^{+}20]. Circulation [CTL^{+}04, CGZ^{+}16, DIM09, DWC06, Fuk91, GCS91, Has06, Hic92, HHP06, Hut95, JS90, KJH^{+}22, LP87, LPS^{+}19, LPF^{+}20, LWL87, Mit91, ON05, OJB99, ÖHÜ89, PMC21, RKS^{+}15, TMPM^{+}16b, TMPM^{+}16a, WBB^{+}01, ASFB^{+}13, ABM^{+}05, AGS10, Arb22, APHGC^{+}22, ABC^{+}99, ARG11, BAARB05, BLT^{+}15, BBPHG^{+}11, BLP^{+}20, BMC05, BSÖ^{+}94, BMG^{+}21a, Bum73, CGL^{+}20, CPG^{+}18, CSW96, Ché14, CD07, CM18a, CCH^{+}12, CGD^{+}22, DML^{+}16, DWH^{+}14, Dea85, DGMM85, DSC^{+}21, EMU21, FWO15, FMP19, GWB14, GSPP^{+}20, GR85, GCZ^{+}00, HMTL05, Ham90, Hau18, Hen85, HHDS02, HGTP^{+}19, HT97, HWB^{+}18, KK20, KRL08, Kes06, KA85, LBC^{+}23, LW85, LC16, LPW^{+}23, LGH^{+}21, LO85, Mac98, MMR^{+}09, MRRC73, MRMD^{+}97, MEMP15, MFS^{+}16a, MFS^{+}16b, MBKS08, MLS^{+}15, MB07, MJC^{+}17, MEMC05, MW96, NDEG22, NTU^{+}14, NGLSSG14, Nof00]. circulation [ORCH^{+}19, OAB^{+}16, OPL^{+}21, Owe91, PGLG^{+}05, PS91, PSK96, PTG95, PLJR22, PP85, PGC^{+}96, PFE10, Rea00, RBS^{+}20, RBS^{+}22, Rei86, Rei89, Rei94, Rei97, Rei03, RCB^{+}20, Ric08, RM89, Roo82, Rud15, SCPN15, STPHM^{+}23, STEB16, SOB^{+}08, SFMT12, SFMT14, SM01, SDS02, SDS22a, Sek86, Sek99, SZG06, SC23, SBLA10, SSM^{+}18, SON^{+}20, Sme93, SNdSR^{+}24, Ste91, SJ02c, SJ02b, SJM^{+}19, Suk88, SMP07, SCB^{+}16, Tal08, TRY^{+}04, TBK^{+}99, TAO05, Tsu86, TAH^{+}11, VNMS91, VAGMDRS22, WLD^{+}15, WSL20, Wen88, WSO01, WBH15, WTH12, WG82, XWL^{+}22, XDG^{+}23, YKH^{+}24, YGC^{+}21, YJS86, ZSW^{+}22, ZBLF23, MSJ^{+}15]. circulation-ice [WSL20]. circulations [HNR^{+}17, IHT^{+}21, MMF^{+}07]. Circum [BF11]. Circum-arctic [BF11]. circumglobal [GBG05]. Circumpolar [BH85, CPO^{+}19, CdD^{+}15, CP07, OYKK^{+}23, TFZS14]. CITHER [GA00]. City [Let87]. cladocerans [KOT^{+}21]. clam [YAK13]. Clarion [SLBH^{+}19, SLPA^{+}20, JSLA^{+}21, RMG90]. Clarion-Clipperton [RMG90]. clarity [RKK^{+}21]. class [DFH^{+}16, DHB^{+}21, SAB^{+}22, YFK21]. classes [CPPPEAG22, LBH^{+}21, LLX^{+}21, Peñ03b, PS98]. classical [vRGW10]. Classification [WR00, CTP^{+}18, DHB^{+}21, LC10, MMF^{+}17, MBH^{+}23, WWW^{+}23]. Claus [GBG05]. Clausocalanus [PMH17]. Clay [ORR^{+}02, Tur65]. Climate [APC^{+}12, BK08, BGL^{+}17, CBC^{+}06, CP02, DAvD^{+}20, DAIS10, Dur09, FHG03, GML^{+}23, HS02, LPF23, LS15, MLL^{+}22, MHTG10, MCB^{+}10, OÁT^{+}05, PL01, PCH08a, RB20, SHT^{+}01, YBS^{+}01, ZK06, BM01, BMG13, Bre82, BDC^{+}08, CDTM^{+}21, Cai95, CALS^{+}23, CAT^{+}08, CB17, DL69, Don94, DMF^{+}09, Dri11, DP13, DAKV99, FFA06, FPS^{+}13, HMRA^{+}03, HMWM00, Has82, HRSM08, HGB^{+}21, HHW01, HHW22, HWLT10, HSC^{+}16, HTV^{+}20, HMH^{+}15, INI^{+}17, IIM^{+}23, IIS^{+}17, JTQ^{+}18, JSA^{+}08, KKB00, KC15, KA94, KYS^{+}17, LAHI10, LMH^{+}13, LNB13, LM14, LCPSMR^{+}10, LS13, MPMFL^{+}23, MDAW^{+}19, MKB00, McI10, MBH^{+}01, McK08, MNM06, MS00, MVV^{+}19, MST^{+}23b, NDEG22, NPO^{+}19, NGPH10, NSE^{+}24, ORPRGIS22, OWH14, OIC^{+}23, PRTC13, PDAM^{+}15, PCR^{+}22, QLW10, RFFL21, RK03a, RAB^{+}11, RKC^{+}10, SLM^{+}16, SJA^{+}23, SBL^{+}23, SSVP00, SMG02, SKSK06, SRFHDH22]. climate [Sha82, STJ^{+}14, SYB^{+}15, SAY^{+}16, SDH^{+}14, SON^{+}20, SHF01, SBM^{+}23, SAB^{+}22, SKH00, SKT01, STGR^{+}14, TKW06, TLH^{+}15, TLP^{+}16, TAO05, TSH^{+}17, TSJC07, THM^{+}14, VMV^{+}23, WAH^{+}20, Was11, WJPHB15, WWSJ07, dlHRA^{+}18, LM10]. Climate-driven [RB20, OIC^{+}23, PRTC13]. Climate-forcing [ZK06]. Climate-induced [GML^{+}23, MLL^{+}22]. Climate-related [CP02, PCH08a]. climates [HMH^{+}15]. Climatic [AVG^{+}19, CD65, CM14b, FHG03, ZLKO00, BGM^{+}99, CMF^{+}09, CS03, Don65, Don87, Kaz17, Kra82, KRL^{+}22, Med87, Ola65b, SNZ^{+}20, SCHD23, WHI^{+}02, XYL^{+}22, YAK^{+}08, dMGS^{+}11b, dMGS^{+}11a]. Climatological [Ola65a, SK18, CPG^{+}18, KS06, Mor91]. climatologically [CGMP14]. climatologically-based [CGMP14]. Climatology [DHL^{+}21, ŠPM^{+}22, TWBC^{+}13, Cai95, CMHM18, HM08, KC15, LOC95, MW96, RZTD17, SKH00, WC15]. CLIOTOP [LM10, Ano10b, LM10].

Clipperton [JSLA^{+}21, SLBH^{+}19, SLPA^{+}20, RMG90]. CLIVAR [STPHM^{+}23]. Closing [MHA^{+}11]. closure [HMH^{+}15]. closures [JOGM^{+}10]. cloud [WM13]. cloud-free [WM13]. Clumsy [KN10, KN11]. Clupea [GBT^{+}19, Nag01, STF^{+}13]. clupeoids [EBvdL^{+}09]. CMIP6 [SKCP23]. CO [LGG18, RCGC^{+}16, SPK^{+}19, ZDG^{+}21, BL02, HLR17, SOWS17, ARELAK24, BK19, BF01, CKP^{+}20, CPG^{+}18, CKM^{+}21, DCD^{+}23, EFC^{+}23, EHSI12, FGS^{+}15, GDSCU09, GSF^{+}15, GLLB22, LM00, LGZ^{+}20, lLdZQ^{+}22, MKOLA20, OYKK^{+}23, OKdA^{+}19, PPKR14, RVS^{+}21, TŠT^{+}17, TAM^{+}15, WD94, WBA^{+}22, WWL^{+}24, WST^{+}16, WR03]. Co-Editors [WR03]. co-limitation [HLR17]. co-occurrence [SOWS17]. co-variability [BL02]. coalescence [WF17]. coast [AYH^{+}23, BFH01, BBSN04, BSC^{+}07, BM86, Bum73, CJMO87, CBB^{+}02, CdTH^{+}16, CCHM02, CLB^{+}14, Con87, CJG88, DLM91, Eme65, FB01, FGGDF^{+}04, HHB^{+}01, HFW^{+}98, HFO90, JJA^{+}13, JJA^{+}17, JJS03, KKB00, KP03, KGJ^{+}10, Lie88, LPARF^{+}20, Mar20, MMMWZ23, Mid69, Mit91, MMPG07, NEI^{+}22, OÁSG^{+}16, PELAA18, PKF02, RBPGJ^{+}20, RN02, SST^{+}17, STR01, TTMM^{+}17, VSGD21, WHT86, BASS^{+}20, FZ88, WLM07]. coast-fjord [SST^{+}17]. Coastal [BLT^{+}15, CWW15, FWO15, HMP^{+}13, HFO^{+}22, Huy83, JS87, KMMC09, KC15, KTW^{+}22, Pai20, RMHL09, SBD^{+}07, SP08, WF17, XLL^{+}20, ZJZ^{+}21, AJA^{+}22, AR18, ASJ^{+}23, ÁSFP^{+}03, AAM^{+}14, AMG^{+}16, AJHC19, BTS^{+}15b, BA04, BEP02, BMC17, BBPHG^{+}11, BCM^{+}02, BFJ18, BPGD^{+}14, BPGC^{+}20, BHLU^{+}07, BMG^{+}21b, Bri83, BLMR^{+}20, BCG^{+}08, CN22, CDH^{+}13, CJMO87, CTF07, CSS^{+}21, CAB^{+}18, CC88, CLB^{+}14, CCS^{+}21, CGD^{+}18, CNSHT15, CM14b, DLM^{+}12, DNNNN16, DWC06, ESTM13, ES07, EHG^{+}07, EM12, EHFD12, EBS^{+}18, EHSI12, FC07, FTSF21, FAAV^{+}15, FELJ16, FTG^{+}18, FLUC08, FAH^{+}13, GMBU12, GCCY^{+}14, GMAMB04, GSVB23, GASV^{+}09, GCB^{+}22, GLF^{+}17, GEO09, HSS^{+}12, HHDS02, HLGA07, HYM^{+}12, HEF^{+}12, HSF02, ILI^{+}12, JTQ^{+}18, JAS^{+}20, JJA^{+}17, JHW^{+}14, KCPM09, KOhL^{+}10, KLC^{+}15, KRL^{+}22, LOG^{+}09, LDB^{+}02, LDD^{+}22, LSB^{+}17]. coastal [LSD^{+}18, LLGS21, MERB12, MCG^{+}02, MBCB88, MDGC^{+}12, MFDH22, MGC^{+}18, Mit83, MMF^{+}12, MKOLA20, MDC^{+}07, MGH^{+}07, MTH^{+}10, MA12, NHS^{+}14, NXT^{+}17, NIF^{+}15, O’B83, OAWAN18, OAD22, PAM^{+}88, PK02, PD15, PAVB^{+}21, RCGC^{+}16, RNBP^{+}19, RTBR^{+}22, RF17, RÁSG^{+}13, RR01, RGM01, SCHBC^{+}22, SRF^{+}19, SAY^{+}16, SHF01, SMPC^{+}12, SAd^{+}17, SPV^{+}15, SCB^{+}16, SCS87, Tol85a, Tol85b, TMKJ^{+}09, USH15b, VSGC21, VDS^{+}18, VVV21, VSPP14, WM13, WD94, WWL^{+}22, WZFW16, WFS^{+}15, WH89, YNTS22, YSY^{+}19, Yos80, ZD17, vFB82]. coastal-offshore [WH89]. coastal-shelf [YSY^{+}19]. coastal-to-open [KRL^{+}22]. coastal-transition [GASV^{+}09]. coastally [HW02]. coastlines [CB17]. coasts [HAA^{+}14, Man69]. cobalt [SWT^{+}17]. Coburg [RZW^{+}23]. Coccolith [PGRP^{+}18]. Coccolith-derived [PGRP^{+}18]. Coccolithophore [KLIRK17, PHCA17, BRG^{+}15, GBB^{+}19, HBV^{+}10, ISM^{+}02, LCGH07, LSW^{+}21, OVG16, VHV^{+}12, Kru19]. coccolithophores [MPD^{+}22]. coccolithophorid [WPW^{+}14]. Coccoliths [MBP65]. Cod [HMA18, AHP19, BF11, BGL^{+}17, BHMS09, CHS^{+}24, Dol09, FKH^{+}13, HLP^{+}16, HvDL^{+}17, HCGK11, KSG^{+}17, LHC^{+}21, LNB13, MHTG10, RKC^{+}10, SLY^{+}15, Tit20, VMH^{+}21]. CODAR [SCB^{+}16]. code [SWH^{+}24]. coexisting [SM16]. Coherence [LO07, FBS22, HMS^{+}22]. Coherent [AVS23, MMF^{+}17, NO14, CDDF11, MBS20]. Coho [AHC^{+}13, BL02, IOGS13]. Cohort [SBFP21]. Cold [CS18, GLV12, YMK^{+}04, ALG^{+}21, BBB^{+}14, BLC23, CVHM^{+}18, CLG^{+}22, FVA^{+}19, GTS^{+}21, HGD22, HHB^{+}00, HDB13, JFG^{+}90, KiL14, LZF^{+}24, LRGV^{+}18, LWT^{+}20, MCB^{+}90, MSGGM18, MRW^{+}14, MHCS^{+}23, MSV^{+}14, OSH^{+}96, QYF^{+}24, SOWS17, SDB^{+}21, TAF^{+}22, VGJ^{+}19, YW22, vHMDL14, Cia22]. cold-pool [GTS^{+}21]. cold-spells [SOWS17, SDB^{+}21, YW22]. cold-water [ALG^{+}21, CVHM^{+}18, CLG^{+}22, FVA^{+}19, LRGV^{+}18, MRW^{+}14, MHCS^{+}23, TAF^{+}22, VGJ^{+}19, vHMDL14]. colder [HMH07]. collaboration [KY15]. collaborative [BN03]. collapse [BHMS09, DTC^{+}06, LNB13, SGLF^{+}13]. collected [MM80, SKF20, UKK^{+}19]. collection [CLdPHL23, MLL^{+}15]. collective [YSD15]. Collins [Ang80]. Collisions [WF17]. Cololabis [OOTA15, WCS^{+}23, XYL^{+}22]. colonial [SIB^{+}06]. colonization [TAF^{+}22]. colonizing [KSKN21]. Color [KKS^{+}19, BPGD^{+}14, McK15, WMB^{+}18, ZL24]. colour [BBE^{+}15, HHMB^{+}09, KPSB17]. Columbia [CMHM18, FDM^{+}13, GKS^{+}13, GDM^{+}15, LC10, RHBS13, STF^{+}13, TSC03, BJ90, CM14b, EKB06, Ham90, JS90, JGS90, JSHB90, LMH^{+}13, MB01, PMC16, SJH^{+}90, SC90, SSM90a, SSM^{+}90b, SMM^{+}90, SL13]. column [CWZ^{+}20, FTSF21, FK86, Her88, HGH^{+}19, JJL^{+}19, LPA92, NMK^{+}03, NKK03, PPCWJ18, SHP^{+}23, SDGVE17, SSTL16, SMP^{+}22b, SBD^{+}07, SYN^{+}21, TDL^{+}17]. Comau [MSV^{+}14, ZHSMM14]. Combined [STEB16, SLY^{+}15, FVLC^{+}23, LSV14]. Combining [CDP14, CRiI^{+}15a, LGL^{+}18, WFS^{+}15, TM13]. cometabolism [Sie88]. Commemorative [BDB^{+}22]. comment [Ano87d]. Comments [dZTG05, GBM^{+}01]. commercial [JRR^{+}24, SEG^{+}22b]. Common [MS02, CCM^{+}14, FDB^{+}21, HBK^{+}24, SGL^{+}18, SYB^{+}15, TTMM^{+}17]. commonalities [LSS^{+}09]. commons [LPF^{+}18]. Communities [HKGH^{+}06, ACE^{+}07, AAM^{+}14, AV23, Ang89, Ano94k, AUE^{+}14, BRR^{+}22, BAM^{+}09, BJMP19, BJMP20, BLP93, CDS90, CWB^{+}22, Car98, CQZ^{+}18, CMHM18, CGS23, CGD^{+}18, DAU22, DCRR^{+}22, DHHP18, DSV^{+}24, FBR^{+}13, FSAO22, FPY^{+}16, GdRGC^{+}14, GAPM16, HM90, HCC02, IG19, JMSB^{+}23, KM10, KZD^{+}19, KGJ^{+}10, LT06, LDB^{+}02, Law04, LMH^{+}13, LMT^{+}19, LEDR^{+}22, LBC^{+}15, Lon85, MCGR07, MSMH19, MSV^{+}14, NHH^{+}23, Nie07, OSH^{+}96, PELAA18, PLEF^{+}23, QOS^{+}22, RLP^{+}18, RÁSG^{+}13, SPSV^{+}20, SH09, TCL20, VPM^{+}19, VKDS^{+}18, gWjNfLyD20, WWW^{+}23, WHS^{+}23, XLL^{+}20, YNTS22, YGMR^{+}23, ZPC^{+}16]. Community [CML^{+}16, CSV^{+}07, CGS23, GPP22, INT14, JSHB90, Kos93, LdCSB^{+}20, MMES16, TSFA22, VDGGD^{+}22, AMFY20, AMG^{+}16, Ang79b, Ang84, AJHC19, ABT^{+}04, BKD^{+}20, BFJ18, BWB^{+}09, BAOM^{+}12, BLES16, CRGA17, CLSD18, CC88, CS03, CBM^{+}21, CKT^{+}13, CLMR23, Dag93, DDK^{+}18, DSR21, Dom84, DBR20, EHFD12, EHG^{+}12, FARRL^{+}13, FFS^{+}20, FLdST98, FMH02, FPS^{+}09, GSFP^{+}09, GFGGD^{+}23, GMDD^{+}22b, GASV^{+}09, GWM^{+}22, GHVG19, GvOSW11, GGQ07, GVKD^{+}13, HHB^{+}01, HGBG20, HLD^{+}21, hHRW^{+}05, HG04, IBW^{+}01, JTQ^{+}18, JRR^{+}24, JYK^{+}14, JFG^{+}90, KP03, KOhL^{+}10, KRHS14, LGK^{+}93, Law04, LdSH^{+}15, LSD^{+}15, MT99, MG02, MERB12, MCG^{+}02, MTC14, MDR20, MTK^{+}22, NHSP23, OPH^{+}24, PGY^{+}22, PCM11, PJH^{+}15, PBS22, Pug84, QSC^{+}15, RPG^{+}18, RBNJ^{+}12, RBPGJ^{+}20, RBHLA04, RAB^{+}84, Roe84a, Roe84b, RB84, RJT84, RF17, RLR^{+}18, SSB19, SNZ^{+}20]. community [SBB^{+}22, STW^{+}15, SWZS^{+}21, SST^{+}17, SPWH21, TAW^{+}15, TKSI08, TVT^{+}23, TKW06, TPP^{+}00, VGM^{+}23, VHV^{+}12, VMC^{+}19, WFH^{+}22, WLL^{+}23, WMWR08, WHK23, XYK^{+}22, XHW^{+}20, YNMY23, YPVP^{+}22, ZCV^{+}19, ZLX^{+}20, ZSBL00]. Comparability [LB20, BCF^{+}03]. Comparative [Ano09h, BHLU^{+}07, DM13, MLM09, MLHM09, THM^{+}14, BHMS09, CRF^{+}10, FBA09, GRDS10, HBL^{+}13, Hey78, HSC^{+}16, Mil88, PMMN^{+}22, VPS09, VCM04]. compare [BBE^{+}15, KHJ^{+}10]. compared [FMH02, ROBRB^{+}22]. compares [RN02]. Comparing [BE99, BAOM^{+}12, GA00, JPM^{+}08, LCBN14, MGE^{+}12, MPN09, TTB^{+}08a, WMWR08, WPA^{+}24, WJPHB15]. Comparison [ABM^{+}05, DCL^{+}13a, DP13, FMC^{+}20, HBV^{+}99, JPBB20, SBLA10, WZBK^{+}21, BMK12, BMO12, BWB^{+}09, BF11, BMG^{+}21a, CGV13a, CMM^{+}04, CM09, CSK^{+}12, CQC15, DHC^{+}20, DDDT99, DAF^{+}22a, DAF^{+}22b, DMF^{+}09, EBvdL^{+}09, FPJ^{+}15, GSFP^{+}09, GRMB18, GA01, GDM^{+}15, HMP^{+}13, HW02, HKPV12, HKGH^{+}06, hHRW^{+}05, HVEF09, LDB^{+}02, LSS^{+}09, MB20, MDB^{+}20, MHS^{+}09, PL09, RK03a, RCS^{+}11, SH09, SCB^{+}16, WLM^{+}22, YNTS22, ZSBL00]. comparisons [HRSM08, HHB^{+}22, MPV12, TKK^{+}05]. compartment [ESGP17]. Compendium [Ano10a, BPW10, PBOW10]. Compensation [BHK^{+}19]. compensatory [RKS01]. competition [ÁLC22, CLB^{+}13]. competitors [RSB^{+}13]. compiler [Koc65b]. Complex [BAP^{+}22, RA15, AGS10, BCB^{+}05, SBLA10, TSL10, WSO^{+}13, YGC^{+}21]. complexity [Den03, HVS10, KKO10, MFS^{+}16a, MFS^{+}16b, MCH^{+}12, RSG06]. complicated [Sie88]. component [Ang84, CDDF11, CP10, LLX^{+}21, MKHO96]. components [FMC^{+}15, HKGH^{+}06, HvdLS^{+}09, KKB00, MLD^{+}03, McC92, MMF^{+}12, ML09, MGH^{+}07, SG91, Tal08]. Composition [GIPG17, JFG^{+}90, PGT^{+}13, ABSDC07, AJHC19, BS90, BHHR15, BJMP19, BJMP20, BAOM^{+}12, BJ90, CSR90, CMF^{+}09, CCW^{+}18, DDE^{+}95, DDP^{+}00, DMD^{+}00, DDD^{+}00, DDCE^{+}23, FMC^{+}15, FMH02, FELJ16, FTHK19, GASV^{+}09, GVKD^{+}13, HCAFD^{+}20, Hey78, HWL^{+}20, HBH^{+}17, HVEF09, IPHW^{+}23, JLRB20, Kam19, KSR^{+}01, KT04, KKS^{+}19, Law04, LSB^{+}17, LSD^{+}18, LXC^{+}22, LdSH^{+}15, MG02, MCG^{+}02, MFDH22, MWS^{+}10, NKK^{+}05, PVM^{+}20, PPHM18, PGY^{+}22, PD15, RCS^{+}11, RCB^{+}20, RBPGJ^{+}20, RF17, SSB19, SM21, SBB^{+}22, SPWH21, SHT^{+}01, TVT^{+}23, TWBC^{+}13, TPM^{+}00, VMC^{+}19, WHK23, YMA^{+}17]. Compositional [RBL^{+}19, VCM04]. compounds [AB90, FK86, GLY23, LFBP^{+}13, NEI^{+}22, SGL^{+}13]. comprehensive [BHH^{+}16, CW06]. compressibility [Due77]. computation [MFM85]. concentration [ABS^{+}20, BMC17, FRV^{+}19, LTJ^{+}15, MSC^{+}15, MVC^{+}11, NM17, SNR^{+}10, UNN^{+}14, VK90, WAH^{+}20, Wil65, YIY^{+}04]. concentrations [ABP15, Ban96, BK19, FGL^{+}23, IHT^{+}21, KTN14, LDD^{+}22, LLL^{+}24, PLHLF05, RNL^{+}13, SM05, SKP99, TBW09, WM13, Whe93, WJE^{+}92, YGMR^{+}23]. Concepción [AMEV07, AMG^{+}16, BM07, CPNL07, DLM^{+}12, GGQ07, GMAB07, IVT^{+}12, MDC^{+}07, MGH^{+}07, MA12, MSL^{+}07, MDL^{+}12, SLG^{+}12]. concept [BRG^{+}15, BMG13, CWW15, LRW^{+}15]. Conception [CLMR23, DW02, FWO15]. concepts [CW06, PSK96, RCM^{+}03]. Conceptual [TR99, HSN^{+}18, WCB^{+}20b]. concerning [WO85]. Concholepas [GKC^{+}14]. concluding [Car97b]. Conclusion [Ano10a]. Conclusions [PBOW10, PAB^{+}87b]. Concomitant [FFS^{+}20]. concurrent [KOhL^{+}10, Min02, PVG^{+}20]. condition [BWMGCB08, BFV^{+}17, CQC15, DMC^{+}18, FMC^{+}20, GCD^{+}18, GDL^{+}15, HSG^{+}15, HBG^{+}21, MLB^{+}20, VDB^{+}20]. Conditions [EBS^{+}18, ASJ^{+}23, AMEV07, AAM^{+}14, AKH^{+}23, BP02, BH07, Ber65b, BL02, BASS^{+}20, Cai95, CBB^{+}02, CCS^{+}21, CPNL07, CCM^{+}14, DPB06, DWFP^{+}19, DBC^{+}23, DL17, DB02, EKB06, EMU^{+}23, EHG^{+}12, FSVL10, FDB^{+}21, GCCY^{+}14, GGQ07, HMP^{+}13, Hey78, HFO^{+}22, IPG^{+}16, JST^{+}24, KMU^{+}12, KGJ^{+}10, LBH^{+}87, LBP15, LPF^{+}18, LLAPG^{+}22, LWBD^{+}17, LGG18, MSd^{+}16, MAFS^{+}22, MMF^{+}12, MGH^{+}07, MKSW^{+}15, NRS^{+}19, Nag01, NRA^{+}21, OMS^{+}09, PGRP^{+}18, RNBP^{+}19, SF02, SS69, SGO^{+}08, SCHD23, SBC^{+}24, TTB^{+}08a, THP21, VKDS^{+}18, WHI^{+}02, XRC^{+}15, ZHBW01]. conducted [KDB95]. conference [MBH^{+}01]. Configuration [WLM^{+}22]. configuring [MCG^{+}14]. Confinement [ZLG17b, ZLG17a]. confirmed [YLY^{+}14]. confluence [RPPM^{+}23, WR00, ZZM^{+}24]. confocal [PKA19]. congeneric [WSH15]. connected [FBS^{+}18]. Connecting [Yu23]. connection [BPGC^{+}20, GIC20, Kaz17, NRS^{+}19, ŠVL^{+}15]. Connections [XC14, Car98, TSS^{+}12]. Connectivity [OACB^{+}15, ALT10, CDTM^{+}21, CZG^{+}21, HLP^{+}16, KAK^{+}22b, LAHI10, RKCH15, SGMVF14, ZLRVB24a, ZLRVB24b]. consensus [BT07]. Consequences [LSH^{+}11, Bea04, BMM01, CPC^{+}02, JLP^{+}20a, JLP^{+}20b, KHD22, Leh01, RBF^{+}09]. conservation [ALG^{+}21, GKC^{+}14, Kno04, LRGV^{+}18, McD81a, SSS^{+}11]. conserved [Nee85]. conserving [HAH^{+}22]. Considerations [SJ18, DHB^{+}21, dHA^{+}04]. considering [CSC^{+}12, His22, MRM^{+}14, MCMT^{+}17, XYL^{+}22]. consistency [BCF^{+}03, PGY^{+}22]. consistent [HHWW20]. consortia [Sie88]. constant [Ång65]. constituent [AK97]. constituents [SBBV04]. constraint [Kiv97]. constraints [BCL^{+}09, FKH^{+}13]. construction [SLBR18, ZZM^{+}24]. Consumer [DSB^{+}24]. consumers [WLP^{+}21]. Consumption [SSM90a, LH89, Lon95, SRT^{+}18]. consumptions [MRM^{+}14]. Contaminant [ARDP14]. contamination [CGD^{+}18, KFC^{+}13]. contemporary [HKGH^{+}06, Was06, Was15]. Content [Ano73b, CWB^{+}22, CM14a, DMT15, FMC^{+}20, Fra69, GvOS^{+}08, GA10, HWPLvW20, KSG^{+}17, OYKK^{+}23, PPSVC^{+}13, SRFHDH22, WLM07, Yas07a]. Contents [Ano64a, Ano65c, Ano65d, Ano69b, Ano85c, Ano86a, Ano87a, Ano89a, Ano92a, TRMV15]. context [BMN19, JFG^{+}90, LFG10, VBL^{+}21]. contiguous [CWW15, WKS^{+}15]. Continental [BHE^{+}98, BHPC06, BPP^{+}98, CTKF^{+}23, FLdST98, FKH^{+}13, Gor92, HM98, KFM^{+}17, LHEB98, MY23, OB98, PS98, SBPGP^{+}23, TvW98, WAH^{+}20, dWDB^{+}98, vWHdS^{+}98, ACK^{+}13, ASJ^{+}23, BHA^{+}14, Bum73, CB06, CSG^{+}15, CS06, Coo65, CFG07, DWH^{+}14, DMD^{+}00, DTOD00, DKRL22, DGGdR02, DCRR^{+}22, DCL^{+}13b, FZY^{+}23, FJA^{+}21, FARRL^{+}13, FMWW14, GRD^{+}23, GGG^{+}18, GEP^{+}08, HHK^{+}22, HVTV22, HWLT10, HHH^{+}12, HG04, HAH^{+}22, HWF^{+}21, Hut81, Hut87, HHK^{+}02, IG19, JW01a, JCIG18, KSVT00, KY23, Kit03, LM18, LLL^{+}11, LCBN14, LDB^{+}02, LSIB23, MH02, MZ14, ON05, ORW^{+}01, PCD^{+}18, PPHM18, RGMPR23, RGPB^{+}23, RAE^{+}05, SCMAR^{+}99, Sel65, SH09, SGO^{+}08, SBD^{+}07, SSL07, Sol00, TCN20, TCDPP^{+}22, TRLA^{+}13, TvG02, TSRF14, TPP^{+}00, TPM^{+}00, VHV^{+}12, WDMMK89, YMI88, YGC^{+}21, YCP^{+}12, ZCA21, dPCS23, vHVAT22]. continentales [Ber65c]. Continued [AvD15, KFM^{+}17]. Continuous [BGMP03, BDE03, HFK03, JJS03, RWJ^{+}06, UCB^{+}18, VBL^{+}21, WH94]. continuously [SMFM^{+}21]. continuum [Coo69, LML^{+}23, RBL^{+}19]. contraction [BCT^{+}09]. Contrasted [ERT^{+}22, VBAC^{+}21, BFV^{+}17, GGA^{+}05]. Contrasting [CDTM^{+}21, SWT^{+}17, TOKLC08, WCS^{+}23, AGD^{+}18, AKH^{+}23, DAF^{+}22a, DAF^{+}22b, DL17, EMU^{+}23, FPY^{+}16, GvOS^{+}08, GRdSS^{+}22, HG04, KKNT23, LT06, LSM^{+}22, LNB13, LdSH^{+}15, MHCS^{+}23, RNBP^{+}19, SWP^{+}13a, SH09, SCHD23]. contrasts [ST10, SQJ^{+}17]. contributes [PMS^{+}15]. contributing [SYN^{+}21]. Contribution [GMAGH^{+}17, GEO09, LSV14, MKS^{+}22, AMEV07, AOMZ^{+}23, BFH01, JJR^{+}08, MMF^{+}12, RKS^{+}15, SMB88, WSC^{+}21, Wüs64]. contributions [CPSM20, GS24, LNB13, SSB^{+}20b, ZGZ19]. control [AHSS22, ACE^{+}07, AUE^{+}14, Dag93, FZY^{+}23, GAF15, HS07, HLS^{+}14a, HS02, HM06, KGdS^{+}08, LLL^{+}11, NGNV12, PPD^{+}12, RGMPR23, WBC^{+}22, YZX^{+}23]. Controlling [MVN^{+}15, CMC^{+}16, CVBG21, HBV^{+}99, HMKF08, LLGS21, SOS^{+}07, SGS^{+}23, VMC^{+}19, XCH^{+}16]. Controls [KSR^{+}01, OMR^{+}22, WAH^{+}20, YSN20, BFV^{+}17, CIL^{+}23, CBPS^{+}22, DWH^{+}14, FBB^{+}21, FRCH15, FPY^{+}16, FDE^{+}22, KDB95, LFBR^{+}18, MLPN06, OPL^{+}21, PAVB^{+}21, Pow06, SCHS^{+}24, SR15, VPM^{+}19, VCM04, dlPPÁB24]. controversial [Cia14]. convection [AJA^{+}22, KFG^{+}03, McD81a, McD81b, PPdM^{+}12, Sch03, TDGY22, YN03b]. convective [DLM^{+}96]. convergence [RASVB^{+}22]. conversion [SNdSR^{+}24]. Converting [FCEZ10]. CONVEX [Rea00]. CONVEX-91 [Rea00]. conveyor [HS07]. conveys [STJ^{+}14]. convolutional [KM22]. cool [BMG^{+}21a, FMM^{+}20]. Cooling [RKS01, SDGVE17]. Cooperative [ALV^{+}21]. coordinated [DLM^{+}96, ÖHÜ89]. Copepod [BHC^{+}18, GHVG19, HCC02, HVEF09, LPHL^{+}05b, LPHL^{+}05a, MTK^{+}22, YHRT22, AVK91, BRE^{+}24, BEP02, BMdMS^{+}21, BD20, BAOM^{+}12, BMG^{+}19, CCG07, Dag93, DWFP^{+}19, DLJ^{+}21, EHFD12, FELJ16, FCEZ10, Fro05, FB05, GBG05, GHC^{+}17, GD91, HTG15, Joh04, MDGC^{+}12, MTH^{+}10, NGPH10, NGNV12, PD15, Reb02, SSTD^{+}95, STS^{+}12, SJJ^{+}03, SJD10, TKSI08, UB10, WPB^{+}08, WSH15, ZGB^{+}20]. Copepoda [FMT15, GS19, GPP22, HLPL05, JC04, Mil88, NMN08, PMG15, PMH17]. copepodites [NMN08]. copepods [BVJE19, BHS^{+}15, CTF07, CÁM06, CSK^{+}12, DLD^{+}19, DAF^{+}22a, DAF^{+}22b, HHB^{+}01, HE07, HEF^{+}12, IS19, JAC^{+}12, KKKS14, KVNT20, KSKN21, KHP^{+}18, LBP15, MMG^{+}13, MSMR93, MFB^{+}84, MC88, Mil93a, Nie07, OPG^{+}10, PK02, Roe84b, SMN^{+}13, SS03, SSV^{+}11, THBA19, VJJ^{+}22, WGG^{+}08, YHLA^{+}04, dPGSHL23, MN88]. Coral [GCF^{+}19, MP04, ALG^{+}21, CPPPEAG22, CALS^{+}23, CVHM^{+}18, CLG^{+}22, CKL^{+}14, CDB^{+}24, FVA^{+}19, KAK^{+}22b, LRGV^{+}18, MRW^{+}14, MHCS^{+}23, NYL^{+}17, PRA^{+}18, TAF^{+}22, TKC^{+}22, VGJ^{+}19, WFJ^{+}15, YW22, RDG^{+}21]. corals [GGSM^{+}20, vHMDL14]. core [CS18, CQO^{+}15, LJ65, LXC^{+}22, NC80, Soh03]. core-use [CQO^{+}15]. corer [JPBB20, LB20]. cores [AS20, CD65, FAAV^{+}15, Ola65b]. correlation [Ber65c, KC15, Par65]. correlations [Con87, Emi65]. Correspondence [CCS^{+}21]. corridor [LRJ^{+}15]. Corrigendum [ÁBMÁS15, BLAM00, BJMP20, Fei04, FDH20, GFB^{+}15b, JLP^{+}20a, KN11, Kru19, KMF^{+}20a, MHS^{+}20a, MFS^{+}16a, RBS^{+}20, RLSF07, SE09, SHC^{+}07, ZLRVB24a, dMGS^{+}11b]. Corsica [GHL15, BMC05, Con87]. CoSiNE [ZCH^{+}17]. Costa [SDS^{+}22b]. Could [DWFP^{+}19, IFC^{+}07]. counter [ÁSFP^{+}03, BTNK13, HMX^{+}23, MCT03]. counterclockwise [TOiF^{+}12]. counting [CTP^{+}18]. Coupled [DWH^{+}14, HKK12, LC22, BBE^{+}15, CGZ^{+}16, GL23, HLK13, HMH07, MMGL^{+}07, PMC16, Pie01, PST^{+}15, RFC^{+}15, SGWF^{+}19, SNV^{+}18, SDJ14, TS10, USH15b, VBM21, WSL20, WSS15, WWZ19, WWSJ07, YWUK15, ZCH^{+}17]. Coupling [FELMGM^{+}22, HHDS02, HE07, STM10, BD20, CW06, DTW^{+}00, HJLLN07, KLC^{+}15, LHC^{+}19, PAM^{+}88, SHC^{+}06, SHC^{+}07]. covariance [WBA^{+}22]. covariation [MHS^{+}09]. Cove [GBB96]. Cover [Ano17i, Ano17j, Ano17k, Ano17e, Ano17f, Ano17g, Ano17h, CDB^{+}24, DWFP^{+}19, FMCG15, GW89, LSH^{+}11, ORR^{+}02, RBR^{+}23]. CPR [BCF^{+}03, BBE^{+}03, BIL03, CL03, VR03]. CPUE [LOBG^{+}10]. crab [AHRT90, BL02, BMN19, DTG^{+}24, MGKW19, RKCH15, YFK21]. crabs [ZK06]. crack [GMD^{+}22]. Cracking [SWH^{+}24]. crawler [CAO^{+}20]. created [BEI^{+}20, WPB05]. Cretan [GCZ^{+}00, GSPMAI99, KSPK99, SKP99, TBK^{+}99, BTK^{+}99, CLG^{+}00, CLA^{+}00, DDDT99, DDP^{+}00, DMD^{+}00, DTOD00, DDD^{+}00, DTW^{+}00, GSPMAI99, KSVT00, KKS^{+}18, KTB^{+}99, PLP99, PTI00, SK21, SCT^{+}00, TZP^{+}00, TPM^{+}00, TVD^{+}99, VCB^{+}00, VKT15, VOT^{+}99, WYT00]. Crete [TPP^{+}00]. Creus [LRGV^{+}18, DCRR^{+}22]. cristatus [MFB^{+}84]. criteria [Ano94k]. criterium [DFD23]. critical [BM01, Kir06, Sim81, WBA^{+}22, vRGW10]. Cross [AFH^{+}11, CCD^{+}13, HPZC21, LMM03, ZAC^{+}23, BASS^{+}20, KFC^{+}23, LZG20, MHS^{+}09, PMC21, RCSHW22, WC15]. cross-ecosystem [MHS^{+}09]. cross-frontal [KFC^{+}23]. Cross-shelf [HPZC21, ZAC^{+}23, BASS^{+}20, LZG20, PMC21, RCSHW22, WC15]. Cross-shore [CCD^{+}13, LMM03]. crucible [SEO13]. cruises [AFH^{+}11]. Crustacea [JZ19, SBS90, Thu90, Roe84a, ACK^{+}13, BBFS19, BHK^{+}19, CHSB^{+}21, GPP22, JPB20, JLRB20, MB20, MKD90, PMG15, PKA19, RBPGJ^{+}20]. crustacean [ACL^{+}18, HWL^{+}20, PCR^{+}22]. crustaceans [ACB^{+}13, CJRÁ^{+}13, SLBVRR^{+}22]. crustal [IU14]. crusts [Gri22, VCM04]. cryopelagic [HKE^{+}10]. Cryptic [GWGR^{+}19]. Cs [AFH^{+}11, IHT^{+}21, KKS^{+}03, PAF^{+}11, SCMAR^{+}99, SCLG^{+}11, TAH^{+}11]. Csanady [PCK^{+}06]. CSK [ALV^{+}21]. CTD [KSK21, KHM^{+}88, PCBA^{+}20]. ctenophores [RJT84]. CTZ [RBHLA04]. culture [PMMN^{+}22]. Cumulative [BHK^{+}16, TCS15]. Current [AH10, AAM^{+}14, ARG11, ABP^{+}23, BBM^{+}14, CCW^{+}02, CBGC^{+}08, CB09, CRT^{+}22, CRHM12, CLdPHL23, DFM^{+}21, DRD^{+}07, DPF^{+}20, EBR^{+}14, FMC^{+}20, GPAB^{+}16, GDI^{+}09, GBC^{+}16, GCED22, Hau84, HLGA07, Hic79, HW02, HMS^{+}22, Huy83, JBB^{+}14, KSD84, KAK^{+}22b, LO07, LQU07, LO21, LPF^{+}18, LABD^{+}24, LB02, MKM86, ML09, MAB^{+}11a, MAB^{+}11b, MFM15, PO15, REG^{+}15, RMK^{+}21, RD03, RZTD17, SFMA20, SDK84, Sim84, SKHD84, SP08, SBG^{+}08, TTB^{+}08a, TTB^{+}08b, TBW09, TMÁGC^{+}21, TSJ^{+}12, XLL^{+}20, ÁSFP^{+}03, AASJ23, BPC^{+}05, BPM^{+}14, BC88, CP17, DWS^{+}24, DDJ^{+}21, FZ88, GDM^{+}20, GGPG^{+}19, Ike88, IHT^{+}21, KMOM88, Ken88, LG22, Lie88, LGG18, LMM03, McC92, MS02, NNM^{+}21, NH88b, NH88a, NPO^{+}19, NC80, PSP^{+}21, Sie69, SÖÜ94b, TWAL^{+}11, TMH^{+}16, TKW06, TKWI08, UCB^{+}18, Web69, WWN^{+}99, Yux88, Zen08, ZYN^{+}24]. Current [ASFB^{+}13, ABE^{+}15, AN04, AH80, Ant09, ABÁS^{+}09, AE09, ABP15, BRB^{+}01, BAT^{+}98, BFR13, BDT^{+}08, BB14, BL02, BLT^{+}08, BH85, CCD^{+}13, CdD^{+}15, CF07, CP07, DSBP15, DLJ^{+}21, ESTM^{+}12, EB08, EBvdL^{+}09, FRCH15, FFA06, FVLC^{+}23, GW91, HvdLS^{+}09, IMW^{+}14, JTQ^{+}18, JJR^{+}08, JAJS08, Jón07, JJ08, KHL12, KL86, KAK^{+}22b, LS20, LJPGC02, Lav09, LG23, LBP15, MPC^{+}17, MJC^{+}17, MCT03, MDR22, NJCD01, NBR^{+}08, NHN^{+}21, OBD^{+}20, ORCH^{+}19, Pai20, Par86, PBS22, RLL^{+}09, RBPGJ^{+}20, RFC^{+}15, RBE^{+}12, SGF^{+}19, SGWF^{+}19, SSS^{+}11, SFS^{+}12, SHL13, Sch83, SHD^{+}21, Sek86, SBM^{+}23, SKRM^{+}95, SHT^{+}01, STGR^{+}14, TCN20, TWAL^{+}11, TFZS14, Ven12, VAGMDRS22, WOW^{+}14, WR00, YKWF21, YJ88, YYhT^{+}17, ZLC^{+}15, dSPF^{+}23, YJW88]. current- [ZYN^{+}24]. current-meter [UCB^{+}18]. Current/Florida [KAK^{+}22b]. Currents [GSA^{+}20, KZSH85, SBPGP^{+}23, BBL^{+}09, CTL^{+}04, CP07, Dea85, DEW^{+}97, FY88, FBS22, HBW17, Hut81, Hut87, Hut92, ICB^{+}19, JJA^{+}13, JSKM02, KAG^{+}19, LG22, LG23, LGH^{+}21, PL89, Pra97, RI86, Rog00, SVU02, STGR^{+}23, TKC^{+}22, WO85, WCS^{+}23, XD95, YMI88, YSS14]. curve [McK08, Oll15]. curves [ED82, FGS^{+}23]. cut [IST^{+}88]. CWC [ALG^{+}21]. cyanobacterium [LMT^{+}19]. cycle [AYH^{+}23, ADS^{+}22, BNC05, BB24, BMM01, BDBJ01, BM07, CPG08, Car97a, CVBG21, CFG07, EALF08, Her97, HTdM^{+}15, KYT^{+}16, LB02, MLB^{+}20, MST^{+}23a, MW96, PRTC13, PST^{+}15, RGC^{+}01, RNBP^{+}19, RG09, SVL^{+}23, SSTD^{+}95, TWMY08, TSFA22, VBA^{+}18, WP91, WSO01, WCN^{+}05, Yu23]. cycles [AH15, FRCH15, HE07, HTV^{+}20, KRHS14, LK13, Lon95, MDC^{+}07, RM93, SBD^{+}07, TAO05, TDL^{+}17]. Cyclic [DBC^{+}23]. cycling [BSMC15, BHE^{+}98, BGS^{+}04, CBPS^{+}22, DHDM22, FC07, HPC^{+}20, MBP^{+}11, MPM^{+}17, MMF^{+}12, OB98, PPCWJ18, RN06, SE16, SMP^{+}22b, WMC^{+}89, WD94, WH20, WRS^{+}92]. Cyclone [PTZ^{+}23, VVV21, BBB^{+}21, KM22, LOO22, VBM21]. cyclones [LYM12, MMR^{+}12, SCY^{+}23]. Cyclonic [LKDL14, MD07, MZK^{+}23, MSS^{+}02, WOW^{+}14, XHC^{+}20]. Cyclorhagida [AM19]. cygnus [WOW^{+}14]. cyst [BPTT19, SPF^{+}23]. cysts [KPM^{+}23, ZHBW01]. cytometry [LSM^{+}22, LGL^{+}18, SHS^{+}05].

D [BHH^{+}16, CNSHT15, HYM^{+}24, MMR^{+}09, SW01, Tho77, XRC^{+}15]. Dabob [FB05, HL05, HPHL^{+}05, PHLL05, PLHLF05]. Daily [PAM^{+}88, CHC^{+}12, WFS^{+}15]. Dall [NHE^{+}13]. Dam [Ore69]. damming [CMF11]. Danois [CSV^{+}07]. dans [Ber65b, Rou65]. dark [BRD^{+}15]. Darwin [BB65]. Data [dHA^{+}04, ARD^{+}03, ARH^{+}00, ATC^{+}19, AHC^{+}13, ASR^{+}20, BMGN15, BBE^{+}03, BIL03, BBM^{+}14, BDL08, BECA22, CFM^{+}18, CL03, CLdPHL23, CGD^{+}22, DTC^{+}06, DHL^{+}21, Egb97, EAB^{+}23, FPJ^{+}15, GCCY^{+}14, GdRGL^{+}01, GJ00, GA00, GAS^{+}22, HMRA^{+}03, HLS^{+}14a, HHH^{+}12, HM00b, KDL^{+}01, Kiv97, KBSB18, LCBN14, LHE^{+}13, LAGM^{+}23, LB20, LHF^{+}16, LM97, MDB^{+}20, Man04, MB05, MAB^{+}11c, MAB^{+}11a, MAB^{+}11b, MZGA^{+}20, OMS^{+}09, ORVES17, OIC^{+}23, OÁT^{+}05, PSP^{+}21, PRC^{+}20, PFW15, PHD^{+}18, REG^{+}15, RBS^{+}20, RBS^{+}22, RDL^{+}91, Ric93, RWJ^{+}06, RBS^{+}09, SPC^{+}23, SO91, SKCP23, TM13, TKWI08, TFZS14, TGR05, UPPS^{+}21, VGJ^{+}19, WM13, WCX^{+}21, WSO01, WLM07, WLM^{+}22, WFS^{+}15, Woo18, WZC20, WWL^{+}24, XLX^{+}20]. data-assimilative [WLM^{+}22]. data-poor [EAB^{+}23, GAS^{+}22]. database [VR03, dlHRA^{+}18, SAA^{+}15]. dataset [DHB^{+}21, KS06, LGZ^{+}20]. datasets [GCP08, PCBA^{+}20]. dated [SCS87]. dates [ST65]. dating [Emi65, OE65, TAF^{+}22]. day [Cai95, DJG^{+}02]. days [HBD^{+}18]. dead [WWL^{+}22, WLL^{+}23]. Dealing [LIH^{+}12]. Dease [CML^{+}16]. Death [GHVG19, Wal83, WKS^{+}15]. debris [SGF^{+}19]. Decadal [JG07, JCIG18, LAD^{+}18, Lev88, LMC^{+}20, SGWF^{+}19, Wu13, CSS^{+}21, DMT15, HDZY15, HHH^{+}12, KLB^{+}21, KKK^{+}04b, LYM12, LDD^{+}22, lLdZQ^{+}22, MTC12, MCD^{+}07, MHCS^{+}23, Mol04, Mol22, NNM^{+}21, OMS^{+}15, QNK^{+}22, SF15, SON^{+}20, TSH^{+}17, WHI^{+}02]. Decadal-scale [SGWF^{+}19]. decade [MWS^{+}10, MMG^{+}11, PHKS17, VLUC^{+}07, WBD^{+}15, dPAJ07]. decades [ABD^{+}17, CBM^{+}21, GS24, HBD^{+}21, KOT^{+}21, LWY07, LDMH09, McK15, MJC^{+}17, PBB^{+}20, PM13, RL23, SEG^{+}22b, SEG22a, TKW06, VBL^{+}21]. decapod [ACB^{+}13, CSV^{+}07, PPPdS20, PPdS21, Roe84a]. Decapoda [CHSB^{+}21, ACK^{+}13, Mar20]. December [Ano20s, Ano22-33, LM10, Ano98b, Ano99e, Ano03f, Ano04f, Ano07q, Ano11i, Ano12p, Ano16n, Ano17o, Ano19a, Ano21c, Ano22c, Ano23c]. Deciphering [RPRCAG^{+}21]. decision [BMN19, JPM^{+}08]. Decline [PBBH^{+}22, BRG^{+}23, Jac10, MPB^{+}23, ONR^{+}14]. Declining [RNL^{+}13, WFR07]. decompositions [MNT14]. Decrease [PMS^{+}15]. decreases [DLD15, VFS^{+}15]. dedicated [BFPS06]. Dedication [vWM02b]. deduced [Soh03]. Deep [BS95, CF20, FBD18, dCFK17, FJ19, GDN^{+}18, GLAHH^{+}22, Hen85, JFEC13, Mun69, RGI05, SKF20, Tit20, VBJ^{+}20, WWSJ07, ZBLF23, vHCY^{+}20, ACK^{+}13, ACB^{+}13, AS20, AP20, AS96, Ang89, Ano94k, AUE^{+}14, BW65, BRC^{+}18, BSF95, BS90, BMNW01, Bil01, BKC15, BHC^{+}18, BAB^{+}19, BBRM20, BGS^{+}04, CDS90, CCM^{+}13, Car98, CGM^{+}02, CMM^{+}04, CSV^{+}07, CPG08, CMF^{+}09, CMF11, CJRÁ^{+}13, CAO^{+}20, CMHM18, CP19, CFML22, CTR^{+}19, DDE^{+}95, DTOD00, DP18, DOS^{+}18, DSB^{+}24, DGMM85, DBJ^{+}15, ECGP01, Eri65, FPD^{+}01, FPIJ85, FARRL^{+}13, FBT^{+}22, FSVL10, FWL^{+}15, GBM^{+}01, GM19, GvOSW11, GBB96, GD85, GJ00, GGA^{+}16, GGSM^{+}20, Gri22, GVKD^{+}13, HM90, Hau18, HS22, HGTP^{+}19, HLCdP19, HCV^{+}20, HWBT03, HG04, HWB^{+}18, JPB20, JM19, JP90, JSLA^{+}21, KSK21, Koc65a, KFC^{+}13, KM22, Kos93, LM18]. deep [LF12, LRNK99, LFI^{+}13, LB20, LSIB23, MGF^{+}13, MGS90, MKD90, MNT14, McC92, MD07, MM80, MMN12, MM90, MSFZ19, NMK^{+}03, NKK03, Ola65b, OT19, ORMR^{+}19, PPdM^{+}12, PPHM18, PMFNGQ21, PVB23, Peñ24, PPdS21, PVA24, PB94, PdMS^{+}13, RGC^{+}01, RLDC^{+}13, RCC^{+}18, RM97, RSW^{+}23, RKM^{+}07, Rot65, RSD^{+}90, SSB20a, San73, SCB^{+}09, SSKA19, Sha82, SSTL16, She65, SFAD^{+}90, SLGI^{+}21, Soh03, Sok90, SV97, SCT^{+}00, SM16, SWH^{+}24, SBS90, Tal08, TRLA^{+}13, TCS15, TTF^{+}22, TVD^{+}99, TCL^{+}15, Tur65, UPPS^{+}21, VKGP^{+}13, WO85, WGCS13, Web69, WWN^{+}99, WHBW03, Wüs64, XHW^{+}20, YTL^{+}19, Zez90, ZKT88, vPRT90, vdS94c]. deep-Mediterranean [CTR^{+}19]. deep-ocean [HWB^{+}18]. deep-pelagic [LSIB23, RSW^{+}23]. Deep-sea [FBD18, FJ19, Mun69, SKF20, VBJ^{+}20, ACB^{+}13, AS20, AP20, AS96, Ano94k, BRC^{+}18, BS90, Bil01, BAB^{+}19, BBRM20, BGS^{+}04, Car98, CGM^{+}02, CMM^{+}04, CPG08, CMF11, CJRÁ^{+}13, CAO^{+}20, CMHM18, CP19, CFML22, DTOD00, DOS^{+}18, DSB^{+}24, ECGP01, Eri65, FARRL^{+}13, GM19, GvOSW11, GBB96, GD85, GVKD^{+}13, HLCdP19, HCV^{+}20, HWBT03, JM19, JP90, JSLA^{+}21, Koc65a, KFC^{+}13, Kos93, LB20, MGS90, MKD90, MM80, MMN12, MM90, MSFZ19, Ola65b, OT19, PPHM18, RGC^{+}01, RCC^{+}18, RSD^{+}90, SSB20a, Sha82, Sok90, SM16, TCS15, TTF^{+}22, Tur65, VKGP^{+}13, Web69, WHBW03, Wüs64, YTL^{+}19, Zez90, vPRT90]. deep-slope [CMF^{+}09]. Deep-water [Tit20, BW65, CSV^{+}07, Gri22, HG04, PPdS21, SCB^{+}09]. deeper [KY14]. Deepest [FTHK19, BHK^{+}19, DSAB20]. deeply [JFEC13, ZLG17a, ZLG17b]. deeply-mixing [ZLG17a, ZLG17b]. Deepwater [DYL^{+}15]. define [CSBL^{+}15]. Defining [MD10, TCL20, ZWP23, HAP^{+}16, PHKS01, SFS^{+}12]. definition [Ban65]. Definitions [ORMB08]. deglaciation [ZLC^{+}15]. degradation [FBR^{+}13]. Degraded [CMC^{+}16, VBJ^{+}20]. Degree [HBD^{+}18]. Degree-days [HBD^{+}18]. delays [MMD^{+}16]. deleteriously [HLPL05]. delimitation [JGB20]. delimited [ARELAK24]. Delimiting [LJM^{+}16]. Delphinus [SGL^{+}18]. delphis [SGL^{+}18]. delta [CFC^{+}18]. demand [GSSWK20, MMKS^{+}21]. demands [WMC^{+}89]. Demersal [KSVT00, ERT^{+}22, FTC^{+}16, GFB^{+}15b, GFB^{+}15a, HM90, PSA^{+}19, VBJ^{+}20, XYK^{+}22]. demographic [BEP02, GTS^{+}21, SGA^{+}19]. demography [HRA^{+}08]. Demonstration [BMG^{+}19, RFC^{+}15]. demosponges [SGA^{+}19]. dendritic [AHD18]. Dendrobranchiata [CHSB^{+}21]. denitrification [BHE^{+}98, PLN^{+}23]. Denmark [JJR^{+}08]. Dense [BBB^{+}14, PNF^{+}21, RTN90, BGV^{+}23, ISH^{+}04, NAH^{+}21, ONR^{+}14, PdMS^{+}13, RCC^{+}18, RKM^{+}07, SGL^{+}17, UNN^{+}14, VKT15, dZTG05]. dense-water [RKM^{+}07]. density [BFB^{+}20, CP02, DDK^{+}18, GDN^{+}18, Her88, JS90, KSVT00, Mol22, RCF^{+}13, Sie69]. deoxygenation [BHK^{+}16, CWS^{+}21]. dependence [BBS21]. dependency [KBF^{+}08, KSK21, LMS93]. dependent [ACK^{+}13, BECA22, OIC^{+}23, ZZM^{+}24]. depicted [VNMS91]. depicting [BDP^{+}06]. depleted [AIHB^{+}07, DGVGR24, MPN09]. depletion [ASB^{+}08, PAVB^{+}21]. deployments [BKD^{+}20]. Deposition [GvOS^{+}08, BJ17, GCD^{+}99, JGO^{+}98, LGG18, Ola65a, RDD^{+}18, SvWRvB02, TAW^{+}15, VOG^{+}08, VBA^{+}18, WDK^{+}01, YSY^{+}19, ZCLS20, dWDB^{+}98]. deposits [LCJ^{+}07, SCS87]. Depth [BLES16, BAB^{+}19, FTC^{+}16, Phl65, UAM05, ATS01, AHRT90, CDS90, Car98, CSMGS19, CSG^{+}15, CCB^{+}20, DW02, ED82, GGJ^{+}10, HLTB^{+}17, LT06, Lev88, PFW15, RSW^{+}23, RLR^{+}18, UKK^{+}19, VFS^{+}15, XY21, vRGW10, BHK^{+}19]. depth-distribution [GGJ^{+}10]. Depth-related [BLES16, FTC^{+}16, UAM05, AHRT90]. depths [BMK12, CBL^{+}19, HFO90, Mar20, OT19]. Derelict [GCF^{+}19]. derelicts [Ric85]. derived [ASR^{+}20, BMC05, BKD^{+}20, BS95, Cai95, DHL^{+}21, HMRB^{+}03, HTdM^{+}15, LCBN14, LC16, LW13, LLX^{+}21, MDR22, PGRP^{+}18, PLK14, RBS^{+}20, RBS^{+}22, RDD^{+}18, SCMAR^{+}99, SW92, SJ02c, SJ02b, SLH^{+}19, SI97, WPA^{+}24, WPB^{+}08, ZYN^{+}24, ZD17]. described [JZ19]. describing [Ano94k, GLF^{+}17, GLS08, PBN13]. description [AHW99, BDP^{+}06, Bum73, FGSA97, GBG05, JGB20, Mil09, Mit83, Owe91, PMG15, Rud15, iSIS02, MFM85]. descriptions [CES^{+}19, CP19, Mar20]. Descriptive [Hof81]. Descriptors [MLM09]. Desert [FAAV^{+}15, RDD^{+}18]. Design [Whi95, Ben85, SPC^{+}23, SCY^{+}23, TDH^{+}95, TIOM16]. desmosomatid [JGB20]. destinations [LGH^{+}21]. detail [AF10]. detailed [VMA^{+}24]. detected [KCL^{+}12]. Detecting [LHF^{+}16, dHA^{+}04, Man04]. detection [BECA22, BPGD^{+}14, BLC23, OP18, ON22, Tho87]. determinants [SSS^{+}11]. Determination [HF65, YKWF21, Koc65a, SC65, XD95]. determine [LJM^{+}16, Pra04, Suk88]. determined [DEW^{+}97, WSO01]. determines [PDAM^{+}15]. Determining [SQJ^{+}17, Dav99, FRCH15, SRAV19, TNY^{+}24, WSS15]. detours [MIN^{+}20]. detrital [LLL^{+}11, SMM^{+}90]. detritus [GMAB07, KBC^{+}22]. develop [JRR^{+}24]. developed [SFAD^{+}90]. Developing [MCH^{+}12, RSK^{+}23, SJP10, XY20]. Development [CBHL07, HFK03, JHDT12, MN88, Mil93a, WCB20a, BMK12, CZL^{+}24, Eme65, EBW^{+}23, FHP83, GPEV20, GMDD^{+}22b, KAH^{+}16, PM13, PFHM10, RAB^{+}11, SAB^{+}21]. developmental [BM76, OAWAN18]. developments [FPJ^{+}15, JAS^{+}20, Szu12]. device [RLC85]. devices [AKAL20, ICB^{+}19]. diabatic [WCN^{+}05]. diadromous [EAB^{+}23]. diagenesis [RGC^{+}01]. diagenetic [GLS08]. diagnosed [APHGC^{+}22]. Diagnosing [CSS^{+}21, PKV18]. diagnostic [OACA20, YJS86, YLY^{+}14]. diagram [Tom81b]. diagrams [Ric08]. diapausing [JC04, KHP^{+}18, WBC^{+}22]. diapiric [OSH^{+}96]. Diapycnal [KY23, RPPM^{+}23, HDB13]. Diatom [Fro05, RCB^{+}20, AAML22, FPJ^{+}15, FB05, HL05, HLPL05, ILA21, JTQ^{+}18, KV13, KV18, KTW^{+}22, MST^{+}23a, OMS^{+}09, PHLL05, PTPY^{+}23, PMS^{+}15, RFKC16, RF17, ST03, ST10, TSFA22, WPB^{+}08]. diatom-derived [WPB^{+}08]. diatomaceous [LvIKB07]. Diatoms [Don65, PMA^{+}14, SYN^{+}21, CLV^{+}19, GDI^{+}09, GGE^{+}65, HHAR23, KV18, KBE^{+}22, LPHL^{+}05b, MDC^{+}07, TKK^{+}05, WPB^{+}08]. DIC [PRL^{+}18]. dicarboxylic [SCCJ^{+}18]. Dichato [RBPGJ^{+}20]. did [JvdLL^{+}15]. DIDSON [GDN^{+}18]. Diel [Her88, UPPS^{+}21, dPGSHL23, Ang84, Ant09, BEP02, CDL^{+}22, Dom84, GBB^{+}20, HGD22, OACA20, OCH^{+}18, Pug84, RAB^{+}84, Roe84a, Roe84b, RB84, RJT84, RG94, SK17, ZGB^{+}20]. Diet [ABE^{+}15, BOMdP15, EBD^{+}20, ESD^{+}21, EBvdL^{+}09, HHB^{+}01, AGD^{+}18, BFH01, BGM^{+}10, BD20, CPG08, CQC15, DIQJ21, GMR^{+}23, HGBG20, SHT^{+}01]. Dietary [NHE^{+}13, SM16]. diets [DWNN04, EBvdL^{+}09, Gif93]. Differences [LBP^{+}21, Tan99, ACK^{+}13, AAMB^{+}24, CP02, KJZ^{+}12, LSS^{+}09, RLR^{+}18, WSH15]. Different [KCL^{+}12, VPM^{+}19, AMEV07, Bak01, BB24, BBL^{+}18, CSMGS19, CGC^{+}20, CCB^{+}20, DDDT99, HLK13, HSK^{+}19, HMH^{+}15, IPG^{+}16, KKO10, KGJ^{+}10, MGH^{+}07, PPKR14, Sie69, WZBK^{+}21, XHC^{+}20, YYK^{+}12, YGMR^{+}23]. Differential [Gar03, Due77]. differentiation [WLP^{+}21]. diffraction [LGL^{+}18]. diffusion [Gar03, RG03a, Wen88]. Diffusive [KFC^{+}23, Fla02, KFG^{+}03, McD81a, McD81b, ÖÜT93, PSGVS^{+}14, WZFW16, YYK88, YN03a]. diffusively [TG81]. digestive [RMB^{+}01]. dimensional [BBPHG^{+}11, BASS^{+}20, Coo69, Dav85, DJ92, GWB14, Kaw86, KDF97, LHC^{+}19, MR03, MAB^{+}11c, MAB^{+}11a, MAB^{+}11b, SGMVF14, WB03, XD95, YHZ^{+}22, YWUK15, YFY05]. dimethyl [ZPY^{+}20]. dimethylated [GLY23]. dimethylsulfide [FLDF22, ZBY^{+}22]. Dimethylsulphide [FPS^{+}09]. Dinoflagellate [BPTT19, ST10, KPM^{+}23, MPMA13, OMS^{+}09, RSM^{+}23, SPF^{+}23, Sma10a, Sma10b, VMV^{+}23, ZHBW01]. dinoflagellates [CM14a, PMA^{+}14]. Dinophysis [DRVMC^{+}22, PMMN^{+}22]. dioxide [FLUC08, KBHML17, XCH^{+}16]. dioxins [CJRÁ^{+}13]. dipole [SW22, RHB23]. Direct [ITO^{+}14, JSKM02, LGK^{+}93, Web69]. directions [MPC12, PV18]. disadvantages [KSY^{+}19]. discharge [AJHC19, CKB^{+}17, CTI^{+}19, FMWW14, GdRGC^{+}14, WBH15]. discharges [LDMH09]. discovered [GKS^{+}13]. Discovery [WTT14]. discrepancy [HL05]. discriminate [LSD^{+}15]. discriminated [BCLD^{+}17]. Discrimination [YGL^{+}10]. discussion [MSFZ19]. Disentangling [HGB^{+}21, HLCdP19]. disequilibrium [AYK^{+}05]. disjunct [NHG19]. Disko [HJLLN07]. disparities [HS07]. Dispersal [DTG^{+}24, HPW10, JGO^{+}98, PHFK14, PPdS21, ALT10, CZG^{+}21, CZL^{+}24, DCM16, GKC^{+}14, GGJ^{+}10, GHC^{+}17, JLRB20, JLP^{+}20a, JLP^{+}20b, LAHI10, LYZ16, LAGM^{+}23, MKSvA^{+}22, OOTA15, QCdS^{+}07, SNV^{+}18, TTMM^{+}17, TKC^{+}22, TMR^{+}21, VMH^{+}21, VOJD02b, ZLRVB24a, ZLRVB24b]. Dispersion [ZLG17b, ARDP14, Gar06, IAN13, MMIB10, OC06, San15, SVIA14, ZLG17a]. Dispersion-Confinement [ZLG17b, ZLG17a]. displacement [SDGVE17]. displacements [LO21]. Dissecting [XHW^{+}20]. dissipation [Car97a, Kat18, LL97]. dissolution [GTR01, TGR05]. Dissolved [DHDM22, MPCNC^{+}19, VFCC^{+}22, BFJ18, CDS90, CLL^{+}18, DGH^{+}20, FZY^{+}23, FTG^{+}18, FGL^{+}23, GLAHH^{+}22, GF19, HOY^{+}21a, JZZY24, JMG^{+}13, KFC^{+}23, MNFY21, OB98, RFSCF19, SHd13, SWT^{+}17, SIS^{+}14, TBW00, WLM^{+}13, XYGJ23, YT06, YTB^{+}21, ZNR^{+}24, ZMW^{+}23, ZKK^{+}16, ZSY^{+}22, CMPNC^{+}22]. distance [JLP^{+}20a, JLP^{+}20b]. distinct [CdD^{+}15, DBRK17, PMM^{+}23]. Distinctive [AJA^{+}22, MCGS^{+}16]. distinguished [YRKC08]. Distinguishing [Pie01]. distributed [THM^{+}14]. Distribution [AYH^{+}23, AE09, CF20, EBM^{+}21, ECFT20, GMAMB04, KT04, KMF^{+}20a, KMF^{+}20b, LDAM^{+}07, LRAE23, LM00, LCR^{+}93, lLdZQ^{+}22, MSA^{+}22, MPTMK22, NMLBCM^{+}01, PL87, PBD^{+}88, SPF^{+}23, SGL^{+}13, SGA^{+}19, STS^{+}12, SBS90, TZP^{+}00, YCP^{+}12, YAI^{+}14, AHSS22, AEP^{+}23, AGD^{+}18, AOMZ^{+}23, AHRT90, BM76, Ban64, BSF95, BLI^{+}99, BCLD^{+}17, BHB^{+}19, BDL08, BGB^{+}08, BECA22, BBFS19, BHC^{+}18, BJ90, BSH^{+}20, BTV^{+}17, CAH^{+}22, CSR90, CGV13a, CB06, CMM^{+}04, CTF07, CMPNC^{+}22, CLB^{+}13, CLG^{+}00, CS06, CFG07, Cra09, CNSHT15, CJG88, DDE^{+}95, DGP^{+}13, DLL^{+}23, DGGdR02, DAU22, DCRR^{+}22, DDCE^{+}23, DAIS10, DIQJ21, EAB^{+}23, EKB06, EHG^{+}12, dCFK17, FTG^{+}18, FTHK19, GCLD19, GHF^{+}21, GPP22, GGJ^{+}10, GSA^{+}20, GBT^{+}19, GBB^{+}20, GGA^{+}16, GMDS20, GAS^{+}22, HGD22, HVTV22, HFNG00, HKPV12, HOY^{+}21a, Hob10, HF10, HCGK11, IMM^{+}22, IHY^{+}01, JS90]. distribution [JC04, JMZ23, JSHB90, JAJS08, JHW^{+}14, Kam19, KSK^{+}15, KFM^{+}17, KGdS^{+}08, KVNT20, KSKN21, KDB95, KHP^{+}18, KFH^{+}15, LJ65, LF12, LOBG^{+}10, LLGS21, LRGV^{+}18, LPARF^{+}20, LSIB23, MCB^{+}90, MSC^{+}15, MGWZ20, MBdM^{+}18, MPCNC^{+}19, McG64, MS17, MPSD15, MBB^{+}96, MFA^{+}15, MWFH02, MGH^{+}07, MKMF^{+}89, NHH^{+}23, NCC^{+}15, NMN08, OSH^{+}96, ORMR^{+}19, PSL87, PPSV^{+}18, PJS^{+}22, PPVG12, PRC^{+}20, PPSVC^{+}13, PWZ^{+}16, PK02, PDAM^{+}15, PCC^{+}19, PLHLF05, RLDC^{+}13, RCB^{+}20, RAP95, RMB^{+}01, RG94, SOS^{+}07, SCHS^{+}24, SSB20a, SIR^{+}07, SM21, SBK^{+}95, Sie69, SAY^{+}16, SCC^{+}19, SSV^{+}11, SDL^{+}19, TNY^{+}24, TPRS10, TBW00, TIOM16, Ver91, WMB^{+}21, WCX^{+}21, WSS15, WQ08, XYL^{+}22, XNT^{+}17, YTB^{+}21, YSY^{+}19, dFHR^{+}24]. Distributions [HHY03, SIS^{+}14, Ang79b, Ang84, ATC^{+}19, ABP15, BSW86, BMG^{+}19, Bri79, BASS^{+}20, CDH^{+}13, CIL^{+}23, CGM^{+}02, CBOP15, CCM^{+}14, CW02, Dom84, DHL^{+}21, DTKvH15, FCN^{+}19, FW91, GBC^{+}00, GMR^{+}23, Hau84, HSLG11, KCPM09, KTIT22, Lev88, MHCR^{+}12, OYKK^{+}23, OCH^{+}18, Par65, PBP^{+}99, Pug84, RB20, RAB^{+}84, Roe84a, Roe84b, RB84, RJT84, SMN^{+}14, SWT^{+}17, SK18, SBE^{+}20, SW21, TCDPP^{+}23, WGG^{+}08, WL16, Yao88, ZHBW01]. disturbance [GWK17, RLR^{+}18, dJSL^{+}20]. disturbances [His22, LL21]. diurnal [CIL^{+}23, HYM^{+}24, PVV23, Rou65, SNS^{+}22, vHMDL14]. diurne [Rou65]. dive [HHP10]. Divergent [FMCG15]. Diverging [HSG^{+}15]. diverse [HGB^{+}21, IG19, MRBS^{+}24, OSH^{+}96, RSW^{+}23, SKH^{+}23, SPH^{+}15a]. diversification [GWGR^{+}19, HS22]. diversities [RTBR^{+}22]. Diversity [BBFS19, BGWP^{+}17, BTV^{+}17, DGVGR24, DCRR^{+}22, EBW^{+}23, GGA^{+}16, Mar20, PJS^{+}22, PAPL15, ABE^{+}15, AS96, ÁLC22, BMdMS^{+}21, BWB^{+}09, CMF^{+}09, CTR^{+}19, DDK^{+}18, DNNNN16, DBJ^{+}15, DBR20, FTC^{+}16, FBT^{+}22, GCC^{+}24, GFB^{+}15b, GFB^{+}15a, GLV12, GRdSS^{+}22, HEF^{+}12, HTG15, JPB20, KSVT00, KKKS14, LWBD^{+}17, MGWZ20, MJD^{+}21, NGNV12, OVG16, PTPY^{+}23, PPD^{+}12, QOS^{+}22, SBS90, VWDF14, WPBG^{+}18, ZWM^{+}15]. Dividing [RSB^{+}13]. diving [FDB^{+}21, STEB16]. DMS [MY23]. DMSP [CM14a, FPS^{+}09, MY23]. DMSP-lyase [FPS^{+}09]. DNA [CBL^{+}19, JM19, RSM^{+}23]. Do [HLD^{+}21, KCPM09, PCR^{+}22, TKC^{+}22, NGLSSG14, Ric22, RMB^{+}01, RSM^{+}23, SMKK21, Zez90]. DOC [OB98, SHd13]. documented [SBMB18]. Does [LPHL^{+}05b, SCC^{+}19, WBC^{+}22, XRC^{+}15, ALT10, CPC88, Dag93, FMWW14, GCCY^{+}14, IL20, MAH^{+}15, SGLF^{+}13, UAM05]. Dohrn [GCLD19]. doliolids [IMM^{+}22]. dolphin [SGL^{+}18]. dolphins [LPF^{+}18]. DOM [CMPNC^{+}22, CDB^{+}22, CMPNC^{+}22, MPCNC^{+}19, MZ14]. domain [TSJ^{+}12, CWW15]. domains [CLV^{+}19, MWFH02, WKS^{+}15]. dome [SDS^{+}22b]. Dominance [JJA^{+}17, ERT^{+}22, PFW15]. Dominant [HDZY15, AVK91, INI^{+}17, MGH^{+}07, SLBVRR^{+}22]. dominate [CBT07]. dominated [BTNK13, BRE^{+}24, TPPG10, Whe06, XLX^{+}20, ZDG^{+}21]. Dongsha [HWF^{+}21]. Dosidicus [ATT^{+}08, BGM^{+}10, RS10]. Double [McD81a, McD81b, PSGVS^{+}14, KFG^{+}03, ÖÜT93, RG03a, RG03b, TG81, YYK88, YN03a]. double-diffusion [RG03a]. Double-diffusive [McD81a, McD81b, PSGVS^{+}14, KFG^{+}03, YN03a]. double-diffusively [TG81]. double-infusion [RG03b]. Douro [MSd^{+}16]. down [AH10, AF10, BVJE19, CGM^{+}02, FZY^{+}23, FFA06, HS22, HM06, MLPN06, OWR^{+}07, PPD^{+}12, RGMPR23, SCHS^{+}24, SIB^{+}06, YNMY23, ZK06]. downcore [SLG^{+}12]. downscaling [LPF23, SCLS10]. downstream [TOiF^{+}12, WWL^{+}22]. Downward [SCT^{+}00, BS02, FUOG^{+}16]. downwelling [ABS^{+}20, BLT^{+}15, LPF^{+}21]. Dr. [Ano20u]. Dragon [GKS^{+}13]. Drake [ARELAK24, CLV^{+}19, CP07, GWGR^{+}19, Spr08, VKDS^{+}18]. Drastic [KOhL^{+}10]. drawdown [TRP^{+}23, TAM^{+}15]. drift [Coo65, HKPV12, HMA18, ICB^{+}19, MK86, SCHD23, SOO^{+}14, VBL^{+}09, YSS14]. drifter [LKDL14, LC16, TM13, TFZS14]. drifters [AKAL20, BBM^{+}14, GL06, ICB^{+}19]. Drifting [Ric85]. driftnet [INI^{+}17]. drive [BMG^{+}21a, CCB^{+}20, GSA^{+}20, GRD^{+}23, LBC^{+}23, RAB^{+}11, WCS^{+}23, YFY^{+}22]. Driven [SBPGP^{+}23, BBS21, BBS23, CLY22, CBM^{+}21, FMP19, GW91, KQP^{+}17, MHGGS19, NNM^{+}21, OIC^{+}23, OÁSG^{+}16, PRTC13, RB20, SDGVE17, Sek86, SC23, SDO^{+}14, TAW^{+}15, TG81, WHBK05, WHBW03, XWL^{+}22, vHCY^{+}20]. driver [MTK^{+}22]. Drivers [BTS^{+}15a, CM18a, GHL15, IMM^{+}22, OVG16, OHH^{+}22, dFHR^{+}24, BBL^{+}18, CLD22, CMPNC^{+}22, EBR^{+}14, FJA^{+}21, GBT^{+}19, HSGJ23, HVTV22, HLD^{+}21, HPW10, KPSB22, KHP^{+}18, LJM^{+}16, LM18, LFCSV^{+}13, LDMH09, MPCNC^{+}19, MCKS17, PVM^{+}20, PGY^{+}22, PSA^{+}19, RCSVGP^{+}16, SHP^{+}23, SBL^{+}23, SLH^{+}19, TAM^{+}15, VSGD21, VDB^{+}20, YPVP^{+}22]. drives [APC^{+}12, SOO^{+}14, ZWM^{+}15]. driving [NRA17, Woo18]. drop [CMF11]. droughts [Has06]. dry [Ken88, KAK^{+}22b]. dryer [JTQ^{+}18]. dual [LH08, TSJ^{+}12, GDN^{+}18]. dual-basin [LH08]. dual-domain [TSJ^{+}12]. Dual-frequency [GDN^{+}18]. due [DLD^{+}19, IST^{+}88, KLC^{+}15, LRJ^{+}15, SAY^{+}16, YHRT22]. Dungeness [BL02]. duration [FP03, GGJ^{+}10, STHM02, WO15]. during [AMFY20, AAML22, ABD^{+}17, ABS^{+}20, AYK^{+}05, AVK91, AVG^{+}19, AKH^{+}23, ABP15, BC91, BGM^{+}01, BIST01, BCM^{+}02, BRD^{+}15, BPC^{+}05, BM07, BASS^{+}20, CAH^{+}22, CCW^{+}02, CTF07, CB91, CJMI^{+}91, CFC^{+}18, CBD^{+}24, DW02, DRVMC^{+}22, DHDM22, EKB06, ESGP17, EMU^{+}23, FB01, FGGDF^{+}04, FWBC02, GWB14, GFGGD^{+}23, GMDD^{+}22b, GCZ^{+}00, GS24, GIHJ23, GA10, GMAB07, GRS08, GAM98a, GAM98b, HBK^{+}24, HMB^{+}86, HGPFN^{+}14, HFW^{+}98, HCC02, HHSR07, HBH^{+}17, HCGK11, HSF02, INO^{+}24, JJA^{+}17, JS87, JC88, JSKM02, JMZ23, JST^{+}24, JW01a, KKB00, KIS^{+}05, KBHML17, Kos02, KC02, KNI^{+}05, KYS^{+}17, LBH^{+}87, LPHL^{+}05b, LPHL^{+}05a, LGR^{+}02, LBH^{+}21, LC22, LXC^{+}22, LOO22, LCGH07, LWBD^{+}17, LDMH09, MSMR93, MG02, MRMD^{+}97, MMR^{+}12, MST^{+}23a, MHA^{+}11, MPD^{+}22, McG64, MBB^{+}96, MWS^{+}10, Mil93a, MSA^{+}22, MIW91, MSS^{+}02, MMF^{+}12, MFM85]. during [MPTMK22, MGH^{+}07, MHCR^{+}12, MA12, MEMC05, MJA^{+}07, NSE^{+}24, NBR^{+}08, NKK^{+}05, NST^{+}23, ORW^{+}01, OELP04, OEL^{+}14, PBB^{+}20, PVG^{+}20, PPdM^{+}12, PTP^{+}22, PELAA18, PAM^{+}88, PK02, PHLL05, PWMIM91, PVV23, PLN^{+}23, PNF^{+}21, RAP95, RVS^{+}21, Sai65, SLG^{+}12, SLBVRR^{+}22, SMdG02, SGMP15, SMP^{+}22a, SWT^{+}17, SZG06, SKF20, SEG^{+}22b, SEG22a, SW01, SPH83, SMGL01, STW^{+}15, STR01, SYN^{+}21, SDL^{+}19, TMN^{+}12, TII^{+}14, TSAM^{+}22, TFM03, TSFA22, VSGC21, VHV^{+}12, VDGGD^{+}22, Ver91, WFH^{+}22, WP91, WLM^{+}13, WLM07, WDMC02, WHK23, YMA^{+}17, YNMY23, YLL19, ZBY^{+}22, ZLC^{+}15, ZDG^{+}21, dPAJ07, dFKdLZTT17]. dusky [LPF^{+}18]. dust [LGG18, Qiu15, RDD^{+}18]. dust-derived [RDD^{+}18]. DWBC [FKZ^{+}15]. dwelling [KFC^{+}13]. DYFAMED [CLL^{+}18, HLM^{+}13, MMG^{+}11]. Dynamic [LB02, TAW^{+}15, TJ90, AMG^{+}16, GCCY^{+}14, KHS^{+}14, Kaw86, Kit03, KMU^{+}12, Leh01, MB01, PHKS01, PBN13, RSMIS03, Ric15, RG94, Sac16, SWH^{+}24, TSH^{+}17, VDS^{+}18, XDG^{+}23, YAK^{+}08]. Dynamical [RHB23, Tho95, HKL^{+}15, PC87]. Dynamics [AYK^{+}05, Car98, DTOD00, DGH^{+}20, KIS^{+}05, Kir06, LSM^{+}22, LSB^{+}17, LSD^{+}18, LZG20, MFS^{+}16a, MFS^{+}16b, MYH^{+}22, MMG^{+}11, SFMT12, SFMT14, SBPGP^{+}23, SNdSR^{+}24, WHIH97, AEP^{+}23, AHP19, AS88, AOMZ^{+}23, Ano94c, AG22, ALM^{+}23, AKH^{+}23, BDP^{+}06, Bak06, BBB^{+}14, BFJ18, BHLU^{+}07, BRR^{+}12, Bri83, BB10, CRGA17, CGL^{+}20, CKB^{+}17, CML^{+}16, CWZ^{+}20, CGMP14, CSV^{+}07, CMS^{+}13, CSS^{+}21, CÁM06, CMHM18, CLG^{+}00, CS16, CCD^{+}13, CRiI^{+}15a, DAvD^{+}20, DAvD^{+}21, DCD^{+}23, DXH^{+}02, DBW^{+}22, DLL^{+}23, DZ04, DL17, DBRK17, DLJ^{+}21, DSV^{+}24, DCL^{+}13b, DK07, EMU^{+}23, FBS^{+}18, FFA06, FK99, Gam14, GCB^{+}22, GCV^{+}24, GCD^{+}13, GBC^{+}16, GRD^{+}23, GCFS06, GC14, GZCL23, HPS^{+}01, HPC^{+}20, HSK^{+}19, HMP^{+}13, HvDL^{+}17, HGH^{+}19, IVR^{+}13, JB15, JvdLL^{+}15, JHM^{+}22, JMG^{+}13, JCF^{+}23, KKS^{+}18, KBC^{+}22, KHC^{+}99, KHD22, KMB01, KAK^{+}22a, KMS^{+}24, KSKN21, KS15]. dynamics [KNI^{+}05, LGK^{+}93, LSM08, LMS10, LG23, LNB13, LFG10, LZCZ05, LGH^{+}21, LS12, LS15, LGD^{+}20, MHGGS19, MGA^{+}23, Mau10, MY92, MKM93, MFDH22, MS00, MMD^{+}16, NMK^{+}03, NKK03, NHSP23, NHN^{+}21, OVR^{+}02, PVC^{+}08, PMMN^{+}22, Peñ03b, PV07, PMC16, PVB23, PST^{+}02, PWMIM91, PCH^{+}08b, PDD^{+}22, Pra91, REG^{+}15, RTF^{+}05, RHB23, RGPB^{+}23, RKC^{+}10, SBMB18, ST03, SHd13, SDK84, SPH^{+}15a, SDH^{+}14, SPN98, SPK^{+}22, SLH^{+}19, SGR^{+}22, TTMM^{+}17, TTB^{+}08b, TG81, TFZS14, VFCC^{+}22, VGM^{+}23, VHV^{+}12, VSC01, VOJD02a, VOJD02b, WCX^{+}21, WZBK^{+}21, YGC^{+}21, Yos80, Yu23, ZLG17a, ZLG17b, ZDG^{+}21, ZPC^{+}16, ZFSV^{+}09, vWM02a]. dynamique [Ber65c]. DYNAMO [Ano01a, MLW01]. Dysfunctionality [AM10, Fly10].

E. [GBG05]. each [CBL^{+}19]. eAP [WCC^{+}20]. Earlier [KGB^{+}23]. Early [BSF^{+}21, FMT15, NNFL21, PSK96, AHP19, AHC^{+}13, BLT^{+}08, CCM^{+}14, DML^{+}16, Dea85, EKB06, FVLC^{+}23, FCN^{+}19, FMW91, GPEV20, GKS^{+}13, GDM^{+}15, GiIKX22, HBG^{+}21, HPW10, IMHL07, KSK^{+}15, KBF^{+}08, KLP^{+}17, LSY^{+}14, LBC^{+}23, MVBC^{+}21, NHG19, PDAM^{+}15, RGC^{+}01, SLY^{+}15, Tho87, TMR^{+}21, TRP^{+}23, VMH^{+}21, WZ04, YJ88, dFKdLZTT17]. earth [Ång65, GSPP^{+}20, CR97, GCD97, Kag97, Bak83]. earthquake [CLB^{+}14]. earthquake-induced [CLB^{+}14]. East [Ang84, AYH^{+}23, ALT10, CSC^{+}12, Dom84, HCV^{+}20, HWS^{+}07, LC16, LCJ^{+}07, LXC^{+}22, NGNV12, ORMR^{+}19, Pug84, RAB^{+}84, Roe84a, Roe84b, RB84, RJT84, TTL^{+}04, VHV^{+}12, ABSDC07, Bum73, DMBHG10, JJS03, KMOM88, Lie88, Man69, MKM86, MM90, STJ^{+}14, TMH^{+}16, Yin88, ZSH^{+}24, CPG^{+}18, CTL^{+}04, CS03, Cho86, DL17, DCS^{+}22, GCCY^{+}14, HMRA^{+}03, HOY^{+}21b, hHCK01, HSH97, ITM86, JJR^{+}08, JS87, JJL^{+}19, JZZY24, JAJS08, Jón07, JJ08, KKK04a, KKK^{+}04b, KPM^{+}23, KHM^{+}88, LSF^{+}17, LGH^{+}21, MLL^{+}22, MY23, MH14, MKS^{+}22, NC80, NBR^{+}08, NHN^{+}21, ORCH^{+}19, PMG15, Pai20, PO00, PL18, RD03, RI86, SN24, SW12, Suk88, SP08, TM13, TKW06, TKWI08, WGZZ19, YIY^{+}04, YKS^{+}12, YMI88, YKH^{+}24, YJS86, YJ88, YPM^{+}10, Yux88, ZKT88, ZLS^{+}04, ZLR^{+}07, ZGZ19, ZSY^{+}22, ZZWL06, vHVAT22]. East-Pacific [vHVAT22]. East/Japan [CTL^{+}04, JJL^{+}19, KKK04a, YPM^{+}10]. Eastern [KHC^{+}99, KTB^{+}99, AMY^{+}23, AJA^{+}22, AALM06, ADS^{+}22, BM76, BPF06, BTK^{+}99, BE99, BCLD^{+}17, Ber65b, BHS^{+}15, BHC^{+}18, Bri79, BDC^{+}08, CARBML^{+}22, CGV13b, CC88, CS16, CMF15, CHS^{+}24, DDDT99, DWS^{+}24, DVL^{+}99, DBC^{+}23, DMT15, DGVGR24, DTG^{+}24, ECGP01, EBvdL^{+}09, FÁFL06, FCMCÁS19, FT06, FL06, FMP19, FK99, GPP22, GTS^{+}21, HSS^{+}12, HMTL05, HPB^{+}09, HFNG00, HLP^{+}16, HHMB^{+}09, HFO90, HVEF09, ISM^{+}02, KKB00, KSV08, Kes06, KJH^{+}22, KH09, KFH^{+}15, LBNBM13, LRNK99, LFA^{+}06, LKDL14, LW85, LSW^{+}21, Luk86, MT99, MGG22, MST^{+}23a, MLK^{+}09, MC15, MNM06, MWFH02, MWO^{+}12, NCH^{+}07, OLH^{+}18, OPG^{+}10, PBB^{+}20, PMK^{+}06, PLEF^{+}23, PDAM^{+}15, PLB^{+}23, PP10, PST^{+}15, RMG90, RKCH15, RGPB^{+}23, RR01, SCLG^{+}11, SCHBC^{+}22, SRAV19, SGS^{+}23, SH09, STHM02, SSN23, SLPA^{+}20, SNMW10, SGR^{+}22, TJ73, Tan99]. eastern [TMKJ^{+}09, TPRS10, THM^{+}06, VOT^{+}99, WF06, WF07, WSL20, WCC^{+}20, WFR07, WHI^{+}02, WLL06, WTH12, XWW^{+}21, ZK06, ZWP23, ZLZ^{+}17, vAB96, Ano09h, BMC^{+}10, CM09, Cia14, DDE^{+}95, DRE^{+}08, DGMM85, EMU21, FGSA97, FBA09, FAB^{+}09, GBC^{+}16, JHM^{+}22, KKS^{+}18, Kat18, KMU^{+}12, KEV10, LBH^{+}87, MLPN06, NDEG22, PJS^{+}22, Phl65, PGC^{+}96, PLN^{+}23, RKM^{+}07, SPH^{+}15a, TPN^{+}18, TVD^{+}99, VKT15, VOT^{+}99, WFBN^{+}13]. eastern-Mediterranean [DDDT99]. eastward [BPSGP^{+}23, FGR^{+}06]. eaux [Rot65]. Echinoderes [AM19]. Echinodermata [MMK19]. Echinoderms [MDG^{+}19, PJS^{+}22]. echinoids [SM16]. Echiura [GS19, MA20]. echiurans [MA20]. echo [ZZPL18, ZYN^{+}24]. echosounder [ON22]. eco [Lav09]. eco-regions [Lav09]. Eco3M [BFPS06]. ecohydrographic [KAAK^{+}16]. Ecological [CQO^{+}15, CLG^{+}22, EAL^{+}07, Fly10, GVBV^{+}21, GHC^{+}17, GRB^{+}08, MP04, SEO13, BRG^{+}15, BMG13, BEH19, CAO^{+}20, CRS04, DFM^{+}15, DFH^{+}16, FVA^{+}19, GMD^{+}22, GA01, HSMLDC^{+}22, HFS^{+}20, HBW17, HLD^{+}21, Igu04, JHDT12, LAGM^{+}23, LM14, LSH^{+}22, MDGC^{+}12, MGC^{+}18, MIH06, MR03, MCH^{+}12, OOTA15, POS^{+}07, PRC^{+}20, PBO10, PM13, PYKF15, RSK^{+}23, SHD^{+}21, TMÁGC^{+}21, VPM^{+}19]. ecologically [FACM^{+}23, SFS^{+}12]. Ecology [McK04, SRK15, Ang89, BFB^{+}20, BLI^{+}99, BRH^{+}05, CCG07, CDL19, CH07b, CSM^{+}15, CLCBB19, DM13, DKRL22, EBM^{+}20, EBM^{+}21, FDB^{+}21, GD85, HS22, HMP^{+}13, HWBT03, JSLA^{+}21, KHS^{+}14, MM80, MPC^{+}17, MHVS19, PGLG^{+}05, PHCA17, SFS^{+}12, THM^{+}14, YGMR^{+}23, dB94, WHG^{+}16]. Economic [McI10, GRDS10, JHDT12, MBF^{+}14]. ecophysiological [THBA19]. ecophysiology [PBB^{+}12a, PBB^{+}12b, PRTC13]. ecoregion [BRR^{+}22]. ECOSystem [Mau10, Ant09, AYK^{+}05, BDG^{+}17, BTS22, FK99, GCFS06, GBC^{+}15, Hea12, JSA^{+}08, JBB^{+}14, KNI^{+}05, LL21, MP13, MOSN^{+}13, ML09, MBD^{+}09, PHD^{+}18, PCH^{+}08b, PL09, SCD^{+}07, SDP^{+}22, SKGS20, YFY05, ABE^{+}15, AN04, AF10, AM10, AGS10, AHC^{+}13, ABÁS^{+}09, BCB^{+}05, Bak06, BW08, BLR^{+}23a, BMB^{+}16, BCM^{+}02, BRD^{+}15, BMM01, BDE03, BHMS09, CARBML^{+}22, CSS^{+}21, CALS^{+}23, CSC^{+}12, CHB02, CAB^{+}99, CRiI^{+}15a, DKRL22, DLL^{+}23, DSR21, DPH^{+}18, DPR^{+}18, DWC06, EGP^{+}18, FFA06, FACM^{+}23, FLDF22, FPS^{+}13, GMBU12, GLF^{+}17, Gir15, GRS08, GCP08, GEP^{+}08, HVS10, HFS^{+}20, HSG^{+}15, HGB^{+}21, HSC09, HBD^{+}18, HvdLS^{+}09, IMM^{+}22, JPM^{+}08, JHDT12, JLS^{+}22, JPBB20, KKB00, KHL12, KTH^{+}21, KIS^{+}05, LR07, LOG^{+}09, LLL^{+}11, LCBN14, LML^{+}23, LDH90, Leh01, LSM08, LSXT01, LFG10, LIH^{+}12, LHW^{+}20, LCANAS^{+}07, LDMH09, MLL^{+}22, MRM^{+}14]. ecosystem [MRA^{+}19, MRSS02, MGS90, MV10, MKHO96, MBH^{+}01, MSI17, MHS^{+}09, MLM09, MLHM09, MS00, MSB^{+}23, NYL^{+}17, ORVES17, OAWAN18, PL01, Peñ03b, PBH^{+}10, PBN13, PPD^{+}12, QCdS^{+}07, REG^{+}15, RCS^{+}11, RBD^{+}07, RLX^{+}24, RGMPR23, RRLS22, RN06, SGWF^{+}19, SSL08, SJP10, SBLA10, SFF^{+}24, SRM^{+}10, SCB^{+}07, STGR^{+}23, TTB^{+}08b, TCL20, TKWI08, Tit20, TFM03, TS10, Tur99, VBL04, WSS15, WFJ^{+}15, WPH^{+}10, WPB05, WJPHB15, YFK21, YZX^{+}23, YWUK15, ZLKO00, ZLS^{+}04, ZWM^{+}15, dLLdAWL^{+}23, dMGS^{+}11b, dMGS^{+}11a, dPCS23, TTB^{+}08a, TTB^{+}08b]. ecosystem-based [HFS^{+}20, JHDT12, WFJ^{+}15]. ecosystem-level [PBN13]. ecosystem-linked [SSL08]. ecosystemic [YPGE^{+}10]. Ecosystems [Ano09h, BCS09, Bel09, CM09, FAB^{+}09, IGG^{+}19, KSE^{+}09, BDP^{+}06, BH07, BWB^{+}09, BHMS09, CSBL^{+}15, CCB^{+}20, CS04, DM13, Den03, DMF^{+}09, Dri11, DP13, DAKV99, GSFP^{+}09, GAF15, GML^{+}23, GLH13, HAA^{+}14, HM06, HPH^{+}08, HDA^{+}16, HBD^{+}18, JAS^{+}20, KCPM09, KGL22, Kno04, LDD^{+}22, LSS^{+}09, Man04, MBCB88, MFS^{+}07, MSF^{+}07, Mau17, MLK^{+}09, MS15, MPN09, MBD^{+}09, PS08, QSC^{+}15, RK03a, RKK^{+}21, Sie88, SPV^{+}15, Tan99, TCS15, TMKJ^{+}09, TAM^{+}13, TSJC07, Was11, Was15, Woo05, YBPS08, dlHRA^{+}18, FBA09]. Ectinosomatidae [GS19]. Eddies [PLEF^{+}23, RFKC16, WLL06, BRB^{+}01, BA04, BNCC15, BF12, BH85, CGG08, CBB^{+}22c, CMG15, CSS11, CBT07, DMBHG10, GLY23, HMX^{+}23, KM10, Kli10, KBSB18, LLS01, MXC^{+}21, MNS^{+}24, MTL05, MSS^{+}02, MHCR^{+}12, NC80, OKdA^{+}19, Ric93, TOiF^{+}12, WOW^{+}14, XWL^{+}22, YHLA^{+}04, ZSH^{+}24, ZL24, HMS^{+}22]. Eddy [CED09, RL85, SOB^{+}08, ZYN^{+}24, BATNP04, BBS21, BBS23, BRR^{+}12, BKB85, CDH^{+}13, CCRS20, CGG08, CLY22, CCD^{+}13, DCD^{+}23, GMAMB04, Har05a, Hau84, HZCZ16, JLB^{+}08, KZSH85, KDL^{+}01, KSD84, LRS^{+}03, LBD11, MZK^{+}23, ORCH^{+}19, OP18, Owe91, QYF^{+}24, RLL^{+}09, RAE^{+}05, RBS^{+}09, SC23, SDK84, Sim84, SBD01, TG05, VBM21, WBA^{+}22, WLM^{+}13, WBB^{+}01, XHC^{+}20, ZZM^{+}24]. eddy-driven [CLY22]. eddy-Ekman [HZCZ16]. eddy-induced [HZCZ16]. eddy-permitting [KDL^{+}01, WBB^{+}01]. Eddy-resolving [SOB^{+}08]. edge [BHK^{+}16, BH85, GGG^{+}18, Hut81, Hut95, HHK^{+}02, KHM^{+}88, LMP22, LH08, ZCA21, ZAC^{+}23]. edifice [GSA^{+}20]. Edited [Ang79a, Ang80, Ang88, SW81]. Edition [Bak83]. Editor [Swa76]. Editorial [AO79, AO84, Ang85, Ang87, AS89, Ang94, Ano87d, Ano99a, Ano17e, Ano17f, Ano17g, Ano17h, Ano18a, CD87, Har05b, OA81, SA90, SA86a, SA86b, SB69, Swa77, War73, Ano63a, Ano64b, Ano73c, Ano76a, Ano79a, Ano80b, Ano81a, Ano82b, Ano83a, Ano84a, Ano85d, Ano85e, Ano86b, Ano86c, Ano87b, Ano87c, Ano88a, Ano88b, Ano89b, Ano89c, Ano90a, Ano90b, Ano91a, Ano91b, Ano91c, Ano92b, Ano92c, Ano93a, Ano93b, Ano94a, Ano94b, Ano95a, Ano95b, Ano96a, Ano96b, Ano97a, Ano97b, Ano07a, Ano07b, Ano07c, Ano07d, Ano07e, Ano07f, Ano07g, Ano07h, Ano07i, Ano07j, Ano08a, Ano08b, Ano08c, Ano08d, Ano08e, Ano08f, Ano08g, Ano08h, Ano08i, Ano08j, Ano08k, Ano08l, Ano08m, Ano09a, Ano09b, Ano09c, Ano09d, Ano09e, Ano09f, Ano09g, Ano10c, Ano10d]. Editorial [Ano10e, Ano10f, Ano10g, Ano10h, Ano10i, Ano11a, Ano11b, Ano11c, Ano11d, Ano11e, Ano11f, Ano12a, Ano12b, Ano12c, Ano12d, Ano12e, Ano12f, Ano12g, Ano12h, Ano12i, Ano12j, Ano13a, Ano13b, Ano13c, Ano13d, Ano13e, Ano13f, Ano17i, Ano17j, Ano17k, Ano18b, Ano18c, Ano18d, Ano18e, Ano18f, Ano18g, Ano18h, Ano18i, Ano19b, Ano19c, Ano19d, Ano19e, Ano19f, Ano19g, Ano19h, Ano19i, Ano20c, Ano20d, Ano20e, Ano20f, Ano20g, Ano20h, Ano20i, Ano20j, Ano20k, Ano20l, Ano21d, Ano21e, Ano21f, Ano21g, Ano21h, Ano21i, Ano21j, Ano21k, Ano21l, Ano21m, Ano22d, Ano22e, Ano22f, Ano22g, Ano22h, Ano22i, Ano22j, Ano22k, Ano22l, Ano22m, Ano22n, Ano22o, Ano22p, Ano22q, Ano22r, Ano22s, Ano22t, Ano23d, Ano23e, Ano23f, Ano23g]. Editorial [Ano23h, Ano23i, Ano23j, Ano23k, Ano23l, Ano23m, Ano24b, Ano24c, Ano24d, Ano24e, Ano24f, Ano24g]. Editors [WR03]. EDM [HSGJ23]. eDNA [WPA^{+}24]. Eds [Ano13h, Ano13i, Ano13j, Ano13k, Ano14a, Ano14b, Ano14c, Ano14d, Ano14e, Ano14f, Ano14g, Ano14h, Ano14i, Ano14j, Ano14k, Ano15a, Ano15b, Ano15c, Ano15d, Ano15e, Ano15f, Ano15g, Ano15h, Ano15i, Ano15j, Ano15k, Ano15l, Ano16a, Ano16b, Ano16c, Ano16d, Ano16e, Ano16f, Ano16g, Ano16h, Ano16i, Ano16j, Ano17b, Ano17c, Ano17d]. Eds/publication [Ano13h, Ano13i, Ano13j, Ano13k, Ano14a, Ano14b, Ano14c, Ano14d, Ano14e, Ano14f, Ano14g, Ano14h, Ano14i, Ano14j, Ano14k, Ano15a, Ano15b, Ano15c, Ano15d, Ano15e, Ano15f, Ano15g, Ano15h, Ano15i, Ano15j, Ano15k, Ano15l, Ano16a, Ano16b, Ano16c, Ano16d, Ano16e, Ano16f, Ano16g, Ano16h, Ano16i, Ano16j, Ano17b, Ano17c, Ano17d]. Edward [Mer65]. eel [GEPC15, PFE10]. eelpouts [SM21]. Effect [BD20, CKP^{+}20, CC88, CB17, FC07, GCD97, MRAP22, RTF^{+}05, TŠT^{+}17, WZC20, BPSGP^{+}23, DCM16, HL05, IMM^{+}22, NHH^{+}23, PK02, RKS01, SVL^{+}23, SE08, SE09, VWDF14, VFS^{+}15, WSH^{+}22, YHLA^{+}04]. effecting [WMC^{+}89]. effectiveness [JSLA^{+}21]. Effects [CKB^{+}17, CMG15, Cia22, DLD15, FJH10, GAPM16, HMH^{+}15, Hut87, IOGS13, LW13, LFBP^{+}13, MBT07, MPM^{+}17, MCG^{+}02, NTU^{+}14, PCSMC12, PRA^{+}18, RLDC^{+}13, SRFHDH22, SPB^{+}02, STGR^{+}23, ACN01, BTS^{+}15a, BHK^{+}16, BFJ18, CDL^{+}22, CCH^{+}12, DAvD^{+}20, FvBA^{+}17, GCD^{+}13, GWK17, HCAFD^{+}20, JOGM^{+}10, LGK^{+}93, LDB^{+}02, LAHI10, LS85, MMGL^{+}07, MFS^{+}07, MSF^{+}07, NGLSSG14, NFMCS^{+}22, OAWAN18, OÁSG^{+}16, PJH^{+}15, PKF02, PDAM^{+}15, RGPB^{+}23, SHL13, SE92, SLY^{+}15, SIB^{+}06, TNC^{+}09, TGJT09, TG81, TRP^{+}23, TSJC07, VMV^{+}23, WPB^{+}08, WC15, WBF^{+}21, WST^{+}16, XHC^{+}20, ZK06, ZHF^{+}24, dLLdAWL^{+}23, dlHRA^{+}18, MPN09]. efficiency [JF13, SHD^{+}21, VBAC^{+}21, XYGJ23]. Efficient [FW91]. effluent [Tol85b]. effluxes [SIS^{+}14]. effort [PAB^{+}21, SOB^{+}08]. Egg [PD15, BD20, HL05, HLP^{+}16, JSdSS^{+}21, LC12, OACB^{+}15, SJJ^{+}03]. eggs [AIA^{+}15, CTP^{+}18, HLS^{+}14a, MMIB10, PHFK14, SMPC^{+}12]. eiders [LRJ^{+}15]. eight [CTKF^{+}23]. eightsii [GWGR^{+}19]. Ekman [CCD^{+}13, DWH^{+}14, HZCZ16]. Elasipodida [GKR20]. electric [Szu12]. electrification [Bla63]. electron [MBCB88]. elegans [FB05]. Eleginus [VMH^{+}21]. Element [LYZ16, CE84, LJ65, LXC^{+}22]. Elemental [CPPPEAG22, BPA^{+}21, PHK^{+}17, PPHM18]. elements [BJ17, CFC^{+}18, FK86, GSPP^{+}20, HF65, IU14, LTJ^{+}15, ORB^{+}18, PLP99, PBP^{+}99, ZNR^{+}24]. elephant [CdD^{+}15, DMBHG10, FGS^{+}23, LVGH^{+}15, LSF^{+}17]. elevation [DOP87]. Elminius [BB65]. elongata [PRA^{+}18]. elucidate [KAAK^{+}16]. Elucidating [TTB^{+}08b, YAK^{+}08]. embayment [DCM16]. embayments [Ang80]. embedded [BRG^{+}15]. emergence [MV10]. emergent [AGS10, SBLA10]. emerging [ÁLC22]. Emiliania [RPG^{+}18, STHM02]. emission [RVS^{+}21, dlPPÁB24]. emissions [BCP09, GLY23]. emphasis [LMT^{+}19, TKW06]. emphasises [LML^{+}23]. Empirical [HM00a, GCCY^{+}14, MDB^{+}20, TCL20]. enclosed [BLT^{+}15]. encounter [HF10, PTF10, PTF12]. encrasicolus [BFB^{+}20, HPW10, TCF^{+}18, TIOM16]. End [SHG12, CP10, FRCH15, Hea12, KHL12, MD10, MVV^{+}19, RFC^{+}15, SGWF^{+}19, STM10, SCB^{+}07, Ste12, TSJC07]. End-to-end [SHG12, FRCH15, Hea12, KHL12, MD10, MVV^{+}19, RFC^{+}15, SGWF^{+}19, STM10, SCB^{+}07, Ste12, TSJC07]. endemic [RBPGJ^{+}20]. Endemicity [SSB19]. endemism [DGVGR24, HTG15]. Enduring [HBD^{+}21]. Energetic [BH85, GAF15, VBAC^{+}21]. Energetics [JGS90, JFUR20, LL97, ZQWP23, LWT^{+}20, WGM^{+}24]. energy [ACL^{+}18, CNT^{+}19, CZL^{+}24, CRT^{+}22, DCM16, DDK^{+}18, GSFP^{+}09, HS02, LZL^{+}22, LBD11, MGS90, MFS^{+}07, MSF^{+}07, MFM85, NGLL^{+}22, QYF^{+}24, QLY^{+}22, RBE^{+}12, SRT^{+}18, SBD01, SDJ14, XD95, XY20]. energy-rich [SRT^{+}18]. engineering [Pir87]. engineers [DLL^{+}23]. England [HHP06, JHDT12, PL09, USH15a]. English [Ang79a, BTS^{+}15b, BLCL14, PMS^{+}15, RMC^{+}15, STW^{+}15, TAW^{+}15, TB15, ALT10, BTS^{+}15a, BHLU^{+}07, GCD^{+}18, LSD^{+}15, MAH^{+}15]. Engraulis [BFB^{+}20, BDL08, CGC^{+}20, DMC^{+}18, EB08, FVLC^{+}23, GRB^{+}08, HSGJ23, HPW10, PCSMC12, PVA24, SAY^{+}16, TCF^{+}18, TIOM16, YPGE^{+}10, dFKdLZTT17]. enhance [NNFL21]. Enhanced [HZD^{+}23, TSP^{+}13, XWL^{+}22, XY20, XY21, LK13, LCZ^{+}24, TDGY22, vHCY^{+}20]. enhancement [MTK^{+}22]. enhances [BGL^{+}17]. Enigmatic [Wal83]. enough [BBLD^{+}11]. Enrichment [SPK^{+}19, AJA^{+}22, Har82, PKV18, RTF^{+}05, TT05, TDGY22, TŠT^{+}17, TSRF14, TKK^{+}05]. enrichments [FYYC05]. Ensemble [Hob10, MZGA^{+}20, TLH^{+}15, VBL^{+}09, Woo05]. ENSO [WF07, BLR^{+}23a, CS03, DNNNN16, HKK12, hHCK01, LYM12, Leh01, LOO22, LSW^{+}21, LWBD^{+}17, PKP14, RR01, RN02, TWMY08, TTB^{+}08b, WF06]. ENSO-induced [LOO22]. enters [Nof00]. Entoprocta [BC19]. Entrainment [BBS23, APN^{+}15, MTL05]. entrance [BC88, NFMCS^{+}22, SCHBC^{+}22, SAB^{+}21]. entre [Ber65c]. ENVIFISH [BN03]. Environment [JSLA^{+}21, AMG^{+}16, BSW86, BHHS83, BAP^{+}22, CNT^{+}19, ERT^{+}22, FFS^{+}20, GCV^{+}24, ILA21, KBE^{+}22, MAH^{+}15, NCC^{+}15, NMLBCM^{+}01, O’B83, PVA24, PBD^{+}88, RLL^{+}09, SCB^{+}09, SRF^{+}19, Sim81, TSH^{+}17, Ven12, ZL01, SW81]. Environmental [CIL^{+}23, CMC^{+}16, CCM^{+}14, HLSX22, KHP^{+}18, LM18, LFBR^{+}18, MRH^{+}14, PVM^{+}20, Pow06, RFS10, dSSDS^{+}20, SR15, SSW^{+}09, STGR^{+}14, ARD^{+}03, ABD^{+}17, Ang80, AT07, AKH^{+}23, BN03, BGB^{+}08, BGM^{+}99, BBL^{+}18, BDE03, BCL^{+}09, BFV^{+}17, CSV^{+}07, CPG08, DDE^{+}95, DBR03, DVL^{+}99, DMC^{+}18, EHG^{+}12, FJA^{+}21, FRCH15, FSVL10, FPY^{+}16, FC05, FDB^{+}21, GAF15, GCCY^{+}14, GiIKX22, HSGJ23, HMP^{+}13, KSE^{+}09, KKKY10, KOhL^{+}10, LSIC12, LBP15, LEDR^{+}22, LLAPG^{+}22, MMG^{+}13, MRAP22, MFS^{+}07, MSF^{+}07, MP04, Nag01, NGNV12, PJH^{+}15, PCC^{+}19, QOS^{+}22, RGB^{+}17, RvBD^{+}22, RSM^{+}23, SGL^{+}18, SPF^{+}23, SBL^{+}23, SPMVP05, SST^{+}17, SLH^{+}19, TNGP22, TGJT09, THP21, VSGD21, VCM04, VHK03, VHK04, WFD^{+}07, WLKM10, WPH^{+}10, WBD^{+}15, XRC^{+}15, YKNO23, YPVP^{+}22, Zav99]. Environmentally [KQP^{+}17, BEP02]. environments [BW65, BVB88, DDDT99, Gal17, GGSM^{+}20, HDA^{+}16, LSMG01, MPV12, MRBS^{+}24, Pas22, PTF12, ZHD^{+}20]. enzymatic [GGQ07]. enzymatically [GGA^{+}05]. enzyme [RMB^{+}01]. enzymes [SPH83]. EOF [Fuk91]. EOF-based [Fuk91]. eOMP [PPVG12]. Ephemeral [Wai21]. Epibenthic [EMBS13, JPBB20, CMHM18, LM18, LB20, RLP^{+}18, YGMR^{+}23]. epibenthic-sledge [LB20]. epibenthos [JSHB90]. epilogue [AS96]. Epipelagic [BMN^{+}99, BLR^{+}23a, BC01, DWNN04, GDI^{+}09, INI^{+}17, MS17, PBS22, YMA^{+}17]. episodes [GEP^{+}08]. episodic [VCSG^{+}01]. equal [BEI^{+}20]. equation [AdAK^{+}18, McD81a, McD81b]. equator [AFH^{+}11, CC23]. Equatorial [LG23, Luk86, MRRC73, MY92, PGG^{+}22, TLF^{+}89, BLR^{+}23a, GR17, GHC^{+}17, LW85, LG22, MAAS^{+}00, MGK^{+}86, Nof96, OHC^{+}17, OCH^{+}18, RFPG15, RPSC22, SO91, TJ73, VMN08, XWW^{+}21, ZSI^{+}05, FGS^{+}15]. Equatorially [CSLJ03]. Equatorward [KHD22, Kos02]. Equilibrium [Fei93]. era [GWB14, JvdLL^{+}15, MMR^{+}09, MHVS19, Was11, CBD^{+}24]. Erignathus [MSC^{+}15]. Erratum [Ano94c, HHW22, RG03a, SDS22a, VH09a, VHK04, WF07, Yas07b]. error [Oll15, RCGC^{+}16]. Errors [SA97, YHRT22, GJ00]. especially [Bri79]. Esperanza [MPTMK22]. Essai [Ber65c]. Establishing [EBD^{+}20]. establishment [JLP^{+}20a, JLP^{+}20b]. esters [PAG^{+}18]. estimate [AIA^{+}15, BBMR19, GTNK21, LL97, Mac98, WFS^{+}15]. estimated [AYK^{+}05, EiIT^{+}22, GSVB23, GDN^{+}18, LCJ^{+}17, NAH^{+}21, SHS^{+}05, Ver91, YMA^{+}17]. Estimates [AHGRAL23, DGMM85, JVJ^{+}17, Tur65, APN^{+}15, AdAK^{+}18, AVK91, CKT^{+}13, CDB^{+}24, DRD^{+}07, HMO^{+}13, HHWW20, HHK^{+}02, HHB^{+}22, LCBN14, LLH^{+}21, MLHE23, MLK^{+}09, MJC^{+}17, PPKR14, SSM^{+}18, Tal08]. Estimating [ARD^{+}03, BBM^{+}14, BDLW14, FEGA^{+}14, MBF^{+}14, OWH14, OIC^{+}23, OPH^{+}24, CPSM20, KPSA17]. Estimation [BBE^{+}15, RM93, RL85, AHW99, CSS^{+}19, HLCdP19, HHMB^{+}09, PVA24, SSL08, SO91]. estimations [TAM^{+}13]. ESTOC [NCH^{+}07]. Estuaries [USH15a, ARDP14, JGO^{+}98]. estuarine [AJV^{+}02, BBS21, BBS23, CDTM^{+}21, LHP^{+}05, PMMN^{+}22, Pra04, TRP^{+}23, ZHF^{+}24]. estuary [CN22, GEPC15, LHE^{+}13, MPSD15, SSM^{+}90b, SPMVP05, SEO13, USH15b, BJ90, Ham90, JS90, JGS90, JSHB90, LZG20, SJH^{+}90, SC90, SSM90a, SMM^{+}90, WL16, ZLR^{+}07, ZCH^{+}17]. EU [BN03]. EU/Southern [BN03]. Eubalaena [GC09]. Eucalanidae [GBG05]. Eucalanus [GBG05, MFB^{+}84, STS^{+}12, THP21]. eukaryotic [XLL^{+}20]. Eukrohnia [MMN12]. Eulerian [HLS^{+}14a]. Euopisthobranch [CES^{+}19]. Euopisthobranchia [KCBS20]. Euphausia [AMY^{+}23, DSBP15, FMC^{+}20, OTNI20, RBPGJ^{+}20, RB20, SBFP21]. Euphausiacea [OTNI20, RBPGJ^{+}20]. Euphausiid [LO21, FP03, GFGGD^{+}23, MCG^{+}02, PELAA18, RBNJ^{+}12]. euphausiids [Ant09, Bri79, GDI^{+}09, LMM03, PCH08a, RJT84, RBE^{+}12, VDGGD^{+}22]. euphotic [HFW^{+}98, ILA21, WP91]. Eurasian [BS95, RKS^{+}15]. European [BHE^{+}98, BPP^{+}98, FLdST98, HM98, LHEB98, OB98, PS98, SHC^{+}07, TvW98, dWDB^{+}98, vWHdS^{+}98, BB65, BFB^{+}20, BSC^{+}07, BRH^{+}05, CMS^{+}13, CCHV^{+}21, DFM^{+}15, DTG^{+}24, GSM^{+}17, GEPC15, HHK^{+}22, HWLT10, HHH^{+}12, HSC^{+}16, HG04, HHB^{+}22, JTGM10, JCIG18, KDF97, LSH^{+}11, MPB^{+}23, NCH^{+}07, NB87, OACB^{+}15, RAB^{+}11, SON^{+}20, SHC^{+}06, TLH^{+}15, TLP^{+}16, WAH^{+}20, YFK21, dlHRA^{+}18]. Eurythenes [HCV^{+}20]. eutrophic [KOhL^{+}10, LZCZ05, MBCB88, RF17, ZLR^{+}07]. eutrophicated [BVB88]. eutrophication [GRS08, MPMFL^{+}23, RBF^{+}09, SÖÜ94b]. eutrophied [GLF^{+}17]. evaluate [BMN19, HSN^{+}18]. Evaluating [CRT^{+}22, DMC^{+}18, FVA^{+}19, GiIKX22, LHW^{+}20, SEW11, XYL^{+}22, ZGZ19]. Evaluation [DLM91, JJA^{+}08, KSY^{+}19, MDL^{+}12, RMG90, RMK^{+}21, AASJ23, CTKF^{+}23, KAH^{+}16, NNO^{+}14, PJH^{+}15, SOA^{+}23, VPH^{+}12, XY21, YYT^{+}14, ZLR^{+}07]. evaporation [MJWK07]. event [BDTC15, CBB^{+}19, CNBD21, FWBC02, JLP^{+}20a, JLP^{+}20b, JIT^{+}01, KiL14, KS15, Lav09, LWBD^{+}17, MG02, MPMFL^{+}23, RCC^{+}18, RN02, SW22, VBVYT05, WLM07, WZ04, YBPS08]. events [ACL^{+}18, ABP15, BLMR^{+}20, CBD^{+}24, FELMGM^{+}22, hHCK01, JHM^{+}22, KM22, MCG^{+}02, PKF02, PBBH^{+}22, PLN^{+}23, SGL^{+}17, SMdG02, STR01, SKT01, VKT15, VDGGD^{+}22, WLCG23]. ever [UKK^{+}19]. Evidence [Bea04, DJW^{+}18, FAH^{+}13, Hog85, ILA21, IS19, KSG^{+}17, LML^{+}23, LXC^{+}22, Ola65b, SM16, SSW^{+}09, SCS87, TWAL^{+}11, TKWI08, UCB^{+}18, VCSG^{+}01, WFD^{+}07, ZNR^{+}24, BB14, HM90, HM00a, HCGK11, IHY^{+}01, IL20, Kag97, KBHML17, MBT07, MTL05, MCGS^{+}16, OV24, PAF^{+}11, RK20, RBL90, SPH^{+}15a, SHF01, TRLA^{+}13, WWL^{+}22, YCP^{+}12, WHBW03]. evidenced [GdRGL^{+}01, UPPS^{+}21]. Evolution [DB02, MFM85, SI97, BVJE19, CHG^{+}18, DW02, DCL^{+}13b, GLPC23, HHB^{+}00, ILI^{+}12, Jer65, JW01a, Kag97, KBSB18, Lon85, LB02, MNS^{+}24, NW87, NC80, ÖÜT93, SN24, STB^{+}92, SMdG02, SK21, SV97, ZHF^{+}24]. evolutionary [BCL^{+}09, HS22, MJD^{+}21]. exacerbates [LHC^{+}21]. examination [AHGRAL23]. examine [PYKF15]. Examining [SRM^{+}10, WSL20]. example [DAKV99, FVA^{+}19, FC05, LHF^{+}16, LPF23, MZF^{+}08]. Examples [CJMO87, CHG^{+}18, Sei63]. exceeds [LHW^{+}20]. exceptional [CBB^{+}19, CBB^{+}22a, FBB^{+}21, LLS01, MPMFL^{+}23]. Excess [Gam14, MRO^{+}08]. Exchange [BCK94, BBF^{+}22, vWMH98, ARDP14, BBS21, BBS23, BS95, BCR^{+}13, Hut95, HHB^{+}22, KAG^{+}19, LTSG13, RDC^{+}21, VGLCS06, WDMMK89, WBA^{+}22, WC15, WH89, WST^{+}16, YN20, ZAC^{+}23, ZMW^{+}23, dlPHF^{+}15]. exchanges [BATNP04, CdMS^{+}18, EMK^{+}17, HØ00, JOBT05, LLH^{+}21, LZG20, PMC21, Rud89]. excitation [SF15]. exclusive [ZMCD11]. excursions [NB87]. exercise [PPPdS20]. exhibiting [BB14]. existing [DTC^{+}06]. exogenous [PVA24]. exopolymer [MPM^{+}17, ORMR^{+}19, Pas22, RTF^{+}05]. Expansion [RSB^{+}15, BCT^{+}09, OBD^{+}20]. expatriates [OWR^{+}07]. expedition [CP19, MPTMK22, NBR^{+}08, SKF20]. expeditions [Wüs64]. expendable [Mol22]. experienced [FG16]. Experiment [AYK^{+}05, KNI^{+}05, MGK^{+}86, BBM^{+}14, FHL^{+}24, GCZ^{+}00, KIS^{+}05, LFCSV^{+}13, MWJ^{+}08, NH88b, NKK^{+}05, PLJR22, RF17, STB^{+}92, TT05, VSA^{+}21, dJSL^{+}20]. Experimental [ZZM^{+}24, SFAD^{+}90, SEW11, SPN98, ZKK^{+}16]. experimentation [CDS90, SFAD^{+}90]. Experiments [IST^{+}88, RD11, BRB^{+}01, FB01, GAM98a, GAM98b, HHP06, JIT^{+}01, Leg91, LGG18, McD81b, NHS^{+}14, ÖHÜ89, SMGL01, SCLS10, SPK^{+}19, YN03a, YN03b]. expert [YRKC08]. explain [BM01, LBP^{+}21, MHGGS19, XWL^{+}18]. explained [SKH^{+}23]. Explaining [NGPH10]. explanation [Dea85, Due77]. explicit [BDL08, CMS^{+}13, LLS01]. exploitable [LAP10]. exploitation [PL09]. exploited [hHRW^{+}05]. exploration [BHMS09]. exploratory [PKP14]. explored [MCGS^{+}16, WZC20]. Explorers [GBB96]. Exploring [ADV^{+}18, CLCBB19, LSF^{+}17, MRH^{+}18, RDG^{+}21, RDP^{+}21, WSS15, ZCA21, BPSGP^{+}23, LFG10]. Export [AYK^{+}05, Law04, AAMB^{+}24, BEI^{+}20, BT07, CWZ^{+}20, DCKB13, GMAB07, HLM^{+}13, HGH^{+}19, HPZC21, KLB^{+}21, KV13, LBNBM13, LvIKB07, MMG^{+}11, NNFL21, PHK^{+}17, RWOA01, Rud89, TLM^{+}17, TDK^{+}16, WGZZ19, WHBW03, WPW^{+}14, YHM^{+}18, YSN20, ZDM^{+}20, JST^{+}24]. EXPORTS [JST^{+}24]. exposed [YGMR^{+}23]. extended [CN22, Fei03, Fei04, HHWW20, PPVG12]. extends [KGB^{+}23]. extension [Tom81b, LMC^{+}20, QNK^{+}22, WCX^{+}21, WWZ19, YTNK00]. Extensive [HTG15, Hut92, INI^{+}17, SDL^{+}19]. extent [DLD^{+}19, GTS^{+}21, HKN^{+}14]. external [ZGZ19]. externe [Ber65c]. extinct [RPRCAG^{+}21]. extinction [CMF11]. Extraction [Coo69]. extraneous [AF10]. Extrapolation [GMDS20]. Extreme [MPMFL^{+}23, BBB^{+}21, FDB^{+}21, GPEV20, KM22, LEDR^{+}22, PKA19, RHB23]. extremes [CBB^{+}22a, OWH14, ŠPM^{+}22]. extremite [Bou65].

F [Ang80, Ang88]. face [LOO22]. facies [Ike88]. facilitating [GBC^{+}15]. facing [ORPRGIS22, TTF^{+}22, Val99b]. factor [BWMGCB08, CGD^{+}18]. Factors [BPP^{+}98, FPIJ85, LZCZ05, STF^{+}13, SNMW10, AH10, CMC^{+}16, CPG08, CVBG21, DDE^{+}95, GiIKX22, HBV^{+}99, HYM^{+}12, HKPV12, HMKF08, HFO90, Li14, LLGS21, MVN^{+}15, MRAP22, MS17, MRH^{+}14, NCC^{+}15, TNY^{+}24]. facts [Hic79]. faecal [RWOA01, WYT00]. faeces [FGL^{+}23]. failures [KMOM88]. Falkland [WR00]. fall [BDC^{+}08]. family [MKD90, MMK19, MSFZ19]. far [RSG06]. Farallones [JSA^{+}08, WDMC02]. farm [CTA16, FvBA^{+}17]. Faroe [HØH^{+}03, HØ07]. fast [CTP^{+}18, DSV^{+}24]. fat [GA10]. fate [BPNB90, JW01a, OWR^{+}07]. Fatty [RBPGJ^{+}20, WOW^{+}14, CSG^{+}15, CDP14, GVKD^{+}13, KO19, KSG^{+}17, LSV14, MRBS^{+}24, MFDH22, PPHM18, SBC^{+}16, WPB^{+}08, YGL^{+}10]. fauna [BS90, CGM^{+}02, CMM^{+}04, DHB^{+}21, FJ19, FTHK19, GHSC19, KSB^{+}22, NRA^{+}21, SSB20a, TJ90]. faunal [GIC20, HM90, PGGG17, dJSL^{+}20]. faunas [HM90, SPB19, SCS87]. favorable [FAAV^{+}15]. Fe [GHG^{+}24, WH20, ZNR^{+}24]. feasible [TSL10]. feature [DCD^{+}23, PHKS01]. features [BAARB05, BSA06, CB06, CFM^{+}18, CdTH^{+}16, CdD^{+}15, Ken88, KSK21, Kra69, KAAK^{+}16, Lie86, LLH^{+}20, STC10, TZP^{+}00, WR00, XYL^{+}22, Zez90]. February [Ano22-27, Ano00d, Ano03e, Ano08s, Ano13l, Ano15n, Ano16p, Ano17m, Ano18j, Ano19m, Ano20m, Ano21n, Ano22u, Ano23n, Ano24h, YYC^{+}18]. Fecal [SE92, Tur15]. Fecundity [AVK91, JSdSS^{+}21]. feed [BBLD^{+}11, Mae88]. feedback [Bak01, KV13]. feedbacks [DFC^{+}21]. Feeding [BHS^{+}15, CSM^{+}15, FLdST98, SCAA07, SMN^{+}13, SPK^{+}22, BOMdP15, BSH^{+}20, BMG^{+}19, BPSN^{+}21, CNT^{+}19, CCG07, CPO^{+}19, CPG08, CCS^{+}21, CCM^{+}14, DM13, DAF^{+}22a, DAF^{+}22b, GBM^{+}01, GA10, GD91, GSC^{+}20, HWBT03, IVR^{+}13, KTIT22, KSB^{+}22, KVNT20, LdSH^{+}15, LRJ^{+}15, OTNI20, Roe84a, RB84, SE92, VWDF14, WHBW03, YHLA^{+}04]. female [Nie07]. females [ATT^{+}08]. Fernández [MPM^{+}18]. Ferrol [CVHM^{+}18]. ferromanganese [Gri22, VCM04]. fertilisation [Qiu15]. fertilization [Peñ03a, SSH^{+}05, SHS^{+}05, TSNO05, YFY05, ZCD08]. fertilized [TNS^{+}05]. FESOM [SKWWGV18]. FGGE [MGK^{+}86, MFM85]. Fidji [Rot65]. Field [HCGK11, YCP^{+}12, BKB85, CCH^{+}12, Emi65, FTG^{+}11, HLPL05, Jac10, JST^{+}24, KSY^{+}19, LRS^{+}03, ORCH^{+}19, RAE^{+}05, STB^{+}92, TM13, WPB^{+}08, XLX^{+}20]. fields [AR18, BJMP19, BJMP20, KM22, MFM85, RL85]. fifty [HDZY15]. Fiji [Rot65]. Filament [ILI^{+}12, ÁSDB^{+}01, ABT^{+}04, BIST01, BATNP04, BFH01, FRK^{+}09, GMAMB04, HHB^{+}01, HPS^{+}01, HCAFD^{+}20, JIT^{+}01, NIF^{+}15, QYF^{+}24]. filament-eddy [BATNP04, GMAMB04]. filament-like [HCAFD^{+}20]. filaments [BA04, KCPM09, SFMA20, YHLA^{+}04]. Filipjev [MSFZ19]. filter [MZGA^{+}20]. final [Ano94k]. find [PPPdS20]. Finding [TRLA^{+}13]. Fine [AGD^{+}18, CTF07, DRVMC^{+}22, DIQJ21, JHW^{+}14, MCB^{+}90, NBLI20, SWP^{+}13a, GBC^{+}16, LL21, Sie69]. Fine-scale [AGD^{+}18, CTF07, DIQJ21, JHW^{+}14, MCB^{+}90, NBLI20, GBC^{+}16, LL21]. finer [BMGN15]. finger [Sch03, YN03b]. fingering [Kun03]. finite [XY21]. finmarchicus [BMK12, BTNK13, BRE^{+}24, DBM17, GHF^{+}21, GPC^{+}03, HMP^{+}13, HRA^{+}08, HBR11, MRH^{+}14, SDH^{+}14, SAB^{+}21, UB10, WBC^{+}22, WSH15]. First [GHSC19, MDR20, PSP^{+}21, PMFNGQ21, War06, AM19, AV23, BC19, CLMR23, MWS^{+}10, PMMN^{+}22, WHG^{+}16, YTL^{+}19, MGK^{+}86]. Fish [ESA^{+}13, GIHJ23, ICB^{+}19, LBC^{+}15, SGF^{+}19, ARD^{+}03, AHP19, AIA^{+}15, ASB^{+}08, Bak01, BBLD^{+}11, BFB^{+}20, BECR^{+}22, BJ90, BCL^{+}09, BAP^{+}22, BFV^{+}17, BB10, BHMS09, CIL^{+}23, CJRÁ^{+}13, CBOP15, CTP^{+}18, CRF^{+}10, CCB^{+}20, DWNN04, DDK^{+}18, Dol09, DPH^{+}18, DHHP18, DPM^{+}09, EAB^{+}23, ERT^{+}22, FTC^{+}16, FARRL^{+}13, FFS^{+}20, FGGDF^{+}04, GMR^{+}23, GD85, GRD^{+}23, GFB^{+}15b, GFB^{+}15a, GCG^{+}14, GDM^{+}15, GAPM16, HM90, HLS^{+}14a, HHW01, HHW22, HAA^{+}14, HSL96, KSVT00, KSS^{+}23, KYS^{+}17, LSY^{+}14, LAP10, LLAPG^{+}22, LSIB23, MRM^{+}14, MPMFL^{+}23, MRAP22, MEST13, MKB00, MLHE23, MHS^{+}09, MAFS^{+}22, MVBC^{+}21, NHH^{+}23, NPO^{+}19, OMR^{+}22, OMK^{+}22, OPH^{+}24, ORB^{+}18, POS^{+}07, PRTC13, PAB^{+}21, PHFK14, PCR^{+}22, QLW10, ROBRB^{+}22, RTBR^{+}22, RBHLA04, RB84, RFC^{+}15, RG94, RBE^{+}12, RSB^{+}13, SOS^{+}07, SKH^{+}23, SCD^{+}07, SCHD23, SLPA^{+}20, SCC^{+}19, SEG^{+}22b, TCL20]. fish [TKW06, TS10, THM^{+}14, UKK^{+}19, WWL^{+}22, XYL^{+}22, YAK^{+}08]. fish-induced [TS10]. fish-killing [MPMFL^{+}23]. fish-mediated [MLHE23]. fisheries [BN03, BBSN04, BECA22, BDE03, CLB^{+}13, CRT^{+}22, CWS^{+}21, CL03, FFA06, FPS^{+}13, Hea12, JPM^{+}08, JOGM^{+}10, JHDT12, JBH20, KN10, KN11, LNB13, LMP22, MEST13, MBF^{+}14, MCB^{+}10, MCG^{+}14, ML09, MPN09, ORPRGIS22, RSK^{+}23, SRAV19, SJP10, SBM^{+}23, TCF^{+}18, THM^{+}14, WWL^{+}22, WLKM10, YBS^{+}01, ZLKO00, ZLS^{+}04]. fisheries-based [FFA06]. fishermen [JBB^{+}14]. fishers [BDT^{+}08]. fishery [ACN01, BW08, BMN19, FBM^{+}08, GRDS10, INI^{+}17, JBB^{+}14, KHL12, MCL^{+}15, OIC^{+}23, PWZ^{+}16, RS04, TNGP22]. fishery-dependent [OIC^{+}23]. fishes [ACB^{+}13, BM76, BLI^{+}99, BOMdP15, CRT^{+}22, DCKB13, EKB06, EBD^{+}20, ESD^{+}21, Hob10, hHRW^{+}05, Kos93, MGS90, MS17, MM80, MMPG07, OHC^{+}17, OCH^{+}18, OT19, PSA^{+}19, SGF^{+}19, SLBVRR^{+}22, SBL^{+}23, SKF20, SPK^{+}22, TCF^{+}18]. Fishing [GCF^{+}19, SENS13, AKAL20, BDL08, ERT^{+}22, GAF15, JRR^{+}24, LML^{+}23, MVV^{+}19, NHS^{+}14, PG10, QLW10, SSI13, TKW06, TSJC07, VBJ^{+}20, WJPHB15]. fit [GCCY^{+}14]. fitness [LLAPG^{+}22]. five [BJMP19, BJMP20, DHC^{+}20, DAF^{+}22a, DAF^{+}22b, DSAB20, GSFP^{+}09, HSG^{+}15, HSC^{+}16, LS15, LWY07]. fixation [LFBR^{+}18, NHH^{+}23, WCC^{+}20, WFBN^{+}13, ZCD08]. Fixed [CNBD21]. Fixed-point [CNBD21]. Fjord [MSV^{+}14, BRR^{+}22, BRE^{+}24, CKB^{+}17, FB05, GHVG19, HWPLvW20, KYT^{+}16, KSKN21, LSV14, LHP^{+}05, LWT^{+}20, MTC14, PTM^{+}22, PVM^{+}20, RBL^{+}19, RPSVLS14, SWZS^{+}21, SST^{+}17, SPW22, VLCCP14, WZBK^{+}21, ZHSMM14, ZCLS20]. Fjords [FF83, IPD14, KMSTK23, SV14, CCM^{+}14, CBPS^{+}22, CTI^{+}19, GHVG19, GCD^{+}13, IPF23, MPMFL^{+}23, PMA^{+}14, PSGVS^{+}14, PTPY^{+}23, QOS^{+}22, SPF^{+}23, VNH^{+}23]. flapping [Ric22, SSB14]. flat [DOP87]. Flatfish [WHI^{+}02]. fleet [GRDS10, PPSV^{+}18]. flemingeri [Mil88]. Flexible [CdD^{+}15, BPA^{+}21]. flexible-stoichiometry [BPA^{+}21]. Flexicover [Hof81]. flight [Sac16, SSB14]. float [RBS^{+}20, RBS^{+}22, Ric93, RLC85]. Floating [PHFK14]. floats [KSK21, LS20, Owe91, PL18, RBZ00, ZBLF23]. flood [Whe06]. flood-dominated [Whe06]. floods [FBB^{+}21]. floor [SW65, TCL^{+}15, WGCS13, vHCY^{+}20]. Florida [KAK^{+}22b, AC85, LW12, MZ14, RLL^{+}09, RLGC10, WWL^{+}22]. Flow [Rei86, Rei89, Rei94, BBS21, BBS23, BHPC06, CSLJ03, CTP^{+}18, Dav85, FAAF88, FMWW14, GL06, GD91, HCAFD^{+}20, HKGH^{+}06, Hol00, HSH97, HS02, JFEC13, LSM^{+}22, LGL^{+}18, MMGL^{+}07, MFS^{+}07, MSF^{+}07, MTL05, Nof03, PLB^{+}23, Rei97, Rei03, RBE^{+}12, Sak86, SHS^{+}05, TMN^{+}12, TM13, TWMY08, TG05, THM^{+}06, TSP^{+}13, Wun24, dJSL^{+}20, vAB96, BBF^{+}22]. flow-through [HSH97]. flowing [FZ88]. flows [FTG^{+}11, GvOSW11, KCL^{+}12, Mil14, NGLL^{+}22, Pra04, SDJ14, TMPM^{+}16a, TMKJ^{+}09, TFM03, VBL04]. fluctuation [FWL^{+}15, RI86]. fluctuations [AMG^{+}16, BBL^{+}18, CSMGS19, CD65, DL69, Don65, LFI^{+}13, NH88a, QLY^{+}22, RCD^{+}94, RFS10, SLG^{+}12, Seg69, SEG22a, SDO^{+}14, TMN^{+}12, TRLA^{+}13, Zen08]. fluid [GD91]. fluids [MSV^{+}14]. fluorescence [MZ14, MSMH19, SAT^{+}22]. fluorescent [GLAHH^{+}22, MPCNC^{+}19, XYGJ23, YTB^{+}21]. fluvial [Pow06]. fluviale [Ber65c]. flux [AGL^{+}15, BSC^{+}19, BSW86, BRG^{+}23, BS02, CKP^{+}20, CHC^{+}12, DDD^{+}00, DOS^{+}18, DRD^{+}07, FUOG^{+}16, HLM^{+}13, HDB13, Jón07, KGB^{+}23, KSP^{+}23, LPA92, Law04, MMG^{+}11, MDC^{+}07, NMK^{+}03, ORW^{+}01, Oll15, PG13, RCS^{+}11, SLBVRR^{+}22, SPG^{+}06, SG91, SW01, SDS^{+}22b, SW65, TRLA^{+}13, TSG^{+}04, TWBC^{+}13, VK90, WZFW16, WCB20a, WYT00, WRH^{+}06, WHBW03, WBH15, WWL^{+}24, XCH^{+}16, YHM^{+}18, ZPY^{+}20]. flux-rich [TRLA^{+}13]. Fluxes [FPD^{+}01, AAML22, ÁSDB^{+}01, AYK^{+}05, ABT^{+}04, AT07, BIST01, CFG07, CF12, DDDT99, DVB^{+}18, EHSI12, FLUC08, GDI^{+}09, GCD^{+}13, GIPG17, GBB^{+}19, GLLB22, Hea12, HMKF08, KKS^{+}18, KNSN^{+}09, KZSH85, KF11, KHC^{+}99, KFC^{+}23, KTB^{+}99, KSPK99, LBNBM13, LBP^{+}21, LGZ^{+}20, LvIKB07, LFCSV^{+}13, LDHW20, MHGP06, MLD^{+}03, MFM85, MSL^{+}07, NFMCS^{+}22, NIF^{+}15, OKdA^{+}19, PGLG^{+}05, RCB^{+}20, RAP95, RGI05, RGE22, RCSVGP^{+}16, SLBR18, SCCJ^{+}18, SBB^{+}22, SCT^{+}00, TFY02, TCDPP^{+}22, TSC03, TDK^{+}16, TDL^{+}17, TSP^{+}13, WJE^{+}92, ZFSV^{+}09, dlPHF^{+}15]. fly [Ric22]. flying [ATT^{+}08, IIS^{+}17]. FOCE [SPK^{+}19]. focus [BZD^{+}21, BB24, DRVMC^{+}22, Rud15, SPF^{+}23, SJM^{+}19, TTB^{+}08b]. follow [VKGP^{+}13]. following [PBBH^{+}22]. Fonera [PGLG^{+}05]. Food [Bon88, CLSD18, CW06, CFML22, GvOSW11, GWK17, IBW^{+}01, JE92, OPG^{+}10, WRH^{+}06, AHW^{+}15, BHA^{+}14, BRE^{+}24, BAOC^{+}07, CLSP17, CP10, CBC^{+}06, CPPPEAG22, CSV^{+}07, CSBL^{+}15, CSC^{+}12, DJW^{+}18, DYO^{+}10, DTOD00, DKRL22, DSB^{+}24, DWC06, FMC^{+}15, FVLC^{+}23, FTG^{+}11, GAF15, GSVB23, GSM^{+}17, Hea12, HKGH^{+}06, HNL14, iIYO^{+}10, LK13, LLL^{+}11, LCBN14, LZF^{+}24, MRA^{+}19, MGC^{+}18, MCH^{+}12, NYL^{+}17, NMC^{+}09, PVG^{+}20, Peñ03a, PG10, Pow06, RMC^{+}15, RBE^{+}12, SBMB18, SH09, SJJ^{+}03, SSM90a, ŠGM^{+}18, SHC^{+}06, SHC^{+}07, SRT^{+}18, SLY^{+}15, SDJ14, TNGP22, TR99, TSS^{+}12, TFM03, VGJ^{+}19, Was06, YAK13, ZHSMM14, ZBRJ23, dJSL^{+}20]. food-mediated [GSM^{+}17]. Food-web [OPG^{+}10, CP10, CSBL^{+}15, MGC^{+}18, dJSL^{+}20]. foot [vHVAT22]. Footnote [SDS22a]. Footprints [MGWZ20]. forage [ASB^{+}08, BBLD^{+}11, BHMS09, CRT^{+}22, DDK^{+}18, EKB06, GMR^{+}23, OMK^{+}22, PSA^{+}19, PHKS01, RBE^{+}12]. Foraging [FDB^{+}21, NRS^{+}19, STC10, WLKM10, YSS14, BMG^{+}21a, CdTH^{+}16, DMBHG10, LSF^{+}17, LBC^{+}23, MHR^{+}10, STEB16, SWP^{+}13a, SWZS^{+}21]. foraminifer [AT07]. foraminifera [DGVGR24, KGdS^{+}08, Par65, Phl65, Sai65]. Foraminiferal [Ban65, CBL^{+}19, DSC^{+}19, GBB96, GA01, LvIKB07, NFMCS^{+}22, SBG16, Wil65]. foraminifers [AOMZ^{+}23]. Forbes [Mer65]. Forbes-Manxman [Mer65]. forced [CPSM20, CD07, DWH^{+}14, Hol00, LPF^{+}21, LO85, SFMT12]. Forcing [LSD^{+}18, SFMT14, WLD^{+}15, AALM06, AC85, AHW99, ASB^{+}08, BLT^{+}15, BDT^{+}08, BGB^{+}08, BSW86, BPSGP^{+}23, Cai95, CBC^{+}06, CPC^{+}15, CS06, CAB^{+}99, DLM^{+}12, DMC^{+}18, DPM^{+}09, DMF^{+}09, DCL^{+}13a, DP13, FZ88, FBS22, GAF15, GPC^{+}03, GEP^{+}08, Kaz17, KSE^{+}09, Kit03, LMH^{+}13, LG23, LBP15, MZH^{+}23, OACA20, OAB^{+}16, PMA^{+}14, PSM^{+}22, RA15, RHM^{+}19, SCPN15, SBD01, SSW^{+}09, SP08, YAK^{+}08, ZK06, ZDM^{+}20]. forcings [TCF^{+}18, dMGS^{+}11b, dMGS^{+}11a]. Forecast [JAS^{+}20, CBB^{+}22a, CP83, RSK^{+}23, YPGE^{+}10]. forecasting [CCS^{+}21, GEPC15, PLK14, Wai21]. Forecasts [SYB^{+}15, HKL^{+}15, LSS^{+}10, TSH^{+}17]. forests [GCF^{+}19]. Foreword [LHM^{+}99, Koc65b, Kra82]. Formation [BGV^{+}23, LPW^{+}23, MSGGM18, NAH^{+}21, PVV23, BS95, CGL^{+}20, CS18, ILI^{+}12, Kat18, KY15, LRNK99, NIC^{+}19, NC80, PSL87, PC87, PZA^{+}15, PMS^{+}15, PGC^{+}96, Rud89, SMFM^{+}21, San73, SASH08, TSAM^{+}22, TG05, VKT15]. formed [ELW06]. formulae [ARDP14]. formulation [MAB^{+}11c, RRLS22, XY20]. formulations [AGS10, BDP^{+}06]. Fortnightly [VGLCS06, BPSGP^{+}23, RCSA01]. forty [CPB^{+}15]. fosse [Rot65]. fossils [ST65]. foster [RMC^{+}15]. found [OT19]. foundations [FvBA^{+}17]. four [BHMS09, CED09, Coo69, DMF^{+}09, GS24, HFNG00, HvdLS^{+}09, LSS^{+}09, MKHO96, MLK^{+}09, MMPG07, PM13, SH09, SBE^{+}20, TKW06, WB03]. four-component [MKHO96]. four-dimensional [Coo69, WB03]. fraction [WSC^{+}21]. fractionated [BTJ^{+}17, FEGA^{+}14, JTD^{+}14, LHC^{+}19, SSN23, XHW^{+}20]. fractioned [SEG22a]. fractions [SRF^{+}19]. Fracture [Nay65, RMG90, SM65]. fragmented [AHC^{+}13]. Fram [GIC20, GSSWK20, HFPS^{+}06, LBNBM13, LF12, MGWZ20, MRO^{+}08, Rud89, RKS^{+}15, SSV^{+}11, TSBS18]. framework [BRG^{+}15, BMG^{+}19, BBF^{+}22, FPS^{+}13, GLH13, HSN^{+}18, JHDT12, KHS^{+}14, SJLW23, TSJ^{+}12]. France [RSM^{+}23, LDD^{+}22]. Fraser [McK08, MI21, SEO13]. free [DS65, MSFZ19, PKA19, WM13, WDC^{+}11, WO15, SPK^{+}19]. free-living [MSFZ19]. freezing [NAH^{+}21]. French [Ber65b, CCB^{+}20, Rot65, Rou65]. frequency [AC85, Bak01, BCLD^{+}17, CM18b, DOS^{+}18, FRV^{+}19, MNT14, NH88a, PZA^{+}15, ŠVL^{+}15, GDN^{+}18]. fresh [JAJS08, Jón07, TAW^{+}15]. Freshening [ONR^{+}14, CRPS^{+}15, WWSJ07]. fresher [HMH07]. Freshwater [BSF95, SSQ19, Tal08, AYH^{+}23, BvdLA^{+}11, BF11, CTI^{+}19, DRD^{+}07, HS07, LLH^{+}21, MTC14, MWS^{+}10, NBR^{+}08, RFSCF19, RGM01, VNH^{+}23, WHS17, Yas07a]. freshwater-influenced [RFSCF19]. friction [LL97, Mun97]. frigatebirds [WLKM10]. fringe [CBB^{+}22c]. Front [Ano17i, Ano17j, Ano17k, Ano17e, Ano17f, Ano17g, Ano17h, PHKS17, STC10, VFS^{+}15, BLR^{+}23b, CMC^{+}16, FPIJ85, Gou85, HHP06, MSI17, OC06, PHKS01, RPPM^{+}23, Ang89, BZD^{+}21, CHSB^{+}21, DCD^{+}23]. Frontal [BSMC15, HMS^{+}22, INI^{+}17, VFCC^{+}22, CLV^{+}19, CPO^{+}19, HBH^{+}17, JOBT05, Kaz17, KFC^{+}23, LL21, LLGS21, NMO^{+}21, RLL^{+}09, Tom81a, WMB^{+}21]. Fronts [BCS09, ARELAK24, BZD^{+}21, BTS22, CLY22, KZSH85, LB14, NXT^{+}17, PLEF^{+}23, PO15, Pra22, SDP^{+}22, Sou94a, VMA^{+}24, XYWY23, XNT^{+}17]. frugality [MHGGS19]. fucoxanthin [CMC^{+}16]. fuegensis [CCM^{+}14]. Fuegian [PMA^{+}14]. full [CN22, LFG10, PST^{+}15]. fully [RFC^{+}15]. fully-coupled [RFC^{+}15]. function [AV23, FM07, HKGH^{+}06, MLM09, Nee85, PBH^{+}10, TAM^{+}13, VBL04, VHK03, Was06, ZWM^{+}15, dLLdAWL^{+}23, VHK04]. Functional [FAB^{+}09, KST03, ANMP15, ASR^{+}20, BPA^{+}21, BMGN15, BMdMS^{+}21, DDK^{+}18, EBM^{+}20, FTC^{+}16, FGS^{+}23, GCC^{+}24, GDM^{+}20, GNH19, GDI^{+}09, Hof10, JG90, LMS10, LBP^{+}21, MPC12, Peñ03b, RLP^{+}18, TSL10, VWDF14, VSGD21, THBA19]. functionally [MHS^{+}09]. functioning [BGM^{+}99, CDL19, CSG^{+}15, CCB^{+}20, EAL^{+}07, ES07, GRS08, MGC^{+}18, TCS15, TDH^{+}95, VPM^{+}19]. functions [TCL20]. fundament [dLLdAWL^{+}23]. fungi [BRC^{+}18, CJG88, GGT^{+}15]. Funiculina [PRA^{+}18]. funnel [MXC^{+}21]. funneling [She65]. Further [BB65, OV24, iUMY86]. fusiformis [GA01]. fusion [VVV21]. Future [Tho87, BW08, BCD^{+}20, CCM^{+}13, CBGC^{+}08, CSS^{+}19, CBOP15, CLdPHL23, Den87, GPAB^{+}16, HM15, Hob10, IHY^{+}01, KKK04a, LDMH09, MPC12, NPO^{+}19, NW87, OVG16, PV18, RDG^{+}21, RSB^{+}15, SDH^{+}14, SAd^{+}17, SBB^{+}14, SPK^{+}19, SJ18, SBC^{+}24, TMÁGC^{+}21, WAH^{+}20, WDC^{+}11, WO15, Was15, WH89, vdS94b]. fuzzy [JPM^{+}08]. fuzzy-logic [JPM^{+}08]. FVCOM [CGZ^{+}16, LC22, SCB^{+}16].

G [Ang79a]. G. [AHRT90]. Gabès [FTG^{+}18]. Gabriel [PCK^{+}06]. gadiforms [DKRL22]. Gadus [AHP19, CHS^{+}24, FKH^{+}13, GTS^{+}21, LSY^{+}14, LNB13, SLY^{+}15]. gained [OELP04]. Galápagos [HMB^{+}86]. Galicia [BAOC^{+}09, DGGdR02, FB01, HPS^{+}01]. Galician [BF01, OÁSG^{+}16, VSPP14]. Gannet [SWP^{+}13a]. gap [DP18, LMS10, MCG^{+}14]. gaps [CLdPHL23, IPF23]. GARCH [PVA24]. GARP [MGK^{+}86]. gastropod [FTHK19]. Gastropoda [CES^{+}19, KCBS20]. gastropods [IPHW^{+}23]. Gateway [ZWP23, JPIP22, PNF^{+}21]. gaudichaudi [AE09]. gauge [HHWW20]. gauges [RM97]. Gaussian [SBFP21]. gayi [BWMGCB08, LC12]. gazetteer [Ang79a]. GCMs [HLK13]. gear [Reb02]. gears [JPBB20, GCF^{+}19]. Gelatinous [MMMWZ23, MGWZ20, dWDB^{+}98]. gender [IVR^{+}13]. gender-life [IVR^{+}13]. genera [MMK19]. General [PGLG^{+}05, TAO05, AGS10, Arb22, MEMP15, SPSV^{+}20, SBLA10, TAH^{+}11, WG82]. generalised [BBMR19]. generality [AF10]. generalization [TCL20]. generated [FG16, Ike88, PdMS^{+}13]. generating [SBD01]. Generation [CBB^{+}22a, HYM^{+}24, BSA06, CCRS20, GLPC23, ORMR^{+}19, PLK14, TCN20]. genesis [OC06]. Genetic [GBG05, ACK^{+}13, BBFS19, HTG15, NMN08, RK20, RAG^{+}19, UB10, WPBG^{+}18]. Genkai [MKM86]. Genovese [Bru88, SMB88, Sie88]. genus [CES^{+}19, GM19, KKKS14, Mar20, MSFZ19, PTPY^{+}23, UKK^{+}19]. geo [RAB^{+}11]. geo-hazard [RAB^{+}11]. Geochemical [AJV^{+}02, MCGS^{+}16, CM11, JFG^{+}90, Mor91, VSGC21]. Geochemistry [MSV^{+}14, AS20, CKB^{+}17, CE84, GSV^{+}01, LYZ16, Soh03]. Geochronology [RANS65]. geoeconomical [Ola65a]. Geographic [HRSM08, iIRM^{+}15, MMPG07, NHH^{+}23, QOS^{+}22, SKH^{+}23, ZSH^{+}24, CALS^{+}23, MRH^{+}18, SAY^{+}16, WCS^{+}23]. Geographical [GCC^{+}24, ZHBW01, HPB^{+}09]. geographically [NHG19]. geoid [NB87]. Geological [RLT^{+}22, Sei63, Gor92]. geology [SC90]. geomagnetic [Emi65]. Geophysical [CBPS^{+}22]. Geophysics [VH09a, VH09b]. George [AYH^{+}23].

Georges [BBL^{+}09, GGJ^{+}10, SCB^{+}07]. Georgia [FDM^{+}13, GKS^{+}13, GDM^{+}15, LMH^{+}13, LC10, PMC16, PM13, RHBS13, STF^{+}13, AW13, AHC^{+}13, ESTM13, GLPC23, Ham87, IIM^{+}23, IOGS13, LAP10, MGF^{+}13, MP13, PBN13, STC10, SLH^{+}19, WMWR08, XCH^{+}16]. GEOSECS [JvdLL^{+}15]. geostationary [HYM^{+}24]. geostrophic [CP07, KAG^{+}19, Rei86, Rei89, Rei94, Rei97, Rei03]. geothermally [YYK88]. GEOTRACES [AH15, BGR^{+}15, HM15, ZNR^{+}24]. German [DOP87]. Germany [PDV12, VPH^{+}12]. Geryon [AHRT90]. Ghir [AAM^{+}14]. giant [JPBB20]. Gibbs [Fei04, FH95, Fei03, FM07]. Gibraltar [BMC05, CGMP14, SGMP15, BPGC^{+}20, BCR^{+}13, BCK94, Dea85, FAAF88, GR85, GWS^{+}23, MMGL^{+}07, MM90, NGLSSG14, Ric94, SBPGP^{+}23, VGLCS06, dlPHF^{+}15]. gigas [ATT^{+}08, BGM^{+}10, RS10]. Gills [JG90]. Gioia [CFM^{+}18, GCLD19, PRA^{+}18]. given [MEMC05]. Glacial [Bro82, HM98, TAF^{+}22, DPB06, DYL^{+}15, DCL^{+}13b, HHR^{+}19, KSKN21]. Glacial-interglacial [HM98]. glacialis [AEP^{+}23]. glacier [CKB^{+}17, GHVG19]. glacier-fjord [GHVG19]. glaciers [BRR^{+}22, VNH^{+}23]. glaciomarine [ZCLS20]. gladius [SYB^{+}15, YGL^{+}10]. Glider [KHD22, MBdM^{+}18, CLY22, PO15, SCY^{+}23]. glider-resolved [PO15]. gliders [RZTD17]. Globa [MGK^{+}86]. Global [Cai95, CSS11, DPCS87, GRMB18, JTGM10, KT97, LSH^{+}22, MBS20, MMN12, OVG16, PBO10, SDJ14, TAM^{+}15, WQ08, AGS10, Arb22, BBC^{+}22, BDLW14, CTKF^{+}23, CWS^{+}21, DEW^{+}97, ERBV21, GJ00, Gri22, GZCL23, HS07, HGB^{+}21, HHDS02, HTG15, HMKF08, HMH^{+}15, HLFL23, IHY^{+}01, JBH20, KV13, KKKY10, KKO10, LGZW22, LM97, Mac98, NGLL^{+}22, OWH14, PAB^{+}21, PSA^{+}19, PHKS01, PPD^{+}12, RG09, SE08, SE09, SBB^{+}22, SJLW23, Tal08, TMH^{+}16, TAM^{+}13, VWDF14, WWW^{+}23, WSS15, WSO01, Whi95, WLM^{+}22, WBH15, XYWY23, YW22, YAK13, ZCD08, CGZ^{+}16]. Global-FVCOM [CGZ^{+}16]. Global-scale [SDJ14]. globally [LRW^{+}15]. GLOBEC [Ano10a, BPW10, BK08, EAL^{+}07, PDV12, PBOW10, PBH^{+}10, SWP^{+}13b, VPH^{+}12]. Goban [BHE^{+}98, FLdST98, LHEB98, OB98, TvW98, FMH02, vWHdS^{+}98]. going [SWP^{+}13b]. golfe [Ber65b]. Golfo [CTF07]. gonad [Nie07]. Gondwanaland [Fai65]. good [GEPC15]. gorbuscha [KBF^{+}08]. Gorda [CSR90]. Gordian [HPNDC15]. gorgonian [GGA^{+}16]. Gorshkov [Ang79a]. governing [BPP^{+}98]. Gower [Ang88]. gracilis [VMH^{+}21]. gradient [BMdMS^{+}21, BGWP^{+}17, CBL^{+}19, DWH^{+}14, GCD^{+}13, KLC^{+}15, LvIKB07, MCD^{+}14, PCH^{+}08b, SST^{+}17, ZZM^{+}24]. gradient-dependent [ZZM^{+}24]. gradients [BLES16, CSG^{+}15, EBW^{+}23, FBD18, GBB^{+}19, JTD^{+}14, SPMVP05, VKGP^{+}13, XWL^{+}18, ZLR^{+}07, dLLdAWL^{+}23]. grain [PPSVC^{+}13]. Gran [Ano09h]. Grande [FBT^{+}22, SNdSR^{+}24]. grandis [RK20]. graph [DYO^{+}10]. graph-theoretic [DYO^{+}10]. gravimetric [LM97]. gravity [Arb22, KCL^{+}12, MZZ^{+}23]. grazer [GBH^{+}20, VPM^{+}19]. Grazing [Dag93, NYL^{+}17, TKSI08, AGS10, Ban96, BM07, CMC^{+}16, HHB^{+}01, LGK^{+}93, LMS93, LOG^{+}09, LHP^{+}05, LPHL^{+}05b, LPHL^{+}05a, MFB^{+}84, MC88, RRLS22, SPB93]. Great [BLAM00, Leg91, WLKM10, BLAM98, BECR^{+}22, DMML88, Her97, HT97, HPHW21]. Greater [ROBRB^{+}22, RLR^{+}18]. greatest [UKK^{+}19]. green [DTG^{+}24, YFK21, TAM^{+}13]. Greenland [JJR^{+}08, JAJS08, JJ08, MFM15, NBR^{+}08, BS95, GS24, HBL^{+}13, HJLLN07, HHR^{+}19, Leg91, MMMWZ23, MLS^{+}15, MWJ^{+}08, MDR22, ON22, RN06, SBK^{+}95, Ste91, SP08, SBS90, VEM^{+}21, VNH^{+}23, WBH15]. Greenland/Norwegian [BS95, SBK^{+}95]. gregaria [CSMGS19]. grenadier [MMPG07]. Grey [NBLI20]. grid [CTA16, SZG06]. gridded [DPF^{+}20, VR03]. grip [Nof03]. groenlandicus [FJH10]. gross [SMP^{+}22a]. ground [BPSN^{+}21]. grounds [CPO^{+}19, KKKY10, LML^{+}23]. Group [Ano94k, Ano03j, WHG^{+}16, CSS^{+}21, FAB^{+}09, GKR20, Hof10, SEG^{+}22b, VSGD21]. groups [GDI^{+}09, LSM^{+}22, LRAE23, LMS10, Peñ03b, XWL^{+}18]. Growth [Hey78, LS12, MPD^{+}22, BFB^{+}20, BSF^{+}21, CBHL07, CCS^{+}21, CWS^{+}21, CHS^{+}24, CP02, DIQJ21, GiIKX22, HLR17, HBL^{+}13, HPNDC15, HCGK11, iIRM^{+}15, KSY^{+}19, KSKN21, KLIRK17, Kru19, LMS93, LOG^{+}09, LHP^{+}05, MN88, NKK^{+}05, RvBD^{+}22, SIR^{+}07, SKSK06, SJD10, VMV^{+}23, VDB^{+}20, WPW^{+}14, YHRT22, YHLA^{+}04, YAK13]. grypus [NBLI20]. Guadalquivir [CN22, GEPC15]. Guadalupe [GPEV20]. Guiana [CdTH^{+}16]. Guinea [Ber65b, KLP^{+}17, NM17, Ver92]. Guinée [Ber65b]. Gulf [AAML22, DHD^{+}23, FTG^{+}11, SRFHDH22, TPTM23, ZLRVB24a, ZLRVB24b, AC85, AJHC19, BDB^{+}04, Ber65b, BHPC06, BF12, BD85, CDH^{+}13, CHC^{+}12, CD07, CCH^{+}12, CBT07, DDDT99, DPM^{+}09, EMU^{+}23, FELMGM^{+}22, FSAO22, FFS^{+}20, FTG^{+}18, Fug63, GFGGD^{+}23, GPE^{+}17, GMD^{+}22, HSMLDC^{+}22, HKL^{+}15, Ham87, Ham09, Hen85, HHDS02, Her88, HMS^{+}22, Hog85, JJA^{+}13, JSA^{+}08, Kli10, KLP^{+}17, KAH^{+}16, LR07, LBC^{+}23, LKDL14, LHW^{+}20, LEDR^{+}22, LdCSB^{+}20, LCANAS^{+}07, LCPSMR^{+}10, LS85, MLB^{+}20, MDAW^{+}19, MPSD15, Mil93a, MPN09, MKSW^{+}15, NFMCS^{+}22, NMLBCM^{+}01, NM17, NHN^{+}21, NAH^{+}21, ON05, OHH^{+}22, PPdM^{+}12, PGT^{+}13, PCH08a, PBBH^{+}22, PCR^{+}22, ROBRB^{+}22, RLP^{+}18, RANS65, RGPB^{+}23, RKC^{+}10, SGL^{+}13, SCHBC^{+}22, San15, SGMVF14, SLBVRR^{+}22, STG^{+}18, SBPGP^{+}23, TPTM23, UPPS^{+}21, Ver92, WM13, WDMMK89, WD94, WFBN^{+}13, WW02, WBC^{+}22, WDMC02]. Gulf [WLM^{+}22, WWL^{+}24]. guts [RMB^{+}01]. Gyre [CMJPH^{+}18, DLC^{+}08, GIPC^{+}15, HPC^{+}20, LBSP01, LSMG01, Ang79b, BDTC15, CRGA17, CBB^{+}22c, DMT15, FCMCÁS19, GTNK21, HPB^{+}09, HLM^{+}16, HHB^{+}00, KJH^{+}22, MCD^{+}14, MD07, MFB^{+}84, NCH^{+}07, PAF^{+}11, Rea00, TŠT^{+}17, VBVYT05, Whi95, CM18b, FGR^{+}06, LLL^{+}24, Pre86, RBS^{+}20, RBS^{+}22, SGLF^{+}13, YHM^{+}18]. gyre-scale [Whi95]. Gyres [SPS^{+}99, BLI^{+}99, FK99, HBV^{+}99, HMX^{+}23, HMPZ11, SWT^{+}17, WQ08, ABD^{+}17, SJ02c, SJ02b].

H [BSF95]. Habitat [BCT^{+}09, DFH^{+}16, PGGG17, SJ18, AQVB^{+}10, BMK12, BFP^{+}18, BGM^{+}10, BSC^{+}07, BHH^{+}16, BMG^{+}19, BMG^{+}21a, CF20, DHD^{+}23, DBJ^{+}15, EiIT^{+}22, FGS^{+}23, FKH^{+}13, HLP^{+}16, HSN^{+}18, KKKY10, LVGH^{+}15, LAD^{+}18, LBC^{+}23, LO21, LMP22, MSMR93, MHR^{+}10, MBH^{+}23, MRH^{+}14, MHVS19, NXT^{+}17, PHKS01, PO15, RHBS13, RVC^{+}13, RASVB^{+}22, RDP^{+}21, SPB^{+}12, SAY^{+}16, TJ90, WMB^{+}21, WWW^{+}23, XYL^{+}22, YWUK15]. Habitat-faunal [PGGG17]. habitats [DHHP18, FVA^{+}19, FJH10, JYK^{+}14, KST^{+}10, KOT^{+}21, LPF^{+}18, LBC^{+}15, MCD^{+}14, RLR^{+}18, TNC^{+}09, TTF^{+}22]. habits [KTIT22, SMN^{+}13]. Hadal [JS21, AM19, BC19, BBRM20, CBL^{+}19, GHSC19, HFO90, JGB20, KO19, SJ18]. hake [BWMGCB08, DFM^{+}15, FCN^{+}19, FB05, LC12]. Half [ALV^{+}21, LSW02]. Half-Century [ALV^{+}21]. Halichoerus [NBLI20]. haline [GS24, GCD^{+}13]. halocline [BvdLA^{+}11]. halocyprid [Ang79b]. hamata [MMN12]. hamatus [DK07]. Hampshire [NW87]. Hans [Ano65e, Ano65f]. Haploniscidae [JLRB20]. harbor [RNL^{+}13]. harbour [JHW^{+}14, LAP10, LH08, NHE^{+}13]. hard [DGVGR24]. hard-shelled [DGVGR24]. Hardcover [Bak83, Hof81]. Harmful [IPF23, PP10, BLMR^{+}20, KSC10, KHJ^{+}10, LPHL^{+}05b, MPMFL^{+}23, PFHM10, SPF^{+}23, Sma10a, Sma10b, TPRS10, WDC^{+}11]. harmonics [XD96]. Harp [HBG^{+}21, FJH10]. harpacticoid [KKKS14]. Harpacticoida [GS19, PMG15]. Harpagifer [NPO^{+}19]. hatchery [IOGS13]. hatchery-origin [IOGS13]. Hatchetfishes [EBM^{+}20]. hatching [BF11, IMHL07]. Hato [DFC^{+}21]. Hatteras [GL06]. Hawaii [GDN^{+}18]. Hawaiian [KST^{+}10, LPF23, PCH^{+}08b]. hazard [RAB^{+}11]. head [CVHM^{+}18, DCL^{+}13b, PGGG17]. Heat [BDTC15, YTNK00, AHW99, BDBJ01, DMT15, EMK^{+}17, HWPLvW20, HTdM^{+}15, HDB13, JVJ^{+}17, LLH^{+}21, LDHW20, MMR^{+}09, MMF^{+}17, RKS^{+}15, SCC14, SRFHDH22, SAH^{+}21, WCN^{+}05, WLM07, YN20, ZYN^{+}24]. heating [GW91]. heatwave [WHK23]. Heatwaves [Cia22, ABP^{+}23, FSAO22, HAP^{+}16, OLH^{+}18, SN24, SOWS17, SJLW23, YW22]. Heavy [PCD^{+}18, aHFS92, JP90, XYK^{+}22]. Heavy-metal [PCD^{+}18]. Hebrides [XD96, PHK^{+}17, Rot65, Rot65]. height [CPSM20, DWH^{+}14, HMRB^{+}03, KDL^{+}01, RG09, SRFHDH22, WLD^{+}15]. heights [BBC^{+}22, MZZ^{+}23]. Helena [FHG03, TFM03]. helgolandicus [BRH^{+}05, CÁM06, MAH^{+}15, WPB^{+}08, WSH15]. helical [MJ88]. helium [SW65]. hemipelagic [HKE^{+}10]. Hemisphere [APSC11, DMF^{+}09, GPA^{+}11, HSLG11, KMWF11, MBD^{+}09, ELW06, SE08, SE09, KF11, LPA^{+}11, TAH^{+}11]. Hensen [WB03]. Herald [LPBM17]. herbivorous [DLD^{+}19, FB01]. Herbst [AHRT90]. herring [GBT^{+}19, HRSM08, HPNDC15, iIRM^{+}15, MHTG10, Nag01, PS08, dSSDS^{+}20, STF^{+}13]. hertwigi [vPRT90]. Heterobranchia [CES^{+}19]. Heterogeneities [Mil14]. heterogeneity [DJW^{+}18, DCS^{+}22, DBJ^{+}15, GLV12, HHK^{+}22, IHR18, RLT^{+}22, SJ18]. heterogeneous [DSB^{+}24, DIQJ21]. heteropods [BGWP^{+}17, WPBG^{+}18]. Heterosigma [MPMFL^{+}23]. Heterotrophic [BVB88, VCB^{+}00, BGM^{+}01, BLP93, DLM^{+}12, GASV^{+}09, JP90, MVN^{+}15, MSMH19, SCHS^{+}24, SHS^{+}05, VDDA^{+}08]. hexactinellid [RTN90]. hexafluoride [OYKK^{+}23, TNS^{+}05]. hexapterus [RHBS13]. HF [BBM^{+}14, Pra91]. hidden [DBJ^{+}15]. Hierarchical [KHS^{+}14, HAP^{+}16, OWH14]. High [BJMP19, BJMP20, Ché14, CTR^{+}19, CRC^{+}19, DOS^{+}18, DAF^{+}22a, DAF^{+}22b, HHAR23, JM19, KV13, MLHE23, SDGVE17, SMP^{+}22a, ŠVL^{+}15, SVIA14, vHMDL14, APC13, APN^{+}15, ASJ^{+}23, ACL^{+}18, AASJ23, AE09, BVJE19, BMC05, Ber65a, CKB^{+}17, CGZ^{+}16, CNBD21, DFD23, DPF^{+}20, FM07, FRV^{+}19, FACM^{+}23, FAAV^{+}15, FYYC05, GWS^{+}23, HL05, HBG^{+}21, HDA^{+}16, INO^{+}24, KS06, KSB^{+}22, KMSTK23, LNB13, LGG18, MFDH22, MRW^{+}14, MJD^{+}21, MBD^{+}09, MSL^{+}07, RMK^{+}21, RN06, SVL^{+}23, SKWWGV18, SCC14, STS^{+}12, STM10, SWZS^{+}21, TDGY22, UNN^{+}14, WZBK^{+}21, ZLC^{+}15, ZL24, dPCS23]. high-Arctic [RN06, SWZS^{+}21]. high-energy [ACL^{+}18]. High-frequency [DOS^{+}18, ŠVL^{+}15, FRV^{+}19]. high-latitude [AE09, BVJE19, MBD^{+}09]. high-quality [KS06]. High-resolution [Ché14, SMP^{+}22a, vHMDL14, APC13, APN^{+}15, ASJ^{+}23, AASJ23, BMC05, CGZ^{+}16, CNBD21, DPF^{+}20, FACM^{+}23, FAAV^{+}15, GWS^{+}23, MFDH22, MRW^{+}14, RMK^{+}21, SCC14, dPCS23]. high-salinity [FM07]. high-water [DFD23]. higher [LB20, XD96]. highlight [NFMCS^{+}22]. highlights [GFB^{+}15b, GFB^{+}15a]. highly [AMG^{+}16, BMG^{+}21b, CSS^{+}21, CHC^{+}12, JOGM^{+}10, LHP^{+}05]. highly-stratified [BMG^{+}21b]. hill [MDR20]. hills [DBJ^{+}15, SBG16]. Hindcast [AW13, Cho86, His22, KRL08]. hindcasting [AEPW93]. hints [GGT^{+}15]. Hirudinea [UKK^{+}19]. historic [RN02]. Historical [DLC^{+}08, SJH^{+}90, CQC15, GJ00, PCBA^{+}20, TKWI08, WMWR08, ZCD08]. histories [MT99, MFB^{+}84, Wil87]. History [CGB^{+}23, vdS94b, BVJE19, BBSN04, BLT^{+}08, CMS^{+}13, CBGC^{+}08, ECFT20, Fre07, FMT15, GLAHH^{+}22, GiIKX22, JAC^{+}12, JJS03, Kaw98, KLP^{+}17, MRH^{+}18, MRH^{+}14, MC88, Nie07, RCM^{+}03, Ric08, SCD^{+}07, dSSDS^{+}20, STS^{+}12, SCHD23, SSM^{+}90b, SSW^{+}09, Wüs64, YAK^{+}08]. hiver [Rou65]. Hjálmar [Ano13g]. HNLC [FYYC05]. Hokkaido [DWNN04, Nag01]. hold [BW08]. holistic [PRC^{+}20]. Holocene [HKN^{+}14, JS87, JKBH87, KiL14, KYS^{+}17, Med87, NB87]. holothurians [GBM^{+}01, HWBT03, RMB^{+}01, WHBW03]. Holothuroidea [GKR20, MMK19, WDK^{+}01]. Holyrood [CLMR23]. home [PPPdS20]. homing [BCL^{+}09]. homogenization [Sak86]. hoplonemerteans [CP19]. Horizontal [BGM^{+}10, IAN13, MWJ^{+}08, PBP^{+}99, RBS^{+}20, RBS^{+}22, SVIA14, GSPMAI99]. Hornsund [SPW22]. horse [ZL01]. hosting [MSV^{+}14]. hot [HNSP^{+}19, ZMCD11]. hot-spot [HNSP^{+}19]. hotspot [FDE^{+}22, JHW^{+}14, PIS13, RVS^{+}21, SC23, SNdSR^{+}24, STGR^{+}23]. hotspots [CCRS20, FJA^{+}21, GFB^{+}15b, GFB^{+}15a, GBC^{+}15, PMM^{+}23, SSS^{+}11]. HPLC [SHS^{+}05]. Hudhud [VBM21]. Hudhud-eddy [VBM21]. Hudson [EHG^{+}12, PGGG17]. Human [ALG^{+}21, KSE^{+}09, NCC^{+}15, OvdSN94]. humans [RSB^{+}13]. Humboldt [BDT^{+}08, CRT^{+}22, GCED22, JBB^{+}14, TTB^{+}08a, TTB^{+}08b, ABE^{+}15, AN04, Ant09, BLT^{+}08, CBGC^{+}08, EB08, EBvdL^{+}09, FVLC^{+}23, GDI^{+}09, JTQ^{+}18, LQU07, LPF^{+}18, ML09, NMC^{+}09, RBPGJ^{+}20, RS10, SGF^{+}19, SHD^{+}21, SBG^{+}08, TBW09, VSGC21]. humpback [AHGRAL23]. hundred [RCD^{+}94]. hurricane [KAH^{+}16, LC22, SZG06]. Hurst [NW87]. huxleyi [RPG^{+}18, STHM02]. Hwanghae [Lie86]. hyalinus [GBG05]. hybrid [PVA24]. hydration [Due77]. hydro [PST^{+}15]. hydro-biogeochemical [PST^{+}15]. HydroBase [KS06]. Hydrobates [SWP^{+}13a]. hydrobiological [DBC^{+}18]. hydrocarbons [FTG^{+}18, GGT^{+}15, SGL^{+}17]. Hydrodynamic [CB06, Con87, ACL^{+}18, BFP^{+}18, FG16, HLP^{+}16, HPW10, KSK^{+}15, LL97, TCL^{+}15, WSS15]. hydrodynamics [JHW^{+}14, MRW^{+}14, MHCS^{+}23, TSG^{+}04]. Hydrogen [BCP09, Ros65]. Hydrographic [ALBP87, HMB^{+}86, KHM^{+}88, LBH^{+}87, PHC^{+}19, Yas07c, AMEV07, BDB^{+}04, FW91, GMD^{+}22, GJ00, GA00, JTD^{+}14, Ken88, KDB95, KGJ^{+}10, Lev88, Lie86, LC16, LWY07, Mac98, PFW15, STB^{+}92, Sva65, TTL^{+}04, TZP^{+}00, WXH07, Yas07b]. hydrographical [FPY^{+}16]. Hydrography [FT06, LWT^{+}20, RBR^{+}23, VMA^{+}24, WDMC02, vAB96, BSÖ^{+}94, BBL^{+}09, CCA^{+}02, DPF^{+}20, FMW91, HHY03, HWLT10, MRMD^{+}97, NCH^{+}07, ÖHÜ89, PKF02, SPSR^{+}14, TBK^{+}99, WSL20]. hydrological [CSG^{+}15, KBSB18, PGRP^{+}18, SSVP00]. Hydrology [GCZ^{+}00, GRD^{+}23, Roo82, GCED22]. hydrolytic [CS89, RMB^{+}01]. hydrolyzable [BPP^{+}98, GGA^{+}05]. hydrostatic [RPG^{+}18]. hydrothermal [GSA^{+}20, JP90, JG90, LDH90, LvBS^{+}24, SF85, TJ90, WLP^{+}21]. Hydrothermalism [DSV99]. hydroxychlorophyll [STW^{+}15]. hydroxychlorophyll-a [STW^{+}15]. hyperboreus [DWFP^{+}19, LBC^{+}23]. hyperiid [BTV^{+}17, gWjNfLyD20]. hyperpycnal [HLS^{+}14b]. hypersaline [LEDR^{+}22]. hypotheses [FFA06, Hic79, SPH^{+}15a, VPH^{+}12]. Hypothesis [CS18, BM01, BF11, DTC^{+}06, HMTL05, MPB^{+}23, MCKS17, MZK^{+}23, YYK88]. Hypoxia [SV14, PJS^{+}22, ZHF^{+}24, ZCH^{+}17].

I.G.Y. [McG64]. I03 [KMWF11]. I03/I04 [KMWF11]. I04 [KMWF11]. Iberia [BIST01, KPSB22, QCdS^{+}07, RBD^{+}07, SCD^{+}07]. Iberian [ÁSDB^{+}01, MSd^{+}16, SVL^{+}23, BGM^{+}01, BCOL^{+}19, BAOC^{+}07, CVHM^{+}18, CNBD21, DXH^{+}02, DJG^{+}02, EAL^{+}07, EvdZSH02, FEGA^{+}14, FMH02, JW01a, JRW01, JIT^{+}01, JWD^{+}02, LDB^{+}02, MH02, MMPG07, ORR^{+}02, Peñ24, RCÁS^{+}15, RWOA01, SNV^{+}18, SvWRvB02, SW01, SMGL01, TTMM^{+}17, TMPM^{+}16b, TMPM^{+}16a, TvG02, dlPPÁB24, vWM02a, vWdSBdH02]. IBM [ANMP15, PST^{+}15]. Ice [BHM^{+}15, OPL^{+}21, AAMB^{+}24, BSC^{+}19, Ber65a, BRG^{+}23, CML^{+}16, CGZ^{+}16, DWFP^{+}19, DLD^{+}19, DLD15, DS65, DSV^{+}24, FPJ^{+}15, FACM^{+}23, FMCG15, FJH10, GDL^{+}15, GMR^{+}23, GW89, HKN^{+}14, HJLLN07, Hol00, Iwa23, JCF^{+}23, KTN14, KON14, KMS^{+}24, KSG^{+}17, KSB^{+}22, KGB^{+}23, LVGH^{+}15, LSF^{+}17, LSH^{+}11, LMA^{+}15, LRJ^{+}15, LDHW20, LSW02, MSC^{+}15, MGA^{+}23, MST^{+}23a, MOSN^{+}13, MMD^{+}16, NNFL21, NAH^{+}21, NST^{+}23, ONR^{+}14, RCB^{+}20, RBR^{+}23, RCSA01, Rud89, SMFM^{+}21, SW92, SOO^{+}14, SEW11, SLGI^{+}21, SNMW10, TRY^{+}04, TSFA22, VH09a, VH09b, WDC^{+}11, WO15, WSL20, WBA^{+}22, WC15, YNTS22, YLL19, YGMR^{+}23, ZKK^{+}16]. ice-associated [DLD15]. ice-band [SMFM^{+}21]. ice-free [DS65, WDC^{+}11]. ice-mediated [MMD^{+}16]. ice-sea [CGZ^{+}16]. ice-sheets [Ber65a]. iceberg [WBH15]. icebergs [NST^{+}23]. Iceland [CGV13a, HØH^{+}03, ZLZ^{+}17]. Icelandic [Jón07, CGV13b, GVBV^{+}21]. ICES [SWP^{+}13b, WHG^{+}16]. ICES/GLOBEC [SWP^{+}13b]. ichthyofauna [OV24]. ichthyoplankton [ABP15, DMBB02, JF13, MDAW^{+}19, RMHL09]. idealized [CS06, LL21]. ideas [Rud15]. IDentification [GDN^{+}18, CGG08, MHR^{+}10]. identified [NMN08, YSS14]. identify [RRS03]. Identifying [GKC^{+}14, KFH^{+}15, RHBS13, HRA00, XYWY23]. if [Ano17i, Ano17j, Ano17k]. IFC [Ano13h, Ano13i, Ano13j, Ano13k, Ano14a, Ano14b, Ano14c, Ano14d, Ano14e, Ano14f, Ano14g, Ano14h, Ano14i, Ano14j, Ano14k, Ano15a, Ano15b, Ano15c, Ano15d, Ano15e, Ano15f, Ano15g, Ano15h, Ano15i, Ano15j, Ano15k, Ano15l, Ano16a, Ano16b, Ano16c, Ano16d, Ano16e, Ano16f, Ano16g, Ano16h, Ano16i, Ano16j, Ano17b, Ano17c, Ano17d]. ignition [USH15a]. II [ÁSÁB^{+}14, BBS23, CP19, HMO^{+}13, KSD84, MPCNC^{+}19, McD81b, MAB^{+}11a, NKK03, NF87, SKF20, TTB^{+}08b, Tol85b, WPB^{+}08]. iii [Ano86j, Ano92i, CKT^{+}13, MAB^{+}11b, Sim84]. iii-xv [Ano86j]. illuminate [CRS04]. illus [Bak83]. illustrations [Ang80]. image [Ano17i, Ano17j, Ano17k, GPAB^{+}16, MERB12]. imagery [CN22, OP18, RRS03, VOT^{+}99]. images [CDB^{+}24, SW92]. imaging [CTP^{+}18, YNMY23]. Imbalance [RSB^{+}01, DHB^{+}21]. Immature [CBB^{+}22c]. Impact [AdAK^{+}18, CZG^{+}21, FDB^{+}21, GCB^{+}22, IIM^{+}23, KNSN^{+}09, LBNBM13, MXC^{+}21, MHS^{+}20a, MHS^{+}20b, MK12, OvdSN94, PPSVC^{+}13, REG^{+}15, SKT01, TMR^{+}21, VTGC19, VNH^{+}23, WBA^{+}22, WW02, ZDM^{+}20, ASFB^{+}13, AH15, BWMGCB08, BRE^{+}24, BF12, CRPS^{+}15, DAvD^{+}21, DHB^{+}21, EBR^{+}14, FB01, FGL^{+}23, HKL^{+}15, HPS^{+}01, HZD^{+}23, HHR^{+}19, HKPV12, HSL96, KTN14, Kat18, KKNT23, KAG^{+}19, KGJ^{+}10, MAB^{+}11b, MJC^{+}17, NDEG22, NRA17, PTM^{+}22, RHB23, RHM^{+}19, SLBR18, SCHBC^{+}22, SNR^{+}10, SCS^{+}18, SAB^{+}22, SST^{+}17, SPB93, TKSI08, TTB^{+}08b, VBL04, XYL^{+}22, YYT^{+}14, ZKK^{+}16]. impacted [MGC^{+}18, PBB^{+}20]. impacting [CGD^{+}18, HPNDC15, SBC^{+}24]. Impacts [BDTC15, BDT^{+}08, GPEV20, HLFL23, HDA^{+}16, INI^{+}17, KKKY10, KMS^{+}24, LM10, RCB^{+}20, XLL^{+}20, ALG^{+}21, CDTM^{+}21, CTA16, CZL^{+}24, CWS^{+}21, DNNNN16, DAIS10, GMD^{+}22, HMWM00, HRSM08, HWLT10, HSC^{+}16, HBW17, HWF^{+}21, KS15, LSH^{+}22, MVN^{+}15, McI10, MBH^{+}01, MS00, MVV^{+}19, MST^{+}23b, OHH^{+}22, PFHM16, PL01, PG10, Pra97, RDD^{+}18, RCSA01, RAB^{+}11, RKC^{+}10, SBM^{+}23, TNGP22, TKW06, TPRS10, THM^{+}14, UBB^{+}23, WSL20, WJPHB15, YZX^{+}23, ZLKO00]. impaired [dJSL^{+}20]. impedes [MLHE23]. imperative [KN10, KN11]. imperfect [BECA22]. implement [CP10]. Implementation [MPC12, SJP10]. implemented [BFPS06]. implication [AYH^{+}23, FZY^{+}23, Gam14, STHM02, TvG02]. Implications [BSA06, CCS^{+}21, Gal17, INI^{+}17, IHT^{+}21, LLL^{+}11, Law04, LRGV^{+}18, NGPH10, OBD^{+}20, OOTA15, Ric01, RS10, RLR^{+}18, SSS^{+}11, SFS^{+}12, SBH^{+}14, SMN^{+}14, THM^{+}14, Ang89, BHM^{+}15, BEP02, BD18, CB06, CGB^{+}23, DVB^{+}18, GAF15, HRSM08, IHY^{+}01, JC04, JCF^{+}23, JG90, LL21, Mor91, PWZ^{+}16, Reb02, RGM01, RCSHW22, SJ18, SPW22, TWBC^{+}13, TCF^{+}18, VB14, WWL^{+}22, WBH15, WL16, ZL01, ZHD^{+}20]. implied [CBB^{+}22a]. Importance [HLK13, INO^{+}24, She65, BCR^{+}13, CLSP17, CMC^{+}16, CLB^{+}13, DWC06, HFS^{+}20, LZF^{+}24, LM14, MDC^{+}07, MCH^{+}12, RGC^{+}01, SRT^{+}18]. important [KFH^{+}15, NXT^{+}17, SFS^{+}12, SSV^{+}11, VBA^{+}18]. improve [CTMV^{+}14, MAFS^{+}22]. Improved [HHWW20, TAM^{+}13, RAB^{+}11]. Improvement [PPVG12]. Improving [CDB^{+}24, KM22, PRC^{+}20, HKL^{+}15]. in- [Mil14, Sek88]. In-flight [Sac16]. in-flow [CTP^{+}18]. in-situ [BGR^{+}15, SNS^{+}22, NIF^{+}15]. incidence [YHRT22]. incidental [HYM^{+}12]. incised [JFEC13]. Including [MCH^{+}12, BFP^{+}18, PMG15, TGJT09, ZHSMM14]. inclusive [CSH^{+}23]. Incoherence [CGW^{+}22]. Incoherent [CZW^{+}22]. incoming [ZKT88]. Incongruous [Wil87]. Incorporating [Arb22, KMB01, FPS^{+}13, PRC^{+}20]. increase [CKB^{+}17, Cra09, LAP10, SDGVE17]. Increased [SPW22, DBJ^{+}15]. Increases [Woo18, AvD15, HLTB^{+}17, PPSV^{+}18]. increasing [RPG^{+}18, VFS^{+}15]. increment [ATT^{+}08]. incubation [SFAD^{+}90]. Incursion [ZJZ^{+}21]. Independence [TWMY08]. independent [CDDF11]. Index [FHG03, SMG02, Ano63b, Ano63d, Ano64c, Ano64e, Ano65a, Ano65b, Ano65j, Ano65k, Ano69a, Ano69d, Ano73a, Ano73e, Ano85b, Ano85k, Ano86a, Ano87i, Ano89l, Ano90d, Ano90e, Ano92a, Ano92j, Ano93g, BCGN^{+}18, CWS^{+}21, KC15, PM13, SMN^{+}14, dMM69]. India [JJA^{+}13, JJA^{+}17]. Indian [Ang79a, ABP^{+}23, JHM^{+}22, KRL08, PFE10, PAF^{+}11, RHB23, SBC^{+}16, SMP07, ZQWP23, AMY^{+}23, ATC^{+}19, Ano94c, CVBG21, CGS23, CRF^{+}10, DAvD^{+}21, Fai65, GCV^{+}24, GCS91, HGT16, HBW17, ICB^{+}19, KS06, LvBS^{+}24, LSW^{+}21, MNS^{+}24, MST^{+}23a, MKM93, MMF^{+}07, MYH^{+}22, MBKS08, MFA^{+}15, NST^{+}23, PTP^{+}22, PS23, Rei03, RGE22, SM01, SDS02, SDS22a, SVU02, SAM^{+}10, SW22, SSN23, SR15, SKCP23, VSA^{+}21, VVV21, VMA^{+}24, Wis65, XWW^{+}21]. Indian-Ocean [MYH^{+}22]. indicate [ECFT20]. indicates [dJSL^{+}20]. indications [AHRT90, WZ04]. Indicator [RCM^{+}03, HSMLDC^{+}22, LML^{+}23, SON^{+}20]. indicators [Don65, DPH^{+}18, GRDS10, PL09, SFF^{+}24, SKGS20, STGR^{+}14, TSS^{+}12, WFD^{+}07, YFK21]. indices [DBR03, GRLS14, GSVB23, HPW10, PCSMC12, SW22, TNC^{+}09, TCL20, WPH^{+}10, YHLA^{+}04]. indigenous [SL13]. indirect [LGK^{+}93]. indirectly [Szu12]. Individual [BEP02, BSH^{+}20, BLT^{+}08, BAP^{+}22, CMS^{+}13, DSBP15, GBH^{+}20, JAC^{+}12, PGS^{+}22, SK17, TMR^{+}21, WPB05]. Individual-based [BEP02, BSH^{+}20, BLT^{+}08, CMS^{+}13, DSBP15, GBH^{+}20, JAC^{+}12, PGS^{+}22, WPB05].

Indonesia [SNR^{+}10]. Indonesian [MMF^{+}07, PCBA^{+}20]. induce [LPHL^{+}05a]. induced [BBB^{+}14, BHPC06, CDH^{+}13, CLB^{+}14, CKL^{+}14, Dav85, FLDF22, GML^{+}23, HZCZ16, HE07, HCGK11, KKB00, LOO22, MLL^{+}22, MMGL^{+}07, NTU^{+}14, NNM^{+}21, OC06, PPSVC^{+}13, SZG06, Szu12, TSAM^{+}22, TS10, VBM21, XD96, YHLA^{+}04, Yux88]. induces [DBR20]. Industrial [CBD^{+}24]. Industrialised [Ang80]. inermis [CNT^{+}19]. Inertial [Hen73, HLFL23, IAN13, KWI20, LZL^{+}22, Lie88, LCZ^{+}24, PTM^{+}22, SSL07, WLCG23]. infaunal [IVR^{+}13]. infer [TSS^{+}12]. inference [Egb97, HMX^{+}23]. inferences [OPG^{+}10]. inferred [ADS^{+}22, BSF95, CDDF11, FCN^{+}19, HLSX22, OMS^{+}09, OÁT^{+}05, PPHM18, PCBA^{+}20, SBC^{+}16, SOB^{+}08, SCC14, TBS^{+}19]. inflated [HLM^{+}16]. inflow [GHF^{+}21, HØH^{+}03, MM01, MIN^{+}20, NBHM01, RBR^{+}23, RKS^{+}15, Woo18, ZBRJ23]. Influence [AGS10, AJHC19, AG22, CSV^{+}07, CPG08, DOP87, FVLC^{+}23, GdRGC^{+}14, KM10, KTW^{+}22, Lav09, MJA^{+}07, NKK^{+}05, PFE10, ST03, SHd13, SJJ^{+}03, SBG16, SJD10, SLH^{+}19, WCX^{+}21, WPB^{+}08, XNT^{+}17, YAK^{+}08, AR18, BRG^{+}23, CLSD18, CGW^{+}22, CPPPEAG22, CdTH^{+}16, CS03, CBD^{+}24, Cra09, Dri11, GSVB23, GGJ^{+}10, GTS^{+}21, GiIKX22, HWL^{+}20, JHW^{+}14, KAK^{+}22b, LSXT01, LPF^{+}18, MBP^{+}11, MHTG10, MRBS^{+}24, MHCR^{+}12, NGNV12, Ore69, PJS^{+}22, PLEF^{+}23, QOS^{+}22, RBNJ^{+}12, RHB23, RBHLA04, SLM^{+}16, SMR^{+}20, STEB16, Sek86, SNZ^{+}20, SCC^{+}19, SBD^{+}07, SK21, SPN98, TSG^{+}04, TDK^{+}16, ZPY^{+}20, ZZWL06, vRGW10]. influenced [BEP02, GCFS06, IVT^{+}12, RFSCF19, SM05]. influences [BPF06, BMG13, BECR^{+}22, CSG^{+}15, HEF^{+}12, MVBC^{+}21, PTP^{+}22, Qiu15, SWZS^{+}21, WHS^{+}23]. influencing [DCD^{+}23, LZCZ05, SNMW10, XD96, vHMDL14]. info [Ano13h, Ano13i, Ano13j, Ano13k, Ano14a, Ano14b, Ano14c, Ano14d, Ano14e, Ano14f, Ano14g, Ano14h, Ano14i, Ano14j, Ano14k, Ano15a, Ano15b, Ano15c, Ano15d, Ano15e, Ano15f, Ano15g, Ano15h, Ano15i, Ano15j, Ano15k, Ano15l, Ano16a, Ano16b, Ano16c, Ano16d, Ano16e, Ano16f, Ano16g, Ano16h, Ano16i, Ano16j, Ano17b, Ano17c, Ano17d]. inform [FACM^{+}23]. Information [Mol22, AGD^{+}18, BTNK13, BBE^{+}03, BMN19, BDE03]. informed [WPA^{+}24]. infragravity [LH08]. infrared [VNMS91]. infusion [RG03b]. ingestion [BD20]. inhabiting [GVBV^{+}21]. Inherent [LC10, SRG^{+}19, HHMB^{+}09, WMB^{+}18]. initial [AHGRAL23, HHMB^{+}09]. initiation [KFM15, RHM^{+}19]. Initiative [MP13]. Inland [YFY^{+}22, Ber65a]. Inlets [TSC03]. Inner [FCN^{+}19, BD18, CAB^{+}18, CP02, MG02, PHC^{+}19, TNC^{+}09]. inner-shelf [TNC^{+}09]. Inorganic [FWBC02, JCF^{+}23, DGH^{+}20, HLR17, Kit03, KFC^{+}23, RBR^{+}23, SRF^{+}19, SAM^{+}04, ZSY^{+}22]. input [XY20]. inputs [aHFS92, MTC14, OAD22, PBB^{+}20, TSRF14]. INSAT [SNS^{+}22]. INSAT-3D [SNS^{+}22]. Inshore [CSG^{+}15, BHAJ12, KLC^{+}15]. Insight [AIHB^{+}07, ALM^{+}23, QLY^{+}22, SF02, Ste12, YTL^{+}19]. Insights [ATC^{+}19, BK19, DOS^{+}18, HPZC21, LSV14, MPMFL^{+}23, McI10, PPCWJ18, SMN^{+}13, VBAC^{+}21, YNMY23, YSY^{+}19, Yas07a, AEP^{+}23, ÁSFP^{+}03, BRB^{+}01, CS16, DFD23, DDK^{+}18, GBT^{+}19, MSd^{+}16, NBHM01, SSKA19, Sch03, Woo18, CDH^{+}13]. instabilities [SÖÜ94b]. Instability [WLL06, Hog85, VMN08]. Institutional [MCB^{+}10]. instrument [Dah69]. Instrumental [Szu12, Jer65]. instrumentation [SFAD^{+}90]. instrumented [NBLI20, PHC^{+}19]. insular [CMM^{+}04, GIPG17]. Integrated [CCM^{+}13, BKC15, DPH^{+}18, DPR^{+}18, FMP19, HFS^{+}20, HMO^{+}13, JHDT12, KSE^{+}09, MCH^{+}12, PM13, PHD^{+}18, PV18, Val99b, WPH^{+}10]. Integration [RKC^{+}10, BMB^{+}16, CFM^{+}18, JTGM10]. Integrative [Ano09h, FBA09, JGB20, HAA^{+}14, JZ19, MCB^{+}10, SMN^{+}13]. integrators [ESTM13]. Intense [ZCLS20, BBB^{+}14, MPN09, PPdM^{+}12, RCC^{+}18]. intensified [MKOLA20]. intensifies [PLN^{+}23]. intensity [AE09, BAOC^{+}09, FP03, GvOS^{+}08, LFI^{+}13, MK12, PLK14, ST03, Whe06]. Inter [DBM17, FCN^{+}19, HS07, LMPB^{+}16, MTC12, SNZ^{+}20, AMFY20, AMG^{+}16, BCOL^{+}19, CLL^{+}18, DZ04, EHG^{+}07, FBS^{+}18, GLLB22, HDZY15, LW13, MB20, MCGR07, NRA17, OACB^{+}15, PVA24, dlPHF^{+}15]. Inter-annual [DBM17, FCN^{+}19, LMPB^{+}16, MTC12, SNZ^{+}20, AMFY20, AMG^{+}16, BCOL^{+}19, CLL^{+}18, DZ04, EHG^{+}07, GLLB22, HDZY15, LW13, MCGR07, NRA17, OACB^{+}15]. Inter-basin [HS07, dlPHF^{+}15]. inter-connected [FBS^{+}18]. inter-ocean [MB20]. Interaction [MSd^{+}16, KBE^{+}22, KAH^{+}16, LC22, LSW^{+}21, SMFM^{+}21]. Interactions [MMF^{+}07, PPY87, Val99a, AUE^{+}14, BLP^{+}20, CDP14, DRVMC^{+}22, Due77, Fly03, GCG^{+}14, HBG^{+}21, HBK^{+}24, MPM^{+}18, Pie01, RBL90, SCD^{+}07, WLKM10, YLL19]. Interannual [AAMB^{+}24, AUE^{+}14, BGMP03, DWNN04, DCS^{+}22, ESTM^{+}12, GFGGD^{+}23, IAFD02, LAA12, LBC^{+}23, LBSP01, MPV12, Min02, MWO^{+}12, NNM^{+}21, PELAA18, PO00, RBE^{+}12, SiSI^{+}02, TBW09, Wu13, YNM^{+}02, BMM01, BPM^{+}14, CHC^{+}12, CD07, CRHM12, FDHT05, GdRGL^{+}01, GA10, GCED22, GTS^{+}21, HBD^{+}18, IIS^{+}17, Iwa23, JAS^{+}20, KST^{+}10, Kat18, KSK^{+}15, LS12, LBD11, MZH^{+}23, MT99, MM99, OLH^{+}18, PV07, PFW15, PD15, PTI00, RWD01, RDP^{+}21, SLM^{+}16, SAM^{+}10, SAB^{+}21, Spr08, SJM^{+}19, TBS^{+}19, XWW^{+}21, YIY^{+}04, YKS^{+}12, YBPS08, PLJR22]. Interannual- [BGMP03]. interbasin [EMK^{+}17]. Intercalibration [JF13, TSG^{+}04]. Intercomparison [BSFM^{+}12, NH88b, SWP^{+}13b, BEI^{+}20, BRB^{+}01, BDBJ01, NP00]. Interdecadal [FBM^{+}08, MNM06, MS00, MM99, MJWK07, RGM01, GCED22, Min02, PO00]. interdisciplinary [WH89]. interface [CDS90, CEF^{+}13, SvWRvB02, SFAD^{+}90, VPW01]. interglacial [Bro82, DPB06, HM98]. interior [HMKF08, ÖÜT93, TSRF14, WFR07, WC15, SDK84]. interleaving [Rud03]. Intermediate [EMU21, MWJ^{+}08, NIC^{+}19, PL18, ZSI^{+}05, BLC23, CS18, DL17, HVS10, HGD22, KON14, LF12, MFS^{+}16a, MFS^{+}16b, MSGGM18, YTB^{+}21]. intermittent [GMR^{+}23, SSB14]. Internal [AHD18, HNSP^{+}19, TYO^{+}14, VOT^{+}99, ZZPL18, vHVAT22, Arb22, BPSGP^{+}23, CGW^{+}22, CZW^{+}22, ESA^{+}13, GXX^{+}22, GLPC23, GCG^{+}14, GC14, Hey78, HLFL23, HHZ^{+}22, HYM^{+}24, HWF^{+}21, ITO^{+}14, JFUR20, KWI20, Kra69, Li14, LCZ^{+}24, MZZ^{+}23, NP00, PM85, RM97, Ric94, RHB23, RPSVLS14, SMFM^{+}21, SNdSR^{+}24, WGM^{+}24, XDG^{+}23, XD96, XHC^{+}20, YHZ^{+}22, ZCA21, ZQWP23]. internal-wave [HWF^{+}21]. International [JLS^{+}22, LM10, BPW10, KY15, ZNR^{+}24, MHA^{+}11, Rud15]. Internet [CAO^{+}20]. interpentadal [MW96]. Interplay [CTI^{+}19, HM06, CRT^{+}22, RGPB^{+}23, RA15]. interpolation [Egb97, MPSS91, SBM91]. interpretation [GAM98b, McD81b]. interrupted [KO19]. interspecific [MDR20]. intertidal [BRR^{+}22, BWB^{+}09]. interzonal [KST03]. Intra [FKZ^{+}15, RCS^{+}11, SGS^{+}23, SGR^{+}22]. Intra-annual [SGS^{+}23, SGR^{+}22]. Intra-regional [RCS^{+}11]. Intra-seasonal [FKZ^{+}15]. intracellular [BPA^{+}21]. Intraseasonal [BPSN^{+}21, RLX^{+}24, QLY^{+}22]. intricate [SBMB18]. Intrinsic [CD07, CPSM20, PMA^{+}14]. Introducing [GLF^{+}17]. Introduction [Ano65g, Ano10a, BPW10, CBB^{+}02, CB91, FL06, ITM86, JM88, LK13, LM10, O’B83, PAB^{+}87a, PBH^{+}10, SRM^{+}10, SBB^{+}14, Tak88, TT05, AB00, BR01, CH07b, HMWM00, JW01a, MBH^{+}01, Peñ03b, RG03a, RG03b, SSM^{+}90b, SB69, TP00, Was06, vWM02a, RAB^{+}84]. introductory [Ang79a]. Intrusion [ZSY^{+}22, BLR^{+}23b, NXY15, WST^{+}21, YYC^{+}18, YYhT^{+}17, YMK^{+}04, YLY^{+}14]. intrusions [ÖÜT93, RK03b, RR03, TG81, VKT15, YN03a]. inundation [CLB^{+}14]. invasion [RSB^{+}15]. invasive [FVA^{+}19, YFK21]. inventories [KKS^{+}03]. inventory [CPG^{+}18, YT06]. Inverse [VBL04, APHGC^{+}22, Ben85, FTG^{+}11, HNL14, Suk88, SMP07]. inversion [BGV^{+}23, Egb97]. invertebrate [AIA^{+}15, BJ90, QCdS^{+}07, RNBP^{+}19, SLPA^{+}20, Wil87]. invertebrates [ALT10, KO19, MKSvA^{+}22, PPHM18, Sok90]. inverted [ZZPL18, ZYN^{+}24]. investigate [LAHI10]. investigated [TFZS14]. Investigating [LOBG^{+}10, BN03, TSJC07, VBL04]. Investigation [BDP^{+}06, BFPS06, LvBS^{+}24, PFE10, ZCH^{+}17, dLLdAWL^{+}23, Ano94c, BH07, GLPC23, JX18, MKM93, Sei63, TCN20, WCX^{+}21]. investigations [KFM15, WTT14]. IOD [LSW^{+}21]. IOD-ENSO [LSW^{+}21]. iodine [TBW00, ZSY^{+}22]. ion [Due77]. ion-water [Due77]. Ionian [SK21, KK20, MRMD^{+}97]. ions [CS89]. IPCC [LSS^{+}10, SYB^{+}15, SAB^{+}22]. IPCC-class [SAB^{+}22]. Ireland [CJMO87, PMM^{+}23]. Irish [DJ92, HHB^{+}00, LML^{+}23]. Irminger [HRA^{+}08, PM22, SIR^{+}07]. Iron [AYK^{+}05, DTKvH15, KTN14, KNI^{+}05, NHN^{+}21, Qiu15, TRP^{+}23, AIHB^{+}07, Ban96, BB14, FYYC05, KIS^{+}05, NNO^{+}14, NKK^{+}05, Peñ03a, RTF^{+}05, SSH^{+}05, SWT^{+}17, SHS^{+}05, TT05, TSNO05, TKK^{+}05, TNS^{+}05, UNN^{+}14, VMN08, WCC^{+}20, YFY05, ZCD08]. iron-depleted [AIHB^{+}07]. iron-enrichment [TT05, TKK^{+}05]. iron-fertilization [TSNO05]. iron-fertilized [TNS^{+}05]. iron-replete [AIHB^{+}07]. irradiance [Ban96, BD18]. Irregular [Par65]. irregularity [OAM00]. irrigation [PMDR06]. ISBN [Ang79a, Ang88, Bak83, Hof81, SW81]. Ischnomesidae [BBFS19]. Isfjorden [SON^{+}20]. Isidella [PRA^{+}18]. Island [DWNN04, SCB^{+}16, VSGD21, GDN^{+}18, RM97, YHLA^{+}04, FHG03, KMOM88, MFM15, WLKM10]. island-induced [YHLA^{+}04]. Islands [CCRS20, PMC21, TSFA22, PMG15, AGL^{+}15, BAM^{+}09, BA04, BATNP04, CPG08, DMBHG10, FWO15, GASV^{+}09, HMB^{+}86, PVB23, RZW^{+}23, TOiF^{+}12, TMH^{+}16]. Isles [HHWW20]. isoirradiance [LRAE23]. isolated [JP90]. isopod [RK20]. Isopoda [BBFS19, MB20, GM19, JLRB20]. isopods [BLES16, CF20, JGB20, SBS90]. isopycnal [CM18b, KY23, RLC85, Tom81b]. isoscape [HOY^{+}21b]. Isoscapes [OMK^{+}22, ALM^{+}23, DSB^{+}24, MRA^{+}19, OPH^{+}24]. isothermal [Due77]. Isotope [CSG^{+}15, AYH^{+}23, BAOC^{+}07, CSK^{+}12, CDP14, EiIT^{+}22, GWM^{+}22, IBW^{+}01, Kli10, KSS^{+}23, KSG^{+}17, KAAK^{+}16, LSV14, MRAP22, OPG^{+}10, PPHM18, QOS^{+}22, SF02, SM16, VKGP^{+}13, YAI^{+}14]. isotopes [BGA^{+}21, CGC^{+}20, ESTM13, GSPP^{+}20, LvBS^{+}24, LCJ^{+}17, PYKF15, SBC^{+}16, SHC^{+}06, SHC^{+}07, WST^{+}21]. Isotopic [WWL^{+}22, WRS^{+}92, WLP^{+}21, CSC^{+}12, DDCE^{+}23, ESTM^{+}12, FMC^{+}15, GDM^{+}20, LCBN14, NAH^{+}21, Ola65b, OPH^{+}24]. Issue [BDB^{+}22, LM10, SBB^{+}14, FJhT^{+}14, MSI17, RW97, SGPdM18]. issues [ORMB08, Zav99]. Istiophorus [RLL^{+}09]. ISW [HYM^{+}24]. ISW-D [HYM^{+}24]. Italia [CPB^{+}15]. Italian [Sel65]. Italy [Pir87, CBB^{+}19, CBB^{+}22b]. itself [UAM05]. IV [Hau84]. Ivanovich [Ano20u].

J [Ang80, Ang88]. J. [Dea85]. jack [ABE^{+}15, ACN01, BHH^{+}16, EiIT^{+}22]. James [Ang80]. January [GSPMAI99, KSPK99, TBK^{+}99, Ano98a, Ano99b, Ano00a, Ano02b, Ano03c, Ano04c, Ano05d, Ano06a, Ano07k, Ano08o, Ano13p, Ano14p, Ano15o, Ano16m, Ano18k, Ano19k, Ano20n, Ano21o, Ano22v, Ano22-27, Ano22-28, Ano23o, Ano24i, KDB95]. Japan [hHCK01, KKK^{+}04b, LXC^{+}22, AB90, CTL^{+}04, CS03, DWNN04, DZ04, Gam14, HFO90, Igu04, Ike88, IHY^{+}01, IMW^{+}14, JJL^{+}19, KKK04a, KKS^{+}03, KWI20, KKKS14, KNS^{+}03, KOT^{+}21, KTIT22, MGKW19, MLD^{+}03, NEI^{+}22, NH88a, NMK^{+}03, NKK03, Nof00, SN24, Sek86, Sek88, SS03, Soh03, Sud86, TMN^{+}12, TTL^{+}04, TKW06, TKWI08, iUMY86, WGM^{+}24, WHT86, YIY^{+}04, YNMY23, YYK88, YPM^{+}10, YFY^{+}22, YAK13, ZLS^{+}04, YJW88]. Japan/East [CS03, TKW06, TKWI08, YIY^{+}04, ZLS^{+}04]. Japanese [CLB^{+}14, SKH^{+}23, YFY^{+}22]. japonica [GC09]. japonicus [GiIKX22, HSGJ23, TMÁGC^{+}21, ZL01]. jashnovi [NMN08]. Jay [GWGR^{+}19]. JCOPE2 [CMG15]. jellyfish [BDC^{+}08, CRT^{+}22, DBC^{+}23, RBE^{+}12]. Jersey [JKBH87]. Jerusalem [NF87]. Jervis [TSC03]. JES [YIY^{+}04]. jet [Kos02]. jets [RFPG15]. JGOFS [CBM^{+}21, JvdLL^{+}15, MB05]. Jiaozhou [LZCZ05]. Johanssonia [UKK^{+}19]. John [Ano85a]. Jones [DGP^{+}13, EMBS13, LTSG13, MEST13, PIS13]. Juan [MPM^{+}18, SZG06]. Juby [ABT^{+}04]. Julian [RLX^{+}24]. July [Ano03a, Ano04d, Ano05c, Ano06d, Ano07m, Ano08n, Ano09i, Ano20o, Ano21p, Ano22w, Ano22-29, Ano22-30, Ano23p, Ano24j, YYC^{+}18]. jumbo [ATT^{+}08, BGM^{+}10]. June [Ano09h, Ano22-31, Ano22-32, YJW88, Ano98d, Ano10k, Ano14q, Ano15m, Ano16q, Ano17q, Ano20p, Ano21q, Ano22x, Ano23q, Ano24k, HBV^{+}10, KTB^{+}99, KMU^{+}12, SLBVRR^{+}22]. juvenile [DIQJ21, EKB06, EiIT^{+}22, FMM^{+}20, GA10, LMM03, SBL^{+}23, YGL^{+}10].

Kalmaegi [WZC20]. Kalman [MZGA^{+}20]. Kamchatka [AM19, AP20, BBFS19, BC19, BBRM20, CES^{+}19, CP19, CBL^{+}19, FTHK19, GKR20, GM19, GHSC19, JPB20, JGB20, Kam19, KCBS20, KKS^{+}19, MA20, MB20, MDG^{+}19, NHN^{+}21, SSKA19, SKF20, SPB19, UKK^{+}19, YTL^{+}19]. Kara [HKGH^{+}06]. Karenia [BTS^{+}15a, BK19, VMV^{+}23]. Kashevarov [RCSA01]. Katsuwonus [KAK^{+}22a, PGS^{+}22]. kelp [BD18]. Kelvin [BDT^{+}08]. Kerch [KAG^{+}19]. Kerguelen [BCLD^{+}17, DMBHG10, PWZ^{+}16]. kernels [HPW10]. keta [SKSK06, YWUK15]. Key [MST^{+}23b, AHSS22, Ang79a, BGM^{+}99, CAO^{+}20, DP18, GDI^{+}09, GBH^{+}20, MTK^{+}22, TCF^{+}18, HPNDC15]. Kharlamenko [Ano20u]. kill [VWDF14]. kill-the-winner [VWDF14]. killing [MPMFL^{+}23]. kinetic [LZL^{+}22, LBD11]. kinetics [GTR01]. King [STC10, BMN19, STEB16]. Kinorhyncha [AM19]. kisutch [AHC^{+}13]. kittiwakes [WFD^{+}07]. knot [HPNDC15]. know [SMKK21]. Knowledge [JPM^{+}08, OELP04, Don87, Eme65, GGT^{+}15]. Knowledge-based [JPM^{+}08]. known [FTHK19]. Knudsen [BBF^{+}22]. kochi [BLCL14]. Kongsfjorden [AKH^{+}23, HFPS^{+}06]. Konrad [SW81]. Korea [BC88, FY88, Lie88, MWO^{+}12, Suk88, YMK^{+}04]. Korean [KKB00, MK12, RK03a, YN20, YMK^{+}04, ZLKO00, ZL01]. Korteweg [NP00]. Krill [OACA20, AMY^{+}23, BLP^{+}20, DSBP15, IIM^{+}23, LAA12, LS15, MPSD15, OTNI20, RBPGJ^{+}20, SSS^{+}11, SFS^{+}12, TNGP22]. Krithe [YTL^{+}19]. Kröyer [CCG07, AHRT90, Ver91]. KuramBio [CP19, SKF20]. Kuril [AM19, AS20, AP20, BBFS19, BC19, BBRM20, CES^{+}19, CP19, CBL^{+}19, FTHK19, GKR20, GM19, GHSC19, JPB20, JGB20, Kam19, KCBS20, KKS^{+}19, MA20, MB20, MDG^{+}19, NTU^{+}14, SSKA19, SPB19, UKK^{+}19, YTL^{+}19]. Kurile [SKF20]. Kuroshio [ALV^{+}21, CGW^{+}22, CMG15, CZW^{+}22, HSH97, JC88, Kaw98, KSS^{+}23, LMC^{+}20, MZH^{+}23, NXY15, QNK^{+}22, RI86, STJ^{+}14, TMN^{+}12, TSJ^{+}12, WCX^{+}21, WST^{+}21, WWZ19, YYC^{+}18, YJ88, YTNK00, YLY^{+}14, YYhT^{+}17, ZLR^{+}07, ZZM^{+}24, ZLC^{+}15, ZSY^{+}22, YJW88]. kyr [AVG^{+}19, SGF^{+}19]. Kyushu [Ike88].

L [WR03]. L4 [TAW^{+}15]. labeled [SCLG^{+}11]. labile [ÁSDB^{+}01, DMD^{+}00, DDD^{+}00, FPD^{+}01, KGdS^{+}08, TRMV15]. lability [BHHR15]. Laboratory [Rud03, FHL^{+}24, Sak86, Sch03, VPS09, YN03a, McD81b]. Labrador [HMP^{+}13, CGV13b, DCL^{+}13a, FPY^{+}16, HBL^{+}13, HHY03, HMP^{+}13, KY15, LRS^{+}03, LNB13, LWY07, LBD11, MJWK07, PCM11, TDGY22, Yas07b, Yas07c]. Labrador/Newfoundland [HMP^{+}13, HBL^{+}13]. Lack [JMSB^{+}23, WSH15]. lagged [dSSDS^{+}20]. Lagoon [CKL^{+}14, AR18, RTBR^{+}22, YAK13, Pir87]. Lagrangian [AKAL20, AVS23, BIST01, BBM^{+}14, FB01, GL06, HCAFD^{+}20, HMS^{+}22, JLB^{+}08, JW01a, JIT^{+}01, LaC08, LOG^{+}09, Men21, ORW^{+}01, Pra22, SMGL01, Woo05]. Laila [MMR^{+}12]. lakes [SvN04]. Laminated [SGO^{+}08]. lampfish [Kli10]. lance [RHBS13]. Land [AYH^{+}23, GCD^{+}13, DSV^{+}24, GHVG19]. land- [GHVG19]. land-fast [DSV^{+}24]. landers [TDH^{+}95]. landfall [MK12]. landscape [MRH^{+}18, SLBH^{+}19]. Lang [PMG15]. lanternfishes [EBM^{+}21]. Laptev [SPG^{+}06]. Large [ASB^{+}08, BCS09, Bel09, GFB^{+}15b, GFB^{+}15a, HPB^{+}09, KSE^{+}09, LBP15, RM89, SH09, TG05, YSY^{+}19, ALM^{+}23, Bak01, CNT^{+}19, CWZ^{+}20, CGC^{+}20, CRT^{+}22, CRF^{+}10, DKRL22, DHL^{+}21, DTKvH15, ELW06, EBW^{+}23, FBB^{+}21, FK86, Har05a, Hob10, HWB^{+}18, IPD14, ITO^{+}14, Kaz17, KGL22, KT04, LMS10, LMH^{+}13, LFBP^{+}13, MSMR93, Man04, MFB^{+}84, MN88, MC88, Mil93a, MSA^{+}22, NBLI20, PK02, PZA^{+}15, RFFL21, RLGC10, RFS10, RPSC22, SSL08, SJ02c, SJ02b, SMP07, TMN^{+}12, YGC^{+}21, ZWM^{+}15]. large-amplitude [ITO^{+}14]. Large-scale [ASB^{+}08, GFB^{+}15b, GFB^{+}15a, LBP15, SH09, YSY^{+}19, ALM^{+}23, CNT^{+}19, FBB^{+}21, HWB^{+}18, IPD14, Kaz17, LMH^{+}13, MSA^{+}22, PZA^{+}15, RFFL21, RPSC22, SJ02b]. large-sized [KT04]. larva [PKA19]. larvae [CZG^{+}21, DMC^{+}18, DTG^{+}24, GGJ^{+}10, HLS^{+}14a, hHRW^{+}05, HCGK11, KSS^{+}23, KTIT22, LDAM^{+}07, LHC^{+}21, MMIB10, MPM^{+}18, OCH^{+}18, OOTA15, PCSMC12, PPdS21, RKCH15, SOS^{+}07, SMPC^{+}12, WOW^{+}14, Wil87, YCP^{+}12]. Larval [CHSB^{+}21, DPM^{+}09, MKSvA^{+}22, OAWAN18, RLGC10, ALT10, CDTM^{+}21, CZG^{+}21, CZL^{+}24, CCM^{+}14, CHS^{+}24, CKL^{+}14, DCM16, DPGC14, FGGDF^{+}04, FB05, GKC^{+}14, GCG^{+}14, GDM^{+}15, HLS^{+}14a, HKPV12, HPNDC15, HCGK11, MGKW19, QCdS^{+}07, RNBP^{+}19, RLL^{+}09, RBHLA04, SNV^{+}18, SMR^{+}20, SCHD23, TTMM^{+}17, TKC^{+}22, TMR^{+}21, VDGGD^{+}22, ZLRVB24a, ZLRVB24b]. laser [BTNK13, LGL^{+}18, PKA19]. last [AVG^{+}19, CFC^{+}18, DYL^{+}15, Emi65, GS24, LXC^{+}22, PBB^{+}20, RSB^{+}01, RCD^{+}94, SGF^{+}19, SF02, TKW06, VSGC21, ZLC^{+}15, TAF^{+}22]. lasting [PDV12]. Late [EHG^{+}12, dWDB^{+}98, DLD^{+}19, GDM^{+}15, HMH07, INI^{+}17, KiL14, KMS^{+}24, KYS^{+}17, SKSK06, ST65, TKWI08, VHV^{+}12, YLL19, dFKdLZTT17]. late-spring [dFKdLZTT17]. Late-summer [EHG^{+}12, dWDB^{+}98]. Lateral [IHT^{+}21, SAH^{+}21, TPPG10, BHPC06, Mar03, VK92]. Latest [RBD^{+}07]. latitude [Ang79b, AE09, BVJE19, HSL96, HDA^{+}16, KA85, MHS^{+}20a, MHS^{+}20b, MBD^{+}09, RGMPR23, Wu13, ZLC^{+}15]. latitudes [ABSDC07]. Latitudinal [BMdMS^{+}21, BHS^{+}15, FWH^{+}17, GVKD^{+}13, YMA^{+}17, BHC^{+}18, BGWP^{+}17, BTV^{+}17, DLL^{+}23, SCHS^{+}24]. Law [War06, dB94]. Lawrence [MPN09, HGD22, LBC^{+}23, MPSD15]. Layer [Car98, SPN98, ATS01, AMEV07, Ano94c, BMK12, BOG20, BTG^{+}03, BLC23, CMHM18, CS18, CDP14, DHC^{+}20, DGMM85, DVB^{+}18, Gam14, HGD22, HFW^{+}98, HTdM^{+}15, HLS^{+}14b, LYS^{+}22, MY92, MKM93, McD81a, MSGGM18, NDEG22, NF06, OVR^{+}02, PMFNGQ21, PFW15, PNF^{+}21, SAT^{+}22, SLGI^{+}21, TvW98, TvG02, Tsu86, VB14, WZFW16, XHW^{+}20, XY21, YTNK00, ZSW^{+}22]. layered [CGL^{+}20, Ros65]. layering [PSGVS^{+}14]. layers [BLMR^{+}20, CLG^{+}00, DL17, FBT^{+}22, GRMB18, GIC20, LRAE23, PVB23, Peñ24, PdMS^{+}13, VOJD02b, WGZZ19, ZPC^{+}16]. lead [VPM^{+}19, SW81]. leading [DYO^{+}10, HHWW20, OOTA15, YMK^{+}04]. leaping [CGW^{+}22]. learn [JvdLL^{+}15]. learned [BK08, MLHM09]. learning [CLX^{+}20, GL23, PVA24, WWL^{+}24]. least [BGL^{+}17]. lecture [Sie88]. lee [HNSP^{+}19]. leeches [UKK^{+}19]. Leeuwin [GW91, MB07, TWAL^{+}11, WOW^{+}14]. legacy [Fly10, PDV12]. legend [Ano17i, Ano17j, Ano17k]. legged [WFD^{+}07]. lemuru [SNR^{+}10]. length [LVGH^{+}15, SEG^{+}22b]. Lepidochelys [CdTH^{+}16]. Lepidodinium [RSM^{+}23]. Lepidopsetta [LDAM^{+}07]. leptocephali [FMC^{+}15, GDM^{+}20, KMF^{+}20a, KMF^{+}20b, MFA^{+}15, MSA^{+}22]. Leptonychotes [NRS^{+}19]. Leptosomatidae [MSFZ19]. less [BFR13]. Lessons [BK08, MLHM09]. Levant [Ore69]. Levantine [ABM^{+}05, ÖHÜ89]. level [BHPC06, CDDF11, CJMO87, CAT^{+}08, Con87, Den87, Dev87, DPCS87, DOP87, FP15, FK99, FG16, FWL^{+}15, GAPM16, HHWW20, HPHW21, JKBH87, KKKS14, LMPB^{+}16, Let87, LSMG01, LB20, MTC12, MVS08, NO14, NH88a, NF87, OPH^{+}24, PS91, Pir87, PBN13, RN02, ŠVL^{+}15, ŠPM^{+}22, SAM^{+}10, SC65, STR01, SIB^{+}06, SSW^{+}09, SCS87, Tho87]. levels [LLL^{+}11, LH08, MPC^{+}17, PHFK14, RD11, STM10, SPMVP05, SNMW10, Tan99, WFR07]. Levitus [Cai95]. lie [CPC88]. Liebig [dB94]. Life [JAC^{+}12, MFB^{+}84, Nie07, PRTC13, SSTD^{+}95, AHP19, BVJE19, BSF^{+}21, BLT^{+}08, CGC^{+}20, CMS^{+}13, CCM^{+}14, ECFT20, FWH^{+}17, FVLC^{+}23, FCN^{+}19, FMT15, GiIKX22, HLM^{+}16, HE07, HPW10, IVR^{+}13, KSK^{+}15, KLP^{+}17, LSY^{+}14, MMG^{+}13, MLB^{+}20, MT99, MST^{+}23a, MRH^{+}14, MC88, MVBC^{+}21, NHG19, PWZ^{+}16, PDAM^{+}15, PST^{+}15, Pra22, RNBP^{+}19, SCD^{+}07, dSSDS^{+}20, STS^{+}12, SCHD23, SLY^{+}15, SSW^{+}09, THBA19, TMR^{+}21, VMH^{+}21, WKS^{+}15, Wil87, YAK^{+}08]. life-histories [Wil87]. life-history [dSSDS^{+}20, SCHD23, YAK^{+}08]. life-stage [MMG^{+}13]. life-stages [PWZ^{+}16]. ligand [HSLG11]. Light [ACK^{+}13, LCB18, Mor91, SK91, ÁLC22, BMGN15, GGPG^{+}19, HOY^{+}21a, JJL^{+}19, KM08, KY15, MBP^{+}11, RHM^{+}19, ST03, SBBV04, JTD^{+}14]. Light-dependent [ACK^{+}13]. Lignin [SMN^{+}14]. Ligurian [BAARB05, BMB^{+}16, BBM^{+}14, CPB^{+}15, CLD22, CBB^{+}22b, CQZ^{+}18, CRC^{+}19, DSC^{+}19, GCF^{+}19, GBB^{+}20, OÁT^{+}05, PCD^{+}18, ZGB^{+}20]. like [HCAFD^{+}20, ICB^{+}19, LSM08, WST^{+}16]. likely [MHR^{+}10]. Limacina [MLB^{+}20]. limit [KAK^{+}22a, SAd^{+}17]. limitation [AFBT^{+}22, BB14, HLR17, ILA21, ZCD08]. limitations [BBE^{+}15, BMG13]. limited [KEV10, LHF^{+}16]. Limits [LRJ^{+}15, ALT10, Hea12, LHW^{+}20, SGL^{+}18]. line [CLdPHL23, DN07, HWS^{+}07, JOGM^{+}10, KH09, CGB07, Fre07, MG02, PV07]. lineage [SPH^{+}15b]. linear [BMGN15, HNL14, KC15, LI10, McD81a, PC87]. lines [Kaw86, KMWF11]. link [AB90, HHSR07, IIS^{+}17, MCG^{+}14, RGI05, SBPGP^{+}23, VMA^{+}24]. Linkage [ZHD^{+}20, BM01, KiL14, KSG^{+}17, NMY^{+}14]. Linkages [YPVP^{+}22, BHA^{+}14, HPS^{+}01, KYS^{+}17, NYH^{+}22, TTB^{+}08a, WLD^{+}15]. linked [CHSB^{+}21, DWS^{+}24, HTV^{+}20, ONR^{+}14, RBR^{+}23, SSL08, YKNO23]. Linking [BFB^{+}20, DMBHG10, FFA06, GF19, MRW^{+}14, SST^{+}17, SPWH21, TSS^{+}12, BB10, STF^{+}13, THP21]. Links [SF15, BC01, GDL^{+}15, HLS^{+}14b]. Linnaeus [YGL^{+}10]. Lion [EMU^{+}23, PGT^{+}13, STG^{+}18, CDH^{+}13, SGL^{+}13]. Lions [DDDT99, FFS^{+}20, PPdM^{+}12]. lipid [CPPPEAG22, FMC^{+}20, PPCWJ18, SGL^{+}17]. Lipids [BC01]. Liquid [NBR^{+}08]. liquidum [Coo69]. Lisbon [KGdS^{+}08, OMS^{+}09]. List [Ano03j]. lit [Bak83]. Lithodes [BMN19]. lithogenic [PBP^{+}99, YAI^{+}14]. litter [CPC88, MDB^{+}20, RLDC^{+}13, TCL^{+}15]. Little [HM90, Nof00, KGJ^{+}10]. littoral [DAKV99, dlHRA^{+}18]. littorals [ST65]. Living [DSC^{+}19, LMP22, BHK^{+}19, LFG10, LIH^{+}12, MM80, MSFZ19, OSH^{+}96, Ric01, SKT01, TSH^{+}17, SAB^{+}22]. LIW [MTL05]. loading [LM97]. lobster [ACK^{+}13, GRB^{+}08, MPM^{+}18, WOW^{+}14]. Local [FBB^{+}21, JOGM^{+}10, ASB^{+}08, BBS21, BBS23, CMM^{+}04, CGC^{+}20, IHT^{+}21, LLAPG^{+}22, NGLSSG14, PMMN^{+}22, PSM^{+}22, RSW^{+}23, RHB23, YCP^{+}12]. local-scale [ASB^{+}08]. localities [Mid69]. locality [VPW01]. localized [LRJ^{+}15]. location [LBC^{+}23]. locations [LC12, MGF^{+}13]. lock [MZK^{+}23]. loco [GKC^{+}14]. Logachev [vHMDL14]. loggerhead [CBB^{+}22c, MHR^{+}10]. logic [JPM^{+}08]. logistics [KNS^{+}03]. Loliginidae [OÁSG^{+}16]. Lombok [GXX^{+}22]. Long [BBR^{+}01, BLCL14, BD18, BBL^{+}09, CMF^{+}09, DLM^{+}96, FMP19, FAH^{+}13, FMSBW13, GMD^{+}22, KON14, KNS^{+}03, LSIC12, LSXT01, MDAW^{+}19, Nag01, PG10, STJ^{+}14, ŠGM^{+}18, Spr08, TFY02, TKW06, VMB^{+}22a, APC13, ABE^{+}15, BPTT19, CSK^{+}12, CB17, DEW^{+}97, FRV^{+}19, GML^{+}23, GHL15, GCD97, HFS^{+}20, hHRW^{+}05, HHZ^{+}22, JOGM^{+}10, JLP^{+}20a, JLP^{+}20b, JLS^{+}22, KRHS14, LO07, LHE^{+}13, LPF^{+}20, MLL^{+}22, MDGC^{+}12, PGY^{+}22, Reb02, SOH21, STHM02, UPPS^{+}21, WFD^{+}07, WHBW03, WHIH97]. long-chain [SOH21]. long-line [JOGM^{+}10]. long-period [GCD97, WHIH97]. long-range [DEW^{+}97]. Long-term [BBR^{+}01, BLCL14, BD18, BBL^{+}09, CMF^{+}09, DLM^{+}96, FMP19, FAH^{+}13, FMSBW13, GMD^{+}22, KON14, KNS^{+}03, LSIC12, LSXT01, MDAW^{+}19, Nag01, PG10, STJ^{+}14, ŠGM^{+}18, Spr08, TKW06, VMB^{+}22a, APC13, ABE^{+}15, CSK^{+}12, CB17, FRV^{+}19, GML^{+}23, GHL15, HFS^{+}20, hHRW^{+}05, HHZ^{+}22, JLS^{+}22, KRHS14, LO07, LHE^{+}13, MLL^{+}22, MDGC^{+}12, PGY^{+}22, Reb02, UPPS^{+}21, WFD^{+}07, WHBW03]. longer [YYhT^{+}17]. Longevity [BMB06]. longicauda [GKR20, RMB^{+}01]. longicaudata [SIR^{+}07]. longipes [SSTD^{+}95]. longirostris [PPdS21]. longitude [YMA^{+}17]. longshore [CB17, Hut87, Yos80]. look [CLMR23, HS22]. loop [MHA^{+}11, dJSL^{+}20, HMS^{+}22, KAK^{+}22b]. loophole [IFC^{+}07]. looping [CGW^{+}22]. LOPC [EBS^{+}18]. lose [DSAB20]. Loss [GMR^{+}23, MBF^{+}14, PLN^{+}23, USH15a]. loss-on-ignition [USH15a]. losses [Wal83]. Louisiana [MCB^{+}90]. Low [Ban96, CM18b, AAML22, AC85, Bak01, DCM16, FYYC05, HL05, HSL96, KA85, KGB^{+}23, Lav09, Mol22, MKS^{+}22, PZA^{+}15, STM10, TSFA22, Wu13]. low-density [Mol22]. Low-frequency [CM18b, PZA^{+}15]. low-latitude [KA85, Wu13]. low-salinity [MKS^{+}22]. Lower [LLL^{+}11, CFM^{+}18, CAT^{+}08, FK99, GAPM16, KAK^{+}22a, LSMG01, MPSD15, MPC^{+}17, PCH^{+}08b, SNMW10, Tan99, TAM^{+}13, WGZZ19, WGG^{+}08, YWUK15]. lower-trophic [CAT^{+}08]. loxosomatid [BC19]. Loxosomella [BC19]. LTER [MDGC^{+}12]. LTER-MC [MDGC^{+}12]. Lunar [RCSA01, UPPS^{+}21]. Lutjanus [ZLRVB24a, ZLRVB24b]. Luzon [CGW^{+}22, MZH^{+}23, YLY^{+}14, YYhT^{+}17, ZYN^{+}24]. lyase [CM14a, FPS^{+}09]. lysis [CRC^{+}19].

M [Ang80, LMPB^{+}16, CDS90, GSSWK20, SS03, LL97, YMI88]. M. [BLCL14, RK20]. M9.0 [CLB^{+}14]. machine [CLX^{+}20, WWL^{+}24, VSPP14]. Machine-based [VSPP14]. machine-learning [CLX^{+}20, WWL^{+}24]. Mackenzie [LPF^{+}20, LPF^{+}21]. mackerel [ACN01, BHH^{+}16, BSH^{+}20, EiIT^{+}22, GiIKX22, JSdSS^{+}21, TMÁGC^{+}21, YKNO23, ZL01]. mackerels [ABE^{+}15]. Macro [CHC^{+}12, KTN14]. Macro- [CHC^{+}12]. macro-nutrient [KTN14]. macroaggregates [KS15]. Macrobenthic [TPP^{+}00, BRR^{+}22, GBC^{+}15, IG19, JMSB^{+}23, Sok90]. macrobenthos [Whe06]. macrocephalus [CHS^{+}24]. Macroecological [BLHB07, BMG13]. macrofauna [BAB^{+}19, FMH02, GSV^{+}01]. Macrofaunal [HDM19, Car98, SH09]. macroinfaunal [CLSD18]. macronutrient [TRP^{+}23]. macronutrients [AIHB^{+}07]. Macrophysiology [HBR11]. macroscopic [PMA^{+}14]. Macrostylis [RK20]. macrozooplankton [CIL^{+}23, DSR21, HGD22, MCT03]. Madagascar [ZQWP23]. Madden [RLX^{+}24]. Madeira [RBL90]. Madeiran [PMG15]. maenas [YFK21]. Magallanes [PTPY^{+}23]. Magellan [AHGRAL23, Gri22, SPF^{+}23]. Magellanic [BRR^{+}22]. magellanicus [GGJ^{+}10]. magnesium [Wis65]. magnetization [Hey78]. magnitude [CCH^{+}12]. Mahlon [CNT03]. main [FZ88]. Maine [BDB^{+}04, MLB^{+}20, RKC^{+}10, TPTM23, WBC^{+}22, WLM^{+}22]. mainland [CMM^{+}04]. maintain [MAH^{+}15]. maintenance [DS65, PC87, PVV23, STS^{+}12, WWL^{+}22]. Major [BTK^{+}99, JTQ^{+}18, LDMH09, VDB^{+}20, AAMB^{+}24, BBSN04, CED09, DWS^{+}24, HS07, Igu04, LTJ^{+}15, MLD^{+}03, MIN^{+}20, NRA17, PLP99, TSAM^{+}22, Wüs64, ZMW^{+}23].

Malaspina [GKS^{+}13]. Malin [XD96]. Malin-Hebrides [XD96]. Mallotus [BSF^{+}21, CGV13a, CGV13b, MMD^{+}16]. Malvinas [PMC21, PSP^{+}21, RPPM^{+}23]. mammal [KFH^{+}15]. mammals [SPS^{+}99]. Management [SHG12, BDE03, CN22, Dev87, EAB^{+}23, FPS^{+}13, GKC^{+}14, GRDS10, HFS^{+}20, INI^{+}17, JOGM^{+}10, JHDT12, JBH20, KHL12, KMB01, LLL^{+}11, LFG10, LIH^{+}12, McI10, MBF^{+}14, Ric01, RS04, SSB^{+}20b, TCF^{+}18, WFJ^{+}15, Zav99, ZL01]. Managing [TSH^{+}17]. mandalas [KV18]. manganese [Hey78]. Mangkhut [DFC^{+}21]. manifestation [KSD84, RM97]. Manila [YAK13]. Mannar [JJA^{+}13]. manta [RHB23]. Manx [SSB14]. Manxman [Mer65]. many [Ric01]. mapping [BPGD^{+}14, LMT^{+}19]. maps [RRS03]. March [Ano20q, Ano21r, Ano22y, Ano22-28, KSPK99, KDB95, TBK^{+}99, GR17, Ano99d, Ano07o, Ano08u, Ano08w, Ano09j, Ano10l, Ano13m, Ano14n, Ano16k, Ano17s, Ano18m, Ano19l, Ano23r, Ano24l, CLB^{+}14, GSPMAI99, KTB^{+}99]. marcusorum [BC19]. Margin [BHE^{+}98, CHG^{+}18, FLdST98, LHEB98, OB98, TMPM^{+}16b, TMPM^{+}16a, TvW98, vWMH98, BGM^{+}01, BCOL^{+}19, CLG^{+}00, CLA^{+}00, CVHM^{+}18, CdMS^{+}18, CBD^{+}24, DXH^{+}02, EvdZSH02, FARRL^{+}13, FMH02, GLV12, HG04, HWF^{+}21, Hut95, JW01a, KKS^{+}18, KiL14, LGR^{+}02, MH02, MCGS^{+}16, ORW^{+}01, PCD^{+}18, PPHM18, PGGG17, RLT^{+}22, RCÁS^{+}15, RLR^{+}18, RCSVGP^{+}16, SCMAR^{+}99, SVHM^{+}13, SvWRvB02, SH09, SW01, SMGL01, TCDPP^{+}22, TvG02, TPP^{+}00, VHV^{+}12, WDMMK89, XLX^{+}20, YMI88, YKH^{+}24, vWM02a, vWdSBdH02, BPP^{+}98, HM98, PS98, dWDB^{+}98, vWHdS^{+}98]. Marginal [BHM^{+}15, FJhT^{+}14, LMW^{+}12, PO00, CWZ^{+}20, DPB06, Gal17, IHT^{+}21, IU14, JCF^{+}23, KJZ^{+}12, MTC12, McK04, MNFY21, NYH^{+}22, PB94, SCC14, SPWH21, LLX^{+}21]. margins [ACK^{+}13, BPGC^{+}20, ESTM13, Kit03, LDB^{+}02, Sol00]. mariculture [Mae88]. Marine [BCS09, Bel09, CC23, CPC88, IGG^{+}19, KSE^{+}09, KFM^{+}17, KA94, LPA^{+}11, OLH^{+}18, Pea02, Pra22, SDB^{+}21, Sim81, SPS^{+}99, SAB^{+}22, TCL^{+}15, YW22, dMGS^{+}11b, dMGS^{+}11a, AHP19, AS88, AF10, AGD^{+}18, AH15, AOMZ^{+}23, AHC^{+}13, ARELAK24, ALM^{+}23, ADV^{+}18, ALT10, ABP^{+}23, BDP^{+}06, Bak06, BW08, BMO12, BVB88, Blu88, BDE03, BCD^{+}20, CLSP17, CSH^{+}23, CM14a, CSBL^{+}15, CSC^{+}12, CTMV^{+}14, CWS^{+}21, CS89, CE84, CJG88, CRiI^{+}15a, DL69, DLD15, DMF^{+}09, DP13, DAKV99, DTKvH15, EVM^{+}15, FSAO22, FMM^{+}20, FPJ^{+}15, FCG88, FAAV^{+}15, FDH20, FDM^{+}13, GSFP^{+}09, GAF15, Gal17, GPEV20, GKC^{+}14, GGT^{+}15, GHVG19, GPAB^{+}16, GMDS20, HVS10, HGD22, HSG^{+}15, HVTV22, HGB^{+}21, HSH^{+}19, HWL^{+}20, HAP^{+}16, HHW01, HHW22, HAA^{+}14, Hop64, HLS^{+}14b, HM06, HDA^{+}16, HBD^{+}18, IL20, JAS^{+}20, KST^{+}10, KTH^{+}21, KGL22, KiL14, KMB01]. marine [KBF^{+}08, Kli10, Kno04, KOT^{+}21, KLC^{+}15, KMSTK23, KKO10, KFH^{+}15, LTJ^{+}15, LK13, LSY^{+}14, Law04, LML^{+}23, LMS10, LAHI10, LDD^{+}22, LFBR^{+}18, LMT^{+}19, LSS^{+}09, LFG10, LIH^{+}12, LHF^{+}16, LSD^{+}15, LRGV^{+}18, LMP22, MPV12, MDB^{+}20, Man04, MV10, MRBS^{+}24, MFS^{+}07, MSF^{+}07, Mau17, MLL^{+}15, MBH^{+}01, MI21, MLM09, MLHM09, MCG^{+}14, MS15, Mor91, MSB^{+}23, MPN09, MVV^{+}19, NCC^{+}15, Nie07, NBLI20, OV24, PL01, PS08, PBO10, PM13, PG10, PHKS01, PPD^{+}12, RLDC^{+}13, RBL^{+}19, RK03a, RCGC^{+}16, RCB^{+}20, RLX^{+}24, RNBP^{+}19, Ric01, RKK^{+}21, RRLS22, RN06, SSB19, SBMB18, SGL^{+}17, SHd13, SN24, SBL^{+}23, SRF^{+}19, SvN04, SOWS17, Sel65, SJJ^{+}03, SMB88, Sie88, SW81, SFF^{+}24, SKGS20, SRM^{+}10, ST65, SPH^{+}15b, SDJ14, SJLW23, SIB^{+}06, SHT^{+}01, TSAM^{+}22, TSH^{+}17, TSJC07, Tur15, VMB^{+}22b, VCM04, Was11, WSS15, WMB^{+}18]. marine [WBF^{+}21, WHK23, YSY^{+}19, YZX^{+}23, ZLKO00, dLLdAWL^{+}23, dPCS23, Cia22]. marine-oil [BCD^{+}20]. marine-terminating [GHVG19]. marins [Bou65]. marker [KSB^{+}22]. markers [NMN08]. market [PG10]. markets [JTGM10]. Marko [SW81]. marlin [JOGM^{+}10]. Marmara [BSÖ^{+}94, FBS^{+}18]. marmoratus [PPPdS20]. marshallae [PD15, THP21]. Martin [Oll15, WR03]. Martínez [RBL^{+}19]. Marukawa [Mil88]. Maryland [Let87]. Mascarene [FMC^{+}15, MFA^{+}15]. Mass [AHW99, MHGP06, AdAK^{+}18, EMK^{+}17, HOY^{+}21a, HPZC21, JG07, JVJ^{+}17, KT04, KMU^{+}12, LVGH^{+}15, LvIKB07, LC10, MRMD^{+}97, MLD^{+}03, Mau10, PTM^{+}22, PZA^{+}15, RGC^{+}01, SGMP15, WWW^{+}23, YRKC08, ZPC^{+}16, dWDB^{+}98]. masses [ÁSÁB^{+}14, ÁBMÁS14, ÁBMÁS15, CPHR98, Fla02, GIPC^{+}15, GR85, GLAHH^{+}22, GSF^{+}15, Hut95, KKK^{+}04b, LPBM17, MRO^{+}08, MAAS^{+}00, PPVG12, PPKR14, Rea00, RGPB^{+}23, Rud15, SWZS^{+}21, SBG^{+}08, TRY^{+}04, TPTM23, YYT^{+}14, YKH^{+}24]. massive [KTW^{+}22, RKM^{+}07]. match [LHC^{+}21, ZLX^{+}20]. match/mismatch [LHC^{+}21]. Material [LBK^{+}01, MBCB88, MWO^{+}12, OE65, ÖÜT93, Tur65, TRMV15, YAI^{+}14]. materials [ÁSDB^{+}01, ÁSÁB^{+}14, ÁBMÁS14, ÁBMÁS15, BCF^{+}03, Kit03, OAD22]. Matter [CMPNC^{+}22, ASC92, BB24, BHHR15, BSW86, BFJ18, BMG^{+}21b, CKP^{+}20, CGM^{+}02, CGC^{+}20, CLG^{+}00, CS89, CE84, DDE^{+}95, DDP^{+}00, DMD^{+}00, DDD^{+}00, DBW^{+}22, DCS^{+}22, FPD^{+}01, FRV^{+}19, GMAMB04, GBM^{+}01, GLAHH^{+}22, GF19, HOY^{+}21a, HSLG11, HOY^{+}21b, HLD^{+}21, IL20, IU14, JZZY24, JW01a, JMG^{+}13, KNSN^{+}09, KLB^{+}21, LBNBM13, LSB^{+}17, LSD^{+}18, LSD^{+}15, LHEB98, LvIKB07, LFBP^{+}13, MPCNC^{+}19, MFDH22, MFY^{+}86, NRA^{+}21, ORW^{+}01, PL87, Par63, PGT^{+}13, PRL^{+}18, PLP99, PBP^{+}99, RFSCF19, SVHM^{+}13, SRF^{+}19, SMN^{+}14, SK18, TPPG10, TvW98, TZP^{+}00, TPM^{+}00, USH15a, VFCC^{+}22, WMC^{+}89, WRS^{+}92, XYGJ23, YTB^{+}21]. Matthias [MC08, YRKC08]. maturation [Nie07]. mature [ATT^{+}08]. Maud [Har05a]. Mauritania [FRK^{+}09, GEO09, NIF^{+}15]. Mauritanian [FPS^{+}09, HNR^{+}17, RF17]. maxima [Yu23]. Maximal [VWDF14]. maximize [SHD^{+}21]. Maximum [BBC^{+}22, APP21, BBB^{+}21, DFD23, DYL^{+}15, FPIJ85, LFI^{+}13, MBP^{+}11, PVV23, XHW^{+}20, TAF^{+}22]. May [Ano21a, Ano22a, Ano23a, Ano24a, LAP10, dSSDS^{+}20, Ano00b, Ano03d, Ano07r, Ano14r, Ano15r, Ano16l, Ano17l, Ano20r, Ano22-31, Ano22-32, GCZ^{+}00, MN88, PGRP^{+}18, YJW88]. MC [MDGC^{+}12]. MCS [Cia22].

Mean [ED82, SJM^{+}19, TM13, DML^{+}16, DGMM85, HLK13, HPHW21, KDL^{+}01, Mol22, Owe91, RD03, RG09, SS69, UAM05, VYGMM^{+}17, Zez90]. meander [TMN^{+}12]. meandering [SSB14]. meanders [dSPF^{+}23]. means [CPSM20]. measurable [Nee85]. measure [Szu12]. measured [SLM^{+}16]. measurement [DPCS87, DDJ^{+}21, HM15, MPMA13, Sac16]. Measurements [HHB^{+}22, TJ73, ADS^{+}22, BBL^{+}09, Due77, FTG^{+}11, GRMB18, GA00, HHK^{+}02, JSKM02, KSK21, MXC^{+}21, NBLI20, RCGC^{+}16, RLSF06, RLSF07, SAT^{+}22, SV97, TSG^{+}04, VSA^{+}21, VBAC^{+}21, WZFW16, Web69, Whe06]. measures [JBH20]. Measuring [MSJ^{+}15, KSY^{+}19]. mechanics [GD91]. Mechanism [KKNT23, SMFM^{+}21, ZLG17b, CGB^{+}23, LYS^{+}22, STHM02, YYC^{+}18, ZLG17a, ZJZ^{+}21]. Mechanisms [BLR^{+}23a, HAH^{+}22, NMC^{+}09, PC87, SvN04, STS^{+}12, THP21, AAMB^{+}24, Bea04, CRS04, CMF15, CS04, DLL^{+}23, DMT15, FBS22, Has06, HLS^{+}14a, JAS^{+}20, MMIB10, NIC^{+}19, NRA17, OMS^{+}15, SGS^{+}23, SD07, TKWI08, VPM^{+}19, VGJ^{+}19, WLD^{+}15, Woo18]. mechanistic [BDP^{+}06, BFPS06, BSH^{+}20, SLY^{+}15]. Med [NDEG22]. Med-ROMS [NDEG22]. Meddies [APC13, RBZ00]. mediate [ALT10]. mediated [DCKB13, GSM^{+}17, MLHE23, MMD^{+}16]. Mediterranean [ACB^{+}13, AQVB^{+}10, ACL^{+}18, BBM^{+}14, CLD22, CQZ^{+}18, CCHV^{+}21, CHSB^{+}21, CLG^{+}22, DDE^{+}95, DDDT99, DAU22, Eri65, EMU^{+}23, FTG^{+}18, GBB^{+}20, GGA^{+}16, GGG^{+}18, GGSM^{+}20, HLM^{+}13, JMG^{+}13, KHC^{+}99, KFC^{+}13, KTB^{+}99, LTG85, LSM^{+}22, LRGV^{+}18, LFBP^{+}13, MGC^{+}18, MAFS^{+}22, NDEG22, NCC^{+}15, POS^{+}07, PCD^{+}18, PGT^{+}13, PPSVC^{+}13, RCC^{+}18, SCAA07, SCMAR^{+}99, SVHM^{+}13, SGA^{+}19, STG^{+}18, TCDPP^{+}22, TPP^{+}00, TCL^{+}15, VGM^{+}23, VBAC^{+}21, WYT00, dMGS^{+}11b, APN^{+}15, AR18, ALG^{+}21, AVS23, ABC^{+}99, AUE^{+}14, AIA^{+}18, BPA^{+}21, BTK^{+}99, BFP^{+}18, BRC^{+}18, BZD^{+}21, BNCC15, BS90, BMC05, BSFM^{+}12, BOMdP15, BGM^{+}99, BD19, BFJ18, BGA^{+}21, BFV^{+}17, CCM^{+}13, CDL19, Car98, CGM^{+}02, CPG08, CMF^{+}09, CMF11, CF20, CJRÁ^{+}13, CMPNC^{+}22, CMS^{+}13, CC88, CLG^{+}00, CLA^{+}00, Cia14, Cia22, CD65, CGD^{+}18, Con87, CFML22, CLL^{+}18, CLG^{+}22]. Mediterranean [CTR^{+}19, CRC^{+}19, CFC^{+}18, CdMS^{+}18, CBD^{+}24, CAB^{+}99, CJG88, DSV99, DDP^{+}00, DTOD00, DVL^{+}99, DSC^{+}19, DVB^{+}18, DCRR^{+}22, DFM^{+}15, DAKV99, DCL^{+}13b, EMU21, FVA^{+}19, FBR^{+}13, FBD18, FTC^{+}16, FARRL^{+}13, FDHT05, FBS^{+}18, Fuk91, GSPP^{+}20, GGT^{+}15, GPE^{+}17, GTB07, GPP22, GHL15, GFB^{+}15b, GFB^{+}15a, GGA^{+}16, GIPG17, GCD^{+}99, HMTL05, HDM19, HTdM^{+}15, IVR^{+}13, JMZ23, JFEC13, JVJ^{+}17, KSVT00, KKS^{+}18, KQP^{+}17, KCL^{+}12, KMU^{+}12, KEV10, LT06, LRNK99, LGL^{+}18, LSIC12, LDMH09, MMGL^{+}07, MHGGS19, MDB^{+}20, MMIB10, MPCNC^{+}19, MHA^{+}11, MCD^{+}07, MDGC^{+}12, MCGR07, Mil09, Mil14, MMG^{+}11, MMPG07, MVV^{+}19, MEMC05, NGLSSG14, ORMR^{+}19, OACB^{+}15, ORB^{+}18, ÖÜT93, PBB^{+}20, PGLG^{+}05, PAG^{+}18, Peñ24, PRA^{+}18, PCC^{+}19, PZA^{+}15, PST^{+}15, PKV18, PRL^{+}18, PBP^{+}99, PTI00, PdMS^{+}13, RLDC^{+}13, RGB^{+}17, RDD^{+}18, RTBR^{+}22, RPG^{+}18]. Mediterranean [RKM^{+}07, RCF^{+}13, RAE^{+}05, SOS^{+}07, SLBR18, SGL^{+}13, San73, SCPN15, SHd13, SCB^{+}09, SCCJ^{+}18, SGPdM18, SFK^{+}99, ŠVL^{+}15, SCC^{+}19, SCT^{+}00, ST65, TCS15, TPN^{+}18, TR99, TAO05, Tol85b, TCF^{+}18, TZP^{+}00, TPM^{+}00, TVD^{+}99, TIOM16, Tur99, Val99b, VCB^{+}00, VYGMM^{+}17, VKT15, VOT^{+}99, VBA^{+}18, YHM^{+}18, YPVP^{+}22, ZGB^{+}20, Zav99, ZMCD11, ZFSV^{+}09, dMGS^{+}11a, dPAJ07, dlPHF^{+}15]. medium [DRE^{+}08, RLGC10]. medium-resolution [DRE^{+}08]. medium-sized [RLGC10]. MEDOC [Kil76]. Medusae [RJT84]. Meeting [FJhT^{+}14, PBH^{+}10]. Megabenthic [GGG^{+}18, MMKS^{+}21, PGGG17, BMNW01, CLMR23, DCRR^{+}22, JLS^{+}22, KOhL^{+}10]. megabenthos [BBR^{+}01, CMF^{+}09, TRLA^{+}13]. Megafauna [PCC^{+}19, ADV^{+}18, BBMR19, CAO^{+}20, FBD18, LJM^{+}16, LM18, LDB^{+}02, MDR20, SLPA^{+}20]. Megafaunal [SLBH^{+}19, Car98, CMHM18, DTC^{+}06, DBJ^{+}15, DBR20, KSVT00, KRHS14]. megalopae [PPPdS20]. Meganyctiphanes [CNT^{+}19, MPSD15, SIR^{+}07]. Megrey [CRiI^{+}15b]. meiobenthic [RCF^{+}13, ZMCD11]. meiobenthos [RMG90, Sol00]. Meiofauna [DDE^{+}95, FMH02, SSKA19, GSV^{+}01, GCLD19, MGA^{+}23, VKGP^{+}13, VKDS^{+}18]. Meiofaunal [BD19, CLD22, DTOD00, LT06]. Mejillones [VOG^{+}08, VSGC21]. Melchior [Bak83]. melt [KGB^{+}23]. meltwater [HHR^{+}19, KMSTK23]. members [DSAB20]. Membership [Ano03j]. membrane [MRBS^{+}24]. Memoriam [Ano04a, CRiI^{+}15b]. memory [Ano20u, Bru88, Sie88]. Mendocino [Nay65]. Menorca [GGA^{+}16, GGG^{+}18, GGSM^{+}20, SGA^{+}19]. mentor [CNT03]. Mercury [AS20, BMC17, CdMS^{+}18, CBD^{+}24]. merging [FHL^{+}24]. Meriadzek [MGS90, VK90]. Meridional [ARH^{+}00, AB00, Ano17a, BJ17, CF12, HLR17, HGPFN^{+}14, HGT16, KBHML17, MSJ^{+}15, PHCA17, RNP^{+}17, SWT^{+}17, ZSBL00, AdAK^{+}18, APHGC^{+}22, BDBJ01, KF11, LM00, MR06, MFS^{+}16a, MFS^{+}16b, MLS^{+}15, Nof03, Ric08, SFMT12, SFMT14, TBW00]. meridionally [CSLJ03]. MERIS [CN22]. Merluccius [BWMGCB08, FCN^{+}19, FB05, LC12]. MERMEX [AIA^{+}18, SGPdM18]. Meroplankton [WZBK^{+}21, MTC14]. Mesima [CFM^{+}18]. meso [CSC^{+}12, Kli10, LLH^{+}20, RÁSG^{+}13, STEB16]. meso- [LLH^{+}20]. meso-scale [CSC^{+}12, Kli10, RÁSG^{+}13, STEB16]. Mesopelagic [OHC^{+}17, Ang84, ATC^{+}19, BOMdP15, BOG20, CLCBB19, DCKB13, Dom84, FCMCÁS19, GBB^{+}20, HLCdP19, KVNT20, MVBC^{+}21, NMN08, PNF^{+}21, Pug84, Roe84b, RB84, RJT84, SMN^{+}13, SLBVRR^{+}22, SLOP^{+}22, SWH^{+}24, ZPC^{+}16, GASV^{+}09, RAB^{+}84, Roe84a]. mesoplankton [DZ04]. Mesoscale [BAM^{+}09, BF12, CGG08, CBT07, FDHT05, MTH^{+}10, MHCR^{+}12, OP18, RD03, RBS^{+}09, SSS^{+}11, WWZ19, BAARB05, BCR^{+}13, CSS11, CD07, CdD^{+}15, DMBHG10, FELMGM^{+}22, GLY23, GBC^{+}16, KM10, KZD^{+}19, LJM^{+}16, LBSP01, LLS01, MXC^{+}21, MNS^{+}24, MPM^{+}18, MZK^{+}23, MTL05, MEMC05, MJA^{+}07, PM22, PLJR22, RBHLA04, STC10, SI97, TG05, UBB^{+}23, WLM^{+}13, WR00, XYL^{+}22, XYWY23]. Mesozooplankton [GCCY^{+}14, HLGA07, MT99, SM05, TSNO05, WTT14, BFH01, CS03, CLdPHL23, DBC^{+}18, DDK^{+}18, EHG^{+}07, GSVB23, GCC^{+}24, HPS^{+}01, HHY03, INO^{+}24, INT14, JJA^{+}13, JJA^{+}17, KJZ^{+}12, KST03, KSS^{+}23, KRL^{+}22, NYL^{+}17, RÁSG^{+}13, SBB^{+}22, VAEP24, Ver91, VHK03, VHK04, YMA^{+}17, YLL19]. meta [BD19]. meta-analysis [BD19]. metabarcodes [CBL^{+}19]. Metabolic [RS10, CBPS^{+}22, XHW^{+}20]. Metabolism [GTB07, ABT^{+}04, GMDD^{+}22b, GMDD^{+}22a, HLCdP19, JTQ^{+}18]. Metabolites [Blu88]. metabolomics [LSD^{+}15]. Metal [TRMV15, CS89, CGD^{+}18, MDL^{+}12, PCD^{+}18]. metals [aHFS92, JP90, KIS^{+}05, TCDPP^{+}23]. metamodel [Woo05]. metamorphosis [RNBP^{+}19]. metaproteomics [XHW^{+}20]. Metazoan [Sol00, HG04]. metazooplankton [CTP^{+}18, KSY^{+}19]. Meteoric [WCC^{+}20]. meteorites [HF65]. meteorological [EMU^{+}23, MFM85, SSVP00]. meter [KMOM88, UCB^{+}18]. meters [NH88b]. Methane [FLDF22, BCP09, FTSF21, WLM^{+}13, dlPHF^{+}15]. methanotrophic [MCB^{+}90]. Method [GAM98a, BBMR19, MCMT^{+}17, VSPP14]. methodological [CP10]. methodologies [BEI^{+}20, KSY^{+}19, Mol04]. Methodology [RLSF06, RLSF07, HRA00]. Methods [LTJ^{+}15, Man04, BCF^{+}03, BIL03, Ben85, BPGD^{+}14, BCG^{+}08, FVA^{+}19, JAS^{+}20, Sei63, WPA^{+}24, WZBK^{+}21]. methylene [KO19]. methylene-interrupted [KO19]. methylmercury [KHBA^{+}24]. metrics [FMC^{+}20, GSFP^{+}09]. Metridia [HL05]. Mexican [Lav09]. Mexico [LM10, AJHC19, BGM^{+}10, Dur09, HKL^{+}15, Ham09, HMS^{+}22, LKDL14, MKSW^{+}15, ON05, ORPRGIS22, OHH^{+}22, PBBH^{+}22, RLP^{+}18, RANS65, RASVB^{+}22, SRFHDH22, UPPS^{+}21, VDGGD^{+}22, WDMMK89, WD94, WWL^{+}24, ZLRVB24a, ZLRVB24b, FGGDF^{+}04]. MHW [Cia22]. MHWs [HZD^{+}23]. Michael [CNT03]. Micro [AIHB^{+}07, AMG^{+}16, Moh15, TLM^{+}17, AMEV07, BTS^{+}15b, CHC^{+}12, DDP^{+}00, YLL19]. Micro- [AIHB^{+}07, Moh15, BTS^{+}15b, YLL19]. micro-nutrient [CHC^{+}12]. micro-particulate [DDP^{+}00]. Micro-phytoplankton [AMG^{+}16, TLM^{+}17]. micro-planktonic [AMEV07]. microalgae [Bon88, DLD^{+}19, LCB18]. microalgal [DSV^{+}24, YNTS22]. microbes [NRA^{+}21, RCC^{+}18]. Microbial [BTG^{+}03, CJ92, FBR^{+}13, GDM^{+}20, MZF^{+}08, VGM^{+}23, VDP^{+}01, VDDA^{+}08, CDS90, DAU22, FBT^{+}22, HSK^{+}19, HMPZ11, KBC^{+}22, LRW^{+}15, LSD^{+}15, MVN^{+}15, MA12, NYL^{+}17, PVG^{+}20, ŠGM^{+}18, SPH^{+}15b, TR99, XYGJ23, XHW^{+}20, XLL^{+}20, dJSL^{+}20]. microbiology [SMB88]. microbiomes [SWH^{+}24]. microcephalus [ON22]. microcosms [SvN04]. Microeukaryotes [LZF^{+}24]. microlayer [Har82, SRG^{+}19]. microlayers [CEF^{+}13]. Micromesistius [MAFS^{+}22]. Micronekton [ATC^{+}19, PBS22, AGL^{+}15, CDL^{+}22, GBB^{+}20, RDG^{+}21, RSW^{+}23, SFS^{+}12]. micronektonic [BCLD^{+}17, NHH^{+}23]. micronutrient [GHG^{+}24]. Microorganisms [Mae88, PCD^{+}18, PAPL15, PPY87, SLG^{+}12, SK91]. microphytoplankton [CCW^{+}18, HLD^{+}21, IPG^{+}16]. Microplankton [LWBD^{+}17, FPS^{+}09, GGQ07, GMAB07, PD15, WYT00]. microplastics [STG^{+}18]. microscopy [LSM^{+}22, LGL^{+}18, PKA19]. Microsetella [KSKN21]. Microstructure [RLSF06, VBAC^{+}21, RLSF07]. Microzooplankton [HPS^{+}01, SDL^{+}19, BFH01, CMC^{+}16, FB01, INO^{+}24, LMS93, LHP^{+}05, MY23, SSH^{+}05, ZBY^{+}22]. Mid [SBK^{+}95, WLM^{+}22, LMS10, Lie88, LS85, MHS^{+}20a, MHS^{+}20b, MBB^{+}96, PS23, PHC^{+}19, RGMPR23, SLPA^{+}20, CSM^{+}15, CB06, KVLA06, SF85, VMB^{+}22a, ZAC^{+}23]. Mid-1980s [SBK^{+}95, MBB^{+}96]. Mid-Atlantic [CSM^{+}15, CB06, KVLA06, SF85, VMB^{+}22a, ZAC^{+}23]. mid-east [Lie88]. mid-eastern [SLPA^{+}20]. mid-latitude [MHS^{+}20a, MHS^{+}20b, RGMPR23]. mid-ocean [LS85]. mid-oceanic [PS23]. mid-shelf [PHC^{+}19]. mid-trophic [LMS10]. Middle [BHPC06, AF10, PVG^{+}20]. middle-out [AF10]. midnight [PNF^{+}21]. Midwater [BM76, BLI^{+}99, HPC^{+}20, HSL96, RBL90]. might [BW08]. Migrant [AGL^{+}15]. migrating [KST03]. migration [ASC07, Ant09, AHRT90, BEP02, CDL^{+}22, EBM^{+}21, FGR^{+}06, FFT^{+}18, GA10, GBB^{+}20, GiIKX22, HGD22, KAK^{+}22a, LRJ^{+}15, MPM^{+}18, OACA20, OAWAN18, PHKS01, RS10, SK17, UCB^{+}18, UPPS^{+}21, VBVYT05, YKNO23, ZGB^{+}20, ZL24, dPGSHL23]. migrations [ACB^{+}13, Ang84, Dom84, DAIS10, Pug84, RAB^{+}84, Roe84a, Roe84b, RB84, RJT84, VLCCP14]. migratory [GHC^{+}17, JOGM^{+}10, SLBVRR^{+}22, Sma10b]. mikimotoi [BTS^{+}15a]. millenium [RSB^{+}01]. millennial [KYS^{+}17]. millennium [PAB^{+}21]. Min [XLL^{+}20]. Mineralization [ÁSÁB^{+}14, ÁBMÁS14, MBCB88, LTSG13, TvG02, ÁBMÁS15]. Mineralogical [IPHW^{+}23]. minerals [ORR^{+}02]. Minho [DGGdR02, MSd^{+}16]. mini [CVHM^{+}18]. mini-mounds [CVHM^{+}18]. miniatuarisation [PKA19]. miniature [BGM^{+}99]. minicosm [LGG18]. Minima [SAM^{+}10, Yu23]. Minimum [Ant09, DLD^{+}19, GEO09, GLV12, HFW^{+}98, KSV08, LGR^{+}02, LQU07, MYH^{+}22, PRP09, RS10, SGR^{+}22, VSGC21, VJJ^{+}22, WGG^{+}08, dB94]. mining [JSLA^{+}21, WFS^{+}15]. minke [FGL^{+}23, HBK^{+}24]. minus [MJWK07]. Mio [GWGR^{+}19]. Mio-Pliocene [GWGR^{+}19]. Miocene [BW65]. Miquelon [BMG^{+}21b]. Mismatch [DLD^{+}19, LHC^{+}21, MMD^{+}16, SIB^{+}06, WBF^{+}21]. mismatched [AHC^{+}13]. mismatching [AHW^{+}15]. missing [KVNT20, VBA^{+}18]. mission [DDJ^{+}21]. mix [Bak01]. mixed [ATS01, Ano94c, BMK12, DHC^{+}20, HTdM^{+}15, LYS^{+}22, MKM93, NF06, PFW15, YTNK00]. mixed-layer [Ano94c, MKM93]. mixing [AUE^{+}14, BIST01, BH07, BCG^{+}08, GvOS^{+}08, GWS^{+}23, HGH^{+}19, HHP06, HLFL23, KY23, Kil76, LFI^{+}13, LCZ^{+}24, MMGL^{+}07, Mar03, MWJ^{+}08, MTK^{+}22, NTU^{+}14, NNM^{+}21, OJB99, ÖÜT93, PTM^{+}22, PIS13, PSGVS^{+}14, PCBA^{+}20, RBF^{+}09, RCSA01, RPPM^{+}23, SÖÜ94b, TYO^{+}14, Tom81a, Tom81b, VMC^{+}19, VBAC^{+}21, WSH^{+}22, XWL^{+}22, XD95, XD96, YYT^{+}14, ZLG17a, ZLG17b, vRGW10]. mixoplankton [LBP^{+}21]. mixotrophic [Sie88]. mixotrophic-cometabolism [Sie88]. mixotrophy [FEGA^{+}14, GLF^{+}17]. mixture [AdAK^{+}18, SBFP21]. mm [Ang79a, Hof81]. mobile [Bak01, FBD18, HF10]. mobility [HHK^{+}02]. MOCI [STGR^{+}14]. MOCNESS [JF13]. modal [Dav85]. Mode [Kat18, ZHD^{+}20, HM08, SW22, SASH08, vFB82]. Model [CCH^{+}12, FPJ^{+}15, KFM15, RSG06, TSJ^{+}12, ABM^{+}05, AKAL20, AGS10, AHC^{+}13, BDP^{+}06, BFPS06, BBC^{+}22, BRB^{+}01, BAARB05, BMGN15, BMB^{+}16, BMC05, BDBJ01, BSH^{+}20, BMG^{+}21a, BLT^{+}08, BCL^{+}09, CDH^{+}13, CMS^{+}13, CLB^{+}14, CHC^{+}12, CJMI^{+}91, Dav85, DJ92, DRE^{+}08, DSBP15, DPF^{+}20, DEW^{+}97, FDHT05, FRCH15, FFA06, FACM^{+}23, Fuk91, GKC^{+}14, GLF^{+}17, Gir15, GWS^{+}23, GBH^{+}20, GiIKX22, Has82, Hea12, HLS^{+}14a, HSC09, HHH^{+}12, HSH97, HMH07, HLFL23, HPW10, iIYO^{+}10, Kag97, KKK04a, KHL12, KDL^{+}01, KLP^{+}17, KDF97, LCBN14, LBC^{+}23, LL97, LSM08, LMS10, LC22, LFG10, LLX^{+}21, LO85, MMGL^{+}07, MHGGS19, MR06, MFS^{+}07, MY92, MKHO96, MFS^{+}16a, MFS^{+}16b, MRW^{+}14, Mor91, MSB^{+}23, MVV^{+}19, MEMC05, MST^{+}23b, NDEG22, NGLL^{+}22, OWH14, PVC^{+}08, PCSMC12, PMC16, PM22, PDAM^{+}15, PGS^{+}22, PVA24, PST^{+}15, PBN13, Pre86, PPD^{+}12, RWD01]. model [RLX^{+}24, RFC^{+}15, RG94, RBS^{+}09, SBMB18, Sak86, SGLF^{+}13, SGWF^{+}19, SGMVF14, SKWWGV18, SPSV^{+}20, SS17, SBFP21, SZG06, SBLA10, SO91, SCLS10, SHC^{+}06, SHC^{+}07, SRM^{+}10, Ste12, Ste91, SDJ14, SOA^{+}23, Suk88, SMP07, TGJT09, TS10, TAH^{+}11, VVV21, VMH^{+}21, VMN08, VBM21, WCB20a, WSL20, WSO^{+}13, WSS15, WLM^{+}22, WPB05, WJPHB15, WWSJ07, XYL^{+}22, XRC^{+}15, XY21, YJS86, YWUK15, YFY05, YLY^{+}14, ZCD08, ZCH^{+}17, dJSL^{+}20, vFB82, Mau10]. model-based [CHC^{+}12, PCSMC12, SOA^{+}23]. model-data [HHH^{+}12]. modeled [REG^{+}15, ZLZ^{+}17]. Modeling [CHS^{+}24, DPGC14, FY88, FUOG^{+}16, JX18, KM08, LSM08, LPF^{+}18, MFS^{+}07, MSF^{+}07, MAB^{+}11c, MAB^{+}11a, MAB^{+}11b, PG13, RDD^{+}18, SMPC^{+}12, TMÁGC^{+}21, VMH^{+}21, ZMW^{+}23, ZLRVB24a, ZLRVB24b, AUE^{+}14, BFP^{+}18, BRR^{+}12, BhTW10, Bre82, BCG^{+}08, CBB^{+}19, CSBL^{+}15, CLX^{+}20, CNBD21, CRiI^{+}15a, DFC^{+}21, DGH^{+}20, FPJ^{+}15, FFA06, FTG^{+}11, GRS08, JAC^{+}12, KSK^{+}15, LL21, LGD^{+}20, MPM^{+}18, MPB^{+}23, PKP14, RCÁS^{+}15, RKC^{+}10, SJA^{+}23, TTB^{+}08a, TTB^{+}08b, VPS09, WH89, YCP^{+}12, ZDG^{+}21, ZHF^{+}24, GAF15]. Modelling [BECA22, BS02, BTJ^{+}17, Den03, DFM^{+}15, DTKvH15, EAB^{+}23, Fly03, GGJ^{+}10, Ham90, HF10, LWY07, MMIB10, OH94, Peñ03a, PWZ^{+}16, RTBR^{+}22, SVL^{+}23, SE16, SW01, TTMM^{+}17, WBH15, APN^{+}15, AK97, BFPS06, BMM97, BMG^{+}19, CFM^{+}18, CPHR98, CAA^{+}07, CAB^{+}18, CTA16, CTMV^{+}14, CKL^{+}14, DLL^{+}23, Den87, EALF08, FRK^{+}09, FVLC^{+}23, Fly10, Fro93, FK99, FPS^{+}13, GCP08, GAPM16, GMDS20, HPS^{+}01, HM15, HT97, HLP^{+}16, HAA^{+}14, ILI^{+}12, Leh01, LAHI10, LWL87, MRM^{+}14, MPMA13, MR03, PBB^{+}20, PRC^{+}20, PPPdS20, SCPN15, SCHD23, SVIA14, SLY^{+}15, TLH^{+}15, THM^{+}14, TGR05, VAEP24, VHK03, VHK04, WCB^{+}20b, Yux88, dlGFM^{+}23]. Models [DXH^{+}02, KPSA17, AF10, AM10, Arb22, APHGC^{+}22, BBE^{+}15, BDP^{+}06, BPA^{+}21, BEP02, BMB^{+}16, BBM^{+}14, BB10, CP10, DHC^{+}20, ECFT20, FCEZ10, Gir15, GBT^{+}19, GAS^{+}22, GL23, HVS10, HMRA^{+}03, HGB^{+}21, HHDS02, Hof10, HLSX22, JB15, KKO10, LI10, LMS10, LAGM^{+}23, LIH^{+}12, LPF23, LM97, MMG^{+}13, MV10, MD10, MR03, MPC12, MCH^{+}12, NBHM01, OWH14, ORVES17, PRC^{+}20, PTF10, PYKF15, PPdS21, RRLS22, SSL08, STM10, SBLA10, SA97, SHG12, SAB^{+}22, TR99, TSL10, TMR^{+}21, TSJC07, WFJ^{+}15, WLM^{+}22, WBB^{+}01, XD95]. modern [GSPP^{+}20, NF87, PRP09, WMWR08]. modes [DAF^{+}22a, DAF^{+}22b, Hut87, LG22, OAM00, RFFL21]. modestus [BB65]. modification [DJW^{+}18]. modified [Iwa23]. MODIS [LHE^{+}13, SNS^{+}22, WM13]. MODIS-Aqua [WM13]. Modular [ZNR^{+}24, BFPS06]. modulate [TDL^{+}17]. modulated [CZW^{+}22, MH14]. modulates [BAM^{+}09, CTI^{+}19, WLL^{+}23]. modulating [HYM^{+}12, JJL^{+}19, TSAM^{+}22]. Modulation [ABS^{+}20, ZLC^{+}15, CSW96, KZD^{+}19, MM99, RCSA01, RR01, VTGC19, Wu13]. Molecular [HCV^{+}20, LAGM^{+}23, JZZY24, KHJ^{+}10, LCB18, WPA^{+}24]. molecular-scale [LCB18]. molecules [SGL^{+}17]. Mollusca [BHB^{+}19, RVC^{+}13]. mollusks [CES^{+}19, Kam19]. Molokai [SM65]. molt [Joh04]. molting [YHRT22]. momentum [CR97]. Monica [CJ92, Hic92, LPA92, SE92, VK92, WRS^{+}92, WJE^{+}92, Gor92]. monitor [RKK^{+}21]. Monitoring [DPH^{+}18, BHLU^{+}07, CAO^{+}20, CDB^{+}24, DPR^{+}18, EGP^{+}18, FC05, GFB^{+}15b, GFB^{+}15a, HWB^{+}18, JRR^{+}24, JLS^{+}22, KNS^{+}03, Man69, MBdM^{+}18, MHA^{+}11, Reb02, TFY02, VMB^{+}22a, WPH^{+}10]. Monod [Fly10]. monodon [GRB^{+}08, YCP^{+}12]. Monsoon [Sch83, WHBK05, ABS^{+}20, CBM^{+}21, DL17, HFW^{+}98, LWBD^{+}17, MJA^{+}07, PVV23, RGE22, RVS^{+}21, SM01, SVU02, gWjNfLyD20, JJA^{+}17]. Monsoon-driven [WHBK05, CBM^{+}21]. Monterey [Bre06, BhTW10, HCC02, MCG^{+}02]. monthly [GAM98a, GAM98b, OÁT^{+}05, SS69, VGLCS06]. Moon [Kag97]. Moored [Kvi69, BBL^{+}09, CNSHT15, LPS^{+}19, MZZ^{+}23]. mooring [HFO^{+}22, HHZ^{+}22, LPF^{+}20, PTZ^{+}23, SSM^{+}18, Woo18, YHZ^{+}22]. moorings [KMOM88, LS20]. morhua [AHP19, FKH^{+}13, LSY^{+}14, LNB13, SLY^{+}15]. Moroccan [ST65]. morpho [CFM^{+}18]. morpho-stratigraphic [CFM^{+}18]. Morphological [GA01, KBE^{+}22, FTC^{+}16, GBG05, RK20, Zez90]. morphology [Nie07, RGPB^{+}23, SW92, USH15a, Sma10a]. mortality [GHVG19, KSKN21, LSY^{+}14, MRM^{+}14, MPD^{+}22, TS10]. Morus [SWP^{+}13a]. MOSSFA [BCD^{+}20]. Motility [Sma10b]. motion [GCD97, UGY^{+}22]. motionally [Szu12]. motionally-induced [Szu12]. motions [Ham87, SSL07]. motorway [CMS^{+}13]. Mound [CLG^{+}22]. mounds [CVHM^{+}18, vHMDL14]. mouth [Ken88]. movement [DXH^{+}02, PGS^{+}22, WSS15]. Movements [BMC^{+}10, BGM^{+}10, BHPC06, Par86, SKH^{+}23, WLKM10, YSS14]. Moving [BMGN15, KKK04a]. moving-boundary [KKK04a]. Mozambique [KM10]. much [GCCY^{+}14, NGLSSG14, Nof00]. mucilage [KS15]. mucronata [RBPGJ^{+}20]. Mud [JJA^{+}17, ZMCD11, LCJ^{+}07]. muddled [KV18]. Muggiaea [BLCL14]. Mullin [CNT03]. Multi [AK97, ALM^{+}23, HHH^{+}12, KC02, LYM12, MVC^{+}11, SLPA^{+}20, BDB^{+}04, BRR^{+}12, CSS^{+}21, Fly03, GRDS10, KSB^{+}22, LSB^{+}17, LSD^{+}18, ROBRB^{+}22, RF17, iSIS02, SSW^{+}09, Tom81b, VVV21, WCX^{+}21, WZC20, ZSW^{+}22]. Multi-decadal [HHH^{+}12, LYM12, CSS^{+}21]. multi-fleet [GRDS10]. multi-layer [ZSW^{+}22]. multi-model [VVV21]. multi-nutrient [Fly03]. multi-parameter [Tom81b]. Multi-platform [KC02, WZC20]. Multi-satellite [AK97]. Multi-scale [SLPA^{+}20, WCX^{+}21]. Multi-sensor [MVC^{+}11, BRR^{+}12, iSIS02]. multi-systems [LSB^{+}17, LSD^{+}18]. Multi-taxa [ALM^{+}23]. multi-trophic [KSB^{+}22, SSW^{+}09]. multi-use [ROBRB^{+}22]. multi-year [BDB^{+}04, RF17]. multicontaminated [CGD^{+}18]. Multidecadal [BGMP03, KYS^{+}17, XYK^{+}22, YS15]. Multidecadal-scale [BGMP03]. multidimensional [RGB^{+}17]. Multidisciplinary [PV18, DVL^{+}99]. multifleet [HSC09]. multifrequency [BOG20, ON22]. multiparameter [ÁBMÁS14, ÁBMÁS15]. Multiple [Kno04, LFCSV^{+}13, BRR^{+}22, BGB^{+}08, CBB^{+}22b, DIM09, FSVL10, GCP08, KWI20, ORVES17, WWL^{+}24, ZCV^{+}19]. multiple-inertial [KWI20]. multiple-scale [DIM09]. multiscale [DDJ^{+}21, LM18, Zav99]. multispecies [HSC09, TMR^{+}21]. Multivariate [PTG95, STGR^{+}14, Dom84, PJH^{+}15, RGB^{+}17, YPGE^{+}10]. multiyear [Lav09, PFW15]. munida [GRB^{+}08, ACK^{+}13, CSMGS19]. Munidopsidae [Mar20]. Munidopsis [Mar20]. Munnopsidae [MB20]. murphyi [BHH^{+}16]. mussel [VMB^{+}22a]. mussels [MCB^{+}90]. mutabilis [RMB^{+}01, WDK^{+}01]. Myctophidae [EBM^{+}21]. myctophids [BM76, SCS^{+}18]. mykiss [WBF^{+}21]. Myriotrochidae [MMK19]. Mysids [Roe84a, VDGGD^{+}22]. mystery [LVGH^{+}15].

N [AAM^{+}14, BPP^{+}98, CBB^{+}02, HM98, RKFD07, WBH15, BC19, Ang79b, BM76, BHC^{+}18, BDBJ01, CBB^{+}02, CGC^{+}20, CSC^{+}12, CFG07, CF12, CMF15, DVB^{+}18, FC07, FMSBW13, GMAGH^{+}17, GBC^{+}00, GLS08, JJA^{+}08, MSJ^{+}15, MM80, OMK^{+}22, PHK^{+}17, SKH^{+}23, SAA^{+}15, WNNI21, WRS^{+}92, ZCD08, vWHdS^{+}98, SWT^{+}17]. N-50^{o} [CBB^{+}02]. N.E. [CLG^{+}00, FC05]. N.W [vWHdS^{+}98]. N.W. [BHE^{+}98, BPP^{+}98, FLdST98, HM98, LDB^{+}02, LHEB98, OB98, PS98, TvW98, dWDB^{+}98]. Name [Ano63b, Ano64c]. Namibian [BCP09]. nannofossil [Gal17]. nannoniscid [JGB20]. nano [AMEV07, BAM^{+}09, BTS^{+}15b, CCW^{+}18, DDP^{+}00]. nano- [AMEV07, BAM^{+}09, CCW^{+}18, DDP^{+}00]. nano-phytoplankton [BTS^{+}15b]. nanophytoplankton [LGL^{+}18]. Nanoplankton [TB15, FEGA^{+}14]. Nanoplanktonic [BM07]. nanotechnologies [Moh15]. NANSEN [vAB96, EFC^{+}23, JMSB^{+}23, SBK^{+}95]. Nargis [MMR^{+}12]. natal [BCL^{+}09]. National [VDGGD^{+}22]. nationwide [JRR^{+}24]. Natural [Iwa23, MKSW^{+}15, Tit20, BM01, BhTW10, LSW^{+}21, RLDC^{+}13, RPG^{+}18, RLSF06, RLSF07]. naturally [TCDPP^{+}23]. Nature [KAH^{+}16, HMWM00, LPF^{+}21, LPBM17, MJ88, MB01, Pra97, SRAV19, TJ90]. naupliar [IMHL07]. Navigating [Bak06]. Nazaré [EvdZSH02]. Nd [GSPP^{+}20, YAI^{+}14]. neap [Car97a, JS90, TDL^{+}17]. neap-spring [JS90, TDL^{+}17]. Near [DSC^{+}21, HHK^{+}02, KWI20, Lie88, PTM^{+}22, SSL07, VK90, BDBJ01, Bri83, CHSB^{+}21, DXH^{+}02, DW02, FWL^{+}15, GL06, Hut81, ICB^{+}19, LW85, LZL^{+}22, LCZ^{+}24, LS12, PS23, Pra91, RDL^{+}91, RCSA01, Tho77, TBW00, TDL^{+}17, VSC01, WAH^{+}20, WLCG23, XWL^{+}22, YMI88, YMK^{+}04, ZQWP23, vHVAT22]. Near-bed [HHK^{+}02, DXH^{+}02, WAH^{+}20]. Near-bottom [VK90, VSC01, XWL^{+}22, YMK^{+}04]. Near-inertial [KWI20, Lie88, PTM^{+}22, SSL07, LZL^{+}22, LCZ^{+}24, WLCG23]. near-seafloor [TDL^{+}17]. near-shore [Pra91]. Near-surface [DSC^{+}21, Bri83, GL06, ICB^{+}19, TBW00]. near-uniform [FWL^{+}15]. nearby [ORMR^{+}19, WLP^{+}21]. nearly [RPRCAG^{+}21]. Nearshore [SOWS17, BEP02, BSA06, DWC06, GLH13, aHFS92, PAM^{+}88, RL23, Sei63, ZCV^{+}19, ZL24]. necrophagous [Thu90]. Need [ALV^{+}21, RAB^{+}11]. needs [KA94]. negative [KWI20, KV13, KSK21]. negentropy [DFD23]. neighboring [KSS^{+}23, YNMY23]. nekton [BMN^{+}99, INI^{+}17, Pea02]. Nematode [LdSH^{+}15, FSVL10, HVTV22, IVR^{+}13, MSFZ19, RLR^{+}18, UAM05]. nematodes [GVKD^{+}13]. Nemerteans [CP19]. NEMURO [iIYO^{+}10]. Neocalanus [MFB^{+}84, CSK^{+}12, DOS^{+}18, FMT15, Gif93, Kli10, LGK^{+}93, MFB^{+}84, Mil88]. neodymium [DDCE^{+}23]. Neoglacial [KiL14]. neon [IIS^{+}17]. Nepheloid [OVR^{+}02, CLG^{+}00, DL17, GRMB18, HLS^{+}14b, PdMS^{+}13]. nephelometer [GRMB18]. nepholometer [VK90]. Nephrops [LML^{+}23]. Neptunism [AB65]. NESSAS [Dri11]. nested [BAARB05, CGZ^{+}16, SZG06]. nested-grid [SZG06]. Net [BKD^{+}20, Whe93, BCOL^{+}19, CKT^{+}13, JTQ^{+}18, MST^{+}23b, OKdA^{+}19, SMP^{+}22a, SWP^{+}13b, TBS^{+}19, VEM^{+}21, WB03, dlGFM^{+}23]. nets [JF13]. network [AHC^{+}13, CGS23, HMRB^{+}03, KLC^{+}15, LPF23, MGC^{+}18, MDR20, RZTD17, YPGE^{+}10]. networks [KM22, TMKJ^{+}09]. neural [HMRB^{+}03, KM22, LPF23, YPGE^{+}10]. Neutral [McD88, MJ88, SBC^{+}16]. Neutral-surface [McD88]. Newfoundland [HMP^{+}13, CGV13b, DM13, HBL^{+}13, LNB13, PCM11]. Newfoundland/Labrador [CGV13b]. newly [GKS^{+}13, LPW^{+}23]. newly-discovered [GKS^{+}13]. newmani [HLPL05, LPHL^{+}05a, PLHLF05]. Newport [HWS^{+}07]. Niche [BGA^{+}21, PMH17, BMG13, DFM^{+}15, DFH^{+}16, EBM^{+}20, GVBV^{+}21, KBE^{+}22, LAGM^{+}23, NHE^{+}13, RPRCAG^{+}21, TMÁGC^{+}21, WLP^{+}21]. niche-environment [KBE^{+}22]. niches [GRB^{+}08, HLD^{+}21, SBC^{+}16, XWL^{+}18, ZLX^{+}20]. night [BRD^{+}15]. Nile [CMF11, Ore69]. Niña [ABP15, BCM^{+}02, FWBC02, HCC02, LJPGC02, LCGH07, LB02, MCG^{+}02, PK02, PKF02, RKS01, SMdG02, TTB^{+}08a, WW02, WDMC02]. Nino [Pea02, FWBC02, LB02, PK02, PKF02, ACN01, ABP15, BWMGCB08, BCM^{+}02, CBB^{+}02, CCW^{+}02, CSW96, CPC^{+}02, CCHM02, CCMS08, CCA^{+}02, CPNL07, CW02, DW02, FGGDF^{+}04, GR17, GdRGL^{+}01, HLK13, HMB^{+}86, HCC02, HSF02, Kos02, KC02, LJPGC02, LBH^{+}87, LGR^{+}02, LO21, LCGH07, LPF^{+}18, LPARF^{+}20, MG02, MCG^{+}02, MBH^{+}01, NFMCS^{+}22, PBBH^{+}22, SMdG02, SPB^{+}02, SJ02b, SJ02a, STR01, SKT01, TTB^{+}08a, WW02, WLM07, WDMC02, YYK^{+}12]. Niño/La [PKF02]. Niños [SHF01]. Nitrate [CKT^{+}13, CCH^{+}12, GGPG^{+}19, LCR^{+}93, RCSHW22, WCX^{+}21, Whe93]. nitrite [MBP^{+}11]. Nitrogen [BGS^{+}04, WP91, WDMMK89, WFBN^{+}13, BHE^{+}98, BNC05, BAOC^{+}07, DHDM22, FUOG^{+}16, FFA09, HW02, HOY^{+}21b, Kli10, KSS^{+}23, KAAK^{+}16, LFBR^{+}18, MBP^{+}11, MRAP22, NHH^{+}23, PKV18, PLN^{+}23, RDD^{+}18, SIS^{+}14, TFM03, VKGP^{+}13, WMC^{+}89, WNNI21, Whe93]. nitrogenous [KNI^{+}05]. Nitrous [dlPHF^{+}15, dlPPÁB24]. nitzschia [PTPY^{+}23, VSPP14]. NL [CLMR23]. No [IL20, ROBRB^{+}22]. no-take [ROBRB^{+}22]. NOAH [NF87]. nocturnal [LSIB23]. nodule [BJMP19, BJMP20]. nodules [Hey78, VCM04]. NOI [SMG02]. noise [Pie01, RFS10]. noltei [AR18]. Non [FCMCÁS19, GSC^{+}20, MHGGS19, Ang79a, CPNL07, FVA^{+}19, LPF^{+}18, McD81a, MMF^{+}12, PSM^{+}22, PCR^{+}22, RWD01, RGC^{+}01, SL13, TMN^{+}12, Tom81b, TIOM16, WLP^{+}21]. Non-carnivorous [GSC^{+}20]. non-El [CPNL07, LPF^{+}18]. non-indigenous [SL13]. non-invasive [FVA^{+}19]. non-isopycnal [Tom81b]. non-large-meander [TMN^{+}12]. non-linear [McD81a]. non-local [PSM^{+}22]. non-normal [TIOM16]. Non-Redfieldian [MHGGS19, FCMCÁS19]. non-Russian [Ang79a]. non-stationary [PCR^{+}22]. non-steady-state [RWD01, RGC^{+}01]. non-upwelling [MMF^{+}12]. non-vent [WLP^{+}21]. Nonlinear [Bre06, CSS11, GXX^{+}22, GLPC23, Li14, McD81b, vHMDL14]. NORCAN [DP13]. Nordic [HØ00, AHSS22, BLP^{+}20, BGV^{+}23, FWL^{+}15, HØH^{+}03, JJA^{+}08, MRO^{+}08, MM01, NBHM01, SH09, SAH^{+}21, YS15]. normal [LWBD^{+}17, TIOM16]. North [ALT10, BLAM00, BBM^{+}14, BMC^{+}10, CSC^{+}12, DGMM85, DMML88, FGSA97, GLF^{+}17, GSF^{+}15, GLLB22, HBL^{+}13, HMP^{+}13, HCV^{+}20, LZCZ05, MHS^{+}20a, PGC^{+}96, SIR^{+}07, Sei63, Tom81a, AAMB^{+}24, BFH01, BAOM^{+}12, CGC^{+}20, GL06, GdRGC^{+}14, HPB^{+}09, HFO90, LG22, LSIC12, MNS^{+}24, MTC12, PLK14, RBR^{+}23, RWOA01, STC10, SAA^{+}15, SVU02, SKCP23, WDC^{+}11, ZGB^{+}20, vAB96, ABD^{+}17, AHP19, APC^{+}12, Ang84, ABSDC07, AFH^{+}11, ALM^{+}23, BM76, BRB^{+}01, BZD^{+}21, BDTC15, BGMP03, Bea04, BLHB07, BLAM98, BGM^{+}99, BBL^{+}18, BDBJ01, BHLU^{+}07, BMG^{+}21a, BMN^{+}99, BKB85, CRGA17, CPG^{+}18, CMJPH^{+}18, CBB^{+}02, CBB^{+}22c, CCHM02, CAT^{+}08, CSK^{+}12, CHSB^{+}21, CF07, DHC^{+}20, DN07, DML^{+}16, DL69, DLM^{+}96, Dom84, DLC^{+}08, Dri06, DK07, ECGP01, ED82, FCMCÁS19, dCFK17, FKZ^{+}15, FvBA^{+}17, FMP19, FJH10, FW91, FMW91, GHF^{+}21, GMDD^{+}22b, GTB07]. North [GKR20, GdRGL^{+}01, GPC^{+}03, GC09, GLS08, HSS^{+}12, HPC^{+}20, HVRR15, HØ00, HM00a, HMWM00, HFNG00, HDZY15, HRSM08, Hea12, HBR11, HMPZ11, HKPV12, HGBG20, His22, HAA^{+}14, HHSR07, HHP10, HBH^{+}17, HMH07, HPNDC15, HMA18, HWS^{+}07, IIS^{+}17, IHT^{+}21, IMM^{+}22, iIRM^{+}15, IU14, JAS^{+}20, JST^{+}24, JJS03, KRL08, KST^{+}10, Kat18, KKNT23, KY15, KDL^{+}01, KJH^{+}22, Kos93, KAH^{+}16, KDB95, KFM15, KTW^{+}22, KYS^{+}17, LS20, LYM12, LBSP01, LLH^{+}21, LG23, LSMG01, LM14, LGZL22, LOC95, LBF^{+}22, MMG^{+}13, MRM^{+}14, MHS^{+}20b, Man04, MEMP15, MMF^{+}17, MHR^{+}10, McC92, MS17, MSI17, MRH^{+}14, MBF^{+}14, MS00, MM99, Min00, Min02, MTK^{+}22, Mol04, Mol22, MR03, MNFY21, MW96, NEI^{+}22, NHH^{+}23, NCH^{+}07, NJCD01, NMO^{+}21, NGNV12, NRA17, OTNI20, Ola65b, OOTA15, ORMR^{+}19, OMS^{+}15, OAD22, OAM00, ORMB08, PMG15, PVC^{+}08, PL01, PBB^{+}12a]. North [PDV12, PBB^{+}12b, PMDR06, PJH^{+}15, Pie01, Pug84, Qiu15, RWD01, RSB^{+}01, RM93, Rei94, RCD^{+}94, Ric85, Ric93, RDC^{+}21, RAB^{+}84, Roe84a, Roe84b, RB84, RJT84, RM89, RHM^{+}19, SGL^{+}18, SCLG^{+}11, SVHM^{+}13, SHK^{+}14, STPHM^{+}23, SAM^{+}04, SBH^{+}14, Sek99, Sme93, SLGI^{+}21, SIS^{+}14, SBD01, SJD10, STR01, SKH00, SASH08, SKT01, SD07, TOKLC08, TŠT^{+}17, Tom81a, THM^{+}06, THM^{+}14, TRMV15, UKM^{+}14, UBB^{+}23, UB10, VBL^{+}21, VHV^{+}12, VFS^{+}15, VR03, WLD^{+}15, WHS17, WFD^{+}07, WSO^{+}13, WO85, WNNI21, WFBN^{+}13, WBB^{+}01, WJPHB15, WZ04, WWSJ07, Wu13, WG82, YYT^{+}14, YMA^{+}17, YTB^{+}21, Yas07a, YSD15, YAK^{+}08, YBPS08, YYK^{+}12, YFY05, YSN20, YMK^{+}04, ZLZ^{+}17, ZBLF23, vAB96]. North-East [ALT10, CSC^{+}12, CPG^{+}18, PMG15, ABSDC07]. north-eastern [HPB^{+}09, vAB96]. north-west [RWOA01, GKR20, Tom81a]. north-western [AAMB^{+}24, LSIC12, MTC12, ZGB^{+}20, SGL^{+}18, SAM^{+}04]. Northeast [FKH^{+}13, HHW22, KFM^{+}17, VSC01, AGL^{+}15, BP02, BKD^{+}20, CZG^{+}21, CGS23, CR20, DCKB13, DMBB02, DAKV99, KJZ^{+}12, KT04, KRHS14, LFG10, PFHM16, PB07, PP85, RTN90, SMdG02, SMG02, SJ02a, Thu90, dPCS23, ARG11, BMNW01, BSH^{+}20, CTKF^{+}23, CPHR98, CZL^{+}24, CBOP15, DBC^{+}18, DSC^{+}21, DP13, FMSBW13, HBD^{+}21, Has06, Hau18, HHW01, JSdSS^{+}21, JLB^{+}08, KYT^{+}16, KSR^{+}01, KRL^{+}22, LBK^{+}01, LMP22, MBT07, MMMWZ23, MFB^{+}84, MHCS^{+}23, NXT^{+}17, RGC^{+}01, Rea00, RSK^{+}23, RFKC16, RGM01, SLOP^{+}22, SBFP21, SLY^{+}15, SBE^{+}20, TWBC^{+}13, Tit20, WHBW03, WDK^{+}01, WXH07, XNT^{+}17]. Northeastern [GD85, BDG^{+}17, HLM^{+}16, PCD^{+}18, SDP^{+}22, VKJ^{+}23, WLP^{+}21, JG07, SSVP00, VPW01]. northeastward [FZ88]. Northern [DMF^{+}09, FAAV^{+}15, Mar20, MBD^{+}09, WO85, ABE^{+}15, AOMZ^{+}23, AJV^{+}02, ABP15, BLT^{+}08, CIL^{+}23, CWZ^{+}20, CSM^{+}15, CSMGS19, CVBG21, CBGC^{+}08, CCM^{+}14, CCH^{+}12, CBT07, CS04, DCD^{+}23, DML^{+}16, DBR03, DWNN04, DJG^{+}02, DSR21, DFC^{+}21, EFC^{+}23, EBvdL^{+}09, FJA^{+}21, FUOG^{+}16, FMC^{+}20, FVLC^{+}23, GIHJ23, GC14, HBV^{+}10, HBK^{+}24, HSC09, HW02, HOY^{+}21a, HPZC21, HMH07, HHZ^{+}22, HSF02, IHY^{+}01, JCM^{+}21, JMSB^{+}23, Kli10, KLP^{+}17, Kos02, KS15, LDAM^{+}07, LBP15, McC92, MZZ^{+}23, MMD^{+}16, MFM15, NEI^{+}22, ON05, OVR^{+}02, PS23, PK02, PBS22, PCH08a, PVA24, PST^{+}15, PCH^{+}08b, QLW10, RCS^{+}11, RGE22, RDC^{+}21, RB20, SE16, SSTL16, SR15, SPB19, SD07, SBC^{+}24, SYN^{+}21, TVT^{+}23, TMÁGC^{+}21, VOG^{+}08, VBM21, VKJ^{+}23, VOJD02a, VOJD02b, WD94, WST^{+}21, WHS^{+}23, WL16, WZC20, XLX^{+}20, YBS^{+}01, YPGE^{+}10, YGC^{+}21, YHZ^{+}22, ZHD^{+}20, ZDG^{+}21]. northern [dFHR^{+}24, ASFB^{+}13, BH85, BPSN^{+}21, CBB^{+}19, CBB^{+}22b, CRT^{+}22, DMML88, DLJ^{+}21, FWO15, FFA06, GCED22, GCFS06, HMRB^{+}03, HWF^{+}21, JBB^{+}14, MPC^{+}17, OBD^{+}20, PL09, RBE^{+}12, SMG02, TTB^{+}08a, TTB^{+}08b, ZLRVB24a, ZLRVB24b]. Northern-Boundary [WO85]. Northward [KAK^{+}22a, Nof03, SE08, SE09]. northwest [AS20, HHK^{+}22, HWLT10, HHH^{+}12, HG04, LPARF^{+}20, MMIB10, MCGR07, MM80, Mit91, SHS^{+}05, TLH^{+}15, TLP^{+}16, YJS86, BIST01, BHB^{+}19, BAB^{+}19, BHMS09, CGV13a, DP13, GM19, JS21, JM19, JLRB20, JCIG18, Kam19, KKS^{+}19, KHP^{+}18, LFC^{+}15, MH02, Mit83, NB87, OMK^{+}22, ORR^{+}02, QPR03, SPK^{+}22, VNMS91, WMB^{+}21, WAH^{+}20]. northwestern [ANH21, AJHC19, BRG^{+}23, BGA^{+}21, Car98, CLL^{+}18, CBD^{+}24, DZ04, FVA^{+}19, FEGA^{+}14, FMWW14, GCC^{+}24, GIPG17, KCL^{+}12, KiL14, KMS^{+}24, LYZ16, lLdZQ^{+}22, LSMG01, Mar20, MGA^{+}23, MZZ^{+}23, MMG^{+}11, OT19, OACB^{+}15, Owe91, PGLG^{+}05, PAG^{+}18, RLP^{+}18, SGL^{+}13, SHP^{+}23, San73, SM21, SvWRvB02, SCC14, STS^{+}12, WCS^{+}23, ZZPL18, dlGFM^{+}23, CGM^{+}02, CNBD21, CFC^{+}18, CdMS^{+}18, HLM^{+}13, JMG^{+}13, LGL^{+}18, SGA^{+}19, TMPM^{+}16b, TMPM^{+}16a, Tol85a, TCL^{+}15, WH20, ZFSV^{+}09]. Northwind [RCB^{+}20]. norvegica [CNT^{+}19, KSKN21, MPSD15, SIR^{+}07]. Norway [FGL^{+}23, Mid69, MPN09, SST^{+}17]. Norwegian [HBL^{+}13, BMK12, BPNB90, BS95, DCL^{+}13a, DBM17, EBD^{+}20, FDE^{+}22, FAH^{+}13, HBG^{+}21, HMP^{+}13, Leg91, Mos69, NF06, SBK^{+}95, Ste91, SBS90]. notably [dSSDS^{+}20]. Note [Ano03h, Ano03i, Ano19o, Ano07s, Ano08x, Ano10a, BSA06, FZ88, Sud86, Swa76]. notes [BM76]. notice [Swa77]. nov [AM19]. Novel [CCM^{+}13, SPH^{+}15b]. November [Ano04g, Ano11h, Ano12o, Ano14o, Ano15p, Ano16s, Ano18n, Ano19j, Ano20s, Ano21s, Ano22z, Ano22-33, Ano23s, FBB^{+}21]. NPP [VEM^{+}21]. NSCS [XLX^{+}20]. NSCS-RiOMar [XLX^{+}20]. Numerical [AR18, AEPW93, GLPC23, GRS08, HT97, Leg91, Li14, MGKW19, SZG06, TCN20, YN03a, YN03b, Yux88, ASFB^{+}13, Ano94c, BFPS06, BCG^{+}08, CDH^{+}13, CCMS08, FDHT05, HSH97, HYM^{+}24, KRL08, MKM93, PTF10, Pre86, RDL^{+}91, SGMVF14, Ste91, WDMMK89, WD94, vFB82]. nurseries [DFM^{+}15]. nursery [CDTM^{+}21, HvDL^{+}17, RASVB^{+}22]. Nutriclines [SGS^{+}23]. Nutrient [CCW^{+}02, FGL^{+}23, HNR^{+}17, LGZL22, LGD^{+}20, PMDR06, SMP^{+}22b, ZLR^{+}07, AJA^{+}22, AFBT^{+}22, CHC^{+}12, CW02, DHL^{+}21, FCMCÁS19, Fly03, FYYC05, HLR17, HAH^{+}22, ILA21, iIYO^{+}10, JJL^{+}19, JX18, JRW01, JCF^{+}23, KTN14, KMS^{+}24, KSPK99, KNI^{+}05, KZD^{+}19, LDD^{+}22, LFBR^{+}18, LZCZ05, MRA^{+}19, MHGGS19, MHCR^{+}12, NHN^{+}21, NKK^{+}05, PV07, PKV18, ST03, SGS^{+}23, SKP99, SM16, TDGY22, TAM^{+}15, TSP^{+}13, WW02]. nutrient-conserving [HAH^{+}22]. Nutrients [CDH^{+}13, TWAL^{+}11, BMK12, FAH^{+}13, KFC^{+}23, KSC10, LMC^{+}20, LGG18, LDMH09, RBR^{+}23, RDD^{+}18, SK18, SK21, TNY^{+}24, TZP^{+}00, WGZZ19, WDMC02, WXH07, ZGZ19]. nutrition [GLV12]. Nutritional [LDH90, DMC^{+}18, KTN14]. NW [AAM^{+}14, ABT^{+}04, BAOC^{+}09, BMG^{+}21b, CLD22, CQZ^{+}18, EMU^{+}23, KFC^{+}13, LRGV^{+}18, PGT^{+}13, SCAA07, STG^{+}18, VFCC^{+}22, VGM^{+}23, ACB^{+}13, ACL^{+}18, ÁSDB^{+}01, ÁSFP^{+}03, AUE^{+}14, BAM^{+}09, BGM^{+}01, BCOL^{+}19, BBPHG^{+}11, BBRM20, CJRÁ^{+}13, CCHV^{+}21, CHSB^{+}21, CVHM^{+}18, DAU22, DCRR^{+}22, DCL^{+}13b, FBR^{+}13, FARRL^{+}13, FB01, GMAMB04, GASV^{+}09, GHL15, GBB^{+}20, HHB^{+}01, HDM19, HBV^{+}99, HHB^{+}22, IVR^{+}13, JW01a, LSM^{+}22, LB20, LFBP^{+}13, LMP22, MAFS^{+}22, ORW^{+}01, POS^{+}07, PPHM18, PPSVC^{+}13, PGGG17, PRL^{+}18, RCC^{+}18, RCF^{+}13, RÁSG^{+}13, RAE^{+}05, SOS^{+}07, SSB19, SSB20a, SCMAR^{+}99, SVHM^{+}13, SCB^{+}09, VSPP14, dlPPÁB24, vWM02a, vWdSBdH02].

O [BSF95, CFG07, CF12, CMF15, FC07, LCJ^{+}17]. oak [FBR^{+}13]. obesus [HHP10, MRAP22]. objective [HRA00, RGB^{+}17, VBL04, WSO^{+}13]. Objectives [PV18]. observable [SFF^{+}24]. Observation [ILI^{+}12, KGL22, LLH^{+}21, MAB^{+}11b, VBVYT05, DDJ^{+}21, SCY^{+}23, UGY^{+}22]. Observation-based [KGL22, LLH^{+}21]. Observational [BCG^{+}08, DFC^{+}21, QYF^{+}24, Sch03, FRK^{+}09, Kag97]. Observations [BhTW10, LRS^{+}03, LCZ^{+}24, MM80, MKMF^{+}89, PL18, Tho77, WSH^{+}22, XHC^{+}20, APN^{+}15, ABM^{+}05, ALBP87, BIST01, BDBJ01, BPC^{+}05, BRR^{+}12, CSS11, CLY22, CPSM20, CCH^{+}12, CRiI^{+}15a, DHC^{+}20, DOS^{+}18, DLC^{+}08, FRV^{+}19, FGR^{+}06, GBM^{+}01, GTNK21, HTdM^{+}15, HHZ^{+}22, KHD22, LaC08, LS20, LHE^{+}13, LC16, MBS20, MZZ^{+}23, MS00, MKM86, PC87, Pra91, RM97, RR03, Rud15, RKC^{+}10, SJA^{+}23, ŠGM^{+}18, Ste04, SOA^{+}23, SNS^{+}22, TBS^{+}19, TSAM^{+}22, VBL^{+}21, VNMS91, WLM^{+}22, WZC20, WWL^{+}24, XYWY23, YHZ^{+}22, YN20, YLY^{+}14, ZYN^{+}24, vHMDL14, vHCY^{+}20]. observatories [RAB^{+}11]. Observatory [CLMR23, HBD^{+}21, VGM^{+}23, CNSHT15, DOS^{+}18, TAW^{+}15, VGJ^{+}19, SAW^{+}15]. Observed [CGD^{+}22, KSK21, MZH^{+}23, MVS08, PTZ^{+}23, ZLZ^{+}17, CS04, HMRA^{+}03, HZD^{+}23, HBD^{+}18, JJA^{+}08, Kos02, KRHS14, LO85, Man69, McG64, MSS^{+}02, Ric93, TYO^{+}14, VK90, YYK88, ZZPL18, ZBLF23]. observing [HBD^{+}21, MJC^{+}17, SFK^{+}99, VSA^{+}21, Whi95]. obtained [PHC^{+}19]. occidental [MEMC05]. Occurrence [CJG88, GLY23, PAG^{+}18, STG^{+}18, SCCJ^{+}18, DWNN04, IMM^{+}22, MB01, RSW^{+}23, SCB^{+}09, SOWS17, SWP^{+}13a]. occurrences [KT04, MRW^{+}14]. Ocean [AC85, Ang79a, APSC11, BMG^{+}21a, Bre82, DDJ^{+}21, FTG^{+}11, HMRA^{+}03, HBV^{+}10, HHB^{+}22, Ike88, JJA^{+}13, JCM^{+}21, Kra82, KPSB17, KMF^{+}20a, MCL^{+}15, MYH^{+}22, MAB^{+}11c, MAB^{+}11a, MAB^{+}11b, OHH^{+}22, RHB23, SCS^{+}18, SAA^{+}15, SPK^{+}19, SJM^{+}19, UGY^{+}22, ASJ^{+}23, AFBT^{+}22, AASJ23, AK97, AGS10, Ang89, AG22, Arb22, AB65, BBE^{+}15, BB24, BHK^{+}16, BHHR15, BGM^{+}99, BPGD^{+}14, Bro82, BB10, BCG^{+}08, Cai95, CKP^{+}20, CTKF^{+}23, CVBG21, CBB^{+}22c, CSS^{+}19, CP83, CWS^{+}21, CGB07, CEF^{+}13, Dea85, Dur09, ERBV21, EHSI12, EGPM^{+}15, FPD^{+}01, FDHT05, FBS22, FACM^{+}23, FWH^{+}17, FGS^{+}23, Fuk91, FWL^{+}15, GPEV20, GSPP^{+}20, Gar06, GR17, GLAHH^{+}22, GCD^{+}13, GBC^{+}16, GCED22, GCD97, GZCL23, HWPLvW20, HKL^{+}15, HHK^{+}22, Has82, HS07, HDZY15, HMX^{+}23, HM15, HLGA07, HHMB^{+}09, HAA^{+}14, HMKF08, HTdM^{+}15, HLFL23, HWB^{+}18]. ocean [HDB13, Hut87, Hut92, Hut95, HSF02, IGG^{+}19, IPF23, IG19, JB15, JRR^{+}24, JvdLL^{+}15, JSKM02, JST^{+}24, JCF^{+}23, KM08, KC15, KBHML17, KFM^{+}17, KY14, KA85, KMS^{+}24, KLIRK17, Kru19, KRL^{+}22, LVGH^{+}15, LK13, LL97, LRW^{+}15, LMS10, LGZW22, LL21, LM97, LB14, LH89, LS85, MLL^{+}22, Mac98, MBT07, MGE^{+}12, MHS^{+}20a, MHS^{+}20b, MB20, MMR^{+}12, MPM^{+}17, MBS20, MFS^{+}16a, MFS^{+}16b, MJ88, McK15, MFDH22, MLHE23, MWS^{+}10, MFB^{+}84, Mit91, MK12, MJC^{+}17, MZGA^{+}20, NHS^{+}14, NNM^{+}21, NH88b, NBG^{+}05, OWH14, PRP09, PCK^{+}06, PM22, PSK96, Pie01, PZA^{+}15, PAVB^{+}21, PLK14, RM97, RSMIS03, RGI05, RG09, RFKC16, Roo82, RLC85, RA15, RAB^{+}11, SMFM^{+}21, SGLF^{+}13, SOB^{+}08, SSB^{+}20b, SMN^{+}14, SF15, SZG06, SE08, SE09, SBLA10, SON^{+}20, SA97, SHF01, SBB^{+}22, SV97, SMKK21, Ste04, SJD10, Ste91]. ocean [SK91, SBBV04, SW65, SSW^{+}09, STGR^{+}14, Szu12, TMH^{+}16, TAM^{+}13, TAH^{+}11, Tur15, VBL^{+}21, Val99b, VGJ^{+}19, VDS^{+}18, WD94, War06, WGCS13, WSO01, WMB^{+}18, Whi94, Whi95, WLM07, WHI^{+}02, WLM^{+}22, WBH15, Wun24, XLX^{+}20, XYWY23, YFK21, YS15, YZX^{+}23, YSS14, YN03a, Yu23, ZCD08, ZKT88, ZL01, ZNR^{+}24, ZHF^{+}24, ZL24, ZHBW01, ZBLF23, dMM69, dHA^{+}04, dPCS23, vHCY^{+}20, AMFY20, AMY^{+}23, AdAK^{+}18, ÁSÁB^{+}14, ÁBMÁS14, ÁBMÁS15, AKAL20, ATC^{+}19, Ano94c, AFH^{+}11, AGL^{+}15, AvD15, APHGC^{+}22, AT07, ARG11, ABP^{+}23, BBE^{+}15, BSC^{+}19, BLR^{+}23a, BGMP03, BP02, BSF95, BvdLA^{+}11, BGM^{+}10, BLHB07, BEH19, BMdMS^{+}21, BKC15, BHS^{+}15, BHC^{+}18, BC16, BS95, BS02, BLES16, BTJ^{+}17, BGWP^{+}17, BTV^{+}17, CSR90, CKM^{+}21, CPC^{+}15, CVBG21, CSLJ03, CGZ^{+}16, CGS23, CKT^{+}13, CDP14, CRPS^{+}15, Dag93, DAvD^{+}21, DYO^{+}10, DML^{+}16, DPB06]. Ocean [DCKB13, DMC^{+}18, Don65, Don87, Don94, DS65, DMBB02, DBRK17, FPD^{+}01, Fai65, FGS^{+}15, dCFK17, FPJ^{+}15, FACM^{+}23, Fre07, Fro93, FWL^{+}15, GLY23, GWB14, GMAGH^{+}17, GTB07, GCV^{+}24, GBC^{+}00, Gif93, GLAHH^{+}22, GGAA^{+}23, GWGR^{+}19, GTNK21, GBH^{+}20, GCS91, GSF^{+}15, GLV12, GVKD^{+}13, GMDS20, HLR17, HSK^{+}19, HPB^{+}09, HVTV22, Hay65, HBR11, HGPFN^{+}14, HGTP^{+}19, HMO^{+}13, HSLG11, HKE^{+}10, HBW17, HCV^{+}20, HFK03, HLTB^{+}17, HHP10, HK65, HPH^{+}08, IMM^{+}22, iIRM^{+}15, ISH^{+}04, JS21, JAC^{+}12, JHM^{+}22, JG07, JCM^{+}21, JPIP22, Kam19, KRL08, KPM^{+}23, KSR^{+}01, Kiv97, KAK^{+}22a, KS06, KH09, KAH^{+}16, KB65, KTW^{+}22, KMF^{+}20b, LJ65, Leh01, Let87, Lev88, LMH^{+}13, LMT^{+}19, LdSH^{+}15, LM14, LSW^{+}21, LLX^{+}21, LS15, LB14, MMR^{+}09, MNS^{+}24, MGWZ20, Mar20, MRO^{+}08, MR06, MRAP22, MOS^{+}13, MST^{+}23a, MMF^{+}17, MKM93]. Ocean [MFS^{+}16a, MFS^{+}16b, MBKS08, McG64, MRH^{+}14, MKSvA^{+}22, MPSS91, MHH^{+}15, MN88, Mil88, MC88, MS00, MFA^{+}15, MTK^{+}22, MFM85, MGK^{+}86, MKOLA20, MJD^{+}21, MCMT^{+}17, MVS08, MCH^{+}12, MW96, NJCD01, NXT^{+}17, NMY^{+}14, NYH^{+}22, Nof03, NKK^{+}05, NRA^{+}21, NST^{+}23, OMK^{+}22, OWR^{+}07, OPG^{+}10, OT19, OKdA^{+}19, ORMR^{+}19, PMG15, PFHM16, PTP^{+}22, PPVG12, PPKR14, PAPL15, PMH17, PLEF^{+}23, PS91, PYKF15, PGS^{+}22, PP85, PHCA17, PFE10, PAF^{+}11, RWD01, RSB^{+}01, RJO^{+}19, RFFL21, Rei86, Rei89, Rei94, Rei97, Rei03, RCB^{+}20, RKFD07, RTN90, RGE22, RM89, RFPG15, Rud15, RKS^{+}15, RHM^{+}19, SSB19, Sai65, SCLG^{+}11, STPHM^{+}23, SBC^{+}16, SLOP^{+}22, SJA^{+}23, SBK^{+}95, SM01, SDS02, SDS22a, SKWWGV18, SVU02, SAM^{+}10, SWT^{+}17, SW22, SSN23, SR15, SEW11, SO91, SQJ^{+}17, SPB93, SMP07, SW21, SKCP23, TFY02]. Ocean [TKSI08, Thu90, TLM^{+}17, TWBC^{+}13, TNY^{+}24, TAM^{+}15, TTF^{+}22, TBW00, TSJ^{+}12, Tsu86, TRMV15, UB10, VSA^{+}21, VDP^{+}01, VDDA^{+}08, VKGP^{+}13, VMA^{+}24, WMC^{+}89, WHS17, WH20, WWW^{+}23, WMWR08, WTT14, WKS^{+}15, Was15, WSC^{+}21, WLM^{+}13, WSG^{+}93, Whe93, WCN^{+}05, Wil65, WC15, Wis65, WHS^{+}23, WG82, XC14, XWW^{+}21, YGMR^{+}23, ZSI^{+}05, ZPY^{+}20, ZBY^{+}22, ZQWP23, ZJZ^{+}21, dZTG05, dPGSHL23, vAB96, vWMH98, PS23]. ocean-climate [STGR^{+}14]. ocean-color [McK15]. Ocean-facing [TTF^{+}22]. ocean-shelf [Hut92]. Oceanic [BM86, CR97, HFO^{+}22, LM10, PMG15, RG03b, RK03b, RR03, BH07, BCLD^{+}17, BT07, CLV^{+}19, CTI^{+}19, DFC^{+}21, Emi65, FELMGM^{+}22, FK86, Fra69, Gif93, HTG15, Hol00, HRA00, Jac10, Joh04, Kaz17, Kun03, LDHW20, MSMR93, MM99, MTH^{+}10, MCMT^{+}17, OCH^{+}18, PS23, PELAA18, PTZ^{+}23, PK02, PFE10, RMK^{+}21, RL85, SMdG02, SJ02a, TBS^{+}19, UBB^{+}23, XHW^{+}20, YT06, dlHRA^{+}18, RG03a]. Oceanogr [ÁBMÁS15, BJMP20, FDH20, GFB^{+}15b, JLP^{+}20a, KN11, KMF^{+}20a, MHS^{+}20a, MFS^{+}16a, RBS^{+}20, SE09, VH09a, WF07, Yas07b, ZLRVB24a, dMGS^{+}11b]. oceanographer [CNT03, YRKC08]. Oceanographic [BPF06, BECR^{+}22, GAM98b, IPD14, MVBC^{+}21, Ore69, QCdS^{+}07, RAG^{+}19, TNY^{+}24, VKDS^{+}18, AKAL20, BL02, BASS^{+}20, BPSN^{+}21, CBB^{+}02, CGC^{+}20, CdTH^{+}16, DMBHG10, DMF^{+}09, DCL^{+}13a, DB02, EKB06, EBR^{+}14, Gar03, HGBG20, ICB^{+}19, JST^{+}24, KJZ^{+}12, Kvi69, LJM^{+}16, MNT14, MLL^{+}15, MMES16, MIH06, Mit91, MWFH02, MGH^{+}07, MBD^{+}09, MKSW^{+}15, MMN^{+}24, NRS^{+}19, NM17, NBLI20, PMM^{+}23, PGY^{+}22, PC87, PRC^{+}20, REG^{+}15, RNBP^{+}19, RBPGJ^{+}20, RHB23, SSS^{+}11, SF02, SMR^{+}20, STC10, SWP^{+}13a, TKC^{+}22, WC15, XYL^{+}22, Yas07a]. oceanographical [Coo69, VKGP^{+}13]. oceanographically [DIQJ21]. Oceanography [Ano94c, BDB^{+}22, BLAM00, Fei04, HHW22, Kru19, RLSF07, RG03a, SDS22a, SAA^{+}15, SHC^{+}07, VHK04, BC91, BTK^{+}99, BMGN15, BD85, Dav99, FF83, FJhT^{+}14, FÁFL06, FL06, Gou85, Jac10, LFA^{+}06, OEL^{+}14, Pai20, PIS13, PFHM10, RBD^{+}07, RSW^{+}23, SGPdM18, SFK^{+}99, SENS13, SMB88, TDH^{+}95, TAO05, Tol85a, Tol85b, WB03, Wüs64, YSD15, Hof81, Ang88, WR03]. Oceanologica [Ang86]. oceanological [AVG^{+}19]. oceans [BPGD^{+}14, DP18, FC05, KF11, KKNT23, LRAE23, LCR^{+}93, McK15, OP18, SvN04, SBM91, VBC^{+}20, CRF^{+}10, GPA^{+}11, HGT16, ICB^{+}19, KSV08, KMWF11, MMF^{+}07, Ang79a]. OCLE [dlHRA^{+}18]. October [Ano21t, Ano22-34, Ano23t, CBB^{+}22b, Ano98e, Ano00c, Ano00h, Ano02c, Ano03b, Ano04e, Ano05b, Ano06b, Ano07l, Ano08q, Ano11g, Ano11j, Ano12l, Ano18l, Ano20t, CBB^{+}19, SLBVRR^{+}22]. octocoral [SW21]. octopus [LAGM^{+}23, OÁSG^{+}16]. ODAS [CPB^{+}15]. Oden [NBR^{+}08]. Off [ÁSDB^{+}01, ÁSFP^{+}03, AMEV07, AAM^{+}14, AMG^{+}16, Ang79b, AYH^{+}23, ACN01, ASB^{+}08, ACHSH08, BW08, BBLD^{+}11, BSMC15, BGM^{+}01, BIST01, BFH01, BGM^{+}10, BBSN04, BSC^{+}07, BAOC^{+}09, BM07, BM86, BASS^{+}20, CPG08, CMF^{+}09, CCW^{+}02, CMS^{+}13, CGG08, CCMS08, CPNL07, CFG07, CF12, CW02, Cra09, DM13, DWS^{+}24, DSC^{+}21, DB02, Dur09, ES07, EHG^{+}07, EM12, EHFD12, FZ88, FC07, FP03, FB01, FGL^{+}23, GMBU12, GEO09, GGQ07, GMAB07, GDI^{+}09, GRB^{+}08, HHB^{+}01, HPS^{+}01, HFW^{+}98, HEF^{+}12, Hob10, HFO90, aHFS92, HWF^{+}21, HSF02, HWS^{+}07, IVT^{+}12, JE92, JW01a, JIT^{+}01, JJS03, JGO^{+}98, KKB00, KP03, Ken88, KC15, Kos02, KPSB22, LC12, LOG^{+}09, LAA12, LQU07, Lie88, LZG20, LPF^{+}18, LS13, LABD^{+}24, Man69, MERB12, MLD^{+}03, MMMWZ23, MB01, MM80, Mid69, Mit83, MMF^{+}12, MDC^{+}07, MGH^{+}07]. off [MTH^{+}10, MHCR^{+}12, MA12, MSL^{+}07, MDL^{+}12, MCGS^{+}16, NEI^{+}22, NIF^{+}15, OLH^{+}18, ORW^{+}01, OÁSG^{+}16, PMM^{+}23, PCSMC12, PELAA18, Pea02, PAM^{+}88, PK02, PKF02, PD15, RBNJ^{+}12, RBPGJ^{+}20, RWOA01, RB20, RZW^{+}23, RASVB^{+}22, SLM^{+}16, SSQ19, SLG^{+}12, SDGVE17, SS69, SYB^{+}15, SAY^{+}16, SHF01, SBD^{+}07, SSL07, SMPC^{+}12, Tom81a, VNMS91, VSGD21, Ven12, WOW^{+}14, WCC^{+}20, WLM07, YPGE^{+}10, ZCH^{+}17]. Off-shelf [ÁSDB^{+}01]. off-shore [JIT^{+}01]. offs [KSY^{+}19]. offset [DWFP^{+}19]. Offshore [FRK^{+}09, SFMA20, ZL24, BDG^{+}17, CTF07, CTA16, CZL^{+}24, CSG^{+}15, FvBA^{+}17, GDN^{+}18, Hau84, HW02, KSD84, KLC^{+}15, LYZ16, MKSW^{+}15, PMMN^{+}22, SOWS17, SDK84, Sim84, SKHD84, WH89]. offshore-estuarine [PMMN^{+}22]. offspring [DWFP^{+}19]. OI [ZZM^{+}24]. oil [BCD^{+}20]. Okhotsk [KKS^{+}19, AVG^{+}19, FJ19, FMWW14, HKN^{+}14, INT14, Iwa23, KTN14, KON14, KCBS20, MMN^{+}24, NTU^{+}14, NO14, NNM^{+}21, NMY^{+}14, NNO^{+}14, ONR^{+}14, Rog00, RCSA01, SM21, SBH^{+}14, SMN^{+}14, SOO^{+}14, SIS^{+}14, SPB19, TII^{+}14, UNN^{+}14, WNNI21, YYT^{+}14, YNTS22, YAI^{+}14]. Okinawa [DYL^{+}15, ZLC^{+}15]. old [CBB^{+}22c]. oligotrophic [DDP^{+}00, DLC^{+}08, DTW^{+}00, EGPM^{+}15, JHM^{+}22, KGB^{+}23, KAAK^{+}16, LRAE23, LL21, LEDR^{+}22, PTI00, QPR03, TWAL^{+}11, WYT00, WQ08, YFY^{+}22, ZLR^{+}07]. oligotrophy [VDS^{+}18]. olivacea [CdTH^{+}16]. olive [CdTH^{+}16]. Oman [HFW^{+}98]. OMEX [vWMH98]. Ommastrephes [IIS^{+}17]. Ommastrephid [RASVB^{+}22]. omorii [SJJ^{+}03]. OMP [PPVG12]. OMZs [PRP09]. onboard [CTP^{+}18]. Oncaea [KSKN21]. Once [Mun97]. Oncorhynchus [AHC^{+}13, KBF^{+}08, SKSK06, WBF^{+}21, YWUK15]. One [MSFZ19, DSAB20, PMG15, RCD^{+}94, RZW^{+}23, YFY05]. one-dimensional [YFY05]. Oneirophanta [RMB^{+}01, WDK^{+}01]. Ongoing [CHG^{+}18, IG19, OV24]. only [SHF01]. onset [KHBA^{+}24, VEM^{+}21]. onshore [YYC^{+}18]. Ontogenetic [CSMGS19, RK20, SKH^{+}23]. ontogeny [IS19]. opal [RGC^{+}01]. Open [PBH^{+}10, BB24, BPGD^{+}14, CMPNC^{+}22, CSC^{+}12, FBR^{+}13, FBD18, Fro93, KKS^{+}18, KFC^{+}13, KRL^{+}22, LSM^{+}22, LFCSV^{+}13, MPCNC^{+}19, MCKS17, PPdM^{+}12, PAVB^{+}21, RFKC16, SGL^{+}17, SK91, SPB93, TCDPP^{+}22, Whi94, ZGB^{+}20, dlHRA^{+}18]. open-ocean [RFKC16, Whi94]. open-sea [PPdM^{+}12]. open-slope [PPdM^{+}12]. operated [CAO^{+}20]. Operational [NHS^{+}14, SGLF^{+}13]. opheliid [VCSG^{+}01]. opilio [MGKW19]. opportunities [FDH20, HSH^{+}19, MLHM09]. opportunity [SPH^{+}15b]. oppositely [LBH^{+}21]. Optical [BGR^{+}15, BDB^{+}04, BTNK13, BMB^{+}16, CDB^{+}22, CTMV^{+}14, FRV^{+}19, GNH19, HHMB^{+}09, JZZY24, KTH^{+}21, LCB18, LGL^{+}18, LC10, MVC^{+}11, MCT03, PTP^{+}22, SHP^{+}23, SRG^{+}19, TII^{+}14, WMB^{+}18, XC14]. optics [Jer65]. optima [TOKLC08]. Optimal [SF15, SCY^{+}23, MPSS91]. optimisation [TIOM16]. Optimization [MV10, HLS^{+}14a, RG94]. Optimizing [SPC^{+}23]. optimum [ÁBMÁS14, ÁBMÁS15]. options [HSC09]. orbit [HYM^{+}24]. order [CES^{+}19]. Oregon [BASS^{+}20, CW02, FP03, HSF02, HWS^{+}07, KP03, Kos02, Pea02, PK02, PKF02, PD15, SHF01]. Organic [CMPNC^{+}22, GBM^{+}01, HW02, LB14, RSD^{+}90, TPM^{+}00, VBA^{+}18, AYK^{+}05, ASC92, BB24, BHHR15, BFJ18, BS02, BMG^{+}21b, CDS90, CKP^{+}20, CGM^{+}02, CGC^{+}20, CJ92, DDE^{+}95, DDP^{+}00, DMD^{+}00, DDD^{+}00, DBW^{+}22, DOS^{+}18, DHDM22, DVB^{+}18, EvdZSH02, FPD^{+}01, FZY^{+}23, FUOG^{+}16, FK86, GMAMB04, GvOS^{+}08, GLAHH^{+}22, GF19, HM98, HOY^{+}21a, HSLG11, HOY^{+}21b, HPZC21, HMKF08, JZZY24, JW01a, JMG^{+}13, KNSN^{+}09, KSR^{+}01, Kit03, KGdS^{+}08, LSV14, LSB^{+}17, LSD^{+}18, LSD^{+}15, LHEB98, LvIKB07, LFBP^{+}13, MPCNC^{+}19, MBCB88, MMKS^{+}21, MDC^{+}07, NRA^{+}21, OB98, Par63, PGT^{+}13, PPSVC^{+}13, PRL^{+}18, RFSCF19, RGE22, SVHM^{+}13, SFMA20, SHd13, SRF^{+}19, SMN^{+}14, SIS^{+}14, SBC^{+}24, TPPG10, TZP^{+}00, TDK^{+}16, VFCC^{+}22, VK92, WMC^{+}89, WRS^{+}92, XYGJ23, YTB^{+}21, ZMW^{+}23, ZKK^{+}16, ZZWL06, ZHBW01, dFHR^{+}24]. organic-walled [ZHBW01]. organism [VFS^{+}15]. organisms [BCLD^{+}17, Bri79, KFC^{+}13, PNF^{+}21]. organization [JFG^{+}90]. organizing [RRS03]. organochlorine [SGL^{+}13]. orientale [Ber65b, Bou65]. orientalis [BMC^{+}10, FFT^{+}18, KKKY10, KTIT22]. orientated [STM10]. oriented [ŠPM^{+}22, VMN08]. Origin [BvdLA^{+}11, PRL^{+}18, AS96, AGD^{+}18, BPNB90, IOGS13, LCJ^{+}07, LPBM17, MM01, NBHM01, PL18, SA97, VKT15, Wil87]. original [MJD^{+}21, SDS22a]. originating [NTU^{+}14]. Origination [BLC23, LGH^{+}21]. Origins [SG91, BhTW10, NMK^{+}03]. Orkney [TSFA22]. orthogonal [TCL20]. Oscillation [BGMP03, FJH10, GPC^{+}03, HHSR07, SMG02, Wu13, LYM12, RLX^{+}24]. oscillations [BhTW10, CDDF11, CMF^{+}09, Hen73, IAN13, KWI20, Lie88, ŠVL^{+}15]. OSCOM [DDJ^{+}21]. OSMOSE [MSB^{+}23]. OSSE [HKL^{+}15, KAH^{+}16]. ostracod [ABSDC07, BHK^{+}19]. Ostracoda [BHK^{+}19]. ostracods [Ang79b, Ang84, YTL^{+}19]. Osumi [Ike88]. other [DDE^{+}95, EBvdL^{+}09, FMC^{+}15, HHY03, HM15, HVEF09, RZW^{+}23, Sma10a, Sma10b, WXH07]. otolith [EiIT^{+}22]. our [SRM^{+}10]. out-flows [Mil14]. outbreak [BBB^{+}14]. Outer [FCN^{+}19]. Outflow [BF12, APN^{+}15, GPE^{+}17, JX18, MHH^{+}15, Mil09, Sek88, TVD^{+}99]. Outflows [PB94]. outlining [IGG^{+}19]. output [RKM^{+}07]. ovata [RK20]. over-generality [AF10]. Overarching [Was15]. overbanking [CFM^{+}18]. overcapacity [FBM^{+}08]. overfishing [BWMGCB08, BMB06]. Overflow [LBF^{+}22, Men21, ZLZ^{+}17, HØ07, JJR^{+}08]. Overlap [LMP22, GRB^{+}08, NHE^{+}13]. overlying [LPA92]. Overrunning [KVLA06]. Overturning [MSJ^{+}15, APHGC^{+}22, GWB14, HGPFN^{+}14, HGT16, HGTP^{+}19, KJH^{+}22, MR06, MFS^{+}16a, MFS^{+}16b, MLS^{+}15, Ric08, SFMT12, SFMT14, SDS02, SDS22a, SSM^{+}18, Tal08]. Overview [CTL^{+}04, ZSW^{+}22, ARH^{+}00, BIL03, BR01, BRH^{+}05, KNS^{+}03, KY14, MP13, Mau10, MAB^{+}11c, SPF^{+}23, WMB^{+}18]. Overwintering [GHF^{+}21, DOS^{+}18, DBM17, HPNDC15, SRK15, SK17]. OVIDE [MLS^{+}15]. oxic [GGQ07]. oxidation [EvdZSH02, STW^{+}15]. oxide [dlPHF^{+}15, dlPPÁB24]. oxides [FC07]. oxoacids [SCCJ^{+}18]. oxycline [WGG^{+}08]. Oxygen [Ant09, KSV08, KSPK99, PRP09, RCÁS^{+}15, SE16, SGR^{+}22, VSGC21, VJJ^{+}22, AYH^{+}23, BKD^{+}20, CLL^{+}18, DGVGR24, EiIT^{+}22, GEO09, GLV12, HFW^{+}98, LGR^{+}02, LQU07, MYH^{+}22, McG64, MSB^{+}23, NAH^{+}21, PAVB^{+}21, RS10, SGF^{+}19, SM05, SKP99, SK21, WAH^{+}20, WFR07, WGG^{+}08, XYGJ23]. oxygen-depleted [DGVGR24]. Oxygenation [GEP^{+}08, FC07]. oxymeter [Føy65]. Oyashio [INO^{+}24, iIYO^{+}10, KTW^{+}22, NNM^{+}21, OTNI20, ST03, TKSI08, ZZM^{+}24].

P [Ang80, Hof81, CGB07, CR20, DVB^{+}18, Fre07, GGQ07, GLS08, JJA^{+}08, MG02, PHK^{+}17, PV07]. P. [vPRT90]. P06 [KMWF11]. Pa [DTKvH15]. Pachygrapsus [PPPdS20]. Pacific [AYK^{+}05, AFH^{+}11, ARG11, BJMP20, CLSD18, CSR90, CF07, GLV12, Hau18, HHW22, HHP10, IMM^{+}22, iIRM^{+}15, JS21, Kam19, KSV08, KTW^{+}22, KMF^{+}20a, KMF^{+}20b, LLL^{+}24, LM14, LLX^{+}21, MS00, MTK^{+}22, MW96, NXT^{+}17, OMK^{+}22, PYKF15, Rei86, RM89, SJ02c, SJ02b, SKH00, SW21, TWBC^{+}13, WF07, WH20, WSC^{+}21, WWN^{+}99, AS20, AALM06, ALM^{+}23, APHGC^{+}22, AT07, BPF06, BE99, BLR^{+}23a, BP02, BGM^{+}10, BLI^{+}99, BHB^{+}19, BHH^{+}16, BJMP19, BLP93, BMC^{+}10, BAB^{+}19, BBRM20, Bri79, BMN^{+}99, BPM^{+}14, BH85, BPSN^{+}21, CARBML^{+}22, CBB^{+}22a, CES^{+}19, CSLJ03, CBOP15, CAT^{+}08, CSK^{+}12, CMF15, CHS^{+}24, CRPS^{+}15, CR20, Dag93, DYO^{+}10, DN07, DWH^{+}14, DCKB13, DOS^{+}18, DRE^{+}08, DGVGR24, Don65, Don87, Don94, DLC^{+}08, DMBB02, DBRK17, DLM91, ED82, Emi65, FELMGM^{+}22, FJhT^{+}14, FMM^{+}20, FÁFL06, FT06, FL06, FGGDF^{+}04]. Pacific [FC05, FMCG15, FLUC08, Fro93, FK99, FMT15, FFT^{+}18, FB05, GLY23, GR17, GCC^{+}24, GKR20, GDM^{+}20, Gif93, GBT^{+}19, GW91, GdRGL^{+}01, GM19, GCFS06, GBC^{+}15, GC09, Gri22, GRdSS^{+}22, HSS^{+}12, HKK12, HPC^{+}20, HM00a, HMWM00, HBV^{+}99, HSK^{+}19, HFNG00, HRSM08, HZCZ16, HGT16, HGBG20, HKY^{+}11, His22, HM08, HHW01, HFO90, HLTB^{+}17, HBH^{+}17, HWS^{+}07, INO^{+}24, IIS^{+}17, IHT^{+}21, IHY^{+}01, IAFD02, iIRM^{+}15, IU14, JM19, JLRB20, JSKM02, JSLA^{+}21, KST^{+}10, Kat18, KYT^{+}16, Kes06, KHD22, KKNT23, KiL14, KKKY10, KIS^{+}05, KAK^{+}22a, KST03, KTIT22, KKS^{+}19, KNI^{+}05, KRHS14, KYS^{+}17, KRL^{+}22, LMPB^{+}16, LJ65, LGK^{+}93, LMS93, LFA^{+}06, LYM12, LMW^{+}12, LW85, LBH^{+}87, Leh01, LSS^{+}10, LBSP01, LMH^{+}13, lLdZQ^{+}22, LG23, LSMG01, LB20, LCGH07, LLS01, LGZL22, Luk86, MMR^{+}09, MSMR93, MT99, MBT07, MRRC73, MCD^{+}14]. Pacific [Man04, MTC12, Mar20, MRAP22, MMF^{+}07, McI10, MS17, MSI17, MNM06, MPSS91, MZZ^{+}23, MFB^{+}84, MN88, Mil88, MC88, Mil93b, MSA^{+}22, MM99, Min00, Min02, MS15, MNFY21, MCGS^{+}16, NEI^{+}22, Nag01, NHH^{+}23, NMLBCM^{+}01, NMO^{+}21, NMY^{+}14, NYH^{+}22, Nof96, Nof00, NKK^{+}05, NRA^{+}21, ORPRGIS22, OTNI20, OPG^{+}10, OOTA15, OT19, OAM00, ORMB08, PGY^{+}22, Peñ03a, PV07, PB07, PMK^{+}06, PLEF^{+}23, PS08, PGS^{+}22, Phl65, PLB^{+}23, Pie01, PVA24, PLK14, Qiu15, RTF^{+}05, Rei97, RMG90, RvBD^{+}22, RDC^{+}21, RHBS13, Rog00, RGM01, RPSC22, SSB19, SSB20a, Sai65, SSH^{+}05, SCLG^{+}11, SF02, SDGVE17, STF^{+}13, SMdG02, SMG02, SBH^{+}14, Sek99, SCC14, SBFP21, STS^{+}12, SO91, SIS^{+}14, SPS^{+}99, SPB93, SJ02a, SASH08, SKT01, SHS^{+}05, TJ73, TFY02, TKSI08, TOKLC08, TT05, Tan99, TMÁGC^{+}21, TSJ^{+}12, TLF^{+}89]. Pacific [TSNO05, TKK^{+}05, Tur65, UKM^{+}14, UBB^{+}23, VCM04, VMH^{+}21, VMN08, WFH^{+}22, WF06, WO85, WNNI21, WCS^{+}23, WLM^{+}13, WSG^{+}93, Whe93, Whi94, WCN^{+}05, WFBN^{+}13, WFR07, Wil65, WLL06, WHK23, WXH07, WBD^{+}15, Woo18, WJPHB15, WZ04, Wu13, XYL^{+}22, XC14, XNT^{+}17, YYT^{+}14, YMA^{+}17, YTB^{+}21, YAK^{+}08, YBPS08, YYK^{+}12, YFY05, ZSI^{+}05, vHVAT22, HHW22, Wu13]. Pacific-Asian [FJhT^{+}14]. Pacific-influenced [GCFS06]. pacifica [DSBP15, FMC^{+}20, HL05, OTNI20, RB20, SBFP21]. pacificus [JC04, KSK^{+}15, LPHL^{+}05b, PHLL05, PLHLF05]. pack [SLGI^{+}21]. paedomorphosis [GM19]. PAEs [PAG^{+}18]. Page [Ano17i, Ano17j, Ano17k, Ano17e, Ano17f, Ano17g, Ano17h]. Page/Cover [Ano17i, Ano17j, Ano17k]. Pages [Ano63c, Ano64d, Ano65h, Ano65i, Ano69c, Ano73d, Ano76b, Ano76c, Ano77b, Ano77a, Ano78, Ano79b, Ano79c, Ano79d, Ano80c, Ano80f, Ano80d, Ano80e, Ano81b, Ano81e, Ano81c, Ano81d, Ano82c, Ano82f, Ano82d, Ano82e, Ano83b, Ano83c, Ano83d, Ano83e, Ano84b, Ano84c, Ano84d, Ano85f, Ano85g, Ano85j, Ano85h, Ano85i, Ano86j, Ano86e, Ano86i, Ano86f, Ano86h, Ano86d, Ano86g, Ano87e, Ano87f, Ano87h, Ano87g, Ano88d, Ano88i, Ano88f, Ano88g, Ano88c, Ano88e, Ano88h, Ano89d, Ano89f, Ano89h, Ano89j, Ano89e, Ano89k, Ano89g, Ano89i, Ano90c, Ano91d, Ano91g, Ano91j, Ano91l, Ano91f, Ano91i, Ano91e, Ano91h, Ano91k, Ano92i, Ano92d, Ano92f, Ano92g, Ano92h, Ano92e, Ano93c, Ano93d, Ano93e, Ano93f, Ano94j, Ano94e, Ano94i, Ano94f, Ano94d, Ano94h, Ano94g, Ano95d]. Pages [Ano95j, Ano95f, Ano95h, Ano95c, Ano95i, Ano95e, Ano95g, Ano96c, Ano96e, Ano96f, Ano96d, Ano96i, Ano96g, Ano96h, Ano97c, Ano97f, Ano97d, Ano97e, Ano98c, Ano98b, Ano98a, Ano98d, Ano98e, Ano99c, Ano99e, Ano99b, Ano99d, Ano00f, Ano00e, Ano00g, Ano00d, Ano00a, Ano00b, Ano00c, Ano00h, Ano01b, Ano01c, Ano02a, Ano02b, Ano02c, Ano03f, Ano03e, Ano03c, Ano03a, Ano03d, Ano03b, Ano04b, Ano04f, Ano04c, Ano04d, Ano04g, Ano04e, Ano05a, Ano05d, Ano05c, Ano05b, Ano06c, Ano06a, Ano06d, Ano06b, Ano07n, Ano07q, Ano07k, Ano07m, Ano07o, Ano07r, Ano07l, Ano07p, Ano08p, Ano08r, Ano08s, Ano08o, Ano08n, Ano08u, Ano08w, Ano08q, Ano08t, Ano08v, Ano09m, Ano09i, Ano09j, Ano09k, Ano09l, Ano10j, Ano10k, Ano10l, Ano10m, Ano11i, Ano11h, Ano11g]. Pages [Ano11j, Ano12m, Ano12n, Ano12p, Ano12o, Ano12l, Ano12k, Ano13n, Ano13l, Ano13p, Ano13m, Ano13o, Ano14m, Ano14p, Ano14q, Ano14n, Ano14r, Ano14o, Ano14l, Ano15n, Ano15o, Ano15m, Ano15r, Ano15p, Ano15q, Ano16t, Ano16o, Ano16n, Ano16p, Ano16m, Ano16q, Ano16k, Ano16l, Ano16s, Ano16r, Ano17r, Ano17n, Ano17o, Ano17m, Ano17q, Ano17s, Ano17l, Ano17p, Ano18j, Ano18k, Ano18m, Ano18n, Ano18l, Ano19n, Ano19m, Ano19k, Ano19l, Ano22-29, Ano22-30, Ano22-27, Ano22-28, Ano22-31, Ano22-32, Ano22-33, SW81]. Pagophilus [FJH10]. Palaeo [DVL^{+}99]. Palaeo-environmental [DVL^{+}99]. palaeoceanography [Gal17]. palaeoclimate [Gal17]. Palamós [CHSB^{+}21, PGLG^{+}05]. palatability [PMS^{+}15]. Paleoceanographic [LXC^{+}22]. paleoceanography [SBH^{+}14]. paleoclimate [SOH21]. paleoenvironment [YSY^{+}19]. paleogeography [Sai65]. paleooceanographical [Ola65a]. paleoproductivity [MDL^{+}12]. paleotemperature [Ola65b]. paleothermometry [SBH^{+}14]. paleoxygenation [MDL^{+}12]. Palk [JJA^{+}13]. pallasi [GBT^{+}19, Nag01, STF^{+}13]. Palmas [Ano09h]. PAMS [FJhT^{+}14, LMW^{+}12]. pan [CW06, PFW15, WCB^{+}20b, CSK^{+}12, HRSM08, MVN^{+}15, VMB^{+}22b]. pan-Arctic [CW06, PFW15, WCB^{+}20b, MVN^{+}15, VMB^{+}22b]. Pan-North [CSK^{+}12]. Pan-Pacific [HRSM08]. Panama [BSW86]. panarctic [CBC^{+}06, CWW15, HKGH^{+}06]. Panulirus [SS69, WOW^{+}14]. PAP [HBD^{+}21, BEI^{+}20, BKD^{+}20, KVNT20]. PAP-SO [HBD^{+}21, BEI^{+}20]. Papa [AEPW93, Fre07]. papers [BPW10]. PAR-light [JTD^{+}14]. Paradigm [Fro05]. paradox [Fro05, LIH^{+}12]. paralarvae [OÁSG^{+}16, RAG^{+}19, RASVB^{+}22]. Parallel [SPMVP05]. parallels [GA01]. paralytic [CAH^{+}22]. Paramesochridae [PMG15]. Parameter [HLS^{+}14a, SSL08, SO91, Tom81b]. parameterisation [BMM97, SRM^{+}10]. parameterisations [GLS08]. parameterization [BBS21, BBS23, HLFL23, SCLS10]. parameterizations [ZCD08]. parameterizing [Den03]. Parameters [Bri79, CÁM06, KPSA17, SSVP00, SW22, SCHD23, SMPC^{+}12, WXH07, XDG^{+}23]. Paramunnidae [GM19]. Parapenaeus [PPdS21]. parasitism [LZF^{+}24]. Paratethyan [Gal17]. parcels [RLC85]. Park [VDGGD^{+}22]. part [AVG^{+}19, Igu04, IMW^{+}14, INT14, KFKO03, KNS^{+}03, KSD84, Sim84, BBS21, BBS23, GAM98a, GAM98b, Hau84, MFS^{+}07, MSF^{+}07, McD81a, McD81b, MR03, MAB^{+}11c, MAB^{+}11a, MAB^{+}11b, RANS65, SFMT12, SFMT14, SDK84, SJ02c, SJ02b, TTB^{+}08a, TTB^{+}08b, VOJD02a, VOJD02b]. Partial [LGZ^{+}20, BF01, LM00, YCP^{+}12]. Particle [DL17, KKS^{+}18, NMK^{+}03, NKK03, RCSVGP^{+}16, TCDPP^{+}22, ZFSV^{+}09, BS02, DTKvH15, FLdST98, GIPG17, HPC^{+}20, JvdLL^{+}15, KNSN^{+}09, KHC^{+}99, LFCSV^{+}13, MBdM^{+}18, MMG^{+}11, NFMCS^{+}22, NIF^{+}15, ORMR^{+}19, SG91, SMM^{+}90, SRG^{+}19, TFY02, VK90, VSC01, ZDM^{+}20]. particles [AH15, BPNB90, BB24, Bla63, BGR^{+}15, BT07, FRK^{+}09, FAAV^{+}15, FK86, FDE^{+}22, HM15, KSR^{+}01, LTJ^{+}15, MPM^{+}17, MLD^{+}03, MLL^{+}15, NMK^{+}03, Pas22, PPCWJ18, RTF^{+}05, SPB^{+}02, SCT^{+}00, TSBS18, VDP^{+}01, WGZZ19]. particular [KDB95, TR99]. Particulate [BSW86, HMKF08, LPA92, AYK^{+}05, CGM^{+}02, CGC^{+}20, CLG^{+}00, CS89, CE84, DDDT99, DDP^{+}00, DOS^{+}18, DCS^{+}22, FUOG^{+}16, FRV^{+}19, HSLG11, HOY^{+}21b, HPZC21, IU14, KLB^{+}21, KBC^{+}22, KSP^{+}23, LSB^{+}17, LSD^{+}18, LSD^{+}15, LFBP^{+}13, MFDH22, PL87, PLP99, PBP^{+}99, SMN^{+}14, SK18, TvW98, TZP^{+}00, TDK^{+}16, TRMV15, USH15a, Whe93, WRS^{+}92, ZZWL06]. particulates [VKJ^{+}23]. partition [BEH19, FP15]. Partitioning [FDM^{+}13, BGA^{+}21, EBM^{+}20, ERBV21, GVBV^{+}21, IS19, KPSB17, MSMR93, NHE^{+}13, PPKR14, SPB^{+}12, SM16]. Partnering [JRR^{+}24]. partnership [KY15]. Pass [STGR^{+}23]. Passage [ARELAK24, CLV^{+}19, CP07, GWGR^{+}19, Spr08, VKDS^{+}18]. Past [BCD^{+}20, DPB06, WH89, BDLW14, Den87, HDZY15, HPHW21, KKK04a, LWY07, LDMH09, NW87, PM13, SLG^{+}12, SBB^{+}14, SKRM^{+}95, XNT^{+}17]. Patagonia [SPF^{+}23, BPSN^{+}21, CSMGS19, CCM^{+}14, CTI^{+}19, FCN^{+}19, IPD14, JTD^{+}14, LSV14, MVBC^{+}21, MSV^{+}14, SSQ19, SPSR^{+}14, TRP^{+}23]. Patagonian [CMC^{+}16, CM18a, CCM^{+}14, CBPS^{+}22, GHVG19, GF97, GCD^{+}13, IPF23, MPMFL^{+}23, MTC14, PMA^{+}14, PSGVS^{+}14, PWZ^{+}16, QOS^{+}22, RPSVLS14, SV14, TSRF14, VMV^{+}23]. patagonicus [STC10, STEB16]. patch [PSL87, TNS^{+}05]. patchiness [Mar03]. path [Cia14, ORCH^{+}19, TMN^{+}12, TSS^{+}12]. paths [IMW^{+}14]. pathway [BLC23, NBHM01, QLY^{+}22]. Pathways [YKWF21, ZSI^{+}05, BNCC15, BGV^{+}23, CRT^{+}22, DTG^{+}24, KSS^{+}23, LGH^{+}21, LBF^{+}22, MRMD^{+}97, Men21, MJD^{+}21, NMO^{+}21, PGG^{+}22, RBL90, RBE^{+}12, SPG^{+}06, STM10, THM^{+}06, WHS17, ZLZ^{+}17]. pathways-orientated [STM10]. pattern [FPIJ85, GKS^{+}13, HMRB^{+}03, SMFM^{+}21, SIR^{+}07, UAM05, XLX^{+}20, XLL^{+}20, ZGB^{+}20]. Patterns [BDL08, BL02, CMM^{+}04, CLD22, CB09, CRF^{+}10, KPSB22, QPR03, RKCH15, SM05, SBB^{+}22, Ant09, ABP^{+}23, BGMP03, BMO12, BWB^{+}09, BHC^{+}18, BRG^{+}23, BD18, BAP^{+}22, BFV^{+}17, CGG08, CRHM12, CCW^{+}18, DGGdR02, FWO15, FWH^{+}17, FPY^{+}16, FMCG15, GSV^{+}01, GHF^{+}21, GL06, GCC^{+}24, GNH19, GBC^{+}16, HLR17, HDZY15, HEF^{+}12, HLP^{+}16, HHW01, HHW22, HVEF09, IGG^{+}19, JYK^{+}14, KKKS14, KFH^{+}15, LJM^{+}16, LDAM^{+}07, LRAE23, LW85, LFC^{+}15, LC16, LW13, MHGGS19, MB20, MDB^{+}20, MVN^{+}15, MEMP15, MMF^{+}17, MHS^{+}09, MCMT^{+}17, NDEG22, OMK^{+}22, OHC^{+}17, OPG^{+}10, OACB^{+}15, PGLG^{+}05, PCM11, PK02, PSA^{+}19, Phl65, PG10, QOS^{+}22, Rei86, Rei89, Rei94, Rei97, Rei03, RRS03, SSB20a, SLM^{+}16, SGA^{+}19, SPG^{+}06, ŠVL^{+}15, SEG22a, SK17, SPH^{+}15b, SWH^{+}24, SI97, SBS90, TSC03, WTT14, WPH^{+}10, YHRT22, YKNO23, ZLX^{+}20]. Paul [Bak83]. Paz [LM10]. Pb [KKS^{+}03, MSL^{+}07, SCMAR^{+}99, ZNR^{+}24]. PBDEs [FDM^{+}13, RNL^{+}13]. PCBs [FDM^{+}13, RNL^{+}13]. PCDEs [RNL^{+}13]. PCNs [RNL^{+}13]. pCO [AEPW93, CVBG21, GCB^{+}22, HSS^{+}12, VSA^{+}21, ZKK^{+}16]. PCR [AIA^{+}15]. PDO [Wu13]. peaks [CC23]. Pearl [LYZ16, LZG20, WL16]. Pedro [Gor92, Hic92, WJE^{+}92]. Pelagic [BPNB90, FvBA^{+}17, Mil93b, Sou94a, WWW^{+}23, ARD^{+}03, AAMB^{+}24, Ang89, AE09, BDP^{+}06, BN03, BAOC^{+}07, Bol94, BCL^{+}09, BAP^{+}22, BFV^{+}17, CKB^{+}17, CSMGS19, CBOP15, CRF^{+}10, CAB^{+}99, DYO^{+}10, DPH^{+}18, DPR^{+}18, DHHP18, DTW^{+}00, EKB06, ESGP17, FFS^{+}20, GRD^{+}23, HKGH^{+}06, Hob10, HFPS^{+}06, IMM^{+}22, JYK^{+}14, JLP^{+}20a, JLP^{+}20b, KHS^{+}14, KLC^{+}15, KFH^{+}15, KAAK^{+}16, LR07, LSY^{+}14, Law04, Leh01, LLAPG^{+}22, Lon95, LSIB23, LSH^{+}22, MRA^{+}19, MHR^{+}10, MMN12, OMR^{+}22, OMK^{+}22, OH94, OPG^{+}10, OvdSN94, POS^{+}07, PRTC13, PAB^{+}21, PLEF^{+}23, PSA^{+}19, PBBH^{+}22, PNF^{+}21, PGG^{+}22, RSW^{+}23, RLGC10, RFC^{+}15, RG94, SGF^{+}19, SCD^{+}07, SSL08, SHC^{+}06, SHC^{+}07, SPK^{+}22, Tan99, TCF^{+}18, THM^{+}14, VFS^{+}15, WPH^{+}10, YBS^{+}01, vdS94a, vdS94b]. pelagics [RFS10]. pelagicus [SWP^{+}13a]. PELAGOS [BC99]. pelamis [KAK^{+}22a, PGS^{+}22]. PELGAS [DPR^{+}18]. pellets [RWOA01, Tur15, WYT00]. pen [PRA^{+}18]. penguin [STEB16]. Penguins [STC10]. Peninsula [FDH20, SVL^{+}23, Peñ24, RK03a, BHA^{+}14, BAOC^{+}07, DCD^{+}23, DSR21, DHDM22, HWPLvW20, HSH^{+}19, HGH^{+}19, IIM^{+}23, KKB00, LS12, LS13, LWT^{+}20, MWS^{+}10, RZW^{+}23, TNGP22, WCC^{+}20, YN20, ZHD^{+}20]. peninsular [MK12]. pentadal [Lev88]. pentadecadal [Min00]. perception [JSdSS^{+}21]. Perciformes [SM21]. Performance [BTNK13, MAB^{+}11a, HCGK11, RSG06, SHT^{+}01]. Pergamon [Ang79a, Ang80, Ang88, Bak83, Hof81, SW81]. period [BM01, FGGDF^{+}04, GFGGD^{+}23, GIHJ23, GCD97, Hen73, HGBG20, INO^{+}24, KGB^{+}23, LAA12, NF06, OEL^{+}14, PPdM^{+}12, Rud15, TFM03, WHIH97, YYhT^{+}17]. periods [CQC15, EMU^{+}23, FMM^{+}20]. periphyton [VPM^{+}19]. permanent [MTL05]. permanently [DGVGR24]. permitting [KDL^{+}01, WBB^{+}01]. Persian [BD85]. Persistent [LNB13, PGY^{+}22, FELJ16, GBC^{+}15, RNL^{+}13]. Persistently [WFR07]. perspective [BCB^{+}05, BKC15, BBFS19, CEF^{+}13, DLC^{+}08, KPSB22, PL18, PL01, Pow06, WCB^{+}20b, WPW^{+}14, Zav99]. Perspectives [SM16, CW06, JVJ^{+}17, MHH^{+}15, PHCA17, Was15]. Perth [EVM^{+}15]. pertinent [BGR^{+}15]. perturbations [WWZ19]. Peru [TWMY08, ASB^{+}08, ACHSH08, BW08, BBLD^{+}11, BDT^{+}08, CGG08, CCMS08, CRHM12, FBM^{+}08, GGPG^{+}19, GRB^{+}08, GEP^{+}08, LGR^{+}02, LPF^{+}18, MS02]. peruanus [BWMGCB08]. Peruvian [ATT^{+}08, BWMGCB08, BDL08, BHHS83, EALF08, EB08, FFA09, FVLC^{+}23, GHG^{+}24, GCP08, HLS^{+}14a, JBB^{+}14, SGO^{+}08, SBG^{+}08, TMÁGC^{+}21, XRC^{+}15]. Peterson [BDB^{+}22, SSB^{+}20b]. Petrace [GCLD19]. petrel [FDB^{+}21, SWP^{+}13a]. Petrology [Nay65]. Pettersson [Ano65e, Ano65f]. PFT [TSL10]. pH [CVBG21, Rot65, SSTL16]. Phaeocystis [SDL^{+}19]. phase [GRS08, HHSR07, IIS^{+}17, MP04, WP91, MRMD^{+}97, NF87]. phaseoliformis [AB90]. phases [Kil76]. phasic [RNBP^{+}19]. Phelliactis [vPRT90]. Phenological [WBF^{+}21]. Phenology [MI21, MCKS17, AHW^{+}15, CAT^{+}08, DAIS10, KPSB22, LBC^{+}23, LMA^{+}15, MGE^{+}12, NSE^{+}24, OIC^{+}23, THP21]. phenols [SMN^{+}14]. phenomena [ANMP15, Mit83]. phenomenon [DFD23, RV17]. phenomics [HHAR23]. phenoregion [KPSB22]. phenotypic [ACK^{+}13]. Pheronema [RTN90, VBJ^{+}20]. Philinidae [CES^{+}19]. Philippine [lLdZQ^{+}22]. Philippines [ZSH^{+}24]. Phoca [LAP10, RNL^{+}13]. Phocoena [RPRCAG^{+}21]. Phosichthyidae [CSM^{+}15]. phosphatase [IPG^{+}16, SAM^{+}04]. phosphate [CKT^{+}13, LCR^{+}93, MHGGS19, NNM^{+}21]. Phosphorus [DBRK17, PKV18, HSK^{+}19, IPG^{+}16, KEV10, McG64, PDD^{+}22, RDD^{+}18, SAM^{+}04, VBA^{+}18]. photic [TR99]. Photoadaptation [SPH83]. photographically [BBMR19]. photographs [DHB^{+}21]. Photoheterotrophy [EGPM^{+}15]. photophysiology [FDE^{+}22]. Photoreactivity [BFJ18]. photosynthesis [BK19, Epp92, KPSA17, Mor91, SPH83, TŠT^{+}17, TLM^{+}17]. Photosynthetic [NHG19, HHMB^{+}09, LCB18]. phthalate [PAG^{+}18]. phycotoxins [MPTMK22]. phyllosoma [WOW^{+}14]. phylogenetic [GMBU12]. phylogenetics [HCV^{+}20]. phylogeny [MSFZ19]. phylogeography [MMN12, YTL^{+}19]. Physical [ACE^{+}07, BP02, CMF15, FDE^{+}22, Gou85, HVRR15, HFPS^{+}06, KAK^{+}22b, RBD^{+}07, SOS^{+}07, SCD^{+}07, TNS^{+}05, UBB^{+}23, USH15b, WC15, BBE^{+}15, BC91, BHM^{+}15, BKC15, BL02, BHHS83, CBC^{+}06, CW06, CM11, CFM^{+}18, CZL^{+}24, CAB^{+}99, DBC^{+}23, DRVMC^{+}22, DPM^{+}09, DMBHG10, DMF^{+}09, EBW^{+}23, FF83, GL23, HKPV12, Hut95, ILA21, JHM^{+}22, LBSP01, LHC^{+}19, MMGL^{+}07, MHS^{+}20a, MHS^{+}20b, MJA^{+}07, MST^{+}23b, NM17, NMY^{+}14, NYH^{+}22, O’B83, OEL^{+}14, PIS13, PMA^{+}14, PKP14, PMC16, PFHM10, REG^{+}15, RBHLA04, RA15, RCSHW22, SWP^{+}13a, SFK^{+}99, SRAV19, SW22, SNMW10, SDJ14, TAO05, TCF^{+}18, VOJD02a, XC14, ZHD^{+}20, ZDM^{+}20]. physical-biological [DRVMC^{+}22]. physically [BCB^{+}05]. physicochemical [CEF^{+}13]. Physics [HAA^{+}14, BDT^{+}08, OELP04, PCK^{+}06, SSI13, SBM^{+}23]. physiognomies [HFO90]. physiological [GNH19, MSB^{+}23]. physiologically [BEP02, Mau10]. physiology [BLP^{+}20, LBP^{+}21, MLB^{+}20, RS10, ST03]. phyto [GRLS14]. phyto- [GRLS14]. phytodetritus [BMNW01, NKK03, TAW^{+}15, Tur15, ZCLS20, dWDB^{+}98]. phytopigments [FPD^{+}01, WDK^{+}01]. Phytoplankton [ÁLC22, CRGA17, CKM^{+}21, DGP^{+}13, HBL^{+}13, JHM^{+}22, LBH^{+}21, Mar03, MIW91, MA12, PWMIM91, RHM^{+}19, VHV^{+}12, Ven12, ZLG17a, ZLG17b, ZLX^{+}20, AJA^{+}22, AW13, AMG^{+}16, AJHC19, BSC^{+}19, BDP^{+}06, BFPS06, BE99, BTS^{+}15b, BGMP03, BPGD^{+}14, BLMR^{+}20, CKB^{+}17, CMC^{+}16, CSS^{+}21, CBM^{+}21, CRPS^{+}15, CBT07, CTI^{+}19, Dag93, DAvD^{+}20, DN07, DLM^{+}12, DNNNN16, DLC^{+}08, ERBV21, FHP83, FEGA^{+}14, Fly03, FPY^{+}16, FDE^{+}22, GBC^{+}00, GdRGC^{+}14, GHL15, GGAA^{+}23, GSPMAI99, GF19, GBB^{+}19, HBV^{+}99, HMX^{+}23, HCAFD^{+}20, HWL^{+}20, HPHL^{+}05, HBH^{+}17, iIYO^{+}10, JX18, KTH^{+}21, KMMC09, KMS^{+}24, KPSB22, LMS93, LOG^{+}09, LSM^{+}22, LRAE23, LHP^{+}05, LMT^{+}19, LLX^{+}21, MY23, MHGGS19, MCKS17, MHCR^{+}12, NEI^{+}22, NHSP23, NSE^{+}24, NGNV12, NKK^{+}05, OBD^{+}20, PFHM16, PVM^{+}20, PV07, PPY87, PPD^{+}12, RM93, RSMIS03, RBPGJ^{+}20, RGMPR23, RSM^{+}23, RPSC22, iSIS02]. phytoplankton [SVL^{+}23, SiSI^{+}02, SSN23, STW^{+}15, SPWH21, SHS^{+}05, TŠT^{+}17, TLM^{+}17, TSFA22, TRMV15, VGM^{+}23, VMN08, VBM21, VMC^{+}19, WFH^{+}22, Wal83, WCC^{+}20, WLL^{+}23, WCS^{+}23, WLM^{+}13, WSH^{+}22, XWL^{+}18, ZWM^{+}15, ZHD^{+}20, ZJZ^{+}21]. phytoplanktonic [BFJ18]. PIC [LCGH07]. PICES [BK08, HMWM00]. PICES/GLOBEC [BK08]. pico [DDP^{+}00]. pico- [DDP^{+}00]. picoeukaryotic [CGS23]. picophytoplankton [CLX^{+}20, PBD^{+}88]. Picoplankton [ZSBL00, BAM^{+}09, TB15, WSC^{+}21]. picture [MVV^{+}19]. pie [RSB^{+}13]. piece [KVNT20, VBA^{+}18]. Pierre [BMG^{+}21b]. Pigment [MKMF^{+}89, GBC^{+}00, LSM^{+}22, SHS^{+}05]. pigmented [FEGA^{+}14]. pigments [CMC^{+}16, DN07]. PII [SDS22a]. pilchardus [BCGN^{+}18, CCHV^{+}21, SNV^{+}18]. pilot [HPB^{+}09]. Pink [KBF^{+}08, Kli10]. Pisces [CSM^{+}15]. Piscicolidae [UKK^{+}19]. piscivorous [EKB06]. pit} [Bak06]. pitfall [AM10]. Pitzer [FM07]. Placopecten [GGJ^{+}10]. Plain [BBR^{+}01, HBD^{+}21, HCV^{+}20, IBW^{+}01, VSC01, VPW01, CES^{+}19, SSKA19, BKD^{+}20, BGS^{+}04, RBL90, VDP^{+}01, WDK^{+}01]. plains [BHB^{+}19]. planet [Bak83]. planktivorous [RG94, dSSDS^{+}20]. Plankton [AFBT^{+}22, BPA^{+}21, BGMP03, BDE03, FGGDF^{+}04, GMDD^{+}22b, Hau84, Hof10, HFK03, JJS03, LJPGC02, MGC^{+}18, NGLL^{+}22, Peñ03b, AV23, ABT^{+}04, AKH^{+}23, AUE^{+}14, BTNK13, BAOC^{+}09, CDH^{+}13, CWB^{+}22, CNT03, CGS23, Col69, CFML22, CNSHT15, FWH^{+}17, Fra69, Fro93, FDM^{+}13, FG16, GMAGH^{+}17, GMDD^{+}22a, GCD^{+}13, HPB^{+}09, IVT^{+}12, JSHB90, KZD^{+}19, LGK^{+}93, LLGS21, Lon85, LH89, MSMH19, Moh15, MPC12, MCT03, OMR^{+}22, PVC^{+}08, PTF12, PMC16, RMC^{+}15, RWJ^{+}06, SHL13, STF^{+}13, SPSV^{+}20, SS17, TSAM^{+}22, TSL10, TS10, TSBS18, VDDA^{+}08, VR03, WMWR08, WH94, Woo05, WPB05, dB94]. Planktonic [CQZ^{+}18, GRdSS^{+}22, HNL14, NYL^{+}17, NFMCS^{+}22, TCF^{+}18, AMEV07, AAM^{+}14, ABSDC07, AOMZ^{+}23, AT07, AHW^{+}15, Bri79, Den03, FTG^{+}11, GMBU12, GGJ^{+}10, GHC^{+}17, JE92, MVN^{+}15, MPM^{+}18, MCGR07, MMF^{+}12, MGH^{+}07, Par65, PAPL15, Peñ03a, RLX^{+}24, RAG^{+}19, SBMB18, STS^{+}12, SJJ^{+}03, SDJ14, SPB93, TFM03, UB10, Wil65, Ang84]. planning [LRGV^{+}18]. Plans [Mos69]. plant [SMM^{+}90]. plasticity [KSB^{+}22]. Plateau [FMC^{+}15, GIHJ23, MFA^{+}15, PWZ^{+}16]. plateformes [Ber65c]. platform [KC02, WZC20]. platforms [RLP^{+}18]. platypterus [RLL^{+}09]. plausible [CMF11]. played [OKdA^{+}19]. Pleistocene [Ban65, BW65, Don65, Emi65, MBP65, Sel65, ST65, Wil65]. plenary [Sie88]. Pleuromamma [HTG15]. Pleuroncodes [GRB^{+}08, YCP^{+}12]. Pliocene [BW65, GWGR^{+}19, MBP65, Sel65]. plumchrus [MFB^{+}84, FMT15, LGK^{+}93, Mil88]. plume [EKB06, HW02, HLS^{+}14b, JX18, Wen88, MSd^{+}16]. plumes [DJW^{+}18, MSd^{+}16, SLM^{+}16, WF17]. plutonium [HKY^{+}11]. Plymouth [USH15b]. POC [LCGH07, LHC^{+}19]. POEM [MRMD^{+}97, ÖHÜ89]. POEM-Phase [MRMD^{+}97]. point [CNBD21, DW02, FWO15]. poisoning [CAH^{+}22]. polaires [Rou65]. Polar [MHA^{+}11, Rud15, BRD^{+}15, BF11, BGL^{+}17, GCD97, HDA^{+}16, KTH^{+}21, KSG^{+}17, Rou65, Rud89, SBH^{+}14, VMH^{+}21, WMB^{+}21, WBA^{+}22, NBR^{+}08, STC10]. Polcevera [CLD22, DSC^{+}19, GBB^{+}20]. Pole [SLGI^{+}21]. poleward [Kos02, SDGVE17]. Policy [HSC09, SSB^{+}20b, Val99a, Val99b]. pollen [GGE^{+}65, YSY^{+}19]. Pollock [PDAM^{+}15, BCB^{+}05, GTS^{+}21, MLPN06, YNM^{+}02]. pollutants [LB14, PAF^{+}11]. pollution [CC88]. polychaete [JP90, LMPB^{+}16, VCSG^{+}01]. polycyclic [FTG^{+}18, SGL^{+}17]. Polycystina [BC16]. polymetallic [BJMP19, BJMP20]. polymorphism [Sma10a]. Polynya [SDL^{+}19, THM^{+}06, BvdLA^{+}11, Hol00, YLL19, HFO^{+}22]. polyunsaturated [KO19, WPB^{+}08, WL16]. polyxystra [LDAM^{+}07]. pompejana [JP90]. pool [GTS^{+}21, HHB^{+}00]. pools [ELW06, FGL^{+}23]. poor [EAB^{+}23, GAS^{+}22]. Population [KSKN21, LS12, AEP^{+}23, AH10, BCGN^{+}18, BAP^{+}22, BB10, DLL^{+}23, DBM17, FRCH15, GGA^{+}16, HMP^{+}13, JLP^{+}20a, JLP^{+}20b, LMS10, LSS^{+}10, LAHI10, MAH^{+}15, Mau10, Reb02, RA15, SIR^{+}07, SSL08, SBFP21, STS^{+}12, SDH^{+}14, SAB^{+}21, SJD10, UB10, WBF^{+}21]. populations [ABSDC07, APC^{+}21, ALT10, Bak01, Bak06, BBLD^{+}11, BCT^{+}09, BEP02, CNSHT15, GRLS14, GPC^{+}03, HRSM08, HFW^{+}98, hHRW^{+}05, KQP^{+}17, LSM08, Mau17, MM90, PRTC13, PCR^{+}22, YAK^{+}08]. Porcupine [VPW01, BBR^{+}01, BKD^{+}20, BGS^{+}04, HBD^{+}21, HCV^{+}20, IBW^{+}01, RTN90, VSC01, VDP^{+}01, VBJ^{+}20, WDK^{+}01]. pore [KGdS^{+}08, PRL^{+}18, Wil65]. porewater [SSTL16]. porpoises [JHW^{+}14, NHE^{+}13]. portion [MG02]. Portland [MB07]. Portugal [ÁSFP^{+}03, EAL^{+}07, JGO^{+}98, KGdS^{+}08, MLS^{+}15]. Portuguese [AJV^{+}02, OVR^{+}02, VOJD02a, VOJD02b]. POSEIDON [DEW^{+}97]. position [GWM^{+}22, HLTB^{+}17]. positive [BBLD^{+}11, HHSR07]. possibility [SW65]. Possible [AB90, IIS^{+}17, Ber65a, Car98, CS03, Coo65, Cra09, CS04, KMF^{+}20a, KMF^{+}20b, RTN90, RVS^{+}21, SNZ^{+}20, SA97, TKWI08]. possibly [ZBY^{+}22]. Post [OTNI20, DCL^{+}13b, INO^{+}24, WMWR08]. post-bloom [INO^{+}24]. post-glacial [DCL^{+}13b]. Post-spring-bloom [OTNI20]. postlarval [VDGGD^{+}22]. posts [PMG15]. Potential [CALS^{+}23, CZL^{+}24, DVB^{+}18, HSC^{+}16, LRW^{+}15, MLK^{+}09, MMF^{+}12, VMV^{+}23, YZX^{+}23, YWUK15, BM07, CPO^{+}19, DRVMC^{+}22, FH95, Fei03, Fei04, FGL^{+}23, GGT^{+}15, GGQ07, GCG^{+}14, HWLT10, HSN^{+}18, JLRB20, JSLA^{+}21, KV13, LYS^{+}22, LPF^{+}18, MVN^{+}15, MPMFL^{+}23, McD88, MDR20, RASVB^{+}22, RDP^{+}21, Sak86, SBL^{+}23, SJD10, SPB93, TLM^{+}17, WAH^{+}20, WWW^{+}23, ZCA21]. potentially [VBJ^{+}20]. poutassou [MAFS^{+}22]. power [Iwa23, YYhT^{+}17]. pp [Ang79a, Ang80, Hof81]. Pre [OMS^{+}09, CMS^{+}13, PCSMC12]. Pre-conditions [OMS^{+}09]. pre-recruitment [CMS^{+}13, PCSMC12]. precipitation [MJWK07]. precision [MLHE23, TIOM16]. precursor [STW^{+}15]. predation [AHSS22, AIA^{+}15, DWFP^{+}19, HSL96, LZF^{+}24, OMR^{+}22, TS10, WBC^{+}22]. Predator [PTF10, PTF12, AGD^{+}18, AUE^{+}14, Bak06, DAF^{+}22a, DAF^{+}22b, GPEV20, KM10, LAD^{+}18, LMP22, NBLI20, STEB16, SDO^{+}14, SIB^{+}06]. predator-driven [SDO^{+}14]. Predators [LM10, Mau10, AHP19, DYO^{+}10, GCG^{+}14, HF10, Jac10, LMS10, SFS^{+}12, SSL08, TSS^{+}12]. predatory [SCS^{+}18]. predict [QSC^{+}15, SWP^{+}13a]. Predictability [BBL^{+}18, JAS^{+}20, OLH^{+}18, TBS^{+}19, WPB05]. predictable [KCPM09, STC10]. predicted [WST^{+}16]. Predicting [CLB^{+}13, LRGV^{+}18, MCD^{+}14, MDB^{+}20, SAY^{+}16, RSMIS03, VSPP14]. Prediction [Ste12, FY88, HSS^{+}12, HKK12, JAS^{+}20, NHS^{+}14, PBH^{+}10, RL85, SPV^{+}15, VBL^{+}09, VVV21]. predictions [GRLS14, SCHD23]. Predictive [BFP^{+}18, BMGN15, PRC^{+}20, PYKF15]. predictor [CARBML^{+}22, LPARF^{+}20]. Preface [Ano03g, CH07a, RW97, Sea63, SGPdM18, Wyr06, SKHD84, War73]. preference [CdD^{+}15]. preferences [ARD^{+}03, LAD^{+}18, LSIC12]. preferred [RVC^{+}13]. prehistoric [DVL^{+}99]. Preindustrial [ZCD08, LHW^{+}20]. prejudice [MRH^{+}18]. Preliminary [LSS^{+}10, ZHSMM14, Ken88, WO85]. premier [STGR^{+}23]. prerecruit [LLS01]. Presence [LAP10, ANMP15, BPSN^{+}21, CSLJ03, DHD^{+}23, LML^{+}23, SPW22]. Present [DJG^{+}02, BCD^{+}20, Cai95, CBGC^{+}08, KKK04a, NW87, SBB^{+}14]. present-day [Cai95]. Preservation [SLG^{+}12, LvIKB07]. Press [Ang79a, Ang80, Ang88, Bak83, Hof81, SW81]. pressure [BF01, ERT^{+}22, FWL^{+}15, HWB^{+}18, KSK21, LM00, LGZ^{+}20, RPG^{+}18, SV97, ZZPL18, ZYN^{+}24]. pressure-recording [ZZPL18, ZYN^{+}24]. pressures [BVJE19, KSE^{+}09]. prevailing [TZP^{+}00]. previous [Ano65c, Ano65d, Ano69b, Ano73b, Ano85c, Ano86a, Ano87a, Ano89a, Ano92a]. Prey [HWL^{+}20, SRT^{+}18, AHGRAL23, AGD^{+}18, AUE^{+}14, BFB^{+}20, BJ90, CCHV^{+}21, Cra09, DRVMC^{+}22, DAF^{+}22a, DAF^{+}22b, GCG^{+}14, HBG^{+}21, HBK^{+}24, HF10, LOBG^{+}10, LPF^{+}18, LRJ^{+}15, PTF10, PTF12, VWDF14, YGL^{+}10]. prey-predator [DAF^{+}22a, DAF^{+}22b]. prey-switching [VWDF14]. prey/predator [AUE^{+}14]. Priceless [PG10]. prices [PG10]. pride [MRH^{+}18]. Primary [CCB^{+}20, FFA09, IVT^{+}12, JRW01, KNI^{+}05, MB05, PMK^{+}06, PS08, PTI00, SMM^{+}90, WFH^{+}22, WSG^{+}93, AHGRAL23, AvD15, BBE^{+}15, BTS^{+}15b, BCOL^{+}19, BTJ^{+}17, CKT^{+}13, DAvD^{+}21, Dem09, FEGA^{+}14, GSPMAI99, GF19, HMO^{+}13, HSC^{+}16, JTD^{+}14, JJL^{+}19, KFKO03, KLB^{+}21, KM08, LSH^{+}11, MBP^{+}11, MOS^{+}13, MC15, MZK^{+}23, MSL^{+}07, MST^{+}23b, NHH^{+}23, PAM^{+}88, STS^{+}12, SSN23, SR15, SEW11, SKCP23, TBS^{+}19, TLM^{+}17, TNY^{+}24, TAM^{+}15, VEM^{+}21, WFS^{+}15, XLX^{+}20, ZGZ19, dlGFM^{+}23, vRGW10]. primer [Gar03]. Principal [YKWF21, KDF97]. principles [AV23, LMA^{+}15]. prior [FGGDF^{+}04, Rud15, SSM^{+}90b, TSFA22]. priorities [FDH20, HSH^{+}19, IPF23]. priority [JAS^{+}20]. prism [JFG^{+}90, OSH^{+}96]. pristine [GHL15, MPN09]. pro [CFC^{+}18]. pro-delta [CFC^{+}18]. probabilistic [AHC^{+}13]. probability [HLP^{+}16]. probe [SPH^{+}15b]. problem [Cox63, CPC88, KN10, KN11, MFH86]. problems [Ang80, Pir87, Sei63, Sva65]. procedures [RAB^{+}84]. process [QSC^{+}15, ŠPM^{+}22, VMN08]. process-oriented [ŠPM^{+}22, VMN08]. Processes [IPD14, TvG02, XD96, AHA^{+}16, Ano94c, BP02, BRD^{+}15, BVB88, BPGC^{+}20, BCR^{+}13, CB09, CZL^{+}24, CLA^{+}00, CVHM^{+}18, DJG^{+}02, ESTM^{+}12, FFA09, Fro93, GBC^{+}15, GTS^{+}21, GLS08, HGBG20, HKE^{+}10, HPNDC15, Hut95, IGG^{+}19, JGS90, JST^{+}24, KSK^{+}15, KNS^{+}03, Kit03, KZD^{+}19, LCB18, LGR^{+}02, LC22, LHC^{+}19, LFBP^{+}13, MLL^{+}22, MY23, MHS^{+}20a, MHS^{+}20b, MVN^{+}15, MKM93, MZZ^{+}23, MIH06, ML09, MJA^{+}07, MST^{+}23b, NGLSSG14, NMY^{+}14, NNO^{+}14, OOTA15, ORB^{+}18, PKP14, QCdS^{+}07, RKS01, RGC^{+}01, RLDC^{+}13, RGI05, RHB23, RBHLA04, RAG^{+}19, Rud15, SOS^{+}07, SRAV19, SW22, SSM90a, Soh03, SÖÜ94b, Tho95, TCL^{+}15, UBB^{+}23, USH15b, VPW01, VOJD02a, VPH^{+}12, WHBK05, WH89, XC14, XCH^{+}16, YAI^{+}14, YFY^{+}22, vWMH98, vWM02a]. processing [BCLD^{+}17, ZCLS20]. Prochlorococcus [RA15]. prodelta [PRL^{+}18]. produced [KSG^{+}17, KMWF11]. producers [LSH^{+}11]. product [STW^{+}15, dPCS23]. Production [RWOA01, STHM02, AAMB^{+}24, AvD15, BBE^{+}15, BTS^{+}15b, BCOL^{+}19, BLP^{+}20, BD20, BKD^{+}20, BD18, BTJ^{+}17, BPTT19, CKT^{+}13, CRPS^{+}15, CCH^{+}12, CCB^{+}20, Dag93, DAvD^{+}21, Dem09, Epp92, FJA^{+}21, FFA09, FEGA^{+}14, FLDF22, Fro93, GdRGL^{+}01, GGPG^{+}19, GGQ07, GWK17, GSSWK20, GSPMAI99, GF19, HL05, HBL^{+}13, HSG^{+}15, HMO^{+}13, HWL^{+}20, HGH^{+}19, HHW01, HHW22, HSC^{+}16, HVEF09, IVT^{+}12, JTD^{+}14, JSdSS^{+}21, JJL^{+}19, JW01a, JRW01, KFKO03, KLB^{+}21, KON14, KV13, KM08, KPSA17, KC02, LQU07, LCGH07, LLS01, LH89, Lon95, MOS^{+}13, MLK^{+}09, MC15, MZK^{+}23, Mil93b, MIW91, MTK^{+}22, MSL^{+}07, MST^{+}23b, NYL^{+}17, OMR^{+}22, OWR^{+}07, OJB99, OACB^{+}15, PMK^{+}06, PAM^{+}88, PD15, PB94, QLW10, RM93, SCAA07, SMP^{+}22a, STS^{+}12, SJJ^{+}03, SSN23, SR15, SDH^{+}14, SEW11, SMM^{+}90, SE92, SBC^{+}24, SYN^{+}21, TBS^{+}19, TGJT09, TLM^{+}17, TRP^{+}23]. production [TNY^{+}24, TAM^{+}15, VCB^{+}00, Ver91, VEM^{+}21, WSG^{+}93, Whe93, WFS^{+}15, XLX^{+}20, YMA^{+}17, YHRT22, YHM^{+}18, ZGZ19, ZHD^{+}20, dlGFM^{+}23]. production/export [HGH^{+}19]. Productive [MIN^{+}20, CSS^{+}21, CHC^{+}12, DMC^{+}18, LHP^{+}05]. Productivity [BHA^{+}14, MDC^{+}07, AHGRAL23, BW08, CMHM18, CQC15, DGP^{+}13, GCV^{+}24, HBV^{+}99, JTQ^{+}18, KTH^{+}21, KNI^{+}05, LT06, LNB13, LdSH^{+}15, MCD^{+}14, MB05, McK08, MMN^{+}24, MJA^{+}07, NHH^{+}23, NGNV12, PS08, PCH^{+}08b, PTI00, RCS^{+}11, RFKC16, RGM01, SGF^{+}19, STF^{+}13, SGO^{+}08, SKCP23, SIB^{+}06, STGR^{+}23, TDGY22, WFH^{+}22, vRGW10]. products [WWL^{+}24, YN20, ZD17]. productus [FB05]. PROFAN [MPTMK22]. professor [Bru88, JW01b, SMB88, Sie88]. profile [BMGN15]. Profiles [LH89, Ang89, ASR^{+}20, Gam14, HKY^{+}11, KPSA17, MMF^{+}17, NKK03, PAF^{+}11, RSMIS03, RMB^{+}01, WOW^{+}14, YYK88]. profiling [BOG20, KSK21, KHM^{+}88, TDH^{+}95]. profondes [Rot65]. Prog [ÁBMÁS15, GFB^{+}15b, KN11, MFS^{+}16a, SE09, VH09a, WF07, Yas07b, ZLRVB24a, dMGS^{+}11b]. Progr [BJMP20, FDH20, JLP^{+}20a, KMF^{+}20a, MHS^{+}20a, RBS^{+}20]. Program [BK08, AH15, EBD^{+}20, JRR^{+}24, PDV12, Reb02, RG09]. programme [Ano17a, BPW10, BR01, RNP^{+}17, ALV^{+}21, AB00, PHCA17]. programmes [HM15, OELP04]. programs [HMKF08]. Progress [Ano94c, BDB^{+}22, BLAM00, CGL^{+}20, Fei04, HHW22, Kru19, LGH^{+}21, OEL^{+}14, RLSF07, RG03a, SDS22a, SHC^{+}07, VHK04, CAT^{+}08, vdS94b, WR03]. progression [KMS^{+}24, STW^{+}15]. Progressive [RMC^{+}15]. Project [WSG^{+}93, APSC11, BN03, Kit03, SJA^{+}23, SGPdM18, TP00, VBL^{+}21, VYGMM^{+}17, vAB96, BC99, Dri11, DP13, NF87]. Projected [WO15, ABP^{+}23, SJD10]. Projecting [CBOP15, SBM^{+}23, TNGP22, MST^{+}23b]. projections [SBC^{+}24, TLH^{+}15, TLP^{+}16]. projects [SPK^{+}19]. prokaryote [CQZ^{+}18, LQU07]. prokaryotes [GLAHH^{+}22, RCC^{+}18, SCHS^{+}24]. prokaryotic [CRC^{+}19, GASV^{+}09, ORMR^{+}19, RPG^{+}18]. prolonged [HZD^{+}23]. promotes [WCC^{+}20]. promoting [ZJZ^{+}21]. promotion [KSC10]. pronounced [RS10]. propagating [UKM^{+}14]. Propagation [KAG^{+}19, ZKT88, BPSGP^{+}23, His22, Oll15, RKS01]. Proper [Sud86]. Properties [BNCC15, Kat18, AGS10, BDB^{+}04, BCR^{+}13, CTMV^{+}14, CR20, GCS91, HGD22, HMX^{+}23, JZZY24, JSKM02, KF11, KMU^{+}12, LF12, LC10, McD81a, MVC^{+}11, Mit91, NF06, NBLI20, NIF^{+}15, PM22, PFW15, SHP^{+}23, SW92, SRG^{+}19, TII^{+}14, TTL^{+}04, WMB^{+}18, WST^{+}16, YNM^{+}02]. property [GJ00, HHMB^{+}09, McK08]. Proportion [WLL^{+}23]. proposed [WGCS13]. prospect [CLdPHL23]. Prospects [CSS^{+}19, ALG^{+}21]. protected [EAB^{+}23, JSLA^{+}21]. protection [BDE03]. proteins [YT06]. protist [CQZ^{+}18, GRdSS^{+}22, MGA^{+}23, ZPC^{+}16]. Protista [BC16]. protistan [SWH^{+}24]. protistan-bacterial [SWH^{+}24]. protobranch [AS96]. Protozoa [Gif93, Mae88, TSFA22]. protozooplankton [SSV^{+}11]. provenance [GGE^{+}65, PGT^{+}13, YSY^{+}19]. provide [ALM^{+}23, GBT^{+}19]. provides [EB08]. Providing [dLLdAWL^{+}23]. Province [CLG^{+}22, SQJ^{+}17, MPTMK22]. provinces [HRA00, LJM^{+}16, MRW^{+}14, SW21]. provisioning [KGJ^{+}10]. proxies [CLV^{+}19, MDL^{+}12]. proxy [FPJ^{+}15, GEPC15, SOH21]. Pseudo [PTPY^{+}23, VSPP14]. Pseudo-nitzschia [PTPY^{+}23, VSPP14]. Pseudocalanus [HLPL05, LPHL^{+}05a, PLHLF05]. Pseudoikedella [GS19]. Pseudostichopus [RMB^{+}01]. Pseudotanaidae [JPB20]. PSP [CAH^{+}22]. Psychropotes [GKR20, RMB^{+}01]. Psychropotidae [GKR20]. pteropod [JMZ23, MLB^{+}20]. Pteropods [BHK^{+}16, HPH^{+}08, KCBS20, LS13, AOMZ^{+}23, BGWP^{+}17]. Pu [GPA^{+}11]. publication [Ano13h, Ano13i, Ano13j, Ano13k, Ano14a, Ano14b, Ano14c, Ano14d, Ano14e, Ano14f, Ano14g, Ano14h, Ano14i, Ano14j, Ano14k, Ano15a, Ano15b, Ano15c, Ano15d, Ano15e, Ano15f, Ano15g, Ano15h, Ano15i, Ano15j, Ano15k, Ano15l, Ano16a, Ano16b, Ano16c, Ano16d, Ano16e, Ano16f, Ano16g, Ano16h, Ano16i, Ano16j, Ano17b, Ano17c, Ano17d]. Publisher [Ano07s, Ano08x, Ano19o]. Publisher‘s [Ano03i, Ano03h]. Puffinus [SSB14]. Puget [CCS^{+}21, SKGS20]. Pulley [KAK^{+}22b]. Pulmo [VDGGD^{+}22]. pump [CLCBB19, FZY^{+}23, GDI^{+}09, HMKF08, HKE^{+}10, HAH^{+}22, LRW^{+}15, LH89, RGI05, RGE22, Tur15, XYGJ23, SHK^{+}14]. pumping [HZCZ16, HNR^{+}17]. purposeful [HHP06]. Puyuhuapi [SPSR^{+}14]. pycnocline [SASH08]. pyramid [XYK^{+}22]. pyrenees [Bou65]. Pyrosoma [OBD^{+}20].

quadrangularis [PRA^{+}18]. qualitative [WHBW03]. quality [BMG^{+}21b, CLSP17, CLSD18, KS06, LSH^{+}11, MFDH22, NRA^{+}21, SJJ^{+}03, ZD17]. quantification [dCFK17]. Quantifying [IHR18, NRA17, SL13, BCG^{+}08]. Quantitative [NNO^{+}14, BTNK13, FAAV^{+}15, KHS^{+}14, RSMIS03]. quantities [Nee85]. quantity [CLSD18, LSH^{+}11]. Quasi [SGMP15, Mol22, dSPF^{+}23]. quasi-decadal [Mol22]. quasi-stationary [dSPF^{+}23]. Quasi-synoptic [SGMP15]. Quaternary [Hay65, HK65, Sai65, Con87]. Questioning [AHW^{+}15]. questions [FAB^{+}09]. quick [RKS01]. quiescent [VBAC^{+}21]. quiscent [WP91]. quotas [BPA^{+}21].

R [Ang88, SKF20]. R/V [SKF20]. Ra [IHT^{+}21, LvBS^{+}24]. Radar [VOT^{+}99, BBM^{+}14, Ric94, Pra91]. RADIALES [VBL^{+}21]. radiated [CGW^{+}22]. radiation [HYM^{+}12, LLL^{+}24]. radiocarbon [KMWF11]. Radiolaria [BC16]. Radiolarians [Coo65]. Radionuclide [WJE^{+}92]. radiotracers [LPA^{+}11]. radium [LCJ^{+}17, WST^{+}21]. RADMED [VYGMM^{+}17]. Radon [CB06, OP18]. Radon-222 [CB06]. raised [SCS87]. Random [RCGC^{+}16]. range [ABSDC07, ALT10, BK19, DOP87, DEW^{+}97, ESTM^{+}12, HTG15, NRA^{+}21, OBD^{+}20, VDS^{+}18]. Rapid [Bel09, EFC^{+}23, KRHS14, ORPRGIS22, RCC^{+}18, SCHBC^{+}22, MHVS19, Was11, ZWM^{+}15, ZCLS20]. rapidly [BHM^{+}15]. Rare [GSPP^{+}20, EAB^{+}23]. raschii [MPSD15]. rate [Hey78, HTdM^{+}15, PPSV^{+}18, Tur65]. Rates [KB65, ÁSÁB^{+}14, ARDP14, BS95, CBHL07, CRC^{+}19, FCEZ10, GvOS^{+}08, HL05, HHMB^{+}09, HF10, KKS^{+}03, KBF^{+}08, Kli10, KSY^{+}19, KSKN21, LS12, OTNI20, PTF10, PTF12, PD15, PZA^{+}15, SCMAR^{+}99, SIR^{+}07, SKSK06, STHM02, VEM^{+}21, Whe93]. ratio [SS03, Soh03]. rationale [KNS^{+}03]. ratios [ÁSÁB^{+}14, BPA^{+}21, CPPPEAG22, DAF^{+}22a, DAF^{+}22b, IVR^{+}13, KSS^{+}23, KG65, KAAK^{+}16, LDD^{+}22, NKK^{+}05, NAH^{+}21, YAI^{+}14]. ray [RHB23]. Re [YYT^{+}14, MB01, McK08, PMG15, VPH^{+}12, ZLR^{+}07]. re- [PMG15]. re-analysis [McK08]. Re-evaluation [YYT^{+}14, VPH^{+}12, ZLR^{+}07]. re-occurrence [MB01]. reach [CFM^{+}18]. real [AIA^{+}15, CP83, QSC^{+}15, TGR05, ZZM^{+}24]. real-time [AIA^{+}15, CP83, ZZM^{+}24]. realism [Fly03]. realities [Val99a]. Realized [XWL^{+}18, ZLX^{+}20]. really [UAM05]. reanalyses [AASJ23, CTKF^{+}23]. Reanalysis [HNL14, ASJ^{+}23, BBC^{+}22, CMG15, SOB^{+}08, SCC14, SKCP23, TMH^{+}16, WLM^{+}22, YN20]. rebuilding [KN10, KN11]. Recirculating [McC92]. recirculation [Hog85]. recognition [CM11, HMRB^{+}03, Kaw98]. recognize [KV18]. REcoM2 [SKWWGV18]. Recommendations [HM15, RRLS22, SPK^{+}19]. reconciliation [GCP08]. Reconciling [OMS^{+}15, Ste04]. reconstruct [RCD^{+}94]. reconstructing [RD11]. reconstruction [FAAV^{+}15]. reconstructions [VBL04, WM13]. Record [PVG^{+}20, BDB^{+}04, Bol94, BC19, BHK^{+}19, DLD^{+}19, Emi65, GHSC19, LHE^{+}13, RN02, Sha82, SGO^{+}08, VOG^{+}08]. Record-breaking [PVG^{+}20]. recorded [UKK^{+}19]. Recorder [BGMP03, BDE03, HFK03, JJS03, BTNK13, RWJ^{+}06, WH94]. recorders [RCM^{+}03]. recording [Dah69, PMFNGQ21, ZZPL18, ZYN^{+}24]. records [ALG^{+}21, BB65, CGG08, HHWW20, MDL^{+}12, RvBD^{+}22, TAF^{+}22, UGY^{+}22, UCB^{+}18, WMWR08]. recoverability [Dol09]. recovery [MMD^{+}16, NMC^{+}09, PBBH^{+}22, dJSL^{+}20]. recreational [JOGM^{+}10]. recruit [IFC^{+}07]. Recruitment [BCB^{+}05, CGV13b, HSGJ23, SEG^{+}22b, VPH^{+}12, ZK06, BGL^{+}17, CDTM^{+}21, CMS^{+}13, FARRL^{+}13, GSM^{+}17, GEPC15, HMRB^{+}03, HLS^{+}14a, HCGK11, IIM^{+}23, KKKY10, MHTG10, MHS^{+}09, MKSvA^{+}22, MAFS^{+}22, NGPH10, OACB^{+}15, OAWAN18, PCSMC12, RKC^{+}10, Tit20, TIOM16, VCSG^{+}01, WHI^{+}02, XRC^{+}15, ZL01]. recruits [OACB^{+}15]. rectified [YYhT^{+}17]. recurrence [ATS01]. Recurrent [VKT15]. Recycling [LHEB98, CRC^{+}19, RWD01]. Red [ZLRVB24a, ZLRVB24b, AHRT90, WWL^{+}22, BF12, KAAK^{+}16, KZD^{+}19, Men21, SCHS^{+}24]. redefinition [SCLS10]. redescription [Mil88]. Redfield [Fly10]. Redfieldian [MHGGS19, FCMCÁS19]. redistribution [SAH^{+}21]. redox [DYL^{+}15, KGdS^{+}08, SGO^{+}08]. reduce [RSK^{+}23]. reduced [DWFP^{+}19, MSV^{+}14]. reducing [CDB^{+}24]. reductase [GGPG^{+}19]. reduction [ONR^{+}14, WSO^{+}13]. Reef [BMM97, BECR^{+}22, CPPPEAG22, CALS^{+}23, CLG^{+}22, CKL^{+}14, NYL^{+}17, ROBRB^{+}22, RHB23, VGJ^{+}19, YW22, BECR^{+}22]. reefed [RLP^{+}18]. reefs [FSAO22, KAK^{+}22b, MP04, WFJ^{+}15, WLP^{+}21]. Reevaluation [RK20]. reference [Dol09, KDB95, MWFH02, NMN08, SYN^{+}21, TR99, Ver91, WFH^{+}22]. refinements [MCKS17]. reflect [CLV^{+}19, GLAHH^{+}22, KBF^{+}08, PGY^{+}22]. reflected [PG10]. reflection [LCZ^{+}24]. refs [Bak83]. refuge [BF11]. regeneration [FUOG^{+}16, FFA09]. Regime [AN04, BHAJ12, CRS04, CS04, RS04, Ste04, WZ04, ANH21, AAMB^{+}24, Bea04, BDC^{+}08, CAT^{+}08, CHB02, Dri06, FC07, FAH^{+}13, FPS^{+}13, HM00a, HMWM00, INI^{+}17, IIM^{+}23, iIRM^{+}15, KFKO03, KJZ^{+}12, KFG^{+}03, Law04, LM14, LHF^{+}16, Man04, MS00, MFY^{+}86, ORMB08, PM13, Qiu15, SvN04, SKH00, SKT01, TOKLC08, TKW06, TKWI08, VBL04, WXH07, YTNK00, ZLKO00, dHA^{+}04]. Regime-shifts [BHAJ12]. Regimes [BBSN04, AIHB^{+}07, AVS23, BDTC15, FJA^{+}21, HSK^{+}19, LdSH^{+}15, MB01, Pow06, ZHF^{+}24]. Region [INI^{+}17, Ang79b, BHA^{+}14, BAT^{+}98, Ber65b, BB14, BHAJ12, CLV^{+}19, CGMP14, Dav85, DL69, DIQJ21, FPIJ85, GMAMB04, GDSCU09, HSMLDC^{+}22, HE07, hHRW^{+}05, INO^{+}24, IMM^{+}22, iIYO^{+}10, LJPGC02, LvBS^{+}24, LMC^{+}20, LSW02, MPV12, MSA^{+}22, MTK^{+}22, NNM^{+}21, OTNI20, PMC21, PHC^{+}19, RCGC^{+}16, RBR^{+}23, ST03, SOB^{+}08, SDGVE17, SGMP15, SBH^{+}14, SJM^{+}19, TKSI08, TKW06, VVV21, WBD^{+}15, ZHD^{+}20, ZZM^{+}24, ZBRJ23, ZHBW01, FMCG15]. Regional [DMBB02, GZCL23, HVTV22, KJZ^{+}12, LHC^{+}21, MZGA^{+}20, RLR^{+}18, SIR^{+}07, ABÁS^{+}09, BGV^{+}23, CB17, DBC^{+}23, DRE^{+}08, FP15, HHDS02, Hof81, HSC^{+}16, HDB13, LPF23, Mac98, MB20, MEMC05, NGPH10, RCS^{+}11, SOA^{+}23, TAM^{+}15, WSL20, YRKC08, MAB^{+}11c, MAB^{+}11a, MAB^{+}11b]. Regional-scale [HVTV22]. Regionalisation [AIA^{+}18]. regionalism [KPM^{+}23]. Regions [ALV^{+}21, AP20, BEP02, BBL^{+}18, BBRM20, DMF^{+}09, DBRK17, ESTM13, FYYC05, Lav09, MKOLA20, O’B83, RGB^{+}17, Rou65, RAG^{+}19, SCB^{+}16, VBAC^{+}21, WMC^{+}89, WWL^{+}22]. regularities [PMA^{+}14]. regularly [Nof96]. regulate [Mau17]. Regulating [NF87, Fro93]. Regulation [SC23, BM01, HBL^{+}13, HCGK11, KSC10, MC15, OWR^{+}07, Tit20]. regulations [HLSX22]. regulators [Law04]. related [AQVB^{+}10, AHRT90, BLES16, CMF^{+}09, CW02, CP02, FTC^{+}16, GSV^{+}01, GdRGL^{+}01, IPG^{+}16, Mit83, MR03, Nag01, OELP04, PCH08a, Pir87, RR01, SSTL16, UAM05, WXH07]. relatedness [BLES16]. relates [DBC^{+}23, LLAPG^{+}22]. Relating [DBR03, HMRB^{+}03, Bri79, Leg91]. Relation [MNT14, ARG11, AE09, BCOL^{+}19, BGB^{+}08, BDBJ01, BMG^{+}19, BDC^{+}08, DDE^{+}95, FMH02, HPC^{+}20, HGD22, HHY03, KON14, MIW91, NMLBCM^{+}01, NM17, PC87, PD15, PLHLF05, RBPGJ^{+}20, STC10, SKSK06, TZP^{+}00, VDGGD^{+}22, YSS14, YN03a]. relations [Don94, HFW^{+}98]. Relationship [EKB06, SCB^{+}09, WLKM10, YFK21, BCGN^{+}18, CPG08, DBC^{+}18, hHCK01, IL20, LSF^{+}17, LLL^{+}24, MSC^{+}15, RD03, SAM^{+}04, Sel65, SBG^{+}08, XYGJ23, ZHBW01]. Relationships [MM90, PCC^{+}19, RFFL21, DTOD00, EHG^{+}12, JJA^{+}08, LOG^{+}09, LOBG^{+}10, MHTG10, MDAW^{+}19, PPHM18, PGGG17, SPC^{+}23, Wai21, WZFW16, ZL01]. Relative [EBR^{+}14, ADV^{+}18, DCM16, JF13, KSE^{+}09, MLB^{+}20, MWFH02, NCC^{+}15]. relaxation [DRVMC^{+}22, TFM03]. Release [OB98, IOGS13, JOGM^{+}10, TNS^{+}05]. released [MK86]. relevance [Dav99, SCPN15]. relevant [FACM^{+}23, FEGA^{+}14, MCMT^{+}17]. Reliability [Kvi69]. Remarks [Koc65a, OE65, SCLS10, Car97b, GM19, Kam19, Ola65b]. remediate [GGT^{+}15]. remineralisation [BHHR15]. remineralization [FCMCÁS19, SLOP^{+}22]. remineralized [PRL^{+}18]. remobilization [PPSVC^{+}13]. RemoTe [JST^{+}24, GEP^{+}08, SMGL01, ARD^{+}03, BBE^{+}15, BPGD^{+}14, BGR^{+}15, BTJ^{+}17, KY14, KPSB17, KC02, McK15, RHB23, iSIS02, Tho87, WMB^{+}18, ZL24, dlGFM^{+}23]. remote-sensing [McK15]. Remotely [UKM^{+}14, CTMV^{+}14, HMO^{+}13, XLX^{+}20]. remotely-sensed [CTMV^{+}14]. removal [SGL^{+}18]. renewal [APC^{+}21, LEDR^{+}22, Tit20]. Reorganization [KLC^{+}15, CHB02, RMC^{+}15]. reorganizations [BDT^{+}08]. repeat [CNBD21, MDR22]. replanting [AR18]. replete [AIHB^{+}07]. report [Ano94k, KMOM88]. representation [FW91, MST^{+}23b]. representative [AM19, SS69]. representativeness [Web69]. representativity [Kvi69]. representing [GLH13]. Reprint [DAF^{+}22b, PBB^{+}12b]. reproducibility [ŠGM^{+}18]. Reproduction [HLPL05, IMHL07, vPRT90, BFB^{+}20, DWFP^{+}19, HGBG20, PMH17, PCH08a, SIB^{+}06, WPB^{+}08]. Reproductive [FARRL^{+}13, PHLL05, BWMGCB08, BCL^{+}09, CÁM06, HWBT03, LS12, Nie07, SS03, SHT^{+}01, YFY^{+}22]. requirements [AHGRAL23, BMG13]. Research [FDH20, HSH^{+}19, MP13, MS15, ACHSH08, BPW10, CL03, EAL^{+}07, MRH^{+}18, MDGC^{+}12, Moh15, PAB^{+}21, PHKS17, Val99a, Wüs64, ALV^{+}21, Ang88]. reserve [ROBRB^{+}22]. reserves [CNT^{+}19]. reservoir [SLBR18]. reshape [ZCA21]. residence [CSMGS19, WST^{+}21, WJE^{+}92]. residual [PL89, Yux88]. residuals [SL13]. resilience [CARBML^{+}22, HSMLDC^{+}22, MDGC^{+}12, ROBRB^{+}22]. resilient [MCB^{+}10]. Resistance [JP90, CARBML^{+}22]. resistant [PCD^{+}18]. Resolution [SPC^{+}23, APC13, APN^{+}15, ASJ^{+}23, AASJ23, BMC05, CN22, CGZ^{+}16, Ché14, CNBD21, DRE^{+}08, DPF^{+}20, FACM^{+}23, FAAV^{+}15, GWS^{+}23, KM08, MFDH22, MRW^{+}14, RMK^{+}21, SKWWGV18, SMP^{+}22a, SCC14, SVIA14, dPCS23, vHMDL14]. resolved [PO15]. Resolving [CMS^{+}13, SOB^{+}08]. resonant [SMFM^{+}21]. resource [IS19, LFG10, LIH^{+}12, OMR^{+}22, VMC^{+}19]. resources [CWS^{+}21, GMR^{+}23, PHKS01, Ric01, SKT01, TSH^{+}17, ZLKO00, ZLS^{+}04, SAB^{+}22]. respiration [AE09, GMAGH^{+}17, GMDD^{+}22b, GMDD^{+}22a, HLCdP19, PG13, YHLA^{+}04, ZKK^{+}16]. respiratory [AGL^{+}15]. respond [LG23]. Responding [JBH20]. Response [BBB^{+}14, CPC^{+}15, HVRR15, LFI^{+}13, ANMP15, ABD^{+}17, AC85, BVJE19, BLR^{+}23a, BSW86, BDE03, CGC^{+}20, CMHM18, CAB^{+}99, DDDT99, DMD^{+}00, ESA^{+}13, GCCY^{+}14, GFGGD^{+}23, GEP^{+}08, iIRM^{+}15, JHM^{+}22, KJZ^{+}12, LJPGC02, Let87, LO21, LH08, MMES16, MOSN^{+}13, MH14, MSB^{+}23, MPC12, Peñ03a, PTZ^{+}23, PSM^{+}22, RCC^{+}18, RLX^{+}24, RKM^{+}07, RN02, SDP^{+}22, SFMT12, Sch83, SS17, VWDF14, VMV^{+}23, WWW^{+}23, WHI^{+}02, WDK^{+}01]. Responses [FB05, MST^{+}23a, QLW10, RK03a, SSH^{+}05, SHS^{+}05, TKK^{+}05, gWjNfLyD20, YAK13, AJA^{+}22, AFBT^{+}22, BTS22, BAP^{+}22, CLMR23, CNBD21, DFC^{+}21, DCL^{+}13a, EHFD12, FMM^{+}20, FDB^{+}21, Fro05, FYYC05, GLY23, GPC^{+}03, HMX^{+}23, HGB^{+}21, HWF^{+}21, JSA^{+}08, JYK^{+}14, LG22, MCB^{+}10, MBD^{+}09, PCR^{+}22, dSSDS^{+}20, SSW^{+}09, TAW^{+}15, TSNO05, VSGD21, YAK^{+}08, YYK^{+}12, dMGS^{+}11b, dMGS^{+}11a]. result [KCPM09]. resulting [NP00, VTGC19]. Results [AUE^{+}14, BBM^{+}14, CGZ^{+}16, FRK^{+}09, HHP06, LFCSV^{+}13, LGG18, MGF^{+}13, RDL^{+}91, RF17, SWP^{+}13b, ZHF^{+}24, ABM^{+}05, ÁBMÁS14, ÁBMÁS15, CCM^{+}13, CMG15, GAM98a, GA00, HNL14, JLS^{+}22, Kag97, KNS^{+}03, KY14, MR03, WO85, NCH^{+}07, ÖHÜ89, SGLF^{+}13]. resuspension [KYT^{+}16, VSC01]. retention [BEP02, CKL^{+}14, MMIB10, MCMT^{+}17, RKCH15, RWOA01, YCP^{+}12]. retentive [SMR^{+}20]. Rethinking [CLSP17]. retreat [MOSN^{+}13, WC15]. retrieving [WMB^{+}18]. Retrospective [CHB02, CAT^{+}08, PZA^{+}15]. retroversa [MLB^{+}20]. returning [CCS^{+}21]. reveal [BD18, CNBD21, DOS^{+}18, DSB^{+}24, GIC20, HLP^{+}16, HCV^{+}20, LAGM^{+}23, OMK^{+}22, SGF^{+}19, SDO^{+}14, TSAM^{+}22, VGJ^{+}19]. revealed [BLP^{+}20, BDL08, BMC^{+}10, FFT^{+}18, HLK13, HFO^{+}22, HHZ^{+}22, JM19, LSM^{+}22, MMF^{+}17, MZZ^{+}23, OYKK^{+}23, PHD^{+}18, SHC^{+}06, SHC^{+}07, SKCP23, TM13, TMH^{+}16, VOT^{+}99, WST^{+}21, XLX^{+}20, YHZ^{+}22, YAI^{+}14]. reveals [AGD^{+}18, KSB^{+}22, MFDH22, MDR20, MJD^{+}21, PMM^{+}23, SRT^{+}18, VMB^{+}22a]. reversal [Emi65, gWjNfLyD20]. Review [AMFY20, Ang79a, Ang80, Ang88, Bak83, BHK^{+}19, Dri11, Hof81, LCANAS^{+}07, MR03, SW81, UBB^{+}23, AALM06, Arb22, ACHSH08, BPF06, BRD^{+}15, BPGD^{+}14, CTL^{+}04, Cow05, DPB06, DMF^{+}09, FÁFL06, FT06, FL06, FK99, GCV^{+}24, Gri22, GC14, Igu04, Kes06, KDB95, KPSB17, Kun03, LR07, LRNK99, LFA^{+}06, LCB18, LNB13, LB14, Man04, MPMA13, McK15, MNM06, MHH^{+}15, MB07, MMN^{+}24, NXY15, POS^{+}07, PMK^{+}06, RBF^{+}09, RMHL09, SJA^{+}23, Sim81, SPH^{+}15a, Sol00, TDH^{+}95, VH09a, VH09b, VNH^{+}23, WF06, WF07, WBA^{+}22, WLL06]. Reviewer [Ano08y]. Revised [MC88]. revisited [BMM01, Fei93, HLP^{+}16, HNL14]. Revisiting [CSC^{+}12, EB08, ORCH^{+}19, YKH^{+}24]. Revisits [KMWF11]. reworking [SvWRvB02]. Rhône [CFC^{+}18, MBdM^{+}18]. Rhone [PAG^{+}18, PRL^{+}18, SCCJ^{+}18]. Ria [BLT^{+}15]. rias [VSPP14]. Rica [SDS^{+}22b]. Rice [Bil01]. rich [IG19, SRT^{+}18, TRLA^{+}13]. richness [BJMP19, BJMP20, CC23, JM19, JPBB20, MB20]. Ricker [McK08]. Ridge [RCB^{+}20, LvBS^{+}24, CSR90, CSM^{+}15, KAK^{+}22b, MPM^{+}18, SF85, VMB^{+}22a]. ridley [CdTH^{+}16]. Right [CPO^{+}19, BMG^{+}21a, GC09, LBC^{+}23]. ring [Ban96]. ringed [CQC15]. ringens [BDL08, CGC^{+}20, EB08, FVLC^{+}23, GRB^{+}08, PCSMC12, PVA24, SAY^{+}16, YPGE^{+}10]. rings [Nof96, PM22]. Rio [FBT^{+}22, SNdSR^{+}24]. RiOMar [XLX^{+}20]. Rise [BDC^{+}08, FBT^{+}22, Har05a, BHE^{+}98, EFC^{+}23, HHWW20, KY23, Let87, MVS08, NF87, Tho87, HKE^{+}10, SNdSR^{+}24]. risk [KFC^{+}13, MPMFL^{+}23, MCL^{+}15]. Risso [SCB^{+}09]. River [BJ90, Ham90, JS90, JGS90, JSHB90, LSV14, LZG20, LDMH09, SJH^{+}90, SC90, SSM90a, SMM^{+}90, WL16, ZLR^{+}07, ZDG^{+}21, AJHC19, CPPPEAG22, HVRR15, HW02, HLS^{+}14b, IVT^{+}12, LCJ^{+}17, Pra04, SLM^{+}16, WF17, XLX^{+}20, EKB06, FMWW14, GdRGC^{+}14, Ken88, LYZ16, McK08, MI21, MSd^{+}16, PAG^{+}18, PRL^{+}18, SMN^{+}14, SCCJ^{+}18, SSM^{+}90b, SEO13, YAI^{+}14]. River-dominated [ZDG^{+}21, XLX^{+}20]. river-influenced [IVT^{+}12]. Riverine [OAD22, SVHM^{+}13, BFJ18, SLBR18, CWW15]. rivers [PBB^{+}20]. Rn [Gam14]. Robert [WR03]. Roberts [SEO13]. robusta [vPRT90]. rochei [KTIT22]. rock [LDAM^{+}07, WOW^{+}14]. Rockall [DDCE^{+}23, GD85]. rockwalls [CHG^{+}18]. rocky [BWB^{+}09, Con87, FSAO22, WLP^{+}21]. ROFI [MBdM^{+}18]. Roland [JW01b, vWM02b]. Role [CPG^{+}18, FK86, Iwa23, JJL^{+}19, LZL^{+}22, SRAV19, AFBT^{+}22, ABC^{+}99, AHW^{+}15, Ber65a, CRT^{+}22, CNSHT15, DP18, DTKvH15, FRCH15, GCG^{+}14, GCD^{+}99, HZCZ16, HGH^{+}19, Hut95, IAN13, KYT^{+}16, KSC10, LDHW20, Mar03, MTC14, NJCD01, OOTA15, OKdA^{+}19, PHK^{+}17, RBE^{+}12, SGF^{+}19, SBD01, SBBV04, SDS^{+}22b, TCN20, TSAM^{+}22, TR99, TWAL^{+}11, TSH^{+}17, TCF^{+}18, TPTM23, TCL^{+}15, Val99b, VMN08, VMC^{+}19, WSS15, XDG^{+}23, YS15, Yos80]. Roles [MY23, NNM^{+}21, EBM^{+}20, KST03, NCC^{+}15, Sie88]. ROMS [MAB^{+}11c, MAB^{+}11a, MAB^{+}11b, NDEG22, ZCH^{+}17]. ROMS-CoSiNE [ZCH^{+}17]. ropes [IST^{+}88]. rose [SCB^{+}09]. Ross [PHC^{+}19, SK18, SAT^{+}22, ZPC^{+}16]. Rossby [ABS^{+}20, CSLJ03, Ham09, Tho77]. rotating [FHL^{+}24, Sak86]. rotation [CR97]. round [DHD^{+}23, LFCSV^{+}13, LPARF^{+}20, MGF^{+}13, Woo18]. route [AG22, TSL10]. rule [JPM^{+}08]. rule-based [JPM^{+}08]. Run [KAH^{+}16, HMX^{+}23]. runoff [HVRR15, SGO^{+}08]. Russian [Ang79a, EBD^{+}20, Mar20, OT19]. Ryukyu [KKKS14, TMH^{+}16].

S [Ang80, Dea85, PSP^{+}21, APHGC^{+}22, BHC^{+}18, BM07, CPNL07, CFG07, CF12, EM12, FC07, GMAGH^{+}17, GBC^{+}00, GMAB07, GCD^{+}13, HGT16, HGTP^{+}19, HYM^{+}12, JTD^{+}14, MBKS08, MSL^{+}07, MDL^{+}12, MSV^{+}14, MCGS^{+}16, PMA^{+}14, PTPY^{+}23, QOS^{+}22, RBL^{+}19, RKFD07, SLG^{+}12, SJM^{+}19, TSRF14, VSGC21, ZHSMM14, dSPF^{+}23]. S-50^{o} [RKFD07]. S-54^{o} [QOS^{+}22]. S. [KLP^{+}17]. s.s [GBG05]. S.W. [JGO^{+}98]. S0079 [SDS22a]. S0079-6611 [SDS22a]. Saanich [TSC03]. sablefish [KMB01]. sac [SMPC^{+}12]. SADCP [CGD^{+}22]. Sado [JGO^{+}98]. safe [SGL^{+}18]. saffron [VMH^{+}21]. Sagami [KFKO03, KKS^{+}03, KNS^{+}03, Kit03, MLD^{+}03, NMK^{+}03, NKK03, NMN08, SMN^{+}13, SS03, Soh03, TMN^{+}12]. sagax [DBR03, EBvdL^{+}09, FELMGM^{+}22, NMLBCM^{+}01, ORPRGIS22, PVA24, YPGE^{+}10]. Sagitta [FB05]. Saharan [LGG18]. saida [BF11, KSG^{+}17, VMH^{+}21]. Sailfish [RLL^{+}09]. sailing [SSB14]. Saint [BMG^{+}21b]. Saint-Pierre-et-Miquelon [BMG^{+}21b]. saira [OOTA15, WCS^{+}23, XYL^{+}22]. saison [Ber65b]. Sakhalin [Nag01]. saline [NBHM01]. salinities [NPO^{+}19, PVG^{+}20]. Salinity [ANH21, BLAM98, BLAM00, DMML88, Leg91, USH15a, AR18, ATS01, APP21, Cox63, CGB07, Don94, ED82, FM07, Ham90, HHSR07, KKNT23, KC15, KSK21, Lav09, LSXT01, Mid69, MKS^{+}22, MJWK07, NRA17, RCD^{+}94, RKFD07, RG09, SDGVE17, SMKK21, SD07, Tom81b, UKM^{+}14, VYGMM^{+}17, Yao88, Yu23]. salinity-depth [ED82]. Salish [DTG^{+}24, NHSP23, NHE^{+}13, RNL^{+}13, SOA^{+}23, WHK23]. Salmon [DIQJ21, AHC^{+}13, BM01, BL02, CCS^{+}21, FMM^{+}20, IOGS13, IHY^{+}01, IAFD02, KBF^{+}08, Kli10, McK08, MI21, SF02, SKSK06, Wai21, YWUK15]. salmonids [EKB06]. Salpa [GBH^{+}20, LS12]. salps [IMM^{+}22]. salt [EMK^{+}17, Kun03, Sch03, SAH^{+}21, YN03b]. salt-finger [YN03b]. salt-fingering [Kun03]. salty [War06]. sampler [SPH^{+}15b]. samplers [Hop64]. samples [LB20, Soh03, TCL20]. Sampling [WH94, BCF^{+}03, JJS03, JPBB20, LPA^{+}11, MGF^{+}13, MSA^{+}22, Mol04, Reb02, RAB^{+}84, SPC^{+}23, SFAD^{+}90, SWP^{+}13b, SVIA14, TIOM16, VLUC^{+}07, WZBK^{+}21, ZNR^{+}24]. sampling-gear [Reb02]. San [Hic92, WJE^{+}92, Gor92]. sand [LTSG13, RHBS13, Sei63]. sand-bank [LTSG13]. sand-transport [Sei63]. sandbank [PC87]. Sandy [LC22]. Santa [AHW99, CJ92, Gor92, Hic92, LPA92, SPB^{+}02, SC23, SE92, VK92, WRS^{+}92, WJE^{+}92]. santolla [BMN19]. SAR [VOT^{+}99]. Sardina [BCGN^{+}18, CCHV^{+}21, SNV^{+}18]. Sardine [BSC^{+}07, SNV^{+}18, BCT^{+}09, BCGN^{+}18, CCHV^{+}21, CCM^{+}14, DBR03, DPGC14, EBvdL^{+}09, FELMGM^{+}22, FRCH15, GSM^{+}17, GCD^{+}18, HMRB^{+}03, HLS^{+}14a, KYS^{+}17, LLS01, MRBS^{+}24, MPB^{+}23, NMLBCM^{+}01, ORPRGIS22, PVA24, Qiu15, RFC^{+}15, SKH^{+}23, SGWF^{+}19, SYB^{+}15, SBG^{+}08, TOKLC08, VOG^{+}08, VDB^{+}20, YPGE^{+}10]. sardinella [LPARF^{+}20, DPGC14, SNR^{+}10, TCF^{+}18]. sardines [APC^{+}12, MB01]. Sardinops [DBR03, EBvdL^{+}09, FELMGM^{+}22, NMLBCM^{+}01, ORPRGIS22, PVA24, YPGE^{+}10]. Sargasso [PPCWJ18]. Sargassum [JLP^{+}20a, JLP^{+}20b, PGG^{+}22]. Sarmatian [Gal17]. sars [GS19]. satelite [iSIS02]. Satellite [HSS^{+}12, SLM^{+}16, SW12, AK97, CED09, CSS^{+}19, FGR^{+}06, GdRGL^{+}01, GA00, HSMLDC^{+}22, HMRA^{+}03, HMRB^{+}03, HZD^{+}23, HMX^{+}23, HHMB^{+}09, KBSB18, LC16, LW13, LLX^{+}21, MBdM^{+}18, MVC^{+}11, OMS^{+}09, OP18, OÁT^{+}05, PLK14, RRS03, RBS^{+}09, SLH^{+}19, SKCP23, SI97, TBS^{+}19, TM13, Tho87, VNMS91, WCX^{+}21, WLM07, WFS^{+}15, WZC20, XYWY23, ZD17]. Satellite-based [HSS^{+}12, HMX^{+}23]. satellite-derived [HMRB^{+}03, LLX^{+}21, PLK14, SLH^{+}19, ZD17]. Satellite-measured [SLM^{+}16]. satellite-observed [HZD^{+}23]. satellite-tracked [LC16]. satellites [HYM^{+}24, KY14]. saury [OOTA15, WCS^{+}23, XYL^{+}22]. SBE [KSK21]. scale [AGD^{+}18, ALM^{+}23, ASB^{+}08, BMK12, BGMP03, BRR^{+}12, CNT^{+}19, CMM^{+}04, CTF07, CGC^{+}20, CSS^{+}19, CS03, CSC^{+}12, CCS^{+}21, CNBD21, DRVMC^{+}22, DIM09, DIQJ21, DTKvH15, FBB^{+}21, GRLS14, GBC^{+}00, GGAA^{+}23, GBC^{+}16, GFB^{+}15b, GFB^{+}15a, GPC^{+}03, GAS^{+}22, HVTV22, HGBG20, HWB^{+}18, IHT^{+}21, IPD14, IMM^{+}22, JLP^{+}20a, JLP^{+}20b, JHW^{+}14, KLB^{+}21, Kaz17, Kli10, LCB18, LMH^{+}13, LSB^{+}17, LSD^{+}18, LBP15, LL21, MCB^{+}90, MBH^{+}23, MSA^{+}22, MHCS^{+}23, MZGA^{+}20, MJA^{+}07, MCH^{+}12, NGLSSG14, NBLI20, OPL^{+}21, PZA^{+}15, RFFL21, RM89, RÁSG^{+}13, RPSC22, SGWF^{+}19, STEB16, SWP^{+}13a, SPC^{+}23, SSL08, SKSK06, SH09, SLPA^{+}20, SDJ14, SJ02c, SJ02b, SMP07, TG05, UB10, VOG^{+}08, VMB^{+}22b, WCX^{+}21, Whi95, YSY^{+}19, YBPS08]. scales [BRR^{+}22, BMO12, BMGN15, BGB^{+}08, BBL^{+}18, CIL^{+}23, CMF11, EM12, FSVL10, GKC^{+}14, GCED22, Kra82, MCD^{+}07, MPSS91, OÁT^{+}05, Ric01, SSW^{+}09, Yos80]. Scaling [WZFW16, KHS^{+}14, KSKN21]. scallop [CZG^{+}21, CZL^{+}24, GGJ^{+}10]. scanning [PKA19]. scarce [BMN19]. scattering [GIC20, PMFNGQ21, PVB23, Peñ24, PO15, PNF^{+}21, SLGI^{+}21, SK91]. scavenging [CS89, GPA^{+}11, HS22, HTV^{+}20]. scenario [KKKY10, LSS^{+}10, SPF^{+}23, SYB^{+}15, WAH^{+}20]. Scenarios [BDT^{+}08, BMG^{+}21a, JBB^{+}14, KHL12, NPO^{+}19, SAd^{+}17, Ste12, TMÁGC^{+}21]. scheduling [HBD^{+}18]. schematic [Ric08]. scheme [SBM91]. schemes [BMM97, YGC^{+}21]. Schoenefeldt [SDS22a]. School [Bak01]. School-mix [Bak01]. Schooling [BGB^{+}08, ANMP15, Mau17]. Science [PBH^{+}10, BCD^{+}20, CSH^{+}23, Car97b, HBD^{+}21, MCB^{+}10, RW97, SSB^{+}20b, SJP10, CSH^{+}23]. sciences [MCG^{+}14]. Scientific [ALV^{+}21, SPK^{+}19, Ric01, Val99b, Zav99]. scleractinian [ALG^{+}21]. Sclerochronological [RvBD^{+}22]. Scomber [GiIKX22, TMÁGC^{+}21]. SCOR [Ano94k]. scorpaenid [ORB^{+}18]. Scotia [SCS^{+}18, TSFA22]. Scotian [SPH^{+}15a, SZG06]. Scotland [ZLZ^{+}17]. Scott [GBG05]. Scottish [Fra69]. Screening [KHL12]. SE [GRDS10, LvIKB07, MCGS^{+}16]. Sea [AQVB^{+}10, BBM^{+}14, BF12, CJMO87, CSV^{+}07, CLD22, CÁM06, CQZ^{+}18, CLG^{+}22, CEF^{+}13, Den87, Dev87, EMU^{+}23, EHSI12, FGR^{+}06, FCN^{+}19, HBL^{+}13, HMP^{+}13, JYK^{+}14, KK20, KGL22, KKK^{+}04b, KKS^{+}19, KS15, lLdZQ^{+}22, LDHW20, Men21, NHN^{+}21, Par86, PGT^{+}13, PLN^{+}23, Pre86, RCSA01, RN02, SCAA07, SIR^{+}07, SCS^{+}18, STG^{+}18, SMP^{+}22b, SNMW10, SDL^{+}19, TCN20, TPP^{+}00, WGZZ19, XLX^{+}20, XCH^{+}16, YSY^{+}19, YHM^{+}18, Yin88, ZLR^{+}07, ZLC^{+}15, ACB^{+}13, AS20, AP20, ATS01, AS96, AAMB^{+}24, Ano94k, AEPW93, BSC^{+}19, Ban64, BRC^{+}18, BTNK13, BCOL^{+}19, BS90, BBB^{+}21, Bil01, Bla63, BHC^{+}18, BRG^{+}23, BHPC06, BAB^{+}19, BBRM20, Bre06, BGS^{+}04, CDS90, CWZ^{+}20, CKP^{+}20, CDB^{+}22, Car98, CGM^{+}02, CMM^{+}04, CPG08, CMF11, CF20, CJRÁ^{+}13, CdTH^{+}16, CAO^{+}20, CMHM18, CGZ^{+}16, CZG^{+}21, CP19, CPSM20, Con87, CFML22, CFG07, CF12, DWH^{+}14]. sea [DTOD00, DWFP^{+}19, Dav85, DBW^{+}22, DOS^{+}18, Dem09, DSB^{+}24, DLD^{+}19, DPCS87, DLD15, DSV^{+}24, DBJ^{+}15, ECGP01, EAB^{+}23, ESA^{+}13, Eri65, FBD18, FARRL^{+}13, FPJ^{+}15, FACM^{+}23, Fly10, FP15, FMCG15, FLUC08, FJH10, FJ19, FWL^{+}15, GMDD^{+}22a, GDL^{+}15, GGJ^{+}10, GBM^{+}01, GDN^{+}18, GMR^{+}23, GM19, GvOSW11, GBB96, GD85, GF19, GVKD^{+}13, GW89, GLLB22, HM90, HKN^{+}14, HMRB^{+}03, Har82, HS22, HJLLN07, HLCdP19, Her97, HT97, HHWW20, HPHW21, Hol00, hHCK01, HCV^{+}20, HWBT03, HAH^{+}22, IAN13, IHT^{+}21, Iwa23, JM19, JP90, JKBH87, JSLA^{+}21, KTN14, KON14, KKNT23, KDL^{+}01, Koc65a, KFC^{+}13, KMS^{+}24, KSG^{+}17, KGB^{+}23, KG65, Kos93, LVGH^{+}15, LSF^{+}17, Let87, LSH^{+}11, LGZ^{+}20, LC22, LB20, LFBP^{+}13, LRJ^{+}15, MSC^{+}15, MBP^{+}11, MGS90, MKD90, MTC12, MGA^{+}23, MEMP15, MST^{+}23a, MHR^{+}10, McK04]. sea [MM80, MOSN^{+}13, Mit91, MMN12, MFM85, MM90, MKS^{+}22, MSFZ19, MJD^{+}21, MNFY21, MVS08, Mun69, NNFL21, NO14, NH88a, NF87, NAH^{+}21, NST^{+}23, ONR^{+}14, Ola65b, OE65, OAB^{+}16, OT19, OKdA^{+}19, PPdM^{+}12, PO00, Par63, PPHM18, PRA^{+}18, Pir87, PDD^{+}22, RGC^{+}01, RCC^{+}18, RD11, RCB^{+}20, RBR^{+}23, RSD^{+}90, RCSHW22, SSB20a, SMFM^{+}21, SDGVE17, SW92, ŠVL^{+}15, ŠPM^{+}22, Sha82, SAM^{+}10, SENS13, SSI13, SC65, She65, SKF20, SFAD^{+}90, SOO^{+}14, SWP^{+}13b, SEW11, SMGL01, Soh03, Sok90, SM16, SPN98, SPV^{+}15, SRG^{+}19, STR01, SCS87, TRY^{+}04, TCS15, Tho87, Tho95, TSRF14, TSFA22, TTF^{+}22, Tur65, TSP^{+}13, VKGP^{+}13, VBJ^{+}20, VH09a, VH09b, VTGC19, Wal83, WO15, WLD^{+}15, WGM^{+}24, WBA^{+}22, Web69, WHBW03, WC15, WPW^{+}14, WWL^{+}24, WST^{+}16, Wüs64, XD95, XY20, YIY^{+}04, YNTS22, YHRT22, YLL19, YHZ^{+}22, YPM^{+}10]. sea [YN20, YFY^{+}22, YTL^{+}19, YYK^{+}12, YGMR^{+}23, ZLG17a, ZLG17b, Zez90, ZPY^{+}20, ZKK^{+}16, dlPHF^{+}15, vHCY^{+}20, vPRT90, AEP^{+}23, ANH21, AIHB^{+}07, ACB^{+}13, AHP19, AJA^{+}22, ACE^{+}07, ACL^{+}18, AAMB^{+}24, ABS^{+}20, ALG^{+}21, AOMZ^{+}23, AVS23, ADS^{+}22, AVG^{+}19, AT07, APC^{+}21, BMK12, BSC^{+}19, BPA^{+}21, BTK^{+}99, BNC05, BE99, BRC^{+}18, BAARB05, BPNB90, Bea04, BBB^{+}14, BMB^{+}16, BSF^{+}21, BSÖ^{+}94, BGM^{+}99, BD19, BTG^{+}03, BDG^{+}17, BFJ18, BBL^{+}18, BRG^{+}23, BPSGP^{+}23, BHLU^{+}07, BD18, BDC^{+}08, BPM^{+}14, BLC23, BBF^{+}22, CGL^{+}20, CDL19, CPB^{+}15, CIL^{+}23, CWZ^{+}20, CDDF11, CGV13a, CGV13b, CGM^{+}02, CFM^{+}18, CMJPH^{+}18, CMPNC^{+}22, CBB^{+}22b, CC88, CTL^{+}04, CSW96, CLX^{+}20, CCHV^{+}21, CS03, Cho86, CBM^{+}21, CLG^{+}00, CLA^{+}00, CS18, Cia22, CS16, CD65, CHSB^{+}21, CFML22, CDP14, CHB02, CLL^{+}18, CRC^{+}19, CHS^{+}24, Cow05, CP02, CGD^{+}22, CJG88]. Sea [DM13, DAvD^{+}20, DSV99, DDE^{+}95, DDDT99, DDP^{+}00, DMD^{+}00, DTOD00, DDD^{+}00, DGP^{+}13, DBC^{+}23, DLD^{+}19, DSC^{+}19, DAU22, DVB^{+}18, DZ04, Dol09, DFC^{+}21, Dri11, DCL^{+}13a, DFM^{+}15, DL17, DGH^{+}20, DHL^{+}21, DTG^{+}24, DCS^{+}22, DTW^{+}00, DWC06, DBM17, EMBS13, ERT^{+}22, EBW^{+}23, EFC^{+}23, ESGP17, EGP^{+}18, EBD^{+}20, ESD^{+}21, Eri65, FVA^{+}19, FZY^{+}23, FTC^{+}16, FHP83, FDHT05, FvBA^{+}17, FTG^{+}18, FPY^{+}16, FJ19, FMWW14, GCLD19, Gam14, GHF^{+}21, GCCY^{+}14, GMDD^{+}22b, GSPP^{+}20, GGT^{+}15, GR85, GTB07, GDL^{+}15, GPP22, GCZ^{+}00, GCB^{+}22, GLF^{+}17, GWM^{+}22, GCF^{+}19, GS24, GMR^{+}23, GSPMAI99, GBB^{+}20, GFB^{+}15b, GFB^{+}15a, GRS08, GGA^{+}16, GIPG17, GGG^{+}18, GGSM^{+}20, GTS^{+}21, GCD^{+}99, GC14, GLS08, GLLB22, HMTL05, HSG^{+}15, HBK^{+}24, HZD^{+}23, HHY03, HMP^{+}13, HRA^{+}08, Hea12, HLM^{+}13, HvDL^{+}17, HKGH^{+}06, HOY^{+}21b, hHCK01, HPZC21, HHB^{+}00, HTdM^{+}15, HSH97, HLSX22]. Sea [HHZ^{+}22, HYM^{+}24, HPNDC15, HWF^{+}21, HS02, HMA18, ITM86, Igu04, ISM^{+}02, IAFD02, IMW^{+}14, INT14, Iwa23, JJL^{+}19, JZZY24, JFUR20, JMZ23, JMG^{+}13, JCF^{+}23, JMSB^{+}23, JVJ^{+}17, KKK04a, KTN14, KON14, KWI20, KCL^{+}12, KY15, KOT^{+}21, KTIT22, KFC^{+}23, KMS^{+}24, KSB^{+}22, KGB^{+}23, KCBS20, KTB^{+}99, KDB95, KBSB18, KAG^{+}19, KHM^{+}88, KAAK^{+}16, KZD^{+}19, LJM^{+}16, LT06, LDAM^{+}07, LTSG13, LRNK99, LSM^{+}22, LML^{+}23, Li14, LYZ16, LPS^{+}19, LGZ^{+}20, LBH^{+}21, LZL^{+}22, LYS^{+}22, LRS^{+}03, LNB13, LXC^{+}22, LSXT01, LLH^{+}20, LPF^{+}21, LPW^{+}23, LPBM17, LGH^{+}21, LBC^{+}15, LFBP^{+}13, LWY07, LGD^{+}20, LSW02, LDMH09, LBD11, LO85, MXC^{+}21, MLL^{+}22, MY23, MMG^{+}13, MRM^{+}14, MHGGS19, MRSS02, MGG22, MRMD^{+}97, MGKW19, MMIB10, MGA^{+}23, MB05, MZF^{+}08, MPCNC^{+}19, MEST13, MEMP15, MPD^{+}22, MCD^{+}07, MKHO96, MGC^{+}18, MVC^{+}11, Men21, MBB^{+}96, MCGR07, MWJ^{+}08, MZZ^{+}23, MCKS17, MOSN^{+}13]. Sea [MSGGM18, MFY^{+}86, Mil09, Mil14, Min02, MMG^{+}11, MAFS^{+}22, MSS^{+}02, MKM86, MR03, MH14, MKS^{+}22, MWFH02, Mos69, MEMC05, MLPN06, MKMF^{+}89, MMN^{+}24, MJWK07, NRS^{+}19, NNFL21, NDEG22, NTU^{+}14, NO14, NNM^{+}21, NH88a, NXY15, NBG^{+}05, NGLSSG14, NCC^{+}15, NHSP23, NHE^{+}13, NSE^{+}24, NF06, NMY^{+}14, NNO^{+}14, Nof00, NRA17, ONR^{+}14, OELP04, OEL^{+}14, OÁT^{+}05, ORMR^{+}19, OPL^{+}21, ÖÜT93, PIS13, POS^{+}07, PAG^{+}18, PCD^{+}18, PJS^{+}22, PL18, PDV12, PPSVC^{+}13, PPCWJ18, PDAM^{+}15, PSM^{+}22, PLB^{+}23, PRA^{+}18, PZA^{+}15, PM85, PHC^{+}19, PTG95, PST^{+}02, PST^{+}15, PDD^{+}22, PKV18, PL09, PLP99, PBP^{+}99, PTI00, PGG^{+}22, QLW10, QLY^{+}22, RM93, RKS01, RLDC^{+}13, RCC^{+}18, RDG^{+}21, RCS^{+}11, RBF^{+}09, RLX^{+}24, RGB^{+}17, RKCH15, RDD^{+}18, RTBR^{+}22, RI86, RDC^{+}21, RKM^{+}07, Rog00, RNL^{+}13, RDP^{+}21, SCHS^{+}24, SLBR18, iSIS02, SGL^{+}13, SCMAR^{+}99]. Sea [SVHM^{+}13, SHP^{+}23, SGA^{+}19, SHd13, SN24, SDP^{+}22, SM21, SPG^{+}06, SPB^{+}12, Seg69, SMP^{+}22a, SBH^{+}14, SMN^{+}14, SOH21, Sek86, Sek88, SCCJ^{+}18, SGPdM18, SFK^{+}99, SRAV19, SSTL16, SGS^{+}23, SW12, SNZ^{+}20, STHM02, SiSI^{+}02, SOO^{+}14, SBPGP^{+}23, SDH^{+}14, SAB^{+}21, SEG^{+}22b, SEG22a, Smi05, SM05, SK18, SAT^{+}22, SIS^{+}14, ŠGM^{+}18, SKP99, SK21, SNMW10, SCT^{+}00, SPB19, SDO^{+}14, SOA^{+}23, SGR^{+}22, Sud86, Suk88, SWH^{+}24, SÖÜ94b, SI97, TFY02, TM13, TII^{+}14, TTL^{+}04, TCDPP^{+}22, TAF^{+}22, TDGY22, TBK^{+}99, TVT^{+}23, TR99, TKW06, TKWI08, Tit20, Tol85a, Tol85b, TCF^{+}18, TZP^{+}00, TPM^{+}00, TCL^{+}15, Tur99, UNN^{+}14, VGM^{+}23, VHV^{+}12, VCB^{+}00, VKDS^{+}18, VKT15, VR03, VJJ^{+}22, VH09a, VH09b, VKJ^{+}23, VBA^{+}18, VBAC^{+}21, VHK03, VHK04, WFH^{+}22, WD94, gWjNfLyD20, WST^{+}21, WLCG23, WFD^{+}07, WYT00, WRH^{+}06, WHT86, WNNI21, WHBK05]. Sea [WHI^{+}02, WHK23, WGG^{+}08, WHS^{+}23, WL16, WZC20, XWL^{+}18, XDG^{+}23, XHC^{+}20, YYT^{+}14, YIY^{+}04, YKS^{+}12, YMI88, YNTS22, YNM^{+}02, YLL19, YGC^{+}21, Yao88, YJS86, YJ88, Yas07b, Yas07c, YAI^{+}14, YYK88, YPM^{+}10, Yux88, ZGB^{+}20, Zav99, ZMCD11, ZLS^{+}04, ZGZ19, ZZPL18, ZDG^{+}21, ZK06, ZMW^{+}23, ZDM^{+}20, ZSY^{+}22, ZZWL06, ZSW^{+}22, ZPC^{+}16, dFHR^{+}24, dPAJ07, dlGFM^{+}23]. sea-air [OKdA^{+}19]. sea-going [SWP^{+}13b]. sea-ice [BRG^{+}23, DWFP^{+}19, Hol00, KGB^{+}23, LSH^{+}11, MGA^{+}23, TRY^{+}04, WC15, YGMR^{+}23]. sea-ice-free [WO15]. Sea-level [CJMO87, Den87, Dev87, Con87, DPCS87, JKBH87, Let87, NF87, Pir87, SAM^{+}10, SCS87, Tho87]. sea-surface [KKNT23, MFM85, SRG^{+}19]. sea-to-air [ZPY^{+}20]. seabed [BMNW01, CDB^{+}24, DJW^{+}18, DHB^{+}21]. Seabight [RTN90, VBJ^{+}20]. Seabird [DWS^{+}24, AH10, FDB^{+}21, HGBG20, KFH^{+}15, MCD^{+}14, SWP^{+}13a, SWZS^{+}21, SIB^{+}06, STGR^{+}23]. seabirds [BPF06, DWNN04, UGY^{+}22, WMB^{+}21]. seafloor [LBK^{+}01, RGPB^{+}23, TDL^{+}17, ZCA21, ZCLS20]. seafood [PG10]. Seaglider [BKD^{+}20]. seal [CQC15, FGS^{+}23, GPEV20, HBG^{+}21]. Seals [BHMS09, CdD^{+}15, DMBHG10, FJH10, LVGH^{+}15, LSF^{+}17, LAP10, MSC^{+}15, NRS^{+}19, NBLI20, PHC^{+}19, RNL^{+}13]. Seamount [CDL^{+}22, HNSP^{+}19, TDK^{+}16, GPP22]. seamounts [ATC^{+}19, HFNG00, PMG15, XWL^{+}22, Gri22]. SEAPODYM [HLS^{+}14a, LSM08]. SEAPODYM-SP [HLS^{+}14a]. search [SF85]. Seas [KPM^{+}23, LW13, PO00, BEH19, CAB^{+}18, CTA16, CBOP15, CQC15, DPB06, DJ92, DRD^{+}07, DCL^{+}13a, EAL^{+}07, FBS^{+}18, FSVL10, HBL^{+}13, HOY^{+}21a, HSC^{+}16, HHB^{+}22, IU14, KJZ^{+}12, Leg91, LBP^{+}21, LSH^{+}11, MTC12, MHVS19, NYH^{+}22, Pra97, PB94, PCBA^{+}20, SCC14, SFF^{+}24, SVIA14, SPWH21, TLH^{+}15, TLP^{+}16, WZFW16, WGZZ19, WTH12, YN20, ZKT88, AHSS22, APC^{+}12, BS95, BGV^{+}23, CM11, CBB^{+}15, CCW^{+}18, FJhT^{+}14, FMCG15, FWL^{+}15, GCFS06, HVRR15, HØ00, HØH^{+}03, HKN^{+}14, HKPV12, JJA^{+}08, KK20, KFH^{+}15, LMW^{+}12, LCJ^{+}17, LC16, LCJ^{+}07, LLX^{+}21, MSC^{+}15, MRO^{+}08, MM01, NBHM01, OMR^{+}22, OACA20, PBB^{+}12a, PBB^{+}12b, RAB^{+}11, SBK^{+}95, Sei63, SH09, SK21, SAH^{+}21, Ste91, SYN^{+}21, SBS90, WMC^{+}89, WDC^{+}11, YS15]. Seascape [FG16, PVB23]. Seascape-level [FG16]. seascapes [KHS^{+}14]. season [Ber65b, BF11, CTF07, Ken88, KMSTK23, LOO22, MHCR^{+}12, MA12, SW01, TII^{+}14, WO15]. Seasonal [ACB^{+}13, ADS^{+}22, AKH^{+}23, BMK12, BSC^{+}19, BCOL^{+}19, BMC05, BRG^{+}23, BDBJ01, BPM^{+}14, CNT^{+}19, CWZ^{+}20, CDB^{+}22, CGMP14, CÁM06, CVBG21, CLL^{+}18, CFG07, DDD^{+}00, DBW^{+}22, DZ04, DK07, ESD^{+}21, EHG^{+}07, GMDD^{+}22a, GSVB23, GWM^{+}22, GA10, GDSCU09, GSPMAI99, GLLB22, HPC^{+}20, HHB^{+}00, HTdM^{+}15, HDB13, JAS^{+}20, Joh04, KFKO03, KLB^{+}21, KC15, KFH^{+}15, KZD^{+}19, LC12, LC16, LW12, LvIKB07, Lon95, MLB^{+}20, MGA^{+}23, MCGR07, MC15, MH14, NHSP23, PV07, PCM11, PFW15, PDD^{+}22, SMR^{+}20, SCS^{+}18, SvWRvB02, SAB^{+}21, SHC^{+}06, SHC^{+}07, SPV^{+}15, SOA^{+}23, TBS^{+}19, TOiF^{+}12, TRLA^{+}13, VGM^{+}23, VAGMDRS22, VBC^{+}20, WGZZ19, XWW^{+}21, YIY^{+}04, YKS^{+}12, YGC^{+}21, YN20, YBPS08, dlPPÁB24, AC85, BAM^{+}09, BRB^{+}01, BMM01, BHLU^{+}07, CIL^{+}23, CGC^{+}20, CSS^{+}21, CMHM18, CLG^{+}00, CLMR23, DMD^{+}00, DTOD00, DOS^{+}18, DRE^{+}08, DAF^{+}22a, DAF^{+}22b, EALF08]. seasonal [EBW^{+}23, EBR^{+}14, FC07, FTSF21, FDHT05, FKZ^{+}15, FCN^{+}19, GR17, GBC^{+}16, GCED22, KSVT00, KP03, KKNT23, KMS^{+}24, KSP^{+}23, KRL^{+}22, LZL^{+}22, LW13, LO85, MT99, MCD^{+}07, MW96, PCSMC12, PJS^{+}22, PD15, PLJR22, PTI00, SKH^{+}23, SSQ19, SSVP00, Sek88, SNZ^{+}20, SPB^{+}02, SBD^{+}07, STW^{+}15, SM16, SJ02c, SJM^{+}19, SWH^{+}24, TSC03, TSH^{+}17, TSP^{+}13, UPPS^{+}21, VAEP24, Ver92, WSL20, WZBK^{+}21, WCN^{+}05, WDK^{+}01, Woo18, XWL^{+}18, XLX^{+}20, XDG^{+}23, ZHF^{+}24, NKK03]. Seasonal-to-decadal [KLB^{+}21, MCD^{+}07]. Seasonal-to-interannual [JAS^{+}20]. Seasonality [AOMZ^{+}23, AVS23, HHK^{+}22, HWBT03, JCF^{+}23, LGZW22, PFHM16, RDC^{+}21, RCSHW22, SHP^{+}23, SAd^{+}17, Ban96, BE99, GGJ^{+}10, JMSB^{+}23, MGE^{+}12, RGC^{+}01, SLBVRR^{+}22, SBE^{+}20, TB15, TMPM^{+}16a, TSS^{+}12, ZBRJ23]. seasonally [CTA16, CBB^{+}22c, HAH^{+}22, MBP^{+}11, WHS^{+}23]. seasons [SM05, ZSBL00]. Seawater [Due77, AdAK^{+}18, AYH^{+}23, Fei93, FH95, Fei03, Fei04, FM07, GTR01, GLLB22, LGG18, SBBV04, TGR05, YNTS22, ZBY^{+}22]. seaweeds [GNH19, NHG19]. SeaWiFS [HM00b, WM13, YKS^{+}12]. Sebastiano [SMB88, Sie88, Bru88]. second [LSW02]. secondary [LSH^{+}11, LQU07, MTK^{+}22, PAM^{+}88, SEW11, YHRT22]. Secrets [JPB20, BW08]. section [MLS^{+}15]. sections [Man69, MDR22, STPHM^{+}23, Sel65]. sector [AMY^{+}23, Don65, JRR^{+}24, MCL^{+}15, MST^{+}23a, NST^{+}23, SSN23, VMA^{+}24]. sectors [LO07, TAF^{+}22]. Secular [CR20]. Sediment [CS06, DGGdR02, DCL^{+}13b, EMU^{+}23, PPdM^{+}12, RMB^{+}01, Sai65, SCMAR^{+}99, AS20, BEI^{+}20, BPTT19, CDS90, CJMO87, CB17, DXH^{+}02, FC07, FMH02, GSV^{+}01, HMKF08, HHK^{+}02, JGO^{+}98, KCL^{+}12, LJ65, LCJ^{+}07, LvIKB07, MXC^{+}21, ML09, MDL^{+}12, NIF^{+}15, PGLG^{+}05, RF17, RLR^{+}18, SLBR18, SvWRvB02, Sha82, SSTL16, SFAD^{+}90, TCDPP^{+}22, TWBC^{+}13, VK90, VPW01, VOJD02b, VB14, WCB20a, Whe06, YYK88, dJSL^{+}20, vWHdS^{+}98, vWdSBdH02]. sediment-starved [TCDPP^{+}22]. sediment-water [SvWRvB02]. Sedimentary [CVHM^{+}18, KGdS^{+}08, SC90, Bol94, DJG^{+}02, Ike88, JGS90, JFG^{+}90, JWD^{+}02, KNS^{+}03, Kit03, ORR^{+}02, RWD01]. sedimentation [Ber65c, DOP87, HK65, JS87, JWD^{+}02, KKS^{+}03, KB65, PPSV^{+}18]. sedimentation-rate [PPSV^{+}18]. sedimenting [BPNB90]. Sedimentological [CLA^{+}00]. sediments [AJV^{+}02, BPP^{+}98, BGS^{+}04, CLV^{+}19, CJRÁ^{+}13, CPNL07, CFC^{+}18, CBD^{+}24, Cow05, CJ92, DMD^{+}00, DTOD00, DVL^{+}99, ECGP01, Emi65, EvdZSH02, Eri65, GMAB07, GGA^{+}05, GGE^{+}65, HVTV22, Hay65, KKS^{+}03, KPM^{+}23, Koc65a, KKS^{+}19, LYZ16, LXC^{+}22, MNFY21, MSL^{+}07, Nay65, OE65, OB98, PCD^{+}18, PPSVC^{+}13, PRL^{+}18, RSB^{+}01, RGC^{+}01, RBL^{+}19, SPF^{+}23, SGL^{+}13, SGF^{+}19, SLG^{+}12, SOH21, She65, SGO^{+}08, TPM^{+}00, Wis65, YSY^{+}19, YAI^{+}14, ZHBW01, vWHdS^{+}98]. Seeding [ST10]. SEEDS [AYK^{+}05, NKK^{+}05, SSH^{+}05, TT05, TKK^{+}05, TNS^{+}05, YFY05]. SEEDS2001 [KIS^{+}05, TSNO05]. seen [KRL08, XYGJ23]. seep [GLV12, JFG^{+}90, LZF^{+}24, MCB^{+}90, OSH^{+}96]. segregation [PNF^{+}21]. SEIBM [CMS^{+}13]. Selected [BHM^{+}15, BC01]. selection [BVJE19, BPW10, CCHV^{+}21, Sma10a, Sma10b, WSO^{+}13]. selective [GBM^{+}01, WHBW03]. selectivity [CNT^{+}19, HWBT03]. Selensky [UKK^{+}19]. self [CDDF11, RRS03]. self-organizing [RRS03]. self-oscillations [CDDF11]. selliformis [VMV^{+}23]. semelparous [BSF^{+}21]. semi [BLT^{+}15]. semi-enclosed [BLT^{+}15]. Semidiurnal [RPSVLS14, WGM^{+}24, JFUR20, ZQWP23]. Senegalese [TGJT09]. Sense [CSH^{+}23]. sensed [CTMV^{+}14, HMO^{+}13, XLX^{+}20]. Sensing [KPSB17, ARD^{+}03, BBE^{+}15, BPGD^{+}14, BTJ^{+}17, JST^{+}24, KY14, KC02, McK15, iSIS02, SMGL01, Tho87, WMB^{+}18, ZL24, dlGFM^{+}23]. sensitive [Tur99]. Sensitivities [LK13]. Sensitivity [HWPLvW20, KM08, MMG^{+}13, OAB^{+}16, SCHD23, BRG^{+}15, BFPS06, BHHR15, HPS^{+}01, MAB^{+}11b, SBMB18]. sensor [BRR^{+}12, MVC^{+}11, iSIS02]. sensors [FRV^{+}19, KSK21, KHM^{+}88, MSMH19]. sentinel [DLJ^{+}21, HGBG20]. Separating [ZL24]. separation [GWGR^{+}19, LMM03, ORCH^{+}19, PMH17, SRF^{+}19, TG05]. September [Ano65i, Ano13g, Ano20b, GCZ^{+}00, Ano07p, Ano08t, Ano08v, Ano09k, Ano09l, Ano10m, Ano12k, Ano13o, Ano14l, Ano15q, Ano16r, Ano17p, Ano19p, Ano21t, Ano22-34, Ano23t]. sequential [DTC^{+}06, ORVES17]. Sergei [Ang79a]. series [AT07, BSFM^{+}12, BAOM^{+}12, Coo69, CNBD21, EHFD12, FTSF21, HFS^{+}20, HLM^{+}13, HFO^{+}22, LQU07, MGE^{+}12, MGF^{+}13, MAH^{+}15, MFDH22, MVC^{+}11, MMG^{+}11, MA12, NCH^{+}07, NIF^{+}15, ORVES17, PB07, SS69, UPPS^{+}21, VLUC^{+}07, VBL^{+}21, VDGGD^{+}22, WSO^{+}13, Whe06, XCH^{+}16, dPAJ07]. Serrano [LSV14]. set [DHL^{+}21, KMOM88]. Seto [YFY^{+}22]. setting [DVL^{+}99, Gor92, WC15]. Settling [TSC03, KHC^{+}99, KSR^{+}01, SCT^{+}00]. Setúbal [KGdS^{+}08]. seven [BHLU^{+}07, YNMY23]. seventies [Leg91]. several [aHFS92]. Severe [MAFS^{+}22]. Sex [SS03, PWZ^{+}16]. shadow [SMR^{+}20]. Shallow [PMG15, Tal08, ATC^{+}19, DP18, GBB96, LAGM^{+}23, MGG22, MSV^{+}14, SCHBC^{+}22, SDS02, SDS22a, SPN98, WLP^{+}21, XD95]. Shallow-water [PMG15, GBB96, LAGM^{+}23, WLP^{+}21]. shape [JBB^{+}14, RSW^{+}23, RSMIS03, RAG^{+}19, TKC^{+}22, ZPC^{+}16]. shaped [LdSH^{+}15, PMA^{+}14]. shapes [BMG^{+}19]. Shaping [PVB23, GBT^{+}19, JJA^{+}17]. shark [HNSP^{+}19, ON22]. sharks [Jac10]. shear [AR18, OC06, SS17, Zen08]. sheared [CSLJ03]. shearing [IST^{+}88]. Shearwater [SSB14]. shearwaters [YSS14]. shed [Nof96]. sheets [Ber65a]. Shelf [GMDD^{+}22b, JOBT05, WAH^{+}20, XD96, ANH21, ÁSDB^{+}01, AJV^{+}02, BHA^{+}14, BHAJ12, BAOC^{+}07, BASS^{+}20, BCP09, Bum73, CKP^{+}20, CDB^{+}22, CMC^{+}16, CTA16, CZG^{+}21, CZL^{+}24, CBOP15, CSG^{+}15, CM18a, CS06, CDP14, CFG07, CP02, CCH^{+}12, DM13, DWH^{+}14, DMD^{+}00, DTOD00, DBW^{+}22, DKRL22, DW02, DJG^{+}02, DGGdR02, DCRR^{+}22, DFC^{+}21, DCL^{+}13b, ESA^{+}13, FZY^{+}23, FJA^{+}21, FMWW14, GMDD^{+}22a, GMAB07, GRD^{+}23, GGG^{+}18, GHG^{+}24, GEP^{+}08, Ham87, HHK^{+}22, HVTV22, HWLT10, HHH^{+}12, HPZC21, HAH^{+}22, Hut81, Hut87, Hut92, Hut95, HHK^{+}02, HHB^{+}22, IAN13, IG19, JCIG18, JMSB^{+}23, JGO^{+}98, JWD^{+}02, KSVT00, KMS^{+}24, KVLA06, KHM^{+}88, KDF97, LM18, LLL^{+}11, LCBN14, LSH^{+}11, LFC^{+}15, LBH^{+}21, LCJ^{+}07, LZG20, LGH^{+}21, LHEB98, LMM03, MZ14, MIN^{+}20, MB07, MR03, MH14, MWFH02, MLPN06, MSL^{+}07, NSE^{+}24, NAH^{+}21, ON05, ONR^{+}14, ORR^{+}02, OVR^{+}02]. shelf [OC06, PMC21, PJS^{+}22, PL89, PHC^{+}19, PDD^{+}22, Pra97, PdMS^{+}13, RCC^{+}18, RWOA01, RGMPR23, RAE^{+}05, RCSHW22, SHP^{+}23, SENS13, SSI13, SMP^{+}22b, SFF^{+}24, SBD^{+}07, SSL07, SNMW10, SVIA14, SPW22, SCB^{+}16, SÖÜ94b, TCN20, TNC^{+}09, TLH^{+}15, TLP^{+}16, TPM^{+}00, TSP^{+}13, UNN^{+}14, VEM^{+}21, VOJD02a, VOJD02b, WWL^{+}22, WSL20, WGM^{+}24, WYT00, Whe06, WC15, YMI88, YSY^{+}19, YGC^{+}21, YCP^{+}12, YJS86, ZCA21, ZAC^{+}23, ZMW^{+}23, ZBRJ23, dPCS23, CTKF^{+}23, FKH^{+}13, GF97, HMP^{+}13, Hic92, KFM^{+}17, LNB13, LW12, LDHW20, MPD^{+}22, PHK^{+}17, PMC21, PCM11, PLN^{+}23, SMP^{+}22b, SZG06, SPH^{+}15a, SBPGP^{+}23, Tol85a, TPTM23, WGZZ19, ZLR^{+}07]. shelf-break [CMC^{+}16, LMM03, OC06]. Shelf-slope [JOBT05, LHEB98, MB07]. shelf-water [ZAC^{+}23]. shelfbreak [HHP06, LPS^{+}19]. shell [MLB^{+}20, OE65]. Shelled [JMZ23, AOMZ^{+}23, DGVGR24]. shellfish [CAH^{+}22]. shells [IPHW^{+}23]. shelves [BD18, CBC^{+}06, CW06, HBL^{+}13, HKGH^{+}06, MHH^{+}15, MB07, Was06, WC15]. Shift [RF17, XYK^{+}22, Bea04, CHB02, DRVMC^{+}22, Dri06, FAH^{+}13, INI^{+}17, IIM^{+}23, IIS^{+}17, PM13, SSVP00, SE08, SE09, SKH00, SKT01, TCN20, TKW06, TKWI08]. shifts [AN04, AHW^{+}15, BHAJ12, BMG^{+}21a, BDC^{+}08, CALS^{+}23, CAT^{+}08, CRS04, CS04, EiIT^{+}22, FPS^{+}13, HM00a, HMWM00, HPB^{+}09, HDZY15, iIRM^{+}15, KJZ^{+}12, KFM^{+}17, Law04, LO21, LM14, LHF^{+}16, Man04, MP04, OIC^{+}23, ORMB08, Qiu15, RS04, SvN04, Ste04, SM16, TOKLC08, VBL04, WXH07, WZ04, YTNK00, ZLKO00, dHA^{+}04]. Ship [DPR^{+}18, GdRGL^{+}01, SGR^{+}22]. Ship-based [DPR^{+}18]. ship-borne [SGR^{+}22]. shipboard [WFS^{+}15]. shipping [DN07]. ships [SPH^{+}15b]. shoaling [PFW15]. shore [CCD^{+}13, JIT^{+}01, LMM03, Pra91]. shoreface [Let87]. Short [AVK91, BC91, BGM^{+}01, BF01, CB91, CJMI^{+}91, MIW91, PWMIM91, RÁSG^{+}13, TMN^{+}12, Ver91, WP91, CSMGS19, DLM^{+}12, FB01, Fre07, JSA^{+}08, SGL^{+}18, Sie69]. short-beaked [SGL^{+}18]. Short-term [AVK91, BC91, BGM^{+}01, BF01, CB91, CJMI^{+}91, MIW91, PWMIM91, RÁSG^{+}13, TMN^{+}12, Ver91, WP91, CSMGS19, DLM^{+}12, FB01, JSA^{+}08]. short-time [Sie69]. SHOTS [APSC11]. show [PCR^{+}22]. shrimp [CPG08, CMF11, CHSB^{+}21, PPdS21, SCB^{+}09, TGJT09]. Si [PHK^{+}17]. Siberian [DCS^{+}22, KPM^{+}23, OV24]. Sicily [BMC05, BFB^{+}20, OAB^{+}16]. side [CBL^{+}19, RZW^{+}23]. sidescan [SW92]. signal [Mol04, RKS01, SJ02a]. signals [BPSGP^{+}23, FG16, SD07, Szu12]. signature [DCS^{+}22, MVS08, VSGC21, YGL^{+}10]. Signatures [KiL14, CZW^{+}22, CSC^{+}12, CDP14, OPH^{+}24, PRL^{+}18, QOS^{+}22, RAG^{+}19, VKGP^{+}13, WRS^{+}92]. significance [AS88, BB24, GDI^{+}09, KSE^{+}09, LRW^{+}15, MGS90, Nie07, OB98]. Significant [WSC^{+}21, DBR20, ZPY^{+}20]. significantly [CBL^{+}19, PBB^{+}20]. silica [CWZ^{+}20, GTR01, KYT^{+}16, RGC^{+}01, TRP^{+}23, TGR05, WSC^{+}21]. silicate [LCR^{+}93]. silicates [Ros65]. Siliceous [SLG^{+}12]. Silicic [TSRF14, TRP^{+}23]. Sill [Zen08]. Similar [GGQ07, VPM^{+}19]. Similarities [KTIT22]. Similarity [HMPZ11, RLP^{+}18]. Simple [GLS08, KKK04a, KKO10, RLC85, Zez90]. simplicity [Fly03]. simplify [RSG06]. Simulated [FYYC05, TFM03, Hea12, RKCH15, SGWF^{+}19, WWSJ07, YFY05]. Simulating [BSH^{+}20, JB15, PST^{+}15, PGG^{+}22, VDS^{+}18, HLK13, RM93, Woo05]. Simulation [UNN^{+}14, APC13, AR18, Ché14, CJMI^{+}91, DRE^{+}08, Har05a, RDL^{+}91, SEW11, TS10, VSA^{+}21, WDC^{+}11]. Simulations [MSF^{+}07, HYM^{+}24, LC22, MHGGS19, Men21, RMK^{+}21, ZCD08]. Simultaneous [DDJ^{+}21]. Since [ALV^{+}21, DYL^{+}15, HHWW20, HMKF08, TAF^{+}22]. single [EGP^{+}18, GASV^{+}09]. single-cell [GASV^{+}09]. Sinica [Ang86]. sinicus [NMN08]. Sink [GGT^{+}15, ARELAK24, CMF15, MHS^{+}20a, MHS^{+}20b, MKOLA20]. sinking [IL20, KSP^{+}23, LFBP^{+}13, MLD^{+}03, MLL^{+}15, NMK^{+}03, SPB^{+}02, VDP^{+}01, VKJ^{+}23, YYK88]. sinks [SE16]. sinuosity [MHR^{+}10]. sinus [RPRCAG^{+}21]. siphonophores [BLCL14, LSIC12, Pug84]. sister [GBG05]. Site [SIB^{+}06, BEI^{+}20, BTS^{+}15b, BKD^{+}20, CLL^{+}18, GSV^{+}01, GHL15, GBB96, HFO90, KVNT20, MMG^{+}11, OAD22, RWD01, TAW^{+}15, Tho77, XCH^{+}16, ZHSMM14]. Site-specific [SIB^{+}06]. sites [HG04, MHCS^{+}23, WSO^{+}13]. situ [BGR^{+}15, FRV^{+}19, GA00, ORMR^{+}19, PSP^{+}21, SSH^{+}05, SPK^{+}19, SHS^{+}05, SNS^{+}22, TT05, WZFW16, WFS^{+}15, dlGFM^{+}23, ASR^{+}20, CDS90, FGR^{+}06, HMO^{+}13, KPSA17, LKDL14, NIF^{+}15, SFAD^{+}90, SAT^{+}22, TSAM^{+}22]. situation [BTNK13]. situations [LSM^{+}22]. Six [YTL^{+}19, BSFM^{+}12, SEG^{+}22b]. Size [ATT^{+}08, IL20, SSN23, SEG22a, WSS15, ARD^{+}03, AJA^{+}22, AV23, BM01, BTJ^{+}17, CNT03, CTI^{+}19, DAF^{+}22a, DAF^{+}22b, DFH^{+}16, DHB^{+}21, FEGA^{+}14, FDE^{+}22, GCCY^{+}14, GDN^{+}18, GGAA^{+}23, GWK17, GSSWK20, GGA^{+}16, HEF^{+}12, HHMB^{+}09, HVEF09, IOGS13, IAFD02, JTD^{+}14, KRL^{+}22, LC12, LBH^{+}21, LHC^{+}19, LLX^{+}21, LLGS21, MBdM^{+}18, MRAP22, MRBS^{+}24, MFS^{+}07, MSF^{+}07, MPB^{+}23, MMPG07, NGLL^{+}22, PPSVC^{+}13, Peñ03b, PS98, QPR03, RSW^{+}23, RB20, SPSV^{+}20, SE92, Sok90, TSBS18, UAM05, VFS^{+}15, WSC^{+}21, WJPHB15, XHW^{+}20]. size- [ARD^{+}03, SPSV^{+}20]. size-abundance [GGAA^{+}23]. Size-based [WSS15, WJPHB15]. Size-fractionated [SSN23, BTJ^{+}17, FEGA^{+}14, JTD^{+}14, LHC^{+}19, XHW^{+}20]. Size-fractioned [SEG22a]. size-sinking [IL20]. size-specific [HHMB^{+}09]. size-structure [CTI^{+}19]. size-structured [HEF^{+}12, MFS^{+}07, MSF^{+}07, NGLL^{+}22]. sized [KT04, RLGC10]. sizes [MCT03]. sizing [CTP^{+}18]. Skagerrak [Sva65]. skate [BECA22]. skates [ECFT20]. skipjack [KAK^{+}22a, PGS^{+}22, SSL08]. sled [JPBB20]. sledge [LB20]. sliding [KHM^{+}88, XYWY23]. sliding-window-threshold [XYWY23]. Slope [DCD^{+}23, TPTM23, ACL^{+}18, BHE^{+}98, BHPC06, BPC^{+}05, CB06, CGM^{+}02, CMF^{+}09, CS06, CFML22, CP17, DW02, FBR^{+}13, FLdST98, GGG^{+}18, HM98, HDM19, HFW^{+}98, Hut87, Hut92, IVR^{+}13, JOBT05, KSVT00, KFC^{+}13, LPS^{+}19, LHEB98, LFCSV^{+}13, LSIB23, MB07, PPdM^{+}12, PCD^{+}18, PL89, RGPB^{+}23, RCSVGP^{+}16, SGO^{+}08, SPB19, TCDPP^{+}22, TRLA^{+}13, dWDB^{+}98, vHVAT22, RDP^{+}21]. slopes [CTR^{+}19, DSB^{+}24, FBD18, GD85, KKS^{+}18]. slurry [SL13]. Small [BCL^{+}09, LLAPG^{+}22, POS^{+}07, PAB^{+}21, PS98, BRR^{+}12, BAP^{+}22, BFV^{+}17, CSS^{+}19, DPH^{+}18, DHHP18, FFS^{+}20, GRD^{+}23, IPD14, PRTC13, RFC^{+}15, SGF^{+}19, SCD^{+}07, SLBVRR^{+}22, SLOP^{+}22, SPC^{+}23, SPK^{+}22, SSV^{+}11, TCF^{+}18, VCSG^{+}01, ZJZ^{+}21]. small-scale [BRR^{+}12, CSS^{+}19]. smart [NHS^{+}14]. Smith [WR03]. smolt [IOGS13]. Snapper [ZLRVB24a, ZLRVB24b]. snow [AS88, BCD^{+}20, IL20, MGKW19, TSAM^{+}22, Tur15]. snowdrift [Ber65a]. snowmelt [NNFL21]. SNU [HLK13]. SOAR [MS15]. soaring [Ric15, Sac16]. SOCCOM [SJA^{+}23]. social [Dav99, Jac10, PBO10, PM13]. Societal [RAB^{+}11, KA94]. Society [CSH^{+}23]. socio [GRDS10]. socio-economic [GRDS10]. sockeye [McK08, MI21]. SOFAR [Owe91, Ric93]. Soft [GGSM^{+}20, QOS^{+}22]. soft-bottoms [QOS^{+}22]. software [GPAB^{+}16]. Sognefjord [SST^{+}17]. solar [Ång65, HYM^{+}12, LLL^{+}24]. sole [LDAM^{+}07, TTMM^{+}17]. Solea [TTMM^{+}17]. Solenogastres [BHB^{+}19]. Solitary [PM85, GC14, HHZ^{+}22, HYM^{+}24, NP00, VB14, XDG^{+}23, XHC^{+}20, YHZ^{+}22]. soliton [BPSGP^{+}23]. Solomon [CGD^{+}22]. soluble [BJ17, DVB^{+}18, WRS^{+}92]. solute [LTSG13]. solutions [HHAR23, KN10, KN11, TDH^{+}95]. Somali [Sch83]. Some [Car97a, Eme65, Eri65, Sva65, WO85, Ano86a, Ano87a, Ano89a, GBM^{+}01, Mid69, Wil87, Wis65]. Somniosus [ON22]. SONar [GDN^{+}18, SW92]. Sonne [SKF20]. Sørfjord [MPN09]. Sound [RN06, PNF^{+}21, CCS^{+}21, SKGS20, SCB^{+}16, USH15b]. sounder [ZYN^{+}24]. sounders [ZZPL18]. soundscape [EVM^{+}15, PS23]. Soundscapes [MGG22]. Source [dCFK17, TLF^{+}89, WMC^{+}89, ARELAK24, DJW^{+}18, DBJ^{+}15, HMS^{+}22, KBF^{+}08, MXC^{+}21, SMN^{+}14, SHC^{+}06, SHC^{+}07, UKM^{+}14, WWL^{+}22]. sourced [TCDPP^{+}23]. Sources [BMG^{+}21b, CdMS^{+}18, CBD^{+}24, FTG^{+}18, JJR^{+}08, JAJS08, BSF95, CFC^{+}18, GLV12, KSS^{+}23, SF85, SE16, TSC03, WRS^{+}92, YAK13, ZGZ19]. sous [Bou65]. sous-marins [Bou65]. South [CAH^{+}22, DRE^{+}08, FBT^{+}22, HWF^{+}21, KMF^{+}20a, PFE10, TSFA22, TPP^{+}00, ATC^{+}19, CWB^{+}22, Hen85, Ike88, MAAS^{+}00, SCLG^{+}11, SDGVE17, SPW22, TMKJ^{+}09, WGM^{+}24, YHZ^{+}22, YCP^{+}12, ARD^{+}03, ABD^{+}17, ÁSÁB^{+}14, ÁBMÁS14, ÁBMÁS15, AFH^{+}11, APP21, BMdMS^{+}21, BWB^{+}09, BTS22, CGL^{+}20, CWZ^{+}20, CKM^{+}21, CSW96, CLX^{+}20, CMF15, Cra09, CKL^{+}14, DFC^{+}21, DGH^{+}20, DHL^{+}21, EMK^{+}17, GDM^{+}20, GC14, HMRA^{+}03, HZD^{+}23, HMPZ11, HKY^{+}11, HPZC21, HLSX22, HHZ^{+}22, IIM^{+}23, JFG^{+}90, KHD22, KMF^{+}20b, LvBS^{+}24, Li14, LYZ16, LGZ^{+}20, LBH^{+}21, LZL^{+}22, LYS^{+}22, LLL^{+}24, LLH^{+}20, LGD^{+}20, MXC^{+}21, MCD^{+}14, MBH^{+}23, MZZ^{+}23, MSA^{+}22, NXY15, OKdA^{+}19, PJS^{+}22, PS91, PG10, PAF^{+}11, QLW10, QLY^{+}22, RFFL21, Rei86, Rei89, RLX^{+}24, RDC^{+}21, STC10, SDGVE17, SJP10, SSTL16, SMP07, SWH^{+}24, TBK^{+}99, Tom81a, VCM04, WWL^{+}22, gWjNfLyD20, WST^{+}21, WMWR08, WL16, WZC20, XWL^{+}18, XLX^{+}20, XHC^{+}20, YGC^{+}21]. South [Yao88, ZZPL18, ZDG^{+}21, ZMW^{+}23, ZDM^{+}20, ZSW^{+}22, YJW88]. south-central [YCP^{+}12]. south-eastern [SCLG^{+}11, TMKJ^{+}09]. south-west [CWB^{+}22, SPW22, CKL^{+}14]. South-Western [PFE10, ATC^{+}19]. southeast [BPSN^{+}21, DBRK17, JJA^{+}13, LHF^{+}16, SJ02a, FZ88, FHG03, FMM^{+}20]. southeastern [AIHB^{+}07, ABS^{+}20, CP02, DBM17, FLUC08, GCB^{+}22, HZCZ16, HS02, LDAM^{+}07, Lie86, MWFH02, WD94, AG22, BHH^{+}16, dFKdLZTT17]. Southern [CPO^{+}19, DMBHG10, FTG^{+}18, SCS^{+}18, SC23, ZHD^{+}20, ALG^{+}21, AVK91, BC91, BHA^{+}14, BTK^{+}99, BMC17, BHAJ12, BCGN^{+}18, BMN19, CFM^{+}18, CB91, CLG^{+}00, CLA^{+}00, CJMI^{+}91, CdD^{+}15, CS04, DBC^{+}18, DYL^{+}15, ES07, EHG^{+}07, EM12, FCN^{+}19, GHVG19, GLPC23, HE07, HEF^{+}12, HOY^{+}21a, Igu04, JTQ^{+}18, KTN14, KKS^{+}18, KTIT22, KGdS^{+}08, KVLA06, LVGH^{+}15, LSF^{+}17, LJPGC02, LO21, LMC^{+}20, LGH^{+}21, LFCSV^{+}13, LB02, MERB12, MMES16, MB07, MIW91, MMF^{+}12, MH14, MTH^{+}10, MHCR^{+}12, MA12, MMN^{+}24, MSL^{+}07, NRS^{+}19, NMC^{+}09, OSH^{+}96, PMMN^{+}22, PRA^{+}18, PWMIM91, PVV23, RFSCF19, RBNJ^{+}12, RSM^{+}23, RBS^{+}09, SLM^{+}16, SCMAR^{+}99, SE08, SE09, SMPC^{+}12, SAd^{+}17, SKP99, SJM^{+}19, SHT^{+}01, TAF^{+}22, TFM03, VAEP24, Ven12, Ver91, WP91, WHK23, YNTS22, ZLC^{+}15, ZSY^{+}22, AMY^{+}23, AH80, APSC11, ASC92, ABP^{+}23, Ban65]. Southern [BN03, BLES16, CBB^{+}22a, CFML22, Don65, DIM09, DPF^{+}20, Epp92, FPJ^{+}15, FACM^{+}23, GWB14, GPA^{+}11, GCZ^{+}00, GLF^{+}17, GWGR^{+}19, GBH^{+}20, GVKD^{+}13, GMDS20, GLLB22, HVTV22, HSLG11, HFK03, hHRW^{+}05, HVEF09, HPH^{+}08, JE92, KF11, KMMC09, KC15, KMWF11, LSV14, LOG^{+}09, LPA^{+}11, LdSH^{+}15, LS15, LLS01, MR06, MST^{+}23a, MFS^{+}16a, MFS^{+}16b, MKOLA20, MJD^{+}21, MCMT^{+}17, MVS08, MCH^{+}12, Nof03, NST^{+}23, PPVG12, PPKR14, PO15, PL09, RVS^{+}21, SJA^{+}23, SSN23, TNY^{+}24, TTF^{+}22, TAH^{+}11, UCB^{+}18, VKGP^{+}13, VMA^{+}24, WWW^{+}23, WMWR08, WTT14, ZPY^{+}20, ZBY^{+}22]. southward [SOO^{+}14]. Southwest [CBM^{+}21, GCS91, GRdSS^{+}22, HFW^{+}98, JJA^{+}17, MZ14, USH15a, WLM^{+}13, BH85, HHR^{+}19, HM08, JJA^{+}17, WWN^{+}99, ZQWP23]. southwestern [AVG^{+}19, ASR^{+}20, CTL^{+}04, IMW^{+}14, MFA^{+}15, PTP^{+}22, TTF^{+}22, DMC^{+}18, EBM^{+}21, LMT^{+}19, SNdSR^{+}24]. sp [AM19, BC19, GA01, HLS^{+}14a]. Space [MPSS91, Ang88, MKMF^{+}89]. Spain [BAOC^{+}09, ÁSFP^{+}03, Ano09h, BFH01, BAOM^{+}12, EAL^{+}07, HHB^{+}01, LMP22, RWOA01, RÁSG^{+}13, VSPP14]. Spanish [DAKV99, JW01a, ORW^{+}01, SGL^{+}18, TIOM16, VBL^{+}21]. spanning [MJC^{+}17]. sparse [SVIA14]. Spatial [BMO12, BAOM^{+}12, CRHM12, CCW^{+}18, ESGP17, EM12, FJA^{+}21, FFT^{+}18, GXX^{+}22, GMAB07, GBC^{+}16, GHG^{+}24, HRA^{+}08, Igu04, IVR^{+}13, IU14, JZZY24, JYK^{+}14, KSP^{+}23, KRL^{+}22, LT06, LSY^{+}14, LZL^{+}22, LYS^{+}22, LCGH07, LLX^{+}21, MERB12, MPSD15, MTC14, OSH^{+}96, PPSV^{+}18, PCBA^{+}20, RCF^{+}13, SLM^{+}16, SFS^{+}12, SPB^{+}12, SSTL16, SEG^{+}22b, SEG22a, SBE^{+}20, TvW98, TSBS18, TIOM16, VEM^{+}21, WM13, WTT14, WL16, YMK^{+}04, AMFY20, AGD^{+}18, BRR^{+}22, BSC^{+}19, BBE^{+}03, BDT^{+}08, BDL08, BGB^{+}08, CNT^{+}19, CDH^{+}13, CDB^{+}22, CMM^{+}04, CMF11, CDL^{+}22, DAvD^{+}20, DAvD^{+}21, DCRR^{+}22, DDCE^{+}23, DAIS10, DK07, ESD^{+}21, FBD18, FHP83, FPIJ85, FSVL10, FTG^{+}18, GKC^{+}14, GLH13, GTS^{+}21, HHK^{+}22, Her88, HEF^{+}12, HKPV12, HPW10, ISM^{+}02, IFC^{+}07, IHR18, JFG^{+}90, JBB^{+}14, KM10, KFH^{+}15, KMWF11, Leh01, LSM08, LO21, MRM^{+}14, MZF^{+}08, MHS^{+}09, MCT03, NCC^{+}15]. spatial [NHSP23, OPG^{+}10, OPH^{+}24, ORB^{+}18, OAM00, PWZ^{+}16, PHD^{+}18, PS98, RKC^{+}10, iSIS02, STB^{+}92, SK18, SMPC^{+}12, TKK^{+}05, WPH^{+}10, XWL^{+}18, XRC^{+}15, XNT^{+}17, YN20, dLLdAWL^{+}23, BDB^{+}04]. Spatial/temporal [Igu04]. spatialized [Mau10]. Spatially [LLS01, BDL08, CMS^{+}13, RBNJ^{+}12]. Spatially-explicit [LLS01, CMS^{+}13]. spatially-structured [RBNJ^{+}12]. Spatio [BFV^{+}17, DSV^{+}24, GBT^{+}19, GGA^{+}05, HvDL^{+}17, HLD^{+}21, HVEF09, LFC^{+}15, MDB^{+}20, Min00, ŠGM^{+}18, TS10, WHK23, YPVP^{+}22, BBB^{+}21, BSC^{+}07, CGG08, CSC^{+}12, ERT^{+}22, GRD^{+}23, GFB^{+}15b, GFB^{+}15a, JFUR20, JHW^{+}14, LSB^{+}17, LSD^{+}18, MSC^{+}15, SLH^{+}19, SJLW23]. Spatio-temporal [BFV^{+}17, DSV^{+}24, GBT^{+}19, GGA^{+}05, HvDL^{+}17, HLD^{+}21, LFC^{+}15, MDB^{+}20, Min00, ŠGM^{+}18, TS10, WHK23, YPVP^{+}22, BBB^{+}21, BSC^{+}07, CGG08, CSC^{+}12, ERT^{+}22, GFB^{+}15b, GFB^{+}15a, JFUR20, JHW^{+}14, LSB^{+}17, LSD^{+}18, MSC^{+}15, SLH^{+}19, SJLW23]. Spatiotemporal [AHP19, AOMZ^{+}23, CAH^{+}22, HHP10, PTP^{+}22, SDO^{+}14, DIQJ21, LG22, LS15]. Spawning [BFB^{+}20, HLP^{+}16, MMD^{+}16, dFKdLZTT17, AQVB^{+}10, BSF^{+}21, BSC^{+}07, DPGC14, FP03, FCN^{+}19, GGJ^{+}10, HLS^{+}14a, JSdSS^{+}21, KKKY10, KMF^{+}20a, KMF^{+}20b, LC12, PDAM^{+}15, PFE10, RLL^{+}09, RDP^{+}21, SMPC^{+}12, TOKLC08]. speaking [Ang79a]. Special [FJhT^{+}14, LM10, MSI17, SBB^{+}14, BZD^{+}21, Dol09, HWB^{+}18, NMN08, RW97, SGPdM18, Ver91]. speciation [GWGR^{+}19, LTJ^{+}15]. specie [MIW91]. Species [Ant09, BJ90, FTHK19, Kam19, SM21, ARD^{+}03, AP20, AQVB^{+}10, ASB^{+}08, BJMP19, BJMP20, BC19, CES^{+}19, CC23, CBOP15, DAF^{+}22a, DAF^{+}22b, EGP^{+}18, FKH^{+}13, GTB07, GKR20, GS19, GPP22, GBG05, GM19, GA01, GMDS20, GAS^{+}22, HSG^{+}15, HCV^{+}20, HLD^{+}21, Igu04, IS19, JM19, JZ19, JGB20, JOGM^{+}10, JLRB20, JG90, JPBB20, KSVT00, KFM^{+}17, KT04, KGdS^{+}08, KDB95, LMPB^{+}16, LML^{+}23, LAGM^{+}23, LS15, Mar20, Mil88, MAFS^{+}22, MSFZ19, NHG19, NKK^{+}05, PMG15, PRC^{+}20, PMH17, PD15, PTPY^{+}23, PVA24, RLGC10, RK20, RVC^{+}13, RF17, SSB19, SPF^{+}23, SWP^{+}13a, SEG^{+}22b, Sma10a, Sma10b, SJD10, SBE^{+}20, SL13, TPRS10, THM^{+}14, TSFA22, UKK^{+}19, VCSG^{+}01, Wil65, WSH15, WJPHB15, YGL^{+}10, ZMCD11, ZSY^{+}22]. species-based [WJPHB15]. species-level [LMPB^{+}16]. Species-specific [Ant09]. specific [Ant09, HHMB^{+}09, SIB^{+}06]. spectra [GCCY^{+}14, GGAA^{+}23, GWK17, GSSWK20, Kra69, KRL^{+}22, QPR03, YYhT^{+}17]. Spectral [STR01, KM08, Mor91]. spectrum [GWM^{+}22, LBC^{+}15]. Spells [Cia22, SOWS17, SDB^{+}21, YW22]. spiciness [Fla02, Kat18]. spinifer [GBG05]. spinifera [FMC^{+}20]. spiny [MPM^{+}18]. spiral [QYF^{+}24]. spirals [Wun24]. Spiraphiline [CES^{+}19]. Spit [NW87]. Spitsbergen [KMSTK23, KGJ^{+}10, SPW22]. split [ON22]. sponge [RTN90, VBJ^{+}20]. sponges [KT04]. sporadic [SSB14]. Spores [GGE^{+}65]. spot [HNSP^{+}19]. spots [YSS14, ZMCD11]. spp [DOS^{+}18, Gif93, HHY03, KSKN21, MMG^{+}13, McK15, VSPP14]. sprat [CCM^{+}14, HKPV12, MHTG10, SRK15, SK17, VPH^{+}12]. Sprattus [CCM^{+}14, PBB^{+}12a, PBB^{+}12b, SRK15, SK17]. spreading [BF12, Kil76, LRNK99]. Spring [DHHP18, GCS91, MSMR93, NSE^{+}24, ZGB^{+}20, AW13, CMC^{+}16, Car97a, DBC^{+}18, DBM17, EKB06, FHP83, FB05, GMDD^{+}22b, GHL15, GKS^{+}13, GDM^{+}15, HHY03, HPHL^{+}05, HBH^{+}17, JS90, JMZ23, KHBA^{+}24, KFM15, KTW^{+}22, LBC^{+}23, LPHL^{+}05b, LPHL^{+}05a, LMA^{+}15, MERB12, MPD^{+}22, MCKS17, Mil93a, MA12, MMD^{+}16, NEI^{+}22, OTNI20, RHM^{+}19, iSIS02, STB^{+}92, STF^{+}13, SMP^{+}22a, SPH83, THP21, TRP^{+}23, TDL^{+}17, VHV^{+}12, VEM^{+}21, YMA^{+}17, ZLG17a, ZLG17b, ZWM^{+}15, dFKdLZTT17, RAP95]. spring-neap [Car97a]. spring/early [EKB06]. Springtime [APC^{+}21, HJLLN07, JTD^{+}14, DMBB02, GRD^{+}23]. Spur [BHE^{+}98, FLdST98, FMH02, LHEB98, OB98, TvW98, vWHdS^{+}98]. squat [ACK^{+}13, GRB^{+}08]. squid [ATT^{+}08, BGM^{+}10, IIS^{+}17, OMK^{+}22, RS10, RASVB^{+}22]. Sr [YAI^{+}14]. SST [MS02, SNS^{+}22, WWZ19]. St [LBC^{+}23, MPSD15, MPN09, TFM03]. St. [HGD22, FHG03]. Sta. [LMPB^{+}16]. Stability [MDGC^{+}12, WPB05, Cai95, CGS23, Fla02, Gir15, HGH^{+}19, JJL^{+}19, LAD^{+}18, Law04, NGNV12, VMB^{+}22a]. Stable [BAOC^{+}07, CGC^{+}20, FMC^{+}15, Kli10, SF02, Ang84, AYH^{+}23, BGA^{+}21, CSK^{+}12, CSC^{+}12, CDP14, ESTM13, EiIT^{+}22, GDM^{+}20, GWM^{+}22, IBW^{+}01, Kno04, KSS^{+}23, KSG^{+}17, KAAK^{+}16, LSV14, MRAP22, OPG^{+}10, PPHM18, PYKF15, QOS^{+}22, SBC^{+}16, SC65, SHC^{+}06, SHC^{+}07, SM16, VKGP^{+}13]. Stage [NMN08, HCGK11, IVR^{+}13, MMG^{+}13]. Stage-V [NMN08]. stages [AHP19, CSMGS19, CGC^{+}20, CCM^{+}14, Eme65, Emi65, FVLC^{+}23, FCN^{+}19, HPW10, KSK^{+}15, LSY^{+}14, MVBC^{+}21, NHG19, OCH^{+}18, PWZ^{+}16, PDAM^{+}15, SLY^{+}15, TCF^{+}18, VMH^{+}21, WFH^{+}22]. staging [PMG15]. Stainforthia [GA01]. stalagmites [SCS87]. standard [MDR22]. standing [ADS^{+}22, Fro93, GSPMAI99, IVR^{+}13, JSHB90, MMKS^{+}21, OBD^{+}20, RLP^{+}18, TRLA^{+}13, WSC^{+}21]. stars [MJD^{+}21]. starvation [HCGK11]. starvation-induced [HCGK11]. starved [TCDPP^{+}22]. State [MPC12, SBB^{+}14, AdAK^{+}18, CAA^{+}07, Fla02, GTNK21, HLK13, JCM^{+}21, MHA^{+}11, McD81a, McD81b, RWD01, RGC^{+}01, SRM^{+}10, TAM^{+}13]. states [Kno04, Bum73, Eme65]. Station [AEPW93, Fre07, AVK91, BC91, CB91, CJMI^{+}91, HLM^{+}13, MIW91, MA12, NCH^{+}07, PWMIM91, Ver91, WP91, CR20, KLB^{+}21, KBC^{+}22, MDGC^{+}12]. stationary [DCD^{+}23, PCR^{+}22, dSPF^{+}23]. stations [BHLU^{+}07, DLJ^{+}21, Lev88, SS69, SH09, YNMY23]. Statistical [FPJ^{+}15, SW92, WPH^{+}10, BIL03, BPGD^{+}14, GAM98a, HLSX22, MNT14, Owe91, SBM91, SDO^{+}14, SLY^{+}15, dMM69]. Statistics [LaC08, BBB^{+}21]. status [CSBL^{+}15, CAO^{+}20, CBGC^{+}08, FVA^{+}19, Fai65, Joh04, JJS03, KTN14, KSE^{+}09, LCANAS^{+}07, SGL^{+}18, SF02, VJJ^{+}22]. Stay [CMS^{+}13]. STCC [CMG15]. steady [RWD01, RGC^{+}01]. steelhead [SKGS20, WBF^{+}21]. steep [JOBT05, vHVAT22]. Steering [LMA^{+}15]. Stephos [SSTD^{+}95]. Stepwise [TSL10]. Steric [MZZ^{+}23, RG09, SRFHDH22]. Sternoptychidae [CSM^{+}15, EBM^{+}20]. stimulate [CRC^{+}19]. stirring [Mar03]. Stock [ZL01, ADS^{+}22, BHMS09, Dol09, Fro93, IAFD02, JSdSS^{+}21, JSHB90, KMB01, MHTG10, SEG22a, TRLA^{+}13, WBC^{+}22]. stocks [CGV13b, GSPMAI99, IVR^{+}13, LNB13, MMKS^{+}21, MHS^{+}09, OBD^{+}20, SEG^{+}22b, TGJT09, WSC^{+}21]. Stoichiometric [ÁSÁB^{+}14]. stoichiometry [BPA^{+}21, Fly10, FAH^{+}13, HWL^{+}20, PHK^{+}17, SGS^{+}23]. stomach [KSG^{+}17]. stomachs [EBD^{+}20]. stomatopod [MKD90]. Stomiiformes [CSM^{+}15, EBM^{+}20]. Stomiiforms [SLOP^{+}22]. Storage [PPKR14, HTdM^{+}15, lLdZQ^{+}22, WCN^{+}05]. storm [CBB^{+}19, CBB^{+}22b, LFBP^{+}13, MFH86, PFHM16, PPSVC^{+}13, SWP^{+}13a, SZG06]. storm-induced [PPSVC^{+}13, SZG06]. Strait [JJR^{+}08, KHC^{+}99, SNR^{+}10, AHGRAL23, AW13, AHC^{+}13, BFB^{+}20, BPGC^{+}20, BCR^{+}13, BCK94, BC88, CML^{+}16, CGW^{+}22, CGMP14, CZW^{+}22, DWH^{+}14, Dea85, DL17, ESTM13, FY88, FAAF88, FDM^{+}13, GIC20, GXX^{+}22, GLPC23, GWS^{+}23, GSSWK20, GKS^{+}13, GDM^{+}15, HFPS^{+}06, HMH^{+}15, Ike88, IOGS13, ITO^{+}14, JC88, KAG^{+}19, LBNBM13, LF12, LAP10, LMH^{+}13, LHC^{+}19, LC10, MMGL^{+}07, MGF^{+}13, MGWZ20, MRO^{+}08, MP13, MK86, MH14, NGLSSG14, PMC16, PM13, PST^{+}02, PBN13, Ric94, RHBS13, Rud89, RKS^{+}15, STF^{+}13, SGMP15, SBPGP^{+}23, SLH^{+}19, SSV^{+}11, TYO^{+}14, TSBS18, VGLCS06, VKDS^{+}18, WDC^{+}11, Woo18, YYT^{+}14, YLY^{+}14, YYhT^{+}17, YMK^{+}04, ZYN^{+}24, ZLX^{+}20, ZWP23, dlPHF^{+}15]. Straits [KTB^{+}99, ABC^{+}99, ZMW^{+}23, AC85, BTK^{+}99, BMC05, GR85, GSPMAI99, KSPK99, MM90, MWO^{+}12, NTU^{+}14, RLGC10, SKP99, TBK^{+}99, VOT^{+}99]. Strangomera [CCM^{+}14, SYB^{+}15]. strategies [BVJE19, BOMdP15, BHS^{+}15, BCL^{+}09, DPGC14, KPSB17, LDH90, LPA^{+}11, MGS90, Nie07, RMB^{+}01, SSW^{+}09, THBA19, VBC^{+}20]. strategy [SSTD^{+}95, SSB14, YAK^{+}08]. Stratification [BBS21, SHL13, GS24, GCD^{+}13, LK13, LFC^{+}15, MZK^{+}23, PFW15, SHd13, SYN^{+}21]. stratified [BMG^{+}21b, CTA16, FvBA^{+}17, Her88, HAH^{+}22, IAN13, KV13, KV18, MBP^{+}11, SMFM^{+}21, SENS13, TSP^{+}13, WPW^{+}14]. stratigraphic [CFM^{+}18, Par65]. stratigraphies [Sel65]. stratigraphy [Med87, Sai65]. streaked [YSS14]. Stream [TPTM23, AC85, BHPC06, Fug63, Ham87, Hen85, Hog85, KAH^{+}16, LS85]. streamer [ZAC^{+}23]. streams [LGZL22]. strength [YFK21]. strengthen [CKM^{+}21]. Strengths [BBE^{+}15, HS07]. Stress [FWH^{+}17, AR18, BCOL^{+}19, HHK^{+}02, IPG^{+}16, MS02, QOS^{+}22, VTGC19, WWZ19, XY21]. stress-SST [WWZ19]. stresses [LEDR^{+}22]. Stressors [SKCP23, dSSDS^{+}20, TCS15]. Striking [XLX^{+}20]. striped [JOGM^{+}10]. Strong [KPM^{+}23, KSG^{+}17, FELJ16, GTS^{+}21, HSLG11, JMSB^{+}23, NFMCS^{+}22, SEG22a, WST^{+}21]. stronger [KLC^{+}15]. Structural [RLP^{+}18, BHHR15]. Structure [GIPC^{+}15, HKGH^{+}06, JG90, MRA^{+}19, MDR22, PP85, SP08, TCS15, Was06, AMFY20, AAM^{+}14, AMG^{+}16, AV23, Ang79b, ABT^{+}04, ASR^{+}20, BAM^{+}09, BIST01, BS90, BWB^{+}09, BLP93, BLES16, CRGA17, CLSD18, CWB^{+}22, CSV^{+}07, CDL^{+}22, CC88, CS03, CBM^{+}21, CFML22, CF12, CCB^{+}20, CTI^{+}19, DBC^{+}18, DW02, DRE^{+}08, DSR21, Dol09, Dom84, DCRR^{+}22, DBR20, EBD^{+}20, ESD^{+}21, ES07, EHG^{+}12, FMH02, FDE^{+}22, FPS^{+}09, FAB^{+}09, GSFP^{+}09, GAF15, GCCY^{+}14, GPP22, GDM^{+}20, GWM^{+}22, GHVG19, GGA^{+}16, GVKD^{+}13, Ham87, Hey78, HGBG20, HSL96, HG04, IBW^{+}01, INT14, JSHB90, KP03, KOhL^{+}10, Kos93, KBSB18, LCBN14, LdCSB^{+}20, Lon85, LS85, MLL^{+}22, MT99, MERB12, MDAW^{+}19, MTC14, MLM09, Min00, MTK^{+}22, Mol22, MK12, MTH^{+}10, MMN^{+}24, NYL^{+}17, PCM11, PJH^{+}15, PHD^{+}18, PBS22, PLK14, PCBA^{+}20, QSC^{+}15, RSW^{+}23]. structure [RD03, RLR^{+}18, RCSHW22, RBE^{+}12, SSQ19, SSS^{+}11, SIR^{+}07, SNZ^{+}20, Sie69, SSM90a, Sim84, ŠGM^{+}18, SPMVP05, SNMW10, SPN98, SST^{+}17, Tan99, TWMY08, TMPM^{+}16a, TKW06, THM^{+}06, TSFA22, TSBS18, TPP^{+}00, UB10, VKDS^{+}18, VDGGD^{+}22, WFH^{+}22, WLL^{+}23, WLM07, XYK^{+}22, XHW^{+}20, YYT^{+}14, YNMY23, YJ88, ZHSMM14, ZLS^{+}04, ZZM^{+}24, ZBRJ23, ZSBL00]. structured [HEF^{+}12, MMG^{+}13, MFS^{+}07, MSF^{+}07, Mau10, NGLL^{+}22, RBNJ^{+}12]. structures [AHSS22, DMBHG10, HCAFD^{+}20, Ken88, MPM^{+}18, OSH^{+}96, SHC^{+}06, SHC^{+}07, YHZ^{+}22, ZSH^{+}24, AVS23]. structuring [CGD^{+}18, JJA^{+}13, KM10]. studied [PKA19, RWD01]. Studies [Ang79b, APSC11, HWS^{+}07, KY15, LJ65, Ros65, BGM^{+}99, BAOC^{+}07, CAT^{+}08, CSC^{+}12, Cow05, DFC^{+}21, Dri11, FACM^{+}23, FRK^{+}09, Has82, Hey78, ITM86, IPD14, JW01a, LOG^{+}09, MGKW19, ML09, Mos69, Rud03, SSM^{+}90b, SJ18, SGR^{+}22, VPS09, WFD^{+}07, WBA^{+}22, WPA^{+}24, WH89, ZNR^{+}24, dLLdAWL^{+}23]. Study [ALV^{+}21, AYK^{+}05, BOG20, KNI^{+}05, ASFB^{+}13, ABE^{+}15, AHA^{+}16, Ang80, AVK91, AUE^{+}14, ALT10, BC91, BEI^{+}20, BLHB07, BCLD^{+}17, BHH^{+}16, BDE03, BPTT19, BLT^{+}08, BCL^{+}09, CCM^{+}13, CCRS20, CGMP14, CB91, CJMI^{+}91, CCMS08, CKL^{+}14, DGH^{+}20, EALF08, ESTM13, FTSF21, Fro93, GHL15, GAPM16, HLPL05, HBV^{+}10, HBL^{+}13, HPNDC15, HPW10, JAC^{+}12, JLB^{+}08, JW01a, Ken88, KIS^{+}05, KA85, KFC^{+}13, KGdS^{+}08, KH09, KS15, LJM^{+}16, LMW^{+}12, LSB^{+}17, LB20, LGD^{+}20, Mau10, MMG^{+}11, MIW91, MRW^{+}14, NHG19, ORW^{+}01, PBB^{+}20, PKP14, PWMIM91, PGC^{+}96, PCH^{+}08b, Pre86, RCÁS^{+}15, RLX^{+}24, RZW^{+}23, RAE^{+}05, SZG06, SKHD84, iUMY86, VMA^{+}24, Ver91, VMN08, WP91, WDMMK89, XRC^{+}15, YKWF21, YFY^{+}22, ZDG^{+}21, vWMH98]. Sub [ABD^{+}17, ABÁS^{+}09, BRR^{+}12, NYH^{+}22, BMG^{+}21b, BCR^{+}13, GvOSW11, HMPZ11, HDA^{+}16, MPM^{+}18, PM22, SBH^{+}14, TŠT^{+}17, VPM^{+}19, WMB^{+}21, WXH07]. sub- [HMPZ11, MPM^{+}18]. sub-Antarctic [VPM^{+}19]. Sub-Arctic [NYH^{+}22, GvOSW11, WXH07, BMG^{+}21b]. sub-mesoscale [BCR^{+}13, PM22]. sub-polar [HDA^{+}16, SBH^{+}14, WMB^{+}21]. Sub-regional [ABÁS^{+}09]. Sub-surface [BRR^{+}12]. Sub-Tropical [ABD^{+}17, TŠT^{+}17]. Subacrtic [Mil93b]. Subantarctic [BRR^{+}22, Ban96, GRdSS^{+}22, IGG^{+}19, NHG19, TSRF14]. subaratic [SPB93]. Subarctic [BLI^{+}99, BMN^{+}99, CNT^{+}19, CM11, CCW^{+}18, KNI^{+}05, MN88, SPS^{+}99, AT07, BE99, BLP93, Dag93, DRD^{+}07, Fro93, FMT15, Gif93, HBV^{+}99, INO^{+}24, KIS^{+}05, KST03, KRL^{+}22, LGK^{+}93, LMS93, LCGH07, LBP15, MSMR93, MT99, MFB^{+}84, Mil88, MC88, MTK^{+}22, MHVS19, MST^{+}23b, NMO^{+}21, NKK^{+}05, Peñ03a, PV07, RTF^{+}05, Rog00, SSH^{+}05, Sek99, SPMVP05, SHS^{+}05, TFY02, TKSI08, TT05, Tan99, TWBC^{+}13, TSNO05, TKK^{+}05, WFH^{+}22, WNNI21, WSG^{+}93, Whe93, WFR07, YAK13, AYK^{+}05, BRE^{+}24, FK99, HBL^{+}13, HMP^{+}13]. subbasin [MTL05]. subinertial [BPSGP^{+}23]. Subject [Ano63d, Ano64e, Ano65j, Ano65k, Ano69d, Ano73e, Ano85k, Ano87i, Ano89l, Ano90d, Ano90e, Ano92j, Ano93g, Ano86a, Ano92a]. subjected [ACL^{+}18]. sublittoral [GA01]. Submarine [RVC^{+}13, AHA^{+}16, AHD18, BRC^{+}18, BD19, CDL19, CHG^{+}18, CJRÁ^{+}13, CQZ^{+}18, CHC^{+}12, CD65, CHSB^{+}21, DP18, DSB^{+}24, DSC^{+}19, DAU22, DCRR^{+}22, DCL^{+}13b, FBR^{+}13, FBD18, GvOS^{+}08, GCF^{+}19, GBB^{+}20, HSN^{+}18, IVR^{+}13, IHR18, JFEC13, JOBT05, KCL^{+}12, KFC^{+}13, LFCSV^{+}13, MRH^{+}18, Nay65, PGLG^{+}05, PCD^{+}18, PGT^{+}13, PRC^{+}20, PCC^{+}19, RCC^{+}18, RCF^{+}13, RCSVGP^{+}16, SCB^{+}09, She65, TCDPP^{+}22, TCDPP^{+}23, TAF^{+}22, TPPG10, TTF^{+}22, TCL^{+}15, ZFSV^{+}09]. Submersible [SFAD^{+}90]. Submesoscale [BPGC^{+}20, CLY22, HNR^{+}17, LLH^{+}20, MBS20, MZZ^{+}23]. suboxic [GGQ07]. Subpolar [GIPC^{+}15, VFS^{+}15, BDTC15, DMT15, FKZ^{+}15, FMP19, GTNK21, HPB^{+}09, HLM^{+}16, KY15, KJH^{+}22, LWT^{+}20, Rea00, SF15, TNY^{+}24, Yas07a, ZBLF23]. Subsea [CLMR23]. subseasonal [RSK^{+}23]. subspecies [PMG15]. substrates [CDS90, GGT^{+}15]. Subsurface [HMX^{+}23, KY14, SN24, ZSH^{+}24, ZSY^{+}22, CLY22, Don94, JRR^{+}24, MVS08, PTZ^{+}23, PVV23, RKK^{+}21, SBM91, YYC^{+}18]. Subtidal [NH88a, BRR^{+}22, RHBS13]. Subtle [DBR20]. Subtropical [CMJPH^{+}18, DLC^{+}08, GRdSS^{+}22, HPC^{+}20, INI^{+}17, Kat18, LSMG01, AGL^{+}15, BHS^{+}15, BDBJ01, CRGA17, CMF15, DBRK17, FHG03, FCMCÁS19, FMSBW13, HSK^{+}19, HMX^{+}23, HM08, KF11, LRAE23, LLH^{+}21, LMT^{+}19, MZH^{+}23, MS17, MSI17, MZZ^{+}23, MTK^{+}22, Mol04, NHH^{+}23, NCH^{+}07, NMO^{+}21, PHCA17, PAF^{+}11, RASVB^{+}22, SHL13, SASH08, WLM^{+}13, Yu23, YSN20, ZLG17a, ZLG17b, dPGSHL23, BSMC15, GSF^{+}15]. success [BWMGCB08, BEP02, IMHL07, MAFS^{+}22, PHLL05, YFY^{+}22]. successful [DWFP^{+}19]. succession [LBP^{+}21, PFHM16]. Suez [Ore69]. Sugars [TPN^{+}18]. Suggested [Ano94k]. suggests [PDAM^{+}15, SLY^{+}15]. suitability [BHH^{+}16, DFH^{+}16, HLP^{+}16, SAY^{+}16, WWW^{+}23]. sulfide [ZPY^{+}20]. sulfur [GLY23, NEI^{+}22, OYKK^{+}23]. sulphide [BCP09, TJ90]. sulphidic [PLN^{+}23]. sulphur [TNS^{+}05]. Sulu [HYM^{+}24, XDG^{+}23]. Summarising [BBE^{+}03]. Summary [LFA^{+}06, PAB^{+}87b, Man69, TT05]. Summer [DLC^{+}08, MDR22, OMS^{+}09, YLY^{+}14, AMY^{+}23, AAMB^{+}24, BTNK13, BvdLA^{+}11, BSH^{+}20, DSR21, DHDM22, DL17, EKB06, EHG^{+}12, FB01, GDL^{+}15, HWPLvW20, HBK^{+}24, HHY03, HEF^{+}12, IMM^{+}22, INT14, KMS^{+}24, KGB^{+}23, LBC^{+}23, LPBM17, LHC^{+}19, LZG20, MST^{+}23a, MSS^{+}02, MKOLA20, MKS^{+}22, MJA^{+}07, NST^{+}23, PVG^{+}20, PTP^{+}22, PELAA18, Par86, PVV23, RBHLA04, RVS^{+}21, SM21, SGMP15, SSN23, SEW11, SYN^{+}21, TII^{+}14, TTL^{+}04, TVT^{+}23, WFH^{+}22, WMB^{+}21, WST^{+}21, WQ08, YKS^{+}12, YNMY23, YLL19, YJS86, YJ88, ZDG^{+}21, dWDB^{+}98, dFKdLZTT17]. summers [AMFY20]. Summertime [Lie86, RLX^{+}24, WMWR08]. sun [PNF^{+}21]. sunken [JZ19]. sunscreens [FWH^{+}17]. Super [VBL^{+}09, WSG^{+}93]. Super-ensemble [VBL^{+}09]. superba [AMY^{+}23]. superficial [GGA^{+}05, SGL^{+}13]. supplies [ZCV^{+}19]. Supply [RNBP^{+}19, CJMO87, DTOD00, LBK^{+}01, QCdS^{+}07, RCSHW22, SH09, TAM^{+}15, VGJ^{+}19, VMN08, WP91, WW02]. support [AHGRAL23, BMN19, BDE03, JPM^{+}08, WFJ^{+}15, VSPP14]. supports [GTS^{+}21]. suppresses [MZK^{+}23]. Suprabenthic [ACL^{+}18]. Surface [ABM^{+}05, CP07, Don94, GBC^{+}00, KBHML17, KSD84, NBR^{+}08, NMO^{+}21, RM97, RCD^{+}94, RKFD07, San15, SBPGP^{+}23, SAT^{+}22, SCB^{+}16, ATS01, AMEV07, AEPW93, APP21, Ban96, BCOL^{+}19, BH07, BHC^{+}18, BF01, BRR^{+}12, Bre06, Bri83, Cai95, CLV^{+}19, CPC^{+}15, CVBG21, CMHM18, CSS^{+}19, Cia22, CPSM20, CPNL07, CMF15, CRHM12, CBT07, CM14b, CEF^{+}13, DHC^{+}20, DWH^{+}14, Dem09, DVB^{+}18, DSC^{+}21, DL17, DDJ^{+}21, EALF08, EGPM^{+}15, EBR^{+}14, FFA09, FBT^{+}22, FWH^{+}17, FLDF22, FTG^{+}18, GPA^{+}11, GL06, GEPC15, HMTL05, HKN^{+}14, HMRB^{+}03, Har82, HHK^{+}22, HZCZ16, HMX^{+}23, HJLLN07, Her97, HT97, HMPZ11, His22, hHCK01, ICB^{+}19, IHT^{+}21, IU14, Iwa23, JW01a, KK20, KTN14, KGL22, KKNT23, KDL^{+}01, KPSB17, LW85, LOO22, LCZ^{+}24, LFBP^{+}13, LBD11, McD88, Mit91, MFM85, MKS^{+}22, MKSW^{+}15, PO00, PFW15]. surface [PTG95, RCB^{+}20, RDL^{+}91, RI86, RGI05, Rud89, SPF^{+}23, SCHBC^{+}22, SLG^{+}12, SN24, STG^{+}18, SDGVE17, SCCJ^{+}18, SSTL16, SC23, SMGL01, Soh03, SPV^{+}15, SRG^{+}19, SJ02c, SJ02b, SYN^{+}21, TM13, Tho95, TSRF14, TBW00, VSA^{+}21, VBL^{+}09, WFH^{+}22, WLD^{+}15, WH20, WST^{+}16, XWL^{+}22, XCH^{+}16, YIY^{+}04, YSY^{+}19, YPM^{+}10, YSS14, YYK^{+}12, ZLC^{+}15, MOS^{+}13]. surface-sediment [SSTL16]. surficial [Cow05, RSB^{+}01, ZHBW01]. Surge [Cho86, MFH86]. surplus [TGJT09]. surprising [WSH15]. surrogacy [HSN^{+}18]. surrounding [RK03a]. Survey [HBD^{+}18, ASJ^{+}23, BGMP03, CNBD21, DPH^{+}18, DPR^{+}18, Hop64, HFK03, KDB95, OIC^{+}23, PHD^{+}18, TIOM16, VR03, WH94, WPA^{+}24]. surveyed [BBMR19]. surveys [CHB02, EGP^{+}18, HKL^{+}15, LPBM17, MSA^{+}22, PV18]. survival [AHP19, AHC^{+}13, CCS^{+}21, CHS^{+}24, DPGC14, HLP^{+}16, HPNDC15, IMHL07, IOGS13, KBF^{+}08, Kli10, MGKW19, MI21, PHFK14, SKGS20, SMPC^{+}12, SLY^{+}15, WBF^{+}21]. Suspended [CLG^{+}00, MFY^{+}86, Par63, DCS^{+}22, FRV^{+}19, FDE^{+}22, IU14, JGO^{+}98, MBdM^{+}18, MLL^{+}15, NKK03, PPCWJ18, PLP99, PBP^{+}99, SRF^{+}19, SPB^{+}02, USH15a, VFCC^{+}22]. Sussex [JS87]. sustainability [GHF^{+}21, LCANAS^{+}07]. sustainable [McI10, Ric01]. Sustained [HBD^{+}21]. sustaining [INO^{+}24]. Svalbard [SON^{+}20, FGL^{+}23, GIHJ23, JCM^{+}21, KMSTK23, MIN^{+}20, ON22, RBR^{+}23, VNH^{+}23]. Sverdrup [MCKS17]. SW [CÁM06, SVL^{+}23, CMS^{+}13, Gal17, KPSB22, LWBD^{+}17, MGC^{+}18, PYKF15, TCDPP^{+}22, YHM^{+}18, YPVP^{+}22]. swallow [RLC85, Ano85a]. Swansea [Ang80]. swell [CBB^{+}22a, VTGC19, ZKT88]. swift [Cra09]. swim [GGT^{+}15]. Swimming [ASC07, FGS^{+}23, BLP^{+}20, SS17, SK17]. Swirling [CBB^{+}22c]. swish [SL13]. switch [GS24]. switching [VWDF14]. Swoddies [TMPM^{+}16b]. swordfish [SYB^{+}15, YGL^{+}10]. symbiosis [AB90]. sympagic [SHC^{+}06, SHC^{+}07]. sympatric [NHE^{+}13, SBC^{+}16]. symposia [CSH^{+}23]. Symposium [LM10, HMWM00, BK08, SRM^{+}10]. synchronicity [PVA24]. Synchronies [LSS^{+}09, KQP^{+}17]. Synchronization [Yu23]. Synchronized [QNK^{+}22]. synchronous [KOhL^{+}10, PCR^{+}22, TOKLC08]. synchrony [BMO12, MDAW^{+}19, SBL^{+}23]. Synechococcus [SDS^{+}22b]. synergistic [Sie88]. synergistic-consortia [Sie88]. Synoptic [GRLS14, LPBM17, OPL^{+}21, CP83, LZL^{+}22, SMR^{+}20, SGMP15, ŠVL^{+}15, iUMY86]. Synoptic-scale [GRLS14, OPL^{+}21]. Synthesis [BK08, CAT^{+}08, CKT^{+}13, GTS^{+}21, HMO^{+}13, LGZ^{+}20, MOS^{+}13, MS15, ARH^{+}00, ABD^{+}17, AIA^{+}18, Dri11, DP13, EAL^{+}07, FTG^{+}11, HHH^{+}12, HMKF08, MRMD^{+}97, MB05, MGK^{+}86, MST^{+}23b, TBK^{+}99]. synthetic [Ric94, VOT^{+}99]. System [AAM^{+}14, GBC^{+}16, MAB^{+}11c, MAB^{+}11a, MAB^{+}11b, PAVB^{+}21, ABE^{+}15, AASJ23, AVK91, BC91, BTK^{+}99, BIST01, BFB^{+}20, BDT^{+}08, BF01, BMG^{+}21b, BHHS83, BLT^{+}08, BLMR^{+}20, CFM^{+}18, CQZ^{+}18, CB91, CGZ^{+}16, CP83, CJMI^{+}91, CTP^{+}18, CKL^{+}14, DLM^{+}12, EALF08, ESTM^{+}12, ES07, EB08, EBvdL^{+}09, EHG^{+}12, FFA09, FVLC^{+}23, FDH20, FWBC02, GSFP^{+}09, GMAMB04, HHK^{+}22, HGD22, HSH^{+}19, HKGH^{+}06, HFPS^{+}06, HNR^{+}17, IVR^{+}13, JE92, JW01a, KAH^{+}16, KAK^{+}22b, KMSTK23, LSV14, LFG10, LBP15, LSW02, MERB12, MCG^{+}02, McD81a, MMES16, MIW91, MR03, MJC^{+}17, PM13, PBS22, PWMIM91, QNK^{+}22, RBL^{+}19, RAP95, RDC^{+}21, RÁSG^{+}13, RBS^{+}09, SNV^{+}18, SHD^{+}21, SAM^{+}04, SGO^{+}08, SCLS10, SKRM^{+}95, SHT^{+}01, TMH^{+}16, TLH^{+}15, USH15b, VSGC21, VSA^{+}21, VAEP24, Ven12, Ver91, WP91, Whi95, WFS^{+}15, YKWF21, YGC^{+}21, ZHF^{+}24, AH10, ÁSDB^{+}01]. System [ARG11, CCW^{+}02, CBGC^{+}08, CB09, CRT^{+}22, CRHM12, CLdPHL23, DFM^{+}21, DPF^{+}20, GDI^{+}09, GCED22, Hau84, HLGA07, Hic79, HW02, Huy83, JIT^{+}01, JBB^{+}14, KSD84, LO07, LQU07, LO21, LPF^{+}18, LABD^{+}24, LB02, ML09, MAB^{+}11a, MAB^{+}11b, PO15, REG^{+}15, RMK^{+}21, RD03, RZTD17, SFMA20, SE16, SDK84, Sim84, SKHD84, SAd^{+}17, SBG^{+}08, TMÁGC^{+}21, TSJ^{+}12]. Systematic [GJ00]. Systematics [BHB^{+}19]. systems [BCL^{+}09, CLV^{+}19, CPO^{+}19, CED09, CTI^{+}19, Dem09, EBvdL^{+}09, GHVG19, GCF^{+}19, HDM19, HVEF09, IPF23, JPM^{+}08, KMB01, KSC10, KHJ^{+}10, LZL^{+}22, LSB^{+}17, LSD^{+}18, LHF^{+}16, MS02, MC15, MAB^{+}11c, MAB^{+}11a, MAB^{+}11b, MZGA^{+}20, MCGS^{+}16, PBO10, PP10, PFHM10, RFSCF19, SWP^{+}13b, Sma10a, Sma10b, ST10, TPRS10, MFM15, TBW09].

T [Ang80, BDB^{+}22, PCK^{+}06]. T. [GBG05, MRAP22]. Tōhoku [CLB^{+}14]. Tagging [PMM^{+}23]. tags [BMC^{+}10, FFT^{+}18]. Tagus [JGO^{+}98]. Taiwan [DL17, JC88, LHC^{+}19, MH14, STJ^{+}14, WLP^{+}21, Yin88, ZLX^{+}20]. take [ROBRB^{+}22]. takes [WJPHB15]. tale [BKC15, SWT^{+}17, Wai21]. tall [TDK^{+}16]. Tanaidacea [JPB20]. Tanaidacean [SPB19, BJMP19, BJMP20, JPBB20]. tank [FHL^{+}24]. Tanner [RKCH15]. Tantulocarida [PKA19]. tantulus [PKA19]. target [CBB^{+}22c]. Tasmania [OLH^{+}18, RL23]. taxa [ALM^{+}23, GSV^{+}01, WPA^{+}24]. taxon [LB20]. Taxonomic [CSR90, CCW^{+}18, Kam19, PVM^{+}20, YMA^{+}17]. taxonomy [JZ19]. Tchernia [Hof81]. technical [BCF^{+}03, SPK^{+}19, TDH^{+}95]. technification [PPSV^{+}18]. technique [Jer65]. techniques [BPGD^{+}14, BGR^{+}15, SNV^{+}18, Tom81b, VBL^{+}09]. Tectonic [JFG^{+}90, Kit03]. teleconnection [Has06]. teleconnections [OÁT^{+}05, ZLC^{+}15]. telemetering [Dah69]. telemetry [MHR^{+}10]. tell [RMB^{+}01]. telling [VLUC^{+}07]. temperate [CDB^{+}22, DBW^{+}22, FB05, GMDD^{+}22a, HAH^{+}22, LM18, LBP^{+}21, MPD^{+}22, MFDH22, MAFS^{+}22, PDD^{+}22, RCSHW22, RHM^{+}19, SSB20a, SMP^{+}22b]. Temperature [GSM^{+}17, Yao88, Zen08, Ång65, ATS01, BCOL^{+}19, BHHR15, Bre06, CGB07, Dem09, DDJ^{+}21, ED82, FVLC^{+}23, GAPM16, GEPC15, HKN^{+}14, HJLLN07, Her97, HT97, HHH^{+}12, hHCK01, JRR^{+}24, KGL22, KC15, LSXT01, LHW^{+}20, LLX^{+}21, MEMP15, MMF^{+}17, MPSS91, Mid69, MKM86, MFM85, Mol22, MSB^{+}23, NGNV12, NKK^{+}05, PO00, RG09, Rou65, SCHBC^{+}22, SDP^{+}22, SMGL01, ŠGM^{+}18, Spr08, SLY^{+}15, SPV^{+}15, STGR^{+}23, TMN^{+}12, TOKLC08, Tom81b, VYGMM^{+}17, WHT86, Whi95, YYK88, YPM^{+}10, Yin88, YYK^{+}12, ZLC^{+}15, dLLdAWL^{+}23, dPCS23, vHMDL14, Rou65]. temperature-depth [ED82]. temperature-salinity [ED82]. temperature/salinity [Tom81b]. temperatures [DW02, NPO^{+}19, YKNO23]. template [DAKV99]. Temporal [BTS^{+}15b, BMNW01, BLP93, CCHV^{+}21, FTSF21, FELJ16, GMBU12, GSV^{+}01, HSMLDC^{+}22, HLM^{+}13, HYM^{+}12, HHZ^{+}22, ILA21, ISM^{+}02, KBC^{+}22, KMWF11, LSMG01, MLD^{+}03, MCD^{+}07, MPC^{+}17, MCT03, MSL^{+}07, PS23, SBL^{+}23, STB^{+}92, SKP99, SK21, VK90, VSC01, VDB^{+}20, WLM^{+}13, BDB^{+}04, BMO12, BBE^{+}03, BBB^{+}21, BSC^{+}07, BFV^{+}17, Car98, CGG08, CSC^{+}12, DAvD^{+}20, DAvD^{+}21, DDCE^{+}23, DSV^{+}24, ERT^{+}22, ESGP17, EM12, FHP83, FFT^{+}18, GBT^{+}19, GMAB07, GRD^{+}23, GFB^{+}15b, GFB^{+}15a, GGA^{+}05, GIPG17, GHG^{+}24, GVKD^{+}13, HvDL^{+}17, HLD^{+}21, HVEF09, Igu04, IVR^{+}13, JFUR20, JW01a, JHW^{+}14, KBE^{+}22, LFC^{+}15, LSB^{+}17, LSD^{+}18, LCGH07, LLX^{+}21, LdCSB^{+}20, MRM^{+}14, MSC^{+}15, MHS^{+}20a, MHS^{+}20b, MDB^{+}20, MFDH22, MPSD15, MHS^{+}09, MZ14, Min00, NTU^{+}14, OAM00, PSP^{+}21, PS98, RCF^{+}13, iSIS02, SPB^{+}12, SEG^{+}22b, SEG22a, SK18, ŠGM^{+}18, SMPC^{+}12]. temporal [SLH^{+}19, SKH00, SJLW23, TvW98, TS10, TKK^{+}05, WM13, WGM^{+}24, WTT14, WHK23, YPVP^{+}22, YMK^{+}04, dLLdAWL^{+}23]. temporary [MOSN^{+}13]. tenuimana [ACK^{+}13]. TEOS [AdAK^{+}18]. TEOS-10 [AdAK^{+}18]. TEP [ORMR^{+}19, Pas22]. term [APC13, ABE^{+}15, AVK91, BC91, BGM^{+}01, BBR^{+}01, BLCL14, BD18, BF01, BBL^{+}09, CMF^{+}09, CSMGS19, CB91, CSK^{+}12, CB17, CJMI^{+}91, DLM^{+}12, DLM^{+}96, FRV^{+}19, FB01, FMP19, FAH^{+}13, FMSBW13, GML^{+}23, GMD^{+}22, GHL15, HFS^{+}20, hHRW^{+}05, HHZ^{+}22, JSA^{+}08, JLS^{+}22, KON14, KNS^{+}03, KRHS14, LO07, LHE^{+}13, LSIC12, LSXT01, MLL^{+}22, MDAW^{+}19, MDGC^{+}12, MIW91, Nag01, PGY^{+}22, PG10, PWMIM91, Reb02, RÁSG^{+}13, STJ^{+}14, ŠGM^{+}18, Spr08, TFY02, TMN^{+}12, TKW06, UPPS^{+}21, VMB^{+}22a, Ver91, WP91, WFD^{+}07, WHBW03]. terminal [MBCB88]. terminating [GHVG19]. termination [Luk86]. terms [Due77]. terns [Cra09]. Terrace [MGS90, VK90]. terrain [DBR20]. terrestrial [CKP^{+}20, DWC06, KiL14, RBL^{+}19, SVHM^{+}13, SMN^{+}14, SGO^{+}08, SCS87, ZHF^{+}24]. terrestrial-marine [RBL^{+}19]. terrigenous [PGT^{+}13]. test [BTNK13, BF11]. Testing [DTC^{+}06, LM97, MPB^{+}23]. tests [PTF10]. Teuri [DWNN04]. TEX [SBH^{+}14]. text [Ang79a]. Th [AYK^{+}05, DTKvH15, HPZC21, ST65, TAF^{+}22]. Th/ [AYK^{+}05]. Thalasseus [Cra09]. Their [VJJ^{+}22, AH15, Ang79b, Ang80, Ang89, AOMZ^{+}23, BB24, BLP^{+}20, BPSGP^{+}23, BF12, BAP^{+}22, Car98, CVBG21, CRT^{+}22, DSAB20, EHG^{+}12, Fla02, Gam14, GLY23, GGT^{+}15, GA10, GMAB07, HPS^{+}01, HKK12, HBK^{+}24, HHSR07, HPW10, KTN14, KiL14, KM22, LSXT01, MPV12, MGS90, MB20, MNS^{+}24, MDAW^{+}19, Mau10, McD81b, MS17, MR03, MST^{+}23b, NMLBCM^{+}01, NMN08, OOTA15, PTM^{+}22, Pra97, Ric22, RGI05, RCSVGP^{+}16, ŠVL^{+}15, SIS^{+}14, TCN20, TvG02, VCM04, WGZZ19, WR03, WL16, YSS14, YN03a, ZLKO00, ZGZ19, ZLX^{+}20, ZHBW01]. thematic [MRH^{+}18]. theme [BHMS09]. Themisto [AE09]. theorem [BBF^{+}22]. theoretic [DYO^{+}10]. theoretical [BH07, GD91, Szu12]. theories [MS00]. Theory [Har05a, CGB^{+}23, CRS04, Dea85, GCCY^{+}14, Kun03, NP00, RK03b, Ste04, vRGW10, vdS94b]. Theragra [MLPN06, YNM^{+}02]. there [BBLD^{+}11, BT07, MKB00, SHF01, PPPdS20]. Thermal [FKH^{+}13, TNC^{+}09, Wun24, CP83, GS24, HGD22, KAK^{+}22a, MK12, PLK14, SBM91, SI97, WLM07]. Thermobaric [Har05a]. thermocline [FMSBW13, HTdM^{+}15, HLTB^{+}17, LS85, TSP^{+}13]. thermodynamic [FH95, Fei03, Fei04]. thermodynamical [HKL^{+}15]. Thermodynamics [War06, Ano94c, Fei93, FM07, MKM93]. Thermohaline [KBSB18, ASR^{+}20, Cai95, NTU^{+}14, RK03b, RR03, SCPN15, Sme93, ZSH^{+}24]. Thermostads [Tsu86]. Thick [PdMS^{+}13]. thickness [WZFW16]. Thin [BLMR^{+}20]. think [Bak01]. thompsoni [GBH^{+}20, LS12]. those [KFC^{+}13]. thoughts [Car97a]. threatened [CLB^{+}13, YZX^{+}23]. threatens [IG19]. Three [BBPHG^{+}11, BASS^{+}20, Dav85, HBD^{+}21, LHC^{+}19, McK15, RL23, SGMVF14, VBL^{+}21, YHZ^{+}22, CES^{+}19, CP19, CBPS^{+}22, DJ92, FYYC05, GWB14, GDM^{+}20, GGA^{+}05, HG04, KDF97, LSM^{+}22, LLX^{+}21, LGG18, Mar20, MRW^{+}14, MR03, MJC^{+}17, PMG15, RFSCF19, RK03a, SN24, SEG22a, SBM^{+}23, SPMVP05, WBB^{+}01, XD95, YWUK15]. three-component [LLX^{+}21]. Three-dimensional [BBPHG^{+}11, BASS^{+}20, LHC^{+}19, SGMVF14, YHZ^{+}22, GWB14, KDF97, MR03, YWUK15]. threshold [XYWY23]. through-flow [vAB96]. throughflow [DWH^{+}14, MH14, Tal08, MMF^{+}07]. throughout [BHS^{+}15, DTG^{+}24]. throughput [HHAR23]. Thunnus [BMC^{+}10, FFT^{+}18, GA10, HHP10, KKKY10, KTIT22, LOBG^{+}10, MRAP22, NXT^{+}17, XNT^{+}17]. thynnus [GA10]. Thysanoessa [CNT^{+}19, FMC^{+}20, MPSD15, SIR^{+}07]. Tidal [Car97b, Egb97, FBS^{+}18, MHCS^{+}23, Pra97, RI86, RGPB^{+}23, TDK^{+}16, ZZWL06, BLR^{+}23b, BMM97, Car97a, CR97, DOP87, DEW^{+}97, ESA^{+}13, FY88, GWS^{+}23, Hen73, Kag97, Mun97, RW97, RCSA01, SVL^{+}23, SCPN15, SPN98, TDL^{+}17, WSL20, XD95, YMI88]. tidally [NNM^{+}21, XD96]. tide [AK97, HHWW20, LM97, NTU^{+}14, SA97, SNdSR^{+}24, XD96, Yux88, vHMDL14]. tide-induced [NTU^{+}14, Yux88]. tides [Arb22, AHD18, Bak83, CGW^{+}22, CZW^{+}22, DJ92, FGSA97, GF97, GCD97, JFUR20, KT97, Kiv97, KDF97, LL97, MMGL^{+}07, Mun69, NGLSSG14, Pra04, RM97, RPSVLS14, SNMW10, TYO^{+}14, WWL^{+}22, WGM^{+}24, WHIH97, ZCA21, ZQWP23, ZZPL18]. tidewater [VNH^{+}23]. tied [TDGY22]. Time [KV18, LMS93, NIF^{+}15, PB07, Whe06, AIA^{+}15, AT07, BSFM^{+}12, BBL^{+}18, BAOM^{+}12, CIL^{+}23, CP83, Coo69, CNBD21, EHFD12, FTSF21, GW89, HFS^{+}20, HLM^{+}13, HFO^{+}22, LQU07, LCJ^{+}17, MGE^{+}12, MGF^{+}13, MNT14, MAH^{+}15, MFDH22, MVC^{+}11, MPSS91, MMG^{+}11, MA12, NCH^{+}07, ORVES17, OÁT^{+}05, OAWAN18, Ric01, RPRCAG^{+}21, RSD^{+}90, Sie69, VLUC^{+}07, VBL^{+}21, VDGGD^{+}22, WST^{+}21, WSO^{+}13, Wun24, XCH^{+}16, ZZM^{+}24, dPAJ07]. time-average [Wun24]. Time-dependency [LMS93]. Time-series [Whe06, AT07, BAOM^{+}12, Coo69, HLM^{+}13, MMG^{+}11, NCH^{+}07, WSO^{+}13, dPAJ07]. times [Men21, WJE^{+}92]. timeseries [LPS^{+}19]. Timing [LSH^{+}11, AW13, BSF^{+}21, CCH^{+}12, HBD^{+}18, IOGS13, LBC^{+}23]. tip [WCC^{+}20]. tissues [AB90]. TIW [NIC^{+}19]. Tiwi [RZW^{+}23]. Todarodes [KSK^{+}15]. together [RK20]. Tokara [CZW^{+}22]. Tomczak [MC08, YRKC08]. tomography [GW89]. Tongue [Fuk91, HDB13]. tonsa [BD20]. Tony [Bil01]. too [WSO^{+}13]. tool [BFPS06, BMN19, CAO^{+}20, FAAV^{+}15, FC05]. toolbox [KHJ^{+}10, MLL^{+}15]. toolkit [WCB^{+}20b]. tools [JPM^{+}08]. toothfish [PWZ^{+}16]. TOp [LM10, AH10, AF10, FZY^{+}23, FFA06, PPD^{+}12, RGMPR23, BVJE19, DYO^{+}10, DSAB20, GPEV20, HS22, HM06, Jac10, KM10, LAD^{+}18, MLPN06, OWR^{+}07, SCHS^{+}24, SFS^{+}12, SIB^{+}06, XYK^{+}22, ZK06]. Top-down [AH10, AF10, FZY^{+}23, FFA06, PPD^{+}12, RGMPR23, BVJE19, HS22, HM06, MLPN06, SCHS^{+}24, SIB^{+}06, ZK06]. top-heavy [XYK^{+}22]. TOPEX [DEW^{+}97]. TOPEX/POSEIDON [DEW^{+}97]. Topographic [Ham09, VKDS^{+}18]. topographies [Bak06]. topography [GSA^{+}20, OAWAN18, RD03, SBG16]. Tortugas [KAK^{+}22b]. TOSCA [BBM^{+}14]. total [BTS^{+}15b, KBC^{+}22, Rei86, Rei89, Rei94, Rei97, Rei03, BBF^{+}22]. totally [GGA^{+}05]. Toxic [PMMN^{+}22, CGD^{+}18, PTPY^{+}23, VMV^{+}23]. toxins [CAH^{+}22]. TPXO.2 [DEW^{+}97]. Trace [CGD^{+}18, ORB^{+}18, BJ17, CE84, CFC^{+}18, HF65, KIS^{+}05, LTJ^{+}15, LJ65, MDL^{+}12, TCDPP^{+}23, ZNR^{+}24]. traced [PRL^{+}18, ZSY^{+}22]. Tracer [APSC11, BS95, HHP06, MWJ^{+}08, TNS^{+}05, VKT15, WNNI21]. tracers [AJV^{+}02, GSPP^{+}20, LvBS^{+}24, Rei86, Rei89, Rei94, Rei97, Rei03, WRS^{+}92]. Trachurus [BHH^{+}16, ZL01]. Tracing [Sme93, ZGZ19]. track [KST^{+}10, LKDL14, LMH^{+}13, MHR^{+}10]. tracked [LC16, RBZ00, ZBLF23]. Tracking [GPE^{+}17, PM22, AGD^{+}18, JIT^{+}01, Ric94, RLC85]. trade [KSY^{+}19]. trade-offs [KSY^{+}19]. traditional [KSY^{+}19, WPA^{+}24]. Trafalgar [SVL^{+}23, BPGC^{+}20]. Trails [WKS^{+}15]. training [DHB^{+}21]. trait [ECFT20, GNH19, NGLL^{+}22, SPSV^{+}20]. trait-based [NGLL^{+}22, SPSV^{+}20]. traits [DLL^{+}23, KBE^{+}22, MRH^{+}14, RNBP^{+}19, SHD^{+}21, THBA19, TMR^{+}21]. Trajectories [MK86, AKAL20, Kir06, LKDL14, LC16, MJ88, PKV18]. Trans [GPC^{+}03, FFT^{+}18, LAGM^{+}23, GWGR^{+}19]. Trans-Atlantic [GPC^{+}03, LAGM^{+}23]. trans-Drake [GWGR^{+}19]. trans-Pacific [FFT^{+}18]. Transatlantic [GBB^{+}19]. Transect [AB00, BJ17, KBHML17, PHCA17, SWT^{+}17, BHC^{+}18, BTV^{+}17, CGM^{+}02, KRL^{+}22, LM00, LGR^{+}02, LHEB98, ARH^{+}00, Ano17a, RNP^{+}17, ZSBL00]. transects [Mol22]. Transfer [SGL^{+}17, VHK03, MRA^{+}19, MGS90, SSS^{+}11, SHD^{+}21, Tho95, VHK04]. Transfer-function [VHK03, VHK04]. transfers [BMG^{+}21b, NBG^{+}05, QYF^{+}24]. transform [OP18]. transformation [ASC92, CCRS20, KHBA^{+}24, MY23, RKS^{+}15, SGMP15]. transformations [BS02, GIPC^{+}15]. transforming [OCH^{+}18]. transgressions [Med87]. Transient [Hol00, RKM^{+}07]. transiently [BGL^{+}17]. Transit [LCJ^{+}17, Men21]. Transition [BMN^{+}99, INI^{+}17, PHKS17, RMHL09, BAM^{+}09, BAT^{+}98, BA04, BBPHG^{+}11, BB14, EM12, GMAMB04, GASV^{+}09, HLGA07, IMM^{+}22, JPIP22, MTK^{+}22, MGH^{+}07, MTH^{+}10, PHKS01, SAM^{+}04, SJM^{+}19]. transitional [VKT15]. transitions [JS90, ORMB08]. translated [Ang79a]. transmissions [DEW^{+}97]. transmissometer [GRMB18]. transparency [GZCL23]. Transparent [MPM^{+}17, ORMR^{+}19, Pas22, RTF^{+}05]. Transport [GPA^{+}11, MNFY21, SCLG^{+}11, Soh03, TCDPP^{+}23, TAH^{+}11, AdAK^{+}18, AC85, AFH^{+}11, ASC92, BBB^{+}14, BMC05, BBM^{+}14, BDBJ01, BMG^{+}19, CB17, CHSB^{+}21, CCD^{+}13, CS06, CBD^{+}24, DWH^{+}14, FLdST98, FK86, GMAMB04, JC04, Jón07, KHD22, KSK^{+}15, KBSB18, LDAM^{+}07, LGD^{+}20, MGKW19, MLHE23, MWJ^{+}08, MWO^{+}12, NBR^{+}08, NNO^{+}14, OOTA15, OAWAN18, ÖÜT93, PPdM^{+}12, PCSMC12, PDAM^{+}15, PGG^{+}22, RBL90, SVHM^{+}13, SFMA20, SGMP15, Sei63, SMN^{+}14, Sek88, SCC14, STJ^{+}14, SMM^{+}90, SMPC^{+}12, Tal08, VK92, VAGMDRS22, VMH^{+}21, VB14, YSY^{+}19, YCP^{+}12, YAI^{+}14, ZSH^{+}24, ZYN^{+}24, ZZWL06, vWdSBdH02]. Transportation [AHSS22]. transported [PGT^{+}13]. Transports [MRO^{+}08, APN^{+}15, GIPC^{+}15, HGPFN^{+}14, HGT16, RBS^{+}20, RBS^{+}22, Rei86, Rei89, Rei94, Rei97, Rei03]. trap [BPTT19, HMKF08, MXC^{+}21, NIF^{+}15, RF17]. trapezoidal [LLL^{+}24]. trapped [CSLJ03, KWI20, vHMDL14]. traps [BEI^{+}20, VK90]. Travel [GW89]. trawl [ATC^{+}19, CHB02, DJW^{+}18]. trawling [DJW^{+}18, PPSV^{+}18, PRA^{+}18]. trawls [RKK^{+}21]. Trench [AM19, AP20, AB90, BBFS19, BC19, BBRM20, CES^{+}19, CP19, CBL^{+}19, FTHK19, GKR20, GM19, GHSC19, JPB20, JGB20, Kam19, KKKS14, KCBS20, KKS^{+}19, MA20, MB20, MDG^{+}19, SSKA19, SKF20, SPB19, UKK^{+}19, YTL^{+}19]. trenches [BHB^{+}19, DSAB20, KKKS14, SJ18]. trend [KC15, LGZW22]. Trends [Dem09, GSF^{+}15, LW13, OLH^{+}18, RJO^{+}19, AH10, BAOC^{+}09, CSR90, CMF^{+}09, CSG^{+}15, CLdPHL23, CM14b, CR20, FTC^{+}16, FAH^{+}13, GPAB^{+}16, GSSWK20, HSG^{+}15, HHH^{+}12, hHRW^{+}05, JSdSS^{+}21, KGL22, LSS^{+}10, LSS^{+}09, MDAW^{+}19, MS02, MMPG07, Reb02, SGL^{+}18, SGA^{+}19, SON^{+}20, SAd^{+}17, Spr08, VYGMM^{+}17, VDB^{+}20, Woo18]. tribute [Ano20u, SMB88, WR03]. Trichodesmium [LMT^{+}19, McK15]. tridens [AHRT90]. Trieste [Her88]. trigger [RFKC16]. triggers [HS22]. trispinosus [AHRT90]. tritium [MBB^{+}96, SBK^{+}95]. Trophic [BWB^{+}09, CARBML^{+}22, CDP14, DKRL22, Dol09, FMM^{+}20, GCP08, HBK^{+}24, HLTB^{+}17, KSS^{+}23, MGS90, PPHM18, SBC^{+}16, SPG^{+}06, TTB^{+}08a, TTB^{+}08b, THM^{+}06, WDC^{+}11, YLL19, YGMR^{+}23, ALM^{+}23, AHW^{+}15, BAM^{+}09, BCGN^{+}18, Car98, CMF^{+}09, CAT^{+}08, CSG^{+}15, CCB^{+}20, DDDT99, EBM^{+}20, EBM^{+}21, EBD^{+}20, ESD^{+}21, FAB^{+}09, FDB^{+}21, FK99, GSFP^{+}09, GSVB23, GWM^{+}22, GCD^{+}13, GAPM16, HBG^{+}21, HSL96, IPG^{+}16, KSB^{+}22, KLC^{+}15, LLL^{+}11, LCBN14, LMS10, LPHL^{+}05a, LSMG01, LLAPG^{+}22, LvIKB07, MPC^{+}17, NYL^{+}17, NRA^{+}21, PCH^{+}08b, RSW^{+}23, RPRCAG^{+}21, SSS^{+}11, SF02, SHD^{+}21, STM10, Sie88, SPH^{+}15a, SPMVP05, SNMW10, SIB^{+}06, SSW^{+}09, Tan99, TWMY08, TSS^{+}12, TMKJ^{+}09, TAM^{+}13, TS10, YZX^{+}23, YWUK15]. trophic-level [SIB^{+}06]. trophodynamic [PL09]. trophodynamics [CFML22, EB08]. Tropical [ABD^{+}17, EBM^{+}21, MPSS91, SKH00, VVV21, WFBN^{+}13, WLL06, AALM06, ASR^{+}20, BPF06, BHS^{+}15, Bri79, CPC^{+}15, DN07, DNNNN16, Don87, Don94, FÁFL06, FT06, FL06, FLUC08, GLY23, GdRGC^{+}14, GAM98a, GAM98b, GA00, Gri22, HKK12, HMPZ11, HNSP^{+}19, JLP^{+}20a, JLP^{+}20b, JSKM02, KSV08, Kes06, LRAE23, LFA^{+}06, LYM12, Leh01, LOO22, LSW^{+}21, MZH^{+}23, MRA^{+}19, MMR^{+}12, MNM06, MFM85, OHC^{+}17, OCH^{+}18, PTP^{+}22, PMK^{+}06, PGS^{+}22, PHCA17, PFE10, RDL^{+}91, RHB23, RASVB^{+}22, SBC^{+}16, SBM91, SCY^{+}23, THBA19, Thu90, TŠT^{+}17, VMN08, WF06, WF07, Yu23, dPGSHL23, AKAL20, EBM^{+}20, PTZ^{+}23]. tropical-subtropical [RASVB^{+}22]. tropicalize [FSAO22]. tropics [MBH^{+}01]. Trough [GD85, DDCE^{+}23, Rot65, ZLC^{+}15]. Trougho [DYL^{+}15]. troughs [MIN^{+}20]. trout [SKGS20, WBF^{+}21]. tsunami [CLB^{+}14, DLM91, RV17]. Tsushima [BC88, IMW^{+}14, KMOM88, KL86, MK86, MKM86, MWO^{+}12, Sek86, TOiF^{+}12, TKW06, TKWI08, YKWF21]. Tsushima/Korea [MWO^{+}12]. tuna [AQVB^{+}10, BMC^{+}10, DFH^{+}16, FFT^{+}18, HLTB^{+}17, HHP10, JTGM10, KKKY10, KAK^{+}22a, KTIT22, LSM08, LSS^{+}10, LOBG^{+}10, MMIB10, Mau10, McI10, NXT^{+}17, OPG^{+}10, PMM^{+}23, PYKF15, PGS^{+}22, RDP^{+}21, SBC^{+}16, SSL08, XNT^{+}17]. tuna-like [LSM08]. tunas [DAIS10, GA10]. tunicate [MM90]. tunicates [LSH^{+}22]. Tunisia [FTG^{+}18]. turbid [LHE^{+}13, SLM^{+}16]. turbiditic [CBD^{+}24]. Turbidity [MH02, PGRP^{+}18]. turbulence [BCG^{+}08, FG16, HHK^{+}02, XD95, vHCY^{+}20]. turbulence- [FG16]. Turbulent [Gar06, BIST01, GD91, PTF10, PTF12, PCBA^{+}20, TYO^{+}14, YYT^{+}14]. Turks [Ché14]. turnover [CC23, CRC^{+}19, CJ92, KKKS14, RSD^{+}90, STHM02]. turtle [CdTH^{+}16]. turtles [CBB^{+}22c, MHR^{+}10]. Twelve [GAM98a, GAM98b]. twentieth [LSW02, RD11, WBH15]. twentieth-century [RD11]. Twilight [VLCCP14, HSLG11, MLHE23]. Two [BAARB05, JIT^{+}01, LOO22, MMK19, WJPHB15, ABD^{+}17, ATC^{+}19, AKH^{+}23, BFP^{+}18, BKC15, CBM^{+}21, DM13, DLL^{+}23, DSC^{+}19, DAU22, EMU^{+}23, FB01, GvOS^{+}08, HGH^{+}19, HBH^{+}17, JG90, Kaw86, KOT^{+}21, LO07, LNB13, McD81a, MSA^{+}22, MMG^{+}11, MHCS^{+}23, MSFZ19, MCGS^{+}16, OMS^{+}15, RAG^{+}19, SWT^{+}17, SBLA10, SHC^{+}06, SHC^{+}07, SJD10, WSH15, YYK^{+}12, vFB82, vPRT90]. two-dimensional [Kaw86]. two-layer [McD81a]. two-mode [vFB82]. two-source [SHC^{+}06, SHC^{+}07]. Two-way [BAARB05]. Tyler [Ang80]. Type [NP00, BPA^{+}21, JSdSS^{+}21, TSL10]. types [BB24, FLdST98, HLK13, IVR^{+}13, KCL^{+}12, LBP^{+}21, Nie07, SN24, YYK^{+}12]. typhoon [DL17, GLY23, HLS^{+}14b, PLK14, vHCY^{+}20, WZC20]. typhoons [MK12, DFC^{+}21]. Typical [Kra69]. typicus [BLHB07, BHLU^{+}07, CCG07, CH07b, CBHL07, DK07, IMHL07, MCD^{+}07]. Tyrrhenian [ALG^{+}21, CFM^{+}18, GCLD19, NIC^{+}19, PRA^{+}18].

U [ST65, AYK^{+}05, TAF^{+}22]. U.S [Ang79a, CTKF^{+}23]. U.S. [CZG^{+}21, KFM^{+}17, RSK^{+}23, WLM^{+}22, WH89, dPCS23]. U.S.$69.00 [SW81]. U.S.A [Ban65]. U.S.A. [Let87]. U/Th [TAF^{+}22]. UAVs [Ric15]. ubiquitous [BD20, EGPM^{+}15]. UK [CWB^{+}22, SV97]. Ulleung [LXC^{+}22]. Última [MPTMK22]. ultraabyssalis [AM19]. ultrahigh [ZBY^{+}22]. ultraphytoplankton [CGD^{+}18]. ultrastructure [JG90]. unaccounted [CMF15]. uncertain [SFF^{+}24]. Uncertainties [FRV^{+}19, ZD17, JVJ^{+}17]. Uncertainty [LS20, TLP^{+}16, BMN19, CSBL^{+}15, LIH^{+}12, MLHE23, MCB^{+}10, MCH^{+}12, Oll15]. uncertainty-based [BMN19]. uncommon [MS02]. Uncorrelated [WPB^{+}08]. under-ice [KSB^{+}22]. Undercurrent [Luk86, JC04, Kos02, TJ73, Yos80, MZH^{+}23, TLF^{+}89]. Underground [NF87]. underlying [BLR^{+}23a]. underpin [HBD^{+}18]. underrated [AM10]. understand [BMG13, CWS^{+}21, CRiI^{+}15a]. Understanding [BT07, GWS^{+}23, HLS^{+}14a, KSK^{+}15, PCK^{+}06, RKC^{+}10, VAEP24, YYK^{+}12, ABC^{+}99, CDL19, FGS^{+}23, MSI17, RAB^{+}11, Smi05]. underwater [Ban96, Jer65, Kaw86, KHM^{+}88, RZTD17]. undescribed [HCV^{+}20]. undesirable [McK08]. undulating [RCM^{+}03]. uneven [YHRT22]. Unexpected [SWH^{+}24, LZF^{+}24]. unexploited [hHRW^{+}05]. unicellular [AV23]. unicorns [Mar20]. unified [CEF^{+}13]. Uniform [VMB^{+}22b, FWL^{+}15]. Unifying [CW06, BRG^{+}15, CWW15, WCB^{+}20b]. Unimak [STGR^{+}23]. unique [BJMP19, BJMP20, vRGW10]. United [Bum73, Eme65]. Uniting [RK20]. units [FCEZ10]. univariate [SBM91]. unknowns [OHH^{+}22]. unprecedented [VMB^{+}22a]. unpredicted [dSSDS^{+}20]. Unraveling [Oll15, SBMB18, ZCV^{+}19]. Unravelling [DLL^{+}23, HPNDC15]. unsaturated [WPB^{+}08]. unselective [DAF^{+}22a, DAF^{+}22b]. unstable [dSPF^{+}23]. Unstructured [CTA16]. unsupervised [MMF^{+}17]. Unusual [AB90, CB06, KO19, RPSC22]. unusually [ZWM^{+}15]. Unveiling [PTPY^{+}23, TTF^{+}22]. Updated [GGT^{+}15, CR20, PHKS17]. Updating [VYGMM^{+}17]. Upper [GCED22, MMR^{+}12, PS91, PP85, WLM07, BHHR15, CWZ^{+}20, Dur09, GR17, GBC^{+}16, GGG^{+}18, HWPLvW20, HKL^{+}15, HGTP^{+}19, HTdM^{+}15, HLSX22, HDB13, JSKM02, KMS^{+}24, KVNT20, LZL^{+}22, LGZW22, LH89, MR06, MWS^{+}10, MK12, NDEG22, NH88b, RGPB^{+}23, SIB^{+}06, Tsu86, WGZZ19, Whi95, ZHBW01, SW21]. Upper-level [PS91]. Upper-ocean [WLM07, GR17, HKL^{+}15, HTdM^{+}15, LGZW22, MK12]. Upstream [LS85, ESTM^{+}12]. Uptake [FCG88, CKM^{+}21, HMPZ11, WCC^{+}20]. upwelled [HHB^{+}01, HW02]. Upwelling [AH80, Ano09h, CM09, FBA09, FAB^{+}09, JIT^{+}01, PSM^{+}22, SCD^{+}07, SE16, VNMS91, AJA^{+}22, ÁSDB^{+}01, AMEV07, AAM^{+}14, AMG^{+}16, AJHC19, ABT^{+}04, ABÁS^{+}09, AVK91, AE09, BC91, BGM^{+}01, BAT^{+}98, BIST01, BA04, BFR13, BLT^{+}15, BFB^{+}20, BEP02, BFH01, BCM^{+}02, BWB^{+}09, BAOC^{+}09, BF01, BM07, Bri83, BHHS83, BCL^{+}09, BLMR^{+}20, CTF07, CED09, CB91, CJMI^{+}91, CCD^{+}13, CNBD21, CS04, DLM^{+}12, Dem09, DRVMC^{+}22, EALF08, ES07, EHG^{+}07, EM12, EHFD12, EBvdL^{+}09, FC07, FTSF21, FFA09, FEGA^{+}14, FAAV^{+}15, FLDF22, FELJ16, FPS^{+}09, FWBC02, GMBU12, GRLS14, GSVB23, GDL^{+}15, GCB^{+}22, GEO09, GGPG^{+}19, GMAB07, GCP08, HPS^{+}01, HYM^{+}12, HE07, HEF^{+}12, HHMB^{+}09, HNR^{+}17, HVEF09, Huy83, IVT^{+}12, JJA^{+}17, JW01a, JRW01, KCPM09, KSC10, KHJ^{+}10, LOG^{+}09, LDB^{+}02, LPF^{+}21, LWBD^{+}17, MRA^{+}19, MERB12]. upwelling [MCG^{+}02, MMES16, MLK^{+}09, MC15, MIW91, Mit83, MMF^{+}12, MDC^{+}07, MHCR^{+}12, MA12, MDL^{+}12, NMC^{+}09, NH83, NIF^{+}15, O’B83, OAWAN18, OÁSG^{+}16, PAM^{+}88, PD15, PBS22, PWMIM91, PP10, PFHM10, QCdS^{+}07, RCGC^{+}16, RBNJ^{+}12, RF17, RÁSG^{+}13, RAG^{+}19, RR01, RBS^{+}09, SLG^{+}12, SFMA20, SNV^{+}18, SMR^{+}20, SDGVE17, Sch83, SHD^{+}21, SAM^{+}04, SGO^{+}08, SC23, SW01, Sma10a, Sma10b, ST10, SM05, SBD^{+}07, SAd^{+}17, SDS^{+}22b, SKRM^{+}95, SÖÜ94b, TNC^{+}09, TSAM^{+}22, TSS^{+}12, TFM03, TPRS10, VSGC21, VDS^{+}18, VAEP24, Ver91, WP91, WFS^{+}15, YHLA^{+}04, Yos80, ZL24, ZHBW01, dlPPÁB24, vFB82]. upwelling-induced [TSAM^{+}22]. upwellings [Ver92]. Upwind [Ric15, Sac16]. Uranium [KG65]. Uranium-234 [KG65]. Uranium-234/uranium-238 [KG65]. uranium-238 [KG65]. urban [XYK^{+}22]. Urup [ITO^{+}14]. US$50.00 [Ang88]. US$90 [Bak83]. USA [FMC^{+}20, PD15]. Use [BDE03, CSC^{+}12, CQO^{+}15, CL03, FGS^{+}23, LVGH^{+}15, LSV14, LIH^{+}12, LMP22, PMM^{+}23, ROBRB^{+}22, SvN04, SSB14, SAB^{+}22]. used [CQO^{+}15, QSC^{+}15]. users [Ang79a]. uses [Ste12]. Using [HLP^{+}16, KHJ^{+}10, LSD^{+}15, MSMH19, PYKF15, RRS03, RWJ^{+}06, RKK^{+}21, Szu12, TSS^{+}12, ARD^{+}03, ASJ^{+}23, AKAL20, AHC^{+}13, BBE^{+}15, BPA^{+}21, BECA22, BGA^{+}21, CLMR23, DPH^{+}18, DLJ^{+}21, EiIT^{+}22, FVA^{+}19, FDHT05, FRV^{+}19, FACM^{+}23, GKC^{+}14, GL06, GDN^{+}18, GTNK21, GiIKX22, HMRB^{+}03, HSC09, HOY^{+}21a, JZ19, JJA^{+}08, KHL12, KSK^{+}15, KM22, KPSB17, LHE^{+}13, LvBS^{+}24, LLX^{+}21, MMG^{+}13, MNT14, NDEG22, NGLL^{+}22, NBLI20, OP18, OIC^{+}23, ON22, Owe91, PM22, PHC^{+}19, RK20, RFC^{+}15, SGL^{+}18, iSIS02, SNV^{+}18, SPC^{+}23, SBFP21, SZG06, Sme93, SNS^{+}22, TCL20, TMR^{+}21, TS10, VMH^{+}21, VBM21, WM13, WSL20, WSS15, WZC20, XRC^{+}15, YSS14, YGL^{+}10, YYhT^{+}17, ZCD08, ZCH^{+}17, dlGFM^{+}23]. utilization [BGM^{+}10, CCH^{+}12, CJ92, NKK^{+}05, XYGJ23]. utilizing [UGY^{+}22]. UV [FWH^{+}17].

V [AYH^{+}23, NMN08, WR03, Ano92i, SKF20]. Validation [TLH^{+}15, TSJ^{+}12, DHB^{+}21, HM00b]. valley [CSR90, JFEC13]. valleys [She65]. value [BBB^{+}21, HWB^{+}18]. values [SSVP00, VYGMM^{+}17]. Vancouver [SL13, VSGD21]. vaquita [RPRCAG^{+}21]. Var [KCL^{+}12, WSO01, MZGA^{+}20]. variabilities [QNK^{+}22, XWW^{+}21]. Variability [ABT^{+}04, APHGC^{+}22, BRR^{+}22, BATNP04, CWB^{+}22, Col69, CF07, DDK^{+}18, FP03, FBS22, Fra69, FDH20, HSH^{+}19, JJL^{+}19, Kaz17, LLGS21, MRSS02, MEMP15, MLS^{+}15, MIH06, Mid69, MA12, NM17, OOTA15, RFSCF19, SKSK06, SON^{+}20, SKRM^{+}95, TKC^{+}22, WLD^{+}15, dPAJ07, ASFB^{+}13, APC^{+}12, ÁSFP^{+}03, ABÁS^{+}09, AVK91, BSC^{+}19, BC91, Bak01, BDB^{+}04, BAM^{+}09, BN03, BGM^{+}01, BTS^{+}15b, BRB^{+}01, BGMP03, BCOL^{+}19, BMC17, BMC05, BSC^{+}07, BMNW01, BLCL14, BDG^{+}17, BHPC06, BPSGP^{+}23, BL02, BM86, BPM^{+}14, CNT^{+}19, CBC^{+}06, CDB^{+}22, CM14a, CCW^{+}02, CLB^{+}13, CB91, CSS^{+}19, CLY22, CHC^{+}12, CS03, CB17, CPSM20, CJMI^{+}91, CD07, CCD^{+}13, CM18a, CF12, CRHM12, CKL^{+}14, CM14b, CM18b, CDB^{+}24, DLM^{+}12, DBC^{+}23, DBC^{+}18, DFM^{+}21, DRE^{+}08, DIM09, Dri11, DBM17, Dur09, ESTM^{+}12, ESTM13, ESA^{+}13, EM12, FTSF21]. variability [FJA^{+}21, FDHT05, FKZ^{+}15, FP15, FBM^{+}08, FFT^{+}18, GMBU12, GSM^{+}17, GR17, GCB^{+}22, GBC^{+}00, GA10, GDSCU09, GGAA^{+}23, GMAB07, GSPMAI99, GJ00, GIPG17, GTS^{+}21, GVKD^{+}13, GLLB22, HSMLDC^{+}22, Has82, HDZY15, HMB^{+}86, HLM^{+}13, HYM^{+}12, HE07, Hol00, HHW01, HHW22, HHH^{+}12, HHP10, hHRW^{+}05, HHZ^{+}22, HDB13, HBD^{+}18, ILA21, ISM^{+}02, IAFD02, Iwa23, JSA^{+}08, JX18, JFUR20, JCIG18, JCM^{+}21, JFEC13, JOBT05, KK20, KKB00, KRL08, KST^{+}10, KLB^{+}21, KTH^{+}21, KY15, KKK^{+}04b, KC15, KSK^{+}15, KKKY10, KMSTK23, Kra82, KYS^{+}17, KRL^{+}22, LT06, LAA12, LBC^{+}23, LBSP01, LSIC12, LSD^{+}18, LNB13, LO21, LMC^{+}20, LCGH07, LHW^{+}20, LW12, LS15, LBD11, MZH^{+}23, MT99, MPV12, MMR^{+}12, Man69, MDB^{+}20, MBdM^{+}18, MZF^{+}08, MRAP22, MCD^{+}07, MBH^{+}01, MNM06, MM99, Min00, MIW91, Mol04, Mol22, MHCR^{+}12, Mos69, MBD^{+}09]. variability [MKSW^{+}15, MSL^{+}07, MW96, MJWK07, MDR22, NNM^{+}21, NGPH10, NRA17, OACA20, OLH^{+}18, OMS^{+}15, OPH^{+}24, OACB^{+}15, Owe91, PTP^{+}22, PSP^{+}21, PS23, PCSMC12, PMA^{+}14, PELAA18, PL01, PV07, PFW15, PAM^{+}88, PDAM^{+}15, PS98, PZA^{+}15, PCH08a, PHC^{+}19, PLJR22, PWMIM91, PVA24, PVV23, PTI00, QLW10, RWD01, RM93, RvBD^{+}22, Ric01, RL23, RBPGJ^{+}20, RB20, RPRCAG^{+}21, Rog00, RCF^{+}13, RÁSG^{+}13, RGM01, RBE^{+}12, RDP^{+}21, iSIS02, SLM^{+}16, SSQ19, SGWF^{+}19, STB^{+}92, SF15, Sha82, SAM^{+}10, SSTL16, STJ^{+}14, SNZ^{+}20, SAB^{+}21, SEG^{+}22b, ST10, SAT^{+}22, SKP99, Spr08, SPB93, SJ02c, SJM^{+}19, STR01, SOA^{+}23, SP08, SNS^{+}22, TBW09, TvW98, TCL20, TAO05, TS10, TSBS18, iUMY86, VSGC21, VSC01, VGLCS06, VAEP24, Ver91, VEM^{+}21, WM13, WP91, WF06, WF07, WSL20, WCS^{+}23, Whi95, WHI^{+}02, XDG^{+}23]. variability [YIY^{+}04, YBS^{+}01, YS15, YPVP^{+}22, YPM^{+}10, YMK^{+}04, Zav99, ZSI^{+}05, dMM69]. Variable [LTG85, AAM^{+}14, Fla02, Fly10, KAG^{+}19, MRM^{+}14, MAH^{+}15, PVA24, SBD01]. variables [BFP^{+}18, BPSN^{+}21, CSV^{+}07, FACM^{+}23, LLAPG^{+}22, MNT14, NM17, PCC^{+}19, TIOM16, WLKM10]. Variation [CS03, EHFD12, FSVL10, HBH^{+}17, OCH^{+}18, VKJ^{+}23, YKWF21, AC85, BMM01, BLP93, BPSN^{+}21, CMHM18, CCHV^{+}21, CSK^{+}12, DMBB02, DIQJ21, DBR20, EHG^{+}07, FELJ16, FG16, GDM^{+}20, GWM^{+}22, HM98, HRSM08, HVEF09, IIS^{+}17, iIRM^{+}15, IU14, JZZY24, Joh04, KSVT00, KFKO03, KON14, Kli10, LC12, LYS^{+}22, LdCSB^{+}20, MLB^{+}20, MLD^{+}03, MST^{+}23a, MZ14, MPC^{+}17, MCT03, NTU^{+}14, NO14, RBNJ^{+}12, Rou65, SIR^{+}07, SCS^{+}18, Sek88, SH09, SiSI^{+}02, SLBH^{+}19, SEG^{+}22b, SGR^{+}22, TM13, TOiF^{+}12, THP21, VHK03, VHK04, WHT86, WLM^{+}13, WHK23, XRC^{+}15, YYC^{+}18, Yin88, YBPS08, YYhT^{+}17, ZLC^{+}15, ZK06, ZSY^{+}22]. Variational [SO91, GAM98a, GAM98b, GA00, MAB^{+}11c, MAB^{+}11a, MAB^{+}11b, Suk88]. Variations [NF06, SNR^{+}10, SCC14, WHS17, WXH07, ZLS^{+}04, BFB^{+}20, BF01, BHMS09, Car98, CGC^{+}20, CR97, CSC^{+}12, Con87, CLL^{+}18, CRF^{+}10, CW02, DWH^{+}14, DVL^{+}99, DWNN04, Don87, ESD^{+}21, FCN^{+}19, GML^{+}23, GdRGL^{+}01, GHG^{+}24, Her88, Hut87, Igu04, IHT^{+}21, JG07, KP03, KKNT23, KOhL^{+}10, KMWF11, KSP^{+}23, LMPB^{+}16, LSY^{+}14, LO07, LYM12, lLdZQ^{+}22, LZL^{+}22, LXC^{+}22, LSMG01, LSXT01, LLX^{+}21, LS12, MLL^{+}22, MHS^{+}20a, MHS^{+}20b, MTC12, MPSD15, MWO^{+}12, Nag01, ORB^{+}18, PO00, PD15, RSB^{+}01, RBL^{+}19, SKH^{+}23, Sek99, SGS^{+}23, SNZ^{+}20, SPB^{+}02, Sie69, SLPA^{+}20, SEG22a, SQJ^{+}17, TMN^{+}12, VK90, VCM04, WGZZ19, WGM^{+}24, YKS^{+}12, YMA^{+}17, YN20]. varies [UKM^{+}14]. various [Cai95, JYK^{+}14, OSH^{+}96]. vary [BBL^{+}18, CBL^{+}19]. Vector [VSPP14, DDJ^{+}21]. vectors [HKK12]. vehicle [KHM^{+}88]. velocities [DW02]. velocity [CSS^{+}19, LGZW22, NH88a, Nee85, PSP^{+}21, Szu12]. Vema [Zen08]. Venice [Pir87, DFD23, FBB^{+}21]. vent [JP90, JG90, TJ90, VMB^{+}22a, WLP^{+}21]. ventilated [LPW^{+}23]. Ventilation [SASH08, NJCD01, PGC^{+}96, UKM^{+}14, dlPPÁB24]. vents [MSV^{+}14, WLP^{+}21]. Verde [VFCC^{+}22, CCRS20]. Verifications [SBC^{+}24]. vernal [LMA^{+}15]. versus [CMF15, JPBB20, JJA^{+}08, NYL^{+}17, Peñ24, PSM^{+}22, VBAC^{+}21, CSG^{+}15, MPN09]. vertebrate [HSG^{+}15]. Vertical [Ang89, BIST01, DRE^{+}08, DAU22, EiIT^{+}22, GBB^{+}20, HGD22, HSLG11, HKY^{+}11, HCGK11, JC04, MSMR93, ORW^{+}01, PLHLF05, RBF^{+}09, SCHS^{+}24, TMPM^{+}16a, VK92, WYT00, WGG^{+}08, YNMY23, YKNO23, ASC07, Ang79b, ABT^{+}04, AHRT90, Ban64, BEP02, BGM^{+}10, BRG^{+}23, CDL^{+}22, CTF07, CFG07, CF12, EBM^{+}20, EBM^{+}21, GNH19, GSPMAI99, HLM^{+}13, KNSN^{+}09, KGB^{+}23, KVNT20, KSKN21, KHP^{+}18, LGZW22, LFI^{+}13, MGWZ20, NTU^{+}14, NNM^{+}21, NMN08, OACA20, OHC^{+}17, OCH^{+}18, Oll15, OAWAN18, PTM^{+}22, PG13, PNF^{+}21, RS10, RG94, SK17, SBB^{+}22, SDS^{+}22b, UCB^{+}18, UPPS^{+}21, VLCCP14, WCX^{+}21, WSH^{+}22, YYT^{+}14, ZGB^{+}20, ZSH^{+}24, dPGSHL23, BM76, Roe84a, RB84]. vertical-habitat [BGM^{+}10]. very [Nof00]. vessel [BOG20]. vessels [Wüs64]. via [HPB^{+}09]. viability [SJJ^{+}03]. Vicariance [Whi94]. vicinity [Ang89, LPF^{+}20, Nay65, VOT^{+}99]. video [BTNK13]. Vietnamese [LWBD^{+}17]. view [CBGC^{+}08, PAB^{+}21, Pra91, SBM^{+}23, SBB^{+}14, SJ02a, YSD15]. views [RBD^{+}07, SW12]. Vigo [BLT^{+}15]. Vilhjálmsson [Ano13g]. Villefranche [LSIC12]. villosus [BSF^{+}21, CGV13a, CGV13b, MMD^{+}16, RMB^{+}01]. viral [CRC^{+}19, HLSX22]. virtual [WPB05]. viruses [RCC^{+}18]. viscosity [BBS21, BBS23]. visible [VNMS91]. vision [EB08]. Visual [CAO^{+}20]. vital [FCEZ10, LRJ^{+}15]. vitro [GTR01, TGR05, PMMN^{+}22]. vitulina [LAP10, RNL^{+}13]. Vivaldi [PGC^{+}96]. Vladimir [Ano20u]. vocalizing [MSC^{+}15]. void [SEW11, Soh03]. Vol [Ang79a]. volcanics [Nay65]. volcanoes [ZMCD11]. Volume [Jón07, Ano64a, Ano85a, RBS^{+}20, RBS^{+}22, Sek88, VAGMDRS22, ZSH^{+}24]. volumes [Ano65c, Ano65d, Ano69b, Ano73b, Ano85c, Ano86a, Ano87a, Ano89a, Ano92a]. volumetric [BBMR19]. vortex [FHL^{+}24]. vortices [MBS20]. vorticity [CSS^{+}19, KWI20, McD88, Sak86]. Vries [NP00]. Vries-Type [NP00]. vs [CMM^{+}04, CCD^{+}13, ERT^{+}22, HBD^{+}18, LPF^{+}18]. vs. [CPNL07]. vulcanism [AB65]. vulgaris [OÁSG^{+}16]. vulnerability [ECFT20, LHC^{+}21, MDR20, ORPRGIS22, RLR^{+}18]. vulnerable [BECA22, LML^{+}23].

W [BM76, CMJPH^{+}18, FMSBW13, Hen85, HBH^{+}17, JG07, STPHM^{+}23]. waist [FAB^{+}09, GAF15, HM06, Bak06]. Waitz [Dea85]. Wakasa [iUMY86]. wake [HLS^{+}14b]. Wakefield [Ang80]. walled [ZHBW01]. Walleye [PDAM^{+}15, BCB^{+}05, GTS^{+}21, MLPN06, YNM^{+}02]. Warm [IMW^{+}14, Par86, TCN20, YKWF21, BMG^{+}21a, ELW06, FMM^{+}20, LBNBM13, MM01, NC80, TKW06, TKWI08, Ber65b]. warm-core [NC80]. warmer [MBT07]. Warming [DAvD^{+}21, FSAO22, TPTM23, ABP^{+}23, BDTC15, BHK^{+}16, Bel09, BD18, BGL^{+}17, Cia22, CWS^{+}21, ESGP17, GPEV20, GFGGD^{+}23, IHY^{+}01, IG19, JBH20, KV13, KKKY10, KFM^{+}17, LHC^{+}21, MVN^{+}15, MMKS^{+}21, MKS^{+}22, NSE^{+}24, Peñ03a, RL23, SCHBC^{+}22, SEG^{+}22b, SEG22a, SAd^{+}17, SJD10, VDGGD^{+}22, WWW^{+}23, YZX^{+}23, YAK13, dLLdAWL^{+}23]. warning [RAB^{+}11, vdS94c]. Warped [YYK88]. Washington [CCS^{+}21, FB05, HL05, HPHL^{+}05, PHLL05, PLHLF05]. Wasp [GAF15, FAB^{+}09, HM06, Bak06]. wasp-waist [FAB^{+}09, HM06, Bak06]. Water [CDDF11, GR85, HOY^{+}21a, Kat18, KY15, KKK^{+}04b, KTB^{+}99, KMU^{+}12, LH08, MWJ^{+}08, MTK^{+}22, NIC^{+}19, OJB99, PPVG12, Par86, PL18, PGC^{+}96, SDGVE17, Sud86, THM^{+}06, YMK^{+}04, ZSI^{+}05, ZSY^{+}22, ZPC^{+}16, AdAK^{+}18, ÁSÁB^{+}14, ÁBMÁS14, ÁBMÁS15, ALG^{+}21, BW65, Ban96, BNCC15, BvdLA^{+}11, BBB^{+}14, BMC05, BS95, BGV^{+}23, BPM^{+}14, BC88, CDS90, CWZ^{+}20, CSV^{+}07, CPHR98, CVHM^{+}18, CLG^{+}22, CMF15, CR20, DFD23, DOP87, Due77, FVA^{+}19, FZY^{+}23, FC07, FTSF21, FAAF88, dCFK17, FBT^{+}22, FWO15, Fla02, FLDF22, FK86, Gam14, GIPC^{+}15, GLAHH^{+}22, GBB96, GJ00, GTNK21, Gri22, GCS91, GSF^{+}15, GAPM16, HØH^{+}03, Her88, HMS^{+}22, HGH^{+}19, HM08, HG04, Hut95, ISH^{+}04, JJL^{+}19, JSKM02, JG07, JAJS08, Jón07, KON14, KF11, KGdS^{+}08, KG65, KVLA06, KTW^{+}22, LVGH^{+}15, LTG85, LPA92, LRNK99, LCJ^{+}17, LAGM^{+}23, LPW^{+}23]. water [LPBM17, LZG20, LRGV^{+}18, LBF^{+}22, LSW02, LDMH09, MGF^{+}13, MGG22, MRMD^{+}97, MRO^{+}08, MEMP15, MAAS^{+}00, MIN^{+}20, MAFS^{+}22, MRW^{+}14, MHCS^{+}23, MH14, MKS^{+}22, NMK^{+}03, NKK03, NMO^{+}21, NAH^{+}21, ONR^{+}14, OPL^{+}21, ÖÜT93, PTM^{+}22, PMG15, PPKR14, Par63, PTF10, PPCWJ18, PLB^{+}23, PZA^{+}15, PPdS21, PAF^{+}11, PRL^{+}18, PB94, PdMS^{+}13, RCC^{+}18, Rea00, RKM^{+}07, RKK^{+}21, RGPB^{+}23, RLC85, Rud89, Rud15, RKS^{+}15, SGL^{+}17, SCHBC^{+}22, SHP^{+}23, San73, SCB^{+}09, SvWRvB02, SGMP15, SSTL16, SMP^{+}22b, SFAD^{+}90, SBD^{+}07, SWZS^{+}21, SPH^{+}15b, SASH08, SYN^{+}21, SBG^{+}08, TMN^{+}12, TRY^{+}04, TAF^{+}22, Tit20, TSRF14, TVD^{+}99, TDL^{+}17, UNN^{+}14, VGJ^{+}19, VPW01, VKT15, WH20, WCC^{+}20, WWW^{+}23, WHT86, Wen88, WPA^{+}24, WLP^{+}21, YYT^{+}14, YTB^{+}21, YKH^{+}24, YYC^{+}18, Yin88, YRKC08, Yu23, ZAC^{+}23, ZD17, dZTG05, vHMDL14, BRG^{+}23, BF12, Cia14, Men21, NBR^{+}08, OYKK^{+}23]. Water [RBR^{+}23, SPW22, SKH00, Tom81a, TPTM23, YS15, Zen08, ZJZ^{+}21, ZLZ^{+}17]. Water-column [SDGVE17, SSTL16]. Water-level [CDDF11, DOP87]. water-mass [YRKC08]. water/sediment [CDS90, SFAD^{+}90]. Waters [EMU21, HMP^{+}13, ATT^{+}08, BSF95, BCLD^{+}17, BCGN^{+}18, BRH^{+}05, BF01, BT07, BLC23, CWB^{+}22, CB06, CC88, CPC^{+}02, CCA^{+}02, CRC^{+}19, CFG07, CdMS^{+}18, CTI^{+}19, CM14b, DGVGR24, DNNNN16, Dri11, EBS^{+}18, FMM^{+}20, FFA09, FTG^{+}18, FAH^{+}13, GVBV^{+}21, GPA^{+}11, GIHJ23, GEO09, GRdSS^{+}22, HSS^{+}12, HBG^{+}21, HW02, HMPZ11, ILI^{+}12, IHT^{+}21, IU14, KTN14, KV13, KV18, KSS^{+}23, KAG^{+}19, LF12, LC10, LMP22, MOS^{+}13, MH02, MGC^{+}18, MPB^{+}23, MTK^{+}22, MTH^{+}10, ON22, PLN^{+}23, QPR03, RL23, RKM^{+}07, RLSF06, RLSF07, Rot65, SGL^{+}18, SGL^{+}17, SCLG^{+}11, SCHBC^{+}22, STG^{+}18, SCCJ^{+}18, SSTL16, SYN^{+}21, TRP^{+}23, TPTM23, TBW00, TLF^{+}89, TIOM16, VBL^{+}21, VKT15, VMV^{+}23, VEM^{+}21, VSPP14, WFH^{+}22, WWL^{+}22, WFR07, XY21, YT06, YNMY23, Yao88, ZLKO00, ZL01, ZLR^{+}07, dlPHF^{+}15]. watershed [HVRR15]. Wave [CAA^{+}07, CAB^{+}18, His22, BBC^{+}22, BBB^{+}14, BDT^{+}08, BSA06, CSH^{+}23, CB17, FG16, GC14, HLFL23, HWF^{+}21, Iwa23, Kra69, LCZ^{+}24, RFFL21, RHB23, SSB14, UGY^{+}22, VB14, XY21, vHVAT22]. wave-generated [FG16]. wave-induced [BBB^{+}14]. wave-meandering [SSB14]. wavelet [YYhT^{+}17]. Waves [Hut81, WLL06, ABS^{+}20, Arb22, BBB^{+}21, CSLJ03, DWH^{+}14, ESA^{+}13, GXX^{+}22, GLPC23, GCG^{+}14, Ham09, HNSP^{+}19, HHZ^{+}22, HYM^{+}24, Hut87, ITO^{+}14, KWI20, Li14, LOO22, LH08, MZZ^{+}23, NP00, PTM^{+}22, PM85, Ric94, SMFM^{+}21, TCN20, Tho77, VOT^{+}99, VMN08, WLCG23, XDG^{+}23, XY20, XHC^{+}20, YHZ^{+}22]. way [Bak01, BAARB05, PPPdS20]. Weak [Kiv97, ZBRJ23, BH07]. weaknesses [HS07]. weather [BDTC15]. Weathership [AEPW93]. web [AHW^{+}15, BHA^{+}14, BRE^{+}24, BAOC^{+}07, CP10, CPPPEAG22, CSBL^{+}15, CFML22, DKRL22, FMC^{+}15, FTG^{+}11, GAF15, GSVB23, GvOSW11, Hea12, IBW^{+}01, iIYO^{+}10, JE92, LLL^{+}11, LCBN14, MRA^{+}19, MGC^{+}18, NMC^{+}09, OPG^{+}10, PVG^{+}20, Peñ03a, Pow06, RMC^{+}15, RBE^{+}12, SBMB18, SSM90a, ŠGM^{+}18, SHC^{+}06, SHC^{+}07, SDJ14, TNGP22, TR99, TSS^{+}12, TFM03, ZHSMM14, ZBRJ23, dJSL^{+}20]. webs [CLSP17, CBC^{+}06, CW06, Car98, CMF^{+}09, CSC^{+}12, DYO^{+}10, DSB^{+}24, DWC06, HKGH^{+}06, HNL14, LK13, LZF^{+}24, MCH^{+}12, NYL^{+}17, PG10, SBMB18, SRT^{+}18, WRH^{+}06, Was06]. Weddell [CS16, FZY^{+}23, MBB^{+}96, NRS^{+}19, RBS^{+}20, RBS^{+}22, VKDS^{+}18]. weddellii [NRS^{+}19]. weekly [VDGGD^{+}22]. well [SBMB18, SMKK21]. well-documented [SBMB18]. West [HJLLN07, LvBS^{+}24, WLM07, BFH01, BBSN04, CWB^{+}22, CBB^{+}02, CCHM02, DHDM22, FMC^{+}15, FDH20, HMB^{+}86, HSH^{+}19, Man69, MWS^{+}10, MFA^{+}15, MM90, RWOA01, SPW22, STR01, VSGD21, WBH15, CKL^{+}14, GKR20, KGJ^{+}10, LW12, Med87, MFM15, MDR22, Tom81a, WWL^{+}22]. westerlies [SE08, SE09]. Western [BBM^{+}14, CMJPH^{+}18, Con87, DHD^{+}23, EBR^{+}14, HLTB^{+}17, SCD^{+}07, AMFY20, AQVB^{+}10, AAMB^{+}24, ALG^{+}21, ATC^{+}19, BSC^{+}19, BHA^{+}14, BE99, BTS^{+}15a, BOMdP15, BPC^{+}05, CMF^{+}09, CMF11, CF20, CAT^{+}08, CLG^{+}22, DDDT99, DWS^{+}24, DSC^{+}21, DWC06, DK07, EBM^{+}20, ELW06, EHSI12, FTC^{+}16, FKZ^{+}15, FK99, FMT15, GLY23, GDM^{+}20, GdRGC^{+}14, Gri22, HHB^{+}00, INO^{+}24, IHT^{+}21, INT14, JWD^{+}02, KHD22, KAK^{+}22a, KST03, KFH^{+}15, KYS^{+}17, LYM12, LG22, LG23, LSIC12, LPW^{+}23, LSD^{+}15, LBC^{+}15, LWT^{+}20, MT99, MRRC73, MTC12, MAAS^{+}00, MSA^{+}22, Mol04, MMPG07, MNFY21, NEI^{+}22, NMO^{+}21, NHN^{+}21, Nof96, NKK^{+}05, NRA^{+}21, OTNI20, OOTA15, PPdM^{+}12, PELAA18, PJH^{+}15, PHFK14, PHC^{+}19, PdMS^{+}13, PLK14, QCdS^{+}07, RKS01, RTF^{+}05, RBD^{+}07, RCB^{+}20, RDC^{+}21, Rog00, SGL^{+}18, SSH^{+}05, SSQ19, SAM^{+}04, SFK^{+}99, SAB^{+}21, Sme93, SIS^{+}14, SST^{+}17, SKT01, TKSI08]. western [TT05, Tan99, TTMM^{+}17, TSNO05, TKK^{+}05, UPPS^{+}21, VBVYT05, WFH^{+}22, WOW^{+}14, WWN^{+}99, Wu13, YMA^{+}17, YHM^{+}18, YYK^{+}12, YFY05, ZGB^{+}20, ZSI^{+}05, ZJZ^{+}21, BTS^{+}15b, BZD^{+}21, BLCL14, CPG08, Eri65, FGR^{+}06, GR85, GGA^{+}16, GGG^{+}18, GGSM^{+}20, HWPLvW20, HGH^{+}19, LTG85, LLX^{+}21, McI10, MSd^{+}16, NCC^{+}15, PCD^{+}18, PRA^{+}18, PMS^{+}15, PFE10, RMC^{+}15, SGLF^{+}13, SBC^{+}16, SBPGP^{+}23, SAW^{+}15, STW^{+}15, TAW^{+}15, TB15, VYGMM^{+}17, VBAC^{+}21, WOW^{+}14, WSC^{+}21, YWUK15, dPAJ07]. western-boundary [WWN^{+}99]. western-Mediterranean [DDDT99]. westward [MTL05, TCN20]. wet [OAD22]. whale [BMG^{+}19, BMG^{+}21a, BPSN^{+}21, DHD^{+}23, FGL^{+}23, GDL^{+}15, GC09, LBC^{+}23, MHVS19]. Whales [RSB^{+}13, AHGRAL23, CPO^{+}19, CQO^{+}15, CBB^{+}15, GVBV^{+}21, HPB^{+}09, HBK^{+}24]. Where [CPC88, MZK^{+}23]. white [TGJT09]. Whiteaves [Mar20]. whiting [HPB^{+}09, MAFS^{+}22]. Whitsand [USH15b]. Whittard [AHA^{+}16, CHG^{+}18, DJW^{+}18]. Who [THBA19]. whole [MRA^{+}19]. WHP [KMWF11, STPHM^{+}23]. wicked [KN10, KN11]. wide [BMO12, CMPNC^{+}22, MPCNC^{+}19, PKV18, SSM^{+}18]. widely [THM^{+}14]. wider [MEST13]. Widespread [YTB^{+}21]. wild [WBF^{+}21]. Wilkes [AYH^{+}23]. Will [DSAB20]. William [BDB^{+}22, SSB^{+}20b]. Wind [CKL^{+}14, DLM^{+}12, OÁSG^{+}16, OC06, Sek86, AR18, AC85, BBS21, BBS23, BCOL^{+}19, CPC^{+}15, CTA16, CZL^{+}24, CS06, DWH^{+}14, Dav85, Dem09, DDJ^{+}21, FvBA^{+}17, HMRA^{+}03, HGH^{+}19, Iwa23, KM22, LPF^{+}21, MZH^{+}23, MS02, MH14, NNM^{+}21, OACA20, OMS^{+}09, OPL^{+}21, SC23, SBD01, VTGC19, WWZ19, Wun24, XY20, XY21, ZHF^{+}24]. Wind-driven [OÁSG^{+}16, Sek86, BBS21, BBS23, NNM^{+}21, SC23]. wind-forced [DWH^{+}14, LPF^{+}21]. Wind-induced [CKL^{+}14, OC06]. wind-modulated [MH14]. wind-sea [VTGC19]. windier [JTQ^{+}18]. window [HWB^{+}18, XYWY23]. winds [BBB^{+}21, DL17, FAAV^{+}15, GW91, GMR^{+}23, KAG^{+}19, LO85, SFMT14, SNMW10]. wing [SSB14]. wing-sailing [SSB14]. wings [Ric22]. winner [VWDF14]. Winter [ATS01, DSR21, HPHL^{+}05, LVGH^{+}15, MOSN^{+}13, VOJD02a, VOJD02b, YNTS22, AJA^{+}22, ABS^{+}20, AUE^{+}14, BPC^{+}05, BF11, EMU^{+}23, FPJ^{+}15, GHL15, HBH^{+}17, Iwa23, KMS^{+}24, LPW^{+}23, MSd^{+}16, MCKS17, MAFS^{+}22, MEMC05, Rou65, SBC^{+}24, TDGY22, VNH^{+}23]. Winter-spring [HPHL^{+}05]. Winter-to-winter [ATS01]. winters [WDMC02]. wintertime [JC88, Suk88]. wishes [Val99a]. Withdrawn [ZLG17b]. Within [MPV12, Ang84, Arb22, BLC23, Car98, CdD^{+}15, Dom84, DDCE^{+}23, FMM^{+}20, FRK^{+}09, GDI^{+}09, KFC^{+}13, MCG^{+}02, PTPY^{+}23, Pug84, QYF^{+}24, RCS^{+}11, RAB^{+}84, Roe84a, Roe84b, RB84, RJT84, STS^{+}12, WMC^{+}89, WDC^{+}11, WWL^{+}22, ZBRJ23]. Within- [MPV12]. without [Ric22, SPN98]. WOCE [GWB14, GJ00, GA00, MMR^{+}09, STPHM^{+}23]. WOCE-era [MMR^{+}09]. WOCE-WHP [STPHM^{+}23]. Wollast [JW01b, vWM02b]. Women [CSH^{+}23]. wood [FBR^{+}13, GGT^{+}15, JZ19, RVC^{+}13]. wood-boring [RVC^{+}13]. work [CNT03]. Working [Ano94k, Ano03j, WHG^{+}16]. workshop [SWP^{+}13b]. world [BMO12, CBB^{+}15, FC05, LCR^{+}93, Ric22, Val99a, VBC^{+}20, BC16, GMD^{+}22, ISH^{+}04, Lev88, McK15, SAA^{+}15, dZTG05, Ang79a]. world-wide [BMO12]. WRF [LC22]. WRF-FVCOM [LC22].

X [Hof81, Ang88]. XBT [LS20]. xiii [Ano92i]. Xiphias [SYB^{+}15, YGL^{+}10]. xv [Ano86j]. Xylophaga [RVC^{+}13].

Yangtze [ZLR^{+}07, Wen88]. Year [MHA^{+}11, Rud15, AT07, AUE^{+}14, BDB^{+}04, BMG^{+}21a, BPTT19, CMG15, CBB^{+}15, CLdPHL23, DHC^{+}20, DBC^{+}18, DHD^{+}23, FTSF21, HGBG20, HFO^{+}22, LGZ^{+}20, LPF^{+}20, LFCSV^{+}13, MGF^{+}13, MAH^{+}15, RF17, SGO^{+}08, TSFA22, WHT86, Woo18, YFK21, EBD^{+}20]. year-long [BPTT19, LPF^{+}20]. year-round [DHD^{+}23, MGF^{+}13, Woo18]. year-to-year [DBC^{+}18, WHT86]. years [AKH^{+}23, CPB^{+}15, CFC^{+}18, GRMB18, HDZY15, HPHW21, KRL08, KRHS14, LXC^{+}22, RCD^{+}94, SLG^{+}12, SF02, SKRM^{+}95, VOG^{+}08, VSGC21, WR03, WHG^{+}16, XNT^{+}17, dJSL^{+}20]. Yellow [LW13, ITM86, JZZY24, LC16, LCJ^{+}07, LGH^{+}21, MLL^{+}22, MFY^{+}86, Par86, SW12, SNZ^{+}20, TCN20, YKS^{+}12]. yellowfin [HLTB^{+}17, OPG^{+}10]. Yermak [GIHJ23]. yields [Hea12]. yolk [SMPC^{+}12]. young [CSMGS19, GMR^{+}23, RN06]. Yutu [GLY23].

Zdenka [SW81]. Zealand [BSMC15, BWB^{+}09, FBS22, JRR^{+}24, RLR^{+}18]. Zenkevitch [GS19]. Zhe [XLL^{+}20]. Zhe-Min [XLL^{+}20]. Zoarcidae [SM21]. Zonal [KP03, RFPG15, STPHM^{+}23]. zonally [KKNT23]. zonation [BAB^{+}19, GNH19, HM90, LdCSB^{+}20, VMB^{+}22b, WGG^{+}08, ZGB^{+}20]. Zone [Ant09, BHM^{+}15, INI^{+}17, JSLA^{+}21, PHKS17, RMHL09, SLBH^{+}19, SLPA^{+}20, VFCC^{+}22, AMG^{+}16, BAM^{+}09, BAT^{+}98, BA04, BMC17, BBPHG^{+}11, BB14, BC01, DLM^{+}12, EHG^{+}07, EHFD12, FELJ16, GMAMB04, GASV^{+}09, GEO09, GDI^{+}09, GBB^{+}20, GLV12, HLGA07, HLCdP19, HFW^{+}98, HSLG11, ILA21, JCF^{+}23, LGR^{+}02, LQU07, LLGS21, MLHE23, MGH^{+}07, MTH^{+}10, NMO^{+}21, NMN08, PAM^{+}88, PD15, PBS22, PHKS01, RS10, SAM^{+}04, SM65, SMPC^{+}12, SGR^{+}22, SWH^{+}24, TR99, Tom81a, VSGC21, VJJ^{+}22, WMB^{+}21, WP91, WR00, WGG^{+}08, WPH^{+}10, YLL19, BSMC15, Nay65, RMG90, STC10]. Zones [BMN^{+}99, ARELAK24, CSMGS19, EM12, FCN^{+}19, FPY^{+}16, GDM^{+}20, HHMB^{+}09, HBH^{+}17, JS21, JPIP22, KSV08, Kaz17, MYH^{+}22, PRP09, YW22]. ZooCAM [CTP^{+}18]. zoogeography [AS96]. Zooplankton [ASB^{+}08, ACHSH08, Ban64, ESTM13, EHFD12, FÁFL06, HFNG00, LMH^{+}13, MRM^{+}14, MG02, MGF^{+}13, SHD^{+}21, SSV^{+}11, Tur15, VSGD21, WHG^{+}16, YHM^{+}18, dPAJ07, AMFY20, ACE^{+}07, ADS^{+}22, AGL^{+}15, APC^{+}21, BBLD^{+}11, BTNK13, BFB^{+}20, BMO12, BMG13, BMC17, BSFM^{+}12, BASS^{+}20, BPSN^{+}21, CP10, CS16, CP02, DAvD^{+}20, EBW^{+}23, EBS^{+}18, EHG^{+}12, FUOG^{+}16, FKH^{+}13, GRLS14, GWM^{+}22, GSPMAI99, GBB^{+}20, GRD^{+}23, HPC^{+}20, HLCdP19, HGBG20, HWL^{+}20, Igu04, iIYO^{+}10, KP03, KH09, KDB95, KAAK^{+}16, KGJ^{+}10, LO07, Lav09, LEDR^{+}22, LdCSB^{+}20, MBT07, MGE^{+}12, MPV12, MGWZ20, MERB12, MCG^{+}02, MMMWZ23, MFDH22, MMES16, MCG^{+}14, MPC12, Nie07, OWR^{+}07, PG13, PSL87, PGY^{+}22, PCM11, PJH^{+}15, PKF02, PPY87, PO15, Reb02, RWOA01, SCAA07, SCS^{+}18, SPB^{+}12, SPC^{+}23, SNZ^{+}20, SWP^{+}13b, SEG22a, SE92, SDO^{+}14, SOA^{+}23, TSAM^{+}22]. zooplankton [UCB^{+}18, VLUC^{+}07, VLCCP14, VMA^{+}24, WCS^{+}23, WHK23, WHS^{+}23, YNMY23, YPVP^{+}22]. ZooScan [YNMY23]. Zostera [AR18].

References

Amador:2006:AFE

[AALM06] Jorge A. Amador, Eric J. Alfaro, Omar G. Lizano, and Victor O. Magaña. Atmospheric forcing of the eastern tropical Pacific: a review. Progress in Oceanography, 69(2–4):101– 142, May 2006. CODEN POCNA8. ISSN 0079-6611 (print),

              1873-4472 (electronic). URL http://www.sciencedirect.
              com/science/article/pii/S0079661106000292.

Anabalon:2014:SPC

[AAM^{+}14] V. Anabalón, J. Arístegui, C. E. Morales, I. Andrade, M. Benavides, M. A. Correa-Ramírez, M. Espino, O. Ettahiri, S. Hormazabal, A. Makaoui, M. F. Montero, and A. Orbi. The structure of planktonic communities under variable coastal upwelling conditions off Cape Ghir (31^{o}n) in the Canary Current System (NW Africa). Progress in Oceanography, 120(??):320–339, January 2014. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

              tronic). URL http://www.sciencedirect.com/science/
              article/pii/S0079661113002097.

Amargant-Arumi:2024:IDS

[AAMB^{+}24] Martí Amargant-Arumí, Oliver Müller, Yasemin V. Bodur, Iliana-V. Ntinou, Tobias Vonnahme, Philipp Assmy, Doreen Kohlbach, Melissa Chierici, Elizabeth Jones, Lasse M. Olsen, Tatiana M. Tsagaraki, Marit Reigstad, Gunnar Bratbak, and Rolf Gradinger. Interannual differences in sea ice regime in the north-western Barents Sea cause major changes in summer pelagic production and export mechanisms. Progress in Oceanography, 220(??):??, January 2024. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661123002215.

Acevedo-Acosta:2022:ALD

[AAML22] Juan David Acevedo-Acosta and Aída Martínez-López. Anomalously low diatom fluxes during 2009–2010 at Alfonso Basin, Gulf of California. Progress in Oceanography, 206(??): ??, August 2022. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661122000982.

Amaya:2023:EHR

[AASJ23] Dillon J. Amaya, Michael A. Alexander, James D. Scott, and Michael G. Jacox. An evaluation of high-resolution ocean reanalyses in the California current system. Progress in Oceanography, 210(??):??, January 2023. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661122002105.

Arrhenius:1965:NVO

[AB65] Gustaf Arrhenius and Enrico Bonatti. Neptunism and vulcanism in the ocean. Progress in Oceanography, 3(??):7–22, ???? 1965. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661165900054.

Alberic:1990:UAC

[AB90] Patrick Alberic and Jacques Boulegue. Unusual amino compounds in the tissues of Calyptogena phaseoliformis (Japan Trench): Possible link to symbiosis. Progress in Oceanography, 24(??):89–101, ???? 1990. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/007966119090022T.

Aiken:2000:IAM

[AB00] J. Aiken and A. J. Bale. An introduction to the Atlantic Meridional Transect (AMT) Programme. Progress in Oceanography, 45(3–4):251–256, April 2000. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661100000045.

Aristegui:2009:SRE

[ABÁS^{+}09] Javier Arístegui, Eric D. Barton, Xosé A. Álvarez-Salgado, A. Miguel P. Santos, Francisco G. Figueiras, Souad Kifani, Santiago Hernández-León, Evan Mason, Eric Machú, and Hervé Demarcq. Sub-regional ecosystem variability in the Canary Current upwelling. Progress in Oceanography, 83(1–4):33–48, December 2009. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661109001001.

Astraldi:1999:RSC

[ABC^{+}99] M. Astraldi, S. Balopoulos, J. Candela, J. Font, M. Gacic, G. P. Gasparini, B. Manca, A. Theocharis, and J. Tintoré. The role of straits and channels in understanding the characteristics of Mediterranean circulation. Progress in Oceanography, 44(1–3):65–108, August 1999. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S007966119900021X.

Aiken:2017:SER

[ABD^{+}17] Jim Aiken, Robert J. W. Brewin, Francois Dufois, Luca Polimene, Nick J. Hardman-Mountford, Thomas Jackson, Ben Loveday, Silvana Mallor Hoya, Giorgio Dall’Olmo, John Stephens, and Takafumi Hirata. A synthesis of the environmental response of the North and South Atlantic Sub-Tropical Gyres during two decades of AMT. Progress in Oceanography, 158(??):236–254, November 2017. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661116300179.

Alegre:2015:DDJ

[ABE^{+}15] Ana Alegre, Arnaud Bertrand, Marco Espino, Pepe Espinoza, Teobaldo Dioses, Miguel Ñiquen, Iván Navarro, Monique Simier, and Frédéric Ménard. Diet diversity of jack and chub mackerels and ecosystem changes in the northern Humboldt Current system: a long-term study. Progress in Oceanography, 137 (part A)(??):299–313, September 2015. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661115001627.

Alhammoud:2005:SCL

[ABM^{+}05] B. Alhammoud, K. Béranger, L. Mortier, M. Crépon, and I. Dekeyser. Surface circulation of the Levantine Basin: Comparison of model results with observations. Progress in Oceanography, 66(2–4):299–320, August 2005. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661105000753.

Alvarez:2014:MBM

[ÁBMÁS14] M. Álvarez, S. Brea, H. Mercier, and X. A. Álvarez-Salgado. Mineralization of biogenic materials in the water masses of the South Atlantic Ocean. i: Assessment and results of an optimum multiparameter analysis. Progress in Oceanography, 123(??):1– 23, April 2014. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661113002450. See corri-

gendum [ÁBMÁS15].

Alvarez:2015:CMB

[ÁBMÁS15] M. Álvarez, S. Brea, H. Mercier, and X. A. Álvarez-Salgado. Corrigendum: Mineralization of biogenic materials in the water masses of the South Atlantic Ocean. I: Assessment and results of an optimum multiparameter Analysis [Prog. Oceanogr. 123 (2014) 1–23]. Progress in Oceanography, 130(??):249, January 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661114001748. See [ÁBMÁS14].

Auth:2015:AID

[ABP15] Toby D. Auth, Richard D. Brodeur, and Jay O. Peterson. Anomalous ichthyoplankton distributions and concentrations in the northern California Current during the 2010 El Niño and La Niña events. Progress in Oceanography, 137 (part A)(??):103–120, September 2015. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661115001305.

Azarian:2023:CPP

[ABP^{+}23] Clara Azarian, Laurent Bopp, Alice Pietri, Jean-Baptiste Sallée, and Francesco d’Ovidio. Current and projected patterns of warming and marine heatwaves in the Southern Indian Ocean. Progress in Oceanography, 215(??):??, July 2023. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661123000794.

Amol:2020:MCC

[ABS^{+}20] P. Amol, Suchandan Bemal, D. Shankar, V. Jain, V. Thushara, V. Vijith, and P. N. Vinayachandran. Modulation of chlorophyll concentration by downwelling Rossby waves during the winter monsoon in the southeastern Arabian Sea. Progress in Oceanography, 186(??):??, July 2020. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S007966112030104X.

Angel:2007:CCP

[ABSDC07] Martin V. Angel, Kasia Blachowiak-Samolyk, Inna Drapun, and Ruth Castillo. Changes in the composition of planktonic ostracod populations across a range of latitudes in the North-east Atlantic. Progress in Oceanography, 73(1):60–78, April 2007. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661106001601.

Aristegui:2004:VPC

[ABT^{+}04] Javier Arístegui, Eric D. Barton, Paul Tett, María F. Montero, Mercedes García-Muñoz, Gotzon Basterretxea, Anne-Sophie Cussatlegras, Alicia Ojeda, and Demetrio de Armas. Variability in plankton community structure, metabolism, and vertical carbon fluxes along an upwelling filament (Cape Juby, NW Africa). Progress in Oceanography, 62(2–4):95–113, August 2004. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S007966110400093X.

Anderson:1985:ORL

[AC85] D. L. T. Anderson and R. A. Corry. Ocean response to low frequency wind forcing with application to the seasonal variation in the Florida Straits–Gulf Stream transport. Progress in Oceanography, 14(??):7–40, ???? 1985. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0079661185900035.

Aguzzi:2013:SBM

[ACB^{+}13] J. Aguzzi, J. B. Company, N. Bahamon, M. M. Flexas, S. Tecchio, U. Fernandez-Arcaya, J. A. García, A. Mechó, S. Koenig, and M. Canals. Seasonal bathymetric migrations of deepsea fishes and decapod crustaceans in the NW Mediterranean Sea. Progress in Oceanography, 118(??):210–221, November 2013. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661113001389.

Alcaraz:2007:PCZ

[ACE^{+}07] M. Alcaraz, A. Calbet, M. Estrada, C. Marrasé, E. Saiz, and I. Trepat. Physical control of zooplankton communities in the Catalan Sea. Progress in Oceanography, 74(2–3):294–312, August 2007. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661107000614.

Ayon:2008:ZRP

[ACHSH08] Patricia Ayón, Maria I. Criales-Hernandez, Ralf Schwamborn, and Hans-Jürgen Hirche. Zooplankton research off Peru: a review. Progress in Oceanography, 79(2–4):238–255, October 2008. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661108001778.

Aguzzi:2013:LDG

[ACK^{+}13] J. Aguzzi, C. Costa, V. Ketmaier, C. Angelini, F. Antonucci, P. Menesatti, and J. B. Company. Light-dependent genetic and phenotypic differences in the squat lobster Munida tenuimana (Crustacea: Decapoda) along deep continental margins. Progress in Oceanography, 118(??):199–209, November 2013. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661113001249.

Almeida:2018:SCA

[ACL^{+}18] Mariana Almeida, Marina R. Cunha, Nikoloaos Lampadariou, Patricia Esquete, and Joan B. Company. Suprabenthic crustacean assemblages subjected to high-energy hydrodynamic events in the Blanes Canyon and adjacent slope (NW Mediterranean Sea). Progress in Oceanography, 169(??):138–150, December 2018. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661117302252.

Arcos:2001:JMF

[ACN01] Dagoberto F. Arcos, Luis A. Cubillos, and Sergio P. Núñez. The jack mackerel fishery and El Niño 1997–98 effects off Chile. Progress in Oceanography, 49(??):597–617, ???? 2001. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S007966110100043X.

Almeida:2018:INE

[AdAK^{+}18] Lucas Almeida, José Luiz Lima de Azevedo, Rodrigo Kerr, Moacyr Araujo, and Mauricio M. Mata. Impact of the new equation of state of seawater (TEOS-10) on the estimates of water mass mixture and meridional transport in the Atlantic Ocean. Progress in Oceanography, 162(??):13–24, March 2018. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661117302513.

Aparna:2022:SCZ

[ADS^{+}22] S. G. Aparna, D. V. Desai, D. Shankar, A. C. Anil, Shrikant Dora, and R. Khedekar. Seasonal cycle of zooplankton standing stock inferred from ADCP backscatter measurements in the eastern Arabian Sea. Progress in Oceanography, 203(??):??, April–May 2022. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661122000271.

Authier:2018:ECR

[ADV^{+}18] Matthieu Authier, Ghislain Dorémus, Olivier Van Canneyt, Jean-Jacques Boubert, Gérard Gautier, Mathieu Doray, Erwan Duhamel, Jacques Massé, Pierre Petitgas, Vincent Ridoux, and Jérôme Spitz. Exploring change in the relative abundance of marine megafauna in the Bay of Biscay, 2004– 2016. Progress in Oceanography, 166(??):159–167, September 2018. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661117300861.

Auel:2009:DRH

[AE09] Holger Auel and Werner Ekau. Distribution and respiration of the high-latitude pelagic amphipod Themisto gaudichaudi in the Benguela Current in relation to upwelling intensity. Progress in Oceanography, 83(1–4):237–241, December 2009. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661109000871.

Aarflot:2023:NIB

[AEP^{+}23] Johanna M. Aarflot, Elena Eriksen, Irina P. Prokopchuk, Camilla Svensen, Janne E. Søreide, Anette Wold, and Morten D. Skogen. New insights into the Barents Sea Calanus glacialis population dynamics and distribution. Progress in Oceanography, 217(??):??, September–October 2023. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661123001490.

Archer:1993:NHS

[AEPW93] D. Archer, S. Emerson, T. Powell, and C. S. Wong. Numerical hindcasting of sea surface pCO_{2} at Weathership Station Papa. Progress in Oceanography, 32(??):319–351, ???? 1993. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/007966119390019A.

Allen:2010:TBM

[AF10] J. I. Allen and E. A. Fulton. Top-down, bottom-up or middleout? Avoiding extraneous detail and over-generality in marine ecosystem models. Progress in Oceanography, 84(1–2):129–133, January 2010. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661109001499.

Alvarez-Fernandez:2022:PRO

[AFBT^{+}22] S. Alvarez-Fernandez, L. T. Bach, J. Taucher, U. Riebesell, U. Sommer, N. Aberle, C. P. D. Brussaard, and M. Boersma. Plankton responses to ocean acidification: the role of nutrient limitation. Progress in Oceanography, 165(??):11–18, ???? 2022. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661117300113.

Aoyama:2011:CET

[AFH^{+}11] M. Aoyama, M. Fukasawa, K. Hirose, Y. Hamajima, T. Kawano, P. P. Povinec, and J. A. Sanchez-Cabeza. Cross equator transport of ^{137}Cs from North Pacific Ocean to South Pacific Ocean (BEAGLE2003 cruises). Progress in Oceanography, 89(1–4):7–16, April 2011. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661110001552.

Araujo:2022:IOD

[AG22] Carolina C. Araújo and Maria A. Gasalla. Influence of ocean dynamics on the route of argonauts in the Southeastern Brazil Bight. Progress in Oceanography, 209(??):??, December 2022. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661122001653.

Alonso:2018:FST

[AGD^{+}18] Hany Alonso, José P. Granadeiro, Maria P. Dias, Teresa Catry, and Paulo Catry. Fine-scale tracking and diet information of a marine predator reveals the origin and contrasting spatial distribution of prey. Progress in Oceanography, 162(??):1–12, March 2018. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S007966111730215X.

Ariza:2015:MBR

[AGL^{+}15] A. Ariza, J. C. Garijo, J. M. Landeira, F. Bordes, and S. Hernández-León. Migrant biomass and respiratory carbon flux by zooplankton and micronekton in the subtropical northeast Atlantic Ocean (Canary Islands). Progress in Oceanography, 134(??):330–342, May 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661115000440.

Anderson:2010:IGF

[AGS10] Thomas R. Anderson, Wendy C. Gentleman, and Bablu Sinha. Influence of grazing formulations on the emergent properties of a complex ecosystem model in a global ocean general circulation model. Progress in Oceanography, 87(1–4):201–213, October 2010. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661110000972.

Andrews:1980:USB

[AH80] W. R. H. Andrews and L. Hutchings. Upwelling in the Southern Benguela Current. Progress in Oceanography, 9 (1):1–8, IN1–IN2, 9–76, IN3–IN4, 77–81, ???? 1980. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0079661180900154.

Ainley:2010:TBF

[AH10] David G. Ainley and K. David Hyrenbach. Top-down and bottom-up factors affecting seabird population trends in the California Current System (1985–2006). Progress in Oceanography, 84(3–4):242–254, March 2010. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661109001608.

Anderson:2015:CMP

[AH15] Robert F. Anderson and Christopher T. Hayes. Characterizing marine particles and their impact on biogeochemical cycles in the GEOTRACES program. Progress in Oceanography, 133(??):1–5, April 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661114001992.

Amaro:2016:WCC

[AHA^{+}16] T. Amaro, V. A. I. Huvenne, A. L. Allcock, T. Aslam, J. S. Davies, R. Danovaro, H. C. De Stigter, G. C. A. Duineveld, C. Gambi, A. J. Gooday, L. M. Gunton, R. Hall, K. L. Howell, J. Ingels, K. Kiriakoulakis, C. E. Kershaw, M. S. S. Lavaleye, K. Robert, H. Stewart, D. Van Rooij, M. White, and A. M. Wilson. The whittard canyon — a case study of submarine canyon processes. Progress in Oceanography, 146(??):38–57, August 2016. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661115300549.

Araujo:2013:BEM

[AHC^{+}13] H. Andres Araujo, Carrie Holt, Janelle M. R. Curtis, R. I. Perry, James R. Irvine, and Catherine G. J. Michielsens. Building an ecosystem model using mismatched and fragmented data: a probabilistic network of early marine survival for Coho salmon Oncorhynchus kisutch in the Strait of Georgia. Progress in Oceanography, 115(??):41–52, August 2013. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661113000736.

Aslam:2018:ITD

[AHD18] Tahmeena Aslam, Rob A. Hall, and Stephen R. Dye. Internal tides in a dendritic submarine canyon. Progress in Oceanography, 169(??):20–32, December 2018. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661117302495.

Huh:1992:AIS

[aHFS92] Chih an Huh, Bruce P. Finney, and Janet K. Stull. Anthropogenic inputs of several heavy metals to nearshore basins off Los Angeles. Progress in Oceanography, 30(1–4):335–351, ???? 1992. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/007966119290018U.

Acevedo:2023:EPR

[AHGRAL23] Jorge Acevedo, Daniela Haro, Eduardo González-Rodríguez, and Anelio Aguayo-Lobo. Estimates of prey requirements and initial examination of the primary productivity to support the Magellan Strait humpback whales. Progress in Oceanography, 211(??):??, February 2023. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661123000149.

Akimova:2019:SDP

[AHP19] Anna Akimova, Marc Hufnagl, and Myron A. Peck. Spatiotemporal dynamics of predators and survival of marine fish early life stages: Atlantic cod (Gadus morhua) in the North Sea. Progress in Oceanography, 176(??):??, September 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661118302490.

Attrill:1990:DRD

[AHRT90] M. J. Attrill, R. G. Hartnoll, A. L. Rice, and M. H. Thurston. A depth-related distribution of the red crab, Geryon trispinosus (Herbst) [= G. tridens Krøyer]: indications of vertical migration. Progress in Oceanography, 24(??):197–206, ???? 1990. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0079661190900306.

Aarflot:2022:TPC

[AHSS22] Johanna M. Aarflot, Solfrid S. Hjøllo, Espen Strand, and Morten D. Skogen. Transportation and predation control structures the distribution of a key calanoid in the Nordic Seas. Progress in Oceanography, 202(??):??, March 2022. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661122000234.

Auad:1999:MHB

[AHW99] Guillermo Auad, Myrl C. Hendershott, and Clinton D. Winant. Mass and heat balances in the Santa Barbara Channel: estimation, description and forcing. Progress in Oceanography, 43(1):111–155, January 1999. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661199000063.

Atkinson:2015:QRP

[AHW^{+}15] Angus Atkinson, Rachel A. Harmer, Claire E. Widdicombe, Andrea J. McEvoy, Tim J. Smyth, Denise G. Cummings, Paul J. Somerfield, Jacqueline L. Maud, and Kristian Mc-Conville. Questioning the role of phenology shifts and trophic mismatching in a planktonic food web. Progress in Oceanography, 137(??):498–512, September 2015. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661115000907.

Albaina:2015:RTP

[AIA^{+}15] Aitor Albaina, Xabier Irigoien, Unai Aldalur, Unai Cotano, María Santos, Guillermo Boyra, and Andone Estonba. A real-time PCR assay to estimate invertebrate and fish predation on anchovy eggs in the Bay of Biscay. Progress in Oceanography, 131(??):82–99, February 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661114002043.

Ayata:2018:RMB

[AIA^{+}18] Sakina-Dorothée Ayata, Jean-Olivier Irisson, Anaïs Aubert, Léo Berline, Jean-Claude Dutay, Nicolas Mayot, Anne-Elise Nieblas, Fabrizio D’Ortenzio, Julien Palmiéri, Gabriel Reygondeau, Vincent Rossi, and Cécile Guieu. Regionalisation of the Mediterranean basin, a MERMEX synthesis. Progress in Oceanography, 163(??):7–20, April 2018. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661117300393.

Aguilar-Islas:2007:MMS

[AIHB^{+}07] Ana M. Aguilar-Islas, Matthew P. Hurst, Kristen N. Buck, Bettina Sohst, Geoffrey J. Smith, Maeve C. Lohan, and Kenneth W. Bruland. Micro- and macronutrients in the southeastern Bering Sea: Insight into iron-replete and irondepleted regimes. Progress in Oceanography, 73(2):99–126, May 2007. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S007966110700016X.

Albin:2022:DPS

[AJA^{+}22] K. J. Albin, R. Jyothibabu, K. T. Alok, S. Santhikrishnan, S. Sarath, V. Sudheesh, C. K. Sherin, K. K. Balachandran, C. R. Asha Devi, and G. V. M. Gupta. Distinctive phytoplankton size responses to the nutrient enrichment of coastal upwelling and winter convection in the eastern Arabian Sea. Progress in Oceanography, 203(??):??, April–May 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661122000404.

Angles:2019:ICU

[AJHC19] Sílvia Anglès, Antoni Jordi, Darren W. Henrichs, and Lisa Campbell. Influence of coastal upwelling and river discharge on the phytoplankton community composition in the northwestern Gulf of Mexico. Progress in Oceanography, 173(??): 26–36, April 2019. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661118302921.

Araujo:2002:GTN

[AJV^{+}02] M. F. Araújo, J.-M. Jouanneau, P. Valério, T. Barbosa, A. Gouveia, O. Weber, A. Oliveira, A. Rodrigues, and J. M. A. Dias. Geochemical tracers of northern Portuguese estuarine sediments on the shelf. Progress in Oceanography, 52(??):277– 297, ???? 2002. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661102000113.

Andersen:1997:MSO

[AK97] Ole Baltazar Andersen and Per Knudsen. Multi-satellite ocean tide modelling — the K_{1} constituent. Progress in Oceanography, 40(??):197–216, ???? 1997. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661198000020.

Amemou:2020:ALM

[AKAL20] H. Amemou, V. Koné, A. Aman, and C. Lett. Assessment of a Lagrangian model using trajectories of oceanographic drifters and fishing devices in the Tropical Atlantic Ocean.

Progress in Oceanography, 188(??):??, October 2020. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661120301658.

Assmy:2023:SPD

[AKH^{+}23] Philipp Assmy, Ane Cecilie Kvernvik, Haakon Hop, Clara J. M. Hoppe, Melissa Chierici, Divya David T., Pedro Duarte, Agneta Fransson, Laura M. García, Weronika Patuła, Sławomir Kwaśniewski, Marion Maturilli, Olga Pavlova, Agnieszka Tatarek, Jozef M. Wiktor, Anette Wold, Klara K. E. Wolf, and Allison Bailey. Seasonal plankton dynamics in Kongsfjorden during two years of contrasting environmental conditions. Progress in Oceanography, 213(??):??, April–May 2023. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661123000393.

Atkinson:1987:HO

[ALBP87] Larry P. Atkinson, Thomas N. Lee, Jackson O. Blanton, and Gustav-Adolf Paffenhöfer. Hydrographic observations. Progress in Oceanography, 19(3–4):231–266, ???? 1987. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0079661187900103.

Alvarez:2022:PDE

[ÁLC22] Eva Álvarez, Paolo Lazzari, and Gianpiero Cossarini. Phytoplankton diversity emerging from chromatic adaptation and competition for light. Progress in Oceanography, 204(??): ??, June 2022. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661122000507.

Angiolillo:2021:NRS

[ALG^{+}21] Michela Angiolillo, Gabriele La Mesa, Michela Giusti, Eva Salvati, Bianca Di Lorenzo, Lorenzo Rossi, Simonepietro Canese, and Leonardo Tunesi. New records of scleractinian cold-water coral (CWC) assemblages in the southern Tyrrhenian Sea (western Mediterranean Sea): Human impacts and conservation prospects. Progress in Oceanography, 197(??):??, September–October 2021. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661121001415.

Arnoldi:2023:MTM

[ALM^{+}23] Natalie S. Arnoldi, Steven Y. Litvin, Daniel J. Madigan, Fiorenza Micheli, and Aaron Carlisle. Multi-taxa marine isoscapes provide insight into large-scale trophic dynamics in the North Pacific. Progress in Oceanography, 213(??):??, April–May 2023. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661123000484.

Ayata:2010:HDC

[ALT10] Sakina-Dorothée Ayata, Pascal Lazure, and Éric Thiébaut. How does the connectivity between populations mediate range limits of marine invertebrates? A case study of larval dispersal between the Bay of Biscay and the English Channel (North-East Atlantic). Progress in Oceanography, 87(1–4):18– 36, October 2010. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661110001345.

Ando:2021:HCS

[ALV^{+}21] Kentaro Ando, Xiaopei Lin, Cesar Villanoy, Mikhail Danchenkov, Jae-Hak Lee, Hui-Jun He, Qian Liu, Yang Liu, Vyacheslav Lobanov, Xiao-Lin Ma, Hanung Agus Mulyadi, Akira Nagano, Jin-Ling Ren, Augy Syahailatua, Yongjun Tian, Lunyu Wu, Jing Zhang, Linlin Zhang, Meixun Zhao, Jingjing Zheng, Shuyang Ma, and Wenxi Zhu. Half-century of scientific advancements since the cooperative study of the Kuroshio and Adjacent Regions (CSK) Programme — need for a new Kuroshio Research. Progress in Oceanography, 193(??):??, April–May 2021. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661121000033.

Anderson:2010:DEM

[AM10] Thomas R. Anderson and Aditee Mitra. Dysfunctionality in ecosystem models: an underrated pitfall? Progress in Oceanography, 84(1–2):66–68, January 2010. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661109001402.

Adrianov:2019:EUS

[AM19] Andrey V. Adrianov and Anastassya S. Maiorova. Echinoderes ultraabyssalis sp. nov. from the Kuril–Kamchatka Trench — the first hadal representative of the Kinorhyncha (Kinorhyncha: Cyclorhagida). Progress in Oceanography, 178(??):??, November 2019. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661119301740.

Anabalon:2007:CNM

[AMEV07] V. Anabalón, C. E. Morales, R. Escribano, and M. Angélica Varas. The contribution of nano- and micro-planktonic assemblages in the surface layer (0–30m) under different hydrographic conditions in the upwelling area off Concepción, central Chile. Progress in Oceanography, 75(3):396–414, November 2007. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661107001590.

Abe:2020:RSI

[AMFY20] Yoshiyuki Abe, Kohei Matsuno, Amane Fujiwara, and Atsushi Yamaguchi. Review of spatial and inter-annual changes in the zooplankton community structure in the western Arctic Ocean during summers of 2008–2017. Progress in Oceanography, 186 (??):??, July 2020. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661120301300.

Anabalon:2016:MPC

[AMG^{+}16] V. Anabalón, C. E. Morales, H. E. González, E. Menschel, W. Schneider, S. Hormazabal, L. Valencia, and R. Escribano. Micro-phytoplankton community structure in the coastal upwelling zone off Concepción (central Chile): Annual and inter-annual fluctuations in a highly dynamic environment. Progress in Oceanography, 149(??):174–188, December 2016. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661115300409.

Abe:2023:BAK

[AMY^{+}23] Koki Abe, Ryuichi Matsukura, Natsuki Yamamoto, Kazuo Amakasu, Reiko Nagata, and Hiroto Murase. Biomass of Antarctic krill (Euphausia superba) in the eastern Indian sector of the Southern Ocean (80–150^{o}e) in the 2018–19 austral summer. Progress in Oceanography, 218(??):??, November 2023. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661123001507.

Alheit:2004:RSH

[AN04] Jürgen Alheit and Miguel Niquen. Regime shifts in the Humboldt Current ecosystem. Progress in Oceanography, 60 (2–4):201–222, February 2004. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S007966110400028X.

Aangstrom:1965:SCT

[Ång65] Anders A. Ångström. The solar constant and the temperature of the earth. Progress in Oceanography, 3(??):1–5, ???? 1965. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0079661165900042.

Angel:1979:BRW

[Ang79a] M. V. Angel. Book review: World Ocean Atlas, Vol. 2 — Atlantic and Indian Oceans: Edited by Sergei G. Gorshkov (with an introductory text and gazetteer translated into English and key for non-Russian speaking users). Pergamon Press. 460 × 330 mm, 352 pp. (1979) U.S. $300.00, £150.00, ISBN 0-08-021953-5. Progress in Oceanography, 8(4): 281, ???? 1979. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/007966117990003X.

Angel:1979:SAH

[Ang79b] Martin V. Angel. Studies on Atlantic halocyprid ostracods: their vertical distributions and community structure in the central gyre region along latitude 30^{o}N from off Africa to Bermuda. Progress in Oceanography, 8(1–2):3–124, ???? 1979. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0079661179900090.

Angel:1980:BRI

[Ang80] Martin Angel. Book review: Industrialised embayments and their environmental problems; a case study of Swansea Bay: Edited by M. B. Collins, F. T. Banner, P. A. Tyler, S. J. Wakefield and A. E. James, Pergamon Press (1980) $84.00 (£35.00) 632 pp., 266 illustrations. Progress in Oceanography, 9(4):245– 247, ???? 1980. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661180900038.

Angel:1984:DMD

[Ang84] M. V. Angel. The diel migrations and distributions within a mesopelagic community in the North East Atlantic. 3. Planktonic ostracods, a stable component in the community. Progress in Oceanography, 13(3–4):319–351, ???? 1984. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0079661184900120.

Angel:1985:E

[Ang85] M. V. Angel. Editorial. Progress in Oceanography, 15(1): i, ???? 1985. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661185900357.

Angel:1986:AOS

[Ang86] M. V. Angel. Acta Oceanologica Sinica. Progress in Oceanography, 16(1):iii, ???? 1986. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0079661186900352.

Angel:1987:E

[Ang87] Martin V. Angel. Editorial. Progress in Oceanography, 19(3– 4):i, ???? 1987. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661187900085.

Angel:1988:BRO

[Ang88] Martin V. Angel. Book review: Oceanography from space, advances in space research: 7(2). Edited by J. F. R. Gower. Pergamon Press. 141pp. (1988) US$50.00, £30.00 ISBN 0-08-036649-X. Progress in Oceanography, 20(2):i–ii, ???? 1988.

CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/007966118890050X.

Angel:1989:VPP

[Ang89] M. V. Angel. Vertical profiles of pelagic communities in the vicinity of the Azores Front and their implications to deep ocean ecology. Progress in Oceanography, 22(1):1–46, ???? 1989. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0079661189900098.

Angel:1994:E

[Ang94] Martin V. Angel. Editorial. Progress in Oceanography, 34(2–3): 79–80, ???? 1994. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661194900019.

Abe:2021:SRN

[ANH21] Hiroto Abe, Daiki Nomura, and Toru Hirawake. Salinity regime of the northwestern Bering Sea shelf. Progress in Oceanography, 198(??):??, November 2021. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661121001592.

Accolla:2015:AFR

[ANMP15] Chiara Accolla, David Nerini, Olivier Maury, and Jean-Christophe Poggiale. Analysis of functional response in presence of schooling phenomena: an IBM approach. Progress in Oceanography, 134(??):232–243, May 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661115000361.

Anonymous:1963:EB

[Ano63a] Anonymous. Editorial Board. Progress in Oceanography, 1(??): v, ???? 1963. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661163900016.

Anonymous:1963:NI

[Ano63b] Anonymous. Name index. Progress in Oceanography, 1 (??):375–383, ???? 1963. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0079661163900089.

Anonymous:1963:P

[Ano63c] Anonymous. Pages 3–383 (1963). Progress in Oceanography, 1(??):3–383, ???? 1963. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1963:SI

[Ano63d] Anonymous. Subject index. Progress in Oceanography, 1(??): 381, ???? 1963. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661163900090.

Anonymous:1964:CV

[Ano64a] Anonymous. Contents of volume 1. Progress in Oceanography, 2(??):272, ???? 1964. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0079661164900084.

Anonymous:1964:EB

[Ano64b] Anonymous. Editorial Board. Progress in Oceanography, 2(??): iii, ???? 1964. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661164900011.

Anonymous:1964:NI

[Ano64c] Anonymous. Name index. Progress in Oceanography, 2 (??):257–262, ???? 1964. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0079661164900060.

Anonymous:1964:P

[Ano64d] Anonymous. Pages 1–272 (1964). Progress in Oceanography, 2(??):1–272, ???? 1964. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1964:SI

[Ano64e] Anonymous. Subject index. Progress in Oceanography, 2 (??):263–271, ???? 1964. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0079661164900072.

Anonymous:1965:AIa

[Ano65a] Anonymous. Author index. Progress in Oceanography, 3 (??):391–398, ???? 1965. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0079661165900340.

Anonymous:1965:AIb

[Ano65b] Anonymous. Author index. Progress in Oceanography, 4 (??):323–328, September 2, 1965. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0079661165900583.

Anonymous:1965:CPVa

[Ano65c] Anonymous. Contents of previous volumes. Progress in Oceanography, 3(??):409, IN5, IN7, IN9, IN11, IN13, IN15, ???? 1965. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661165900364.

Anonymous:1965:CPVb

[Ano65d] Anonymous. Contents of previous volumes. Progress in Oceanography, 4(??):345–346, September 2, 1965. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0079661165900601.

Anonymous:1965:HPA

[Ano65e] Anonymous. Hans Pettersson — an appreciation. Progress in Oceanography, 3(??):xi–xiii, ???? 1965. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0079661165900029.

Anonymous:1965:HPB

[Ano65f] Anonymous. Hans Pettersson’s bibliography. Progress in Oceanography, 3(??):xv–xxiii, ???? 1965. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0079661165900030.

Anonymous:1965:I

[Ano65g] Anonymous. Introduction. Progress in Oceanography, 4 (??):ix–xi, September 2, 1965. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0079661165900376.

Anonymous:1965:P

[Ano65h] Anonymous. Pages 1–409 (1965). Progress in Oceanography, 3(??):1–409, ???? 1965. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1965:PS

[Ano65i] Anonymous. Pages 3–346 (2 September 1965). Progress in Oceanography, 4(??):3–346, September 2, 1965. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1965:SIa

[Ano65j] Anonymous. Subject index. Progress in Oceanography, 3 (??):399–407, ???? 1965. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0079661165900352.

Anonymous:1965:SIb

[Ano65k] Anonymous. Subject index. Progress in Oceanography, 4 (??):329–344, September 2, 1965. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0079661165900595.

Anonymous:1969:AI

[Ano69a] Anonymous. Author index. Progress in Oceanography, 5 (??):185–186, ???? 1969. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0079661169900408.

Anonymous:1969:CPV

[Ano69b] Anonymous. Contents of previous volumes. Progress in Oceanography, 5(??):193–196, ???? 1969. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0079661169900421.

Anonymous:1969:P

[Ano69c] Anonymous. Pages 3–196 (1969). Progress in Oceanography, 5(??):3–196, ???? 1969. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1969:SI

[Ano69d] Anonymous. Subject index. Progress in Oceanography, 5 (??):187–191, ???? 1969. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/007966116990041X.

Anonymous:1973:AI

[Ano73a] Anonymous. Author index. Progress in Oceanography, 6 (??):181–183, ???? 1973. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0079661173900086.

Anonymous:1973:CPV

[Ano73b] Anonymous. Content of previous volumes. Progress in Oceanography, 6(??):189–193, ???? 1973. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0079661173900104.

Anonymous:1973:EB

[Ano73c] Anonymous. Editorial Board. Progress in Oceanography, 6(??): ifc, ???? 1973. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661173900013.

Anonymous:1973:P

[Ano73d] Anonymous. Pages 1–193 (1973). Progress in Oceanography, 6(??):1–193, ???? 1973. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1973:SI

[Ano73e] Anonymous. Subject index. Progress in Oceanography, 6 (??):185–188, ???? 1973. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0079661173900098.

Anonymous:1976:EB

[Ano76a] Anonymous. Editorial Board. Progress in Oceanography, 7(1): ifc, ???? 1976. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/007966117690001X.

Anonymous:1976:Pa

[Ano76b] Anonymous. Pages 1–58 (1976). Progress in Oceanography, 7 (1):??, ???? 1976. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1976:Pb

[Ano76c] Anonymous. Pages 59–90 (1976). Progress in Oceanography, 7 (2):??, ???? 1976. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1977:Pb

[Ano77a] Anonymous. Pages 135–162 (1977). Progress in Oceanography, 7(4):??, ???? 1977. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1977:Pa

[Ano77b] Anonymous. Pages 91–133 (1977). Progress in Oceanography, 7 (3):??, ???? 1977. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1978:P

[Ano78] Anonymous. Pages 163–239 (1978). Progress in Oceanography, 7(5–6):??, ???? 1978. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1979:EB

[Ano79a] Anonymous. Editorial Board. Progress in Oceanography, 8(1– 2):ifc, ???? 1979. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661179900077.

Anonymous:1979:Pa

[Ano79b] Anonymous. Pages 1–124 (1979). Progress in Oceanography, 8(1–2):??, ???? 1979. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1979:Pb

[Ano79c] Anonymous. Pages 125–189 (1979). Progress in Oceanography, 8(3):??, ???? 1979. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1979:Pc

[Ano79d] Anonymous. Pages 191–281 (1979). Progress in Oceanography, 8(4):??, ???? 1979. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1980:A

[Ano80a] Anonymous. Announcement. Progress in Oceanography, 9(3): i, ???? 1980. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661180900099.

Anonymous:1980:EB

[Ano80b] Anonymous. Editorial Board. Progress in Oceanography, 9(1): ifc, ???? 1980. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661180900142.

Anonymous:1980:Pa

[Ano80c] Anonymous. Pages 1–81 (1980). Progress in Oceanography, 9 (1):??, ???? 1980. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1980:Pc

[Ano80d] Anonymous. Pages 133–183 (1980). Progress in Oceanography, 9(3):??, ???? 1980. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1980:Pd

[Ano80e] Anonymous. Pages 185–247 (1980). Progress in Oceanography, 9(4):??, ???? 1980. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1980:Pb

[Ano80f] Anonymous. Pages 83–131 (1980). Progress in Oceanography, 9 (2):??, ???? 1980. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1981:EB

[Ano81a] Anonymous. Editorial Board. Progress in Oceanography, 10(1): ifc, ???? 1981. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661181900057.

Anonymous:1981:Pa

[Ano81b] Anonymous. Pages 1–70 (1981). Progress in Oceanography, 10 (1):??, ???? 1981. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1981:Pc

[Ano81c] Anonymous. Pages 147–192 (1981). Progress in Oceanography, 10(3):??, ???? 1981. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1981:Pd

[Ano81d] Anonymous. Pages 193–226 (1981). Progress in Oceanography, 10(4):??, ???? 1981. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1981:Pb

[Ano81e] Anonymous. Pages 71–145 (1981). Progress in Oceanography, 10(2):??, ???? 1981. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1982:A

[Ano82a] Anonymous. Announcement. Progress in Oceanography, 11 (1):I, ???? 1982. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661182900118.

Anonymous:1982:EB

[Ano82b] Anonymous. Editorial Board. Progress in Oceanography, 11(1): ifc, ???? 1982. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/007966118290009X.

Anonymous:1982:Pa

[Ano82c] Anonymous. Pages 1–59 (1982). Progress in Oceanography, 11 (1):??, ???? 1982. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1982:Pc

[Ano82d] Anonymous. Pages 219–305 (1982). Progress in Oceanography, 11(3):??, ???? 1982. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1982:Pd

[Ano82e] Anonymous. Pages 307–392 (1982). Progress in Oceanography, 11(4):??, ???? 1982. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1982:Pb

[Ano82f] Anonymous. Pages 61–218 (1982). Progress in Oceanography, 11(2):??, ???? 1982. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1983:EB

[Ano83a] Anonymous. Editorial Board. Progress in Oceanography, 12(1): ifc, ???? 1983. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661183900058.

Anonymous:1983:Pa

[Ano83b] Anonymous. Pages 1–145 (1983). Progress in Oceanography, 12 (1):??, ???? 1983. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1983:Pb

[Ano83c] Anonymous. Pages 147–219 (1983). Progress in Oceanography, 12(2):??, ???? 1983. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1983:Pc

[Ano83d] Anonymous. Pages 221–381 (1983). Progress in Oceanography, 12(3):??, ???? 1983. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1983:Pd

[Ano83e] Anonymous. Pages 383–459 (1983). Progress in Oceanography, 12(4):??, ???? 1983. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1984:EB

[Ano84a] Anonymous. Editorial Board. Progress in Oceanography, 13(1): ifc, ???? 1984. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661184900016.

Anonymous:1984:Pa

[Ano84b] Anonymous. Pages 1–111 (1984). Progress in Oceanography, 13 (1):??, ???? 1984. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1984:Pb

[Ano84c] Anonymous. Pages 113–243 (1984). Progress in Oceanography, 13(2):??, ???? 1984. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1984:Pc

[Ano84d] Anonymous. Pages 245–511 (1984). Progress in Oceanography, 13(3–4):??, ???? 1984. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1985:AVJ

[Ano85a] Anonymous. An anniversary volume for John Swallow. Progress in Oceanography, 14(??):1–5, ???? 1985. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0079661185900023.

Anonymous:1985:AI

[Ano85b] Anonymous. Author index. Progress in Oceanography, 14(??): 561, ???? 1985. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661185900291.

Anonymous:1985:CPV

[Ano85c] Anonymous. Contents of previous volumes. Progress in Oceanography, 14(??):563–569, ???? 1985. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0079661185900308.

Anonymous:1985:EBa

[Ano85d] Anonymous. Editorial Board. Progress in Oceanography, 14 (??):iii, ???? 1985. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661185900011.

Anonymous:1985:EBb

[Ano85e] Anonymous. Editorial Board. Progress in Oceanography, 15(1): ifc, ???? 1985. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661185900345.

Anonymous:1985:Pa

[Ano85f] Anonymous. Pages 1–569 (1985). Progress in Oceanography, 14(??):1–569, ???? 1985. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1985:Pb

[Ano85g] Anonymous. Pages 1–69 (1985). Progress in Oceanography, 15 (1):??, ???? 1985. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1985:Pd

[Ano85h] Anonymous. Pages 157–216 (1985). Progress in Oceanography, 15(3):??, ???? 1985. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1985:Pe

[Ano85i] Anonymous. Pages 217–316 (1985). Progress in Oceanography, 15(4):??, ???? 1985. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1985:Pc

[Ano85j] Anonymous. Pages 71–156 (1985). Progress in Oceanography, 15(2):??, ???? 1985. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1985:SI

[Ano85k] Anonymous. Subject index. Progress in Oceanography, 14 (??):553–559, ???? 1985. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/007966118590028X.

Anonymous:1986:CSI

[Ano86a] Anonymous. Contents and subject index and contents of some previous volumes. Progress in Oceanography, 17(??):iii–xv, ???? 1986. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661190900454.

Anonymous:1986:EBa

[Ano86b] Anonymous. Editorial Board. Progress in Oceanography, 16(1): ifc, ???? 1986. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661186900339.

Anonymous:1986:EBb

[Ano86c] Anonymous. Editorial Board. Progress in Oceanography, 17(1– 2):ifc, ???? 1986. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661186900224.

Anonymous:1986:Pe

[Ano86d] Anonymous. Pages 1–162 (1986). Progress in Oceanography, 17(1–2):??, ???? 1986. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1986:Pa

[Ano86e] Anonymous. Pages 1–61 (1986). Progress in Oceanography, 16 (1):??, ???? 1986. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1986:Pc

[Ano86f] Anonymous. Pages 113–194 (1986). Progress in Oceanography, 16(3):??, ???? 1986. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1986:Pf

[Ano86g] Anonymous. Pages 163–399 (1986). Progress in Oceanography, 17(3–4):??, ???? 1986. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1986:Pd

[Ano86h] Anonymous. Pages 195–267 (1986). Progress in Oceanography, 16(4):??, ???? 1986. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1986:Pb

[Ano86i] Anonymous. Pages 63–112 (1986). Progress in Oceanography, 16(2):??, ???? 1986. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1986:PIX

[Ano86j] Anonymous. Pages iii-xv (1986). Progress in Oceanography, 17(??):iii–+, ???? 1986. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1987:CSP

[Ano87a] Anonymous. Contents of some previous volumes. Progress in Oceanography, 18(1–4):351–358, ???? 1987. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0079661187900401.

Anonymous:1987:EBa

[Ano87b] Anonymous. Editorial Board. Progress in Oceanography, 18(1– 4):ifc, ???? 1987. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661187900218.

Anonymous:1987:EBb

[Ano87c] Anonymous. Editorial Board. Progress in Oceanography, 19(1): ifc, ???? 1987. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661187900012.

Anonymous:1987:EC

[Ano87d] Anonymous. Editorial comment. Progress in Oceanography, 19 (2):i–ii, ???? 1987. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661187900048.

Anonymous:1987:Pa

[Ano87e] Anonymous. Pages 1–358 (1987). Progress in Oceanography, 18(1–4):??, ???? 1987. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1987:Pb

[Ano87f] Anonymous. Pages 1–85 (1987). Progress in Oceanography, 19 (1):??, ???? 1987. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1987:Pd

[Ano87g] Anonymous. Pages 221–441 (1987). Progress in Oceanography, 19(3–4):??, ???? 1987. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1987:Pc

[Ano87h] Anonymous. Pages 87–220 (1987). Progress in Oceanography, 19(2):??, ???? 1987. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1987:SI

[Ano87i] Anonymous. Subject index. Progress in Oceanography, 18 (1–4):347–350, ???? 1987. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0079661187900395.

Anonymous:1988:EBa

[Ano88a] Anonymous. Editorial Board. Progress in Oceanography, 20(1): ifc, ???? 1988. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661188900511.

Anonymous:1988:EBb

[Ano88b] Anonymous. Editorial Board. Progress in Oceanography, 21(1): ifc, ???? 1988. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661188900547.

Anonymous:1988:Pe

[Ano88c] Anonymous. Pages 1–105 (1988). Progress in Oceanography, 21 (1):??, ???? 1988. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1988:Pa

[Ano88d] Anonymous. Pages 1–82 (1988). Progress in Oceanography, 20 (1):??, ???? 1988. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1988:Pf

[Ano88e] Anonymous. Pages 107–226 (1988). Progress in Oceanography, 21(2):??, ???? 1988. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1988:Pc

[Ano88f] Anonymous. Pages 153–221 (1988). Progress in Oceanography, 20(3):??, ???? 1988. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1988:Pd

[Ano88g] Anonymous. Pages 223–313 (1988). Progress in Oceanography, 20(4):??, ???? 1988. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1988:Pg

[Ano88h] Anonymous. Pages 227–536 (1988). Progress in Oceanography, 21(3–4):??, ???? 1988. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1988:Pb

[Ano88i] Anonymous. Pages 83–151 (1988). Progress in Oceanography, 20(2):??, ???? 1988. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1989:CSP

[Ano89a] Anonymous. Contents of some previous volumes. Progress in Oceanography, 23(1):ix–xiv, ???? 1989. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0079661189900220.

Anonymous:1989:EBa

[Ano89b] Anonymous. Editorial Board. Progress in Oceanography, 22(1): ifc, ???? 1989. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661189900074.

Anonymous:1989:EBb

[Ano89c] Anonymous. Editorial Board. Progress in Oceanography, 23 (1):ii, ???? 1989. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661189900207.

Anonymous:1989:Pa

[Ano89d] Anonymous. Pages 1–123 (1989). Progress in Oceanography, 22 (1):??, ???? 1989. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1989:Pe

[Ano89e] Anonymous. Pages 1–64 (1989). Progress in Oceanography, 23 (1):??, ???? 1989. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1989:Pb

[Ano89f] Anonymous. Pages 125–204 (1989). Progress in Oceanography, 22(2):??, ???? 1989. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1989:Pg

[Ano89g] Anonymous. Pages 149–244 (1989). Progress in Oceanography, 23(3):??, ???? 1989. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1989:Pc

[Ano89h] Anonymous. Pages 205–275 (1989). Progress in Oceanography, 22(3):??, ???? 1989. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1989:Ph

[Ano89i] Anonymous. Pages 245–338 (1989). Progress in Oceanography, 23(4):??, ???? 1989. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1989:Pd

[Ano89j] Anonymous. Pages 277–359 (1989). Progress in Oceanography, 22(4):??, ???? 1989. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1989:Pf

[Ano89k] Anonymous. Pages 65–147 (1989). Progress in Oceanography, 23(2):??, ???? 1989. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1989:SI

[Ano89l] Anonymous. Subject index. Progress in Oceanography, 23(1): v–viii, ???? 1989. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661189900219.

Anonymous:1990:EBa

[Ano90a] Anonymous. Editorial Board. Progress in Oceanography, 24(??):ifc, ???? 1990. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/007966119090013R.

Anonymous:1990:EBb

[Ano90b] Anonymous. Editorial Board. Progress in Oceanography, 25 (1–4):IFC, i, ???? 1990. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/007966119090001I.

Anonymous:1990:P

[Ano90c] Anonymous. Pages 1–358 (1990). Progress in Oceanography, 25(1–4):??, ???? 1990. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1990:SIa

[Ano90d] Anonymous. Subject index. Progress in Oceanography, 24 (??):341–343, ???? 1990. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0079661190900432.

Anonymous:1990:SIb

[Ano90e] Anonymous. Subject index. Progress in Oceanography, 25 (1–4):353–358, ???? 1990. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/007966119090012Q.

Anonymous:1991:EBa

[Ano91a] Anonymous. Editorial Board. Progress in Oceanography, 26(1): ifc, ???? 1991. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661191900057.

Anonymous:1991:EBb

[Ano91b] Anonymous. Editorial Board. Progress in Oceanography, 27(1– 2):ifc, ???? 1991. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/007966119190013C.

Anonymous:1991:EBc

[Ano91c] Anonymous. Editorial Board. Progress in Oceanography, 28(1– 2):ifc, ???? 1991. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/007966119190017G.

Anonymous:1991:Pa

[Ano91d] Anonymous. Pages 1–113 (1991). Progress in Oceanography, 26 (1):??, ???? 1991. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1991:Pg

[Ano91e] Anonymous. Pages 1–188 (1991). Progress in Oceanography, 28(1–2):??, ???? 1991. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1991:Pe

[Ano91f] Anonymous. Pages 1–224 (1991). Progress in Oceanography, 27(1–2):??, ???? 1991. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1991:Pb

[Ano91g] Anonymous. Pages 115–241 (1991). Progress in Oceanography, 26(2):??, ???? 1991. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1991:Ph

[Ano91h] Anonymous. Pages 189–303 (1991). Progress in Oceanography, 28(3):??, ???? 1991. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1991:Pf

[Ano91i] Anonymous. Pages 225–438 (1991). Progress in Oceanography, 27(3–4):??, ???? 1991. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1991:Pc

[Ano91j] Anonymous. Pages 243–306 (1991). Progress in Oceanography, 26(3):??, ???? 1991. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1991:Pi

[Ano91k] Anonymous. Pages 305–383 (1991). Progress in Oceanography, 28(4):??, ???? 1991. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1991:Pd

[Ano91l] Anonymous. Pages 307–402 (1991). Progress in Oceanography, 26(4):??, ???? 1991. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1992:CSI

[Ano92a] Anonymous. Contents and subject index and contents of previous volumes. Progress in Oceanography, 20(??):v–xiii, ???? 1992. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/007966119290024T.

Anonymous:1992:EBa

[Ano92b] Anonymous. Editorial Board. Progress in Oceanography, 29(1): ifc, ???? 1992. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/007966119290001G.

Anonymous:1992:EBb

[Ano92c] Anonymous. Editorial Board. Progress in Oceanography, 30(1– 4):ifc, ???? 1992. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/007966119290007M.

Anonymous:1992:Pa

[Ano92d] Anonymous. Pages 1–132 (1992). Progress in Oceanography, 29 (1):??, ???? 1992. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1992:Pe

[Ano92e] Anonymous. Pages 1–398 (1992). Progress in Oceanography, 30(1–4):??, ???? 1992. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1992:Pb

[Ano92f] Anonymous. Pages 133–196 (1992). Progress in Oceanography, 29(2):??, ???? 1992. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1992:Pc

[Ano92g] Anonymous. Pages 197–281 (1992). Progress in Oceanography, 29(3):??, ???? 1992. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1992:Pd

[Ano92h] Anonymous. Pages 283–383 (1992). Progress in Oceanography, 29(4):??, ???? 1992. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1992:PIV

[Ano92i] Anonymous. Pages iii, v–xiii (1992). Progress in Oceanography, 20(??):??, ???? 1992. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1992:SI

[Ano92j] Anonymous. Subject index. Progress in Oceanography, 30 (1–4):393–398, ???? 1992. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/007966119290020Z.

Anonymous:1993:EBa

[Ano93a] Anonymous. Editorial Board. Progress in Oceanography, 31(1): ifc, ???? 1993. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661193900215.

Anonymous:1993:EBb

[Ano93b] Anonymous. Editorial Board. Progress in Oceanography, 32(??):ifc, ???? 1993. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/007966119390006Y.

Anonymous:1993:Pa

[Ano93c] Anonymous. Pages 1–99 (1993). Progress in Oceanography, 31 (1):??, ???? 1993. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1993:Pb

[Ano93d] Anonymous. Pages 101–179 (1993). Progress in Oceanography, 31(2):??, ???? 1993. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1993:Pc

[Ano93e] Anonymous. Pages 181–338 (1993). Progress in Oceanography, 31(3):??, ???? 1993. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1993:Pd

[Ano93f] Anonymous. Pages 339–419 (1993). Progress in Oceanography, 31(4):??, ???? 1993. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1993:SI

[Ano93g] Anonymous. Subject index. Progress in Oceanography, 32 (??):353–358, ???? 1993. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/007966119390020E.

Anonymous:1994:EBa

[Ano94a] Anonymous. Editorial Board. Progress in Oceanography, 33(1): ifc, ???? 1994. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661194900132.

Anonymous:1994:EBb

[Ano94b] Anonymous. Editorial Board. Progress in Oceanography, 34(1): ifc, ???? 1994. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661194900248.

Anonymous:1994:ENI

[Ano94c] Anonymous. Erratum: A numerical investigation of dynamics, thermodynamics and mixed-layer processes in the Indian Ocean [Progress in Oceanography 31(4) 181–244]. Progress in Oceanography, 33(3):I, A204, A209, A219, ???? 1994. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0079661194900299.

Anonymous:1994:Pd

[Ano94d] Anonymous. Pages 1–78 (1994). Progress in Oceanography, 34 (1):??, ???? 1994. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1994:Pa

[Ano94e] Anonymous. Pages 1–92 (1994). Progress in Oceanography, 33 (1):??, ???? 1994. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1994:Pc

[Ano94f] Anonymous. Pages 249–386 (1994). Progress in Oceanography, 33(4):??, ???? 1994. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1994:Pf

[Ano94g] Anonymous. Pages 285–392 (1994). Progress in Oceanography, 34(4):??, ???? 1994. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1994:Pe

[Ano94h] Anonymous. Pages 79–284 (1994). Progress in Oceanography, 34(2–3):??, ???? 1994. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1994:Pb

[Ano94i] Anonymous. Pages 93–159 (1994). Progress in Oceanography, 33(2):??, ???? 1994. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1994:PAA

[Ano94j] Anonymous. Pages A204, A209, A219, 161–248 (1994). Progress in Oceanography, 33(3):??, ???? 1994. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1994:SCD

[Ano94k] Anonymous. Suggested criteria for describing deep-sea benthic communities; the final report of SCOR Working Group 76. Progress in Oceanography, 34(2–3):81–100, ???? 1994. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0079661194900027.

Anonymous:1995:EBa

[Ano95a] Anonymous. Editorial Board. Progress in Oceanography, 35(1): ifc, ???? 1995. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661195900047.

Anonymous:1995:EBb

[Ano95b] Anonymous. Editorial Board. Progress in Oceanography, 36(1): ifc, ???? 1995. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661195900071.

Anonymous:1995:Pe

[Ano95c] Anonymous. Pages 1–75 (1995). Progress in Oceanography, 36 (1):??, ???? 1995. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1995:Pa

[Ano95d] Anonymous. Pages 1–80 (1995). Progress in Oceanography, 35 (1):??, ???? 1995. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1995:Pg

[Ano95e] Anonymous. Pages 169–247 (1995). Progress in Oceanography, 36(3):??, ???? 1995. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1995:Pc

[Ano95f] Anonymous. Pages 207–294 (1995). Progress in Oceanography, 35(3):??, ???? 1995. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1995:Ph

[Ano95g] Anonymous. Pages 249–376 (1995). Progress in Oceanography, 36(4):??, ???? 1995. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1995:Pd

[Ano95h] Anonymous. Pages 295–431 (1995). Progress in Oceanography, 35(4):??, ???? 1995. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1995:Pf

[Ano95i] Anonymous. Pages 77–167 (1995). Progress in Oceanography, 36(2):??, ???? 1995. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1995:Pb

[Ano95j] Anonymous. Pages 81–205 (1995). Progress in Oceanography, 35(2):??, ???? 1995. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1996:EBa

[Ano96a] Anonymous. Editorial Board. Progress in Oceanography, 37(1): ifc, ???? 1996. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661196800012.

Anonymous:1996:EBb

[Ano96b] Anonymous. Editorial Board. Progress in Oceanography, 38(1): ifc, ???? 1996. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661196900063.

Anonymous:1996:Pa

[Ano96c] Anonymous. Pages 1–115 (1996). Progress in Oceanography, 37 (1):??, ???? 1996. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1996:Pd

[Ano96d] Anonymous. Pages 1–93 (1996). Progress in Oceanography, 38 (1):??, ???? 1996. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1996:Pb

[Ano96e] Anonymous. Pages 117–192 (1996). Progress in Oceanography, 37(2):??, ???? 1996. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1996:Pc

[Ano96f] Anonymous. Pages 193–291 (1996). Progress in Oceanography, 37(3–4):??, ???? 1996. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1996:Pf

[Ano96g] Anonymous. Pages 205–295 (1996). Progress in Oceanography, 38(3):??, ???? 1996. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1996:Pg

[Ano96h] Anonymous. Pages 297–451 (1996). Progress in Oceanography, 38(4):??, ???? 1996. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1996:Pe

[Ano96i] Anonymous. Pages 95–204 (1996). Progress in Oceanography, 38(2):??, ???? 1996. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1997:EBa

[Ano97a] Anonymous. Editorial Board. Progress in Oceanography, 39(1): ifc, ???? 1997. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661197900008.

Anonymous:1997:EBb

[Ano97b] Anonymous. Editorial Board. Progress in Oceanography, 40(??):ifc, ???? 1997. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S007966119790001X.

Anonymous:1997:Pa

[Ano97c] Anonymous. Pages 1–78 (1997). Progress in Oceanography, 39 (1):??, ???? 1997. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1997:Pc

[Ano97d] Anonymous. Pages 153–262 (1997). Progress in Oceanography, 39(3):??, ???? 1997. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1997:Pd

[Ano97e] Anonymous. Pages 263–352 (1997). Progress in Oceanography, 39(4):??, ???? 1997. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1997:Pb

[Ano97f] Anonymous. Pages 79–151 (1997). Progress in Oceanography, 39(2):??, ???? 1997. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1998:PJa

[Ano98a] Anonymous. Pages 1–139 (January 1998). Progress in Oceanography, 41(1):??, January 1998. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1998:PD

[Ano98b] Anonymous. Pages 1–258 (December 1998). Progress in Oceanography, 42(1–4):??, December 1998. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1998:PA

[Ano98c] Anonymous. Pages 141–247 (August 1998). Progress in Oceanography, 41(2):??, August 1998. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1998:PJb

[Ano98d] Anonymous. Pages 249–382 (June 1998). Progress in Oceanography, 41(3):??, June 1998. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1998:PO

[Ano98e] Anonymous. Pages 383–578 (October 1998). Progress in Oceanography, 41(4):??, October 1998. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1999:E

[Ano99a] Anonymous. Editorial. Progress in Oceanography, 43(2–4):157– 161, March 1999. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661199000075.

Anonymous:1999:PJ

[Ano99b] Anonymous. Pages 1–156 (January 1999). Progress in Oceanography, 43(1):??, January 1999. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1999:PA

[Ano99c] Anonymous. Pages 1–468 (August 1999). Progress in Oceanography, 44(1–3):??, August 1999. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1999:PM

[Ano99d] Anonymous. Pages 157–488 (March 1999). Progress in Oceanography, 43(2–4):??, March 1999. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:1999:PD

[Ano99e] Anonymous. Pages 469–700 (December 1999). Progress in Oceanography, 44(4):??, December 1999. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2000:PJ

[Ano00a] Anonymous. Pages 1–108 (January 2000). Progress in Oceanography, 45(1):??, January 2000. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2000:PM

[Ano00b] Anonymous. Pages 1–84 (May 2000). Progress in Oceanography, 46(1):??, May 2000. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2000:POa

[Ano00c] Anonymous. Pages 1–98 (October 2000). Progress in Oceanography, 47(1):??, October 2000. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2000:PF

[Ano00d] Anonymous. Pages 109–250 (February 2000). Progress in Oceanography, 45(2):??, February 2000. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2000:PAa

[Ano00e] Anonymous. Pages 251–466 (April 2000). Progress in Oceanography, 45(3–4):??, April 2000. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2000:P

[Ano00f] Anonymous. Pages 337–510 (2000). Progress in Oceanography, 48(4):??, ???? 2000. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2000:PAb

[Ano00g] Anonymous. Pages 85–482 (August 2000). Progress in Oceanography, 46(2–4):??, August 2000. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2000:POb

[Ano00h] Anonymous. Pages 99–408 (October 2000). Progress in Oceanography, 47(2–4):??, October 2000. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2001:D

[Ano01a] Anonymous. (DYNAMO). Progress in Oceanography, 48(??): ??, ???? 2001. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2001:Pa

[Ano01b] Anonymous. Pages 1–120 (2001). Progress in Oceanography, 48 (1):??, ???? 2001. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2001:Pb

[Ano01c] Anonymous. Pages 1–214 (2001). Progress in Oceanography, 51 (1):??, ???? 2001. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2002:PA

[Ano02a] Anonymous. Pages 1–103 (April 2002). Progress in Oceanography, 53(1):??, April 2002. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2002:PJ

[Ano02b] Anonymous. Pages 1–120 (January 2002). Progress in Oceanography, 52(1):??, January 2002. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2002:PO

[Ano02c] Anonymous. Pages 1–261 (October 2002). Progress in Oceanography, 55(1–2):??, October 2002. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2003:PJb

[Ano03a] Anonymous. Pages 1–114 (July 2003). Progress in Oceanography, 58(1):??, July 2003. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2003:PO

[Ano03b] Anonymous. Pages 1–176 (October 2003). Progress in Oceanography, 59(1):??, October 2003. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2003:PJa

[Ano03c] Anonymous. Pages 1–186 (January 2003). Progress in Oceanography, 56(1):??, January 2003. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2003:PM

[Ano03d] Anonymous. Pages 125–237 (May 2003). Progress in Oceanography, 57(2):??, May 2003. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2003:PF

[Ano03e] Anonymous. Pages 187–380 (February 2003). Progress in Oceanography, 56(2):??, February 2003. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2003:PD

[Ano03f] Anonymous. Pages 339–474 (December 2003). Progress in Oceanography, 59(4):??, December 2003. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2003:P

[Ano03g] Anonymous. Preface. Progress in Oceanography, 58(2–4):115– 116, August 2003. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661103001265.

Anonymous:2003:PNa

[Ano03h] Anonymous. Publishers note. Progress in Oceanography, 56(1): 1, January 2003. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661103000028.

Anonymous:2003:PNb

[Ano03i] Anonymous. Publisher‘s note. Progress in Oceanography, 56(2):187, February 2003. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661103000235.

Anonymous:2003:WGL

[Ano03j] Anonymous. Working group list of membership. Progress in Oceanography, 56(3–4):395–397, March 2003. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661103000351.

Anonymous:2004:M

[Ano04a] Anonymous. In memoriam. Progress in Oceanography, 60 (2–4):133, February 2004. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661104000230.

Anonymous:2004:PA

[Ano04b] Anonymous. Pages 1–102 (April 2004). Progress in Oceanography, 61(1):??, April 2004. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2004:PJa

[Ano04c] Anonymous. Pages 1–132 (January 2004). Progress in Oceanography, 60(1):??, January 2004. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2004:PJb

[Ano04d] Anonymous. Pages 1–66 (July 2004). Progress in Oceanography, 62(1):??, July 2004. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2004:PO

[Ano04e] Anonymous. Pages 1–98 (October 2004). Progress in Oceanography, 63(1–2):??, October 2004. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2004:PD

[Ano04f] Anonymous. Pages 159–228 (December 2004). Progress in Oceanography, 63(4):??, December 2004. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2004:PN

[Ano04g] Anonymous. Pages 99–158 (November 2004). Progress in Oceanography, 63(3):??, November 2004. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2005:PA

[Ano05a] Anonymous. Pages 1–112 (April 2005). Progress in Oceanography, 65(1):??, April 2005. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2005:PO

[Ano05b] Anonymous. Pages 1–282 (October 2005). Progress in Oceanography, 67(1–2):??, October 2005. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2005:PJb

[Ano05c] Anonymous. Pages 1–86 (July 2005). Progress in Oceanography, 66(1):??, July 2005. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2005:PJa

[Ano05d] Anonymous. Pages 1–90 (January 2005). Progress in Oceanography, 64(1):??, January 2005. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2006:PJa

[Ano06a] Anonymous. Pages 1–114 (January 2006). Progress in Oceanography, 68(1):??, January 2006. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2006:PO

[Ano06b] Anonymous. Pages 1–122 (October 2006). Progress in Oceanography, 71(1):??, October 2006. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2006:PA

[Ano06c] Anonymous. Pages 1–90 (April 2006). Progress in Oceanography, 69(1):??, April 2006. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2006:PJb

[Ano06d] Anonymous. Pages 1–90 (July 2006). Progress in Oceanography, 70(1):??, July 2006. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2007:EBa

[Ano07a] Anonymous. Editorial Board. Progress in Oceanography, 73(1): ifc, April 2007. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661107000511.

Anonymous:2007:EBb

[Ano07b] Anonymous. Editorial Board. Progress in Oceanography, 73 (2):ifc, May 2007. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661107000882.

Anonymous:2007:EBc

[Ano07c] Anonymous. Editorial Board. Progress in Oceanography, 73(3– 4):ifc, May 2007. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661107001036.

Anonymous:2007:EBd

[Ano07d] Anonymous. Editorial Board. Progress in Oceanography, 74 (1):ifc, July 2007. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661107001152.

Anonymous:2007:EBe

[Ano07e] Anonymous. Editorial Board. Progress in Oceanography, 74(2–3):ifc, August 2007. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661107001255.

Anonymous:2007:EBf

[Ano07f] Anonymous. Editorial Board. Progress in Oceanography, 74 (4):ifc, September 2007. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661107001371.

Anonymous:2007:EBg

[Ano07g] Anonymous. Editorial Board. Progress in Oceanography, 75(1): ifc, October 2007. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661107001772.

Anonymous:2007:EBh

[Ano07h] Anonymous. Editorial Board. Progress in Oceanography, 75(2): ifc, October 2007. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661107001887.

Anonymous:2007:EBi

[Ano07i] Anonymous. Editorial Board. Progress in Oceanography, 75(3):ifc, November 2007. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661107002017.

Anonymous:2007:EBj

[Ano07j] Anonymous. Editorial Board. Progress in Oceanography, 75(4):ifc, December 2007. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661107002157.

Anonymous:2007:PJa

[Ano07k] Anonymous. Pages 1–114 (January 2007). Progress in Oceanography, 72(1):??, January 2007. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2007:PO

[Ano07l] Anonymous. Pages 1–114 (October 2007). Progress in Oceanography, 75(1):??, October 2007. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2007:PJb

[Ano07m] Anonymous. Pages 1–94 (July 2007). Progress in Oceanography, 74(1):??, July 2007. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2007:PA

[Ano07n] Anonymous. Pages 1–98 (April 2007). Progress in Oceanography, 73(1):??, April 2007. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2007:PMa

[Ano07o] Anonymous. Pages 275–384 (March 2007). Progress in Oceanography, 72(4):??, March 2007. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2007:PS

[Ano07p] Anonymous. Pages 423–572 (September 2007). Progress in Oceanography, 74(4):??, September 2007. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2007:PD

[Ano07q] Anonymous. Pages 603–860 (December 2007). Progress in Oceanography, 75(4):??, December 2007. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2007:PMb

[Ano07r] Anonymous. Pages 99–202 (May 2007). Progress in Oceanography, 73(2):??, May 2007. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2007:PN

[Ano07s] Anonymous. Publisher’s note. Progress in Oceanography, 72(4):275, March 2007. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661106001595.

Anonymous:2008:EBa

[Ano08a] Anonymous. Editorial Board. Progress in Oceanography, 76(1): ifc, January 2008. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661107002248.

Anonymous:2008:EBb

[Ano08b] Anonymous. Editorial Board. Progress in Oceanography, 76(2): ifc, February 2008. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661108000165.

Anonymous:2008:EBc

[Ano08c] Anonymous. Editorial Board. Progress in Oceanography, 76(3): ifc, March 2008. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661108000268.

Anonymous:2008:EBd

[Ano08d] Anonymous. Editorial Board. Progress in Oceanography, 76(4): ifc, March 2008. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661108000372.

Anonymous:2008:EBe

[Ano08e] Anonymous. Editorial Board. Progress in Oceanography, 77(1): ifc, April 2008. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S007966110800075X.

Anonymous:2008:EBf

[Ano08f] Anonymous. Editorial Board. Progress in Oceanography, 77(2– 3):ifc, May 2008. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S007966110800092X.

Anonymous:2008:EBg

[Ano08g] Anonymous. Editorial Board. Progress in Oceanography, 77(4): ifc, June 2008. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661108000992.

Anonymous:2008:EBh

[Ano08h] Anonymous. Editorial Board. Progress in Oceanography, 78 (1):ifc, July 2008. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661108001110.

Anonymous:2008:EBi

[Ano08i] Anonymous. Editorial Board. Progress in Oceanography, 78(2): ifc, August 2008. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661108001250.

Anonymous:2008:EBj

[Ano08j] Anonymous. Editorial Board. Progress in Oceanography, 78 (3):ifc, September 2008. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661108001389.

Anonymous:2008:EBk

[Ano08k] Anonymous. Editorial Board. Progress in Oceanography, 78 (4):ifc, September 2008. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S007966110800150X.

Anonymous:2008:EBl

[Ano08l] Anonymous. Editorial Board. Progress in Oceanography, 79(1): ifc, October 2008. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661108001602.

Anonymous:2008:EBm

[Ano08m] Anonymous. Editorial Board. Progress in Oceanography, 79 (2–4):ifc, October 2008. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S007966110800195X.

Anonymous:2008:PJb

[Ano08n] Anonymous. Pages 1–106 (July 2008). Progress in Oceanography, 78(1):??, July 2008. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2008:PJa

[Ano08o] Anonymous. Pages 1–148 (January 2008). Progress in Oceanography, 76(1):??, January 2008. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2008:PAa

[Ano08p] Anonymous. Pages 1–82 (April 2008). Progress in Oceanography, 77(1):??, April 2008. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2008:PO

[Ano08q] Anonymous. Pages 1–94 (October 2008). Progress in Oceanography, 79(1):??, October 2008. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2008:PAb

[Ano08r] Anonymous. Pages 107–192 (August 2008). Progress in Oceanography, 78(2):??, August 2008. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2008:PF

[Ano08s] Anonymous. Pages 149–216 (February 2008). Progress in Oceanography, 76(2):??, February 2008. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2008:PSa

[Ano08t] Anonymous. Pages 193–256 (September 2008). Progress in Oceanography, 78(3):??, September 2008. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2008:PMa

[Ano08u] Anonymous. Pages 217–398 (March 2008). Progress in Oceanography, 76(3):??, March 2008. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2008:PSb

[Ano08v] Anonymous. Pages 257–336 (September 2008). Progress in Oceanography, 78(4):??, September 2008. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2008:PMb

[Ano08w] Anonymous. Pages 399–486 (March 2008). Progress in Oceanography, 76(4):??, March 2008. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2008:PN

[Ano08x] Anonymous. Publisher’s note. Progress in Oceanography, 76(1):1, January 2008. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661107002303.

Anonymous:2008:RA

[Ano08y] Anonymous. Reviewer acknowledgement 2007. Progress in Oceanography, 76(2):149–150, February 2008. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661107002212.

Anonymous:2009:EBa

[Ano09a] Anonymous. Editorial Board. Progress in Oceanography, 80(3– 4):ifc, March 2009. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661109000391.

Anonymous:2009:EBb

[Ano09b] Anonymous. Editorial Board. Progress in Oceanography, 81(1– 4):ifc, April 2009. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661109000615.

Anonymous:2009:EBc

[Ano09c] Anonymous. Editorial Board. Progress in Oceanography, 82 (1):ifc, July 2009. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661109001141.

Anonymous:2009:EBd

[Ano09d] Anonymous. Editorial Board. Progress in Oceanography, 82(2): ifc, August 2009. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661109001232.

Anonymous:2009:EBe

[Ano09e] Anonymous. Editorial Board. Progress in Oceanography, 82 (3):ifc, September 2009. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661109001530.

Anonymous:2009:EBf

[Ano09f] Anonymous. Editorial Board. Progress in Oceanography, 82 (4):ifc, September 2009. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661109001633.

Anonymous:2009:EBg

[Ano09g] Anonymous. Editorial Board. Progress in Oceanography, 83 (1–4):ifc, December 2009. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661109001761.

Anonymous:2009:EBU

[Ano09h] Anonymous. Eastern boundary upwelling ecosystems: Integrative and comparative approaches: Integrative and comparative approaches, 2–8 June 2008, Las Palmas, Gran Canaria, Spain. Progress in Oceanography, 83(1–4):??, December 2009. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2009:PJ

[Ano09i] Anonymous. Pages 1–80 (July 2009). Progress in Oceanography, 82(1):??, July 2009. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2009:PM

[Ano09j] Anonymous. Pages 113–218 (March 2009). Progress in Oceanography, 80(3–4):??, March 2009. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2009:PSa

[Ano09k] Anonymous. Pages 149–226 (September 2009). Progress in Oceanography, 82(3):??, September 2009. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2009:PSb

[Ano09l] Anonymous. Pages 227–326 (September 2009). Progress in Oceanography, 82(4):??, September 2009. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2009:PA

[Ano09m] Anonymous. Pages 81–148 (August 2009). Progress in Oceanography, 82(2):??, August 2009. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2010:BNI

[Ano10a] Anonymous. A brief note on introduction and conclusion chapters in GLOBEC compendium. Progress in Oceanography, 87(S1):S1, ???? 2010. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661110001692.

Anonymous:2010:C

[Ano10b] Anonymous. CLIOTOP. Progress in Oceanography, 86(1–2): ??, July 2010. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2010:EBa

[Ano10c] Anonymous. Editorial Board. Progress in Oceanography, 84 (1–2):ifc, January 2010. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661110000054.

Anonymous:2010:EBb

[Ano10d] Anonymous. Editorial Board. Progress in Oceanography, 84(3– 4):ifc, March 2010. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661110000236.

Anonymous:2010:EBc

[Ano10e] Anonymous. Editorial Board. Progress in Oceanography, 85(1– 2):ifc, April 2010. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661110000662.

Anonymous:2010:EBd

[Ano10f] Anonymous. Editorial Board. Progress in Oceanography, 85(3– 4):ifc, June 2010. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661110000765.

Anonymous:2010:EBe

[Ano10g] Anonymous. Editorial Board. Progress in Oceanography, 86(1– 2):ifc, July 2010. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661110000881.

Anonymous:2010:EBf

[Ano10h] Anonymous. Editorial Board. Progress in Oceanography, 86(3–4):ifc, September 2010. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661110001011.

Anonymous:2010:EBg

[Ano10i] Anonymous. Editorial Board. Progress in Oceanography, 87 (1–4):ifc, October 2010. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661110001412.

Anonymous:2010:P

[Ano10j] Anonymous. Pages 1–8 (2010). Progress in Oceanography, 87(S1):??, ???? 2010. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2010:PJ

[Ano10k] Anonymous. Pages 137–244 (June 2010). Progress in Oceanography, 85(3–4):??, June 2010. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2010:PM

[Ano10l] Anonymous. Pages 139–254 (March 2010). Progress in Oceanography, 84(3–4):??, March 2010. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2010:PS

[Ano10m] Anonymous. Pages 317–414 (September 2010). Progress in Oceanography, 86(3–4):??, September 2010. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2011:EBa

[Ano11a] Anonymous. Editorial Board. Progress in Oceanography, 89(1– 4):ifc, April 2011. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661111000309.

Anonymous:2011:EBb

[Ano11b] Anonymous. Editorial Board. Progress in Oceanography, 90(1– 4):ifc, July 2011. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661111000413.

Anonymous:2011:EBc

[Ano11c] Anonymous. Editorial Board. Progress in Oceanography, 91(1): ifc, October 2011. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661111000784.

Anonymous:2011:EBd

[Ano11d] Anonymous. Editorial Board. Progress in Oceanography, 91(2): ifc, October 2011. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661111000917.

Anonymous:2011:EBe

[Ano11e] Anonymous. Editorial Board. Progress in Oceanography, 91(3):ifc, November 2011. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661111001029.

Anonymous:2011:EBf

[Ano11f] Anonymous. Editorial Board. Progress in Oceanography, 91(4):ifc, December 2011. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661111001340.

Anonymous:2011:POa

[Ano11g] Anonymous. Pages 1–96 (October 2011). Progress in Oceanography, 91(1):??, October 2011. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2011:PN

[Ano11h] Anonymous. Pages 217–344 (November 2011). Progress in Oceanography, 91(3):??, November 2011. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2011:PD

[Ano11i] Anonymous. Pages 345–594 (December 2011). Progress in Oceanography, 91(4):??, December 2011. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2011:POb

[Ano11j] Anonymous. Pages 97–216 (October 2011). Progress in Oceanography, 91(2):??, October 2011. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2012:EBa

[Ano12a] Anonymous. Editorial Board. Progress in Oceanography, 92–95(??):ifc, January 2012. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661111001467.

Anonymous:2012:EBb

[Ano12b] Anonymous. Editorial Board. Progress in Oceanography, 96(1): ifc, April 2012. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661112000122.

Anonymous:2012:EBc

[Ano12c] Anonymous. Editorial Board. Progress in Oceanography, 97–100(??):ifc, May 2012. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661112000298.

Anonymous:2012:EBd

[Ano12d] Anonymous. Editorial Board. Progress in Oceanography, 101(1):ifc, August 2012. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661112000523.

Anonymous:2012:EBe

[Ano12e] Anonymous. Editorial Board. Progress in Oceanography, 102(??):ifc, September 2012. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661112000730.

Anonymous:2012:EBf

[Ano12f] Anonymous. Editorial Board. Progress in Oceanography, 103(??):ifc, September 2012. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661112000821.

Anonymous:2012:EBg

[Ano12g] Anonymous. Editorial Board. Progress in Oceanography, 104(??):ifc, October 2012. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S007966111200105X.

Anonymous:2012:EBh

[Ano12h] Anonymous. Editorial Board. Progress in Oceanography, 105(??):ifc, October 2012. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661112001218.

Anonymous:2012:EBi

[Ano12i] Anonymous. Editorial Board. Progress in Oceanography, 106 (??):ifc, November 2012. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S007966111200136X.

Anonymous:2012:EBj

[Ano12j] Anonymous. Editorial Board. Progress in Oceanography, 107 (??):ifc, December 2012. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S007966111200153X.

Anonymous:2012:PS

[Ano12k] Anonymous. Pages 1–106 (September 2012). Progress in Oceanography, 103(??):1–106, September 2012. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2012:PO

[Ano12l] Anonymous. Pages 1–110 (October 2012). Progress in Oceanography, 104(??):1–110, October 2012. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2012:PAa

[Ano12m] Anonymous. Pages 1–140 (April 2012). Progress in Oceanography, 96(1):??, April 2012. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2012:PAb

[Ano12n] Anonymous. Pages 1–146 (August 2012). Progress in Oceanography, 101(1):??, August 2012. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2012:PN

[Ano12o] Anonymous. Pages 1–174 (November 2012). Progress in Oceanography, 106(??):1–174, November 2012. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2012:PD

[Ano12p] Anonymous. Pages 1–80 (December 2012). Progress in Oceanography, 107(??):1–80, December 2012. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2013:EBa

[Ano13a] Anonymous. Editorial Board. Progress in Oceanography, 108(??):ifc, January 2013. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661112001760.

Anonymous:2013:EBb

[Ano13b] Anonymous. Editorial Board. Progress in Oceanography, 109(??):ifc, February 2013. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661113000086.

Anonymous:2013:EBc

[Ano13c] Anonymous. Editorial Board. Progress in Oceanography, 110(??):ifc, March 2013. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661113000219.

Anonymous:2013:EBd

[Ano13d] Anonymous. Editorial Board. Progress in Oceanography, 111(??):ifc, April 2013. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661113000359.

Anonymous:2013:EBe

[Ano13e] Anonymous. Editorial Board. Progress in Oceanography, 114 (??):ifc, July 2013. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S007966111300089X.

Anonymous:2013:EBf

[Ano13f] Anonymous. Editorial Board. Progress in Oceanography, 115(??):ifc, August 2013. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661113001183.

Anonymous:2013:HVS

[Ano13g] Anonymous. Hjálmar Vilhjálmsson (September 25, 1937– August 20, 2011). Progress in Oceanography, 114(??):1–2, July 2013. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661113000438.

Anonymous:2013:IEPa

[Ano13h]  Anonymous.  IFC — eds/publication info.  Progress

in Oceanography, 116(??):ifc, September 2013. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661113001547.

Anonymous:2013:IEPb

[Ano13i]  Anonymous.  IFC — eds/publication info.  Progress

in Oceanography, 117(??):ifc, October 2013. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661113001766.

Anonymous:2013:IEPc

[Ano13j]  Anonymous.  IFC — eds/publication info.  Progress

in Oceanography, 118(??):ifc, November 2013. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661113001985.

Anonymous:2013:IEPd

[Ano13k]  Anonymous.  IFC — eds/publication info.  Progress

in Oceanography, 119(??):ifc, December 2013. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661113002139.

Anonymous:2013:PF

[Ano13l] Anonymous. Pages 1–130 (February 2013). Progress in Oceanography, 109(??):1–130, February 2013. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2013:PM

[Ano13m] Anonymous. Pages 1–150 (March 2013). Progress in Oceanography, 110(??):1–150, March 2013. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2013:PA

[Ano13n] Anonymous. Pages 1–154 (April 2013). Progress in Oceanography, 111(??):1–154, April 2013. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2013:PS

[Ano13o] Anonymous. Pages 1–260 (September 2013). Progress in Oceanography, 116(??):1–260, September 2013. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2013:PJ

[Ano13p] Anonymous. Pages 1–80 (January 2013). Progress in Oceanography, 108(??):1–80, January 2013. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2014:IEPa

[Ano14a]  Anonymous.  IFC — eds/publication info.  Progress

in Oceanography, 120(??):ifc, January 2014. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661113002334.

Anonymous:2014:IEPb

[Ano14b]  Anonymous.  IFC — eds/publication info.  Progress

in Oceanography, 121(??):ifc, February 2014. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661114000068.

Anonymous:2014:IEPc

[Ano14c] Anonymous. IFC — eds/publication info. Progress in Oceanography, 122(??):ifc, March 2014. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661114000214.

Anonymous:2014:IEPd

[Ano14d] Anonymous. IFC — eds/publication info. Progress in Oceanography, 123(??):ifc, April 2014. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661114000445.

Anonymous:2014:IEPe

[Ano14e] Anonymous. IFC — eds/publication info. Progress in Oceanography, 124(??):ifc, May 2014. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661114000792.

Anonymous:2014:IEPf

[Ano14f] Anonymous. IFC — eds/publication info. Progress in Oceanography, 125(??):ifc, June 2014. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661114001025.

Anonymous:2014:IEPg

[Ano14g] Anonymous. IFC — eds/publication info. Progress in Oceanography, 126(??):ifc, August 2014. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661114001177.

Anonymous:2014:IEPh

[Ano14h]  Anonymous.  IFC — eds/publication info.  Progress

in Oceanography, 127(??):ifc, September 2014. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661114001360.

Anonymous:2014:IEPi

[Ano14i]  Anonymous.  IFC — eds/publication info.  Progress

in Oceanography, 128(??):ifc, November 2014. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661114001566.

Anonymous:2014:IEPj

[Ano14j]  Anonymous.  IFC — eds/publication info.  Progress

in Oceanography, 129(??):ifc, December 2014. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661114001839.

Anonymous:2014:IEPk

[Ano14k]  Anonymous.  IFC — eds/publication info.  Progress

in Oceanography, 129(??):ifc, December 2014. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S007966111400189X.

Anonymous:2014:PS

[Ano14l] Anonymous. Pages 1–128 (September 2014). Progress in Oceanography, 127(??):1–128, September 2014. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2014:PA

[Ano14m] Anonymous. Pages 1–144 (April 2014). Progress in Oceanography, 123(??):1–144, April 2014. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2014:PMa

[Ano14n] Anonymous. Pages 1–174 (March 2014). Progress in Oceanography, 122(??):1–174, March 2014. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2014:PN

[Ano14o] Anonymous. Pages 1–184 (November 2014). Progress in Oceanography, 128(??):1–184, November 2014. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2014:PJa

[Ano14p] Anonymous. Pages 1–424 (January 2014). Progress in Oceanography, 120(??):1–424, January 2014. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2014:PJb

[Ano14q] Anonymous. Pages 1–74 (June 2014). Progress in Oceanography, 125(??):1–74, June 2014. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2014:PMb

[Ano14r] Anonymous. Pages 1–92 (May 2014). Progress in Oceanography, 124(??):1–92, May 2014. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2015:IEPa

[Ano15a]  Anonymous.  IFC — eds/publication info.  Progress

in Oceanography, 130(??):ifc, January 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661114002080.

Anonymous:2015:IEPb

[Ano15b]  Anonymous.  IFC — eds/publication info.  Progress

in Oceanography, 131(??):ifc, February 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661115000142.

Anonymous:2015:IEPc

[Ano15c] Anonymous. IFC — eds/publication info. Progress in Oceanography, 132(??):ifc, March 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661115000300.

Anonymous:2015:IEPd

[Ano15d] Anonymous. IFC — eds/publication info. Progress in Oceanography, 133(??):ifc, April 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661115000580.

Anonymous:2015:IEPe

[Ano15e] Anonymous. IFC — eds/publication info. Progress in Oceanography, 134(??):ifc, May 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661115000762.

Anonymous:2015:IEPf

[Ano15f] Anonymous. IFC — eds/publication info. Progress in Oceanography, 135(??):ifc, June 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661115001214.

Anonymous:2015:IEPg

[Ano15g] Anonymous. IFC — eds/publication info. Progress in Oceanography, 136(??):ifc, August 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661115001421.

Anonymous:2015:IEPh

[Ano15h] Anonymous. IFC — eds/publication info. Progress in Oceanography, 137 (part A)(??):ifc, September 2015. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S007966111500169X.

Anonymous:2015:IEPi

[Ano15i]  Anonymous.  IFC — eds/publication info.  Progress

in Oceanography, 137(??):ifc, September 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661115001780.

Anonymous:2015:IEPj

[Ano15j] Anonymous. IFC — eds/publication info. Progress in Oceanography, 138 (part A)(??):ifc, November 2015. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661115002062.

Anonymous:2015:IEPk

[Ano15k]  Anonymous.  IFC — eds/publication info.  Progress

in Oceanography, 138(??):ifc, November 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661115002207.

Anonymous:2015:IEPl

[Ano15l]  Anonymous.  IFC — eds/publication info.  Progress

in Oceanography, 139(??):ifc, December 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S007966111500230X.

Anonymous:2015:PJb

[Ano15m] Anonymous. Pages 1–176 (June 2015). Progress in Oceanography, 135(??):1–176, June 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2015:PF

[Ano15n] Anonymous. Pages 1–200 (February 2015). Progress in Oceanography, 131(??):1–200, February 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2015:PJa

[Ano15o] Anonymous. Pages 1–250 (January 2015). Progress in Oceanography, 130(??):1–250, January 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2015:PN

[Ano15p] Anonymous. Pages 1–324 (November 2015). Progress in Oceanography, 138 (part A)(??):1–324, November 2015. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2015:PS

[Ano15q] Anonymous. Pages 1–334 (September 2015). Progress in Oceanography, 137 (part A)(??):1–334, September 2015. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2015:PM

[Ano15r] Anonymous. Pages 1–474 (May 2015). Progress in Oceanography, 134(??):1–474, May 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2016:IEPa

[Ano16a]  Anonymous.  IFC — eds/publication info.  Progress

in Oceanography, 140(??):ifc, January 2016. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661115002438.

Anonymous:2016:IEPb

[Ano16b]  Anonymous.  IFC — eds/publication info.  Progress

in Oceanography, 141(??):ifc, February 2016. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661116000124.

Anonymous:2016:IEPc

[Ano16c] Anonymous. IFC — eds/publication info. Progress in Oceanography, 142(??):ifc, March 2016. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661116000227.

Anonymous:2016:IEPd

[Ano16d] Anonymous. IFC — eds/publication info. Progress in Oceanography, 143(??):ifc, April 2016. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661116300593.

Anonymous:2016:IEPe

[Ano16e] Anonymous. IFC — eds/publication info. Progress in Oceanography, 144(??):ifc, May 2016. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661116300751.

Anonymous:2016:IEPf

[Ano16f] Anonymous. IFC — eds/publication info. Progress in Oceanography, 145(??):ifc, June 2016. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661116301008.

Anonymous:2016:IEPg

[Ano16g] Anonymous. IFC — eds/publication info. Progress in Oceanography, 146(??):ifc, August 2016. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661116301379.

Anonymous:2016:IEPh

[Ano16h]  Anonymous.  IFC — eds/publication info.  Progress

in Oceanography, 147(??):ifc, September 2016. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661116301677.

Anonymous:2016:IEPi

[Ano16i]  Anonymous.  IFC — eds/publication info.  Progress

in Oceanography, 148(??):ifc, November 2016. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661116302142.

Anonymous:2016:IEPj

[Ano16j]  Anonymous.  IFC — eds/publication info.  Progress

in Oceanography, 149(??):ifc, December 2016. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661116302464.

Anonymous:2016:PMa

[Ano16k] Anonymous. Pages 1–116 (March 2016). Progress in Oceanography, 142(??):1–116, March 2016. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2016:PMb

[Ano16l] Anonymous. Pages 1–118 (May 2016). Progress in Oceanography, 144(??):1–118, May 2016. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2016:PJa

[Ano16m] Anonymous. Pages 1–154 (January 2016). Progress in Oceanography, 140(??):1–154, January 2016. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2016:PD

[Ano16n] Anonymous. Pages 1–214 (December 2016). Progress in Oceanography, 149(??):1–214, December 2016. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2016:PAb

[Ano16o] Anonymous. Pages 1–222 (August 2016). Progress in Oceanography, 146(??):1–222, August 2016. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2016:PF

[Ano16p] Anonymous. Pages 1–258 (February 2016). Progress in Oceanography, 141(??):1–258, February 2016. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2016:PJb

[Ano16q] Anonymous. Pages 1–56 (June 2016). Progress in Oceanography, 145(??):1–56, June 2016. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2016:PS

[Ano16r] Anonymous. Pages 1–68 (September 2016). Progress in Oceanography, 147(??):1–68, September 2016. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2016:PN

[Ano16s] Anonymous. Pages 1–74 (November 2016). Progress in Oceanography, 148(??):1–74, November 2016. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2016:PAa

[Ano16t] Anonymous. Pages 1–82 (April 2016). Progress in Oceanography, 143(??):1–82, April 2016. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2017:AMT

[Ano17a] Anonymous. The Atlantic Meridional Transect programme (1995–2016). Progress in Oceanography, 158(??):1–2, November 2017. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661117302884.

Anonymous:2017:IEPa

[Ano17b]  Anonymous.  IFC — eds/publication info.  Progress

in Oceanography, 150(??):ifc, January 2017. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661117300174.

Anonymous:2017:IEPb

[Ano17c]  Anonymous.  IFC — eds/publication info.  Progress

in Oceanography, 151(??):ifc, February 2017. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661117300411.

Anonymous:2017:IEPc

[Ano17d] Anonymous. IFC — eds/publication info. Progress in Oceanography, 152(??):ifc, March 2017. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661117301040.

Anonymous:2017:IFCd

[Ano17e] Anonymous. Inside front cover — editorial board page. Progress in Oceanography, 156(??):ifc, August 2017. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661117302628.

Anonymous:2017:IFCe

[Ano17f] Anonymous. Inside front cover — editorial board page. Progress in Oceanography, 157(??):ifc, September 2017. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661117303014.

Anonymous:2017:IFCf

[Ano17g] Anonymous. Inside front cover — editorial board page. Progress in Oceanography, 158(??):ifc, November 2017. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S007966111730318X.

Anonymous:2017:IFCg

[Ano17h] Anonymous. Inside front cover — editorial board page. Progress in Oceanography, 159(??):ifc, December 2017. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S007966111730349X.

Anonymous:2017:IFCa

[Ano17i] Anonymous. Inside front cover — Editorial Board page/cover image legend if applicable. Progress in Oceanography, 153(??): ifc, April 2017. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661117301441.

Anonymous:2017:IFCb

[Ano17j] Anonymous. Inside front cover — Editorial Board page/cover image legend if applicable. Progress in Oceanography, 154(??): ifc, May 2017. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661117301751.

Anonymous:2017:IFCc

[Ano17k] Anonymous. Inside front cover — Editorial Board page/cover image legend if applicable. Progress in Oceanography, 155(??): ifc, June 2017. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661117301970.

Anonymous:2017:PMb

[Ano17l] Anonymous. Pages 1–106 (May 2017). Progress in Oceanography, 154(??):1–106, May 2017. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2017:PF

[Ano17m] Anonymous. Pages 1–292 (February 2017). Progress in Oceanography, 151(??):1–292, February 2017. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2017:PAb

[Ano17n] Anonymous. Pages 1–326 (August 2017). Progress in Oceanography, 156(??):1–326, August 2017. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2017:PD

[Ano17o] Anonymous. Pages 1–340 (December 2017). Progress in Oceanography, 159(??):1–340, December 2017. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2017:PS

[Ano17p] Anonymous. Pages 1–72 (September 2017). Progress in Oceanography, 157(??):1–72, September 2017. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2017:PJ

[Ano17q] Anonymous. Pages 1–74 (June 2017). Progress in Oceanography, 155(??):1–74, June 2017. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2017:PAa

[Ano17r] Anonymous. Pages 1–84 (April 2017). Progress in Oceanography, 153(??):1–84, April 2017. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2017:PMa

[Ano17s] Anonymous. Pages 1–86 (March 2017). Progress in Oceanography, 152(??):1–86, March 2017. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2018:E

[Ano18a] Anonymous. Editorial. Progress in Oceanography, 166(??):1, September 2018. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661118302301.

Anonymous:2018:EBa

[Ano18b] Anonymous. Editorial Board. Progress in Oceanography, 160(??):ii, January 2018. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S007966111830017X.

Anonymous:2018:EBb

[Ano18c] Anonymous. Editorial Board. Progress in Oceanography, 161(??):ii, February 2018. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661118300430.

Anonymous:2018:EBc

[Ano18d] Anonymous. Editorial Board. Progress in Oceanography, 162(??):ii, March 2018. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661118300764.

Anonymous:2018:EBd

[Ano18e] Anonymous. Editorial Board. Progress in Oceanography, 163 (??):ii, April 2018. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661118301423.

Anonymous:2018:EBe

[Ano18f] Anonymous. Editorial Board. Progress in Oceanography, 166 (??):ii, September 2018. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661118302374.

Anonymous:2018:EBf

[Ano18g] Anonymous. Editorial Board. Progress in Oceanography, 167(??):ii, October 2018. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661118302726.

Anonymous:2018:EBg

[Ano18h] Anonymous. Editorial Board. Progress in Oceanography, 168(??):ii, November 2018. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661118303082.

Anonymous:2018:EBh

[Ano18i] Anonymous. Editorial Board. Progress in Oceanography, 169(??):ii, December 2018. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661118303355.

Anonymous:2018:PF

[Ano18j] Anonymous. Pages 1–152 (February 2018). Progress in Oceanography, 161(??):1–152, February 2018. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2018:PJ

[Ano18k] Anonymous. Pages 1–224 (January 2018). Progress in Oceanography, 160(??):1–224, January 2018. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2018:PO

[Ano18l] Anonymous. Pages 1–234 (October 2018). Progress in Oceanography, 167(??):1–234, October 2018. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2018:PM

[Ano18m] Anonymous. Pages 1–302 (March 2018). Progress in Oceanography, 162(??):1–302, March 2018. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2018:PN

[Ano18n] Anonymous. Pages 1–324 (November 2018). Progress in Oceanography, 168(??):1–324, November 2018. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2019:D

[Ano19a] Anonymous. December 2019. Progress in Oceanography, 179 (??):??, December 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2019:EBa

[Ano19b] Anonymous. Editorial Board. Progress in Oceanography, 170(??):ii, January 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661118303604.

Anonymous:2019:EBb

[Ano19c] Anonymous. Editorial Board. Progress in Oceanography, 171(??):ii, February 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661119300230.

Anonymous:2019:EBc

[Ano19d] Anonymous. Editorial Board. Progress in Oceanography, 172(??):ii, March 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661119300618.

Anonymous:2019:EBd

[Ano19e] Anonymous. Editorial Board. Progress in Oceanography, 173 (??):ii, April 2019. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661119301430.

Anonymous:2019:EBe

[Ano19f] Anonymous. Editorial Board. Progress in Oceanography, 174 (??):ii, May 2019. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661119302174.

Anonymous:2019:EBf

[Ano19g] Anonymous. Editorial Board. Progress in Oceanography, 176(??):??, September 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661119303374.

Anonymous:2019:EBg

[Ano19h] Anonymous. Editorial Board. Progress in Oceanography, 178 (??):??, November 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S007966111930388X.

Anonymous:2019:EBh

[Ano19i] Anonymous. Editorial Board. Progress in Oceanography, 179 (??):??, December 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661119304094.

Anonymous:2019:N

[Ano19j] Anonymous. November 2019. Progress in Oceanography, 178 (??):??, November 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2019:PJ

[Ano19k] Anonymous. Pages 1–146 (January 2019). Progress in Oceanography, 170(??):1–146, January 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2019:PM

[Ano19l] Anonymous. Pages 1–198 (March 2019). Progress in Oceanography, 172(??):1–198, March 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2019:PF

[Ano19m] Anonymous. Pages 1–252 (February 2019). Progress in Oceanography, 171(??):1–252, February 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2019:PA

[Ano19n] Anonymous. Pages 1–350 (April 2019). Progress in Oceanography, 173(??):1–350, April 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2019:PN

[Ano19o] Anonymous. Publisher’s note. Progress in Oceanography, 176(??):??, September 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661119302691.

Anonymous:2019:S

[Ano19p] Anonymous. September 2019. Progress in Oceanography, 176 (??):??, September 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2020:A

[Ano20a] Anonymous. April 2020. Progress in Oceanography, 183(??): ??, April 2020. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2020:AS

[Ano20b] Anonymous. August–September 2020. Progress in Oceanography, 187(??):??, August–September 2020. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2020:EBa

[Ano20c] Anonymous. Editorial Board. Progress in Oceanography, 180(??):??, January 2020. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661119304318.

Anonymous:2020:EBb

[Ano20d] Anonymous. Editorial Board. Progress in Oceanography, 181(??):??, February 2020. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661120300173.

Anonymous:2020:EBc

[Ano20e] Anonymous. Editorial Board. Progress in Oceanography, 182(??):??, March 2020. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661120300434.

Anonymous:2020:EBd

[Ano20f] Anonymous. Editorial Board. Progress in Oceanography, 183(??):??, April 2020. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661120300653.

Anonymous:2020:EBe

[Ano20g] Anonymous. Editorial Board. Progress in Oceanography, 184 (??):??, May 2020. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661120301002.

Anonymous:2020:EBf

[Ano20h] Anonymous. Editorial Board. Progress in Oceanography, 185(??):??, June 2020. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661120301245.

Anonymous:2020:EBg

[Ano20i] Anonymous. Editorial Board. Progress in Oceanography, 186 (??):??, July 2020. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661120301488.

Anonymous:2020:EBh

[Ano20j] Anonymous. Editorial Board. Progress in Oceanography, 187 (??):??, August–September 2020. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661120301695.

Anonymous:2020:EBi

[Ano20k] Anonymous. Editorial Board. Progress in Oceanography, 188(??):??, October 2020. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661120301981.

Anonymous:2020:EBj

[Ano20l] Anonymous. Editorial Board. Progress in Oceanography, 189(??):??, November–December 2020. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661120302238.

Anonymous:2020:F

[Ano20m] Anonymous. February 2020. Progress in Oceanography, 181 (??):??, February 2020. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2020:Ja

[Ano20n] Anonymous. January 2020. Progress in Oceanography, 180(??): ??, January 2020. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2020:Jc

[Ano20o] Anonymous. July 2020. Progress in Oceanography, 186(??): ??, July 2020. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2020:Jb

[Ano20p] Anonymous. June 2020. Progress in Oceanography, 185(??): ??, June 2020. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2020:Ma

[Ano20q] Anonymous. March 2020. Progress in Oceanography, 182(??): ??, March 2020. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2020:Mb

[Ano20r] Anonymous. May 2020. Progress in Oceanography, 184(??): ??, May 2020. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2020:ND

[Ano20s] Anonymous. November–December 2020. Progress in Oceanography, 189(??):??, November–December 2020. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2020:O

[Ano20t] Anonymous. October 2020. Progress in Oceanography, 188(??): ??, October 2020. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2020:TMD

[Ano20u] Anonymous. A tribute to the memory of Dr. Vladimir Ivanovich Kharlamenko. Progress in Oceanography, 183(??): ??, April 2020. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661120300501.

Anonymous:2021:AM

[Ano21a] Anonymous. April–May 2021. Progress in Oceanography, 193 (??):??, April–May 2021. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2021:A

[Ano21b] Anonymous. August 2021. Progress in Oceanography, 196(??): ??, August 2021. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2021:D

[Ano21c] Anonymous. December 2021. Progress in Oceanography, 199 (??):??, December 2021. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2021:EBa

[Ano21d] Anonymous. Editorial Board. Progress in Oceanography, 190(??):??, January 2021. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661120302391.

Anonymous:2021:EBb

[Ano21e] Anonymous. Editorial Board. Progress in Oceanography, 191(??):??, February 2021. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661121000161.

Anonymous:2021:EBc

[Ano21f] Anonymous. Editorial Board. Progress in Oceanography, 192(??):??, March 2021. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661121000276.

Anonymous:2021:EBd

[Ano21g] Anonymous. Editorial Board. Progress in Oceanography, 193(??):??, April–May 2021. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661121000677.

Anonymous:2021:EBe

[Ano21h] Anonymous. Editorial Board. Progress in Oceanography, 194(??):??, June 2021. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661121000884.

Anonymous:2021:EBf

[Ano21i] Anonymous. Editorial Board. Progress in Oceanography, 195 (??):??, July 2021. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661121001117.

Anonymous:2021:EBg

[Ano21j] Anonymous. Editorial Board. Progress in Oceanography, 196(??):??, August 2021. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661121001336.

Anonymous:2021:EBh

[Ano21k] Anonymous. Editorial Board. Progress in Oceanography, 197(??):??, September–October 2021. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661121001646.

Anonymous:2021:EBi

[Ano21l] Anonymous. Editorial Board. Progress in Oceanography, 198 (??):??, November 2021. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661121001853.

Anonymous:2021:EBj

[Ano21m] Anonymous. Editorial Board. Progress in Oceanography, 199 (??):??, December 2021. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661121002044.

Anonymous:2021:F

[Ano21n] Anonymous. February 2021. Progress in Oceanography, 191 (??):??, February 2021. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2021:Ja

[Ano21o] Anonymous. January 2021. Progress in Oceanography, 190(??): ??, January 2021. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2021:Jc

[Ano21p] Anonymous. July 2021. Progress in Oceanography, 195(??): ??, July 2021. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2021:Jb

[Ano21q] Anonymous. June 2021. Progress in Oceanography, 194(??): ??, June 2021. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2021:M

[Ano21r] Anonymous. March 2021. Progress in Oceanography, 192(??): ??, March 2021. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2021:N

[Ano21s] Anonymous. November 2021. Progress in Oceanography, 198 (??):??, November 2021. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2021:SO

[Ano21t] Anonymous. September–October 2021. Progress in Oceanography, 197(??):??, September–October 2021. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2022:AM

[Ano22a] Anonymous. April–May 2022. Progress in Oceanography, 203 (??):??, April–May 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2022:A

[Ano22b] Anonymous. August 2022. Progress in Oceanography, 206(??): ??, August 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2022:D

[Ano22c] Anonymous. December 2022. Progress in Oceanography, 209 (??):??, December 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2022:EBa

[Ano22d] Anonymous. Editorial Board. Progress in Oceanography, 80(1– 2):ifc, ???? 2022. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661109000056.

Anonymous:2022:EBb

[Ano22e] Anonymous. Editorial Board. Progress in Oceanography, 88(1– 4):ifc, ???? 2022. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661111000073.

Anonymous:2022:EBc

[Ano22f] Anonymous. Editorial Board. Progress in Oceanography, 112–113(??):ifc, ???? 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661113000669.

Anonymous:2022:EBd

[Ano22g] Anonymous. Editorial Board. Progress in Oceanography, 164 (??):ii, ???? 2022. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661118301502.

Anonymous:2022:EBe

[Ano22h] Anonymous. Editorial Board. Progress in Oceanography, 165 (??):ii, ???? 2022. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661118302131.

Anonymous:2022:EBf

[Ano22i] Anonymous. Editorial Board. Progress in Oceanography, 175 (??):ii, ???? 2022. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661119302630.

Anonymous:2022:EBg

[Ano22j] Anonymous. Editorial Board. Progress in Oceanography, 177(??):??, October 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661119303544.

Anonymous:2022:EBh

[Ano22k] Anonymous. Editorial Board. Progress in Oceanography, 200(??):??, January 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661121002159.

Anonymous:2022:EBi

[Ano22l] Anonymous. Editorial Board. Progress in Oceanography, 201(??):??, February 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661122000131.

Anonymous:2022:EBj

[Ano22m] Anonymous. Editorial Board. Progress in Oceanography, 202(??):??, March 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661122000337.

Anonymous:2022:EBk

[Ano22n] Anonymous. Editorial Board. Progress in Oceanography, 203(??):??, April–May 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S007966112200060X.

Anonymous:2022:EBl

[Ano22o] Anonymous. Editorial Board. Progress in Oceanography, 204(??):??, June 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661122000751.

Anonymous:2022:EBm

[Ano22p] Anonymous. Editorial Board. Progress in Oceanography, 205 (??):??, July 2022. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661122000945.

Anonymous:2022:EBn

[Ano22q] Anonymous. Editorial Board. Progress in Oceanography, 206(??):??, August 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL https://www.

sciencedirect.com/science/article/pii/S007966112200129X.

Anonymous:2022:EBo

[Ano22r] Anonymous. Editorial Board. Progress in Oceanography, 207(??):??, September–October 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661122001471.

Anonymous:2022:EBp

[Ano22s] Anonymous. Editorial Board. Progress in Oceanography, 208 (??):??, November 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661122001720.

Anonymous:2022:EBq

[Ano22t] Anonymous. Editorial Board. Progress in Oceanography, 209 (??):??, December 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S007966112200194X.

Anonymous:2022:F

[Ano22u] Anonymous. February 2022. Progress in Oceanography, 201 (??):??, February 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2022:Ja

[Ano22v] Anonymous. January 2022. Progress in Oceanography, 200(??): ??, January 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2022:Jc

[Ano22w] Anonymous. July 2022. Progress in Oceanography, 205(??): ??, July 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2022:Jb

[Ano22x] Anonymous. June 2022. Progress in Oceanography, 204(??): ??, June 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2022:M

[Ano22y] Anonymous. March 2022. Progress in Oceanography, 202(??): ??, March 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2022:N

[Ano22z] Anonymous. November 2022. Progress in Oceanography, 208 (??):??, November 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2022:PJF

[Ano22-27] Anonymous. Pages 1–112 (January–February 2009). Progress in Oceanography, 80(1–2):??, ???? 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2022:PJM

[Ano22-28] Anonymous. Pages 1–150 (January–March 2011). Progress in Oceanography, 88(1–4):??, ???? 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2022:PJAa

[Ano22-29] Anonymous. Pages 1–348 (July–August 2018). Progress in Oceanography, 165(??):1–348, ???? 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2022:PJAb

[Ano22-30] Anonymous. Pages 1–358 (July–August 2019). Progress in Oceanography, 175(??):1–358, ???? 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2022:PMJa

[Ano22-31] Anonymous. Pages 1–60 (May–June 2013). Progress in Oceanography, 112–113(??):1–60, ???? 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2022:PMJb

[Ano22-32] Anonymous. Pages 1–88 (May–June 2018). Progress in Oceanography, 164(??):1–88, ???? 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2022:PND

[Ano22-33] Anonymous. Pages 263–384 (November–December 2002). Progress in Oceanography, 55(3–4):??, ???? 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2022:SO

[Ano22-34] Anonymous. September–October 2022. Progress in Oceanography, 207(??):??, September–October 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2023:AM

[Ano23a] Anonymous. April–May 2023. Progress in Oceanography, 213 (??):??, April–May 2023. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2023:A

[Ano23b] Anonymous. August 2023. Progress in Oceanography, 216(??): ??, August 2023. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2023:D

[Ano23c] Anonymous. December 2023. Progress in Oceanography, 219 (??):??, December 2023. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2023:EBa

[Ano23d] Anonymous. Editorial Board. Progress in Oceanography, 210(??):??, January 2023. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661123000046.

Anonymous:2023:EBb

[Ano23e] Anonymous. Editorial Board. Progress in Oceanography, 211(??):??, February 2023. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661123000332.

Anonymous:2023:EBc

[Ano23f] Anonymous. Editorial Board. Progress in Oceanography, 212(??):??, March 2023. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661123000538.

Anonymous:2023:EBd

[Ano23g] Anonymous. Editorial Board. Progress in Oceanography, 213(??):??, April–May 2023. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661123000708.

Anonymous:2023:EBe

[Ano23h] Anonymous. Editorial Board. Progress in Oceanography, 214(??):??, June 2023. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661123000836.

Anonymous:2023:EBf

[Ano23i] Anonymous. Editorial Board. Progress in Oceanography, 215 (??):??, July 2023. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661123001040.

Anonymous:2023:EBg

[Ano23j] Anonymous. Editorial Board. Progress in Oceanography, 216(??):??, August 2023. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661123001349.

Anonymous:2023:EBh

[Ano23k] Anonymous. Editorial Board. Progress in Oceanography, 217(??):??, September–October 2023. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S007966112300157X.

Anonymous:2023:EBi

[Ano23l] Anonymous. Editorial Board. Progress in Oceanography, 218 (??):??, November 2023. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661123001829.

Anonymous:2023:EBj

[Ano23m] Anonymous. Editorial Board. Progress in Oceanography, 219 (??):??, December 2023. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661123002252.

Anonymous:2023:F

[Ano23n] Anonymous. February 2023. Progress in Oceanography, 211 (??):??, February 2023. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2023:Ja

[Ano23o] Anonymous. January 2023. Progress in Oceanography, 210(??): ??, January 2023. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2023:Jc

[Ano23p] Anonymous. July 2023. Progress in Oceanography, 215(??): ??, July 2023. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2023:Jb

[Ano23q] Anonymous. June 2023. Progress in Oceanography, 214(??): ??, June 2023. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2023:Ma

[Ano23r] Anonymous. March 2023. Progress in Oceanography, 212(??): ??, March 2023. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2023:N

[Ano23s] Anonymous. November 2023. Progress in Oceanography, 218 (??):??, November 2023. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2023:SO

[Ano23t] Anonymous. September–October 2023. Progress in Oceanography, 217(??):??, September–October 2023. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2024:AM

[Ano24a] Anonymous. April–May 2024. Progress in Oceanography, 223 (??):??, April–May 2024. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2024:EBa

[Ano24b] Anonymous. Editorial Board. Progress in Oceanography, 220(??):??, January 2024. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661124000053.

Anonymous:2024:EBb

[Ano24c] Anonymous. Editorial Board. Progress in Oceanography, 221(??):??, February 2024. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661124000260.

Anonymous:2024:EBc

[Ano24d] Anonymous. Editorial Board. Progress in Oceanography, 222(??):??, March 2024. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661124000442.

Anonymous:2024:EBd

[Ano24e] Anonymous. Editorial Board. Progress in Oceanography, 223(??):??, April–May 2024. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661124000612.

Anonymous:2024:EBe

[Ano24f] Anonymous. Editorial Board. Progress in Oceanography, 224(??):??, June 2024. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661124000788.

Anonymous:2024:EBf

[Ano24g] Anonymous. Editorial Board. Progress in Oceanography, 225 (??):??, July 2024. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661124000995.

Anonymous:2024:F

[Ano24h] Anonymous. February 2024. Progress in Oceanography, 221 (??):??, February 2024. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2024:Ja

[Ano24i] Anonymous. January 2024. Progress in Oceanography, 220(??): ??, January 2024. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2024:Jc

[Ano24j] Anonymous. July 2024. Progress in Oceanography, 225(??): ??, July 2024. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2024:Jb

[Ano24k] Anonymous. June 2024. Progress in Oceanography, 224(??): ??, June 2024. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Anonymous:2024:M

[Ano24l] Anonymous. March 2024. Progress in Oceanography, 222(??): ??, March 2024. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

Antezana:2009:SSP

[Ant09] Tarsicio Antezana. Species-specific patterns of diel migration into the oxygen minimum zone by euphausiids in the Humboldt Current ecosystem. Progress in Oceanography, 83 (1–4):228–236, December 2009. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661109000834.

Angel:1979:E

[AO79] Martin Angel and James O’Brien. Editorial. Progress in Oceanography, 8(1–2):1, ???? 1979. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0079661179900089.

Angel:1984:E

[AO84] Martin Angel and James J. O’Brien. Editorial. Progress in Oceanography, 13(1):i, ???? 1984. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0079661184900028.

Anglada-Ortiz:2023:SMC

[AOMZ^{+}23] Griselda Anglada-Ortiz, Julie Meilland, Patrizia Ziveri, Melissa Chierici, Agneta Fransson, Elizabeth Jones, and Tine L. Rasmussen. Seasonality of marine calcifiers in the northern Barents Sea: Spatiotemporal distribution of planktonic foraminifers and shelled pteropods and their contribution to carbon dynamics. Progress in Oceanography, 218(??):??, November 2023. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661123001647.

Alalykina:2020:NDS

[AP20] Inna L. Alalykina and Neonila E. Polyakova. New deepsea species of Anobothrus (Annelida: Ampharetidae) from the Kuril–Kamchatka Trench and adjacent abyssal regions. Progress in Oceanography, 182(??):??, March 2020. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661119304173.

Alheit:2012:CVD

[APC^{+}12] Jürgen Alheit, Thomas Pohlmann, Michele Casini, Wulf Greve, Rosemarie Hinrichs, Moritz Mathis, Kieran O’Driscoll, Ralf Vorberg, and Carola Wagner. Climate variability drives anchovies and sardines into the North and Baltic Seas. Progress in Oceanography, 96(1):128–139, April 2012. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661111001418.

Aguiar:2013:CML

[APC13] Ana Cláudia Barbosa Aguiar, Álvaro Peliz, and Xavier Carton. A census of meddies in a long-term high-resolution simulation. Progress in Oceanography, 116(??):80–94, September 2013. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661113001043.

Ashjian:2021:SRZ

[APC^{+}21] Carin J. Ashjian, Robert S. Pickart, Robert G. Campbell, Zhixuan Feng, Cecilia Gelfman, Philip Alatalo, and Jinlun Zhang. Springtime renewal of zooplankton populations in the Chukchi Sea. Progress in Oceanography, 197(??):??, September– October 2021. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661121001208.

Arumi-Planas:2022:VMO

[APHGC^{+}22] Cristina Arumí-Planas, Alonso Hernández-Guerra, Verónica Caínzos, Pedro Vélez-Belchí, Riccardo Farneti, Matthew R. Mazloff, Sabine Mecking, Isabella Rosso, Lena M. Schulze Chretien, Kevin G. Speer, and Lynne D. Talley. Variability in the meridional overturning circulation at 32^{o}S in the Pacific Ocean diagnosed by inverse box models. Progress in Oceanography, 203(??):??, April–May 2022. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661122000416.

Aguiar:2015:MOT

[APN^{+}15] A. C. Barbosa Aguiar, A. Peliz, F. Neves, I. Bashmachnikov, and X. Carton. Mediterranean outflow transports and entrainment estimates from observations and high-resolution modelling. Progress in Oceanography, 131(??):33–45, February 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661114001979.

Aubone:2021:SSM

[APP21] N. Aubone, E. D. Palma, and A. R. Piola. The surface salinity maximum of the South Atlantic. Progress in Oceanography, 191(??):??, February 2021. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661120302342.

Aoyama:2011:SHO

[APSC11] Michio Aoyama, Pavel P. Povinec, and Joan-Albert Sanchez-Cabeza. The Southern Hemisphere Ocean Tracer Studies (SHOTS) project. Progress in Oceanography, 89(1–4):1–6, April 2011. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661110001540.

Alemany:2010:CSH

[AQVB^{+}10] F. Alemany, L. Quintanilla, P. Velez-Belchí, A. García, D. Cortés, J. M. Rodríguez, M. L. Fernández de Puelles, C. González-Pola, and J. L. López-Jurado. Characterization of the spawning habitat of Atlantic bluefin tuna and related species in the Balearic Sea (western Mediterranean). Progress in Oceanography, 86(1–2):21–38, July 2010. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661110000480.

Alekseenko:2018:NSW

[AR18] E. Alekseenko and B. Roux. Numerical simulation of the wind influence on bottom shear stress and salinity fields in areas of zostera noltei replanting in a Mediterranean coastal lagoon. Progress in Oceanography, 163(??):147–160, April 2018. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661116302397.

Arbic:2022:ITI

[Arb22] Brian K. Arbic. Incorporating tides and internal gravity waves within global ocean general circulation models: a review. Progress in Oceanography, 206(??):??, August 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661122000854.

Agenbag:2003:EEP

[ARD^{+}03] J. J. Agenbag, A. J. Richardson, H. Demarcq, P. Fréon, S. Weeks, and F. A. Shillington. Estimating environmental preferences of South African pelagic fish species using catch size- and remote sensing data. Progress in Oceanography, 59(2–3):275–300, October 2003. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661103001708.

Andutta:2014:CER

[ARDP14] Fernando P. Andutta, Peter V. Ridd, Eric Deleersnijder, and David Prandle. Contaminant exchange rates in estuaries — new formulae accounting for advection and dispersion. Progress in Oceanography, 120(??):139–153, January 2014. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661113001699.

Arbilla:2024:CSS

[ARELAK24] Lisandro A. Arbilla, Laura A. Ruiz-Etcheverry, Celeste López-Abbate, and Lucía C. Kahl. CO_{2} sink and source zones delimited by marine fronts in the Drake Passage. Progress in Oceanography, 223(??):??, April–May 2024. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661124000521.

Auad:2011:CCS

[ARG11] Guillermo Auad, Dean Roemmich, and John Gilson. The California Current System in relation to the Northeast Pacific Ocean circulation. Progress in Oceanography, 91(4):576–592, December 2011. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661111001157.

Aiken:2000:AMT

[ARH^{+}00] J. Aiken, N. Rees, S. Hooker, P. Holligan, A. Bale, D. Robins, G. Moore, R. Harris, and D. Pilgrim. The Atlantic Meridional Transect: overview and synthesis of data. Progress in Oceanography, 45(3–4):257–312, April 2000. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661100000057.

Alldredge:1988:CDS

[AS88] Alice L. Alldredge and Mary W. Silver. Characteristics, dynamics and significance of marine snow. Progress in Oceanography, 20(1):41–82, ???? 1988. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0079661188900535.

Angel:1989:E

[AS89] M. V. Angel and R. L. Smith. Editorial. Progress in Oceanography, 22(1):i, ???? 1989. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0079661189900086.

Allen:1996:ZDO

[AS96] J. A. Allen and H. L. Sanders. The zoogeography, diversity and origin of the deep-sea protobranch bivalves of the Atlantic: the epilogue. Progress in Oceanography, 38(2):95– 153, ???? 1996. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661196000110.

Aksentov:2020:MGD

[AS20] Kirill I. Aksentov and Valentina V. Sattarova. Mercury geochemistry of deep-sea sediment cores from the Kuril area, northwest Pacific. Progress in Oceanography, 180(??):??, January 2020. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S007966111930415X.

Alvarez-Salgado:2014:MBM

[ÁSÁB^{+}14] X. A. Álvarez-Salgado, M. Álvarez, S. Brea, L. Mèmery, and M. J. Messias. Mineralization of biogenic materials in the water masses of the South Atlantic Ocean. II: Stoichiometric ratios and mineralization rates. Progress in Oceanography, 123(??): 24–37, April 2014. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661114000032.

Ayon:2008:ZFF

[ASB^{+}08] Patricia Ayón, Gordon Swartzman, Arnaud Bertrand, Mariano Gutiérrez, and Sophie Bertrand. Zooplankton and forage fish species off Peru: Large-scale bottom-up forcing and localscale depletion. Progress in Oceanography, 79(2–4):208–214, October 2008. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661108001742.

Azam:1992:BTT

[ASC92] Farooq Azam, David C. Smith, and Angelo F. Carlucci. Bacterial transformation and transport of organic matter in the Southern California Bight. Progress in Oceanography, 30(1–4):151–166, ???? 1992. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/007966119290011N.

Alcaraz:2007:CBS

[ASC07] Miguel Alcaraz, Enric Saiz, and Albert Calbet. Centropages behaviour: Swimming and vertical migration. Progress in Oceanography, 72(2–3):121–136, February 2007. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S007966110700002X.

Alvarez-Salgado:2001:SFL

[ÁSDB^{+}01] X. A. Álvarez-Salgado, M. D. Doval, A. V. Borges, I. Joint, M. Frankignoulle, E. M. S. Woodward, and F. G. Figueiras. Off-shelf fluxes of labile materials by an upwelling filament in the NW Iberian Upwelling System. Progress in Oceanography, 51(2–4):321–337, November 2001. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661101000738.

Ahumada-Sempoal:2013:NCV

[ASFB^{+}13] M.-A. Ahumada-Sempoal, M. M. Flexas, R. Bernardello, N. Bahamon, and A. Cruzado. Northern Current variability and its impact on the Blanes Canyon circulation: a numerical study. Progress in Oceanography, 118(??):61–70, November 2013. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661113001432.

Alvarez-Salgado:2003:PCC

[ÁSFP^{+}03] X. A. Álvarez-Salgado, F. G. Figueiras, F. F. Pérez, S. Groom, E. Nogueira, A. V. Borges, L. Chou, C. G. Castro, G. Moncoiffé, A. F. Ríos, A. E. J. Miller, M. Frankignoulle, G. Savidge, and R. Wollast. The Portugal coastal counter current off NW Spain: new insights on its biogeochemical variability. Progress in Oceanography, 56(2):281–321, February 2003. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661103000077.

Alexander:2023:SCC

[ASJ^{+}23] Michael A. Alexander, James D. Scott, Michael G. Jacox, Clara Deser, Dillon J. Amaya, Antonietta Capotondi, and Adam S. Phillips. A survey of coastal conditions around the continental US using a high-resolution ocean reanalysis. Progress in Oceanography, 216(??):??, August 2023. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661123000988.

Assuncao:2020:CTS

[ASR^{+}20] Ramilla V. Assunção, Alex C. Silva, Amédée Roy, Bernard Bourlès, Carlos Henrique S. Silva, Jean-François Ternon, Moacyr Araujo, and Arnaud Bertrand. 3D characterisation of the thermohaline structure in the southwestern tropical Atlantic derived from functional data analysis of in situ profiles. Progress in Oceanography, 187(??):??, August–September 2020. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661120301385.

Asahi:2007:YTS

[AT07] Hirofumi Asahi and Kozo Takahashi. A 9-year time-series of planktonic foraminifer fluxes and environmental change in the Bering Sea and the central subarctic Pacific Ocean, 1990–1999. Progress in Oceanography, 72(4):343–363, March 2007. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661106001339.

Annasawmy:2019:MDA

[ATC^{+}19] Pavanee Annasawmy, Jean-François Ternon, Pascal Cotel, Yves Cherel, Evgeny V. Romanov, Gildas Roudaut, Anne Lebourges-Dhaussy, Fréderic Ménard, and Francis Marsac. Micronekton distributions and assemblages at two shallow seamounts of the south-western Indian Ocean: Insights from acoustics and mesopelagic trawl data. Progress in Oceanography, 178(??):??, November 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661118302313.

Alexander:2001:WWR

[ATS01] Michael A. Alexander, Michael S. Timlin, and James D. Scott. Winter-to-winter recurrence of sea surface temperature, salinity and mixed layer depth anomalies. Progress in Oceanography, 49(??):41–61, ???? 2001. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661101000155.

Arguelles:2008:SIJ

[ATT^{+}08] Juan Argüelles, Ricardo Tafur, Anatolio Taipe, Piero Villegas, Friedeman Keyl, Noel Dominguez, and Martín Salazar. Size increment of jumbo flying squid Dosidicus gigas mature females in Peruvian waters, 1989–2004. Progress in Oceanography, 79(2–4):308–312, October 2008. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661108001833.

Auger:2014:ICP

[AUE^{+}14] P. A. Auger, C. Ulses, C. Estournel, L. Stemmann, S. Somot, and F. Diaz. Interannual control of plankton communities by deep winter mixing and prey/predator interactions in the NW Mediterranean: Results from a 30-year 3D modeling study. Progress in Oceanography, 124(??):12–27, May 2014. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661114000524.

Andersen:2023:CSF

[AV23] K. H. Andersen and A. W. Visser. From cell size and first principles to structure and function of unicellular plankton communities. Progress in Oceanography, 213(??):??, April– May 2023. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661123000381.

Arrigo:2015:CIA

[AvD15] Kevin R. Arrigo and Gert L. van Dijken. Continued increases in Arctic Ocean primary production. Progress in Oceanography, 136(??):60–70, August 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661115000993.

Artemova:2019:COC

[AVG^{+}19] Antonina V. Artemova, Yuriy P. Vasilenko, Sergey A. Gorbarenko, Aleksandr A. Bosin, and Valentina V. Sattarova. Climatic and oceanological changes in the southwestern part of the Sea of Okhotsk during the last 94 kyr.

Progress in Oceanography, 179(??):??, December 2019. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661119303957.

Armstrong:1991:STV

[AVK91] D. A. Armstrong, H. M. Verheye, and A. D. Kemp. Shortterm variability during an anchor station study in the southern Benguela upwelling system: Fecundity estimates of the dominant copepod, Calanoides carinatus. Progress in Oceanography, 28(1–2):167–188, ???? 1991. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/007966119190025H.

Antivachis:2023:LCS

[AVS23] Dimitrios Antivachis, Vassilios Vervatis, and Sarantis Sofianos. Lagrangian Coherent Structures in the Mediterranean Sea: Seasonality and basin regimes. Progress in Oceanography, 215 (??):??, July 2023. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661123000940.

Allen:2013:HTS

[AW13] S. E. Allen and M. A. Wolfe. Hindcast of the timing of the spring phytoplankton bloom in the Strait of Georgia, 1968–2010. Progress in Oceanography, 115(??):6–13, August 2013. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661113000773.

Aoki:2023:DSO

[AYH^{+}23] Shigeru Aoki, Kaihe Yamazaki, Daisuke Hirano, Daiki Nomura, and Hiroto Murase. Distribution of stable oxygen isotope in seawater and implication on freshwater cycle off the coast from Wilkes to George V Land, East Antarctica. Progress in Oceanography, 217(??):??, September–October 2023. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661123001441.

Aono:2005:EFP

[AYK^{+}05] Tatsuo Aono, Masatoshi Yamada, Isao Kudo, Keiri Imai, Yukihiro Nojiri, and Atsushi Tsuda. Export fluxes of particulate organic carbon estimated from ^{234}Th/^{238}U disequilibrium during the Subarctic Pacific Iron Experiment for Ecosystem Dynamics Study (SEEDS 2001). Progress in Oceanography, 64(2–4):263–282, February 2005. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661105000340.

Barton:2004:CIC

[BA04] E. D. Barton and J. Arístegui. The Canary Islands coastal transition zone — upwelling, eddies and filaments. Progress in Oceanography, 62(2–4):67–69, August 2004. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661104000916.

Barth:2005:TWN

[BAARB05] A. Barth, A. Alvera-Azcárate, M. Rixen, and J.-M. Beckers. Two-way nested model of mesoscale circulation features in the Ligurian Sea. Progress in Oceanography, 66(2–4):171–189, August 2005. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661105000698.

Brandt:2019:DZN

[BAB^{+}19] Angelika Brandt, Inna Alalykina, Saskia Brix, Nils Brenke, Magdalena Błażewicz, Olga A. Golovan, Nele Johannsen, Anna M. Hrinko, Anna M. Jażdżewska, Karen Jeskulke, Gennady M. Kamenev, Anna V. Lavrenteva, Marina V. Malyutina, Torben Riehl, and Lidia Lins. Depth zonation of Northwest Pacific deep-sea macrofauna. Progress in Oceanography, 176(??):??, September 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S007966111830301X.

Baker:1983:BRT

[Bak83] Trevor Baker. Book review: The tides of the planet earth: 2nd Edition, by Paul Melchior, Pergamon Press, 1983, 648pp., 142 illus, 3467 lit refs. £45.00 US$90 — ISBN 0-08-026248-1 (Hardcover). Progress in Oceanography, 12(4):i–ii, ???? 1983. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0079661183900034.

Bakun:2001:LSM

[Bak01] Andrew Bakun. ‘School-mix feedback’: a different way to think about low frequency variability in large mobile fish populations. Progress in Oceanography, 49(??):485–511, ???? 2001. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661101000374.

Bakun:2006:WWP

[Bak06] Andrew Bakun. Wasp-waist populations and marine ecosystem dynamics: Navigating the “predator pit” topographies. Progress in Oceanography, 68(2–4):271–288, February 2006. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661106000097.

Baltar:2009:MVM

[BAM^{+}09] Federico Baltar, Javier Arístegui, María F. Montero, Minerva Espino, Josep M. Gasol, and Gerhard J. Herndl. Mesoscale variability modulates seasonal changes in the trophic structure of nano- and picoplankton communities across the NW Africa–Canary Islands transition zone. Progress in Oceanography, 83(1–4):180–188, December 2009. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661109000706.

Banse:1964:VDZ

[Ban64] K. Banse. On the vertical distribution of zooplankton in the sea. Progress in Oceanography, 2(??):53, 55–125, ???? 1964. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0079661164900035.

Bandy:1965:FDB

[Ban65] Orville L. Bandy. Foraminiferal definition of the boundaries of the Pleistocene in Southern California, U.S.A. Progress in Oceanography, 4(??):27–49, September 2, 1965. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/007966116590039X.

Banse:1996:LSL

[Ban96] Karl Banse. Low seasonality of low concentrations of surface chlorophyll in the subantarctic water ring: underwater irradiance, iron, or grazing? Progress in Oceanography, 37(3–4):241– 291, ???? 1996. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661196000067.

Bode:2007:SNI

[BAOC^{+}07] Antonio Bode, M. Teresa Alvarez-Ossorio, M. Emilia Cunha, Susana Garrido, J. Benito Peleteiro, Carmela Porteiro, Luis Valdés, and Manuel Varela. Stable nitrogen isotope studies of the pelagic food web on the Atlantic shelf of the Iberian Peninsula. Progress in Oceanography, 74(2–3):115–131, August 2007. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S007966110700064X.

Bode:2009:RTP

[BAOC^{+}09] A. Bode, M. T. Alvarez-Ossorio, J. M. Cabanas, A. Miranda, and M. Varela. Recent trends in plankton and upwelling intensity off Galicia (NW Spain). Progress in Oceanography, 83(1–4):342–350, December 2009. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661109000810.

Bode:2012:CCT

[BAOM^{+}12] Antonio Bode, Maria Teresa Alvarez-Ossorio, Ana Miranda, Angel López-Urrutia, and Luis Valdés. Comparing copepod time-series in the north of Spain: Spatial autocorrelation of community composition. Progress in Oceanography, 97–100(??):108–119, May 2012. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661111001315.

Brochier:2022:CSP

[BAP^{+}22] Timothée Brochier, Pierre-Amaël Auger, Laure Pecquerie, Eric Machu, Xavier Capet, Modou Thiaw, Baye Cheikh Mbaye, Cheikh-Baye Braham, Omar Ettahiri, Najib Charouki, Ousseynou Ndaw Sène, Francisco Werner, and Patrice Brehmer. Complex small pelagic fish population patterns arising from individual behavioral responses to their environment. Progress in Oceanography, 164(??):12–27, ???? 2022. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661117303737.

Briseno-Avena:2020:TDC

[BASS^{+}20] Christian Briseño-Avena, Moritz S. Schmid, Kelsey Swieca, Su Sponaugle, Richard D. Brodeur, and Robert K. Cowen. Three-dimensional cross-shelf zooplankton distributions off the Central Oregon Coast during anomalous oceanographic conditions. Progress in Oceanography, 188(??):??, October 2020. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661120301750.

Barton:1998:TZC

[BAT^{+}98] E. D. Barton, J. Arıśtegui, P. Tett, M. Cantón, J. Garcıá-Braun, S. Hernández-León, L. Nykjaer, C. Almeida, J. Almunia, S. Ballesteros, G. Basterretxea, J. Escánez, L. García-Weill, A. Hernández-Guerra, F. López-Laatzen, R. Molina, M. F. Montero, E. Navarro-Pérez, J. M. Rodríguez, K. van Lenning, H. Vélez, and K. Wild. The transition zone of the Canary Current upwelling region. Progress in Oceanography, 41(4):455–504, October 1998. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661198000238.

Barton:2004:VCI

[BATNP04] Eric D. Barton, Javier Arístegui, Paul Tett, and Eleuteria Navarro-Pérez. Variability in the Canary Islands area of filament-eddy exchanges. Progress in Oceanography, 62(2– 4):71–94, August 2004. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661104000928.

Barnes:1965:EMD

[BB65] H. Barnes and Margaret Barnes. Elminius modestus Darwin: Further European records. Progress in Oceanography, 3(??): 23–30, ???? 1965. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661165900066.

Buckley:2010:TLO

[BB10] Lawrence J. Buckley and Lauren B. Buckley. Toward linking ocean models to fish population dynamics. Progress in Oceanography, 84(1–2):85–88, January 2010. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661109001426.

Biller:2014:CCC

[BB14] Dondra V. Biller and Kenneth W. Bruland. The central California Current transition zone: a broad region exhibiting evidence for iron limitation. Progress in Oceanography, 120(??):370–382, January 2014. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661113002218.

Baumas:2024:FDT

[BB24] Chloé Baumas and Mina Bizic. A focus on different types of organic matter particles and their significance in the open ocean carbon cycle. Progress in Oceanography, 224(??):??, June 2024. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661124000399.

Benetazzo:2014:RAS

[BBB^{+}14] A. Benetazzo, A. Bergamasco, D. Bonaldo, F. M. Falcieri, M. Sclavo, L. Langone, and S. Carniel. Response of the Adriatic Sea to an intense cold air outbreak: Dense water dynamics and wave-induced transport. Progress in Oceanography, 128(??):115–138, November 2014. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661114001487.

Benetazzo:2021:EVS

[BBB^{+}21] Alvise Benetazzo, Francesco Barbariol, Filippo Bergamasco, Luciana Bertotti, Jeseon Yoo, Jae-Seol Shim, and Luigi Cavaleri. On the extreme value statistics of spatio-temporal maximum sea waves under cyclone winds. Progress in Oceanography, 197(??):??, September–October 2021. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661121001270.

Barbariol:2022:MWH

[BBC^{+}22] Francesco Barbariol, Jean-Raymond Bidlot, Luigi Cavaleri, Mauro Sclavo, Jim Thomson, and Alvise Benetazzo. Maximum wave heights from global model reanalysis. Progress in Oceanography, 175(??):139–160, ???? 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661118302763.

Beare:2003:SST

[BBE^{+}03] D. J. Beare, S. D. Batten, M. Edwards, E. McKenzie, P. C. Reid, and D. G. Reid. Summarising spatial and temporal information in CPR data. Progress in Oceanography, 58(2–4):217– 233, August 2003. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661103001307.

Babin:2015:EPP

[BBE^{+}15] M. Babin, S. Bélanger, I. Ellingsen, A. Forest, V. Le Fouest, T. Lacour, M. Ardyna, and D. Slagstad. Estimation of primary production in the Arctic Ocean using ocean colour remote sensing and coupled physical–biological models: Strengths, limitations and how they compare. Progress in Oceanography, 139(??):197–220, December 2015. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661115001883.

Burchard:2022:KTT

[BBF^{+}22] Hans Burchard, Karsten Bolding, Rainer Feistel, Ulf Gräwe, Knut Klingbeil, Parker MacCready, Volker Mohrholz, Lars Umlauf, and Eefke M. van der Lee. The Knudsen theorem and the Total Exchange Flow analysis framework applied to the Baltic Sea. Progress in Oceanography, 165(??):268–286, ???? 2022. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661117303774.

Bober:2019:DDI

[BBFS19] Johanna Bober, Angelika Brandt, Inmaculada Frutos, and Martin Schwentner. Diversity and distribution of Ischnomesidae (Crustacea: Isopoda: Asellota) along the Kuril– Kamchatka Trench — a genetic perspective. Progress in Oceanography, 178(??):??, November 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661119301727.

Brink:2009:LTM

[BBL^{+}09] K. H. Brink, R. C. Beardsley, R. Limeburner, J. D. Irish, and M. Caruso. Long-term moored array measurements of currents and hydrography over Georges Bank: 1994– 1999. Progress in Oceanography, 82(3):191–223, September 2009. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661109000585.

Blauw:2018:PED

[BBL^{+}18] Anouk N. Blauw, Elisa Benincà, Remi W. P. M. Laane, Naomi Greenwood, and Jef Huisman. Predictability and environmental drivers of chlorophyll fluctuations vary across different time scales and regions of the North Sea. Progress in Oceanography, 161(??):1–18, February 2018. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661117303324.

Ballon:2011:TEZ

[BBLD^{+}11] Michael Ballón, Arnaud Bertrand, Anne Lebourges-Dhaussy, Mariano Gutiérrez, Patricia Ayón, Daniel Grados, and François Gerlotto. Is there enough zooplankton to feed forage fish populations off Peru? An acoustic (positive) answer. Progress in Oceanography, 91(4):360–381, December 2011. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661111000279.

Berta:2014:ELT

[BBM^{+}14] Maristella Berta, Lucio Bellomo, Marcello G. Magaldi, Annalisa Griffa, Anne Molcard, Julien Marmain, Mireno Borghini, and Vincent Taillandier. Estimating Lagrangian transport blending drifters with HF radar data and models: Results from the TOSCA experiment in the Ligurian Current (North Western Mediterranean Sea). Progress in Oceanography, 128(??):15–29, November 2014. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S007966111400127X.

Benoist:2019:GVM

[BBMR19] Noëlie M. A. Benoist, Brian J. Bett, Kirsty J. Morris, and Henry A. Ruhl. A generalised volumetric method to estimate the biomass of photographically surveyed benthic megafauna. Progress in Oceanography, 178(??):??, November 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661119302204.

Benitez-Barrios:2011:TDC

[BBPHG^{+}11] V. M. Benítez-Barrios, J. L. Pelegrí, A. Hernández-Guerra, K. M. M. Lwiza, D. Gomis, P. Vélez-Belchí, and S. Hernández-León. Three-dimensional circulation in the NW Africa coastal transition zone. Progress in Oceanography, 91(4):516–533, December 2011. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661111000887.

Billett:2001:LTC

[BBR^{+}01] D. S. M. Billett, B. J. Bett, A. L. Rice, M. H. Thurston, J. Galéron, M. Sibuet, and G. A. Wolff. Long-term change in the megabenthos of the Porcupine Abyssal Plain (NE Atlantic). Progress in Oceanography, 50(??):325–348, ???? 2001. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S007966110100060X.

Brandt:2020:BBA

[BBRM20] Angelika Brandt, Saskia Brix, Torben Riehl, and Marina Malyutina. Biodiversity and biogeography of the abyssal and hadal Kuril–Kamchatka Trench and adjacent NW Pacific deepsea regions. Progress in Oceanography, 181(??):??, February 2020. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661119304124.

Basdurak:2021:LEV

[BBS21] N. Berkay Basdurak, Hans Burchard, and Henk M. Schuttelaars. A local eddy viscosity parameterization for winddriven estuarine exchange flow. Part I: Stratification dependence. Progress in Oceanography, 193(??):??, April–May 2021. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661121000367.

Basdurak:2023:LEV

[BBS23] N. Berkay Basdurak, Hans Burchard, and Henk M. Schuttelaars. A local eddy viscosity parameterization for winddriven estuarine exchange flow, Part II: Entrainment. Progress in Oceanography, 219(??):??, December 2023. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661123002094.

Beamish:2004:RHM

[BBSN04] R. J. Beamish, A. J. Benson, R. M. Sweeting, and C. M. Neville. Regimes and the history of the major fisheries off Canada’s west coast. Progress in Oceanography, 60(2–4): 355–385, February 2004. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661104000357.

Byun:1988:TCW

[BC88] Sang-Kyung Byun and Kyung-Il Chang. Tsushima current water at the entrance of the Korea Strait in autumn. Progress in Oceanography, 21(3–4):295–306, ???? 1988. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0079661188900092.

Bailey:1991:STV

[BC91] G. W. Bailey and P. Chapman. Short-term variability during an anchor station study in the southern Benguela upwelling system: Chemical and physical oceanography. Progress in Oceanography, 28(1–2):9–37, ???? 1991. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/007966119190019I.

Balopoulos:1999:PP

[BC99] Efstathios Balopoulos and Michael Collins. The PELAGOS Project. Progress in Oceanography, 44(4):v–ix, December 1999. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661100000021.

Buhring:2001:LSA

[BC01] Solveig I. Bühring and Bernd Christiansen. Lipids in selected abyssal benthopelagic animals: links to the epipelagic zone? Progress in Oceanography, 50(??):369–382, ???? 2001. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661101000611.

Boltovskoy:2016:BRP

[BC16] Demetrio Boltovskoy and Nancy Correa. Biogeography of Radiolaria Polycystina (Protista) in the World Ocean. Progress in Oceanography, 149(??):82–105, December 2016. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661116301501.

Borisanova:2019:NLS

[BC19] Anastasia O. Borisanova and Alexei V. Chernyshev. A new loxosomatid species from the Kuril–Kamchatka Trench area: Loxosomella marcusorum sp. n., the first record of hadal Entoprocta. Progress in Oceanography, 178(??):??, November 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S007966111930179X.

Bailey:2005:RWP

[BCB^{+}05] Kevin M. Bailey, Lorenzo Ciannelli, Nicholas A. Bond, Andrea Belgrano, and Nils Chr. Stenseth. Recruitment of walleye pollock in a physically and biologically complex ecosystem: a new perspective. Progress in Oceanography, 67(1–2):24–42, October 2005. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661105001102.

Burd:2020:SBM

[BCD^{+}20] Adrian B. Burd, Jeffrey P. Chanton, Kendra L. Daly, Sherryl Gilbert, Uta Passow, and Antonietta Quigg. The science behind marine-oil snow and MOSSFA: Past, present, and future. Progress in Oceanography, 187(??):??, August– September 2020. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661120301373.

Batten:2003:CST

[BCF^{+}03] S. D. Batten, R. Clark, J. Flinkman, G. Hays, E. John, A. W. G. John, T. Jonas, J. A. Lindley, D. P. Stevens, and A. Walne. CPR sampling: the technical background, materials and methods, consistency and comparability. Progress in Oceanography, 58(2–4):193–215, August 2003. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661103001290.

Burchard:2008:ONM

[BCG^{+}08] Hans Burchard, Peter D. Craig, Johannes R. Gemmrich, Hans van Haren, Pierre-Philippe Mathieu, H. E. Markus Meier, W. Alex M. Nimmo Smith, Hartmut Prandke, Tom P. Rippeth, Eric D. Skyllingstad, William D. Smyth, David J. S. Welsh, and Hemantha W. Wijesekera. Observational and numerical modeling methods for quantifying coastal ocean turbulence and mixing. Progress in Oceanography, 76(4):399–442, March 2008. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661108000050.

Bode:2018:TIS

[BCGN^{+}18] Antonio Bode, Pablo Carrera, Gonzalo González-Nuevo, Enrique Nogueira, Isabel Riveiro, and M. Begoña Santos. A trophic index for sardine (Sardina pilchardus) and its relationship to population abundance in the southern Bay of Biscay and adjacent waters of the NE Atlantic. Progress in Oceanography, 166(??):139–147, September 2018. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661117300897.

Bryden:1994:ETS

[BCK94] Harry L. Bryden, Julio Candela, and Thomas H. Kinder. Exchange through the Strait of Gibraltar. Progress in Oceanography, 33(3):201–248, ???? 1994. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0079661194900280.

Brochier:2009:SPF

[BCL^{+}09] T. Brochier, F. Colas, C. Lett, V. Echevin, L. A. Cubillos, J. Tam, M. Chlaida, C. Mullon, and P. Fréon. Small pelagic fish reproductive strategies in upwelling systems: a natal homing evolutionary model to study environmental constraints. Progress in Oceanography, 83(1–4):261–269, December 2009. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661109000901.

Behagle:2017:ADD

[BCLD^{+}17] Nolwenn Béhagle, Cédric Cotté, Anne Lebourges-Dhaussy, Gildas Roudaut, Guy Duhamel, Patrice Brehmer, Erwan Josse, and Yves Cherel. Acoustic distribution of discriminated micronektonic organisms from a bi-frequency processing: the case study of eastern Kerguelen oceanic waters. Progress in Oceanography, 156(??):276–289, August 2017. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661115300239.

Benson:2002:CCA

[BCM^{+}02] Scott R. Benson, Donald A. Croll, Baldo B. Marinovic, Francisco P. Chavez, and James T. Harvey. Changes in the cetacean assemblage of a coastal upwelling ecosystem during El Niño 1997–98 and La Niña 1999. Progress in Oceanography, 54(1–4):279–291, July 2002. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S007966110200054X.

Beca-Carretero:2019:SIA

[BCOL^{+}19] P. P. Beca-Carretero, J. Otero, P. E. Land, S. Groom, and X. A. Álvarez-Salgado. Seasonal and inter-annual variability of net primary production in the NW Iberian margin (1998–2016) in relation to wind stress and sea surface temperature. Progress in Oceanography, 178(??):??, November 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661118302830.

Bruchert:2009:HSM

[BCP09] Volker Brüchert, Bronwen Currie, and Kathleen R. Peard. Hydrogen sulphide and methane emissions on the central Namibian shelf. Progress in Oceanography, 83(1–4):169–179, December 2009. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S007966110900072X.

Bruno:2013:ISM

[BCR^{+}13] M. Bruno, J. Chioua, J. Romero, A. Vázquez, D. Macías, C. Dastis, E. Ramírez-Romero, F. Echevarria, J. Reyes, and C. M. García. The importance of sub-mesoscale processes for the exchange of properties through the Strait of Gibraltar. Progress in Oceanography, 116(??):66–79, September 2013. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661113000943.

Belkin:2009:FLM

[BCS09] Igor M. Belkin, Peter C. Cornillon, and Kenneth Sherman. Fronts in large marine ecosystems. Progress in Oceanography, 81(1–4):223–236, April 2009. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661109000330.

Barange:2009:HEC

[BCT^{+}09] Manuel Barange, Janet Coetzee, Akinori Takasuka, Kevin Hill, Mariano Gutierrez, Yoshioki Oozeki, Carl van der Lingen, and Vera Agostini. Habitat expansion and contraction in anchovy and sardine populations. Progress in Oceanography, 83(1–4):251–260, December 2009. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S007966110900086X.

Brewer:1985:COP

[BD85] Peter G. Brewer and D. Dyrssen. Chemical oceanography of the Persian Gulf. Progress in Oceanography, 14(??):41– 55, ???? 1985. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661185900047.

Bonsell:2018:LTP

[BD18] Christina Bonsell and Kenneth H. Dunton. Long-term patterns of benthic irradiance and kelp production in the central Beaufort Sea reveal implications of warming for Arctic inner shelves. Progress in Oceanography, 162(??):160–170, March 2018. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661117302288.

Bianchelli:2019:MBS

[BD19] Silvia Bianchelli and Roberto Danovaro. Meiofaunal biodiversity in submarine canyons of the Mediterranean Sea: a meta-analysis. Progress in Oceanography, 170(??):69–80, January 2019. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661118300806.

Besiktepe:2020:EDC

[BD20] Sengul Besiktepe and Hans G. Dam. Effect of diet on the coupling of ingestion and egg production in the ubiquitous copepod, Acartia tonsa. Progress in Oceanography, 186(??): ??, July 2020. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661120300859.

Balch:2004:MYR

[BDB^{+}04] W. M. Balch, D. T. Drapeau, B. C. Bowler, E. S. Booth, J. I. Goes, A. Ashe, and J. M. Frye. A multi-year record of hydrographic and bio-optical properties in the Gulf of Maine: I. Spatial and temporal variability. Progress in Oceanography, 63(1–2):57–98, October 2004. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661104001004.

Batchelder:2022:WPC

[BDB^{+}22] Hal Batchelder, Hans Dam, Eric Bjorkstedt, Skip McKinnell, Sam McClatchie, and So Kawaguchi. William T. Peterson commemorative issue: Progress in oceanography. Progress in Oceanography, 209(??):??, December 2022. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661122001689.

Boning:2001:SCM

[BDBJ01] Claus W. Böning, Christian Dieterich, Bernard Barnier, and Yanli Jia. Seasonal cycle of meridional heat transport in the subtropical North Atlantic: a model intercomparison in relation to observations near 25^{o}N. Progress in Oceanography, 48(??):231–253, ???? 2001. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661101000064.

Brodeur:2008:RFJ

[BDC^{+}08] Richard D. Brodeur, Mary Beth Decker, Lorenzo Ciannelli, Jennifer E. Purcell, Nicholas A. Bond, Phyllis J. Stabeno, Erika Acuna, and George L. Hunt. Rise and fall of jellyfish in the eastern Bering Sea in relation to climate regime shifts. Progress in Oceanography, 77(2–3):103–111, May 2008. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661108000487.

Brander:2003:UCP

[BDE03] K. M. Brander, R. R. Dickson, and M. Edwards. Use of continuous plankton recorder information in support of marine management: applications in fisheries, environmental protection, and in the study of ecosystem response to environmental change. Progress in Oceanography, 58(2–4):175–191, August 2003. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661103001289.

Blanchard:2017:EVO

[BDG^{+}17] Arny L. Blanchard, Robert H. Day, Adrian E. Gall, Lisanne A. M. Aerts, Julien Delarue, Elizabeth L. Dobbins, Russell R. Hopcroft, Jennifer M. Questel, Thomas J. Weingartner, and Sheyna S. Wisdom. Ecosystem variability in the offshore northeastern Chukchi Sea. Progress in Oceanography, 159(??):130–153, December 2017. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661117300952.

Bertrand:2008:PSD

[BDL08] Sophie Bertrand, Erich Díaz, and Matthieu Lengaigne. Patterns in the spatial distribution of Peruvian anchovy (Engraulis ringens) revealed by spatially explicit fishing data. Progress in Oceanography, 79(2–4):379–389, October 2008. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661108001894.

Boyce:2014:EGC

[BDLW14] Daniel G. Boyce, Michael Dowd, Marlon R. Lewis, and Boris Worm. Estimating global chlorophyll changes over the past century. Progress in Oceanography, 122(??):163–173, March 2014. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661114000135.

Baklouti:2006:IMF

[BDP^{+}06] M. Baklouti, F. Diaz, C. Pinazo, V. Faure, and B. Quéguiner. Investigation of mechanistic formulations depicting phytoplankton dynamics for models of marine pelagic ecosystems and description of a new model. Progress in Oceanography, 71(1):1–33, October 2006. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661106000760.

Bertrand:2008:IKW

[BDT^{+}08] Sophie Bertrand, Boris Dewitte, Jorge Tam, Erich Díaz, and Arnaud Bertrand. Impacts of Kelvin wave forcing in the Peru Humboldt Current system: Scenarios of spatial reorganizations from physics to fishers. Progress in Oceanography, 79(2–4):278–289, October 2008. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661108001808.

Barrier:2015:HBN

[BDTC15] Nicolas Barrier, Julie Deshayes, Anne-Marie Treguier, and Christophe Cassou. Heat budget in the North Atlantic subpolar gyre: Impacts of atmospheric weather regimes on the 1995 warming event. Progress in Oceanography, 130(??):75– 90, January 2015. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661114001645.

Banse:1999:CPS

[BE99] Karl Banse and David C. English. Comparing phytoplankton seasonality in the eastern and western subarctic Pacific and the western Bering Sea. Progress in Oceanography, 43(2–4):235– 288, March 1999. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661199000105.

Beaugrand:2004:NSR

[Bea04] Grégory Beaugrand. The North Sea regime shift: Evidence, causes, mechanisms and consequences. Progress in Oceanography, 60(2–4):245–262, February 2004. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661104000308.

Bisch:2022:MDV

[BECA22] Amaëlle Bisch, Sophie A. M. Elliott, Alexandre Carpentier, and Anthony Acou. Modelling the distribution of vulnerable skate from fisheries dependent data using imperfect detection. Progress in Oceanography, 206(??):??, August 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661122001185.

Benthuysen:2022:OIR

[BECR^{+}22] Jessica A. Benthuysen, Michael J. Emslie, Leanne M. Currey-Randall, Alistair J. Cheal, and Michelle R. Heupel. Oceanographic influences on reef fish assemblages along the Great Barrier Reef. Progress in Oceanography, 208(??):??, November 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661122001604.

Beaugrand:2019:EPA

[BEH19] Gregory Beaugrand, Martin Edwards, and Pierre Hélaouët. An ecological partition of the Atlantic Ocean and its adjacent seas. Progress in Oceanography, 173(??):86–102, April 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661118302027.

Baker:2020:AST

[BEI^{+}20] Chelsey A. Baker, Margaret L. Estapa, Morten Iversen, Richard Lampitt, and Ken Buesseler. Are all sediment traps created equal? An intercomparison study of carbon export methodologies at the PAP-SO site. Progress in Oceanography, 184(??):??, May 2020. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661120300562.

Belkin:2009:RWL

[Bel09] Igor M. Belkin. Rapid warming of large marine ecosystems. Progress in Oceanography, 81(1–4):207–213, April 2009. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661109000317.

Bennett:1985:ADI

[Ben85] A. F. Bennett. Array design by inverse methods. Progress in Oceanography, 15(2):129–156, ???? 1985. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0079661185900333.

Batchelder:2002:IBM

[BEP02] Harold P. Batchelder, Christopher A. Edwards, and Thomas M. Powell. Individual-based models of copepod populations in coastal upwelling regions: implications of physiologically and environmentally influenced diel vertical migration on demographic success and nearshore retention. Progress in Oceanography, 53(??):307–333, ???? 2002. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661102000356.

Bergeron:1965:PRS

[Ber65a] Tor Bergeron. The possible role of snowdrift in building up high inland ice-sheets. Progress in Oceanography, 3(??):385– 390, ???? 1965. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661165900339.

Berrit:1965:CSC

[Ber65b] G. R. Berrit. Les conditions de saison chaude dans la region orientale du golfe de Guinée. (French) [Warm season conditions in the eastern region of the Gulf of Guinea]. Progress in Oceanography, 3(??):31–47, ???? 1965. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0079661165900078.

Berthois:1965:ECS

[Ber65c] Léopold Berthois. Essai de correlation entre la sedimentation actuelle sur le bord externe des plateformes continentales et la dynamique fluviale. Progress in Oceanography, 3(??):49– 62, ???? 1965. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/007966116590008X.

Borges:2001:STV

[BF01]  Alberto V. Borges and Michel Frankignoulle.  Short-

term variations of the partial pressure of CO_{2} in surface waters of the Galician upwelling system. Progress in Oceanography, 51(2–4):283–302, November 2001. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661101000714.

Bouchard:2011:CAC

[BF11] Caroline Bouchard and Louis Fortier. Circum-arctic comparison of the hatching season of polar cod Boreogadus saida: a test of the freshwater winter refuge hypothesis. Progress in Oceanography, 90(1–4):105–116, July 2011. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661111000231.

Bower:2012:MEG

[BF12] Amy S. Bower and Heather H. Furey. Mesoscale eddies in the Gulf of Aden and their impact on the spreading of Red Sea Outflow Water. Progress in Oceanography, 96(1):14–39, April 2012. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661111001145.

Basilone:2020:SEE

[BFB^{+}20] Gualtiero Basilone, Rosalia Ferreri, Marco Barra, Angelo Bonanno, Maurizio Pulizzi, Antonella Gargano, Ignazio Fontana, Giovanni Giacalone, Paola Rumolo, Salvatore Mazzola, Simona Genovese, Richard McBride, and Salvatore Aronica. Spawning ecology of the European anchovy (Engraulis encrasicolus) in the Strait of Sicily: Linking variations of zooplankton prey, fish density, growth, and reproduction in an upwelling system. Progress in Oceanography, 184(??):??, May 2020. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661120300690.

Batten:2001:CMD

[BFH01] Sonia D. Batten, Elaine S. Fileman, and Elisabeth Halvorsen. The contribution of microzooplankton to the diet of mesozooplankton in an upwelling filament off the north west coast of Spain. Progress in Oceanography, 51(2–4):385–398, November 2001. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661101000763.

Blanchet:2018:PRP

[BFJ18] Marine Blanchet, Camila Fernandez, and Fabien Joux. Photoreactivity of riverine and phytoplanktonic dissolved organic matter and its effects on the dynamics of a bacterial community from the coastal Mediterranean Sea. Progress in Oceanography, 163(??):82–93, April 2018. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661116301999.

Bargain:2018:PHM

[BFP^{+}18] A. Bargain, F. Foglini, I. Pairaud, D. Bonaldo, S. Carniel, L. Angeletti, M. Taviani, S. Rochette, and M. C. Fabri. Predictive habitat modeling in two Mediterranean canyons including hydrodynamic variables. Progress in Oceanography, 169(??):151–168, December 2018. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S007966111730280X.

Baklouti:2006:ISA

[BFPS06] Melika Baklouti, Vincent Faure, Lionel Pawlowski, and Antoine Sciandra. Investigation and sensitivity analysis of a mechanistic phytoplankton model implemented in a new modular numerical tool (Eco3M) dedicated to biogeochemical modelling. Progress in Oceanography, 71(1):34–58, October 2006. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661106000772.

Barton:2013:CCU

[BFR13] E. D. Barton, D. B. Field, and C. Roy. Canary Current upwelling: More or less? Progress in Oceanography, 116 (??):167–178, September 2013. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661113001134.

Brosset:2017:STP

[BFV^{+}17] Pablo Brosset, Jean-Marc Fromentin, Elisabeth Van Beveren, Josep Lloret, Virginie Marques, Gualtiero Basilone, Angelo Bonanno, Piera Carpi, Fortunata Donato, Vanja Čikeš Keč, Andrea De Felice, Rosalia Ferreri, Denis Gašparević, Ana Giráldez, Ali Gücü, Magdalena Iglesias, Iole Leonori, Isabel Palomera, Stylianos Somarakis, Vjekoslav Tičina, Pedro Torres, Ana Ventero, Barbara Zorica, Frédéric Ménard, and Claire Saraux. Spatio-temporal patterns and environmental controls of small pelagic fish body condition from contrasted Mediterranean areas. Progress in Oceanography, 151(??):149–162, February 2017. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661116301276.

Borrell:2021:NPA

[BGA^{+}21] Asunción Borrell, Manel Gazo, Alex Aguilar, Juan A. Raga, Eduard Degollada, Patricia Gozalbes, and Raquel García-Vernet. Niche partitioning amongst northwestern Mediterranean cetaceans using stable isotopes. Progress in Oceanography, 193(??):??, April–May 2021. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S007966112100046X.

Bertrand:2008:SBE

[BGB^{+}08] Arnaud Bertrand, François Gerlotto, Sophie Bertrand, Mariano Gutiérrez, Luis Alza, Andres Chipollini, Erich Díaz, Pepe Espinoza, Jesús Ledesma, Roberto Quesquén, Salvador Peraltilla, and Francisco Chavez. Schooling behaviour and environmental forcing in relation to anchoveta distribution: an analysis across multiple spatial scales. Progress in Oceanography, 79(2–4):264–277, October 2008. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661108001791.

Bouchard:2017:CWE

[BGL^{+}17] Caroline Bouchard, Maxime Geoffroy, Mathieu LeBlanc, Andrew Majewski, Stéphane Gauthier, Wojciech Walkusz, James D. Reist, and Louis Fortier. Climate warming enhances polar cod recruitment, at least transiently. Progress in Oceanography, 156(??):121–129, August 2017. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661117301131.

Bethoux:1999:MSM

[BGM^{+}99] J. P. Bethoux, B. Gentili, P. Morin, E. Nicolas, C. Pierre, and D. Ruiz-Pino. The Mediterranean Sea: a miniature ocean for climatic and environmental studies and a key for the climatic functioning of the North Atlantic. Progress in Oceanography, 44(1–3):131–146, August 1999. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661199000233.

Barbosa:2001:STV

[BGM^{+}01] A. B. Barbosa, H. M. Galvão, P. A. Mendes, X. A. Álvarez-Salgado, F. G. Figueiras, and I. Joint. Short-term variability of heterotrophic bacterioplankton during upwelling off the NW Iberian margin. Progress in Oceanography, 51(2–4):339–359, November 2001. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S007966110100074X.

Bazzino:2010:HMV

[BGM^{+}10] Gastón Bazzino, William F. Gilly, Unai Markaida, César A. Salinas-Zavala, and Jorge Ramos-Castillejos. Horizontal movements, vertical-habitat utilization and diet of the jumbo squid (Dosidicus gigas) in the Pacific Ocean off Baja California Sur, Mexico. Progress in Oceanography, 86(1–2):59–71, July 2010. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661110000510.

Barton:2003:CPR

[BGMP03] A. D. Barton, C. H. Greene, B. C. Monger, and A. J. Pershing. The continuous plankton recorder survey and the North Atlantic oscillation: Interannual- to multidecadal-scale patterns of phytoplankton variability in the North Atlantic Ocean. Progress in Oceanography, 58(2–4):337–358, August 2003. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S007966110300137X.

Boss:2015:OTR

[BGR^{+}15] Emmanuel Boss, Lionel Guidi, Mary Jo Richardson, Lars Stemmann, Wilford Gardner, James K. B. Bishop, Robert F. Anderson, and Robert M. Sherrell. Optical techniques for remote and in-situ characterization of particles pertinent to GEOTRACES. Progress in Oceanography, 133(??):43–54, April 2015. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661114001542.

Brunnegaard:2004:NCD

[BGS^{+}04] Jenny Brunnegård, Sibylle Grandel, Henrik Ståhl, Anders Tengberg, and Per O. J. Hall. Nitrogen cycling in deepsea sediments of the Porcupine Abyssal Plain, NE Atlantic. Progress in Oceanography, 63(4):159–181, December 2004. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661104001235.

Brakstad:2023:FPD

[BGV^{+}23] Ailin Brakstad, Geoffrey Gebbie, Kjetil Våge, Emil Jeansson, and Sólveig Rósa Ólafsdóttir. Formation and pathways of dense water in the Nordic Seas based on a regional inversion. Progress in Oceanography, 212(??):??, March 2023. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661123000241.

Burridge:2017:DAP

[BGWP^{+}17] Alice K. Burridge, Erica Goetze, Deborah Wall-Palmer, Serena L. Le Double, Jef Huisman, and Katja T. C. A. Peijnenburg. Diversity and abundance of pteropods and heteropods along a latitudinal gradient across the Atlantic Ocean. Progress in Oceanography, 158(??):213–223, November 2017. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661116300040.

Bryden:1985:EEN

[BH85] Harry L. Bryden and Ron A. Heath. Energetic eddies at the Northern edge of the Antarctic Circumpolar Current in the Southwest Pacific. Progress in Oceanography, 14(??):65– 87, ???? 1985. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661185900060.

Beckmann:2007:BSC

[BH07] Aike Beckmann and Inga Hense. Beneath the surface: Characteristics of oceanic ecosystems under weak mixing conditions — a theoretical investigation. Progress in Oceanography, 75(4):771–796, December 2007. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661107001826.

Ballerini:2014:PLF

[BHA^{+}14] Tosca Ballerini, Eileen E. Hofmann, David G. Ainley, Kendra Daly, Marina Marrari, Christine A. Ribic, Walker O. Smith, and John H. Steele. Productivity and linkages of the food web of the southern region of the western Antarctic Peninsula continental shelf. Progress in Oceanography, 122(??):10– 29, March 2014. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661113002267.

Blamey:2012:RSS

[BHAJ12] Laura K. Blamey, James A. E. Howard, Jacobus Agenbag, and Astrid Jarre. Regime-shifts in the southern Benguela shelf and inshore region. Progress in Oceanography, 106(??):80–95, November 2012. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S007966111200078X.

Bergmeier:2019:BPT

[BHB^{+}19] Franziska S. Bergmeier, Gerhard Haszprunar, Angelika Brandt, Hiroshi Saito, Yasunori Kano, and Katharina M. Jörger. Of basins, plains, and trenches: Systematics and distribution of Solenogastres (Mollusca, Aplacophora) in the Northwest Pacific. Progress in Oceanography, 178(??):??, November 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661119302265.

Bode:2018:CDB

[BHC^{+}18] Maya Bode, Wilhelm Hagen, Astrid Cornils, Patricia Kaiser, and Holger Auel. Copepod distribution and biodiversity patterns from the surface to the deep sea along a latitudinal transect in the eastern Atlantic Ocean (24^{o}N to 21^{o}S). Progress in Oceanography, 161(??):66–77, February 2018. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S007966111730188X.

Balzer:1998:BDN

[BHE^{+}98] W. Balzer, W. Helder, E. Epping, L. Lohse, and S. Otto. Benthic denitrification and nitrogen cycling at the slope and rise of the N.W. European Continental Margin (Goban Spur). Progress in Oceanography, 42(1–4):111–126, December 1998. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661198000305.

Bertrand:2016:DHS

[BHH^{+}16] Arnaud Bertrand, Jérémie Habasque, Tarek Hattab, Niels T. Hintzen, Ricardo Oliveros-Ramos, Mariano Gutiérrez, Hervé Demarcq, and François Gerlotto. 3-D habitat suitability of jack mackerel Trachurus murphyi in the Southeastern Pacific, a comprehensive study. Progress in Oceanography, 146(??):199– 211, August 2016. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661116300404.

Bendtsen:2015:ARL

[BHHR15] Jørgen Bendtsen, Karen Marie Hilligsøe, Jørgen L. S. Hansen, and Katherine Richardson. Analysis of remineralisation, lability, temperature sensitivity and structural composition of organic matter from the upper ocean. Progress in Oceanography, 130(??):125–145, January 2015. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661114001724.

Brink:1983:PEP

[BHHS83] K. H. Brink, D. Halpern, A. Huyer, and R. L. Smith. The physical environment of the Peruvian upwelling system. Progress in Oceanography, 12(3):285–305, ???? 1983. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0079661183900113.

Bednarsek:2016:PEC

[BHK^{+}16] Nina Bednaršek, Chris J. Harvey, Isaac C. Kaplan, Richard A. Feely, and Jasna Možina. Pteropods on the edge: Cumulative effects of ocean acidification, warming, and deoxygenation. Progress in Oceanography, 145(??):1–24, June 2016. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661115300112.

Brandao:2019:ROC

[BHK^{+}19] Simone N. Brandão, Mario Hoppema, Gennady M. Kamenev, Ivana Karanovic, Torben Riehl, Hayato Tanaka, Helenice Vital, Hyunsu Yoo, and Angelika Brandt. Review of Ostracoda (Crustacea) living below the Carbonate Compensation Depth and the deepest record of a calcified ostracod. Progress in Oceanography, 178(??):??, November 2019. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661119300655.

Bonnet:2007:CSD

[BHLU^{+}07] Delphine Bonnet, Roger Harris, Angel Lopez-Urrutia, Claudia Halsband-Lenk, Wulf Greve, Luis Valdes, Hans-Jürgen Hirche, Marcus Engel, Maria Teresa Alvarez-Ossorio, and Karen Wiltshire. Comparative seasonal dynamics of Centropages typicus at seven coastal monitoring stations in the North Sea, English Channel and Bay of Biscay. Progress in Oceanography, 72(2–3):233–248, February 2007. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661107000043.

Barber:2015:SPB

[BHM^{+}15] David G. Barber, Haakon Hop, Christopher J. Mundy, Brent Else, Igor A. Dmitrenko, Jean-Eric Tremblay, Jens K. Ehn, Philipp Assmy, Malin Daase, Lauren M. Candlish, and Søren Rysgaard. Selected physical, biological and biogeochemical implications of a rapidly changing Arctic marginal ice zone. Progress in Oceanography, 139(??):122–150, December 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661115300148.

Bundy:2009:SCF

[BHMS09] Alida Bundy, Johanna J. Heymans, Lyne Morissette, and Claude Savenkoff. Seals, cod and forage fish: a comparative exploration of variations in the theme of stock collapse and ecosystem change in four Northwest Atlantic ecosystems. Progress in Oceanography, 81(1–4):188–206, April 2009. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661109000305.

Bohm:2006:CSS

[BHPC06] E. Böhm, T. S. Hopkins, L. J. Pietrafesa, and J. H. Churchill. Continental slope sea level and flow variability induced by lateral movements of the Gulf Stream in the middle Atlantic Bight. Progress in Oceanography, 70(2–4):196–212, August 2006. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661106000875.

Bode:2015:FST

[BHS^{+}15] Maya Bode, Wilhelm Hagen, Anna Schukat, Lena Teuber, Debany Fonseca-Batista, Frank Dehairs, and Holger Auel. Feeding strategies of tropical and subtropical calanoid copepods throughout the eastern Atlantic Ocean — latitudinal and bathymetric aspects. Progress in Oceanography, 138 (part A)(??):268–282, November 2015. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661115002116.

Breaker:2010:NOM

[BhTW10] Laurence C. Breaker, Yu heng Tseng, and Xiaochun Wang. On the natural oscillations of Monterey Bay: Observations, modeling, and origins. Progress in Oceanography, 86(3–4):380–395, September 2010. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661110000935.

Billett:2001:TRD

[Bil01] D. S. M. Billett. Tony Rice — deep-sea biologist. Progress in Oceanography, 50(??):1–11, ???? 2001. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661101000453.

Beaugrand:2003:OSM

[BIL03] G. Beaugrand, F. Ibañez, and J. A. Lindley. An overview of statistical methods applied to CPR data. Progress in Oceanography, 58(2–4):235–262, August 2003. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661103001319.

Barton:2001:VST

[BIST01] E. D. Barton, M. E. Inall, T. J. Sherwin, and R. Torres. Vertical structure, turbulent mixing and fluxes during Lagrangian observations of an upwelling filament system off Northwest Iberia. Progress in Oceanography, 51(2–4):249–267, November 2001. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661101000696.

Bottom:1990:SCD

[BJ90] Daniel L. Bottom and Kim K. Jones. Species composition, distribution, and invertebrate prey of fish assemblages in the Columbia River Estuary. Progress in Oceanography, 25(1–4):243–270, ???? 1990. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/007966119090009Q.

Baker:2017:ADS

[BJ17] Alex R. Baker and Tim D. Jickells. Atmospheric deposition of soluble trace elements along the Atlantic Meridional Transect (AMT). Progress in Oceanography, 158(??):41–51, November 2017. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661116300362.

Blazewicz:2019:HSR

[BJMP19] Magdalena Błażewicz, Piotr Jóźwiak, Lenaick Menot, and Krzysztof Pabis. High species richness and unique composition of the tanaidacean communities associated with five areas in the Pacific polymetallic nodule fields. Progress in Oceanography, 176(??):??, September 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S0079661119301399. See corrigendum [BJMP20].

Blazewicz:2020:CHS

[BJMP20] Magdalena Błażewicz, Piotr Jóźwiak, Lenaick Menot, and Krzysztof Pabis. Corrigendum to “High species richness and unique composition of the tanaidacean communities associated with five areas in the Pacific polymetallic nodule fields” [Progr. Oceanogr. 176 (2019) 102141]. Progress in Oceanography, 181(??):??, February 2020. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S007966111930401X. See [BJMP19].

Batchelder:2008:LLP

[BK08] Harold P. Batchelder and Suam Kim. Lessons learned from the PICES/GLOBEC Climate Change and Carrying Capacity (CCCC) Program and Synthesis Symposium. Progress in Oceanography, 77(2–3):83–91, May 2008. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661108000463.

Bercel:2019:ICA

[BK19] T. L. Bercel and S. A. Kranz. Insights into carbon acquisition and photosynthesis in Karenia brevis under a range of CO_{2} concentrations. Progress in Oceanography, 172(??):65– 76, March 2019. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661118302593.

Bruce:1985:NBE

[BKB85] J. G. Bruce, J. L. Kerling, and W. H. Beatty. On the North Brazilian eddy field. Progress in Oceanography, 14(??):57– 63, ???? 1985. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661185900059.

Bluhm:2015:TTB

[BKC15] B. A. Bluhm, K. N. Kosobokova, and E. C. Carmack. A tale of two basins: an integrated physical and biological perspective of the deep Arctic Ocean. Progress in Oceanography, 139(??):89–121, December 2015. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661115001639.

Binetti:2020:NCO

[BKD^{+}20] U. Binetti, J. Kaiser, G. M. Damerell, A. Rumyantseva, A. P. Martin, S. Henson, and K. J. Heywood. Net community oxygen production derived from Seaglider deployments at the Porcupine Abyssal Plain site (PAP; northeast Atlantic) in 2012–13. Progress in Oceanography, 183(??):??, April 2020. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661120300318.

Botsford:2002:PCV

[BL02] L. W. Botsford and C. A. Lawrence. Patterns of covariability among California Current Chinook salmon, Coho salmon, Dungeness crab, and physical oceanographic conditions. Progress in Oceanography, 53(??):283–305, ???? 2002. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661102000344.

Blanchard:1963:EAP

[Bla63] Duncan C. Blanchard. The electrification of the atmosphere by particles from bubbles in the sea. Progress in Oceanography, 1(??):73–112, IN7, 113–202, ???? 1963. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0079661163900041.

Belkin:1998:GSA

[BLAM98] Igor M. Belkin, Sydney Levitus, John Antonov, and Svend-Aage Malmberg. “Great Salinity Anomalies” in the North Atlantic. Progress in Oceanography, 41(1):1–68, January 1998. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661198000159.

Belkin:2000:CGS

[BLAM00] I. M. Belkin, S. Levitus, J. Antonov, and S.-A. Malmberg. Corrigendum to “Great Salinity Anomalies” in the North Atlantic’: [Progress in Oceanography, 41 (1998) 1– 68]. Progress in Oceanography, 45(1):107–108, January 2000. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661100000033.

Bukanova:2023:BIW

[BLC23] Tatiana Bukanova, Olga Lobchuk, and Irina Chubarenko. The Bornholm intermediate waters: Origination, pathway, and detection within the cold intermediate layer of the Baltic Sea. Progress in Oceanography, 210(??):??, January 2023. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661122001859.

Blackett:2014:LTV

[BLCL14] Michael Blackett, Priscilla Licandro, Steve H. Coombs, and Cathy H. Lucas. Long-term variability of the siphonophores Muggiaea atlantica and M. kochi in the Western English Channel. Progress in Oceanography, 128(??):1–14, November 2014. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661114001232.

Brandt:2016:DRG

[BLES16] Angelika Brandt, Katrin Linse, Kari E. Ellingsen, and Paul J. Somerfield. Depth-related gradients in community structure and relatedness of bivalves and isopods in the Southern Ocean. Progress in Oceanography, 144(??):25–38, May 2016. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661116300507.

Beaugrand:2007:MSC

[BLHB07] G. Beaugrand, J. A. Lindley, P. Helaouet, and D. Bonnet. Macroecological study of Centropages typicus in the North Atlantic Ocean. Progress in Oceanography, 72(2–3):259–273, February 2007. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661107000031.

Beamish:1999:EDA

[BLI^{+}99] R. J. Beamish, K. D. Leask, O. A. Ivanov, A. A. Balanov, A. M. Orlov, and B. Sinclair. The ecology, distribution, and abundance of midwater fishes of the subarctic Pacific gyres. Progress in Oceanography, 43(2–4):399–442, March 1999. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661199000178.

Broullon:2020:TLP

[BLMR^{+}20] Esperanza Broullón, Marta López-Mozos, Beatriz Reguera, Paloma Chouciño, María Dolores Doval, Bieito Fernández-Castro, Miguel Gilcoto, Enrique Nogueira, Carlos Souto, and Beatriz Mouriño-Carballido. Thin layers of phytoplankton and harmful algae events in a coastal upwelling system. Progress in Oceanography, 189(??):??, November–December 2020. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661120301877.

Booth:1993:TVS

[BLP93] Beatrice C. Booth, Joyce Lewin, and James R. Postel. Temporal variation in the structure of autotrophic and heterotrophic communities in the subarctic Pacific. Progress in Oceanography, 32(??):57–99, ???? 1993. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0079661193900093.

Benkort:2020:ANK

[BLP^{+}20] Déborah Benkort, Diane Lavoie, Stéphane Plourde, Christiane Dufresne, and Frédéric Maps. Arctic and Nordic krill circuits of production revealed by the interactions between their physiology, swimming behaviour and circulation. Progress in Oceanography, 182(??):??, March 2020. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661120300082.

Barrier:2023:MUE

[BLR^{+}23a] Nicolas Barrier, Matthieu Lengaigne, Jonathan Rault, Renaud Person, Christian Ethé, Olivier Aumont, and Olivier Maury. Mechanisms underlying the epipelagic ecosystem response to ENSO in the equatorial Pacific Ocean. Progress in Oceanography, 213(??):??, April–May 2023. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661123000459.

Bassett:2023:ABT

[BLR^{+}23b] Christopher Bassett, Andone C. Lavery, David Ralston, W. Rockwell Geyer, Joseph T. Jurisa, Jim Thomson, David A. Honegger, Alexandra Simpson, Malcom E. Scully, and Merrick C. Haller. Acoustic backscattering at a tidal intrusion front. Progress in Oceanography, 219(??):??, December 2023. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661123002100.

Brochier:2008:IBM

[BLT^{+}08] Timothée Brochier, Christophe Lett, Jorge Tam, Pierre Fréon, François Colas, and Patricia Ayón. An individual-based model study of anchovy early life history in the northern Humboldt Current system. Progress in Oceanography, 79(2–4):313–325, October 2008. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661108001845.

Barton:2015:CUD

[BLT^{+}15] E. D. Barton, J. L. Largier, R. Torres, M. Sheridan, A. Trasviña, A. Souza, Y. Pazos, and A. Valle-Levinson. Coastal upwelling and downwelling forcing of circulation in a semi-enclosed bay: Ria de Vigo. Progress in Oceanography, 134(??):173–189, May 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661115000257.

Blunden:1988:MMA

[Blu88] Gerald Blunden. Metabolites from marine algae. Progress in Oceanography, 21(2):217–226, ???? 1988. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0079661188900419.

Badcock:1976:MFE

[BM76] Julian Badcock and N. R. Merrett. Midwater fishes in the eastern North Atlantic — I. Vertical distribution and associated biology in 30^{o}N, 23^{o}W, with developmental notes on certain myctophids. Progress in Oceanography, 7(1):3–58, ???? 1976. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661176900033.

Breaker:1986:OVC

[BM86] Laurence C. Breaker and Christopher N. K. Mooers. Oceanic variability off the Central California coast. Progress in Oceanography, 17(1–2):61–135, ???? 1986. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/007966118690025X.

Beamish:2001:CSP

[BM01] R. J. Beamish and Conrad Mahnken. A critical size and period hypothesis to explain natural regulation of salmon abundance and the linkage to climate and climate change. Progress in Oceanography, 49(??):423–437, ???? 2001. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661101000349.

Bottjer:2007:NAU

[BM07] Daniela Böttjer and Carmen E. Morales. Nanoplanktonic assemblages in the upwelling area off Concepción (≈ 36^{o}S), central Chile: Abundance, biomass, and grazing potential during the annual cycle. Progress in Oceanography, 75(3):415–434, November 2007. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661107001607.

Beamish:2006:LO

[BMB06] R. J. Beamish, G. A. McFarlane, and A. Benson. Longevity overfishing. Progress in Oceanography, 68(2–4):289–302, February 2006. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661106000103.

Bengil:2016:BOM

[BMB^{+}16] Fethi Bengil, David McKee, Sükrü T. Beşiktepe, Violeta Sanjuan Calzado, and Charles Trees. A bio-optical model for integration into ecosystem models for the Ligurian Sea. Progress in Oceanography, 149(??):1–15, December 2016. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661116300982.

Beranger:2005:SVW

[BMC05] K. Béranger, L. Mortier, and M. Crépon. Seasonal variability of water transport through the Straits of Gibraltar, Sicily and Corsica, derived from a high-resolution model of the Mediterranean circulation. Progress in Oceanography, 66(2–4):341– 364, August 2005. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661105000777.

Boustany:2010:MPB

[BMC^{+}10] Andre M. Boustany, Robyn Matteson, Michael Castleton, Charles Farwell, and Barbara A. Block. Movements of Pacific bluefin tuna (Thunnus orientalis) in the Eastern North Pacific revealed with archival tags. Progress in Oceanography, 86(1–2): 94–104, July 2010. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661110000492.

Beldowska:2017:MCV

[BMC17] Magdalena Bełdowska and Stella Mudrak-Cegiołka. Mercury concentration variability in the zooplankton of the southern Baltic coastal zone. Progress in Oceanography, 159(??):73–85, December 2017. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661117300058.

Becker:2021:LGC

[BMdMS^{+}21] Érica Caroline Becker, Maria Grazia Mazzocchi, Luis Carlos Pinto de Macedo-Soares, Manoela Costa Brandão, and Andrea Santarosa Freire. Latitudinal gradient of copepod functional diversity in the South Atlantic Ocean. Progress in Oceanography, 199(??):??, December 2021. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661121001932.

Beaugrand:2013:ACE

[BMG13] Gregory Beaugrand, Dave Mackas, and Eric Goberville. Applying the concept of the ecological niche and a macroecological approach to understand how climate influences zooplankton: Advantages, assumptions, limitations and requirements. Progress in Oceanography, 111(??):75–90, April 2013.

CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661112001681.

Brennan:2019:HTS

[BMG^{+}19] Catherine E. Brennan, Frédéric Maps, Wendy C. Gentleman, Stéphane Plourde, Diane Lavoie, Joël Chassé, Caroline Lehoux, Kira A. Krumhansl, and Catherine L. Johnson. How transport shapes copepod distributions in relation to whale feeding habitat: Demonstration of a new modelling framework. Progress in Oceanography, 171(??):1–21, February 2019. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661118302507.

Brennan:2021:OCC

[BMG^{+}21a] Catherine E. Brennan, Frédéric Maps, Wendy C. Gentleman, Diane Lavoie, Joël Chassé, Stéphane Plourde, and Catherine L. Johnson. Ocean circulation changes drive shifts in Calanus abundance in North Atlantic right whale foraging habitat: a model comparison of cool and warm year scenarios. Progress in Oceanography, 197(??):??, September– October 2021. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661121001154.

Bridier:2021:SQT

[BMG^{+}21b] Guillaume Bridier, Tarik Meziane, Jacques Grall, Laurent Chauvaud, Sébastien Donnet, Pascal Lazure, and Frédéric Olivier. Sources, quality and transfers of organic matter in a highly-stratified sub-Arctic coastal system (Saint-Pierre-et-Miquelon, NW Atlantic). Progress in Oceanography, 190(??): ??, January 2021. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661120302184.

Bayle:2015:MTF

[BMGN15] Séverine Bayle, Pascal Monestiez, Christophe Guinet, and David Nerini. Moving toward finer scales in oceanography: Predictive linear functional model of chlorophyll-a profile from light data. Progress in Oceanography, 134(??):221–231, May 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S007966111500035X.

Bagoien:2012:SDM

[BMK12] Espen Bagøien, Webjørn Melle, and Stein Kaartvedt. Seasonal development of mixed layer depths, nutrients, chlorophyll and Calanus finmarchicus in the Norwegian Sea — a basin-scale habitat comparison. Progress in Oceanography, 103(??):58–79, September 2012. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S007966111200047X.

Bode:1997:RPS

[BMM97] Lance Bode, Luciano B. Mason, and Jason H. Middleton. Reef parameterisation schemes with applications to tidal modelling. Progress in Oceanography, 40(??):285–324, ???? 1997. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661198000068.

Bertram:2001:SCR

[BMM01] Douglas F. Bertram, David L. Mackas, and Stewart M. McKinnell. The seasonal cycle revisited: interannual variation and ecosystem consequences. Progress in Oceanography, 49(??):283–307, ???? 2001. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661101000271.

Brodeur:1999:ENN

[BMN^{+}99] Richard Brodeur, Skip McKinnell, Kazuya Nagasawa, William Pearcy, Vladimir Radchenko, and Shogo Takagi. Epipelagic nekton of the North Pacific subarctic and transition zones. Progress in Oceanography, 43(2–4):365–397, March 1999. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661199000130.

Bozzeda:2019:UBD

[BMN19] Fabio Bozzeda, Sandra L. Marín, and Laura Nahuelhual. An uncertainty-based decision support tool to evaluate the southern king crab (Lithodes santolla) fishery in a scarce information context. Progress in Oceanography, 174(??):64–71, May 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661118300715.

Bett:2001:TVP

[BMNW01] Brian J. Bett, M. Gabriella Malzone, Bhavani E. Narayanaswamy, and Benjamin D. Wigham. Temporal variability in phytodetritus and megabenthic activity at the seabed in the deep Northeast Atlantic. Progress in Oceanography, 50(??):349– 368, ???? 2001. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661101000660.

Batchelder:2012:STS

[BMO12] Harold P. Batchelder, David L. Mackas, and Todd D. O’Brien. Spatial–temporal scales of synchrony in marine zooplankton biomass and abundance patterns: a world-wide comparison. Progress in Oceanography, 97–100(??):15–30, May 2012. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661111001285.

Barange:2003:EES

[BN03] Manuel Barange and Leo Nykjaer. ENVIFISH: an EU/ Southern African collaborative project investigating environmental causes of pelagic fisheries variability. Progress in Oceanography, 59(2–3):177–179, October 2003. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661103001666.

Bange:2005:NCA

[BNC05] Hermann W. Bange, S. Wajih A. Naqvi, and L. A. Codispoti. The nitrogen cycle in the Arabian Sea. Progress in Oceanography, 65(2–4):145–158, May 2005. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661105000406.

Bashmachnikov:2015:PPM

[BNCC15] I. Bashmachnikov, F. Neves, T. Calheiros, and X. Carton. Properties and pathways of Mediterranean water eddies in the Atlantic. Progress in Oceanography, 137 (part A)(??):149–172, September 2015. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661115001342.

Bernardes:2020:SAM

[BOG20] Inês Dias Bernardes, Egil Ona, and Harald Gjøsæter. Study of the Arctic mesopelagic layer with vessel and profiling multifrequency acoustics. Progress in Oceanography, 182(??):??, March 2020. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661119304409.

Boltovskoy:1994:SRP

[Bol94] Demetrio Boltovskoy. The sedimentary record of pelagic biogeography. Progress in Oceanography, 34(2–3):135–160, ???? 1994. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/007966119490006X.

Bernal:2015:DFS

[BOMdP15] Ainhoa Bernal, M. Pilar Olivar, Francesc Maynou, and M. Luz Fernández de Puelles. Diet and feeding strategies of mesopelagic fishes in the western Mediterranean. Progress in Oceanography, 135(??):1–17, June 2015. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661115000464.

Bonotto:1988:FCM

[Bon88] S. Bonotto. Food and chemicals from microalgae. Progress in Oceanography, 21(2):207–215, ???? 1988. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0079661188900407.

Bourcart:1965:CSM

[Bou65] Jacques Bourcart. Les canyons sous-marins de l’extremite orientale des pyrenees. Progress in Oceanography, 3(??):63–69, ???? 1965. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661165900091.

Batchelder:2002:PBC

[BP02] Harold P. Batchelder and Thomas M. Powell. Physical and biological conditions and processes in the northeast Pacific Ocean. Progress in Oceanography, 53(??):105–114, ???? 2002. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661102000265.

Baklouti:2021:BUC

[BPA^{+}21] Melika Baklouti, Rémi Pagès, Elena Alekseenko, Arnaud Guyennon, and Gérald Grégori. On the benefits of using cell quotas in addition to intracellular elemental ratios in flexible-stoichiometry plankton functional type models. Application to the Mediterranean Sea. Progress in Oceanography, 197(??):??, September–October 2021. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661121001191.

Book:2005:AOW

[BPC^{+}05] J. W. Book, H. T. Perkins, L. Cavaleri, J. D. Doyle, and J. D. Pullen. ADCP observations of the western Adriatic slope current during winter of 2001. Progress in Oceanography, 66(2–4):270–286, August 2005. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661105000741.

Ballance:2006:OIS

[BPF06] Lisa T. Ballance, Robert L. Pitman, and Paul C. Fiedler. Oceanographic influences on seabirds and cetaceans of the eastern tropical Pacific: a review. Progress in Oceanography, 69(2–4):360–390, May 2006. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661106000371.

Bolado-Penagos:2020:SPC

[BPGC^{+}20] Marina Bolado-Penagos, Carlos J. González, Jamal Chioua, Iria Sala, Juan Jesús Gomiz-Pascual, Águeda Vázquez, and Miguel Bruno. Submesoscale processes in the coastal margins of the Strait of Gibraltar. The Trafalgar–Alboran connection. Progress in Oceanography, 181(??):??, February 2020. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661119303994.

Blondeau-Patissier:2014:ROC

[BPGD^{+}14] David Blondeau-Patissier, James F. R. Gower, Arnold G. Dekker, Stuart R. Phinn, and Vittorio E. Brando. A review of ocean color remote sensing methods and statistical techniques for the detection, mapping and analysis of phytoplankton blooms in coastal and open oceans. Progress in Oceanography, 123(??):123–144, April 2014. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661114000020.

Brugler:2014:SIV

[BPM^{+}14] Eric T. Brugler, Robert S. Pickart, G. W. K. Moore, Steven Roberts, Thomas J. Weingartner, and Hank Statscewich. Seasonal to interannual variability of the Pacific water boundary current in the Beaufort Sea. Progress in Oceanography, 127(??):1–20, September 2014. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661114000871.

Bathmann:1990:POF

[BPNB90] Ulrich V. Bathmann, Rolf Peinert, Thomas T. Noji, and Bodo V. Bodungen. Pelagic origin and fate of sedimenting particles in the Norwegian Sea. Progress in Oceanography, 24(??):117–125, ???? 1990. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/007966119090024V.

Boski:1998:FGA

[BPP^{+}98] T. Boski, J. Pessoa, P. Pedro, J. Thorez, J. M. A. Dias, and I. R. Hall. Factors governing abundance of hydrolyzable amino acids in the sediments from the N.W. European Continental Margin (47–50^{o}n). Progress in Oceanography, 42(1–4):145–164, December 1998. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661198000329.

Bolado-Penagos:2023:AIS

[BPSGP^{+}23] Marina Bolado-Penagos, Iria Sala, Juan Jesús Gomiz-Pascual, Carlos J. González, Alfredo Izquierdo, Óscar Álvarez, Águeda Vázquez, Miguel Bruno, and Hans van Haren. Analysis of internal soliton signals and their eastward propagation in the Alboran Sea: exploring the effect of subinertial forcing and fortnightly variability. Progress in Oceanography, 217(??):??, September–October 2023. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661123001209.

Buchan:2021:IVS

[BPSN^{+}21] Susannah J. Buchan, Iván Pérez-Santos, Diego Narváez, Leonardo Castro, Kathleen M. Stafford, Mark F. Baumgartner, Arnoldo Valle-Levinson, Paulina Montero, Laura Gutiérrez, Constanza Rojas, Giovanni Daneri, and Sergio Neira. Intraseasonal variation in southeast Pacific blue whale acoustic presence, zooplankton backscatter, and oceanographic variables on a feeding ground in Northern Chilean Patagonia. Progress in Oceanography, 199(??):??, December 2021. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661121001920.

Bringue:2019:DCP

[BPTT19] Manuel Bringué, Vera Pospelova, Eric J. Tappa, and Robert C. Thunell. Dinoflagellate cyst production in the Cariaco Basin: a 12.5 year-long sediment trap study. Progress in Oceanography, 171(??):175–211, February 2019. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661118302052.

Barange:2010:IGC

[BPW10] Manuel Barange, R. Ian Perry, and F. E. Werner. Introduction to the GLOBEC compendium: a selection of papers from the GLOBEC international research programme. Progress in Oceanography, 87(S1):S2–S5, ???? 2010. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661110001497.

Billett:2001:BPI

[BR01] D. S. M. Billett and A. L. Rice. The BENGAL programme: introduction and overview. Progress in Oceanography, 50(??): 13–25, ???? 2001. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661101000465.

Barnier:2001:SVE

[BRB^{+}01] Bernard Barnier, Thierry Reynaud, Aike Beckmann, Claus Böning, Jean-Marc Molines, Sally Barnard, and Yanli Jia. On the seasonal variability and eddies in the North Brazil Current: insights from model intercomparison experiments. Progress in Oceanography, 48(??):195–230, ???? 2001. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661101000052.

Barone:2018:BDS

[BRC^{+}18] Giulio Barone, Eugenio Rastelli, Cinzia Corinaldesi, Michael Tangherlini, Roberto Danovaro, and Antonio Dell’Anno. Benthic deep-sea fungi in submarine canyons of the Mediterranean Sea. Progress in Oceanography, 168(??):57–64, November 2018. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661118301587.

Berge:2015:DRE

[BRD^{+}15] Jørgen Berge, Paul E. Renaud, Gerald Darnis, Finlo Cottier, Kim Last, Tove M. Gabrielsen, Geir Johnsen, Lena Seuthe, Jan Marcin Weslawski, Eva Leu, Mark Moline, Jasmine Nahrgang, Janne E. Søreide, Øystein Varpe, Ole Jørgen Lønne, Malin Daase, and Stig Falk-Petersen. In the dark: a review of ecosystem processes during the Arctic polar night. Progress in Oceanography, 139(??):258–271, December 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661115001858.

Bretherton:1982:OCM

[Bre82] Francis P. Bretherton. Ocean climate modeling. Progress in Oceanography, 11(2):93–129, ???? 1982. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0079661182900052.

Breaker:2006:NAS

[Bre06] Laurence C. Breaker. Nonlinear aspects of sea surface temperature in Monterey Bay. Progress in Oceanography, 69(1): 61–89, April 2006. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661106000140.

Basedow:2024:IAG

[BRE^{+}24] S. L. Basedow, A. H. H. Renner, B. Espinasse, S. Falk-Petersen, M. Graeve, K. Bandara, K. Sørensen, K. Eiane, and W. Hagen. The impact of advection on a Subarctic fjord food web dominated by the copepod Calanus finmarchicus. Progress in Oceanography, 224(??):??, June 2024. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661124000740.

Bach:2015:UCC

[BRG^{+}15] Lennart Thomas Bach, Ulf Riebesell, Magdalena A. Gutowska, Luisa Federwisch, and Kai Georg Schulz. A unifying concept of coccolithophore sensitivity to changing carbonate chemistry embedded in an ecological framework. Progress in Oceanography, 135(??):125–138, June 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661115000725.

Bodur:2023:SPV

[BRG^{+}23] Yasemin V. Bodur, Paul E. Renaud, Lucie Goraguer, Martí Amargant-Arumí, Philipp Assmy, Anna Maria Dabrowska, Miriam Marquardt, Angelika H. H. Renner, Agnieszka Tatarek, and Marit Reigstad. Seasonal patterns of vertical flux in the northwestern Barents Sea under Atlantic Water influence and sea-ice decline. Progress in Oceanography, 219(??):??, December 2023. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661123001751.

Bonnet:2005:OCH

[BRH^{+}05] Delphine Bonnet, Anthony Richardson, Roger Harris, Andrew Hirst, Gregory Beaugrand, Martin Edwards, Sara Ceballos, Rabea Diekman, Angel López-Urrutia, Luis Valdes, François Carlotti, Juan Carlos Molinero, Horst Weikert, Wulf Greve, Davor Lucic, Aitor Albaina, Nejib Daly Yahia, Serena Fonda Umani, Ana Miranda, Antonina dos Santos, Kathryn Cook, Susan Robinson, and Marie Luz Fernandez de Puelles. An overview of Calanus helgolandicus ecology in European waters. Progress in Oceanography, 65(1):1–53, April 2005. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661105000212.

Brinton:1979:PRD

[Bri79] Edward Brinton. Parameters relating to the distributions of planktonic organisms, especially euphausiids in the eastern tropical Pacific. Progress in Oceanography, 8 (3):125–168, IN1–IN2, 169–189, ???? 1979. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0079661179900016.

Brink:1983:NSD

[Bri83] K. H. Brink. The near-surface dynamics of coastal upwelling. Progress in Oceanography, 12(3):223–257, ???? 1983. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0079661183900095.

Broecker:1982:GIC

[Bro82] Wallace S. Broecker. Glacial to interglacial changes in ocean chemistry. Progress in Oceanography, 11(2):151–197, ???? 1982. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0079661182900076.

Bouffard:2012:SSS

[BRR^{+}12] J. Bouffard, L. Renault, S. Ruiz, A. Pascual, C. Dufau, and J. Tintoré. Sub-surface small-scale eddy dynamics from multi-sensor observations and modeling. Progress in Oceanography, 106(??):62–79, November 2012. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661112000705.

Bahamonde:2022:VMS

[BRR^{+}22] Francisco Bahamonde, Juan Pablo Rodríguez, Sebastián Rosenfeld, Fabio Méndez, Zambra López, Karin Gerard, and Andrés Mansilla. Variability at multiple spatial scales in intertidal and subtidal macrobenthic communities in a fjord with glaciers, Magellanic Subantarctic ecoregion, Chile. Progress in Oceanography, 208(??):??, November 2022. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661122001380.

Bruni:1988:MPS

[Bru88] Vivia Bruni. In memory of professor Sebastiano Genovese. Progress in Oceanography, 21(2):107–108, ???? 1988. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0079661188900286.

Bellan-Santini:1990:MDS

[BS90] Denise Bellan-Santini. Mediterranean deep-sea amphipods: composition, structure and affinities of the fauna. Progress in Oceanography, 24(??):275–287, ???? 1990. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0079661190900373.

Bonisch:1995:DWF

[BS95] Gerhard Bönisch and Peter Schlosser. Deep water formation and exchange rates in the Greenland/Norwegian Seas and the Eurasian Basin of the Arctic Ocean derived from tracer balances. Progress in Oceanography, 35(1):29–52, ???? 1995. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/007966119500004Z.

Boyd:2002:MPT

[BS02] P. W. Boyd and C. L. Stevens. Modelling particle transformations and the downward organic carbon flux in the NE Atlantic Ocean. Progress in Oceanography, 52(1):1–29, January 2002. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661102000204.

Bolanos:2006:NNW

[BSA06] Rodolfo Bolaños and Agustín Sánchez-Arcilla. A note on nearshore wave features: Implications for wave generation. Progress in Oceanography, 70(2–4):168–180, August 2006. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661106000802.

Bernal:2007:SSE

[BSC^{+}07] M. Bernal, Y. Stratoudakis, S. Coombs, M. M. Angelico, A. Lago de Lanzós, C. Porteiro, Y. Sagarminaga, M. Santos, A. Uriarte, E. Cunha, L. Valdés, and D. Borchers. Sardine spawning off the European Atlantic coast: Characterization of and spatio-temporal variability in spawning habitat. Progress in Oceanography, 74(2–3):210–227, August 2007. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661107000687.

Bai:2019:SSV

[BSC^{+}19] Youcheng Bai, Marie-Alexandrine Sicre, Jianfang Chen, Vincent Klein, Haiyan Jin, Jian Ren, Hongliang Li, Bin Xue, Zhongqiang Ji, Yanpei Zhuang, and Meixun Zhao. Seasonal and spatial variability of sea ice and phytoplankton biomarker flux in the Chukchi Sea (western Arctic Ocean). Progress in Oceanography, 171(??):22–37, February 2019. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661118300727.

Bauch:1995:FBS

[BSF95] Dorothea Bauch, Peter Schlosser, and Richard G. Fairbanks. Freshwater balance and the sources of deep and bottom waters in the Arctic Ocean inferred from the distribution of H_{2} ^{18}O. Progress in Oceanography, 35(1):53–80, ???? 1995. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0079661195000052.

Berg:2021:ELG

[BSF^{+}21] Florian Berg, Samina Shirajee, Arild Folkvord, Jane Aanestad Godiksen, Georg Skaret, and Aril Slotte. Early life growth is affecting timing of spawning in the semelparous Barents Sea capelin (Mallotus villosus). Progress in Oceanography, 196(??): ??, August 2021. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661121001014.

Berline:2012:ISM

[BSFM^{+}12] Léo Berline, Ioanna Siokou-Frangou, Ivona Marasović, Olja Vidjak, Ma Luz Fernández de Puelles, Maria Grazia Mazzocchi, Georgia Assimakopoulou, Soultana Zervoudaki, Serena Fonda-Umani, Alessandra Conversi, Carmen Garcia-Comas, Frédéric Ibanez, Stéphane Gasparini, Lars Stemmann, and Gabriel Gorsky. Intercomparison of six Mediterranean zooplankton time series. Progress in Oceanography, 97–100(??): 76–91, May 2012. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661111001297.

Boyd:2020:SSF

[BSH^{+}20] R. J. Boyd, R. Sibly, K. Hyder, N. Walker, R. Thorpe, and S. Roy. Simulating the summer feeding distribution of Northeast Atlantic mackerel with a mechanistic individualbased model. Progress in Oceanography, 183(??):??, April 2020. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661120300380.

Baltar:2015:BCC

[BSMC15] Federico Baltar, Esther Stuck, Sergio Morales, and Kim Currie. Bacterioplankton carbon cycling along the Subtropical Frontal Zone off New Zealand. Progress in Oceanography, 135(??):168– 175, June 2015. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661115001160.

Besiktepe:1994:CHM

[BSÖ^{+}94]  Şükrü T. Beşiktepe, Halil İ. Sur, Emin Özsoy, M. Abdul Latif,

Temel Oğuz, and Ümit Ünlüata. The circulation and hydrography of the Marmara Sea. Progress in Oceanography, 34(4):285– 334, ???? 1994. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661194900183.

Bishop:1986:PMD

[BSW86] James K. B. Bishop, Jeanne C. Stepien, and Peter H. Wiebe. Particulate matter distributions, chemistry and flux in the Panama Basin: response to environment forcing. Progress in Oceanography, 17(1–2):1–59, ???? 1986. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0079661186900248.

Boyd:2007:UEB

[BT07] P. W. Boyd and T. W. Trull. Understanding the export of biogenic particles in oceanic waters: Is there consensus? Progress in Oceanography, 72(4):276–312, March 2007. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661106001340.

Bianchi:2003:MAB

[BTG^{+}03] A. Bianchi, O. Tholosan, J. Garcin, T. Polychronaki, A. Tselepides, R. Buscail, and G. Duineveld. Microbial activities at the benthic boundary layer in the Aegean Sea. Progress in Oceanography, 57(2):219–236, May 2003. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S007966110300034X.

Brewin:2017:MSF

[BTJ^{+}17] Robert J. W. Brewin, Gavin H. Tilstone, Thomas Jackson, Terry Cain, Peter I. Miller, Priscila K. Lange, Ankita Misra, and Ruth L. Airs. Modelling size-fractionated primary production in the Atlantic Ocean from remote sensing. Progress in Oceanography, 158(??):130–149, November 2017. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661116300350.

Balopoulos:1999:MAO

[BTK^{+}99] E. Th Balopoulos, A. Theocharis, H. Kontoyiannis, S. Varnavas, F. Voutsinou-Taliadouri, A. Iona, A. Souvermezoglou, L. Ignatiades, O. Gotsis-Skretas, and A. Pavlidou. Major advances in the oceanography of the southern Aegean Sea– Cretan Straits system (eastern Mediterranean). Progress in Oceanography, 44(1–3):109–130, August 1999. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661199000221.

Basedow:2013:CQZ

[BTNK13] Sünnje L. Basedow, Kurt S. Tande, M. Fredrika Norrbin, and Stian A. Kristiansen. Capturing quantitative zooplankton information in the sea: Performance test of laser optical plankton counter and video plankton recorder in a Calanus finmarchicus dominated summer situation. Progress in Oceanography, 108(??):72–80, January 2013. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661112001292.

Barnes:2015:DEK

[BTS^{+}15a] Morvan K. Barnes, Gavin H. Tilstone, Timothy J. Smyth, Claire E. Widdicombe, Johanna Gloël, Carol Robinson, Jan Kaiser, and David J. Suggett. Drivers and effects of Karenia mikimotoi blooms in the western English Channel. Progress in Oceanography, 137(??):456–469, September 2015. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661115000853.

Barnes:2015:TVT

[BTS^{+}15b] Morvan K. Barnes, Gavin H. Tilstone, David J. Suggett, Claire E. Widdicombe, John Bruun, Victor Martinez-Vicente, and Timothy J. Smyth. Temporal variability in total, micro- and nano-phytoplankton primary production at a coastal site in the Western English Channel. Progress in Oceanography, 137(??):470–483, September 2015. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661115000841.

Brandini:2022:ERB

[BTS22] Frederico P. Brandini, Pedro M. Tura, and Pedro P. G. M. Santos. Ecosystem responses to biogeochemical fronts in the South Brazil Bight. Progress in Oceanography, 164(??):52– 62, ???? 2022. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661117302781.

Burridge:2017:DDH

[BTV^{+}17] Alice K. Burridge, Marloes Tump, Ronald Vonk, Erica Goetze, and Katja T. C. A. Peijnenburg. Diversity and distribution of hyperiid amphipods along a latitudinal transect in the Atlantic Ocean. Progress in Oceanography, 158(??):224–235, November 2017. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661116300386.

Bumpus:1973:DCC

[Bum73] Dean F. Bumpus. A description of the circulation on the continental shelf of the east coast of the United States. Progress in Oceanography, 6(??):111–157, ???? 1973. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0079661173900062.

Bianchi:1988:HPE

[BVB88] Micheline Bianchi, France Van Wambeke, and Armand Bianchi. Heterotrophic processes in eutrophicated aquuatic marine environments. Progress in Oceanography, 21(2):159– 166, ???? 1988. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661188900341.

Bauch:2011:OFP

[BvdLA^{+}11] Dorothea Bauch, Michiel Rutgers van der Loeff, Nils Andersen, Sinhue Torres-Valdes, Karel Bakker, and E. Povl Abrahamsen. Origin of freshwater and polynya water in the Arctic Ocean halocline in summer 2007. Progress in Oceanography, 91(4):482–495, December 2011. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661111000711.

Bandara:2019:AEB

[BVJE19] Kanchana Bandara, Øystein Varpe, Rubao Ji, and Ketil Eiane. Artificial evolution of behavioral and life history strategies of high-latitude copepods in response to bottom-up and top-down selection pressures. Progress in Oceanography, 173(??):134– 164, April 2019. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661118301204.

Bandy:1965:MPP

[BW65] Orville L. Bandy and Mary E. Wade. Miocene–Pliocene– Pleistocene boundaries in deep-water environments. Progress in Oceanography, 4(??):51–62, IN7, 63–66, September 2, 1965. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0079661165900406.

Bakun:2008:MEP

[BW08] Andrew Bakun and Scarla J. Weeks. The marine ecosystem off Peru: What are the secrets of its fishery productivity and what might its future hold? Progress in Oceanography, 79(2–4):290–299, October 2008. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S007966110800181X.

Blanchette:2009:TSD

[BWB^{+}09] C. A. Blanchette, E. A. Wieters, B. R. Broitman, B. P. Kinlan, and D. R. Schiel. Trophic structure and diversity in rocky intertidal upwelling ecosystems: a comparison of community patterns across California, Chile, South Africa and New Zealand. Progress in Oceanography, 83(1–4):107–116, December 2009. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661109001037.

Ballon:2008:ION

[BWMGCB08] Michael Ballón, Claudia Wosnitza-Mendo, Renato Guevara-Carrasco, and Arnaud Bertrand. The impact of overfishing and El Niño on the condition factor and reproductive success of Peruvian hake, Merluccius gayi peruanus. Progress in Oceanography, 79(2–4):300–307, October 2008. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661108001821.

Barral:2021:CFW

[BZD^{+}21] Quentin-Boris Barral, Bruno Zakardjian, Franck Dumas, Pierre Garreau, Pierre Testor, and Jonathan Beuvier. Characterization of fronts in the Western Mediterranean with a special focus on the North Balearic Front. Progress in Oceanography, 197(??):??, September–October 2021. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S007966112100121X.

Cavaleri:2007:WMS

[CAA^{+}07] L. Cavaleri, J.-H. G. M. Alves, F. Ardhuin, A. Babanin, M. Banner, K. Belibassakis, M. Benoit, M. Donelan, J. Groeneweg, T. H. C. Herbers, P. Hwang, P. A. E. M. Janssen, T. Janssen, I. V. Lavrenov, R. Magne, J. Monbaliu, M. Onorato, V. Polnikov, D. Resio, W. E. Rogers, A. Sheremet, J. McKee Smith, H. L. Tolman, G. van Vledder, J. Wolf, and I. Young. Wave modelling — the state of the art. Progress in Oceanography, 75(4):603–674, December 2007. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661107001206.

Crise:1999:MPE

[CAB^{+}99] A. Crise, J. I. Allen, J. Baretta, G. Crispi, R. Mosetti, and C. Solidoro. The Mediterranean pelagic ecosystem response to physical forcing. Progress in Oceanography, 44(1–3):219–243, August 1999. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661199000270.

Cavaleri:2018:WMC

[CAB^{+}18] L. Cavaleri, S. Abdalla, A. Benetazzo, L. Bertotti, J.-R. Bidlot, Ø. Breivik, S. Carniel, R. E. Jensen, J. Portilla-Yandun, W. E. Rogers, A. Roland, A. Sanchez-Arcilla, J. M. Smith, J. Staneva, Y. Toledo, G. Ph. van Vledder, and A. J. van der Westhuysen. Wave modelling in coastal and inner seas. Progress in Oceanography, 167(??):164–233, October 2018. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661117302975.

Cadaillon:2022:SDP

[CAH^{+}22] A. M. Cadaillon, G. O. Almandoz, M. P. Hernando, L. Saravia, S. Maldonado, and I. R. Schloss. Spatiotemporal distribution of paralytic shellfish poisoning (PSP) toxins in Beagle Channel (South America) during 2005–2017. Progress in Oceanography, 204(??):??, June 2022. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661122000192.

Cai:1995:GPD

[Cai95] Wenju Cai. Global present-day ocean climate and its stability under various surface thermohaline forcing conditions derived from levitus climatology. Progress in Oceanography, 36(3):219– 247, ???? 1995. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/007966119500016X.

Chaudhary:2023:PGS

[CALS^{+}23] Chhaya Chaudhary, Joan M. Alfaro-Lucas, Marianna V. P. Simões, Angelika Brandt, and Hanieh Saeedi. Potential geographic shifts in the coral reef ecosystem under climate change. Progress in Oceanography, 213(??):??, April–May 2023. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661123000447.

Ceballos:2006:SDR

[CÁM06] Sara Ceballos and Florentina Álvarez-Marqués. Seasonal dynamics of reproductive parameters of the calanoid copepods Calanus helgolandicus and Calanoides carinatus in the Cantabrian Sea (SW Bay of Biscay). Progress in Oceanography, 70(1):1–26, July 2006. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661106000176.

Chatzievangelou:2020:VMK

[CAO^{+}20] Damianos Chatzievangelou, Jacopo Aguzzi, Andrea Ogston, Alejandro Suárez, and Laurenz Thomsen. Visual monitoring of key deep-sea megafauna with an Internet operated crawler as a tool for ecological status assessment.

Progress in Oceanography, 184(??):??, May 2020. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661120300604.

Cartwright:1997:STS

[Car97a] David E. Cartwright. Some thoughts on the spring-neap cycle of tidal dissipation. Progress in Oceanography, 40(??):125– 133, ???? 1997. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661197000268.

Cartwright:1997:TSC

[Car97b] David E. Cartwright. Tidal science 1996: concluding remarks. Progress in Oceanography, 40(??):437–438, ???? 1997. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661198000123.

Cartes:1998:DBB

[Car98] Joan E. Cartes. Dynamics of the bathyal benthic boundary layer in the northwestern Mediterranean: depth and temporal variations in macrofaunal–megafaunal communities and their possible connections within deep-sea trophic webs. Progress in Oceanography, 41(1):111–139, January 1998. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661198000184.

Calderon-Aguilera:2022:TAP

[CARBML^{+}22] Luis E. Calderon-Aguilera, Héctor Reyes-Bonilla, Hem Nalini Morzaria-Luna, Juan Carlos Perusquía-Ardón, Manuel Olán-González, and M. Fernanda Méndez-Martínez. Trophic architecture as a predictor of ecosystem resilience and resistance in the eastern Pacific. Progress in Oceanography, 209(??):??, December 2022. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661122001811.

Chiba:2008:CRS

[CAT^{+}08] Sanae Chiba, Maki N. Aita, Kazuaki Tadokoro, Toshiro Saino, Hiroya Sugisaki, and Kaoru Nakata. From climate regime shifts to lower-trophic level phenology: Synthesis of recent progress in retrospective studies of the western North Pacific. Progress in Oceanography, 77(2–3):112–126, May 2008. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661108000499.

Chapman:1991:STV

[CB91] P. Chapman and G. W. Bailey. Short-term variability during an anchor station study in the southern Benguela upwelling system: Introduction. Progress in Oceanography, 28(1–2):1– 7, ???? 1991. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/007966119190018H.

Carson:2006:UFD

[CB06] Steven R. Carson and Pierre E. Biscaye. Unusual features of the distribution of radon-222 in the bottom waters of the continental slope of the Mid-Atlantic Bight: Hydrodynamic implications. Progress in Oceanography, 70(2–4):271–288, August 2006. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661106000553.

Checkley:2009:PPC

[CB09] David M. Checkley and John A. Barth. Patterns and processes in the California Current System. Progress in Oceanography, 83(1–4):49–64, December 2009. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661109001098.

Chowdhury:2017:ELT

[CB17] Piyali Chowdhury and Manasa Ranjan Behera. Effect of long-term wave climate variability on longshore sediment transport along regional coastlines. Progress in Oceanography, 156(??):145–153, August 2017. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661117300149.

Castro:2002:ILN

[CBB^{+}02] C. G. Castro, T. R. Baumgartner, S. Bograd, R. Castro, F. P. Chavez, C. A. Collins, R. Durazo, J. García, G. Gaxiola-Castro, T. Hayward, A. Huyer, R. Lynn, A. S. Mascarenhas, M. R. D. Robert, R. L. Smith, P. A. Wheeler, and F. A. Whitney. Introduction to ‘the 1997–8 El Niño atlas of oceanographic conditions along the west coast of North America (23^{o}N-50^{o} n)’. Progress in Oceanography, 54(1–4):503– 511, July 2002. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661102000666.

Clark:2015:YAW

[CBB^{+}15] Christopher W. Clark, Catherine L. Berchok, Susanna B. Blackwell, David E. Hannay, Josh Jones, Dimitri Ponirakis, and Kathleen M. Stafford. A year in the acoustic world of bowhead whales in the Bering, Chukchi and Beaufort Seas. Progress in Oceanography, 136(??):223–240, August 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661115001044.

Cavaleri:2019:OSN

[CBB^{+}19] L. Cavaleri, M. Bajo, F. Barbariol, M. Bastianini, A. Benetazzo, L. Bertotti, J. Chiggiato, S. Davolio, C. Ferrarin, L. Magnusson, A. Papa, P. Pezzutto, A. Pomaro, and G. Umgiesser. The October 29, 2018 storm in Northern Italy — an exceptional event and its modeling. Progress in Oceanography, 178(??):??, November 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661119301089.

Cavaleri:2022:ESS

[CBB^{+}22a] L. Cavaleri, A. Benetazzo, L. Bertotti, J.-R. Bidlot, A. Pomaro, and J. Portilla-Yandun. The 2015 exceptional swell in the Southern Pacific: Generation, advection, forecast and implied extremes. Progress in Oceanography, 206(??):??, August 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S007966112200101X.

Cavaleri:2022:OSN

[CBB^{+}22b] Luigi Cavaleri, Francesco Barbariol, Luciana Bertotti, Giovanni Besio, and Francesco Ferrari. The 29 October 2018 storm in Northern Italy: Its multiple actions in the Ligurian Sea. Progress in Oceanography, 201(??):??, February 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661121001981.

Chambault:2022:SOI

[CBB^{+}22c] Philippine Chambault, Alberto Baudena, Karen A. Bjorndal, Marco A. R. Santos, Alan B. Bolten, and Frederic Vandeperre. Swirling in the ocean: Immature loggerhead turtles seasonally target old anticyclonic eddies at the fringe of the North Atlantic gyre. Progress in Oceanography, 175(??):345– 358, ???? 2022. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661118303458.

Carmack:2006:CVP

[CBC^{+}06] Eddy Carmack, David Barber, Jens Christensen, Robie Macdonald, Bert Rudels, and Egil Sakshaug. Climate variability and physical forcing of the food webs and the carbon budget on panarctic shelves. Progress in Oceanography, 71(2–4):145–181, October 2006. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661106001285.

Cossa:2024:STA

[CBD^{+}24] Daniel Cossa, Roselyne Buscail, Bernard Dennielou, Olivier Radakovitch, Pere Puig, Alexis Khripounoff, Bernard Boutier, and Serge Berné. Sources, transport, and accumulation of mercury in the northwestern Mediterranean margin sediments during the Industrial Era and influence of turbiditic events. Progress in Oceanography, 220(??):??, January 2024. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S007966112300229X.

Chavez:2008:NHC

[CBGC^{+}08] Francisco P. Chavez, Arnaud Bertrand, Renato Guevara-Carrasco, Pierre Soler, and Jorge Csirke. The northern Humboldt Current System: Brief history, present status and a view towards the future. Progress in Oceanography, 79(2–4):95–105, October 2008. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661108001651.

Carlotti:2007:DGR

[CBHL07] François Carlotti, Delphine Bonnet, and Claudia Halsband-Lenk. Development and growth rates of Centropages typicus. Progress in Oceanography, 72(2–3):164–194, February 2007. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661107000055.

Cordier:2019:BFD

[CBL^{+}19] Tristan Cordier, Inès Barrenechea, Franck Lejzerowicz, Emanuela Reo, and Jan Pawlowski. Benthic foraminiferal DNA metabarcodes significantly vary along a gradient from abyssal to hadal depths and between each side of the Kuril– Kamchatka Trench. Progress in Oceanography, 178(??):??, November 2019. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661119301818.

Chowdhury:2021:SMD

[CBM^{+}21] Mintu Chowdhury, Haimanti Biswas, Aditi Mitra, Saumya Silori, Diksha Sharma, Debasmita Bandyopadhyay, Aziz Ur Rahman Shaik, Veronica Fernandes, and Jayu Narvekar. Southwest monsoon-driven changes in the phytoplankton community structure in the central Arabian Sea (2017– 2018): After two decades of JGOFS. Progress in Oceanography, 197(??):??, September–October 2021. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661121001397.

Cheung:2015:PFC

[CBOP15] William W. L. Cheung, Richard D. Brodeur, Thomas A. Okey, and Daniel Pauly. Projecting future changes in distributions of pelagic fish species of Northeast Pacific shelf seas. Progress in Oceanography, 130(??):19–31, January 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661114001505.

Crosswell:2022:GCM

[CBPS^{+}22] Joseph R. Crosswell, Francisco Bravo, Iván Pérez-Santos, Geoffrey Carlin, Nagur Cherukuru, Cassie Schwanger, Rob Gregor, and Andrew D. L. Steven. Geophysical controls on metabolic cycling in three Patagonian fjords. Progress in Oceanography, 207(??):??, September–October 2022. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661122001252.

Crawford:2007:MED

[CBT07] William R. Crawford, Peter J. Brickley, and Andrew C. Thomas. Mesoscale eddies dominate surface phytoplankton in northern Gulf of Alaska. Progress in Oceanography, 75(2):287–303, October 2007. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661107001504.

Cavari:1988:EPB

[CC88] Benzion Cavari and Rita R. Colwell. Effect of pollution on the bacterial community structure in the coastal waters of the eastern Mediterranean Sea. Progress in Oceanography, 21(2):147– 157, ???? 1988. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/007966118890033X.

Chaudhary:2023:MST

[CC23] Chhaya Chaudhary and Mark John Costello. Marine species turnover but not richness, peaks at the Equator. Progress in Oceanography, 210(??):??, January 2023. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661122002002.

Collins:2002:CHC

[CCA^{+}02] C. A. Collins, C. G. Castro, H. Asanuma, T. A. Rago, S.- K. Han, R. Durazo, and F. P. Chavez. Changes in the hydrography of central California waters associated with the 1997–98 El Niño. Progress in Oceanography, 54(1–4):129– 147, July 2002. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661102000460.

Cresson:2020:PPD

[CCB^{+}20] Pierre Cresson, Tiphaine Chouvelon, Paco Bustamante, Daniela Bănaru, Jérome Baudrier, François Le Loc’h, Aourell Mauffret, Benoit Mialet, Jérome Spitz, Nathalie Wessel, Marine J. Briand, Margaux Denamiel, Matthieu Doray, Gaël Guillou, Angélique Jadaud, Coline Lazard, Solène Prieur, Manuel Rouquette, Claire Saraux, Sandrine Serre, Charles-André Timmerman, Yves Verin, and Mireille Harmelin-Vivien. Primary production and depth drive different trophic structure and functioning of fish assemblages in French marine ecosystems. Progress in Oceanography, 186(??):??, July 2020. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661120300823.

Combes:2013:CST

[CCD^{+}13] V. Combes, F. Chenillat, E. Di Lorenzo, P. Rivière, M. D. Ohman, and S. J. Bograd. Cross-shore transport variability in the California Current: Ekman upwelling vs. eddy dynamics. Progress in Oceanography, 109(??):78–89, February 2013. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661112001164.

Calbet:2007:FEC

[CCG07] Albert Calbet, François Carlotti, and Raymond Gaudy. The feeding ecology of the copepod Centropages typicus (Kröyer). Progress in Oceanography, 72(2–3):137–150, February 2007. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661107000067.

Coyle:2012:MFO

[CCH^{+}12] Kenneth O. Coyle, Wei Cheng, Sarah L. Hinckley, Evelyn J. Lessard, Terry Whitledge, Albert J. Hermann, and Kate Hedstrom. Model and field observations of effects of circulation on the timing and magnitude of nitrate utilization and production on the northern Gulf of Alaska shelf. Progress in Oceanography, 103(??):16–41, September 2012. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661112000195.

Chavez:2002:NAW

[CCHM02] Francisco P. Chavez, Curtis A. Collins, Adriana Huyer, and David L. Mackas. El Niño along the west coast of North America. Progress in Oceanography, 54(1–4):1–5, July 2002. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S007966110200040X.

Chen:2021:TVP

[CCHV^{+}21] Chia-Ting Chen, François Carlotti, Mireille Harmelin-Vivien, Loic Guilloux, and Daniela Bănaru. Temporal variation in prey selection by adult European sardine (Sardina pilchardus) in the NW Mediterranean Sea. Progress in Oceanography, 196(??):??, August 2021. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S007966112100104X.

Canals:2013:ISM

[CCM^{+}13] M. Canals, J. B. Company, D. Martín, A. Sànchez-Vidal, and E. Ramírez-Llodrà. Integrated study of Mediterranean deep canyons: Novel results and future challenges. Progress in Oceanography, 118(??):1–27, November 2013. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661113001742.

Contreras:2014:ECE

[CCM^{+}14] Tabit Contreras, Leonardo R. Castro, Sandra Montecinos, Humberto E. Gonzalez, Samuel Soto, Maria I. Muñoz, and Sergio Palma. Environmental conditions, early life stages distributions and larval feeding of Patagonian sprat Sprattus fuegensis and common sardine Strangomera bentincki in fjords and channels of the northern Chilean Patagonia. Progress in Oceanography, 129(??):136–148, December 2014. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661114001682.

Colas:2008:NPN

[CCMS08] F. Colas, X. Capet, J. C. McWilliams, and A. Shchepetkin. 1997–1998 El Niño off Peru: a numerical study. Progress in Oceanography, 79(2–4):138–155, October 2008.

CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661108001687.

Cardoso:2020:IET

[CCRS20] Cláudio Cardoso, Rui M. A. Caldeira, Paulo Relvas, and Alexandre Stegner. Islands as eddy transformation and generation hotspots: Cabo Verde case study. Progress in Oceanography, 184(??):??, May 2020. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661120300094.

Claiborne:2021:CBS

[CCS^{+}21] Andrew M. Claiborne, Lance Campbell, Bethany Stevick, Todd Sandell, James P. Losee, Marisa Litz, and Joseph H. Anderson. Correspondence between scale growth, feeding conditions, and survival of adult Chinook salmon returning to Puget Sound and coastal Washington: Implications for forecasting. Progress in Oceanography, 198(??):??, November 2021. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661120301828.

Castro:2002:NVD

[CCW^{+}02] C. G. Castro, C. A. Collins, P. Walz, J. T. Pennington, R. P. Michisaki, G. Friederich, and F. P. Chavez. Nutrient variability during El Niño 1997–98 in the California Current System off central California. Progress in Oceanography, 54(1–4):171– 184, July 2002. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661102000484.

Crawford:2018:SPA

[CCW^{+}18] David W. Crawford, Adrián O. Cefarelli, Ian A. Wrohan, Shea N. Wyatt, and Diana E. Varela. Spatial patterns in abundance, taxonomic composition and carbon biomass of nano- and microphytoplankton in subarctic and Arctic Seas. Progress in Oceanography, 162(??):132–159, March 2018. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S007966111630221X.

Cita:1965:CFS

[CD65] Maria Bianca Cita and Sara D’Onofrio. Climatic fluctuations in submarine cores from the Adriatic Sea (Mediterranean). Progress in Oceanography, 4(??):161–178, September 2, 1965. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661165900492.

Carter:1987:E

[CD87] Bill Carter and Robert Devoy. Editorial. Progress in Oceanography, 18(1–4):v, ???? 1987. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/007966118790022X.

Combes:2007:IFI

[CD07] Vincent Combes and Emanuele Di Lorenzo. Intrinsic and forced interannual variability of the Gulf of Alaska mesoscale circulation. Progress in Oceanography, 75(2):266–286, October 2007. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661107001486.

Carr:2022:SSV

[CDB^{+}22] Nealy Carr, Clare E. Davis, Sabena Blackbird, Lucie R. Daniels, Calum Preece, Malcolm Woodward, and Claire Mahaffey. Seasonal and spatial variability in the optical characteristics of DOM in a temperate shelf sea. Progress in Oceanography, 177(??):??, October 2022. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661117302112.

Curtis:2024:ICM

[CDB^{+}24] Emma J. Curtis, Jennifer M. Durden, Brian J. Bett, Veerle A. I. Huvenne, Nils Piechaud, Jenny Walker, James Albrecht, Miquel Massot-Campos, Takaki Yamada, Adrian Bodenmann, Jose Cappelletto, James A. Strong, and Blair Thornton. Improving coral monitoring by reducing variability and bias in cover estimates from seabed images. Progress in Oceanography, 222(??):??, March 2024. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S007966112400020X.

Cotte:2015:FPS

[CdD^{+}15] Cédric Cotté, Francesco d’Ovidio, Anne-Cécile Dragon, Christophe Guinet, and Marina Lévy. Flexible preference of southern elephant seals for distinct mesoscale features within the Antarctic Circumpolar Current. Progress in Oceanography, 131(??):46–58, February 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661114002006.

Capuano:2011:WLO

[CDDF11] P. Capuano, E. De Lauro, S. De Martino, and M. Falanga. Water-level oscillations in the Adriatic Sea as coherent self-oscillations inferred by independent component analysis. Progress in Oceanography, 91(4):447–460, December 2011. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661111000541.

Campbell:2013:NPS

[CDH^{+}13] Rose Campbell, Frédéric Diaz, ZiYuan Hu, Andrea Doglioli, Anne Petrenko, and Ivan Dekeyser. Nutrients and plankton spatial distributions induced by a coastal eddy in the Gulf of Lion. Insights from a numerical model. Progress in Oceanography, 109(??):47–69, February 2013. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661112001140.

Canals:2019:RAU

[CDL19] Miquel Canals, Roberto Danovaro, and Gian Marco Luna. Recent advances in understanding the ecology and functioning of submarine canyons in the Mediterranean Sea. Progress in Oceanography, 179(??):??, December 2019. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S007966111930357X.

Cascao:2022:SED

[CDL^{+}22] Irma Cascão, Réka Domokos, Marc O. Lammers, Ricardo S. Santos, and Mónica A. Silva. Seamount effects on the diel vertical migration and spatial structure of micronekton.

Progress in Oceanography, 175(??):1–13, ???? 2022. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661116301185.

Cossa:2018:SEM

[CdMS^{+}18] Daniel Cossa, Xavier Durrieu de Madron, Jörg Schäfer, Stéphane Guédron, Nicolas Marusczak, Sabine Castelle, and Jean-Jacques Naudin. Sources and exchanges of mercury in the waters of the Northwestern Mediterranean margin. Progress in Oceanography, 163(??):172–183, April 2018. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661116301963.

Connelly:2014:TIB

[CDP14] Tara L. Connelly, Don Deibel, and Christopher C. Parrish. Trophic interactions in the benthic boundary layer of the Beaufort Sea shelf, Arctic Ocean: Combining bulk stable isotope and fatty acid signatures. Progress in Oceanography, 120(??):79–92, January 2014. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661113001456.

Cahet:1990:SEW

[CDS90] Guy Cahet, Raoul Daumas, and Myriam Sibuet. In situ experimentation at the water/sediment interface in the deep sea:

2. Biotransformation of dissolved organic substrates by microbial communities at 2000 m depth in the Bay of Biscay. Progress in Oceanography, 24(??):169–178, ???? 1990. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/007966119090028Z.

Chambault:2016:IOF

[CdTH^{+}16] Philippine Chambault, Benoît de Thoisy, Karine Heerah, Anna Conchon, Sébastien Barrioz, Virginie Dos Reis, Rachel Berzins, Laurent Kelle, Baptiste Picard, Fabien Roquet, Yvon Le Maho, and Damien Chevallier. The influence of oceanographic features on the foraging behavior of the olive ridley sea turtle Lepidochelys olivacea along the Guiana coast.

Progress in Oceanography, 142(??):58–71, March 2016. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661115300094.

Cabral:2021:CIC

[CDTM^{+}21] Henrique Cabral, Hilaire Drouineau, Ana Teles-Machado, Maud Pierre, Mario Lepage, Jérémy Lobry, Patrick Reis-Santos, and Susanne E. Tanner. Contrasting impacts of climate change on connectivity and larval recruitment to estuarine nursery areas. Progress in Oceanography, 196(??):??, August 2021. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661121000951.

Collier:1984:TEG

[CE84] Robert Collier and John Edmond. The trace element geochemistry of marine biogenic particulate matter. Progress in Oceanography, 13(2):113–199, ???? 1984. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0079661184900089.

Chaigneau:2009:EAF

[CED09] Alexis Chaigneau, Gérard Eldin, and Boris Dewitte. Eddy activity in the four major upwelling systems from satellite altimetry (1992–2007). Progress in Oceanography, 83(1–4):117–123, December 2009. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661109001013.

Cunliffe:2013:SSM

[CEF^{+}13] Michael Cunliffe, Anja Engel, Sanja Frka, Blaženka Gašparović, Carlos Guitart, J. Colin Murrell, Matthew Salter, Christian Stolle, Robert Upstill-Goddard, and Oliver Wurl. Sea surface microlayers: a unified physicochemical and biological perspective of the air–ocean interface. Progress in Oceanography, 109(??):104–116, February 2013. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661112000924.

Chaban:2019:EMO

[CES^{+}19] Elena M. Chaban, Irina A. Ekimova, Dimitry M. Schepetov, Peter C. Kohnert, Michael Schrödl, and Alexei V. Chernyshev. Euopisthobranch mollusks of the order Cephalaspidea (Gastropoda: Heterobranchia) of the Kuril–Kamchatka Trench and the adjacent Pacific abyssal plain with descriptions of three new species of the genus Spiraphiline (Philinidae). Progress in Oceanography, 178(??):??, November 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S007966111930182X.

Cummins:2007:VNP

[CF07] Patrick F. Cummins and Howard J. Freeland. Variability of the North Pacific Current and its bifurcation. Progress in Oceanography, 75(2):253–265, October 2007. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661107001462.

Cornejo:2012:MVV

[CF12] Marcela Cornejo and Laura Farías. Meridional variability of the vertical structure and air–sea fluxes of N_{2}O off central Chile (30–40^{o}S). Progress in Oceanography, 92–95(??):33–42, January 2012. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S007966111100070X.

Cartes:2020:DSI

[CF20] Joan E. Cartes and Diego F. Figueroa. Deep sea isopods from the western Mediterranean: Distribution and habitat. Progress in Oceanography, 188(??):??, October 2020. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661120301543.

Cossa:2018:CST

[CFC^{+}18] D. Cossa, A.-S. Fanget, J.-F. Chiffoleau, M.-A. Bassetti, R. Buscail, B. Dennielou, K. Briggs, M. Arnaud, S. Guédron, and S. Berné. Chronology and sources of trace elements accumulation in the Rhône pro-delta sediments (Northwestern Mediterranean) during the last 400 years. Progress in Oceanography, 163(??):161–171, April 2018. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661116301070.

Cornejo:2007:SCV

[CFG07] Marcela Cornejo, Laura Farías, and Mauricio Gallegos. Seasonal cycle of N_{2}O vertical distribution and air–sea fluxes over the continental shelf waters off central Chile (≈ 36^{o}S). Progress in Oceanography, 75(3):383–395, November 2007. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661107001589.

Casalbore:2018:COF

[CFM^{+}18] Daniele Casalbore, Federico Falcini, Eleonora Martorelli, Eleonora Morelli, Alessandro Bosman, Marilena Calarco, and Francesco Latino Chiocci. Characterization of overbanking features on the lower reach of the Gioia–Mesima canyon– channel system (southern Tyrrhenian Sea) through integration of morpho-stratigraphic data and physical modelling. Progress in Oceanography, 169(??):66–78, December 2018. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661117302574.

Conese:2022:FWS

[CFML22] I. Conese, E. Fanelli, S. Miserocchi, and L. Langone. Food web structure and trophodynamics of deep-sea plankton from the Bari Canyon and adjacent slope (Southern Adriatic, central Mediterranean Sea). Progress in Oceanography, 175(??):92– 104, ???? 2022. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S007966111830096X.

Crawford:2007:LPO

[CGB07] William Crawford, Jake Galbraith, and Nick Bolingbroke. Line P ocean temperature and salinity, 1956–2005. Progress in Oceanography, 75(2):161–178, October 2007. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661107001498.

Civitarese:2023:BMH

[CGB^{+}23] Giuseppe Civitarese, Miroslav Gačić, Mirna Batistić, Manuel Bensi, Vanessa Cardin, Jakov Dulčić, Rade Garić, and Milena Menna. The BiOS mechanism: History, theory, implications. Progress in Oceanography, 216(??):??, August 2023. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S007966112300099X.

Castro:2020:SIS

[CGC^{+}20] Leonardo R. Castro, Valeria González, Gabriel Claramunt, Pamela Barrientos, and Samuel Soto. Stable isotopes (δ ^{13}C, δ ^{15}N) seasonal changes in particulate organic matter and in different life stages of anchoveta (Engraulis ringens) in response to local and large scale oceanographic variations in north and central Chile. Progress in Oceanography, 186(??): ??, July 2020. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661120300811.

Coclet:2018:TMC

[CGD^{+}18] Clément Coclet, Cédric Garnier, Floriane Delpy, Dominique Jamet, Gaël Durrieu, Christophe Le Poupon, Marie Mayer, and Benjamin Misson. Trace metal contamination as a toxic and structuring factor impacting ultraphytoplankton communities in a multicontaminated Mediterranean coastal area. Progress in Oceanography, 163(??):196–213, April 2018. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661116302361.

Cravatte:2022:OCS

[CGD^{+}22] Sophie Cravatte, Alexandre Ganachaud, Quoc-Phi Duong, William S. Kessler, Gérard Eldin, and Pierre Dutrieux. Observed circulation in the Solomon Sea from SADCP data. Progress in Oceanography, 88(1–4):116–130, ???? 2022. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661110001679.

Chaigneau:2008:MEP

[CGG08] Alexis Chaigneau, Arnaud Gizolme, and Carmen Grados. Mesoscale eddies off Peru in altimeter records: Identification algorithms and eddy spatio-temporal patterns. Progress in Oceanography, 79(2–4):106–119, October 2008. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661108001663.

Cai:2020:PFD

[CGL^{+}20] Zhongya Cai, Jianping Gan, Zhiqiang Liu, Chiwing Rex Hui, and Junlu Li. Progress on the formation dynamics of the layered circulation in the South China Sea. Progress in Oceanography, 181(??):??, February 2020. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661119304264.

Cartes:2002:BCD

[CGM^{+}02] J. E. Cartes, A. Grémare, F. Maynou, S. Villora-Moreno, and A. Dinet. Bathymetric changes in the distributions of particulate organic matter and associated fauna along a deep-sea transect down the Catalan Sea slope (Northwestern Mediterranean). Progress in Oceanography, 53(1):29–56, April 2002. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S007966110200023X.

Carracedo:2014:SDA

[CGMP14] L. I. Carracedo, M. Gilcoto, H. Mercier, and F. F. Pérez. Seasonal dynamics in the Azores–Gibraltar Strait region: a climatologically-based study. Progress in Oceanography, 122(??):116–130, March 2014. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661113002437.

Chen:2023:CAN

[CGS23] Zhuo Chen, Ting Gu, and Jun Sun. Community assembly and network stability of picoeukaryotic plankton communities in the northeast Indian Ocean. Progress in Oceanography, 219(??):??, December 2023. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661123002008.

Carscadden:2013:CRC

[CGV13a] James E. Carscadden, Harald Gjøsæter, and Hjálmar Vilhjálmsson. A comparison of recent changes in distribution of capelin (Mallotus villosus) in the Barents Sea, around Iceland and in the Northwest Atlantic. Progress in Oceanography, 114(??):64– 83, July 2013. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661113000475.

Carscadden:2013:RBS

[CGV13b] James E. Carscadden, Harald Gjøsæter, and Hjálmar Vilhjálmsson. Recruitment in the Barents Sea, Icelandic, and eastern Newfoundland/Labrador capelin (Mallotus villosus) stocks. Progress in Oceanography, 114(??):84–96, July 2013. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661113000487.

Cao:2022:IIT

[CGW^{+}22] Anzhou Cao, Zheng Guo, Shuya Wang, Xinyu Guo, and Jinbao Song. Incoherence of the M_{2} and K_{1} internal tides radiated from the Luzon Strait under the influence of looping and leaping Kuroshio. Progress in Oceanography, 206(??):??, August 2022. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661122001094.

Chen:2016:CAO

[CGZ^{+}16] Changsheng Chen, Guoping Gao, Yu Zhang, Robert C. Beardsley, Zhigang Lai, Jianhua Qi, and Huichan Lin. Circulation in the Arctic Ocean: Results from a high-resolution coupled ice-sea nested Global-FVCOM and Arctic-FVCOM system. Progress in Oceanography, 141(??):60–80, February 2016. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661115002499.

Carlotti:2007:P

[CH07a] F. Carlotti and R. Harris. Preface. Progress in Oceanography, 72(2–3):115–116, February 2007. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661107000122.

Carlotti:2007:BEC

[CH07b] Francois Carlotti and Roger Harris. The biology and ecology of Centropages typicus: an introduction. Progress in Oceanography, 72(2–3):117–120, February 2007. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661107000146.

Conners:2002:RAB

[CHB02] M. E. Conners, A. B. Hollowed, and E. Brown. Retrospective analysis of Bering Sea bottom trawl surveys: regime shift and ecosystem reorganization. Progress in Oceanography, 55(1–2):209–222, October 2002. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661102000794.

Cheng:2012:MMN

[CHC^{+}12] Wei Cheng, Albert J. Hermann, Ken O. Coyle, Elizabeth L. Dobbins, Nancy B. Kachel, and Phyllis J. Stabeno. Macroand micro-nutrient flux to a highly productive submarine bank in the Gulf of Alaska: a model-based analysis of daily and interannual variability. Progress in Oceanography, 101(1):63–77, August 2012. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S007966111200002X.

Cherubin:2014:HRS

[Ché14] Laurent M. Chérubin. High-resolution simulation of the circulation in the Bahamas and Turks and Caicos Archipelagos. Progress in Oceanography, 127(??):21–46, September 2014. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661114000925.

Carter:2018:OES

[CHG^{+}18] Gareth D. O. Carter, Veerle A. I. Huvenne, Jennifer A. Gales, Claudio Lo Iacono, Leigh Marsh, Audrey Ougier-Simonin, Katleen Robert, and Russell B. Wynn. Ongoing evolution of submarine canyon rockwalls; examples from the whittard canyon, Celtic margin (NE Atlantic). Progress in Oceanography, 169(??):79–88, December 2018. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S007966111730277X.

Choi:1986:SHE

[Cho86] B. H. Choi. Surge hindcast in the East China Sea. Progress in Oceanography, 17(3–4):177–192, ???? 1986. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0079661186900431.

Correa:2024:MLG

[CHS^{+}24] Giancarlo M. Correa, Thomas P. Hurst, William T. Stockhausen, Lorenzo Ciannelli, Trond Kristiansen, and Darren J. Pilcher. Modeling the larval growth and survival of Pacific cod (Gadus macrocephalus) in the eastern Bering Sea. Progress in Oceanography, 225(??):??, July 2024. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661124000880.

Clavel-Henry:2021:LTA

[CHSB^{+}21] Morane Clavel-Henry, Jordi Solé, Nixon Bahamon, Marta Carretón, and Joan Baptista Company. Larval transport of Aristeus antennatus shrimp (Crustacea: Decapoda: Dendrobranchiata: Aristeidae) near the Palamós submarine canyon (NW Mediterranean Sea) linked to the North Balearic Front. Progress in Oceanography, 192(??):??, March 2021. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661121000057.

Ciappa:2014:CPA

[Cia14] Achille Carlo Ciappa. The controversial path of Atlantic Water in the Eastern Mediterranean. Progress in Oceanography, 123(??):74–83, April 2014. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661114000147.

Ciappa:2022:EMH

[Cia22] Achille Carlo Ciappa. Effects of Marine Heatwaves (MHW) and Cold Spells (MCS) on the surface warming of the Mediterranean Sea from 1989 to 2018. Progress in Oceanography, 205 (??):??, July 2022. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661122000891.

Cannaby:2023:ECM

[CIL^{+}23] Heather Cannaby, Randi B. Ingvaldsen, Øyvind Lundesgaard, Angelika H. H. Renner, Georg Skaret, Serdar S. Sakinan, Terje Hovland, Melissa Chierici, and Harald Gjøsæter. Environmental controls on macrozooplankton and fish distributions over diurnal to seasonal time scales in the northern Barents Sea. Progress in Oceanography, 219(??):??, December 2023. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661123002021.

Craven:1992:MUT

[CJ92] D. B. Craven and R. A. Jahnke. Microbial utilization and turnover of organic carbon in Santa Monica Basin sediments. Progress in Oceanography, 30(1–4):313–333, ???? 1992. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/007966119290017T.

Cuomo:1988:ODM

[CJG88] V. Cuomo, E. B. Gareth Jones, and S. Grasso. Occurrence and distribution of marine fungi along the coast of the Mediterranean Sea. Progress in Oceanography, 21(2):189–200, ???? 1988. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0079661188900389.

Cochrane:1991:STV

[CJMI^{+}91] K. L. Cochrane, A. G. James, B. A. Mitchell-Innes, G. C. Pitcher, H. M. Verheye, and D. R. Walker. Short-term variability during an anchor station study in the southern Benguela upwelling system: a simulation model. Progress in Oceanography, 28(1–2):121–152, ???? 1991. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/007966119190023F.

Carter:1987:SLS

[CJMO87] R. W. G. Carter, T. W. Johnston, J. McKenna, and J. D. Orford. Sea-level, sediment supply and coastal changes: Examples from the coast of Ireland. Progress in Oceanography, 18(1–4):79–101, ???? 1987. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0079661187900279.

Castro-Jimenez:2013:ADD

[CJRÁ^{+}13] Javier Castro-Jiménez, Guiomar Rotllant, Manuela Ábalos, Jordi Parera, Jordi Dachs, Joan B. Company, Antoni Calafat, and Esteban Abad. Accumulation of dioxins in deep-sea crustaceans, fish and sediments from a submarine canyon (NW Mediterranean). Progress in Oceanography, 118(??):260–272, November 2013. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661113001304.

Calleja:2017:EIG

[CKB^{+}17] Maria Ll. Calleja, P. Kerhervé, S. Bourgeois, M. Kedra, A. Leynaert, E. Devred, M. Babin, and N. Morata. Effects of increase glacier discharge on phytoplankton bloom dynamics and pelagic geochemistry in a high Arctic fjord. Progress in Oceanography, 159(??):195–210, December 2017. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S007966111630180X.

Cuif:2014:WIV

[CKL^{+}14] Marion Cuif, David Michael Kaplan, Jérôme Lefèvre, Vincent Martin Faure, Matthieu Caillaud, Philippe Verley, Laurent Vigliola, and Christophe Lett. Wind-induced variability in larval retention in a coral reef system: a biophysical modelling study in the South-West Lagoon of New Caledonia. Progress in Oceanography, 122(??):105–115, March 2014. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661113002449.

Carvalho:2021:PSC

[CKM^{+}21] A. C. O. Carvalho, R. Kerr, C. R. B. Mendes, J. L. L. Azevedo, and V. M. Tavano. Phytoplankton strengthen CO_{2} uptake in the South Atlantic Ocean. Progress in Oceanography, 190(??): ??, January 2021. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661120302111.

Capelle:2020:ETO

[CKP^{+}20] David W. Capelle, Zou Zou A. Kuzyk, Tim Papakyriakou, Céline Guéguen, Lisa A. Miller, and Robie W. Macdonald. Effect of terrestrial organic matter on ocean acidification and CO_{2} flux in an Arctic shelf sea. Progress in Oceanography, 185(??):??, June 2020. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661120300586.

Codispoti:2013:SPP

[CKT^{+}13] L. A. Codispoti, V. Kelly, A. Thessen, P. Matrai, S. Suttles, V. Hill, M. Steele, and B. Light. Synthesis of primary production in the Arctic Ocean: III. nitrate and phosphate based estimates of net community production. Progress in Oceanography, 110(??):126–150, March 2013. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661112001723.

Corten:2003:UCD

[CL03] A. Corten and J. A. Lindley. The use of CPR data in fisheries research. Progress in Oceanography, 58(2–4):285–300, August 2003. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661103001332.

Chronis:2000:SPS

[CLA^{+}00] G. Chronis, V. Lykousis, C. Anagnostou, A. Karageorgis, S. Stavrakakis, and S. Poulos. Sedimentological processes in the southern margin of the Cretan Sea (NE Mediterranean). Progress in Oceanography, 46(2–4):143–162, August 2000. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661100000161.

Catry:2013:PDT

[CLB^{+}13] P. Catry, R. T. Lemos, P. Brickle, R. A. Phillips, R. Matias, and J. P. Granadeiro. Predicting the distribution of a threatened albatross: the importance of competition, fisheries and annual variability. Progress in Oceanography, 110(??):1– 10, March 2013. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661113000153.

Chen:2014:MTM

[CLB^{+}14] Changsheng Chen, Zhigang Lai, Robert C. Beardsley, Jun Sasaki, Jian Lin, Huichan Lin, Rubao Ji, and Yunfang Sun. The March 11, 2011 Tōhoku M9.0 earthquake-induced tsunami and coastal inundation along the Japanese coast: a model assessment. Progress in Oceanography, 123(??):84–104, April 2014. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661114000111.

Cavan:2019:EEM

[CLCBB19] E. L. Cavan, E. C. Laurenceau-Cornec, M. Bressac, and P. W. Boyd. Exploring the ecology of the mesopelagic biological pump. Progress in Oceanography, 176(??):??, September 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661119300370.

Carugati:2022:PDM

[CLD22] L. Carugati, M. Lo Martire, and R. Danovaro. Patterns and drivers of meiofaunal assemblages in the canyons Polcevera and Bisagno of the Ligurian Sea (NW Mediterranean Sea). Progress in Oceanography, 175(??):81–91, ???? 2022. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661118300934.

Couret:2023:YMB

[CLdPHL23] María Couret, José María Landeira, Ángelo Santana del Pino, and Santiago Hernández-León. A 50-year (1971–2021) mesozooplankton biomass data collection in the Canary Current System: Base line, gaps, trends, and future prospect.

Progress in Oceanography, 216(??):??, August 2023. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661123001167.

Chronis:2000:SPM

[CLG^{+}00] G. Chronis, V. Lykousis, D. Georgopoulos, V. Zervakis, S. Stavrakakis, and S. Poulos. Suspended particulate matter and nepheloid layers over the southern margin of the Cretan Sea (N.E. Mediterranean): seasonal distribution and dynamics. Progress in Oceanography, 46(2–4):163–185, August 2000. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661100000173.

Corbera:2022:ECM

[CLG^{+}22] Guillem Corbera, Claudio Lo Iacono, Eulàlia Gràcia, Jordi Grinyó, Martina Pierdomenico, Veerle A. I. Huvenne, Ricardo Aguilar, and Josep Maria Gili. Ecological characterisation of a Mediterranean cold-water coral reef: Cabliers Coral Mound Province (Alboran Sea, western Mediterranean). Progress in Oceanography, 175(??):245–262, ???? 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661118303677.

Coppola:2018:SIA

[CLL^{+}18] Laurent Coppola, Louis Legendre, Dominique Lefevre, Louis Prieur, Vincent Taillandier, and Emilie Diamond Riquier. Seasonal and inter-annual variations of dissolved oxygen in the northwestern Mediterranean Sea (DYFAMED site). Progress in Oceanography, 162(??):187–201, March 2018. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661117302434.

Command:2023:FLM

[CLMR23] Rylan J. Command, Fabio C. De Leo, Cynthia H. McKenzie, and Katleen Robert. A first look at megabenthic community responses to seasonal change using the new Holyrood Subsea Observatory in Conception Bay, NL. Progress in Oceanography, 216(??):??, August 2023. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661123001143.

Campanya-Llovet:2018:FQQ

[CLSD18] Neus Campanyà-Llovet, Paul V. R. Snelgrove, and Fabio C. De Leo. Food quantity and quality in Barkley Canyon (NE Pacific) and its influence on macroinfaunal community structure. Progress in Oceanography, 169(??):106–119, December 2018. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661117301842.

Campanya-Llovet:2017:RIF

[CLSP17] Neus Campanyà-Llovet, Paul V. R. Snelgrove, and Christopher C. Parrish. Rethinking the importance of food quality in marine benthic food webs. Progress in Oceanography, 156(??):240–251, August 2017. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661115300471.

Cardenas:2019:BPD

[CLV^{+}19] Paola Cárdenas, Carina B. Lange, Maria Vernet, Oliver Esper, Benjamin Srain, Maria-Elena Vorrath, Sophie Ehrhardt, Juliane Müller, Gerhard Kuhn, Helge W. Arz, Lester Lembke-Jene, and Frank Lamy. Biogeochemical proxies and diatoms in surface sediments across the Drake Passage reflect oceanic domains and frontal systems in the region. Progress in Oceanography, 174(??):72–88, May 2019. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661118301903.

Chen:2020:MLA

[CLX^{+}20] Bingzhang Chen, Hongbin Liu, Wupeng Xiao, Lei Wang, and Bangqin Huang. A machine-learning approach to modeling picophytoplankton abundances in the South China Sea. Progress in Oceanography, 189(??):??, November–December 2020. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661120301932.

Chen:2022:SVS

[CLY22] Yinchao Chen, Qian P. Li, and Jiancheng Yu. Submesoscale variability of subsurface chlorophyll-a across eddy-driven fronts by glider observations. Progress in Oceanography, 209(??):??, December 2022. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661122001641.

Chavez:2009:CEB

[CM09] Francisco P. Chavez and Monique Messié. A comparison of Eastern boundary upwelling ecosystems. Progress in Oceanography, 83(1–4):80–96, December 2009. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661109000998.

Carmack:2011:TRP

[CM11] Eddy Carmack and Fiona McLaughlin. Towards recognition of physical and geochemical change in subarctic and Arctic Seas. Progress in Oceanography, 90(1–4):90–104, July 2011. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S007966111100022X.

Caruana:2014:VDC

[CM14a] Amandine M. N. Caruana and Gill Malin. The variability in DMSP content and DMSP lyase activity in marine dinoflagellates. Progress in Oceanography, 120(??):410–424, January 2014. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661113002061.

Cummins:2014:CVT

[CM14b] Patrick F. Cummins and Diane Masson. Climatic variability and trends in the surface waters of coastal British Columbia. Progress in Oceanography, 120(??):279–290, January 2014. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661113001882.

Combes:2018:PSC

[CM18a] Vincent Combes and Ricardo P. Matano. The Patagonian shelf circulation: Drivers and variability. Progress in Oceanography, 167(??):24–43, October 2018. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661117303130.

Cummins:2018:LFI

[CM18b] Patrick F. Cummins and Diane Masson. Low-frequency isopycnal variability in the Alaska Gyre from Argo. Progress in Oceanography, 168(??):310–324, November 2018. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661118301642.

Carreto:2016:EBF

[CMC^{+}16] José I. Carreto, Nora G. Montoya, Mario O. Carignan, Rut Akselman, E. Marcelo Acha, and Carla Derisio. Environmental and biological factors controlling the spring phytoplankton bloom at the Patagonian shelf-break front — degraded fucoxanthin pigments and the importance of microzooplankton grazing. Progress in Oceanography, 146(??):1–21, August 2016. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661115300215.

Cartes:2009:LTC

[CMF^{+}09] J. E. Cartes, F. Maynou, E. Fanelli, V. Papiol, and D. Lloris. Long-term changes in the composition and diversity of deepslope megabenthos and trophic webs off Catalonia (western Mediterranean): Are trends related to climatic oscillations? Progress in Oceanography, 82(1):32–46, July 2009. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661109000172.

Cartes:2011:NDP

[CMF11] J. E. Cartes, F. Maynou, and E. Fanelli. Nile damming as plausible cause of extinction and drop in abundance of deepsea shrimp in the western Mediterranean over broad spatial scales. Progress in Oceanography, 91(3):286–294, November 2011. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661111000139.

Cornejo:2015:USS

[CMF15] Marcela Cornejo, Alejandro A. Murillo, and Laura Farías. An unaccounted for N_{2}O sink in the surface water of the eastern subtropical South Pacific: Physical versus biological mechanisms. Progress in Oceanography, 137 (part A)(??):12–23, September 2015. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661114002262.

Chang:2015:ESE

[CMG15] Yu-Lin Chang, Yasumasa Miyazawa, and Xinyu Guo. Effects of the STCC eddies on the Kuroshio based on the 20-year JCOPE2 reanalysis results. Progress in Oceanography, 135(??): 64–76, June 2015. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S007966111500066X.

Chauvet:2018:ASD

[CMHM18] Pauline Chauvet, Anna Metaxas, Alex E. Hay, and Marjolaine Matabos. Annual and seasonal dynamics of deepsea megafaunal epibenthic communities in Barkley Canyon (British Columbia, Canada): a response to climatology, surface productivity and benthic boundary layer variation. Progress in Oceanography, 169(??):89–105, December 2018. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661117302501.

Casanova-Masjoan:2018:CAW

[CMJPH^{+}18] M. Casanova-Masjoan, T. M. Joyce, M. D. Pérez-Hernández, P. Vélez-Belchí, and A. Hernández-Guerra. Changes across 66^{o}W, the Caribbean Sea and the western boundaries of the North Atlantic subtropical gyre. Progress in Oceanography, 168(??):296–309, November 2018. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S007966111830034X.

Campbell:2016:CDB

[CML^{+}16] K. Campbell, C. J. Mundy, J. C. Landy, A. Delaforge, C. Michel, and S. Rysgaard. Community dynamics of bottom-ice algae in Dease Strait of the Canadian Arctic.

Progress in Oceanography, 149(??):27–39, December 2016. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661116300878.

Cartes:2004:PBD

[CMM^{+}04] Joan E. Cartes, Francesc Maynou, Joan Moranta, Enric Massutí, Domènec Lloris, and Beatriz Morales-Nin. Patterns of bathymetric distribution among deep-sea fauna at local spatial scale: comparison of mainland vs. insular areas. Progress in Oceanography, 60(1):29–45, January 2004. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661104000205.

Catala:2022:DOM

[CMPNC^{+}22] Teresa S. Catalá, Alba María Martínez-Pérez, Mar Nieto-Cid, Marta Álvarez, Jaime Otero, Mikhail Emelianov, Isabel Reche, Javier Arístegui, and Xosé Antón Álvarez-Salgado. Dissolved Organic Matter (DOM) in the open Mediterranean Sea. I. Basin-wide distribution and drivers of chromophoric DOM. Progress in Oceanography, 165(??):35–51, ???? 2022. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661117303609.

Catalan:2013:SMR

[CMS^{+}13] Ignacio A. Catalán, Diego Macías, Jordi Solé, Andrés Ospina-Álvarez, and Javier Ruiz. Stay off the motorway: Resolving the pre-recruitment life history dynamics of the European anchovy in the SW Mediterranean through a spatially-explicit individual-based model (SEIBM). Progress in Oceanography, 111(??):140–153, April 2013. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661113000189.

Caballero:2022:AEF

[CN22] Isabel Caballero and Gabriel Navarro. Application of extended full resolution MERIS imagery to assist coastal management of the area adjacent to the Guadalquivir estuary. Progress in Oceanography, 165(??):215–232, ???? 2022. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661117302367.

Cordeiro:2021:FPT

[CNBD21] Nuno G. F. Cordeiro, Rita Nolasco, Eric D. Barton, and Jesus Dubert. Fixed-point time series, repeat survey and highresolution modeling reveal event scale responses of the Northwestern Iberian upwelling. Progress in Oceanography, 190(??): ??, January 2021. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661120302159.

Cross:2015:RAD

[CNSHT15] Jaimie Cross, W. Alex M. Nimmo-Smith, Philip J. Hosegood, and Ricardo Torres. The role of advection in the distribution of plankton populations at a moored 1-D coastal observatory. Progress in Oceanography, 137(??):342–359, September 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S007966111500083X.

Checkley:2003:MMM

[CNT03] Dave Checkley, Jeffrey M. Napp, and Cynthia T. Tynan. Michael Mahlon Mullin, oceanographer: his work on plankton size and as a mentor. Progress in Oceanography, 57(3–4):243– 249, June 2003. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661103000995.

Cabrol:2019:SLS

[CNT^{+}19] Jory Cabrol, Jean-Bruno Nadalini, Réjean Tremblay, Peter S. Galbraith, Christian Nozais, Michel Starr, Stéphane Plourde, and Gesche Winkler. Seasonal and large-scale spatial variability of the energy reserves and the feeding selectivity of Meganyctiphanes norvegica and Thysanoessa inermis in a subarctic environment. Progress in Oceanography, 179(??):??, December 2019. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661119303830.

Colebrook:1969:VP

[Col69] J. M. Colebrook. Variability in the plankton. Progress in Oceanography, 5(??):115–125, ???? 1969. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0079661169900354.

Conchon:1987:HVC

[Con87] Odette Conchon. Hydrodynamic variations and correlations with Quaternary sea-level change on a rocky coast in Corsica, Western Mediterranean. Progress in Oceanography, 18(1–4):103–117, ???? 1987. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0079661187900280.

Cooper:1965:RPC

[Coo65] L. H. N. Cooper. Radiolarians as possible chronometers of continental drift. Progress in Oceanography, 3(??):71–82, ???? 1965. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0079661165900108.

Cooper:1969:EOT

[Coo69] L. H. N. Cooper. Extraction of oceanographical timeseries from the four-dimensional “liquidum” (or continuum). Progress in Oceanography, 5(??):127–148, ???? 1969. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0079661169900366.

Cowie:2005:BAS

[Cow05] Greg Cowie. The biogeochemistry of Arabian Sea surficial sediments: a review of recent studies. Progress in Oceanography, 65(2–4):260–289, May 2005. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661105000418.

Cox:1963:SP

[Cox63] Roland A. Cox. The salinity problem. Progress in Oceanography, 1(??):243–261, ???? 1963. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0079661163900065.

Clancy:1983:RTS

[CP83] R. Michael Clancy and Kenneth D. Pollak. A real-time synoptic ocean thermal analysis/forecast system. Progress in Oceanography, 12(4):383–424, ???? 1983. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0079661183900010.

Coyle:2002:CRD

[CP02] K. O. Coyle and A. I. Pinchuk. Climate-related differences in zooplankton density and growth on the inner shelf of the southeastern Bering Sea. Progress in Oceanography, 55(1–2):177–194, October 2002. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661102000770.

Cunningham:2007:SGC

[CP07] Stuart Cunningham and Marko Pavic. Surface geostrophic currents across the Antarctic Circumpolar Current in Drake Passage from 1992 to 2004. Progress in Oceanography, 73 (3–4):296–310, May 2007. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661107000778.

Carlotti:2010:TMA

[CP10] F. Carlotti and J. C. Poggiale. Towards methodological approaches to implement the zooplankton component in “end to end” food-web models. Progress in Oceanography, 84(1–2):20– 38, January 2010. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661109001360.

Corlett:2017:CSC

[CP17] W. Bryce Corlett and Robert S. Pickart. The Chukchi slope current. Progress in Oceanography, 153(??):50–65, April 2017. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661116301434.

Chernyshev:2019:NDS

[CP19] Alexei V. Chernyshev and Neonila E. Polyakova. Nemerteans from the deep-sea expedition KuramBio II with descriptions of three new hoplonemerteans from the Kuril–Kamchatka Trench. Progress in Oceanography, 178(??):??, November 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661119301867.

Canepa:2015:OIB

[CPB^{+}15] Elisa Canepa, Sara Pensieri, Roberto Bozzano, Marco Faimali, Pierluigi Traverso, and Luigi Cavaleri. The ODAS Italia 1 buoy: More than forty years of activity in the Ligurian Sea. Progress in Oceanography, 135(??):48–63, June 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661115000658.

Cuomo:1988:MLW

[CPC88] V. Cuomo, V. Pane, and P. Contesso. Marine litter: Where does the problem lie? Progress in Oceanography, 21(2):177– 180, ???? 1988. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661188900365.

Chavez:2002:BCC

[CPC^{+}02] F. P. Chavez, J. T. Pennington, C. G. Castro, J. P. Ryan, R. P. Michisaki, B. Schlining, P. Walz, K. R. Buck, A. Mc-Fadyen, and C. A. Collins. Biological and chemical consequences of the 1997–1998 El Niño in central California waters. Progress in Oceanography, 54(1–4):205–232, July 2002. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661102000502.

Castellanos:2015:RST

[CPC^{+}15] Paola Castellanos, Josep L. Pelegrí, Edmo J. D. Campos, Miquel Rosell-Fieschi, and Marc Gasser. Response of the surface tropical Atlantic Ocean to wind forcing. Progress in Oceanography, 134(??):271–292, May 2015. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661115000403.

Cartes:2008:FDD

[CPG08] Joan E. Cartes, Vanesa Papiol, and Beatriz Guijarro. The feeding and diet of the deep-sea shrimp Aristeus antennatus off the Balearic Islands (Western Mediterranean): Influence of environmental factors and relationship with the biological cycle. Progress in Oceanography, 79(1):37–54, October 2008. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661108001328.

Carracedo:2018:RCA

[CPG^{+}18] L. I. Carracedo, F. F. Pérez, M. Gilcoto, A. Velo, A. Padín, and G. Rosón. Role of the circulation on the anthropogenic CO_{2} inventory in the North-East Atlantic: a climatological analysis. Progress in Oceanography, 161(??):78–86, February 2018. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661117301544.

Castro:1998:CCM

[CPHR98] C. G. Castro, F. F. Pérez, S. E. Holley, and A. F. Ríos. Chemical characterisation and modelling of water masses in the Northeast Atlantic. Progress in Oceanography, 41(3):249– 279, June 1998. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661198000214.

Contreras:2007:BSS

[CPNL07] Sergio Contreras, Silvio Pantoja, Carlos Neira, and Carina B. Lange. Biogeochemistry of surface sediments off Concepción (≈ 36^{o}S), Chile: El Niño vs. non-El Niño conditions. Progress in Oceanography, 75(3):576–585, November 2007. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S007966110700170X.

Carman:2019:CFS

[CPO^{+}19] Victoria González Carman, Alberto Piola, Todd D. O’Brien, Dmitriy D. Tormosov, and E. Marcelo Acha. Circumpolar frontal systems as potential feeding grounds of southern right whales. Progress in Oceanography, 176(??):??, September 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661118302180.

Carreon-Palau:2022:ERL

[CPPPEAG22] Laura Carreón-Palau, Christopher C. Parrish, Horacio Pérez-España, and Sergio Aguiñiga-Garcia. Elemental ratios and lipid classes in a coral reef food web under river influence. Progress in Oceanography, 164(??):1–11, ???? 2022. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661117300022.

Close:2020:MEI

[CPSM20] S. Close, T. Penduff, S. Speich, and J.-M. Molines. A means of estimating the intrinsic and atmospherically-forced contributions to sea surface height variability applied to altimetric observations. Progress in Oceanography, 184(??):??, May 2020. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661120300537.

Crawford:2015:CRB

[CQC15] Justin A. Crawford, Lori T. Quakenbush, and John J. Citta. A comparison of ringed and bearded seal diet, condition and productivity between historical (1975–1984) and recent (2003–2012) periods in the Alaskan Bering and Chukchi seas. Progress in Oceanography, 136(??):133–150, August 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661115001081.

Citta:2015:ECC

[CQO^{+}15] John J. Citta, Lori T. Quakenbush, Stephen R. Okkonen, Matthew L. Druckenmiller, Wieslaw Maslowski, Jaclyn Clement-Kinney, John C. George, Harry Brower, Robert J. Small, Carin J. Ashjian, Lois A. Harwood, and Mads Peter Heide-Jørgensen. Ecological characteristics of core-use areas used by Bering–chukchi–beaufort (BCB) bowhead whales, 2006–2012. Progress in Oceanography, 136(??):201–222, August 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661114001438.

Celussi:2018:PPP

[CQZ^{+}18] Mauro Celussi, Grazia Marina Quero, Luca Zoccarato, Annalisa Franzo, Cinzia Corinaldesi, Eugenio Rastelli, Marco Lo Martire, Pierre E. Galand, Jean-François Ghiglione, Jacopo Chiggiato, Alessandro Coluccelli, Aniello Russo, Alberto Pallavicini, Serena Fonda Umani, Paola Del Negro, and Gian Marco Luna. Planktonic prokaryote and protist communities in a submarine canyon system in the Ligurian Sea (NW Mediterranean). Progress in Oceanography, 168(??):210–221, November 2018. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661118300582.

Chao:1997:OTA

[CR97] B. F. Chao and R. D. Ray. Oceanic tidal angular momentum and Earth’s rotation variations. Progress in Oceanography, 40(??):399–421, ???? 1997. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S007966119800010X.

Cummins:2020:STW

[CR20] Patrick F. Cummins and Tetjana Ross. Secular trends in water properties at Station P in the northeast Pacific: an updated analysis. Progress in Oceanography, 186(??):??, July 2020. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661120300689.

Crawford:2009:RIS

[Cra09] Robert J. M. Crawford. A recent increase of swift terns Thalasseus bergii off South Africa — the possible influence of an altered abundance and distribution of prey. Progress in Oceanography, 83(1–4):398–403, December 2009. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661109000767.

Corinaldesi:2019:HRV

[CRC^{+}19] Cinzia Corinaldesi, Eugenio Rastelli, Sara Canensi, Michael Tangherlini, Roberto Danovaro, and Antonio Dell’Anno. High rates of viral lysis stimulate prokaryotic turnover and C recycling in bathypelagic waters of a Ligurian canyon (Mediterranean Sea). Progress in Oceanography, 171(??):70–75, February 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661118301009.

Corbineau:2010:PVL

[CRF^{+}10] A. Corbineau, T. Rouyer, J.-M. Fromentin, B. Cazelles, A. Fonteneau, and F. Ménard. Patterns of variations in large pelagic fish: a comparative approach between the Indian and the Atlantic Oceans. Progress in Oceanography, 86(1–2):276– 282, July 2010. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661110000534.

Caceres:2017:PCS

[CRGA17] Carlos Cáceres, Antonella Rivera, Sonia González, and Ricardo Anadón. Phytoplankton community structure and dynamics in the North Atlantic subtropical gyre. Progress in Oceanography, 151(??):177–188, February 2017. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661116300155.

Correa-Ramirez:2012:SPA

[CRHM12] Marco A. Correa-Ramirez, Samuel E. Hormazabal, and Carmen E. Morales. Spatial patterns of annual and interannual surface chlorophyll-a variability in the Peru–Chile Current System. Progress in Oceanography, 92–95(??):8–17, January 2012. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661111000620.

Curchitser:2015:CMO

[CRiI^{+}15a] Enrique N. Curchitser, Kenneth A. Rose, Shin ichi Ito, Myron A. Peck, and Michio J. Kishi. Combining modeling and observations to better understand marine ecosystem dynamics. Progress in Oceanography, 138(??):327–330, November 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661115002177.

Curchitser:2015:MBM

[CRiI^{+}15b] Enrique N. Curchitser, Kenneth A. Rose, Shin ichi Ito, Myron A. Peck, and Michio J. Kishi. In memoriam: Bernard A. Megrey. Progress in Oceanography, 138(??):325–326, November 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661115002189.

Coupel:2015:IFP

[CRPS^{+}15] P. Coupel, D. Ruiz-Pino, M. A. Sicre, J. F. Chen, S. H. Lee, N. Schiffrine, H. L. Li, and J. C. Gascard. The impact of freshening on phytoplankton production in the Pacific Arctic Ocean. Progress in Oceanography, 131(??):113–125, February 2015. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661114002055.

Collie:2004:RSC

[CRS04] Jeremy S. Collie, Katherine Richardson, and John H. Steele. Regime shifts: Can ecological theory illuminate the mechanisms? Progress in Oceanography, 60(2–4):281–302, February 2004. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661104000321.

Chiaverano:2022:ERL

[CRT^{+}22] Luciano M. Chiaverano, Kelly L. Robinson, Jorge Tam, James J. Ruzicka, Javier Quiñones, Katrina T. Aleksa, Frank J. Hernandez, Richard D. Brodeur, Robert Leaf, Shin ichi Uye, Mary Beth Decker, Marcelo Acha, Hermes W. Mianzan, and William M. Graham. Evaluating the role of large jellyfish and forage fishes as energy pathways, and their interplay with fisheries, in the Northern Humboldt Current System. Progress in Oceanography, 164(??):28–36, ???? 2022. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661117303312.

Clegg:1989:HSM

[CS89] Simon L. Clegg and Jorge L. Sarmiento. The hydrolytic scavenging of metal ions by marine particulate matter.

Progress in Oceanography, 23(1):1–21, ???? 1989. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0079661189900232.

Chiba:2003:VMC

[CS03] Sanae Chiba and Toshiro Saino. Variation in mesozooplankton community structure in the Japan/East Sea (1991– 1999) with possible influence of the ENSO scale climatic variability. Progress in Oceanography, 57(3–4):317–339, June 2003. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661103001046.

Cury:2004:RSU

[CS04] Philippe Cury and Lynne Shannon. Regime shifts in upwelling ecosystems: observed changes and possible mechanisms in the northern and southern Benguela. Progress in Oceanography, 60(2–4):223–243, February 2004. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661104000291.

Condie:2006:SDT

[CS06] Scott A. Condie and Christopher R. Sherwood. Sediment distribution and transport across the continental shelf and slope under idealized wind forcing. Progress in Oceanography, 70(2–4):255–270, August 2006. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661106000528.

Cisewski:2016:AIZ

[CS16] Boris Cisewski and Volker H. Strass. Acoustic insights into the zooplankton dynamics of the eastern Weddell Sea. Progress in Oceanography, 144(??):62–92, May 2016. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S007966111630057X.

Chubarenko:2018:CIL

[CS18] Irina Chubarenko and Natalia Stepanova. Cold intermediate layer of the Baltic Sea: Hypothesis of the formation of its core. Progress in Oceanography, 167(??):1–10, October 2018. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661117301179.

Chaalali:2015:NMA

[CSBL^{+}15] Aurélie Chaalali, Blanche Saint-Béat, Géraldine Lassalle, François Le Loc’h, Samuele Tecchio, Georges Safi, Claude Savenkoff, Jérémy Lobry, and Nathalie Niquil. A new modeling approach to define marine ecosystems food-web status with uncertainty assessment. Progress in Oceanography, 135(??):37– 47, June 2015. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661115000531.

Chouvelon:2012:RUM

[CSC^{+}12] T. Chouvelon, J. Spitz, F. Caurant, P. Mèndez-Fernandez, A. Chappuis, F. Laugier, E. Le Goff, and P. Bustamante. Revisiting the use of δ ^{15}N in meso-scale studies of marine food webs by considering spatio-temporal variations in stable isotopic signatures — the case of an open ecosystem: the Bay of Biscay (North-East Atlantic). Progress in Oceanography, 101(1):92–105, August 2012. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661112000067.

Chouvelon:2015:IFA

[CSG^{+}15] T. Chouvelon, G. Schaal, J. Grall, F. Pernet, M. Perdriau, E. J. A-Pernet, and H. Le Bris. Isotope and fatty acid trends along continental shelf depth gradients: Inshore versus offshore hydrological influences on benthic trophic functioning. Progress in Oceanography, 138 (part A)(??):158–175, November 2015. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661115001652.

Canfield:2023:BIW

[CSH^{+}23] Katherine N. Canfield, Alexa R. Sterling, Christina M. Hernández, Sophie N. Chu, Bethanie R. Edwards, Diana N. Fontaine, Jillian M. Freese, Marissa S. Giroux, Aubree E. Jones, Alexandra J. McCarty, Hannah K. Morrissette, Hilary I.

Palevsky, Catherine E. Raker, Anna R. Robuck, Gabriela Serrato Marks, Patricia S. Thibodeau, and Anna E. Windle. Building an inclusive wave in marine science: Sense of belonging and Society for Women in Marine Science symposia. Progress in Oceanography, 218(??):??, November 2023. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661123001532.

Chiba:2012:PNP

[CSK^{+}12] Sanae Chiba, Hiroya Sugisaki, Akira Kuwata, Kazuaki Tadokoro, Toru Kobari, Atsushi Yamaguchi, and David L. Mackas. Pan-North Pacific comparison of long-term variation in Neocalanus copepods based on stable isotope analysis. Progress in Oceanography, 97–100(??):63–75, May 2012. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S007966111100125X.

Chelton:2003:ETR

[CSLJ03] D. B. Chelton, M. G. Schlax, J. M. Lyman, and G. C. Johnson. Equatorially trapped Rossby waves in the presence of meridionally sheared baroclinic flow in the Pacific Ocean. Progress in Oceanography, 56(2):323–380, February 2003. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661103000089.

Carmo:2015:FES

[CSM^{+}15] Vanda Carmo, Tracey Sutton, Gui Menezes, Tone Falkenhaug, and Odd Aksel Bergstad. Feeding ecology of the Stomiiformes (Pisces) of the northern Mid-Atlantic Ridge.

1. The Sternoptychidae and Phosichthyidae. Progress in Oceanography, 130(??):172–187, January 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661114001785.

Castro:2019:OST

[CSMGS19] Leonardo R. Castro, Samuel Soto-Mendoza, and Fernanda González-Saldías. Ontogenetic and short-term fluctuations in the residence depth of young pelagic stages of Munida gregaria in different zones of northern Patagonia. Progress in Oceanography, 174(??):173–184, May 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661118301113.

Carey:1990:BGR

[CSR90] A. G. Carey, D. L. Stein, and P. L. Rona. Benthos of the Gorda Ridge axial valley (NE Pacific Ocean): Taxonomic composition and trends in distribution. Progress in Oceanography, 24(??): 47–57, ???? 1990. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/007966119090018W.

Chelton:2011:GON

[CSS11] Dudley B. Chelton, Michael G. Schlax, and Roger M. Samelson. Global observations of nonlinear mesoscale eddies. Progress in Oceanography, 91(2):167–216, October 2011. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661111000036.

Chelton:2019:PFS

[CSS^{+}19] Dudley B. Chelton, Michael G. Schlax, Roger M. Samelson, J. Thomas Farrar, M. Jeroen Molemaker, James C. McWilliams, and Jonathan Gula. Prospects for future satellite estimation of small-scale variability of ocean surface velocity and vorticity. Progress in Oceanography, 173(??):256– 350, April 2019. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661117303610.

Catlett:2021:DSM

[CSS^{+}21] D. Catlett, D. A. Siegel, R. D. Simons, N. Guillocheau, F. Henderikx-Freitas, and C. S. Thomas. Diagnosing seasonal to multi-decadal phytoplankton group dynamics in a highly productive coastal ecosystem. Progress in Oceanography, 197(??):??, September–October 2021. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661121001221.

Cartes:2007:CSD

[CSV^{+}07] Joan E. Cartes, Alberto Serrano, Francisco Velasco, Santiago Parra, and Francisco Sánchez. Community structure and dynamics of deep-water decapod assemblages from Le Danois Bank (Cantabrian Sea, NE Atlantic): Influence of environmental variables and food availability. Progress in Oceanography, 75(4):797–816, December 2007. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661107001838.

Chao:1996:NMS

[CSW96] Shenn-Yu Chao, Ping-Tung Shaw, and Sunny Y. Wu. El Niño modulation of the South China Sea circulation. Progress in Oceanography, 38(1):51–93, ???? 1996. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661196000109.

Cazenave:2016:UGM

[CTA16] Pierre William Cazenave, Ricardo Torres, and J. Icarus Allen. Unstructured grid modelling of offshore wind farm impacts on seasonally stratified shelf seas. Progress in Oceanography, 145(??):25–41, June 2016. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661115300379.

Castro:2007:FSV

[CTF07] Leonardo R. Castro, Victor A. Troncoso, and Dante R. Figueroa. Fine-scale vertical distribution of coastal and offshore copepods in the golfo de arauco, central Chile, during the upwelling season. Progress in Oceanography, 75(3):486–500, November 2007. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661107001644.

Cuevas:2019:IBF

[CTI^{+}19] L. Antonio Cuevas, Fabián J. Tapia, José Luis Iriarte, Humberto E. González, Nelson Silva, and Cristian A. Vargas. Interplay between freshwater discharge and oceanic waters modulates phytoplankton size-structure in fjords and channel systems of the Chilean Patagonia. Progress in Oceanography, 173(??):103–113, April 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661118301836.

Castillo-Trujillo:2023:EEG

[CTKF^{+}23] Alma Carolina Castillo-Trujillo, Young-Oh Kwon, Paula Fratantoni, Ke Chen, Hyodae Seo, Michael A. Alexander, and Vincent S. Saba. An evaluation of eight global ocean reanalyses for the Northeast U.S. Continental Shelf. Progress in Oceanography, 219(??):??, December 2023. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661123001696.

Chang:2004:CCS

[CTL^{+}04] K.-I. Chang, W. J. Teague, S. J. Lyu, H. T. Perkins, D.- K. Lee, D. R. Watts, Y.-B. Kim, D. A. Mitchell, C. M. Lee, and K. Kim. Circulation and currents in the southwestern East/Japan Sea: Overview and review. Progress in Oceanography, 61(2–4):105–156, May 2004. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S007966110400059X.

Ciavatta:2014:ARS

[CTMV^{+}14] Stefano Ciavatta, Ricardo Torres, Victor Martinez-Vicente, Timothy Smyth, Giorgio Dall’Olmo, Luca Polimene, and J. Icarus Allen. Assimilation of remotely-sensed optical properties to improve marine biogeochemistry modelling. Progress in Oceanography, 127(??):74–95, September 2014. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661114001098.

Colas:2018:ZNF

[CTP^{+}18] F. Colas, M. Tardivel, J. Perchoc, M. Lunven, B. Forest, G. Guyader, M. M. Danielou, S. Le Mestre, P. Bourriau, E. Antajan, M. Sourisseau, M. Huret, P. Petitgas, and J. B. Romagnan. The ZooCAM, a new in-flow imaging system for fast onboard counting, sizing and classification of fish eggs and metazooplankton. Progress in Oceanography, 166(??):54–65, September 2018. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661117300873.

Corinaldesi:2019:HDB

[CTR^{+}19] Cinzia Corinaldesi, Michael Tangherlini, Eugenio Rastelli, Emanuela Buschi, Marco Lo Martire, Roberto Danovaro, and Antonio Dell’Anno. High diversity of benthic bacterial and archaeal assemblages in deep-Mediterranean canyons and adjacent slopes. Progress in Oceanography, 171(??):154–161, February 2019. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661118301174.

Chakraborty:2021:SCS

[CVBG21] Kunal Chakraborty, Vinu Valsala, Trishneeta Bhattacharya, and Jayashree Ghosh. Seasonal cycle of surface ocean pCO_{2} and pH in the northern Indian Ocean and their controlling factors. Progress in Oceanography, 198(??):??, November 2021. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661121001671.

Collart:2018:SPC

[CVHM^{+}18] Tim Collart, Wencke Verreydt, F. Javier Hernández-Molina, Estefanía Llave, Ricardo León, María Gómez-Ballesteros, Edwige Pons-Branchu, Heather Stewart, and David Van Rooij. Sedimentary processes and cold-water coral minimounds at the Ferrol canyon head, NW Iberian margin. Progress in Oceanography, 169(??):48–65, December 2018. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661117302410.

Corwith:2002:NRV

[CW02] H. L. Corwith and P. A. Wheeler. El Niño related variations in nutrient and chlorophyll distributions off Oregon. Progress in Oceanography, 54(1–4):361–380, July 2002. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661102000587.

Carmack:2006:FWP

[CW06] Eddy Carmack and Paul Wassmann. Food webs and physical–biological coupling on pan-Arctic shelves: Unifying concepts and comprehensive perspectives. Progress in Oceanography, 71(2–4):446–477, October 2006. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661106001297.

Capuzzo:2022:VSC

[CWB^{+}22] Elisa Capuzzo, Serena Wright, Paul Bouch, Kate Collingridge, Veronique Creach, Sophie Pitois, David Stephens, and Jeroen van der Kooij. Variability in structure and carbon content of plankton communities in autumn in the waters southwest of the UK. Progress in Oceanography, 204(??):??, June 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661122000660.

Clarke:2021:AGI

[CWS^{+}21] Tayler M. Clarke, Colette C. C. Wabnitz, Sandra Striegel, Thomas L. Frölicher, Gabriel Reygondeau, and William W. L. Cheung. Aerobic growth index (AGI): an index to understand the impacts of ocean warming and deoxygenation on global marine fisheries resources. Progress in Oceanography, 195(??): ??, July 2021. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661121000756.

Carmack:2015:CPR

[CWW15] Eddy Carmack, Peter Winsor, and William Williams. The contiguous panarctic Riverine Coastal Domain: a unifying concept. Progress in Oceanography, 139(??):13–23, December 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661115300069.

Cao:2020:SDE

[CWZ^{+}20] Zhimian Cao, Danna Wang, Zhouling Zhang, Kuanbo Zhou, Xin Liu, Lei Wang, Bangqin Huang, Pinghe Cai, and Minhan Dai. Seasonal dynamics and export of biogenic silica in the upper water column of a large marginal sea, the northern South China Sea. Progress in Oceanography, 188(??):??, October 2020. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661120301609.

Chen:2021:ILB

[CZG^{+}21] Changsheng Chen, Liuzhi Zhao, Scott Gallager, Rubao Ji, Pingguo He, Cabell Davis, Robert C. Beardsley, Deborah Hart, Wendy C. Gentleman, Lu Wang, Siqi Li, Huichan Lin, Kevin Stokesbury, and David Bethoney. Impact of larval behaviors on dispersal and connectivity of sea scallop larvae over the northeast U.S. shelf. Progress in Oceanography, 195(??): ??, July 2021. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661121000914.

Chen:2024:PIO

[CZL^{+}24] Changsheng Chen, Liuzhi Zhao, Huichan Lin, Pingguo He, Siqi Li, Zhongxiang Wu, Jianhua Qi, Qichun Xu, Kevin Stokesbury, and Lu Wang. Potential impacts of offshore wind energy development on physical processes and scallop larval dispersal over the US Northeast shelf. Progress in Oceanography, 224(??): ??, June 2024. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661124000697.

Chen:2022:ISI

[CZW^{+}22] Juntian Chen, Xiao-Hua Zhu, Min Wang, Hua Zheng, Ruixiang Zhao, Hirohiko Nakamura, and Toru Yamashiro. Incoherent signatures of internal tides in the Tokara Strait modulated by the Kuroshio. Progress in Oceanography, 206(??):??, August 2022. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661122001227.

Djeghri:2022:HPP

[DAF^{+}22a] Nicolas Djeghri, Angus Atkinson, Elaine S. Fileman, Rachel A. Harmer, Claire E. Widdicombe, Andrea J. McEvoy, Louise Cornwell, and Daniel J. Mayor. High prey-predator size ratios and unselective feeding in copepods: a seasonal comparison of five species with contrasting feeding modes.

Progress in Oceanography, 165(??):63–74, ???? 2022. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661117301258.

Djeghri:2022:RHP

[DAF^{+}22b] Nicolas Djeghri, Angus Atkinson, Elaine S. Fileman, Rachel A. Harmer, Claire E. Widdicombe, Andrea J. McEvoy, Louise Cornwell, and Daniel J. Mayor. Reprint of: High preypredator size ratios and unselective feeding in copepods: a seasonal comparison of five species with contrasting feeding modes. Progress in Oceanography, 177(??):??, October 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661118303409.

Dagg:1993:GCC

[Dag93] Michael Dagg. Grazing by the copepod community does not control phytoplankton production in the subarctic Pacific Ocean. Progress in Oceanography, 32(??):163–183, ???? 1993. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0079661193900123.

Dahl:1969:CAR

[Dah69] Odd Dahl. The capability of the aanderaa recording and telemetering instrument. Progress in Oceanography, 5(??):103– 106, ???? 1969. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661169900329.

Dufour:2010:CIA

[DAIS10] Florence Dufour, Haritz Arrizabalaga, Xabier Irigoien, and Josu Santiago. Climate impacts on albacore and bluefin tunas migrations phenology and spatial distribution. Progress in Oceanography, 86(1–2):283–290, July 2010. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661110000418.

Duarte:1999:MCT

[DAKV99] Carlos M. Duarte, Susana Agustí, Hilary Kennedy, and Dolors Vaqué. The Mediterranean climate as a template for Mediterranean marine ecosystems: the example of the northeast Spanish littoral. Progress in Oceanography, 44(1–3):245– 270, August 1999. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661199000282.

Diociaiuti:2022:VDM

[DAU22] Tommaso Diociaiuti, Fabrizio Bernardi Aubry, and Serena Fonda Umani. Vertical distribution of microbial communities abundance and biomass in two NW Mediterranean Sea submarine canyons. Progress in Oceanography, 175(??):14– 23, ???? 2022. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661118300375.

Davies:1985:TDM

[Dav85] A. M. Davies. Three dimensional modal model of wind induced flow in a sea region. Progress in Oceanography, 15(2):71– 128, ???? 1985. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661185900321.

Davos:1999:DSR

[Dav99] Climis A. Davos. On determining the social relevance of oceanography. Progress in Oceanography, 44(1–3):457–468, August 1999. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661199000361.

Dalpadado:2020:CET

[DAvD^{+}20] Padmini Dalpadado, Kevin R. Arrigo, Gert L. van Dijken, Hein Rune Skjoldal, Espen Bagøien, Andrey V. Dolgov, Irina P. Prokopchuk, and Erik Sperfeld. Climate effects on temporal and spatial dynamics of phytoplankton and zooplankton in the Barents Sea. Progress in Oceanography, 185(??):??, June 2020. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661120300598.

Dalpadado:2021:WIO

[DAvD^{+}21] Padmini Dalpadado, Kevin R. Arrigo, Gert L. van Dijken, Sudheera S. Gunasekara, Marek Ostrowski, Gabriella Bianchi, and Erik Sperfeld. Warming of the Indian Ocean and its impact on temporal and spatial dynamics of primary production. Progress in Oceanography, 198(??):??, November 2021. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661121001725.

deBaar:1994:LLM

[dB94] H. J. W. de Baar. von Liebig’s law of the minimum and plankton ecology (1899–1991). Progress in Oceanography, 33(4):347– 386, ???? 1994. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661194900221.

Durazo:2002:EOC

[DB02] R. Durazo and T. R. Baumgartner. Evolution of oceanographic conditions off Baja California: 1997–1999. Progress in Oceanography, 54(1–4):7–31, July 2002. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0079661102000411.

Dessier:2018:SMV

[DBC^{+}18] Aurélie Dessier, Paco Bustamante, Tiphaine Chouvelon, Martin Huret, Marc Pagano, Elise Marquis, Frédéric Rousseaux, Cécilia Pignon-Mussaud, Françoise Mornet, Martine Bréret, and Christine Dupuy. The spring mesozooplankton variability and its relationship with hydrobiological structure over year-to-year changes (2003–2013) in the southern Bay of Biscay (Northeast Atlantic). Progress in Oceanography, 166(??):76–87, September 2018. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661117300265.

Decker:2023:CVE

[DBC^{+}23] Mary Beth Decker, Richard D. Brodeur, Lorenzo Ciannelli, Lyle L. Britt, Nicholas A. Bond, Bartholomew P. Di-Fiore, and George L. Hunt. Cyclic variability of eastern Bering Sea jellyfish relates to regional physical conditions. Progress in Oceanography, 210(??):??, January 2023. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661122001823.

Durden:2015:AHH

[DBJ^{+}15] Jennifer M. Durden, Brian J. Bett, Daniel O. B. Jones, Veerle A. I. Huvenne, and Henry A. Ruhl. Abyssal hills — hidden source of increased habitat heterogeneity, benthic megafaunal biomass and diversity in the deep sea. Progress in Oceanography, 137 (part A)(??):209–218, September 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661115001391.

Dupont:2017:IAV

[DBM17] Nicolas Dupont, Espen Bagøien, and Webjørn Melle. Interannual variability in spring abundance of adult Calanus finmarchicus from the overwintering population in the southeastern Norwegian Sea. Progress in Oceanography, 152(??):75–85, March 2017. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661115300550.

Daskalov:2003:RSS

[DBR03] G. M. Daskalov, D. C. Boyer, and J. P. Roux. Relating sardine Sardinops sagax abundance to environmental indices in northern Benguela. Progress in Oceanography, 59(2–3):257–274, October 2003. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661103001733.

Durden:2020:SVA

[DBR20] Jennifer M. Durden, Brian J. Bett, and Henry A. Ruhl. Subtle variation in abyssal terrain induces significant change in benthic megafaunal abundance, diversity, and community structure. Progress in Oceanography, 186(??):??, July 2020. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661120301348.

Duhamel:2017:PDB

[DBRK17] Solange Duhamel, Karin M. Björkman, Daniel J. Repeta, and David M. Karl. Phosphorus dynamics in biogeochemically distinct regions of the southeast subtropical Pacific Ocean. Progress in Oceanography, 151(??):261–274, February 2017. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661115300707.

Davis:2022:SOM

[DBW^{+}22] Clare E. Davis, Sabena Blackbird, George Wolff, Malcolm Woodward, and Claire Mahaffey. Seasonal organic matter dynamics in a temperate shelf sea. Progress in Oceanography, 177(??):??, October 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661117302136.

Damini:2023:ASF

[DCD^{+}23] Brendon Yuri Damini, Raul Rodrigo Costa, Tiago S. Dotto, Carlos Rafael Borges Mendes, Juan Camilo Torres-Lasso, Marina do V. C. Azaneu, Mauricio M. Mata, and Rodrigo Kerr. Antarctica Slope Front bifurcation eddy: a stationary feature influencing CO_{2} dynamics in the northern Antarctic Peninsula. Progress in Oceanography, 212(??):??, March 2023. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661123000289.

Ferreira:2017:SWD

[dCFK17] Maria Luiza de Carvalho Ferreira and Rodrigo Kerr. Source water distribution and quantification of North Atlantic deep water and Antarctic bottom water in the Atlantic Ocean. Progress in Oceanography, 153(??):66–83, April 2017. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661116301227.

Davison:2013:CEM

[DCKB13] P. C. Davison, D. M. Checkley, J. A. Koslow, and J. Barlow. Carbon export mediated by mesopelagic fishes in the northeast Pacific Ocean. Progress in Oceanography, 116(??):14–30, September 2013. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661113000554.

Drinkwater:2013:CAF

[DCL^{+}13a] K. Drinkwater, E. Colbourne, H. Loeng, S. Sundby, and T. Kristiansen. Comparison of the atmospheric forcing and oceanographic responses between the Labrador Sea and the Norwegian and Barents seas. Progress in Oceanography, 114(??):11–25, July 2013. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661113000323.

Duran:2013:SDP

[DCL^{+}13b] Ruth Durán, Miquel Canals, Galderic Lastras, Aaron Micallef, David Amblas, Rut Pedrosa-Pàmies, and José Luis Sanz. Sediment dynamics and post-glacial evolution of the continental shelf around the Blanes submarine canyon head (NW Mediterranean). Progress in Oceanography, 118(??):28–46, November 2013. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661113001444.

Daigle:2016:REB

[DCM16] Rémi M. Daigle, Joël Chassé, and Anna Metaxas. The relative effect of behaviour in larval dispersal in a low energy embayment. Progress in Oceanography, 144(??):93–117, May 2016. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661116300684.

Dominguez-Carrio:2022:DSS

[DCRR^{+}22] Carlos Dominguez-Carrió, Joan Lluís Riera, Katleen Robert, Mikel Zabala, Susana Requena, Andrea Gori, Covadonga Orejas, Claudio Lo Iacono, Claude Estournel, Guillem Corbera, Stefano Ambroso, Maria Jesús Uriz, Pablo J. López-González, Rafael Sardá, and Josep-Maria Gili. Diversity, structure and spatial distribution of megabenthic communities in Cap de Creus continental shelf and submarine canyon (NW Mediterranean). Progress in Oceanography, 208(??):??, November 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661122001367.

Dudarev:2022:ESS

[DCS^{+}22] Oleg Dudarev, Alexander Charkin, Natalia Shakhova, Aleksey Ruban, Denis Chernykh, Jorien Vonk, Tommaso Tesi, Jannik Martens, Irina Pipko, Svetlana Pugach, Elena Gershelis, Andrey Leusov, Andrey Grinko, Örjan Gustafsson, and Igor Semiletov. East Siberian Sea: Interannual heterogeneity of the suspended particulate matter and its biogeochemical signature. Progress in Oceanography, 208(??):??, November 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661122001628.

Dubois-Dauphin:2023:STD

[DDCE^{+}23] Quentin Dubois-Dauphin, Christophe Colin, Mary Elliot, Julius Förstel, Frederic Haurine, Rosella Pinna, Eric Douville, and Norbert Frank. The spatial and temporal distribution of neodymium isotopic composition within the Rockall Trough. Progress in Oceanography, 218(??):??, November 2023. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661123001702.

Danovaro:2000:SCB

[DDD^{+}00] Roberto Danovaro, Norberto Della Croce, Antonio Dell’Anno, Mauro Fabiano, Daniela Marrale, and Daniela Martorano. Seasonal changes and biochemical composition of the labile organic matter flux in the Cretan Sea. Progress in Oceanography, 46(2–4):259–278, August 2000. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661100000227.

Danovaro:1999:BRP

[DDDT99] R. Danovaro, A. Dinet, G. Duineveld, and A. Tselepides. Benthic response to particulate fluxes in different trophic environments: a comparison between the Gulf of Lions–Catalan Sea (western-mediterranean) and the Cretan Sea (easternmediterranean). Progress in Oceanography, 44(1–3):287–312, August 1999. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661199000300.

Danovaro:1995:MDE

[DDE^{+}95] Roberto Danovaro, Norberto Della Croce, Anastasios Eleftheriou, Mauro Fabiano, Nadia Papadopoulou, Chris Smith, and Anastasios Tselepides. Meiofauna of the deep Eastern Mediterranean Sea: distribution and abundance in relation to bacterial biomass, organic matter composition and other environmental factors. Progress in Oceanography, 36(4):329– 341, ???? 1995. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/007966119600002X.

Du:2021:OSC

[DDJ^{+}21] Yan Du, Xiaolong Dong, Xingwei Jiang, Yuhong Zhang, Di Zhu, Qiwei Sun, Zhenzhan Wang, Xinhua Niu, Wen Chen, Cheng Zhu, Zhiyou Jing, Shilin Tang, Yineng Li, Ju Chen, Xiaoqing Chu, Chi Xu, Tianyu Wang, Yinghui He, Bing Han, Ying Zhang, Minyang Wang, Wei Wu, Yifan Xia, Kun Chen, Yu-Kun Qian, Ping Shi, Haigang Zhan, and Shiqiu Peng. Ocean surface current multiscale observation mission (OSCOM): Simultaneous measurement of ocean surface current, vector wind, and temperature. Progress in Oceanography, 193(??):??, April–May 2021. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661121000215.

Dessier:2018:VED

[DDK^{+}18] Aurélie Dessier, Christine Dupuy, Anaïs Kerric, Françoise Mornet, Matthieu Authier, Paco Bustamante, and Jérôme Spitz. Variability of energy density among mesozooplankton community: New insights in functional diversity to forage fish. Progress in Oceanography, 166(??):121–128, September 2018. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661116302105.

Danovaro:2000:BCP

[DDP^{+}00] Roberto Danovaro, Antonio Dell’Anno, Antonio Pusceddu, Daniela Marrale, Norberto Della Croce, Mauro Fabiano, and Anastasios Tselepides. Biochemical composition of pico-, nano- and micro-particulate organic matter and bacterioplankton biomass in the oligotrophic Cretan Sea (NE Mediterranean). Progress in Oceanography, 46(2–4):279–310, August 2000. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661100000239.

Deacon:1985:ETO

[Dea85] Margaret Deacon. An early theory of ocean circulation: J. S. Von Waitz and his explanation of the currents in the Strait of Gibraltar. Progress in Oceanography, 14(??):89–101, ???? 1985. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0079661185900072.

Demarcq:2009:TPP

[Dem09] Hervé Demarcq. Trends in primary production, sea surface temperature and wind in upwelling systems (1998–2007). Progress in Oceanography, 83(1–4):376–385, December 2009. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661109000779.

Denness:1987:SLM

[Den87] Bruce Denness. Sea-level modelling: the past and the future. Progress in Oceanography, 18(1–4):41–59, ???? 1987. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0079661187900255.

Denman:2003:MPE

[Den03] K. L. Denman. Modelling planktonic ecosystems: parameterizing complexity. Progress in Oceanography, 57(3–4):429– 452, June 2003. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661103001095.

Devoy:1987:SLA

[Dev87] R. J. N. Devoy. Sea-level applications and management. Progress in Oceanography, 18(1–4):273–286, ???? 1987. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/007966118790036X.

Dushaw:1997:TPG

[DEW^{+}97] Brian D. Dushaw, Gary D. Egbert, Peter F. Worcester, Bruce D. Cornuelle, Bruce M. Howe, and Kurt Metzger. A TOPEX/POSEIDON global tidal model (TPXO.2) and barotropic tidal currents determined from long-range acoustic transmissions. Progress in Oceanography, 40(??):337–367, ???? 1997. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661198000081.

Dong:2021:OMS

[DFC^{+}21] Wenjie Dong, Yanqing Feng, Changsheng Chen, Zhongxiang Wu, Danya Xu, Siqi Li, Qichun Xu, Lu Wang, Robert C. Beardsley, Huichan Lin, Ruixiang Li, Junkun Chen, and Jiahui Li. Observational and modeling studies of oceanic responses and feedbacks to Typhoons Hato and Mangkhut over the northern shelf of the South China Sea. Progress in Oceanography, 191(??):??, February 2021. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661120302421.

DeLauro:2023:NIV

[DFD23] E. De Lauro, M. Falanga, and S. De Martino. New insights on the Venice high-water phenomenon through a maximum negentropy criterium. Progress in Oceanography, 211(??):??, February 2023. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661123000216.

Druon:2016:HSA

[DFH^{+}16] Jean-Noël Druon, Jean-Marc Fromentin, Alex R. Hanke, Haritz Arrizabalaga, Dimitrios Damalas, Vjekoslav Tičina, Gemma Quílez-Badia, Karina Ramirez, Igor Arregui, George Tserpes, Patricia Reglero, Michele Deflorio, Isik Oray, F. Saadet Karakulak, Persefoni Megalofonou, Tevfik Ceyhan, Leon Grubišić, Brian R. MacKenzie, John Lamkin, Pedro Afonso, and Piero Addis. Habitat suitability of the Atlantic bluefin tuna by size class: an ecological niche approach. Progress in Oceanography, 142(??):30–46, March 2016. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661116000070.

deFreitas:2024:DOC

[dFHR^{+}24] Thaise Ricardo de Freitas, Silvia Hess, Paul E. Renaud, Peter Appleby, and Elisabeth Alve. Drivers of organic carbon distribution and accumulation in the northern Barents Sea. Progress in Oceanography, 225(??):??, July 2024. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661124000922.

delFavero:2017:SAE

[dFKdLZTT17] Jana M. del Favero, Mario Katsuragawa, Maria de Lourdes Zani-Teixeira, and Jefferson T. Turner. Spawning areas of Engraulis anchoita in the Southeastern Brazilian Bight during late-spring and early summer. Progress in Oceanography, 153(??):37–49, April 2017. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661116302294.

Druon:2015:MEH

[DFM^{+}15] Jean-Noël Druon, Fabio Fiorentino, Matteo Murenu, Leyla Knittweis, Francesco Colloca, Chato Osio, Bastien Mérigot, Germana Garofalo, Alessandro Mannini, Angélique Jadaud, Mario Sbrana, Giuseppe Scarcella, George Tserpes, Panagiota Peristeraki, Roberto Carlucci, and Jukka Heikkonen. Modelling of European hake nurseries in the Mediterranean Sea: an ecological niche approach. Progress in Oceanography, 130(??):188–204, January 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661114001803.

Deutsch:2021:BVC

[DFM^{+}21] Curtis Deutsch, Hartmut Frenzel, James C. McWilliams, Lionel Renault, Faycal Kessouri, Evan Howard, Jun-Hong Liang, Daniele Bianchi, and Simon Yang. Biogeochemical variability in the California Current System. Progress in Oceanography, 196(??):??, August 2021. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661121000525.

Dias:2002:SDP

[DGGdR02] J. M. A. Dias, R. Gonzalez, C. Garcia, and V. Diaz del Rio. Sediment distribution patterns on the Galicia–Minho continental shelf. Progress in Oceanography, 52(??):215–231, ???? 2002. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661102000071.

Du:2020:DDI

[DGH^{+}20] Chuanjun Du, Jianping Gan, Chiwing Rex Hui, Zhongming Lu, Xiaozheng Zhao, Elliott Roberts, and Minhan Dai. Dynamics of dissolved inorganic carbon in the South China Sea: a modeling study. Progress in Oceanography, 186(??):??, July 2020. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661120301063.

Dickson:1985:EMC

[DGMM85] R. R. Dickson, W. J. Gould, T. J. Müller, and C. Maillard. Estimates of the mean circulation in the deep (>2,000m) layer of the Eastern North Atlantic. Progress in Oceanography, 14(??):103–127, ???? 1985. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0079661185900084.

Davidson:2013:PBD

[DGP^{+}13] Keith Davidson, Linda C. Gilpin, Romain Pete, Debra Brennan, Sharon McNeill, Grigorios Moschonas, and Jonathon Sharples. Phytoplankton and bacterial distribution and productivity on and around Jones Bank in the Celtic Sea. Progress in Oceanography, 117(??):48–63, October 2013. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661113000402.

Diz:2024:DEH

[DGVGR24] Paula Diz, Rita González-Villanueva, and Emilio García-Roselló. Diversity and endemism of hard-shelled benthic foraminifera in permanently oxygen-depleted bottom waters: an analysis from the eastern Pacific. Progress in Oceanography, 225(??):??, July 2024. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661124000831.

deYoung:2004:DRS

[dHA^{+}04] B. deYoung, R. Harris, J. Alheit, G. Beaugrand, N. Mantua, and L. Shannon. Detecting regime shifts in the ocean: Data considerations. Progress in Oceanography, 60(2–4):143–164, February 2004. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661104000254.

Durden:2021:ACF

[DHB^{+}21] Jennifer M. Durden, Brett Hosking, Brian J. Bett, Danelle Cline, and Henry A. Ruhl. Automated classification of fauna in seabed photographs: the impact of training and validation dataset size, with considerations for the class imbalance. Progress in Oceanography, 196(??):??, August 2021. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661121000999.

Damerell:2020:CFS

[DHC^{+}20] Gillian M. Damerell, Karen J. Heywood, Daley Calvert, Alan L. M. Grant, Michael J. Bell, and Stephen E. Belcher. A comparison of five surface mixed layer models with a year of observations in the North Atlantic. Progress in Oceanography, 187(??):??, August–September 2020. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661120300550.

Diogou:2023:BWY

[DHD^{+}23] Nikoletta Diogou, William D. Halliday, Stan E. Dosso, Xavier Mouy, Andrea Niemi, and Stephen J. Insley. Bowhead whale year-round acoustic presence and habitat associations in the Amundsen Gulf, Western Canadian Arctic, 2018–2019. Progress in Oceanography, 213(??):??, April–May 2023. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661123000472.

Dittrich:2022:DOC

[DHDM22] Ribanna Dittrich, Sian F. Henley, Hugh W. Ducklow, and Michael P. Meredith. Dissolved organic carbon and nitrogen cycling along the west Antarctic Peninsula during summer. Progress in Oceanography, 206(??):??, August 2022. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661122001136.

Doray:2018:SHS

[DHHP18] Mathieu Doray, Camille Hervy, Martin Huret, and Pierre Petitgas. Spring habitats of small pelagic fish communities in the Bay of Biscay. Progress in Oceanography, 166(??):88–108, September 2018. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661117300927.

Du:2021:CND

[DHL^{+}21] Chuanjun Du, Ruoying He, Zhiyu Liu, Tao Huang, Lifang Wang, Zhongwei Yuan, Yanping Xu, Zhe Wang, and Minhan Dai. Climatology of nutrient distributions in the South China Sea based on a large data set derived from a new algorithm. Progress in Oceanography, 195(??):??, July 2021. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661121000732.

Dong:2009:CMS

[DIM09] Changming Dong, Eileen Y. Idica, and James C. McWilliams. Circulation and multiple-scale variability in the Southern California Bight. Progress in Oceanography, 82(3):168–190, September 2009. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661109000573.

Duguid:2021:FSS

[DIQJ21] William D. P. Duguid, Thomas W. Iwanicki, Jessica Qualley, and Francis Juanes. Fine-scale spatiotemporal variation in juvenile Chinook salmon distribution, diet and growth in an oceanographically heterogeneous region. Progress in Oceanography, 193(??):??, April–May 2021. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661121000021.

Davies:1992:TDM

[DJ92] A. M. Davies and J. E. Jones. A three dimensional model of the M_{2}, S_{2}, N_{2}, K_{1} and O_{1} tides in the Celtic and Irish seas. Progress in Oceanography, 29(3):197–234, ???? 1992. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/007966119290004J.

Dias:2002:PDS

[DJG^{+}02] J. M. A. Dias, J. M. Jouanneau, R. Gonzalez, M. F. Araújo, T. Drago, C. Garcia, A. Oliveira, A. Rodrigues, J. Vitorino, and O. Weber. Present day sedimentary processes on the northern Iberian shelf. Progress in Oceanography, 52(??):249– 259, ???? 2002. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661102000095.

deJonge:2020:AFW

[dJSL^{+}20] Daniëlle S. W. de Jonge, Tanja Stratmann, Lidia Lins, Ann Vanreusel, Autun Purser, Yann Marcon, Clara F. Rodrigues, Ascensão Ravara, Patricia Esquete, Marina R. Cunha, Erik Simon-Lledó, Peter van Breugel, Andrew K. Sweetman, Karline Soetaert, and Dick van Oevelen. Abyssal food-web model indicates faunal carbon flow recovery and impaired microbial loop 26 years after a sediment disturbance experiment. Progress in Oceanography, 189(??):??, November–December 2020. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661120301841.

Daly:2018:BTW

[DJW^{+}18] Eoghan Daly, Mark P. Johnson, Annette M. Wilson, Hans D. Gerritsen, Konstadinos Kiriakoulakis, A. Louise Allcock, and Martin White. Bottom trawling at whittard canyon: Evidence for seabed modification, trawl plumes and food source heterogeneity. Progress in Oceanography, 169(??):227–240, December 2018. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661117302549.

Durbin:2007:SSD

[DK07] Edward Durbin and Joseph Kane. Seasonal and spatial dynamics of Centropages typicus and C. hamatus in the western North Atlantic. Progress in Oceanography, 72(2–3):249–258, February 2007. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661107000079.

Day:2022:TEL

[DKRL22] Louise Day, Dorothée Kopp, Marianne Robert, and Hervé Le Bris. Trophic ecology of large gadiforms in the food web of a continental shelf ecosystem. Progress in Oceanography, 175(??):105–114, ???? 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661118302428.

Dickson:1969:AMC

[DL69] Robert Dickson and Arthur Lee. Atmospheric and marine climate fluctuations in the North Atlantic region. Progress in Oceanography, 5(??):55–65, ???? 1969. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0079661169900275.

Du:2017:PDS

[DL17] Xiaoqin Du and James T. Liu. Particle dynamics of the surface, intermediate, and benthic nepheloid layers under contrasting conditions of summer monsoon and typhoon winds on the boundary between the Taiwan Strait and East China Sea. Progress in Oceanography, 156(??):130–144, August 2017. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661117302185.

Dore:2008:SPB

[DLC^{+}08] John E. Dore, Ricardo M. Letelier, Matthew J. Church, Roger Lukas, and David M. Karl. Summer phytoplankton blooms in the oligotrophic North Pacific subtropical gyre: Historical perspective and recent observations. Progress in Oceanography, 76(1):2–38, January 2008. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661107001954.

Divoky:2015:ERD

[DLD15] George J. Divoky, Paul M. Lukacs, and Matthew L. Druckenmiller. Effects of recent decreases in Arctic sea ice on an ice-associated marine bird. Progress in Oceanography, 136(??):151–161, August 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S007966111500107X.

Dezutter:2019:MBM

[DLD^{+}19] Thibaud Dezutter, Catherine Lalande, Christiane Dufresne, Gérald Darnis, and Louis Fortier. Mismatch between microalgae and herbivorous copepods due to the record sea ice minimum extent of 2012 and the late sea ice break-up of 2013 in the Beaufort Sea. Progress in Oceanography, 173(??):66– 77, April 2019. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661118301071.

delaGuardia:2023:ANP

[dlGFM^{+}23] Laura Castro de la Guardia, Tania Hernández Fariñas, Christian Marchese, Martí Amargant-Arumí, Paul G. Myers, Simon Bélanger, Philipp Assmy, Rolf Gradinger, and Pedro Duarte. Assessing net primary production in the northwestern Barents Sea using in situ, remote sensing and modelling approaches. Progress in Oceanography, 219(??):??, December 2023. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661123002033.

delaHoz:2018:OEO

[dlHRA^{+}18] Camino F. de la Hoz, Elvira Ramos, Adrián Acevedo, Araceli Puente, Íñigo J. Losada, and José A. Juanes. OCLE: a European open access database on climate change effects on littoral and oceanic ecosystems. Progress in Oceanography, 168(??):222–231, November 2018. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661118300570.

Dumelle:2021:CCD

[DLJ^{+}21] Michael Dumelle, Jesse F. Lamb, Kym C. Jacobson, Mary Hunsicker, Cheryl A. Morgan, Brian J. Burke, and William T. Peterson. Capturing copepod dynamics in the Northern California Current using sentinel stations. Progress in Oceanography, 193(??):??, April–May 2021. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661121000379.

DeCubber:2023:UMB

[DLL^{+}23] Lola De Cubber, Sébastien Lefebvre, Théo Lancelot, Daniel Schaffer Ferreira Jorge, and Sylvie Marylène Gaudron. Unravelling mechanisms behind population dynamics, biological traits and latitudinal distribution in two benthic ecosystem engineers: a modelling approach. Progress in Oceanography, 219(??):??, December 2023. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661123001970.

deAmorim:2023:IMT

[dLLdAWL^{+}23] Felipe de Luca Lopes de Amorim, Karen Helen Wiltshire, Peter Lemke, Kristine Carstens, Silvia Peters, Johannes Rick, Luis Gimenez, and Mirco Scharfe. Investigation of marine temperature changes across temporal and spatial gradients: Providing a fundament for studies on the effects of warming on marine ecosystem function and biodiversity. Progress in Oceanography, 216(??):??, August 2023. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661123001234.

Dunbar:1991:ETA

[DLM91] D. Dunbar, P. LeBlond, and T. S. Murty. Evaluation of tsunami amplitudes for the Pacific coast of Canada. Progress in Oceanography, 26(2):115–177, ???? 1991. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0079661191900013.

Dickson:1996:LTC

[DLM^{+}96] Robert Dickson, John Lazier, Jens Meincke, Peter Rhines, and James Swift. Long-term coordinated changes in the convective activity of the North Atlantic. Progress in Oceanography, 38(3):241–295, ???? 1996. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661197000025.

Daneri:2012:WFS

[DLM^{+}12] Giovanni Daneri, Lorena Lizárraga, Paulina Montero, Humberto E. González, and Fabián J. Tapia. Wind forcing and short-term variability of phytoplankton and heterotrophic bacterioplankton in the coastal zone of the Concepción upwelling system (Central Chile). Progress in Oceanography, 92–95(??):92–96, January 2012. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S007966111100067X.

delaPaz:2015:NOM

[dlPHF^{+}15] M. de la Paz, I. E. Huertas, S. Flecha, A. F. Ríos, and F. F. Pérez. Nitrous oxide and methane in Atlantic and Mediterranean waters in the Strait of Gibraltar: Air-sea fluxes and inter-basin exchange. Progress in Oceanography, 138 (part A)(??):18–31, November 2015. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661115001998.

delaPaz:2024:SVC

[dlPPÁB24] Mercedes de la Paz, Fiz F. Pérez, Marta Álvarez, and Antonio Bode. Seasonal ventilation controls nitrous oxide emission in the NW Iberian upwelling. Progress in Oceanography, 224(??): ??, June 2024. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661124000673.

Dalpadado:2013:CAF

[DM13] Padmini Dalpadado and Fran Mowbray. Comparative analysis of feeding ecology of capelin from two shelf ecosystems, off Newfoundland and in the Barents Sea. Progress in Oceanography, 114(??):97–105, July 2013. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661113000499.

Doyle:2002:RVS

[DMBB02] Miriam J. Doyle, Kathryn L. Mier, Morgan S. Busby, and Richard D. Brodeur. Regional variation in springtime ichthyoplankton assemblages in the northeast Pacific Ocean.

Progress in Oceanography, 53(??):247–281, ???? 2002. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661102000332.

Dragon:2010:LFB

[DMBHG10] Anne-Cecile Dragon, P. Monestiez, A. Bar-Hen, and C. Guinet. Linking foraging behaviour to physical oceanographic structures: Southern elephant seals and mesoscale eddies east of Kerguelen Islands. Progress in Oceanography, 87(1–4):61–71, October 2010. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661110001370.

Diaz:2018:EEF

[DMC^{+}18] Marina V. Diaz, Marina Marrari, Valeria Casa, Florencia Gattás, Marcelo Pájaro, and Gustavo J. Macchi. Evaluating environmental forcing on nutritional condition of Engraulis anchoita larvae in a productive area of the Southwestern Atlantic Ocean. Progress in Oceanography, 168(??):13–22, November 2018. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661117303701.

Danovaro:2000:BRS

[DMD^{+}00] Roberto Danovaro, Daniela Marrale, Antonio Dell’Anno, Norberto Della Croce, Anastasios Tselepides, and Mauro Fabiano. Bacterial response to seasonal changes in labile organic matter composition on the continental shelf and bathyal sediments of the Cretan Sea. Progress in Oceanography, 46(2–4):345–366, August 2000. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661100000252.

Drinkwater:2009:RCF

[DMF^{+}09] K. F. Drinkwater, F. Mueter, K. D. Friedland, M. Taylor, G. L. Hunt, J. Hare, and W. Melle. Recent climate forcing and physical oceanographic changes in northern hemisphere regions: a review and comparison of four marine ecosystems. Progress in Oceanography, 81(1–4):10–28, April 2009. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661109000202.

deMadron:2011:MER

[dMGS^{+}11a] X. Durrieu de Madron, C. Guieu, R. Sempéré, P. Conan, D. Cossa, F. D’Ortenzio, C. Estournel, F. Gazeau, C. Rabouille, L. Stemmann, S. Bonnet, F. Diaz, P. Koubbi, O. Radakovitch, M. Babin, M. Baklouti, C. Bancon-Montigny, S. Belviso, N. Bensoussan, B. Bonsang, I. Bouloubassi, C. Brunet, J.-F. Cadiou, F. Carlotti, M. Chami, S. Charmasson, B. Charrière, J. Dachs, D. Doxaran, J.-C. Dutay, F. Elbaz-Poulichet, M. Eléaume, F. Eyrolles, C. Fernandez, S. Fowler, P. Francour, J. C. Gaertner, R. Galzin, S. Gasparini, J.-F. Ghiglione, J.-L. Gonzalez, C. Goyet, L. Guidi, K. Guizien, L.-E. Heimbürger, S. H. M. Jacquet, W. H. Jeffrey, F. Joux, P. Le Hir, K. Leblanc, D. Lefèvre, C. Lejeusne, R. Lemé, M.- D. Loÿe-Pilot, M. Mallet, L. Méjanelle, F. Mélin, C. Mellon, B. Mérigot, P.-L. Merle, C. Migon, W. L. Miller, L. Mortier, B. Mostajir, L. Mousseau, T. Moutin, J. Para, T. Pérez, A. Petrenko, J.-C. Poggiale, L. Prieur, M. Pujo-Pay, Pulido-Villena, P. Raimbault, A. P. Rees, C. Ridame, J.-F. Rontani, D. Ruiz Pino, M. A. Sicre, V. Taillandier, C. Tamburini, T. Tanaka, I. Taupier-Letage, M. Tedetti, P. Testor, H. Thébault, B. Thouvenin, F. Touratier, J. Tronczynski, C. Ulses, F. Van Wambeke, V. Vantrepotte, S. Vaz, and R. Verney. Marine ecosystems’ responses to climatic and anthropogenic forcings in the Mediterranean. Progress in Oceanography, 91(2):97–166, October 2011. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661111000188.

deMadron:2011:CME

[dMGS^{+}11b] X. Durrieu de Madron, C. Guieu, R. Sempéré, P. Conan, D. Cossa, F. D’Ortenzio, C. Estournel, F. Gazeau, C. Rabouille, L. Stemmann, S. Bonnet, F. Diaz, P. Koubbi, O. Radakovitch, M. Babin, M. Baklouti, C. Bancon-Montigny, S. Belviso, N. Bensoussan, B. Bonsang, I. Bouloubassi, C. Brunet, J.-F. Cadiou, F. Carlotti, M. Chami, S. Charmasson, B. Charrière, J. Dachs, D. Doxaran, J.-C. Dutay, F. Elbaz-Poulichet, M. Eléaume, F. Eyrolles, C. Fernandez, S. Fowler, P. Francour, J. C. Gaertner, R. Galzin, S. Gasparini, J.-F. Ghiglione, J.-L. Gonzalez, C. Goyet, L. Guidi, K. Guizien, L.-E. Heimbürger, S. H. M. Jacquet, W. H. Jeffrey, F. Joux, P. Le Hir, K. Leblanc, D. Lefèvre, C. Lejeusne, R. Lemé, M.- D. Loÿe-Pilot, M. Mallet, L. Méjanelle, F. Mélin, C. Mellon, B. Mérigot, P.-L. Merle, C. Migon, W. L. Miller, L. Mortier, B. Mostajir, L. Mousseau, T. Moutin, J. Para, T. Pérez, A. Petrenko, J.-C. Poggiale, L. Prieur, M. Pujo-Pay, Pulido-Villena, P. Raimbault, A. P. Rees, C. Ridame, J.-F. Rontani, D. Ruiz Pino, M. A. Sicre, V. Taillandier, C. Tamburini, T. Tanaka, I. Taupier-Letage, M. Tedetti, P. Testor, H. Thébault, B. Thouvenin, F. Touratier, J. Tronczynski, C. Ulses, F. Van Wambeke, V. Vantrepotte, S. Vaz, and R. Verney. Corrigendum to “Marine ecosystems’ responses to climatic and anthropogenic forcings in the Mediterranean” [Prog. Oceanogr. 91 (2011) 97– 166]. Progress in Oceanography, 91(4):593–594, December 2011. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S007966111100111X.

Daniault:2016:NNA

[DML^{+}16] Nathalie Daniault, Herlé Mercier, Pascale Lherminier, Artem Sarafanov, Anastasia Falina, Patricia Zunino, Fiz F. Pérez, Aida F. Ríos, Bruno Ferron, Thierry Huck, Virginie Thierry, and Sergey Gladyshev. The northern North Atlantic Ocean mean circulation in the early 21st century. Progress in Oceanography, 146(??):142–158, August 2016. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661116300659.

deMaio:1969:SIO

[dMM69] A. de Maio and M. Moretti. On a statistical index of ocean variability. Progress in Oceanography, 5(??):71–79, ???? 1969. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0079661169900299.

Dickson:1988:GSA

[DMML88] Robert R. Dickson, Jens Meincke, Svend-Aage Malmberg, and Arthur J. Lee. The “Great Salinity Anomaly” in the Northern North Atlantic 1968–1982. Progress in Oceanography, 20(2):103–151, ???? 1988. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0079661188900493.

Desbruyeres:2015:MBD

[DMT15] Damien Desbruyères, Herlé Mercier, and Virginie Thierry. On the mechanisms behind decadal heat content changes in the eastern subpolar gyre. Progress in Oceanography, 132(??):262– 272, March 2015. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661114000184.

Dandonneau:2007:APP

[DN07] Yves Dandonneau and Awa Niang. Assemblages of phytoplankton pigments along a shipping line through the North Atlantic and tropical Pacific. Progress in Oceanography, 73(2):127– 144, May 2007. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661107000420.

Doan-Nhu:2016:EAI

[DNNNN16] Hai Doan-Nhu, Lam Nguyen-Ngoc, and Chi-Thoi Nguyen. ENSO and anthropogenic impacts on phytoplankton diversity in tropical coastal waters. Progress in Oceanography, 140(??):1–13, January 2016. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S007966111500213X.

Dolgov:2009:TSB

[Dol09] Andrey V. Dolgov. Trophic structure of the Barents Sea fish assemblage with special reference to the cod stock recoverability. Progress in Oceanography, 81(1–4):165–173, April 2009. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661109000287.

Domanski:1984:DMD

[Dom84] Peter Domanski. The diel migrations and distributions within a mesopelagic community in the North East Atlantic. 8. a multivariate analysis of community structure. Progress in Oceanography, 13(3–4):491–511, ???? 1984. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/007966118490017X.

Donahue:1965:DIP

[Don65] Jessie G. Donahue. Diatoms as indicators of Pleistocene climatic fluctuations in the Pacific sector of the Southern Ocean. Progress in Oceanography, 4(??):133–134, IN9, 135–138, IN11–IN14, 139–140, September 2, 1965. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0079661165900467.

Donguy:1987:RAK

[Don87] Jean René Donguy. Recent advances in the knowledge of the climatic variations in the tropical Pacific Ocean. Progress in Oceanography, 19(1):49–85, ???? 1987. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0079661187900036.

Donguy:1994:SSS

[Don94] Jean-René Donguy. Surface and subsurface salinity in the tropical Pacific Ocean relations with climate. Progress in Oceanography, 34(1):45–78, ???? 1994. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0079661194900264.

Dieckmann:1987:IWL

[DOP87] R. Dieckmann, M. Osterthun, and H.-W. Partenscky. Influence of water-level elevation and tidal range on the sedimentation in a German tidal flat area. Progress in Oceanography, 18(1–4):151–166, ???? 1987. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0079661187900310.

DeLeo:2018:HFO

[DOS^{+}18] Fabio Cabrera De Leo, Bruno Ogata, Akash R. Sastri, Martin Heesemann, Steven Mihály, Moira Galbraith, and Michael G. Morley. High-frequency observations from a deep-sea cabled observatory reveal seasonal overwintering of Neocalanus spp. in Barkley Canyon, NE Pacific: Insights into particulate organic carbon flux. Progress in Oceanography, 169(??):120–137, December 2018. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661118300132.

Drinkwater:2013:CCF

[DP13] Ken Drinkwater and Pierre Pepin. Comparison of climate forcing on the marine ecosystems of the Northeast and Northwest Atlantic: a synthesis of the NORCAN Project. Progress in Oceanography, 114(??):3–10, July 2013. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S007966111300044X.

DeLeo:2018:BGB

[DP18] Fabio C. De Leo and Pere Puig. Bridging the gap between the shallow and deep oceans: the key role of submarine canyons. Progress in Oceanography, 169(??):1–5, December 2018. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661118302520.

dePuelles:2007:ZTS

[dPAJ07] Maria Luz Fernández de Puelles, Francisco Alemany, and Javier Jansá. Zooplankton time-series in the Balearic Sea (Western Mediterranean): Variability during the decade 1994– 2003. Progress in Oceanography, 74(2–3):329–354, August 2007. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661107000705.

Darby:2006:PGI

[DPB06] Dennis A. Darby, Leonid Polyak, and Henning A. Bauch. Past glacial and interglacial conditions in the Arctic Ocean and marginal seas — a review. Progress in Oceanography, 71(2–4):129–144, October 2006. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661106001200.

Diamante:1987:GCM

[DPCS87] John M. Diamante, Thomas E. Pyle, William E. Carter, and Wolfgang Scherer. Global change and the measurement of absolute sea-level. Progress in Oceanography, 18(1–4):1–21, ???? 1987. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661187900231.

duPontavice:2023:HRO

[dPCS23] Hubert du Pontavice, Zhuomin Chen, and Vincent S. Saba. A high-resolution ocean bottom temperature product for the northeast U.S. continental shelf marine ecosystem. Progress in Oceanography, 210(??):??, January 2023. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661122002075.

Duran:2020:SAC

[DPF^{+}20] Earl R. Duran, Helen E. Phillips, Ryo Furue, Paul Spence, and Nathaniel L. Bindoff. Southern Australia Current System based on a gridded hydrography and a high-resolution model. Progress in Oceanography, 181(??):??, February 2020. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661119304343.

Dias:2014:MSS

[DPGC14] Daniela Faggiani Dias, Luciano Ponzi Pezzi, Douglas Francisco Marcolino Gherardi, and Ricardo Camargo. Modeling the spawning strategies and larval survival of the Brazilian sardine (Sardinella brasiliensis). Progress in Oceanography, 123(??): 38–53, April 2014. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661114000354.

dePuelles:2023:DVM

[dPGSHL23] Maria Luz Fernandez de Puelles, Magdalena Gazá, Mar Santandreu, and Santiago Hernández-León. Diel vertical migration of copepods in the tropical and subtropical Atlantic Ocean. Progress in Oceanography, 219(??):??, December 2023. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661123001908.

Doray:2018:MSP

[DPH^{+}18] Mathieu Doray, Pierre Petitgas, Martin Huret, Erwan Duhamel, Jean Baptiste Romagnan, Matthieu Authier, Christine Dupuy, and Jérome Spitz. Monitoring small pelagic fish in the Bay of Biscay ecosystem, using indicators from an integrated survey. Progress in Oceanography, 166(??):168–188, September 2018. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661117300915.

Doyle:2009:LFA

[DPM^{+}09] Miriam J. Doyle, Susan J. Picquelle, Kathryn L. Mier, Michael C. Spillane, and Nicholas A. Bond. Larval fish abundance and physical forcing in the Gulf of Alaska, 1981– 2003. Progress in Oceanography, 80(3–4):163–187, March 2009. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661109000147.

Doray:2018:PSS

[DPR^{+}18] Mathieu Doray, Pierre Petitgas, Jean Baptiste Romagnan, Martin Huret, Erwan Duhamel, Christine Dupuy, Jérome Spitz, Matthieu Authier, Florence Sanchez, Laurent Berger, Ghislain Dorémus, Paul Bourriau, Patrick Grellier, and Jacques Massé. The PELGAS survey: Ship-based integrated monitoring of the Bay of Biscay pelagic ecosystem. Progress in Oceanography, 166(??):15–29, September 2018. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661117300939.

Dickson:2007:CEF

[DRD^{+}07] Robert Dickson, Bert Rudels, Stephen Dye, Michael Karcher, Jens Meincke, and Igor Yashayaev. Current estimates of freshwater flux through Arctic and subarctic seas. Progress in Oceanography, 73(3–4):210–230, May 2007. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S007966110700081X.

Dewitte:2008:VSV

[DRE^{+}08] Boris Dewitte, Marcel Ramos, Vincent Echevin, Oscar Pizarro, and Yves duPenhoat. Vertical structure variability in a seasonal simulation of a medium-resolution regional model of the Eastern South Pacific. Progress in Oceanography, 79(2–4):120–137, October 2008. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661108001675.

Drinkwater:2006:RSN

[Dri06] Kenneth F. Drinkwater. The regime shift of the 1920s and 1930s in the North Atlantic. Progress in Oceanography, 68 (2–4):134–151, February 2006. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661106000036.

Drinkwater:2011:ICV

[Dri11] Kenneth F. Drinkwater. The influence of climate variability and change on the ecosystems of the Barents Sea and adjacent waters: Review and synthesis of recent studies from the NESSAS Project. Progress in Oceanography, 90(1–4):47– 61, July 2011. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661111000218.

Diaz:2022:FSP

[DRVMC^{+}22] Patricio A. Díaz, Manuel Ruiz-Villarreal, Beatriz Mouriño-Carballido, Concepción Fernández-Pena, Pilar Riobó, and Beatriz Reguera. Fine scale physical-biological interactions during a shift from relaxation to upwelling with a focus on Dinophysis acuminata and its potential ciliate prey. Progress in Oceanography, 175(??):309–327, ???? 2022. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661117303087.

Donn:1965:MIF

[DS65] William L. Donn and David M. Shaw. The maintenance of an ice-free Arctic Ocean. Progress in Oceanography, 4 (??):105–113, September 2, 1965. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0079661165900443.

Dreutter:2020:WTF

[DSAB20] Simon Dreutter, Melanie Steffen, Pedro Martínez Arbizu, and Angelika Brandt. Will the “top five” deepest trenches lose one of their members? Progress in Oceanography, 181(??):??, February 2020. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661119304380.

Demopoulos:2024:CIR

[DSB^{+}24] Amanda W. J. Demopoulos, Brian J. Smith, Jill R. Bourque, Jason D. Chaytor, Jennifer McClain-Counts, Nancy Prouty, Steve W. Ross, Sandra Brooke, Gerard Duineveld, and Furu Mienis. Consumer isoscapes reveal heterogeneous food webs in deep-sea submarine canyons and adjacent slopes. Progress in Oceanography, 223(??):??, April–May 2024. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661124000375.

Dorman:2015:IBM

[DSBP15] Jeffrey G. Dorman, William J. Sydeman, Steven J. Bograd, and Thomas M. Powell. An individual-based model of the krill Euphausia pacifica in the California Current. Progress in Oceanography, 138(??):504–520, November 2015. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661115000439.

DiBella:2019:LFA

[DSC^{+}19] L. Di Bella, A. Sabbatini, L. Carugati, M. Lo Martire, G. M. Luna, M. Pierdomenico, R. Danovaro, and A. Negri. Living foraminiferal assemblages in two submarine canyons (Polcevera and Bisagno) of the Ligurian basin (Mediterranean Sea). Progress in Oceanography, 173(??):114–133, April 2019. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661118300867.

Dossa:2021:NSW

[DSC^{+}21] Alina N. Dossa, Alex C. Silva, Alexis Chaigneau, Gérard Eldin, Moacyr Araujo, and Arnaud Bertrand. Nearsurface western boundary circulation off Northeast Brazil. Progress in Oceanography, 190(??):??, January 2021. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S007966112030210X.

daSilveira:2023:BCQ

[dSPF^{+}23] Ilson C. A. da Silveira, Filipe Pereira, Glenn R. Flierl, Iury T. Simoes-Sousa, André Palóczy, Milton Borges-Silva, and César B. Rocha. The Brazil Current quasi-stationary unstable meanders at 22^{o}S–23^{o}S. Progress in Oceanography, 210(??):??, January 2023. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661122001847.

Dietrich:2021:WSB

[DSR21] Kimberly S. Dietrich, Jarrod A. Santora, and Christian S. Reiss. Winter and summer biogeography of macrozooplankton community structure in the northern Antarctic Peninsula ecosystem. Progress in Oceanography, 196(??):??, August 2021. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661121000975.

Schmidt:2020:ESM

[dSSDS^{+}20] Thassya C. dos Santos Schmidt, Jennifer A. Devine, Aril Slotte, Marion Claireaux, Arne Johannessen, Katja Enberg, Gudmundur J. Óskarsson, James Kennedy, Yutaka Kurita, and Olav Sigurd Kjesbu. Environmental stressors may cause unpredicted, notably lagged life-history responses in adults of the planktivorous Atlantic herring. Progress in Oceanography, 181(??):??, February 2020. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661119304379.

Dando:1999:HMS

[DSV99] P. R. Dando, D. Stüben, and S. P. Varnavas. Hydrothermalism in the Mediterranean Sea. Progress in Oceanography, 44(1–3):333–367, August 1999. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661199000324.

Duncan:2024:STD

[DSV^{+}24] Rebecca J. Duncan, Janne E. Søreide, Øystein Varpe, Józef Wiktor, Vanessa Pitusi, Elaine Runge, and Katherina Petrou.

Spatio-temporal dynamics in microalgal communities in Arctic land-fast sea ice. Progress in Oceanography, 224(??):??, June 2024. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661124000545.

DeMaster:2006:SMC

[DTC^{+}06] Douglas P. DeMaster, Andrew W. Trites, Phillip Clapham, Sally Mizroch, Paul Wade, Robert J. Small, and Jay Ver Hoef. The sequential megafaunal collapse hypothesis: Testing with existing data. Progress in Oceanography, 68(2–4):329–342, February 2006. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661106000127.

Du:2024:DPG

[DTG^{+}24] Jiabi Du, Carolyn K. Tepolt, Emily W. Grason, P. Sean Mc-Donald, Yan Jia, and Weifeng G. Zhang. Dispersal pathways of European green crab larvae into and throughout the eastern Salish Sea. Progress in Oceanography, 223(??):??, April–May 2024. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S007966112400051X.

Dutay:2015:MRM

[DTKvH15] J.-C. Dutay, A. Tagliabue, I. Kriest, and M. M. P. van Hulten. Modelling the role of marine particle on large scale 231Pa, ^{230}Th, iron and aluminium distributions. Progress in Oceanography, 133(??):66–72, April 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S007966111500021X.

Danovaro:2000:DMA

[DTOD00] Roberto Danovaro, Anastasios Tselepides, Ana Otegui, and Norberto Della Croce. Dynamics of meiofaunal assemblages on the continental shelf and deep-sea sediments of the Cretan Sea (NE Mediterranean): relationships with seasonal changes in food supply. Progress in Oceanography, 46(2–4):367–400, August 2000. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661100000264.

Duineveld:2000:BPC

[DTW^{+}00] G. C. A. Duineveld, A. Tselepides, R. Witbaard, R. P. M. Bak, E. M. Berghuis, G. Nieuwland, J. van der Weele, and A. Kok. Benthic–pelagic coupling in the oligotrophic Cretan Sea. Progress in Oceanography, 46(2–4):457–481, August 2000. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S007966110000029X.

Duedall:1977:SED

[Due77] Iver W. Duedall. Seawater: an explanation of differential isothermal compressibility measurements in terms of hydration and ion-water interactions. Progress in Oceanography, 7(3):91– 133, ???? 1977. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661177900015.

Durazo:2009:CUO

[Dur09] Reginaldo Durazo. Climate and upper ocean variability off Baja California, Mexico: 1997–2008. Progress in Oceanography, 83(1–4):361–368, December 2009. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661109000974.

Djaoudi:2018:AFS

[DVB^{+}18] K. Djaoudi, F. Van Wambeke, A. Barani, S. Hélias-Nunige, R. Sempéré, and E. Pulido-Villena. Atmospheric fluxes of soluble organic C, N, and P to the Mediterranean Sea: Potential biogeochemical implications in the surface layer. Progress in Oceanography, 163(??):59–69, April 2018. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661117300587.

DeLange:1999:PEV

[DVL^{+}99] Gert J. De Lange, P. J. M. Van Santvoort, C. Langereis, J. Thomson, C. Corselli, A. Michard, M. Rossignol-Strick, M. Paterne, and G. Anastasakis. Palaeo-environmental variations in eastern Mediterranean sediments: a multidisciplinary approach in a prehistoric setting. Progress in Oceanography, 44(1–3):369–386, August 1999. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661199000373.

Dever:2002:EDS

[DW02] E. P. Dever and C. D. Winant. The evolution and depth structure of shelf and slope temperatures and velocities during the 1997–1998 El Niño near Point Conception, California. Progress in Oceanography, 54(1–4):77–103, July 2002. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661102000447.

Dunton:2006:NWB

[DWC06] Kenneth H. Dunton, Thomas Weingartner, and Eddy C. Carmack. The nearshore western Beaufort Sea ecosystem: Circulation and importance of terrestrial carbon in Arctic coastal food webs. Progress in Oceanography, 71(2–4):362–378, October 2006. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661106001248.

deWilde:1998:LSM

[dWDB^{+}98] P. A. W. J. de Wilde, G. C. A. Duineveld, E. M. Berghuis, M. S. S. Lavaleye, and A. Kok. Late-summer mass deposition of gelatinous phytodetritus along the slope of the N.W. European Continental Margin. Progress in Oceanography, 42(1–4):165–187, December 1998. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661198000330.

Darnis:2019:COP

[DWFP^{+}19] Gérald Darnis, Anette Wold, Stig Falk-Petersen, Martin Graeve, and Louis Fortier. Could offspring predation offset the successful reproduction of the Arctic copepod Calanus hyperboreus under reduced sea-ice cover conditions? Progress in Oceanography, 170(??):107–118, January 2019. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661117303415.

Danielson:2014:CWF

[DWH^{+}14] Seth L. Danielson, Thomas J. Weingartner, Katherine S. Hedstrom, Knut Aagaard, Rebecca Woodgate, Enrique Curchitser, and Phyllis J. Stabeno. Coupled wind-forced controls of the Bering–Chukchi shelf circulation and the Bering Strait throughflow: Ekman transport, continental shelf waves, and variations of the Pacific–Arctic sea surface height gradient. Progress in Oceanography, 125(??):40–61, June 2014. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661114000548.

Deguchi:2004:IVO

[DWNN04] Tomohiro Deguchi, Yutaka Watanuki, Yasuaki Niizuma, and Akifumi Nakata. Interannual variations of the occurrence of epipelagic fish in the diets of the seabirds breeding on Teuri Island, northern Hokkaido, Japan. Progress in Oceanography, 61(2–4):267–275, May 2004. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661104000667.

Daudt:2024:SAL

[DWS^{+}24] Nicholas W. Daudt, Eric J. Woehler, Matthew R. Schofield, Robert O. Smith, Leandro Bugoni, and William J. Rayment. Seabird assemblages are linked to the major western boundary current off eastern Australia. Progress in Oceanography, 223(??):??, April–May 2024. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661124000211.

Davies:2002:MNB

[DXH^{+}02] Alan M. Davies, Jiuxing Xing, John M. Huthnance, Philip Hall, and Laurenz Thomsen. Models of near-bed dynamics and sediment movement at the Iberian margin. Progress in Oceanography, 52(??):373–397, ???? 2002. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661102000162.

Dou:2015:DRC

[DYL^{+}15] Yanguang Dou, Shouye Yang, Chao Li, Xuefa Shi, Jihua Liu, and Lei Bi. Deepwater redox changes in the southern Okinawa Trougho since the last glacial maximum. Progress in Oceanography, 135(??):77–90, June 2015. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661115000671.

Dambacher:2010:APF

[DYO^{+}10] Jeffrey M. Dambacher, Jock W. Young, Robert J. Olson, Valérie Allain, Felipe Galván-Magaña, Matthew J. Lansdell, Noemí Bocanegra-Castillo, Vanessa Alatorre-Ramírez, Scott P. Cooper, and Leanne M. Duffy. Analyzing pelagic food webs leading to top predators in the Pacific Ocean: a graph-theoretic approach. Progress in Oceanography, 86(1–2):152–165, July 2010. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661110000455.

Dolganova:2004:SIA

[DZ04] Natalia T. Dolganova and Yury I. Zuenko. Seasonal and inter-annual dynamics of mesoplankton in the northwestern Japan Sea. Progress in Oceanography, 61(2–4):227–243, May 2004. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661104000643.

deMadron:2005:CCD

[dZTG05] X. Durrieu de Madron, V. Zervakis, A. Theocharis, and D. Georgopoulos. Comments on “Cascades of dense water around the World Ocean”. Progress in Oceanography, 64(1):83– 90, January 2005. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661105000200. See [ISH^{+}04].

Elliott:2023:MDR

[EAB^{+}23] Sophie A. M. Elliott, Anthony Acou, Laurent Beaulaton, Jérôme Guitton, Elodie Réveillac, and Etienne Rivot. Modelling the distribution of rare and data-poor diadromous fish at sea for protected area management. Progress in Oceanography, 210(??):??, January 2023. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661122001835.

Echevarria:2007:EFI

[EAL^{+}07] Fidel Echevarría, Jose Luis Acuña, Santiago Hernández León, Enric Saiz, A. Miguel P. Santos, and Manuel Barangé. Ecological functioning of the Iberian seas: a synthesis of GLOBEC research in Spain and Portugal. Progress in Oceanography, 74(2–3):95–97, August 2007. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661107000584.

Echevin:2008:SCS

[EALF08] V. Echevin, O. Aumont, J. Ledesma, and G. Flores. The seasonal cycle of surface chlorophyll in the Peruvian upwelling system: a modelling study. Progress in Oceanography, 79(2–4):167–176, October 2008. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661108001705.

Espinoza:2008:RPA

[EB08] Pepe Espinoza and Arnaud Bertrand. Revisiting Peruvian anchovy (Engraulis ringens) trophodynamics provides a new vision of the Humboldt Current system. Progress in Oceanography, 79(2–4):215–227, October 2008. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661108001754.

Eriksen:2020:DTS

[EBD^{+}20] Elena Eriksen, Aleksander N. Benzik, Andrey V. Dolgov, Hein Rune Skjoldal, Mikko Vihtakari, Edda Johannesen, Tatiana A. Prokhorova, Felicia Keulder-Stenevik, Irina Prokopchuk, and Espen Strand. Diet and trophic structure of fishes in the Barents Sea: the Norwegian–Russian program “Year of stomachs” 2015 — establishing a baseline. Progress in Oceanography, 183(??):??, April 2020. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661119304422.

Eduardo:2020:HSS

[EBM^{+}20] Leandro Nolé Eduardo, Arnaud Bertrand, Michael Maia Mincarone, Lucas V. Santos, Thierry Frédou, Ramilla V. Assunção, Alex Silva, Frédéric Ménard, Ralf Schwamborn, François Le Loc’h, and Flávia Lucena-Frédou. Hatchetfishes (Stomiiformes: Sternoptychidae) biodiversity, trophic ecology, vertical niche partitioning and functional roles in the western Tropical Atlantic. Progress in Oceanography, 187(??):??, August– September 2020. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661120301282.

Eduardo:2021:DVM

[EBM^{+}21] Leandro Nolé Eduardo, Arnaud Bertrand, Michael Maia Mincarone, Júlia Rodrigues Martins, Thierry Frédou, Ramilla Vieira Assunção, Rayssa Siqueira Lima, Frédéric Ménard, François Le Loc’h, and Flávia Lucena-Frédou. Distribution, vertical migration, and trophic ecology of lanternfishes (Myctophidae) in the Southwestern Tropical Atlantic. Progress in Oceanography, 199(??):??, December 2021. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661121001798.

Everett:2014:RIS

[EBR^{+}14] Jason D. Everett, Mark E. Baird, Moninya Roughan, Iain M. Suthers, and Martina A. Doblin. Relative impact of seasonal and oceanographic drivers on surface chlorophyll-a along a western boundary current. Progress in Oceanography, 120(??):340–351, January 2014. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661113002103.

Espinasse:2018:CAZ

[EBS^{+}18] B. Espinasse, S. Basedow, S. Schultes, M. Zhou, L. Berline, and F. Carlotti. Conditions for assessing zooplankton abundance with LOPC in coastal waters. Progress in Oceanography, 163(??):260–270, April 2018. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661117300630.

Espinoza:2009:DSS

[EBvdL^{+}09] Pepe Espinoza, Arnaud Bertrand, Carl D. van der Lingen, Susana Garrido, and Blanca Rojas de Mendiola. Diet of sardine (Sardinops sagax) in the northern Humboldt Current system and comparison with the diets of clupeoids in this and other eastern boundary upwelling systems. Progress in Oceanography, 83(1–4):242–250, December 2009. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661109000883.

Engeland:2023:DSD

[EBW^{+}23] Tom Van Engeland, Espen Bagøien, Anette Wold, Heather A. Cannaby, Sanna Majaneva, Anna Vader, Jon Rønning, Nils Olav Handegard, Padmini Dalpadado, and Randi B. Ingvaldsen. Diversity and seasonal development of large zooplankton along physical gradients in the Arctic Barents Sea. Progress in Oceanography, 216(??):??, August 2023. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661123001088.

Elliott:2020:DLH

[ECFT20] Sophie A. M. Elliott, Alexandre Carpentier, Eric Feunteun, and Thomas Trancart. Distribution and life history trait models indicate vulnerability of skates. Progress in Oceanography, 181(??):??, February 2020. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661119304367.

Eardly:2001:BAA

[ECGP01] D. F. Eardly, M. W. Carton, J. M. Gallagher, and J. W. Patching. Bacterial abundance and activity in deepsea sediments from the eastern North Atlantic. Progress in Oceanography, 50(??):245–259, ???? 2001. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661101000568.

Emery:1982:MTS

[ED82] W. J. Emery and J. S. Dewar. Mean temperature-salinity, salinity-depth and temperature-depth curves for the North Atlantic and the North Pacific. Progress in Oceanography, 11(3):219–256, IN1–IN4, 257–305, ???? 1982. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0079661182900155.

Ericson:2023:RCR

[EFC^{+}23] Ylva Ericson, Agneta Fransson, Melissa Chierici, Elizabeth M. Jones, Ingunn Skjelvan, Abdirahman Omar, Are Olsen, and Meike Becker. Rapid f CO _{2} rise in the northern Barents Sea and Nansen Basin. Progress in Oceanography, 217(??):??, September–October 2023. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661123001222.

Egbert:1997:TDI

[Egb97] Gary D. Egbert. Tidal data inversion: interpolation and inference. Progress in Oceanography, 40(??):53–80, ???? 1997. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661197000232.

Eriksen:2018:SSS

[EGP^{+}18] E. Eriksen, H. Gjøsæter, D. Prozorkevich, E. Shamray, A. Dolgov, M. Skern-Mauritzen, J. E. Stiansen, Yu. Kovalev, and K. Sunnanå. From single species surveys towards monitoring of the Barents Sea ecosystem. Progress in Oceanography, 166(??):4–14, September 2018. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661117300642.

Evans:2015:PBU

[EGPM^{+}15] Claire Evans, Paola R. Gómez-Pereira, Adrian P. Martin, David J. Scanlan, and Mikhail V. Zubkov. Photoheterotrophy of bacterioplankton is ubiquitous in the surface oligotrophic ocean. Progress in Oceanography, 135(??):139–145, June 2015. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661115000749.

Escribano:2012:ZTS

[EHFD12] Ruben Escribano, Pamela Hidalgo, Marcelo Fuentes, and Katty Donoso. Zooplankton time series in the coastal zone off Chile: Variation in upwelling and responses of the copepod community. Progress in Oceanography, 97–100(??):174– 186, May 2012. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661111001248.

Escribano:2007:SIA

[EHG^{+}07] Ruben Escribano, Pamela Hidalgo, Humberto González, Ricardo Giesecke, Ramiro Riquelme-Bugueño, and Karen Manríquez. Seasonal and inter-annual variation of mesozooplankton in the coastal upwelling zone off central-southern Chile. Progress in Oceanography, 75(3):470–485, November 2007. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661107001632.

Estrada:2012:LSZ

[EHG^{+}12] Rafael Estrada, Michel Harvey, Michel Gosselin, Michel Starr, Peter S. Galbraith, and Fiammetta Straneo. Late-summer zooplankton community structure, abundance, and distribution in the Hudson Bay system (Canada) and their relationships with environmental conditions, 2003–2006. Progress in Oceanography, 101(1):121–145, August 2012. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661112000092.

Evans:2012:SAC

[EHSI12] Wiley Evans, Burke Hales, Peter G. Strutton, and Debby Ianson. Sea–air CO_{2} fluxes in the western Canadian coastal ocean. Progress in Oceanography, 101(1):78–91, August 2012. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661112000043.

Enomoto:2022:VHS

[EiIT^{+}22] Megumi Enomoto, Shin ichi Ito, Motomitsu Takahashi, Chiyuki Sassa, Tomihiko Higuchi, and Kotaro Shirai. Vertical habitat shifts of juvenile jack mackerel estimated using otolith oxygen stable isotope. Progress in Oceanography, 208(??):??, November 2022. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661122001562.

Emmett:2006:ADP

[EKB06] Robert L. Emmett, Gregory K. Krutzikowsky, and Paul Bentley. Abundance and distribution of pelagic piscivorous fishes in the Columbia River plume during spring/early summer 1998– 2003: Relationship to oceanographic conditions, forage fishes, and juvenile salmonids. Progress in Oceanography, 68(1):1–26, January 2006. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661105001412.

Enfield:2006:HLW

[ELW06] David B. Enfield, Sang-Ki Lee, and Chunzai Wang. How are large western hemisphere warm pools formed? Progress in Oceanography, 70(2–4):346–365, August 2006. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661106000486.

Escribano:2012:STS

[EM12] Ruben Escribano and Carmen E. Morales. Spatial and temporal scales of variability in the coastal upwelling and coastal transition zones off central-southern Chile (35–40^{o}S). Progress in Oceanography, 92–95(??):1–7, January 2012. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661111000735.

Ellis:2013:EAC

[EMBS13] J. R. Ellis, I. Martinez, G. J. Burt, and B. E. Scott. Epibenthic assemblages in the Celtic Sea and associated with the Jones Bank. Progress in Oceanography, 117(??):76–88, October 2013. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661113001006.

Emery:1965:SSD

[Eme65] K. O. Emery. Some stages in the development of knowledge of the Atlantic coast of the United States. Progress in Oceanography, 4(??):307–322, September 2, 1965. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0079661165900571.

Emiliani:1965:PRA

[Emi65] Cesare Emiliani. The Pleistocene record of the Atlantic and Pacific oceanic sediments; correlations with the Alaskan stages by absolute dating; and the age of the last reversal of the geomagnetic field. Progress in Oceanography, 4(??):219–224, September 2, 1965. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661165900510.

Evans:2017:SAI

[EMK^{+}17] G. R. Evans, E. L. McDonagh, B. A. King, H. L. Bryden, D. C. E. Bakker, P. J. Brown, U. Schuster, K. G. Speer, and S. M. A. C. van Heuven. South Atlantic interbasin exchanges of mass, heat, salt and anthropogenic carbon. Progress in Oceanography, 151(??):62–82, February 2017. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661115300380.

Estournel:2021:NAC

[EMU21] Claude Estournel, Patrick Marsaleix, and Caroline Ulses. A new assessment of the circulation of Atlantic and intermediate waters in the Eastern Mediterranean. Progress in Oceanography, 198(??):??, November 2021. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661121001579.

Estournel:2023:SDG

[EMU^{+}23] Claude Estournel, Guillaume Mikolajczak, Caroline Ulses, François Bourrin, Miquel Canals, Sabine Charmasson, David Doxaran, Thomas Duhaut, Xavier Durrieu de Madron, Patrick Marsaleix, Albert Palanques, Pere Puig, Olivier Radakovitch, Anna Sanchez-Vidal, and Romaric Verney. Sediment dynamics in the Gulf of Lion (NW Mediterranean Sea) during two autumn–winter periods with contrasting meteorological conditions. Progress in Oceanography, 210(??):??, January 2023. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661122002014.

Eppley:1992:CPN

[Epp92] Richard W. Eppley. Chlorophyll, photosynthesis and new production in the Southern California Bight. Progress in Oceanography, 30(1–4):117–150, ???? 1992. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/007966119290010W.

Elizondo:2021:BPG

[ERBV21] Urs Hofmann Elizondo, Damiano Righetti, Fabio Benedetti, and Meike Vogt. Biome partitioning of the global ocean based on phytoplankton biogeography. Progress in Oceanography, 194(??):??, June 2021. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661121000203.

Eriksson:1965:SDS

[Eri65] K. Gösta Eriksson. Some deep-sea sediments in the Western Mediterranean Sea. Progress in Oceanography, 4(??): 267–278, IN15–IN18, 279–280, September 2, 1965. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0079661165900546.

Eme:2022:CST

[ERT^{+}22] David Eme, Marta M. Rufino, Verena M. Trenkel, Youen Vermard, Pascal Laffargue, Pierre Petitgas, Loïc Pellissier, and Camille Albouy. Contrasted spatio-temporal changes in the demersal fish assemblages and the dominance of the environment vs fishing pressure, in the Bay of Biscay and Celtic Sea. Progress in Oceanography, 204(??):??, June 2022. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661122000490.

Escribano:2007:SFC

[ES07] Ruben Escribano and Wolfgang Schneider. The structure and functioning of the coastal upwelling system off central/southern Chile. Progress in Oceanography, 75(3):343–347, November 2007. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661107001553.

Embling:2013:FBR

[ESA^{+}13] C. B. Embling, J. Sharples, E. Armstrong, M. R. Palmer, and B. E. Scott. Fish behaviour in response to tidal variability and internal waves over a shelf sea bank. Progress in Oceanography, 117(??):106–117, October 2013. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661113001018.

Eriksen:2021:DTS

[ESD^{+}21] Elena Eriksen, Hein Rune Skjoldal, Andrey V. Dolgov, Espen Strand, Felicia Keulder-Stenevik, Irina P. Prokopchuk, Tatiana A. Prokhorova, Dmitry Prozorkevich, and Aleksander N. Benzik. Diet and trophic structure of fishes in the Barents Sea: Seasonal and spatial variations. Progress in Oceanography, 197(??):??, September–October 2021. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661121001488.

Eriksen:2017:STC

[ESGP17] Elena Eriksen, Hein Rune Skjoldal, Harald Gjøsæter, and Raul Primicerio. Spatial and temporal changes in the Barents Sea pelagic compartment during the recent warming. Progress in Oceanography, 151(??):206–226, February 2017. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661116300854.

El-Sabaawi:2012:IVB

[ESTM^{+}12] Rana W. El-Sabaawi, Marc Trudel, David L. Mackas, John F. Dower, and Asit Mazumder. Interannual variability in bottom-up processes in the upstream range of the California Current system: an isotopic approach. Progress in Oceanography, 106(??):16–27, November 2012. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661112000675.

El-Sabaawi:2013:ZSI

[ESTM13] Rana El-Sabaawi, Marc Trudel, and Asit Mazumder. Zooplankton stable isotopes as integrators of bottom-up variability in coastal margins: a case study from the Strait of Georgia and adjacent coastal regions. Progress in Oceanography, 115(??):76–89, August 2013. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661113000529.

Epping:2002:OBO

[EvdZSH02] Eric Epping, Claar van der Zee, Karline Soetaert, and Willem Helder. On the oxidation and burial of organic carbon in sediments of the Iberian margin and Nazaré Canyon (NE Atlantic). Progress in Oceanography, 52(??):399–431, ???? 2002. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661102000174.

Erbe:2015:MSP

[EVM^{+}15] Christine Erbe, Arti Verma, Robert McCauley, Alexander Gavrilov, and Iain Parnum. The marine soundscape of the Perth Canyon. Progress in Oceanography, 137 (part A)(??):38–51, September 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661115001123.

Farmer:1988:FAW

[FAAF88] David M. Farmer, Laurence Armi, Laurence Armi, and David M. Farmer. The flow of Atlantic water through the Strait of Gibraltar. Progress in Oceanography, 21(1):1, 3– 41, 43–47, 49, 51, 53, 55, 57–63, 65–67, 69, 71, 73, 75– 77, 79, 81–89, 91–93, 95, 97–101, 103, ???? 1988. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0079661188900559.

Flores-Aqueveque:2015:APM

[FAAV^{+}15] Valentina Flores-Aqueveque, Stéphane Alfaro, Gabriel Vargas, José A. Rutllant, and Sandrine Caquineau. Aeolian particles in marine cores as a tool for quantitative high-resolution reconstruction of upwelling favorable winds along coastal Atacama Desert, Northern Chile. Progress in Oceanography, 134(??):244–255, May 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661115000373.

Freon:2009:FGB

[FAB^{+}09] Pierre Fréon, Javier Arístegui, Arnaud Bertrand, Robert J. M. Crawford, John C. Field, Mark J. Gibbons, Jorge Tam, Larry Hutchings, Hicham Masski, Christian Mullon, Mohamed Ramdani, Bernard Seret, and Monique Simier. Functional group biodiversity in Eastern boundary upwelling ecosystems questions the wasp-waist trophic structure. Progress in Oceanography, 83(1–4):97–106, December 2009. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661109001025.

Fierro-Arcos:2023:AER

[FACM^{+}23] Denisse Fierro-Arcos, Stuart Corney, Amelie Meyer, Hakase Hayashida, Andrew E. Kiss, and Petra Heil. Analysis of ecologically relevant sea ice and ocean variables for the Southern Ocean using a high-resolution model to inform ecosystem studies. Progress in Oceanography, 215(??):??, July 2023. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661123000927.

Fernandez-Alamo:2006:ZOE

[FÁFL06] María Ana Fernández-Álamo and Jaime Färber-Lorda. Zooplankton and the oceanography of the eastern tropical Pacific: a review. Progress in Oceanography, 69(2–4):318–359, May 2006. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S007966110600036X.

Frigstad:2013:LTT

[FAH^{+}13] Helene Frigstad, Tom Andersen, Dag O. Hessen, Emil Jeansson, Morten Skogen, Lars-Johan Naustvoll, Martin W. Miles, Truls Johannessen, and Richard G. J. Bellerby. Long-term trends in carbon, nutrients and stoichiometry in Norwegian coastal waters: Evidence of a regime shift. Progress in Oceanography, 111(??):113–124, April 2013. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661113000165.

Fairbridge:1965:IOS

[Fai65] Rhodes W. Fairbridge. The Indian Ocean and the status of Gondwanaland. Progress in Oceanography, 3(??):83–136, ???? 1965. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/007966116590011X.

Fernandez-Arcaya:2013:RBR

[FARRL^{+}13] U. Fernandez-Arcaya, G. Rotllant, E. Ramirez-Llodra, L. Recasens, J. Aguzzi, M. M. Flexas, A. Sanchez-Vidal, P. López-Fernández, J. A. García, and J. B. Company. Reproductive biology and recruitment of the deep-sea fish community from the NW Mediterranean continental margin. Progress in Oceanography, 118(??):222–234, November 2013. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661113001328.

Fileman:2001:HIM

[FB01] Elaine Fileman and Peter Burkill. The herbivorous impact of microzooplankton during two short-term Lagrangian experiments off the NW coast of Galicia in summer 1998. Progress in Oceanography, 51(2–4):361–383, November 2001. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661101000751.

Fulmer:2005:RCS

[FB05] Julia H. Fulmer and Stephen M. Bollens. Responses of the chaetognath, Sagitta elegans, and larval Pacific hake, Merluccius productus, to spring diatom and copepod blooms in a temperate fjord (Dabob Bay, Washington). Progress in Oceanography, 67(3–4):442–461, November 2005. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661105001527.

Freon:2009:EBU

[FBA09] Pierre Fréon, Manuel Barange, and Javier Arístegui. Eastern Boundary Upwelling Ecosystems: Integrative and comparative approaches. Progress in Oceanography, 83(1–4):1–14, December 2009. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661109001323.

Ferrarin:2021:LLS

[FBB^{+}21] Christian Ferrarin, Marco Bajo, Alvise Benetazzo, Luigi Cavaleri, Jacopo Chiggiato, Silvio Davison, Silvio Davolio, Piero Lionello, Mirko Orlić, and Georg Umgiesser. Local and largescale controls of the exceptional Venice floods of November 2019. Progress in Oceanography, 197(??):??, September– October 2021. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661121001142.

Fanelli:2018:DSM

[FBD18] Emanuela Fanelli, Silvia Bianchelli, and Roberto Danovaro. Deep-sea mobile megafauna of Mediterranean submarine canyons and open slopes: Analysis of spatial and bathymetric gradients. Progress in Oceanography, 168(??):23–34, November 2018. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661118300995.

Freon:2008:IVA

[FBM^{+}08] Pierre Fréon, Marilú Bouchon, Christian Mullon, Christian García, and Miguel Ñiquen. Interdecadal variability of anchoveta abundance and overcapacity of the fishery in Peru. Progress in Oceanography, 79(2–4):401–412, October 2008. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661108001912.

Fagervold:2013:MCA

[FBR^{+}13] S. K. Fagervold, S. Bessette, C. Romano, D. Martin, M. Plyuscheva, N. Le Bris, and P. E. Galand. Microbial communities associated with the degradation of oak wood in the Blanes submarine canyon and its adjacent open slope (NW Mediterranean). Progress in Oceanography, 118(??):137–143, November 2013. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661113001250.

Ferrarin:2018:TDI

[FBS^{+}18] Christian Ferrarin, Debora Bellafiore, Gianmaria Sannino, Marco Bajo, and Georg Umgiesser. Tidal dynamics in the inter-connected Mediterranean, Marmara, Black and Azov seas. Progress in Oceanography, 161(??):102–115, February 2018. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661117303567.

Fernandez:2022:VCF

[FBS22] Denise Fernandez, Melissa Bowen, and Philip Sutton. Variability, coherence and forcing mechanisms in the New Zealand ocean boundary currents. Progress in Oceanography, 165 (??):168–188, ???? 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661117303440.

Ferreira:2022:AMD

[FBT^{+}22] Juliana Correa Neiva Ferreira, Natascha Menezes Bergo, Pedro Marone Tura, Mateus Gustavo Chuqui, Frederico P. Brandini, Luigi Jovane, and Vivian H. Pellizari. Abundance and microbial diversity from surface to deep water layers over the Rio Grande Rise, South Atlantic. Progress in Oceanography, 201(??):??, February 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661121002196.

Freeland:2005:ANT

[FC05] Howard J. Freeland and Patrick F. Cummins. Argo: a new tool for environmental monitoring and assessment of the world’s oceans, an example from the N.E. Pacific. Progress in Oceanography, 64(1):31–44, January 2005. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661105000194.

Farias:2007:ESC

[FC07] Laura Farías and Marcela Cornejo. Effect of seasonal changes in bottom water oxygenation on sediment N oxides and N_{2}O cycling in the coastal upwelling regime off central Chile (36.5^{o}S). Progress in Oceanography, 75(3):561–575, November 2007. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661107001693.

Frangoulis:2010:CCV

[FCEZ10] C. Frangoulis, F. Carlotti, L. Eisenhauer, and S. Zervoudaki. Converting copepod vital rates into units appropriate for biogeochemical models. Progress in Oceanography, 84(1–2):43–51, January 2010. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661109001384.

Flatau:1988:UCM

[FCG88] G. N. Flatau, R. L. Clement, and M. J. Gauthier. Uptake of cadmium by marine bacteria. Progress in Oceanography, 21(2):181–188, ???? 1988. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0079661188900377.

Fernandez-Castro:2019:NRM

[FCMCÁS19] Bieito Fernández-Castro, Beatriz Mouriño-Carballido, and Xosé Antón Álvarez-Salgado. Non-redfieldian mesopelagic nutrient remineralization in the eastern North Atlantic subtropical gyre. Progress in Oceanography, 171(??):136–153, February 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661118300338.

Flores:2019:IAS

[FCN^{+}19] Eduardo A. Flores, Leonardo R. Castro, Diego A. Narváez, Sergio Lillo, Fernando Balbontín, and Francisca Osorio-Zúñiga. Inter-annual and seasonal variations in the outer and inner sea spawning zones of southern hake, Merluccius australis, inferred from early life stages distributions in Chilean Patagonia. Progress in Oceanography, 171(??):93–107, February 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661118301666.

Fromant:2021:IEE

[FDB^{+}21] Aymeric Fromant, Karine Delord, Charles-André Bost, Yonina H. Eizenberg, Jonathan A. Botha, Yves Cherel, Paco Bustamante, Brett R. Gardner, Maud Brault-Favrou, Arnaud Lec’hvien, and John P. Y. Arnould. Impact of extreme environmental conditions: Foraging behaviour and trophic ecology responses of a diving seabird, the common diving petrel. Progress in Oceanography, 198(??):??, November 2021. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661121001609.

Fragoso:2022:PCP

[FDE^{+}22] Glaucia M. Fragoso, Emlyn J. Davies, Ingrid Ellingsen, Matilde S. Chauton, Trygve Fossum, Martin Ludvigsen, Kristine B. Steinhovden, Kanna Rajan, and Geir Johnsen. Physical controls on phytoplankton size structure, photophysiology and suspended particles in a Norwegian biological hotspot. Progress in Oceanography, 175(??):284–299, ???? 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661119300485.

Fraser:2020:CVC

[FDH20] William R. Fraser, Hugh W. Ducklow, and Sian F. Henley. Corrigendum to “Variability and change in the west Antarctic Peninsula marine system: Research priorities and opportunities” [Progr. Oceanogr. 173 (2019) 208–237]. Progress in Oceanography, 186(??):??, July 2020. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661120300896. See [HSH^{+}19].

Fernandez:2005:MSI

[FDHT05] Vicente Fernández, David E. Dietrich, Robert L. Haney, and Joaquín Tintoré. Mesoscale, seasonal and interannual variability in the Mediterranean Sea using a numerical ocean model. Progress in Oceanography, 66(2–4):321–340, August 2005. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661105000765.

Frouin:2013:PBP

[FDM^{+}13] Heloise Frouin, Neil Dangerfield, Robie W. Macdonald, M. Galbraith, Norman Crewe, Patrick Shaw, David Mackas, and Peter S. Ross. Partitioning and bioaccumulation of PCBs and PBDEs in marine plankton from the Strait of Georgia, British Columbia, Canada. Progress in Oceanography, 115(??):65–75, August 2013. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661113000748.

Figueiras:2014:EPS

[FEGA^{+}14] F. G. Figueiras, O. Espinoza-González, B. Arbones, J. L. Garrido, I. G. Teixeira, and C. G. Castro. Estimating phytoplankton size-fractionated primary production in the northwestern Iberian upwelling: Is mixotrophy relevant in pigmented nanoplankton? Progress in Oceanography, 128(??):88–97, November 2014. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661114001347.

Feistel:1993:ETS

[Fei93] Rainer Feistel. Equilibrium thermodynamics of seawater revisited. Progress in Oceanography, 31(2):101–179, ???? 1993. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0079661193900248.

Feistel:2003:NEG

[Fei03] Rainer Feistel. A new extended Gibbs thermodynamic potential of seawater. Progress in Oceanography, 58(1):43–114, July 2003. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661103000880. See corrigendum

[Fei04].

Feistel:2004:CNE

[Fei04] Rainer Feistel. Corrigendum to: A new extended Gibbs thermodynamic potential of seawater [Progress in Oceanography 58 (2003) 43–114]. Progress in Oceanography, 61(1): 99, April 2004. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661104000564. See [Fei03].

Fontana:2016:TVC

[FELJ16] Rachel E. Fontana, Meredith L. Elliott, John L. Largier, and Jaime Jahncke. Temporal variation in copepod abundance and composition in a strong, persistent coastal upwelling zone. Progress in Oceanography, 142(??):1–16, March 2016. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661116000094.

Farach-Espinoza:2022:COM

[FELMGM^{+}22] Edgardo B. Farach-Espinoza, Juana López-Martínez, Ricardo García-Morales, Manuel O. Nevárez-Martínez, Sofía Ortega-García, and Daniel B. Lluch-Cota. Coupling oceanic mesoscale events with catches of the Pacific sardine (Sardinops sagax) in the Gulf of California. Progress in Oceanography, 206(??):??, August 2022. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661122001173.

Farmer:1983:POF

[FF83] David M. Farmer and Howard J. Freeland. The physical oceanography of fjords. Progress in Oceanography, 12(2):147– 194, IN1, 195–219, ???? 1983. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0079661183900046.

Field:2006:TMB

[FFA06] J. C. Field, R. C. Francis, and K. Aydin. Top-down modeling and bottom-up dynamics: Linking a fisheries-based ecosystem model with climate hypotheses in the Northern California Current. Progress in Oceanography, 68(2–4):238–270, February 2006. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661106000085.

Fernandez:2009:PPN

[FFA09] C. Fernández, L. Farías, and M. E. Alcaman. Primary production and nitrogen regeneration processes in surface waters of the Peruvian upwelling system. Progress in Oceanography, 83(1–4):159–168, December 2009. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661109001086.

Feuilloley:2020:CCE

[FFS^{+}20] Guillaume Feuilloley, Jean-Marc Fromentin, Lars Stemmann, Hervé Demarcq, Claude Estournel, and Claire Saraux. Concomitant changes in the environment and small pelagic fish community of the Gulf of Lions. Progress in Oceanography, 186 (??):??, July 2020. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661120301142.

Fujioka:2018:STV

[FFT^{+}18] Ko Fujioka, Hiromu Fukuda, Yaoki Tei, Suguru Okamoto, Hidetada Kiyofuji, Seishiro Furukawa, Junichi Takagi, Ethan Estess, Charles J. Farwell, Daniel W. Fuller, Nobuaki Suzuki, Seiji Ohshimo, and Takashi Kitagawa. Spatial and temporal variability in the trans-Pacific migration of Pacific bluefin tuna (Thunnus orientalis) revealed by archival tags. Progress in Oceanography, 162(??):52–65, March 2018. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661117300988.

Fuchs:2016:SLV

[FG16] Heidi L. Fuchs and Gregory P. Gerbi. Seascape-level variation in turbulence- and wave-generated hydrodynamic signals experienced by plankton. Progress in Oceanography, 141(??):109–129, February 2016. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661115002566.

Franco-Gordo:2004:PBL

[FGGDF^{+}04] C. Franco-Gordo, E. Godínez-Domínguez, A. E. Filonov, I. E. Tereshchenko, and J. Freire. Plankton biomass and larval fish abundance prior to and during the El Niño period of 1997–1998 along the central Pacific coast of México. Progress in Oceanography, 63(3):99–123, November 2004. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661104001259.

Freitas:2023:NCM

[FGL^{+}23] Carla Freitas, Kjell Gundersen, Lotta Lindblom, Martin Biuw, and Tore Haug. Nutrient concentrations in minke whale faeces and the potential impact on dissolved nutrient pools off Svalbard, Norway. Progress in Oceanography, 210(??):??, January 2023. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661122001860.

Flexas:2006:SSO

[FGR^{+}06] M. M. Flexas, D. Gomis, S. Ruiz, A. Pascual, and P. León. In situ and satellite observations of the eastward migration of the Western Alboran Sea Gyre. Progress in Oceanography, 70(2–4):486–509, August 2006. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661106000590.

Fajar:2015:ACC

[FGS^{+}15] N. M. Fajar, E. F. Guallart, R. Steinfeldt, A. F. Ríos, J. L. Pelegrí, C. Pelejero, E. Calvo, and F. F. Pérez. Anthropogenic CO_{2} changes in the Equatorial Atlantic Ocean. Progress in Oceanography, 134(??):256–270, May 2015. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661115000385.

Fonvieille:2023:SOC

[FGS^{+}23] Nadège Fonvieille, Christophe Guinet, Martin Saraceno, Baptiste Picard, Martin Tournier, Pauline Goulet, Claudio Campagna, Julieta Campagna, and David Nerini. Swimming in an ocean of curves: a functional approach to understanding elephant seal habitat use in the Argentine Basin. Progress in Oceanography, 218(??):??, November 2023. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661123001635.

Fanjul:1997:DTE

[FGSA97] Enrique Alvarez Fanjul, Begoña Pérez Gómez, and Ignacio Rodríguez Sánchez-Arévalo. A description of the tides in the Eastern North Atlantic. Progress in Oceanography, 40(??):217–244, ???? 1997. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661198000032.

Feistel:1995:GTP

[FH95] Rainer Feistel and Eberhard Hagen. On the GIBBS thermodynamic potential of seawater. Progress in Oceanography, 36(4):249–327, ???? 1995. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0079661196000018.

Feistel:2003:CCS

[FHG03] Rainer Feistel, Eberhard Hagen, and Keith Grant. Climatic changes in the subtropical Southeast Atlantic: the St. Helena Island Climate Index (1893–1999). Progress in Oceanography, 59(2–3):321–337, October 2003. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661103001678.

Fu:2024:AVM

[FHL^{+}24] Minghan Fu, Guoqing Han, Xiaojie Lu, Wenjin Sun, Joel Sommeria, Alexandre Stegner, Rui M. A. Caldeira, and Changming Dong. Analysis of vortex merging from a rotating tank laboratory experiment. Progress in Oceanography, 222(??):??, March 2024. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661124000338.

Fasham:1983:STD

[FHP83] M. J. R. Fasham, P. M. Holligan, and P. R. Pugh. The spatial and temporal development of the spring phytoplankton bloom in the Celtic Sea, April 1979. Progress in Oceanography, 12(1):87–145, ???? 1983. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0079661183900071.

Frutos:2019:DSA

[FJ19] Inmaculada Frutos and Anna Maria Jażdżewska. Deepsea amphipod fauna of the Sea of Okhotsk. Progress in Oceanography, 178(??):??, November 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661119301946.

Feng:2021:BHN

[FJA^{+}21] Zhixuan Feng, Rubao Ji, Carin Ashjian, Jinlun Zhang, Robert Campbell, and Jacqueline M. Grebmeier. Benthic hotspots on the northern Bering and Chukchi continental shelf: Spatial variability in production regimes and environmental drivers. Progress in Oceanography, 191(??):??, February 2021. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661120302329.

Friedlaender:2010:ENA

[FJH10] Ari S. Friedlaender, David W. Johnston, and Patrick N. Halpin. Effects of the North Atlantic oscillation on sea ice breeding habitats of harp seals (Pagophilus groenlandicus) across the North Atlantic. Progress in Oceanography, 86(1–2):261– 266, July 2010. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661110000364.

Farmer:2014:SIP

[FJhT^{+}14] David Farmer, Sen Jan, Yu heng Tseng, Dongliang Yuan, and SangHo Lee. Special issue of 16th PAMS Meeting: Recent advances in the oceanography of Pacific-Asian Marginal Seas. Progress in Oceanography, 121(??):1–6, February 2014. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661113001924.

Fowler:1986:RLP

[FK86] Scott W. Fowler and George A. Knauer. Role of large particles in the transport of elements and organic compounds through the oceanic water column. Progress in Oceanography, 16(3):147–194, ???? 1986. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0079661186900327.

Frost:1999:EDE

[FK99] Bruce W. Frost and Michio J. Kishi. Ecosystem dynamics in the eastern and western gyres of the Subarctic Pacific — a review of lower trophic level modelling. Progress in Oceanography, 43(2–4):317–333, March 1999. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661199000166.

Friedland:2013:THC

[FKH^{+}13] Kevin D. Friedland, Joe Kane, Jonathan A. Hare, R. Gregory Lough, Paula S. Fratantoni, Michael J. Fogarty, and Janet A. Nye. Thermal habitat constraints on zooplankton species associated with Atlantic cod (Gadus morhua) on the US Northeast Continental Shelf. Progress in Oceanography, 116(??):1–13, September 2013. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661113000530.

Fischer:2015:ISV

[FKZ^{+}15] J. Fischer, J. Karstensen, R. Zantopp, M. Visbeck, A. Biastoch, E. Behrens, C. W. Böning, D. Quadfasel, K. Jochumsen, H. Valdimarsson, S. Jónsson, S. Bacon, N. P. Holliday, S. Dye, M. Rhein, and C. Mertens. Intra-seasonal variability of the DWBC in the western subpolar North Atlantic. Progress in Oceanography, 132(??):233–249, March 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S007966111400041X.

Fiedler:2006:IRE

[FL06] Paul C. Fiedler and Miguel F. Lavín. Introduction: a review of eastern tropical Pacific oceanography. Progress in Oceanography, 69(2–4):94–100, May 2006. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661106000280.

Flament:2002:SVC

[Fla02] P. Flament. A state variable for characterizing water masses and their diffusive stability: spiciness. Progress in Oceanography, 54(1–4):493–501, July 2002. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661102000654.

Florez-Leiva:2022:MPI

[FLDF22] Lennin Florez-Leiva, Ellen Damm, and Laura Farías. Methane production induced by dimethylsulfide in surface water of an upwelling ecosystem. Progress in Oceanography, 112–113(??): 38–48, ???? 2022. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S007966111300030X.

Flach:1998:FTB

[FLdST98] E. Flach, M. Lavaleye, H. de Stigter, and L. Thomsen. Feeding types of the benthic community and particle transport across the slope of the N.W. European Continental Margin (Goban Spur). Progress in Oceanography, 42(1–4):209–231, December 1998. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661198000354.

Friederich:2008:ASC

[FLUC08] Gernot E. Friederich, Jesus Ledesma, Osvaldo Ulloa, and Francisco P. Chavez. Air–sea carbon dioxide fluxes in the coastal southeastern tropical Pacific. Progress in Oceanography, 79(2–4):156–166, October 2008. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661108001699.

Flynn:2003:MMN

[Fly03] Kevin J. Flynn. Modelling multi-nutrient interactions in phytoplankton; balancing simplicity and realism. Progress in Oceanography, 56(2):249–279, February 2003. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661103000065.

Flynn:2010:EMS

[Fly10] Kevin J. Flynn. Ecological modelling in a sea of variable stoichiometry: Dysfunctionality and the legacy of Redfield and Monod. Progress in Oceanography, 84(1–2):52–65, January 2010. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661109001396.

Feistel:2007:GPF

[FM07] Rainer Feistel and Giles M. Marion. A Gibbs–Pitzer function for high-salinity seawater thermodynamics. Progress in Oceanography, 74(4):515–539, September 2007. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661107000973.

Feunteun:2015:SIC

[FMC^{+}15] Eric Feunteun, Michael J. Miller, Alexandre Carpentier, Jun Aoyama, Christine Dupuy, Mari Kuroki, Marc Pagano, Elodie Réveillac, Daniel Sellos, Shun Watanabe, Katsumi Tsukamoto, and Tsuguo Otake. Stable isotopic composition of anguilliform leptocephali and other food web components from west of the Mascarene Plateau. Progress in Oceanography, 137 (part A)(??):69–83, September 2015. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661115001299.

Fisher:2020:CCM

[FMC^{+}20] Jennifer L. Fisher, Jennifer Menkel, Louise Copeman, C. Tracy Shaw, Leah R. Feinberg, and William T. Peterson. Comparison of condition metrics and lipid content between Euphausia pacifica and Thysanoessa spinifera in the northern California Current, USA. Progress in Oceanography, 188(??):??, October 2020. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661120301567.

Frey:2015:DPR

[FMCG15] Karen E. Frey, G. W. K. Moore, Lee W. Cooper, and Jacqueline M. Grebmeier. Divergent patterns of recent sea ice cover across the Bering, Chukchi, and Beaufort Seas of the Pacific Arctic Region. Progress in Oceanography, 136(??):32–49, August 2015. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661115001068.

Flach:2002:MMC

[FMH02] Els Flach, Agnes Muthumbi, and Carlo Heip. Meiofauna and macrofauna community structure in relation to sediment composition at the Iberian margin compared to the Goban Spur (NE Atlantic). Progress in Oceanography, 52(??):433– 457, ???? 2002. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661102000186.

Fergusson:2020:TRJ

[FMM^{+}20] Emily Fergusson, Todd Miller, Megan V. McPhee, Corey Fugate, and Haila Schultz. Trophic responses of juvenile Pacific salmon to warm and cool periods within inside marine waters of Southeast Alaska. Progress in Oceanography, 186(??): ??, July 2020. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661120301178.

Fontela:2019:LTI

[FMP19] Marcos Fontela, Herlé Mercier, and Fiz F. Pérez. Longterm integrated biogeochemical budget driven by circulation in the eastern subpolar North Atlantic. Progress in Oceanography, 173(??):51–65, April 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661118302660.

Frundt:2013:LTC

[FMSBW13] B. Fründt, T. J. Müller, D. E. Schulz-Bull, and J. J. Waniek. Long-term changes in the thermocline of the subtropical Northeast Atlantic (33^{o}N, 22^{o}W). Progress in Oceanography, 116(??):246–260, September 2013. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661113001109.

Fujioka:2015:ELH

[FMT15] Hidefumi A. Fujioka, Ryuji J. Machida, and Atsushi Tsuda. Early life history of Neocalanus plumchrus (Calanoida: Copepoda) in the western subarctic Pacific. Progress in Oceanography, 137 (part A)(??):196–208, September 2015. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661115001378.

Fukumori:1991:HNA

[FMW91] Ichiro Fukumori, Françoise Martel, and Carl Wunsch. The hydrography of the North Atlantic in the early 1980s. An atlas. Progress in Oceanography, 27(1–2):1–110, ???? 1991. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/007966119190014D.

Fujisaki:2014:HDA

[FMWW14] Ayumi Fujisaki, Humio Mitsudera, Jia Wang, and Masaaki Wakatsuchi. How does the Amur River discharge flow over the northwestern continental shelf in the Sea of Okhotsk? Progress in Oceanography, 126(??):8–20, August 2014. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661114000767.

Foyn:1965:O

[Føy65] Ernst Føyn. The oxymeter. Progress in Oceanography, 3 (??):137–144, ???? 1965. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0079661165900121.

Feinberg:2003:VDI

[FP03] L. R. Feinberg and W. T. Peterson. Variability in duration and intensity of euphausiid spawning off central Oregon, 1996–2001. Progress in Oceanography, 57(3–4):363–379, June 2003. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S007966110300106X.

Forget:2015:PRS

[FP15] Gaël Forget and Rui M. Ponte. The partition of regional sea level variability. Progress in Oceanography, 137 (part A)(??):173–195, September 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661115001354.

Fabiano:2001:FPL

[FPD^{+}01] M. Fabiano, A. Pusceddu, A. Dell’Anno, M. Armeni, S. Vanucci, R. S. Lampitt, G. A. Wolff, and R. Danovaro. Fluxes of phytopigments and labile organic matter to the deep ocean in the NE Atlantic Ocean. Progress in Oceanography, 50(??):89–104, ???? 2001. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661101000490.

Fasham:1985:FAS

[FPIJ85] M. J. R. Fasham, T. Platt, B. Irwin, and K. Jones. Factors affecting the spatial pattern of the deep chlorophyll maximum in the region of the Azores front. Progress in Oceanography, 14(??):129–165, ???? 1985. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0079661185900096.

Ferry:2015:SMS

[FPJ^{+}15] Alexander J. Ferry, Tania Prvan, Brian Jersky, Xavier Crosta, and Leanne K. Armand. Statistical modeling of Southern Ocean marine diatom proxy and winter sea ice data: Model comparison and developments. Progress in Oceanography, 131(??):100–112, February 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661114002031.

Franklin:2009:DDL

[FPS^{+}09] Daniel J. Franklin, Alex J. Poulton, Michael Steinke, Jeremy Young, Ilka Peeken, and Gill Malin. Dimethylsulphide, DMSPlyase activity and microplankton community structure inside and outside of the Mauritanian upwelling. Progress in Oceanography, 83(1–4):134–142, December 2009. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661109001074.

Fu:2013:EMF

[FPS^{+}13] Caihong Fu, R. Ian Perry, Yunne-Jai Shin, Jake Schweigert, and Huizhu Liu. An ecosystem modelling framework for incorporating climate regime shifts into fisheries management. Progress in Oceanography, 115(??):53–64, August 2013. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661113000281.

Fragoso:2016:BPE

[FPY^{+}16] Glaucia M. Fragoso, Alex J. Poulton, Igor M. Yashayaev, Erica J. H. Head, Mark C. Stinchcombe, and Duncan A. Purdie. Biogeographical patterns and environmental controls of phytoplankton communities from contrasting hydrographical zones of the Labrador Sea. Progress in Oceanography, 141(??):212–226, February 2016. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661115002530.

Fraser:1969:VOC

[Fra69] J. H. Fraser. Variability in the oceanic content of plankton in the Scottish area. Progress in Oceanography, 5(??):149– 159, ???? 1969. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661169900378.

Fiechter:2015:REC

[FRCH15] Jerome Fiechter, Kenneth A. Rose, Enrique N. Curchitser, and Katherine S. Hedstrom. The role of environmental controls in determining sardine and anchovy population cycles in the California Current: Analysis of an end-to-end model. Progress in Oceanography, 138(??):381–398, November 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S007966111400202X.

Freeland:2007:SHO

[Fre07] Howard Freeland. A short history of Ocean Station Papa and Line P. Progress in Oceanography, 75(2):120–125, October 2007. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661107001450.

Fischer:2009:OAP

[FRK^{+}09] G. Fischer, C. Reuter, G. Karakas, N. Nowald, and G. Wefer. Offshore advection of particles within the Cape Blanc filament, Mauritania: Results from observational and modelling studies. Progress in Oceanography, 83(1–4):322–330, December 2009. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661109000792.

Frost:1993:MSP

[Fro93] Bruce W. Frost. A modelling study of processes regulating plankton standing stock and production in the open subarctic Pacific Ocean. Progress in Oceanography, 32(??):17–56, ???? 1993. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661193900082.

Frost:2005:DBC

[Fro05] Bruce W. Frost. Diatom blooms and copepod responses: Paradigm or paradox? Progress in Oceanography, 67(3–4): 283–285, November 2005. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661105001436.

Fettweis:2019:UAS

[FRV^{+}19] Michael Fettweis, Rolf Riethmüller, Romaric Verney, Marius Becker, Joan Backers, Matthias Baeye, Marion Chapalain, Styn Claeys, Jan Claus, Tom Cox, Julien Deloffre, Davy Depreiter, Flavie Druine, Götz Flöser, Steffen Grünler, Frédéric Jourdin, Robert Lafite, Janine Nauw, Bouchra Nechad, Rüdiger Röttgers, Aldo Sottolichio, Tom Van Engeland, Wim Vanhaverbeke, and Hans Vereecken. Uncertainties associated with in situ high-frequency long-term observations of suspended particulate matter concentration using optical and acoustic sensors. Progress in Oceanography, 178(??):??, November 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661118302003.

Favoretto:2022:WMH

[FSAO22] Fabio Favoretto, Carlos Sánchez, and Octavio Aburto-Oropeza. Warming and marine heatwaves tropicalize rocky reefs communities in the Gulf of California. Progress in Oceanography, 206(??):??, August 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0079661122000994.

Fonseca:2010:VNA

[FSVL10] Gustavo Fonseca, Thomas Soltwedel, Ann Vanreusel, and Mats Lindegarth. Variation in nematode assemblages over multiple spatial scales and environmental conditions in Arctic deep seas. Progress in Oceanography, 84(3–4):174–184, March 2010. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661109001888.

Fiedler:2006:HET

[FT06] Paul C. Fiedler and Lynne D. Talley. Hydrography of the eastern tropical Pacific: a review. Progress in Oceanography, 69(2–4):143–180, May 2006. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661106000309.

Farre:2016:DRT

[FTC^{+}16] Marc Farré, Víctor M. Tuset, Joan E. Cartes, Enric Massutí, and Antoni Lombarte. Depth-related trends in morphological and functional diversity of demersal fish assemblages in the western Mediterranean Sea. Progress in Oceanography, 147(??):22–37, September 2016. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S007966111530046X.

Forest:2011:BCF

[FTG^{+}11] Alexandre Forest, Jean-Éric Tremblay, Yves Gratton, Johannie Martin, Jonathan Gagnon, Gérald Darnis, Makoto Sampei, Louis Fortier, Mathieu Ardyna, Michel Gosselin, Hiroshi Hattori, Dan Nguyen, Roxane Maranger, Dolors Vaqué, Cèlia Marrasé, Carlos Pedrós-Alió, Amélie Sallon, Christine Michel, Colleen Kellogg, Jody Deming, Elizabeth Shadwick, Helmuth Thomas, Heike Link, Philippe Archambault, and Dieter Piepenburg. Biogenic carbon flows through the planktonic food web of the Amundsen Gulf (Arctic Ocean): a synthesis of field measurements and inverse modeling analyses. Progress in Oceanography, 91(4):410–436, December 2011. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661111000498.

Fourati:2018:SSD

[FTG^{+}18] Rania Fourati, Marc Tedetti, Catherine Guigue, Madeleine Goutx, Nicole Garcia, Hatem Zaghden, Sami Sayadi, and Boubaker Elleuch. Sources and spatial distribution of dissolved aliphatic and polycyclic aromatic hydrocarbons in surface coastal waters of the Gulf of Gabès (Tunisia, Southern Mediterranean Sea). Progress in Oceanography, 163(??):232– 247, April 2018. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661116301094.

Fukumori:2019:DKG

[FTHK19] Hiroaki Fukumori, Tsuyoshi Takano, Kazunori Hasegawa, and Yasunori Kano. Deepest known gastropod fauna: Species composition and distribution in the Kuril–Kamchatka Trench. Progress in Oceanography, 178(??):??, November 2019. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661119302307.

Farias:2021:TMV

[FTSF21] L. Farías, S. Tenorio, K. Sanzana, and J. Faundez. Temporal methane variability in the water column of an area of seasonal coastal upwelling: a study based on a 12 year time series. Progress in Oceanography, 195(??):??, July 2021. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661121000768.

Fuglister:1963:GS

[Fug63] F. C. Fuglister. Gulf Stream ’60. Progress in Oceanography, 1(??):265–272, IN9, 273–274, IN11–IN14, 275–373, ???? 1963. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0079661163900077.

Fukumori:1991:CAM

[Fuk91] Ichiro Fukumori. Circulation about the Mediterranean tongue: an analysis of an EOF-based model ocean. Progress in Oceanography, 27(1–2):197–224, ???? 1991. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/007966119190016F.

Fernandez-Urruzola:2016:MDP

[FUOG^{+}16] I. Fernández-Urruzola, N. Osma, M. Gómez, F. Pollehne, L. Postel, and T. T. Packard. Modeling downward particulate organic nitrogen flux from zooplankton ammonium regeneration in the northern Benguela. Progress in Oceanography, 149(??):121–133, December 2016. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661115300392.

Fabri:2019:EES

[FVA^{+}19] Marie-Claire Fabri, Beatriz Vinha, Anne-Gaëlle Allais, Marie-Edith Bouhier, Olivier Dugornay, Arnaud Gaillot, and Aurelien Arnaubec. Evaluating the ecological status of coldwater coral habitats using non-invasive methods: an example from cassidaigne canyon, northwestern Mediterranean Sea. Progress in Oceanography, 178(??):??, November 2019. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661119301302.

Floeter:2017:PEO

[FvBA^{+}17] Jens Floeter, Justus E. E. van Beusekom, Dominik Auch, Ulrich Callies, Jeffrey Carpenter, Tim Dudeck, Sabine Eberle, André Eckhardt, Dominik Gloe, Kristin Hänselmann, Marc Hufnagl, Silke Janßen, Hermann Lenhart, Klas Ove Möller, Ryan P. North, Thomas Pohlmann, Rolf Riethmüller, Sabrina Schulz, Stefan Spreizenbarth, Axel Temming, Bettina Walter, Oliver Zielinski, and Christian Möllmann. Pelagic effects of offshore wind farm foundations in the stratified North Sea. Progress in Oceanography, 156(??):154–173, August 2017. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661117300381.

Flores-Valiente:2023:ICT

[FVLC^{+}23] Jorge Flores-Valiente, Christophe Lett, François Colas, Laure Pecquerie, Arturo Aguirre-Velarde, Fanny Rioual, Jorge Tam, Arnaud Bertrand, Patricia Ayón, Saidou Sall, Nicolas Barrier, and Timothée Brochier. Influence of combined temperature and food availability on Peruvian anchovy (Engraulis ringens) early life stages in the northern Humboldt Current system: a modelling approach. Progress in Oceanography, 215 (??):??, July 2023. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661123000770.

Fukumori:1991:ERN

[FW91] Ichiro Fukumori and Carl Wunsch. Efficient representation of the North Atlantic hydrographic and chemical distributions. Progress in Oceanography, 27(1–2):111–195, ???? 1991. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/007966119190015E.

Friederich:2002:ICC

[FWBC02] G. E. Friederich, P. M. Walz, M. G. Burczynski, and F. P. Chavez. Inorganic carbon in the central California upwelling system during the 1997–1999 El Niño–La Niña event. Progress in Oceanography, 54(1–4):185–203, July 2002. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661102000496.

Fileman:2017:SLO

[FWH^{+}17] Elaine S. Fileman, Daniel A. White, Rachel A. Harmer, Ülgen Aytan, Glen A. Tarran, Tim Smyth, and Angus Atkinson. Stress of life at the ocean’s surface: Latitudinal patterns of UV sunscreens in plankton across the Atlantic. Progress in Oceanography, 158(??):171–184, November 2017. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661116300246.

Fukumori:2015:NUF

[FWL^{+}15] Ichiro Fukumori, Ou Wang, William Llovel, Ian Fenty, and Gael Forget. A near-uniform fluctuation of ocean bottom pressure and sea level across the deep ocean basins of the Arctic Ocean and the Nordic Seas. Progress in Oceanography, 134(??):152–172, May 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661115000245.

Fewings:2015:CWC

[FWO15] Melanie R. Fewings, Libe Washburn, and J. Carter Ohlmann. Coastal water circulation patterns around the Northern Channel Islands and Point Conception, California. Progress in Oceanography, 138 (part A)(??):283–304, November 2015. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661115002025.

Fang:1988:MPT

[FY88] Guohong Fang and Jingfei Yang. Modeling and prediction of tidal currents in the Korea Strait. Progress in Oceanography, 21(3–4):307–318, ???? 1988. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0079661188900109.

Fujii:2005:SBR

[FYYC05] Masahiko Fujii, Naoki Yoshie, Yasuhiro Yamanaka, and Fei Chai. Simulated biogeochemical responses to iron enrichments in three high nutrient, low chlorophyll (HNLC) regions. Progress in Oceanography, 64(2–4):307–324, February 2005. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661105000364.

Fang:1988:NMF

[FZ88] Guohong Fang and Baoren Zhao. A note on the main forcing of the northeastward flowing current off the Southeast China Coast. Progress in Oceanography, 21(3–4):363–372, ???? 1988. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661188900146.

Fang:2023:TCD

[FZY^{+}23] Ziming Fang, Kun Zhang, Weifeng Yang, Min Chen, Aron Stubbins, and Huina Hu. Top-down control over dissolved organic carbon in the bottom water of the Weddell Sea and its implication for the continental shelf pump. Progress in Oceanography, 219(??):??, December 2023. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S007966112300188X.

Greiner:2000:CRV

[GA00] Eric Greiner and Sabine Arnault. Comparing the results of a 4D-variational assimilation of satellite and in situ data with WOCE CITHER hydrographic measurements in the tropical Atlantic. Progress in Oceanography, 47(1):1–68, October 2000. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661100000318.

Gooday:2001:MEP

[GA01] Andrew J. Gooday and Elisabeth Alve. Morphological and ecological parallels between sublittoral and abyssal foraminiferal species in the NE Atlantic: a comparison of Stainforthia fusiformis and Stainforthia sp. Progress in Oceanography, 50(??):261–283, ???? 2001. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S007966110100057X.

Goni:2010:SIV

[GA10] Nicolas Goñi and Haritz Arrizabalaga. Seasonal and interannual variability of fat content of juvenile albacore (Thunnus alalunga) and bluefin (Thunnus thynnus) tunas during their feeding migration to the Bay of Biscay. Progress in Oceanography, 86(1–2):115–123, July 2010. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661110000509.

Gaichas:2015:WWB

[GAF15] Sarah Gaichas, Kerim Aydin, and Robert C. Francis. Wasp waist or beer belly? Modeling food web structure and energetic control in Alaskan marine ecosystems, with implications for fishing and environmental forcing. Progress in Oceanography, 138 (part A)(??):1–17, November 2015. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661115002001.

Galovic:2017:SCN

[Gal17] Ines Galović. Sarmatian calcareous nannofossil assemblages in the SW Paratethyan marginal marine environments: Implications for palaeoceanography and the palaeoclimate. Progress in Oceanography, 156(??):209–220, August 2017. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661116300970.

Greiner:1998:TMEa

[GAM98a] Eric Greiner, Sabine Arnault, and Alain Morlière. Twelve monthly experiments of 4D-variational assimilation in the tropical Atlantic during 1987: Part 1: Method and statistical results. Progress in Oceanography, 41(2):141–202, August 1998. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661198000160.

Greiner:1998:TMEb

[GAM98b] Eric Greiner, Sabine Arnault, and Alain Morlière. Twelve monthly experiments of 4D-variational assimilation in the tropical Atlantic during 1987:Part 2: Oceanographic interpretation. Progress in Oceanography, 41(2):203–247, August 1998. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661198000172.

Gamo:2014:ERP

[Gam14] Toshitaka Gamo. Excess ^{222}Rn profiles in the bottom layer of the Japan Sea and their implication for bottom water dynamics. Progress in Oceanography, 121(??):94–97, February 2014. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661113002048.

Guiet:2016:ELT

[GAPM16] Jérôme Guiet, Olivier Aumont, Jean-Christophe Poggiale, and Olivier Maury. Effects of lower trophic level biomass and water temperature on fish communities: a modelling study. Progress in Oceanography, 146(??):22–37, August 2016. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661115300367.

Gargett:2003:DDO

[Gar03] A. E. Gargett. Differential diffusion: an oceanographic primer. Progress in Oceanography, 56(3–4):559–570, March 2003. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661103000259.

Garrett:2006:TDO

[Gar06] Chris Garrett. Turbulent dispersion in the ocean. Progress in Oceanography, 70(2–4):113–125, August 2006. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661106000577.

Guillaumot:2022:BSS

[GAS^{+}22] Charlène Guillaumot, Jean Artois, Thomas Saucède, Laura Demoustier, Camille Moreau, Marc Eléaume, Antonio Agüera, and Bruno Danis. Broad-scale species distribution models applied to data-poor areas. Progress in Oceanography, 175(??):198–207, ???? 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661118301939.

Gasol:2009:MPB

[GASV^{+}09] Josep M. Gasol, Laura Alonso-Sáez, Dolors Vaqué, Federico Baltar, Maria Ll. Calleja, Carlos M. Duarte, and Javier Arístegui. Mesopelagic prokaryotic bulk and singlecell heterotrophic activity and community composition in the NW Africa–Canary Islands coastal-transition zone. Progress in Oceanography, 83(1–4):189–196, December 2009. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661109000822.

Gooday:1996:BFA

[GBB96] Andrew J. Gooday, Samuel S. Bowser, and Joan M. Bernhard. Benthic foraminiferal assemblages in explorers cove, Antarctica: a shallow-water site with deep-sea characteristics. Progress in Oceanography, 37(2):117–166, ???? 1996. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661196000079.

Guerreiro:2019:TGC

[GBB^{+}19] Catarina V. Guerreiro, Karl-Heinz Baumann, Geert-Jan A. Brummer, Laura F. Korte, Carolina Sá, and Jan-Berend W. Stuut. Transatlantic gradients in calcifying phytoplankton (coccolithophore) fluxes. Progress in Oceanography, 176(??):??, September 2019. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661118302477.

Granata:2020:VDD

[GBB^{+}20] Antonia Granata, Alessandro Bergamasco, Pietro Battaglia, Giacomo Milisenda, Marco Pansera, Vincenzo Bonanzinga, Giuseppe Arena, Franco Andaloro, Salvatore Giacobbe, Silvestro Greco, Rosanna Guglielmo, Nunziacarla Spanò, Giacomo Zagami, and Letterio Guglielmo. Vertical distribution and diel migration of zooplankton and micronekton in Polcevera submarine canyon of the Ligurian mesopelagic zone (NW Mediterranean Sea). Progress in Oceanography, 183(??): ??, April 2020. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661120300379.

Gibb:2000:SPP

[GBC^{+}00] S. W. Gibb, R. G. Barlow, D. G. Cummings, N. W. Rees, C. C. Trees, P. Holligan, and D. Suggett. Surface phytoplankton pigment distributions in the Atlantic Ocean: an assessment of basin scale variability between 50^{o}N and 50^{o}S. Progress in Oceanography, 45(3–4):339–368, April 2000. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661100000070.

Grebmeier:2015:ECP

[GBC^{+}15] Jacqueline M. Grebmeier, Bodil A. Bluhm, Lee W. Cooper, Seth L. Danielson, Kevin R. Arrigo, Arny L. Blanchard, Janet T. Clarke, Robert H. Day, Karen E. Frey, Rolf R. Gradinger, Monika Kedra, Brenda Konar, Kathy J. Kuletz, Sang H. Lee, James R. Lovvorn, Brenda L. Norcross, and Stephen R. Okkonen. Ecosystem characteristics and processes facilitating persistent macrobenthic biomass hotspots and associated benthivory in the Pacific Arctic. Progress in Oceanography, 136(??):92–114, August 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661115001032.

Grados:2016:SSP

[GBC^{+}16] Daniel Grados, Arnaud Bertrand, François Colas, Vincent Echevin, Alexis Chaigneau, Dimitri Gutiérrez, Gary Vargas, and Ronan Fablet. Spatial and seasonal patterns of fine-scale to mesoscale upper ocean dynamics in an Eastern boundary current system. Progress in Oceanography, 142(??):105–116, March 2016. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661115300057.

Goetze:2005:GMD

[GBG05] Erica Goetze and Janet Bradford-Grieve. Genetic and morphological description of Eucalanus spinifer T. Scott, 1894 (Calanoida: Eucalanidae), a circumglobal sister species of the copepod E. hyalinus s.s. (Claus, 1866). Progress in Oceanography, 65(1):55–87, April 2005. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661105000376.

Groeneveld:2020:BKG

[GBH^{+}20] Jürgen Groeneveld, Uta Berger, Natasha Henschke, Evgeny A. Pakhomov, Christian S. Reiss, and Bettina Meyer. Blooms of a key grazer in the Southern Ocean — an individual-based model of Salpa thompsoni. Progress in Oceanography, 185(??): ??, June 2020. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661120300781.

Ginger:2001:OMA

[GBM^{+}01] Michael L. Ginger, David S. M. Billett, Karen L. Mackenzie, Konstandinos Kiriakoulakis, Renato R. Neto, Daniel K. Boardman, Vera L. C. S. Santos, Ian M. Horsfall, and George A. Wolff. Organic matter assimilation and selective feeding by holothurians in the deep sea: some observations and comments. Progress in Oceanography, 50(??):407–421, ???? 2001. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661101000635.

Godefroid:2019:STM

[GBT^{+}19] Martin Godefroid, Jennifer L. Boldt, James T. Thorson, Robyn Forrest, Stéphane Gauthier, Linnea Flostrand, R. Ian Perry, Andrew R. S. Ross, and Moira Galbraith. Spatiotemporal models provide new insights on the biotic and abiotic drivers shaping Pacific herring (Clupea pallasi) distribution. Progress in Oceanography, 178(??):??, November 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661119300850.

Gregr:2009:BNP

[GC09] Edward J. Gregr and Kenneth O. Coyle. The biogeography of the North Pacific right whale (Eubalaena japonica). Progress in Oceanography, 80(3–4):188–198, March 2009. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661108002115.

Guo:2014:RIS

[GC14] C. Guo and X. Chen. A review of internal solitary wave dynamics in the northern South China Sea. Progress in Oceanography, 121(??):7–23, February 2014. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661113000414.

Ghosh:2022:ICU

[GCB^{+}22] Jayashree Ghosh, Kunal Chakraborty, Trishneeta Bhattacharya, Vinu Valsala, and Balaji Baduru. Impact of coastal upwelling dynamics on the pCO_{2} variability in the southeastern Arabian Sea. Progress in Oceanography, 203(??):??, April–May 2022. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661122000465.

Ge:2024:GPM

[GCC^{+}24] Ruping Ge, Hongju Chen, Tianying Chen, Yu Zang, Weimin Wang, Yunyun Zhuang, and Guangxing Liu. Geographical patterns of mesozooplankton functional diversity in the northwestern Pacific. Progress in Oceanography, 221(??):??, February 2024. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661124000016.

Garcia-Comas:2014:MSS

[GCCY^{+}14] Carmen García-Comas, Chun-Yi Chang, Lin Ye, Akash R. Sastri, Yu-Ching Lee, Gwo-Ching Gong, and Chih hao Hsieh. Mesozooplankton size structure in response to environmental conditions in the East China Sea: How much does size spectra theory fit empirical data of a dynamic coastal area? Progress in Oceanography, 121(??):141–157, February 2014. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661113001961.

Gross:1997:ELP

[GCD97] Richard S. Gross, Ben F. Chao, and Shailen D. Desai. Effect of long-period ocean tides on the Earth’s polar motion. Progress in Oceanography, 40(??):385–397, ???? 1997. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661198000093.

Guerzoni:1999:RAD

[GCD^{+}99] Stefano Guerzoni, Roy Chester, François Dulac, Barak Herut, Marie-Dominique Loÿe-Pilot, Chris Measures, Christophe Migon, Emanuela Molinaroli, Cyril Moulin, Paolo Rossini, Cemal Saydam, Alexandre Soudine, and Patrizia Ziveri. The role of atmospheric deposition in the biogeochemistry of the Mediterranean Sea. Progress in Oceanography, 44(1–3):147– 190, August 1999. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661199000245.

Gonzalez:2013:LOG

[GCD^{+}13] H. E. González, L. R. Castro, G. Daneri, J. L. Iriarte, N. Silva, F. Tapia, E. Teca, and C. A. Vargas. Land–ocean gradient in haline stratification and its effects on plankton dynamics and trophic carbon fluxes in Chilean Patagonian fjords (47– 50^{o}S). Progress in Oceanography, 119(??):32–47, December 2013. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661113000840.

Gatti:2018:BCA

[GCD^{+}18] Paul Gatti, Louise Cominassi, Erwan Duhamel, Patrick Grellier, Hervé Le Delliou, Sophie Le Mestre, Pierre Petitgas, Manuella Rabiller, Jérôme Spitz, and Martin Huret. Bioenergetic condition of anchovy and sardine in the Bay of Biscay and English Channel. Progress in Oceanography, 166(??):129–138, September 2018. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661116302634.

Grados:2022:UOH

[GCED22] Carmen Grados, Alexis Chaigneau, Vincent Echevin, and Noel Dominguez. Upper ocean hydrology of the Northern Humboldt Current System at seasonal, interannual and interdecadal scales. Progress in Oceanography, 165(??):123–144, ???? 2022. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661116302026.

Giusti:2019:CFD

[GCF^{+}19] M. Giusti, S. Canese, M. Fourt, M. Bo, C. Innocenti, A. Goujard, B. Daniel, L. Angeletti, M. Taviani, L. Aquilina, and L. Tunesi. Coral forests and Derelict Fishing Gears in submarine canyon systems of the Ligurian Sea. Progress in Oceanography, 178(??):??, November 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661118300685.

Grebmeier:2006:EDP

[GCFS06] Jacqueline M. Grebmeier, Lee W. Cooper, Howard M. Feder, and Boris I. Sirenko. Ecosystem dynamics of the Pacificinfluenced Northern Bering and Chukchi Seas in the amerasian Arctic. Progress in Oceanography, 71(2–4):331–361, October 2006. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661106001303.

Greer:2014:RIW

[GCG^{+}14] Adam T. Greer, Robert K. Cowen, Cedric M. Guigand, Jonathan A. Hare, and Dorothy Tang. The role of internal waves in larval fish interactions with potential predators and prey. Progress in Oceanography, 127(??):47–61, September 2014. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661114000962.

Gambi:2019:BDM

[GCLD19] Cristina Gambi, Laura Carugati, Marco Lo Martire, and Roberto Danovaro. Biodiversity and distribution of meiofauna in the Gioia, Petrace and Dohrn Canyons (Tyrrhenian Sea). Progress in Oceanography, 171(??):162–174, February 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661118300880.

Guenette:2008:TMP

[GCP08] Sylvie Guénette, Villy Christensen, and Daniel Pauly. Trophic modelling of the Peruvian upwelling ecosystem: Towards reconciliation of multiple datasets. Progress in Oceanography, 79(2–4):326–335, October 2008. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661108001857.

Grundlingh:1991:CWP

[GCS91] M. L. Gründlingh, R. A. Carter, and R. C. Stanton. Circulation and water properties of the southwest Indian Ocean, spring 1987. Progress in Oceanography, 28(4):305–342, ???? 1991. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/007966119190031G.

Ghosh:2024:RIO

[GCV^{+}24] Jayashree Ghosh, Kunal Chakraborty, Vinu Valsala, Trishneeta Bhattacharya, and Prasanna Kanti Ghoshal. A review of the Indian Ocean carbon dynamics, acidity, and productivity in a changing environment. Progress in Oceanography, 221(??): ??, February 2024. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661124000168.

Georgopoulos:2000:HCS

[GCZ^{+}00] D. Georgopoulos, G. Chronis, V. Zervakis, V. Lykousis, S. Poulos, and A. Iona. Hydrology and circulation in the Southern Cretan Sea during the CINCS experiment (May 1994– September 1995). Progress in Oceanography, 46(2–4):89–112, August 2000. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661100000148.

Gordon:1985:EDS

[GD85] J. D. M. Gordon and J. A. R. Duncan. The ecology of the deep-sea benthic and benthopelagic fish on the slopes of the Rockall Trough, Northeastern Atlantic. Progress in Oceanography, 15(1):37–69, ???? 1985. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0079661185900370.

Granata:1991:FMC

[GD91] T. C. Granata and T. D. Dickey. The fluid mechanics of copepod feeding in a turbulent flow: a theoretical approach. Progress in Oceanography, 26(3):243–261, ???? 1991. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0079661191900035.

Gonzalez:2009:CFW

[GDI^{+}09] Humberto E. González, Giovanni Daneri, José L. Iriarte, Beatriz Yannicelli, Eduardo Menschel, Claudio Barría, Silvio Pantoja, and Lorena Lizárraga. Carbon fluxes within the epipelagic zone of the Humboldt Current System off Chile: the significance of euphausiids and diatoms as key functional groups for the biological pump. Progress in Oceanography, 83(1–4):217–227, December 2009. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661109000895.

George:2015:BWB

[GDL^{+}15] John C. George, Matthew L. Druckenmiller, Kristin L. Laidre, Robert Suydam, and Brian Person. Bowhead whale body condition and links to summer sea ice and upwelling in the Beaufort Sea. Progress in Oceanography, 136(??):250–262, August 2015. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661115000981.

Guan:2015:CSL

[GDM^{+}15] Lu Guan, John F. Dower, Skip M. McKinnell, Pierre Pepin, Evgeny A. Pakhomov, and Brian P. V. Hunt. A comparison of spring larval fish assemblages in the Strait of Georgia (British Columbia, Canada) between the early 1980s and late 2000s. Progress in Oceanography, 138 (part A)(??):45–57, November 2015. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661115001962.

Ghinter:2020:MFS

[GDM^{+}20] Leopold Ghinter, Christine Dupuy, Michael J. Miller, Alexander Carpentier, Christel Lefrançois, Anthony Acou, Jun Aoyama, Mari Kuroki, Camilla Liénart, Shun Watanabe, Katsumi Tsukamoto, Tsuguo Otake, and Eric Feunteun. Microbial functional structure and stable isotopic variation of leptocephali across three current zones in the western South Pacific. Progress in Oceanography, 182(??):??, March 2020. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661120300021.

Giorli:2018:DSA

[GDN^{+}18] Giacomo Giorli, Jeffrey C. Drazen, Anna B. Neuheimer, Adrienne Copeland, and Whitlow W. L. Au. Deep sea animal density and size estimated using a Dual-frequency IDentification SONar (DIDSON) offshore the island of Hawaii. Progress in Oceanography, 160(??):155–166, January 2018. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661116301136.

Goes:2014:IAR

[GdRGC^{+}14] Joaquim I. Goes, Helga do Rosario Gomes, Alexander M. Chekalyuk, Edward J. Carpenter, Joseph P. Montoya, Victoria J. Coles, Patricia L. Yager, William M. Berelson, Douglas G. Capone, Rachel A. Foster, Deborah K. Steinberg, Ajit Subramaniam, and Mark A. Hafez. Influence of the Amazon River discharge on the biogeography of phytoplankton communities in the western tropical north Atlantic. Progress in Oceanography, 120(??):29–40, January 2014. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661113001237.

Goes:2001:NRI

[GdRGL^{+}01] Joaquim I. Goes, Helga do R. Gomes, Atsamon Limsakul, William M. Balch, and Toshiro Saino. El Niño related interannual variations in biological production in the North Pacific as evidenced by satellite and ship data. Progress in Oceanography, 49(??):211–225, ???? 2001. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661101000234.

Gonzalez-Davila:2009:SVC

[GDSCU09] Melchor González-Dávila, J. Magdalena Santana-Casiano, and Ivan R. Ucha. Seasonal variability of f CO_{2} in the Angola– Benguela region. Progress in Oceanography, 83(1–4):124–133, December 2009. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661109001062.

Glessmer:2009:COM

[GEO09] Mirjam S. Glessmer, Carsten Eden, and Andreas Oschlies. Contribution of oxygen minimum zone waters to the coastal upwelling off Mauritania. Progress in Oceanography, 83(1–4):143–150, December 2009. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661109001050.

Gutierrez:2008:OEC

[GEP^{+}08] Dimitri Gutiérrez, E. Enríquez, S. Purca, L. Quipúzcoa, R. Marquina, G. Flores, and M. Graco. Oxygenation episodes on the continental shelf of central Peru: Remote forcing and benthic ecosystem response. Progress in Oceanography, 79(2–4):177–189, October 2008. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661108001717.

Gutierrez-Estrada:2015:AST

[GEPC15] Juan Carlos Gutiérrez-Estrada and Inmaculada Pulido-Calvo. Is the Atlantic surface temperature a good proxy for forecasting the recruitment of European eel in the Guadalquivir estuary? Progress in Oceanography, 130(??):112–124, January 2015. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661114001700.

Glorioso:1997:PST

[GF97] Pablo D. Glorioso and Roger A. Flather. The Patagonian Shelf tides. Progress in Oceanography, 40(??):263–283, ???? 1997. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661198000044.

Guallar:2019:LPP

[GF19] Carles Guallar and Jordi Flos. Linking phytoplankton primary production and chromophoric dissolved organic matter in the sea. Progress in Oceanography, 176(??):??, September 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661117303439.

Granger:2015:LSS

[GFB^{+}15a] Victoria Granger, Jean-Marc Fromentin, Nicolas Bez, Giulio Relini, Christine N. Meynard, Jean-Claude Gaertner, Porzia Maiorano, Cristina Garcia Ruiz, Cristina Follesa, Michele Gristina, Panagiota Peristeraki, Anik Brind’Amour, Pierluigi Carbonara, Charis Charilaou, Antonio Esteban, Angélique Jadaud, Aleksandar Joksimovic, Argyris Kallianiotis, Jerina Kolitari, Chiara Manfredi, Enric Massuti, Roberta Mifsud, Antoni Quetglas, Wahid Refes, Mario Sbrana, Nedo Vrgoc, Maria Teresa Spedicato, and Bastien Mérigot. Largescale spatio-temporal monitoring highlights hotspots of demersal fish diversity in the Mediterranean Sea. Progress in Oceanography, 130(??):65–74, January 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S0079661114001657. See corrigendum [GFB^{+}15b].

Granger:2015:CLS

[GFB^{+}15b] Victoria Granger, Jean-Marc Fromentin, Nicolas Bez, Giulio Relini, Christine N. Meynard, Jean-Claude Gaertner, Porzia Maiorano, Cristina Garcia Ruiz, Cristina Follesa, Michele Gristina, Panagiota Peristeraki, Anik Brind’Amour, Pierluigi Carbonara, Charis Charilaou, Antonio Esteban, Angélique Jadaud, Aleksandar Joksimovic, Argyris Kallianiotis, Jerina Kolitari, Chiara Manfredi, Enric Massuti, Roberta Mifsud, Antoni Quetglas, Wahid Refes, Mario Sbrana, Nedo Vrgoc, Maria Teresa Spedicato, and Bastien Mérigot. Corrigendum to “Large-scale spatio-temporal monitoring highlights hotspots of demersal fish diversity in the Mediterranean Sea” [Prog. Oceanogr. 130 (2015) 65–74]. Progress in Oceanography, 138 (part A)(??):322–323, November 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661115001676. See [GFB^{+}15a].

Garcia-Fernandez:2023:IRE

[GFGGD^{+}23] Franklin García-Fernández, Jaime Gómez-Gutiérrez, Roxana De Silva-Dávila, Cristian Hakspiel-Segura, Israel Ambriz-Arreola, Aída Martínez-López, Arturo R. Sánchez-Uvera, Martín E. Hernández-Rivas, and Carlos J. Robinson. Interannual response of euphausiid community abundance during the anomalous warming period (2014–2016) in the Gulf of California. Progress in Oceanography, 212(??):??, March 2023. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S007966112300037X.

Gremare:2005:STC

[GGA^{+}05] Antoine Grémare, Dimitri Gutiérrez, Pierre Anschutz, Jean Michel Amouroux, Bruno Deflandre, and Gilles Vétion. Spatiotemporal changes in totally and enzymatically hydrolyzable amino acids of superficial sediments from three contrasted areas. Progress in Oceanography, 65(1):89–111, April 2005. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661105000388.

Grinyo:2016:DDP

[GGA^{+}16] Jordi Grinyó, Andrea Gori, Stefano Ambroso, Ariadna Purroy, Clara Calatayud, Carlos Dominguez-Carrió, Martina Coppari, Claudio Lo Iacono, Pablo J. López-González, and Josep-Maria Gili. Diversity, distribution and population size structure of deep Mediterranean gorgonian assemblages (Menorca Channel, Western Mediterranean Sea). Progress in Oceanography, 145(??):42–56, June 2016. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661115300446.

Gonzalez-Garcia:2023:BSV

[GGAA^{+}23] Cristina González-García, Susana Agustí, Jim Aiken, Arnaud Bertrand, Gabriel Bittencourt Farias, Antonio Bode, Claire Carré, Rafael Gonçalves-Araujo, Derek S. Harbour, María Huete-Ortega, Pedro A. M. C. Melo, Enrique Moreno-Ostos, Andrew P. Rees, Jaime Rodríguez, Sonia da Silva, Mikhail Zubkov, and Emilio Marañón. Basin-scale variability in phytoplankton size-abundance spectra across the Atlantic Ocean. Progress in Oceanography, 217(??):??, September– October 2023. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661123001477.

Groot:1965:SPD

[GGE^{+}65] J. J. Groot, C. R. Groot, M. Ewing, L. Burckle, and J. R. Conolly. Spores, pollen, diatoms and provenance of the Argentine Basin sediments. Progress in Oceanography, 4(??): 179–217, September 2, 1965. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0079661165900509.

Grinyo:2018:MAC

[GGG^{+}18] Jordi Grinyó, Andrea Gori, Michael Greenacre, Susana Requena, Antonio Canepa, Claudio Lo Iacono, Stefano Ambroso, Ariadna Purroy, and Josep-Maria Gili. Megabenthic assemblages in the continental shelf edge and upper slope of the Menorca Channel, Western Mediterranean Sea. Progress in Oceanography, 162(??):40–51, March 2018. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661116302336.

Gilbert:2010:MDS

[GGJ^{+}10] C. S. Gilbert, W. C. Gentleman, C. L. Johnson, C. DiBacco, J. M. Pringle, and C. Chen. Modelling dispersal of sea scallop (Placopecten magellanicus) larvae on Georges Bank: the influence of depth-distribution, planktonic duration and spawning seasonality. Progress in Oceanography, 87(1–4):37–48, October 2010. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661110001333.

Gonzalez-Galisteo:2019:CNP

[GGPG^{+}19] S. González-Galisteo, T. T. Packard, M. Gómez, A. Herrera, R. C. Dugdale, F. P. Wilkerson, R. T. Barber, D. Blasco, J. P. Christensen, and L. A. Codispoti. Calculating new production from nitrate reductase activity and light in the Peru current upwelling. Progress in Oceanography, 173(??):78–85, April 2019. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661117302732.

Gonzalez:2007:SPA

[GGQ07] Rodrigo R. González, Marcelo H. Gutiérrez, and Renato A. Quiñones. Similar potential ATP-P production and enzymatic activities in the microplankton community off Concepción (Chile) under oxic and suboxic conditions. Progress in Oceanography, 75(3):550–560, November 2007. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661107001681.

Grinyo:2020:SCA

[GGSM^{+}20] Jordi Grinyó, Alba Garriga, Anna Soler-Membrives, Andreu Santín, Stefano Ambroso, Pablo J. López-González, and David Díaz. Soft corals assemblages in deep environments of the Menorca Channel (Western Mediterranean Sea). Progress in Oceanography, 188(??):??, October 2020. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661120301749.

Garzoli:2015:SSU

[GGT^{+}15] Laura Garzoli, Giorgio Gnavi, Federica Tamma, Solveig Tosi, Giovanna C. Varese, and Anna M. Picco. Sink or swim: Updated knowledge on marine fungi associated with wood substrates in the Mediterranean Sea and hints about their potential to remediate hydrocarbons. Progress in Oceanography, 137 (part A)(??):140–148, September 2015. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661115001330.

Goetze:2017:EDB

[GHC^{+}17] Erica Goetze, Patricia T. Hüdepohl, Chantel Chang, Lauren Van Woudenberg, Matthew Iacchei, and Katja T. C. A. Peijnenburg. Ecological dispersal barrier across the equatorial Atlantic in a migratory planktonic copepod. Progress in Oceanography, 158(??):203–212, November 2017. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661115300525.

Gao:2021:ODI

[GHF^{+}21] Shuang Gao, Solfrid Sætre Hjøllo, Tone Falkenhaug, Espen Strand, Martin Edwards, and Morten D. Skogen. Overwintering distribution, inflow patterns and sustainability of Calanus finmarchicus in the North Sea. Progress in Oceanography, 194(??):??, June 2021. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661121000549.

Gu:2024:STV

[GHG^{+}24] Yuanyuan Gu, Mark James Hopwood, Martha Gledhill, Insa Rapp, Kathrin Wuttig, and Eric Pieter Achterberg. Spatial and temporal variations in the micronutrient Fe across the Peruvian shelf from 1984 to 2017. Progress in Oceanography, 221(??):??, February 2024. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661124000144.

Goffart:2015:DWS

[GHL15] Anne Goffart, Jean-Henri Hecq, and Louis Legendre. Drivers of the winter–spring phytoplankton bloom in a pristine NW Mediterranean site, the Bay of Calvi (Corsica): a long-term study (1979–2011). Progress in Oceanography, 137 (part A)(??):121–139, September 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661115001329.

Grischenko:2019:FRA

[GHSC19] Andrei V. Grischenko, Masato Hirose, Thomas Schwaha, and Alexei V. Chernyshev. First record of an abyssal and hadal bryozoan fauna from the Kuril–Kamchatka Trench.

Progress in Oceanography, 176(??):??, September 2019. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661119301545.

Giesecke:2019:DSP

[GHVG19] R. Giesecke, J. Höfer, T. Vallejos, and H. E. González. Death in southern Patagonian fjords: Copepod community structure and mortality in land- and marine-terminating glacierfjord systems. Progress in Oceanography, 174(??):162–172, May 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S007966111830065X.

Gjosaeter:2020:ASL

[GIC20] Harald Gjøsæter, Randi Ingvaldsen, and Jørgen S. Christiansen. Acoustic scattering layers reveal a faunal connection across the Fram Strait. Progress in Oceanography, 185(??): ??, June 2020. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661120300872.

Gifford:1993:PDN

[Gif93] Dian J. Gifford. Protozoa in the diets of Neocalanus spp. in the oceanic subarctic Pacific Ocean. Progress in Oceanography, 32(??):223–237, ???? 1993. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0079661193900156.

Gjosaeter:2023:FAY

[GIHJ23] Harald Gjøsæter, Randi B. Ingvaldsen, Elvar H. Hallfredsson, and Edda Johannesen. Fish assemblages at the Yermak Plateau and in northern Svalbard waters during the period 2012–2020. Progress in Oceanography, 219(??):??, December 2023. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661123001994.

Guo:2022:EIE

[GiIKX22] Chenying Guo, Shin ichi Ito, Yasuhiro Kamimura, and Peng Xiu. Evaluating the influence of environmental factors on the early life history growth of chub mackerel (Scomber japonicus) using a growth and migration model. Progress in Oceanography, 206(??):??, August 2022. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL https://www.sciencedirect.com/science/
article/pii/S0079661122000829.

Garcia-Ibanez:2015:STT

[GIPC^{+}15] Maribel I. García-Ibáñez, Paula C. Pardo, Lidia I. Carracedo, Herlé Mercier, Pascale Lherminier, Aida F. Ríos, and Fiz F. Pérez. Structure, transports and transformations of the water masses in the Atlantic subpolar gyre. Progress in Oceanography, 135(??):18–36, June 2015. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661115000506.

Grinyo:2017:CTV

[GIPG17] Jordi Grinyó, Enrique Isla, Laura Peral, and Josep-Maria Gili. Composition and temporal variability of particle fluxes in an insular canyon of the northwestern Mediterranean Sea. Progress in Oceanography, 159(??):323–339, December 2017. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661116301823.

Giricheva:2015:AEM

[Gir15] Evgeniya Giricheva. Aggregation in ecosystem models and model stability. Progress in Oceanography, 134(??):190–196, May 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661115000270.

Gouretski:2000:SEC

[GJ00] V. V. Gouretski and K. Jancke. Systematic errors as the cause for an apparent deep water property variability: global analysis of the WOCE and historical hydrographic data. Progress in Oceanography, 48(4):337–402, ???? 2000. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661100000495.

Garavelli:2014:IAS

[GKC^{+}14] Lysel Garavelli, David Michael Kaplan, François Colas, Wolfgang Stotz, Beatriz Yannicelli, and Christophe Lett. Identifying appropriate spatial scales for marine conservation and management using a larval dispersal model: the case of Concholepas concholepas (loco) in Chile. Progress in Oceanography, 124(??): 42–53, May 2014. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661114000378.

Gebruk:2020:GSP

[GKR20] Andrey V. Gebruk, Antonina Kremenetskaia, and Greg W. Rouse. A group of species “ Psychropotes longicauda ” (Psychropotidae, Elasipodida, Holothuroidea) from the Kuril– Kamchatka Trench area (North-West Pacific). Progress in Oceanography, 180(??):??, January 2020. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661119304021.

Gower:2013:MDN

[GKS^{+}13] J. Gower, S. King, S. Statham, R. Fox, and E. Young. The Malaspina Dragon: a newly-discovered pattern of the early spring bloom in the Strait of Georgia, British Columbia, Canada. Progress in Oceanography, 115(??):181–188, August 2013. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S007966111300075X.

Gawarkiewicz:2006:LFP

[GL06] Glen Gawarkiewicz and Christopher A. Linder. Lagrangian flow patterns north of Cape Hatteras using near-surface drifters. Progress in Oceanography, 70(2–4):181–195, August 2006. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661106000851.

Gupta:2023:BLC

[GL23] Abhinav Gupta and Pierre F. J. Lermusiaux. Bayesian learning of coupled biogeochemical–physical models. Progress in Oceanography, 216(??):??, August 2023. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0079661123000939.

Gomez-Letona:2022:DOP

[GLAHH^{+}22] Markel Gómez-Letona, Javier Arístegui, Nauzet Hernández-Hernández, Xosé Antón Álvarez-Salgado, Marta Álvarez, Erick Delgadillo, María Pérez-Lorenzo, Eva Teira, Santiago Hernández-León, and Marta Sebastián. Deep ocean prokaryotes and fluorescent dissolved organic matter reflect the history of the water masses across the Atlantic Ocean. Progress in Oceanography, 205(??):??, July 2022. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661122000805.

Ghyoot:2017:IMB

[GLF^{+}17] Caroline Ghyoot, Christiane Lancelot, Kevin J. Flynn, Aditee Mitra, and Nathalie Gypens. Introducing mixotrophy into a biogeochemical model describing an eutrophied coastal ecosystem: the Southern North Sea. Progress in Oceanography, 157(??):1–11, September 2017. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S007966111630235X.

Gregr:2013:SFR

[GLH13] Edward J. Gregr, Joanne Lessard, and John Harper. A spatial framework for representing nearshore ecosystems. Progress in Oceanography, 115(??):189–201, August 2013. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661113000797.

Gypens:2022:SIA

[GLLB22] N. Gypens, G. Lacroix, C. Lancelot, and A. V. Borges. Seasonal and inter-annual variability of air–sea CO_{2} fluxes and seawater carbonate chemistry in the Southern North Sea. Progress in Oceanography, 88(1–4):59–77, ???? 2022. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661110001515.

Gong:2023:NIG

[GLPC23] Yankun Gong, Lan Li, Rich Pawlowicz, and Shuqun Cai. Numerical investigation on the generation and evolution of nonlinear internal waves in the southern Strait of Georgia. Progress in Oceanography, 215(??):??, July 2023. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661123000964.

Gypens:2008:SPD

[GLS08] Nathalie Gypens, Christiane Lancelot, and Karline Soetaert. Simple parameterisations for describing N and P diagenetic processes: Application in the North Sea. Progress in Oceanography, 76(1):89–110, January 2008. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S007966110700198X.

Guilini:2012:CSO

[GLV12] Katja Guilini, Lisa A. Levin, and Ann Vanreusel. Cold seep and oxygen minimum zone associated sources of margin heterogeneity affect benthic assemblages, diversity and nutrition at the Cascadian margin (NE Pacific Ocean). Progress in Oceanography, 96(1):77–92, April 2012. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661111001182.

Gao:2023:OEB

[GLY23] Xu-Xu Gao, Heng-Chang Liu, and Gui-Peng Yang. Occurrence and emissions of biogenic dimethylated sulfur compounds and their responses to mesoscale eddies and typhoon Yutu in the western tropical Pacific Ocean. Progress in Oceanography, 219(??):??, December 2023. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661123001945.

Golovan:2019:NDS

[GM19] Olga A. Golovan and Marina V. Malyutina. A new deep-sea genus of Paramunnidae (Isopoda: Asellota) from the Northwest Pacific adjacent to the Kuril–Kamchatka Trench with remarks on paedomorphosis in deep-sea species. Progress in Oceanography, 178(??):??, November 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661119301971.

Gonzalez:2007:STV

[GMAB07] Humberto E. González, Eduardo Menschel, Carola Aparicio, and Claudio Barría. Spatial and temporal variability of microplankton and detritus, and their export to the shelf sediments in the upwelling area off Concepción, Chile (≈ 36^{o}S), during 2002–2005. Progress in Oceanography, 75(3):435–451, November 2007. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661107001619.

Garcia-Martin:2017:CBR

[GMAGH^{+}17] E. E. García-Martín, M. Aranguren-Gassis, M. Hartmann, M. V. Zubkov, and P. Serret. Contribution of bacterial respiration to plankton respiration from 50^{o}N to 44^{o}S in the Atlantic Ocean. Progress in Oceanography, 158(??):99–108, November 2017. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661115300513.

Garcia-Munoz:2004:DTO

[GMAMB04] Mercedes García-Muñoz, Javier Arístegui, María F. Montero, and Eric D. Barton. Distribution and transport of organic matter along a filament-eddy system in the canaries — NW Africa coastal transition zone region. Progress in Oceanography, 62(2–4):115–129, August 2004. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661104000941.

Galan:2012:TVP

[GMBU12] Alexander Galán, Verónica Molina, Lucy Belmar, and Osvaldo Ulloa. Temporal variability and phylogenetic characterization of planktonic anammox bacteria in the coastal upwelling ecosystem off central Chile. Progress in Oceanography, 92–95(??):110–120, January 2012. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661111000619.

Gilly:2022:LTH

[GMD^{+}22] William Gilly, Unai Markaida, Patrick Daniel, Tim Frawley, Carlos Robinson, Jaime Gómez-Gutiérrez, Dylan Hyun, Jacob Soliman, Puneeta Pandey, and Lorenzo Rosenzweig. Long-term hydrographic changes in the Gulf of California and ecological impacts: a crack in the World’s Aquarium? Progress in Oceanography, 206(??):??, August 2022. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661122001161.

Garcia-Martin:2022:SCP

[GMDD^{+}22a] E. Elena García-Martín, Chris J. Daniels, Keith Davidson, Clare E. Davis, Claire Mahaffey, Kyle M. J. Mayers, Sharon McNeill, Alex J. Poulton, Duncan A. Purdie, Glen A. Tarran, and Carol Robinson. Seasonal changes in plankton respiration and bacterial metabolism in a temperate shelf sea. Progress in Oceanography, 177(??):??, October 2022. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661117301283.

Garcia-Martin:2022:PCR

[GMDD^{+}22b] E. Elena García-Martín, Chris J. Daniels, Keith Davidson, Jose Lozano, Kyle M. J. Mayers, Sharon McNeill, Elaine Mitchell, Alex J. Poulton, Duncan A. Purdie, Glen A. Tarran, Callum Whyte, and Carol Robinson. Plankton community respiration and bacterial metabolism in a North Atlantic Shelf Sea during spring bloom development (April 2015). Progress in Oceanography, 177(??):??, October 2022. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661117301234.

Guillaumot:2020:ESD

[GMDS20] Charlène Guillaumot, Camille Moreau, Bruno Danis, and Thomas Saucède. Extrapolation in species distribution modelling. Application to Southern Ocean marine species. Progress in Oceanography, 188(??):??, October 2020. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661120301774.

Gao:2023:CIL

[GML^{+}23] Zihui Gao, Shuyang Ma, Jianchao Li, Peng Sun, Yang Liu, Qinwang Xing, Qingsong He, and Yongjun Tian. Climateinduced long-term variations of the Arctic ecosystems. Progress in Oceanography, 213(??):??, April–May 2023. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661123000496.

Goldstein:2023:LSI

[GMR^{+}23] Esther D. Goldstein, Ryan M. McCabe, Matthew C. Rogers, Alison L. Deary, and Janet T. Duffy-Anderson. Loss of sea ice and intermittent winds alter distributions and diet resources of young forage fish in the Chukchi Sea. Progress in Oceanography, 217(??):??, September–October 2023. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661123001404.

Gomez:2019:BOP

[GNH19] Iván Gómez, Nelso P. Navarro, and Pirjo Huovinen. Biooptical and physiological patterns in Antarctic seaweeds: a functional trait based approach to characterize vertical zonation. Progress in Oceanography, 174(??):17–27, May 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661117303798.

Gorsline:1992:GSS

[Gor92] D. S. Gorsline. The geological setting of Santa Monica and San Pedro Basins, California Continental Borderland. Progress in Oceanography, 30(1–4):1–36, ???? 1992. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/007966119290008N.

Gould:1985:POA

[Gou85] W. J. Gould. Physical oceanography of the Azores front. Progress in Oceanography, 14(??):167–190, ???? 1985. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0079661185900102.

Gastaud:2011:TSP

[GPA^{+}11] J. Gastaud, P. P. Povinec, M. Aoyama, K. Hirose, J. A. Sanchez-Cabeza, I. Levy, P. Roos, M. Eriksson, E. Bosc, and S. Rezzoug. Transport and scavenging of Pu in surface waters of the Southern Hemisphere Oceans. Progress in Oceanography, 89(1–4):92–100, April 2011. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661110001618.

Gomes-Pereira:2016:CFT

[GPAB^{+}16] Jose Nuno Gomes-Pereira, Vincent Auger, Kolja Beisiegel, Robert Benjamin, Melanie Bergmann, David Bowden, Pal Buhl-Mortensen, Fabio C. De Leo, Gisela Dionísio, Jennifer M. Durden, Luke Edwards, Ariell Friedman, Jens Greinert, Nancy Jacobsen-Stout, Steve Lerner, Murray Leslie, Tim W. Nattkemper, Jessica A. Sameoto, Timm Schoening, Ronald Schouten, James Seager, Hanumant Singh, Olivier Soubigou, Inês Tojeira, Inge van den Beld, Frederico Dias, Fernando Tempera, and Ricardo S. Santos. Current and future trends in marine image annotation software. Progress in Oceanography, 149(??):106–120, December 2016. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661116301240.

Greene:2003:TAR

[GPC^{+}03] C. H. Greene, A. J. Pershing, A. Conversi, B. Planque, C. Hannah, D. Sameoto, E. Head, P. C. Smith, P. C. Reid, J. Jossi, D. Mountain, M. C. Benfield, P. H. Wiebe, and E. Durbin. Trans-Atlantic responses of Calanus finmarchicus populations to basin-scale forcing associated with the North Atlantic oscillation. Progress in Oceanography, 58(2–4):301–312, August 2003. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661103001344.

Gasser:2017:TMO

[GPE^{+}17] Marc Gasser, Josep L. Pelegrí, Mikhail Emelianov, Miguel Bruno, Eulàlia Gràcia, Marcos Pastor, Hartmut Peters, Ángel Rodríguez-Santana, Joaquín Salvador, and Ricardo F. Sánchez-Leal. Tracking the Mediterranean outflow in the Gulf of Cadiz. Progress in Oceanography, 157(??):47–71, September 2017. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661116301525.

Galvez:2020:IEO

[GPEV20] Casandra Gálvez, Mario A. Pardo, and Fernando R. Elorriaga-Verplancken. Impacts of extreme ocean warming on the early development of a marine top predator: the Guadalupe fur seal. Progress in Oceanography, 180(??):??, January 2020. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661119304008.

George:2022:BCC

[GPP22] Kai Horst George, Karin Pointner, and Jana Packmor. The benthic Copepoda (Crustacea) of Anaximenes Seamount (eastern Mediterranean Sea) — community structure and species distribution. Progress in Oceanography, 165(??):299–316, ???? 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661117302458.

Gascard:1985:WMC

[GR85] J. C. Gascard and C. Richez. Water masses and circulation in the Western Alboran Sea and in the Straits of Gibraltar. Progress in Oceanography, 15(3):157–216, ???? 1985. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/007966118590031X.

Gasparin:2017:SME

[GR17] Florent Gasparin and Dean Roemmich. The seasonal march of the equatorial Pacific upper-ocean and its El Niño variability. Progress in Oceanography, 156(??):1–16, August 2017. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661116301628.

Gutierrez:2008:ENA

[GRB^{+}08] Mariano Gutiérrez, Argiro Ramirez, Sophie Bertrand, Octavio Móron, and Arnaud Bertrand. Ecological niches and areas of overlap of the squat lobster ‘munida’ (Pleuroncodes monodon) and anchoveta (Engraulis ringens) off Peru. Progress in Oceanography, 79(2–4):256–263, October 2008. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S007966110800178X.

Grandremy:2023:HSP

[GRD^{+}23] Nina Grandremy, Jean-Baptiste Romagnan, Christine Dupuy, Mathieu Doray, Martin Huret, and Pierre Petitgas. Hydrology and small pelagic fish drive the spatio–temporal dynamics of springtime zooplankton assemblages over the Bay of Biscay continental shelf. Progress in Oceanography, 210(??):??, January 2023. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661122002087.

Gasalla:2010:CMF

[GRDS10] Maria A. Gasalla, Amanda R. Rodrigues, Luis F. A. Duarte, and U. Rashid Sumaila. A comparative multi-fleet analysis of socio-economic indicators for fishery management in SE Brazil. Progress in Oceanography, 87(1–4):304–319, October 2010. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661110001229.

Gutierrez-Rodriguez:2022:PPD

[GRdSS^{+}22] Andres Gutiérrez-Rodríguez, Adriana Lopes dos Santos, Karl Safi, Ian Probert, Fabrice Not, Denise Fernández, Priscillia Gourvil, Jaret Bilewitch, Debbie Hulston, Matt Pinkerton, and Scott D. Nodder. Planktonic protist diversity across contrasting subtropical and subantarctic waters of the southwest Pacific. Progress in Oceanography, 206(??):??, August 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661122000702.

Grischenko:2022:NDW

[Gri22] Andrei V. Grischenko. New deep-water Bryozoa from the Magellan Seamounts, tropical western Pacific, with a global review of bryozoans associated with ferromanganese crusts. Progress in Oceanography, 205(??):??, July 2022. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S007966112200088X.

Garcia-Reyes:2014:SSU

[GRLS14] Marisol García-Reyes, John L. Largier, and William J. Sydeman. Synoptic-scale upwelling indices and predictions of phyto- and zooplankton populations. Progress in Oceanography, 120(??):177–188, January 2014. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S007966111300164X.

Gardner:2018:GCB

[GRMB18] Wilford D. Gardner, Mary Jo Richardson, Alexey V. Mishonov, and Pierre E. Biscaye. Global comparison of benthic nepheloid layers based on 52 years of nephelometer and transmissometer measurements. Progress in Oceanography, 168(??):100–111, November 2018. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661118301290.

Gregoire:2008:NMC

[GRS08] M. Grégoire, C. Raick, and K. Soetaert. Numerical modeling of the central Black Sea ecosystem functioning during the eutrophication phase. Progress in Oceanography, 76(3):286– 333, March 2008. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661108000037.

George:2019:NSE

[GS19] Kai Horst George and Enrico Schwabe. A new species of Ectinosomatidae sars (Copepoda, Harpacticoida) associated with Pseudoikedella achaeta (Zenkevitch, 1958) (Echiura, Bonelliida). Progress in Oceanography, 173(??):1–8, April 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661118303422.

Gjelstrup:2024:STH

[GS24] Caroline V. B. Gjelstrup and Colin A. Stedmon. A switch in thermal and haline contributions to stratification in the Greenland Sea during the last four decades.

Progress in Oceanography, 225(??):??, July 2024. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661124000892.

Girard:2020:CTD

[GSA^{+}20] Fanny Girard, Jozée Sarrazin, Aurélien Arnaubec, Mathilde Cannat, Pierre-Marie Sarradin, Benjamin Wheeler, and Marjolaine Matabos. Currents and topography drive assemblage distribution on an active hydrothermal edifice. Progress in Oceanography, 187(??):??, August–September 2020. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661120301361.

Grigor:2020:NCF

[GSC^{+}20] Jordan J. Grigor, Moritz S. Schmid, Marianne Caouette, Vicky St.-Onge, Thomas A. Brown, and Roxane-M. Barthélémy. Non-carnivorous feeding in Arctic chaetognaths. Progress in Oceanography, 186(??):??, July 2020. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661120301270.

Guallart:2015:TAC

[GSF^{+}15] Elisa F. Guallart, Ute Schuster, Noelia M. Fajar, Oliver Legge, Peter Brown, Carles Pelejero, Marie-Jose Messias, Eva Calvo, Andrew Watson, Aida F. Ríos, and Fiz F. Pérez. Trends in anthropogenic CO_{2} in water masses of the Subtropical North Atlantic Ocean. Progress in Oceanography, 131(??):21–32, February 2015. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661114001815.

Gaichas:2009:CCT

[GSFP^{+}09] Sarah Gaichas, Georg Skaret, Jannike Falk-Petersen, Jason S. Link, William Overholtz, Bernard A. Megrey, Harald Gjøsæter, William T. Stockhausen, Are Dommasnes, Kevin D. Friedland, and Kerim Aydin. A comparison of community and trophic structure in five marine ecosystems based on energy budgets and system metrics. Progress in Oceanography, 81(1–4):47– 62, April 2009. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661109000226.

Garrido:2017:TFM

[GSM^{+}17] Susana Garrido, Alexandra Silva, Vitor Marques, Ivone Figueiredo, Philippe Bryère, Antoine Mangin, and A. Miguel P. Santos. Temperature and food-mediated variability of European Atlantic sardine recruitment. Progress in Oceanography, 159(??):267–275, December 2017. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661116301768.

Gotsis-Skretas:1999:SHV

[GSPMAI99] Olympia Gotsis-Skretas, Kalliopi Pagou, Maria Moraitou-Apostolopoulou, and Lydia Ignatiades. Seasonal horizontal and vertical variability in primary production and standing stocks of phytoplankton and zooplankton in the Cretan Sea and the Straits of the Cretan Arc (March 1994–January 1995). Progress in Oceanography, 44(4):625–649, December 1999. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661199000488.

Garcia-Solsona:2020:REE

[GSPP^{+}20] Ester Garcia-Solsona, Leopoldo D. Pena, Eduardo Paredes, José N. Pérez-Asensio, Lucía Quirós-Collazos, Fabrizio Lirer, and Isabel Cacho. Rare earth elements and Nd isotopes as tracers of modern ocean circulation in the central Mediterranean Sea. Progress in Oceanography, 185(??):??, June 2020. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661120300793.

Gorska:2020:BTB

[GSSWK20] Barbara Górska, Thomas Soltwedel, Ingo Schewe, and Maria Włodarska-Kowalczuk. Bathymetric trends in biomass size spectra, carbon demand, and production of Arctic benthos (76–5561 m, Fram Strait). Progress in Oceanography, 186(??): ??, July 2020. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661120301099.

Galeron:2001:TPA

[GSV^{+}01] J. Galéron, M. Sibuet, A. Vanreusel, K. Mackenzie, A. J. Gooday, A. Dinet, and G. A. Wolff. Temporal patterns among meiofauna and macrofauna taxa related to changes in sediment geochemistry at an abyssal NE Atlantic site. Progress in Oceanography, 50(??):303–324, ???? 2001. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661101000593.

Garcia-Seoane:2023:SUI

[GSVB23] R. García-Seoane, I. G. Viana, and A. Bode. Seasonal upwelling influence on trophic indices of mesozooplankton in a coastal food web estimated from δ ^{15}N in amino acids. Progress in Oceanography, 219(??):??, December 2023. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661123001921.

Gaudy:2007:MCS

[GTB07] Raymond Gaudy and Delphine Thibault-Botha. Metabolism of Centropages species in the Mediterranean Sea and the North Atlantic Ocean. Progress in Oceanography, 72(2–3):151–163, February 2007. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661107000109.

Grabon:2021:AAW

[GTNK21] Jeffrey S. Grabon, John M. Toole, An T. Nguyen, and Richard A. Krishfield. An analysis of Atlantic water in the Arctic Ocean using the Arctic subpolar gyre state estimate and observations. Progress in Oceanography, 198(??):??, November 2021. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661121001695.

Greenwood:2001:BSD

[GTR01] J. E. Greenwood, V. W. Truesdale, and A. R. Rendell. Biogenic silica dissolution in seawater — in vitro chemical kinetics. Progress in Oceanography, 48(1):1–23, ???? 2001. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S007966110000046X.

Gruss:2021:SIV

[GTS^{+}21] Arnaud Grüss, James T. Thorson, Christine C. Stawitz, Jonathan C. P. Reum, Sean K. Rohan, and Cheryl L. Barnes. Synthesis of interannual variability in spatial demographic processes supports the strong influence of coldpool extent on eastern Bering Sea walleye pollock (Gadus chalcogrammus). Progress in Oceanography, 194(??):??, June 2021. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661121000562.

Garcia-Vernet:2021:ENP

[GVBV^{+}21] Raquel García-Vernet, Asunción Borrell, Gísli Víkingsson, Sverrir D. Halldórsson, and Alex Aguilar. Ecological niche partitioning between baleen whales inhabiting Icelandic waters. Progress in Oceanography, 199(??):??, December 2021. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661121001749.

Guilini:2013:LTV

[GVKD^{+}13] Katja Guilini, Gritta Veit-Köhler, Marleen De Troch, Dirk Van Gansbeke, and Ann Vanreusel. Latitudinal and temporal variability in the community structure and fatty acid composition of deep-sea nematodes in the Southern Ocean. Progress in Oceanography, 110(??):80–92, March 2013. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661113000037.

Garcia:2008:DRM

[GvOS^{+}08] R. García, D. van Oevelen, K. Soetaert, L. Thomsen, H. C. De Stigter, and E. Epping. Deposition rates, mixing intensity and organic content in two contrasting submarine canyons. Progress in Oceanography, 76(2):192–215, February 2008. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661108000062.

Gontikaki:2011:FWF

[GvOSW11] E. Gontikaki, D. van Oevelen, K. Soetaert, and U. Witte. Food web flows through a sub-arctic deep-sea benthic community. Progress in Oceanography, 91(3):245–259, November 2011. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661110001667.

Guoliang:1989:TTC

[GW89] Jin Guoliang and Peter Wadhams. Travel time changes in a tomography array caused by a sea ice cover. Progress in Oceanography, 22(3):249–275, ???? 1989. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0079661189900141.

Godfrey:1991:LCD

[GW91] J. Stuart Godfrey and Andrew J. Weaver. Is the Leeuwin Current driven by Pacific heating and winds? Progress in Oceanography, 27(3–4):225–272, ???? 1991. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/007966119190026I.

Garabato:2014:TDO

[GWB14] Alberto C. Naveira Garabato, Adam P. Williams, and Sheldon Bacon. The three-dimensional overturning circulation of the Southern Ocean during the WOCE era. Progress in Oceanography, 120(??):41–78, January 2014. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661113001316.

Gonzalez-Wevar:2019:CSS

[GWGR^{+}19] Claudio A. González-Wevar, Karin Gérard, Sebastian Rosenfeld, Thomas Saucède, Javier Naretto, Angie Díaz, Simon A. Morley, Paul Brickle, and Elie Poulin. Cryptic speciation in Southern Ocean Aequiyoldia eightsii (Jay, 1839): Mio-Pliocene trans-Drake Passage separation and diversification. Progress in Oceanography, 174(??):44–54, May 2019. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661118300508.

Wang:2020:RHA

[gWjNfLyD20] Liang gen Wang, Jia jia Ning, Ya fang Li, and Fei yan Du. Responses of hyperiid (Amphipoda) communities to monsoon reversal in the central South China Sea. Progress in Oceanography, 189(??):??, November–December 2020. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661120301798.

Gorska:2017:FDE

[GWK17] Barbara Górska and Maria Włodarska-Kowalczuk. Food and disturbance effects on Arctic benthic biomass and production size spectra. Progress in Oceanography, 152(??):50–61, March 2017. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661116300349.

Giering:2022:SVZ

[GWM^{+}22] Sarah L. C. Giering, Seona R. Wells, Kyle M. J. Mayers, Hanna Schuster, Louise Cornwell, Elaine S. Fileman, Angus Atkinson, Kathryn B. Cook, Calum Preece, and Daniel J. Mayor. Seasonal variation of zooplankton community structure and trophic position in the Celtic Sea: a stable isotope and biovolume spectrum approach. Progress in Oceanography, 177(??): ??, October 2022. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661118300399.

Gonzalez:2023:UTM

[GWS^{+}23] Nicolas M. Gonzalez, Robin Waldman, Gianmaria Sannino, Hervé Giordani, and Samuel Somot. Understanding tidal mixing at the Strait of Gibraltar: a high-resolution model approach. Progress in Oceanography, 212(??):??, March 2023. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S007966112300023X.

Gong:2022:SAN

[GXX^{+}22] Yankun Gong, Jieshuo Xie, Jiexin Xu, Zhiwu Chen, Yinghui He, and Shuqun Cai. Spatial asymmetry of nonlinear internal waves in the Lombok Strait. Progress in Oceanography, 202(??):??, March 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661122000210.

Guo:2023:RGA

[GZCL23] Junting Guo, Zihan Zhao, Ruichen Cao, and Xianqing Lv. Regional to global assessments of ocean transparency dynamics from 1997 to 2019. Progress in Oceanography, 219(??):??, December 2023. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661123002082.

Holt:2014:CIA

[HAA^{+}14] Jason Holt, J. Icarus Allen, Thomas R. Anderson, Robert Brewin, Momme Butenschön, James Harle, Geir Huse, Patrick Lehodey, Christian Lindemann, Laurent Memery, Baris Salihoglu, Inna Senina, and Andrew Yool. Challenges in integrative approaches to modelling the marine ecosystems of the North Atlantic: Physics to fish and coasts to ocean. Progress in Oceanography, 129(??):285–313, December 2014. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S007966111400072X.

Humphreys:2022:MNC

[HAH^{+}22] Matthew P. Humphreys, Eric P. Achterberg, Joanne E. Hopkins, Mohammed Z. H. Chowdhury, Alex M. Griffiths, Susan E. Hartman, Tom Hull, Angelina Smilenova, Juliane U. Wihsgott, E. Malcolm S. Woodward, and C. Mark Moore. Mechanisms for a nutrient-conserving carbon pump in a seasonally stratified, temperate continental shelf sea. Progress in Oceanography, 177(??):??, October 2022. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661117301374.

Hamilton:1987:SSG

[Ham87] Peter Hamilton. The structure of shelf and Gulf Stream motions in the Georgia Bight. Progress in Oceanography, 19(3–4):329–351, ???? 1987. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0079661187900139.

Hamilton:1990:MSC

[Ham90] Peter Hamilton. Modelling salinity and circulation for the Columbia River Estuary. Progress in Oceanography, 25(1– 4):113–156, ???? 1990. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/007966119090005M.

Hamilton:2009:TRW

[Ham09] Peter Hamilton. Topographic Rossby waves in the Gulf of Mexico. Progress in Oceanography, 82(1):1–31, July 2009. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661109000342.

Hobday:2016:HAD

[HAP^{+}16] Alistair J. Hobday, Lisa V. Alexander, Sarah E. Perkins, Dan A. Smale, Sandra C. Straub, Eric C. J. Oliver, Jessica A. Benthuysen, Michael T. Burrows, Markus G. Donat, Ming Feng, Neil J. Holbrook, Pippa J. Moore, Hillary A. Scannell, Alex Sen Gupta, and Thomas Wernberg. A hierarchical approach to defining marine heatwaves. Progress in Oceanography, 141(??):227–238, February 2016. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661116000057.

Hardy:1982:SSM

[Har82] J. T. Hardy. The sea surface microlayer: Biology, chemistry and anthropogenic enrichment. Progress in Oceanography, 11(4):307–328, ???? 1982. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0079661182900015.

Harcourt:2005:TCM

[Har05a] Ramsey R. Harcourt. Thermobaric cabbeling over maud rise: Theory and large eddy simulation. Progress in Oceanography, 67(1–2):186–244, October 2005. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661105001072.

Harrison:2005:E

[Har05b] Paul J. Harrison. Editorial. Progress in Oceanography, 64(2–4):91–93, February 2005. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661105000236.

Hasselmann:1982:OMC

[Has82] K. Hasselmann. An ocean model for climate variability studies. Progress in Oceanography, 11(2):69–92, ???? 1982. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0079661182900040.

Hastenrath:2006:CTM

[Has06] Stefan Hastenrath. Circulation and teleconnection mechanisms of Northeast Brazil droughts. Progress in Oceanography, 70(2–4):407–415, August 2006. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661106000565.

Haury:1984:OEC

[Hau84] L. R. Haury. An offshore eddy in the California Current System Part IV: Plankton distributions. Progress in Oceanography, 13(1):95–111, ???? 1984. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0079661184900077.

Hautala:2018:ADC

[Hau18] Susan L. Hautala. The abyssal and deep circulation of the Northeast Pacific Basin. Progress in Oceanography, 160(??):68– 82, January 2018. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661117300046.

Hays:1965:QSA

[Hay65] James D. Hays. Quaternary sediments of the Antarctic Ocean. Progress in Oceanography, 4(??):117–131, September 2, 1965. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661165900455.

Huret:2018:STV

[HBD^{+}18] Martin Huret, Paul Bourriau, Mathieu Doray, Francis Gohin, and Pierre Petitgas. Survey timing vs. ecosystem scheduling: Degree-days to underpin observed interannual variability in marine ecosystems. Progress in Oceanography, 166(??):30–40, September 2018. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661117300824.

Hartman:2021:EST

[HBD^{+}21] Susan E. Hartman, Brian J. Bett, Jennifer M. Durden, Stephanie A. Henson, Morten Iversen, Rachel M. Jeffreys, Tammy Horton, Richard Lampitt, and Andrew R. Gates. Enduring science: Three decades of observing the Northeast Atlantic from the Porcupine Abyssal Plain Sustained Observatory (PAP-SO). Progress in Oceanography, 191(??):??, February 2021. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661120302433.

Haug:2021:HSB

[HBG^{+}21] Tore Haug, Martin Biuw, Harald Gjøsæter, Tor Knutsen, Ulf Lindstrøm, Kirsteen M. MacKenzie, Sonnich Meier, and Kjell T. Nilssen. Harp seal body condition and trophic interactions with prey in Norwegian high Arctic waters in early autumn. Progress in Oceanography, 191(??):??, February 2021. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661120302330.

Howell:2017:VPC

[HBH^{+}17] Evan A. Howell, Steven J. Bograd, Aimee L. Hoover, Michael P. Seki, and Jeffrey J. Polovina. Variation in phytoplankton composition between two North Pacific frontal zones along 158^{o}W during winter–spring 2008–2011. Progress in Oceanography, 150(??):3–12, January 2017. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661115001366.

Haug:2024:TIB

[HBK^{+}24] T. Haug, M. Biuw, K. M. Kovacs, L. Lindblom, U. Lindstrøm, C. Lydersen, K. M. MacKenzie, and S. Meier. Trophic interactions between common minke whales (Balaenoptera acutorostrata) and their prey during summer in the northern Barents Sea. Progress in Oceanography, 224(??):??, June 2024. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661124000739.

Harrison:2013:PPG

[HBL^{+}13] W. Glen Harrison, K. Yngve Børsheim, William K. W. Li, Gary L. Maillet, Pierre Pepin, Egil Sakshaug, Morten D. Skogen, and Philip A. Yeats. Phytoplankton production and growth regulation in the Subarctic North Atlantic: a comparative study of the Labrador Sea–Labrador/Newfoundland shelves and Barents/Norwegian/Greenland seas and shelves. Progress in Oceanography, 114(??):26–45, July 2013. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661113000451.

Helaouet:2011:MCF

[HBR11] Pierre Helaouët, Gregory Beaugrand, and Philip Chris Reid. Macrophysiology of Calanus finmarchicus in the North Atlantic Ocean. Progress in Oceanography, 91(3):217–228, November 2011. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661110001503.

Harrison:1999:CFC

[HBV^{+}99] P. J. Harrison, P. W. Boyda, D. E. Varela, S. Takeda, A. Shiomoto, and T. Odate. Comparison of factors controlling phytoplankton productivity in the NE and NW subarctic Pacific gyres. Progress in Oceanography, 43(2–4):205–234, March 1999. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661199000154.

Harlay:2010:BSC

[HBV^{+}10] J. Harlay, A. V. Borges, C. Van Der Zee, B. Delille, R. H. M. Godoi, L.-S. Schiettecatte, N. Roevros, K. Aerts, P.-E. Lapernat, L. Rebreanu, S. Groom, M.-H. Daro, R. Van Grieken, and L. Chou. Biogeochemical study of a coccolithophore bloom in the northern Bay of Biscay (NE Atlantic Ocean) in June 2004. Progress in Oceanography, 86(3–4):317–336, September 2010. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661110000637.

Hood:2017:BEI

[HBW17] Raleigh R. Hood, Lynnath E. Beckley, and Jerry D. Wiggert. Biogeochemical and ecological impacts of boundary currents in the Indian Ocean. Progress in Oceanography, 156(??):290–325, August 2017. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661117301507.

Hernandez-Carrasco:2020:LFE

[HCAFD^{+}20] Ismael Hernández-Carrasco, Eva Alou-Font, Pierre-Antoine Dumont, Andrea Cabornero, John Allen, and Alejandro Orfila. Lagrangian flow effects on phytoplankton abundance and composition along filament-like structures. Progress in Oceanography, 189(??):??, November–December 2020. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661120302056.

Hopcroft:2002:CCM

[HCC02] R. R. Hopcroft, C. Clarke, and F. P. Chavez. Copepod communities in Monterey Bay during the 1997–1999 El Niño and La Niña. Progress in Oceanography, 54(1–4):251–264, July 2002. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661102000526.

Huwer:2011:VDG

[HCGK11] Bastian Huwer, Catriona Clemmesen, Peter Grønkjær, and Friedrich W. Köster. Vertical distribution and growth performance of Baltic cod larvae — field evidence for starvationinduced recruitment regulation during the larval stage? Progress in Oceanography, 91(4):382–396, December 2011. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S007966111100036X.

Horton:2020:MPD

[HCV^{+}20] Tammy Horton, Harry Cooper, Rianna Vlierboom, Michael Thurston, Chris Hauton, and C. Robert Young. Molecular phylogenetics of deep-sea amphipods (Eurythenes) reveal a new undescribed species at the Porcupine Abyssal Plain, North East Atlantic Ocean. Progress in Oceanography, 183(??): ??, April 2020. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661120300306.

Hunt:2016:APS

[HDA^{+}16] George L. Hunt, Kenneth F. Drinkwater, Kevin Arrigo, Jørgen Berge, Kendra L. Daly, Seth Danielson, Malin Daase, Haakon Hop, Enrique Isla, Nina Karnovsky, Kristin Laidre, Franz J. Mueter, Eugene J. Murphy, Paul E. Renaud, Walker O. Smith, Philip Trathan, John Turner, and Dieter Wolf-Gladrow. Advection in polar and sub-polar environments: Impacts on high latitude marine ecosystems. Progress in Oceanography, 149(??):40–81, December 2016. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661116302051.

Hummels:2013:SRV

[HDB13] Rebecca Hummels, Marcus Dengler, and Bernard Bourlès. Seasonal and regional variability of upper ocean diapycnal heat flux in the Atlantic cold tongue. Progress in Oceanography, 111(??):52–74, April 2013. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661112001607.

Harriague:2019:MAC

[HDM19] Anabella Covazzi Harriague, Roberto Danovaro, and Cristina Misic. Macrofaunal assemblages in canyon and adjacent slope of the NW and central Mediterranean systems. Progress in Oceanography, 171(??):38–48, February 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661118301010.

Hauser:2015:NAA

[HDZY15] Tristan Hauser, Entcho Demirov, Jieshun Zhu, and Igor Yashayaev. North Atlantic atmospheric and ocean inter-annual variability over the past fifty years — dominant patterns and decadal shifts. Progress in Oceanography, 132(??):197–219, March 2015. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661114001712.

Hidalgo:2007:CLC

[HE07] Pamela Hidalgo and Ruben Escribano. Coupling of life cycles of the copepods Calanus chilensis and Centropages brachiatus to upwelling induced variability in the central-southern region of Chile. Progress in Oceanography, 75(3):501–517, November 2007. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661107001656.

Heath:2012:ELF

[Hea12] Michael R. Heath. Ecosystem limits to food web fluxes and fisheries yields in the North Sea simulated with an end-to-end food web model. Progress in Oceanography, 102(??):42–66, September 2012. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661112000213.

Hidalgo:2012:HCU

[HEF^{+}12] Pamela Hidalgo, Ruben Escribano, Marcelo Fuentes, Erika Jorquera, and Odette Vergara. How coastal upwelling influences spatial patterns of size-structured diversity of copepods off central-southern Chile (summer 2009). Progress in Oceanography, 92–95(??):134–145, January 2012. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661111000668.

Hendershott:1973:IOT

[Hen73] M. C. Hendershott. Inertial oscillations of tidal period. Progress in Oceanography, 6(??):1–27, ???? 1973. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0079661173900037.

Hendry:1985:DCS

[Hen85] R. M. Hendry. Deep circulation south of the Gulf Stream at 50^{o}W. Progress in Oceanography, 14(??):191–207, ???? 1985. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0079661185900114.

Herndl:1988:DSV

[Her88] Gerhard J. Herndl. Diel and spatial variations in bacterial density in a stratified water column of the Gulf of Trieste. Progress in Oceanography, 21(2):139–146, ???? 1988. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0079661188900328.

Herzfeld:1997:ACS

[Her97] Michael Herzfeld. The annual cycle of sea surface temperature in the Great Australian Bight. Progress in Oceanography, 39(1):1–27, ???? 1997. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661197000104.

Heye:1978:GCM

[Hey78] Dietrich Heye. Growth conditions of manganese nodules comparative studies of growth rate, magnetization, chemical composition and internal structure. Progress in Oceanography, 7(5–6):163–239, ???? 1978. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0079661178900010.

Hecht:1965:DTE

[HF65] Friedrich Hecht and Helmut Fenninger. Determination of trace elements in meteorites. Progress in Oceanography, 3(??):145– 148, ???? 1965. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661165900133.

Huse:2010:MER

[HF10] Geir Huse and Øyvind Fiksen. Modelling encounter rates and distribution of mobile predators and prey. Progress in Oceanography, 84(1–2):93–104, January 2010. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S007966110900144X.

Hosie:2003:DSO

[HFK03] G. W. Hosie, M. Fukuchi, and S. Kawaguchi. Development of the Southern Ocean continuous plankton recorder survey. Progress in Oceanography, 58(2–4):263–283, August 2003. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661103001320.

Haury:2000:ZDA

[HFNG00] Loren Haury, Connie Fey, Carol Newland, and Amatzia Genin. Zooplankton distribution around four eastern North Pacific seamounts. Progress in Oceanography, 45(1):69–105, January 2000. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661199000518.

Horikoshi:1990:BAB

[HFO90] Masuoki Horikoshi, Toshihiko Fujita, and Suguru Ohta. Benthic associations in bathyal and hadal depths off the Pacific coast of north eastern Japan: physiognomies and site factors. Progress in Oceanography, 24(??):331–339, ???? 1990. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/007966119090042Z.

Hirano:2022:OCB

[HFO^{+}22] Daisuke Hirano, Yasushi Fukamachi, Kay I. Ohshima, Masato Ito, Takeshi Tamura, Daisuke Simizu, Toru Takatsuka, Andrew R. Mahoney, Joshua Jones, Craig George, Billy Adams, and Hajo Eicken. Oceanic conditions in the Barrow Coastal Polynya revealed by a 10-year mooring time series. Progress in Oceanography, 203(??):??, April–May 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661122000428.

Hop:2006:PBC

[HFPS^{+}06] Haakon Hop, Stig Falk-Petersen, Harald Svendsen, Slawek Kwasniewski, Vladimir Pavlov, Olga Pavlova, and Janne E.

Søreide. Physical and biological characteristics of the pelagic system across Fram Strait to Kongsfjorden. Progress in Oceanography, 71(2–4):182–231, October 2006. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661106001212.

Harvey:2020:ILT

[HFS^{+}20] Chris J. Harvey, Jennifer L. Fisher, Jameal F. Samhouri, Gregory D. Williams, Tessa B. Francis, Kym C. Jacobson, Yvonne L. deReynier, Mary E. Hunsicker, and Newell Garfield. The importance of long-term ecological time series for integrated ecosystem assessment and ecosystem-based management. Progress in Oceanography, 188(??):??, October 2020. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661120301579.

Herring:1998:ASR

[HFW^{+}98] P. J. Herring, M. J. R. Fasham, A. R. Weeks, J. C. P. Hemmings, H. S. J. Roe, P. R. Pugh, S. Holley, N. A. Crisp, and M. V. Angel. Across-slope relations between the biological populations, the euphotic zone and the oxygen minimum layer off the coast of Oman during the southwest monsoon (August, 1994). Progress in Oceanography, 41(1):69–109, January 1998. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661198000196.

Hughes:2004:BMB

[HG04] D. J. Hughes and J. D. Gage. Benthic metazoan biomass, community structure and bioturbation at three contrasting deep-water sites on the northwest European continental margin. Progress in Oceanography, 63(1–2):29–55, October 2004. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661104000990.

Heneghan:2021:DDR

[HGB^{+}21] Ryan F. Heneghan, Eric Galbraith, Julia L. Blanchard, Cheryl Harrison, Nicolas Barrier, Catherine Bulman, William Cheung, Marta Coll, Tyler D. Eddy, Maite Erauskin-Extramiana, Jason D. Everett, Jose A. Fernandes-Salvador, Didier Gascuel, Jerome Guiet, Olivier Maury, Juliano Palacios-Abrantes, Colleen M. Petrik, Hubert du Pontavice, Anthony J. Richardson, Jeroen Steenbeek, Travis C. Tai, Jan Volkholz, Phoebe A. Woodworth-Jefcoats, and Derek P. Tittensor. Disentangling diverse responses to climate change among global marine ecosystem models. Progress in Oceanography, 198(??):??, November 2021. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661121001440.

Hipfner:2020:BSO

[HGBG20] J. Mark Hipfner, Moira Galbraith, Douglas F. Bertram, and David J. Green. Basin-scale oceanographic processes, zooplankton community structure, and diet and reproduction of a sentinel North Pacific seabird over a 22-year period. Progress in Oceanography, 182(??):??, March 2020. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661120300288.

Harvey:2022:VDD

[HGD22] Michel Harvey, Peter S. Galbraith, and Aurélie Descroix. Vertical distribution and diel migration of macrozooplankton in the St. Lawrence marine system (Canada) in relation with the cold intermediate layer thermal properties. Progress in Oceanography, 80(1–2):1–21, ???? 2022. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661108001924.

Hofer:2019:RWC

[HGH^{+}19] Juan Höfer, Ricardo Giesecke, Mark J. Hopwood, Vania Carrera, Emilio Alarcón, and Humberto E. González. The role of water column stability and wind mixing in the production/export dynamics of two bays in the Western Antarctic Peninsula. Progress in Oceanography, 174(??):105–116, May 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661118300661.

Hernandez-Guerra:2014:MOT

[HGPFN^{+}14] Alonso Hernández-Guerra, Josep L. Pelegrí, Eugenio Fraile-Nuez, Verónica Benítez-Barrios, Mikhail Emelianov, María Dolores Pérez-Hernández, and Pedro Vélez-Belchí. Meridional overturning transports at 7.5N and 24.5N in the Atlantic Ocean during 1992–93 and 2010–11. Progress in Oceanography, 128(??):98–114, November 2014. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661114001499.

Hernandez-Guerra:2016:MOT

[HGT16] Alonso Hernández-Guerra and Lynne D. Talley. Meridional overturning transports at 30^{o}S in the Indian and Pacific Oceans in 2002–2003 and 2009. Progress in Oceanography, 146(??):89– 120, August 2016. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661115300689.

Hernandez-Guerra:2019:UDA

[HGTP^{+}19] Alonso Hernández-Guerra, Lynne D. Talley, José Luis Pelegrí, Pedro Vélez-Belchí, Molly O. Baringer, Alison M. Macdonald, and Elaine L. McDonagh. The upper, deep, abyssal and overturning circulation in the Atlantic Ocean at 30^{o}S in 2003 and 2011. Progress in Oceanography, 176(??):??, September 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661118303501.

Hoch:2023:HTP

[HHAR23] Lilian Hoch, Andrei Herdean, Phoebe A. Argyle, and Peter J. Ralph. High throughput phenomics for diatoms: Challenges and solutions. Progress in Oceanography, 216(??):??, August 2023. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661123001179.

Horsburgh:2000:SEC

[HHB^{+}00] K. J. Horsburgh, A. E. Hill, J. Brown, L. Fernand, R. W. Garvine, and M. M. P. Angelico. Seasonal evolution of the cold pool gyre in the western Irish Sea.

Progress in Oceanography, 46(1):1–58, May 2000. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661199000543.

Halvorsen:2001:DCG

[HHB^{+}01] E. Halvorsen, A. G. Hirst, S. D. Batten, K. S. Tande, and R. S. Lampitt. Diet and community grazing by copepods in an upwelled filament off the NW coast of Spain. Progress in Oceanography, 51(2–4):399–421, November 2001. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661101000775.

Huthnance:2022:OSE

[HHB^{+}22] John Huthnance, Jo Hopkins, Bee Berx, Andy Dale, Jason Holt, Philip Hosegood, Mark Inall, Sam Jones, Benjamin R. Loveday, Peter I. Miller, Jeff Polton, Marie Porter, and Carl Spingys. Ocean shelf exchange, NW European shelf seas: Measurements, estimates and comparisons. Progress in Oceanography, 202(??):??, March 2022. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661122000222.

Hong:2001:RBE

[hHCK01] Chul hoon Hong, Kyu-Dae Cho, and Hee-Joon Kim. The relationship between ENSO events and sea surface temperature in the East (Japan) Sea. Progress in Oceanography, 49(??):21– 40, ???? 2001. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661101000143.

Hermann:2002:CGR

[HHDS02] A. J. Hermann, D. B. Haidvogel, E. L. Dobbins, and P. J. Stabeno. Coupling global and regional circulation models in the coastal Gulf of Alaska. Progress in Oceanography, 53(??):335– 367, ???? 2002. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661102000368.

Holt:2012:MDV

[HHH^{+}12] Jason Holt, Sarah Hughes, Joanne Hopkins, Sarah L. Wakelin, N. Penny Holliday, Stephen Dye, César González-Pola, Solfrid Sætre Hjøllo, Kjell Arne Mork, Glen Nolan, Roger Proctor, Jane Read, Theresa Shammon, Toby Sherwin, Tim Smyth, Graham Tattersall, Ben Ward, and Karen Helen Wiltshire. Multi-decadal variability and trends in the temperature of the northwest European continental shelf: a modeldata synthesis. Progress in Oceanography, 106(??):96–117, November 2012. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661112000894.

Huthnance:2002:NBT

[HHK^{+}02] John M. Huthnance, John D. Humphery, Philip J. Knight, Paul G. Chatwin, Laurenz Thomsen, and Martin White. Near-bed turbulence measurements, stress estimates and sediment mobility at the continental shelf edge. Progress in Oceanography, 52(??):171–194, ???? 2002. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661102000058.

Hartman:2022:SSH

[HHK^{+}22] S. E. Hartman, M. P. Humphreys, C. Kivimäe, E. M. S. Woodward, V. Kitidis, T. McGrath, D. J. Hydes, N. Greenwood, T. Hull, C. Ostle, D. J. Pearce, D. Sivyer, B. M. Stewart, P. Walsham, S. C. Painter, E. McGovern, C. Harris, A. Griffiths, A. Smilenova, J. Clarke, C. Davis, R. Sanders, and P. Nightingale. Seasonality and spatial heterogeneity of the surface ocean carbonate system in the northwest European continental shelf. Progress in Oceanography, 177(??):??, October 2022. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S007966111730099X.

Hirata:2009:IOP

[HHMB^{+}09] Takafumi Hirata, Nick J. Hardman-Mountford, Ray Barlow, Tarron Lamont, Robert Brewin, Tim Smyth, and James Aiken. An inherent optical property approach to the estimation of size-specific photosynthetic rates in eastern boundary upwelling zones from satellite ocean colour: an initial assessment. Progress in Oceanography, 83(1–4):393–397, December 2009. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661109000743.

Houghton:2006:CMN

[HHP06] Robert W. Houghton, Dave Hebert, and Mark Prater. Circulation and mixing at the New England shelfbreak front: Results of purposeful tracer experiments. Progress in Oceanography, 70(2–4):289–312, August 2006. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661106000541.

Howell:2010:SVB

[HHP10] Evan A. Howell, Donald R. Hawn, and Jeffrey J. Polovina. Spatiotemporal variability in bigeye tuna (Thunnus obesus) dive behavior in the central North Pacific Ocean. Progress in Oceanography, 86(1–2):81–93, July 2010. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661110000479.

Hendry:2019:BIG

[HHR^{+}19] Katharine R. Hendry, Veerle A. I. Huvenne, Laura F. Robinson, Amber Annett, Marcus Badger, Allison W. Jacobel, Hong Chin Ng, Jacob Opher, Rebecca A. Pickering, Michelle L. Taylor, Stephanie L. Bates, Adam Cooper, Grace G. Cushman, Claire Goodwin, Shannon Hoy, George Rowland, Ana Samperiz, James A. Williams, Eric P. Achterberg, Carol Arrowsmith, J. Alexander Brearley, Sian F. Henley, Jeffrey W. Krause, Melanie J. Leng, Tao Li, Jerry F. McManus, Michael P. Meredith, Rupert Perkins, and E. Malcolm S. Woodward. The biogeochemical impact of glacial meltwater from Southwest Greenland. Progress in Oceanography, 176(??):??, September 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661118302817.

Hsieh:2005:CLT

[hHRW^{+}05] Chih hao Hsieh, Christian Reiss, William Watson, M. James Allen, John R. Hunter, Robert N. Lea, Richard H. Rosenblatt, Paul E. Smith, and George Sugihara. A comparison of long-term trends and variability in populations of larvae of exploited and unexploited fishes in the Southern California region: a community approach. Progress in Oceanography, 67(1–2):160–185, October 2005. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661105001084.

Houssais:2007:ASA

[HHSR07] Marie-Noelle Houssais, Christophe Herbaut, Pawel Schlichtholz, and Clément Rousset. Arctic salinity anomalies and their link to the North Atlantic during a positive phase of the Arctic oscillation. Progress in Oceanography, 73(2):160–189, May 2007. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661107000377.

Hollowed:2001:PBC

[HHW01] Anne Babcock Hollowed, Steven R. Hare, and Warren S. Wooster. Pacific Basin climate variability and patterns of Northeast Pacific marine fish production. Progress in Oceanography, 49(??):257–282, ???? 2001. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S007966110100026X. See erratum [HHW22].

Hollowed:2022:EPB

[HHW22] Anne Babcock Hollowed, Steven R. Hare, and Warren S. Wooster. Erratum to: Pacific Basin climate variability and patterns of Northeast Pacific marine fish production [Progress on Oceanography 49 (2001) 257–282]. Progress in Oceanography, 55(3–4):371, ???? 2022. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661102001167. See [HHW01].

Hogarth:2020:IET

[HHWW20] P. Hogarth, C. W. Hughes, S. D. P. Williams, and C. Wilson. Improved and extended tide gauge records for the British Isles leading to more consistent estimates of sea level rise and acceleration since 1958. Progress in Oceanography, 184(??): ??, May 2020. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661120300720.

Head:2003:DCS

[HHY03] E. J. H. Head, L. R. Harris, and I. Yashayaev. Distributions of Calanus spp. and other mesozooplankton in the Labrador Sea in relation to hydrography in spring and summer (1995–2000). Progress in Oceanography, 59(1):1–30, October 2003. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661103001113.

Huang:2022:TVI

[HHZ^{+}22] Xiaodong Huang, Siwei Huang, Wei Zhao, Zhiwei Zhang, Chun Zhou, and Jiwei Tian. Temporal variability of internal solitary waves in the northern South China Sea revealed by long-term mooring observations. Progress in Oceanography, 201(??):??, February 2022. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661121001993.

Hickey:1979:CCS

[Hic79] Barbara M. Hickey. The California Current System — hypotheses and facts. Progress in Oceanography, 8(4):191–279, ???? 1979. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661179900028.

Hickey:1992:CSM

[Hic92] Barbara M. Hickey. Circulation over the Santa Monica–San Pedro Basin and Shelf. Progress in Oceanography, 30(1–4):37– 115, ???? 1992. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/007966119290009O.

Hisaki:2022:WHN

[His22] Yukiharu Hisaki. Wave hindcast in the North Pacific area considering the propagation of surface disturbances. Progress in Oceanography, 165(??):332–347, ???? 2022. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661117302306.

Heide-Jorgensen:2007:SCB

[HJLLN07] M. P. Heide-Jørgensen, K. L. Laidre, M. L. Logsdon, and T. G. Nielsen. Springtime coupling between chlorophylla, sea ice and sea surface temperature in Disko Bay, West Greenland. Progress in Oceanography, 73(1):79–95, April 2007. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661107000110.

Hunkins:1965:QSA

[HK65] Kenneth Hunkins and Henry Kutschale. Quaternary sedimentation in the Arctic Ocean. Progress in Oceanography, 4(??):89–94, September 2, 1965. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/007966116590042X.

Honjo:2010:BPP

[HKE^{+}10] Susumu Honjo, Richard A. Krishfield, Timothy I. Eglinton, Steven J. Manganini, John N. Kemp, Kenneth Doherty, Jeomshik Hwang, Theresa K. McKee, and Takatoshi Takizawa. Biological pump processes in the cryopelagic and hemipelagic Arctic Ocean: Canada Basin and Chukchi Rise. Progress in Oceanography, 85(3–4):137–170, June 2010. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661110000200.

Hirche:2006:SFC

[HKGH^{+}06] H. J. Hirche, K. N. Kosobokova, B. Gaye-Haake, I. Harms, B. Meon, and E.-M. Nöthig. Structure and function of contemporary food webs on Arctic shelves: a panarctic comparison: the pelagic system of the Kara Sea — communities and components of carbon flow. Progress in Oceanography, 71(2–4):288–313, October 2006. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661106001224.

Ham:2012:CBV

[HKK12] Yoo-Geun Ham, In-Sik Kang, and Jong-Seong Kug. Coupled bred vectors in the tropical Pacific and their application to ENSO prediction. Progress in Oceanography, 105(??):90–101, October 2012. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661112000389.

Halliwell:2015:OIA

[HKL^{+}15] George R. Halliwell, Vassiliki Kourafalou, Matthieu Le Hénaff, Lynn K. Shay, and Robert Atlas. OSSE impact analysis of airborne ocean surveys for improving upper-ocean dynamical and thermodynamical forecasts in the Gulf of Mexico. Progress in Oceanography, 130(??):32–46, January 2015. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661114001517.

Harada:2014:HSS

[HKN^{+}14] Naomi Harada, Kota Katsuki, Mitsuhiro Nakagawa, Akiko Matsumoto, Osamu Seki, Jason A. Addison, Bruce P. Finney, and Miyako Sato. Holocene sea surface temperature and sea ice extent in the Okhotsk and Bering Seas. Progress in Oceanography, 126(??):242–253, August 2014. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661114000652.

Hinrichsen:2012:IPB

[HKPV12] Hans-Harald Hinrichsen, Wilfried Kühn, Myron A. Peck, and Rudi Voss. The impact of physical and biological factors on the drift and spatial distribution of larval sprat: a comparison of the Baltic and North Seas. Progress in Oceanography, 107(??):47–60, December 2012. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661112000614.

Hirose:2011:VPP

[HKY^{+}11] K. Hirose, C. S. Kim, S. A. Yim, M. Aoyama, M. Fukasawa, K. Komura, P. P. Povinec, and J. A. Sanchez-Cabeza. Vertical profiles of plutonium in the central South Pacific. Progress in Oceanography, 89(1–4):101–107, April 2011. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S007966111000162X.

Halsband-Lenk:2005:MPD

[HL05] Claudia Halsband-Lenk. Metridia pacifica in Dabob Bay, Washington: the diatom effect and the discrepancy between high abundance and low egg production rates. Progress in Oceanography, 67(3–4):422–441, November 2005. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661105001515.

Hernandez-Leon:2019:EMM

[HLCdP19] Santiago Hernández-León, Susana Calles, and María Luz Fernández de Puelles. The estimation of metabolism in the mesopelagic zone: Disentangling deep-sea zooplankton respiration. Progress in Oceanography, 178(??):??, November 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661118303161.

Houliez:2021:STD

[HLD^{+}21] Emilie Houliez, Sébastien Lefebvre, Aurélie Dessier, Martin Huret, Elise Marquis, Martine Bréret, and Christine Dupuy. Spatio-temporal drivers of microphytoplankton community in the Bay of Biscay: Do species ecological niches matter? Progress in Oceanography, 194(??):??, June 2021. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661121000458.

Hu:2023:III

[HLFL23] XiaoKai Hu, GuangHong Liao, ZhiSong Fan, and XiYing Liu. Impacts of inertial internal wave breaking mixing parameterization on a global ocean model. Progress in Oceanography, 212(??):??, March 2023. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S007966112300040X.

Hernandez-Leon:2007:MCC

[HLGA07] Santiago Hernández-León, May Gómez, and Javier Arístegui. Mesozooplankton in the Canary Current System: the coastal– ocean transition zone. Progress in Oceanography, 74(2–3):397– 421, August 2007. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661107000729.

Ham:2013:IMS

[HLK13] Yoo-Geun Ham, Mijung Lim, and Jong-Seong Kug. Importance of mean state in simulating different types of El Niño revealed by SNU coupled GCMs. Progress in Oceanography, 116(??):130–141, September 2013. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661113001110.

Heimburger:2013:TVV

[HLM^{+}13] Lars-Eric Heimbürger, Héloïse Lavigne, Christophe Migon, Fabrizio D’Ortenzio, Claude Estournel, Laurent Coppola, and Juan-Carlos Miquel. Temporal variability of vertical export flux at the DYFAMED time-series station (Northwestern Mediterranean Sea). Progress in Oceanography, 119(??):59–67, December 2013. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661113001651.

Hatun:2016:ISG

[HLM^{+}16] H. Hátún, K. Lohmann, D. Matei, J. H. Jungclaus, S. Pacariz, M. Bersch, A. Gislason, J. Ólafsson, and P. C. Reid. An inflated subpolar gyre blows life toward the northeastern Atlantic. Progress in Oceanography, 147(??):49–66, September 2016. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S007966111630129X.

Hinrichsen:2016:SAE

[HLP^{+}16] H.-H. Hinrichsen, A. Lehmann, C. Petereit, A. Nissling, D. Ustups, U. Bergström, and K. Hüssy. Spawning areas of eastern Baltic cod revisited: Using hydrodynamic modelling to reveal spawning habitat suitability, egg survival probability, and connectivity patterns. Progress in Oceanography, 143(??):13– 25, April 2016. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661116000197.

Halsband-Lenk:2005:RPN

[HLPL05] Claudia Halsband-Lenk, James J. Pierson, and Andrew W. Leising. Reproduction of Pseudocalanus newmani (Copepoda: Calanoida) is deleteriously affected by diatom blooms — a field study. Progress in Oceanography, 67(3–4):332–348, November 2005. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661105001461.

Hale:2017:MPI

[HLR17] Michelle S. Hale, William K. W. Li, and Richard B. Rivkin. Meridional patterns of inorganic nutrient limitation and co-limitation of bacterial growth in the Atlantic Ocean. Progress in Oceanography, 158(??):90–98, November 2017. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S007966111630009X.

Hernandez:2014:UMC

[HLS^{+}14a] Olga Hernandez, Patrick Lehodey, Inna Senina, Vincent Echevin, Patricia Ayón, Arnaud Bertrand, and Philippe Gaspar. Understanding mechanisms that control fish spawning and larval recruitment: Parameter optimization of an Eulerian model (SEAPODYM-SP) with Peruvian anchovy and sardine eggs and larvae data. Progress in Oceanography, 123(??):105– 122, April 2014. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661114000275.

Hsu:2014:LBR

[HLS^{+}14b] Ray T. Hsu, James T. Liu, Chih-Chieh Su, Shuh-Ji Kao, Shih-Nan Chen, Fang-Hsu Kuo, and Jeff C. Huang. On the links between a river’s hyperpycnal plume and marine benthic nepheloid layer in the wake of a typhoon. Progress in Oceanography, 127(??):62–73, September 2014. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661114001074.

Hu:2022:ERV

[HLSX22] Caiqin Hu, Xiangfu Li, Zhen Shi, and Jie Xu. Environmental regulations on viral abundance in the upper South China Sea inferred from statistical models. Progress in Oceanography, 208(??):??, November 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661122001598.

Houssard:2017:TPI

[HLTB^{+}17] Patrick Houssard, Anne Lorrain, Laura Tremblay-Boyer, Valérie Allain, Brittany S. Graham, Christophe E. Menkes, Heidi Pethybridge, Lydie I. E. Couturier, David Point, Bruno Leroy, Aurore Receveur, Brian P. V. Hunt, Elodie Vourey, Sophie Bonnet, Martine Rodier, Patrick Raimbault, Eric Feunteun, Petra M. Kuhnert, Jean-Marie Munaron, Benoit Lebreton, Tsuguo Otake, and Yves Letourneur. Trophic position increases with thermocline depth in yellowfin and bigeye tuna across the western and central Pacific Ocean. Progress in Oceanography, 154(??):49–63, May 2017. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S007966111630177X.

Haedrich:1990:LEF

[HM90] Richard L. Haedrich and Nigel R. Merrett. Little evidence for faunal zonation or communities in deep sea demersal fish faunas. Progress in Oceanography, 24(??):239–250, ???? 1990. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/007966119090033X.

Hall:1998:GIV

[HM98] I. R. Hall and I. N. McCave. Glacial-interglacial variation in organic carbon burial on the slope of the N.W. European Continental Margin (48^{o}–50^{o}n). Progress in Oceanography, 42(1–4):37–60, December 1998. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661198000275.

Hare:2000:EEN

[HM00a] Steven R. Hare and Nathan J. Mantua. Empirical evidence for North Pacific regime shifts in 1977 and 1989. Progress in Oceanography, 47(2–4):103–145, October 2000. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661100000331.

Hooker:2000:CVS

[HM00b] S. B. Hooker and C. R. McClain. The calibration and validation of SeaWiFS data. Progress in Oceanography, 45(3–4):427– 465, April 2000. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661100000124.

Hunt:2006:IBT

[HM06] George L. Hunt and Skip McKinnell. Interplay between topdown, bottom-up, and wasp-waist control in marine ecosystems. Progress in Oceanography, 68(2–4):115–124, February 2006. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661106000139.

Holbrook:2008:SPS

[HM08] Neil J. Holbrook and Angela M. Maharaj. Southwest Pacific subtropical mode water: a climatology. Progress in Oceanography, 77(4):298–315, June 2008. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661108000657.

Henderson:2015:RFM

[HM15] Gideon M. Henderson and Olivier Marchal. Recommendations for future measurement and modelling of particles in GEOTRACES and other ocean biogeochemistry programmes. Progress in Oceanography, 133(??):73–78, April 2015. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661115000269.

Huserbraaten:2018:CDN

[HMA18] M. B. O. Huserbråten, E. Moland, and J. Albretsen. Cod at drift in the North Sea. Progress in Oceanography, 167 (??):116–124, October 2018. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661117303282.

Hayes:1986:HVW

[HMB^{+}86] S. P. Hayes, L. J. Mangum, R. T. Barber, A. Huyer, and R. L. Smith. Hydrographic variability west of the Galápagos Islands during the 1982–1983 El Niño. Progress in Oceanography, 17(1–2):137–162, ???? 1986. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0079661186900261.

Hu:2007:CFC

[HMH07] Aixue Hu, Gerald A. Meehl, and Weiqing Han. Causes of a fresher, colder northern North Atlantic in late 20th century in a coupled model. Progress in Oceanography, 73(3–4):384– 405, May 2007. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661107000766.

Hu:2015:EBS

[HMH^{+}15] Aixue Hu, Gerald A. Meehl, Weiqing Han, Bette Otto-Bliestner, Ayako Abe-Ouchi, and Nan Rosenbloom. Effects of the Bering Strait closure on AMOC and global climate under different background climates. Progress in Oceanography, 132(??):174–196, March 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661114000172.

Honjo:2008:POC

[HMKF08] Susumu Honjo, Steven J. Manganini, Richard A. Krishfield, and Roger Francois. Particulate organic carbon fluxes to the ocean interior and factors controlling the biological pump: a synthesis of global sediment trap programs since 1983. Progress in Oceanography, 76(3):217–285, March 2008. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661108000025.

Hill:2013:SIP

[HMO^{+}13] Victoria J. Hill, Patricia A. Matrai, Elise Olson, S. Suttles, Mike Steele, L. A. Codispoti, and Richard C. Zimmerman. Synthesis of integrated primary production in the Arctic Ocean: II. In situ and remotely sensed estimates. Progress in Oceanography, 110(??):107–125, March 2013. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661112001711.

Head:2013:ECF

[HMP^{+}13] Erica J. H. Head, Webjørn Melle, Pierre Pepin, Espen Bagøien, and Cecilie Broms. On the ecology of Calanus finmarchicus in the Subarctic North Atlantic: a comparison of population dynamics and environmental conditions in areas of the Labrador Sea–Labrador/Newfoundland Shelf and Norwegian Sea Atlantic and coastal waters. Progress in Oceanography, 114(??): 46–63, July 2013. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661113000463.

Hill:2011:SMA

[HMPZ11] Polly G. Hill, Isabelle Mary, Duncan A. Purdie, and Mikhail V. Zubkov. Similarity in microbial amino acid uptake in surface waters of the North and South Atlantic (sub-)tropical gyres. Progress in Oceanography, 91(4):437–446, December 2011. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S007966111100053X.

Hardman-Mountford:2003:OCS

[HMRA^{+}03] N. J. Hardman-Mountford, A. J. Richardson, J. J. Agenbag, E. Hagen, L. Nykjaer, F. A. Shillington, and C. Villacastin. Ocean climate of the South East Atlantic observed from satellite data and wind models. Progress in Oceanography, 59(2–3):181–221, October 2003. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S007966110300168X. See [MKM93].

Hardman-Mountford:2003:RSR

[HMRB^{+}03] N. J. Hardman-Mountford, A. J. Richardson, D. C. Boyer, A. Kreiner, and H. J. Boyer. Relating sardine recruitment in the Northern Benguela to satellite-derived sea surface height using a neural network pattern recognition approach. Progress in Oceanography, 59(2–3):241–255, October 2003. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661103001721.

Hiron:2022:LCS

[HMS^{+}22] Luna Hiron, Philippe Miron, Lynn K. Shay, William E. Johns, Eric P. Chassignet, and Alexandra Bozec. Lagrangian coherence and source of water of Loop Current Frontal Eddies in the Gulf of Mexico. Progress in Oceanography, 208(??):??, November 2022. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661122001355.

Hamad:2005:NHA

[HMTL05] Najwa Hamad, Claude Millot, and Isabelle Taupier-Letage. A new hypothesis about the surface circulation in the eastern basin of the Mediterranean Sea. Progress in Oceanography, 66(2–4):287–298, August 2005. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661105000984.

Hare:2000:IPS

[HMWM00] Steven R. Hare, Shoshiro Minobe, Warren S. Wooster, and Skip McKinnell. An introduction to the PICES symposium on the nature and impacts of North Pacific climate regime shifts. Progress in Oceanography, 47(2–4):99–102, October 2000. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661100000458.

He:2023:SPR

[HMX^{+}23] Qingyou He, Dennis J. McGillicuddy, Xiaogang Xing, Shuqun Cai, Weikang Zhan, Yinghui He, Jiexin Xu, and Haigang Zhan. Subsurface phytoplankton responses to ocean eddies can run counter to satellite-based inference from surface properties in subtropical gyres. Progress in Oceanography, 218(??):??, November 2023. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661123001611.

Hlaili:2014:PFW

[HNL14] Asma Sakka Hlaili, Nathalie Niquil, and Louis Legendre. Planktonic food webs revisited: Reanalysis of results from the linear inverse approach. Progress in Oceanography, 120 (??):216–229, January 2014. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661113001730.

Hosegood:2017:NPS

[HNR^{+}17] P. J. Hosegood, P. D. Nightingale, A. P. Rees, C. E. Widdicombe, E. M. S. Woodward, D. R. Clark, and R. J. Torres. Nutrient pumping by submesoscale circulations in the Mauritanian upwelling system. Progress in Oceanography, 159(??):223–236, December 2017. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661117301003.

Hosegood:2019:ILW

[HNSP^{+}19] P. J. Hosegood, W. A. M. Nimmo-Smith, R. Proud, K. Adams, and A. S. Brierley. Internal lee waves and baroclinic bores over a tropical seamount shark ‘hot-spot’. Progress in Oceanography, 172(??):34–50, March 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661118301964.

Hansen:2000:NAN

[HØ00] B. Hansen and S. Østerhus. North Atlantic–Nordic Seas exchanges. Progress in Oceanography, 45(2):109–208, February 2000. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S007966119900052X.

Hansen:2007:FBC

[HØ07] Bogi Hansen and Svein Østerhus. Faroe Bank Channel overflow 1995–2005. Progress in Oceanography, 75(4):817–856, December 2007. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661107001851.

Hobday:2010:EAF

[Hob10] Alistair J. Hobday. Ensemble analysis of the future distribution of large pelagic fishes off Australia. Progress in Oceanography, 86(1–2):291–301, July 2010. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661110000571.

Hofmann:1981:BRD

[Hof81] Eileen E. Hofmann. Book review: Descriptive regional oceanography: P. Tchernia. Pergamon Press (1980). 19.5 × 13.5 mm, 253 pp. US $52.00 (£23.00), ISBN 0 08 020925 4 (Hardcover). US $19.00 (£9.50), ISBN 0-08-020919-X (Flexicover). Progress in Oceanography, 10(1):iii–iv, ???? 1981. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0079661181900082.

Hofmann:2010:PFG

[Hof10] Eileen E. Hofmann. Plankton functional group models — an assessment. Progress in Oceanography, 84(1–2):16–19, January 2010. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661109001359.

Hogg:1985:EBI

[Hog85] Nelson G. Hogg. Evidence for baroclinic instability in the Gulf Stream recirculation. Progress in Oceanography, 14(??):209– 229, ???? 1985. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661185900126.

Hansen:2003:IFI

[HØH^{+}03] B. Hansen, S. Østerhus, H. Hátún, R. Kristiansen, and K. M. H. Larsen. The Iceland–Faroe inflow of Atlantic water to the Nordic Seas. Progress in Oceanography, 59(4):443–474, December 2003. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661103001770.

Holland:2000:TSI

[Hol00] David M. Holland. Transient sea-ice polynya forced by oceanic flow variability. Progress in Oceanography, 48(4):403–460, ???? 2000. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661101000106.

Hopkins:1964:SMB

[Hop64] Thomas L. Hopkins. A survey of marine bottom samplers. Progress in Oceanography, 2(??):213, 215–256, ????

1964. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0079661164900059.

Hirawake:2021:WMD

[HOY^{+}21a] Toru Hirawake, Joji Oida, Youhei Yamashita, Hisatomo Waga, Hiroto Abe, Jun Nishioka, Daiki Nomura, Hiromichi Ueno, and Atsushi Ooki. Water mass distribution in the northern Bering and southern Chukchi seas using light absorption of chromophoric dissolved organic matter. Progress in Oceanography, 197(??):??, September–October 2021. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661121001269.

Ho:2021:CNI

[HOY^{+}21b] Pei-Chi Ho, Noboru Okuda, Chih-Fu Yeh, Pei-Ling Wang, Gwo-Ching Gong, and Chih hao Hsieh. Carbon and nitrogen isoscape of particulate organic matter in the East China Sea. Progress in Oceanography, 197(??):??, September– October 2021. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S007966112100152X.

Hatun:2009:LBG

[HPB^{+}09] H. Hátún, M. R. Payne, G. Beaugrand, P. C. Reid, A. B. Sandø, H. Drange, B. Hansen, J. A. Jacobsen, and D. Bloch. Large bio-geographical shifts in the north-eastern Atlantic Ocean: From the subpolar gyre, via plankton, to blue whiting and pilot whales. Progress in Oceanography, 80(3–4):149– 162, March 2009. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661109000135.

Hannides:2020:SDM

[HPC^{+}20] Cecelia C. S. Hannides, Brian N. Popp, Hilary G. Close, Claudia R. Benitez-Nelson, Cassie A. Ka’apu-Lyons, Kristen Gloeckler, Natalie Wallsgrove, Blaire Umhau, Emily Palmer, and Jeffrey C. Drazen. Seasonal dynamics of midwater zooplankton and relation to particle cycling in the North Pacific subtropical gyre. Progress in Oceanography, 182(??):??, March 2020. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661120300045.

Hunt:2008:PSO

[HPH^{+}08] B. P. V. Hunt, E. A. Pakhomov, G. W. Hosie, V. Siegel, P. Ward, and K. Bernard. Pteropods in Southern Ocean ecosystems. Progress in Oceanography, 78(3):193–221, September 2008. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661108001183.

Horner:2005:WSP

[HPHL^{+}05] Rita A. Horner, James R. Postel, Claudia Halsband-Lenk, James J. Pierson, Georg Pohnert, and Thomas Wichard. Winter-spring phytoplankton blooms in Dabob Bay, Washington. Progress in Oceanography, 67(3–4):286–313, November 2005. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661105001448.

Hogarth:2021:CMS

[HPHW21] P. Hogarth, D. T. Pugh, C. W. Hughes, and S. D. P. Williams. Changes in mean sea level around Great Britain over the past 200 years. Progress in Oceanography, 192(??):??, March 2021. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661121000112.

Hufnagl:2015:UGK

[HPNDC15] Marc Hufnagl, Myron A. Peck, Richard D. M. Nash, and Mark Dickey-Collas. Unravelling the Gordian knot! Key processes impacting overwintering larval survival and growth: a North Sea herring case study. Progress in Oceanography, 138(??):486–503, November 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661114000779.

Halvorsen:2001:MMU

[HPS^{+}01] E. Halvorsen, O. P. Pedersen, D. Slagstad, K. S. Tande, E. S. Fileman, and S. D. Batten. Microzooplankton and mesozooplankton in an upwelling filament off Galicia: modelling and sensitivity analysis of the linkages and their impact on the carbon dynamics. Progress in Oceanography, 51(2–4):499–513, November 2001. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661101000817.

Huret:2010:DKT

[HPW10] M. Huret, P. Petitgas, and M. Woillez. Dispersal kernels and their drivers captured with a hydrodynamic model and spatial indices: a case study on anchovy (Engraulis encrasicolus) early life stages in the Bay of Biscay. Progress in Oceanography, 87(1–4):6–17, October 2010. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661110001357.

Hong:2021:CSE

[HPZC21] Qingquan Hong, Shiyun Peng, Daochen Zhao, and Pinghe Cai. Cross-shelf export of particulate organic carbon in the northern South China Sea: Insights from a ^{234}Th mass balance. Progress in Oceanography, 193(??):??, April–May 2021. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661121000227.

Hooker:2000:OMI

[HRA00] S. B. Hooker, N. W. Rees, and J. Aiken. An objective methodology for identifying oceanic provinces. Progress in Oceanography, 45(3–4):313–338, April 2000. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661100000069.

Heath:2008:SDC

[HRA^{+}08] M. R. Heath, J. Rasmussen, Y. Ahmed, J. Allen, C. I. H. Anderson, A. S. Brierley, L. Brown, A. Bunker, K. Cook, R. Davidson, S. Fielding, W. S. C. Gurney, R. Harris, S. Hay, S. Henson, A. G. Hirst, N. P. Holliday, A. Ingvarsdottir, X. Irigoien, P. Lindeque, D. J. Mayor, D. Montagnes, C. Moffat, R. Pollard, S. Richards, R. A. Saunders, J. Sidey, G. Smerdon, D. Speirs, P. Walsham, J. Waniek, L. Webster, and D. Wilson. Spatial demography of Calanus finmarchicus in the Irminger Sea. Progress in Oceanography, 76(1):39–88, January 2008. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661107001966.

Hay:2008:GVN

[HRSM08] D. E. Hay, K. A. Rose, J. Schweigert, and B. A. Megrey. Geographic variation in North Pacific herring populations: Pan-Pacific comparisons and implications for climate change impacts. Progress in Oceanography, 77(2–3):233–240, May 2008. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S007966110800058X.

Hunt:2002:CCC

[HS02] George L. Hunt and Phyllis J. Stabeno. Climate change and the control of energy flow in the southeastern Bering Sea. Progress in Oceanography, 55(1–2):5–22, October 2002. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661102000678.

Haupt:2007:SWG

[HS07] Bernd J. Haupt and Dan Seidov. Strengths and weaknesses of the global ocean conveyor: Inter-basin freshwater disparities as the major control. Progress in Oceanography, 73(3–4):358– 369, May 2007. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661107000791.

Havermans:2022:BTT

[HS22] Charlotte Havermans and Victor Smetacek. Bottom-up and top-down triggers of diversification: a new look at the evolutionary ecology of scavenging amphipods in the deep sea. Progress in Oceanography, 164(??):37–51, ???? 2022. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S007966111730229X.

Heymans:2009:PON

[HSC09] Johanna Jacomina Heymans, Ussif Rashid Sumaila, and Villy Christensen. Policy options for the northern Benguela ecosystem using a multispecies, multifleet ecosystem model. Progress in Oceanography, 83(1–4):417–425, December 2009. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S007966110900069X.

Holt:2016:PIC

[HSC^{+}16] Jason Holt, Corinna Schrum, Heather Cannaby, Ute Daewel, Icarus Allen, Yuri Artioli, Laurent Bopp, Momme Butenschon, Bettina A. Fach, James Harle, Dhanya Pushpadas, Baris Salihoglu, and Sarah Wakelin. Potential impacts of climate change on the primary production of regional seas: a comparative analysis of five European seas. Progress in Oceanography, 140(??):91–115, January 2016. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661115002372.

Huyer:2002:COO

[HSF02] A. Huyer, R. L. Smith, and J. Fleischbein. The coastal ocean off Oregon and northern California during the 1997– 8 El Niño. Progress in Oceanography, 54(1–4):311–341, July 2002. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661102000563.

Harwood:2015:CBS

[HSG^{+}15] L. A. Harwood, T. G. Smith, J. C. George, S. J. Sandstrom, W. Walkusz, and G. J. Divoky. Change in the Beaufort Sea ecosystem: Diverging trends in body condition and/or production in five marine vertebrate species. Progress in Oceanography, 136(??):263–273, August 2015. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661115001007.

Han:2023:REJ

[HSGJ23] Qingpeng Han, Xiujuan Shan, Harry Gorfine, and Xianshi Jin. Recruitment of Engraulis japonicus and its environmental drivers: an EDM analysis. Progress in Oceanography, 219(??):??, December 2023. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661123001866.

Hsueh:1997:KFT

[HSH97] Y. Hsueh, John R. Schultz, and William R. Holland. The Kuroshio flow-through in the East China Sea: a numerical model. Progress in Oceanography, 39(2):79–108, ???? 1997. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661197000086.

Henley:2019:VCW

[HSH^{+}19] Sian F. Henley, Oscar M. Schofield, Katharine R. Hendry, Irene R. Schloss, Deborah K. Steinberg, Carlos Moffat, Lloyd S. Peck, Daniel P. Costa, Dorothee C. E. Bakker, Claire Hughes, Patrick D. Rozema, Hugh W. Ducklow, Doris Abele, Jacqueline Stefels, Maria A. Van Leeuwe, Corina P. D. Brussaard, Anita G. J. Buma, Josh Kohut, Ricardo Sahade, Ari S. Friedlaender, Sharon E. Stammerjohn, Hugh J. Venables, and Michael P. Meredith. Variability and change in the west Antarctic Peninsula marine system: Research priorities and opportunities. Progress in Oceanography, 173(??):208–237, April 2019. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661118300818. See corri-

gendum [FDH20].

Hashihama:2019:AMD

[HSK^{+}19] Fuminori Hashihama, Shuhei Suwa, Jota Kanda, Makoto Ehama, Ryousuke Sakuraba, Shinko Kinouchi, Mitsuhide Sato, Tamaha Yamaguchi, Hiroaki Saito, Yoshitoshi Ogura, Tetsuya Hayashi, Hiroshi Mori, Ken Kurokawa, Shotaro Suzuki, and Koji Hamasaki. Arsenate and microbial dynamics in different phosphorus regimes of the subtropical Pacific Ocean. Progress in Oceanography, 176(??):??, September 2019. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661118303033.

Hopkins:1996:TSP

[HSL96] Thomas L. Hopkins, Tracey T. Sutton, and Thomas M. Lancraft. The trophic structure and predation impact of a low latitude midwater fish assemblage. Progress in Oceanography, 38(3):205–239, ???? 1996. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661197000037.

Hirose:2011:VDS

[HSLG11] K. Hirose, T. Saito, S. H. Lee, and J. Gastaud. Vertical distributions of the strong organic ligand in the twilight zone of Southern Hemisphere Ocean particulate matter. Progress in Oceanography, 89(1–4):108–119, April 2011. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661110001631.

Hakspiel-Segura:2022:TVS

[HSMLDC^{+}22] Cristian Hakspiel-Segura, Aída Martínez-López, Juan Antonio Delgado-Contreras, Carlos J. Robinson, and Jaime Gómez-Gutiérrez. Temporal variability of satellite chlorophyll-a as an ecological resilience indicator in the central region of the Gulf of California. Progress in Oceanography, 205(??):??, July 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661122000866.

Huang:2018:CSF

[HSN^{+}18] Zhi Huang, Thomas A. Schlacher, Scott Nichol, Alan Williams, Franziska Althaus, and Rudy Kloser. A conceptual surrogacy framework to evaluate the habitat potential of submarine canyons. Progress in Oceanography, 169(??):199–213, December 2018. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661117301817.

Hales:2012:SBP

[HSS^{+}12] Burke Hales, Peter G. Strutton, Martin Saraceno, Ricardo Letelier, Taro Takahashi, Richard Feely, Christopher Sabine, and Francisco Chavez. Satellite-based prediction of pCO_{2} in coastal waters of the eastern North Pacific. Progress in Oceanography, 103(??):1–15, September 2012. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661112000183.

Herzfeld:1997:NMS

[HT97] Michael Herzfeld and Matthias Tomczak. Numerical modelling of sea surface temperature and circulation in the Great Australian Bight. Progress in Oceanography, 39(1):29–78, ???? 1997. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661197000116.

Houpert:2015:SCM

[HTdM^{+}15] L. Houpert, P. Testor, X. Durrieu de Madron, S. Somot, F. D’Ortenzio, C. Estournel, and H. Lavigne. Seasonal cycle of the mixed layer, the seasonal thermocline and the upperocean heat storage rate in the Mediterranean Sea derived from observations. Progress in Oceanography, 132(??):333–352, March 2015. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661114001797.

Hirai:2015:EGD

[HTG15] Junya Hirai, Atsushi Tsuda, and Erica Goetze. Extensive genetic diversity and endemism across the global range of the oceanic copepod Pleuromamma abdominalis. Progress in Oceanography, 138 (part A)(??):77–90, November 2015. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661115001913.

Horton:2020:ASA

[HTV^{+}20] Tammy Horton, Michael H. Thurston, Rianna Vlierboom, Zoe Gutteridge, Corinne A. Pebody, Andrew R. Gates, and Brian J. Bett. Are abyssal scavenging amphipod assemblages linked to climate cycles? Progress in Oceanography, 184(??): ??, May 2020. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661120300574.

Huthnance:1981:WCN

[Hut81] J. M. Huthnance. Waves and currents near the continental shelf edge. Progress in Oceanography, 10(4):193–226, ???? 1981. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0079661181900045.

Huthnance:1987:ELS

[Hut87] J. M. Huthnance. Effects of longshore shelf variations on barotropic continental shelf waves, slope currents and ocean modes. Progress in Oceanography, 19(2):177–220, ???? 1987. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0079661187900073.

Huthnance:1992:ESC

[Hut92] J. M. Huthnance. Extensive slope currents and the oceanshelf boundary. Progress in Oceanography, 29(2):161–196, ???? 1992. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/007966119290023S.

Huthnance:1995:CEW

[Hut95] John M. Huthnance. Circulation, exchange and water masses at the ocean margin: the role of physical processes at the shelf edge. Progress in Oceanography, 35(4):353–431, ???? 1995. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/007966119580003C.

Huyer:1983:CUC

[Huy83] Adriana Huyer. Coastal upwelling in the California Current System. Progress in Oceanography, 12(3):259–284, ???? 1983. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0079661183900101.

Hinrichsen:2017:STD

[HvDL^{+}17] H.-H. Hinrichsen, B. von Dewitz, A. Lehmann, U. Bergström, and K. Hüssy. Spatio-temporal dynamics of cod nursery areas in the Baltic Sea. Progress in Oceanography, 155(??):28– 40, June 2017. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661116302580.

Hutchings:2009:BCE

[HvdLS^{+}09] L. Hutchings, C. D. van der Lingen, L. J. Shannon, R. J. M. Crawford, H. M. S. Verheye, C. H. Bartholomae, A. K. van der Plas, D. Louw, A. Kreiner, M. Ostrowski, Q. Fidel, R. G. Barlow, T. Lamont, J. Coetzee, F. Shillington, J. Veitch, J. C. Currie, and P. M. S. Monteiro. The Benguela Current: an ecosystem of four components. Progress in Oceanography, 83(1–4):15–32, December 2009. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661109001104.

Huggett:2009:CBS

[HVEF09] Jenny Huggett, Hans Verheye, Rubén Escribano, and Tracey Fairweather. Copepod biomass, size composition and production in the Southern Benguela: Spatio–temporal patterns of variation, and comparison with other eastern boundary upwelling systems. Progress in Oceanography, 83(1–4):197–207, December 2009. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661109000858.

Hanninen:2015:RBN

[HVRR15] Jari Hänninen, Ilppo Vuorinen, Marjut Rajasilta, and Philip C. Reid. Response of the Baltic and North Seas to river runoff from the Baltic watershed — physical and biological changes. Progress in Oceanography, 138 (part A)(??):91–104, November 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661115001901.

Hannah:2010:CME

[HVS10] Charles Hannah, Alain Vezina, and Mike St. John. The case for marine ecosystem models of intermediate complexity. Progress in Oceanography, 84(1–2):121–128, January 2010. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661109001487.

Hauquier:2022:RSD

[HVTV22] Freija Hauquier, Elie Verleyen, Bjorn Tytgat, and Ann Vanreusel. Regional-scale drivers of marine nematode distribution in Southern Ocean continental shelf sediments. Progress in Oceanography, 165(??):1–10, ???? 2022. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661117301866.

Hill:2002:OCN

[HW02] J. K. Hill and P. A. Wheeler. Organic carbon and nitrogen in the northern California Current System: comparison of offshore, river plume, and coastally upwelled waters. Progress in Oceanography, 53(??):369–387, ???? 2002. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S007966110200037X.

Hughes:2018:WDO

[HWB^{+}18] Chris W. Hughes, Joanne Williams, Adam Blaker, Andrew Coward, and Vladimir Stepanov. A window on the deep ocean: the special value of ocean bottom pressure for monitoring the large-scale, deep-ocean circulation. Progress in Oceanography, 161(??):19–46, February 2018. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661117303154.

Hudson:2003:SSF

[HWBT03] Ian R. Hudson, Benjamin D. Wigham, David S. M. Billett, and Paul A. Tyler. Seasonality and selectivity in the feeding ecology and reproductive biology of deep-sea bathyal holothurians. Progress in Oceanography, 59(4):381–407, December 2003. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661103001782.

Hung:2021:BRI

[HWF^{+}21] J.-J. Hung, Y.-H. Wang, K.-H. Fu, I-H. Lee, S.-S. Tsai, C.-Y. Lee, W.-T. Lu, Y.-J. Shen, and Y.-H. Lin. Biogeochemical responses to internal-wave impacts in the continental margin off dongsha atoll in the Northern South China Sea. Progress in Oceanography, 199(??):??, December 2021. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661121001737.

Ho:2020:PSP

[HWL^{+}20] Pei-Chi Ho, Esther Wong, Fan-Sian Lin, Akash R. Sastri, Carmen García-Comas, Noboru Okuda, Fuh-Kwo Shiah, Gwo-Ching Gong, Rita S. W. Yam, and Chih hao Hsieh. Prey stoichiometry and phytoplankton and zooplankton composition influence the production of marine crustacean zooplankton. Progress in Oceanography, 186(??):??, July 2020. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661120301087.

Holt:2010:PIC

[HWLT10] Jason Holt, Sarah Wakelin, Jason Lowe, and Jonathan Tinker. The potential impacts of climate change on the hydrography of the northwest European continental shelf. Progress in Oceanography, 86(3–4):361–379, September 2010. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661110000856.

Hahn-Woernle:2020:SSU

[HWPLvW20] Lisa Hahn-Woernle, Brian Powell, Øyvind Lundesgaard, and Melchior van Wessem. Sensitivity of the summer upper ocean heat content in a Western Antarctic Peninsula fjord. Progress in Oceanography, 183(??):??, April 2020. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661120300252.

Huyer:2007:NLO

[HWS^{+}07] Adriana Huyer, Patricia A. Wheeler, P. Ted Strub, Robert L. Smith, Ricardo Letelier, and P. Michael Kosro. The Newport line off Oregon — studies in the North East Pacific. Progress in Oceanography, 75(2):126–160, October 2007. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661107001474.

Hernandez:2012:TVI

[HYM^{+}12] Klaudia Hernández, Beatriz Yannicelli, Aldo Montecinos, Marcel Ramos, Humberto E. González, and Giovanni Daneri.

Temporal variability of incidental solar radiation and modulating factors in a coastal upwelling area (36^{o}S). Progress in Oceanography, 92–95(??):18–32, January 2012. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661111000656.

Huang:2024:GDI

[HYM^{+}24] Longyu Huang, Jingsong Yang, Zetai Ma, Bingqing Liu, Lin Ren, Antony K. Liu, and Peng Chen. Generation of diurnal internal solitary waves (ISW-D) in the Sulu Sea: From geostationary orbit satellites and numerical simulations. Progress in Oceanography, 225(??):??, July 2024. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661124000855.

He:2016:AEI

[HZCZ16] Qingyou He, Haigang Zhan, Shuqun Cai, and Guozhen Zha. On the asymmetry of eddy-induced surface chlorophyll anomalies in the southeastern Pacific: the role of eddy-Ekman pumping. Progress in Oceanography, 141(??):202–211, February 2016. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661116000033.

He:2023:EIP

[HZD^{+}23] Wenbo He, Xiyuan Zeng, Lin Deng, Qi Lin Chun Pi, and Jun Zhao. Enhanced impact of prolonged MHWs on satellite-observed chlorophyll in the South China Sea. Progress in Oceanography, 218(??):??, November 2023. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661123001660.

Ishida:2002:IVS

[IAFD02] Y. Ishida, T. Azumaya, M. Fukuwaka, and N. Davis. Interannual variability in stock abundance and body size of Pacific salmon in the central Bering Sea. Progress in Oceanography, 55(1–2):223–234, October 2002. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661102000800.

Inall:2013:HAD

[IAN13] Mark E. Inall, Dmitry Aleynik, and Clare Neil. Horizontal advection and dispersion in a stratified shelf sea: the role of inertial oscillations. Progress in Oceanography, 117(??):25–36, October 2013. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661113000967.

Iken:2001:FWS

[IBW^{+}01] K. Iken, T. Brey, U. Wand, J. Voigt, and P. Junghans. Food web structure of the benthic community at the Porcupine Abyssal Plain (NE Atlantic): a stable isotope analysis. Progress in Oceanography, 50(??):383–405, ???? 2001. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661101000623.

Imzilen:2019:FAD

[ICB^{+}19] T. Imzilen, E. Chassot, J. Barde, H. Demarcq, A. Maufroy, L. Roa-Pascuali, J-F. Ternon, and C. Lett. Fish aggregating devices drift like oceanographic drifters in the nearsurface currents of the Atlantic and Indian Oceans. Progress in Oceanography, 171(??):108–127, February 2019. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661118301563.

Irigoien:2007:CBB

[IFC^{+}07] X. Irigoien, Ø. Fiksen, U. Cotano, A. Uriarte, P. Alvarez, H. Arrizabalaga, G. Boyra, M. Santos, Y. Sagarminaga, P. Otheguy, E. Etxebeste, L. Zarauz, I. Artetxe, and L. Motos. Could Biscay Bay anchovy recruit through a spatial loophole? Progress in Oceanography, 74(2–3):132–148, August 2007. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661107000651.

Isla:2019:OOW

[IG19] Enrique Isla and Dieter Gerdes. Ongoing ocean warming threatens the rich and diverse macrobenthic communities of the Antarctic continental shelf. Progress in Oceanography, 178(??):??, November 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661118301794.

Iriarte:2019:SAM

[IGG^{+}19] J. L. Iriarte, I. Gómez, H. E. González, L. Nahuelhual, and J. M. Navarro. Subantarctic and Antarctic marine ecosystems: outlining patterns and processes in a changing ocean. Progress in Oceanography, 174(??):1–6, May 2019. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661119302137.

Iguchi:2004:STV

[Igu04] Naoki Iguchi. Spatial/temporal variations in zooplankton biomass and ecological characteristics of major species in the southern part of the Japan Sea: a review. Progress in Oceanography, 61(2–4):213–225, May 2004. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661104000631.

Ismail:2018:QSH

[IHR18] Khaira Ismail, Veerle Huvenne, and Katleen Robert. Quantifying spatial heterogeneity in submarine canyons. Progress in Oceanography, 169(??):181–198, December 2018. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661117302835.

Inoue:2021:LVC

[IHT^{+}21] Mutsuo Inoue, Shotaro Hanaki, Ryosei Takehara, Hisaki Kofuji, Tetsuya Matsunaka, Hiroshi Kuroda, Yukiko Taniuchi, Hiromi Kasai, Takami Morita, Shizuho Miki, and Seiya Nagao. Lateral variations of ^{134}Cs and ^{228}Ra concentrations in surface waters in the western North Pacific and its marginal sea (2018–2019): Implications for basin-scale and local current circulations. Progress in Oceanography, 195(??):??, July 2021. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661121000744.

Ishida:2001:AEP

[IHY^{+}01] Yukimasa Ishida, Tsutomu Hariu, Junichi Yamashiro, Skip McKinnell, Takeshi Matsuda, and Hiromasa Kaneko. Archeological evidence of Pacific salmon distribution in northern Japan and implications for future global warming. Progress in Oceanography, 49(??):539–550, ???? 2001. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661101000398.

Ichii:2023:ICR

[IIM^{+}23] Taro Ichii, Hiromichi Igarashi, Mao Mori, Kedarnath Mahapatra, Haruka Nishikawa, and Takehiro Okuda. Impact of the climate regime shift around 2000 on recruitment of Antarctic krill at the Antarctic Peninsula and South Georgia. Progress in Oceanography, 213(??):??, April–May 2023. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661123000630.

Ito:2015:GVP

[iIRM^{+}15] Shin ichi Ito, Kenneth A. Rose, Bernard A. Megrey, Jake Schweigert, Douglas Hay, Francisco E. Werner, and Maki Noguchi Aita. Geographic variation in Pacific herring growth in response to regime shifts in the North Pacific Ocean. Progress in Oceanography, 138(??):331–347, November 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661115001196.

Igarashi:2017:PLB

[IIS^{+}17] Hiromichi Igarashi, Taro Ichii, Mitsuo Sakai, Yoichi Ishikawa, Takahiro Toyoda, Shuhei Masuda, Nozomi Sugiura, Kedarnath Mahapatra, and Toshiyuki Awaji. Possible link between interannual variation of neon flying squid (Ommastrephes bartramii) abundance in the North Pacific and the climate phase shift in 1998/1999. Progress in Oceanography, 150(??):20–34, January 2017. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S007966111500049X.

Ito:2010:AAA

[iIYO^{+}10] Shin ichi Ito, Naoki Yoshie, Takeshi Okunishi, Tsuneo Ono, Yuji Okazaki, Akira Kuwata, Taketo Hashioka, Kenneth A. Rose, Bernard A. Megrey, Michio J. Kishi, Miwa Nakamachi, Yugo Shimizu, Shigeho Kakehi, Hiroaki Saito, Kazutaka Takahashi, Kazuaki Tadokoro, Akira Kusaka, and Hiromi Kasai. Application of an automatic approach to calibrate the NEMURO nutrient–phytoplankton–zooplankton food web model in the Oyashio region. Progress in Oceanography, 87(1–4):186–200, October 2010. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S007966111000114X.

Ikehara:1988:OCG

[Ike88] Ken Ikehara. Ocean current generated sedimentary facies in the Osumi Strait, south of Kyushu, Japan. Progress in Oceanography, 21(3–4):515–524, ???? 1988. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0079661188900250.

Iversen:2020:SDM

[IL20] Morten H. Iversen and Richard S. Lampitt. Size does not matter after all: No evidence for a size-sinking relationship for marine snow. Progress in Oceanography, 189(??):??, November– December 2020. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S007966112030183X.

Iachetti:2021:TVP

[ILA21] Clara M. Iachetti, Gustavo Lovrich, and Viviana A. Alder. Temporal variability of the physical and chemical environment, chlorophyll and diatom biomass in the euphotic zone of the Beagle Channel (Argentina): Evidence of nutrient limitation. Progress in Oceanography, 195(??):??, July 2021. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S007966112100063X.

Iermano:2012:FFE

[ILI^{+}12] Ilaria Iermano, Giovanni Liguori, Daniele Iudicone, Bruno Buongiorno Nardelli, Simone Colella, Adriana Zingone, Vincenzo Saggiomo, and Maurizio Ribera d’Alcalà. Filament formation and evolution in buoyant coastal waters: Observation and modelling. Progress in Oceanography, 106(??):118–137, November 2012. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661112000912.

Ianora:2007:RHS

[IMHL07] A. Ianora, A. Miralto, and C. Halsband-Lenk. Reproduction, hatching success, and early naupliar survival in Centropages typicus. Progress in Oceanography, 72(2–3):195–213, February 2007. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661107000080.

Ishak:2022:BSD

[IMM^{+}22] Nurul Huda Ahmad Ishak, Kentaro Motoki, Hiroomi Miyamoto, Taiki Fuji, Yukiko Taniuchi, Shigeho Kakehi, Hiroshi Kuroda, Takashi Setou, and Kazutaka Takahashi. Basin-scale distribution of salps and doliolids in the transition region of the North Pacific Ocean in summer: Drivers of bloom occurrence and effect on the pelagic ecosystem. Progress in Oceanography, 204(??):??, June 2022. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661122000544.

Ito:2014:TWC

[IMW^{+}14] Masashi Ito, Akihiko Morimoto, Tatsuro Watanabe, Osamu Katoh, and Tetsutaro Takikawa. Tsushima Warm Current paths in the southwestern part of the Japan Sea. Progress in Oceanography, 121(??):83–93, February 2014. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661113001936.

Ichii:2017:IED

[INI^{+}17] T. Ichii, H. Nishikawa, H. Igarashi, H. Okamura, K. Mahapatra, M. Sakai, T. Wakabayashi, D. Inagake, and Y. Okada. Impacts of extensive driftnet fishery and late 1990s climate regime shift on dominant epipelagic nekton in the transition region and subtropical frontal zone: Implications for fishery management. Progress in Oceanography, 150(??):35–47, January 2017. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661115000488.

Ichinomiya:2024:IMS

[INO^{+}24] Mutsuo Ichinomiya, Yuichiro Nishibe, Yuji Okazaki, Mitsuhide Sato, and Kazutaka Takahashi. Importance of microzooplankton for sustaining high mesozooplankton biomass during post-bloom period in the Oyashio region of the western subarctic Pacific. Progress in Oceanography, 224(??):??, June 2024. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661124000569.

Itoh:2014:CSM

[INT14] Hiroshi Itoh, Jun Nishioka, and Atsushi Tsuda. Community structure of mesozooplankton in the western part of the Sea of Okhotsk in summer. Progress in Oceanography, 126(??):224– 232, August 2014. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661114000949.

Irvine:2013:ESR

[IOGS13] J. R. Irvine, M. O’Neill, L. Godbout, and J. Schnute. Effects of smolt release timing and size on the survival of hatchery-origin Coho salmon in the Strait of Georgia. Progress in Oceanography, 115(??):111–118, August 2013. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661113000566.

Iriarte:2014:OPC

[IPD14] José L. Iriarte, Silvio Pantoja, and Giovanni Daneri. Oceanographic processes in Chilean fjords of Patagonia: From small to large-scale studies. Progress in Oceanography, 129(??):1–7, December 2014. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661114001670.

Iriarte:2023:HAB

[IPF23] José Luis Iriarte, Gemita Pizarro, and Máximo Frangopulos. Harmful algal blooms in Patagonian fjords and channels systems: Recent advances, gaps, and priorities in a changing ocean. Progress in Oceanography, 217(??):??, September– October 2023. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661123001301.

Ivancic:2016:APA

[IPG^{+}16] Ingrid Ivančić, Martin Pfannkuchen, Jelena Godrijan, Tamara Djakovac, Daniela Marić Pfannkuchen, Marino Korlević, Blaženka Gašparović, and Mirjana Najdek. Alkaline phosphatase activity related to phosphorus stress of microphytoplankton in different trophic conditions. Progress in Oceanography, 146(??):175–186, August 2016. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661116300295.

Iglikowska:2023:MCC

[IPHW^{+}23] Anna Iglikowska, Joanna Przytarska, Emma Humphreys-Williams, Jens Najorka, Maciej Chełchowski, Anna Sowa, Haakon Hop, Maria Włodarska-Kowalczuk, and Piotr Kukliński. Mineralogical and chemical composition of Arctic gastropods shells. Progress in Oceanography, 218(??):??, November 2023. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661123001775.

Im:2019:ERP

[IS19] Dong-Hoon Im and Hae-Lip Suh. Evidence for resource partitioning by ontogeny and species in calanoid copepods. Progress in Oceanography, 176(??):??, September 2019. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661117300629.

Ivanov:2004:CDW

[ISH^{+}04] V. V. Ivanov, G. I. Shapiro, J. M. Huthnance, D. L. Aleynik, and P. N. Golovin. Cascades of dense water around the World Ocean. Progress in Oceanography, 60(1):47–98, January 2004. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661104000035. See com-

ment [dZTG05].

Saitoh:2002:DTS

[iSIS02] Sei ichi Saitoh, Takahiro Iida, and Kosei Sasaoka. A description of temporal and spatial variability in the Bering Sea spring phytoplankton blooms (1997–1999) using satelite multi-sensor remote sensing. Progress in Oceanography, 55(1–2):131–146, October 2002. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661102000745.

Iida:2002:TSV

[ISM^{+}02] T. Iida, S. I. Saitoh, T. Miyamura, M. Toratani, H. Fukushima, and N. Shiga. Temporal and spatial variability of coccolithophore blooms in the eastern Bering Sea, 1998–2001. Progress in Oceanography, 55(1–2):165–175, October 2002. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661102000769.

Ishii:1988:ECR

[IST^{+}88] Hideo Ishii, Takashige Shinozaki, Akimasa Tashiro, Tomoki Nagahama, Kazumi Matsushige, Hiroshi Sasaki, and Kazuo Kawatate. Experiments on cut ropes due to the shearing action of bodies. Progress in Oceanography, 21(3–4):329–341, ???? 1988. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661188900122.

Ichiye:1986:ECY

[ITM86] Takashi Ichiye, Kenzo Takano, and John D. Milliman. East China and Yellow Sea studies — introduction. Progress in Oceanography, 17(3–4):v, ???? 1986. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0079661186900418.

Itoh:2014:DBL

[ITO^{+}14] Sachihiko Itoh, Yuki Tanaka, Satoshi Osafune, Ichiro Yasuda, Masahiro Yagi, Hitoshi Kaneko, Shunsuke Konda, Jun Nishioka, and Yuri. N. Volkov. Direct breaking of large-amplitude internal waves in the Urup Strait. Progress in Oceanography, 126(??):109–120, August 2014. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661114000627.

Iwamoto:2014:SVB

[IU14] Yoko Iwamoto and Mitsuo Uematsu. Spatial variation of biogenic and crustal elements in suspended particulate matter from surface waters of the North Pacific and its marginal seas. Progress in Oceanography, 126(??):211–223, August 2014. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661114000676.

Umatani:1986:FSS

[iUMY86] Shin ichiro Umatani, Noriaki Masunaga, and Toshio Yamagata. Further study of synoptic variability in Wakasa Bay, Japan. Progress in Oceanography, 17(3–4):359–373, ???? 1986. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0079661186900546.

Ingels:2013:STI

[IVR^{+}13] Jeroen Ingels, Ann Vanreusel, Chiara Romano, Johan Coenjaerts, M. Mar Flexas, Diana Zúñiga, and Daniel Martin. Spatial and temporal infaunal dynamics of the Blanes submarine canyon-slope system (NW Mediterranean); changes in nematode standing stocks, feeding types and gender-life stage ratios. Progress in Oceanography, 118(??):159–174, November 2013. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661113001341.

Iriarte:2012:PPP

[IVT^{+}12] Jose Luis Iriarte, Cristian A. Vargas, Fabián J. Tapia, Rafael Bermúdez, and Roberto E. Urrutia. Primary production and plankton carbon biomass in a river-influenced upwelling area off Concepción, Chile. Progress in Oceanography, 92– 95(??):97–109, January 2012. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661111000632.

Iwasaki:2023:RSI

[Iwa23] S. Iwasaki. Role of sea ice on winter wave power and its interannual variability in the Sea of Okhotsk: Natural breakwater modified by surface wind changes. Progress in Oceanography, 210(??):??, January 2023. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661122002038.

Jacques:2010:SOT

[Jac10] Peter J. Jacques. The social oceanography of top oceanic predators and the decline of sharks: a call for a new field. Progress in Oceanography, 86(1–2):192–203, July 2010. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661110000352.

Ji:2012:LHB

[JAC^{+}12] Rubao Ji, Carin J. Ashjian, Robert G. Campbell, Changsheng Chen, Guoping Gao, Cabell S. Davis, Geoffrey W. Cowles, and Robert C. Beardsley. Life history and biogeography of Calanus copepods in the Arctic Ocean: an individualbased modeling study. Progress in Oceanography, 96(1):40– 56, April 2012. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661111001169.

Jones:2008:SDF

[JAJS08] E. P. Jones, L. G. Anderson, S. Jutterström, and J. H. Swift. Sources and distribution of fresh water in the East Greenland Current. Progress in Oceanography, 78(1):37–44, July 2008. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661108000839.

Jacox:2020:SIP

[JAS^{+}20] Michael G. Jacox, Michael A. Alexander, Samantha Siedlecki, Ke Chen, Young-Oh Kwon, Stephanie Brodie, Ivonne Ortiz, Desiree Tommasi, Matthew J. Widlansky, Daniel Barrie, Antonietta Capotondi, Wei Cheng, Emanuele Di Lorenzo, Christopher Edwards, Jerome Fiechter, Paula Fratantoni, Elliott L. Hazen, Albert J. Hermann, Arun Kumar, Arthur J. Miller, Douglas Pirhalla, Mercedes Pozo Buil, Sulagna Ray, Scott C. Sheridan, Aneesh Subramanian, Philip Thompson, Lesley Thorne, Hariharasubramanian Annamalai, Kerim Aydin, Steven J. Bograd, Roger B. Griffis, Kelly Kearney, Hyemi Kim, Annarita Mariotti, Mark Merrifield, and Ryan Rykaczewski. Seasonal-to-interannual prediction of North American coastal marine ecosystems: Forecast methods, mechanisms of predictability, and priority developments. Progress in Oceanography, 183(??):??, April 2020. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S007966112030046X.

Jackson:2015:SAD

[JB15] George A. Jackson and Adrian B. Burd. Simulating aggregate dynamics in ocean biogeochemical models. Progress in Oceanography, 133(??):55–65, April 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661114001451.

Joo:2014:ESS

[JBB^{+}14] Rocio Joo, Arnaud Bertrand, Marilu Bouchon, Alexis Chaigneau, Hervé Demarcq, Jorge Tam, Monique Simier, Dimitri Gutiérrez, Mariano Gutiérrez, Marceliano Segura, Ronan Fablet, and Sophie Bertrand. Ecosystem scenarios shape fishermen spatial behavior. The case of the Peruvian anchovy fishery in the Northern Humboldt Current System. Progress in Oceanography, 128(??):60–73, November 2014. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661114001323.

Jorgensen:2020:RGW

[JBH20] L. L. Jørgensen, G. Bakke, and A. H. Hoel. Responding to global warming: New fisheries management measures in the Arctic. Progress in Oceanography, 188(??):??, October 2020. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661120301622.

Joe:1988:KBT

[JC88] Wang Joe and Ching-Sheng Chern. On the Kuroshio branch in the Taiwan Strait during wintertime. Progress in Oceanography, 21(3–4):469–491, ???? 1988. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0079661188900225.

Johnson:2004:VDD

[JC04] Catherine Lynn Johnson and David M. Checkley. Vertical distribution of diapausing Calanus pacificus (Copepoda) and implications for transport in the California undercurrent. Progress in Oceanography, 62(1):1–13, July 2004. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661104000898.

Jones:2023:ICN

[JCF^{+}23] Elizabeth M. Jones, Melissa Chierici, Agneta Fransson, Karen M. Assmann, Angelika H. H. Renner, and Helene Hodal Lødemel. Inorganic carbon and nutrient dynamics in the marginal ice zone of the Barents Sea: Seasonality and implications for ocean acidification. Progress in Oceanography, 219(??):??, December 2023. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S007966112300174X.

Jones:2018:DVN

[JCIG18] Sam Jones, Finlo Cottier, Mark Inall, and Colin Griffiths. Decadal variability on the Northwest European continental shelf. Progress in Oceanography, 161(??):131–151, February 2018. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661117302094.

Jones:2021:OAS

[JCM^{+}21] Elizabeth M. Jones, Melissa Chierici, Sebastian Menze, Agneta Fransson, Randi B. Ingvaldsen, and Helene Hodal Lødemel. Ocean acidification state variability of the Atlantic Arctic Ocean around northern Svalbard. Progress in Oceanography, 199(??):??, December 2021. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661121001919.

Jackson:1992:FWA

[JE92] George A. Jackson and Peter M. Eldridge. Food web analysis of a planktonic system off Southern California.

Progress in Oceanography, 30(1–4):223–251, ???? 1992. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/007966119290014Q.

Jerlov:1965:EIT

[Jer65] Nils Jerlov. The evolution of the instrumental technique in underwater optics. Progress in Oceanography, 3(??):149–154, ???? 1965. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661165900145.

Johnson:2013:IMB

[JF13] Donna L. Johnson and Michael J. Fogarty. Intercalibration of MOCNESS and bongo nets: Assessing relative efficiency for ichthyoplankton. Progress in Oceanography, 108(??):43–71, January 2013. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661112001310.

Jorda:2013:DFV

[JFEC13] G. Jorda, M. M. Flexas, M. Espino, and A. Calafat. Deep flow variability in a deeply incised Mediterranean submarine valley (Blanes Canyon). Progress in Oceanography, 118(??):47–60, November 2013. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661113001377.

Jollivet:1990:CSO

[JFG^{+}90] Didier Jollivet, Jean-Claude Faugeres, Roger Griboulard, Daniel Desbruyers, and Gerard Blanc. Composition and spatial organization of a cold seep community on the South Barbados accretionary prism: Tectonic, geochemical and sedimentary context. Progress in Oceanography, 24(??):25–45, ???? 1990. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/007966119090017V.

Jithin:2020:EST

[JFUR20] A. K. Jithin, P. A. Francis, A. S. Unnikrishnan, and S. S. V. S. Ramakrishna. Energetics and spatio-temporal variability of semidiurnal internal tides in the Bay of Bengal and Andaman Sea. Progress in Oceanography, 189(??):??, November– December 2020. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661120301816.

Jouin:1990:GHV

[JG90] Claude Jouin and Françoise Gaill. Gills of hydrothermal vent annelids: Structure, ultrastructure and functional implications in two alvinellid species. Progress in Oceanography, 24(??):59– 69, ???? 1990. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/007966119090019X.

Johnson:2007:DWM

[JG07] Gregory C. Johnson and Nicolas Gruber. Decadal water mass variations along 20^{o}W in the Northeastern Atlantic Ocean. Progress in Oceanography, 73(3–4):277–295, May 2007. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661107000742.

Jennings:2020:ISD

[JGB20] Robert M. Jennings, Olga Golovan, and Saskia Brix. Integrative species delimitation of desmosomatid and nannoniscid isopods from the Kuril–Kamchatka Trench, with description of a hadal species. Progress in Oceanography, 182(??): ??, March 2020. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661119304161.

Jouanneau:1998:DDS

[JGO^{+}98] J. M. Jouanneau, C. Garcia, A. Oliveira, A. Rodrigues, J. A. Dias, and O. Weber. Dispersal and deposition of suspended sediment on the shelf off the Tagus and Sado estuaries, S.W. Portugal. Progress in Oceanography, 42(1–4):233–257, December 1998. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661198000366.

Jay:1990:ESP

[JGS90] David A. Jay, Benjamin S. Giese, and Christopher R. Sherwood. Energetics and sedimentary processes in the Columbia River Estuary. Progress in Oceanography, 25(1–4):157–174, ???? 1990. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/007966119090006N.

Jin:2012:DIE

[JHDT12] D. Jin, P. Hoagland, T. M. Dalton, and E. M. Thunberg. Development of an integrated economic and ecological framework for ecosystem-based fisheries management in New England. Progress in Oceanography, 102(??):93–101, September 2012. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661112000249.

Jiang:2022:PDR

[JHM^{+}22] Siyu Jiang, Fuminori Hashihama, Yukio Masumoto, Hongbin Liu, Hiroshi Ogawa, and Hiroaki Saito. Phytoplankton dynamics as a response to physical events in the oligotrophic Eastern Indian Ocean. Progress in Oceanography, 203(??):??, April–May 2022. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661122000453.

Jones:2014:FSH

[JHW^{+}14] A. R. Jones, P. Hosegood, R. B. Wynn, M. N. De Boer, S. Butler-Cowdry, and C. B. Embling. Fine-scale hydrodynamics influence the spatio-temporal distribution of harbour porpoises at a coastal hotspot. Progress in Oceanography, 128(??):30–48, November 2014. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661114001256.

Joint:2001:TLE

[JIT^{+}01] Ian Joint, Mark Inall, Ricardo Torres, Francisco G. Figueiras, Xosé A. Álvarez-Salgado, Andrew P. Rees, and E. Malcolm S. Woodward. Two Lagrangian experiments in the Iberian Upwelling System: tracking an upwelling event and an off-shore filament. Progress in Oceanography, 51(2–4):221–248, November 2001. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661101000684.

Jutterstrom:2008:ACE

[JJ08] Sara Jutterström and Emil Jeansson. Anthropogenic carbon in the East Greenland Current. Progress in Oceanography, 78(1): 29–36, July 2008. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661108000876.

Jutterstrom:2008:EAC

[JJA^{+}08] Sara Jutterström, Emil Jeansson, Leif G. Anderson, Richard Bellerby, E. Peter Jones, William M. Smethie, and James H. Swift. Evaluation of anthropogenic carbon in the Nordic Seas using observed relationships of N, P and C versus CFCs. Progress in Oceanography, 78(1):78–84, July 2008. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661108000815.

Jagadeesan:2013:OCS

[JJA^{+}13] L. Jagadeesan, R. Jyothibabu, A. Anjusha, Arya P. Mohan, N. V. Madhu, K. R. Muraleedharan, and K. Sudheesh. Ocean currents structuring the mesozooplankton in the Gulf of Mannar and the Palk Bay, southeast coast of India. Progress in Oceanography, 110(??):27–48, March 2013. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661112001838.

Jagadeesan:2017:DCU

[JJA^{+}17] L. Jagadeesan, R. Jyothibabu, N. Arunpandi, C. Karnan, and K. K. Balachandran. Dominance of coastal upwelling over mud bank in shaping the mesozooplankton along the southwest coast of India during the Southwest Monsoon. Progress in Oceanography, 156(??):252–275, August 2017. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661116300258.

Ji:2019:VPP

[JJL^{+}19] Rubao Ji, Meibing Jin, Yun Li, Yun-Ho Kang, and Chang-Keun Kang. Variability of primary production among basins in the East/Japan Sea: Role of water column stability in modulating nutrient and light availability. Progress in Oceanography, 178(??):??, November 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661118303203.

Jeansson:2008:SEG

[JJR^{+}08] Emil Jeansson, Sara Jutterström, Bert Rudels, Leif G. Anderson, K. Anders Olsson, E. Peter Jones, William M. Smethie, and James H. Swift. Sources to the East Greenland Current and its contribution to the Denmark Strait Overflow. Progress in Oceanography, 78(1):12–28, July 2008. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661108000803.

Jossi:2003:CPR

[JJS03] J. W. Jossi, A. W. G. John, and D. Sameoto. Continuous plankton recorder sampling off the east coast of North America: history and status. Progress in Oceanography, 58(2–4):313–325, August 2003. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661103001356.

Jones:1987:HSL

[JKBH87] R. L. Jones, D. H. Keen, J. F. Birnie, and D. T. Holyoak. Holocene sea-level changes on Jersey. Progress in Oceanography, 18(1–4):177–204, ???? 1987. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0079661187900334.

Jickells:2008:LBS

[JLB^{+}08] T. D. Jickells, P. S. Liss, W. Broadgate, S. Turner, A. J. Kettle, J. Read, J. Baker, L. M. Cardenas, F. Carse, M. Hamren-Larssen, L. Spokes, M. Steinke, A. Thompson, A. Watson, S. D. Archer, R. G. J. Bellerby, C. S. Law, P. D. Nightingale, M. I. Liddicoat, C. E. Widdicombe, A. Bowie, L. C. Gilpin, G. Moncoiffé, G. Savidge, T. Preston, P. Hadziabdic, T. Frost, R. Upstill-Goddard, C. Pedrós-Alió, R. Simó, A. Jackson, A. Allen, and M. D. DeGrandpre. A Lagrangian biogeochemical study of an eddy in the Northeast Atlantic. Progress in Oceanography, 76(3):366–398, March 2008. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661108000098.

Johns:2020:CEP

[JLP^{+}20a] Elizabeth M. Johns, Rick Lumpkin, Nathan F. Putman, Ryan H. Smith, Frank E. Muller-Karger, Digna T. Rueda-Roa, Chuanmin Hu, Mengqiu Wang, Maureen T. Brooks, Lewis J. Gramer, and Francisco E. Werner. Corrigendum to “The establishment of a pelagic Sargassum population in the tropical Atlantic: Biological consequences of a basinscale long distance dispersal event” [Progr. Oceanogr. 182 (2020) 102269]. Progress in Oceanography, 188(??):??, October 2020. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661120301762. See [JLP^{+}20b].

Johns:2020:EPS

[JLP^{+}20b] Elizabeth M. Johns, Rick Lumpkin, Nathan F. Putman, Ryan H. Smith, Frank E. Muller-Karger, Digna T. Rueda-Roa, Chuanmin Hu, Mengqiu Wang, Maureen T. Brooks, Lewis J. Gramer, and Francisco E. Werner. The establishment of a pelagic Sargassum population in the tropical Atlantic: Biological consequences of a basin-scale long distance dispersal event. Progress in Oceanography, 182(??):??, March 2020. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661120300070. See corrigendum

[JLP^{+}20a].

Johannsen:2020:CSC

[JLRB20] Nele Johannsen, Lidia Lins, Torben Riehl, and Angelika Brandt. Changes in species composition of haploniscidae (Crustacea: Isopoda) across potential barriers to dispersal in the Northwest Pacific. Progress in Oceanography, 180(??):??, January 2020. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661119304136.

Jorgensen:2022:IML

[JLS^{+}22] Lis L. Jørgensen, Elizabeth A. Logerwell, Natalia Strelkova, Denis Zakharov, Virginie Roy, Claude Nozères, Bodil A. Bluhm, Steinunn Hilma Ólafsdóttir, Julian M. Burgos, Jan Sørensen, Olga Zimina, and Kimberly Rand. International megabenthic long-term monitoring of a changing Arctic ecosystem: Baseline results. Progress in Oceanography, 200(??):??, January 2022. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661121001956.

Jones:1988:I

[JM88] E. B. Gareth Jones and J. David Miller. Introduction. Progress in Oceanography, 21(2):v–vii, ???? 1988. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0079661188900274.

Jazdzewska:2019:HSR

[JM19] Anna Maria Jażdżewska and Tomasz Mamos. High species richness of Northwest Pacific deep-sea amphipods revealed through DNA barcoding. Progress in Oceanography, 178 (??):??, November 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661119300643.

Jones:2013:CDD

[JMG^{+}13] V. Jones, T. B. Meador, A. Gogou, C. Migon, K. E. H. Penkman, M. J. Collins, and D. J. Repeta. Characterisation and dynamics of dissolved organic matter in the Northwestern Mediterranean Sea. Progress in Oceanography, 119(??):78–89, December 2013. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661113000955.

Jorda-Molina:2023:LSS

[JMSB^{+}23] Èric Jordà-Molina, Arunima Sen, Bodil A. Bluhm, Paul E. Renaud, Maria Włodarska-Kowalczuk, Joanna Legeżyńska, Barbara Oleszczuk, and Henning Reiss. Lack of strong seasonality in macrobenthic communities from the northern Barents Sea shelf and Nansen Basin. Progress in Oceanography, 219(??):??, December 2023. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661123001933.

Johnson:2023:SPA

[JMZ23] Roberta Johnson, Clara Manno, and Patrizia Ziveri. Shelled pteropod abundance and distribution across the Mediterranean Sea during spring. Progress in Oceanography, 210(??):??, January 2023. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661122001896.

Jordi:2005:SSE

[JOBT05] A. Jordi, A. Orfila, G. Basterretxea, and J. Tintoré. Shelfslope exchanges by frontal variability in a steep submarine canyon. Progress in Oceanography, 66(2–4):120–141, August 2005. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661105000662.

Jensen:2010:LMH

[JOGM^{+}10] Olaf P. Jensen, Sofia Ortega-Garcia, Steven J. D. Martell, Robert N. M. Ahrens, Michael L. Domeier, Carl J. Walters, and James F. Kitchell. Local management of a “highly migratory species”: the effects of long-line closures and recreational catch-and-release for Baja California striped marlin fisheries. Progress in Oceanography, 86(1–2):176–186, July 2010. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661110000546.

Johnson:2004:SVM

[Joh04] Catherine Lynn Johnson. Seasonal variation in the molt status of an oceanic copepod. Progress in Oceanography, 62(1): 15–32, July 2004. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661104000904.

Jonsson:2007:VFF

[Jón07] Steingrímur Jónsson. Volume flux and fresh water transport associated with the East Icelandic Current. Progress in Oceanography, 73(3–4):231–241, May 2007. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S007966110700078X.

Jeanthon:1990:RHM

[JP90] Christian Jeanthon and Daniel Prieur. Resistance to heavy metals of heterotrophic bacteria isolated from the deepsea hydrothermal vent polychaete, Alvinella pompejana. Progress in Oceanography, 24(??):81–88, ???? 1990. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/007966119090021S.

Jakiel:2020:SDP

[JPB20] Aleksandra Jakiel, Ferran Palero, and Magdalena Błażewicz. Secrets from the deep: Pseudotanaidae (Crustacea: Tanaidacea) diversity from the Kuril–Kamchatka Trench. Progress in Oceanography, 183(??):??, April 2020. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0079661120300264.

Jozwiak:2020:ESV

[JPBB20] Piotr Jóźwiak, Krzysztof Pabis, Angelika Brandt, and Magdalena Błażewicz. Epibenthic sled versus giant box corer — comparison of sampling gears for tanaidacean species richness assessment in the abyssal benthic ecosystem. Progress in Oceanography, 181(??):??, February 2020. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661119304355.

Jorgensen:2022:BTZ

[JPIP22] Lis L. Jørgensen, Laurene Pecuchet, Randi B. Ingvaldsen, and Raul Primicerio. Benthic transition zones in the Atlantic gateway to a changing Arctic Ocean. Progress in Oceanography, 204(??):??, June 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661122000532.

Jarre:2008:KBS

[JPM^{+}08] Astrid Jarre, Barbara Paterson, Coleen L. Moloney, David C. M. Miller, John G. Field, and Anthony M. Starfield. Knowledge-based systems as decision support tools in an ecosystem approach to fisheries: Comparing a fuzzy-logic and a rule-based approach. Progress in Oceanography, 79 (2–4):390–400, October 2008. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661108001900.

Jakoboski:2024:PCF

[JRR^{+}24] Julie Jakoboski, Moninya Roughan, John Radford, João Marcos Azevedo Correia de Souza, Malene Felsing, Robert Smith, Naomi Puketapu-Waite, Mireya Montaño Orozco, Kimberley H. Maxwell, and Cooper Van Vranken. Partnering with the commercial fishing sector and Aotearoa New Zealand’s ocean community to develop a nationwide subsurface temperature monitoring program. Progress in Oceanography, 225(??): ??, July 2024. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661124000843.

Joint:2001:PPN

[JRW01] Ian Joint, Andrew P. Rees, and E. Malcolm S. Woodward. Primary production and nutrient assimilation in the Iberian upwelling in August 1998. Progress in Oceanography, 51 (2–4):303–320, November 2001. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661101000726.

Jennings:1987:CSE

[JS87] S. Jennings and C. Smyth. Coastal sedimentation in East Sussex during the Holocene. Progress in Oceanography, 18 (1–4):205–241, ???? 1987. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0079661187900346.

Jay:1990:CDD

[JS90] David A. Jay and J. Dungan Smith. Circulation, density distribution and neap-spring transitions in the Columbia River Estuary. Progress in Oceanography, 25(1–4):81–112, ???? 1990. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/007966119090004L.

Jamieson:2021:HZN

[JS21] Alan J. Jamieson and Heather A. Stewart. Hadal zones of the Northwest Pacific Ocean. Progress in Oceanography, 190(??): ??, January 2021. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661120302123.

Jahncke:2008:ERS

[JSA^{+}08] J. Jahncke, B. L. Saenz, C. L. Abraham, C. Rintoul, R. W. Bradley, and W. J. Sydeman. Ecosystem responses to shortterm climate variability in the Gulf of the Farallones, California. Progress in Oceanography, 77(2–3):182–193, May 2008. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661108000530.

Jansen:2021:BEP

[JSdSS^{+}21] Teunis Jansen, Aril Slotte, Thassya Christina dos Santos Schmidt, Claus Reedtz Sparrevohn, Jan Arge Jacobsen, and Olav Sigurd Kjesbu. Bioenergetics of egg production in Northeast Atlantic mackerel changes the perception of fecundity type and annual trends in spawning stock biomass. Progress in Oceanography, 198(??):??, November 2021. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661121001439.

Jones:1990:CSD

[JSHB90] Kim K. Jones, Charles A. Simenstad, Duane L. Higley, and Daniel L. Bottom. Community structure, distribution, and standing stock of benthos, epibenthos, and plankton in the Columbia River Estuary. Progress in Oceanography, 25(1–4):211–241, ???? 1990. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/007966119090008P.

Johnson:2002:DMU

[JSKM02] Gregory C. Johnson, Bernadette M. Sloyan, William S. Kessler, and Kristene E. McTaggart. Direct measurements of upper ocean currents and water properties across the tropical Pacific during the 1990s. Progress in Oceanography, 52(1):31–61, January 2002. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661102000216.

Jones:2021:EEP

[JSLA^{+}21] Daniel O. B. Jones, Erik Simon-Lledó, Diva J. Amon, Brian J. Bett, Clémence Caulle, Louis Clément, Douglas P. Connelly, Thomas G. Dahlgren, Jennifer M. Durden, Jeffrey C. Drazen, Janine Felden, Andrew R. Gates, Magdalena N. Georgieva, Adrian G. Glover, Andrew J. Gooday, Anita L. Hollingsworth, Tammy Horton, Rachael H. James, Rachel M. Jeffreys, Claire Laguionie-Marchais, Astrid B. Leitner, Anna Lichtschlag, Amaya Menendez, Gordon L. J. Paterson, Kate Peel, Katleen Robert, Timm Schoening, Natalia A. Shulga, Craig R. Smith, Sergio Taboada, Andreas M. Thurnherr, Helena Wiklund, C. Robert Young, and Veerle A. I. Huvenne. Environment, ecology, and potential effectiveness of an area protected from deep-sea mining (Clarion Clipperton Zone, abyssal Pacific). Progress in Oceanography, 197(??):??, September– October 2021. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661121001385.

Johnson:2024:AOC

[JST^{+}24] Leah Johnson, David A. Siegel, Andrew F. Thompson, Erik Fields, Zachary K. Erickson, Ivona Cetinic, Craig M. Lee, Eric A. D’Asaro, Norman B. Nelson, Melissa M. Omand, Michaela Sten, Shawnee Traylor, David P. Nicholson, Jason R. Graff, Deborah K. Steinberg, Heidi M. Sosik, Ken O. Buesseler, Mark A. Brzezinski, Inia Soto Ramos, Filipa Carvalho, and Stephanie A. Henson. Assessment of oceanographic conditions during the North Atlantic EXport processes in the ocean from RemoTe sensing (EXPORTS) field campaign. Progress in Oceanography, 220(??):??, January 2024. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661123002136.

Jacob:2014:SSF

[JTD^{+}14] Bárbara G. Jacob, Fabián J. Tapia, Giovanni Daneri, Jose L. Iriarte, Paulina Montero, Marcus Sobarzo, and Renato A. Quiñones. Springtime size-fractionated primary production across hydrographic and PAR-light gradients in Chilean Patagonia (41–50^{o}S). Progress in Oceanography, 129(??):75–84, December 2014. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661114001268.

Jimenez-Toribio:2010:GIE

[JTGM10] Ramòn Jiménez-Toribio, Patrice Guillotreau, and Rémi Mongruel. Global integration of European tuna markets. Progress in Oceanography, 86(1–2):166–175, July 2010. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S007966111000056X.

Jacob:2018:MCD

[JTQ^{+}18] Bárbara G. Jacob, Fabián J. Tapia, Renato A. Quiñones, Rodrigo Montes, Marcus Sobarzo, Wolfgang Schneider, Giovanni Daneri, Carmen E. Morales, Paulina Montero, and Humberto E. González. Major changes in diatom abundance, productivity, and net community metabolism in a windier and dryer coastal climate in the southern Humboldt Current. Progress in Oceanography, 168(??):196–209, November 2018. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S007966111730006X.

Jeandel:2015:WDW

[JvdLL^{+}15] Catherine Jeandel, Michiel Rutgers van der Loeff, Phoebe J. Lam, Matthieu Roy-Barman, Robert M. Sherrell, Sven Kretschmer, Chris German, and Frank Dehairs. What did we learn about ocean particle dynamics in the GEOSECS– JGOFS era? Progress in Oceanography, 133(??):6–16, April 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661115000026.

Jorda:2017:MSH

[JVJ^{+}17] G. Jordà, K. Von Schuckmann, S. A. Josey, G. Caniaux, J. García-Lafuente, S. Sammartino, E. Özsoy, J. Polcher, G. Notarstefano, P.-M. Poulain, F. Adloff, J. Salat, C. Naranjo, K. Schroeder, J. Chiggiato, G. Sannino, and D. Macías. The Mediterranean Sea heat and mass budgets: Estimates, uncertainties and perspectives. Progress in Oceanography, 156(??):174–208, August 2017. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661116300817.

Joint:2001:LSI

[JW01a] I. Joint and P. Wassmann. Lagrangian studies of the Iberian upwelling system — an introduction. a study of the temporal evolution of surface production and fate of organic matter during upwelling on and off the NW Spanish continental margin. Progress in Oceanography, 51(2–4):217–220, November 2001. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661101000672.

Joint:2001:PRW

[JW01b] I. Joint and P. Wassmann. Professor Roland Wollast. Progress in Oceanography, 51(2–4):215–216, November 2001. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661101000829.

Jouanneau:2002:RSS

[JWD^{+}02] J. M. Jouanneau, O. Weber, T. Drago, A. Rodrigues, A. Oliveira, J. M. A. Dias, C. Garcia, S. Schmidt, and J. L. Reyss. Recent sedimentation and sedimentary budgets on the western Iberian shelf. Progress in Oceanography, 52(??):261– 275, ???? 2002. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661102000101.

Jiang:2018:MIN

[JX18] Long Jiang and Meng Xia. Modeling investigation of the nutrient and phytoplankton variability in the Chesapeake Bay outflow plume. Progress in Oceanography, 162(??):290–302, March 2018. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661117300344.

Jiang:2014:SPP

[JYK^{+}14] Yong Jiang, Eun Jin Yang, Sun Yong Kim, Young-Nam Kim, and SangHoon Lee. Spatial patterns in pelagic ciliate community responses to various habitats in the Amundsen Sea (Antarctica). Progress in Oceanography, 128(??):49–59, November 2014. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661114001293.

Jazdzewska:2019:NAA

[JZ19] Anna Maria Jażdżewska and Aleksandra Ziemkiewicz. A new abyssal amphipod species (crustacea) from sunken wood described using integrative taxonomy. Progress in Oceanography, 172(??):1–13, March 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661118302581.

Ji:2024:SVO

[JZZY24] Xuan Ji, Ming-Liang Zhao, Jing Zhang, and Gui-Peng Yang. Spatial variation in the optical and molecular properties of dissolved organic matter in the Yellow Sea and East China Sea. Progress in Oceanography, 220(??):??, January 2024. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661123002355.

Knox:1985:RAS

[KA85] R. A. Knox and D. L. T. Anderson. Recent advances in the study of the low-latitude ocean circulation. Progress in Oceanography, 14(??):259–300, IN1, 301–317, ???? 1985. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/007966118590014X.

Krause:1994:MBC

[KA94] Dale C. Krause and Martin V. Angel. Marine biogeography, climate change and societal needs. Progress in Oceanography, 34(2–3):221–235, ???? 1994. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0079661194900108.

Kurten:2016:CNS

[KAAK^{+}16] Benjamin Kürten, Ali M. Al-Aidaroos, Saskia Kürten, Mohsen M. El-Sherbiny, Reny P. Devassy, Ulrich Struck, Nikolaos Zarokanellos, Burton H. Jones, Thomas Hansen, Gerd Bruss, and Ulrich Sommer. Carbon and nitrogen stable isotope ratios of pelagic zooplankton elucidate ecohydrographic features in the oligotrophic Red Sea. Progress in Oceanography, 140(??):69–90, January 2016. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661115002281.

Kagan:1997:EMT

[Kag97] Boris A. Kagan. Earth–Moon tidal evolution: model results and observational evidence. Progress in Oceanography, 40(??):109–124, ???? 1997. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S007966119700027X.

Kubryakov:2019:PAS

[KAG^{+}19] A. A. Kubryakov, A. A. Aleskerova, Yu. N. Goryachkin, S. V. Stanichny, A. A. Latushkin, and A. V. Fedirko. Propagation of the Azov Sea waters in the Black Sea under impact of variable winds, geostrophic currents and exchange in the Kerch Strait. Progress in Oceanography, 176(??):??, September 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661118302957.

Kourafalou:2016:NAO

[KAH^{+}16] Vassiliki H. Kourafalou, Yannis S. Androulidakis, George R. Halliwell, HeeSook Kang, Michael M. Mehari, Matthieu Le Hénaff, Robert Atlas, and Rick Lumpkin. North Atlantic Ocean OSSE system development: Nature run evaluation and application to hurricane interaction with the Gulf Stream. Progress in Oceanography, 148(??):1–25, November 2016. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661115300653.

Kiyofuji:2022:NMD

[KAK^{+}22a] H. Kiyofuji, Y. Aoki, J. Kinoshita, S. Okamoto, M. Masujima, T. Matsumoto, K. Fujioka, R. Ogata, T. Nakao, N. Sugimoto, and T. Kitagawa. Northward migration dynamics of skipjack tuna (Katsuwonus pelamis) associated with the lower thermal limit in the western Pacific Ocean. Progress in Oceanography, 175(??):55–67, ???? 2022. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S007966111830185X.

Kourafalou:2022:PCB

[KAK^{+}22b] V. H. Kourafalou, Y. S. Androulidakis, H. Kang, R. H. Smith, and A. Valle-Levinson. Physical connectivity between Pulley Ridge and Dry Tortugas coral reefs under the influence of the Loop Current/Florida Current system. Progress in Oceanography, 165(??):75–99, ???? 2022. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661117303385.

Kamenev:2019:BMK

[Kam19] Gennady M. Kamenev. Bivalve mollusks of the Kuril– Kamchatka Trench, Northwest Pacific Ocean: Species composition, distribution and taxonomic remarks. Progress in Oceanography, 176(??):??, September 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661119301247.

Katsura:2018:PFD

[Kat18] Shota Katsura. Properties, formation, and dissipation of the North Pacific Eastern subtropical mode water and its impact on interannual spiciness anomalies. Progress in Oceanography, 162(??):120–131, March 2018. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661117302707.

Kawatate:1986:DTD

[Kaw86] Kazuo Kawatate. A dynamic two-dimensional calculation of underwater lines. Progress in Oceanography, 17(3–4):375–387, ???? 1986. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661186900558.

Kawai:1998:BHR

[Kaw98] Hideo Kawai. A brief history of recognition of the Kuroshio. Progress in Oceanography, 41(4):505–578, October 1998. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S007966119800024X.

Kazmin:2017:VCO

[Kaz17] Alexander S. Kazmin. Variability of the climatic oceanic frontal zones and its connection with the large-scale atmospheric forcing. Progress in Oceanography, 154(??):38–48, May 2017. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661116301124.

Ku:1965:RSA

[KB65] Teh-Lung Ku and Wallace S. Broecker. Rates of sedimentation in the Arctic Ocean. Progress in Oceanography, 4 (??):95–104, September 2, 1965. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0079661165900431.

Karl:2022:TDT

[KBC^{+}22] David M. Karl, Karin M. Björkman, Matthew J. Church, Lance A. Fujieki, Eric M. Grabowski, and Ricardo M. Letelier. Temporal dynamics of total microbial biomass and particulate detritus at Station ALOHA. Progress in Oceanography, 205 (??):??, July 2022. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661122000647.

Kleparski:2022:MTN

[KBE^{+}22] Loïck Kléparski, Grégory Beaugrand, Martin Edwards, François G. Schmitt, Richard R. Kirby, Elsa Breton, François Gevaert, and Emeline Maniez. Morphological traits, nicheenvironment interaction and temporal changes in diatoms. Progress in Oceanography, 201(??):??, February 2022. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S007966112200009X.

Kline:2008:PSO

[KBF^{+}08] Thomas C. Kline, Jennifer L. Boldt, Edward V. Farley, Lewis J. Haldorson, and John H. Helle. Pink salmon (Oncorhynchus gorbuscha) marine survival rates reflect early marine carbon source dependency. Progress in Oceanography, 77(2–3):194– 202, May 2008. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661108000542.

Kitidis:2017:SOC

[KBHML17] Vassilis Kitidis, Ian Brown, Nicholas Hardman-Mountford, and Nathalie Lefèvre. Surface ocean carbon dioxide during the Atlantic meridional transect (1995–2013); evidence of ocean acidification. Progress in Oceanography, 158(??):65–75, November 2017. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661116300039.

Kubryakov:2018:TST

[KBSB18] A. A. Kubryakov, A. V. Bagaev, S. V. Stanichny, and V. N. Belokopytov. Thermohaline structure, transport and evolution of the Black Sea eddies from hydrological and satellite data. Progress in Oceanography, 167(??):44–63, October 2018. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661117303695.

Kudela:2002:MPR

[KC02] R. M. Kudela and F. P. Chavez. Multi-platform remote sensing of new production in central California during the 1997–1998 El Niño. Progress in Oceanography, 54(1–4):233– 249, July 2002. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661102000514.

Kim:2015:COC

[KC15] Sung Yong Kim and Bruce D. Cornuelle. Coastal ocean climatology of temperature and salinity off the Southern California Bight: Seasonal variability, climate index correlation, and linear trend. Progress in Oceanography, 138 (part A)(??):136–157, November 2015. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661115001664.

Kohnert:2020:PKK

[KCBS20] P. C. Kohnert, A. F. Cerwenka, A. Brandt, and M. Schrödl. Pteropods from the Kuril–Kamchatka Trench and the Sea of Okhotsk (Euopisthobranchia; Gastropoda). Progress in Oceanography, 181(??):??, February 2020. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661119304392.

Khripounoff:2012:DTS

[KCL^{+}12] A. Khripounoff, P. Crassous, N. Lo Bue, B. Dennielou, and R. Silva Jacinto. Different types of sediment gravity flows detected in the var submarine canyon (northwestern Mediterranean Sea). Progress in Oceanography, 106(??):138–153, November 2012. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661112000936.

Keister:2009:DUF

[KCPM09] J. E. Keister, T. J. Cowles, W. T. Peterson, and C. A. Morgan. Do upwelling filaments result in predictable biological distributions in coastal upwelling ecosystems? Progress in Oceanography, 83(1–4):303–313, December 2009. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661109000950.

Krause:1995:RHC

[KDB95] Michael Krause, Joachim W. Dippner, and Jutta Beil. A review of hydrographic controls on the distribution of zooplankton biomass and species in the North Sea with particular reference to a survey conducted in January–March 1987. Progress in Oceanography, 35(2):81–152, ???? 1995. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0079661195000063.

Kwong:1997:TDM

[KDF97] Simon C. M. Kwong, Alan M. Davies, and Roger A. Flather. A three-dimensional model of the principal tides on the European shelf. Progress in Oceanography, 39(3):205–262, ???? 1997. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661197000141.

Killworth:2001:AAD

[KDL^{+}01] Peter D. Killworth, C. Dieterich, C. Le Provost, A. Oschlies, and J. Willebrand. Assimilation of altimetric data and mean sea surface height into an eddy-permitting model of the North Atlantic. Progress in Oceanography, 48(??):313– 335, ???? 2001. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S007966110100009X.

Kentang:1988:PSB

[Ken88] Le Kentang. A preliminary study of the basic hydrographic features and the current structures off the Changjiang River mouth in the dry season. Progress in Oceanography, 21(3– 4):387–400, ???? 1988. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/007966118890016X.

Kessler:2006:CET

[Kes06] William S. Kessler. The circulation of the eastern tropical Pacific: a review. Progress in Oceanography, 69(2–4):181– 217, May 2006. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661106000310.

Krom:2010:WEM

[KEV10] M. D. Krom, K-C. Emeis, and P. Van Cappellen. Why is the Eastern Mediterranean phosphorus limited? Progress in Oceanography, 85(3–4):236–244, June 2010. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661110000340.

Katsumata:2011:CMF

[KF11] K. Katsumata and M. Fukasawa. Changes in meridional fluxes and water properties in the Southern Hemisphere subtropical oceans between 1992/1995 and 2003/2004. Progress in Oceanography, 89(1–4):61–91, April 2011. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661110001606.

Koenig:2013:DSO

[KFC^{+}13] Samuel Koenig, Pilar Fernández, Joan B. Company, David Huertas, and Montserrat Solé. Are deep-sea organisms dwelling within a submarine canyon more at risk from anthropogenic contamination than those from the adjacent open slope? A case study of Blanes canyon (NW Mediterranean). Progress in Oceanography, 118(??):249–259, November 2013. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661113001298.

Koenig:2023:DAC

[KFC^{+}23] Zoe Koenig, Ilker Fer, Melissa Chierici, Agneta Fransson, Elizabeth Jones, and Eivind H. Kolås. Diffusive and advective cross-frontal fluxes of inorganic nutrients and dissolved inorganic carbon in the Barents Sea in autumn. Progress in Oceanography, 219(??):??, December 2023. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661123002045.

Kelley:2003:DRD

[KFG^{+}03] D. E. Kelley, H. J. S. Fernando, A. E. Gargett, J. Tanny, and E. Özsoy. The diffusive regime of double-diffusive convection. Progress in Oceanography, 56(3–4):461–481, March 2003. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661103000260.

Kuletz:2015:SSP

[KFH^{+}15] Kathy J. Kuletz, Megan C. Ferguson, Brendan Hurley, Adrian E. Gall, Elizabeth A. Labunski, and Tawna C. Morgan. Seasonal spatial patterns in seabird and marine mammal distribution in the eastern Chukchi and western Beaufort Seas: Identifying biologically important pelagic areas. Progress in Oceanography, 136(??):175–200, August 2015. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661115001093.

Kanda:2003:SAV

[KFKO03] Jota Kanda, Shiho Fujiwara, Hiroshi Kitazato, and Yoshihiro Okada. Seasonal and annual variation in the primary production regime in the central part of Sagami Bay. Progress in Oceanography, 57(1):17–29, April 2003. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S007966110300048X.

Kuhn:2015:MIN

[KFM15] Angela M. Kuhn, Katja Fennel, and Jann Paul Mattern. Model investigations of the North Atlantic spring bloom initiation. Progress in Oceanography, 138 (part A)(??):176–193, November 2015. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661115001561.

Kleisner:2017:MSD

[KFM^{+}17] Kristin M. Kleisner, Michael J. Fogarty, Sally McGee, Jonathan A. Hare, Skye Moret, Charles T. Perretti, and Vincent S. Saba. Marine species distribution shifts on the U.S. Northeast Continental Shelf under continued ocean warming. Progress in Oceanography, 153(??):24–36, April 2017. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661116301896.

Koide:1965:UUR

[KG65] Minoru Koide and Edward D. Goldberg. Uranium-234/ uranium-238 ratios in sea water. Progress in Oceanography, 3(??):173–177, ???? 1965. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0079661165900169.

Kohlbach:2023:ESI

[KGB^{+}23] Doreen Kohlbach, Lucie Goraguer, Yasemin V. Bodur, Oliver Müller, Martí Amargant-Arumí, Katalin Blix, Gunnar Bratbak, Melissa Chierici, Anna Maria Dabrowska, Ulrike Dietrich, Bente Edvardsen, Laura M. García, Rolf Gradinger, Haakon Hop, Elizabeth Jones, Øyvind Lundesgaard, Lasse M. Olsen, Marit Reigstad, Karoline Saubrekka, Agnieszka Tatarek, Józef Maria Wiktor, Anette Wold, and Philipp Assmy. Earlier sea-ice melt extends the oligotrophic summer period in the Barents Sea with low algal biomass and associated low vertical flux. Progress in Oceanography, 213(??):??, April–May 2023. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661123000617.

Koho:2008:SLO

[KGdS^{+}08] K. A. Koho, R. García, H. C. de Stigter, E. Epping, E. Koning, T. J. Kouwenhoven, and G. J. van der Zwaan.

Sedimentary labile organic carbon and pore water redox control on species distribution of benthic foraminifera: a case study from Lisbon–Setúbal Canyon (southern Portugal). Progress in Oceanography, 79(1):55–82, October 2008. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661108001456.

Kwasniewski:2010:IDH

[KGJ^{+}10] Slawomir Kwasniewski, Marta Gluchowska, Dariusz Jakubas, Katarzyna Wojczulanis-Jakubas, Wojciech Walkusz, Nina Karnovsky, Katarzyna Blachowiak-Samolyk, Malgorzata Cisek, and Lech Stempniewicz. The impact of different hydrographic conditions and zooplankton communities on provisioning Little Auks along the West coast of Spitsbergen. Progress in Oceanography, 87(1–4):72–82, October 2010. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661110000984.

Kessler:2022:OBS

[KGL22] Augustin Kessler, Nadine Goris, and Siv Kari Lauvset. Observation-based sea surface temperature trends in Atlantic large marine ecosystems. Progress in Oceanography, 208(??):??, November 2022. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661122001616.

Kosobokova:2009:BZE

[KH09] Ksenia Kosobokova and Hans-Jürgen Hirche. Biomass of zooplankton in the eastern Arctic Ocean — a base line study. Progress in Oceanography, 82(4):265–280, September 2009. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661109000561.

Kohler:2024:BTM

[KHBA^{+}24] Stephen G. Kohler, Lars-Eric Heimbürger-Boavida, Philipp Assmy, Oliver Müller, Stefan Thiele, Maria G. Digernes, Kuria Ndungu, and Murat V. Ardelan. Biotic transformation of methylmercury at the onset of the Arctic spring bloom. Progress in Oceanography, 222(??):??, March 2024.

CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661124000302.

Kerherve:1999:BDS

[KHC^{+}99] Philippe Kerhervé, Serge Heussner, Bruno Charrière, Spyros Stavrakakis, Jean-Luc Ferrand, André Monaco, and Nicole Delsaut. Biogeochemistry and dynamics of settling particle fluxes at the Antikythira Strait (Eastern Mediterranean). Progress in Oceanography, 44(4):651–675, December 1999. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661199000403.

Kessler:2022:EWB

[KHD22] William S. Kessler, Hristina G. Hristova, and Russ E. Davis. Equatorward western boundary transport from the South Pacific: Glider observations, dynamics and consequences. Progress in Oceanography, 175(??):208–225, ???? 2022. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661118303021.

Kudela:2010:UMT

[KHJ^{+}10] R. M. Kudela, M. D. A. Howard, B. D. Jenkins, P. E. Miller, and G. J. Smith. Using the molecular toolbox to compare harmful algal blooms in upwelling systems. Progress in Oceanography, 85(1–2):108–121, April 2010. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661110000182.

Kaplan:2012:SCC

[KHL12] Isaac C. Kaplan, Peter J. Horne, and Phillip S. Levin. Screening California Current fishery management scenarios using the Atlantis end-to-end ecosystem model. Progress in Oceanography, 102(??):5–18, September 2012. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661112000262.

Kuroda:1988:HPA

[KHM^{+}88] Yoshifumi Kuroda, Toshiaki Hara, Atsunobu Misumi, Gentaro Kai, and Jun Suwa. Hydrographic profiling across the East China Sea shelf edge by an underwater sliding vehicle with CTD sensors. Progress in Oceanography, 21(3–4):401– 416, ???? 1988. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661188900171.

Krumhansl:2018:EDV

[KHP^{+}18] Kira A. Krumhansl, Erica J. H. Head, Pierre Pepin, Stéphane Plourde, Nicholas R. Record, Jeffrey A. Runge, and Catherine L. Johnson. Environmental drivers of vertical distribution in diapausing Calanus copepods in the Northwest Atlantic. Progress in Oceanography, 162(??):202–222, March 2018. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661117303270.

Kavanaugh:2014:HDS

[KHS^{+}14] Maria T. Kavanaugh, Burke Hales, Martin Saraceno, Yvette H. Spitz, Angelicque E. White, and Ricardo M. Letelier. Hierarchical and dynamic seascapes: a quantitative framework for scaling pelagic biogeochemistry and ecology. Progress in Oceanography, 120(??):291–304, January 2014. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S007966111300205X.

Killworth:1976:MSP

[Kil76] Peter D. Killworth. The mixing and spreading phases of MEDOC. I. Progress in Oceanography, 7(2):59–90, ???? 1976. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0079661176900057.

Kim:2014:SLH

[KiL14] So-Young Kim and Dhong il Lim. Signatures of the late Holocene Neoglacial cold event and their marine–terrestrial linkage in the northwestern Pacific margin. Progress in Oceanography, 124(??):54–65, May 2014. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661114000366.

Kirwan:2006:DCT

[Kir06] A. D. Kirwan. Dynamics of “critical” trajectories. Progress in Oceanography, 70(2–4):448–465, August 2006. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661106000516.

Kinugasa:2005:DTM

[KIS^{+}05] Masatoshi Kinugasa, Tsunemi Ishita, Yoshiki Sohrin, Kei Okamura, Shigenobu Takeda, Jun Nishioka, and Atsushi Tsuda. Dynamics of trace metals during the subarctic Pacific iron experiment for ecosystem dynamics study (SEEDS2001). Progress in Oceanography, 64(2–4):129–147, February 2005. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661105000261.

Kitazato:2003:PSD

[Kit03] Hiroshi Kitazato. “The project Sagami” — dynamic sedimentary processes of both organic and inorganic materials at continental margins with active tectonic forcing. Progress in Oceanography, 57(1):1–2, April 2003. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661103000661.

Kivman:1997:WCD

[Kiv97] Gennady A. Kivman. Weak constraint data assimilation for tides in the Arctic Ocean. Progress in Oceanography, 40 (??):179–196, ???? 1997. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661198000056.

Koman:2022:COE

[KJH^{+}22] G. Koman, W. E. Johns, A. Houk, L. Houpert, and F. Li. Circulation and overturning in the eastern North Atlantic subpolar gyre. Progress in Oceanography, 208(??):??, November 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661122001434.

Kang:2012:RDR

[KJZ^{+}12] Young Shil Kang, Sukgeun Jung, Yury Zuenko, Ilsu Choi, and Natalia Dolganova. Regional differences in the response of mesozooplankton to oceanographic regime shifts in the northeast Asian marginal seas. Progress in Oceanography, 97–100(??):120–134, May 2012. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661111001303.

Kalimeris:2020:SSC

[KK20] Anastasios Kalimeris and Dimitris Kassis. Sea surface circulation variability in the Ionian–Adriatic Seas. Progress in Oceanography, 189(??):??, November–December 2020. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661120301919.

Kang:2000:CEC

[KKB00] Sukyung Kang, Suam Kim, and Sang-Won Bae. Changes in ecosystem components induced by climate variability off the eastern coast of the Korean Peninsula during 1960– 1990. Progress in Oceanography, 47(2–4):205–222, October 2000. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661100000434.

Kang:2004:PPF

[KKK04a] Dong-Jin Kang, Kuh Kim, and Kyung-Ryul Kim. The past, present and future of the East/Japan Sea in change: a simple moving-boundary box model approach. Progress in Oceanography, 61(2–4):175–191, May 2004. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661104000618.

Kim:2004:WMD

[KKK^{+}04b] Kuh Kim, Kyung-Ryul Kim, Young-Gyu Kim, Yang-Ki Cho, Dong-Jin Kang, Masaki Takematsu, and Yuri Volkov. Water masses and decadal variability in the East Sea (Sea of Japan). Progress in Oceanography, 61(2–4):157–174, May 2004. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661104000606.

Kitahashi:2014:BPD

[KKKS14] Tomo Kitahashi, Kiichiro Kawamura, Shigeaki Kojima, and Motohiro Shimanaga. Bathymetric patterns of α and β diversity of harpacticoid copepods at the genus level around the Ryukyu Trench, and turnover diversity between trenches around Japan. Progress in Oceanography, 123(??):54–63, April 2014. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661114000263.

Kimura:2010:IEV

[KKKY10] Shingo Kimura, Yoshiki Kato, Takashi Kitagawa, and Naoki Yamaoka. Impacts of environmental variability and global warming scenario on Pacific bluefin tuna (Thunnus orientalis) spawning grounds and recruitment habitat. Progress in Oceanography, 86(1–2):39–44, July 2010. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661110000522.

Kido:2023:MIZ

[KKNT23] Shoichiro Kido, Shota Katsura, Masami Nonaka, and Youichi Tanimoto. Mechanism and impact of zonally contrasting seasonal variations in sea-surface salinity in the North Pacific and North Atlantic oceans. Progress in Oceanography, 219(??):??, December 2023. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661123001672.

Kriest:2010:TAS

[KKO10] I. Kriest, S. Khatiwala, and A. Oschlies. Towards an assessment of simple global marine biogeochemical models of different complexity. Progress in Oceanography, 86(3–4):337–360, September 2010. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661110000844.

Kato:2003:PCS

[KKS^{+}03] Y. Kato, H. Kitazato, M. Shimanaga, T. Nakatsuka, Y. Shirayama, and T. Masuzawa. ^{210}Pb and ^{137}Cs in sediments from Sagami Bay, Japan: sedimentation rates and inventories. Progress in Oceanography, 57(1):77–95, April 2003. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661103000521.

Karageorgis:2018:PDF

[KKS^{+}18] A. P. Karageorgis, H. Kontoyiannis, S. Stavrakakis, E. Krasakopoulou, A. Gogou, A. Papadopoulos, Th. D. Kanellopoulos, G. Rousakis, E. Malinverno, M. V. Triantaphyllou, and V. Lykousis. Particle dynamics and fluxes in canyons and open slopes of the southern Cretan margin (Eastern Mediterranean). Progress in Oceanography, 169(??):33–47, December 2018. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661117302057.

Kolesnik:2019:CCC

[KKS^{+}19] Alexander N. Kolesnik, Olga N. Kolesnik, Valentina V. Sattarova, Alexander A. Karabtsov, and Elena I. Yaroshchuk. Color and chemical composition of bottom sediments from the Kuril Basin (Sea of Okhotsk) and the Kuril–Kamchatka Trench area (Northwest Pacific). Progress in Oceanography, 178(??):??, November 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661119301752.

Kim:1986:BTC

[KL86] Kuh Kim and Richard Legeckis. Branching of the Tsushima Current in 1981–83. Progress in Oceanography, 17(3–4):265– 276, ???? 1986. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661186900492.

Karl:2021:SDS

[KLB^{+}21] David M. Karl, Ricardo M. Letelier, Robert R. Bidigare, Karin M. Björkman, Matthew J. Church, John E. Dore, and Angelicque E. White. Seasonal-to-decadal scale variability in primary production and particulate matter export at Station ALOHA. Progress in Oceanography, 195(??):??, July 2021. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661121000501.

Kopp:2015:RMT

[KLC^{+}15] Dorothée Kopp, Sébastien Lefebvre, Marie Cachera, Maria Ching Villanueva, and Bruno Ernande. Reorganization of a marine trophic network along an inshore–offshore gradient due to stronger pelagic–benthic coupling in coastal areas. Progress in Oceanography, 130(??):157–171, January 2015. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S007966111400175X.

Kline:2010:SCN

[Kli10] Thomas C. Kline. Stable carbon and nitrogen isotope variation in the northern lampfish and Neocalanus, marine survival rates of pink salmon, and meso-scale eddies in the Gulf of Alaska. Progress in Oceanography, 87(1–4):49–60, October 2010. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661110001369.

Krumhardt:2017:CGC

[KLIRK17] Kristen M. Krumhardt, Nicole S. Lovenduski, M. Debora Iglesias-Rodriguez, and Joan A. Kleypas. Coccolithophore growth and calcification in a changing ocean. Progress in Oceanography, 159(??):276–295, December 2017. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661117302148. See corrigendum [Kru19].

Kone:2017:BMA

[KLP^{+}17] Vamara Koné, Christophe Lett, Pierrick Penven, Bernard Bourlès, and Sandrine Djakouré. A biophysical model of S. aurita early life history in the northern Gulf of Guinea. Progress in Oceanography, 151(??):83–96, February 2017. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661116300477.

Kettle:2008:MOP

[KM08] Helen Kettle and Chris J. Merchant. Modeling ocean primary production: Sensitivity to spectral resolution of attenuation and absorption of light. Progress in Oceanography, 78(2):135– 146, August 2008. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S007966110800089X.

Kai:2010:IME

[KM10] Emilie Tew Kai and Francis Marsac. Influence of mesoscale eddies on spatial structuring of top predators’ communities in the Mozambique Channel. Progress in Oceanography, 86(1–2):214– 223, July 2010. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661110000443.

Kolukula:2022:ICW

[KM22] Siva Srinivas Kolukula and P. L. N. Murty. Improving cyclone wind fields using deep convolutional neural networks and their application in extreme events. Progress in Oceanography, 202(??):??, March 2022. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661122000258.

King:2001:IDM

[KMB01] Jacquelynne R. King, Gordon A. McFarlane, and Richard J. Beamish. Incorporating the dynamics of marine systems into the stock assessment and management of sablefish. Progress in Oceanography, 49(??):619–639, ???? 2001. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661101000441.

Kuroki:2020:CDA

[KMF^{+}20a] Mari Kuroki, Michael J. Miller, Eric Feunteun, Pierre Sasal, Timothy Pikering, Yu-San Han, Elisabeth Faliex, Anthony Acou, Aurélie Dessier, Robert Schabetsberger, Shun Watanabe, Tatsuya Kawakami, Hiroaki Onda, Takatoshi Higuchi, Aya Takeuchi, Madoka Shimizu, Chinthaka A. Hewavitharane, Seishi Hagihara, Terumasa Taka, Shingo Kimura, Noritaka Mochioka, Tsuguo Otake, and Katsumi Tsukamoto. Corrigendum to “Distribution of anguillid leptocephali and possible spawning areas in the South Pacific Ocean” [Progr. Oceanogr. 180 (2020) 102234]. Progress in Oceanography, 183(??):??, April 2020. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661120300495. See [KMF^{+}20b].

Kuroki:2020:DAL

[KMF^{+}20b] Mari Kuroki, Michael J. Miller, Eric Feunteun, Pierre Sasal, Timothy Pikering, Yu-San Han, Elisabeth Faliex, Anthony Acou, Aurélie Dessier, Robert Schabetsberger, Shun Watanabe, Tatsuya Kawakami, Hiroaki Onda, Takatoshi Higuchi, Aya Takeuchi, Madoka Shimizu, Chinthaka A. Hewavitharane, Seishi Hagihara, Terumasa Taka, Shingo Kimura, Noritaka Mochioka, Tsuguo Otake, and Katsumi Tsukamoto. Distribution of anguillid leptocephali and possible spawning areas in the South Pacific Ocean. Progress in Oceanography, 180(??):??, January 2020. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661119304148. See corri-

gendum [KMF^{+}20a].

Kim:2009:CPB

[KMMC09] Hey-Jin Kim, Arthur J. Miller, John McGowan, and Melissa L. Carter. Coastal phytoplankton blooms in the Southern California Bight. Progress in Oceanography, 82(2):137–147, August 2009. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661109000494.

Kawatate:1988:RFC

[KMOM88] Kazuo Kawatate, Tsunehiro Miita, Yasunori Ouchi, and Shinjiro Mizuno. A report on failures of current meter moorings set east of Tsushima Island from 1983 to 1987. Progress in Oceanography, 21(3–4):319–327, ???? 1988. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0079661188900110.

Koenig:2024:WLS

[KMS^{+}24] Zoe Koenig, Morven Muilwijk, Håkon Sandven, Øyvind Lundesgaard, Philipp Assmy, Sigrid Lind, Karen M. Assmann, Melissa Chierici, Agneta Fransson, Sebastian Gerland, Elizabeth Jones, Angelika H. H. Renner, and Mats A. Granskog. From winter to late summer in the northwestern Barents Sea shelf: Impacts of seasonal progression of sea ice and upper ocean on nutrient and phytoplankton dynamics. Progress in Oceanography, 220(??):??, January 2024. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661123002173.

Koziorowska-Makuch:2023:MCS

[KMSTK23] Katarzyna Koziorowska-Makuch, Beata Szymczycha, Helmuth Thomas, and Karol Kuliński. The marine carbonate system variability in high meltwater season (Spitsbergen Fjords, Svalbard). Progress in Oceanography, 211(??):??, February 2023. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661123000204.

Kovacevic:2012:WMP

[KMU^{+}12] Vedrana Kovačević, Beniamino Bruno Manca, Laura Ursella, Katrin Schroeder, Stefano Cozzi, Mihai Burca, Elena Mauri, Riccardo Gerin, Giulio Notarstefano, and Davide Deponte. Water mass properties and dynamic conditions of the Eastern Mediterranean in June 2007. Progress in Oceanography, 104(??):59–79, October 2012. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661112000638.

Kumamoto:2011:TSV

[KMWF11] Yuichiro Kumamoto, Akihiko Murata, Shuichi Watanabe, and Masao Fukasawa. Temporal and spatial variations in bombproduced radiocarbon along BEAGLE2003 lines — revisits of WHP P06, A10, and I03/I04 in the Southern Hemisphere Oceans. Progress in Oceanography, 89(1–4):49–60, April 2011. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S007966111000159X.

Khan:2010:RIF

[KN10] Ahmed S. Khan and Barb Neis. The rebuilding imperative in fisheries: Clumsy solutions for a wicked problem? Progress in Oceanography, 87(1–4):347–356, October 2010. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661110001230. See erratum [KN11].

Khan:2011:CRI

[KN11] Ahmed S. Khan and Barb Neis. Corrigendum to “The rebuilding imperative in fisheries: Clumsy solutions for a wicked problem?” [Prog. Oceanogr. 87 (2010) 347–356]. Progress in Oceanography, 91(1):95, October 2011. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661111000140. See [KN10].

Kudo:2005:PPN

[KNI^{+}05] Isao Kudo, Yoshifumi Noiri, Keiri Imai, Yukihiro Nojiri, Jun Nishioka, and Atsushi Tsuda. Primary productivity and nitrogenous nutrient assimilation dynamics during the subarctic Pacific iron experiment for ecosystem dynamics study. Progress in Oceanography, 64(2–4):207–221, February 2005. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661105000303.

Knowlton:2004:MSS

[Kno04] Nancy Knowlton. Multiple “stable” states and the conservation of marine ecosystems. Progress in Oceanography, 60 (2–4):387–396, February 2004. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661104000369.

Kitazato:2003:LTM

[KNS^{+}03] Hiroshi Kitazato, Takeshi Nakatsuka, Motohiro Shimanaga, Jota Kanda, Wonn Soh, Yoshihisa Kato, Yoshihiro Okada, Akio Yamaoka, Toshiyuki Masuzawa, Koji Suzuki, and Yoshihisa Shirayama. Long-term monitoring of the sedimentary processes in the central part of Sagami Bay, Japan: rationale, logistics and overview of results. Progress in Oceanography, 57(1):3–16, April 2003. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661103000478.

Karakas:2009:IPA

[KNSN^{+}09] G. Karakaş, N. Nowald, C. Schäfer-Neth, M. Iversen, W. Barkmann, G. Fischer, P. Marchesiello, and R. Schlitzer. Impact of particle aggregation on vertical fluxes of organic matter. Progress in Oceanography, 83(1–4):331–341, December 2009. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661109000986.

Kharlamenko:2019:UMI

[KO19] Vladimir I. Kharlamenko and Nelly A. Odintsova. Unusual methylene-interrupted polyunsaturated fatty acids of abyssal and hadal invertebrates. Progress in Oceanography, 178(??):??, November 2019. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661119302046.

Koczy:1965:RAD

[Koc65a] F. F. Koczy. Remarks on age determination in deep-sea sediments. Progress in Oceanography, 3(??):155–171, ???? 1965. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0079661165900157.

Koczy:1965:CF

[Koc65b] Fritz Koczy. The compiler’s foreword. Progress in Oceanography, 3(??):ix, ???? 1965. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0079661165900017.

Kodama:2010:DSC

[KOhL^{+}10] Keita Kodama, Masaaki Oyama, Jeong hoon Lee, Gen Kume, Atsuko Yamaguchi, Yasuyuki Shibata, Hiroaki Shiraishi, Masatoshi Morita, Makoto Shimizu, and Toshihiro Horiguchi. Drastic and synchronous changes in megabenthic community structure concurrent with environmental variations in a eutrophic coastal bay. Progress in Oceanography, 87(1–4):157–167, October 2010. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661110001114.

Kashiwase:2014:LTV

[KON14] Haruhiko Kashiwase, Kay I. Ohshima, and Sohey Nihashi. Long-term variation in sea ice production and its relation to the intermediate water in the Sea of Okhotsk. Progress in Oceanography, 126(??):21–32, August 2014. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661114000895.

Koslow:1993:CSN

[Kos93] J. Anthony Koslow. Community structure in North Atlantic deep-sea fishes. Progress in Oceanography, 31(3):321–338, ???? 1993. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/007966119390005X.

Kosro:2002:PJE

[Kos02] P. M. Kosro. A poleward jet and an equatorward undercurrent observed off Oregon and northern California, during the 1997–98 El Niño. Progress in Oceanography, 54(1–4):343– 360, July 2002. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661102000575.

Kodama:2021:AHM

[KOT^{+}21] Taketoshi Kodama, Seiji Ohshimo, Hiroshige Tanaka, Hiroshi Ashida, Takahiko Kameda, Toshiyuki Tanabe, Makoto Okazaki, Tsuneo Ono, and Yosuke Tanaka. Abundance and habitats of marine cladocerans in the Sea of Japan over two decades. Progress in Oceanography, 194(??):??, June 2021. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661121000483.

Keister:2003:ZSV

[KP03] Julie E. Keister and William T. Peterson. Zonal and seasonal variations in zooplankton community structure off the central Oregon coast, 1998–2000. Progress in Oceanography, 57(3–4):341–361, June 2003. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661103001058.

Kim:2023:SRD

[KPM^{+}23] So-Young Kim, Taewook Park, Fabienne Marret, Éric Potvin, Kyoung-Ho Cho, Jinyoung Jung, Youngju Lee, Hyoung Sul La, Sun-Yong Ha, Seung-Kyu Kim, Sung-Ho Kang, Eun Jin Yang, Seung-Il Nam, and Jong Kuk Hong. Strong regionalism in dinoflagellate cysts in recent sediments from the Chukchi-East Siberian Seas, Arctic Ocean. Progress in Oceanography, 211(??):??, February 2023. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661123000137.

Kovac:2017:MEP

[KPSA17] Žarko Kovač, Trevor Platt, Shubha Sathyendranath, and Suzana Antunović. Models for estimating photosynthesis parameters from in situ production profiles. Progress in Oceanography, 159(??):255–266, December 2017. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661117300095.

Krug:2017:OSP

[KPSB17] Lilian Anne Krug, Trevor Platt, Shubha Sathyendranath, and Ana B. Barbosa. Ocean surface partitioning strategies using ocean colour remote sensing: a review. Progress in Oceanography, 155(??):41–53, June 2017. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661116301562.

Krug:2022:PDP

[KPSB22] Lilian Anne Krug, Trevor Platt, Shubha Sathyendranath, and Ana B. Barbosa. Patterns and drivers of phytoplankton phenology off SW Iberia: a phenoregion based perspective. Progress in Oceanography, 165(??):233–256, ???? 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S007966111830082X.

Keller:2017:EDS

[KQP^{+}17] Stefanie Keller, Antoni Quetglas, Patricia Puerta, Isabella Bitetto, Loredana Casciaro, Danila Cuccu, Antonio Esteban, Cristina Garcia, Germana Garofalo, Beatriz Guijarro, Marios Josephides, Angelique Jadaud, Evgenia Lefkaditou, Porzia Maiorano, Chiara Manfredi, Bojan Marceta, Reno Micallef, Panagiota Peristeraki, Giulio Relini, Paolo Sartor, Maria Teresa Spedicato, George Tserpes, and Manuel Hidalgo. Environmentally driven synchronies of Mediterranean cephalopod populations. Progress in Oceanography, 152(??):1–14, March 2017. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661116302038.

Krauss:1969:TFI

[Kra69] W. Krauss. Typical features of internal wave spectra. Progress in Oceanography, 5(??):95–101, ???? 1969. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0079661169900317.

Kraus:1982:OVC

[Kra82] E. B. Kraus. Ocean variability on climatic scales (foreword). Progress in Oceanography, 11(2):61–68, ???? 1982. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0079661182900039.

Kuhnz:2014:RCL

[KRHS14] Linda A. Kuhnz, Henry A. Ruhl, Christine L. Huffard, and Kenneth L. Smith. Rapid changes and long-term cycles in the benthic megafaunal community observed over 24 years in the abyssal northeast Pacific. Progress in Oceanography, 124(??): 1–11, May 2014. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S007966111400055X.

Kantha:2008:NIO

[KRL08] Lakshmi Kantha, Thaned Rojsiraphisal, and Joseph Lopez. The North Indian Ocean circulation and its variability as seen in a numerical hindcast of the years 1993–2004. Progress in Oceanography, 76(1):111–147, January 2008. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661107001978.

Kwong:2022:SSC

[KRL^{+}22] Lian E. Kwong, Tetjana Ross, Florian Lüskow, Katie R. N. Florko, and Evgeny A. Pakhomov. Spatial, seasonal, and climatic variability in mesozooplankton size spectra along a coastal-to-open ocean transect in the subarctic Northeast Pacific. Progress in Oceanography, 201(??):??, February 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661121002111.

Krumhardt:2019:CCG

[Kru19] Kristen M. Krumhardt. Corrigendum to Coccolithophore growth and calcification in a changing ocean’ [Progress in Oceanography 159 (2017) 276–295]. Progress in Oceanography, 171(??):251, February 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661119300199.

See [KLIRK17].

Kobayashi:2006:IOH

[KS06] Taiyo Kobayashi and Toshio Suga. The Indian Ocean HydroBase: a high-quality climatological dataset for the Indian Ocean. Progress in Oceanography, 68(1):75–114, January 2006. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661105001242.

Kraus:2015:SDI

[KS15] Romina Kraus and Nastjenjka Supić. Sea dynamics impacts on the macroaggregates: a case study of the 1997 mucilage event in the northern Adriatic. Progress in Oceanography, 138 (part A)(??):249–267, November 2015. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S007966111500138X.

Kohlbach:2022:MTM

[KSB^{+}22] Doreen Kohlbach, Lukas Smik, Simon T. Belt, Haakon Hop, Anette Wold, Martin Graeve, and Philipp Assmy. A multitrophic marker approach reveals high feeding plasticity in Barents Sea under-ice fauna. Progress in Oceanography, 208(??):??, November 2022. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661122001549.

Kudela:2010:RNR

[KSC10] R. M. Kudela, S. Seeyave, and W. P. Cochlan. The role of nutrients in regulation and promotion of harmful algal blooms in upwelling systems. Progress in Oceanography, 85(1–2):122– 135, April 2010. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661110000194.

Koblinsky:1984:OEC

[KSD84] C. J. Koblinsky, J. J. Simpson, and T. D. Dickey. An offshore eddy in the California Current System part II: Surface manifestation. Progress in Oceanography, 13(1):51–69, ???? 1984. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661184900053.

Kenny:2009:IAA

[KSE^{+}09] Andrew J. Kenny, Hein Rune Skjoldal, Georg H. Engelhard, Peter J. Kershaw, and James B. Reid. An integrated approach for assessing the relative significance of human pressures and environmental forcing on the status of large marine ecosystems. Progress in Oceanography, 81(1–4):132–148, April 2009. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661109000263.

Kohlbach:2017:SLP

[KSG^{+}17] Doreen Kohlbach, Fokje L. Schaafsma, Martin Graeve, Benoit Lebreton, Benjamin Allen Lange, Carmen David, Martina Vortkamp, and Hauke Flores. Strong linkage of polar cod (Boreogadus saida) to sea ice algae-produced carbon: Evidence from stomach content, fatty acid and stable isotope analyses. Progress in Oceanography, 152(??):62–74, March 2017. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661116301495.

Kim:2015:UIV

[KSK^{+}15] Jung Jin Kim, William Stockhausen, Suam Kim, Yang-Ki Cho, Gwang-Ho Seo, and Joon-Soo Lee. Understanding interannual variability in the distribution of, and transport processes affecting, the early life stages of Todarodes pacificus using behavioral-hydrodynamic modeling approaches. Progress in Oceanography, 138(??):571–583, November 2015. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661115000634.

Kobayashi:2021:OFS

[KSK21] Taiyo Kobayashi, Kanako Sato, and Brian A. King. Observed features of salinity bias with negative pressure dependency for measurements by SBE 41CP and SBE 61 CTD sensors on deep profiling floats. Progress in Oceanography, 198(??):??, November 2021. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661121001701.

Koski:2021:ACC

[KSKN21] Marja Koski, Rasmus Swalethorp, Sanne Kjellerup, and Torkel Gissel Nielsen. Aggregate-colonizing copepods in a glacial fjord: Population dynamics, vertical distribution and allometric scaling of growth and mortality rates of Microsetella norvegica and Oncaea spp. Progress in Oceanography, 197(??):??, September–October 2021. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661121001543.

Kurian:2023:SSV

[KSP^{+}23] Siby Kurian, Damodar M. Shenoy, Akhi V. P, Pratima M. Kessarkar, Mangesh Gauns, Suhas S. Shetye, Muhamed Kabeer, Apsara P. Vijayan, Anand Methar, Supriya Karapurkar, A. V. Chndrasekhararao, and S. W. A. Naqvi. Spatial and seasonal variations in the particulate sinking flux in the Bay of Bengal. Progress in Oceanography, 211(??):??, February 2023. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661123000265.

Krasakopoulou:1999:ONF

[KSPK99] E. Krasakopoulou, E. Souvermezoglou, A. Pavlidou, and H. Kontoyiannis. Oxygen and nutrient fluxes through the Straits of the Cretan Arc (March 1994–January 1995). Progress in Oceanography, 44(4):601–624, December 1999. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661199000476.

Kiriakoulakis:2001:COC

[KSR^{+}01] Konstadinos Kiriakoulakis, Edward Stutt, Steven J. Rowland, Annick Vangriesheim, Richard S. Lampitt, and George A. Wolff. Controls on the organic chemical composition of settling particles in the Northeast Atlantic Ocean. Progress in Oceanography, 50(??):65–87, ???? 2001. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661101000489.

Kobari:2023:TSP

[KSS^{+}23] Toru Kobari, Maharu Shinyashiki, Kanako Saito, Gen Kume, Mutsuo Ichinomiya, Tomohiro Komorita, Hiroomi Miyamoto, Yuji Okazaki, Fumihiro Makino, Ryuji Fukuda, Fujio Hyodo, and Maki Noguchi-Aita. Trophic sources and pathways of mesozooplankton and fish larvae in the Kuroshio and its neighboring waters based on stable isotope ratios of carbon and nitrogen. Progress in Oceanography, 210(??):??, January 2023. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661122002117.

Kobari:2003:FRI

[KST03] Toru Kobari, Akiyoshi Shinada, and Atsushi Tsuda. Functional roles of interzonal migrating mesozooplankton in the western subarctic Pacific. Progress in Oceanography, 57(3–4):279– 298, June 2003. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661103001022.

Kappes:2010:HAT

[KST^{+}10] Michelle A. Kappes, Scott A. Shaffer, Yann Tremblay, David G. Foley, Daniel M. Palacios, Patrick W. Robinson, Steven J. Bograd, and Daniel P. Costa. Hawaiian albatrosses track interannual variability of marine habitats in the North Pacific. Progress in Oceanography, 86(1–2):246–260, July 2010. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661110000467.

Karstensen:2008:OMZ

[KSV08] Johannes Karstensen, Lothar Stramma, and Martin Visbeck. Oxygen minimum zones in the eastern tropical Atlantic and Pacific Oceans. Progress in Oceanography, 77(4):331–350, June 2008. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661108000670.

Kallianiotis:2000:DFM

[KSVT00] A. Kallianiotis, K. Sophronidis, P. Vidoris, and A. Tselepides. Demersal fish and megafaunal assemblages on the Cretan continental shelf and slope (NE Mediterranean): seasonal variation in species density, biomass and diversity. Progress in Oceanography, 46(2–4):429–455, August 2000. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661100000288.

Kobari:2019:ETO

[KSY^{+}19] Toru Kobari, Akash R. Sastri, Lidia Yebra, Hui Liu, and Russell R. Hopcroft. Evaluation of trade-offs in traditional methodologies for measuring metazooplankton growth rates: Assumptions, advantages and disadvantages for field applications. Progress in Oceanography, 178(??):??, November 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S007966111930120X.

Kantha:1997:GBT

[KT97] Lakshmi H. Kantha and Craig C. Tierney. Global baroclinic tides. Progress in Oceanography, 40(??):163–178, ???? 1997. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661197000281.

Klitgaard:2004:DSC

[KT04] A. B. Klitgaard and O. S. Tendal. Distribution and species composition of mass occurrences of large-sized sponges in the northeast Atlantic. Progress in Oceanography, 61(1):57–98, April 2004. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661104000680.

Kontoyiannis:1999:WFT

[KTB^{+}99] Harilaos Kontoyiannis, Alexander Theocharis, Efstathios Balopoulos, Soterios Kioroglou, Vassilios Papadopoulos, Michael Collins, Antonios F. Velegrakis, and Athanasia Iona. Water fluxes through the Cretan Arc Straits, Eastern Mediterranean Sea: March 1994 to June 1995. Progress in Oceanography, 44(4):511–529, December 1999. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661199000440.

Kerkar:2021:CPP

[KTH^{+}21] Anvita U. Kerkar, Sarat C. Tripathy, David J. Hughes, P. Sabu, Sudarsana Rao Pandi, Amit Sarkar, and Manish Tiwari. Characterization of phytoplankton productivity and bio-optical variability in a polar marine ecosystem. Progress in Oceanography, 195(??):??, July 2021. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661121000604.

Kodama:2022:SDF

[KTIT22] Taketoshi Kodama, Atsushi Tawa, Taiki Ishihara, and Yosuke Tanaka. Similarities of distributions and feeding habits between bullet tuna, Auxis rochei, and Pacific bluefin tuna, Thunnus orientalis, larvae in the southern Sea of Japan. Progress in Oceanography, 202(??):??, March 2022. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661122000209.

Kanna:2014:IMN

[KTN14] Naoya Kanna, Takenobu Toyota, and Jun Nishioka. Iron and macro-nutrient concentrations in sea ice and their impact on the nutritional status of surface waters in the southern Okhotsk Sea. Progress in Oceanography, 126(??):44–57, August 2014. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661114000603.

Kuroda:2022:ICO

[KTW^{+}22] Hiroshi Kuroda, Yuko Toya, Tsuyoshi Watanabe, Jun Nishioka, Daisuke Hasegawa, Yukiko Taniuchi, and Akira Kuwata. Influence of coastal Oyashio water on massive spring diatom blooms in the Oyashio area of the North Pacific Ocean. Progress in Oceanography, 175(??):328–344, ???? 2022. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661118302465.

Kunze:2003:ROS

[Kun03] Eric Kunze. A review of oceanic salt-fingering theory. Progress in Oceanography, 56(3–4):399–417, March 2003. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661103000272.

Kemp:2013:HDP

[KV13] Alan E. S. Kemp and Tracy A. Villareal. High diatom production and export in stratified waters — a potential negative feedback to global warming. Progress in Oceanography, 119(??):4–23, December 2013. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661113000852.

Kemp:2018:CDM

[KV18] Alan E. S. Kemp and Tracy A. Villareal. The case of the diatoms and the muddled mandalas: Time to recognize diatom adaptations to stratified waters. Progress in Oceanography, 167(??):138–149, October 2018. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661118301095.

Kvinge:1969:MOB

[Kvi69] Thor Kvinge. Moored oceanographic buoys: Reliability, representativity and application. Progress in Oceanography, 5 (??):107–109, ???? 1969. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0079661169900330.

Kumar:2006:OSW

[KVLA06] Ajoy Kumar, A. Valle-Levinson, and L. P. Atkinson. Overrunning of shelf water in the southern Mid-Atlantic Bight. Progress in Oceanography, 70(2–4):213–232, August 2006. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661106000887.

Koski:2020:MPU

[KVNT20] Marja Koski, Bellineth Valencia, Rebekah Newstead, and Christine Thiele. The missing piece of the upper mesopelagic carbon budget? Biomass, vertical distribution and feeding of aggregate-associated copepods at the PAP site. Progress in Oceanography, 181(??):??, February 2020. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661119304239.

Kawaguchi:2020:NII

[KWI20] Yusuke Kawaguchi, Taku Wagawa, and Yosuke Igeta. Nearinertial internal waves and multiple-inertial oscillations trapped by negative vorticity anomaly in the central Sea of Japan. Progress in Oceanography, 181(??):??, February 2020. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661119304203.

Klemas:2014:SDO

[KY14] Victor Klemas and Xiao-Hai Yan. Subsurface and deeper ocean remote sensing from satellites: an overview and new results. Progress in Oceanography, 122(??):1–9, March 2014. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661113002310.

Kieke:2015:SLS

[KY15] Dagmar Kieke and Igor Yashayaev. Studies of Labrador Sea water formation and variability in the subpolar North Atlantic in the light of international partnership and collaboration. Progress in Oceanography, 132(??):220–232, March 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661114002201.

Katsumata:2023:DIM

[KY23] Katsuro Katsumata and Kaihe Yamazaki. Diapycnal and isopycnal mixing along the continental rise in the Australian– Antarctic Basin. Progress in Oceanography, 211(??):??, February 2023. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661123000228.

Kuwae:2017:MCM

[KYS^{+}17] Michinobu Kuwae, Masanobu Yamamoto, Takuya Sagawa, Ken Ikehara, Tomohisa Irino, Keiji Takemura, Hidetaka Takeoka, and Takashige Sugimoto. Multidecadal, centennial, and millennial variability in sardine and anchovy abundances in the western North Pacific and climate–fish linkages during the late Holocene. Progress in Oceanography, 159(??):86–98, December 2017. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661116301604.

Katz:2016:SCN

[KYT^{+}16] Timor Katz, Gitai Yahel, Verena Tunnicliffe, Barak Herut, Frank Whitney, Paul V. R. Snelgrove, and Boaz Lazar. The silica cycle in a Northeast Pacific fjord; the role of biological resuspension. Progress in Oceanography, 147(??):10–21, September 2016. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661116301239.

Kurten:2019:SMM

[KZD^{+}19] Benjamin Kürten, Nikolaos D. Zarokanellos, Reny P. Devassy, Mohsen M. El-Sherbiny, Ulrich Struck, Douglas G. Capone, Isabelle K. Schulz, Ali M. Al-Aidaroos, Xabier Irigoien, and Burton H. Jones. Seasonal modulation of mesoscale processes alters nutrient availability and plankton communities in the Red Sea. Progress in Oceanography, 173(??):238–255, April 2019. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661117301829.

Kase:1985:CFE

[KZSH85] R. H. Käse, W. Zenk, T. B. Sanford, and W. Hiller. Currents, fronts and eddy fluxes in the Canary Basin. Progress in Oceanography, 14(??):231–257, ???? 1985. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0079661185900138.

Lavaniegos:2012:IVK

[LAA12] Bertha E. Lavaniegos and Israel Ambriz-Arreola. Interannual variability in krill off Baja California in the period 1997– 2005. Progress in Oceanography, 97–100(??):164–173, May 2012. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661111001261.

Lopez-Aviles:2024:CCS

[LABD^{+}24] Brandon López-Aviles, Emilio Beier, Rodrigo Duran, José Gómez-Valdés, Rubén Castro, and Laura Sánchez-Velasco. The California Current System off Baja California Sur. Progress in Oceanography, 222(??):??, March 2024. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661124000314.

LaCasce:2008:SLO

[LaC08] J. H. LaCasce. Statistics from Lagrangian observations. Progress in Oceanography, 77(1):1–29, April 2008. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661108000232.

Lambert:2018:DST

[LAD^{+}18] C. Lambert, M. Authier, M. Doray, G. Dorémus, J. Spitz, and V. Ridoux. Decadal stability in top predator habitat preferences in the Bay of Biscay. Progress in Oceanography, 166(??):109–120, September 2018. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661117300289.

Lima:2023:MDE

[LAGM^{+}23] Françoise D. Lima, Luis Enrique Angeles-Gonzalez, Hugulay Maia, Tatiana S. Leite, Miguel Cahuich-López, Ismael Mariño-Tapia, Mariana L. Santana-Cisneros, Pedro-Luis Ardisson, and Sergio M. Q. Lima. Molecular data, ecological niche, and dispersal models reveal a trans-Atlantic shallow-water octopus species. Progress in Oceanography, 213(??):??, April–May 2023. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661123000629.

Lett:2010:BMI

[LAHI10] Christophe Lett, Sakina-Dorothée Ayata, Martin Huret, and Jean-Olivier Irisson. Biophysical modelling to investigate the effects of climate change on marine population dispersal and connectivity. Progress in Oceanography, 87(1–4):106–113, October 2010. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661110001151.

Li:2010:PHS

[LAP10] Lingbo Li, Cameron Ainsworth, and Tony Pitcher. Presence of harbour seals (Phoca vitulina) may increase exploitable fish biomass in the Strait of Georgia. Progress in Oceanography, 87(1–4):235–241, October 2010. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661110001163.

Lavaniegos:2009:IME

[Lav09] Bertha E. Lavaniegos. Influence of a multiyear event of low salinity on the zooplankton from Mexican eco-regions of the California Current. Progress in Oceanography, 83 (1–4):369–375, December 2009. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661109000780.

Laws:2004:EFS

[Law04] Edward Laws. Export flux and stability as regulators of community composition in pelagic marine biological communities: Implications for regime shifts. Progress in Oceanography, 60(2–4):343–354, February 2004. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661104000345.

Lynn:2002:DEN

[LB02] R. J. Lynn and S. J. Bograd. Dynamic evolution of the 1997–1999 El Niño–La Niña cycle in the southern California Current System. Progress in Oceanography, 54(1–4):59– 75, July 2002. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661102000435.

Lohmann:2014:OPO

[LB14] Rainer Lohmann and Igor M. Belkin. Organic pollutants and ocean fronts across the Atlantic Ocean: a review. Progress in Oceanography, 128(??):172–184, November 2014. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S007966111400144X.

Lins:2020:CBB

[LB20] Lidia Lins and Angelika Brandt. Comparability between box-corer and epibenthic-sledge data on higher taxon level: a case study based on deep-sea samples from the NW Pacific. Progress in Oceanography, 182(??):??, March 2020. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661120300112.

Logerwell:2015:FCA

[LBC^{+}15] E. Logerwell, M. Busby, C. Carothers, S. Cotton, J. Duffy-Anderson, E. Farley, P. Goddard, R. Heintz, B. Holladay, J. Horne, S. Johnson, B. Lauth, L. Moulton, D. Neff, B. Norcross, S. Parker-Stetter, J. Seigle, and T. Sformo. Fish communities across a spectrum of habitats in the western Beaufort Sea and Chukchi Sea. Progress in Oceanography, 136(??):115– 132, August 2015. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S007966111500110X.

LeCorre:2023:BMC

[LBC^{+}23] Nicolas Le Corre, Catherine E. Brennan, Joël Chassé, Catherine L. Johnson, Diane Lavoie, Jean-Philippe Paquin, Nancy Soontiens, and Stéphane Plourde. A biophysical model of Calanus hyperboreus in the Gulf of St. Lawrence: Interannual variability in phenology and circulation drive the timing and location of right whale foraging habitat in spring and early summer. Progress in Oceanography, 219(??):??, December 2023. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661123001957.

Luo:2011:IVS

[LBD11] Hao Luo, Annalisa Bracco, and Emanuele Di Lorenzo. The interannual variability of the surface eddy kinetic energy in the Labrador Sea. Progress in Oceanography, 91(3):295–311, November 2011. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661111000152.

Lozier:2022:OWP

[LBF^{+}22] M. Susan Lozier, Amy S. Bower, Heather H. Furey, Kimberley L. Drouin, Xiaobiao Xu, and Sijia Zou. Overflow water pathways in the North Atlantic. Progress in Oceanography, 208(??):??, November 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661122001331.

Leetmaa:1987:HCE

[LBH^{+}87] A. Leetmaa, D. W. Behringer, A. Huyer, R. L. Smith, and J. Toole. Hydrographic conditions in the Eastern Pacific before, during and after the 1982/83 El Niño. Progress in Oceanography, 19(1):1–47, ???? 1987. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0079661187900024.

Li:2021:PSC

[LBH^{+}21] Teng Li, Yan Bai, Xianqiang He, Bangyi Tao, Xiaoyan Chen, Fang Gong, and Tianyu Wang. Phytoplankton size classes changed oppositely over shelf and basin areas of the South China Sea during 2003–2018. Progress in Oceanography, 191(??):??, February 2021. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661120302317.

Lampitt:2001:MSA

[LBK^{+}01] R. S. Lampitt, B. J. Bett, K. Kiriakoulakis, E. E. Popova, O. Ragueneau, A. Vangriesheim, and G. A. Wolff. Material supply to the abyssal seafloor in the Northeast Atlantic. Progress in Oceanography, 50(??):27–63, ???? 2001. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661101000477.

Lalande:2013:IWA

[LBNBM13] Catherine Lalande, Eduard Bauerfeind, Eva-Maria Nöthig, and Agnieszka Beszczynska-Möller. Impact of a warm anomaly on export fluxes of biogenic matter in the eastern Fram Strait. Progress in Oceanography, 109(??):70–77, February 2013. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661112001152.

Liu:2015:LSF

[LBP15] Hui Liu, Hongsheng Bi, and William T. Peterson. Largescale forcing of environmental conditions on subarctic copepods in the northern California Current system. Progress in Oceanography, 134(??):404–412, May 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661115000555.

Leles:2021:DPE

[LBP^{+}21] Suzana Gonçalves Leles, Jorn Bruggeman, Luca Polimene, Jerry Blackford, Kevin J. Flynn, and Aditee Mitra. Differences in physiology explain succession of mixoplankton functional types and affect carbon fluxes in temperate seas. Progress in Oceanography, 190(??):??, January 2021. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661120302160.

Leonard:2001:IMP

[LBSP01] Carrie L. Leonard, Robert R. Bidigare, Michael P. Seki, and Jeffrey J. Polovina. Interannual mesoscale physical and biological variability in the North Pacific central gyre. Progress in Oceanography, 49(??):227–244, ???? 2001. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661101000246.

Loos:2010:IOP

[LC10] Eduardo A. Loos and Maycira Costa. Inherent optical properties and optical mass classification of the waters of the Strait of Georgia, British Columbia, Canada. Progress in Oceanography, 87(1–4):144–156, October 2010. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661110001138.

Landaeta:2012:SAV

[LC12] Mauricio F. Landaeta and Leonardo R. Castro. Seasonal and annual variation in Chilean hake Merluccius gayi spawning locations and egg size off central Chile. Progress in Oceanography, 92–95(??):166–177, January 2012. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661111000565.

Lie:2016:SCP

[LC16] Heung-Jae Lie and Cheol-Ho Cho. Seasonal circulation patterns of the Yellow and East China Seas derived from satellite-tracked drifter trajectories and hydrographic observations. Progress in Oceanography, 146(??):121–141, August 2016. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661116300076.

Li:2022:ASI

[LC22] Siqi Li and Changsheng Chen. Air–sea interaction processes during Hurricane Sandy: Coupled WRF-FVCOM model simulations. Progress in Oceanography, 206(??):??, August 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661122001148.

Lluch-Cota:2007:GCR

[LCANAS^{+}07] Salvador E. Lluch-Cota, Eugenio A. Aragón-Noriega, Francisco Arreguín-Sánchez, David Aurioles-Gamboa, J. Jesús Bautista-Romero, Richard C. Brusca, Rafael Cervantes-Duarte, Roberto Cortés-Altamirano, Pablo Del-Monte-Luna, Alfonso Esquivel-Herrera, Guillermo Fernández, Michel E. Hendrickx, Sergio Hernández-Vázquez, Hugo Herrera-Cervantes, Mati Kahru, Miguel Lavín, Daniel Lluch-Belda, Daniel B. Lluch-Cota, Juana López-Martínez, Silvio G. Marinone, Manuel O. Nevárez-Martínez, Sofia Ortega-García, Eduardo Palacios-Castro, Alejandro Parés-Sierra, Germán Ponce-Díaz, Mauricio Ramírez-Rodríguez, Cesar A. Salinas-Zavala, Richard A. Schwartzlose, and Arturo P. Sierra-Beltrán. The Gulf of California: Review of ecosystem status and sustainability challenges. Progress in Oceanography, 73(1):1–26, April 2007. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661107000134.

Lehmuskero:2018:LPM

[LCB18] Anni Lehmuskero, Matilde Skogen Chauton, and Tobias Boström. Light and photosynthetic microalgae: a review of cellular- and molecular-scale optical processes. Progress in Oceanography, 168(??):43–56, November 2018. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661118301058.

Lassalle:2014:ATS

[LCBN14] G. Lassalle, T. Chouvelon, P. Bustamante, and N. Niquil. An assessment of the trophic structure of the Bay of Biscay continental shelf food web: Comparing estimates derived from an ecosystem model and isotopic data. Progress in Oceanography, 120(??):205–215, January 2014. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661113001729.

Lipsen:2007:STV

[LCGH07] M. S. Lipsen, D. W. Crawford, J. Gower, and P. J. Harrison. Spatial and temporal variability in coccolithophore abundance and production of PIC and POC in the NE subarctic Pacific during El Niño (1998), La Niña (1999) and 2000. Progress in Oceanography, 75(2):304–325, October 2007. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S007966110700153X.

Lim:2007:RSA

[LCJ^{+}07] D. I. Lim, J. Y. Choi, H. S. Jung, K. C. Rho, and K. S. Ahn. Recent sediment accumulation and origin of shelf mud deposits in the Yellow and East China Seas. Progress in Oceanography, 73(2):145–159, May 2007. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661107000444.

Li:2017:TTR

[LCJ^{+}17] Qi Li, Min Chen, Renming Jia, Jian Zeng, Hui Lin, Minfang Zheng, and Yusheng Qiu. Transit time of river water in the Bering and Chukchi Seas estimated from δ ^{18}O and radium isotopes. Progress in Oceanography, 159(??):115–129, December 2017. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S007966111730037X.

Lluch-Cota:2010:CCG

[LCPSMR^{+}10] Salvador E. Lluch-Cota, Alejandro Parés-Sierra, Víctor O. Magaña-Rueda, Francisco Arreguín-Sánchez, Gastón Bazzino, Hugo Herrera-Cervantes, and Daniel Lluch-Belda. Changing climate in the Gulf of California. Progress in Oceanography, 87(1–4):114–126, October 2010. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661110001175.

Levitus:1993:DNP

[LCR^{+}93] Sydney Levitus, Margarita E. Conkright, Joseph L. Reid, Raymond G. Najjar, and Arnold Mantyla. Distribution of nitrate, phosphate and silicate in the world oceans. Progress in Oceanography, 31(3):245–273, ???? 1993. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/007966119390003V.

Liu:2024:ONI

[LCZ^{+}24] Kun Liu, Xu Chen, Peng Zhan, Lianglong Da, Hui Wang, Wuhong Guo, Jingyi Liu, Lei Chen, Bing Liu, Guandong Gao, and Deyan Tian. Observations of near-inertial internal wave amplification and enhanced mixing after surface reflection. Progress in Oceanography, 220(??):??, January 2024. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661123002203.

Lanksbury:2007:DTP

[LDAM^{+}07] Jennifer A. Lanksbury, Janet T. Duffy-Anderson, Kathryn L. Mier, Morgan S. Busby, and Phyllis J. Stabeno. Distribution and transport patterns of northern rock sole, Lepidopsetta polyxystra, larvae in the southeastern Bering Sea.

Progress in Oceanography, 72(1):39–62, January 2007. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661106001145.

Lavaleye:2002:CBM

[LDB^{+}02] M. S. S. Lavaleye, G. C. A. Duineveld, E. M. Berghuis, A. Kok, and R. Witbaard. A comparison between the megafauna communities on the N.W. Iberian and Celtic continental margins — effects of coastal upwelling? Progress in Oceanography, 52(??):459–476, ???? 2002. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661102000198.

Llope:2020:GCZ

[LdCSB^{+}20] Marcos Llope, Gustavo F. de Carvalho-Souza, Francisco Baldó, Carmen González-Cabrera, María Paz Jiménez, Priscilla Licandro, and César Vilas. Gulf of Cadiz zooplankton: Community structure, zonation and temporal variation. Progress in Oceanography, 186(??):??, July 2020. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S007966112030118X.

Lheureux:2022:BDC

[LDD^{+}22] A. Lheureux, V. David, Y. Del Amo, D. Soudant, I. Auby, F. Ganthy, H. Blanchet, M-A. Cordier, L. Costes, S. Ferreira, L. Mornet, A. Nowaczyk, M. Parra, F. D’Amico, L. Gouriou, C. Meteigner, H. Oger-Jeanneret, L. Rigouin, M. Rumebe, M-P. Tournaire, F. Trut, G. Trut, and N. Savoye. Bidecadal changes in nutrient concentrations and ratios in marine coastal ecosystems: the case of the Arcachon bay, France. Progress in Oceanography, 201(??):??, February 2022. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661122000027.

LePennec:1990:NSH

[LDH90] Marcel Le Pennec, Anne Donval, and Angèle Herry. Nutritional strategies of the hydrothermal ecosystem bivalves. Progress in Oceanography, 24(??):71–80, ???? 1990. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0079661190900203.

Lu:2020:ARO

[LDHW20] Kofan Lu, Seth Danielson, Katherine Hedstrom, and Thomas Weingartner. Assessing the role of oceanic heat fluxes on ice ablation of the central Chukchi Sea Shelf. Progress in Oceanography, 184(??):??, May 2020. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661120300525.

Ludwig:2009:RDW

[LDMH09] Wolfgang Ludwig, Egon Dumont, Michel Meybeck, and Serge Heussner. River discharges of water and nutrients to the Mediterranean and Black Sea: Major drivers for ecosystem changes during past and future decades? Progress in Oceanography, 80(3–4):199–217, March 2009. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661109000020.

Lins:2015:NCC

[LdSH^{+}15] Lidia Lins, Maria Cristina da Silva, Freija Hauquier, André Morgado Esteves, and Ann Vanreusel. Nematode community composition and feeding shaped by contrasting productivity regimes in the Southern Ocean. Progress in Oceanography, 134(??):356–369, May 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661115000476.

Liu:2022:CRZ

[LEDR^{+}22] Hui Liu, Nehad Nour El-Din, Gilbert Rowe, Mohsin Al-Ansi, Chih-Lin Wei, Yousria Soliman, Clifton Nunnally, Antonietta Quigg, Ibrahim S. Al-Ansari, Ibrahim Al-Maslamani, and Mohamed A. Abdel-Moati. Characteristics and renewal of zooplankton communities under extreme environmental stresses in the oligotrophic hypersaline Arabian Gulf. Progress in Oceanography, 201(??):??, February 2022. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661121001282.

Legutke:1991:NER

[Leg91] Stephanie Legutke. Numerical experiments relating to the “Great Salinity Anomaly” of the seventies in the Greenland and Norwegian seas. Progress in Oceanography, 27(3–4):341– 363, ???? 1991. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/007966119190028K.

Lehodey:2001:PET

[Leh01] Patrick Lehodey. The pelagic ecosystem of the tropical Pacific Ocean: dynamic spatial modelling and biological consequences of ENSO. Progress in Oceanography, 49(??):439– 468, ???? 2001. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661101000350.

Letherman:1987:BSR

[Let87] Stephen P. Letherman. Beach and shoreface response to sea-level rise: Ocean City, Maryland, U.S.A. Progress in Oceanography, 18(1–4):139–149, ???? 1987. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0079661187900309.

Levitus:1988:DPD

[Lev88] Sydney Levitus. Decadal and pentadal distributions of hydrographic stations at 1000m depth for the World Ocean. Progress in Oceanography, 20(2):83–101, ???? 1988. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0079661188900481.

Langehaug:2012:CPD

[LF12] Helene R. Langehaug and Eva Falck. Changes in the properties and distribution of the intermediate and deep waters in the Fram Strait. Progress in Oceanography, 96(1):57–76, April 2012. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661111001170.

Lavin:2006:RET

[LFA^{+}06] M. F. Lavín, P. C. Fiedler, J. A. Amador, L. T. Ballance, J. Färber-Lorda, and A. M. Mestas-Nuñez. A review of eastern tropical Pacific oceanography: Summary. Progress in Oceanography, 69(2–4):391–398, May 2006. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661106000395.

Lopez-Fernandez:2013:BCS

[LFBP^{+}13] Pilar Lopez-Fernandez, Silvia Bianchelli, Antonio Pusceddu, Antoni Calafat, Roberto Danovaro, and Miquel Canals. Bioavailable compounds in sinking particulate organic matter, Blanes Canyon, NW Mediterranean Sea: Effects of a large storm and sea surface biological processes. Progress in Oceanography, 118(??):108–121, November 2013. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661113001353.

Li:2018:ENC

[LFBR^{+}18] Xuefeng Li, Debany Fonseca-Batista, Nathalie Roevros, Frank Dehairs, and Lei Chou. Environmental and nutrient controls of marine nitrogen fixation. Progress in Oceanography, 167(??):125–137, October 2018. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661117303592.

Li:2015:STP

[LFC^{+}15] Yun Li, Paula S. Fratantoni, Changsheng Chen, Jonathan A. Hare, Yunfang Sun, Robert C. Beardsley, and Rubao Ji. Spatio-temporal patterns of stratification on the Northwest Atlantic shelf. Progress in Oceanography, 134(??):123–137, May 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661115000051.

Lopez-Fernandez:2013:MDP

[LFCSV^{+}13] Pilar Lopez-Fernandez, Antoni Calafat, Anna Sanchez-Vidal, Miquel Canals, M. Mar Flexas, Jordi Cateura, and Joan B. Company. Multiple drivers of particle fluxes in the Blanes submarine canyon and southern open slope: Results of a year round experiment. Progress in Oceanography, 118(??):95–107, November 2013. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661113001420.

Link:2010:NUA

[LFG10] Jason S. Link, Elizabeth A. Fulton, and Robert J. Gamble. The northeast US application of ATLANTIS: a full system model exploring marine ecosystem dynamics in a living marine resource management context. Progress in Oceanography, 87(1–4):214–234, October 2010. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661110001321.

Liccardo:2013:RDC

[LFI^{+}13] Antonella Liccardo, Annalisa Fierro, Daniele Iudicone, Pascale Bouruet-Aubertot, and Laurent Dubroca. Response of the deep chlorophyll maximum to fluctuations in vertical mixing intensity. Progress in Oceanography, 109(??):33–46, February 2013. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661112001139.

Li:2022:NEC

[LG22] Junlu Li and Jianping Gan. On the north equatorial current spatiotemporal modes and responses in the western boundary currents. Progress in Oceanography, 205(??):??, July 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661122000817.

Li:2023:HFD

[LG23] Junlu Li and Jianping Gan. How the forcing dynamics of the western boundary currents in the Pacific respond to the North Equatorial Current. Progress in Oceanography, 210(??): ??, January 2023. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661122002099.

Lu:2020:NTD

[LGD^{+}20] Zhongming Lu, Jianping Gan, Minhan Dai, Xiaozheng Zhao, and Chiwing Rex Hui. Nutrient transport and dynamics in the South China Sea: a modeling study.

Progress in Oceanography, 183(??):??, April 2020. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661120300471.

Louis:2018:ANS

[LGG18] J. Louis, F. Gazeau, and C. Guieu. Atmospheric nutrients in seawater under current and high pCO_{2} conditions after Saharan dust deposition: Results from three minicosm experiments. Progress in Oceanography, 163(??):40–49, April 2018. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661117300666.

Liu:2021:PCD

[LGH^{+}21] Zhiqiang Liu, Jianping Gan, Jianyu Hu, Hui Wu, Zhongya Cai, and Yongfei Deng. Progress on circulation dynamics in the East China Sea and southern Yellow Sea: Origination, pathways, and destinations of shelf currents. Progress in Oceanography, 193(??):??, April–May 2021. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661121000410.

Landry:1993:DIE

[LGK^{+}93] Michael R. Landry, Dian J. Gifford, David L. Kirchman, Patricia A. Wheeler, and Bruce C. Monger. Direct and indirect effects of grazing by Neocalanus plumchrus on plankton community dynamics in the subarctic Pacific. Progress in Oceanography, 32(??):239–258, ???? 1993. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0079661193900167.

Leroux:2018:CLD

[LGL^{+}18] R. Leroux, G. Gregori, K. Leblanc, F. Carlotti, M. Thyssen, M. Dugenne, M. Pujo-Pay, P. Conan, M.-P. Jouandet, N. Bhairy, and L. Berline. Combining laser diffraction, flow cytometry and optical microscopy to characterize a nanophytoplankton bloom in the Northwestern Mediterranean. Progress in Oceanography, 163(??):248–259, April 2018. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661117300976.

Levin:2002:BPP

[LGR^{+}02] L. Levin, D. Gutiérrez, A. Rathburn, C. Neira, J. Sellanes, P. Muñoz, V. Gallardo, and M. Salamanca. Benthic processes on the Peru margin: a transect across the oxygen minimum zone during the 1997–98 El Niño. Progress in Oceanography, 53(1):1–27, April 2002. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661102000228.

Li:2020:PPC

[LGZ^{+}20] Qian Li, Xianghui Guo, Weidong Zhai, Yi Xu, and Minhan Dai. Partial pressure of CO_{2} and air–sea CO_{2} fluxes in the South China Sea: Synthesis of an 18-year dataset. Progress in Oceanography, 182(??):??, March 2020. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661120300100.

Long:2022:NSN

[LGZL22] Yu Long, Xinyu Guo, Xiao-Hua Zhu, and Zhiyuan Li. Nutrient streams in the North Pacific. Progress in Oceanography, 202(??):??, March 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661122000180.

Liao:2022:STG

[LGZW22] Fanglou Liao, Guandong Gao, Peng Zhan, and Yan Wang. Seasonality and trend of the global upper-ocean vertical velocity over 1998–2017. Progress in Oceanography, 204(??): ??, June 2022. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661122000659.

Longhurst:1989:BPP

[LH89] Alan R. Longhurst and W. Glen Harrison. The biological pump: Profiles of plankton production and consumption in the upper ocean. Progress in Oceanography, 22(1):47–123, ???? 1989. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661189900104.

Luick:2008:WLD

[LH08] John L. Luick and Jon B. Hinwood. Water levels in a dualbasin harbour in response to infragravity and edge waves. Progress in Oceanography, 77(4):367–375, June 2008. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661108000694.

Liu:2019:TDC

[LHC^{+}19] James T. Liu, Bangqin Huang, Yi Chang, Xiaoqin Du, Xin Liu, Rick J. Yang, Ray T. Hsu, Saulwood Lin, Jia-Jang Hung, Jay Lee, Chih-Chieh Su, and Yuan-Pin Chang. Threedimensional coupling between size-fractionated chlorophylla, POC and physical processes in the Taiwan Strait in summer. Progress in Oceanography, 176(??):??, September 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661118301599.

Laurel:2021:RWE

[LHC^{+}21] Benjamin J. Laurel, Mary E. Hunsicker, Lorenzo Ciannelli, Thomas P. Hurst, Janet Duffy-Anderson, Robert O’Malley, and Michael Behrenfeld. Regional warming exacerbates match/ mismatch vulnerability for cod larvae in Alaska. Progress in Oceanography, 193(??):??, April–May 2021. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661121000434.

Le:2013:TLT

[LHE^{+}13] Chengfeng Le, Chuanmin Hu, David English, Jennifer Cannizzaro, Zhiqiang Chen, Lian Feng, Richard Boler, and Charles Kovach. Towards a long-term chlorophyll-a data record in a turbid estuary using MODIS observations. Progress in Oceanography, 109(??):90–103, February 2013. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661112001176.

Lohse:1998:ROM

[LHEB98] L. Lohse, W. Helder, E. H. G. Epping, and W. Balzer. Recycling of organic matter along a shelf-slope transect across the N.W. European Continental Margin (Goban Spur). Progress in Oceanography, 42(1–4):77–110, December 1998. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661198000299.

Litzow:2016:DRS

[LHF^{+}16] Michael A. Litzow, Alistair J. Hobday, Stewart D. Frusher, Peter Dann, and Geoffrey N. Tuck. Detecting regime shifts in marine systems with limited biological data: an example from southeast Australia. Progress in Oceanography, 141(??):96–108, February 2016. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661115002487.

Lipiatou:1999:F

[LHM^{+}99] E. Lipiatou, S. Heussner, R. Mosetti, J. Tintoré, and A. Tselepides. Foreword. Progress in Oceanography, 44(1–3):3–4, August 1999. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S007966119900018X.

Leising:2005:BBM

[LHP^{+}05] Andrew W. Leising, Rita Horner, James J. Pierson, James Postel, and Claudia Halsband-Lenk. The balance between microzooplankton grazing and phytoplankton growth in a highly productive estuarine fjord. Progress in Oceanography, 67(3–4):366–383, November 2005. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661105001485.

Litzow:2020:EEC

[LHW^{+}20] Michael A. Litzow, Mary E. Hunsicker, Eric J. Ward, Sean C. Anderson, Jin Gao, Stephani G. Zador, Sonia Batten, Sherri C. Dressel, Janet Duffy-Anderson, Emily Fergusson, Russell R. Hopcroft, Benjamin J. Laurel, and Robert O’Malley. Evaluating ecosystem change as Gulf of Alaska temperature exceeds the limits of preindustrial variability.

Progress in Oceanography, 186(??):??, July 2020. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661120301324.

LaCasce:2010:LMA

[LI10] J. H. LaCasce and P. E. Isachsen. The linear models of the ACC. Progress in Oceanography, 84(3–4):139–157, March 2010. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S007966110900189X.

Li:2014:NAF

[Li14] Qiang Li. Numerical assessment of factors affecting nonlinear internal waves in the South China Sea. Progress in Oceanography, 121(??):24–43, February 2014. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661113000311.

Lie:1986:SHF

[Lie86] Heung-Jae Lie. Summertime hydrographic features in the southeastern Hwanghae. Progress in Oceanography, 17(3– 4):229–242, ???? 1986. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0079661186900467.

Lie:1988:NIC

[Lie88] Heung-Jae Lie. Near-inertial current oscillations off the mideast coast of Korea. Progress in Oceanography, 21(3–4):241– 253, ???? 1988. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661188900055.

Link:2012:DUE

[LIH^{+}12] J. S. Link, T. F. Ihde, C. J. Harvey, S. K. Gaichas, J. C. Field, J. K. T. Brodziak, H. M. Townsend, and R. M. Peterman. Dealing with uncertainty in ecosystem models: the paradox of use for living marine resource management. Progress in Oceanography, 102(??):102–114, September 2012. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661112000250.

Landergren:1965:STE

[LJ65] Sture Landergren and Oiva Joensuu. Studies on trace element distribution in a sediment core from the Pacific Ocean. Progress in Oceanography, 3(??):179–189, ???? 1965. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0079661165900170.

Lacharite:2016:DOP

[LJM^{+}16] Myriam Lacharité, Lis Lindal Jørgensen, Anna Metaxas, Vidar S. Lien, and Hein Rune Skjoldal. Delimiting oceanographic provinces to determine drivers of mesoscale patterns in benthic megafauna: a case study in the Barents Sea. Progress in Oceanography, 146(??):187–198, August 2016. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661116300015.

Lavaniegos:2002:PRN

[LJPGC02] B. E. Lavaniegos, L. C. Jiménez-Pérez, and G. Gaxiola-Castro. Plankton response to El Niño 1997–1998 and La Niña 1999 in the southern region of the California Current. Progress in Oceanography, 54(1–4):33–58, July 2002. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661102000423.

Landry:2013:ISM

[LK13] Michael R. Landry and Michio J. Kishi. Introduction: Sensitivities of marine food webs and biogeochemical cycles to enhanced ocean stratification. Progress in Oceanography, 119(??):1–3, December 2013. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661113001869.

LeHenaff:2014:CAE

[LKDL14] Matthieu Le Hénaff, Vassiliki H. Kourafalou, Renaud Dussurget, and Rick Lumpkin. Cyclonic activity in the eastern Gulf of Mexico: Characterization from along-track altimetry and in situ drifter trajectories. Progress in Oceanography, 120(??):120–138, January 2014. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661113001626.

LeProvost:1997:EMB

[LL97] C. Le Provost and Florent Lyard. Energetics of the M_{2} barotropic ocean tides: an estimate of bottom friction dissipation from a hydrodynamic model. Progress in Oceanography, 40(??):37–52, ???? 1997. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661197000220.

Liu:2021:EIF

[LL21] Xiao Liu and Naomi M. Levine. Ecosystem implications of finescale frontal disturbances in the oligotrophic ocean — an idealized modeling approach. Progress in Oceanography, 192(??): ??, March 2021. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661121000094.

Lloret-Lloret:2022:SPF

[LLAPG^{+}22] E. Lloret-Lloret, M. Albo-Puigserver, J. Giménez, J. Navarro, M. G. Pennino, J. Steenbeek, J. M. Bellido, and M. Coll. Small pelagic fish fitness relates to local environmental conditions and trophic variables. Progress in Oceanography, 202(??):??, March 2022. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661122000076.

Li:2022:ACP

[lLdZQ^{+}22] Cheng long Li, Wei dong Zhai, Di Qi, Lin lin Zhang, Li wen Zheng, and Yong Zhang. Anthropogenic CO_{2} in the Philippine Sea, northwestern Pacific: Distribution, storage, and decadal variations. Progress in Oceanography, 207(??):??, September– October 2022. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S007966112200132X.

Liu:2021:VPS

[LLGS21] Zijia Liu, Qian P. Li, Zaiming Ge, and Yiping Shuai. Variability of plankton size distribution and controlling factors across a coastal frontal zone. Progress in Oceanography, 197(??):??, September–October 2021. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661121001506.

Lin:2020:CMS

[LLH^{+}20] Hongyang Lin, Zhiyu Liu, Jianyu Hu, Dimitris Menemenlis, and Yongxiang Huang. Characterizing meso- to submesoscale features in the South China Sea. Progress in Oceanography, 188(??):??, October 2020. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661120301592.

Li:2021:OBE

[LLH^{+}21] Feili Li, M. Susan Lozier, N. Penny Holliday, William E. Johns, Isabela A. Le Bras, Ben I. Moat, Stuart A. Cunningham, and M. Femke de Jong. Observation-based estimates of heat and freshwater exchanges from the subtropical North Atlantic to the Arctic. Progress in Oceanography, 197(??):??, September– October 2021. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661121001257.

Lassalle:2011:LTL

[LLL^{+}11] G. Lassalle, J. Lobry, F. Le Loc’h, P. Bustamante, G. Certain, D. Delmas, C. Dupuy, C. Hily, C. Labry, O. Le Pape, E. Marquis, P. Petitgas, C. Pusineri, V. Ridoux, J. Spitz, and N. Niquil. Lower trophic levels and detrital biomass control the Bay of Biscay continental shelf food web: Implications for ecosystem management. Progress in Oceanography, 91(4):561–575, December 2011. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661111001133.

Lian:2024:TRB

[LLL^{+}24] Dongmei Lian, Xin Liu, Edward A. Laws, Tongtong Liu, Jingxiao Wang, Shaoling Shang, and Zhongping Lee. A “trapezoidal” relationship between solar radiation and chlorophyll concentrations at the center of the South Pacific Gyre. Progress in Oceanography, 225(??):??, July 2024. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661124000879.

Logerwell:2001:SEB

[LLS01] Elizabeth A. Logerwell, Bertha Lavaniegos, and Paul E. Smith. Spatially-explicit bioenergetics of Pacific sardine in the Southern California Bight: are mesoscale eddies areas of exceptional prerecruit production? Progress in Oceanography, 49(??):391– 406, ???? 2001. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661101000325.

Liu:2021:STV

[LLX^{+}21] Haoran Liu, Xin Liu, Wupeng Xiao, Edward A. Laws, and Bangqin Huang. Spatial and temporal variations of satellitederived phytoplankton size classes using a three-component model bridged with temperature in Marginal Seas of the Western Pacific Ocean. Progress in Oceanography, 191(??):??, February 2021. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S007966112100001X.

Llubes:1997:TRG

[LM97] M. Llubes and P. Mazzega. Testing recent global ocean tide models with loading gravimetric data. Progress in Oceanography, 40(??):369–383, ???? 1997. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661198000147.

Lefevre:2000:DCP

[LM00] N. Lefèvre and G. F. Moore. Distribution of the CO_{2} partial pressure along an Atlantic meridional transect. Progress in Oceanography, 45(3–4):401–413, April 2000. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661100000100.

Lehodey:2010:CIO

[LM10] Patrick Lehodey and Olivier Maury. CLimate Impacts on Oceanic TOp Predators (CLIOTOP): Introduction to the special issue of the CLIOTOP International Symposium, La Paz, Mexico, 3–7 December 2007. Progress in Oceanography, 86(1– 2):1–7, July 2010. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661110000832.

Litzow:2014:AEI

[LM14] Michael A. Litzow and Franz J. Mueter. Assessing the ecological importance of climate regime shifts: an approach from the North Pacific Ocean. Progress in Oceanography, 120(??):110–119, January 2014. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661113001638.

Lacharite:2018:EDE

[LM18] Myriam Lacharité and Anna Metaxas. Environmental drivers of epibenthic megafauna on a deep temperate continental shelf: a multiscale approach. Progress in Oceanography, 162(??):171– 186, March 2018. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661117302173.

Leu:2015:ASA

[LMA^{+}15] E. Leu, C. J. Mundy, P. Assmy, K. Campbell, T. M. Gabrielsen, M. Gosselin, T. Juul-Pedersen, and R. Gradinger. Arctic spring awakening — steering principles behind the phenology of vernal ice algal blooms. Progress in Oceanography, 139(??):151–170, December 2015. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661115001640.

Lin:2020:DVN

[LMC^{+}20] Pengfei Lin, Jinfeng Ma, Fei Chai, Peng Xiu, and Hailong Liu. Decadal variability of nutrients and biomass in the southern region of Kuroshio Extension. Progress in Oceanography, 188(??): ??, October 2020. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661120301804.

Li:2013:ZCS

[LMH^{+}13] Lingbo Li, Dave Mackas, Brian Hunt, Jake Schweigert, Evgeny Pakhomov, R. Ian Perry, Moira Galbraith, and Tony J. Pitcher. Zooplankton communities in the Strait of Georgia, British Columbia, track large-scale climate forcing over the Pacific Ocean. Progress in Oceanography, 115(??):90–102, August 2013. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661113000761.

LeJoncour:2023:WBE

[LML^{+}23] Anna Le Joncour, Maud Mouchet, Boris Leroy, Jennifer Doyle, Mathieu Lundy, Marianne Robert, and Dorothée Kopp. When bioregionalisation emphasises the presence of vulnerable marine ecosystem indicator species in Nephrops fishing grounds: Evidence in the Bay of Biscay/Irish Sea continuum. Progress in Oceanography, 214(??):??, June 2023. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661123000757.

Lu:2003:CSS

[LMM03] Beiwei Lu, David L. Mackas, and Douglas F. Moore. Crossshore separation of adult and juvenile euphausiids in a shelfbreak alongshore current. Progress in Oceanography, 57(3– 4):381–404, June 2003. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661103001071.

Lopez:2022:LEO

[LMP22] Bruno Díaz López, Séverine Methion, and Oriol Giralt Paradell. Living on the edge: Overlap between a marine predator’s habitat use and fisheries in the Northeast Atlantic waters (NW Spain). Progress in Oceanography, 175(??):115–123, ???? 2022. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661118302568.

Laguionie-Marchais:2016:IAS

[LMPB^{+}16] Claire Laguionie-Marchais, Gordon L. J. Paterson, Brian J. Bett, Kenneth L. Smith, and Henry A. Ruhl. Interannual species-level variations in an abyssal polychaete assemblage (Sta. M, NE Pacific, 4000 m). Progress in Oceanography, 140(??):43–53, January 2016. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661115002153.

Landry:1993:TDM

[LMS93] Michael R. Landry, Bruce C. Monger, and Karen E. Selph. Time-dependency of microzooplankton grazing and phytoplankton growth in the subarctic Pacific. Progress in Oceanography, 32(??):205–222, ???? 1993. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0079661193900145.

Lehodey:2010:BGO

[LMS10] Patrick Lehodey, Raghu Murtugudde, and Inna Senina. Bridging the gap from ocean models to population dynamics of large marine predators: a model of mid-trophic functional groups. Progress in Oceanography, 84(1–2):69–84, January 2010. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661109001414.

Lima:2019:CBM

[LMT^{+}19] Camila Rodrigues Lima, Carlos Rafael Borges Mendes, Virginia Maria Tavano, Amália Maria Sacilotto Detoni, and Eduardo Resende Secchi. Chemotaxonomy-based mapping of phytoplankton communities in the subtropical Southwestern Atlantic Ocean, with emphasis on the marine cyanobacterium Trichodesmium. Progress in Oceanography, 172(??):77–88, March 2019. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661118301770.

Lee:2012:PPA

[LMW^{+}12] Jae Hak Lee, Takeshi Matsuno, Mark Wimbush, Dongliang Yuan, and Kyung-Il Chang. The PAMS (Pacific–Asian Marginal Seas) study. Progress in Oceanography, 105(??):1–3, October 2012. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661112000353.

Lilly:2013:RCF

[LNB13] George R. Lilly, Odd Nakken, and John Brattey. A review of the contributions of fisheries and climate variability to contrasting dynamics in two arcto-boreal Atlantic cod (Gadus morhua) stocks: Persistent high productivity in the Barents Sea and collapse on the Newfoundland and Labrador Shelf. Progress in Oceanography, 114(??):106–125, July 2013. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661113000505.

Luther:1985:MSC

[LO85] Mark E. Luther and James J. O’Brien. A model of the seasonal circulation in the Arabian Sea forced by observed winds. Progress in Oceanography, 14(??):353–385, ???? 1985. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0079661185900175.

Lavaniegos:2007:CLT

[LO07] Bertha E. Lavaniegos and Mark D. Ohman. Coherence of longterm variations of zooplankton in two sectors of the California Current System. Progress in Oceanography, 75(1):42–69, October 2007. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661107001334.

Lilly:2021:ESD

[LO21] Laura E. Lilly and Mark D. Ohman. Euphausiid spatial displacements and habitat shifts in the southern California Current System in response to El Niño variability. Progress in Oceanography, 193(??):??, April–May 2021. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661121000343.

Lezama-Ochoa:2010:IRB

[LOBG^{+}10] Ainhoa Lezama-Ochoa, Guillermo Boyra, Nicolas Goñi, Haritz Arrizabalaga, and Arnaud Bertrand. Investigating relationships between albacore tuna (Thunnus alalunga) CPUE and prey distribution in the Bay of Biscay. Progress in Oceanography, 86(1–2):105–114, July 2010. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661110000406.

Lozier:1995:CNA

[LOC95] M. Susan Lozier, W. Brechner Owens, and Ruth G. Curry. The climatology of the North Atlantic. Progress in Oceanography, 36(1):1–44, ???? 1995. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0079661195000135.

Landry:2009:LSP

[LOG^{+}09] Michael R. Landry, Mark D. Ohman, Ralf Goericke, Michael R. Stukel, and Kate Tsyrklevich. Lagrangian studies of phytoplankton growth and grazing relationships in a coastal upwelling ecosystem off Southern California. Progress in Oceanography, 83(1–4):208–216, December 2009. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661109000846.

Longhurst:1985:SEP

[Lon85] Alan R. Longhurst. The structure and evolution of plankton communities. Progress in Oceanography, 15(1):1–35, ???? 1985. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0079661185900369.

Longhurst:1995:SCP

[Lon95] Alan Longhurst. Seasonal cycles of pelagic production and consumption. Progress in Oceanography, 36(2):77–167, ???? 1995. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0079661195000151.

Lin:2022:TLF

[LOO22] Yuchun Lin, Lie-Yauw Oey, and Alejandro Orfila. Two ‘faces’ of ENSO-induced surface waves during the tropical cyclone season. Progress in Oceanography, 175(??):40–54, ???? 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661118301605.

Lee:1987:C

[LP87] Thomas N. Lee and Leonard J. Pietrafesa. Circulation. Progress in Oceanography, 19(3–4):267–312, ???? 1987. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0079661187900115.

Landry:1992:PFW

[LPA92] Michael R. Landry, Willis K. Peterson, and Christine C. Andrews. Particulate flux in the water column overlying Santa Monica Basin. Progress in Oceanography, 30(1–4):167–195, ???? 1992. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/007966119290012O.

Levy:2011:MAR

[LPA^{+}11] I. Levy, P. P. Povinec, M. Aoyama, K. Hirose, J. A. Sanchez-Cabeza, J.-F. Comanducci, J. Gastaud, M. Eriksson, Y. Hamajima, C. S. Kim, K. Komura, I. Osvath, P. Roos, and S. A. Yim. Marine anthropogenic radiotracers in the Southern Hemisphere: New sampling and analytical strategies. Progress in Oceanography, 89(1–4):120–133, April 2011. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661110001643.

Lopez-Parages:2020:NPR

[LPARF^{+}20] Jorge López-Parages, Pierre-Amaël Auger, Belén Rodríguez-Fonseca, Noel Keenlyside, Carlo Gaetan, Angelo Rubino, Maeregu Woldeyes Arisido, and Timothée Brochier. El Niño as a predictor of round sardinella distribution along the northwest African coast. Progress in Oceanography, 186(??):??, July 2020. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S007966112030080X.

Linders:2017:NOW

[LPBM17] Johanna Linders, Robert. S. Pickart, Göran Björk, and G. W. K. Moore. On the nature and origin of water masses in herald canyon, Chukchi Sea: Synoptic surveys in summer 2004, 2008, and 2009. Progress in Oceanography, 159(??):99–114, December 2017. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661117302069.

Llapapasca:2018:MPH

[LPF^{+}18] Miguel A. Llapapasca, Aldo S. Pacheco, Paul Fiedler, Elisa Goya, Jesús Ledesma, Cecilia Peña, and Luis Vásquez. Modeling the potential habitats of dusky, commons and bottlenose dolphins in the Humboldt Current System off Peru: the influence of non-El Niño vs. El Niño 1997– 98 conditions and potential prey availability. Progress in Oceanography, 168(??):169–181, November 2018. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661117301611.

Lin:2020:CVM

[LPF^{+}20] Peigen Lin, Robert S. Pickart, David Fissel, Ed Ross, Jeremy Kasper, Frank Bahr, Daniel J. Torres, Jeff O’Brien, Keath Borg, Humfrey Melling, and Francis K. Wiese. Circulation in the vicinity of Mackenzie Canyon from a year-long mooring array. Progress in Oceanography, 187(??):??, August– September 2020. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S007966112030135X.

Lin:2021:NWF

[LPF^{+}21] Peigen Lin, Robert S. Pickart, David B. Fissel, Keath Borg, Humfrey Melling, and Francis K. Wiese. On the nature of wind-forced upwelling and downwelling in Mackenzie Canyon, Beaufort Sea. Progress in Oceanography, 198(??):??, November 2021. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661121001580.

Liu:2023:CDR

[LPF23] Guangpeng Liu, Brian Powell, and Tobias Friedrich. Climate downscaling for regional models with a neural network: a Hawaiian example. Progress in Oceanography, 215(??):??, July 2023. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661123000903.

Leising:2005:CGDb

[LPHL^{+}05a] Andrew W. Leising, James J. Pierson, Claudia Halsband-Lenk, Rita Horner, and James Postel. Copepod grazing during spring blooms: Can Pseudocalanus newmani induce trophic cascades? Progress in Oceanography, 67(3–4):406–421, November 2005. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661105001503.

Leising:2005:CGDa

[LPHL^{+}05b] Andrew W. Leising, James J. Pierson, Claudia Halsband-Lenk, Rita Horner, and James Postel. Copepod grazing during spring blooms: Does Calanus pacificus avoid harmful diatoms? Progress in Oceanography, 67(3–4):384–405, November 2005. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661105001497.

Li:2019:CCS

[LPS^{+}19] Min Li, Robert S. Pickart, Michael A. Spall, Thomas J. Weingartner, Peigen Lin, G. W. K. Moore, and Yiquan Qi. Circulation of the Chukchi Sea shelfbreak and slope from moored timeseries. Progress in Oceanography, 172(??):14–33, March 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661118301757.

Lin:2023:FCN

[LPW^{+}23] Peigen Lin, Robert S. Pickart, Thomas J. Weingartner, Harper L. Simmons, Motoyo Itoh, and Takashi Kikuchi. Formation and circulation of newly ventilated winter water in the western Beaufort Sea. Progress in Oceanography, 216(??):??, August 2023. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661123001118.

Levipan:2007:TSP

[LQU07] H. A. Levipan, R. A. Quiñones, and H. Urrutia. A time series of prokaryote secondary production in the oxygen minimum zone of the Humboldt Current System, off central Chile. Progress in Oceanography, 75(3):531–549, November 2007. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S007966110700167X.

Lafuente:2007:GCP

[LR07] Jesús García Lafuente and Javier Ruiz. The Gulf of Cádiz pelagic ecosystem: a review. Progress in Oceanography, 74(2–3):228–251, August 2007. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661107000602.

Latasa:2023:DPP

[LRAE23] Mikel Latasa, Francisco Rodríguez, Susana Agustí, and Marta Estrada. Distribution patterns of phytoplankton groups along isoirradiance layers in oligotrophic tropical and subtropical oceans. Progress in Oceanography, 217(??):??, September– October 2023. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661123001416.

LoIacono:2018:PCW

[LRGV^{+}18] Claudio Lo Iacono, Katleen Robert, Rita Gonzalez-Villanueva, Andrea Gori, Josep-Maria Gili, and Covadonga Orejas. Predicting cold-water coral distribution in the Cap de Creus Canyon (NW Mediterranean): Implications for marine conservation planning. Progress in Oceanography, 169(??):169–180, December 2018. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661117302860.

Lovvorn:2015:LBF

[LRJ^{+}15] James R. Lovvorn, Aariel R. Rocha, Stephen C. Jewett, Douglas Dasher, Steffen Oppel, and Abby N. Powell. Limits to benthic feeding by eiders in a vital Arctic migration corridor due to localized prey and changing sea ice. Progress in Oceanography, 136(??):162–174, August 2015. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661115001111.

Lascaratos:1999:RCD

[LRNK99] Alex Lascaratos, Wolfgang Roether, Kostas Nittis, and Birgit Klein. Recent changes in deep water formation and spreading in the eastern Mediterranean Sea: a review. Progress in Oceanography, 44(1–3):5–36, August 1999. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661199000191.

Lilly:2003:OLS

[LRS^{+}03] Jonathan M. Lilly, Peter B. Rhines, Friedrich Schott, Kara Lavender, John Lazier, Uwe Send, and Eric D’Asaro. Observations of the Labrador Sea eddy field. Progress in Oceanography, 59(1):75–176, October 2003. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661103001460.

Legendre:2015:MCP

[LRW^{+}15] Louis Legendre, Richard B. Rivkin, Markus G. Weinbauer, Lionel Guidi, and Julia Uitz. The microbial carbon pump concept: Potential biogeochemical significance in the globally changing ocean. Progress in Oceanography, 134(??):432– 450, May 2015. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661115000105.

Luyten:1985:UEG

[LS85] James Luyten and Henry Stommel. Upstream effects of the Gulf Stream on the structure of the mid-ocean thermocline. Progress in Oceanography, 14(??):387–399, ???? 1985. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0079661185900187.

Loeb:2012:PDS

[LS12] V. J. Loeb and J. A. Santora. Population dynamics of Salpa thompsoni near the Antarctic Peninsula: Growth rates and interannual variations in reproductive activity (1993–2009). Progress in Oceanography, 96(1):93–107, April 2012. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661111001194.

Loeb:2013:PCA

[LS13] Valerie J. Loeb and Jarrod A. Santora. Pteropods and climate off the Antarctic Peninsula. Progress in Oceanography, 116(??):31–48, September 2013. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661113000815.

Loeb:2015:CVS

[LS15] Valerie J. Loeb and Jarrod A. Santora. Climate variability and spatiotemporal dynamics of five Southern Ocean krill species. Progress in Oceanography, 134(??):93–122, May 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S007966111500004X.

Lankhorst:2020:UNA

[LS20] Matthias Lankhorst and Uwe Send. Uncertainty of North Atlantic Current observations from altimetry, floats, moorings, and XBT. Progress in Oceanography, 187(??):??, August– September 2020. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661120301415.

Lienart:2017:DPO

[LSB^{+}17] Camilla Liénart, Nicolas Savoye, Yann Bozec, Elsa Breton, Pascal Conan, Valérie David, Eric Feunteun, Karine Grangeré, Philippe Kerhervé, Benoît Lebreton, Sébastien Lefebvre, Stéphane L’Helguen, Laure Mousseau, Patrick Raimbault, Pierre Richard, Pascal Riera, Pierre-Guy Sauriau, Gauthier Schaal, Fabien Aubert, Sébastien Aubin, Sabrina Bichon, Christophe Boinet, Line Bourasseau, Martine Bréret, Jocelyne Caparros, Thierry Cariou, Karine Charlier, Pascal Claquin, Vincent Cornille, Anne-Marie Corre, Laurence Costes, Olivier Crispi, Muriel Crouvoisier, Marie Czamanski, Yolanda Del Amo, Hervé Derriennic, François Dindinaud, Maïa Durozier, Vincent Hanquiez, Antoine Nowaczyk, Jérémy Devesa, Sophie Ferreira, Michel Fornier, Fabrice Garcia, Nicole Garcia, Sandrine Geslin, Emilie Grossteffan, Aurore Gueux, Julien Guillaudeau, Gaël Guillou, Orianne Joly, Nicolas Lachaussée, Michel Lafont, Jézabel Lamoureux, Eric Lecuyer, Jean-Paul Lehodey, David Lemeille, Cédric Leroux, Eric Macé, Eric Maria, Philippe Pineau, Franck Petit, Mireille Pujo-Pay, Peggy Rimelin-Maury, and Emmanuelle Sultan. Dynamics of particulate organic matter composition in coastal systems: a spatio-temporal study at multisystems scale. Progress in Oceanography, 156(??):221–239, August 2017. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661116301914.

Llewellyn:2015:UCM

[LSD^{+}15] Carole A. Llewellyn, Ulf Sommer, Chris L. Dupont, Andrew E. Allen, and Mark R. Viant. Using community metabolomics as a new approach to discriminate marine microbial particulate organic matter in the western English Channel. Progress in Oceanography, 137(??):421–433, September 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661115000890.

Lienart:2018:DPO

[LSD^{+}18] Camilla Liénart, Nicolas Savoye, Valérie David, Pierre Ramond, Paco Rodriguez Tress, Vincent Hanquiez, Vincent Marieu, Fabien Aubert, Sébastien Aubin, Sabrina Bichon, Christophe Boinet, Line Bourasseau, Yann Bozec, Martine Bréret, Elsa Breton, Jocelyne Caparros, Thierry Cariou, Pascal Claquin, Pascal Conan, Anne-Marie Corre, Laurence Costes, Muriel Crouvoisier, Yolanda Del Amo, Hervé Derriennic, François Dindinaud, Robert Duran, Maïa Durozier, Jérémy Devesa, Sophie Ferreira, Eric Feunteun, Nicole Garcia, Sandrine Geslin, Emilie Grossteffan, Aurore Gueux, Julien Guillaudeau, Gaël Guillou, Orianne Jolly, Nicolas Lachaussée, Michel Lafont, Véronique Lagadec, Jézabel Lamoureux, Béatrice Lauga, Benoît Lebreton, Eric Lecuyer, Jean-Paul Lehodey, Cédric Leroux, Stéphane L’Helguen, Eric Macé, Eric Maria, Laure Mousseau, Antoine Nowaczyk, Philippe Pineau, Franck Petit, Mireille Pujo-Pay, Patrick Raimbault, Peggy Rimmelin-Maury, Vanessa Rouaud, Pierre-Guy Sauriau, Emmanuelle Sultan, and Nicolas Susperregui. Dynamics of particulate organic matter composition in coastal systems: Forcing of spatio-temporal variability at multi-systems scale. Progress in Oceanography, 162(??):271–289, March 2018. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661117302100.

Labrousse:2017:USI

[LSF^{+}17] Sara Labrousse, Jean-Baptiste Sallée, Alexander D. Fraser, Robert A. Massom, Phillip Reid, Michael Sumner, Christophe Guinet, Robert Harcourt, Clive McMahon, Frédéric Bailleul, Mark A. Hindell, and Jean-Benoit Charrassin. Under the sea ice: Exploring the relationship between sea ice and the foraging behaviour of southern elephant seals in East Antarctica. Progress in Oceanography, 156(??):17–40, August 2017. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661116300441.

Leu:2011:CCS

[LSH^{+}11] E. Leu, J. E. Søreide, D. O. Hessen, S. Falk-Petersen, and J. Berge. Consequences of changing sea-ice cover for primary and secondary producers in the European Arctic shelf seas: Timing, quantity, and quality. Progress in Oceanography, 90(1–4):18–32, July 2011. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S007966111100019X.

Luo:2022:GEB

[LSH^{+}22] Jessica Y. Luo, Charles A. Stock, Natasha Henschke, John P. Dunne, and Todd D. O’Brien. Global ecological and biogeochemical impacts of pelagic tunicates. Progress in Oceanography, 205(??):??, July 2022. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661122000830.

Loutrage:2023:NDD

[LSIB23] Liz Loutrage, Jérôme Spitz, Samuel P. Iglésias, and Anik Brind’Amour. The nocturnal distribution of deep-pelagic fish on the continental slope of the Bay of Biscay. Progress in Oceanography, 216(??):??, August 2023. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0079661123001131.

Licandro:2012:LTV

[LSIC12] P. Licandro, S. Souissi, F. Ibanez, and C. Carré. Longterm variability and environmental preferences of calycophoran siphonophores in the Bay of Villefranche (north-western Mediterranean). Progress in Oceanography, 97–100(??):152– 163, May 2012. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661111001224.

Lehodey:2008:SEP

[LSM08] Patrick Lehodey, Inna Senina, and Raghu Murtugudde. A spatial ecosystem and populations dynamics model (SEAPODYM)

— modeling of tuna and tuna-like populations. Progress in Oceanography, 78(4):304–318, September 2008. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S007966110800116X.

Latasa:2022:DPG

[LSM^{+}22] Mikel Latasa, Renate Scharek, Xosé Anxelu G. Morán, Andrés Gutiérrez-Rodríguez, Mikhail Emelianov, Jordi Salat, Montserrat Vidal, and Marta Estrada. Dynamics of phytoplankton groups in three contrasting situations of the open NW Mediterranean Sea revealed by pigment, microscopy, and flow cytometry analyses. Progress in Oceanography, 201(??):??, February 2022. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661121002202.

Limsakul:2001:TVL

[LSMG01] Atsamon Limsakul, Toshiro Saino, Takashi Midorikawa, and Joaquim I. Goes. Temporal variations in lower trophic level biological environments in the northwestern North Pacific subtropical gyre from 1950 to 1997. Progress in Oceanography, 49(??):129–149, ???? 2001. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661101000192.

Link:2009:CBT

[LSS^{+}09] Jason S. Link, William T. Stockhausen, Georg Skaret, William Overholtz, Bernard A. Megrey, Harald Gjøsæter, Sarah Gaichas, Are Dommasnes, Jannike Falk-Petersen, Joseph Kane, Franz J. Mueter, Kevin D. Friedland, and Jonathan A. Hare. A comparison of biological trends from four marine ecosystems: Synchronies, differences, and commonalities. Progress in Oceanography, 81(1–4):29–46, April 2009. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661109000214.

Lehodey:2010:PFP

[LSS^{+}10] P. Lehodey, I. Senina, J. Sibert, L. Bopp, B. Calmettes, J. Hampton, and R. Murtugudde. Preliminary forecasts of Pacific bigeye tuna population trends under the A2 IPCC scenario. Progress in Oceanography, 86(1–2):302–315, July 2010. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661110000558.

Lafon:2014:CAO

[LSV14] Alejandra Lafon, Nelson Silva, and Cristian A. Vargas. Contribution of allochthonous organic carbon across the Serrano River Basin and the adjacent fjord system in Southern Chilean Patagonia: Insights from the combined use of stable isotope and fatty acid biomarkers. Progress in Oceanography, 129(??):98–113, December 2014. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661114000305.

Luchin:2002:CBS

[LSW02] V. A. Luchin, I. P. Semiletov, and G. E. Weller. Changes in the Bering Sea region: atmosphere — ice — water system in the second half of the twentieth century. Progress in Oceanography, 55(1–2):23–44, October 2002. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S007966110200068X.

Liu:2021:IEI

[LSW^{+}21] Haijiao Liu, Jun Sun, Dongxiao Wang, Misun Yun, Dhiraj Dhondiram Narale, Guicheng Zhang, Xiaodong Zhang, and Satheeswaran Thangaraj. IOD-ENSO interaction with natural coccolithophore assemblages in the tropical eastern Indian Ocean. Progress in Oceanography, 193(??):??, April–May 2021. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661121000355.

Lin:2001:LTV

[LSXT01] Chuanlan Lin, Jilan Su, Bingrong Xu, and Qisheng Tang. Long-term variations of temperature and salinity of the Bohai Sea and their influence on its ecosystem. Progress

in Oceanography, 49(??):7–19, ????  2001.  CODEN

POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661101000131.

Langangen:2014:SVM

[LSY^{+}14] Øystein Langangen, Leif C. Stige, Natalia A. Yaragina, Geir Ottersen, Frode B. Vikebø, and Nils Chr. Stenseth. Spatial variations in mortality in pelagic early life stages of a marine fish (Gadus morhua). Progress in Oceanography, 127 (??):96–107, September 2014. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661114001116.

Lampadariou:2006:SVM

[LT06] N. Lampadariou and A. Tselepides. Spatial variability of meiofaunal communities at areas of contrasting depth and productivity in the Aegean Sea (NE Mediterranean). Progress in Oceanography, 69(1):19–36, April 2006. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661106000152.

Lacombe:1985:VBW

[LTG85] H. Lacombe, P. Tchernia, and L. Gamberoni. Variable bottom water in the Western Mediterranean Basin. Progress in Oceanography, 14(??):319–338, ???? 1985. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0079661185900151.

Lam:2015:MAC

[LTJ^{+}15] Phoebe J. Lam, Benjamin S. Twining, Catherine Jeandel, Alakendra Roychoudhury, Joseph A. Resing, Peter H. Santschi, and Robert F. Anderson. Methods for analyzing the concentration and speciation of major and trace elements in marine particles. Progress in Oceanography, 133(??):32–42, April 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661115000075.

Larsen:2013:BMS

[LTSG13] Morten Larsen, Bo Thamdrup, Tracy Shimmield, and Ronnie N. Glud. Benthic mineralization and solute exchange on a Celtic Sea sand-bank (Jones Bank). Progress in Oceanography, 117(??):64–75, October 2013. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661113000980.

Lukas:1986:TEU

[Luk86] Roger Lukas. The termination of the equatorial undercurrent in the eastern Pacific. Progress in Oceanography, 16(2):63– 90, ???? 1986. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661186900078.

Leon:2024:IHA

[LvBS^{+}24] Morgane Léon, Pieter van Beek, Virginie Sanial, Corentin Baudet, Matthew A. Charette, Marc Souhaut, Frédéric Vivier, Elodie Kestenare, Catherine Jeandel, and Hélène Planquette. Investigation of hydrothermal activity in the South West Indian ridge region using Ra isotopes and ^{227}Ac as tracers. Progress in Oceanography, 221(??):??, February 2024. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661123002343.

Labrousse:2015:WUS

[LVGH^{+}15] Sara Labrousse, Jade Vacquié-Garcia, Karine Heerah, Christophe Guinet, Jean-Baptiste Sallée, Matthieu Authier, Baptiste Picard, Fabien Roquet, Frédéric Bailleul, Mark Hindell, and Jean-Benoit Charrassin. Winter use of sea ice and ocean water mass habitat by southern elephant seals: the length and breadth of the mystery. Progress in Oceanography, 137 (part A)(??):52–68, September 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661115001287.

Loncaric:2007:SES

[LvIKB07] Neven Lončarić, Jolanda van Iperen, Dick Kroon, and Geert-Jan A. Brummer. Seasonal export and sediment preservation of diatomaceous, foraminiferal and organic matter mass fluxes in a trophic gradient across the SE Atlantic. Progress in Oceanography, 73(1):27–59, April 2007. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661106001443.

Leetmaa:1985:CNS

[LW85] Ants Leetmaa and Doug Wilson. Characteristics of near surface circulation patterns in the eastern equatorial Pacific. Progress in Oceanography, 14(??):339–352, ???? 1985. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0079661185900163.

Liu:2012:SVW

[LW12] Yonggang Liu and Robert H. Weisberg. Seasonal variability on the West Florida Shelf. Progress in Oceanography, 104(??):80– 98, October 2012. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S007966111200064X.

Liu:2013:TSD

[LW13] Dongyan Liu and Yueqi Wang. Trends of satellite derived chlorophyll-a (1997–2011) in the Bohai and Yellow Seas, China: Effects of bathymetry on seasonal and inter-annual patterns. Progress in Oceanography, 116(??):154–166, September 2013. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661113001092.

Loick-Wilde:2017:MBD

[LWBD^{+}17] Natalie Loick-Wilde, Deniz Bombar, Hai Nhu Doan, Lam Ngoc Nguyen, Anh Mai Nguyen-Thi, Maren Voss, and Joachim W. Dippner. Microplankton biomass and diversity in the Vietnamese upwelling area during SW monsoon under normal conditions and after an ENSO event. Progress in Oceanography, 153(??):1–15, April 2017. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661116300283.

Lorenzzetti:1987:CM

[LWL87] Joāo Lorenzzetti, John D. Wang, and Thomas N. Lee. Circulation modelling. Progress in Oceanography, 19(3–4):313– 327, ???? 1987. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661187900127.

Lundesgaard:2020:HEC

[LWT^{+}20] Øyvind Lundesgaard, Peter Winsor, Martin Truffer, Mark Merrifield, Brian Powell, Hank Statscewich, Emily Eidam, and Craig R. Smith. Hydrography and energetics of a cold subpolar fjord: Andvord Bay, western Antarctic Peninsula. Progress in Oceanography, 181(??):??, February 2020. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661119304045.

Lu:2007:MHC

[LWY07] Youyu Lu, Daniel G. Wright, and Igor Yashayaev. Modelling hydrographic changes in the Labrador Sea over the past five decades. Progress in Oceanography, 73(3–4):406– 426, May 2007. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661107000808.

Lim:2022:PCU

[LXC^{+}22] Dhongil Lim, Zhaokai Xu, Jinyong Choi, Soyoung Kim, Eunhee Kim, Sora Kang, and Hoisoo Jung. Paleoceanographic changes in the Ulleung Basin, East (Japan) Sea, during the last 20,000 years: Evidence from variations in element composition of core sediments. Progress in Oceanography, 88(1–4):101– 115, ???? 2022. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661110001680.

Lee:2012:MDV

[LYM12] Han Soo Lee, Takao Yamashita, and Toyoaki Mishima. Multidecadal variations of ENSO, the Pacific Decadal Oscillation and tropical cyclones in the western North Pacific. Progress in Oceanography, 105(??):67–80, October 2012. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661112000420.

Li:2022:SVB

[LYS^{+}22] Jianing Li, Qingxuan Yang, Hui Sun, Wei Zhao, and Jiwei Tian. Spatial variation of bottom mixed layer in the South China Sea and a potential mechanism. Progress in Oceanography, 206(??):??, August 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S007966112200115X.

Li:2016:EGO

[LYZ16] Gang Li, Wen Yan, and Lifeng Zhong. Element geochemistry of offshore sediments in the northwestern South China Sea and the dispersal of Pearl River sediments. Progress in Oceanography, 141(??):17–29, February 2016. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661115002402.

Liu:2005:FIN

[LZCZ05] Su Mei Liu, Jing Zhang, Hong Tao Chen, and Guo Sen Zhang. Factors influencing nutrient dynamics in the eutrophic Jiaozhou Bay, North China. Progress in Oceanography, 66(1): 66–85, July 2005. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S007966110500087X.

Li:2024:MUI

[LZF^{+}24] Longzhao Li, Feng Zhao, Sabine Filker, Weiyue Liu, Rongjie Zhao, Yuanyuan Wan, and Kuidong Xu. Microeukaryotes have unexpected importance in cold seep food webs through predation and parasitism. Progress in Oceanography, 222(??):??, March 2024. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661124000223.

Liu:2020:DCS

[LZG20] Zhiqiang Liu, Tingting Zu, and Jianping Gan. Dynamics of cross-shelf water exchanges off Pearl River Estuary in summer. Progress in Oceanography, 189(??):??, November–December 2020. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661120302020.

Li:2022:SSV

[LZL^{+}22] Juan Li, Xiaoming Zhai, Junliang Liu, Tong Yan, Yinghui He, Zhiwu Chen, and Shuqun Cai. Spatial and seasonal variations of near-inertial kinetic energy in the upper South China Sea: Role of synoptic atmospheric systems. Progress in Oceanography, 208(??):??, November 2022. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661122001586.

Morales:2012:PBM

[MA12] Carmen E. Morales and Valeria Anabalón. Phytoplankton biomass and microbial abundances during the spring upwelling season in the coastal area off Concepción, centralsouthern Chile: Variability around a time series station. Progress in Oceanography, 92–95(??):81–91, January 2012. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661111000589.

Maiorova:2020:BEE

[MA20] Anastassya S. Maiorova and Andrey V. Adrianov. Biodiversity of echiurans (Echiura) of the Kuril–Kamchatka Trench area. Progress in Oceanography, 180(??):??, January 2020. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661119303969.

Memery:2000:WMA

[MAAS^{+}00] L. Mémery, M. Arhan, X. A. Alvarez-Salgado, M.-J. Messias, H. Mercier, C. G. Castro, and A. F. Rios. The water masses along the western boundary of the south and equatorial Atlantic. Progress in Oceanography, 47(1):69–98, October 2000. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S007966110000032X.

Moore:2011:ROMb

[MAB^{+}11a] Andrew M. Moore, Hernan G. Arango, Gregoire Broquet, Chris Edwards, Milena Veneziani, Brian Powell, Dave Foley, James D. Doyle, Dan Costa, and Patrick Robinson. The Regional Ocean Modeling System (ROMS) 4-dimensional variational data assimilation systems: Part II — performance and application to the California Current System. Progress in Oceanography, 91(1):50–73, October 2011. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661111000504.

Moore:2011:ROMc

[MAB^{+}11b] Andrew M. Moore, Hernan G. Arango, Gregoire Broquet, Chris Edwards, Milena Veneziani, Brian Powell, Dave Foley, James D. Doyle, Dan Costa, and Patrick Robinson. The Regional Ocean Modeling System (ROMS) 4-dimensional variational data assimilation systems: Part III — observation impact and observation sensitivity in the California Current System. Progress in Oceanography, 91(1):74–94, October 2011. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661111000528.

Moore:2011:ROMa

[MAB^{+}11c] Andrew M. Moore, Hernan G. Arango, Gregoire Broquet, Brian S. Powell, Anthony T. Weaver, and Javier Zavala-Garay. The Regional Ocean Modeling System (ROMS) 4-dimensional variational data assimilation systems: Part I — system overview and formulation. Progress in Oceanography, 91(1):34– 49, October 2011. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661111000516.

Macdonald:1998:GOC

[Mac98] Alison M. Macdonald. The global ocean circulation: a hydrographic estimate and regional analysis. Progress in Oceanography, 41(3):281–382, June 1998. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661198000202.

Maeda:1988:MPF

[Mae88] Masachika Maeda. Microorganisms and protozoa as feed in mariculture. Progress in Oceanography, 21(2):201–206, ???? 1988. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0079661188900390.

Mir-Arguimbau:2022:SWC

[MAFS^{+}22] Joan Mir-Arguimbau, M. Mar Flexas, Jordi Salat, Paloma Martín, Marc Balcells, Nuria Raventós, and Ana Sabatés. Severe winter conditions improve recruitment success of blue whiting (Micromesistius poutassou), a temperate water fish species, in the NW Mediterranean Sea. Progress in Oceanography, 205(??):??, July 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661122000799.

Maud:2015:HDC

[MAH^{+}15] Jacqueline L. Maud, Angus Atkinson, Andrew G. Hirst, Penelope K. Lindeque, Claire E. Widdicombe, Rachel A. Harmer, Andrea J. McEvoy, and Denise G. Cummings. How does Calanus helgolandicus maintain its population in a variable environment? Analysis of a 25-year time series from the English Channel. Progress in Oceanography, 137(??):513–523, September 2015. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661115000956.

Mann:1969:SVO

[Man69] C. R. Mann. A summary of variability observed with monitoring sections off the east and west coasts of Canada. Progress in Oceanography, 5(??):17–30, ???? 1969. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/007966116990024X.

Mantua:2004:MDR

[Man04] Nathan Mantua. Methods for detecting regime shifts in large marine ecosystems: a review with approaches applied to North Pacific data. Progress in Oceanography, 60(2–4):165–182, February 2004. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661104000266.

Martin:2003:PPR

[Mar03] A. P. Martin. Phytoplankton patchiness: the role of lateral stirring and mixing. Progress in Oceanography, 57(2):125– 174, May 2003. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661103000855.

Marin:2020:NUD

[Mar20] Ivan Marin. Northern unicorns of the depths: Diversity of the genus Munidopsis Whiteaves, 1874 (Decapoda: Anomura: Munidopsidae) in the northwestern Pacific Ocean, with descriptions of three new species along the Russian coast. Progress in Oceanography, 183(??):??, April 2020. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S007966112030001X.

Maury:2010:OAS

[Mau10] Olivier Maury. An overview of APECOSM, a spatialized mass balanced “Apex Predators ECOSystem Model” to study physiologically structured tuna population dynamics in their ecosystem. Progress in Oceanography, 84(1–2):113–117, January 2010. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661109001463.

Maury:2017:CSR

[Mau17] Olivier Maury. Can schooling regulate marine populations and ecosystems? Progress in Oceanography, 156(??):91–103, August 2017. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661116301471.

McFarlane:2001:ROS

[MB01] Gordon A. McFarlane and Richard J. Beamish. The reoccurrence of sardines off British Columbia characterises the dynamic nature of regimes. Progress in Oceanography, 49 (??):151–165, ???? 2001. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661101000209.

Marra:2005:PPA

[MB05] John Marra and Richard T. Barber. Primary productivity in the Arabian Sea: a synthesis of JGOFS data. Progress in Oceanography, 65(2–4):159–175, May 2005. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S007966110500042X.

Middleton:2007:RSS

[MB07] John F. Middleton and John A. T. Bye. A review of the shelf-slope circulation along Australia’s southern shelves: Cape Leeuwin to Portland. Progress in Oceanography, 75(1):1–41, October 2007. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661107001322.

Malyutina:2020:MCI

[MB20] Marina V. Malyutina and Angelika Brandt. Munnopsidae (Crustacea, Isopoda, Asellota) from the Kuril–Kamchatka Trench with a regional and inter-ocean comparison of their biogeographic and richness patterns. Progress in Oceanography, 183(??):??, April 2020. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661120300276.

Mensch:1996:DTC

[MBB^{+}96] Manfred Mensch, Reinhold Bayer, John L. Bullister, Peter Schlosser, and Ray F. Weiss. The distribution of tritium and CFCs in the Weddell Sea during the mid-1980s. Progress in Oceanography, 38(4):377–388, IN5– IN9, 389–392, IN11–IN16, 393–415, ???? 1996. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661197000074.

Marty:1988:MOM

[MBCB88] D. Marty, J. C. Bertrand, P. Caumette, and A. Bianchi. Mineralization of organic material in coastal eutrophic ecosystems according to available terminal electron acceptors. Progress in Oceanography, 21(2):167–176, ???? 1988. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0079661188900353.

Mueter:2009:ERR

[MBD^{+}09] Franz J. Mueter, Cecilie Broms, Kenneth F. Drinkwater, Kevin D. Friedland, Jonathan A. Hare, George L. Hunt, Webjørn Melle, and Maureen Taylor. Ecosystem responses to recent oceanographic variability in high-latitude northern hemisphere ecosystems. Progress in Oceanography, 81(1–4):93– 110, April 2009. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S007966110900024X.

Many:2018:GSM

[MBdM^{+}18] Gaël Many, François Bourrin, Xavier Durrieu de Madron, Anouck Ody, David Doxaran, and Pierre Cauchy. Glider and satellite monitoring of the variability of the suspended particle distribution and size in the Rhône ROFI. Progress in Oceanography, 163(??):123–135, April 2018. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661117300502.

Merino:2014:EEL

[MBF^{+}14] Gorka Merino, Manuel Barange, Jose A. Fernandes, Christian Mullon, William Cheung, Verena Trenkel, and Vicky Lam. Estimating the economic loss of recent North Atlantic fisheries management. Progress in Oceanography, 129(??):314–323, December 2014. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661114000706.

McKinnell:2001:IBN

[MBH^{+}01] S. M. McKinnell, R. D. Brodeur, K. Hanawa, A. B. Hollowed, J. J. Polovina, and C.-I. Zhang. An introduction to the Beyond El Niño conference: climate variability and marine ecosystem impacts from the tropics to the Arctic. Progress in Oceanography, 49(??):1–6, ???? 2001. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S007966110100012X.

McQuaid:2023:BSB

[MBH^{+}23] Kirsty A. McQuaid, Amelia E. H. Bridges, Kerry L. Howell, Tiago B. R. Gandra, Vitor de Souza, Jock C. Currie, Oliver T. Hogg, Tabitha R. R. Pearman, James B. Bell, Lara J. Atkinson, Diane Baum, Jarbas Bonetti, Alvar Carranza, Omar Defeo, Thomas Furey, Maria A. Gasalla, Neil Golding, Shannon L. Hampton, Sebastián Horta, Daniel O. B. Jones, Amanda T. Lombard, Eleonora Manca, Yamandú Marin, Stephanie Martin, Pål Mortensen, Cecilia Passadore, Nils Piechaud, Kerry J.

Sink, and Andrew Yool. Broad-scale benthic habitat classification of the South Atlantic. Progress in Oceanography, 214(??): ??, June 2023. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661123000599.

McDonagh:2008:CIO

[MBKS08] Elaine L. McDonagh, Harry L. Bryden, Brian A. King, and Richard J. Sanders. The circulation of the Indian Ocean at 32^{o}S. Progress in Oceanography, 79(1):20–36, October 2008. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661108001316.

McIntyre:1965:CPP

[MBP65] Andrew McIntyre, Allan W. H. Bé, and Resa Preikstas. Coccoliths and the Pliocene–Pleistocene boundary. Progress in Oceanography, 4(??):3–8, IN1–IN6, 9–25, September 2, 1965. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0079661165900388.

Mackey:2011:ILN

[MBP^{+}11] Katherine R. M. Mackey, Laura Bristow, David R. Parks, Mark A. Altabet, Anton F. Post, and Adina Paytan. The influence of light on nitrogen cycling and the primary nitrite maximum in a seasonally stratified sea. Progress in Oceanography, 91(4):545–560, December 2011. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661111001121.

McCoy:2020:GOS

[MBS20] Daniel McCoy, Daniele Bianchi, and Andrew L. Stewart. Global observations of submesoscale coherent vortices in the ocean. Progress in Oceanography, 189(??):??, November– December 2020. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661120301890.

Mackas:2007:EZW

[MBT07] David L. Mackas, Sonia Batten, and Marc Trudel. Effects on zooplankton of a warmer ocean: Recent evidence from the Northeast Pacific. Progress in Oceanography, 75(2):223–252, October 2007. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661107001449.

Miller:1988:RLH

[MC88] Charles B. Miller and Martha J. Clemons. Revised life history analysis for large grazing copepods in the subarctic Pacific Ocean. Progress in Oceanography, 20(4):293–313, ???? 1988. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0079661188900444.

Meincke:2008:MT

[MC08] Jens Meincke and John Church. Matthias tomczak. Progress in Oceanography, 77(4):273–275, June 2008. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661108000621.

Messie:2015:SRP

[MC15] Monique Messié and Francisco P. Chavez. Seasonal regulation of primary production in eastern boundary upwelling systems. Progress in Oceanography, 134(??):1–18, May 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661114001773.

MacDonald:1990:FSD

[MCB^{+}90] Ian R. MacDonald, W. Russell Callender, Roger A. Burke, Susanne J. McDonald, and Robert S. Carney. Fine-scale distribution of methanotrophic mussels at a Louisiana cold seep. Progress in Oceanography, 24(??):15–24, ???? 1990. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/007966119090016U.

Miller:2010:CCU

[MCB^{+}10] Kathleen Miller, Anthony Charles, Manuel Barange, Keith Brander, Vincent F. Gallucci, Maria A. Gasalla, Ahmed Khan, Gordon Munro, Raghu Murtugudde, Rosemary E. Ommer, and R. Ian Perry. Climate change, uncertainty, and resilient fisheries: Institutional responses through integrative science. Progress in Oceanography, 87(1–4):338–346, October 2010. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661110001266.

McCartney:1992:RCD

[McC92] M. S. McCartney. Recirculating components to the deep boundary current of the northern North Atlantic. Progress in Oceanography, 29(4):283–383, ???? 1992. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/007966119290006L.

McDougall:1981:DDCa

[McD81a] Trevor J. McDougall. Double-diffusive convection with a non-linear equation of state: Part I. The accurate conservation of properties in a two-layer system. Progress in Oceanography, 10(2):71–89, ???? 1981. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/007966118190001X.

McDougall:1981:DDCb

[McD81b] Trevor J. McDougall. Double-diffusive convection with a nonlinear equation of state: Part II. Laboratory experiments and their interpretation. Progress in Oceanography, 10(2):91–121, ???? 1981. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661181900021.

McDougall:1988:NSP

[McD88] Trevor J. McDougall. Neutral-surface potential vorticity. Progress in Oceanography, 20(3):185–221, ???? 1988. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/007966118890002X.

Mazzocchi:2007:TVC

[MCD^{+}07] Maria Grazia Mazzocchi, Epaminondas D. Christou, Iole Di Capua, MaLuz Fernández de Puelles, Serena Fonda-Umani, Juan Carlos Molinero, Paul Nival, and Ioanna Siokou-Frangou. Temporal variability of Centropages typicus in the Mediterranean Sea over seasonal-to-decadal scales. Progress in Oceanography, 72(2–3):214–232, February 2007. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661107000092.

Mannocci:2014:PCS

[MCD^{+}14] Laura Mannocci, Maxime Catalogna, Ghislain Dorémus, Sophie Laran, Patrick Lehodey, Wendy Massart, Pascal Monestiez, Olivier Van Canneyt, Pierre Watremez, and Vincent Ridoux. Predicting cetacean and seabird habitats across a productivity gradient in the South Pacific gyre. Progress in Oceanography, 120(??):383–398, January 2014. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661113002243.

McGill:1964:DPO

[McG64] David A. McGill. The distribution of phosphorus and oxygen in the Atlantic Ocean, as observed during the I.G.Y., 1957–1958. Progress in Oceanography, 2(??):127, 129–211, ???? 1964. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661164900047.

Marinovic:2002:ENN

[MCG^{+}02] B. B. Marinovic, D. A. Croll, N. Gong, S. R. Benson, and F. P. Chavez. Effects of the 1997–1999 El Niño and La Niña events on zooplankton abundance and euphausiid community composition within the Monterey Bay coastal upwelling system. Progress in Oceanography, 54(1–4):265–277, July 2002. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661102000538.

Mitra:2014:BGB

[MCG^{+}14] Aditee Mitra, Claudia Castellani, Wendy C. Gentleman, Sigrún H. Jónasdóttir, Kevin J. Flynn, Antonio Bode, Claudia Halsband, Penelope Kuhn, Priscilla Licandro, Mette D. Agersted, Albert Calbet, Penelope K. Lindeque, Rolf Koppelmann, Eva F. Møller, Astthor Gislason, Torkel Gissel Nielsen, and Michael St. John. Bridging the gap between marine biogeochemical and fisheries sciences; configuring the zooplankton link. Progress in Oceanography, 129(??):176–199, December 2014. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661114000731.

Mercado:2007:SIA

[MCGR07] Jesús M. Mercado, Dolores Cortés, Alberto García, and Teodoro Ramírez. Seasonal and inter-annual changes in the planktonic communities of the northwest Alboran Sea (Mediterranean Sea). Progress in Oceanography, 74(2–3):273– 293, August 2007. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661107000845.

Munoz:2016:GCT

[MCGS^{+}16] Práxedes Muñoz, Lissette J. Cárdenas, Dieter Garbe-Schönberg, Javier Sellanes, Laurent Dezileau, Ives Melville, and Stephanie D. Mendes. Geochemical characterization of two distinctive systems with evidence of chemosynthetic activity, explored at the SE Pacific margin off Chile (46^{o}S and 33^{o}S). Progress in Oceanography, 148(??):26–43, November 2016. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661116301835.

Murphy:2012:DIM

[MCH^{+}12] E. J. Murphy, R. D. Cavanagh, E. E. Hofmann, S. L. Hill, A. J. Constable, D. P. Costa, M. H. Pinkerton, N. M. Johnston, P. N. Trathan, J. M. Klinck, D. A. Wolf-Gladrow, K. L. Daly, O. Maury, and S. C. Doney. Developing integrated models of Southern Ocean food webs: Including ecological complexity, accounting for uncertainty and the importance of scale. Progress in Oceanography, 102(??):74–92, September 2012. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661112000237.

McIlgorm:2010:EIC

[McI10] Alistair McIlgorm. Economic impacts of climate change on sustainable tuna and billfish management: Insights from the Western Pacific. Progress in Oceanography, 86(1–2):187–191, July 2010. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661110000583.

McKinnell:2004:EMS

[McK04] Skip McKinnell. Ecology of a marginal sea. Progress in Oceanography, 61(2–4):103–104, May 2004. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661104000588.

McKinnell:2008:FRS

[McK08] Skip McKinnell. Fraser River sockeye salmon productivity and climate: a re-analysis that avoids an undesirable property of Ricker’s curve. Progress in Oceanography, 77(2–3):146– 154, May 2008. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661108000517.

McKinna:2015:TDO

[McK15] Lachlan I. W. McKinna. Three decades of ocean-color remotesensing Trichodesmium spp. in the World’s oceans: a review. Progress in Oceanography, 131(??):177–199, February 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661114002237.

Mikaelyan:2017:PDW

[MCKS17] Alexander S. Mikaelyan, Valeriy K. Chasovnikov, Arseny A. Kubryakov, and Sergey V. Stanichny. Phenology and drivers of the winter–spring phytoplankton bloom in the open Black Sea: the application of Sverdrup’s hypothesis and its refinements. Progress in Oceanography, 151(??):163–176, February 2017. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661116300490.

Mathis:2015:OAR

[MCL^{+}15] J. T. Mathis, S. R. Cooley, N. Lucey, S. Colt, J. Ekstrom, T. Hurst, C. Hauri, W. Evans, J. N. Cross, and R. A. Feely. Ocean acidification risk assessment for Alaska’s fishery sector. Progress in Oceanography, 136(??):71–91, August 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661114001141.

Mori:2017:BRM

[MCMT^{+}17] Mao Mori, Stuart P. Corney, Jessica Melbourne-Thomas, Andreas Klocker, Michael Sumner, and Andrew Constable. A biologically relevant method for considering patterns of oceanic retention in the Southern Ocean. Progress in Oceanography, 159(??):1–12, December 2017. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661116302646.

Mullin:2003:TSV

[MCT03] Michael M. Mullin, David M. Checkley, and Michael P. Thimgan. Temporal and spatial variation in the sizes of California Current macrozooplankton: analysis by optical plankton counter. Progress in Oceanography, 57(3–4):299–316, June 2003. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661103001034.

McCartney:2007:DCG

[MD07] Michael S. McCartney and Kathleen A. Donohue. A deep cyclonic gyre in the Australian–Antarctic Basin. Progress in Oceanography, 75(4):675–750, December 2007. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S007966110700122X.

Mitra:2010:DEE

[MD10] Aditee Mitra and Cabell Davis. Defining the “to” in end-to-end models. Progress in Oceanography, 84(1–2):39–42, January 2010. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661109001372.

Marshall:2019:LTT

[MDAW^{+}19] Kristin N. Marshall, Janet T. Duffy-Anderson, Eric J. Ward, Sean C. Anderson, Mary E. Hunsicker, and Benjamin C. Williams. Long-term trends in ichthyoplankton assemblage structure, biodiversity, and synchrony in the Gulf of Alaska and their relationships to climate. Progress in Oceanography, 170(??):134–145, January 2019. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661117303348.

Mansui:2020:PML

[MDB^{+}20] J. Mansui, G. Darmon, T. Ballerini, O. van Canneyt, Y. Ourmieres, and C. Miaud. Predicting marine litter accumulation patterns in the Mediterranean basin: Spatiotemporal variability and comparison with empirical data. Progress in Oceanography, 182(??):??, March 2020. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661120300069.

Montero:2007:PCC

[MDC^{+}07] Paulina Montero, Giovanni Daneri, L. Antonio Cuevas, Humberto E. González, Bárbara Jacob, Lorena Lizárraga, and Eduardo Menschel. Productivity cycles in the coastal upwelling area off Concepción: the importance of diatoms and bacterioplankton in the organic carbon flux. Progress in Oceanography, 75(3):518–530, November 2007. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661107001668.

Mironov:2019:EKK

[MDG^{+}19] A. N. Mironov, A. B. Dilman, A. V. Gebruk, A. V. Kremenetskaia, K. V. Minin, and I. S. Smirnov. Echinoderms of the Kuril–Kamchatka Trench. Progress in Oceanography, 179 (??):??, December 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661119303970.

Mazzocchi:2012:SRC

[MDGC^{+}12] Maria Grazia Mazzocchi, Laurent Dubroca, Carmen García-Comas, Iole Di Capua, and Maurizio Ribera d’Alcalà. Stability and resilience in coastal copepod assemblages: the case of the Mediterranean long-term ecological research at Station MC (LTER-MC). Progress in Oceanography, 97–100(??):135– 151, May 2012. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661111001212.

Munoz:2012:EST

[MDL^{+}12] Praxedes Muñoz, Laurent Dezileau, Carina Lange, Lissette Cardenas, Javier Sellanes, Marco A. Salamanca, and Antonio Maldonado. Evaluation of sediment trace metal records as paleoproductivity and paleoxygenation proxies in the upwelling center off Concepción, Chile (36^{o}S). Progress in Oceanography, 92–95(??):66–80, January 2012. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661111000644.

Mitchell:2020:FNA

[MDR20] Emily G. Mitchell, Jennifer M. Durden, and Henry A. Ruhl. First network analysis of interspecific associations of abyssal benthic megafauna reveals potential vulnerability of abyssal hill community. Progress in Oceanography, 187(??):??, August– September 2020. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661120301403.

Myers:2022:SVW

[MDR22] Paul G. Myers, Chris Donnelly, and Mads H. Ribergaard. Structure and variability of the West Greenland Current in summer derived from 6 repeat standard sections. Progress in Oceanography, 80(1–2):93–112, ???? 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661108002103.

Medus:1987:WAH

[Med87] J. Medus. West African Holocene stratigraphy, transgressions and climatic changes. Progress in Oceanography, 18(1–4):167– 175, ???? 1987. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661187900322.

Mounier:2005:AMC

[MEMC05] Flore Mounier, Vincent Echevin, Laurent Mortier, and Michel Crepon. Analysis of the mesoscale circulation in the occidental Mediterranean Sea during winter 1999–2000 given by a regional circulation model. Progress in Oceanography, 66(2–4):251–269, August 2005. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S007966110500073X.

Mathis:2015:VPG

[MEMP15] M. Mathis, A. Elizalde, U. Mikolajewicz, and T. Pohlmann. Variability patterns of the general circulation and sea water temperature in the North Sea. Progress in Oceanography, 135(??):91–112, June 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661115000695.

Menezes:2021:APT

[Men21] Viviane V. Menezes. Advective pathways and transit times of the Red Sea Overflow Water in the Arabian Sea from Lagrangian simulations. Progress in Oceanography, 199(??):??, December 2021. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661121001804.

Merriman:1965:EFM

[Mer65] Daniel Merriman. Edward Forbes-Manxman. Progress in Oceanography, 3(??):191–206, ???? 1965. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0079661165900182.

Manriquez:2012:SSZ

[MERB12] Karen Manríquez, Ruben Escribano, and Ramiro Riquelme-Bugueño. Spatial structure of the zooplankton community in the coastal upwelling system off central-southern Chile in spring 2004 as assessed by automated image analysis. Progress in Oceanography, 92–95(??):121–133, January 2012. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661111000747.

Martinez:2013:FFJ

[MEST13] I. Martinez, J. R. Ellis, B. Scott, and A. Tidd. The fish and fisheries of Jones Bank and the wider Celtic Sea. Progress in Oceanography, 117(??):89–105, October 2013. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661113000293.

Miller:2015:BDL

[MFA^{+}15] Michael J. Miller, Eric Feunteun, Jun Aoyama, Shun Watanabe, Mari Kuroki, Raymonde Lecomte-Finiger, Yuki Minegishi, Tony Robinet, Elodie Réveillac, Pierre-Alexandre Gagnaire, Patrick Berrebi, Katsumi Tsukamoto, and Tsuguo Otake. Biodiversity and distribution of leptocephali west of the Mascarene Plateau in the southwestern Indian Ocean. Progress in Oceanography, 137 (part A)(??):84–102, September 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661115001317.

Miller:1984:LHL

[MFB^{+}84] Charles B. Miller, Bruce W. Frost, Harold P. Batchelder, Martha J. Clemons, and Richard E. Conway. Life histories of large, grazing copepods in a subarctic ocean gyre: Neocalanus plumchrus, Neocalanus cristatus, and Eucalanus bungii in the Northeast Pacific. Progress in Oceanography, 13(2):201– 243, ???? 1984. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661184900090.

McLaskey:2022:HRT

[MFDH22] Anna K. McLaskey, Ian Forster, Justin Del Bel Belluz, and Brian P. V. Hunt. A high-resolution time series of particulate matter fatty acids reveals temporal dynamics of the composition and quality available to zooplankton in a temperate coastal ocean. Progress in Oceanography, 206(??):??, August 2022. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661122001045.

Murty:1986:SSP

[MFH86] T. S. Murty, R. A. Flather, and R. F. Henry. The storm surge problem in the Bay of Bengal. Progress in Oceanography, 16(4):195–233, ???? 1986. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/007966118690039X.

Molinari:1985:ESS

[MFM85] Robert L. Molinari, John F. Festa, and Eric Marmolejo. Evolution of sea-surface temperature and surface meteorological fields in the tropical Atlantic Ocean during FGGE, 1979: I. Description of surface fields and computation of surface energy fluxes. Progress in Oceanography, 14(??):401–420, ???? 1985. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0079661185900199.

Munchow:2015:BIW

[MFM15] Andreas Münchow, Kelly K. Falkner, and Humfrey Melling. Baffin Island and West Greenland Current Systems in northern Baffin Bay. Progress in Oceanography, 132(??):305–317, March 2015. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661114000408.

Maury:2007:MEEa

[MFS^{+}07] Olivier Maury, Blaise Faugeras, Yunne-Jai Shin, Jean-Christophe Poggiale, Tamara Ben Ari, and Francis Marsac. Modeling environmental effects on the size-structured energy flow through marine ecosystems. Part 1: the model. Progress in Oceanography, 74(4):479–499, September 2007. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661107000985.

McCreary:2016:CDA

[MFS^{+}16a] Julian P. McCreary, Ryo Furue, Fabian Schloesser, Theodore W. Burkhardt, and Masami Nonaka. Corrigendum to “Dynamics of the Atlantic meridional overturning circulation and Southern Ocean in an ocean model of intermediate complexity” [Prog. Oceanogr. 143 (2016) 46–81]. Progress in Oceanography, 147(??):67, September 2016. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661116301483. See [MFS^{+}16b].

McCreary:2016:DAM

[MFS^{+}16b] Julian P. McCreary, Ryo Furue, Fabian Schloesser, Theodore W. Burkhardt, and Masami Nonaka. Dynamics of the Atlantic meridional overturning circulation and Southern Ocean in an ocean model of intermediate complexity. Progress in Oceanography, 143(??):46–81, April 2016. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/

pii/S0079661116000069. See corrigendum [MFS^{+}16a].

Milliman:1986:SMR

[MFY^{+}86] John D. Milliman, Li Fan, Zhao Yiyang, Zheng Tiemin, and Richard Limeburner. Suspended matter regime in the Yellow Sea. Progress in Oceanography, 17(3–4):215–227, ???? 1986. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0079661186900455.

Mackas:2002:ZCC

[MG02] D. L. Mackas and M. Galbraith. Zooplankton community composition along the inner portion of Line P during the 1997– 1998 El Niño event. Progress in Oceanography, 54(1–4):423– 437, July 2002. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661102000629.

Marquardt:2023:SDS

[MGA^{+}23] Miriam Marquardt, Lucie Goraguer, Philipp Assmy, Bodil A. Bluhm, Signe Aaboe, Emily Down, Evan Patrohay, Bente Edvardsen, Agnieszka Tatarek, Zofia Smoła, Jozef Wiktor, and Rolf Gradinger. Seasonal dynamics of sea-ice protist and meiofauna in the northwestern Barents Sea. Progress in Oceanography, 218(??):??, November 2023. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661123001714.

Meddeb:2018:PFW

[MGC^{+}18] Marouan Meddeb, Boutheïna Grami, Aurélie Chaalali, Matilda Haraldsson, Nathalie Niquil, Olivier Pringault, and Asma Sakka Hlaili. Plankton food-web functioning in anthropogenically impacted coastal waters (SW Mediterranean Sea): an ecological network analysis. Progress in Oceanography, 162(??):66–82, March 2018. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661117300782.

Mackas:2012:CZS

[MGE^{+}12] D. L. Mackas, W. Greve, M. Edwards, S. Chiba, K. Tadokoro, D. Eloire, M. G. Mazzocchi, S. Batten, A. J. Richardson, C. Johnson, E. Head, A. Conversi, and T. Peluso. Changing zooplankton seasonality in a changing ocean: Comparing time series of zooplankton phenology. Progress in Oceanography, 97–100(??):31–62, May 2012. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661111001236.

Mackas:2013:ZTS

[MGF^{+}13] David Mackas, Moira Galbraith, Deborah Faust, Diane Masson, Kelly Young, William Shaw, Stephen Romaine, Marc Trudel, John Dower, Rob Campbell, Akash Sastri, Elizabeth A. Bornhold Pechter, Evgeny Pakhomov, and Rana El-Sabaawi. Zooplankton time series from the Strait of Georgia: Results from year-round sampling at deep water locations, 1990–2010. Progress in Oceanography, 115(??):129–159, August 2013. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S007966111300061X.

Mahanty:2022:SSW

[MGG22] Madan Mohan Mahanty, Latha Ganeshan, and Raguraman Govindan. Soundscapes in shallow water of the eastern Arabian Sea. Progress in Oceanography, 165(??):158–167, ???? 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661117300964.

Morales:2007:DCD

[MGH^{+}07] Carmen E. Morales, Humberto E. González, Samuel E. Hormazabal, Gabriel Yuras, Jaime Letelier, and Leonardo R. Castro. The distribution of chlorophyll-a and dominant planktonic components in the coastal transition zone off Concepción, central Chile, during different oceanographic conditions. Progress in Oceanography, 75(3):452–469, November 2007. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661107001620.

Molinari:1986:SFG

[MGK^{+}86] R. L. Molinari, S. L. Garzoli, E. J. Katz, D. E. Harrison, P. L. Richardson, and G. Reverdin. A synthesis of the first GARP Globa Experiment (FGGE) in the equatorial Atlantic Ocean. Progress in Oceanography, 16(2):91–112, ???? 1986. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/007966118690008X.

Mao:2019:NSS

[MGKW19] Xinyan Mao, Xinyu Guo, Taishi Kubota, and Yucheng Wang. Numerical studies on snow crab (Chionoecetes opilio) larval survival and transport in the Sea of Japan. Progress in Oceanography, 179(??):??, December 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661119303842.

Mahaut:1990:TSC

[MGS90] Marie-Laure Mahaut, Patrick Geistdoerfer, and Myriam Sibuet. Trophic strategies in carnivorous fishes: their significance in energy transfer in the deep-sea benthic ecosystem (Meriadzek Terrace — Bay of Biscay). Progress in Oceanography, 24(??):223–237, ???? 1990. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/007966119090032W.

Manko:2020:FAV

[MGWZ20] Maciej K. Mańko, Marta Gluchowska, and Agata Weydmann-Zwolicka. Footprints of Atlantification in the vertical distribution and diversity of gelatinous zooplankton in the Fram Strait (Arctic Ocean). Progress in Oceanography, 189(??):??, November–December 2020. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661120301531.

McCave:2002:TWN

[MH02]  I. N. McCave and I. R. Hall.  Turbidity of waters

over the Northwest Iberian continental margin. Progress in Oceanography, 52(??):299–313, ???? 2002. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661102000125.

Moon:2014:SRS

[MH14] Jae-Hong Moon and Naoki Hirose. Seasonal response of the southern East China Sea shelf water to windmodulated throughflow in the Taiwan Strait. Progress in Oceanography, 121(??):74–82, February 2014. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661113002280.

Mauritzen:2011:CLA

[MHA^{+}11] C. Mauritzen, E. Hansen, M. Andersson, B. Berx, A. Beszczynska-Möller, I. Burud, K. H. Christensen, J. Debernard, L. de Steur, P. Dodd, S. Gerland, Ø. Godøy, B. Hansen, S. Hudson, F. Høydalsvik, R. Ingvaldsen, P. E. Isachsen, Y. Kasajima, I. Koszalka, K. M. Kovacs, M. Køltzow, J. LaCasce, C. M. Lee, T. Lavergne, C. Lydersen, M. Nicolaus, F. Nilsen, O. A. Nøst, K. A. Orvik, M. Reigstad, H. Schyberg, L. Seuthe, Ø. Skagseth, J. Skardhamar, R. Skogseth, A. Sperrevik, C. Svensen, H. Søiland, S. H. Teigen, V. Tverberg, and C. Wexels Riser. Closing the loop — approaches to monitoring the state of the Arctic Mediterranean during the International Polar Year 2007–2008. Progress in Oceanography, 90(1–4):62–89, July 2011. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661111000255.

Morales:2012:MVN

[MHCR^{+}12] Carmen E. Morales, Samuel Hormazabal, Marco Correa-Ramirez, Oscar Pizarro, Nelson Silva, Camila Fernandez, Valeria Anabalón, and M. Loreto Torreblanca. Mesoscale variability and nutrient–phytoplankton distributions off centralsouthern Chile during the upwelling season: the influence of mesoscale eddies. Progress in Oceanography, 104(??):17–29, October 2012. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661112000481.

Mohn:2023:TDS

[MHCS^{+}23] Christian Mohn, Jørgen L. S. Hansen, Marina Carreiro-Silva, Stuart A. Cunningham, Evert de Froe, Carlos Dominguez-Carrió, Stefan Gary, Ronnie N. Glud, Cordula Göke, Clare Johnson, Telmo Morato, Eva Friis Møller, Lorenzo Rovelli, Kirstin Schulz, Karline Soetaert, Anna van der Kaaden, and Dick van Oevelen. Tidal to decadal scale hydrodynamics at two contrasting cold-water coral sites in the Northeast Atlantic. Progress in Oceanography, 214(??):??, June 2023. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661123000745.

Macias:2019:NRD

[MHGGS19] Diego Macias, I. Emma Huertas, Elisa Garcia-Gorriz, and Adolf Stips. Non-redfieldian dynamics driven by phytoplankton phosphate frugality explain nutrient and chlorophyll patterns in model simulations for the Mediterranean Sea. Progress in Oceanography, 173(??):37–50, April 2019. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661118301678.

Machin:2006:MFC

[MHGP06] F. Machín, A. Hernández-Guerra, and J. L. Pelegrí. Mass fluxes in the Canary Basin. Progress in Oceanography, 70(2–4):416– 447, August 2006. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S007966110600053X.

Michel:2015:AOO

[MHH^{+}15] Christine Michel, Jim Hamilton, Edmond Hansen, David Barber, Marit Reigstad, John Iacozza, Lena Seuthe, and Andrea Niemi. Arctic Ocean outflow shelves in the changing Arctic: a review and perspectives. Progress in Oceanography, 139(??):66–88, December 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661115001871.

McCarthy:2010:ILF

[MHR^{+}10] Abigail L. McCarthy, Selina Heppell, Francois Royer, Carla Freitas, and Thomas Dellinger. Identification of likely foraging habitat of pelagic loggerhead sea turtles (Caretta caretta) in the North Atlantic through analysis of telemetry track sinuosity. Progress in Oceanography, 86(1–2):224–231, July 2010. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661110000431.

Megrey:2009:CEC

[MHS^{+}09] Bernard A. Megrey, Jonathan A. Hare, William T. Stockhausen, Are Dommasnes, Harald Gjøsæter, William Overholtz, Sarah Gaichas, Georg Skaret, Jannike Falk-Petersen, Jason S. Link, and Kevin D. Friedland. A cross-ecosystem comparison of spatial and temporal patterns of covariation in the recruitment of functionally analogous fish stocks. Progress in Oceanography, 81(1–4):63–92, April 2009. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661109000238.

Macovei:2020:CIP

[MHS^{+}20a] Vlad A. Macovei, Susan E. Hartman, Ute Schuster, Sinhué Torres-Valdés, C. Mark Moore, and Richard J. Sanders. Corrigendum to “Impact of physical and biological processes on temporal variations of the ocean carbon sink in the midlatitude North Atlantic (2002–2016)” [Progr. Oceanogr. 180 (2020) 102223]. Progress in Oceanography, 186(??):??, July 2020. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661120301294. See [MHS^{+}20b].

Macovei:2020:IPB

[MHS^{+}20b] Vlad A. Macovei, Susan E. Hartman, Ute Schuster, Sinhué Torres-Valdés, C. Mark Moore, and Richard J. Sanders. Impact of physical and biological processes on temporal variations of the ocean carbon sink in the mid-latitude North Atlantic (2002–2016). Progress in Oceanography, 180(??):??, January 2020. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661119304033. See corri-

gendum [MHS^{+}20a].

Margonski:2010:CIB

[MHTG10] Piotr Margonski, Sture Hansson, Maciej T. Tomczak, and Ryszard Grzebielec. Climate influence on Baltic cod, sprat, and herring stock–recruitment relationships. Progress in Oceanography, 87(1–4):277–288, October 2010. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661110001126.

Moore:2019:BWE

[MHVS19] Sue E. Moore, Tore Haug, Gísli A. Víkingsson, and Garry B. Stenson. Baleen whale ecology in Arctic and subarctic seas in an era of rapid habitat alteration. Progress in Oceanography, 176(??):??, September 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661118302775.

McKinnell:2021:PFR

[MI21] Skip McKinnell and James R. Irvine. Phenology and Fraser River sockeye salmon marine survival. Progress in Oceanography, 197(??):??, September–October 2021. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S007966112100118X.

Midttun:1969:VTS

[Mid69] Lars Midttun. Variability of temperature and salinity at some localities off the coast of Norway. Progress in Oceanography, 5(??):41–54, ???? 1969. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0079661169900263.

Michel:2006:VOE

[MIH06] C. Michel, R. G. Ingram, and L. R. Harris. Variability in oceanographic and ecological processes in the Canadian Arctic Archipelago. Progress in Oceanography, 71(2–4):379–401, October 2006. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S007966110600125X.

Miller:1988:NFN

[Mil88] Charles B. Miller. Neocalanus flemingeri, a new species of Calanidae (Copepoda: Calanoida) from the subarctic Pacific Ocean, with a comparative redescription of Neocalanus plumchrus (Marukawa) 1921. Progress in Oceanography, 20(4):223– 273, ???? 1988. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661188900420.

Miller:1993:DLC

[Mil93a] Charles B. Miller. Development of large copepods during spring in the Gulf of Alaska. Progress in Oceanography, 32(??):295– 317, ???? 1993. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661193900189.

Miller:1993:PPP

[Mil93b] Charles B. Miller. Pelagic production production in the subacrtic Pacific. Progress in Oceanography, 32(??):1–15, ???? 1993. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/007966119390007Z.

Millot:2009:ADM

[Mil09] Claude Millot. Another description of the Mediterranean Sea outflow. Progress in Oceanography, 82(2):101–124, August 2009. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661109000366.

Millot:2014:HFM

[Mil14] Claude Millot. Heterogeneities of in- and out-flows in the Mediterranean Sea. Progress in Oceanography, 120(??): 254–278, January 2014. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661113001857.

Minobe:2000:STS

[Min00] Shoshiro Minobe. Spatio-temporal structure of the pentadecadal variability over the North Pacific. Progress in Oceanography, 47(2–4):381–408, October 2000. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661100000422.

Minobe:2002:IIC

[Min02] S. Minobe. Interannual to interdecadal changes in the Bering Sea and concurrent 1998/99 changes over the North Pacific. Progress in Oceanography, 55(1–2):45–64, October 2002. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661102000691.

Menze:2020:PDA

[MIN^{+}20] Sebastian Menze, Randi B. Ingvaldsen, Anna Nikolopoulos, Tore Hattermann, Jon Albretsen, and Harald Gjøsæter. Productive detours — Atlantic water inflow and acoustic backscatter in the major troughs along the Svalbard shelf. Progress in Oceanography, 188(??):??, October 2020. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661120301853.

Mittelstaedt:1983:UAN

[Mit83] Ekkehard Mittelstaedt. The upwelling area off Northwest Africa — a description of phenomena related to coastal upwelling. Progress in Oceanography, 12(3):307–331, ???? 1983. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0079661183900125.

Mittelstaedt:1991:OBA

[Mit91] Ekkehard Mittelstaedt. The ocean boundary along the northwest African coast: Circulation and oceanographic properties at the sea surface. Progress in Oceanography, 26(4):307–355, ???? 1991. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/007966119190011A.

Mitchell-Innes:1991:STV

[MIW91] B. A. Mitchell-Innes and D. R. Walker. Short-term variability during an anchor station study in the southern Benguela upwelling system: Phytoplankton production and biomass in relation to specie changes. Progress in Oceanography, 28(1–2): 65–89, ???? 1991. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/007966119190021D.

McDougall:1988:HNN

[MJ88] Trevor J. McDougall and David R. Jackett. On the helical nature of neutral trajectories in the ocean. Progress in Oceanography, 20(3):153–183, ???? 1988. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0079661188900018.

Muraleedharan:2007:IBS

[MJA^{+}07] K. R. Muraleedharan, P. Jasmine, C. T. Achuthankutty, C. Revichandran, P. K. Dinesh Kumar, P. Anand, and G. Rejomon. Influence of basin-scale and mesoscale physical processes on biological productivity in the Bay of Bengal during the summer monsoon. Progress in Oceanography, 72(4):364– 383, March 2007. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661106001431.

Moore:2017:IOO

[MJC^{+}17] Andrew M. Moore, Michael G. Jacox, William J. Crawford, Bruce Laughlin, Christopher A. Edwards, and Jérôme Fiechter. The impact of the ocean observing system on estimates of the California Current circulation spanning three decades. Progress in Oceanography, 156(??):41–60, August 2017. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661116301653.

Moreau:2021:HDS

[MJD^{+}21] Camille Moreau, Quentin Jossart, Bruno Danis, Marc Eléaume, Henrik Christiansen, Charlène Guillaumot, Rachel Downey, and Thomas Saucède. The high diversity of Southern Ocean sea stars (Asteroidea) reveals original evolutionary pathways. Progress in Oceanography, 190(??):??, January 2021. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661120302081.

Myers:2007:IVL

[MJWK07] Paul G. Myers, Simon A. Josey, Brett Wheler, and Nilgun Kulan. Interdecadal variability in Labrador Sea precipitation minus evaporation and salinity. Progress in Oceanography, 73(3–4):341–357, May 2007. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661107000754.

Miita:1986:TDB

[MK86] Tsunehiro Miita and Kazuo Kawatate. Trajectories of drift bottles released in the Tsushima Strait. Progress in Oceanography, 17(3–4):255–263, ???? 1986. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0079661186900480.

Moon:2012:IUO

[MK12] Il-Ju Moon and Seok Jae Kwon. Impact of upper-ocean thermal structure on the intensity of Korean peninsular landfall typhoons. Progress in Oceanography, 105(??):61–66, October 2012. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661112000419.

McFarlane:2000:TBR

[MKB00] Gordon A. McFarlane, Jacquelynne R. King, and Richard J. Beamish. Have there been recent changes in climate? Ask the fish. Progress in Oceanography, 47(2–4):147–169, October 2000. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661100000343.

Manning:1990:BDS

[MKD90] Raymond B. Manning, Roy K. Kropp, and Jane Dominguez. Biogeography of deep-sea stomatopod Crustacea, family Bathysquillidae. Progress in Oceanography, 24(??):311–316, ???? 1990. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661190900409.

McCreary:1996:FCE

[MKHO96] Julian P. McCreary, Kevin E. Kohler, Raleigh R. Hood, and Donald B. Olson. A four-component ecosystem model of biological activity in the Arabian Sea. Progress in Oceanography, 37(3–4):193–240, ???? 1996. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661196000055.

Mizuno:1986:CTO

[MKM86] S. Mizuno, K. Kawatate, and T. Miita. Current and temperature observations in the east Tsushima channel and the Sea of Genkai. Progress in Oceanography, 17(3–4):277–295, ???? 1986. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661186900509.

McCreary:1993:NID

[MKM93] Julian P. McCreary, Pijush K. Kundu, and Robert L. Molinari. A numerical investigation of dynamics, thermodynamics and mixed-layer processes in the Indian Ocean. Progress in Oceanography, 31(3):181–244, ???? 1993. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/007966119390002U. See erratum [Ano94c].

Muller-Karger:1989:PDC

[MKMF^{+}89] F. E. Müller-Karger, C. R. McClain, T. R. Fisher, W. E. Esaias, and R. Varela. Pigment distribution in the Caribbean Sea: Observations from space. Progress in Oceanography, 23(1):23– 64, ???? 1989. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661189900244.

Monteiro:2020:TIS

[MKOLA20] Thiago Monteiro, Rodrigo Kerr, Iole B. M. Orselli, and Jannine M. Lencina-Avila. Towards an intensified summer CO_{2} sink behaviour in the Southern Ocean coastal regions. Progress in Oceanography, 183(??):??, April 2020. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661120300057.

Moon:2022:CLS

[MKS^{+}22] Jae-Hong Moon, Taekyun Kim, Young Baek Son, Ji-Seok Hong, Joon-Ho Lee, Pil-Hun Chang, and Soo-Kang Kim. Contribution of low-salinity water to sea surface warming of the East China Sea in the summer of 2016. Progress in Oceanography, 175(??):68–80, ???? 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661118301976.

Meyer-Kaiser:2022:LDR

[MKSvA^{+}22] Kirstin S. Meyer-Kaiser, Kharis R. Schrage, Wilken-Jon von Appen, Mario Hoppmann, Normen Lochthofen, Arild Sundfjord, and Thomas Soltwedel. Larval dispersal and recruitment of benthic invertebrates in the Arctic Ocean. Progress in Oceanography, 203(??):??, April–May 2022. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661122000374.

Muller-Karger:2015:NVS

[MKSW^{+}15] Frank E. Muller-Karger, Joseph P. Smith, Sandra Werner, Robert Chen, Mitchell Roffer, Yanyun Liu, Barbara Muhling, David Lindo-Atichati, John Lamkin, Sergio Cerdeira-Estrada, and David B. Enfield. Natural variability of surface oceanographic conditions in the offshore Gulf of Mexico. Progress in Oceanography, 134(??):54–76, May 2015. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661114002171.

Montecino:2009:HCS

[ML09] Vivian Montecino and Carina B. Lange. The Humboldt Current System: Ecosystem components and processes, fisheries, and sediment studies. Progress in Oceanography, 83 (1–4):65–79, December 2009. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661109001049.

Maas:2020:SVP

[MLB^{+}20] Amy E. Maas, Gareth L. Lawson, Alexander J. Bergan, Zhaohui Aleck Wang, and Ann M. Tarrant. Seasonal variation in physiology and shell condition of the pteropod limacina retroversa in the Gulf of Maine relative to life cycle and carbonate chemistry. Progress in Oceanography, 186(??):??, July 2020. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661120301105.

Masuzawa:2003:TVM

[MLD^{+}03] Toshiyuki Masuzawa, Tie Li, Yun Duan, Mineko Yamamoto, Yuriko Hibi, Takeshi Nakatsuka, Hiroshi Kitazato, and Yoshihisa Kato. Temporal variation in mass fluxes and the major components of sinking particles in Sagami Bay, off Japan. Progress in Oceanography, 57(1):59–75, April 2003. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S007966110300051X.

McMonagle:2023:HUF

[MLHE23] Helena McMonagle, Joel K. Llopiz, Ray Hilborn, and Timothy E. Essington. High uncertainty in fish bioenergetics impedes precision of fish-mediated carbon transport estimates into the ocean’s twilight zone. Progress in Oceanography, 217(??):??, September–October 2023. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661123001210.

Megrey:2009:CMEb

[MLHM09] Bernard A. Megrey, Jason S. Link, George L. Hunt, and Erlend Moksness. Comparative marine ecosystem analysis: Applications, opportunities, and lessons learned. Progress in Oceanography, 81(1–4):2–9, April 2009. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661109000196.

Messie:2009:PNP

[MLK^{+}09] Monique Messié, Jesus Ledesma, Dorota D. Kolber, Reiko P. Michisaki, David G. Foley, and Francisco P. Chavez. Potential new production estimates in four eastern boundary upwelling ecosystems. Progress in Oceanography, 83(1–4):151–158, December 2009. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661109000731.

McDonnell:2015:OTC

[MLL^{+}15] Andrew M. P. McDonnell, Phoebe J. Lam, Carl H. Lamborg, Ken O. Buesseler, Richard Sanders, Jennifer S. Riley, Chris Marsay, Helen E. K. Smith, Elizabeth C. Sargent, Richard S. Lampitt, and James K. B. Bishop. The oceanographic toolbox for the collection of sinking and suspended marine particles. Progress in Oceanography, 133(??):17–31, April 2015. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661115000099.

Ma:2022:CIL

[MLL^{+}22] Shuyang Ma, Yang Liu, Jianchao Li, Caihong Fu, Zhenjiang Ye, Peng Sun, Haiqing Yu, Jiahua Cheng, and Yongjun Tian. Climate-induced long-term variations in ecosystem structure and atmosphere-ocean-ecosystem processes in the Yellow Sea and East China Sea. Progress in Oceanography, 175(??):183– 197, ???? 2022. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661118302489.

Megrey:2009:CMEa

[MLM09] Bernard A. Megrey, Jason S. Link, and Erlend Moksness. Comparative marine ecosystem structure and function: Descriptors and characteristics. Progress in Oceanography, 81(1–4): 1, April 2009. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661109000184.

Mueter:2006:BTC

[MLPN06] Franz J. Mueter, Carol Ladd, Michael C. Palmer, and Brenda L. Norcross. Bottom-up and top-down controls of walleye pollock (Theragra chalcogramma) on the Eastern Bering Sea shelf. Progress in Oceanography, 68(2–4):152–183, February 2006. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661106000048.

Mercier:2015:VMO

[MLS^{+}15] Herlé Mercier, Pascale Lherminier, Artem Sarafanov, Fabienne Gaillard, Nathalie Daniault, Damien Desbruyères, Anastasia Falina, Bruno Ferron, Claire Gourcuff, Thierry Huck, and Virginie Thierry. Variability of the meridional overturning circulation at the Greenland–Portugal OVIDE section from 1993 to 2010. Progress in Oceanography, 132(??):250–261, March 2015. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661113002206.

Meincke:2001:D

[MLW01] Jens Meincke, Christian Le Provost, and Jürgen Willebrand. DYNAMO. Progress in Oceanography, 48(??):121–122, ???? 2001. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661101000027.

Merrett:1980:OED

[MM80] N. R. Merrett and N. B. Marshall. Observations on the ecology of deep-sea bottom-living fishes collected off northwest Africa (08^{o}–27^{o}N). Progress in Oceanography, 9(4):185–244, ???? 1980. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661180900026.

Monniot:1990:RBD

[MM90] Claude Monniot and Françoise Monniot. Relationships between deep-sea tunicate populations west and east of the Straits of Gibraltar. Progress in Oceanography, 24(??):289– 296, ???? 1990. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661190900384.

Minobe:1999:IMI

[MM99] Shoshiro Minobe and Nathan Mantua. Interdecadal modulation of interannual atmospheric and oceanic variability over the North Pacific. Progress in Oceanography, 43(2–4):163–192, March 1999. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661199000087.

McCartney:2001:OWI

[MM01] Michael S. McCartney and Cecilie Mauritzen. On the origin of the warm inflow to the Nordic Seas. Progress in Oceanography, 51(1):125–214, ???? 2001. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661101000842.

Mullowney:2016:SDN

[MMD^{+}16] Darrell Mullowney, Gary Maillet, Earl Dawe, George Rose, and Sherrylynn Rowe. Spawning delays of northern capelin (Mallotus villosus) and recovery dynamics: a mismatch with ice-mediated spring bloom? Progress in Oceanography, 141 (??):144–152, February 2016. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661115002529.

Medellin-Mora:2016:CRZ

[MMES16] Johanna Medellín-Mora, Ruben Escribano, and Wolfgang Schneider. Community response of zooplankton to oceanographic changes (2002–2012) in the central/southern upwelling system of Chile. Progress in Oceanography, 142(??):17–29, March 2016. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661116000100.

McCreary:2007:IBI

[MMF^{+}07] Julian P. McCreary, Toru Miyama, Ryo Furue, Tommy Jensen, Hyoun-Woo Kang, Bohyun Bang, and Tangdong Qu. Interactions between the Indonesian Throughflow and circulations in the Indian and Pacific Oceans. Progress in Oceanography, 75(1):70–114, October 2007. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661107001218.

Molina:2012:PCP

[MMF^{+}12] Veronica Molina, Carmen E. Morales, Laura Farías, Marcela Cornejo, Michelle Graco, Yoanna Eissler, and Luis A. Cuevas. Potential contribution of planktonic components to ammonium cycling in the coastal area off centralsouthern Chile during non-upwelling conditions. Progress in Oceanography, 92–95(??):43–49, January 2012. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661111000607.

Maze:2017:CHP

[MMF^{+}17] Guillaume Maze, Herlé Mercier, Ronan Fablet, Pierre Tandeo, Manuel Lopez Radcenco, Philippe Lenca, Charlène Feucher, and Clément Le Goff. Coherent heat patterns revealed by unsupervised classification of Argo temperature profiles in the North Atlantic Ocean. Progress in Oceanography, 151(??):275–292, February 2017. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661116300714.

Miquel:2011:DPF

[MMG^{+}11] Juan-Carlos Miquel, Jacobo Martín, Beat Gasser, Alessia Rodriguez y Baena, Tarik Toubal, and Scott W. Fowler. Dynamics of particle flux and carbon export in the northwestern Mediterranean Sea: a two decade time-series study at the DY-FAMED site. Progress in Oceanography, 91(4):461–481, December 2011. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661111000723.

Maar:2013:SCS

[MMG^{+}13] Marie Maar, Eva Friis Møller, Zeren Gürkan, Sigrún H. Jónasdóttir, and Torkel Gissel Nielsen. Sensitivity of Calanus spp. copepods to environmental changes in the North Sea using life-stage structured models. Progress in Oceanography, 111(??):24–37, April 2013. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661112001280.

Macias:2007:AMB

[MMGL^{+}07] D. Macías, A. P. Martin, J. García-Lafuente, C. M. García, A. Yool, M. Bruno, A. Vázquez-Escobar, A. Izquierdo, D. V. Sein, and F. Echevarría. Analysis of mixing and biogeochemical effects induced by tides on the Atlantic–Mediterranean flow in the Strait of Gibraltar through a physical–biological coupled model. Progress in Oceanography, 74(2–3):252–272, August 2007. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661107000699.

Mariani:2010:MRD

[MMIB10] Patrizio Mariani, Brian R. MacKenzie, Daniele Iudicone, and Alexandra Bozec. Modelling retention and dispersion mechanisms of bluefin tuna eggs and larvae in the northwest Mediterranean Sea. Progress in Oceanography, 86(1–2):45–58, July 2010. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661110000613.

Mironov:2019:TNG

[MMK19] A. N. Mironov, K. V. Minin, and A. V. Kremenetskaia. Two new genera of the family Myriotrochidae (Echinodermata, Holothuroidea). Progress in Oceanography, 178(??):??, November 2019. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S007966111930196X.

Mazurkiewicz:2021:MSS

[MMKS^{+}21] Mikołaj Mazurkiewicz, Kirstin Meyer-Kaiser, Andrew K. Sweetman, Paul E. Renaud, and Maria Włodarska-Kowalczuk. Megabenthic standing stocks and organic carbon demand in a warming Arctic. Progress in Oceanography, 196(??):??, August 2021. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661121001038.

Mazanowski:2023:GZN

[MMMWZ23] Karol Mazanowski, Maciej K. Mańko, Eva F. Møller, and Agata Weydmann-Zwolicka. Gelatinous zooplankton off the Northeast Greenland coast. Progress in Oceanography, 219 (??):??, December 2023. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661123002161.

Miyamoto:2012:GPD

[MMN12] Hiroomi Miyamoto, Ryuji J. Machida, and Shuhei Nishida. Global phylogeography of the deep-sea pelagic chaetognath Eukrohnia hamata. Progress in Oceanography, 104(??):99–109, October 2012. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661112000663.

Mundo:2024:ROS

[MMN^{+}24] Rodrigo Mundo, Tetsuya Matsunaka, Takuya Nakanowatari, Yukiko Taniuchi, Mutsuo Inoue, Hiromi Kasai, Kaisei Mashita, Hayata Mitsunushi, and Seiya Nagao. A review of the oceanographic structure and biological productivity in the southern Okhotsk Sea. Progress in Oceanography, 221(??):??, February 2024. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661123002379.

Moranta:2007:GBT

[MMPG07] Joan Moranta, Enric Massutí, Miquel Palmer, and John D. M. Gordon. Geographic and bathymetric trends in abundance, biomass and body size of four grenadier fishes along the Iberian coast in the western Mediterranean. Progress in Oceanography, 72(1):63–83, January 2007. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661106001170.

Macdonald:2009:WED

[MMR^{+}09] A. M. Macdonald, S. Mecking, P. E. Robbins, J. M. Toole, G. C. Johnson, L. Talley, M. Cook, and S. E. Wijffels. The WOCE-era 3-D Pacific Ocean circulation and heat budget. Progress in Oceanography, 82(4):281–325, September 2009. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661109001335.

Maneesha:2012:UOV

[MMR^{+}12] K. Maneesha, V. S. N. Murty, M. Ravichandran, T. Lee, Weidong Yu, and M. J. McPhaden. Upper ocean variability in the Bay of Bengal during the tropical cyclones Nargis and Laila. Progress in Oceanography, 106(??):49–61, November 2012. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661112000699.

Miller:1988:DGL

[MN88] Charles B. Miller and R. Dreas Nielsen. Development and growth of large, Calanid Copepods in the Ocean Subarctic Pacific, May 1984. Progress in Oceanography, 20(4):275–292, ???? 1988. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661188900432.

Mori:2021:TDB

[MNFY21] Yutaro Mori, Jun Nishioka, Shinzou Fujio, and Youhei Yamashita. Transport of dissolved black carbon from marginal sea sediments to the western North Pacific. Progress in Oceanography, 193(??):??, April–May 2021. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661121000409.

Mestas-Nunez:2006:IVC

[MNM06] Alberto M. Mestas-Nuñez and Arthur J. Miller. Interdecadal variability and climate change in the eastern tropical Pacific: a review. Progress in Oceanography, 69(2–4):267–284, May 2006. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661106000346.

Manche:2024:CME

[MNS^{+}24] Shiva Shankar Manche, Rabindra K. Nayak, Rajesh Sikhakolli, Rajashree V. Bothale, and Prakash Chauhan. Characteristics of mesoscale eddies and their evolution in the north Indian Ocean. Progress in Oceanography, 221(??):??, February 2024. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661124000193.

Martini:2014:RBD

[MNT14] S. Martini, D. Nerini, and C. Tamburini. Relation between deep bioluminescence and oceanographic variables: a statistical analysis using time–frequency decompositions. Progress in Oceanography, 127(??):117–128, September 2014. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661114001220.

Mohammed:2015:MNP

[Moh15] Javeed Shaikh Mohammed. Micro- and nanotechnologies in plankton research. Progress in Oceanography, 134(??):451– 473, May 2015. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661115000518.

Molinari:2004:ADV

[Mol04] Robert L. Molinari. Annual and decadal variability in the western subtropical North Atlantic: signal characteristics and sampling methodologies. Progress in Oceanography, 62(1):33– 66, July 2004. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661104000709.

Molinari:2022:ILD

[Mol22] Robert L. Molinari. Information from low-density expendable bathythermograph transects: North Atlantic mean temperature structure and quasi-decadal variability. Progress in Oceanography, 88(1–4):131–149, ???? 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661110001655.

Morel:1991:LMP

[Mor91] André Morel. Light and marine photosynthesis: a spectral model with geochemical and climatological implications. Progress in Oceanography, 26(3):263–306, ???? 1991. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0079661191900046.

Mosby:1969:PVS

[Mos69] Håkon Mosby. Plans for variability studies in the Norwegian Sea. Progress in Oceanography, 5(??):111–112, ???? 1969. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0079661169900342.

Matrai:2013:SPP

[MOS^{+}13] P. A. Matrai, E. Olson, S. Suttles, V. Hill, L. A. Codispoti, B. Light, and M. Steele. Synthesis of primary production in the Arctic Ocean: I. Surface waters, 1954–2007.

Progress in Oceanography, 110(??):93–106, March 2013. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S007966111200170X.

Miksis-Olds:2013:ERT

[MOSN^{+}13] Jennifer L. Miksis-Olds, Phyllis J. Stabeno, Jeffery M. Napp, Alexei I. Pinchuk, Jeffrey A. Nystuen, Joseph D. Warren, and Samuel L. Denes. Ecosystem response to a temporary sea ice retreat in the Bering Sea: Winter 2009. Progress in Oceanography, 111(??):38–51, April 2013. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661112001590.

McManus:2004:CAP

[MP04] John W. McManus and Johanna F. Polsenberg. Coral–algal phase shifts on coral reefs: Ecological and environmental aspects. Progress in Oceanography, 60(2–4):263–279, February 2004. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S007966110400031X.

Masson:2013:SGE

[MP13] Diane Masson and R. Ian Perry. The Strait of Georgia ecosystem research initiative: an overview. Progress in Oceanography, 115(??):1–5, August 2013. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661113000517.

Menu:2023:TBH

[MPB^{+}23] Clara Menu, Laure Pecquerie, Cedric Bacher, Mathieu Doray, Tarek Hattab, Jeroen van der Kooij, and Martin Huret. Testing the bottom-up hypothesis for the decline in size of anchovy and sardine across European waters through a bioenergetic modeling approach. Progress in Oceanography, 210(??): ??, January 2023. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661122002026.

Morozov:2012:IZF

[MPC12] Andrew Morozov, Jean-Christophe Poggiale, and Flora Cordoleani. Implementation of the zooplankton functional response in plankton models: State of the art, recent challenges and future directions. Progress in Oceanography, 103 (??):80–91, September 2012. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661112000341.

Miller:2017:TVB

[MPC^{+}17] J. A. Miller, W. T. Peterson, L. A. Copeman, X. Du, C. A. Morgan, and M. N. C. Litz. Temporal variation in the biochemical ecology of lower trophic levels in the Northern California Current. Progress in Oceanography, 155(??):1–12, June 2017. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S007966111630252X.

Martinez–Perez:2019:DOM

[MPCNC^{+}19] Alba María Martínez-Pérez, Teresa S. Catalá, Mar Nieto-Cid, Jaime Otero, Marta Álvarez, Mikhail Emelianov, Isabel Reche, Xosé Antón Álvarez-Salgado, and Javier Arístegui. Dissolved organic matter (DOM) in the open Mediterranean Sea. II: Basin–wide distribution and drivers of fluorescent DOM. Progress in Oceanography, 170(??):93–106, January 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661117303579.

Mayers:2022:GMC

[MPD^{+}22] K. M. J. Mayers, A. J. Poulton, C. J. Daniels, S. R. Wells, E. M. S. Woodward, G. A. Tarran, C. E. Widdicombe, D. J. Mayor, A. Atkinson, and S. L. C. Giering. Growth and mortality of coccolithophores during spring in a temperate Shelf Sea (Celtic Sea, April 2015). Progress in Oceanography, 177(??): ??, October 2022. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661117301593.

Mari:2017:TEP

[MPM^{+}17] Xavier Mari, Uta Passow, Christophe Migon, Adrian B. Burd, and Louis Legendre. Transparent exopolymer particles: Effects on carbon cycling in the ocean. Progress in Oceanography, 151(??):13–37, February 2017. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661116300696.

Medel:2018:HBI

[MPM^{+}18] Carolina Medel, Carolina Parada, Carmen E. Morales, Oscar Pizarro, Billy Ernst, and Carlos Conejero. How biophysical interactions associated with sub- and mesoscale structures and migration behavior affect planktonic larvae of the spiny lobster in the Juan Fernández Ridge: a modeling approach. Progress in Oceanography, 162(??):98–119, March 2018. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661116302373.

Marcinko:2013:RMM

[MPMA13] Charlotte L. J. Marcinko, Stuart C. Painter, Adrian P. Martin, and John T. Allen. A review of the measurement and modelling of dinoflagellate bioluminescence. Progress in Oceanography, 109(??):117–129, February 2013. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661112001322.

Mardones:2023:EHA

[MPMFL^{+}23] Jorge I. Mardones, Javier Paredes-Mella, Ana Flores-Leñero, Kyoko Yarimizu, Marcos Godoy, Osvaldo Artal, Andrea Corredor-Acosta, Lara Marcus, Emma Cascales, Juan Pablo Espinoza, Luis Norambuena, René D. Garreaud, Humberto E. González, and José L. Iriarte. Extreme harmful algal blooms, climate change, and potential risk of eutrophication in Patagonian fjords: Insights from an exceptional Heterosigma akashiwo fish-killing event. Progress in Oceanography, 210(??):??, January 2023. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S007966112200180X.

Morissette:2009:CPD

[MPN09] Lyne Morissette, Torstein Pedersen, and Marianne Nilsen. Comparing pristine and depleted ecosystems: the Sørfjord, Norway versus the Gulf of St. Lawrence, Canada. Effects of intense fisheries on marine ecosystems. Progress in Oceanography, 81(1–4):174–187, April 2009. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661109000299.

McQuinn:2015:STV

[MPSD15] Ian H. McQuinn, Stéphane Plourde, Jean-François St. Pierre, and Maxime Dion. Spatial and temporal variations in the abundance, distribution, and aggregation of krill (Thysanoessa raschii and Meganyctiphanes norvegica) in the lower estuary and Gulf of St. Lawrence. Progress in Oceanography, 131(??):159–176, February 2015. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661114002249.

Meyers:1991:STS

[MPSS91] Gary Meyers, Helen Phillips, Neville Smith, and Janet Sprintall. Space and time scales for optimal interpolation of temperature — tropical Pacific Ocean. Progress in Oceanography, 28(3):189–218, ???? 1991. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/007966119190008A.

Moller:2022:DPU

[MPTMK22] Kristof Möller, Marco Pinto-Torres, Jorge I. Mardones, and Bernd Krock. Distribution of phycotoxins in Última Esperanza Province during the PROFAN expedition 2019. Progress in Oceanography, 206(??):??, August 2022. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661122001100.

Mackas:2012:IVM

[MPV12] David L. Mackas, Pierre Pepin, and Hans Verheye. Interannual variability of marine zooplankton and their environments: Within- and between-region comparisons. Progress in Oceanography, 97–100(??):1–14, May 2012. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661111001200.

Moll:2003:RTD

[MR03] Andreas Moll and Günther Radach. Review of threedimensional ecological modelling related to the North Sea shelf system:Part 1: models and their results. Progress in Oceanography, 57(2):175–217, May 2003. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661103000673.

Marshall:2006:MUB

[MR06] John Marshall and Timour Radko. A model of the upper branch of the meridional overturning of the Southern Ocean. Progress in Oceanography, 70(2–4):331–345, August 2006. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S007966110600084X.

MacKenzie:2019:SNT

[MRA^{+}19] K. M. MacKenzie, D. R. Robertson, J. N. Adams, A. H. Altieri, and B. L. Turner. Structure and nutrient transfer in a tropical pelagic upwelling food web: From isoscapes to the whole ecosystem. Progress in Oceanography, 178(??):??, November 2019. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661119300412.

Martinez-Rincon:2022:EEF

[MRAP22] Raúl O. Martínez-Rincón and Tatiana A. Acosta-Pachón. Effect of environmental factors, fish size, and baseline on carbon and nitrogen stable isotope variability in Thunnus alalunga, T. albacares, and T. obesus in the Pacific Ocean. Progress in Oceanography, 203(??):??, April–May 2022. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661122000477.

Mathieu-Resuge:2024:HMF

[MRBS^{+}24] Margaux Mathieu-Resuge, Pablo Brosset, Fany Sardenne, Philippe Soudant, Fabienne Le Grand, Quentin Schull, and Christophe Lebigre. How membrane fatty acids influence sardine size across diverse marine environments. Progress in Oceanography, 221(??):??, February 2024. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661124000156.

Melle:2014:NAO

[MRH^{+}14] Webjørn Melle, Jeffrey Runge, Erica Head, Stéphane Plourde, Claudia Castellani, Priscilla Licandro, James Pierson, Sigrun Jonasdottir, Catherine Johnson, Cecilie Broms, Høgni Debes, Tone Falkenhaug, Eilif Gaard, Astthor Gislason, Michael Heath, Barbara Niehoff, Torkel Gissel Nielsen, Pierre Pepin, Erling Kaare Stenevik, and Guillem Chust. The North Atlantic Ocean as habitat for Calanus finmarchicus: Environmental factors and life history traits. Progress in Oceanography, 129(??):244–284, December 2014. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661114000743.

Matos:2018:CPP

[MRH^{+}18] F. L. Matos, S. W. Ross, V. A. I. Huvenne, J. S. Davies, and M. R. Cunha. Canyons pride and prejudice: Exploring the submarine canyon research landscape, a history of geographic and thematic bias. Progress in Oceanography, 169(??):6–19, December 2018. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661117302744.

Maar:2014:ZME

[MRM^{+}14] Marie Maar, Anna Rindorf, Eva Friis Møller, Asbjørn Christensen, Kristine S. Madsen, and Mikael van Deurs. Zooplankton mortality in 3D ecosystem modelling considering variable spatial–temporal fish consumptions in the North Sea. Progress in Oceanography, 124(??):78–91, May 2014. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661114000287.

Malanotte-Rizzoli:1997:SIS

[MRMD^{+}97] Paola Malanotte-Rizzoli, Beniamino B. Manca, Maurizio Ribera D’Alcalà, Alexander Theocharis, Andrea Bergamasco, Davide Bregant, Giorgio Budillon, Giuseppe Civitarese, Dimitris Georgopoulos, Antonio Michelato, Emilio Sansone, Paolo Scarazzato, and Ekaterini Souvermezoglou. A synthesis of the Ionian Sea hydrography, circulation and water mass pathways during POEM-Phase I. Progress in Oceanography, 39(3):153– 204, ???? 1997. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S007966119700013X.

Marnela:2008:TNS

[MRO^{+}08] Marika Marnela, Bert Rudels, K. Anders Olsson, Leif G. Anderson, Emil Jeansson, Daniel J. Torres, Marie-José Messias, James H. Swift, and Andrew J. Watson. Transports of Nordic Seas water masses and excess SF_{6} through Fram Strait to the Arctic Ocean. Progress in Oceanography, 78(1): 1–11, July 2008. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661108000827.

Magnier:1973:ECW

[MRRC73] Y. Magnier, H. Rotschi, P. Rual, and C. Colin. Equatorial circulation in the western Pacific (170^{o}e). Progress in Oceanography, 6(??):29–30, IN1–IN8, 31–32, IN9–IN24, 33– 34, IN25–IN26, 35–36, IN27–IN34, 37–42, IN35–IN36, 43–44, IN37–IN52, 45–46, ???? 1973. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0079661173900049.

Macklin:2002:VBS

[MRSS02] S. A. Macklin, V. I. Radchenko, S. Saitoh, and P. J. Stabeno. Variability in the Bering Sea ecosystem. Progress in Oceanography, 55(1–2):1–4, October 2002. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661102001155.

Mohn:2014:LBH

[MRW^{+}14] Christian Mohn, Anna Rengstorf, Martin White, Gerard Duineveld, Furu Mienis, Karline Soetaert, and Anthony Grehan. Linking benthic hydrodynamics and cold-water coral occurrences: a high-resolution model study at three cold-water coral provinces in the NE Atlantic. Progress in Oceanography, 122(??):92–104, March 2014. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661113002413.

Miller:2000:ICR

[MS00] Arthur J. Miller and Niklas Schneider. Interdecadal climate regime dynamics in the North Pacific Ocean: theories, observations and ecosystem impacts. Progress in Oceanography, 47(2–4):355–379, October 2000. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661100000446.

Mendelssohn:2002:CUT

[MS02] R. Mendelssohn and F. B. Schwing. Common and uncommon trends in SST and wind stress in the California and Peru– Chile current systems. Progress in Oceanography, 53(??):141– 162, ???? 2002. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661102000289.

Moore:2015:SAR

[MS15] Sue E. Moore and Phyllis J. Stabeno. Synthesis of Arctic research (SOAR) in marine ecosystems of the Pacific Arctic. Progress in Oceanography, 136(??):1–11, August 2015. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661115001147.

McKinnell:2017:AEF

[MS17] Skip McKinnell and Michael P. Seki. Arcane epipelagic fishes of the subtropical North Pacific and factors associated with their distribution. Progress in Oceanography, 150(??):48–61, January 2017. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661116301252.

Miller:2022:DAL

[MSA^{+}22] Michael J. Miller, Madoka Shimizu, Jun Aoyama, Shun Watanabe, Mari Kuroki, Eric Feunteun, Takatoshi Higuchi, Aya Takeuchi, Yu-San Han, Pierre Sasal, Christine Dupuy, Donald Jellyman, Robert Schabetsberger, Shingo Kimura, Noritaka Mochioka, Tsuguo Otake, and Katsumi Tsukamoto.

Distribution and abundance of leptocephali in the western South Pacific region during two large-scale sampling surveys. Progress in Oceanography, 206(??):??, August 2022. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661122001124.

Morell:2023:BOB

[MSB^{+}23] Alaia Morell, Yunne-Jai Shin, Nicolas Barrier, Morgane Travers-Trolet, Ghassen Halouani, and Bruno Ernande. Bioen-OSMOSE: a bioenergetic marine ecosystem model with physiological response to temperature and oxygen. Progress in Oceanography, 216(??):??, August 2023. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661123001076.

MacIntyre:2015:RBS

[MSC^{+}15] Kalyn Q. MacIntyre, Kathleen M. Stafford, Paul B. Conn, Kristin L. Laidre, and Peter L. Boveng. The relationship between sea ice concentration and the spatio-temporal distribution of vocalizing bearded seals (Erignathus barbatus) in the Bering, Chukchi, and Beaufort Seas from 2008 to 2011. Progress in Oceanography, 136(??):241–249, August 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661115001056.

Mendes:2016:NIW

[MSd^{+}16] Renato Mendes, Magda Catarina Sousa, Maite deCastro, Moncho Gómez-Gesteira, and João Miguel Dias. New insights into the Western Iberian Buoyant Plume: Interaction between the Douro and Minho River plumes under winter conditions. Progress in Oceanography, 141(??):30–43, February 2016. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661115002414.

Maury:2007:MEEb

[MSF^{+}07] Olivier Maury, Yunne-Jai Shin, Blaise Faugeras, Tamara Ben Ari, and Francis Marsac. Modeling environmental effects on the size-structured energy flow through marine ecosystems. Part 2: Simulations. Progress in Oceanography, 74(4):500–514, September 2007. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661107000997.

Mordukhovich:2019:ONG

[MSFZ19] Vladimir V. Mordukhovich, Alexandr A. Semenchenko, Natalya P. Fadeeva, and Julia K. Zograf. One new genus and two new free-living deep-sea nematode species with discussion of phylogeny of the family leptosomatidae filipjev, 1916. Progress in Oceanography, 178(??):??, November 2019. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661119302356.

Miladinova:2018:FCB

[MSGGM18] S. Miladinova, A. Stips, E. Garcia-Gorriz, and D. Macias Moy. Formation and changes of the Black Sea cold intermediate layer. Progress in Oceanography, 167(??):11–23, October 2018. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661117303117.

McKinnell:2017:SIA

[MSI17] Skip McKinnell, Michael P. Seki, and Taro Ichii. Special issue on the advances in understanding of the North Pacific subtropical front ecosystem. Progress in Oceanography, 150(??):1–2, January 2017. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661117300150.

McCarthy:2015:MAM

[MSJ^{+}15] G. D. McCarthy, D. A. Smeed, W. E. Johns, E. Frajka-Williams, B. I. Moat, D. Rayner, M. O. Baringer, C. S. Meinen, J. Collins, and H. L. Bryden. Measuring the Atlantic Meridional Overturning Circulation at 26^{o}N. Progress in Oceanography, 130(??):91–111, January 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661114001694.

Munoz:2007:TVP

[MSL^{+}07] Praxedes Muñoz, Javier Sellanes, Carina Lange, Maritza Palma, and Marco A. Salamanca. Temporal variability of^{210}Pb fluxes and bioturbation in shelf sediments beneath the high primary production area off Concepción, central-southern Chile (36^{o}S). Progress in Oceanography, 75(3):586–602, November 2007. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661107001711.

Messie:2019:UFB

[MSMH19] Monique Messié, Igor Shulman, Séverine Martini, and Steven H. D. Haddock. Using fluorescence and bioluminescence sensors to characterize auto- and heterotrophic plankton communities. Progress in Oceanography, 171(??):76–92, February 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661118300478.

Mackas:1993:VHP

[MSMR93] David L. Mackas, Hugh Sefton, Charles B. Miller, and Anne Raich. Vertical habitat partitioning by large calanoid copepods in the oceanic subarctic Pacific during spring. Progress in Oceanography, 32(??):259–294, ???? 1993. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0079661193900178.

Mizobata:2002:BSC

[MSS^{+}02] K. Mizobata, S. I. Saitoh, A. Shiomoto, T. Miyamura, N. Shiga, K. Imai, M. Toratani, Y. Kajiwara, and K. Sasaoka. Bering Sea cyclonic and anticyclonic eddies observed during summer 2000 and 2001. Progress in Oceanography, 55(1–2):65–75, October 2002. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661102000708.

Matsuno:2023:RDA

[MST^{+}23a] Kohei Matsuno, Kohei Sumiya, Manami Tozawa, Daiki Nomura, Hiroko Sasaki, Atsushi Yamaguchi, and Hiroto Murase. Responses of diatom assemblages and life cycle to sea ice variation in the eastern Indian sector of the Southern Ocean during austral summer 2018/2019. Progress in Oceanography, 218(??):??, November 2023. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S007966112300160X.

Myksvoll:2023:KPP

[MST^{+}23b] Mari S. Myksvoll, Anne Britt Sandø, Jerry Tjiputra, Annette Samuelsen, Veli Çağlar Yumruktepe, Camille Li, Erik A. Mousing, Joao P. H. Bettencourt, and Geir Ottersen. Key physical processes and their model representation for projecting climate impacts on subarctic Atlantic net primary production: a synthesis. Progress in Oceanography, 217(??):??, September– October 2023. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661123001271.

Munoz:2014:GRF

[MSV^{+}14] Práxedes Muñoz, Javier Sellanes, Katherine Villalobos, Germán Zapata-Hernández, Christoph Mayr, and Karen Araya. Geochemistry of reduced fluids from shallow cold vents hosting chemosynthetic communities (Comau Fjord, Chilean Patagonia, ≈ 42^{o}S). Progress in Oceanography, 129 (??):159–169, December 2014. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661114001669.

Mackas:1999:MEW

[MT99] D. L. Mackas and A. Tsuda. Mesozooplankton in the eastern and western subarctic Pacific: community structure, seasonal life histories, and interannual variability. Progress in Oceanography, 43(2–4):335–363, March 1999. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661199000129.

Marcos:2012:IAD

[MTC12] Marta Marcos, Michael N. Tsimplis, and Francesc M. Calafat. Inter-annual and decadal sea level variations in the north-western Pacific marginal seas. Progress in Oceanography, 105(??):4–21, October 2012. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661112000432.

Meerhoff:2014:SSM

[MTC14] Erika Meerhoff, Fabián J. Tapia, and Leonardo R. Castro. Spatial structure of the meroplankton community along a Patagonian fjord — the role of changing freshwater inputs. Progress in Oceanography, 129(??):125–135, December 2014. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661114001086.

Morales:2010:MSC

[MTH^{+}10] Carmen E. Morales, M. Loreto Torreblanca, Samuel Hormazabal, Marco Correa-Ramírez, Sergio Nuñez, and Pamela Hidalgo. Mesoscale structure of copepod assemblages in the coastal transition zone and oceanic waters off central-southern Chile. Progress in Oceanography, 84(3–4):158–173, March 2010. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661109001906.

Miyamoto:2022:CCS

[MTK^{+}22] Hiroomi Miyamoto, Kazutaka Takahashi, Hiroshi Kuroda, Tsuyoshi Watanabe, Yukiko Taniuchi, Akira Kuwata, Hiromi Kasai, Shigeho Kakehi, Taiki Fuji, Satoshi Suyama, and Kazuaki Tadokoro. Copepod community structure in the transition region of the North Pacific Ocean: Water mixing as a key driver of secondary production enhancement in subarctic and subtropical waters. Progress in Oceanography, 207(??):??, September–October 2022. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661122001240.

Millot:2005:AEL

[MTL05] C. Millot and I. Taupier-Letage. Additional evidence of LIW entrainment across the Algerian subbasin by mesoscale eddies and not by a permanent westward flow. Progress in Oceanography, 66(2–4):231–250, August 2005. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661105000728.

Munk:1969:DST

[Mun69] W. H. Munk. Deep-sea tides. Progress in Oceanography, 5(??): 67–69, ???? 1969. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661169900287.

Munk:1997:OAO

[Mun97] Walter Munk. Once again: once again — tidal friction. Progress in Oceanography, 40(??):7–35, ???? 1997. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661197000219.

Mariani:2010:OEM

[MV10] Patrizio Mariani and André W. Visser. Optimization and emergence in marine ecosystem models. Progress in Oceanography, 84(1–2):89–92, January 2010. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661109001438.

Molina-Valdivia:2021:OIE

[MVBC^{+}21] Víctor Molina-Valdivia, Claudia A. Bustos, Manuel I. Castillo, Francesca V. Search, Guido Plaza, and Mauricio F. Landaeta. Oceanographic influences on the early life stages of a mesopelagic fish across the Chilean Patagonia. Progress in Oceanography, 195(??):??, July 2021. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661121000598.

Melin:2011:MSS

[MVC^{+}11] F. Mélin, V. Vantrepotte, M. Clerici, D. D’Alimonte, G. Zibordi, J.-F. Berthon, and E. Canuti. Multi-sensor satellite time series of optical properties and chlorophyll-a concentration in the Adriatic Sea. Progress in Oceanography, 91(3):229–244, November 2011. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661110001539.

Maranger:2015:PAP

[MVN^{+}15] Roxane Maranger, Dolors Vaqué, Dan Nguyen, Marie-Pier Hébert, and Elena Lara. Pan-Arctic patterns of planktonic heterotrophic microbial abundance and processes: Controlling factors and potential impacts of warming. Progress in Oceanography, 139(??):221–232, December 2015. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661115001585.

Morrow:2008:OSS

[MVS08] Rosemary Morrow, Guillaume Valladeau, and Jean-Baptiste Sallee. Observed subsurface signature of Southern Ocean sea level rise. Progress in Oceanography, 77(4):351–366, June 2008. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661108000682.

Moullec:2019:CBP

[MVV^{+}19] Fabien Moullec, Laure Velez, Philippe Verley, Nicolas Barrier, Caroline Ulses, Pierluigi Carbonara, Antonio Esteban, Cristina Follesa, Michele Gristina, Angélique Jadaud, Alessandro Ligas, Eduardo López Díaz, Porzia Maiorano, Panagiota Peristeraki, Maria Teresa Spedicato, Ioannis Thasitis, Maria Valls, François Guilhaumon, and Yunne-Jai Shin. Capturing the big picture of Mediterranean marine biodiversity with an end-to-end model of climate and fishing impacts. Progress in Oceanography, 178(??):??, November 2019. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661118303537.

Myers:1996:CNP

[MW96] Paul G. Myers and Andrew J. Weaver. On the circulation of the North Pacific Ocean: climatology, seasonal cycle and interpentadal variability. Progress in Oceanography, 38(1):1– 49, ???? 1996. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661196000092.

Moore:2002:CDR

[MWFH02] S. E. Moore, J. M. Waite, N. A. Friday, and T. Honkalehto. Cetacean distribution and relative abundance on the central–eastern and the southeastern Bering Sea shelf with reference to oceanographic domains. Progress in Oceanography, 55(1–2):249–261, October 2002. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661102000824.

Messias:2008:GST

[MWJ^{+}08] M.-J. Messias, A. J. Watson, T. Johannessen, K. I. C. Oliver, K. A. Olsson, E. Fogelqvist, J. Olafsson, S. Bacon, J. Balle, N. Bergman, G. Budéus, M. Danielsen, J.-C. Gascard, E. Jeansson, S. R. Olafsdottir, K. Simonsen, T. Tanhua, K. Van Scoy, and J. R. Ledwell. The Greenland Sea tracer experiment 1996–2002: Horizontal mixing and transport of Greenland Sea intermediate water. Progress in Oceanography, 78(1): 85–105, July 2008. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661108000852.

Morimoto:2012:IVM

[MWO^{+}12] Akihiko Morimoto, Atsushi Watanabe, Goh Onitsuka, Tetsutaro Takikawa, Masatoshi Moku, and Tetsuo Yanagi. Interannual variations in material transport through the eastern channel of the Tsushima/Korea Straits. Progress in Oceanography, 105(??):38–46, October 2012. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661112000444.

Meredith:2010:CFC

[MWS^{+}10] Michael P. Meredith, Margaret I. Wallace, Sharon E. Stammerjohn, Ian A. Renfrew, Andrew Clarke, Hugh J. Venables, Deborah R. Shoosmith, Terri Souster, and Melanie J. Leng. Changes in the freshwater composition of the upper ocean west of the Antarctic Peninsula during the first decade of the 21st century. Progress in Oceanography, 87(1–4):127–143, October 2010. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S007966111000131X.

Ma:2021:IME

[MXC^{+}21] Wentao Ma, Peng Xiu, Fei Chai, Lihua Ran, Martin G. Wiesner, Jingyuan Xi, Yunwei Yan, and Erick Fredj. Impact of mesoscale eddies on the source funnel of sediment trap measurements in the South China Sea. Progress in Oceanography, 194(??):??, June 2021. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661121000537.

McCreary:1992:EDL

[MY92] Julian P. McCreary and Zuojun Yu. Equatorial dynamics in a 212-layer model. Progress in Oceanography, 29(1):61–132, ???? 1992. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/007966119290003I.

Ma:2023:RPM

[MY23] Qian-Yao Ma and Gui-Peng Yang. Roles of phytoplankton, microzooplankton, and bacteria in DMSP and DMS transformation processes in the East China Continental Sea. Progress in Oceanography, 213(??):??, April–May 2023. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661123000460.

McCreary:2022:DIO

[MYH^{+}22] Julian P. McCreary, Zuojun Yu, Raleigh R. Hood, P. N. Vinaychandran, Ryo Furue, Akio Ishida, and Kelvin J. Richards. Dynamics of the Indian-ocean oxygen minimum zones. Progress in Oceanography, 112–113(??):15–37, ???? 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S007966111300027X.

Mendoza:2014:TVD

[MZ14] Wilson G. Mendoza and Rod G. Zika. On the temporal variation of DOM fluorescence on the southwest Florida continental shelf. Progress in Oceanography, 120(??):189–204, January 2014. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661113001705.

Martin:2008:MSV

[MZF^{+}08] Adrian P. Martin, Mikhail V. Zubkov, Michael J. Fasham, Peter H. Burkill, and Ross J. Holland. Microbial spatial variability: an example from the Celtic Sea. Progress in Oceanography, 76(4):443–465, March 2008. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661108000049.

Moore:2020:RBS

[MZGA^{+}20] Andrew Moore, Javier Zavala-Garay, Hernan G. Arango, Christopher A. Edwards, Jeffrey Anderson, and Tim Hoar. Regional and basin scale applications of ensemble adjustment Kalman filter and 4D-Var ocean data assimilation systems. Progress in Oceanography, 189(??):??, November–December 2020. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661120301889.

Ma:2023:OIV

[MZH^{+}23] Jie Ma, Zhanqing Zhang, Shijian Hu, Cesar Villanoy, Xinying Guo, Fan Wang, and Dunxin Hu. Observed interannual variability of the Kuroshio and Luzon Undercurrent associated with tropical and subtropical wind forcing. Progress in Oceanography, 219(??):??, December 2023. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661123001891.

Mikaelyan:2023:CWM

[MZK^{+}23] Alexander S. Mikaelyan, Andrey G. Zatsepin, Arseny A. Kubryakov, Oleg I. Podymov, Sergey A. Mosharov, Larisa A. Pautova, Alexey V. Fedorov, and Oksana A. Ocherednik. Case where a mesoscale cyclonic eddy suppresses primary production: a stratification–lock hypothesis. Progress in Oceanography, 212(??):??, March 2023. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661123000277.

Miao:2023:SHS

[MZZ^{+}23] Mingfang Miao, Zhiwei Zhang, Jinchao Zhang, Yue Wang, Wei Zhao, and Jiwei Tian. Steric heights of submesoscale processes and internal gravity waves in the subtropical northwestern Pacific and northern South China Sea as revealed by moored observations. Progress in Oceanography, 219(??):??, December 2023. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S007966112300201X.

Nagasawa:2001:LTV

[Nag01] Kazuya Nagasawa. Long-term variations in abundance of Pacific herring (Clupea pallasi) in Hokkaido and Sakhalin related to changes in environmental conditions. Progress in Oceanography, 49(??):551–564, ???? 2001. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661101000404.

Nomura:2021:FDS

[NAH^{+}21] Daiki Nomura, Hiroto Abe, Toru Hirawake, Atsushi Ooki, Youhei Yamashita, Aiko Murayama, Kazuya Ono, and Jun Nishioka. Formation of dense shelf water associated with sea ice freezing in the Gulf of Anadyr estimated with oxygen isotopic ratios. Progress in Oceanography, 196(??):??, August 2021. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661121000823.

Nayudu:1965:PSV

[Nay65] Y. Rammohanroy Nayudu. Petrology of submarine volcanics and sediments in the vicinity of the Mendocino Fracture Zone. Progress in Oceanography, 3(??):207–211, IN1–IN2, 212–220, ???? 1965. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661165900194.

Newman:1987:HEN

[NB87] Walter S. Newman and Cecile Baeteman. Holocene excursions of the Northwest European geoid. Progress in Oceanography, 18(1–4):287–322, ???? 1987. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0079661187900371.

Naqvi:2005:BOA

[NBG^{+}05] S. Wajih A. Naqvi, Hermann W. Bange, Stuart W. Gibb, Catherine Goyet, Angela D. Hatton, and Robert C. Upstill-Goddard. Biogeochemical ocean–atmosphere transfers in the Arabian Sea. Progress in Oceanography, 65(2–4):116–144, May 2005. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661105000431.

New:2001:OPS

[NBHM01] A. L. New, S. Barnard, P. Herrmann, and J.-M. Molines. On the origin and pathway of the saline inflow to the Nordic Seas: insights from models. Progress in Oceanography, 48(??):255– 287, ???? 2001. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661101000076.

Nowak:2020:GSH

[NBLI20] B. V. R. Nowak, W. D. Bowen, D. C. Lidgard, and S. J. Iverson. Grey seals (Halichoerus grypus) as bioprobes: Fine-scale measurements of oceanographic properties using an instrumented large marine predator. Progress in Oceanography, 189(??):??, November–December 2020. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661120301907.

Nilsson:2008:LFT

[NBR^{+}08] Johan Nilsson, Göran Björk, Bert Rudels, Peter Winsor, and Daniel Torres. Liquid freshwater transport and Polar Surface Water characteristics in the East Greenland Current during the AO-02 Oden expedition. Progress in Oceanography, 78(1):45– 57, July 2008. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661108000840.

Nilsson:1980:FEE

[NC80] C. S. Nilsson and G. R. Cresswell. The formation and evolution of East Australian current warm-core eddies. Progress in Oceanography, 9(3):133–183, ???? 1980. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0079661180900087.

Navarro:2015:RRE

[NCC^{+}15] Joan Navarro, Marta Coll, Laura Cardador, Ángel M. Fernández, and José M. Bellido. The relative roles of the environment, human activities and spatial factors in the spatial distribution of marine biodiversity in the Western Mediterranean Sea. Progress in Oceanography, 131(??):126–137, February 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661114002067.

Neuer:2007:BHE

[NCH^{+}07] Susanne Neuer, Andres Cianca, Peer Helmke, Tim Freudenthal, Robert Davenport, Helge Meggers, Michaela Knoll, J. Magdalena Santana-Casiano, Melchor González-Davila, Maria-José Rueda, and Octavio Llinás. Biogeochemistry and hydrography in the eastern subtropical North Atlantic gyre. Results from the European time-series station ESTOC. Progress in Oceanography, 72(1):1–29, January 2007. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661106001157.

Nagy:2022:ICC

[NDEG22] Hazem Nagy, Emanuele Di Lorenzo, and Ahmed El-Gindy. The impact of climate change on circulation patterns in the Eastern Mediterranean Sea upper layer using Med-ROMS model. Progress in Oceanography, 175(??):226–244, ???? 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661118300958.

Needler:1985:AVF

[Nee85] G. T. Needler. The absolute velocity as a function of conserved measurable quantities. Progress in Oceanography, 14(??):421– 429, ???? 1985. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661185900205.

Nagao:2022:BSC

[NEI^{+}22] Ippei Nagao, Young-Joon Eum, Yoko Iwamoto, Yuya Tada, Koji Suzuki, Atsushi Tsuda, Mitsuhiro Toratani, Koji Hamasaki, and Mitsuo Uematsu. Biogenic sulfur compounds in spring phytoplankton bloom in the western North Pacific off the coast of northern Japan. Progress in Oceanography, 165(??):145–157, ???? 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661117301891.

Newman:1987:PNR

[NF87] Walter S. Newman and Rhodes W. Fairbridge. Project NOAH: Regulating modern sea-level rise. Phase II: Jerusalem underground. Progress in Oceanography, 18(1–4):61–78, ???? 1987. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0079661187900267.

Nilsen:2006:VML

[NF06] J. Even Ø. Nilsen and Eva Falck. Variations of mixed layer properties in the Norwegian Sea for the period 1948– 1999. Progress in Oceanography, 70(1):58–90, July 2006. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661106000383.

Nava-Fernandez:2022:PFA

[NFMCS^{+}22] Xinantecatl Antonio Nava-Fernández, María Luisa Machain-Castillo, Abdelfettah Sifeddine, Mercedes Mendez-Millan, Sandrine Caquineau, Kelly A. Gibson, Eric Tappa, Ana Carolina Ruiz-Fernández, Joan-Albert Sanchez-Cabeza, Alejandro Rodríguez-Ramírez, and Andrea Rebeca Lara-Cera. Planktonic foraminiferal assemblages and particle fluxes at the entrance of the Gulf of California highlight the effects of the strong El Niño 2015–2016. Progress in Oceanography, 208 (??):??, November 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661122001392.

Negrete-Garcia:2022:PEF

[NGLL^{+}22] Gabriela Negrete-García, Jessica Y. Luo, Matthew C. Long, Keith Lindsay, Michael Levy, and Andrew D. Barton. Plankton energy flows using a global size-structured and traitbased model. Progress in Oceanography, 209(??):??, December 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661122001574.

Naranjo:2014:HMD

[NGLSSG14] C. Naranjo, J. Garcia-Lafuente, G. Sannino, and J. C. Sanchez-Garrido. How much do tides affect the circulation of the Mediterranean Sea? From local processes in the Strait of Gibraltar to basin-scale effects. Progress in Oceanography, 127(??):108–116, September 2014. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S007966111400113X.

Nogueira:2012:IPP

[NGNV12] Enrique Nogueira, Gonzalo González-Nuevo, and Luis Valdés. The influence of phytoplankton productivity, temperature and environmental stability on the control of copepod diversity in the North East Atlantic. Progress in Oceanography, 97–100(??):92–107, May 2012. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661111001273.

Neuheimer:2010:ERV

[NGPH10] A. B. Neuheimer, W. C. Gentleman, P. Pepin, and E. J. H. Head. Explaining regional variability in copepod recruitment: Implications for a changing climate. Progress in Oceanography, 87(1–4):94–105, October 2010. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661110001187.

Nelson:1983:BUA

[NH83] G. Nelson and L. Hutchings. The Benguela upwelling area. Progress in Oceanography, 12(3):333–356, ???? 1983. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0079661183900137.

Nakata:1988:SFF

[NH88a] Kisaburo Nakata and Kyoko Hata. Subtidal frequency fluctuations of current velocity and sea level in the Japan Sea. Progress in Oceanography, 21(3–4):227–240, ???? 1988. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0079661188900043.

Nakata:1988:IEC

[NH88b] Kisaburo Nakata and Fumio Horiguchi. Intercomparison experiment of current meters in the upper ocean. Progress in Oceanography, 21(3–4):525–536, ???? 1988. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0079661188900262.

Nichol:2013:DON

[NHE^{+}13] Linda M. Nichol, Anna M. Hall, Graeme M. Ellis, Eva Stredulinsky, Melissa Boogaards, and John K. B. Ford. Dietary overlap and niche partitioning of sympatric harbour porpoises and Dall’s porpoises in the Salish Sea. Progress in Oceanography, 115(??):202–210, August 2013. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S007966111300058X.

Navarro:2019:PCG

[NHG19] Nelso P. Navarro, Pirjo Huovinen, and Iván Gómez. Photosynthetic characteristics of geographically disjunct seaweeds: a case study on the early life stages of Antarctic and subantarctic species. Progress in Oceanography, 174(??):28–36, May 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661118300673.

Nagatomo:2023:GDM

[NHH^{+}23] Yutaro Nagatomo, Sachiko Horii, Junya Hirai, Fuminori Hashihama, Tetsuya Sado, Takehiko Fukuchi, Masaki Miya, and Kazutaka Takahashi. Geographic distribution of micronektonic fish communities in the subtropical North Pacific: the effect of primary productivity and nitrogen fixation. Progress in Oceanography, 217(??):??, September– October 2023. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661123001295.

Nishioka:2021:IND

[NHN^{+}21] Jun Nishioka, Toru Hirawake, Daiki Nomura, Youhei Yamashita, Kazuya Ono, Aiko Murayama, Alexey Shcherbinin, Yuri N. Volkov, Humio Mitsudera, Naoto Ebuchi, Masaaki Wakatsuchi, and Ichiro Yasuda. Iron and nutrient dynamics along the East Kamchatka Current, western Bering Sea Basin and Gulf of Anadyr. Progress in Oceanography, 198(??):??, November 2021. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661121001476.

Nakada:2014:OOP

[NHS^{+}14] Satoshi Nakada, Naoki Hirose, Tomoharu Senjyu, Ken ichi Fukudome, Toshihiro Tsuji, and Noriyuki Okei. Operational ocean prediction experiments for smart coastal fishing. Progress in Oceanography, 121(??):125–140, February 2014. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661113001948.

Nemcek:2023:SSD

[NHSP23] Nina Nemcek, Melissa Hennekes, Akash Sastri, and R. Ian Perry. Seasonal and spatial dynamics of the phytoplankton community in the Salish Sea, 2015–2019. Progress in Oceanography, 217(??):??, September–October 2023. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661123001519.

Napolitano:2019:TIW

[NIC^{+}19] Ernesto Napolitano, Roberto Iacono, Tiziana Ciuffardi, Franco Reseghetti, Pierre-Marie Poulain, and Giulio Notarstefano. The Tyrrhenian Intermediate Water (TIW): Characterization and formation mechanisms. Progress in Oceanography, 170(??):53–68, January 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661118300351.

Niehoff:2007:LHS

[Nie07] Barbara Niehoff. Life history strategies in zooplankton communities: the significance of female gonad morphology and maturation types for the reproductive biology of marine calanoid copepods. Progress in Oceanography, 74(1):1–47, July 2007. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661107000432.

Nowald:2015:TSS

[NIF^{+}15] N. Nowald, M. H. Iversen, G. Fischer, V. Ratmeyer, and G. Wefer. Time series of in-situ particle properties and sediment trap fluxes in the coastal upwelling filament off Cape Blanc, Mauritania. Progress in Oceanography, 137 (part A)(??):1–11, September 2015. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661114002250.

New:2001:RAC

[NJCD01] A. L. New, Y. Jia, M. Coulibaly, and J. Dengg. On the role of the Azores Current in the ventilation of the North Atlantic Ocean. Progress in Oceanography, 48(??):163–194, ???? 2001. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661101000040.

Nakatsuka:2003:PDDb

[NKK03] T. Nakatsuka, J. Kanda, and H. Kitazato. Particle dynamics in the deep water column of Sagami Bay, Japan. II: Seasonal change in profiles of suspended phytodetritus. Progress in Oceanography, 57(1):47–57, April 2003. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661103000508.

Noiri:2005:IIT

[NKK^{+}05] Yoshifumi Noiri, Isao Kudo, Hiroshi Kiyosawa, Jun Nishioka, and Atsushi Tsuda. Influence of iron and temperature on growth, nutrient utilization ratios and phytoplankton species composition in the western subarctic Pacific Ocean during the SEEDS experiment. Progress in Oceanography, 64(2–4):149–166, February 2005. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661105000273.

Nieto:2017:VCC

[NM17] Karen Nieto and Frédéric Mélin. Variability of chlorophylla concentration in the Gulf of Guinea and its relation to physical oceanographic variables. Progress in Oceanography, 151(??):97–115, February 2017. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661116301446.

Neira:2009:MAR

[NMC^{+}09] Sergio Neira, Coleen L. Moloney, Philippe Cury, Christian Mullon, and Villy Christensen. Mechanisms affecting recovery in an upwelling food web: the case of the southern Humboldt. Progress in Oceanography, 83(1–4):404–416, December 2009. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661109000718.

Nakatsuka:2003:PDDa

[NMK^{+}03] T. Nakatsuka, T. Masuzawa, J. Kanda, H. Kitazato, Y. Shirayama, M. Shimanaga, and A. Yamaoka. Particle dynamics in the deep water column of Sagami Bay, Japan. I: Origins of apparent flux of sinking particles. Progress in Oceanography, 57(1):31–45, April 2003. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661103000491.

Nevarez-Martinez:2001:DAP

[NMLBCM^{+}01] Manuel O. Nevárez-Martínez, Daniel Lluch-Belda, Miguel A. Cisneros-Mata, J. Pablo Santos-Molina, Ma De los Angeles Martínez-Zavala, and Salvador E. Lluch-Cota. Distribution and abundance of the Pacific sardine (Sardinops sagax) in the Gulf of California and their relation with the environment. Progress in Oceanography, 49(??):565–580, ???? 2001. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661101000416.

Nonomura:2008:SVC

[NMN08] Takumi Nonomura, Ryuji J. Machida, and Shuhei Nishida. Stage-v copepodites of Calanus sinicus and Calanus jashnovi (Copepoda: Calanoida) in mesopelagic zone of Sagami Bay as identified with genetic markers, with special reference to their vertical distribution. Progress in Oceanography, 77(1): 45–55, April 2008. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661108000311.

Nishikawa:2021:SWP

[NMO^{+}21] Hatsumi Nishikawa, Humio Mitsudera, Takeshi Okunishi, Shin ichi Ito, Taku Wagawa, Daisuke Hasegawa, Toru Miyama, Hitoshi Kaneko, and Ren-Chieh Lien. Surface water pathways in the subtropical–subarctic frontal zone of the western North Pacific. Progress in Oceanography, 199(??):??, December 2021. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661121001750.

Nishioka:2014:BPP

[NMY^{+}14] Jun Nishioka, Humio Mitsudera, Ichiro Yasuda, Hongbin Liu, Takeshi Nakatsuka, and Yuri N. Volkov. Biogeochemical and physical processes in the Sea of Okhotsk and the linkage to the Pacific Ocean. Progress in Oceanography, 126(??):1–7, August 2014. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661114000755.

Nadai:2021:ESS

[NNFL21] Gabrielle Nadaï, Eva-Maria Nöthig, Louis Fortier, and Catherine Lalande. Early snowmelt and sea ice breakup enhance algal export in the Beaufort Sea. Progress in Oceanography, 190(??): ??, January 2021. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661120302147.

Nakanowatari:2021:IDV

[NNM^{+}21] Takuya Nakanowatari, Tomohiro Nakamura, Humio Mitsudera, Jun Nishioka, Hiroshi Kuroda, and Keisuke Uchimoto. Interannual to decadal variability of phosphate in the Oyashio region: Roles of wind-driven ocean current and tidally induced vertical mixing in the Sea of Okhotsk. Progress in Oceanography, 197(??):??, September–October 2021. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661121001026.

Nishioka:2014:QEI

[NNO^{+}14] Jun Nishioka, Takeshi Nakatsuka, Kazuya Ono, Yu. N. Volkov, Alexey Scherbinin, and Takayuki Shiraiwa. Quantitative evaluation of iron transport processes in the Sea of Okhotsk.

Progress in Oceanography, 126(??):180–193, August 2014. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661114000597.

Nakanowatari:2014:CSL

[NO14] Takuya Nakanowatari and Kay I. Ohshima. Coherent sea level variation in and around the Sea of Okhotsk. Progress in Oceanography, 126(??):58–70, August 2014. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661114000950.

Nof:1996:WRR

[Nof96] Doron Nof. Why are rings regularly shed in the western equatorial Atlantic but not in the western Pacific? Progress in Oceanography, 38(4):417–451, ???? 1996. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661197000049.

Nof:2000:WMA

[Nof00] Doron Nof. Why much of the Atlantic circulation enters the Caribbean Sea and very little of the Pacific circulation enters the Sea of Japan. Progress in Oceanography, 45(1):39–67, January 2000. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661199000506.

Nof:2003:SOG

[Nof03] D. Nof. The Southern Ocean’s grip on the northward meridional flow. Progress in Oceanography, 56(2):223–247, February 2003. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661103000053.

New:2000:IIS

[NP00] A. L. New and R. D. Pingree. An intercomparison of internal solitary waves in the Bay of Biscay and resulting from Korteweg–de Vries-type theory. Progress in Oceanography, 45(1):1–38, January 2000. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S007966119900049X.

Navarro:2019:AFH

[NPO^{+}19] Jorge M. Navarro, Kurt Paschke, Alejandro Ortiz, Luis Vargas-Chacoff, Luis Miguel Pardo, and Nelson Valdivia. The Antarctic fish Harpagifer antarcticus under current temperatures and salinities and future scenarios of climate change. Progress in Oceanography, 174(??):37–43, May 2019. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661118300703.

Nunez-Riboni:2017:QIM

[NRA17] Ismael Núñez-Riboni and Anna Akimova. Quantifying the impact of the major driving mechanisms of interannual variability of salinity in the North Sea. Progress in Oceanography, 154(??):25–37, May 2017. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S007966111630088X.

Nomaki:2021:AFB

[NRA^{+}21] Hidetaka Nomaki, Eugenio Rastelli, Andreia Alves, Hisami Suga, Sandra Ramos, Tomo Kitahashi, Masashi Tsuchiya, Nanako O. Ogawa, Yohei Matsui, Koji Seike, Norio Miyamoto, Cinzia Corinaldesi, Elisabetta Manea, Naohiko Ohkouchi, Roberto Danovaro, Takuro Nunoura, and Teresa Amaro. Abyssal fauna, benthic microbes, and organic matter quality across a range of trophic conditions in the western Pacific Ocean. Progress in Oceanography, 195(??):??, July 2021. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661121000781.

Nachtsheim:2019:FBW

[NRS^{+}19] Dominik A. Nachtsheim, Svenja Ryan, Michael Schröder, Laura Jensen, W. Chris Oosthuizen, Marthán N. Bester, Wilhelm Hagen, and Horst Bornemann. Foraging behaviour of Weddell seals (Leptonychotes weddellii) in connection to oceanographic conditions in the southern Weddell Sea. Progress in Oceanography, 173(??):165–179, April 2019. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661118301216.

Nielsen:2024:SPB

[NSE^{+}24] Jens M. Nielsen, Michael F. Sigler, Lisa B. Eisner, Jordan T. Watson, Lauren A. Rogers, Shaun W. Bell, Noel Pelland, Calvin W. Mordy, Wei Cheng, Kirill Kivva, Sage Osborne, and Phyllis Stabeno. Spring phytoplankton bloom phenology during recent climate warming on the Bering Sea shelf. Progress in Oceanography, 220(??):??, January 2024. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661123002197.

Nomura:2023:BCB

[NST^{+}23] Daiki Nomura, Reishi Sahashi, Keigo D. Takahashi, Ryosuke Makabe, Masato Ito, Manami Tozawa, Pat Wongpan, Ryo Matsuda, Masayoshi Sano, Michiyo Yamamoto-Kawai, Natsumi Nojiro, Aiko Tachibana, Norio Kurosawa, Masato Moteki, Takeshi Tamura, Shigeru Aoki, and Hiroto Murase. Biogeochemical characteristics of brash sea ice and icebergs during summer and autumn in the Indian sector of the Southern Ocean. Progress in Oceanography, 214(??):??, June 2023. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661123000666.

Nakamura:2014:ETV

[NTU^{+}14] Tomohiro Nakamura, Yuuki Takeuchi, Keisuke Uchimoto, Humio Mitsudera, and Masaaki Wakatsuchi. Effects of temporal variation in tide-induced vertical mixing in the Kuril Straits on the thermohaline circulation originating in the Okhotsk Sea. Progress in Oceanography, 126(??):135–145, August 2014. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661114000937.

Nicholls:1987:PPF

[NW87] R. J. Nicholls and N. B. Webber. The past, present and future evolution of Hurst Castle Spit, Hampshire. Progress in Oceanography, 18(1–4):119–137, ???? 1987. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0079661187900292.

Nieto:2017:HIC

[NXT^{+}17] Karen Nieto, Yi Xu, Steven L. H. Teo, Sam McClatchie, and John Holmes. How important are coastal fronts to albacore tuna (Thunnus alalunga) habitat in the Northeast Pacific Ocean? Progress in Oceanography, 150(??):62–71, January 2017. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661115001019.

Nan:2015:KIS

[NXY15] Feng Nan, Huijie Xue, and Fei Yu. Kuroshio intrusion into the South China Sea: a review. Progress in Oceanography, 137 (part A)(??):314–333, September 2015. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661114000986.

Nishioka:2022:BPL

[NYH^{+}22] Jun Nishioka, Ichiro Yasuda, Toru Hirawake, Tomohiro Nakamura, Yoshiko Kondo, and Yuri N. Volkov. Biogeochemical and physical linkages between the Arctic Ocean and sub-Arctic Pacific through marginal seas. Progress in Oceanography, 203(??):??, April–May 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661122000295.

Nakajima:2017:PTS

[NYL^{+}17] Ryota Nakajima, Haruka Yamazaki, Levi S. Lewis, Adi Khen, Jennifer E. Smith, Nobuyuki Nakatomi, and Haruko Kurihara. Planktonic trophic structure in a coral reef ecosystem — grazing versus microbial food webs and the production of mesozooplankton. Progress in Oceanography, 156(??):104–120, August 2017. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661116302282.

OBrien:1981:E

[OA81] James J. O’Brien and Martin V. Angel. Editorial. Progress in Oceanography, 10(1):i, ???? 1981. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0079661181900069.

Omrani:2016:SSC

[OAB^{+}16] Hiba Omrani, Thomas Arsouze, Karine Béranger, Moncef Boukthir, Philippe Drobinski, Cindy Lebeaupin-Brossier, and Hanen Mairech. Sensitivity of the sea circulation to the atmospheric forcing in the Sicily Channel. Progress in Oceanography, 140(??):54–68, January 2016. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661115002165.

Okkonen:2020:KDV

[OACA20] Stephen Okkonen, Carin Ashjian, Robert G. Campbell, and Philip Alatalo. Krill diel vertical migration: a diagnostic for variability of wind forcing over the Beaufort and Chukchi Seas. Progress in Oceanography, 181(??):??, February 2020. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661120300033.

Ospina-Alvarez:2015:EPR

[OACB^{+}15] Andres Ospina-Alvarez, Ignacio A. Catalán, Miguel Bernal, David Roos, and Isabel Palomera. From egg production to recruits: Connectivity and inter-annual variability in the recruitment patterns of European anchovy in the northwestern Mediterranean. Progress in Oceanography, 138(??):431–447, November 2015. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661115000221.

Ounissi:2022:RWA

[OAD22] Makhlouf Ounissi, Aicha Beya Amira, and François Dulac. Riverine and wet atmospheric inputs of materials to a North Africa coastal site (Annaba Bay, Algeria). Progress in Oceanography, 165(??):19–34, ???? 2022. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661117301325.

Overland:2000:CNP

[OAM00] J. E. Overland, J. M. Adams, and H. O. Mofjeld. Chaos in the North Pacific: spatial modes and temporal irregularity. Progress in Oceanography, 47(2–4):337–354, October 2000. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661100000410.

Otero:2016:WDU

[OÁSG^{+}16] Jaime Otero, X. Antón Álvarez-Salgado, Ángel F. González, Carlos Souto, Miguel Gilcoto, and Ángel Guerra. Winddriven upwelling effects on cephalopod paralarvae: Octopus vulgaris and Loliginidae off the Galician coast (NE Atlantic). Progress in Oceanography, 141(??):130–143, February 2016. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661115002542.

Orfila:2005:CTM

[OÁT^{+}05] A. Orfila, A. Álvarez, J. Tintoré, A. Jordi, and G. Basterretxea. Climate teleconnections at monthly time scales in the Ligurian Sea inferred from satellite data. Progress in Oceanography, 66(2–4):157–170, August 2005. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661105000686.

Ospina-Alvarez:2018:LTU

[OAWAN18] Andrés Ospina-Alvarez, Nicolas Weidberg, Christopher M. Aiken, and Sergio A. Navarrete. Larval transport in the upwelling ecosystem of central Chile: the effects of vertical migration, developmental time and coastal topography on recruitment. Progress in Oceanography, 168(??):82–99, November 2018. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661117300800.

OBrien:1983:PEC

[O’B83] James J. O’Brien. The physical environment in coastal upwelling regions: Introduction. Progress in Oceanography, 12(3): 221, ???? 1983. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661183900083.

Otto:1998:RDO

[OB98] S. Otto and W. Balzer. Release of dissolved organic carbon (DOC) from sediments of the N.W. European Continental Margin (Goban Spur) and its significance for benthic carbon cycling. Progress in Oceanography, 42(1–4):127–144, December 1998. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661198000317.

OLoughlin:2020:IPA

[OBD^{+}20] Jessica H. O’Loughlin, Kim S. Bernard, Elizabeth A. Daly, Samantha Zeman, Jennifer L. Fisher, Richard D. Brodeur, and Thomas P. Hurst. Implications of Pyrosoma atlanticum range expansion on phytoplankton standing stocks in the Northern California Current. Progress in Oceanography, 188(??):??, October 2020. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661120301634.

Ou:2006:WIS

[OC06] Hsien-Wang Ou and Dake Chen. Wind-induced shear dispersion and genesis of the shelf-break front. Progress in Oceanography, 70(2–4):313–330, August 2006. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661106000498.

Olivar:2018:VDV

[OCH^{+}18] M. Pilar Olivar, Tabit Contreras, P. Alexander Hulley, Mikhail Emelianov, Cristina López-Pérez, Víctor Tuset, and Arturo Castellón. Variation in the diel vertical distributions of larvae and transforming stages of oceanic fishes across the tropical and equatorial Atlantic. Progress in Oceanography, 160(??):83–100, January 2018. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661117301908.

Olsson:1965:RCD

[OE65] Ingrid U. Olsson and K. Gösta Eriksson. Remarks on C^{14} dating of shell material in sea sediments. Progress in Oceanography, 3(??):253–266, ???? 1965. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0079661165900212.

Omstedt:2014:PPO

[OEL^{+}14] A. Omstedt, J. Elken, A. Lehmann, M. Leppäranta, H. E. M. Meier, K. Myrberg, and A. Rutgersson. Progress in physical oceanography of the Baltic Sea during the 2003–2014 period. Progress in Oceanography, 128(??):139–171, November 2014. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661114001335.

Omstedt:2004:KBS

[OELP04] Anders Omstedt, Jüri Elken, Andreas Lehmann, and Jan Piechura. Knowledge of the Baltic Sea physics gained during the BALTEX and related programmes. Progress in Oceanography, 63(1–2):1–28, October 2004. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661104000989.

Olson:1994:MPB

[OH94] Donald B. Olson and Raleigh R. Hood. Modelling pelagic biogeography. Progress in Oceanography, 34(2–3):161–205, ???? 1994. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661194900078.

Olivar:2017:MFA

[OHC^{+}17] M. Pilar Olivar, P. Alexander Hulley, Arturo Castellón, Mikhail Emelianov, Cristina López, Víctor M. Tuset, Tabit Contreras, and Balbina Molí. Mesopelagic fishes across the tropical and equatorial Atlantic: Biogeographical and vertical patterns. Progress in Oceanography, 151(??):116–137, February 2017. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661116301458.

Osborne:2022:OAG

[OHH^{+}22] Emily Osborne, Xinping Hu, Emily R. Hall, Kimberly Yates, Jennifer Vreeland-Dawson, Katie Shamberger, Leticia Barbero, J. Martin Hernandez-Ayon, Fabian A. Gomez, Tacey Hicks, Yuan-Yuan Xu, Melissa R. McCutcheon, Michael Acquafredda, Cecilia Chapa-Balcorta, Orion Norzagaray, Denis Pierrot, Alain Munoz-Caravaca, Kerri L. Dobson, Nancy Williams, Nancy Rabalais, and Padmanava Dash. Ocean acidification in the Gulf of Mexico: Drivers, impacts, and unknowns. Progress in Oceanography, 209(??):??, December 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661122001410.

Ozsoy:1989:CHL

[ÖHÜ89] E. Özsoy, A. Hecht, and Ü. Ünlüata. Circulation and hydrography of the Levantine Basin. Results of POEM coordinated experiments 1985–1986. Progress in Oceanography, 22(2):125– 170, ???? 1989. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661189900049.

Olmos:2023:ECD

[OIC^{+}23] Maxime Olmos, James Ianelli, Lorenzo Ciannelli, Ingrid Spies, Carey R. McGilliard, and James T. Thorson. Estimating climate-driven phenology shifts and survey availability using fishery-dependent data. Progress in Oceanography, 215(??): ??, July 2023. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661123000782.

Orsi:1999:CMP

[OJB99] A. H. Orsi, G. C. Johnson, and J. L. Bullister. Circulation, mixing, and production of Antarctic bottom water. Progress in Oceanography, 43(1):55–109, January 1999. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S007966119900004X.

Orselli:2019:SAC

[OKdA^{+}19] Iole B. M. Orselli, Rodrigo Kerr, José L. L. de Azevedo, Felippe Galdino, Moacyr Araujo, and Carlos A. E. Garcia. The sea-air CO_{2} net fluxes in the South Atlantic Ocean and the role played by Agulhas eddies. Progress in Oceanography, 170(??):40–52, January 2019. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661118300600.

Olausson:1965:CGP

[Ola65a] Eric Olausson. Climatological, geoeconomical and paleooceanographical aspects on carbonate deposition. Progress in Oceanography, 4(??):245–265, September 2, 1965. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0079661165900534.

Olausson:1965:ECC

[Ola65b] Eric Olausson. Evidence of climatic changes in North Atlantic deep-sea cores, with remarks on isotopic paleotemperature analysis. Progress in Oceanography, 3(??):221–252, ???? 1965. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0079661165900200.

Oliver:2018:MHE

[OLH^{+}18] Eric C. J. Oliver, Véronique Lago, Alistair J. Hobday, Neil J. Holbrook, Scott D. Ling, and Craig N. Mundy. Marine heatwaves off eastern Tasmania: Trends, interannual variability, and predictability. Progress in Oceanography, 161(??):116–130, February 2018. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661117303336.

Olli:2015:UUE

[Oll15] Kalle Olli. Unraveling the uncertainty and error propagation in the vertical flux Martin curve. Progress in Oceanography, 135(??):146–155, June 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661115001135.

Ohshimo:2022:IRP

[OMK^{+}22] Seiji Ohshimo, Daniel J. Madigan, Taketoshi Kodama, Hiroshige Tanaka, Kaoru Komoto, Satoshi Suyama, Tsuneo Ono, and Takashi Yamakawa. Isoscapes reveal patterns of δ ^{13}C and δ ^{15}N of pelagic forage fish and squid in the Northwest Pacific Ocean. Progress in Oceanography, 175(??):124–138, ???? 2022. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S007966111830260X.

Oguz:2022:CPP

[OMR^{+}22] Temel Oguz, Diego Macias, Lionel Renault, Javier Ruiz, and Joaquin Tintore. Controls of plankton production by pelagic fish predation and resource availability in the alboran and Balearic Seas. Progress in Oceanography, 112–113(??):1–14, ???? 2022. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661113000268.

Oliveira:2009:SDD

[OMS^{+}09] Paulo B. Oliveira, Teresa Moita, Alexandra Silva, Isabel T. Monteiro, and A. Sofia Palma. Summer diatom and dinoflagellate blooms in Lisbon Bay from 2002 to 2005: Preconditions inferred from wind and satellite data. Progress in Oceanography, 83(1–4):270–277, December 2009. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661109000937.

Ortega:2015:RTA

[OMS^{+}15] Pablo Ortega, Juliette Mignot, Didier Swingedouw, Florian Sévellec, and Eric Guilyardi. Reconciling two alternative mechanisms behind bi-decadal variability in the North Atlantic. Progress in Oceanography, 137 (part A)(??):237–249, September 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661115001482.

Ohlmann:2005:CCS

[ON05] J. Carter Ohlmann and P. Peter Niiler. Circulation over the continental shelf in the northern Gulf of Mexico. Progress in Oceanography, 64(1):45–81, January 2005. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661105000224.

Ona:2022:ADG

[ON22] Egil Ona and Julius Nielsen. Acoustic detection of the Greenland shark (Somniosus microcephalus) using multifrequency split beam echosounder in Svalbard waters. Progress in Oceanography, 206(??):??, August 2022. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661122001033.

Ohshima:2014:FDS

[ONR^{+}14] Kay I. Ohshima, Takuya Nakanowatari, Stephen Riser, Yuri Volkov, and Masaaki Wakatsuchi. Freshening and dense shelf water reduction in the Okhotsk Sea linked with sea ice decline. Progress in Oceanography, 126(??):71–79, August 2014. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661114000688.

Oozeki:2015:VTP

[OOTA15] Yoshioki Oozeki, Takeshi Okunishi, Akinori Takasuka, and Daisuke Ambe. Variability in transport processes of Pacific saury Cololabis saira larvae leading to their broad dispersal: Implications for their ecological role in the western North Pacific. Progress in Oceanography, 138(??):448–458, November 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661114000974.

Oliveira:2018:MED

[OP18] Fabricio S. C. Oliveira and Paulo S. Polito. Mesoscale eddy detection in satellite imagery of the oceans using the Radon transform. Progress in Oceanography, 167(??):150–163, October 2018. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661118300247.

Olson:2010:FWI

[OPG^{+}10] Robert J. Olson, Brian N. Popp, Brittany S. Graham, Gladis A. López-Ibarra, Felipe Galván-Magaña, Cleridy E. Lennert-Cody, Noemi Bocanegra-Castillo, Natalie J. Wallsgrove, Elizabeth Gier, Vanessa Alatorre-Ramírez, Lisa T. Ballance, and Brian Fry. Food-web inferences of stable isotope spatial patterns in copepods and yellowfin tuna in the pelagic eastern Pacific Ocean. Progress in Oceanography, 86(1–2):124– 138, July 2010. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661110000601.

Ortiz:2024:ESV

[OPH^{+}24] J. J. Ortiz, I. Preciado, M. Hidalgo, J. M. González-Irusta, I. M. Rabanal, and L. López-López. Estimating spatial variability of baseline isoscapes from fish isotopic signatures at the community level. Progress in Oceanography, 221(??):??, February 2024. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661124000119.

Ovall:2021:IWW

[OPL^{+}21] Bridget Ovall, Robert S. Pickart, Peigen Lin, Phyllis Stabeno, Thomas Weingartner, Motoyo Itoh, Takashi Kikuchi, Elizabeth Dobbins, and Shaun Bell. Ice, wind, and water: Synoptic-scale controls of circulation in the Chukchi Sea. Progress in Oceanography, 199(??):??, December 2021. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661121001907.

Ourgaud:2018:TEM

[ORB^{+}18] Mélanie Ourgaud, Sandrine Ruitton, Hugo Bourgogne, Paco Bustamante, Carine Churlaud, Gaël Guillou, Benoît Lebreton, and Mireille L. Harmelin-Vivien. Trace elements in a Mediterranean scorpaenid fish: Bioaccumulation processes and spatial variations. Progress in Oceanography, 163(??):184–195, April 2018. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661116301355.

Oke:2019:RCE

[ORCH^{+}19] Peter R. Oke, Moninya Roughan, Paulina Cetina-Heredia, Gabriela S. Pilo, Kenneth R. Ridgway, Tatiana Rykova, Matthew R. Archer, Richard C. Coleman, Colette G. Kerry, Carlos Rocha, Amandine Schaeffer, and Eduardo Vitarelli. Revisiting the circulation of the East Australian Current: Its path, separation, and eddy field. Progress in Oceanography, 176(??):??, September 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661118301745.

Oren:1969:OBI

[Ore69] O. H. Oren. Oceanographic and biological influence of the Suez Canal, the Nile and the Aswan Dam on the Levant Basin. Progress in Oceanography, 5(??):161–167, ???? 1969. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/007966116990038X.

Overland:2008:NPR

[ORMB08] James Overland, Sergei Rodionov, Shoshiro Minobe, and Nicholas Bond. North Pacific regime shifts: Definitions, issues and recent transitions. Progress in Oceanography, 77(2–3):92– 102, May 2008. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661108000475.

Ortega-Retuerta:2019:TEP

[ORMR^{+}19] Eva Ortega-Retuerta, Ignacio P. Mazuecos, Isabel Reche, Josep M. Gasol, Xosé A. Álvarez-Salgado, Marta Álvarez, María F. Montero, and Javier Arístegui. Transparent exopolymer particle (TEP) distribution and in situ prokaryotic generation across the deep Mediterranean Sea and nearby North East Atlantic Ocean. Progress in Oceanography, 173(??):180– 191, April 2019. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661118303239.

Ojeda-Ruiz:2022:RVA

[ORPRGIS22] Miguel Ángel Ojeda-Ruiz, David Petatán-Ramírez, Triana Guerrero-Izquierdo, and Christian Salvadeo. Rapid vulnerability assessment of Pacific sardine (Sardinops sagax) fisheries facing climate change in Mexico. Progress in Oceanography, 206(??):??, August 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661122000878.

Oliveira:2002:CMS

[ORR^{+}02] A. Oliveira, F. Rocha, A. Rodrigues, J. Jouanneau, A. Dias, O. Weber, and C. Gomes. Clay minerals from the sedimentary cover from the Northwest Iberian shelf. Progress in Oceanography, 52(??):233–247, ???? 2002. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661102000083.

Oliveros-Ramos:2017:SAC

[ORVES17] Ricardo Oliveros-Ramos, Philippe Verley, Vincent Echevin, and Yunne-Jai Shin. A sequential approach to calibrate ecosystem models with multiple time series data. Progress in Oceanography, 151(??):227–244, February 2017. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661116301926.

Olli:2001:VFB

[ORW^{+}01] Kalle Olli, Christian Wexels Riser, Paul Wassmann, Tatjana Ratkova, Elena Arashkevich, and Anna Pasternak. Vertical flux of biogenic matter during a Lagrangian study off the NW Spanish continental margin. Progress in Oceanography, 51(2–4):443–466, November 2001. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661101000799.

Olu:1996:SDD

[OSH^{+}96] K. Olu, M. Sibuet, F. Harmegnies, J.-P. Foucher, and A. Fiala-Médioni. Spatial distribution of diverse cold seep communities living on various diapiric structures of the southern Barbados prism. Progress in Oceanography, 38(4): 347–356, IN1, 357–360, IN3, 361–376, ???? 1996. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661197000062.

Orlov:2019:CDS

[OT19] Alexei M. Orlov and Alexei M. Tokranov. Checklist of deepsea fishes of the Russian northwestern Pacific Ocean found at depths below 1000m. Progress in Oceanography, 176(??):??, September 2019. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661119300941.

Okazaki:2020:PSB

[OTNI20] Yuji Okazaki, Kazutaka Takahashi, Yuichiro Nishibe, and Mutsuo Ichinomiya. Post-spring-bloom feeding rates of krill Euphausia pacifica (Euphausiacea) in the Oyashio region, western North Pacific. Progress in Oceanography, 181(??):??, February 2020. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661119304252.

Ozsoy:1993:EMW

[ÖÜT93] Emin Özsoy, Ümit Ünlüata, and Zafer Top. The evolution of Mediterranean water in the Black Sea: interior mixing and material transport by double diffusive intrusions. Progress in Oceanography, 31(3):275–320, ???? 1993. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/007966119390004W.

Orlov:2024:OBG

[OV24] Alexei M. Orlov and Igor V. Volvenko. Ongoing borealization of Siberian Arctic marine ichthyofauna: Further evidence. Progress in Oceanography, 225(??):??, July 2024. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661124000946.

Omori:1994:IHA

[OvdSN94] M. Omori, S. van der Spoel, and C. P. Norman. Impact of human activities on pelagic biogeography. Progress in Oceanography, 34(2–3):211–219, ???? 1994. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0079661194900094.

OBrien:2016:GCD

[OVG16] Colleen J. O’Brien, Meike Vogt, and Nicolas Gruber. Global coccolithophore diversity: Drivers and future change. Progress in Oceanography, 140(??):27–42, January 2016. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661115002128.

Oliveira:2002:NLD

[OVR^{+}02] A. Oliveira, J. Vitorino, A. Rodrigues, J. M. Jouanneau, J. A. Dias, and O. Weber. Nepheloid layer dynamics in the northern Portuguese shelf. Progress in Oceanography, 52(??):195– 213, ???? 2002. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S007966110200006X.

Owens:1991:SDM

[Owe91] W. Brechner Owens. A statistical description of the mean circulation and eddy variability in the northwestern Atlantic using SOFAR floats. Progress in Oceanography, 28(3):257– 303, ???? 1991. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/007966119190010J.

Oliver:2014:EEG

[OWH14] Eric C. J. Oliver, Simon J. Wotherspoon, and Neil J. Holbrook. Estimating extremes from global ocean and climate models: a Bayesian hierarchical model approach. Progress in Oceanography, 122(??):77–91, March 2014. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661113002425.

Olli:2007:FPC

[OWR^{+}07] Kalle Olli, Paul Wassmann, Marit Reigstad, Tatjana N. Ratkova, Elena Arashkevich, Anna Pasternak, Patricia A. Matrai, Johan Knulst, Lars Tranvik, Riina Klais, and A. Jacobsen. The fate of production in the central Arctic Ocean — top–down regulation by zooplankton expatriates? Progress in Oceanography, 72(1):84–113, January 2007. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661106001182.

Ohashi:2023:CDA

[OYKK^{+}23] Yoshihiko Ohashi, Michiyo Yamamoto-Kawai, Kazuya Kusahara, Ken’ichi Sasaki, and Kay I. Ohshima. Circumpolar distributions of age and anthropogenic CO_{2} content of Antarctic Bottom Water revealed by chlorofluorocarbon and sulfur hexafluoride. Progress in Oceanography, 219(??):??, December 2023. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661123001969.

Paffenhofer:1987:I

[PAB^{+}87a] Gustav-Adolf Paffenhöfer, Larry P. Atkinson, Jackson O. Blanton, Thomas N. Lee, Lawrence R. Pomeroy, and James A. Yoder. Introduction. Progress in Oceanography, 19(3–4):221– 230, ???? 1987. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661187900097.

Paffenhofer:1987:SC

[PAB^{+}87b] Gustav-Adolf Paffenhöfer, Larry P. Atkinson, Jackson O. Blanton, Thomas N. Lee, Lawrence R. Pomeroy, and James A. Yoder. Summary and conclusions. Progress in Oceanography, 19(3–4):437–441, ???? 1987. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0079661187900176.

Peck:2021:SPF

[PAB^{+}21] Myron A. Peck, Jürgen Alheit, Arnaud Bertrand, Ignacio A. Catalán, Susana Garrido, Marta Moyano, Ryan R. Rykaczewski, Akinori Takasuka, and Carl D. van der Lingen. Small pelagic fish in the new millennium: a bottom-up view of global research effort. Progress in Oceanography, 191(??):??, February 2021. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661120302299.

Povinec:2011:CWP

[PAF^{+}11] P. P. Povinec, M. Aoyama, M. Fukasawa, K. Hirose, K. Komura, J.-A. Sanchez-Cabeza, J. Gastaud, M. Ješkovský, I. Levy, and I. Sýkora. ^{137}Cs water profiles in the South Indian Ocean — an evidence for accumulation of pollutants in the subtropical gyre. Progress in Oceanography, 89(1–4):17– 30, April 2011. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661110001564.

Paluselli:2018:OPA

[PAG^{+}18] Andrea Paluselli, Yann Aminot, François Galgani, Sopheak Net, and Richard Sempéré. Occurrence of phthalate acid esters (PAEs) in the northwestern Mediterranean Sea and the Rhone River. Progress in Oceanography, 163(??):221–231, April 2018. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661117300496.

Painter:2020:BOE

[Pai20] Stuart C. Painter. The biogeochemistry and oceanography of the East African Coastal Current. Progress in Oceanography, 186(??):??, July 2020. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661120301130.

Peterson:1988:NZD

[PAM^{+}88] William T. Peterson, Dagoberto F. Arcos, George B. Mc-Manus, Hans Dam, Diane Bellantoni, Thomas Johnson, and Peter Tiselius. The nearshore zone during coastal upwelling: Daily variability and coupling between primary and secondary production off central Chile. Progress in Oceanography, 20(1): 1–40, ???? 1988. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661188900523.

Pedros-Alio:2015:DPM

[PAPL15] Carlos Pedrós-Alió, Marianne Potvin, and Connie Lovejoy. Diversity of planktonic microorganisms in the Arctic Ocean. Progress in Oceanography, 139(??):233–243, December 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661115001615.

Parsons:1963:SOM

[Par63] T. R. Parsons. Suspended organic matter in sea water. Progress in Oceanography, 1(??):205–239, ???? 1963. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0079661163900053.

Parker:1965:IDP

[Par65] Frances L. Parker. Irregular distributions of planktonic Foraminifera and stratigraphic correlation. Progress in Oceanography, 3(??):267–272, ???? 1965. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://

www.sciencedirect.com/science/article/pii/0079661165900224.

Park:1986:WCM

[Par86] Young-Hyang Park. Water characteristics and movements of the Yellow Sea Warm Current in summer. Progress in Oceanography, 17(3–4):243–254, ???? 1986. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0079661186900479.

Passow:2022:TEP

[Pas22] U. Passow. Transparent exopolymer particles (TEP) in aquatic environments. Progress in Oceanography, 55(3–4):287–333, ???? 2022. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661102001386.

Pitcher:2021:SCC

[PAVB^{+}21] Grant C. Pitcher, Arturo Aguirre-Velarde, Denise Breitburg, Jorge Cardich, Jacob Carstensen, Daniel J. Conley, Boris Dewitte, Anja Engel, Dante Espinoza-Morriberón, Georgina Flores, Véronique Garçon, Michelle Graco, Marilaure Grégoire, Dimitri Gutiérrez, José Martin Hernandez-Ayon, Huai-Hsuan May Huang, Kirsten Isensee, María Elena Jacinto, Lisa Levin, Alberto Lorenzo, Eric Machu, Lander Merma, Ivonne Montes, Naqvi S. W. A., Aurelien Paulmier, Michael Roman, Kenneth Rose, Raleigh Hood, Nancy N. Rabalais, Anne Gro V. Salvanes, Renato Salvatteci, Sonia Sánchez, Abdelfettah Sifeddine, Abdoul Wahab Tall, Anja K. van der Plas, Moriaki Yasuhara, Jing Zhang, and Z. Y. Zhu. System controls of coastal and open ocean oxygen depletion. Progress in Oceanography, 197(??):??, September– October 2021. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661121001002.

Price:1994:ODW

[PB94] James F. Price and Molly O’Neil Baringer. Outflows and deep water production by marginal seas. Progress in Oceanography, 33(3):161–200, ???? 1994. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0079661194900272.

Pena:2007:TSN

[PB07] M. Angelica Peña and Steven J. Bograd. Time series of the northeast Pacific. Progress in Oceanography, 75(2):115–119, October 2007. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661107001437.

Peck:2012:ESS

[PBB^{+}12a] Myron A. Peck, Hannes Baumann, Matthias Bernreuther, Catriona Clemmesen, Jens-Peter Herrmann, Holger Haslob, Bastian Huwer, Philipp Kanstinger, Fritz W. Köster, Christoph Petereit, Axel Temming, and Rudi Voss. The ecophysiology of Sprattus sprattus in the Baltic and North Seas. Progress in Oceanography, 103(??):42–57, September 2012. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661112000468.

Peck:2012:RES

[PBB^{+}12b] Myron A. Peck, Hannes Baumann, Matthias Bernreuther, Catriona Clemmesen, Jens-Peter Herrmann, Holger Haslob, Bastian Huwer, Philipp Kanstinger, Fritz W. Köster, Christoph Petereit, Axel Temming, and Rudi Voss. Reprint of: the ecophysiology of Sprattus sprattus in the Baltic and North Seas. Progress in Oceanography, 107(??):31–46, December 2012. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661112001334.

Pages:2020:CRI

[PBB^{+}20] R. Pagès, M. Baklouti, N. Barrier, C. Richon, J.-C. Dutay, and T. Moutin. Changes in rivers inputs during the last decades significantly impacted the biogeochemistry of the eastern Mediterranean basin: a modelling study.

Progress in Oceanography, 181(??):??, February 2020. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661119304227.

Portner:2022:DRP

[PBBH^{+}22] Elan J. Portner, Kelly J. Benoit-Bird, Elliott L. Hazen, Chad M. Waluk, Carlos J. Robinson, Jaime Gómez-Gutiérrez, and William F. Gilly. Decline and recovery of pelagic acoustic backscatter following El Niño events in the Gulf of California, Mexico. Progress in Oceanography, 206(??):??, August 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661122000842.

Pomar:1988:DAP

[PBD^{+}88] L. Acosta Pomar, V. Bruni, F. Decembrini, G. Giuffré, and T. L. Maugeri. Distribution and activity of picophytoplankton in a brackish environment. Progress in Oceanography, 21(2):129–138, ???? 1988. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0079661188900316.

Perry:2010:IGO

[PBH^{+}10] R. Ian Perry, Manuel Barange, Eileen Hofmann, Coleen Moloney, Geir Ottersen, and Yasunori Sakurai. Introduction to the GLOBEC 3rd Open Science Meeting: From ecosystem function to ecosystem prediction. Progress in Oceanography, 87(1–4):1–5, October 2010. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661110001308.

Preikshot:2013:DMD

[PBN13] Dave Preikshot, Richard J. Beamish, and Chrys M. Neville. A dynamic model describing ecosystem-level changes in the Strait of Georgia from 1960 to 2010. Progress in Oceanography, 115(??):28–40, August 2013. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661113000621.

Perry:2010:GCM

[PBO10] R. Ian Perry, Manuel Barange, and Rosemary E. Ommer. Global changes in marine systems: a social–ecological approach. Progress in Oceanography, 87(1–4):331–337, October 2010. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661110001217.

Perry:2010:CGC

[PBOW10] R. Ian Perry, Manuel Barange, Rosemary E. Ommer, and Francisco E. Werner. Conclusions to the GLOBEC compendium. Progress in Oceanography, 87(S1):S6–S8, ???? 2010. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661110001485.

Price:1999:HDB

[PBP^{+}99] N. B. Price, T. Brand, J. M. Pates, S. Mowbray, A. Theocharis, G. Civitarese, S. Miserocchi, S. Heussner, and F. Lindsay. Horizontal distributions of biogenic and lithogenic elements of suspended particulate matter in the Mediterranean Sea. Progress in Oceanography, 44(1–3):191–218, August 1999. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661199000257.

Phillips:2022:MCS

[PBS22] A. Jason Phillips, Richard D. Brodeur, and Andrey V. Suntsov. Micronekton community structure in the epipelagic zone of the northern California Current upwelling system. Progress in Oceanography, 80(1–2):74–92, ???? 2022. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661108002097.

Pattiaratchi:1987:MLS

[PC87] Charitha Pattiaratchi and Michael Collins. Mechanisms for linear sandbank formation and maintenance in relation to dynamical oceanographic observations. Progress in Oceanography, 19(2):117–176, ???? 1987. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0079661187900061.

Purwandana:2020:SST

[PCBA^{+}20] Adi Purwandana, Yannis Cuypers, Pascale Bouruet-Aubertot, Taira Nagai, Toshiyuki Hibiya, and Agus S. Atmadipoera. Spatial structure of turbulent mixing inferred from historical CTD datasets in the Indonesian seas. Progress in Oceanography, 184 (??):??, May 2020. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661120300513.

Pierdomenico:2019:MDA

[PCC^{+}19] M. Pierdomenico, F. Cardone, A. Carluccio, D. Casalbore, F. Chiocci, P. Maiorano, and G. D’Onghia. Megafauna distribution along active submarine canyons of the central Mediterranean: Relationships with environmental variables. Progress in Oceanography, 171(??):49–69, February 2019. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661118301125.

Papale:2018:HMR

[PCD^{+}18] Maria Papale, Antonella Conte, Marianna Del Core, Elisa Zito, Mario Sprovieri, Filomena De Leo, Carmen Rizzo, Clara Urzì, Emilio De Domenico, Gian Marco Luna, Luigi Michaud, and Angelina Lo Giudice. Heavy-metal resistant microorganisms in sediments from submarine canyons and the adjacent continental slope in the northeastern Ligurian margin (Western Mediterranean Sea). Progress in Oceanography, 168(??):155–168, November 2018. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661118300922.

Pinchuk:2008:CRV

[PCH08a] Alexei I. Pinchuk, Kenneth O. Coyle, and Russell R. Hopcroft. Climate-related variability in abundance and reproduction of euphausiids in the northern Gulf of Alaska in 1998–2003. Progress in Oceanography, 77(2–3):203–216, May 2008. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661108000554.

Polovina:2008:EDP

[PCH^{+}08b] Jeffrey J. Polovina, Fei Chai, Evan A. Howell, Donald R. Kobayashi, Lei Shi, and Yi Chao. Ecosystem dynamics at a productivity gradient: a study of the lower trophic dynamics around the northern atolls in the Hawaiian Archipelago. Progress in Oceanography, 77(2–3):217–224, May 2008. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661108000566.

Pelegri:2006:GCU

[PCK^{+}06] J. L. Pelegrí, J. H. Churchill, A. D. Kirwan, S.-K. Lee, R. E. Munn, and N. R. Pettigrew. Gabriel T. Csanady: Understanding the physics of the ocean. Progress in Oceanography, 70(2–4):91–112, August 2006. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661106000826.

Pepin:2011:SPZ

[PCM11] Pierre Pepin, Eugene Colbourne, and Gary Maillet. Seasonal patterns in zooplankton community structure on the Newfoundland and Labrador Shelf. Progress in Oceanography, 91(3):273–285, November 2011. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661111000048.

Puerta:2022:DGA

[PCR^{+}22] Patricia Puerta, Lorenzo Ciannelli, Ryan R. Rykaczewski, Michael Opiekun, and Michael A. Litzow. Do Gulf of Alaska fish and crustacean populations show synchronous nonstationary responses to climate? Progress in Oceanography, 175(??):161–170, ???? 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661118300028.

Parada:2012:ESV

[PCSMC12] Carolina Parada, Francois Colas, Samuel Soto-Mendoza, and Leonardo Castro. Effects of seasonal variability in acrossand alongshore transport of anchoveta (Engraulis ringens) larvae on model-based pre-recruitment indices off central Chile. Progress in Oceanography, 92–95(??):192–205, January 2012. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661111000693.

Peterson:2015:EPR

[PD15] William T. Peterson and Xiuning Du. Egg production rates of the copepod Calanus marshallae in relation to seasonal and interannual variations in microplankton biomass and species composition in the coastal upwelling zone off Oregon, USA. Progress in Oceanography, 138 (part A)(??):32–44, November 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661115001974.

Petrik:2015:BTM

[PDAM^{+}15] Colleen M. Petrik, Janet T. Duffy-Anderson, Franz Mueter, Katherine Hedstrom, and Enrique N. Curchitser. Biophysical transport model suggests climate variability determines distribution of walleye pollock early life stages in the eastern Bering Sea through effects on spawning. Progress in Oceanography, 138(??):459–474, November 2015. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661114001128.

Poulton:2022:SPC

[PDD^{+}22] Alex J. Poulton, Clare E. Davis, Chris J. Daniels, Kyle M. J. Mayers, Carolyn Harris, Glen A. Tarran, Claire E. Widdicombe, and E. Malcolm S. Woodward. Seasonal phosphorus and carbon dynamics in a temperate shelf sea (Celtic Sea). Progress in Oceanography, 177(??):??, October 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661117301581.

Puig:2013:TBN

[PdMS^{+}13] Pere Puig, Xavier Durrieu de Madron, Jordi Salat, Katrin Schroeder, Jacobo Martín, Aristomenis P. Karageorgis, Albert Palanques, François Roullier, José Luis Lopez-Jurado, Mikhail Emelianov, Thierry Moutin, and Loïc Houpert. Thick bottom nepheloid layers in the western Mediterranean generated by deep dense shelf water cascading. Progress in Oceanography, 111(??):1–23, April 2013. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661112001188.

Peck:2012:LLB

[PDV12] Myron A. Peck, Jörg Dutz, and Rudi Voss. A lasting legacy for the Baltic and North Sea GLOBEC Germany program. Progress in Oceanography, 107(??):1–2, December 2012. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661112000900.

Pearcy:2002:MNO

[Pea02] W. G. Pearcy. Marine nekton off Oregon and the 1997–98 El nino. Progress in Oceanography, 54(1–4):399–403, July 2002. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661102000605.

Pares-Escobar:2018:ISV

[PELAA18] Fernanda Parés-Escobar, Bertha E. Lavaniegos, and Israel Ambriz-Arreola. Interannual summer variability in oceanic euphausiid communities off the Baja California western coast during 1998–2008. Progress in Oceanography, 160(??):53–67, January 2018. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661117300253.

Pena:2003:MRP

[Peñ03a] M. Angélica Peña. Modelling the response of the planktonic food web to iron fertilization and warming in the NE subarctic Pacific. Progress in Oceanography, 57(3–4):453–479, June 2003. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661103001101.

Pena:2003:PSC

[Peñ03b] M. Angelica Peña. Plankton size classes, functional groups and ecosystem dynamics: an introduction. Progress in Oceanography, 57(3–4):239–242, June 2003. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661103000983.

Pena:2024:AVM

[Peñ24] Marian Peña. Atlantic versus Mediterranean deep scattering layers around the Iberian peninsula. Progress in Oceanography, 221(??):??, February 2024. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S007966112400017X.

Pous:2010:ITE

[PFE10] S. Pous, E. Feunteun, and C. Ellien. Investigation of tropical eel spawning area in the south-Western Indian Ocean: Influence of the oceanic circulation. Progress in Oceanography, 86(3–4):396–413, September 2010. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661110000947.

Pitcher:2010:POU

[PFHM10] G. C. Pitcher, F. G. Figueiras, B. M. Hickey, and M. T. Moita. The physical oceanography of upwelling systems and the development of harmful algal blooms. Progress in Oceanography, 85(1–2):5–32, April 2010. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661110000133.

Painter:2016:SPS

[PFHM16] Stuart C. Painter, Madelaine Finlay, Victoria S. Hemsley, and Adrian P. Martin. Seasonality, phytoplankton succession and the biogeochemical impacts of an autumn storm in the northeast Atlantic Ocean. Progress in Oceanography, 142(??):72– 104, March 2016. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661116000173.

Peralta-Ferriz:2015:SIV

[PFW15] Cecilia Peralta-Ferriz and Rebecca A. Woodgate. Seasonal and interannual variability of pan-Arctic surface mixed layer properties from 1979 to 2012 from hydrographic data, and the dominance of stratification for multiyear mixed layer depth shoaling. Progress in Oceanography, 134(??):19–53, May 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661114002158.

Pincinato:2010:PPM

[PG10] R. B. M. Pincinato and M. A. Gasalla. Priceless prices and marine food webs: Long-term patterns of change and fishing impacts in the South Brazil Bight as reflected by the seafood market. Progress in Oceanography, 87(1–4):320–330, October 2010. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661110001242.

Packard:2013:MVC

[PG13] Theodore T. Packard and May Gómez. Modeling vertical carbon flux from zooplankton respiration. Progress in Oceanography, 110(??):59–68, March 2013. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661113000049.

Pollard:1996:VSF

[PGC^{+}96] R. T. Pollard, M. J. Grifftths, S. A. Cunningham, J. F. Read, F. F. Pérez, and A. F. Ríos. Vivaldi 1991 — a study of the formation, circulation and ventilation of Eastern North Atlantic central water. Progress in Oceanography, 37(2):167– 172, IN1–IN3, 173–176, IN5, 177–180, IN7, 181–192, ???? 1996. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661196000080.

Putman:2022:STP

[PGG^{+}22] Nathan F. Putman, Gustavo J. Goni, Lewis J. Gramer, Chuanmin Hu, Elizabeth M. Johns, Joaquin Trinanes, and Mengqiu Wang. Simulating transport pathways of pelagic Sargassum from the equatorial Atlantic into the Caribbean Sea. Progress in Oceanography, 165(??):205–214, ???? 2022. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661118300259.

Pierdomenico:2017:MAH

[PGGG17] Martina Pierdomenico, Andrea Gori, Vincent G. Guida, and Josep-Maria Gili. Megabenthic assemblages at the Hudson Canyon head (NW Atlantic margin): Habitat-faunal relationships. Progress in Oceanography, 157(??):12–26, September 2017. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661116302671.

Palanques:2005:GPC

[PGLG^{+}05] Albert Palanques, Emilio García-Ladona, Damià Gomis, Jacobo Martín, Marta Marcos, Ananda Pascual, Pere Puig, Josep-Maria Gili, Mikhail Emelianov, Sebastià Monserrat, Jorge Guillén, Joaquín Tintoré, Mariona Segura, Antoni Jordi, Simón Ruiz, Gotzon Basterretxea, Jordi Font, Dolors Blasco, and Francesc Pagès. General patterns of circulation, sediment fluxes and ecology of the Palamós (La Fonera) submarine canyon, northwestern Mediterranean. Progress in Oceanography, 66(2–4):89–119, August 2005. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661105000650.

Perrot:2018:CDT

[PGRP^{+}18] Laurie Perrot, Francis Gohin, Diana Ruiz-Pino, Luis Lampert, Martin Huret, Aurélie Dessier, Pascale Malestroit, Christine Dupuy, and Paul Bourriau. Coccolith-derived turbidity and hydrological conditions in May in the Bay of Biscay. Progress in Oceanography, 166(??):41–53, September 2018. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661117302276.

Phillips:2022:IBM

[PGS^{+}22] Joe Scutt Phillips, Alex Sen Gupta, Inna Senina, Erik van Sebille, Michael Lange, Patrick Lehodey, John Hampton, and Simon Nicol. An individual-based model of skipjack tuna (Katsuwonus pelamis) movement in the tropical Pacific Ocean. Progress in Oceanography, 164(??):63–74, ???? 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661117302896.

Pasqual:2013:CPT

[PGT^{+}13] Catalina Pasqual, Miguel A. Goñi, Tommaso Tesi, Anna Sanchez-Vidal, Antoni Calafat, and Miquel Canals. Composition and provenance of terrigenous organic matter transported along submarine canyons in the Gulf of Lion (NW Mediterranean Sea). Progress in Oceanography, 118(??):81–94, November 2013. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661113001262.

Pata:2022:PZB

[PGY^{+}22] Patrick R. Pata, Moira Galbraith, Kelly Young, Andrew R. Margolin, R. Ian Perry, and Brian P. V. Hunt. Persistent zooplankton bioregions reflect long-term consistency of community composition and oceanographic drivers in the NE Pacific. Progress in Oceanography, 206(??):??, August 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661122001082.

Pinones:2019:HVA

[PHC^{+}19] Andrea Piñones, Eileen E. Hofmann, Daniel P. Costa, Kimberly Goetz, Jennifer M. Burns, Fabien Roquet, Michael S. Dinniman, and John M. Klinck. Hydrographic variability along the inner and mid-shelf region of the western Ross Sea obtained using instrumented seals. Progress in Oceanography, 174(??):131–142, May 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S007966111830003X.

Poulton:2017:CET

[PHCA17] Alex J. Poulton, Patrick M. Holligan, Anastasia Charalampopoulou, and Tim R. Adey. Coccolithophore ecology in the tropical and subtropical Atlantic Ocean: New perspectives from the Atlantic Meridional Transect (AMT) Programme. Progress in Oceanography, 158(??):150–170, November 2017. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661116300234.

Petitgas:2018:ESS

[PHD^{+}18] Pierre Petitgas, Martin Huret, Christine Dupuy, Jérôme Spitz, Matthieu Authier, Jean Baptiste Romagnan, and Mathieu Doray. Ecosystem spatial structure revealed by integrated survey data. Progress in Oceanography, 166(??):189–198, September 2018. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661117300885.

Petereit:2014:FAB

[PHFK14] C. Petereit, H.-H. Hinrichsen, A. Franke, and F. W. Köster. Floating along buoyancy levels: Dispersal and survival of western Baltic fish eggs. Progress in Oceanography, 122(??):131– 152, March 2014. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661114000044.

Painter:2017:ESC

[PHK^{+}17] Stuart C. Painter, Susan E. Hartman, Caroline Kivimäe, Lesley A. Salt, Nicola M. Clargo, Chris J. Daniels, Yann Bozec, Lucie Daniels, Stephanie Allen, Victoria S. Hemsley, Grigorios Moschonas, and Keith Davidson. The elemental stoichiometry (C, Si, N, P) of the Hebrides Shelf and its role in carbon export. Progress in Oceanography, 159(??):154–177, December 2017. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661116302610.

Polovina:2001:TZC

[PHKS01] Jeffrey J. Polovina, Evan Howell, Donald R. Kobayashi, and Michael P. Seki. The transition zone chlorophyll front, a dynamic global feature defining migration and forage habitat for marine resources. Progress in Oceanography, 49(??):469– 483, ???? 2001. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661101000362.

Polovina:2017:TZC

[PHKS17] Jeffrey J. Polovina, Evan A. Howell, Donald R. Kobayashi, and Michael P. Seki. The transition zone chlorophyll front updated: Advances from a decade of research. Progress in Oceanography, 150(??):79–85, January 2017. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661115000087.

Phleger:1965:DPB

[Phl65] Fred B. Phleger. Depth patterns of benthonic Foraminifera in the Eastern Pacific. Progress in Oceanography, 3(??):273– 287, ???? 1965. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661165900236.

Pierson:2005:RSC

[PHLL05] James J. Pierson, Claudia Halsband-Lenk, and Andrew W. Leising. Reproductive success of Calanus pacificus during diatom blooms in Dabob Bay, Washington. Progress in Oceanography, 67(3–4):314–331, November 2005. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S007966110500145X.

Pierce:2001:DCO

[Pie01] David W. Pierce. Distinguishing coupled ocean–atmosphere interactions from background noise in the North Pacific. Progress in Oceanography, 49(??):331–352, ???? 2001. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661101000295.

Pirazzoli:1987:RSL

[Pir87] Paolo Antonio Pirazzoli. Recent sea-level changes and related engineering problems in the Lagoon of Venice (Italy). Progress in Oceanography, 18(1–4):323–346, ???? 1987. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0079661187900383.

Palmer:2013:POJ

[PIS13] Matthew R. Palmer, Mark E. Inall, and Jonathan Sharples. The physical oceanography of Jones Bank: a mixing hotspot in the Celtic Sea. Progress in Oceanography, 117(??):9–24, October 2013. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661113000979.

Pepin:2015:MEE

[PJH^{+}15] Pierre Pepin, Catherine L. Johnson, Michel Harvey, Benoit Casault, Joël Chassé, Eugene B. Colbourne, Peter S. Galbraith, Dave Hebert, Gordana Lazin, Gary Maillet, Stéphane Plourde, and Michel Starr. A multivariate evaluation of environmental effects on zooplankton community structure in the western North Atlantic. Progress in Oceanography, 134(??):197– 220, May 2015. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661115000282.

Parameswaran:2022:DDE

[PJS^{+}22] Usha V. Parameswaran, K. U. Abdul Jaleel, V. N. Sanjeevan, Aiswarya Gopal, Anil Kumar Vijayan, G. V. M. Gupta, and M. Sudhakar. Diversity and distribution of echinoderms in the South Eastern Arabian Sea shelf under the influence of seasonal hypoxia. Progress in Oceanography, 165(??):189– 204, ???? 2022. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661117302963.

Peterson:2002:ELC

[PK02] William T. Peterson and Julie E. Keister. The effect of a large cape on distribution patterns of coastal and oceanic copepods off Oregon and northern California during the 1998–1999 El Niño–La Niña. Progress in Oceanography, 53(??):389–411, ???? 2002. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661102000381.

Petrunina:2019:AFT

[PKA19] Alexandra S. Petrunina, Gregory A. Kolbasov, and Pedro Martinez Arbizu. Anatomy of the free tantulus larva (Crustacea: Tantulocarida) studied with confocal laser scanning microscopy: an extreme case of miniatuarisation in the Arthropoda. Progress in Oceanography, 178(??):??, November 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661119301910.

Peterson:2002:ENN

[PKF02] W. T. Peterson, J. E. Keister, and L. R. Feinberg. The effects of the 1997–99 El Niño/La Niña events on hydrography and zooplankton off the central Oregon coast. Progress in Oceanography, 54(1–4):381–398, July 2002. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661102000599.

Park:2014:EMS

[PKP14] Jong-Yeon Park, Jong-Seong Kug, and Young-Gyu Park. An exploratory modeling study on bio-physical processes associated with ENSO. Progress in Oceanography, 124(??):28– 41, May 2014. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661114000391.

Powley:2018:PNT

[PKV18] Helen R. Powley, Michael D. Krom, and Philippe Van Cappellen. Phosphorus and nitrogen trajectories in the Mediterranean Sea (1950–2030): Diagnosing basin-wide anthropogenic nutrient enrichment. Progress in Oceanography, 162(??):257– 270, March 2018. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661117302987.

Paffenhofer:1987:DAP

[PL87] Gustav-Adolf Paffenhöfer and Thomas N. Lee. Distribution and abundance of particulate matter. Progress in Oceanography, 19(3–4):373–401, ???? 1987. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0079661187900152.

Pingree:1989:CAS

[PL89] R. D. Pingree and B. Le Cann. Celtic and Armorican slope and shelf residual currents. Progress in Oceanography, 23(4):303– 338, ???? 1989. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661189900037.

Parsons:2001:CVM

[PL01] L. S. Parsons and W. H. Lear. Climate variability and marine ecosystem impacts: a North Atlantic perspective. Progress in Oceanography, 49(??):167–188, ???? 2001. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661101000210.

Pranovi:2009:EET

[PL09] Fabio Pranovi and Jason S. Link. Ecosystem exploitation and trophodynamic indicators: a comparison between the Northern Adriatic Sea and Southern New England. Progress in Oceanography, 81(1–4):149–164, April 2009. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661109000275.

Park:2018:NPO

[PL18] JongJin Park and Byunghwan Lim. A new perspective on origin of the East Sea intermediate water: Observations of Argo floats. Progress in Oceanography, 160(??):213–224, January 2018. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661116300052.

Pickart:2023:PWF

[PLB^{+}23] Robert S. Pickart, Peigen Lin, Frank Bahr, Leah T. McRaven, Jie Huang, Astrid Pacini, Kevin R. Arrigo, Carin J. Ashjian, Catherine Berchok, Mark F. Baumgartner, Kyoung-Ho Cho, Lee W. Cooper, Seth L. Danielson, Douglas Dasher, Amane Fuiwara, Jeanette Gann, Jacqueline M. Grebmeier, Jianfeng He, Toru Hirawake, Motoyo Itoh, Lauren Juranek, Takashi Kikuchi, G. W. K. Moore, Jeff Napp, R. John Nelson, Shigeto Nishino, Hank Statscewich, Phyllis Stabeno, Kathleen M. Stafford, Hiromichi Ueno, Svein Vagle, Thomas J. Weingartner, Bill Williams, and Sarah Zimmermann. The Pacific water flow branches in the eastern Chukchi Sea. Progress in Oceanography, 219(??):??, December 2023. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661123002124.

Perelman:2023:EFI

[PLEF^{+}23] Jessica N. Perelman, Yoann Ladroit, Pablo Escobar-Flores, E. Firing, and Jeffrey C. Drazen. Eddies and fronts influence pelagic communities across the eastern Pacific Ocean. Progress in Oceanography, 211(??):??, February 2023. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661123000101.

Pierson:2005:VDA

[PLHLF05] James J. Pierson, Andrew W. Leising, Claudia Halsband-Lenk, and Nissa Ferm. Vertical distribution and abundance of Calanus pacificus and Pseudocalanus newmani in relation to chlorophyll-a concentrations in Dabob Bay, Washington. Progress in Oceanography, 67(3–4):349–365, November 2005. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661105001473.

Pinot:2022:CEI

[PLJR22] J.-M. Pinot, J. L. López-Jurado, and M. Riera. The CANALES experiment (1996–1998). Interannual, seasonal, and mesoscale variability of the circulation in the Balearic Channels. Progress in Oceanography, 55(3–4):335–370, ???? 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661102001398.

Pun:2014:NGS

[PLK14] Iam-Fei Pun, I.-I. Lin, and Dong S. Ko. New generation of satellite-derived ocean thermal structure for the western north Pacific typhoon intensity forecasting. Progress in Oceanography, 121(??):109–124, February 2014. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661113001900.

Pratihary:2023:CDI

[PLN^{+}23] Anil Pratihary, Gaute Lavik, S. W. A. Naqvi, Gayatri Shirodkar, Amit Sarkar, Hannah Marchant, Thomas Ohde, Damodar Shenoy, Siby Kurian, Hema Uskaikar, and Marcel M. M. Kuypers. Chemolithoautotrophic denitrification intensifies nitrogen loss in the Eastern Arabian Sea Shelf waters during sulphidic events. Progress in Oceanography, 217(??):??, September–October 2023. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661123001180.

Price:1999:BME

[PLP99] N. B. Price, F. S. Lindsay, and J. M. Pates. The biogeochemistry of major elements of the suspended particulate matter of the Cretan Sea. Progress in Oceanography, 44(4):677–699, December 1999. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661199000439.

Pingree:1985:SIW

[PM85] R. D. Pingree and G. T. Mardell. Solitary internal waves in the Celtic Sea. Progress in Oceanography, 14(??):431–441, ???? 1985. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661185900217.

Perry:2013:IAM

[PM13] R. Ian Perry and Diane Masson. An integrated analysis of the marine social–ecological system of the Strait of Georgia, Canada, over the past four decades, and development of a regime shift index. Progress in Oceanography, 115(??):14–27, August 2013. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661113000633.

Pennelly:2022:TIR

[PM22] Clark Pennelly and Paul G. Myers. Tracking Irminger Rings’ properties using a sub-mesoscale ocean model. Progress in Oceanography, 201(??):??, February 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661121002184.

Paredes:2014:DDM

[PMA^{+}14] María Alejandra Paredes, Vivian Montecino, Vinka Anic, Miguel Egaña, and Leonardo Guzmán. Diatoms and dinoflagellates macroscopic regularities shaped by intrinsic physical forcing variability in Patagonian and Fuegian fjords and channels (48^{o}–56^{o}S). Progress in Oceanography, 129(??):85–97, December 2014. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661114001153.

Pena:2016:APD

[PMC16] M. Angelica Peña, Diane Masson, and Wendy Callendar. Annual plankton dynamics in a coupled physical–biological model of the Strait of Georgia, British Columbia. Progress in Oceanography, 146(??):58–74, August 2016. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661115300197.

Palma:2021:CCS

[PMC21] Elbio D. Palma, Ricardo P. Matano, and Vincent Combes. Circulation and cross-shelf exchanges in the Malvinas Islands Shelf region. Progress in Oceanography, 198(??):??, November 2021. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661121001518.

Pelegri:2006:NIN

[PMDR06] J. L. Pelegrí, A. Marrero-Díaz, and A. W. Ratsimandresy. Nutrient irrigation of the North Atlantic. Progress in Oceanography, 70(2–4):366–406, August 2006. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661106000504.

Pena:2021:FRB

[PMFNGQ21] Marian Peña, Itziar Munuera-Fernández, Enrique Nogueira, and Rafael González-Quirós. First recording of a bathypelagic deep scattering layer in the Bay of Biscay. Progress in Oceanography, 198(??):??, November 2021. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661121001531.

Packmor:2015:OIS

[PMG15] Jana Packmor, Florian Müller, and Kai Horst George. Oceanic islands and seamounts as staging posts for Copepoda Harpacticoida (Crustacea) — shallow-water Paramesochridae Lang, 1944 from the North-East Atlantic Ocean, including the (re- )description of three species and one subspecies from the Madeiran Archipelago. Progress in Oceanography, 131(??):59– 81, February 2015. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661114002018.

Peralba:2017:NSR

[PMH17] Àurea Peralba, Maria Grazia Mazzocchi, and Roger P. Harris. Niche separation and reproduction of Clausocalanus species (Copepoda, Calanoida) in the Atlantic Ocean. Progress in Oceanography, 158(??):185–202, November 2017. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661115300677.

Pennington:2006:PPE

[PMK^{+}06] J. Timothy Pennington, Kevin L. Mahoney, Victor S. Kuwahara, Dorota D. Kolber, Ruth Calienes, and Francisco P. Chavez. Primary production in the eastern tropical Pacific: a review. Progress in Oceanography, 69(2–4):285–317, May 2006. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661106000358.

Pagniello:2023:TAB

[PMM^{+}23] Camille M. L. S. Pagniello, Niall Ó Maoiléidigh, Hugo Maxwell, Michael R. Castleton, Emilius A. Aalto, Jonathan J. Dale, Robert J. Schallert, Michael J. W. Stokesbury, Ronán Cosgrove, Simon Dedman, Alan Drumm, Ross O’Neill, and Barbara A. Block. Tagging of Atlantic bluefin tuna off Ireland reveals use of distinct oceanographic hotspots. Progress in Oceanography, 219(??):??, December 2023. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661123001787.

Paredes-Mella:2022:TDA

[PMMN^{+}22] Javier Paredes-Mella, Jorge I. Mardones, Luis Norambuena, Gonzalo Fuenzalida, Gissela Labra, and Satoshi Nagai. Toxic Dinophysis acuminata in southern Chile: a comparative approach based on its first local in vitro culture and offshore-estuarine bloom dynamics. Progress in Oceanography, 209(??):??, December 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S007966112200177X.

Polimene:2015:DDP

[PMS^{+}15] L. Polimene, A. Mitra, S. F. Sailley, S. Ciavatta, C. E. Widdicombe, A. Atkinson, and J. I. Allen. Decrease in diatom palatability contributes to bloom formation in the Western English Channel. Progress in Oceanography, 137(??):484–497, September 2015. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661115000932.

Priou:2021:DMS

[PNF^{+}21] Pierre Priou, Anna Nikolopoulos, Hauke Flores, Rolf Gradinger, Erin Kunisch, Christian Katlein, Giulia Castellani, Torsten Linders, Jørgen Berge, Jonathan A. D. Fisher, and Maxime Geoffroy. Dense mesopelagic sound scattering layer and vertical segregation of pelagic organisms at the Arctic-Atlantic gateway during the midnight sun. Progress in Oceanography, 196(??): ??, August 2021. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661121000987.

Park:2000:IIV

[PO00] Won-Sun Park and Im Sang Oh. Interannual and interdecadal variations of sea surface temperature in the East Asian marginal seas. Progress in Oceanography, 47(2–4):191–204, October 2000. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661100000367.

Powell:2015:CZH

[PO15] Jesse R. Powell and Mark D. Ohman. Changes in zooplankton habitat, behavior, and acoustic scattering characteristics across glider-resolved fronts in the Southern California Current System. Progress in Oceanography, 134(??):77–92, May 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661114002213.

Palomera:2007:SPF

[POS^{+}07] I. Palomera, M. P. Olivar, J. Salat, A. Sabatés, M. Coll, A. García, and B. Morales-Nin. Small pelagic fish in the NW Mediterranean Sea: an ecological review. Progress in Oceanography, 74(2–3):377–396, August 2007. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661107000717.

Power:2006:ECF

[Pow06] Mary E. Power. Environmental controls on food web regimes: a fluvial perspective. Progress in Oceanography, 68(2–4): 125–133, February 2006. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661106000024.

Pollard:1985:SCU

[PP85] R. T. Pollard and S. Pu. Structure and circulation of the upper Atlantic Ocean northeast of the Azores. Progress in Oceanography, 14(??):443–462, ???? 1985. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0079661185900229.

Pitcher:2010:HAB

[PP10] Grant Pitcher and Stan Pillar. Harmful algal blooms in eastern boundary upwelling systems. Progress in Oceanography, 85(1– 2):1–4, April 2010. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661110000121.

Pedrosa-Pamies:2018:CCS

[PPCWJ18] R. Pedrosa-Pàmies, M. H. Conte, J. C. Weber, and R. Johnson. Carbon cycling in the Sargasso Sea water column: Insights from lipid biomarkers in suspended particles. Progress in Oceanography, 168(??):248–278, November 2018. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661118301241.

Prowe:2012:TCM

[PPD^{+}12] A. E. Friederike Prowe, Markus Pahlow, Stephanie Dutkiewicz, Michael Follows, and Andreas Oschlies. Top-down control of marine phytoplankton diversity in a global ecosystem model. Progress in Oceanography, 101(1):1–13, August 2012. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S007966111100142X.

Palanques:2012:STD

[PPdM^{+}12] A. Palanques, P. Puig, X. Durrieu de Madron, A. Sanchez-Vidal, C. Pasqual, J. Martín, A. Calafat, S. Heussner, and M. Canals. Sediment transport to the deep canyons and open-slope of the western Gulf of Lions during the 2006 intense cascading and open-sea convection period. Progress in Oceanography, 106(??):1–15, November 2012. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661112000596.

Pires:2021:DDM

[PPdS21] Rita F. T. Pires, Álvaro Peliz, and Antonina dos Santos. Into the deep — dispersal models for deep-water decapod shrimp larvae: the case of Parapenaeus longirostris. Progress in Oceanography, 194(??):??, June 2021. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661121000550.

Parzanini:2018:TRD

[PPHM18] Camilla Parzanini, Christopher C. Parrish, Jean-François Hamel, and Annie Mercier. Trophic relationships of deep-sea benthic invertebrates on a continental margin in the NW Atlantic inferred by stable isotope, elemental, and fatty acid composition. Progress in Oceanography, 168(??):279–295, November 2018. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661117303713.

Pardo:2014:ACE

[PPKR14] Paula C. Pardo, F. F. Pérez, S. Khatiwala, and A. F. Ríos. Anthropogenic CO_{2} estimates in the Southern Ocean: Storage partitioning in the different water masses. Progress in Oceanography, 120(??):230–242, January 2014. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661113001833.

Pires:2020:TBA

[PPPdS20] Rita F. T. Pires, Álvaro Peliz, Maria Pan, and Antonina dos Santos. “There and back again” — how decapod megalopae find the way home: a modelling exercise for Pachygrapsus marmoratus. Progress in Oceanography, 184(??):??, May 2020. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661120300707.

Paradis:2018:SDS

[PPSV^{+}18] Sarah Paradis, Pere Puig, Anna Sanchez-Vidal, Pere Masqué, Jordi Garcia-Orellana, Antoni Calafat, and Miquel Canals. Spatial distribution of sedimentation-rate increases in Blanes Canyon caused by technification of bottom trawling fleet. Progress in Oceanography, 169(??):241–252, December 2018. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661117303373.

Pedrosa-Pamies:2013:ISI

[PPSVC^{+}13] R. Pedrosa-Pàmies, A. Sanchez-Vidal, A. Calafat, M. Canals, and R. Durán. Impact of storm-induced remobilization on grain size distribution and organic carbon content in sediments from the Blanes Canyon area, NW Mediterranean Sea. Progress in Oceanography, 118(??):122–136, November 2013. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661113001365.

Pardo:2012:WMD

[PPVG12] Paula C. Pardo, Fiz F. Pérez, Antón Velo, and Miguel Gilcoto. Water masses distribution in the Southern Ocean: Improvement of an extended OMP (eOMP) analysis. Progress in Oceanography, 103(??):92–105, September 2012. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661112000651.

Pomeroy:1987:IPZ

[PPY87] L. R. Pomeroy, G.-A. Paffenhöfer, and J. A. Yoder. Interactions of phytoplankton, zooplankton and microorganisms. Progress in Oceanography, 19(3–4):353–372, ???? 1987. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0079661187900140.

Prandle:1991:NVN

[Pra91] David Prandle. A new view of near-shore dynamics based on observations from HF Radar. Progress in Oceanography, 27(3–4):403–438, ???? 1991. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/007966119190030P.

Prandle:1997:TCS

[Pra97] D. Prandle. Tidal currents in shelf seas — their nature and impacts. Progress in Oceanography, 40(??):245–261, ???? 1997. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661198000135.

Prandle:2004:HTR

[Pra04] D. Prandle. How tides and river flows determine estuarine bathymetries. Progress in Oceanography, 61(1):1–26, April 2004. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661104000576.

Pierdomenico:2018:ETA

[PRA^{+}18] Martina Pierdomenico, Tommaso Russo, Stefano Ambroso, Andrea Gori, Eleonora Martorelli, Lorenzo D’Andrea, Josep-Maria Gili, and Francesco L. Chiocci. Effects of trawling activity on the bamboo-coral Isidella elongata and the sea pen Funiculina quadrangularis along the Gioia Canyon (Western Mediterranean, southern Tyrrhenian Sea). Progress in Oceanography, 169(??):214–226, December 2018. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661117302525.

Prants:2022:MLL

[Pra22] S. V. Prants. Marine life at Lagrangian fronts. Progress in Oceanography, 204(??):??, June 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://

www.sciencedirect.com/science/article/pii/S0079661122000519.

Pearman:2020:IPC

[PRC^{+}20] T. R. R. Pearman, K. Robert, A. Callaway, R. Hall, C. Lo Iacono, and V. A. I. Huvenne. Improving the predictive capability of benthic species distribution models by incorporating oceanographic data — towards holistic ecological modelling of a submarine canyon. Progress in Oceanography, 184(??): ??, May 2020. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S007966112030077X.

Preller:1986:NMS

[Pre86] Ruth H. Preller. A numerical model study of the Alboran Sea Gyre. Progress in Oceanography, 16(3):113–146, ???? 1986. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0079661186900315.

Pozzato:2018:ORO

[PRL^{+}18] Lara Pozzato, Jens Rassmann, Bruno Lansard, Jean-Pascal Dumoulin, Peter van Breugel, and Christophe Rabouille. Origin of remineralized organic matter in sediments from the Rhone River prodelta (NW Mediterranean) traced by Δ ^{14}C and δ ^{13}C signatures of pore water DIC. Progress in Oceanography, 163(??):112–122, April 2018. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661117300526.

Paulmier:2009:OMZ

[PRP09] A. Paulmier and D. Ruiz-Pino. Oxygen minimum zones (OMZs) in the modern ocean. Progress in Oceanography, 80(3–4):113–128, March 2009. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661108001468.

Peck:2013:LCE

[PRTC13] Myron A. Peck, Patricia Reglero, Motomitsu Takahashi, and Ignacio A. Catalán. Life cycle ecophysiology of small pelagic fish and climate-driven changes in populations. Progress in Oceanography, 116(??):220–245, September 2013. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661113000542.

Peterson:1991:ULC

[PS91] Ray G. Peterson and Lothar Stramma. Upper-level circulation in the South Atlantic Ocean. Progress in Oceanography, 26(1): 1–73, ???? 1991. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661191900068.

Pfannkuche:1998:SBS

[PS98] O. Pfannkuche and T. Soltwedel. Small benthic size classes along the N.W. European Continental Margin: spatial and temporal variability in activity and biomass. Progress in Oceanography, 42(1–4):189–207, December 1998. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661198000342.

Perry:2008:PPC

[PS08] R. Ian Perry and Jacob F. Schweigert. Primary productivity and the carrying capacity for herring in NE Pacific marine ecosystems. Progress in Oceanography, 77(2–3):241–251, May 2008. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661108000591.

Panicker:2023:TVS

[PS23] Divya Panicker and Kathleen M. Stafford. Temporal variability of a soundscape near a mid-oceanic atoll in the northern Indian ocean. Progress in Oceanography, 214(??):??, June 2023. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661123000769.

Petrik:2019:BDG

[PSA^{+}19] Colleen M. Petrik, Charles A. Stock, Ken H. Andersen, P. Daniël van Denderen, and James R. Watson. Bottomup drivers of global patterns of demersal, forage, and pelagic fishes. Progress in Oceanography, 176(??):??, September 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661118301952.

Perez-Santos:2014:DDL

[PSGVS^{+}14] Iván Pérez-Santos, José Garcés-Vargas, Wolfgang Schneider, Lauren Ross, Sabrina Parra, and Arnoldo Valle-Levinson. Double-diffusive layering and mixing in Patagonian fjords. Progress in Oceanography, 129(??):35–49, December 2014. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S007966111400038X.

Peterson:1996:ECC

[PSK96] R. G. Peterson, L. Stramma, and G. Kortum. Early concepts and charts of ocean circulation. Progress in Oceanography, 37(1):1–115, ???? 1996. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661196800024.

Paffenhofer:1987:ADP

[PSL87] Gustav-Adolf Paffenhöfer, Byron K. Sherman, and Thomas N. Lee. Abundance, distribution and patch formation of zooplankton. Progress in Oceanography, 19(3–4):403–436, ???? 1987. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0079661187900164.

Pickart:2022:UAB

[PSM^{+}22] Robert S. Pickart, Michael A. Spall, G. W. K. Moore, Thomas J. Weingartner, Rebecca A. Woodgate, Knut Aagaard, and Koji Shimada. Upwelling in the Alaskan Beaufort Sea: Atmospheric forcing and local versus non-local response. Progress in Oceanography, 88(1–4):78–100, ???? 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661110001527.

Paniagua:2021:MCF

[PSP^{+}21] Guillermina F. Paniagua, Martin Saraceno, Alberto R. Piola, Marcela Charo, Ramiro Ferrari, Camila Artana, and Christine Provost. Malvinas current at 44.7^{o}s: First assessment of velocity temporal variability from in situ data. Progress in Oceanography, 195(??):??, July 2021. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661121000793.

Pipko:2002:CCD

[PST^{+}02] I. I. Pipko, I. P. Semiletov, P. Ya Tishchenko, S. P. Pugach, and J. P. Christensen. Carbonate chemistry dynamics in Bering Strait and the Chukchi Sea. Progress in Oceanography, 55(1–2):77–94, October 2002. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S007966110200071X.

Politikos:2015:SAF

[PST^{+}15] D. Politikos, S. Somarakis, K. P. Tsiaras, M. Giannoulaki, G. Petihakis, A. Machias, and G. Triantafyllou. Simulating anchovy’s full life cycle in the northern Aegean Sea (eastern Mediterranean): a coupled hydro-biogeochemical–IBM model. Progress in Oceanography, 138(??):399–416, November 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661114001475.

Pecseli:2010:PPE

[PTF10] Hans L. Pécseli, Jan K. Trulsen, and Øyvind Fiksen. Predator– prey encounter rates in turbulent water: Analytical models and numerical tests. Progress in Oceanography, 85(3–4):171– 179, June 2010. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661110000108.

Pecseli:2012:PPE

[PTF12] H. L. Pécseli, J. Trulsen, and Ø. Fiksen. Predator–prey encounter and capture rates for plankton in turbulent environments. Progress in Oceanography, 101(1):14–32, August 2012. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661111001431.

Pinot:1995:MAS

[PTG95] Jean-Michel Pinot, Joaquín Tintoré, and Damià Gomis. Multivariate analysis of the surface circulation in the Balearic Sea. Progress in Oceanography, 36(4):343–344, IN1, 345–376, ???? 1995. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661196000031.

Psarra:2000:PPO

[PTI00] S. Psarra, A. Tselepides, and L. Ignatiades. Primary productivity in the oligotrophic Cretan Sea (NE Mediterranean): seasonal and interannual variability. Progress in Oceanography, 46(2–4):187–204, August 2000. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661100000185.

P:2022:NIW

[PTM^{+}22] Subeesh M. P., Divya David T., Ravichandran M., Sourav Chatterjee, Ankit Pramanik, and M. Nuncio. Near-inertial waves in an Arctic fjord and their impact on vertical mixing of Atlantic water mass. Progress in Oceanography, 206(??):??, August 2022. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661122001021.

Pandi:2022:SVB

[PTP^{+}22] Sudarsana Rao Pandi, Sarat C. Tripathy, C. Parida, Aneesh A. Lotliker, R. Chandrasekhar Naik, R. K. Naik, R. K. Mishra, and N. Anilkumar. Spatiotemporal variability in bio-optical characteristics of the southwestern tropical Indian Ocean during boreal summer: Biophysical influences. Progress in Oceanography, 208(??):??, November 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661122001422.

Pinto-Torres:2023:USD

[PTPY^{+}23] Marco Pinto-Torres, Gemita Pizarro, Robert B. York, Catharina Alves de Souza, Nina Lundholm, Jorge I. Mardones, María J. Nariño, and José L. Iriarte. Unveiling species diversity within the toxic diatom genus Pseudo-nitzschia from the fjords and channels of Magallanes (48^{o}–51^{o}S), Chile. Progress in Oceanography, 211(??):??, February 2023. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661122002166.

Peng:2023:OOR

[PTZ^{+}23] Yingyu Peng, Di Tian, Feng Zhou, Han Zhang, Xiao Ma, Dingyong Zeng, Qicheng Meng, Beifeng Zhou, Ruijie Ye, Yeqing Chen, Feilong Lin, Zhentao Hu, and Xueming Zhu. Observed oceanic response to Tropical Cyclone Amphan (2020) from a subsurface mooring in the Bay of Bengal. Progress in Oceanography, 219(??):??, December 2023. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S007966112300191X.

Pugh:1984:DMD

[Pug84] P. R. Pugh. The diel migrations and distributions within a mesopelagic community in the North East Atlantic. 7. Siphonophores. Progress in Oceanography, 13(3–4):461–489, ???? 1984. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661184900168.

Pena:2007:SIV

[PV07] M. Angelica Peña and Diana E. Varela. Seasonal and interannual variability in phytoplankton and nutrient dynamics along Line P in the NE subarctic Pacific. Progress in Oceanography, 75(2):200–222, October 2007. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661107001528.

Petitgas:2018:MIS

[PV18] Pierre Petitgas and Begoña Santos Vásquez. Multidisciplinary integrated surveys: Objectives, achievements and future directions. Progress in Oceanography, 166(??):2–3, September 2018. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661118300636.

Plaza-Vega:2024:AER

[PVA24] Francisco Plaza-Vega and Héctor Araya. Anchovy (Engraulis ringens) and Pacific sardine (Sardinops sagax) variability changes in northern Chile associated with the environment and inter species synchronicity: GARCH model with exogenous variable and hybrid Bayesian deep learning estimation approach. Progress in Oceanography, 221(??):??, February 2024. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661123002331.

Pena:2023:SDS

[PVB23] Marian Peña and Pedro Vélez-Belchí. Shaping of deep scattering layers by the seascape dynamics in the Canary Islands. Progress in Oceanography, 210(??):??, January 2023. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661122002129.

Pahlow:2008:AMP

[PVC^{+}08] Markus Pahlow, Alain F. Vézina, Benoit Casault, Heidi Maass, Louise Malloch, Daniel G. Wright, and Youyu Lu. Adaptive model of plankton dynamics for the North Atlantic. Progress in Oceanography, 76(2):151–191, February 2008. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S007966110700208X.

Paklar:2020:RBS

[PVG^{+}20] Gordana Beg Paklar, Ivica Vilibić, Branka Grbec, Frano Matić, Hrvoje Mihanović, Tomislav Džoić, Danijela Šantić, Stefanija Šestanović, Mladen Šolić, Stjepan Ivatek-Šahdan, and Grozdan Kušpilić. Record-breaking salinities in the middle Adriatic during summer 2017 and concurrent changes in the microbial food web. Progress in Oceanography, 185(??):??, June 2020. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661120300847.

Pan:2020:EDP

[PVM^{+}20] B. Jack Pan, Maria Vernet, Lauren Manck, Kiefer Forsch, Lindsey Ekern, Martina Mascioni, Katherine A. Barbeau, Gastón O. Almandoz, and Alexander J. Orona. Environmental drivers of phytoplankton taxonomic composition in an Antarctic fjord. Progress in Oceanography, 183(??):??, April 2020. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661120300343.

Prasanth:2023:FMD

[PVV23] R. Prasanth, V. Vijith, and P. N. Vinayachandran. Formation, maintenance and diurnal variability of subsurface chlorophyll maximum during the summer monsoon in the southern Bay of Bengal. Progress in Oceanography, 212(??):??, March 2023. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661123000174.

Pitcher:1991:STV

[PWMIM91] G. C. Pitcher, D. R. Walker, B. A. Mitchell-Innes, and C. L. Moloney. Short-term variability during an anchor station study in the southern Benguela upwelling system: Phytoplankton dynamics. Progress in Oceanography, 28(1–2):39– 64, ???? 1991. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/007966119190020M.

Peron:2016:MSD

[PWZ^{+}16] Clara Péron, Dirk C. Welsford, Philippe Ziegler, Timothy D. Lamb, Nicolas Gasco, Charlotte Chazeau, Romain Sinègre, and Guy Duhamel. Modelling spatial distribution of Patagonian toothfish through life-stages and sex and its implications for the fishery on the Kerguelen Plateau. Progress in Oceanography, 141(??):81–95, February 2016. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S007966111530015X.

Pethybridge:2015:USI

[PYKF15] Heidi R. Pethybridge, Jock W. Young, Petra M. Kuhnert, and Jessica H. Farley. Using stable isotopes of albacore tuna and predictive models to characterize bioregions and examine ecological change in the SW Pacific Ocean. Progress in Oceanography, 134(??):293–303, May 2015. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661115000415.

Pinardi:2015:MSL

[PZA^{+}15] Nadia Pinardi, Marco Zavatarelli, Mario Adani, Giovanni Coppini, Claudia Fratianni, Paolo Oddo, Simona Simoncelli, Marina Tonani, Vladislav Lyubartsev, Srdjan Dobricic, and Antonio Bonaduce. Mediterranean Sea largescale low-frequency ocean variability and water mass formation rates from 1987 to 2007: a retrospective analysis. Progress in Oceanography, 132(??):318–332, March 2015. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S007966111300222X.

Queiroga:2007:OBP

[QCdS^{+}07] Henrique Queiroga, Teresa Cruz, Antonina dos Santos, Jesus Dubert, Juan Ignácio González-Gordillo, José Paula, Álvaro Peliz, and A. Miguel P. Santos. Oceanographic and behavioural processes affecting invertebrate larval dispersal and supply in the western Iberia upwelling ecosystem. Progress in Oceanography, 74(2–3):174–191, August 2007. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661107000663.

Qiu:2015:IFA

[Qiu15] Yongsong Qiu. Iron fertilisation by Asian dust influences North Pacific sardine regime shifts. Progress in Oceanography, 134(??):370–378, May 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S007966111500052X.

Qiu:2010:RFP

[QLW10] Yongsong Qiu, Zhaojin Lin, and Yuezhong Wang. Responses of fish production to fishing and climate variability in the northern South China Sea. Progress in Oceanography, 85(3–4):197– 212, June 2010. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661110000315.

Quan:2022:CIF

[QLY^{+}22] Qi Quan, Zhiqiang Liu, Yang Yang, Zhongya Cai, Han Zhang, and Xiaohui Liu. Characterization of intraseasonal fluctuations in the abyssal South China Sea: an insight into the energy pathway. Progress in Oceanography, 206(??):??, August 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661122000908.

Qiao:2022:SDV

[QNK^{+}22] Yu-Xiang Qiao, Hirohiko Nakamura, Shinichiro Kako, Ayako Nishina, and Tomohiko Tomita. Synchronized decadal variabilities in the Kuroshio and Kuroshio Extension system. Progress in Oceanography, 204(??):??, June 2022. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661122000696.

Quiroga:2022:GPS

[QOS^{+}22] Eduardo Quiroga, Paula Ortiz, Eulogio H. Soto, Nicole Salinas, Nicole Olguín, and Chester Sands. Geographic patterns of soft-bottoms benthic communities in Chilean Patagonian fjords (47^{o}S-54^{o}S) — influence of environmental stress on diversity patterns and stable isotope signatures. Progress in Oceanography, 204(??):??, June 2022. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661122000714.

Quinones:2003:PBS

[QPR03] Renato A. Quinones, Trevor Platt, and Jaime Rodríguez. Patterns of biomass-size spectra from oligotrophic waters of the Northwest Atlantic. Progress in Oceanography, 57(3–4):405– 427, June 2003. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661103001083.

Queiros:2015:CBC

[QSC^{+}15] Ana M. Queirós, Nicholas Stephens, Richard Cook, Chiara Ravaglioli, Joana Nunes, Sarah Dashfield, Carolyn Harris, Gavin H. Tilstone, James Fishwick, Ulrike Braeckman, Paul J. Somerfield, and Stephen Widdicombe. Can benthic community structure be used to predict the process of bioturbation in real ecosystems? Progress in Oceanography, 137(??):559–569, September 2015. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661115000944.

Qiu:2024:OET

[QYF^{+}24] Chunhua Qiu, Zihao Yang, Ming Feng, Jun Yang, Tom P. Rippeth, Xiaodong Shang, Zhenyu Sun, Chunsheng Jing, and Dongxiao Wang. Observational energy transfers of a spiral cold filament within an anticyclonic eddy. Progress in Oceanography, 220(??):??, January 2024. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661123002306.

Roy:2015:CIP

[RA15] Rajdeep Roy and A. C. Anil. Complex interplay of physical forcing and Prochlorococcus population in ocean. Progress in Oceanography, 137 (part A)(??):250–260, September 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661115001494.

Roe:1984:DMDa

[RAB^{+}84] H. S. J. Roe, M. V. Angel, J. Badcock, P. Domanski, P. T. James, P. R. Pugh, and M. H. Thurston. The diel migrations and distributions within a Mesopelagic community in the North East Atlantic. 1. Introduction and sampling procedures. Progress in Oceanography, 13(3–4):245–268, ???? 1984. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0079661184900107.

Ruhl:2011:SNI

[RAB^{+}11] Henry A. Ruhl, Michel André, Laura Beranzoli, M. Namik Çağatay, Ana Colaço, Mathilde Cannat, Juanjo J. Dañobeitia, Paolo Favali, Louis Géli, Michael Gillooly, Jens Greinert, Per O. J. Hall, Robert Huber, Johannes Karstensen, Richard S. Lampitt, Kate E. Larkin, Vasilios Lykousis, Jürgen Mienert, J. Miguel Miranda, Roland Person, Imants G. Priede, Ingrid Puillat, Laurenz Thomsen, and Christoph Waldmann. Societal need for improved understanding of climate change, anthropogenic impacts, and geo-hazard warning drive development of ocean observatories in European Seas. Progress in Oceanography, 91(1):1–33, October 2011. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661111000383.

Rubio:2005:FSB

[RAE^{+}05] Anna Rubio, Pedro A. Arnau, Manuel Espino, María del Mar Flexas, Gabriel Jordà, Jordi Salat, Joan Puigdefàbregas, and Agustín S.-Arcilla. A field study of the behaviour of an anticyclonic eddy on the Catalan continental shelf (NW Mediterranean). Progress in Oceanography, 66(2–4):142–156, August 2005. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661105000674.

Roura:2019:OPS

[RAG^{+}19] Álvaro Roura, Michael Amor, Ángel F. González, Ángel Guerra, Eric D. Barton, and Jan M. Strugnell. Oceanographic processes shape genetic signatures of planktonic cephalopod paralarvae in two upwelling regions. Progress in Oceanography, 170(??):11–27, January 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661118300545.

Rona:1965:GGM

[RANS65] Elizabeth Rona, L. K. Akers, John E. Noakes, and Irwin Supernaw. Geochronology in the Gulf of Mexico. Part I. Progress in Oceanography, 3(??):289–295, ???? 1965. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0079661165900248.

Rios:1995:CSD

[RAP95] A. F. Ríos, T. R. Anderson, and F. F. Pérez. The carbonic system distribution and fluxes in the NE Atlantic during Spring 1991. Progress in Oceanography, 35(4):295–314, ???? 1995. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/007966119580001A.

Roura:2013:STM

[RÁSG^{+}13] Álvaro Roura, Xosé A. Álvarez-Salgado, Ángel F. González, María Gregori, Gabriel Rosón, and Ángel Guerra. Shortterm meso-scale variability of mesozooplankton communities in a coastal upwelling system (NW Spain). Progress in Oceanography, 109(??):18–32, February 2013. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661112001127.

Ruvalcaba-Aroche:2022:OSP

[RASVB^{+}22] Erick D. Ruvalcaba-Aroche, Laura Sánchez-Velasco, Emilio Beier, Eric D. Barton, Víctor M. Godínez, Jaime Gómez-Gutiérrez, and Raúl O. Martínez-Rincón. Ommastrephid squid paralarvae potential nursery habitat in the tropical-subtropical convergence off Mexico. Progress in Oceanography, 202(??): ??, March 2022. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661122000246.

Roe:1984:DMDd

[RB84] H. S. J. Roe and J. Badcock. The diel migrations and distributions within a mesopelagic community in the North East Atlantic. 5. Vertical migrations and feeding of fish. Progress in Oceanography, 13(3–4):389–424, ???? 1984. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0079661184900144.

Robertson:2020:CDV

[RB20] Roxanne R. Robertson and Eric P. Bjorkstedt. Climate-driven variability in Euphausia pacifica size distributions off northern California. Progress in Oceanography, 188(??):??, October 2020. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661120301518.

Relvas:2007:POW

[RBD^{+}07] Paulo Relvas, E. D. Barton, Jesús Dubert, Paulo B. Oliveira, Álvaro Peliz, J. C. B. da Silva, and A. Miguel P. Santos. Physical oceanography of the western Iberia ecosystem: Latest views and challenges. Progress in Oceanography, 74(2–3):149–173, August 2007. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S007966110700095X.

Ruzicka:2012:IVN

[RBE^{+}12] James J. Ruzicka, Richard D. Brodeur, Robert L. Emmett, John H. Steele, Jeannette E. Zamon, Cheryl A. Morgan, Andrew C. Thomas, and Thomas C. Wainwright. Interannual variability in the Northern California Current food web structure: Changes in energy flow pathways and the role of forage fish, euphausiids, and jellyfish. Progress in Oceanography, 102(??):19–41, September 2012. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661112000080.

Reissmann:2009:VMB

[RBF^{+}09] Jan Hinrich Reissmann, Hans Burchard, Rainer Feistel, Eberhard Hagen, Hans Ulrich Lass, Volker Mohrholz, Günther Nausch, Lars Umlauf, and Gunda Wieczorek. Vertical mixing in the Baltic Sea and consequences for eutrophication — a review. Progress in Oceanography, 82(1):47–80, July 2009. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661109000159.

Rodriguez:2004:IMP

[RBHLA04] J. M. Rodríguez, E. D. Barton, S. Hernández-León, and J. Arístegui. The influence of mesoscale physical processes on the larval fish community in the Canaries CTZ, in summer. Progress in Oceanography, 62(2–4):171–188, August 2004. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661104000953.

Roe:1990:BMI

[RBL90] H. S. J. Roe, D. S. M. Billett, and R. S. Lampitt. Benthic/midwater interactions on the Madeira Abyssal Plain; evidence for biological transport pathways. Progress in Oceanography, 24(??):127–140, ???? 1990. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/007966119090025W.

Rebolledo:2019:CBV

[RBL^{+}19] Lorena Rebolledo, Sebastien Bertrand, Carina B. Lange, Fabian J. Tapia, Eduardo Quiroga, Matthias Troch, Nelson Silva, Paola Cárdenas, and Silvio Pantoja. Compositional and biogeochemical variations of sediments across the terrestrial-marine continuum of the Baker–Martínez fjord system (Chile, 48^{o}S). Progress in Oceanography, 174(??):89– 104, May 2019. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S007966111830048X.

Riquelme-Bugueno:2012:IUV

[RBNJ^{+}12] Ramiro Riquelme-Bugueño, Sergio Núñez, Erika Jorquera, Lesly Valenzuela, Rubén Escribano, and Samuel Hormazábal. The influence of upwelling variation on the spatially-structured euphausiid community off central-southern Chile in 2007– 2008. Progress in Oceanography, 92–95(??):146–165, January 2012. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661111000577.

Riquelme-Bugueno:2020:FAC

[RBPGJ^{+}20] Ramiro Riquelme-Bugueño, Silvio Pantoja-Gutiérrez, Erika Jorquera, Valeria Anabalón, Benjamín Srain, and Wolfgang Schneider. Fatty acid composition in the endemic Humboldt Current krill, Euphausia mucronata (Crustacea, Euphausiacea) in relation to the phytoplankton community and oceanographic variability off Dichato coast in central Chile. Progress in Oceanography, 188(??):??, October 2020. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661120301646.

Renner:2023:HIN

[RBR^{+}23] Angelika H. H. Renner, Allison Bailey, Marit Reigstad, Arild Sundfjord, Melissa Chierici, and Elizabeth M. Jones. Hydrography, inorganic nutrients and chlorophyll a linked to sea ice cover in the Atlantic Water inflow region north of Svalbard. Progress in Oceanography, 219(??):??, December 2023. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661123002057.

Rubio:2009:MEA

[RBS^{+}09] Anna Rubio, Bruno Blanke, Sabrina Speich, Nicolas Grima, and Claude Roy. Mesoscale eddy activity in the southern Benguela upwelling system from satellite altimetry and model data. Progress in Oceanography, 83(1–4):288–295, December 2009. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661109000925.

Reeve:2020:CHC

[RBS^{+}20] Krissy Anne Reeve, Olaf Boebel, Volker Strass, Torsten Kanzow, and Rüdiger Gerdes. Corrigendum to “Horizontal circulation and volume transports in the Weddell Gyre derived from Argo float data” [Progr. Oceanogr. 175 (2019) 263– 283]. Progress in Oceanography, 184(??):??, May 2020. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661120300331. See [RBS^{+}22].

Reeve:2022:HCV

[RBS^{+}22] Krissy Anne Reeve, Olaf Boebel, Volker Strass, Torsten Kanzow, and Rüdiger Gerdes. Horizontal circulation and volume transports in the Weddell Gyre derived from Argo float data. Progress in Oceanography, 175(??):263–283, ???? 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661117302756. See corrigendum

[RBS^{+}20].

Richardson:2000:CMT

[RBZ00] P. L. Richardson, A. S. Bower, and W. Zenk. A census of meddies tracked by floats. Progress in Oceanography, 45(2):209–250, February 2000. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661199000531.

Reboreda:2015:OIM

[RCÁS^{+}15] Rosa Reboreda, Carmen G. Castro, Xosé A. Álvarez-Salgado, Rita Nolasco, Nuno G. F. Cordeiro, Henrique Queiroga, and Jesus Dubert. Oxygen in the Iberian margin: a modeling study. Progress in Oceanography, 131(??):1–20, February 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661114001529.

Ren:2020:DCF

[RCB^{+}20] Jian Ren, Jianfang Chen, Youcheng Bai, Marie-Alexandrine Sicre, Zhixiong Yao, Long Lin, Jingjing Zhang, Hongliang Li, Bin Wu, Haiyan Jin, Zhongqiang Ji, Yanpei Zhuang, and Yangjie Li. Diatom composition and fluxes over the northwind ridge, western Arctic Ocean: Impacts of marine surface circulation and sea ice distribution. Progress in Oceanography, 186 (??):??, July 2020. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661120301166.

Rastelli:2018:RRB

[RCC^{+}18] Eugenio Rastelli, Cinzia Corinaldesi, Miquel Canals, Roberto Danovaro, and Antonio Dell’Anno. Rapid response of benthic deep-sea microbes (viruses and prokaryotes) to an intense dense shelf water cascading event in a submarine canyon of the NW Mediterranean Sea. Progress in Oceanography, 168(??):35–42, November 2018. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661118301149.

Reverdin:1994:SSN

[RCD^{+}94] Gilles Reverdin, D. Cayan, H. D. Dooley, D. J. Ellett, S. Levitus, Y. du Penhoat, and A. Dessier. Surface salinity of the North Atlantic: Can we reconstruct its fluctuations over the last one hundred years? Progress in Oceanography, 33(4):303– 346, ???? 1994. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661194900213.

Romano:2013:STV

[RCF^{+}13] C. Romano, J. Coenjaerts, M. M. Flexas, D. Zúñiga, A. Vanreusel, J. B. Company, and D. Martin. Spatial and temporal variability of meiobenthic density in the Blanes submarine canyon (NW Mediterranean). Progress in Oceanography, 118(??):144–158, November 2013. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661113001390.

Reimer:2016:REA

[RCGC^{+}16] Janet J. Reimer, Alejandro Cueva, Gilberto Gaxiola-Castro, Ruben Lara-Lara, and Rodrigo Vargas. Random error analysis of marine xCO_{2} measurements in a coastal upwelling region. Progress in Oceanography, 143(??):1–12, April 2016.

CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661116000185.

Reid:2003:CPR

[RCM^{+}03] P. C. Reid, J. M. Colebrook, J. B. L. Matthews, J. Aiken, and Continuous Plankton Recorder Team. The continuous plankton recorder: concepts and history, from plankton indicator to undulating recorders. Progress in Oceanography, 58(2–4):117– 173, August 2003. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661103001277.

Reigstad:2011:IRC

[RCS^{+}11] Marit Reigstad, JoLynn Carroll, Dag Slagstad, Ingrid Ellingsen, and Paul Wassmann. Intra-regional comparison of productivity, carbon flux and ecosystem composition within the northern Barents Sea. Progress in Oceanography, 90(1–4): 33–46, July 2011. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661111000206.

Rogachev:2001:LFM

[RCSA01] Konstantin A. Rogachev, Eddy C. Carmack, Alexander S. Salomatin, and Marina G. Alexanina. Lunar fortnightly modulation of tidal mixing near Kashevarov Bank, Sea of Okhotsk, and its impacts on biota and sea ice. Progress in Oceanography, 49(??):373–390, ???? 2001. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661101000313.

Ruiz-Castillo:2022:SCS

[RCSHW22] Eugenio Ruiz-Castillo, Jonathan Sharples, Jo Hopkins, and Malcolm Woodward. Seasonality in the cross-shelf physical structure of a temperate shelf sea and the implications for nitrate supply. Progress in Oceanography, 177(??):??, October 2022. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661117302562.

Rumin-Caparros:2016:PFT

[RCSVGP^{+}16] A. Rumín-Caparrós, A. Sanchez-Vidal, C. González-Pola, G. Lastras, A. Calafat, and M. Canals. Particle fluxes and their drivers in the Avilés submarine canyon and adjacent slope, central Cantabrian margin, Bay of Biscay. Progress in Oceanography, 144(??):39–61, May 2016. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661116300519.

Ridgway:2003:MSM

[RD03] K. R. Ridgway and J. R. Dunn. Mesoscale structure of the mean East Australian Current System and its relationship with topography. Progress in Oceanography, 56(2):189–222, February 2003. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661103000041.

Ray:2011:ERT

[RD11] Richard D. Ray and Bruce C. Douglas. Experiments in reconstructing twentieth-century sea levels. Progress in Oceanography, 91(4):496–515, December 2011. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661111000759.

Roberts:2021:CSN

[RDC^{+}21] Elliott G. Roberts, Minhan Dai, Zhimian Cao, Weidong Zhai, Liguo Guo, Samuel S. P. Shen, and Chuanjun Du. The carbonate system of the northern South China Sea: Seasonality and exchange with the western North Pacific. Progress in Oceanography, 191(??):??, February 2021. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661120302019.

Richon:2018:MIA

[RDD^{+}18] Camille Richon, Jean-Claude Dutay, François Dulac, Rong Wang, Yves Balkanski, Pierre Nabat, Olivier Aumont, Karine Desboeufs, Benoît Laurent, Cécile Guieu, Patrick Raimbault, and Jonathan Beuvier. Modeling the impacts of atmospheric deposition of nitrogen and desert dust-derived phosphorus on nutrients and biological budgets of the Mediterranean Sea. Progress in Oceanography, 163(??):21–39, April 2018. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661116301811.

Receveur:2021:EFC

[RDG^{+}21] Aurore Receveur, Cyril Dutheil, Thomas Gorgues, Christophe Menkes, Matthieu Lengaigne, Simon Nicol, Patrick Lehodey, Valerie Allain, Frederic Menard, and Anne Lebourges-Dhaussy. Exploring the future of the Coral Sea micronekton. Progress in Oceanography, 195(??):??, July 2021. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S007966112100080X.

Reverdin:1991:NST

[RDL^{+}91] G. Reverdin, P. Delécluse, C. Lévy, P. Andrich, A. Morlière, and J. M. Verstraete. The near surface tropical Atlantic in 1982–1984: Results from a numerical simulation and a data analysis. Progress in Oceanography, 27(3–4):273–340, ???? 1991. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/007966119190027J.

Rypina:2021:EIV

[RDP^{+}21] Irina I. Rypina, Michael M. Dotzel, Lawrence J. Pratt, Christina M. Hernandez, and Joel K. Llopiz. Exploring interannual variability in potential spawning habitat for Atlantic bluefin tuna in the Slope Sea. Progress in Oceanography, 192(??):??, March 2021. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661121000045.

Read:2000:CWM

[Rea00] J. F. Read. CONVEX-91: water masses and circulation of the Northeast Atlantic subpolar gyre. Progress in Oceanography, 48(4):461–510, ???? 2000. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661101000118.

Rebstock:2002:AZS

[Reb02] Ginger A. Rebstock. An analysis of a zooplankton samplinggear change in the CalCOFI long-term monitoring program, with implications for copepod population abundance trends. Progress in Oceanography, 53(??):215–229, ???? 2002. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661102000319.

Raghukumar:2015:IAP

[REG^{+}15] Kaustubha Raghukumar, Christopher A. Edwards, Nicole L. Goebel, Gregoire Broquet, Milena Veneziani, Andrew M. Moore, and Jon P. Zehr. Impact of assimilating physical oceanographic data on modeled ecosystem dynamics in the California Current System. Progress in Oceanography, 138 (??):546–558, November 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661115000063.

Reid:1986:TGC

[Rei86] Joseph L. Reid. On the total geostrophic circulation of the South Pacific Ocean: Flow patterns, tracers and transports. Progress in Oceanography, 16(1):1–61, ???? 1986. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0079661186900364.

Reid:1989:TGC

[Rei89] Joseph L. Reid. On the total geostrophic circulation of the South Atlantic Ocean: Flow patterns, tracers, and transports. Progress in Oceanography, 23(3):149–244, ???? 1989. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0079661189900013.

Reid:1994:TGC

[Rei94] Joseph L. Reid. On the total geostrophic circulation of the North Atlantic Ocean: Flow patterns, tracers, and transports. Progress in Oceanography, 33(1):1–92, ???? 1994. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0079661194900140.

Reid:1997:TGC

[Rei97] Joseph L. Reid. On the total geostrophic circulation of the Pacific Ocean: flow patterns, tracers, and transports. Progress in Oceanography, 39(4):263–352, ???? 1997. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661197000128.

Reid:2003:TGC

[Rei03] Joseph L. Reid. On the total geostrophic circulation of the Indian Ocean: flow patterns, tracers, and transports. Progress in Oceanography, 56(1):137–186, January 2003. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661102001416.

Romero:2017:SSC

[RF17] Oscar E. Romero and Gerhard Fischer. Shift in the species composition of the diatom community in the eutrophic Mauritanian coastal upwelling: Results from a multiyear sediment trap experiment (2003–2010). Progress in Oceanography, 159(??):31–44, December 2017. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661117300319.

Rose:2015:DFC

[RFC^{+}15] Kenneth A. Rose, Jerome Fiechter, Enrique N. Curchitser, Kate Hedstrom, Miguel Bernal, Sean Creekmore, Alan Haynie, Shin ichi Ito, Salvador Lluch-Cota, Bernard A. Megrey, Chris A. Edwards, Dave Checkley, Tony Koslow, Sam Mc-Clatchie, Francisco Werner, Alec MacCall, and Vera Agostini. Demonstration of a fully-coupled end-to-end model for small pelagic fish using sardine and anchovy in the California Current. Progress in Oceanography, 138(??):348–380, November 2015. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661115000233.

Ramos:2021:RBL

[RFFL21] Marília S. Ramos, Leandro Farina, Sérgio Henrique Faria, and Chen Li. Relationships between large-scale climate modes and the South Atlantic Ocean wave climate. Progress in Oceanography, 197(??):??, September–October 2021. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661121001452.

Romero:2016:ETD

[RFKC16] Oscar E. Romero, Gerhard Fischer, Johannes Karstensen, and Pedro Cermeño. Eddies as trigger for diatom productivity in the open-ocean Northeast Atlantic. Progress in Oceanography, 147(??):38–48, September 2016. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661116300180.

Rosell-Fieschi:2015:ZJE

[RFPG15] Miquel Rosell-Fieschi, Josep L. Pelegrí, and Jéröme Gourrion. Zonal jets in the equatorial Atlantic Ocean. Progress in Oceanography, 130(??):1–18, January 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661114001311.

Rouyer:2010:ENA

[RFS10] Tristan Rouyer, Jean-Marc Fromentin, and Nils Chr. Stenseth. Environmental noise affects the fluctuations of Atlantic large pelagics. Progress in Oceanography, 86(1–2):267–275, July 2010. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661110000595.

Rain-Franco:2019:VCD

[RFSCF19] Angel Rain-Franco, Marcus Sobarzo, Jocelyne Caparros, and Camila Fernandez. Variability of chromophoric dissolved organic matter in three freshwater-influenced systems along central-southern Chile. Progress in Oceanography, 174(??):154– 161, May 2019. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661118300879.

Rosland:1994:DOM

[RG94] Rune Rosland and Jarl Giske. A dynamic optimization model of the diel vertical distribution of a pelagic planktivorous fish. Progress in Oceanography, 34(1):1–43, ???? 1994. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0079661194900256.

Ruddick:2003:EOD

[RG03a] Barry Ruddick and Ann E. Gargett. Erratum: Oceanic double-diffusion: introduction [Progress in Oceanography 56 (2003) 381–393]. Progress in Oceanography, 57(2):237, May 2003. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661103000867. See [RG03b].

Ruddick:2003:ODI

[RG03b] Barry Ruddick and Ann E. Gargett. Oceanic double-infusion: introduction. Progress in Oceanography, 56(3–4):381–393, March 2003. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661103000247. See erra-

tum [RG03a].

Roemmich:2009:MAC

[RG09] Dean Roemmich and John Gilson. The 2004–2008 mean and annual cycle of temperature, salinity, and steric height in the global ocean from the Argo Program. Progress in Oceanography, 82(2):81–100, August 2009. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661109000160.

Reygondeau:2017:BRM

[RGB^{+}17] Gabriel Reygondeau, Cécile Guieu, Fabio Benedetti, Jean-Olivier Irisson, Sakina-Dorothée Ayata, Stéphane Gasparini, and Philippe Koubbi. Biogeochemical regions of the Mediterranean Sea: an objective multidimensional and multivariate environmental approach. Progress in Oceanography, 151(??):138–148, February 2017. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661116302270.

Ragueneau:2001:BSC

[RGC^{+}01] O. Ragueneau, M. Gallinari, L. Corrin, S. Grandel, P. Hall, A. Hauvespre, R. S. Lampitt, D. Rickert, H. Stahl, A. Tengberg, and R. Witbaard. The benthic silica cycle in the Northeast Atlantic: annual mass balance, seasonality, and importance of non-steady-state processes for the early diagenesis of biogenic opal in deep-sea sediments. Progress in Oceanography, 50(??):171–200, ???? 2001. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661101000532.

Rixen:2022:MCF

[RGE22] Tim Rixen, Birgit Gaye, and Kay-Christian Emeis. The monsoon, carbon fluxes, and the organic carbon pump in the northern Indian Ocean. Progress in Oceanography, 175(??):24–39, ???? 2022. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661117302161.

Rixen:2005:DOF

[RGI05] T. Rixen, M. V. S. Guptha, and V. Ittekkot. Deep ocean fluxes and their link to surface ocean processes and the biological pump. Progress in Oceanography, 65(2–4):240–259, May 2005. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661105000443.

Royer:2001:IVN

[RGM01] Thomas C. Royer, Chester E. Grosch, and Lawrence A. Mysak. Interdecadal variability of Northeast Pacific coastal freshwater and its implications on biological productivity. Progress in Oceanography, 49(??):95–111, ???? 2001. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661101000179.

Rodriguez-Galvez:2023:TBC

[RGMPR23] Susana Rodríguez-Gálvez, Diego Macías, Laura Prieto, and Javier Ruiz. Top-down and bottom-up control of phytoplankton in a mid-latitude continental shelf ecosystem. Progress in Oceanography, 217(??):??, September–October 2023. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S007966112300126X.

Roque:2023:TDU

[RGPB^{+}23] D. Roque, J. J. Gomiz-Pascual, M. Bruno, R. Sánchez-Leal, C. J. González, M. García, L. M. Fernández-Salas, and F. J. Hernández-Molina. Tidal dynamics on the upper continental slope of the eastern Gulf of Cádiz: the interplay between water masses and its effects on seafloor morphology. Progress in Oceanography, 212(??):??, March 2023. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661122002130.

Robinson:2023:DOP

[RHB23] Edward Robinson, Phil Hosegood, and Adam Bolton. Dynamical oceanographic processes impact on reef manta ray behaviour: Extreme Indian Ocean Dipole influence on local internal wave dynamics at a remote tropical atoll. Progress in Oceanography, 218(??):??, November 2023. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661123001726.

Robinson:2013:ISB

[RHBS13] Clifford L. K. Robinson, Doug Hrynyk, J. Vaughn Barrie, and Jake Schweigert. Identifying subtidal burying habitat of Pacific sand lance (Ammodytes hexapterus) in the Strait of Georgia, British Columbia, Canada. Progress in Oceanography, 115(??):119–128, August 2013. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661113000803.

Rumyantseva:2019:PSB

[RHM^{+}19] Anna Rumyantseva, Stephanie Henson, Adrian Martin, Andrew F. Thompson, Gillian M. Damerell, Jan Kaiser, and Karen J. Heywood. Phytoplankton spring bloom initiation: the impact of atmospheric forcing and light in the temperate North Atlantic Ocean. Progress in Oceanography, 178(??):??, November 2019. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661119303829.

Rikiishi:1986:TFS

[RI86] Kunio Rikiishi and Takashi Ichiye. Tidal fluctuation of the surface currents of the Kuroshio in the East China Sea. Progress in Oceanography, 17(3–4):193–213, ???? 1986. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0079661186900443.

Richardson:1985:DDN

[Ric85] P. L. Richardson. Drifting derelicts in the North Atlantic 1883–1902. Progress in Oceanography, 14(??):463–483, ???? 1985. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0079661185900230.

Richardson:1993:CEO

[Ric93] P. L. Richardson. A census of eddies observed in North Atlantic SOFAR float data. Progress in Oceanography, 31(1):1– 50, ???? 1993. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661193900226.

Richez:1994:ASA

[Ric94] Claude Richez. Airborne synthetic aperture radar tracking of internal waves in the Strait of Gibraltar. Progress in Oceanography, 33(2):93–97, 99–102, IN1–IN10, 103–116, IN11, 117–119, 121–128, IN13–IN14, 129–149, 151–159, ???? 1994. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/007966119490023X.

Rice:2001:IVM

[Ric01] Jake Rice. Implications of variability on many time scales for scientific advice on sustainable management of living marine resources. Progress in Oceanography, 49(??):189–209, ???? 2001. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661101000222.

Richardson:2008:HMO

[Ric08] Philip L. Richardson. On the history of meridional overturning circulation schematic diagrams. Progress in Oceanography, 76(4):466–486, March 2008. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661108000086.

Richardson:2015:UDS

[Ric15] Philip L. Richardson. Upwind dynamic soaring of albatrosses and UAVs. Progress in Oceanography, 130(??):146–156, January 2015. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661114001761.

Richardson:2022:HDA

[Ric22] Philip L. Richardson. How do albatrosses fly around the world without flapping their wings? Progress in Oceanography, 88(1–4):46–58, ???? 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661110001072.

Rajasakaren:2019:TAC

[RJO^{+}19] Balamuralli Rajasakaren, Emil Jeansson, Are Olsen, Toste Tanhua, Truls Johannessen, and W. M. Smethie. Trends in anthropogenic carbon in the Arctic Ocean. Progress in Oceanography, 178(??):??, November 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S007966111830168X.

Roe:1984:DMDe

[RJT84] H. S. J. Roe, P. T. James, and M. H. Thurston. The diel migrations and distributions within a mesopelagic community in the North East Atlantic. 6. Medusae, ctenophores, amphipods and euphausiids. Progress in Oceanography, 13(3–4):425–460, ???? 1984. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661184900156.

Rebstock:2003:CTM

[RK03a] Ginger A. Rebstock and Young Shil Kang. A comparison of three marine ecosystems surrounding the Korean peninsula: Responses to climate change. Progress in Oceanography, 59(4):357–379, December 2003. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661103001769.

Ruddick:2003:OTIa

[RK03b] Barry Ruddick and Oliver Kerr. Oceanic thermohaline intrusions: theory. Progress in Oceanography, 56(3–4):483–497, March 2003. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661103000296.

Riehl:2020:UWB

[RK20] Torben Riehl and Mona A. L. Kühn. Uniting what belongs together — reevaluation of the isopod species Macrostylis grandis and M. ovata using ontogenetic, morphological and genetic evidence. Progress in Oceanography, 181(??):??, February 2020. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661119304185.

Runge:2010:UCI

[RKC^{+}10] Jeffrey A. Runge, Adrienne I. Kovach, James H. Churchill, Lisa A. Kerr, John R. Morrison, Robert C. Beardsley, David L. Berlinsky, Changsheng Chen, Steven X. Cadrin, Cabell S. Davis, Kathryn H. Ford, Jonathan H. Grabowski, W. Huntting Howell, Rubao Ji, Rebecca J. Jones, Andrew J. Pershing, Nicholas R. Record, Andrew C. Thomas, Graham D. Sherwood, Shelly M. L. Tallack, and David W. Townsend. Understanding climate impacts on recruitment and spatial dynamics of Atlantic cod in the Gulf of Maine: Integration of observations and modeling. Progress in Oceanography, 87(1–4):251–263, October 2010. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S007966111000128X.

Richar:2015:PCR

[RKCH15] Jonathan I. Richar, Gordon H. Kruse, Enrique Curchitser, and Albert J. Hermann. Patterns in connectivity and retention of simulated Tanner crab (Chionoecetes bairdi) larvae in the eastern Bering Sea. Progress in Oceanography, 138(??):475–485, November 2015. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661114001244.

Reverdin:2007:SSA

[RKFD07] G. Reverdin, E. Kestenare, C. Frankignoul, and T. Delcroix. Surface salinity in the Atlantic Ocean (30^{o}S-50^{o} n). Progress in Oceanography, 73(3–4):311–340, May 2007. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661107000821.

Rohan:2021:UBT

[RKK^{+}21] Sean K. Rohan, Stan Kotwicki, Kelly A. Kearney, Jennifer A. Schulien, Edward A. Laman, Edward D. Cokelet, David A. Beauchamp, Lyle L. Britt, Kerim Y. Aydin, and Stephani G. Zador. Using bottom trawls to monitor subsurface water clarity in marine ecosystems. Progress in Oceanography, 194(??): ??, June 2021. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661121000422.

Roether:2007:TEM

[RKM^{+}07] Wolfgang Roether, Birgit Klein, Beniamino Bruno Manca, Alexander Theocharis, and Sotiris Kioroglou. Transient Eastern Mediterranean deep waters in response to the massive dense-water output of the Aegean Sea in the 1990s. Progress in Oceanography, 74(4):540–571, September 2007. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661107000572.

Radchenko:2001:CWB

[RKS01] V. I. Radchenko, G. V. Khen, and A. M. Slabinsky. Cooling in the western Bering Sea in 1999: quick propagation of La Niña signal or compensatory processes effect? Progress in Oceanography, 49(??):407–422, ???? 2001. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661101000337.

Rudels:2015:CTA

[RKS^{+}15] Bert Rudels, Meri Korhonen, Ursula Schauer, Sergey Pisarev, Benjamin Rabe, and Andreas Wisotzki. Circulation and transformation of Atlantic water in the Eurasian Basin and the contribution of the Fram Strait inflow branch to the Arctic Ocean heat budget. Progress in Oceanography, 132(??):128– 152, March 2015. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661114000494.

Robinson:1985:EPO

[RL85] Allan R. Robinson and Wayne G. Leslie. Estimation and prediction of oceanic eddy fields. Progress in Oceanography, 14(??):485–510, ???? 1985. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0079661185900242.

Ridgway:2023:TDV

[RL23] K. R. Ridgway and S. D. Ling. Three decades of variability and warming of nearshore waters around Tasmania. Progress in Oceanography, 215(??):??, July 2023. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661123000897.

Rossby:1985:ISF

[RLC85] H. T. Rossby, E. R. Levine, and D. N. Connors. The isopycnal swallow float — a simple device for tracking water parcels in the ocean. Progress in Oceanography, 14(??):511–525, ???? 1985. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0079661185900254.

Ramirez-Llodra:2013:ENA

[RLDC^{+}13] Eva Ramirez-Llodra, Ben De Mol, Joan B. Company, Marta Coll, and Francesc Sardà. Effects of natural and anthropogenic processes in the distribution of marine litter in the deep Mediterranean Sea. Progress in Oceanography, 118(??):273–287, November 2013. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661113001407.

Richardson:2010:LAL

[RLGC10] David E. Richardson, Joel K. Llopiz, Cedric M. Guigand, and Robert K. Cowen. Larval assemblages of large and medium-sized pelagic species in the Straits of Florida. Progress in Oceanography, 86(1–2):8–20, July 2010. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S007966111000039X.

Richardson:2009:SIP

[RLL^{+}09] David E. Richardson, Joel K. Llopiz, Kevin D. Leaman, Peter S. Vertes, Frank E. Muller-Karger, and Robert K. Cowen. Sailfish (Istiophorus platypterus) spawning and larval environment in a Florida Current frontal eddy. Progress in Oceanography, 82(4):252–264, September 2009. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S007966110900055X.

Rezek:2018:SFS

[RLP^{+}18] Ryan J. Rezek, Benoit Lebreton, Terence A. Palmer, Gregory W. Stunz, and Jennifer Beseres Pollack. Structural and functional similarity of epibenthic communities on standing and reefed platforms in the northwestern Gulf of Mexico. Progress in Oceanography, 168(??):145–154, November 2018. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661118301630.

Rosli:2018:RSD

[RLR^{+}18] Norliana Rosli, Daniel Leduc, Ashley A. Rowden, P. Keith Probert, and Malcolm R. Clark. Regional and sediment depth differences in nematode community structure greater than between habitats on the New Zealand margin: Implications for vulnerability to anthropogenic disturbance. Progress in Oceanography, 160(??):26–52, January 2018. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661116302683.

Roget:2006:MMN

[RLSF06] Elena Roget, Iossif Lozovatsky, Xavier Sanchez, and Manuel Figueroa. Microstructure measurements in natural waters: Methodology and applications. Progress in Oceanography, 70(2–4):126–148, August 2006. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661106000838.

See corrigendum [RLSF07].

Roget:2007:CMM

[RLSF07] Elena Roget, Iossif Lozovatsky, Xavier Sanchez, and Manuel Figueroa. Corrigendum to Microstructure measurements in natural waters: Methodology and applications [Progress in Oceanography 70 (2006) 126–148]. Progress in Oceanography, 72(1):114, January 2007. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661106001327.

See [RLSF06].

Rathburn:2022:GBH

[RLT^{+}22] A. E. Rathburn, L. A. Levin, M. Tryon, J. M. Gieskes, J. B. Martin, M. E. Pérez, F. J. Fodrie, C. Neira, G. J. Fryer, G. Mendoza, P. A. McMillan, J. Kluesner, J. Adamic, and W. Ziebis. Geological and biological heterogeneity of the Aleutian margin (1965–4822m). Progress in Oceanography, 80(1–2): 22–50, ???? 2022. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661108002073.

Ren:2024:IRM

[RLX^{+}24] Hengye Ren, Wenfang Lu, Wupeng Xiao, Qing Zhu, Canbo Xiao, and Zhigang Lai. Intraseasonal response of marine planktonic ecosystem to summertime Madden–Julian Oscillation in the South China Sea: a model study. Progress in Oceanography, 224(??):??, June 2024. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661124000570.

Roemmich:1989:LSC

[RM89] Dean Roemmich and Tracy McCallister. Large scale circulation of the North Pacific Ocean. Progress in Oceanography, 22(2):171–204, ???? 1989. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0079661189900050.

Radach:1993:EVP

[RM93] Günther Radach and Andreas Moll. Estimation of the variability of production by simulating annual cycles of phytoplankton in the central North Sea. Progress in Oceanography, 31(4):339– 419, ???? 1993. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/007966119390001T.

Ray:1997:SMI

[RM97] Richard D. Ray and Gary T. Mitchum. Surface manifestation of internal tides in the deep ocean: observations from altimetry and island gauges. Progress in Oceanography, 40(??):135– 162, ???? 1997. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661197000256.

Roberts:2001:SDH

[RMB^{+}01] D. Roberts, H. M. Moore, J. Berges, J. W. Patching, M. W. Carton, and D. F. Eardly. Sediment distribution, hydrolytic enzyme profiles and bacterial activities in the guts of Oneirophanta mutabilis, Psychropotes longicauda and Pseudostichopus villosus: what do they tell us about digestive strategies of abyssal holothurians? Progress in Oceanography, 50(??):443– 458, ???? 2001. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661101000659.

Reygondeau:2015:PCW

[RMC^{+}15] Gabriel Reygondeau, Juan Carlos Molinero, Steve Coombs, Brian R. MacKenzie, and Delphine Bonnet. Progressive changes in the Western English Channel foster a reorganization in the plankton food web. Progress in Oceanography, 137(??):524–532, September 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661115000920.

Renaud-Mornant:1990:EAM

[RMG90] Jeanne Renaud-Mornant and Nicole Gourbault. Evaluation of abyssal meiobenthos in the eastern central Pacific (Clarion-Clipperton Fracture Zone). Progress in Oceanography, 24(??):317–329, ???? 1990. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/007966119090041Y.

Rodriguez:2009:IAC

[RMHL09] J. M. Rodriguez, M. Moyano, and S. Hernandez-Leon. The ichthyoplankton assemblage of the canaries–African coastal transition zone: a review. Progress in Oceanography, 83 (1–4):314–321, December 2009. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661109000962.

Renault:2021:EHR

[RMK^{+}21] Lionel Renault, James C. McWilliams, Faycal Kessouri, Alexandre Jousse, Hartmut Frenzel, Ru Chen, and Curtis Deutsch. Evaluation of high-resolution atmospheric and oceanic simulations of the California Current System. Progress in Oceanography, 195(??):??, July 2021. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661121000513.

Ryan:2002:SLR

[RN02] H. F. Ryan and M. Noble. Sea level response to ENSO along the central California coast: how the 1997–1998 event compares with the historic record. Progress in Oceanography, 54(1–4):149–169, July 2002. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661102000472.

Rysgaard:2006:CCH

[RN06] Søren Rysgaard and Torkel Gissel Nielsen. Carbon cycling in a high-Arctic marine ecosystem — Young Sound, NE Greenland. Progress in Oceanography, 71(2–4):426–445, October 2006. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661106001261.

Rey:2019:SLT

[RNBP^{+}19] Felisa Rey, Gina M. Silva Neto, Juan Bueno-Pardo, Regina Bispo, and Ricardo Calado. Supply and larval traits at metamorphosis of a coastal marine invertebrate with a biphasic life cycle under contrasting oceanographic conditions.

Progress in Oceanography, 178(??):??, November 2019. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661118302659.

Ross:2013:DCP

[RNL^{+}13] Peter S. Ross, Marie Noël, Dyanna Lambourn, Neil Dangerfield, John Calambokidis, and Steven Jeffries. Declining concentrations of persistent PCBs, PBDEs, PCDEs, and PCNs in harbor seals (Phoca vitulina) from the Salish Sea. Progress in Oceanography, 115(??):160–170, August 2013. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661113000785.

Rees:2017:AMT

[RNP^{+}17] Andrew P. Rees, Philip D. Nightingale, Alex J. Poulton, Tim J. Smyth, Glen A. Tarran, and Gavin H. Tilstone. The Atlantic Meridional Transect programme (1995–2016). Progress in Oceanography, 158(??):3–18, November 2017. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661117301635.

Ramirez-Ortiz:2022:GRR

[ROBRB^{+}22] Georgina Ramírez-Ortiz, Eduardo F. Balart, Héctor Reyes-Bonilla, Leonardo Huato-Soberanis, Christian Cortés-Fuentes, and Fiorenza Micheli. Greater resilience of reef fish assemblages in a no-take reserve compared to multi-use areas of the Gulf of California. Progress in Oceanography, 204(??): ??, June 2022. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661122000556.

Roe:1984:DMDb

[Roe84a] H. S. J. Roe. The diel migrations and distributions within a Mesopelagic community in the North East Atlantic. 2. Vertical migrations and feeding of mysids and decapod crustacea. Progress in Oceanography, 13(3–4):269–318, ???? 1984. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0079661184900119.

Roe:1984:DMDc

[Roe84b] H. S. J. Roe. The diel migrations and distributions within a mesopelagic community in the North East Atlantic. 4. The copepods. Progress in Oceanography, 13(3–4):353–388, ???? 1984. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0079661184900132.

Rogachev:2000:RVP

[Rog00] Konstantin A. Rogachev. Recent variability in the Pacific western subarctic boundary currents and Sea of Okhotsk. Progress in Oceanography, 47(2–4):299–336, October 2000. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661100000409.

Rooth:1982:HOC

[Roo82] Claes Rooth. Hydrology and ocean circulation. Progress in Oceanography, 11(2):131–149, ???? 1982. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0079661182900064.

Rosenqvist:1965:SHA

[Ros65] I. Th. Rosenqvist. Studies on the hydrogen atoms of layered silicates. Progress in Oceanography, 3(??):297–300, ???? 1965. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/007966116590025X.

Rotschi:1965:PAE

[Rot65] Henri Rotschi. Le pH et l’alcalinité des eaux profondes de la fosse des Hébrides et du bassin des Fidji. (French) [The pH and alkalinity of the deep waters of the Hebrides Trough and the Fiji Basin]. Progress in Oceanography, 3(??):301– 310, ???? 1965. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661165900261.

Rouch:1965:VDT

[Rou65] Jules Rouch. La variation diurne de la température de l’air en hiver dans les regions polaires. (French) [The diurnal variation of air temperature in winter in the polar regions].

Progress in Oceanography, 3(??):311–320, ???? 1965. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0079661165900273.

Riou:2018:BEH

[RPG^{+}18] Virginie Riou, Julien Para, Marc Garel, Catherine Guigue, Badr Al Ali, Chiara Santinelli, Dominique Lefèvre, Jean-Pierre Gattuso, Madeleine Goutx, Stéphanie Jacquet, Frédéric A. C. Le Moigne, Kazuyo Tachikawa, and Christian Tamburini. Biodegradation of Emiliania huxleyi aggregates by a natural Mediterranean prokaryotic community under increasing hydrostatic pressure. Progress in Oceanography, 163(??):271– 281, April 2018. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661116301859.

Roget:2023:DMB

[RPPM^{+}23] Elena Roget, Josep L. Pelegrí, Jesús Planella-Morato, Joan Puigdefàbregas, Mikhail Emelianov, Ignasi Vallès-Casanova, and Dorleta Orúe-Echevarria. Diapycnal mixing in the Brazil– Malvinas confluence front. Progress in Oceanography, 211(??): ??, February 2023. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661123000113.

Rodriguez-Perez:2021:DTN

[RPRCAG^{+}21] Mónica Y. Rodríguez-Pérez, Rocío. I. Ruiz-Cooley, David Aurioles-Gamboa, Laura Sánchez-Velasco, Miguel F. Lavín, and Juan-Pablo Gallo-Reynoso. Deciphering the trophic niche of the nearly extinct vaquita (Phocoena sinus) and its variability through time. Progress in Oceanography, 199(??):??, December 2021. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661121001786.

Ryan:2022:ULS

[RPSC22] John P. Ryan, Paulo S. Polito, Peter G. Strutton, and Francisco P. Chavez. Unusual large-scale phytoplankton blooms in the equatorial Pacific. Progress in Oceanography, 55(3–4):263– 285, ???? 2022. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661102001374.

Ross:2014:SIT

[RPSVLS14] L. Ross, I. Pérez-Santos, A. Valle-Levinson, and W. Schneider. Semidiurnal internal tides in a Patagonian fjord. Progress in Oceanography, 129(??):19–34, December 2014. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661114000329.

Roy:2001:ERM

[RR01] Claude Roy and Chris Reason. ENSO related modulation of coastal upwelling in the eastern Atlantic. Progress in Oceanography, 49(??):245–255, ???? 2001. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661101000258.

Ruddick:2003:OTIb

[RR03] B. Ruddick and K. Richards. Oceanic thermohaline intrusions: observations. Progress in Oceanography, 56(3–4):499– 527, March 2003. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661103000284.

Rohr:2022:RFG

[RRLS22] Tyler Rohr, Anthony J. Richardson, Andrew Lenton, and Elizabeth Shadwick. Recommendations for the formulation of grazing in marine biogeochemical and ecosystem models. Progress in Oceanography, 208(??):??, November 2022. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661122001379.

Richardson:2003:USO

[RRS03] A. J. Richardson, C. Risien, and F. A. Shillington. Using self-organizing maps to identify patterns in satellite imagery. Progress in Oceanography, 59(2–3):223–239, October 2003. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S007966110300171X.

Rothschild:2004:RSF

[RS04] B. J. Rothschild and L. J. Shannon. Regime shifts and fishery management. Progress in Oceanography, 60(2–4):397–402, February 2004. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661104000370.

Rosa:2010:MPH

[RS10] Rui Rosa and Brad A. Seibel. Metabolic physiology of the Humboldt squid, Dosidicus gigas: Implications for vertical migration in a pronounced oxygen minimum zone. Progress in Oceanography, 86(1–2):72–80, July 2010. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661110000388.

Rabouille:2001:ICB

[RSB^{+}01] C. Rabouille, H. Stahl, F. Bassinot, A. Tengberg, J. Brunnegard, P. Hall, K. Kiriakoulakis, J.-L. Reyss, L. Dezileau, P. Crassous, P. Roos, and R. S. Lampitt. Imbalance in the carbonate budget of surficial sediments in the North Atlantic Ocean: variations over the last millenium? Progress in Oceanography, 50(??):201–221, ???? 2001. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661101000544.

Ruzicka:2013:DPW

[RSB^{+}13] James J. Ruzicka, John H. Steele, Tosca Ballerini, Sarah K. Gaichas, and David G. Ainley. Dividing up the pie: Whales, fish, and humans as competitors. Progress in Oceanography, 116(??):207–219, September 2013. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661113001158.

Renaud:2015:FAB

[RSB^{+}15] Paul E. Renaud, Mikael K. Sejr, Bodil A. Bluhm, Boris Sirenko, and Ingrid H. Ellingsen. The future of Arctic benthos: Expansion, invasion, and biodiversity. Progress in Oceanography, 139(??):244–257, December 2015. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661115001597.

Rowe:1990:OCT

[RSD^{+}90] Gilbert T. Rowe, Myriam Sibuet, Jody Deming, John Tietjen, and Alexis Khripounoff. Organic carbon turnover time in deep-sea benthos. Progress in Oceanography, 24(??):141– 160, ???? 1990. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/007966119090026X.

Raick:2006:MCP

[RSG06] C. Raick, K. Soetaert, and M. Grégoire. Model complexity and performance: How far can we simplify? Progress in Oceanography, 70(1):27–57, July 2006. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661106000188.

Roberts:2023:DSE

[RSK^{+}23] K. E. Roberts, J. E. F. Stepanuk, H. Kim, L. H. Thorne, C. Chong-Montenegro, and J. A. Nye. Developing a subseasonal ecological forecast to reduce fisheries bycatch in the Northeast U.S. Progress in Oceanography, 213(??):??, April– May 2023. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661123000642.

Roux:2023:WPD

[RSM^{+}23] Pauline Roux, Mathilde Schapira, Kenneth Neil Mertens, Coralie André, Aouregan Terre-Terrillon, Anne Schmitt, Soazig Manach, Karine Collin, Joelle Serghine, Cyril Noel, and Raffaele Siano. When phytoplankton do not bloom: the case of the dinoflagellate Lepidodinium chlorophorum in southern Brittany (France) assessed by environmental DNA. Progress in Oceanography, 212(??):??, March 2023. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661123000423.

Richardson:2003:DQA

[RSMIS03] A. J. Richardson, N. F. Silulwane, B. A. Mitchell-Innes, and F. A. Shillington. A dynamic quantitative approach for predicting the shape of phytoplankton profiles in the ocean. Progress in Oceanography, 59(2–3):301–319, October 2003. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661103001691.

Richards:2023:BSD

[RSW^{+}23] Travis M. Richards, Tracey T. Sutton, Matthew S. Woodstock, Heather Judkins, and R. J. David Wells. Body size, depth of occurrence, and local oceanography shape trophic structure in a diverse deep-pelagic micronekton assemblage. Progress in Oceanography, 213(??):??, April–May 2023. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661123000411.

Riera:2022:MAD

[RTBR^{+}22] R. Riera, V. M. Tuset, R. Betancur-R., A. Lombarte, C. Marcos, and A. Pérez-Ruzafa. Modelling alpha-diversities of coastal lagoon fish assemblages from the Mediterranean Sea. Progress in Oceanography, 165(??):100–109, ???? 2022. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661116302579.

Ramaiah:2005:EIE

[RTF^{+}05] Neelam Ramaiah, Shigenobu Takeda, Ken Furuya, Takeshi Yoshimura, Jun Nishioka, Tatsuo Aono, Yukihiro Nojiri, Keiri Imai, Isao Kudo, Hiroaki Saito, and Atsushi Tsuda. Effect of iron enrichment on the dynamics of transparent exopolymer particles in the western subarctic Pacific. Progress in Oceanography, 64(2–4):253–261, February 2005. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661105000339.

Rice:1990:DAH

[RTN90] A. L. Rice, M. H. Thurston, and A. L. New. Dense aggregations of a hexactinellid sponge, Pheronema carpenteri, in the Porcupine Seabight (northeast Atlantic Ocean), and possible causes. Progress in Oceanography, 24(??):179–196, ???? 1990. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661190900292.

Rudels:1989:FPS

[Rud89] Bert Rudels. The formation of polar surface water, the ice export and the exchanges through the Fram Strait. Progress in Oceanography, 22(3):205–248, ???? 1989. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/007966118990013X.

Ruddick:2003:LSI

[Rud03] B. Ruddick. Laboratory studies of interleaving. Progress in Oceanography, 56(3–4):529–547, March 2003. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661103000314.

Rudels:2015:AOC

[Rud15] Bert Rudels. Arctic Ocean circulation, processes and water masses: a description of observations and ideas with focus on the period prior to the International Polar Year 2007–2009. Progress in Oceanography, 132(??):22–67, March 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661113002255.

Robke:2017:TP

[RV17] B. R. Röbke and A. Vött. The tsunami phenomenon. Progress in Oceanography, 159(??):296–322, December 2017. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661116302221.

Reynolds:2022:SRE

[RvBD^{+}22] David J. Reynolds, Vanessa R. von Biela, Kenneth H. Dunton, David C. Douglas, and Bryan A. Black. Sclerochronological records of environmental variability and bivalve growth in the Pacific Arctic. Progress in Oceanography, 206(??):??, August 2022. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661122001239.

Romano:2013:SCP

[RVC^{+}13] C. Romano, J. R. Voight, J. B. Company, M. Plyuscheva, and D. Martin. Submarine canyons as the preferred habitat for wood-boring species of Xylophaga (Mollusca, Bivalvia). Progress in Oceanography, 118(??):175–187, November 2013. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661113001419.

Roy:2021:SBB

[RVS^{+}21] Rajdeep Roy, P. N. Vinayachandran, Amit Sarkar, Jenson George, Chandanlal Parida, Aneesh Lotliker, Satya Prakash, and Saroj Bondhu Choudhury. Southern Bay of Bengal: a possible hotspot for CO_{2} emission during the summer monsoon. Progress in Oceanography, 197(??):??, September– October 2021. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661121001233.

Ray:1997:PSI

[RW97] Richard D. Ray and Philip L. Woodworth. Preface to the special issue on tidal science. Progress in Oceanography, 40(??): 1–6, ???? 1997. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661197000207.

Rabouille:2001:AIV

[RWD01] C. Rabouille, R. Witbaard, and G. C. A. Duineveld. Annual and interannual variability of sedimentary recycling studied with a non-steady-state model: application to the North Atlantic Ocean (BENGAL site). Progress in Oceanography, 50(??):147–170, ???? 2001. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661101000520.

Richardson:2006:UCP

[RWJ^{+}06] A. J. Richardson, A. W. Walne, A. W. G. John, T. D. Jonas, J. A. Lindley, D. W. Sims, D. Stevens, and M. Witt. Using continuous plankton recorder data. Progress in Oceanography, 68(1):27–74, January 2006. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661105001424.

Riser:2001:PRE

[RWOA01] Christian Wexels Riser, Paul Wassmann, Kalle Olli, and Elena Arashkevich. Production, retention and export of zooplankton faecal pellets on and off the Iberian shelf, north-west Spain. Progress in Oceanography, 51(2–4):423–441, November 2001. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661101000787.

Rudnick:2017:CCC

[RZTD17] Daniel L. Rudnick, Katherine D. Zaba, Robert E. Todd, and Russ E. Davis. A climatology of the California Current System from a network of underwater gliders. Progress in Oceanography, 154(??):64–106, May 2017. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661116302075.

Robertson:2023:CST

[RZW^{+}23] R. Robertson, C. Zhao, W. Wang, Z. Xu, and Z. Liu. A case study off the Tiwi Islands and the Coburg Peninsula: Baroclinic on one side and barotropic on the other. Progress in Oceanography, 216(??):??, August 2023. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661123001003.

Smith:1986:Ea

[SA86a] Bob Smith and Martin Angel. Editorial. Progress in Oceanography, 16(1):i–ii, ???? 1986. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0079661186900340.

Smith:1986:Eb

[SA86b] Bob Smith and Martin Angel. Editorial. Progress in Oceanography, 17(1–2):i, ???? 1986. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0079661186900236.

Sibuet:1990:E

[SA90] Myriam Sibuet and Martin V. Angel. Editorial. Progress in Oceanography, 24(??):vii–ix, ???? 1990. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/007966119090014S.

Smith:1997:ERO

[SA97] A. J. E. Smith and O. B. Andersen. Errors in recent ocean tide models: possible origin and cause. Progress in Oceanography, 40(??):325–336, ???? 1997. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S007966119800007X.

Seidov:2015:ONW

[SAA^{+}15] D. Seidov, J. I. Antonov, K. M. Arzayus, O. K. Baranova, M. Biddle, T. P. Boyer, D. R. Johnson, A. V. Mishonov, C. Paver, and M. M. Zweng. Oceanography north of 60^{o}N from World Ocean Database. Progress in Oceanography, 132(??):153–173, March 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661114000160.

Skjoldal:2021:SIV

[SAB^{+}21] Hein Rune Skjoldal, Johanna M. Aarflot, Espen Bagøien, Øystein Skagseth, Jon Rønning, and Vidar S. Lien. Seasonal and interannual variability in abundance and population development of Calanus finmarchicus at the western entrance to the Barents Sea, 1995–2019. Progress in Oceanography, 195(??): ??, July 2021. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661121000616.

Stock:2022:UIC

[SAB^{+}22] Charles A. Stock, Michael A. Alexander, Nicholas A. Bond, Keith M. Brander, William W. L. Cheung, Enrique N. Curchitser, Thomas L. Delworth, John P. Dunne, Stephen M. Griffies, Melissa A. Haltuch, Jonathan A. Hare, Anne B. Hollowed, Patrick Lehodey, Simon A. Levin, Jason S. Link, Kenneth A. Rose, Ryan R. Rykaczewski, Jorge L. Sarmiento, Ronald J. Stouffer, Franklin B. Schwing, Gabriel A. Vecchi, and Francisco E. Werner. On the use of IPCC-class models to assess the impact of climate on Living Marine Resources. Progress in Oceanography, 88(1–4):1–27, ????

2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661110001096.

Sachs:2016:FMU

[Sac16] Gottfried Sachs. In-flight measurement of upwind dynamic soaring in albatrosses. Progress in Oceanography, 142(??):47– 57, March 2016. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661116000082.

Sousa:2017:SCU

[SAd^{+}17] Magda Catarina Sousa, Ines Alvarez, Maite deCastro, Moncho Gomez-Gesteira, and João Miguel Dias. Seasonality of coastal upwelling trends under future warming scenarios along the southern limit of the Canary Upwelling System. Progress in Oceanography, 153(??):16–23, April 2017. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S007966111630132X.

Spall:2021:LRH

[SAH^{+}21] Michael A. Spall, Mattia Almansi, Jie Huang, Thomas W. N. Haine, and Robert S. Pickart. Lateral redistribution of heat and salt in the Nordic Seas. Progress in Oceanography, 196(??): ??, August 2021. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661121000963.

Saidova:1965:SSP

[Sai65] H. M. Saidova. Sediment stratigraphy and paleogeography of the Pacific Ocean by benthonic Foraminifera during the quaternary. Progress in Oceanography, 4(??):143–151, September 2, 1965. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661165900479.

Sakai:1986:BBF

[Sak86] Satoshi Sakai. Bifurcation of barotropic flow and homogenization of potential vorticity in a rotating laboratory model. Progress in Oceanography, 17(3–4):389–399, ???? 1986. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/007966118690056X.

Sebastian:2004:APA

[SAM^{+}04] Marta Sebastián, Javier Arístegui, María F. Montero, Jose Escanez, and F. Xavier Niell. Alkaline phosphatase activity and its relationship to inorganic phosphorus in the transition zone of the North-western African upwelling system. Progress in Oceanography, 62(2–4):131–150, August 2004. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661104000965.

Shankar:2010:MIS

[SAM^{+}10] D. Shankar, S. G. Aparna, J. P. McCreary, I. Suresh, S. Neetu, F. Durand, S. S. C. Shenoi, and M. A. Al Saafani. Minima of interannual sea-level variability in the Indian Ocean. Progress in Oceanography, 84(3–4):225–241, March 2010. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661109001712.

Sankey:1973:FDW

[San73] T. Sankey. The formation of deep water in the northwestern Mediterranean. Progress in Oceanography, 6(??):159– 179, IN53–IN54, ???? 1973. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0079661173900074.

Sanson:2015:SDG

[San15] L. Zavala Sansón. Surface dispersion in the Gulf of California. Progress in Oceanography, 137 (part A)(??):24–37, September 2015. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661115000683.

Suga:2008:VNP

[SASH08] Toshio Suga, Yoshikazu Aoki, Hiroko Saito, and Kimio Hanawa. Ventilation of the North Pacific subtropical pycnocline and mode water formation. Progress in Oceanography, 77(4):285–297, June 2008. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661108000645.

Smith:2022:SLV

[SAT^{+}22] Walker O. Smith, Vernon Asper, Sasha Tozzi, Xiao Liu, and Sharon E. Stammerjohn. Surface layer variability in the Ross Sea, Antarctica as assessed by in situ fluorescence measurements. Progress in Oceanography, 88(1–4):28–45, ???? 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661110001084.

Smyth:2015:WCO

[SAW^{+}15] Tim Smyth, Angus Atkinson, Stephen Widdicombe, Matthew Frost, Icarus Allen, James Fishwick, Ana Queiros, David Sims, and Manuel Barange. The Western Channel Observatory. Progress in Oceanography, 137(??):335–341, September 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661115001172.

Silva:2016:PHS

[SAY^{+}16] Claudio Silva, Isabel Andrade, Eleuterio Yáñez, Samuel Hormazabal, María Ángela Barbieri, Antonio Aranis, and Gabriela Böhm. Predicting habitat suitability and geographic distribution of anchovy (Engraulis ringens) due to climate change in the coastal areas off Chile. Progress in Oceanography, 146(??):159–174, August 2016. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661115300495.

Stewart:1969:EI

[SB69] R. W. Stewart and T. Braarud. Editorial introduction. Progress in Oceanography, 5(??):vii–xi, ???? 1969. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0079661169900226.

StJohn:2014:IBS

[SBB^{+}14] M. A. St. John, M. Barange, H. Benway, K. J. Flynn, J. Holt, G. Merino, A. Martin, A. Mitra, W. Melle, R. Sanders, V. M. Trenkel, I. Grigorov, and E. Hoffman. Introduction to the BASIN special issue: State of art, past present a view to the future. Progress in Oceanography, 129(??):171–175, December 2014. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661114001943.

Soviadan:2022:PMC

[SBB^{+}22] Yawouvi Dodji Soviadan, Fabio Benedetti, Manoela C. Brandão, Sakina-Dorothée Ayata, Jean-Olivier Irisson, Jean Louis Jamet, Rainer Kiko, Fabien Lombard, Kissao Gnandi, and Lars Stemmann. Patterns of mesozooplankton community composition and vertical fluxes in the global ocean. Progress in Oceanography, 200(??):??, January 2022. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661121002007.

Stramski:2004:RSC

[SBBV04] Dariusz Stramski, Emmanuel Boss, Darek Bogucki, and Kenneth J. Voss. The role of seawater constituents in light backscattering in the ocean. Progress in Oceanography, 61(1): 27–56, April 2004. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661104000692.

Sardenne:2016:TNS

[SBC^{+}16] Fany Sardenne, Nathalie Bodin, Emmanuel Chassot, Aurélien Amiel, Edwin Fouché, Maxime Degroote, Stéphanie Hollanda, Heidi Pethybridge, Benoit Lebreton, Gaël Guillou, and Frédéric Ménard. Trophic niches of sympatric tropical tuna in the Western Indian Ocean inferred by stable isotopes and neutral fatty acids. Progress in Oceanography, 146(??):75–88, August 2016. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661115300641.

Supic:2024:WCI

[SBC^{+}24] Nastjenjka Supić, Andrea Budiša, Irena Ciglenečki, Milan Čanković, Jelena Dautović, Tamara Djakovac, Natalija Dunić, Mathieu Dutour-Sikirić, Ingrid Ivančić, Matea Kalac, Romina Kraus, Nataša Kužat, Davor Lučić, Daniela Marić Pfannkuchen, Boris Mifka, Hrvoje Mihanović, Jakica Njire, Paolo Paliaga, Miroslava Pasarić, Zoran Pasarić, Niki Simonović, Maja Telišman Prtenjak, and Ivica Vilibić. Are winter conditions impacting annual organic production in the northern Adriatic? Verifications and future projections. Progress in Oceanography, 224(??):??, June 2024. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661124000533.

Stammer:2001:RVW

[SBD01] Detlef Stammer, Claus Böning, and Christian Dieterich. The role of variable wind forcing in generating eddy energy in the North Atlantic. Progress in Oceanography, 48(??):289–311, ???? 2001. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661101000088.

Sobarzo:2007:CUS

[SBD^{+}07] Marcus Sobarzo, Luis Bravo, David Donoso, José Garcés-Vargas, and Wolfgang Schneider. Coastal upwelling and seasonal cycles that influence the water column over the continental shelf off central Chile. Progress in Oceanography, 75(3):363–382, November 2007. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661107001577.

Strand:2020:SDS

[SBE^{+}20] Espen Strand, Espen Bagøien, Martin Edwards, Cecilie Broms, and Thor Klevjer. Spatial distributions and seasonality of four Calanus species in the Northeast Atlantic. Progress in Oceanography, 185(??):??, June 2020. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661120300835.

Shaw:2021:CAE

[SBFP21] C. Tracy Shaw, Hongsheng Bi, Leah R. Feinberg, and William T. Peterson. Cohort analysis of Euphausia pacifica from the Northeast Pacific population using a Gaussian mixture model. Progress in Oceanography, 191(??):??, February 2021. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661120302305.

Swartzman:2008:RAS

[SBG^{+}08] Gordon Swartzman, Arnaud Bertrand, Mariano Gutiérrez, Sophie Bertrand, and Luis Vasquez. The relationship of anchovy and sardine to water masses in the Peruvian Humboldt Current System from 1983 to 2005. Progress in Oceanography, 79(2–4):228–237, October 2008. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661108001766.

Stefanoudis:2016:AHI

[SBG16] Paris V. Stefanoudis, Brian J. Bett, and Andrew J. Gooday. Abyssal hills: Influence of topography on benthic foraminiferal assemblages. Progress in Oceanography, 148(??):44–55, November 2016. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S007966111630026X.

Seki:2014:ACT

[SBH^{+}14] Osamu Seki, James A. Bendle, Naomi Harada, Madoka Kobayashi, Ken Sawada, Heiko Moossen, Gordon N. Inglis, Seiya Nagao, and Tatsuhiko Sakamoto. Assessment and calibration of TEX_{86} paleothermometry in the Sea of Okhotsk and sub-polar North Pacific region: Implications for paleoceanography. Progress in Oceanography, 126(??):254–266, August 2014. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661114000615.

Schlosser:1995:MDT

[SBK^{+}95] Peter Schlosser, Gerhard Bönisch, Bernd Kromer, H. Hugo Loosli, Renedikt Bühler, Reinhold Bayer, Georges Bonani, and Klaus Peter Koltermann. Mid-1980s distribution of tritium, ^{3}He, ^{14}C and ^{39}Ar in the Greenland/ Norwegian Seas and the Nansen Basin of the Arctic Ocean. Progress in Oceanography, 35(1):1–28, ???? 1995. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/007966119500003Y.

Saulnier:2023:TSA

[SBL^{+}23] Erwan Saulnier, Anik Brind’Amour, Jean-Baptiste Lecomte, Eloïse Piette-Semeril, and Verena M. Trenkel. Temporal synchrony among juvenile marine fishes and potential climate and environmental drivers in the Bay of Biscay. Progress in Oceanography, 211(??):??, February 2023. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661123000125.

Sinha:2010:CEB

[SBLA10] Bablu Sinha, Erik T. Buitenhuis, Corinne Le Quéré, and Thomas R. Anderson. Comparison of the emergent behavior of a complex ecosystem model in two ocean general circulation models. Progress in Oceanography, 84(3–4):204–224, March 2010. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661109001724.

Smith:1991:USI

[SBM91] Neville R. Smith, Joseph E. Blomley, and Gary Meyers. A univariate statistical interpolation scheme for subsurface thermal analyses in the tropical oceans. Progress in Oceanography, 28(3):219–256, ???? 1991. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/007966119190009B.

Smith:2023:PCC

[SBM^{+}23] James A. Smith, Mercedes Pozo Buil, Barbara Muhling, Desiree Tommasi, Stephanie Brodie, Timothy H. Frawley, Jerome Fiechter, Stefan Koenigstein, Amber Himes-Cornell, Michael A. Alexander, Steven J. Bograd, Nathalí Cordero Quirós, Larry B. Crowder, Enrique Curchitser, Stephanie J. Green, Natasha A. Hardy, Alan C. Haynie, Elliott L. Hazen, Kirstin Holsman, Gwendal Le Fol, Nerea Lezama-Ochoa, Ryan R. Rykaczewski, Charles A. Stock, Stephen Stohs, Jonathan Sweeney, Heather Welch, and Michael G. Jacox. Projecting climate change impacts from physics to fisheries: a view from three California Current fisheries. Progress in Oceanography, 211(??):??, February 2023. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661123000162.

Saint-Beat:2018:UID

[SBMB18] Blanche Saint-Béat, Frédéric Maps, and Marcel Babin. Unraveling the intricate dynamics of planktonic Arctic marine food webs. a sensitivity analysis of a well-documented food web model. Progress in Oceanography, 160(??):167–185, January 2018. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S007966111730246X.

Sirviente:2023:DAD

[SBPGP^{+}23] S. Sirviente, M. Bolado-Penagos, J. J. Gomiz-Pascual, J. Romero-Cózar, A. Vázquez, and M Bruno. Dynamics of atmospheric-driven surface currents on the Gulf of Cadiz Continental Shelf and its link with the Strait of Gibraltar and the Western Alboran Sea. Progress in Oceanography, 219(??):??, December 2023. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661123002185.

Svavarsson:1990:DDP

[SBS90] Jörundur Svavarsson, Torleiv Brattegard, and Jarl-Ove Strömberg. Distribution and diversity patterns of asellote isopods (crustacea) in the deep Norwegian and Greenland Seas. Progress in Oceanography, 24(??):297–310, ???? 1990. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0079661190900395.

Shepard:1965:CDS

[SC65] F. P. Shepard and J. R. Curray. Carbon-14 determination of sea level changes in stable areas. Progress in Oceanography, 4(??):283–291, September 2, 1965. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0079661165900558.

Sherwood:1990:SGC

[SC90] Christopher R. Sherwood and Joe S. Creager. Sedimentary geology of the Columbia River Estuary. Progress in Oceanography, 25(1–4):15–79, ???? 1990. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/007966119090003K.

Simons:2023:RSC

[SC23] Rachel D. Simons and Dylan Catlett. Regulation of a surface chlorophyll hotspot by wind-driven upwelling and eddy circulation in the Santa Barbara Channel, Southern California. Progress in Oceanography, 217(??):??, September–October 2023. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661123001398.

Saiz:2007:FPZ

[SCAA07] Enric Saiz, Albert Calbet, Dacha Atienza, and Miquel Alcaraz. Feeding and production of zooplankton in the Catalan Sea (NW Mediterranean). Progress in Oceanography, 74(2–3):313– 328, August 2007. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661107000626.

Steele:2007:BEE

[SCB^{+}07] J. H. Steele, J. S. Collie, J. J. Bisagni, D. J. Gifford, M. J. Fogarty, J. S. Link, B. K. Sullivan, M. E. Sieracki, A. R. Beet, D. G. Mountain, E. G. Durbin, D. Palka, and W. T. Stockhausen. Balancing end-to-end budgets of the Georges Bank ecosystem. Progress in Oceanography, 74(4):423–448, September 2007. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661107001000.

Sarda:2009:RBE

[SCB^{+}09] F. Sardà, J. B. Company, N. Bahamón, G. Rotllant, M. M. Flexas, J. D. Sánchez, D. Zúñiga, J. Coenjaerts, D. Orellana, G. Jordà, J. Puigdefábregas, A. Sánchez-Vidal, A. Calafat, D. Martín, and M. Espino. Relationship between environment and the occurrence of the deep-water rose shrimp Aristeus antennatus (Risso, 1816) in the Blanes submarine canyon (NW Mediterranean). Progress in Oceanography, 82(4):227–238, September 2009. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661109000536.

Sun:2016:SCB

[SCB^{+}16] Yunfang Sun, Changsheng Chen, Robert C. Beardsley, Dave Ullman, Bradford Butman, and Huichan Lin. Surface circulation in Block Island Sound and adjacent coastal and shelf regions: a FVCOM–CODAR comparison. Progress in Oceanography, 143(??):26–45, April 2016. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661116000203.

Seo:2014:VHT

[SCC14] Gwang-Ho Seo, Yang-Ki Cho, and Byoung-Ju Choi. Variations of heat transport in the northwestern Pacific marginal seas inferred from high-resolution reanalysis. Progress in Oceanography, 121(??):98–108, February 2014. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661113001912.

Sion:2019:DBC

[SCC^{+}19] Letizia Sion, Crescenza Calculli, Francesca Capezzuto, Roberto Carlucci, Angela Carluccio, Laura Cornacchia, Porzia Maiorano, Alessio Pollice, Pasquale Ricci, Angelo Tursi, and Gianfranco D’Onghia. Does the Bari Canyon (Central Mediterranean) influence the fish distribution and abundance? Progress in Oceanography, 170(??):81–92, January 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661118300892.

Sempere:2018:ODA

[SCCJ^{+}18] Richard Sempéré, Bruno Charrière, Javier Castro-Jiménez, Kimitaka Kawamura, and Christos Panagiotopoulos. Occurrence of α, ω-dicarboxylic acids and ω-oxoacids in surface waters of the Rhone River and fluxes into the Mediterranean Sea. Progress in Oceanography, 163(??):136–146, April 2018. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661117300514.

Santos:2007:PBI

[SCD^{+}07] A. Miguel P. Santos, Alexandra Chícharo, Antonina Dos Santos, Teresa Moita, Paulo B. Oliveira, Álvaro Peliz, and Pedro Ré. Physical–biological interactions in the life history of small pelagic fish in the western Iberia upwelling ecosystem. Progress in Oceanography, 74(2–3):192–209, August 2007. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661107000675.

Schott:1983:MRS

[Sch83] Friedrich Schott. Monsoon response of the Somali Current and associated upwelling. Progress in Oceanography, 12(3):357– 381, ???? 1983. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661183900149.

Schmitt:2003:OLI

[Sch03] R. W. Schmitt. Observational and laboratory insights into salt finger convection. Progress in Oceanography, 56(3–4):419– 433, March 2003. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661103000338.

Sanchez-Cabeza:2022:RSW

[SCHBC^{+}22] Joan-Albert Sanchez-Cabeza, Carlos Alberto Herrera-Becerril, José Luis Carballo, Benjamín Yáñez, León Felipe Álvarez-Sánchez, José-Gilberto Cardoso-Mohedano, and Ana Carolina Ruiz-Fernández. Rapid surface water warming and impact of the recent (2013–2016) temperature anomaly in shallow coastal waters at the eastern entrance of the Gulf of California. Progress in Oceanography, 202(??):??, March 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661122000088.

Silve:2023:SLH

[SCHD23] Violette Silve, Henrique Cabral, Martin Huret, and Hilaire Drouineau. Sensitivity to life-history parameters in larval fish drift modelling predictions for contrasting climatic conditions. Progress in Oceanography, 217(??):??, September– October 2023. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661123001453.

Sabbagh:2024:VLD

[SCHS^{+}24] Eman I. Sabbagh, Maria Ll. Calleja, Tamara M. Huete-Stauffer, Daniele Daffonchio, and Xosé Anxelu G. Morán. Vertical and latitudinal distribution of bottom-up and topdown controls of heterotrophic prokaryotes in the Red Sea. Progress in Oceanography, 224(??):??, June 2024. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661124000661.

Sanchez-Cabeza:2011:TNP

[SCLG^{+}11] J. A. Sanchez-Cabeza, I. Levy, J. Gastaud, M. Eriksson, I. Osvath, M. Aoyama, P. P. Povinec, and K. Komura. Transport of North Pacific ^{137}Cs labeled waters to the south-eastern Atlantic Ocean. Progress in Oceanography, 89(1–4):31–37, April 2011. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661110001576.

Solidoro:2010:RRS

[SCLS10] C. Solidoro, G. Cossarini, S. Libralato, and S. Salon. Remarks on the redefinition of system boundaries and model parameterization for downscaling experiments. Progress in Oceanography, 84(1–2):134–137, January 2010. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661109001505.

Sanchez-Cabeza:1999:SAR

[SCMAR^{+}99] J. A. Sanchez-Cabeza, P. Masqué, I. Ani-Ragolta, J. Merino, M. Frignani, F. Alvisi, A. Palanques, and P. Puig. Sediment accumulation rates in the southern Barcelona continental margin (NW Mediterranean Sea) derived from ^{210}Pb and 137Cs chronology. Progress in Oceanography, 44(1–3):313–332, August 1999. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661199000312.

Sannino:2015:RTF

[SCPN15] G. Sannino, A. Carillo, G. Pisacane, and C. Naranjo. On the relevance of tidal forcing in modelling the Mediterranean thermohaline circulation. Progress in Oceanography, 134(??):304– 329, May 2015. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661115000427.

Sutcliffe:1987:ESL

[SCS87] A. J. Sutcliffe, A. P. Currant, and C. B. Stringer. Evidence of sea-level change from coastal caves with raised beach deposits, terrestrial faunas and dated stalagmites. Progress in Oceanography, 18(1–4):243–271, ???? 1987. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0079661187900358.

Saunders:2018:SVP

[SCS^{+}18] Ryan A. Saunders, Martin A. Collins, Rachael Shreeve, Peter Ward, Gabriele Stowasser, Simeon L. Hill, and Geraint A. Tarling. Seasonal variation in the predatory impact of myctophids on zooplankton in the Scotia Sea (Southern Ocean). Progress in Oceanography, 168(??):123–144, November 2018. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661117302926.

Stavrakakis:2000:DFS

[SCT^{+}00] S. Stavrakakis, G. Chronis, A. Tselepides, S. Heussner, A. Monaco, and A. Abassi. Downward fluxes of settling particles in the deep Cretan Sea (NE Mediterranean). Progress in Oceanography, 46(2–4):217–240, August 2000. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661100000203.

Sun:2023:ODG

[SCY^{+}23] Yutong Sun, Dan Cheng, Shaoqiong Yang, Chenyi Luo, Di Tian, Han Zhang, and Yanhui Wang. Optimal design of a glider array for the observation of tropical cyclones. Progress in Oceanography, 217(??):??, September– October 2023. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661123001465.

Sundby:2007:MBS

[SD07] Svein Sundby and Kenneth Drinkwater. On the mechanisms behind salinity anomaly signals of the northern North Atlantic. Progress in Oceanography, 73(2):190–202, May 2007. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661107000390.

Schlegel:2021:MCS

[SDB^{+}21] Robert W. Schlegel, Sofia Darmaraki, Jessica A. Benthuysen, Karen Filbee-Dexter, and Eric C. J. Oliver. Marine cold-spells. Progress in Oceanography, 198(??):??, November 2021. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661121001683.

Schneider:2017:WCC

[SDGVE17] Wolfgang Schneider, David Donoso, José Garcés-Vargas, and Rubén Escribano. Water-column cooling and sea surface salinity increase in the upwelling region off central-south Chile driven by a poleward displacement of the South Pacific high. Progress in Oceanography, 151(??):38–48, February 2017. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661116300325.

Skaret:2014:CFA

[SDH^{+}14] G. Skaret, P. Dalpadado, S. S. Hjøllo, M. D. Skogen, and E. Strand. Calanus finmarchicus abundance, production and population dynamics in the Barents Sea in a future climate. Progress in Oceanography, 125(??):26–39, June 2014. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661114000561.

Stock:2014:GSC

[SDJ14] Charles A. Stock, John P. Dunne, and Jasmin G. John. Global-scale carbon and energy flows through the marine planktonic food web: an analysis with a coupled physical– biological model. Progress in Oceanography, 120(??):1–28, January 2014. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661113001079.

Simpson:1984:OECa

[SDK84] J. J. Simpson, T. D. Dickey, and C. J. Koblinsky. An offshore eddy in the California Current System Part I: Interior dynamics. Progress in Oceanography, 13(1):5–49, ???? 1984. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0079661184900041.

Swalethorp:2019:MDA

[SDL^{+}19] Rasmus Swalethorp, Julie Dinasquet, Ramiro Logares, Stefan Bertilsson, Sanne Kjellerup, Anders K. Krabberød, Per-Olav Moksnes, Torkel G. Nielsen, and Lasse Riemann. Microzooplankton distribution in the Amundsen Sea Polynya (Antarctica) during an extensive Phaeocystis antarctica bloom. Progress in Oceanography, 170(??):1–10, January 2019. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661118300491.

Stige:2014:SSA

[SDO^{+}14] Leif Christian Stige, Padmini Dalpadado, Emma Orlova, Anne-Cécile Boulay, Joël M. Durant, Geir Ottersen, and Nils Chr. Stenseth. Spatiotemporal statistical analyses reveal predator-driven zooplankton fluctuations in the Barents Sea. Progress in Oceanography, 120(??):243–253, January 2014. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661113001845.

Sarma:2022:ERT

[SDP^{+}22] V. V. S. S. Sarma, D. V. Desai, J. S. Patil, L. Khandeparker, S. G. Aparna, D. Shankar, Selrina D’Souza, H. B. Dalabehera, J. Mukherjee, P. Sudharani, and A. C. Anil. Ecosystem response in temperature fronts in the northeastern Arabian Sea. Progress in Oceanography, 165(??):317–331, ???? 2022. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661117301647.

Schott:2002:SOC

[SDS02] Friedrich A. Schott, Marcus Dengler, and Rena Schoenefeldt. The shallow overturning circulation of the Indian Ocean. Progress in Oceanography, 53(1):57–103, April 2002. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661102000393. See erratum [SDS22a].

Schott:2022:ESO

[SDS22a] Friedrich A. Schott, Marcus Dengler, and Rena Schoenefeldt. Erratum to: The shallow overturning circulation of the Indian Ocean: [Progress in Oceanography (2002) 53(1), 57–103] Footnote: PII of the original article: S0079-6611(02)00039-3: FOotnote: Friedrich A. Schott, Marcus Dengler, Rena Schoenefeldt. Progress in Oceanography, 55(3–4):373–384, ???? 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661102001179. See [SDS02].

Stukel:2022:RSV

[SDS^{+}22b] Michael R. Stukel, Moira Décima, Karen E. Selph, Darcy A. A. Taniguchi, and Michael R. Landry. The role of Synechococcus in vertical flux in the Costa Rica upwelling dome. Progress in Oceanography, 112–113(??):49–59, ???? 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661113000426.

Small:1992:FCP

[SE92] Lawrence F. Small and Steven G. Ellis. Fecal carbon production by zooplankton in Santa Monica Basin: the effects of body size and carnivorous feeding. Progress in Oceanography, 30(1–4):197–221, ???? 1992. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/007966119290013P.

Sijp:2008:ENS

[SE08] Willem P. Sijp and Matthew H. England. The effect of a northward shift in the southern hemisphere westerlies on the global ocean. Progress in Oceanography, 79(1):1–19, October 2008. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661108001304. See corrigendum

[SE09].

Sijp:2009:CEN

[SE09] Willem P. Sijp and Matthew H. England. Corrigendum to “The effect of a northward shift in the southern hemisphere westerlies on the global ocean” [Prog Oceanogr 79 (2008) 1–19]. Progress in Oceanography, 82(3):225, September 2009. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661109000524. See [SE08].

Schmidt:2016:OCN

[SE16] Martin Schmidt and Anja Eggert. Oxygen cycling in the northern Benguela Upwelling System: Modelling oxygen sources and sinks. Progress in Oceanography, 149(??):145–173, December 2016. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661116301975.

Sears:1963:P

[Sea63] Mary Sears. Preface. Progress in Oceanography, 1(??):vii, ???? 1963. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661163900028.

Segerstraale:1969:BFB

[Seg69] Sven G. Segerstråle. Biological fluctuations in the Baltic Sea. Progress in Oceanography, 5(??):169–184, ???? 1969. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0079661169900391.

Skjoldal:2022:SFZ

[SEG22a] Hein Rune Skjoldal, Elena Eriksen, and Harald Gjøsæter. Size-fractioned zooplankton biomass in the Barents Sea: Spatial patterns and temporal variations during three decades of warming and strong fluctuations of the capelin stock (1989–2020). Progress in Oceanography, 206(??):??, August 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661122001112.

Skjoldal:2022:RVF

[SEG^{+}22b] Hein Rune Skjoldal, Elena Eriksen, Harald Gjøsæter, Øystein Skagseth, Dmitry Prozorkevich, and Vidar S. Lien. Recruitment variability of fish stocks in the Barents Sea: Spatial and temporal variation in 0-group fish length of six commercial species during recent decades of warming (1980–2017). Progress in Oceanography, 206(??):??, August 2022. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661122001057.

Seibold:1963:GIN

[Sei63] Eugen Seibold. Geological investigation of nearshore sandtransport — examples of methods and problems from the Baltic and North Seas. Progress in Oceanography, 1(??):3– 6, IN1, 7–22, IN3, 23–28, IN5–IN6, 29–70, ???? 1963. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/007966116390003X.

Sekine:1986:WDC

[Sek86] Yoshihiko Sekine. Wind-driven circulation in the Japan Sea and its influence on the branching of the Tsushima Current. Progress in Oceanography, 17(3–4):297–312, ???? 1986. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0079661186900510.

Sekine:1988:SVO

[Sek88] Yoshihiko Sekine. On the seasonal variation in in- and outflow volume transport of the Japan Sea. Progress in Oceanography, 21(3–4):269–279, ???? 1988. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0079661188900079.

Sekine:1999:VSC

[Sek99] Yoshihiko Sekine. On variations in the subarctic circulation in the North Pacific. Progress in Oceanography, 43(2–4):193– 203, March 1999. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661199000099.

Selli:1965:PPB

[Sel65] Raimondo Selli. The Pliocene–Pleistocene boundary in Italian marine sections and its relationship to continental stratigraphies. Progress in Oceanography, 4(??):67–86, September 2, 1965. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661165900418.

Sharples:2013:FOS

[SENS13] Jonathan Sharples, Jim R. Ellis, Glenn Nolan, and Beth E. Scott. Fishing and the oceanography of a stratified shelf sea. Progress in Oceanography, 117(??):130–139, October 2013. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S007966111300102X.

Sutherland:2013:RBE

[SEO13] Terri F. Sutherland, Robert W. Elner, and Jennifer D. O’Neill. Roberts Bank: Ecological crucible of the Fraser River estuary. Progress in Oceanography, 115(??):171–180, August 2013. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661113000608.

Slagstad:2011:EPS

[SEW11] D. Slagstad, I. H. Ellingsen, and P. Wassmann. Evaluating primary and secondary production in an Arctic Ocean void of summer sea ice: an experimental simulation approach. Progress in Oceanography, 90(1–4):117–131, July 2011. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661111000243.

Saunders:1985:SHS

[SF85] P. M. Saunders and T. J. G. Francis. The search for hydrothermal sources on the Mid-Atlantic Ridge. Progress in Oceanography, 14(??):527–536, ???? 1985. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0079661185900266.

Satterfield:2002:SIA

[SF02] Franklin R. Satterfield and Bruce P. Finney. Stable isotope analysis of Pacific salmon: insight into trophic status and oceanographic conditions over the last 30 years. Progress in Oceanography, 53(??):231–246, ???? 2002. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661102000320.

Sevellec:2015:OEA

[SF15] Florian Sévellec and Alexey V. Fedorov. Optimal excitation of AMOC decadal variability: Links to the subpolar ocean. Progress in Oceanography, 132(??):287–304, March 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661114000196.

Sibuet:1990:SEW

[SFAD^{+}90] Myriam Sibuet, Luc Floury, Anne-Marie Alayse-Danet, André Echardour, Tanguy LeMoign, and Roger Perron. In situ experimentation at the water/sediment interface in the deep sea: 1. Submersible experimental instrumentation developed for sampling and incubation. Progress in Oceanography, 24(??):161– 167, ???? 1990. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/007966119090027Y.

Skakala:2024:HUO

[SFF^{+}24] Jozef Skákala, David Ford, Alison Fowler, Dan Lea, Matthew J. Martin, and Stefano Ciavatta. How uncertain and observable are marine ecosystem indicators in shelf seas? Progress in Oceanography, 224(??):??, June 2024. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661124000557.

Send:1999:RAO

[SFK^{+}99] Uwe Send, Jordi Font, Gerd Krahmann, Claude Millot, Monika Rhein, and Joaquin Tintoré. Recent advances in observing the physical oceanography of the western Mediterranean Sea. Progress in Oceanography, 44(1–3):37–64, August 1999. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661199000208.

Santana-Falcon:2020:OTO

[SFMA20] Yeray Santana-Falcón, Evan Mason, and Javier Arístegui. Offshore transport of organic carbon by upwelling filaments in the Canary Current System. Progress in Oceanography, 186(??): ??, July 2020. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661120300616.

Schloesser:2012:DAM

[SFMT12] Fabian Schloesser, Ryo Furue, Julian P. McCreary, and Axel Timmermann. Dynamics of the Atlantic meridional overturning circulation. Part 1: Buoyancy-forced response. Progress in Oceanography, 101(1):33–62, August 2012. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661112000031.

Schloesser:2014:DAM

[SFMT14] F. Schloesser, R. Furue, J. P. McCreary, and A. Timmermann. Dynamics of the Atlantic meridional overturning circulation. Part 2: Forcing by winds and buoyancy. Progress in Oceanography, 120(??):154–176, January 2014. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661113001675.

Santora:2012:SEK

[SFS^{+}12] Jarrod A. Santora, John C. Field, Isaac D. Schroeder, Keith M. Sakuma, Brian K. Wells, and William J. Sydeman. Spatial ecology of krill, micronekton and top predators in the central California Current: Implications for defining ecologically important areas. Progress in Oceanography, 106(??):154–174, November 2012. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661112000948.

Silver:1991:PFO

[SG91] Mary W. Silver and Marcia M. Gowing. The “particle” flux: Origins and biological components. Progress in Oceanography, 26(1):75–113, ???? 1991. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0079661191900079.

Santin:2019:DPD

[SGA^{+}19] A. Santín, J. Grinyó, S. Ambroso, M. J. Uriz, C. Dominguez-Carrió, and J. M. Gili. Distribution patterns and demographic trends of demosponges at the Menorca Channel (Northwestern Mediterranean Sea). Progress in Oceanography, 173(??): 9–25, April 2019. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S007966111830199X.

Salvatteci:2019:FDS

[SGF^{+}19] Renato Salvatteci, Dimitri Gutierrez, David Field, Abdel Sifeddine, Luc Ortlieb, Sandrine Caquineau, Tim Baumgartner, Vicente Ferreira, and Arnaud Bertrand. Fish debris in sediments from the last 25 kyr in the Humboldt Current reveal the role of productivity and oxygen on small pelagic fishes. Progress in Oceanography, 176(??):??, September 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S007966111830243X.

Salvado:2013:DOC

[SGL^{+}13] Joan A. Salvadó, Joan O. Grimalt, Jordi F. López, Xavier Durrieu de Madron, Catalina Pasqual, and Miquel Canals. Distribution of organochlorine compounds in superficial sediments from the Gulf of Lion, northwestern Mediterranean Sea. Progress in Oceanography, 118(??):235–248, November 2013. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661113001274.

Salvado:2017:TLM

[SGL^{+}17] Joan A. Salvadó, Joan O. Grimalt, Jordi F. López, Albert Palanques, Serge Heussner, Catalina Pasqual, Anna Sanchez-Vidal, and Miquel Canals. Transfer of lipid molecules and polycyclic aromatic hydrocarbons to open marine waters by dense water cascading events. Progress in Oceanography, 159(??):178–194, December 2017. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661117301428.

Saavedra:2018:AES

[SGL^{+}18] Camilo Saavedra, Tim Gerrodette, Maite Louzao, Julio Valeiras, Salvador García, Santiago Cerviño, Graham J. Pierce, and M. Begoña Santos. Assessing the environmental status of the short-beaked common dolphin (Delphinus delphis) in North-western Spanish waters using abundance trends and safe removal limits. Progress in Oceanography, 166(??):66–75, September 2018. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661117300836.

Sanchez-Garrido:2013:WDC

[SGLF^{+}13] José C. Sánchez-Garrido, Jesús García Lafuente, Enrique Álvarez Fanjul, Marcos García Sotillo, and Francisco J. de los Santos. What does cause the collapse of the Western Alboran Gyre? Results of an operational ocean model. Progress in Oceanography, 116(??):142–153, September 2013. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661113001080.

Solic:2018:STR

[ŠGM^{+}18] Mladen Šolić, Branka Grbec, Frano Matić, Danijela Šantić, Stefanija Šestanović, Živana Ninčević Gladan, Natalia Bojanić, Marin Ordulj, Slaven Jozić, and Ana Vrdoljak. Spatiotemporal reproducibility of the microbial food web structure associated with the change in temperature: Long-term observations in the Adriatic Sea. Progress in Oceanography, 161(??):87–101, February 2018. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661117301192.

Segade:2015:QST

[SGMP15] Lidia Isabel Carracedo Segade, Miguel Gilcoto, Herlé Mercier, and Fiz Fernández Pérez. Quasi-synoptic transport, budgets and water mass transformation in the Azores–Gibraltar Strait region during summer 2009. Progress in Oceanography, 130(??):47–64, January 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661114001530.

Santiago-Garcia:2014:TDC

[SGMVF14] M. W. Santiago-García, S. G. Marinone, and O. U. Velasco-Fuentes. Three-dimensional connectivity in the Gulf of California based on a numerical model. Progress in Oceanography, 123(??):64–73, April 2014. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661114000159.

Sifeddine:2008:LSC

[SGO^{+}08] A. Sifeddine, D. Gutiérrez, L. Ortlieb, H. Boucher, F. Velazco, D. Field, G. Vargas, M. Boussafir, R. Salvatteci, V. Ferreira, M. García, J. Valdés, S. Caquineau, M. Mandeng Yogo, F. Cetin, J. Solis, P. Soler, and T. Baumgartner. Laminated sediments from the central Peruvian continental slope: a 500 year record of upwelling system productivity, terrestrial runoff and redox conditions. Progress in Oceanography, 79(2–4):190–197, October 2008. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661108001729.

Sempere:2018:PSI

[SGPdM18] Richard Sempéré, Cécile Guieu, Ivane Pairaud, and Xavier Durrieu de Madron. Preface of special issue of MERMEX project: Recent advances in the oceanography of the Mediterranean Sea. Progress in Oceanography, 163(??):1–6, April 2018. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661118300284.

Sudheesh:2022:OMZ

[SGR^{+}22] V. Sudheesh, G. V. M. Gupta, Yudhishtir Reddy, Kausar F. Bepari, N. V. H. K. Chari, C. K. Sherin, S. S. Shaju, Ch. V. Ramu, and Anil Kumar Vijayan. Oxygen minimum zone along the eastern Arabian Sea: Intra-annual variation and dynamics based on ship-borne studies. Progress in Oceanography, 201(??):??, February 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661122000040.

Sherin:2023:NNS

[SGS^{+}23] C. K. Sherin, G. V. M. Gupta, V. Sudheesh, Ch.V. Ramu, Bikram Reddy, N. V. Harikrishnachari, and Anil Kumar Vijayan. Nutriclines and nutrient stoichiometry in the eastern Arabian Sea: Intra-annual variations and controlling mechanisms. Progress in Oceanography, 215(??):??, July 2023. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661123000915.

Sanchez-Garrido:2019:DSV

[SGWF^{+}19] J. C. Sánchez-Garrido, F. E. Werner, J. Fiechter, K. A. Rose, E. N. Curchitser, A. Ramos, J. García Lafuente, J. Arístegui, S. Hernández-León, and A. Rodríguez Santana. Decadalscale variability of sardine and anchovy simulated with an end-to-end coupled model of the Canary Current ecosystem. Progress in Oceanography, 171(??):212–230, February 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661118300387.

Shields:2009:LSV

[SH09] Mark A. Shields and David J. Hughes. Large-scale variation in macrofaunal communities along the eastern Nordic Seas continental margin: a comparison of four stations with contrasting food supply. Progress in Oceanography, 82(2):125–136, August 2009. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661109000482.

Shackleton:1982:DSS

[Sha82] N. J. Shackleton. The deep-sea sediment record of climate variability. Progress in Oceanography, 11(2):199–218, ???? 1982. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0079661182900088.

Soreide:2006:SFW

[SHC^{+}06] Janne E. Søreide, Haakon Hop, Michael L. Carroll, Stig Falk-Petersen, and Else Nøst Hegseth. Seasonal food web structures and sympagic–pelagic coupling in the European Arctic revealed by stable isotopes and a two-source food web model. Progress in Oceanography, 71(1):59–87, October 2006. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661106000784. See corrigendum

[SHC^{+}07].

Soreide:2007:CSF

[SHC^{+}07] Janne E. Søreide, Haakon Hop, Michael L. Carroll, Stig Falk-Petersen, and Else Nøst Hegseth. Corrigendum to “Seasonal food web structures and sympagic–pelagic coupling in the European Arctic revealed by stable isotopes and a two-source food web model” [Progress in Oceanography 71 (2006), 59–87]. Progress in Oceanography, 73(1):96–98, April 2007. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661107000158. See [SHC^{+}06].

Santinelli:2013:ISM

[SHd13] C. Santinelli, D. A. Hansell, and M. Ribera d’Alcalà. Influence of stratification on marine dissolved organic carbon (DOC) dynamics: the Mediterranean Sea case. Progress in Oceanography, 119(??):68–77, December 2013. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661113000827.

Schukat:2021:ZET

[SHD^{+}21] Anna Schukat, Wilhelm Hagen, Sabrina Dorschner, Jonathan Correa Acosta, Elda Luz Pinedo Arteaga, Patricia Ayón, and Holger Auel. Zooplankton ecological traits maximize the trophic transfer efficiency of the Humboldt Current upwelling system. Progress in Oceanography, 193(??):??, April–May 2021. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661121000392.

Shepard:1965:ISV

[She65] Francis P. Shepard. Importance of submarine valleys in funneling sediments to the deep sea. Progress in Oceanography, 3(??):321–332, ???? 1965. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0079661165900285.

Smith:2001:COO

[SHF01] Robert L. Smith, Adriana Huyer, and Jane Fleischbein. The coastal ocean off Oregon from 1961 to 2000: is there evidence of climate change or only of Los Niños? Progress in Oceanography, 49(??):63–93, ???? 2001. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661101000167.

Steele:2012:EEM

[SHG12] John H. Steele, Eileen E. Hofmann, and Dian J. Gifford. End-to-end models: Management applications. Progress in Oceanography, 102(??):1–4, September 2012. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661112000201.

Sanders:2014:BCP

[SHK^{+}14] Richard Sanders, Stephanie A. Henson, Marja Koski, Christina L. De La Rocha, Stuart C. Painter, Alex J. Poulton, Jennifer Riley, Baris Salihoglu, Andre Visser, Andrew Yool, Richard Bellerby, and Adrian P. Martin. The Biological Carbon Pump in the North Atlantic. Progress in Oceanography, 129(??):200–218, December 2014. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661114000901.

Schmoker:2013:SEP

[SHL13] Claire Schmoker and Santiago Hernández-León. Stratification effects on the plankton of the subtropical Canary Current. Progress in Oceanography, 119(??):24–31, December 2013. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661113001663.

Sandven:2023:SDW

[SHP^{+}23] Håkon Sandven, Børge Hamre, Tristan Petit, Rüdiger Röttgers, Hongbo Liu, and Mats A. Granskog. Seasonality and drivers of water column optical properties on the northwestern Barents Sea shelf. Progress in Oceanography, 217(??):??, September–October 2023. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661123001192.

Suzuki:2005:RPH

[SHS^{+}05] Koji Suzuki, Akira Hinuma, Hiroaki Saito, Hiroshi Kiyosawa, Hongbin Liu, Toshiro Saino, and Atsushi Tsuda. Responses of phytoplankton and heterotrophic bacteria in the northwest subarctic Pacific to in situ iron fertilization as estimated by HPLC pigment analysis and flow cytometry. Progress in Oceanography, 64(2–4):167–187, February 2005. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661105000285.

Sydeman:2001:CCR

[SHT^{+}01] William J. Sydeman, Michelle M. Hester, Julie A. Thayer, Franklin Gress, Paige Martin, and Joelle Buffa. Climate change, reproductive performance and diet composition of marine birds in the southern California Current system, 1969–1997. Progress in Oceanography, 49(??):309–329, ???? 2001. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661101000283.

Sur:1997:ESD

[SI97] Halil I. Sur and Yuri P. Ilyin. Evolution of satellite derived mesoscale thermal patterns in the Black Sea. Progress in Oceanography, 39(2):109–151, ???? 1997. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661197000098.

Suryan:2006:SSE

[SIB^{+}06] Robert M. Suryan, David B. Irons, Evelyn D. Brown, Patrick G. R. Jodice, and Daniel D. Roby. Site-specific effects on productivity of an upper trophic-level marine predator: Bottomup, top-down, and mismatch effects on reproduction in a colonial seabird. Progress in Oceanography, 68(2–4):303–328, February 2006. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661106000115.

Siedler:1969:FSD

[Sie69] G. Siedler. On the fine structure of density and current distribution and its short-time variations in different areas. Progress in Oceanography, 5(??):81–94, ???? 1969. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0079661169900305.

Sieburth:1988:TRB

[Sie88] John McN. Sieburth. The trophic roles of bacteria in marine ecosystems are complicated by synergistic-consortia and mixotrophic-cometabolism: the plenary lecture in memory of Professor Sebastiano Genovese. Progress in Oceanography, 21(2):117–128, ???? 1988. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0079661188900304.

Simpson:1981:CRC

[Sim81] W. R. Simpson. A critical review of cadmium in the marine environment. Progress in Oceanography, 10(1):1–70, ???? 1981. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0079661181900070.

Simpson:1984:OECb

[Sim84] J. J. Simpson. An offshore eddy in the California Current System part III: Chemical structure. Progress in Oceanography, 13(1):71–93, ???? 1984. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0079661184900065.

Saunders:2007:RVD

[SIR^{+}07] R. A. Saunders, A. Ingvarsdóttir, J. Rasmussen, S. J. Hay, and A. S. Brierley. Regional variation in distribution pattern, population structure and growth rates of Meganyctiphanes norvegica and Thysanoessa longicaudata in the Irminger Sea, North Atlantic. Progress in Oceanography, 72(4):313–342, March 2007. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661106001273.

Sohrin:2014:DDO

[SIS^{+}14] Rumi Sohrin, Kunimatsu Imanishi, Yoshimi Suzuki, Kenshi Kuma, Ichiro Yasuda, Koji Suzuki, and Takeshi Nakatsuka. Distributions of dissolved organic carbon and nitrogen in the western Okhotsk Sea and their effluxes to the North Pacific. Progress in Oceanography, 126(??):168–179, August 2014. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661114001001.

Shiomoto:2002:IVP

[SiSI^{+}02] Akihiro Shiomoto, Sei ichi Saitoh, Keiri Imai, Mitsuhiro Toratani, Yukimasa Ishida, and Kosei Sasaoka. Interannual variation in phytoplankton biomass in the Bering Sea basin in the 1990s. Progress in Oceanography, 55(1–2):147–163, October 2002. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661102000757.

Strub:2002:ONS

[SJ02a] P. T. Strub and C. James. The 1997–1998 oceanic El Niño signal along the southeast and northeast Pacific boundaries — an altimetric view. Progress in Oceanography, 54(1–4):439– 458, July 2002. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661102000630.

Strub:2002:ADSb

[SJ02b] P. T. Strub and C. James. Altimeter-derived surface circulation in the large-scale NE Pacific Gyres.:Part 2: 1997–1998 El Niño anomalies. Progress in Oceanography, 53(??):185– 214, ???? 2002. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661102000307.

Strub:2002:ADSa

[SJ02c] P. Ted Strub and Corinne James. Altimeter-derived surface circulation in the large–scale NE Pacific Gyres.:Part 1. seasonal variability. Progress in Oceanography, 53(??):163–183, ???? 2002. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661102000290.

Stewart:2018:HHH

[SJ18] Heather A. Stewart and Alan J. Jamieson. Habitat heterogeneity of hadal trenches: Considerations and implications for future studies. Progress in Oceanography, 161(??):47–65, February 2018. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661117302768.

Sarmiento:2023:SOC

[SJA^{+}23] Jorge L. Sarmiento, Kenneth S. Johnson, Lionel A. Arteaga, Seth M. Bushinsky, Heidi M. Cullen, Alison R. Gray, Roberta M. Hotinski, Tanya L. Maurer, Matthew R. Mazloff, Stephen C. Riser, Joellen L. Russell, Oscar M. Schofield, and Lynne D. Talley. The Southern Ocean carbon and climate observations and modeling (SOCCOM) project: a review. Progress in Oceanography, 219(??):??, December 2023. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661123001738.

Stegert:2010:IPO

[SJD10] Christoph Stegert, Rubao Ji, and Cabell S. Davis. Influence of projected ocean warming on population growth potential in two North Atlantic copepod species. Progress in Oceanography, 87(1–4):264–276, October 2010. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661110001254.

Sherwood:1990:HCC

[SJH^{+}90] Christopher R. Sherwood, David A. Jay, R. Bradford Harvey, Peter Hamilton, and Charles A. Simenstad. Historical changes in the Columbia River Estuary. Progress in Oceanography, 25(1–4):299–352, ???? 1990. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/007966119090011P.

Shin:2003:IFQ

[SJJ^{+}03] Kyoungsoon Shin, Min-Chul Jang, Pung-Kuk Jang, Se-Jong Ju, Tea-Kyun Lee, and Man Chang. Influence of food quality on egg production and viability of the marine planktonic copepod Acartia omorii. Progress in Oceanography, 57(3–4):265– 277, June 2003. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661103001010.

Sun:2023:CGM

[SJLW23] Di Sun, Zhao Jing, Furong Li, and Lixin Wu. Characterizing global marine heatwaves under a spatio-temporal framework. Progress in Oceanography, 211(??):??, February 2023. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661122002063.

Strub:2019:OCA

[SJM^{+}19] P. Ted Strub, Corinne James, Vivian Montecino, José A. Rutllant, and José Luis Blanco. Ocean circulation along the southern Chile transition region (38^{o}–46^{o}S): Mean, seasonal and interannual variability, with a focus on 2014– 2016. Progress in Oceanography, 172(??):159–198, March 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661118301551.

Shannon:2010:DSB

[SJP10] Lynne J. Shannon, Astrid C. Jarre, and Samantha L. Petersen. Developing a science base for implementation of the ecosystem approach to fisheries in South Africa. Progress in Oceanography, 87(1–4):289–303, October 2010. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661110001205.

Stramski:1991:LSM

[SK91] Dariusz Stramski and Dale A. Kiefer. Light scattering by microorganisms in the open ocean. Progress in Oceanography, 28(4):343–383, ???? 1991. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/007966119190032H.

Solberg:2017:DVM

[SK17] Ingrid Solberg and Stein Kaartvedt. The diel vertical migration patterns and individual swimming behavior of overwintering sprat Sprattus sprattus. Progress in Oceanography, 151(??):49– 61, February 2017. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661116301197.

Smith:2018:CTS

[SK18] Walker O. Smith and Daniel E. Kaufman. Climatological temporal and spatial distributions of nutrients and particulate matter in the Ross Sea. Progress in Oceanography, 168 (??):182–195, November 2018. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661118301940.

Souvermezoglou:2021:TEO

[SK21] Ekaterini Souvermezoglou and Evangelia Krasakopoulou. Temporal evolution of oxygen and nutrients in the NE Ionian Sea from 1987 to 2008 under the influence of the Cretan and Adriatic Seas. Progress in Oceanography, 199(??):??, December 2021. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661121001944.

Sunanda:2023:SPP

[SKCP23] N. Sunanda, J. Kuttippurath, A. Chakraborty, and R. Peter. Stressors of primary productivity in the north Indian Ocean revealed by satellite, reanalysis and CMIP6 data. Progress in Oceanography, 219(??):??, December 2023. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661123002070.

Shinohara:2020:DSF

[SKF20] Gento Shinohara, Shigeaki Kojima, and Hiroaki Fukumori. Deep-sea fishes from the Kurile–Kamchatka Trench collected by the R/V Sonne during the KuramBio II expedition in 2016. Progress in Oceanography, 183(??):??, April 2020. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661120300136.

Sobocinski:2020:EIM

[SKGS20] Kathryn L. Sobocinski, Neala W. Kendall, Correigh M. Greene, and Michael W. Schmidt. Ecosystem indicators of marine survival in Puget Sound steelhead trout. Progress in Oceanography, 188(??):??, October 2020. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661120301580.

Suga:2000:NPT

[SKH00] Toshio Suga, Ayato Kato, and Kimio Hanawa. North Pacific Tropical Water: its climatology and temporal changes associated with the climate regime shift in the 1970s. Progress in Oceanography, 47(2–4):223–256, October 2000. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661100000379.

Sakamoto:2023:GSO

[SKH^{+}23] Tatsuya Sakamoto, Taketoshi Kodama, Sachiko Horii, Kazutaka Takahashi, Atsushi Tawa, Yosuke Tanaka, and Seiji Ohshmio. Geographic, seasonal and ontogenetic variations of δ ^{15}N and δ ^{13}C of Japanese sardine explained by baseline variations and diverse fish movements. Progress in Oceanography, 219(??):??, December 2023. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661123002069.

Simpson:1984:OSC

[SKHD84] J. J. Simpson, C. J. Koblinsky, L. R. Haury, and T. D. Dickey. An offshore study in the California Current System preface. Progress in Oceanography, 13(1):1–4, ???? 1984. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/007966118490003X.

Souvermezoglou:1999:TVO

[SKP99] Ekaterini Souvermezoglou, Evangelia Krasakopoulou, and Alexandra Pavlidou. Temporal variability in oxygen and nutrient concentrations in the southern Aegean Sea and the Straits of the Cretan Arc. Progress in Oceanography, 44(4):573–600, December 1999. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661199000452.

Summerhayes:1995:VBC

[SKRM^{+}95] C. P. Summerhayes, D. Kroon, A. Rosell-Melé, R. W. Jordan, H-J. Schrader, R. Hearn, J. Villanueva, J. O. Grimalt, and G. Eglinton. Variability in the Benguela Current upwelling system over the past 70,000 years. Progress in Oceanography, 35(3):207–251, ???? 1995. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0079661195000085.

Seo:2006:VSG

[SKSK06] Hyunju Seo, Suam Kim, Kibeik Seong, and Sukyung Kang. Variability in scale growth rates of chum salmon (Oncorhynchus keta) in relation to climate changes in the late 1980s. Progress in Oceanography, 68(2–4):205–216, February 2006. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661106000061.

Sugimoto:2001:INE

[SKT01] Takashige Sugimoto, Shingo Kimura, and Kazuaki Tadokoro. Impact of El Niño events and climate regime shift on living resources in the western North Pacific. Progress in Oceanography, 49(??):113–127, ???? 2001. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661101000180.

Schourup-Kristensen:2018:AOB

[SKWWGV18] Vibe Schourup-Kristensen, Claudia Wekerle, Dieter A. Wolf-Gladrow, and Christoph Völker. Arctic Ocean biogeochemistry in the high resolution FESOM 1.4-REcoM2 model. Progress in Oceanography, 168(??):65–81, November 2018. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661117303634.

Sutherland:2013:QNI

[SL13] T. F. Sutherland and C. D. Levings. Quantifying nonindigenous species in accumulated ballast slurry residuals (swish) arriving at Vancouver, British Columbia. Progress in Oceanography, 115(??):211–218, August 2013. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661113000578.

Simon-Lledo:2019:MVA

[SLBH^{+}19] Erik Simon-Lledó, Brian J. Bett, Veerle A. I. Huvenne, Timm Schoening, Noelie M. A. Benoist, Rachel M. Jeffreys, Jennifer M. Durden, and Daniel O. B. Jones. Megafaunal variation in the abyssal landscape of the clarion clipperton zone. Progress in Oceanography, 170(??):119–133, January 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661118301824.

Sadaoui:2018:IRC

[SLBR18] Mahrez Sadaoui, Wolfgang Ludwig, François Bourrin, and Estela Romero. The impact of reservoir construction on riverine sediment and carbon fluxes to the Mediterranean Sea. Progress in Oceanography, 163(??):94–111, April 2018. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661117300605.

Sarmiento-Lezcano:2022:AFS

[SLBVRR^{+}22] Airam N. Sarmiento-Lezcano, Geraldine Busquets-Vass, Uriel Rubio-Rodríguez, M. Pilar Olivar, Marian Peña, Ione Medina-Suárez, Eduardo González-Rodríguez, Jaime Gómez-Gutiérrez, Carlos J. Robinson, and Santiago Hernández-León. Active flux seasonality of the small dominant migratory crustaceans and mesopelagic fishes in the Gulf of California during June and October. Progress in Oceanography, 208(??):??, November 2022. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661122001537.

Sanchez:2012:SMU

[SLG^{+}12] Gloria E. Sánchez, Carina B. Lange, Humberto E. González, Gabriel Vargas, Praxedes Muñoz, Carolina Cisternas, and Silvio Pantoja. Siliceous microorganisms in the upwelling center off Concepción, Chile (36^{o}S): Preservation in surface sediments and downcore fluctuations during the past ≈ 150 years. Progress in Oceanography, 92–95(??):50–65, January 2012. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661111000681.

Snoeijs-Leijonmalm:2021:DSL

[SLGI^{+}21] Pauline Snoeijs-Leijonmalm, Harald Gjøsæter, Randi B. Ingvaldsen, Tor Knutsen, Rolf Korneliussen, Egil Ona, Hein Rune Skjoldal, Christian Stranne, Larry Mayer, Martin Jakobsson, and Katarina Gårdfeldt. A deep scattering layer under the North pole pack ice. Progress in Oceanography, 194(??):??, June 2021. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661121000471.

Suchy:2019:IED

[SLH^{+}19] Karyn D. Suchy, Nicole Le Baron, Andrea Hilborn, R. Ian Perry, and Maycira Costa. Influence of environmental drivers on spatio-temporal dynamics of satellite-derived chlorophyll-a in the Strait of Georgia. Progress in Oceanography, 176(??):??, September 2019. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661118303136.

Saldias:2016:SMI

[SLM^{+}16] Gonzalo S. Saldías, John L. Largier, Renato Mendes, Iván Pérez-Santos, Cristian A. Vargas, and Marcus Sobarzo. Satellite-measured interannual variability of turbid river plumes off central-southern Chile: Spatial patterns and the influence of climate variability. Progress in Oceanography, 146(??):212–222, August 2016. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661116300210.

Sarmiento-Lezcano:2022:CRS

[SLOP^{+}22] Airam N. Sarmiento-Lezcano, M. Pilar Olivar, Marian Peña, José M. Landeira, Laia Armengol, Ione Medina-Suárez, Arturo Castellón, and Santiago Hernández-León. Carbon remineralization by small mesopelagic and bathypelagic stomiiforms in the Northeast Atlantic Ocean. Progress in Oceanography, 203(??):??, April–May 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661122000489.

Simon-Lledo:2020:MSV

[SLPA^{+}20] Erik Simon-Lledó, Christina Pomee, Akesa Ahokava, Jeffrey C. Drazen, Astrid B. Leitner, Adrian Flynn, John Parianos, and Daniel O. B. Jones. Multi-scale variations in invertebrate and fish megafauna in the mid-eastern clarion clipperton zone. Progress in Oceanography, 187(??):??, August– September 2020. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661120301440.

Stige:2015:CSM

[SLY^{+}15] Leif Chr. Stige, Øystein Langangen, Natalia A. Yaragina, Frode B. Vikebø, Bjarte Bogstad, Geir Ottersen, Nils Chr. Stenseth, and Dag Ø. Hjermann. Combined statistical and mechanistic modelling suggests food and temperature effects on survival of early life stages of Northeast Arctic cod (Gadus morhua). Progress in Oceanography, 134(??):138–151, May 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661115000117.

Smith:1965:MFZ

[SM65] Stuart M. Smith and H. W. Menard. The Molokai fracture zone. Progress in Oceanography, 3(??):333–345, ???? 1965. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0079661165900297.

Schott:2001:MCI

[SM01] Friedrich A. Schott and Julian P. McCreary. The monsoon circulation of the Indian Ocean. Progress in Oceanography, 51(1): 1–123, ???? 2001. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661101000830.

Smith:2005:MAS

[SM05] S. L. Smith and M. Madhupratap. Mesozooplankton of the Arabian Sea: Patterns influenced by seasons, upwelling, and oxygen concentrations. Progress in Oceanography, 65 (2–4):214–239, May 2005. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661105000455.

Stevenson:2016:ENP

[SM16] Angela Stevenson and Fraser J. G. Mitchell. Evidence of nutrient partitioning in coexisting deep-sea echinoids, and seasonal dietary shifts in seasonal breeders: Perspectives from stable isotope analyses. Progress in Oceanography, 141(??):44–59, February 2016. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661115002505.

Saveliev:2021:SCD

[SM21] Pavel A. Saveliev and Evgeny A. Metelyov. Species composition and distribution of eelpouts (Zoarcidae, Perciformes, Actinopterygii) in the northwestern Sea of Okhotsk in summer. Progress in Oceanography, 196(??):??, August 2021. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661121000926.

Smayda:2010:ASHa

[Sma10a] T. J. Smayda. Adaptations and selection of harmful and other dinoflagellate species in upwelling systems. 1. Morphology and adaptive polymorphism. Progress in Oceanography, 85(1–2): 53–70, April 2010. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661110000157.

Smayda:2010:ASHb

[Sma10b] T. J. Smayda. Adaptations and selection of harmful and other dinoflagellate species in upwelling systems. 2. Motility and migratory behaviour. Progress in Oceanography, 85(1–2):71– 91, April 2010. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661110000169.

Sieburth:1988:BPS

[SMB88] John McN. Sieburth, Kenneth T. Morse, and Vivia Bruni. The bibliography of Professor Sebastiano Genovese, in tribute of his contribution to marine microbiology and biological oceanography. Progress in Oceanography, 21(2):109–115, ???? 1988. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661188900298.

Schwing:2002:EOA

[SMdG02] F. B. Schwing, T. Murphree, L. deWitt, and P. M. Green. The evolution of oceanic and atmospheric anomalies in the northeast Pacific during the El Niño and La Niña events of 1995–2001. Progress in Oceanography, 54(1–4):459–491, July 2002. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661102000642.

Smethie:1993:TTC

[Sme93] William M. Smethie. Tracing the thermohaline circulation in the western North Atlantic using chlorofluorocarbons. Progress in Oceanography, 31(1):51–99, ???? 1993. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0079661193900237.

Saiki:2021:MIB

[SMFM^{+}21] Ryu Saiki, Humio Mitsudera, Ayumi Fujisaki-Manome, Noriaki Kimura, Jinro Ukita, Takenobu Toyota, and Tomohiro Nakamura. Mechanism of ice-band pattern formation caused by resonant interaction between sea ice and internal waves in a continuously stratified ocean. Progress in Oceanography, 190(??):??, January 2021. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661120302093.

Schwing:2002:NOI

[SMG02] F. B. Schwing, T. Murphree, and P. M. Green. The Northern Oscillation Index (NOI): a new climate index for the northeast Pacific. Progress in Oceanography, 53(??):115–139, ???? 2002. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661102000277.

Smyth:2001:RSS

[SMGL01] Tim J. Smyth, Peter I. Miller, Steve B. Groom, and Samantha J. Lavender. Remote sensing of sea surface temperature and chlorophyll during Lagrangian experiments at the Iberian margin. Progress in Oceanography, 51(2–4):269–281, November 2001. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661101000702.

Smith:2005:ASN

[Smi05] S. L. Smith. The Arabian Sea of the 1990s: New biogeochemical understanding. Progress in Oceanography, 65(2–4):113– 115, May 2005. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S007966110500039X.

Stammer:2021:HWD

[SMKK21] D. Stammer, M. Sena Martins, J. Köhler, and A. Köhl. How well do we know ocean salinity and its changes? Progress in Oceanography, 190(??):??, January 2021. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661120302135.

Small:1990:PPP

[SMM^{+}90] Lawrence F. Small, C. David McIntire, Keith B. MacDonald, J. Ruben Lara-Lara, Bruce E. Frey, Michael C. Amspoker, and Ted Winfield. Primary production, plant and detrital biomass, and particle transport in the Columbia River Estuary. Progress in Oceanography, 25(1–4):175–210, ???? 1990. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/007966119090007O.

Sano:2013:FHM

[SMN^{+}13] Masayoshi Sano, Koh Maki, Yuichiro Nishibe, Toshi Nagata, and Shuhei Nishida. Feeding habits of mesopelagic copepods in Sagami Bay: Insights from integrative analysis. Progress in Oceanography, 110(??):11–26, March 2013. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661113000050.

Seki:2014:LPB

[SMN^{+}14] O. Seki, Y. Mikami, S. Nagao, J. A. Bendle, T. Nakatsuka, V. I. Kim, V. P. Shesterkin, A. N. Makinov, M. Fukushima, H. M. Moossen, and S. Schouten. Lignin phenols and BIT index distributions in the Amur River and the Sea of Okhotsk: Implications for the source and transport of particulate terrestrial organic matter to the ocean. Progress in Oceanography, 126(??):146–154, August 2014. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661114000883.

Sultan:2007:IML

[SMP07] E. Sultan, H. Mercier, and R. T. Pollard. An inverse model of the large scale circulation in the South Indian Ocean. Progress in Oceanography, 74(1):71–94, July 2007. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661107000407.

Seguro:2022:HRN

[SMP^{+}22a] Isabel Seguro, Alina D. Marca, Suzanne J. Painting, Jamie D. Shutler, David J. Suggett, and Jan Kaiser. High-resolution net and gross biological production during a Celtic Sea spring bloom. Progress in Oceanography, 177(??):??, October 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661117300812.

Sharples:2022:SSB

[SMP^{+}22b] Jonathan Sharples, Daniel J. Mayor, Alex J. Poulton, Andrew P. Rees, and Carol Robinson. Shelf Sea Biogeochemistry: Nutrient and carbon cycling in a temperate shelf sea water column. Progress in Oceanography, 177(??):??, October 2022. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661119303696.

Soto-Mendoza:2012:MTS

[SMPC^{+}12] Samuel Soto-Mendoza, Carolina Parada, Leonardo Castro, Francois Colas, and Wolfgang Schneider. Modeling transport and survival of anchoveta eggs and yolk–sac larvae in the coastal zone off central-southern Chile: Assessing spatial and temporal spawning parameters. Progress in Oceanography, 92–95(??):178–191, January 2012. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661111000553.

Satterthwaite:2020:SSO

[SMR^{+}20] Erin V. Satterthwaite, Steven G. Morgan, John P. Ryan, Julio B. J. Harvey, and Robert C. Vrijenhoek. Seasonal and synoptic oceanographic changes influence the larval biodiversity of a retentive upwelling shadow. Progress in Oceanography, 182(??): ??, March 2020. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661119304410.

Saranya:2024:SET

[SN24] J. S. Saranya and SungHyun Nam. Subsurface evolution of three types of surface marine heatwaves over the East Sea (Japan Sea). Progress in Oceanography, 222(??):??, March 2024. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661124000326.

Souza-Neto:2024:GRG

[SNdSR^{+}24] Pedro W. M. Souza-Neto, Ilson C. A. da Silveira, Cesar B. Rocha, Cauê Z. Lazaneo, and Paulo H. R. Calil. The Rio Grande Rise circulation: Dynamics of an internal tide conversion hotspot in the Southwestern Atlantic. Progress in Oceanography, 224(??):??, June 2024. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661124000703.

Stabeno:2010:FIP

[SNMW10] Phyllis Stabeno, Jeffrey Napp, Calvin Mordy, and Terry Whitledge. Factors influencing physical structure and lower trophic levels of the eastern Bering Sea shelf in 2005: Sea ice, tides and winds. Progress in Oceanography, 85(3–4):180– 196, June 2010. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661110000303.

Sartimbul:2010:VCC

[SNR^{+}10] Aida Sartimbul, Hideaki Nakata, Erfan Rohadi, Beni Yusuf, and Hanggar Prasetyo Kadarisman. Variations in chlorophylla concentration and the impact on Sardinella lemuru catches in Bali Strait, Indonesia. Progress in Oceanography, 87 (1–4):168–174, October 2010. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661110001102.

Swapna:2022:ISD

[SNS^{+}22] M. Swapna, R. K. Nayak, T. Santhoshi, M. V. R. Sesha Sai, and S. S. Rajashekhar. INSAT-3D SST and its diurnal variability assessment using in-situ and MODIS observations. Progress in Oceanography, 201(??):??, February 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661122000015.

Santos:2018:SSP

[SNV^{+}18] A. M. P. Santos, A.-E. Nieblas, P. Verley, A. Teles-Machado, S. Bonhommeau, C. Lett, S. Garrido, and A. Peliz. Sardine (Sardina pilchardus) larval dispersal in the Iberian upwelling system, using coupled biophysical techniques. Progress in Oceanography, 162(??):83–97, March 2018. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661117300125.

Shi:2020:IAS

[SNZ^{+}20] Yongqiang Shi, Mingxiang Niu, Tao Zuo, Jun Wang, Qingshan Luan, Jianqiang Sun, Wei Yuan, Xiujuan Shan, and Evgeny A. Pakhomov. Inter-annual and seasonal variations in zooplankton community structure in the Yellow Sea with possible influence of climatic variability. Progress in Oceanography, 185(??): ??, June 2020. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661120300884.

Smedstad:1991:VDA

[SO91] Ole Martin Smedstad and James J. O’Brien. Variational data assimilation and parameter estimation in an equatorial Pacific Ocean model. Progress in Oceanography, 26(2):179–241, ???? 1991. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661191900024.

Suchy:2023:SRV

[SOA^{+}23] Karyn D. Suchy, Elise Olson, Susan E. Allen, Moira Galbraith, BethElLee Herrmann, Julie E. Keister, R. Ian Perry, Akash R. Sastri, and Kelly Young. Seasonal and regional variability of model-based zooplankton biomass in the Salish Sea and evaluation against observations. Progress in Oceanography, 219(??):??, December 2023. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661123002148.

Schiller:2008:ERO

[SOB^{+}08] A. Schiller, P. R. Oke, G. Brassington, M. Entel, R. Fiedler, D. A. Griffin, and J. V. Mansbridge. Eddy-resolving ocean circulation in the Asian–Australian region inferred from an ocean reanalysis effort. Progress in Oceanography, 76(3):334– 365, March 2008. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661108000074.

Soh:2003:TPD

[Soh03] Wonn Soh. Transport processes deduced from geochemistry and the void ratio of surface core samples, deep sea Sagami Bay, central Japan. Progress in Oceanography, 57(1):109– 124, April 2003. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661103000545.

Seki:2021:ALC

[SOH21] Osamu Seki, Yusuke Okazaki, and Naomi Harada. Assessment of long-chain n-alkanes as a paleoclimate proxy in the Bering Sea sediments. Progress in Oceanography, 198(??):??, November 2021. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661121001713.

Sokolova:1990:SDS

[Sok90] N. N. Sokolova. On size of the deep-sea macrobenthic invertebrates. Progress in Oceanography, 24(??):251–252, ???? 1990. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/007966119090034Y.

Soltwedel:2000:MMA

[Sol00] T. Soltwedel. Metazoan meiobenthos along continental margins: a review. Progress in Oceanography, 46(1):59–84, May 2000. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661100000306.

Skogseth:2020:VDT

[SON^{+}20] R. Skogseth, L. L. A. Olivier, F. Nilsen, E. Falck, N. Fraser, V. Tverberg, A. B. Ledang, A. Vader, M. O. Jonassen, J. Søreide, F. Cottier, J. Berge, B. V. Ivanov, and S. Falk-Petersen. Variability and decadal trends in the Isfjorden (Svalbard) ocean climate and circulation — an indicator for climate change in the European Arctic. Progress in Oceanography, 187(??):??, August–September 2020. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661120301336.

Simizu:2014:WDS

[SOO^{+}14] Daisuke Simizu, Kay I. Ohshima, Jun Ono, Yasushi Fukamachi, and Genta Mizuta. What drives the southward drift of sea ice in the Sea of Okhotsk? Progress in Oceanography, 126(??):33–43, August 2014. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661114000998.

Sabates:2007:PBP

[SOS^{+}07] A. Sabatés, M. P. Olivar, J. Salat, I. Palomera, and F. Alemany. Physical and biological processes controlling the distribution of fish larvae in the NW Mediterranean. Progress in Oceanography, 74(2–3):355–376, August 2007. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661107000638.

Sournia:1994:PBF

[Sou94a] Alain Sournia. Pelagic biogeography and fronts. Progress in Oceanography, 34(2–3):109–120, ???? 1994. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0079661194900043.

Sur:1994:BCI

[SÖÜ94b] Halil İ Sur, Emin Özsoy, and Ümit Ünlüata. Boundary current instabilities, upwelling, shelf mixing and eutrophication processes in the Black Sea. Progress in Oceanography, 33(4):249– 302, ???? 1994. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661194900205.

Schlegel:2017:NOC

[SOWS17] Robert W. Schlegel, Eric C. J. Oliver, Thomas Wernberg, and Albertus J. Smit. Nearshore and offshore cooccurrence of marine heatwaves and cold-spells. Progress in Oceanography, 151(??):189–205, February 2017. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661116300830.

Sutherland:2008:EGC

[SP08] David A. Sutherland and Robert S. Pickart. The East Greenland coastal current: Structure, variability, and forcing. Progress in Oceanography, 78(1):58–77, July 2008. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661108000864.

Strom:1993:AVP

[SPB93] Suzanne L. Strom, James R. Postel, and Beatrice C. Booth. Abundance, variability, and potential grazing impact of planktonic ciliates in the open subaratic Pacific Ocean. Progress in Oceanography, 32(??):185–203, ???? 1993. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0079661193900134.

Shipe:2002:ENS

[SPB^{+}02] R. F. Shipe, U. Passow, M. A. Brzezinski, W. M. Graham, D. K. Pak, D. A. Siegel, and A. L. Alldredge. Effects of the 1997–98 El Niño on seasonal variations in suspended and sinking particles in the Santa Barbara basin. Progress in Oceanography, 54(1–4):105–127, July 2002. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661102000459.

Schulz:2012:STH

[SPB^{+}12] Jan Schulz, Myron A. Peck, Kristina Barz, Jörn Oliver Schmidt, Frank C. Hansen, Janna Peters, Jasmin Renz, Miriam Dickmann, Volker Mohrholz, Jörg Dutz, and Hans-Jürgen Hirche. Spatial and temporal habitat partitioning by zooplankton in the Bornholm Basin (central Baltic Sea). Progress in Oceanography, 107(??):3–30, December 2012. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661112000791.

Stepien:2019:TFS

[SPB19] Anna Stepień, Krzysztof Pabis, and Magdalena Błażewicz. Tanaidacean faunas of the Sea of Okhotsk and northern slope of the Kuril–Kamchatka Trench. Progress in Oceanography, 178(??):??, November 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661119301478.

Scott:2023:RCR

[SPC^{+}23] James Scott, Sophie Pitois, Veronique Creach, Gill Malin, Phil Culverhouse, and Julian Tilbury. Resolution changes relationships: Optimizing sampling design using small scale zooplankton data. Progress in Oceanography, 210(??):??, January 2023. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661122002051.

Salgado:2023:DDC

[SPF^{+}23] Pablo Salgado, Gemita Pizarro, Máximo Frangopulos, Marco Pinto-Torres, Carolina Toro, Rodrigo Torres, Emilio Alarcón, Leonardo Guzmán, Karen Manríquez, Roberto Raimapo, and Emma Cascales. Distribution of dinoflagellate cyst assemblages in surface sediments of Magellan fjords and channels (Patagonia, Chile) with a focus on harmful species: an overview on environmental scenario. Progress in Oceanography, 213(??):??, April–May 2023. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661123000435.

Schmid:2006:TPC

[SPG^{+}06] Michael K. Schmid, Dieter Piepenburg, Alexander A. Golikov, Karen von Juterzenka, Victor V. Petryashov, and Michael Spindler. Trophic pathways and carbon flux patterns in the Laptev Sea. Progress in Oceanography, 71(2–4):314–330, October 2006. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661106001236.

Smith:1983:PCE

[SPH83] John Clegg Smith, Trevor Platt, and William G. Harrison. Photoadaptation of carboxylating enzymes and photosynthesis during a spring bloom. Progress in Oceanography, 12(4):425– 459, ???? 1983. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661183900022.

Sinclair:2015:ESS

[SPH^{+}15a] Michael Sinclair, Michael Power, Erica Head, William K. W. Li, Michael McMahon, Robert Mohn, Robert O’Boyle, Douglas Swain, and John Tremblay. Eastern Scotian Shelf trophic dynamics: a review of the evidence for diverse hypotheses. Progress in Oceanography, 138 (part A)(??):305–321, November 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661115001950.

Stern:2015:NLP

[SPH^{+}15b] Rowena F. Stern, Kathryn T. Picard, Kristina M. Hamilton, Antony Walne, Glen A. Tarran, David Mills, Abigail McQuatters-Gollop, and Martin Edwards. Novel lineage patterns from an automated water sampler to probe marine microbial biodiversity with ships of opportunity. Progress in Oceanography, 137(??):409–420, September 2015. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661115000828.

Stark:2019:FOC

[SPK^{+}19] Jonathan S. Stark, Edward T. Peltzer, David I. Kline, Ana M. Queirós, T. Erin Cox, Kent Headley, James Barry, Frédéric Gazeau, John W. Runcie, Stephen Widdicombe, Mark Milnes, Nicholas P. Roden, James Black, Steven Whiteside, Glenn Johnstone, Jeroen Ingels, Emily Shaw, Levente Bodrossy, Juan Diego Gaitan-Espitia, William Kirkwood, and JeanPierre Gattuso. Free Ocean CO_{2} Enrichment (FOCE) experiments: Scientific and technical recommendations for future in situ ocean acidification projects. Progress in Oceanography, 172(??):89–107, March 2019. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661117302598.

Suca:2022:FDN

[SPK^{+}22] Justin J. Suca, Julie W. Pringle, Zofia R. Knorek, Sara L. Hamilton, David E. Richardson, and Joel K. Llopiz. Feeding dynamics of Northwest Atlantic small pelagic fishes. Progress in Oceanography, 165(??):52–62, ???? 2022. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661118300041.

Sepic:2022:CPO

[ŠPM^{+}22] Jadranka Šepić, Miroslava Pasarić, Iva Meugorac, Ivica Vilibić, Maja Karlović, and Marko Mlinar. Climatology and process-oriented analysis of the Adriatic sea level extremes. Progress in Oceanography, 209(??):??, December 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661122001677.

Speckman:2005:PSA

[SPMVP05] Suzann G. Speckman, John F. Piatt, Carolina V. Minte-Vera, and Julia K. Parrish. Parallel structure among environmental gradients and three trophic levels in a subarctic estuary. Progress in Oceanography, 66(1):25–65, July 2005. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661105000789.

Stips:1998:SDB

[SPN98] Adolf Stips, Hartmut Prandke, and Thomas Neumann. The structure and dynamics of the bottom boundary layer in shallow sea areas without tidal influence: an experimental approach. Progress in Oceanography, 41(4):383–453, October 1998. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661198000226.

Sprintall:2008:LTT

[Spr08] Janet Sprintall. Long-term trends and interannual variability of temperature in Drake Passage. Progress in Oceanography, 77(4):316–330, June 2008. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661108000669.

Springer:1999:MBM

[SPS^{+}99] A. M. Springer, J. F. Piatt, V. P. Shuntov, G. B. Van Vliet, V. L. Vladimirov, A. E. Kuzin, and A. S. Perlov. Marine birds and mammals of the Pacific subarctic gyres. Progress in Oceanography, 43(2–4):443–487, March 1999. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661199000142.

Schneider:2014:HPC

[SPSR^{+}14] Wolfgang Schneider, Iván Pérez-Santos, Lauren Ross, Luis Bravo, Romanet Seguel, and Freddy Hernández. On the hydrography of Puyuhuapi Channel, Chilean Patagonia. Progress in Oceanography, 129(??):8–18, December 2014. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661114000330.

Serra-Pompei:2020:GST

[SPSV^{+}20] Camila Serra-Pompei, Floor Soudijn, André W. Visser, Thomas Kiørboe, and Ken H. Andersen. A general sizeand trait-based model of plankton communities. Progress in Oceanography, 189(??):??, November–December 2020. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S007966112030207X.

Stock:2015:SSS

[SPV^{+}15] Charles A. Stock, Kathy Pegion, Gabriel A. Vecchi, Michael A. Alexander, Desiree Tommasi, Nicholas A. Bond, Paula S. Fratantoni, Richard G. Gudgel, Trond Kristiansen, Todd D. O’Brien, Yan Xue, and Xiasong Yang. Seasonal sea surface temperature anomaly prediction for coastal ecosystems. Progress in Oceanography, 137 (part A)(??):219–236, September 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661115001408.

Strzelewicz:2022:IPA

[SPW22] Agnieszka Strzelewicz, Anna Przyborska, and Waldemar Walczowski. Increased presence of Atlantic Water on the shelf south-west of Spitsbergen with implications for the Arctic fjord Hornsund. Progress in Oceanography, 200(??):??, January 2022. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S007966112100197X.

Sun:2021:LPA

[SPWH21] Deyong Sun, Tianfeng Pan, Shengqiang Wang, and Chuanmin Hu. Linking phytoplankton absorption to community composition in Chinese marginal seas. Progress in Oceanography, 192(??):??, March 2021. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661121000070.

Smyth:2017:DAO

[SQJ^{+}17] Tim Smyth, Graham Quartly, Thomas Jackson, Glen Tarran, Malcolm Woodward, Carolyn Harris, Chris Gallienne, Rob Thomas, Ruth Airs, Denise Cummings, Robert Brewin, Vassilis Kitidis, John Stephens, Mike Zubkov, and Andrew Rees. Determining Atlantic Ocean province contrasts and variations. Progress in Oceanography, 158(??):19–40, November 2017. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661116300167.

Singh:2015:ECN

[SR15] Arvind Singh and R. Ramesh. Environmental controls on new and primary production in the northern Indian Ocean. Progress in Oceanography, 131(??):138–145, February 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S007966111400216X.

Shankar:2019:RPP

[SRAV19] D. Shankar, R. Remya, A. C. Anil, and V. Vijith. Role of physical processes in determining the nature of fisheries in the eastern Arabian Sea. Progress in Oceanography, 172(??):124– 158, March 2019. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661117303476.

Schartau:2019:SBI

[SRF^{+}19] M. Schartau, R. Riethmüller, G. Flöser, J. E. E. van Beusekom, H. Krasemann, R. Hofmeister, and K. Wirtz. On the separation between inorganic and organic fractions of suspended matter in a marine coastal environment. Progress in Oceanography, 171(??):231–250, February 2019. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661118300557.

Serrano:2022:ECC

[SRFHDH22] David Serrano, Evlin Ramírez-Félix, Michel Hendrickx, and Francisco Delgadillo-Hinojosa. Effects of climate change on the steric height and heat content in the Ballenas Channel, Gulf of California, Mexico (1939–2011). Progress in Oceanography, 207(??):??, September–October 2022. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661122001343.

Stramski:2019:IOP

[SRG^{+}19] Dariusz Stramski, Rick A. Reynolds, Pierre Gernez, Rüdiger Röttgers, and Oliver Wurl. Inherent optical properties and particle characteristics of the sea-surface microlayer. Progress in Oceanography, 176(??):??, September 2019. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661119300084.

Solberg:2015:EOS

[SRK15] Ingrid Solberg, Anders Røstad, and Stein Kaartvedt. Ecology of overwintering sprat (Sprattus sprattus). Progress in Oceanography, 138 (part A)(??):116–135, November 2015. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661115001755.

StJohn:2010:ICS

[SRM^{+}10] Michael A. St. John, Javier Ruiz, Patrick Monfray, Ivo Grigorov, and Charles G. Hannah. Introduction to the Cadiz Symposium on marine ecosystem model parameterisation: Examining the state of our art. Progress in Oceanography, 84(1–2):1–5, January 2010. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661110000029.

Spitz:2018:PCC

[SRT^{+}18] J. Spitz, V. Ridoux, A. W. Trites, S. Laran, and M. Authier. Prey consumption by cetaceans reveals the importance of energy-rich food webs in the Bay of Biscay. Progress in Oceanography, 166(??):148–158, September 2018. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661117300903.

Schroeder:1969:HRS

[SS69] Elizabeth Schroeder and Henry Stommel. How representative is the series of Panulirus stations of monthly mean conditions off Bermuda? Progress in Oceanography, 5(??):31– 40, ???? 1969. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661169900251.

Shimanaga:2003:SRR

[SS03] Motohiro Shimanaga and Yoshihisa Shirayama. Sex ratio and reproductive activity of benthic copepods in bathyal Sagami Bay (1430 m), central Japan. Progress in Oceanography, 57(1):97–107, April 2003. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661103000533.

Shaw:2017:MSR

[SS17] Justin Shaw and Marek Stastna. A model for shear response in swimming plankton. Progress in Oceanography, 151(??):1–12, February 2017. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661116300131.

Spivey:2014:AIF

[SSB14] R. J. Spivey, S. Stansfield, and C. M. Bishop. Analysing the intermittent flapping flight of a Manx Shearwater, Puffinus puffinus, and its sporadic use of a wave-meandering wingsailing flight strategy. Progress in Oceanography, 125(??):62– 73, June 2014. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661114000536.

Saeedi:2019:ECC

[SSB19] Hanieh Saeedi, Marianna Simões, and Angelika Brandt. Endemicity and community composition of marine species along the NW Pacific and the adjacent Arctic Ocean. Progress in Oceanography, 178(??):??, November 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661119301557.

Saeedi:2020:BDP

[SSB20a] Hanieh Saeedi, Marianna Simões, and Angelika Brandt. Biodiversity and distribution patterns of deep-sea fauna along the temperate NW Pacific. Progress in Oceanography, 183(??): ??, April 2020. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661120300355.

Schwing:2020:WBP

[SSB^{+}20b] Franklin B. Schwing, Michael J. Sissenwine, Harold Batchelder, Hans G. Dam, Jaime Gómez-Gutiérrez, Julie E. Keister, Hui Liu, and Jay O. Peterson. William (Bill) Peterson’s contributions to ocean science, management, and policy. Progress in Oceanography, 182(??):??, March 2020. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661119304215.

Saito:2005:RMS

[SSH^{+}05] Hiroaki Saito, Koji Suzuki, Akira Hinuma, Takashi Ota, Kimio Fukami, Hiroshi Kiyosawa, Toshiro Saino, and Atsushi Tsuda. Responses of microzooplankton to in situ iron fertilization in the western subarctic Pacific (SEEDS). Progress in Oceanography, 64(2–4):223–236, February 2005. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661105000315.

Sharples:2013:PFS

[SSI13] Jonathan Sharples, Beth E. Scott, and Mark E. Inall. From physics to fishing over a shelf sea bank. Progress in Oceanography, 117(??):1–8, October 2013. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661113001031.

Schmidt:2019:NID

[SSKA19] Christina Schmidt, Valentina V. Sattarova, Laura Katrynski, and Pedro Martínez Arbizu. New insights from the deep: Meiofauna in the Kuril–Kamchatka Trench and adjacent abyssal plain. Progress in Oceanography, 173(??):192– 207, April 2019. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661118302751.

Sobarzo:2007:NIM

[SSL07] Marcus Sobarzo, R. Kipp Shearman, and Steve Lentz. Nearinertial motions over the continental shelf off 4Concepción, central Chile. Progress in Oceanography, 75(3):348–362, November 2007. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661107001565.

Senina:2008:PEB

[SSL08] Inna Senina, John Sibert, and Patrick Lehodey. Parameter estimation for basin-scale ecosystem-linked population models of large pelagic predators: Application to skipjack tuna. Progress in Oceanography, 78(4):319–335, September 2008. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661108001171.

Simenstad:1990:CPF

[SSM90a] Charles A. Simenstad, Lawrence F. Small, and C. David McIntire. Consumption processes and food web structure in the Columbia River Estuary. Progress in Oceanography, 25(1–4):271–297, ???? 1990. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/007966119090010Y.

Simenstad:1990:CRE

[SSM^{+}90b] Charles A. Simenstad, Lawrence F. Small, C. David McIntire, David A. Jay, and Christopher Sherwood. Columbia River estuary studies: an introduction to the estuary, a brief history, and prior studies. Progress in Oceanography, 25(1–4):1– 13, ???? 1990. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/007966119090002J.

Sinha:2018:AEO

[SSM^{+}18] B. Sinha, D. A. Smeed, G. McCarthy, B. I. Moat, S. A. Josey, J. J.-M. Hirschi, E. Frajka-Williams, A. T. Blaker, D. Rayner, and G. Madec. The accuracy of estimates of the overturning circulation from basin-wide mooring arrays. Progress in Oceanography, 160(??):101–123, January 2018. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661116302439.

Shiomoto:2023:SFP

[SSN23] Akihiro Shiomoto, Hiroko Sasaki, and Daiki Nomura. Sizefractionated phytoplankton biomass and primary production in the eastern Indian sector of the Southern Ocean in the austral summer 2018/2019. Progress in Oceanography, 218(??):??, November 2023. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661123001623.

Saldias:2019:FSS

[SSQ19] Gonzalo S. Saldías, Marcus Sobarzo, and Renato Quiñones. Freshwater structure and its seasonal variability off western Patagonia. Progress in Oceanography, 174(??):143–153, May 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661118301022.

Santora:2011:MSO

[SSS^{+}11] Jarrod A. Santora, William J. Sydeman, Isaac D. Schroeder, Brian K. Wells, and John C. Field. Mesoscale structure and oceanographic determinants of krill hotspots in the California Current: Implications for trophic transfer and conservation. Progress in Oceanography, 91(4):397–409, December 2011. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661111000371.

Storesund:2017:LBC

[SST^{+}17] Julia E. Storesund, Ruth-Anne Sandaa, T. Frede Thingstad, Lars Asplin, Jon Albretsen, and Svein Rune Erga. Linking bacterial community structure to advection and environmental impact along a coast-fjord gradient of the sognefjord, western Norway. Progress in Oceanography, 159(??):13–30, December 2017. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661115300598.

Schnack-Schiel:1995:LCS

[SSTD^{+}95] Sigrid B. Schnack-Schiel, David Thomas, Gerhard S. Dieckmann, Hajo Eicken, Rolf Gradinger, Michael Spindler, Jürgen Weissenberger, Elke Mizdalski, and Kerstin Beyer. Life cycle strategy of the Antarctic calanoid copepod Stephos longipes. Progress in Oceanography, 36(1):45–75, ???? 1995. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0079661195000143.

Shao:2016:SVS

[SSTL16] Changgao Shao, Yi Sui, Danling Tang, and Louis Legendre. Spatial variability of surface-sediment porewater pH and related water-column characteristics in deep waters of the northern South China Sea. Progress in Oceanography, 149(??):134–144, December 2016. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661116300672.

Svensen:2011:ZDA

[SSV^{+}11] Camilla Svensen, Lena Seuthe, Yulia Vasilyeva, Anna Pasternak, and Edmond Hansen. Zooplankton distribution across Fram Strait in autumn: Are small copepods and protozooplankton important? Progress in Oceanography, 91(4):534–544, December 2011. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S007966111100098X.

Savelieva:2000:CSS

[SSVP00] N. I. Savelieva, I. P. Semiletov, L. N. Vasilevskaya, and S. P. Pugach. A climate shift in seasonal values of meteorological and hydrological parameters for Northeastern Asia. Progress in Oceanography, 47(2–4):279–297, October 2000. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661100000392.

Suryan:2009:EFL

[SSW^{+}09] Robert M. Suryan, Vincent S. Saba, Bryan P. Wallace, Scott A. Hatch, Morten Frederiksen, and Sarah Wanless. Environmental forcing on life history strategies: Evidence for multi-trophic level responses at ocean basin scales. Progress in Oceanography, 81(1–4):214–222, April 2009. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661109000329.

Stearns:1965:TUD

[ST65] Charles E. Stearns and David L. Thurber. Th^{230} /U^{234} dates of late Pleistocene marine fossils from the Mediterranean and Moroccan littorals. Progress in Oceanography, 4 (??):293–305, September 2, 1965. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/007966116590056X.

Saito:2003:ILI

[ST03] Hiroaki Saito and Atsushi Tsuda. Influence of light intensity on diatom physiology and nutrient dynamics in the Oyashio region. Progress in Oceanography, 57(3–4):251–263, June 2003. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661103001009.

Smayda:2010:DBU

[ST10] T. J. Smayda and V. L. Trainer. Dinoflagellate blooms in upwelling systems: Seeding, variability, and contrasts with diatom bloom behaviour. Progress in Oceanography, 85(1–2):92– 107, April 2010. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661110000170.

Savidge:1992:BSB

[STB^{+}92] G. Savidge, D. R. Turner, P. H. Burkill, A. J. Watson, M. V. Angel, R. D. Pingree, H. Leach, and K. J. Richards. The BOFS 1990 spring bloom experiment: Temporal evolution and spatial variability of the hydrographic field. Progress in Oceanography, 29(3):235–281, ???? 1992. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/007966119290005K.

Scheffer:2010:FBK

[STC10] Annette Scheffer, Philip N. Trathan, and Martin Collins. Foraging behaviour of king penguins (Aptenodytes patagonicus) in relation to predictable mesoscale oceanographic features in the Polar Front Zone to the north of South Georgia. Progress in Oceanography, 86(1–2):232–245, July 2010. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S007966111000042X.

Stevens:1991:NOC

[Ste91] David P. Stevens. A numerical ocean circulation model of the Norwegian and Greenland Seas. Progress in Oceanography, 27(3–4):365–402, ???? 1991. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/007966119190029L.

Steele:2004:RSO

[Ste04] John H. Steele. Regime shifts in the ocean: reconciling observations and theory. Progress in Oceanography, 60(2–4):135–141, February 2004. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661104000242.

Steele:2012:PSI

[Ste12] John H. Steele. Prediction, scenarios and insight: the uses of an end-to-end model. Progress in Oceanography, 102(??):67–73, September 2012. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661112000225.

Scheffer:2016:CIM

[STEB16] Annette Scheffer, Philip N. Trathan, Johnnie G. Edmonston, and Charles-André Bost. Combined influence of mesoscale circulation and bathymetry on the foraging behaviour of a diving predator, the king penguin (Aptenodytes patagonicus). Progress in Oceanography, 141(??):1–16, February 2016. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661115002141.

Schweigert:2013:FLP

[STF^{+}13] J. F. Schweigert, M. Thompson, C. Fort, D. E. Hay, T. W. Therriault, and L. N. Brown. Factors linking Pacific herring (Clupea pallasi) productivity and the spring plankton bloom in the Strait of Georgia, British Columbia, Canada. Progress in Oceanography, 115(??):103–110, August 2013. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661113000591.

Schmidt:2018:OMS

[STG^{+}18] Natascha Schmidt, Delphine Thibault, François Galgani, Andrea Paluselli, and Richard Sempéré. Occurrence of microplastics in surface waters of the Gulf of Lion (NW Mediterranean Sea). Progress in Oceanography, 163(??):214–220, April 2018. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661117300484.

Sydeman:2014:MOC

[STGR^{+}14] William J. Sydeman, Sarah Ann Thompson, Marisol García-Reyes, Mati Kahru, William T. Peterson, and John L.

Largier. Multivariate ocean-climate indicators (MOCI) for the central California Current: Environmental change, 1990– 2010. Progress in Oceanography, 120(??):352–369, January 2014. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661113002115.

Sydeman:2023:ECT

[STGR^{+}23] William J. Sydeman, Sarah Ann Thompson, Marisol García-Reyes, Caitlin Kroeger, Brian Hoover, Sonia D. Batten, and Nora A. Rojek. Effects of currents and temperature on ecosystem productivity in Unimak Pass, Alaska, a premier seabird and biodiversity hotspot. Progress in Oceanography, 216(??): ??, August 2023. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661123001258.

Shin:2002:PTR

[STHM02] Kyung-Hoon Shin, Noriyuki Tanaka, Naomi Harada, and Jean-Claude Marty. Production and turnover rates of C_{37} alkenones in the eastern Bering Sea: implication for the mechanism of a long duration of Emiliania huxleyi bloom. Progress in Oceanography, 55(1–2):113–129, October 2002. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661102000733.

Shen:2014:LTV

[STJ^{+}14] Mao-Lin Shen, Yu-Heng Tseng, Sen Jan, Chih-Chieh Young, and Ming-Da Chiou. Long-term variability of the Kuroshio transport east of Taiwan and the climate it conveys. Progress in Oceanography, 121(??):60–73, February 2014. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S007966111300195X.

Shin:2010:CLH

[STM10] Yunne-Jai Shin, Morgane Travers, and Olivier Maury. Coupling low and high trophic levels models: Towards a pathwaysorientated approach for end-to-end models. Progress in Oceanography, 84(1–2):105–112, January 2010. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661109001451.

Santana-Toscano:2023:ZCN

[STPHM^{+}23] Daniel Santana-Toscano, M. Dolores Pérez-Hernández, Alison M. Macdonald, Cristina Arumí-Planas, Verónica Caínzos, and Alonso Hernández-Guerra. Zonal circulation in the North Atlantic Ocean at 52^{o}W from WOCE-WHP and CLIVAR sections: 1997, 2003 and 2012. Progress in Oceanography, 216(??): ??, August 2023. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S007966112300112X.

Subbotina:2001:SCS

[STR01] Marina M. Subbotina, Richard E. Thomson, and Alexander B. Rabinovich. Spectral characteristics of sea level variability along the west coast of North America during the 1982–83 and 1997–98 El Niño events. Progress in Oceanography, 49(??):353– 372, ???? 2001. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661101000301.

Shimode:2012:DLH

[STS^{+}12] Shinji Shimode, Kazutaka Takahashi, Yugo Shimizu, Takumi Nonomura, and Atsushi Tsuda. Distribution and life history of the planktonic copepod, Eucalanus californicus, in the northwestern Pacific: Mechanisms for population maintenance within a high primary production area. Progress in Oceanography, 96(1):1–13, April 2012. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661111000991.

Steele:2015:ACP

[STW^{+}15] Deborah J. Steele, Glen A. Tarran, Claire E. Widdicombe, E. Malcolm S. Woodward, Susan A. Kimmance, Daniel J. Franklin, and Ruth L. Airs. Abundance of a chlorophyll-a precursor and the oxidation product hydroxychlorophyll-a during seasonal phytoplankton community progression in the Western English Channel. Progress in Oceanography, 137(??):434–445, September 2015. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661115000889.

Sudo:1986:NJS

[Sud86] Hideo Sudo. A note on the Japan Sea proper water. Progress in Oceanography, 17(3–4):313–336, ???? 1986. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0079661186900522.

Suk:1988:AVI

[Suk88] Moon-Sik Suk. Application of a variational inverse model to determine a wintertime circulation in the East Sea of Korea. Progress in Oceanography, 21(3–4):281–293, ???? 1988. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0079661188900080.

Spencer:1997:EDO

[SV97] R. Spencer and J. M. Vassie. The evolution of deep ocean pressure measurements in the UK. Progress in Oceanography, 40(??):423–435, ???? 1997. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661198000111.

Silva:2014:HCP

[SV14] Nelson Silva and Cristian A. Vargas. Hypoxia in Chilean Patagonian fjords. Progress in Oceanography, 129(??):62–74, December 2014. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661114001104.

Svansson:1965:SHP

[Sva65] Artur Svansson. Some hydrographic problems of the Skagerrak. Progress in Oceanography, 3(??):355–372, ???? 1965. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0079661165900315.

Sanchez-Vidal:2013:RTT

[SVHM^{+}13] Anna Sanchez-Vidal, Marina Higueras, Eugènia Martí, Camino Liquete, Antoni Calafat, Philippe Kerhervé, and Miquel Canals. Riverine transport of terrestrial organic matter to the North Catalan margin, NW Mediterranean Sea.

Progress in Oceanography, 118(??):71–80, November 2013. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S007966111300133X.

Stashchuk:2014:HDS

[SVIA14] Nataliya Stashchuk, Vasiliy Vlasenko, Mark E. Inall, and Dmitry Aleynik. Horizontal dispersion in shelf seas: High resolution modelling as an aid to sparse sampling. Progress in Oceanography, 128(??):74–87, November 2014. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S007966111400130X.

Sepic:2015:HFS

[ŠVL^{+}15] Jadranka Šepić, Ivica Vilibić, Amaury Lafon, Loïc Macheboeuf, and Zvonko Ivanović. High-frequency sea level oscillations in the Mediterranean and their connection to synoptic patterns. Progress in Oceanography, 137 (part A)(??):284–298, September 2015. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661115001573.

Sala:2023:MET

[SVL^{+}23] Iria Sala, Sergio M. Vallina, Marina Lévy, Marina Bolado-Penagos, Carlos M. García, Fidel Echevarría, and José C. Sánchez-Garrido. Modelling the effect of the tidal cycle on the high phytoplankton biomass area of Cape Trafalgar (SW Iberian Peninsula). Progress in Oceanography, 217(??):??, September–October 2023. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661123001283.

Scheffer:2004:MMR

[SvN04] Marten Scheffer and Egbert H. van Nes. Mechanisms for marine regime shifts: Can we use lakes as microcosms for oceans? Progress in Oceanography, 60(2–4):303–319, February 2004. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661104000333.

Shankar:2002:MCN

[SVU02] D. Shankar, P. N. Vinayachandran, and A. S. Unnikrishnan. The monsoon currents in the north Indian Ocean. Progress in Oceanography, 52(1):63–120, January 2002. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661102000241.

Schmidt:2002:SDR

[SvWRvB02] S. Schmidt, Tj. C. E. van Weering, J.-L. Reyss, and P. van Beek. Seasonal deposition and reworking at the sediment-water interface on the northwestern Iberian margin. Progress in Oceanography, 52(??):331–348, ???? 2002. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661102000149.

Suess:1965:PHF

[SW65] Hans E. Suess and Heinrich Wänke. On the possibility of a helium flux through the ocean floor. Progress in Oceanography, 3(??):347–353, ???? 1965. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0079661165900303.

Simpson:1981:BRL

[SW81] W. R. Simpson and I. O. S. Wormley. Book review: Lead in the marine environment: Edited by Marko Branica and Zdenka Konrad. Pergamon Press (1980) U.S.$69.00 (£30.00) 363 pages. ISBN 0-08-022960-3. Progress in Oceanography, 10(1):iv–vi, ???? 1981. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661181900094.

Sear:1992:SPA

[SW92] C. B. Sear and P. Wadhams. Statistical properties of Arctic sea ice morphology derived from sidescan sonar images. Progress in Oceanography, 29(2):133–160, ???? 1992. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/007966119290022R.

Slagstad:2001:MDC

[SW01] Dag Slagstad and Paul Wassmann. Modelling the 3-D carbon flux across the Iberian margin during the upwelling season in 1998. Progress in Oceanography, 51(2–4):467–497, November 2001. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661101000805.

Shi:2012:SVB

[SW12] Wei Shi and Menghua Wang. Satellite views of the Bohai Sea, Yellow Sea, and East China Sea. Progress in Oceanography, 104(??):30–45, October 2012. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661112000584.

Summers:2021:UBP

[SW21] Natalie Summers and Les Watling. Upper Bathyal Pacific Ocean biogeographic provinces from octocoral distributions. Progress in Oceanography, 191(??):??, February 2021. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661120302445.

Shi:2022:BDM

[SW22] Wei Shi and Menghua Wang. Biological dipole mode indices: New parameters to characterize the physical and biological processes of the Indian Ocean dipole event. Progress in Oceanography, 206(??):??, August 2022. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661122001070.

Swallow:1976:EN

[Swa76] Mary Swallow. Editor’s note 1976. Progress in Oceanography, 7(1):1, ???? 1976. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661176900021.

Swallow:1977:EN

[Swa77] Mary Swallow. Editorial notice. Progress in Oceanography, 7 (4):i, ???? 1977. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661177900027.

Sun:2024:CDC

[SWH^{+}24] Ping Sun, Ying Wang, Xin Huang, SuSu Xu, Ramiro Logares, Yibin Huang, Dapeng Xu, and Bangqin Huang. Cracking the dynamic code of the deep: Unexpected seasonal patterns of active protistan-bacterial microbiomes in the mesopelagic zone of the South China Sea. Progress in Oceanography, 225(??): ??, July 2024. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661124000867.

Scott:2013:FSB

[SWP^{+}13a] B. E. Scott, A. Webb, M. R. Palmer, C. B. Embling, and J. Sharples. Fine scale bio-physical oceanographic characteristics predict the foraging occurrence of contrasting seabird species; gannet (Morus bassanus) and storm petrel (Hydrobates pelagicus). Progress in Oceanography, 117(??):118–129, October 2013. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661113000992.

Skjoldal:2013:IZN

[SWP^{+}13b] Hein Rune Skjoldal, Peter H. Wiebe, Lutz Postel, Tor Knutsen, Stein Kaartvedt, and Douglas D. Sameoto. Intercomparison of zooplankton (net) sampling systems: Results from the ICES/GLOBEC sea-going workshop. Progress in Oceanography, 108(??):1–42, January 2013. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661112001309.

Shelley:2017:TTG

[SWT^{+}17] Rachel U. Shelley, Neil J. Wyatt, Glenn A. Tarran, Andrew P. Rees, Paul J. Worsfold, and Maeve C. Lohan. A tale of two gyres: Contrasting distributions of dissolved cobalt and iron in the Atlantic Ocean during an Atlantic Meridional Transect (AMT-19)n. Progress in Oceanography, 158(??):52–64, November 2017. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661115300604.

Stempniewicz:2021:AAW

[SWZS^{+}21] Lech Stempniewicz, Agata Weydmann-Zwolicka, Agnieszka Strzelewicz, Michał Goc, Marta Głuchowska, Dorota Kidawa, Waldemar Walczowski, Jan Marcin Wesławski, and Adrian Zwolicki. Advection of Atlantic water masses influences seabird community foraging in a high-Arctic fjord. Progress in Oceanography, 193(??):??, April–May 2021. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661121000380.

Silva:2015:FSX

[SYB^{+}15] Claudio Silva, Eleuterio Yáñez, María Angela Barbieri, Claudio Bernal, and Antonio Aranis. Forecasts of swordfish (Xiphias gladius) and common sardine (Strangomera bentincki) off Chile under the A2 IPCC climate change scenario. Progress in Oceanography, 134(??):343–355, May 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661115000452.

Suzuki:2021:DCN

[SYN^{+}21] Koji Suzuki, Yuta Yoshino, Yuichi Nosaka, Jun Nishioka, Stanford B. Hooker, and Toru Hirawake. Diatoms contributing to new production in surface waters of the northern Bering and Chukchi Seas during summer with reference to water column stratification. Progress in Oceanography, 199(??):??, December 2021. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661121001762.

Sheng:2006:NSS

[SZG06] Jinyu Sheng, Xiaoming Zhai, and Richard J. Greatbatch. Numerical study of the storm-induced circulation on the Scotian Shelf during Hurricane Juan using a nested-grid ocean model. Progress in Oceanography, 70(2–4):233–254, August 2006. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661106000863.

Szuts:2012:UMI

[Szu12] Zoltan B. Szuts. Using motionally-induced electric signals to indirectly measure ocean velocity: Instrumental and theoretical developments. Progress in Oceanography, 96(1):108– 127, April 2012. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661111001406.

Taviani:2022:UTD

[TAF^{+}22] Marco Taviani, Lorenzo Angeletti, Federica Foglini, Cesare Corselli, Ina Nasto, Edwige Pons-Branchu, and Paolo Montagna. U/Th dating records of cold-water coral colonization in submarine canyons and adjacent sectors of the southern Adriatic Sea since the Last Glacial Maximum. Progress in Oceanography, 175(??):300–308, ???? 2022. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661118300971.

Tsumune:2011:TCS

[TAH^{+}11] Daisuke Tsumune, Michio Aoyama, Katsumi Hirose, Frank O. Bryan, Keith Lindsay, and Gokhan Danabasoglu. Transport of ^{137}Cs to the Southern Hemisphere in an ocean general circulation model. Progress in Oceanography, 89(1–4):38–48, April 2011. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661110001588.

Takano:1988:I

[Tak88] Kenzo Takano. Introduction. Progress in Oceanography, 21(3– 4):v, ???? 1988. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661188900031.

Talley:2008:FTE

[Tal08] Lynne D. Talley. Freshwater transport estimates and the global overturning circulation: Shallow, deep and throughflow components. Progress in Oceanography, 78(4):257–303, September 2008. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661108001080.

Toyoda:2013:ISE

[TAM^{+}13] T. Toyoda, T. Awaji, S. Masuda, N. Sugiura, H. Igarashi, Y. Sasaki, Y. Hiyoshi, Y. Ishikawa, S.-I. Saitoh, S. Yoon, T. In, and M. J. Kishi. Improved state estimations of lower trophic ecosystems in the global ocean based on a Green’s function approach. Progress in Oceanography, 119(??):90–107, December 2013. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661113001687.

Tremblay:2015:GRD

[TAM^{+}15] Jean-Éric Tremblay, Leif G. Anderson, Patricia Matrai, Pierre Coupel, Simon Bélanger, Christine Michel, and Marit Reigstad. Global and regional drivers of nutrient supply, primary production and CO_{2} drawdown in the changing Arctic Ocean. Progress in Oceanography, 139(??):171–196, December 2015. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661115001895.

Taniguchi:1999:DSL

[Tan99] Akira Taniguchi. Differences in the structure of the lower trophic levels of pelagic ecosystems in the eastern and western subarctic Pacific. Progress in Oceanography, 43(2–4):289– 315, March 1999. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661199000117.

Tintore:2005:MPO

[TAO05] Joaquín Tintoré, Alberto Alvarez, and Reiner Onken. Mediterranean physical oceanography and biogeochemical cycles: General circulation and climate variability. Progress in Oceanography, 66(2–4):87–88, August 2005. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661105000996.

Tait:2015:DRB

[TAW^{+}15] Karen Tait, Ruth L. Airs, Claire E. Widdicombe, Glen A. Tarran, Mark R. Jones, and Stephen Widdicombe. Dynamic responses of the benthic bacterial community at the Western English Channel observatory site L4 are driven by deposition of fresh phytodetritus. Progress in Oceanography, 137(??):546–558, September 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661115000877.

Tarran:2015:NPW

[TB15] Glen A. Tarran and John T. Bruun. Nanoplankton and picoplankton in the Western English Channel: abundance and seasonality from 2007–2013. Progress in Oceanography, 137 (??):446–455, September 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661115000919.

Theocharis:1999:SCH

[TBK^{+}99] Alexander Theocharis, Efstathios Balopoulos, Soterios Kioroglou, Harilaos Kontoyiannis, and Athanassia Iona. A synthesis of the circulation and hydrography of the South Aegean Sea and the Straits of the Cretan Arc (March 1994–January 1995). Progress in Oceanography, 44(4):469–509, December 1999. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661199000415.

Taboada:2019:SIP

[TBS^{+}19] Fernando González Taboada, Andrew D. Barton, Charles A. Stock, John Dunne, and Jasmin G. John. Seasonal to interannual predictability of oceanic net primary production inferred from satellite observations. Progress in Oceanography, 170(??):28–39, January 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661117300733.

Truesdale:2000:MDD

[TBW00] V. W. Truesdale, A. J. Bale, and E. M. S. Woodward. The meridional distribution of dissolved iodine in near-surface waters of the Atlantic Ocean. Progress in Oceanography, 45(3–4):387–400, April 2000. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661100000094.

Thomas:2009:IVC

[TBW09] Andrew C. Thomas, Peter Brickley, and Ryan Weatherbee. Interannual variability in chlorophyll concentrations in the Humboldt and California Current Systems. Progress in Oceanography, 83(1–4):386–392, December 2009. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661109000755.

Tarres:2022:PFS

[TCDPP^{+}22] Marta Tarrés, Marc Cerdà-Domènech, Rut Pedrosa-Pàmies, Aitor Rumín-Caparrós, Antoni Calafat, Miquel Canals, and Anna Sanchez-Vidal. Particle fluxes in submarine canyons along a sediment-starved continental margin and in the adjacent open slope and basin in the SW Mediterranean Sea. Progress in Oceanography, 203(??):??, April–May 2022. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661122000441.

Tarres:2023:TDN

[TCDPP^{+}23] Marta Tarrés, Marc Cerdà-Domènech, Rut Pedrosa-Pàmies, Andrea Baza-Varas, Antoni Calafat, Anna Sanchez-Vidal, and Miquel Canals. Transport and distributions of naturally and anthropogenically sourced trace metals and arsenic in submarine canyons. Progress in Oceanography, 218(??):??, November 2023. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661123001659.

Torri:2018:PSS

[TCF^{+}18] Marco Torri, Raffaele Corrado, Federico Falcini, Angela Cuttitta, Luigi Palatella, Guglielmo Lacorata, Bernardo Patti, Marco Arculeo, Roberta Mifsud, Salvatore Mazzola, and Rosalia Santoleri. Planktonic stages of small pelagic fishes (Sardinella aurita and Engraulis encrasicolus) in the central Mediterranean Sea: the key role of physical forcings and implications for fisheries management. Progress in Oceanography, 162(??):25–39, March 2018. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661117300241.

Tubau:2015:MLF

[TCL^{+}15] Xavier Tubau, Miquel Canals, Galderic Lastras, Xavier Rayo, Jesus Rivera, and David Amblas. Marine litter on the floor of deep submarine canyons of the Northwestern Mediterranean Sea: the role of hydrodynamic processes. Progress in Oceanography, 134(??):379–403, May 2015. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661115000543.

Thorson:2020:DIE

[TCL20] James T. Thorson, Lorenzo Ciannelli, and Michael A. Litzow. Defining indices of ecosystem variability using biological samples of fish communities: a generalization of empirical orthogonal functions. Progress in Oceanography, 181(??):??, February 2020. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661119304240.

Tak:2020:NIG

[TCN20] Yong-Jin Tak, Yang-Ki Cho, and SungHyun Nam. Numerical investigation of the generation of continental shelf waves and their role in the westward shift of the Yellow Sea Warm Current. Progress in Oceanography, 187(??):??, August–September 2020. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661120301439.

Tecchio:2015:SFC

[TCS15] Samuele Tecchio, Marta Coll, and Francisco Sardà. Structure, functioning, and cumulative stressors of Mediterranean deep-sea ecosystems. Progress in Oceanography, 135(??):156– 167, June 2015. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661115001159.

Tesdal:2022:RNE

[TDGY22] Jan-Erik Tesdal, Hugh W. Ducklow, Joaquim I. Goes, and Igor Yashayaev. Recent nutrient enrichment and high biological productivity in the Labrador Sea is tied to enhanced winter convection. Progress in Oceanography, 206(??):??, August 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661122001069.

Tengberg:1995:BCP

[TDH^{+}95] A. Tengberg, F. De Bovee, P. Hall, W. Berelson, D. Chadwick, G. Ciceri, P. Crassous, A. Devol, S. Emerson, J. Gage, R. Glud, F. Graziottini, J. Gundersen, D. Hammond, W. Helder, K. Hinga, O. Holby, R. Jahnke, A. Khripounoff, S. Lieberman, V. Nuppenau, O. Pfannkuche, C. Reimers, G. Rowe, A. Sahami, F. Sayles, M. Schurter, D. Smallman, B. Wehrli, and P. De Wilde. Benthic chamber and profiling landers in oceanography — a review of design, technical solutions and functioning. Progress in Oceanography, 35(3):253–294, ???? 1995. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661195000096.

Turnewitsch:2016:TIP

[TDK^{+}16] Robert Turnewitsch, Matthew Dumont, Kostas Kiriakoulakis, Sonya Legg, Christian Mohn, Florian Peine, and George Wolff. Tidal influence on particulate organic carbon export fluxes around a tall seamount. Progress in Oceanography, 149(??):189–213, December 2016. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661116300702.

Turnewitsch:2017:CNS

[TDL^{+}17] Robert Turnewitsch, Andrew Dale, Niko Lahajnar, Richard S. Lampitt, and Kei Sakamoto. Can neap-spring tidal cycles modulate biogeochemical fluxes in the abyssal near-seafloor water column? Progress in Oceanography, 154(??):1–24, May 2017. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661116301860.

Touratier:2003:SCN

[TFM03] Franck Touratier, John G. Field, and Coleen L. Moloney. Simulated carbon and nitrogen flows of the planktonic food web during an upwelling relaxation period in St Helena Bay (southern Benguela ecosystem). Progress in Oceanography, 58(1):1– 41, July 2003. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661103000879.

Takahashi:2002:LTM

[TFY02] K. Takahashi, N. Fujitani, and M. Yanada. Long term monitoring of particle fluxes in the Bering Sea and the central subarctic Pacific Ocean, 1990–2000. Progress in Oceanography, 55(1–2):95–112, October 2002. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661102000721.

Trani:2014:AAC

[TFZS14] M. Trani, P. Falco, E. Zambianchi, and J. B. Sallée. Aspects of the Antarctic Circumpolar Current dynamics investigated with drifter data. Progress in Oceanography, 125(??): 1–15, June 2014. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S007966111400086X.

Toole:1981:DED

[TG81] John M. Toole and Daniel T. Georgi. On the dynamics and effects of double-diffusively driven intrusions. Progress in Oceanography, 10(2):123–145, ???? 1981. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0079661181900033.

Testor:2005:LSF

[TG05] P. Testor and J.-C. Gascard. Large scale flow separation and mesoscale eddy formation in the Algerian Basin. Progress in Oceanography, 66(2–4):211–230, August 2005. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661105000716.

Thiaw:2009:SPM

[TGJT09] Modou Thiaw, Didier Gascuel, Didier Jouffre, and Omar Thiom Thiaw. A surplus production model including environmental effects: Application to the Senegalese white shrimp stocks. Progress in Oceanography, 83(1–4):351–360, December 2009. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661109000809.

Truesdale:2005:VBD

[TGR05] Victor W. Truesdale, Jim E. Greenwood, and Andrew R. Rendell. In vitro, batch-dissolution of biogenic silica in seawater — the application of recent modelling to real data. Progress in Oceanography, 66(1):1–24, July 2005. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661105000546.

Teuber:2019:WWT

[THBA19] Lena Teuber, Wilhelm Hagen, Maya Bode, and Holger Auel. Who is who in the tropical Atlantic? Functional traits, ecophysiological adaptations and life strategies in tropical calanoid copepods. Progress in Oceanography, 171(??):128–135, February 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661118301988.

Tremblay:2006:TSP

[THM^{+}06] Jean-Éric Tremblay, Hiroshi Hattori, Christine Michel, Marc Ringuette, Zhi-Ping Mei, Connie Lovejoy, Louis Fortier, Keith A. Hobson, David Amiel, and Kirk Cochran. Trophic structure and pathways of biogenic carbon flow in the eastern North water polynya. Progress in Oceanography, 71 (2–4):402–425, October 2006. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661106001352.

Trenkel:2014:CEW

[THM^{+}14] V. M. Trenkel, G. Huse, B. R. MacKenzie, P. Alvarez, H. Arrizabalaga, M. Castonguay, N. Goñi, F. Grégoire, H. Hátún, T. Jansen, J. A. Jacobsen, P. Lehodey, M. Lutcavage, P. Mariani, G. D. Melvin, J. D. Neilson, L. Nøttestad, G. J. Óskarsson, M. R. Payne, D. E. Richardson, I. Senina, and D. C. Speirs. Comparative ecology of widely distributed pelagic fish species in the North Atlantic: Implications for modelling climate and fisheries impacts. Progress in Oceanography, 129(??):219–243, December 2014. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661114000913.

Thompson:1977:ORW

[Tho77] Rory O. R. Y. Thompson. Observations of Rossby waves near site D^{1}. Progress in Oceanography, 7(4):135–162, ???? 1977. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0079661177900039.

Thomas:1987:FSL

[Tho87] Robert H. Thomas. Future sea-level rise and its early detection by satellite remote sensing. Progress in Oceanography, 18(1–4): 23–40, ???? 1987. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661187900243.

Thorpe:1995:DPT

[Tho95] S. A. Thorpe. Dynamical processes of transfer at the sea surface. Progress in Oceanography, 35(4):315–352, ???? 1995. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/007966119580002B.

Tommasi:2021:MLV

[THP21] Desiree Tommasi, Brian P. V. Hunt, and Evgeny A. Pakhomov. Mechanisms linking variation in spring environmental conditions to changes in the phenology and abundance of Calanus marshallae and Eucalanus bungii. Progress in Oceanography, 199(??):??, December 2021. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661121001816.

Thurston:1990:ANA

[Thu90] Michael H. Thurston. Abyssal necrophagous amphipods (crustacea: Amphipoda) in the northeast and tropical Atlantic Ocean. Progress in Oceanography, 24(??):257–274, ???? 1990. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0079661190900362.

Takao:2014:BOP

[TII^{+}14] Shintaro Takao, Takahiro Iida, Tomonori Isada, Sei-Ichi Saitoh, Takafumi Hirata, and Koji Suzuki. Bio-optical properties during the summer season in the Sea of Okhotsk.

Progress in Oceanography, 126(??):233–241, August 2014. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661114000585.

Tugores:2016:SDS

[TIOM16] M. Pilar Tugores, Magdalena Iglesias, Dolores Oñate, and Joan Miquel. Spatial distribution, sampling precision and survey design optimisation with non-normal variables: the case of anchovy (Engraulis encrasicolus) recruitment in Spanish Mediterranean waters. Progress in Oceanography, 141(??):168–178, February 2016. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661116000021.

Titov:2020:DWA

[Tit20] Oleg Titov. Deep-water aeration and the Barents Sea ecosystem renewal: Natural regulation of the Northeast Arctic cod recruitment. Progress in Oceanography, 189(??):??, November– December 2020. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661120302044.

Taft:1973:MEU

[TJ73] Bruce A. Taft and James H. Jones. Measurements of the equatorial undercurrent in the eastern Pacific. Progress in Oceanography, 6(??):47–110, ???? 1973. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0079661173900050.

Tunnicliffe:1990:DCH

[TJ90] Verena Tunnicliffe and S. Kim Juniper. Dynamic character of the hydrothermal vent habitat and the nature of sulphide chimney fauna. Progress in Oceanography, 24(??):1– 13, ???? 1990. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/007966119090015T.

Thompson:2022:VOB

[TKC^{+}22] D. M. Thompson, J. Kleypas, F. Castruccio, E. N. Curchitser, M. L. Pinsky, B. Jönsson, and J. R. Watson. Variability in oceanographic barriers to coral larval dispersal: Do currents shape biodiversity? Progress in Oceanography, 165(??):110– 122, ???? 2022. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S007966111730191X.

Tsuda:2005:RDI

[TKK^{+}05] Atsushi Tsuda, Hiroshi Kiyosawa, Akira Kuwata, Mamiko Mochizuki, Naonobu Shiga, Hiroaki Saito, Sanae Chiba, Keiri Imai, Jun Nishioka, and Tsuneo Ono. Responses of diatoms to iron-enrichment (SEEDS) in the western subarctic Pacific, temporal and spatial comparisons. Progress in Oceanography, 64(2–4):189–205, February 2005. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661105000297.

Takahashi:2008:GIC

[TKSI08] Kazutaka Takahashi, Akira Kuwata, Hiroaki Saito, and Keiichiro Ide. Grazing impact of the copepod community in the Oyashio region of the western subarctic Pacific Ocean. Progress in Oceanography, 78(3):222–240, September 2008. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661108001195.

Tian:2006:LTC

[TKW06] Yongjun Tian, Hideaki Kidokoro, and Tatsuro Watanabe. Long-term changes in the fish community structure from the Tsushima warm current region of the Japan/East Sea with an emphasis on the impacts of fishing and climate regime shift over the last four decades. Progress in Oceanography, 68(2–4):217–237, February 2006. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661106000073.

Tian:2008:LRS

[TKWI08] Yongjun Tian, Hideaki Kidokoro, Tatsuro Watanabe, and Naoki Iguchi. The late 1980s regime shift in the ecosystem of Tsushima warm current in the Japan/East Sea: Evidence from historical data and possible mechanisms. Progress in Oceanography, 77(2–3):127–145, May 2008. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661108000505.

Tsuchiya:1989:SWP

[TLF^{+}89] Mizuki Tsuchiya, Roger Lukas, Rana A. Fine, Eric Firing, and Eric Lindstrom. Source waters of the Pacific Equatorial Undercurrent. Progress in Oceanography, 23(2):101–147, ???? 1989. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661189900128.

Tinker:2015:VEM

[TLH^{+}15] Jonathan Tinker, Jason Lowe, Jason Holt, Anne Pardaens, and Andy Wiltshire. Validation of an ensemble modelling system for climate projections for the northwest European shelf seas. Progress in Oceanography, 138 (part A)(??):211–237, November 2015. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661115001536.

Tilstone:2017:MPP

[TLM^{+}17] Gavin H. Tilstone, Priscila K. Lange, Ankita Misra, Robert J. W. Brewin, and Terry Cain. Micro-phytoplankton photosynthesis, primary production and potential export production in the Atlantic Ocean. Progress in Oceanography, 158 (??):109–129, November 2017. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661116300374.

Tinker:2016:UCP

[TLP^{+}16] Jonathan Tinker, Jason Lowe, Anne Pardaens, Jason Holt, and Rosa Barciela. Uncertainty in climate projections for the 21st century northwest European shelf seas. Progress in Oceanography, 148(??):56–73, November 2016. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661116301847.

Takahashi:2013:MFA

[TM13] Daisuke Takahashi and Akihiko Morimoto. Mean field and annual variation of surface flow in the East China Sea as revealed by combining satellite altimeter and drifter data.

Progress in Oceanography, 111(??):125–139, April 2013. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661113000177.

Torrejon-Magallanes:2021:MPC

[TMÁGC^{+}21] Josymar Torrejón-Magallanes, Luis Enrique Ángeles-González, Jorge Csirke, Marilú Bouchon, Enrique Morales-Bojórquez, and Francisco Arreguín-Sánchez. Modeling the Pacific chub mackerel (Scomber japonicus) ecological niche and future scenarios in the northern Peruvian Current System. Progress in Oceanography, 197(??):??, September–October 2021. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661121001567.

Thoppil:2016:CSE

[TMH^{+}16] Prasad G. Thoppil, E. Joseph Metzger, Harley E. Hurlburt, Ole Martin Smedstad, and Hiroshi Ichikawa. The current system east of the Ryukyu Islands as revealed by a global ocean reanalysis. Progress in Oceanography, 141(??):239–258, February 2016. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661116000045.

Tomczak:2009:ATN

[TMKJ^{+}09] Maciej T. Tomczak, Bärbel Müller-Karulis, Leili Järv, Jonne Kotta, Georg Martin, Atis Minde, Arno Põllumäe, Arturas Razinkovas, Solvita Strake, Martynas Bucas, and Thorsten Blenckner. Analysis of trophic networks and carbon flows in south-eastern Baltic coastal ecosystems. Progress in Oceanography, 81(1–4):111–131, April 2009. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661109000251.

Takahashi:2012:STF

[TMN^{+}12] Daisuke Takahashi, Akihiko Morimoto, Tetsuya Nakamura, Takuji Hosaka, Yoshihisa Mino, Vu Hai Dang, and Toshiro Saino. Short-term flow and water temperature fluctuations in Sagami Bay, Japan, associated with variations of the Kuroshio during the non-large-meander path. Progress in Oceanography, 105(??):47–60, October 2012. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661112000456.

Teles-Machado:2016:CNIb

[TMPM^{+}16a] Ana Teles-Machado, Álvaro Peliz, James C. McWilliams, Xavier Couvelard, and Isabel Ambar. Circulation on the Northwestern Iberian margin: Vertical structure and seasonality of the alongshore flows. Progress in Oceanography, 140(??):134–153, January 2016. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661115001184.

Teles-Machado:2016:CNIa

[TMPM^{+}16b] Ana Teles-Machado, Álvaro Peliz, James C. McWilliams, Jesus Dubert, and Bernard Le Cann. Circulation on the Northwestern Iberian margin: Swoddies. Progress in Oceanography, 140(??):116–133, January 2016. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661115002013.

Torrado:2021:IIE

[TMR^{+}21] Héctor Torrado, Baptiste Mourre, Núria Raventos, Carlos Carreras, Joaquín Tintoré, Marta Pascual, and Enrique Macpherson. Impact of individual early life traits in larval dispersal: a multispecies approach using backtracking models. Progress in Oceanography, 192(??):??, March 2021. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661121000082.

Tapia:2009:TIU

[TNC^{+}09] Fabián J. Tapia, Sergio A. Navarrete, Manuel Castillo, Bruce A. Menge, Juan C. Castilla, John Largier, Evie A. Wieters, Bernardo L. Broitman, and John A. Barth. Thermal indices of upwelling effects on inner-shelf habitats. Progress in Oceanography, 83(1–4):278–287, December 2009. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661109000949.

Testa:2022:PEK

[TNGP22] G. Testa, S. Neira, R. Giesecke, and A. Piñones. Projecting environmental and krill fishery impacts on the Antarctic Peninsula food web in 2100. Progress in Oceanography, 206(??):??, August 2022. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661122001215.

Tsumune:2005:PBS

[TNS^{+}05] Daisuke Tsumune, Jun Nishioka, Akifumi Shimamoto, Shigenobu Takeda, and Atsushi Tsuda. Physical behavior of the SEEDS iron-fertilized patch by sulphur hexafluoride tracer release. Progress in Oceanography, 64(2–4):111–127, February 2005. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S007966110500025X.

Tozawa:2024:OFD

[TNY^{+}24] Manami Tozawa, Daiki Nomura, Kaihe Yamazaki, Masaaki Kiuchi, Daisuke Hirano, Shigeru Aoki, Hiroko Sasaki, and Hiroto Murase. Oceanographic factors determining the distribution of nutrients and primary production in the subpolar Southern Ocean. Progress in Oceanography, 225(??): ??, July 2024. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661124000727.

Takikawa:2012:SVC

[TOiF^{+}12] Tetsutaro Takikawa, Goh Onitsuka, Ken ichi Fukudome, Jong-Hwan Yoon, and Akihiko Morimoto. Seasonal variation of counterclockwise eddies downstream of the Tsushima Islands. Progress in Oceanography, 105(??):30–37, October 2012. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661112000390.

Takasuka:2008:CST

[TOKLC08] Akinori Takasuka, Yoshioki Oozeki, Hiroshi Kubota, and Salvador E. Lluch-Cota. Contrasting spawning temperature optima: Why are anchovy and sardine regime shifts synchronous across the North Pacific? Progress in Oceanography, 77(2–3):225–232, May 2008. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661108000578.

Tolmazin:1985:CCOa

[Tol85a] D. Tolmazin. Changing coastal oceanography of the Black Sea. I: Northwestern Shelf. Progress in Oceanography, 15(4):217– 276, ???? 1985. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661185900382.

Tolmazin:1985:CCOb

[Tol85b] D. Tolmazin. Changing coastal oceanography of the Black Sea. II: Mediterranean effluent. Progress in Oceanography, 15(4):277–316, ???? 1985. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0079661185900394.

Tomczak:1981:AMF

[Tom81a] Matthias Tomczak. An analysis of mixing in the frontal zone of South and North Atlantic Central Water off North-West Africa. Progress in Oceanography, 10(3):173–192, ???? 1981. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0079661181900112.

Tomczak:1981:MPE

[Tom81b] Matthias Tomczak. A multi-parameter extension of temperature/salinity diagram techniques for the analysis of nonisopycnal mixing. Progress in Oceanography, 10(3):147–171, ???? 1981. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661181900100.

Tselepides:2000:CPI

[TP00] Anastasios Tselepides and Thalia Polychronaki. The CINCS project: introduction. Progress in Oceanography, 46(2–4):85– 88, August 2000. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661100000136.

Tselepides:2000:OMC

[TPM^{+}00] Anastasios Tselepides, Thalia Polychronaki, Daniela Marrale, Ioanna Akoumianaki, Antonio Dell’Anno, Antonio Pusceddu, and Roberto Danovaro. Organic matter composition of the continental shelf and bathyal sediments of the Cretan Sea (NE Mediterranean). Progress in Oceanography, 46(2–4):311–344, August 2000. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661100000240.

Theodosi:2018:SAA

[TPN^{+}18] C. Theodosi, C. Panagiotopoulos, A. Nouara, P. Zarmpas, P. Nicolaou, K. Violaki, M. Kanakidou, R. Sempéré, and N. Mihalopoulos. Sugars in atmospheric aerosols over the Eastern Mediterranean. Progress in Oceanography, 163(??):70– 81, April 2018. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661116301264.

Tselepides:2000:MCS

[TPP^{+}00] Anastasios Tselepides, Konstantia-N. Papadopoulou, Dimitris Podaras, Wanda Plaiti, and Drosos Koutsoubas. Macrobenthic community structure over the continental margin of Crete (South Aegean Sea, NE Mediterranean). Progress in Oceanography, 46(2–4):401–428, August 2000. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661100000276.

Tesi:2010:LAO

[TPPG10] T. Tesi, P. Puig, A. Palanques, and M. A. Goñi. Lateral advection of organic matter in cascading-dominated submarine canyons. Progress in Oceanography, 84(3–4):185–203, March 2010. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661109001736.

Trainer:2010:DIH

[TPRS10] V. L. Trainer, G. C. Pitcher, B. Reguera, and T. J. Smayda. The distribution and impacts of harmful algal bloom species in eastern boundary upwelling systems. Progress in Oceanography, 85(1–2):33–52, April 2010. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661110000145.

Townsend:2023:WWG

[TPTM23] David W. Townsend, Neal R. Pettigrew, Maura A. Thomas, and Stephen Moore. Warming waters of the Gulf of Maine: the role of Shelf, Slope and Gulf Stream Water masses. Progress in Oceanography, 215(??):??, July 2023. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661123000733.

Thingstad:1999:CMB

[TR99] T. Frede Thingstad and Fereidoun Rassoulzadegan. Conceptual models for the biogeochemical role of the photic zone microbial food web, with particular reference to the Mediterranean Sea. Progress in Oceanography, 44(1–3):271–286, August 1999. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661199000294.

Tecchio:2013:SFD

[TRLA^{+}13] Samuele Tecchio, Eva Ramírez-Llodra, Jacopo Aguzzi, Anna Sanchez-Vidal, M. Mar Flexas, Francisco Sardà, and Joan B. Company. Seasonal fluctuations of deep megabenthos: Finding evidence of standing stock accumulation in a flux-rich continental slope. Progress in Oceanography, 118(??):188–198, November 2013. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661113001286.

Twining:2015:MCP

[TRMV15] Benjamin S. Twining, Sara Rauschenberg, Peter L. Morton, and Stefan Vogt. Metal contents of phytoplankton and labile particulate material in the North Atlantic Ocean. Progress in Oceanography, 137 (part A)(??):261–283, September 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661115001524.

Torres:2023:ISA

[TRP^{+}23] Rodrigo Torres, Brian Reid, Gemita Pizarro, Máximo Frangópulos, Emilio Alarcón, Magdalena Márquez, Francisco Díaz-Rosas, Eduardo Menschel, Humberto E. González, Paulo Moreno-Meynard, Paulina Montero, Hernán Pacheco, Marco Pinto-Torres, Cesar Alarcón, Rodrigo Ibañez, and Jon Hawkings. Iron and silicic acid addition effects on early spring macronutrient drawdown and biogenic silica production of Patagonia estuarine waters. Progress in Oceanography, 214(??): ??, June 2023. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661123000253.

Tang:2004:CWM

[TRY^{+}04] Charles C. L. Tang, Charles K. Ross, Tom Yao, Brian Petrie, Brendan M. DeTracey, and Ewa Dunlap. The circulation, water masses and sea-ice of Baffin Bay. Progress in Oceanography, 63(4):183–228, December 2004. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661104001247.

Travers:2010:STV

[TS10] M. Travers and Y.-J. Shin. Spatio-temporal variability in fishinduced predation mortality on plankton: a simulation approach using a coupled trophic model of the Benguela ecosystem. Progress in Oceanography, 84(1–2):118–120, January 2010. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661109001475.

Taucher:2022:SCO

[TSAM^{+}22] Jan Taucher, Paul Stange, María Algueró-Muñiz, Lennart T. Bach, Alice Nauendorf, Regina Kolzenburg, Jan Büdenbender, and Ulf Riebesell. In situ camera observations reveal major role of zooplankton in modulating marine snow formation during an upwelling-induced plankton bloom. Progress in Oceanography, 164(??):75–88, ???? 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661117302240.

Trudnowska:2018:SVS

[TSBS18] E. Trudnowska, S. Sagan, and K. Błachowiak-Samołyk. Spatial variability and size structure of particles and plankton in the Fram Strait. Progress in Oceanography, 168(??):1–12, November 2018. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661117302446.

Timothy:2003:SFS

[TSC03] David A. Timothy, Maureen Y. S. Soon, and Stephen E. Calvert. Settling fluxes in Saanich and Jervis Inlets, British Columbia, Canada: sources and seasonal patterns. Progress in Oceanography, 59(1):31–73, October 2003. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661103001253.

Trifoglio:2022:ACP

[TSFA22] Noelia L. Trifoglio, Héctor F. Olguín Salinas, Claudio A. Franzosi, and Viviana A. Alder. Annual cycle of phytoplankton, protozoa and diatom species from Scotia Bay (South Orkney Islands, Antarctica): Community structure prior to, during and after an anomalously low sea ice year. Progress in Oceanography, 204(??):??, June 2022. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661122000684.

Tengberg:2004:IBF

[TSG^{+}04] A. Tengberg, H. Stahl, G. Gust, V. Müller, U. Arning, H. Andersson, and P. O. J. Hall. Intercalibration of benthic flux chambers I. Accuracy of flux measurements and influence of chamber hydrodynamics. Progress in Oceanography, 60(1):1– 28, January 2004. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661104000023.

Tommasi:2017:MLM

[TSH^{+}17] Desiree Tommasi, Charles A. Stock, Alistair J. Hobday, Rick Methot, Isaac C. Kaplan, J. Paige Eveson, Kirstin Holsman, Timothy J. Miller, Sarah Gaichas, Marion Gehlen, Andrew Pershing, Gabriel A. Vecchi, Rym Msadek, Tom Delworth, C. Mark Eakin, Melissa A. Haltuch, Roland Séférian, Claire M. Spillman, Jason R. Hartog, Samantha Siedlecki, Jameal F. Samhouri, Barbara Muhling, Rebecca G. Asch, Malin L. Pinsky, Vincent S. Saba, Sarah B. Kapnick, Carlos F. Gaitan, Ryan R. Rykaczewski, Michael A. Alexander, Yan Xue, Kathleen V. Pegion, Patrick Lynch, Mark R.

Payne, Trond Kristiansen, Patrick Lehodey, and Francisco E. Werner. Managing living marine resources in a dynamic environment: the role of seasonal to decadal climate forecasts. Progress in Oceanography, 152(??):15–49, March 2017. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661116301586.

Tseng:2012:VKC

[TSJ^{+}12] Yu-Heng Tseng, Mao-Lin Shen, Sen Jan, David E. Dietrich, and Chia-Ping Chiang. Validation of the Kuroshio Current System in the dual-domain Pacific Ocean model framework. Progress in Oceanography, 105(??):102–124, October 2012. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661112000365.

Travers:2007:TEE

[TSJC07] M. Travers, Y.-J. Shin, S. Jennings, and P. Cury. Towards end-to-end models for investigating the effects of climate and fishing in marine ecosystems. Progress in Oceanography, 75(4):751–770, December 2007. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661107001425.

Thingstad:2010:SBP

[TSL10] T. Frede Thingstad, Espen Strand, and Aud Larsen. Stepwise building of plankton functional type (PFT) models: a feasible route to complex models? Progress in Oceanography, 84(1–2):6–15, January 2010. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661109001347.

Tsuda:2005:MRI

[TSNO05] Atsushi Tsuda, Hiroaki Saito, Jun Nishioka, and Tsuneo Ono. Mesozooplankton responses to iron-fertilization in the western subarctic Pacific (SEEDS2001). Progress in Oceanography, 64(2–4):237–251, February 2005. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661105000327.

Tweddle:2013:ENF

[TSP^{+}13] Jacqueline F. Tweddle, Jonathan Sharples, Matthew R. Palmer, Keith Davidson, and Sharon McNeill. Enhanced nutrient fluxes at the shelf sea seasonal thermocline caused by stratified flow over a bank. Progress in Oceanography, 117(??):37–47, October 2013. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661113001067.

Torres:2014:SAE

[TSRF14] Rodrigo Torres, Nelson Silva, Brian Reid, and Máximo Frangopulos. Silicic acid enrichment of subantarctic surface water from continental inputs along the Patagonian archipelago interior sea (41–56^{o}S). Progress in Oceanography, 129(??):50–61, December 2014. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661114001736.

Thompson:2012:LPS

[TSS^{+}12] Sarah Ann Thompson, William J. Sydeman, Jarrod A. Santora, Bryan A. Black, Robert M. Suryan, John Calambokidis, William T. Peterson, and Steven J. Bograd. Linking predators to seasonality of upwelling: Using food web indicators and path analysis to infer trophic connections. Progress in Oceanography, 101(1):106–120, August 2012. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661112000079.

Tilstone:2017:ECE

[TŠT^{+}17] Gavin Tilstone, Barbora Šedivá, Glen Tarran, Radek Kaňa, and Ondřej Prášil. Effect of CO_{2} enrichment on phytoplankton photosynthesis in the North Atlantic sub-tropical gyre. Progress in Oceanography, 158(??):76–89, November 2017. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661116300064.

Tsuchiya:1986:TCU

[Tsu86] Mizuki Tsuchiya. Thermostads and circulation in the upper layer of the Atlantic Ocean. Progress in Oceanography, 16(4):235–267, ???? 1986. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0079661186900406.

Takeda:2005:SIE

[TT05] Shigenobu Takeda and Atsushi Tsuda. An in situ ironenrichment experiment in the western subarctic Pacific (SEEDS): Introduction and summary. Progress in Oceanography, 64(2–4):95–109, February 2005. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661105000248.

Tam:2008:TMN

[TTB^{+}08a] Jorge Tam, Marc H. Taylor, Verónica Blaskovic, Pepe Espinoza, R. Michael Ballón, Erich Díaz, Claudia Wosnitza-Mendo, Juan Argüelles, Sara Purca, Patricia Ayón, Luis Quipuzcoa, Dimitri Gutiérrez, Elisa Goya, Noemí Ochoa, and Matthias Wolff. Trophic modeling of the Northern Humboldt Current Ecosystem, Part I: Comparing trophic linkages under La Niña and El Niño conditions. Progress in Oceanography, 79(2–4):352–365, October 2008. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661108001870.

Taylor:2008:TMN

[TTB^{+}08b] Marc H. Taylor, Jorge Tam, Verónica Blaskovic, Pepe Espinoza, R. Michael Ballón, Claudia Wosnitza-Mendo, Juan Argüelles, Erich Díaz, Sara Purca, Noemi Ochoa, Patricia Ayón, Elisa Goya, Dimitri Gutiérrez, Luis Quipuzcoa, and Matthias Wolff. Trophic modeling of the Northern Humboldt Current Ecosystem, Part II: Elucidating ecosystem dynamics from 1995 to 2004 with a focus on the impact of ENSO. Progress in Oceanography, 79(2–4):366–378, October 2008. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661108001882.

Trotter:2022:UDS

[TTF^{+}22] Julie A Trotter, Marco Taviani, Federica Foglini, Aleksey Sadekov, Grzegorz Skrzypek, Claudio Mazzoli, Alessandro Remia, Nadia Santodomingo, Giorgio Castellan, Malcolm McCulloch, Charitha Pattiaratchi, and Paolo Montagna. Unveiling deep-sea habitats of the Southern Oceanfacing submarine canyons of southwestern Australia. Progress in Oceanography, 209(??):??, December 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S007966112200163X.

Talley:2004:AJE

[TTL^{+}04] Lynne D. Talley, Pavel Tishchenko, Vladimir Luchin, Alexander Nedashkovskiy, Sergey Sagalaev, Dong-Jin Kang, Mark Warner, and Dong-Ha Min. Atlas of Japan (East) Sea hydrographic properties in summer, 1999. Progress in Oceanography, 61(2–4):277–348, May 2004. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661104000679.

Tanner:2017:MLD

[TTMM^{+}17] Susanne E. Tanner, Ana Teles-Machado, Filipe Martinho, Álvaro Peliz, and Henrique N. Cabral. Modelling larval dispersal dynamics of common sole (Solea solea) along the western Iberian coast. Progress in Oceanography, 156(??):78–90, August 2017. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661116302518.

Turekian:1965:EAP

[Tur65] Karl K. Turekian. Estimates of the average Pacific deepsea clay accumulation rate from material balance calculations. Progress in Oceanography, 4(??):227–244, September 2, 1965. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0079661165900522.

Turley:1999:CMS

[Tur99] Carol M. Turley. The changing Mediterranean Sea — a sensitive ecosystem? Progress in Oceanography, 44(1–3):387–400, August 1999. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661199000336.

Turner:2015:ZFP

[Tur15] Jefferson T. Turner. Zooplankton fecal pellets, marine snow, phytodetritus and the ocean’s biological pump. Progress in Oceanography, 130(??):205–248, January 2015. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661114001281.

Tsimplis:1999:CDW

[TVD^{+}99] M. N. Tsimplis, A. F. Velegrakis, P. Drakopoulos, A. Theocharis, and M. B. Collins. Cretan deep water outflow into the Eastern Mediterranean. Progress in Oceanography, 44(4):531–551, December 1999. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661199000427.

Thomsen:2002:PBB

[TvG02] L. Thomsen, T. vanWeering, and G. Gust. Processes in the benthic boundary layer at the Iberian continental margin and their implication for carbon mineralization. Progress in Oceanography, 52(??):315–329, ???? 2002. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661102000137.

Thiele:2023:SBA

[TVT^{+}23] Stefan Thiele, Anna Vader, Stuart Thomson, Karoline Saubrekka, Elzbieta Petelenz, Hilde Rief Armo, Oliver Müller, Lasse Olsen, Gunnar Bratbak, and Lise Øvreås. The summer bacterial and archaeal community composition of the northern Barents Sea. Progress in Oceanography, 215(??): ??, July 2023. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661123000976.

Thomsen:1998:STV

[TvW98] L. Thomsen and Tj. C. E. van Weering. Spatial and temporal variability of particulate matter in the benthic boundary layer at the N.W. European Continental Margin (Goban Spur). Progress in Oceanography, 42(1–4):61–76, December 1998. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661198000287.

Thompson:2011:NOB

[TWAL^{+}11] P. A. Thompson, K. Wild-Allen, M. Lourey, C. Rousseaux, A. M. Waite, M. Feng, and L. E. Beckley. Nutrients in an oligotrophic boundary current: Evidence of a new role for the Leeuwin Current. Progress in Oceanography, 91(4):345–359, December 2011. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661111000267.

Timothy:2013:CSF

[TWBC^{+}13] D. A. Timothy, C. S. Wong, J. E. Barwell-Clarke, J. S. Page, L. A. White, and R. W. Macdonald. Climatology of sediment flux and composition in the subarctic Northeast Pacific Ocean with biogeochemical implications. Progress in Oceanography, 116(??):95–129, September 2013. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661113001055.

Taylor:2008:CTF

[TWMY08] Marc H. Taylor, Matthias Wolff, Jaime Mendo, and Carmen Yamashiro. Changes in trophic flow structure of Independence Bay (Peru) over an ENSO cycle. Progress in Oceanography, 79(2–4):336–351, October 2008. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661108001869.

Tanaka:2014:ITT

[TYO^{+}14] Yuki Tanaka, Ichiro Yasuda, Satoshi Osafune, Takahiro Tanaka, Jun Nishioka, and Yuri N. Volkov. Internal tides and turbulent mixing observed in the Bussol Strait. Progress in Oceanography, 126(??):98–108, August 2014. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661114000573.

Tselepides:2000:DNP

[TZP^{+}00] A. Tselepides, V. Zervakis, T. Polychronaki, R. Danovaro, and G. Chronis. Distribution of nutrients and particulate organic matter in relation to the prevailing hydrographic features of the Cretan Sea (NE Mediterranean).

Progress in Oceanography, 46(2–4):113–142, August 2000. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S007966110000015X.

Udalov:2005:DRP

[UAM05] A. A. Udalov, A. I. Azovsky, and V. O. Mokievsky. Depthrelated pattern in nematode size: What does the depth itself really mean? Progress in Oceanography, 67(1–2):1–23, October 2005. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661105000972.

Unal:2010:BSP

[UB10] Ebru Unal and Ann Bucklin. Basin-scale population genetic structure of the planktonic copepod Calanus finmarchicus in the North Atlantic Ocean. Progress in Oceanography, 87(1–4):175–185, October 2010. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661110001291.

Ueno:2023:ROM

[UBB^{+}23] Hiromichi Ueno, Annalisa Bracco, John A. Barth, Maxim V. Budyansky, Daisuke Hasegawa, Sachihiko Itoh, Sung Yong Kim, Carol Ladd, Xiaopei Lin, Young-Gyu Park, Sergey Prants, Tetjana Ross, Irina I. Rypina, Yoshikazu Sasai, Olga O. Trusenkova, Elena I. Ustinova, and Yisen Zhong. Review of oceanic mesoscale processes in the North Pacific: Physical and biogeochemical impacts. Progress in Oceanography, 212(??): ??, March 2023. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661122002142.

Ursella:2018:EZV

[UCB^{+}18] Laura Ursella, Vanessa Cardin, Mirna Batistić, Rade Garić, and Miroslav Gačić. Evidence of zooplankton vertical migration from continuous Southern Adriatic buoy current-meter records. Progress in Oceanography, 167(??):78–96, October 2018. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661118300211.

Uesaka:2022:OWO

[UGY^{+}22] Leo Uesaka, Yusuke Goto, Yoshinari Yonehara, Kosei Komatsu, Masaru Naruoka, Henri Weimerskirch, Katsufumi Sato, and Kentaro Q. Sakamoto. Ocean wave observation utilizing motion records of seabirds. Progress in Oceanography, 200(??): ??, January 2022. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661121001968.

Utevsky:2019:NSG

[UKK^{+}19] Serge Utevsky, Anastasiia Kovalchuk, Nataliia Kovalchuk, Andriy Utevsky, and Alexei V. Chernyshev. A new species of the genus Johanssonia Selensky, 1914 (Hirudinea: Piscicolidae) collected in the Kuril–Kamchatka Trench at the greatest depth ever recorded for fish leeches. Progress in Oceanography, 176(??):??, September 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661119302010.

Uehara:2014:RPS

[UKM^{+}14] Hiroki Uehara, Andrey A. Kruts, Humio Mitsudera, Tomohiro Nakamura, Yuri N. Volkov, and Masaaki Wakatsuchi. Remotely propagating salinity anomaly varies the source of North Pacific ventilation. Progress in Oceanography, 126(??):80–97, August 2014. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661114000640.

Uchimoto:2014:SHC

[UNN^{+}14] Keisuke Uchimoto, Tomohiro Nakamura, Jun Nishioka, Humio Mitsudera, Kazuhiro Misumi, Daisuke Tsumune, and Masaaki Wakatsuchi. Simulation of high concentration of iron in dense shelf water in the Okhotsk Sea. Progress in Oceanography, 126(??):194–210, August 2014. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661114000664.

Ursella:2021:DLS

[UPPS^{+}21] Laura Ursella, Sara Pensieri, Enric Pallàs-Sanz, Sharon Z. Herzka, Roberto Bozzano, Miguel Tenreiro, Vanessa Cardin, Julio Candela, and Julio Sheinbaum. Diel, lunar and seasonal vertical migration in the deep western Gulf of Mexico evidenced from a long-term data series of acoustic backscatter. Progress in Oceanography, 195(??):??, July 2021. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661121000495.

Uncles:2015:ESE

[USH15a] R. J. Uncles, J. A. Stephens, and C. Harris. Estuaries of southwest England: Salinity, suspended particulate matter, loss-on-ignition and morphology. Progress in Oceanography, 137(??):385–408, September 2015. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S007966111500097X.

Uncles:2015:PPC

[USH15b] R. J. Uncles, J. A. Stephens, and C. Harris. Physical processes in a coupled bay–estuary coastal system: Whitsand Bay and Plymouth Sound. Progress in Oceanography, 137(??):360–384, September 2015. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661115000865.

vanAken:1996:HTF

[vAB96] H. M. van Aken and G. Becker. Hydrography and throughflow in the north-eastern North Atlantic Ocean: the NANSEN project. Progress in Oceanography, 38(4):297–346, ???? 1996. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661197000050.

Venegas:2024:USV

[VAEP24] Ana Venegas, Pierre-Amaël Auger, Ruben Escribano, and Carolina Parada. Understanding seasonal variability of mesozooplankton biomass in the upwelling system of centralsouthern Chile: a modelling approach. Progress in Oceanography, 220(??):??, January 2024. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661123002367.

Veny:2022:SCV

[VAGMDRS22] Marta Veny, Borja Aguiar-González, Ángeles Marrero-Díaz, and Ángel Rodríguez-Santana. Seasonal circulation and volume transport of the Bransfield Current. Progress in Oceanography, 204(??):??, June 2022. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661122000568.

Valette:1999:IBR

[Val99a] François Valette. Interactions between research and policy.: From wishes to realities, in a changing world. Progress in Oceanography, 44(1–3):433–456, August 1999. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661199000385.

Vallega:1999:RIS

[Val99b] Adalberto Vallega. The role of integrated scientific approach facing the changing ocean policy. The case of the Mediterranean. Progress in Oceanography, 44(1–3):411–431, August 1999. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S007966119900035X.

Vittori:2014:BLB

[VB14] Giovanna Vittori and Paolo Blondeaux. The boundary layer at the bottom of a solitary wave and implications for sediment transport. Progress in Oceanography, 120(??):399–409, January 2014. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661113001717.

Violaki:2018:OPA

[VBA^{+}18] Kalliopi Violaki, François Bourrin, Dominique Aubert, Giorgos Kouvarakis, Nicole Delsaut, and Nikos Mihalopoulos. Organic phosphorus in atmospheric deposition over the Mediterranean Sea: an important missing piece of the phosphorus cycle. Progress in Oceanography, 163(??):50–58, April 2018. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661117300356.

Vladoiu:2021:CME

[VBAC^{+}21] Anda Vladoiu, Pascale Bouruet-Aubertot, Yannis Cuypers, Bruno Ferron, Katrin Schroeder, Mireno Borghini, and Stephane Leizour. Contrasted mixing efficiency in energetic versus quiescent regions: Insights from microstructure measurements in the Western Mediterranean Sea. Progress in Oceanography, 195(??):??, July 2021. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661121000811.

Visser:2020:SSW

[VBC^{+}20] André W. Visser, Philipp Brun, Subhendu Chakraborty, Tim Spaanheden Dencker, P. Daniël van Denderen, Rob van Gemert, Hans van Someren Gréve, Irene Heilmann, Mark Wejlemann Holm, Sigrún Huld Jónasdóttir, Kasia M. Kenitz, Thomas Kiørboe, Martin Lindegren, Patrizio Mariani, Lasse Tor Nielsen, Marina Pancic, Mark Payne, Lauréne Pécuchet, Nicolas Azaña Schnedler-Meyer, Uffe Høgsbro Thygesen, Anna Törnroos, and Ken H. Andersen. Seasonal strategies in the world’s oceans. Progress in Oceanography, 189(??):??, November–December 2020. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661120302032.

Vieira:2020:DSS

[VBJ^{+}20] Rui P. Vieira, Brian J. Bett, Daniel O. B. Jones, Jennifer M. Durden, Kirsty J. Morris, Marina R. Cunha, Clive N. Trueman, and Henry A. Ruhl. Deep-sea sponge aggregations (Pheronema carpenteri) in the Porcupine Seabight (NE Atlantic) potentially degraded by demersal fishing. Progress in Oceanography, 183(??):??, April 2020. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661119302563.

Vezina:2004:IRE

[VBL04] Alain F. Vézina, Florence Berreville, and Svetlana Loza. Inverse reconstructions of ecosystem flows in investigating regime shifts: impact of the choice of objective function. Progress in Oceanography, 60(2–4):321–341, February 2004.

CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661104000229.

Vandenbulcke:2009:SET

[VBL^{+}09] L. Vandenbulcke, J.-M. Beckers, F. Lenartz, A. Barth, P.- M. Poulain, M. Aidonidis, J. Meyrat, F. Ardhuin, M. Tonani, C. Fratianni, L. Torrisi, D. Pallela, J. Chiggiato, M. Tudor, J. W. Book, P. Martin, G. Peggion, and M. Rixen. Super-ensemble techniques: Application to surface drift prediction. Progress in Oceanography, 82(3):149–167, September 2009. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661109000500.

Valdes:2021:TDC

[VBL^{+}21] Luis Valdés, Antonio Bode, Mikel Latasa, Enrique Nogueira, Raquel Somavilla, Marta M. Varela, César González-Pola, and Gerardo Casas. Three decades of continuous ocean observations in North Atlantic Spanish waters: the RADI-ALES time series project, context, achievements and challenges. Progress in Oceanography, 198(??):??, November 2021. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661121001555.

Vidya:2021:CHE

[VBM21] P. J. Vidya, M. Balaji, and R. Mani Murali. Cyclone Hudhud-eddy induced phytoplankton bloom in the northern Bay of Bengal using a coupled model. Progress in Oceanography, 197(??):??, September–October 2021. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661121001178.

Velez-Belchi:2005:OWA

[VBVYT05] P. Vélez-Belchí, M. Vargas-Yáñez, and J. Tintoré. Observation of a western Alborán gyre migration event. Progress in Oceanography, 66(2–4):190–210, August 2005. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661105000704.

VanWambeke:2000:HBP

[VCB^{+}00] France Van Wambeke, Urania Christaki, Micheline Bianchi, Stella Psarra, and Anastasios Tselepides. Heterotrophic bacterial production in the Cretan Sea (NE Mediterranean). Progress in Oceanography, 46(2–4):205–216, August 2000. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661100000197.

Verlaan:2004:CAC

[VCM04] Philomène A. Verlaan, David S. Cronan, and Charles L. Morgan. A comparative analysis of compositional variations in and between marine ferromanganese nodules and crusts in the South Pacific and their environmental controls. Progress in Oceanography, 63(3):125–158, November 2004. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661104001260.

Vanreusel:2001:EER

[VCSG^{+}01] A. Vanreusel, N. Cosson-Sarradin, A. J. Gooday, G. L. J. Paterson, J. Galéron, M. Sibuet, and M. Vincx. Evidence for episodic recruitment in a small opheliid polychaete species from the abyssal NE Atlantic. Progress in Oceanography, 50(??):285–301, ???? 2001. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661101000581.

Veron:2020:MCS

[VDB^{+}20] Matthieu Véron, Erwan Duhamel, Michel Bertignac, Lionel Pawlowski, and Martin Huret. Major changes in sardine growth and body condition in the Bay of Biscay between 2003 and 2016: Temporal trends and drivers. Progress in Oceanography, 182(??):??, March 2020. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661120300124.

Vazquez-Dominguez:2008:MPA

[VDDA^{+}08] Evaristo Vázquez-Domínguez, Carlos M. Duarte, Susana Agustí, Klaus Jürgens, Dolors Vaqué, and Josep M. Gasol.

Microbial plankton abundance and heterotrophic activity across the central Atlantic Ocean. Progress in Oceanography, 79(1):83–94, October 2008. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S007966110800147X.

Ventura-Dominguez:2022:CSB

[VDGGD^{+}22] Phalom D. Ventura-Domínguez, Jaime Gómez-Gutiérrez, Roxana De Silva-Dávila, Eduardo González-Rodríguez, Arturo R. Sánchez-Uvera, Ángel A. Silveyra-Bustamante, Armando Trasviña-Castro, Octavio Aburto-Oropeza, and Carlos Sánchez. Community structure and biomass of larval euphausiids and postlarval mysids in relation to anomalous warming events during a weekly time series (2014–2017) at Cabo Pulmo National Park, Mexico. Progress in Oceanography, 208(??):??, November 2022. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661122001409.

Vanucci:2001:MAA

[VDP^{+}01] S. Vanucci, A. Dell’Anno, A. Pusceddu, M. Fabiano, R. S. Lampitt, and R. Danovaro. Microbial assemblages associated with sinking particles in the Porcupine Abyssal Plain (NE Atlantic Ocean). Progress in Oceanography, 50(??):105–121, ???? 2001. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661101000507.

vanderSpoel:1994:BBP

[vdS94a] S. van der Spoel. The basis for boundaries in pelagic biogeography. Progress in Oceanography, 34(2–3):121–133, ???? 1994. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0079661194900051.

vanderSpoel:1994:HPF

[vdS94b] S. van der Spoel. History, progress and future of theory in pelagic biogeography. Progress in Oceanography, 34(2–3):101– 107, ???? 1994. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661194900035.

vanderSpoel:1994:WD

[vdS94c] S. van der Spoel. A warning from the deep. Progress in Oceanography, 34(2–3):207–210, ???? 1994. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0079661194900086.

VanOostende:2018:SOC

[VDS^{+}18] N. Van Oostende, R. Dussin, C. A. Stock, A. D. Barton, E. Curchitser, J. P. Dunne, and B. B. Ward. Simulating the ocean’s chlorophyll dynamic range from coastal upwelling to oligotrophy. Progress in Oceanography, 168(??):232–247, November 2018. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661117302586.

Vernet:2021:SVR

[VEM^{+}21] Maria Vernet, Ingrid Ellingsen, Christian Marchese, Simon Bélanger, Mattias Cape, Dag Slagstad, and Patricia A. Matrai. Spatial variability in rates of net primary production (NPP) and onset of the spring bloom in Greenland shelf waters. Progress in Oceanography, 198(??):??, November 2021. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661121001403.

Venrick:2012:PCC

[Ven12] Elizabeth L. Venrick. Phytoplankton in the California Current system off southern California: Changes in a changing environment. Progress in Oceanography, 104(??):46–58, October 2012. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661112000626.

Verheye:1991:STV

[Ver91] Hans M. Verheye. Short-term variability during an anchor station study in the southern Benguela upwelling system: Abundance, distribution and estimated production of mesozooplankton with special reference to Calanoides carinatus (Krøyer, 1849). Progress in Oceanography, 28(1–2):91–119, ???? 1991. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/007966119190022E.

Verstraete:1992:SUG

[Ver92] Jean-Marc Verstraete. The seasonal upwellings in the Gulf of Guinea. Progress in Oceanography, 29(1):1–60, ???? 1992. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/007966119290002H.

vanForeest:1982:TMN

[vFB82] D. van Foreest and G. B. Brundrit. A two-mode numerical model with applications to coastal upwelling. Progress in Oceanography, 11(4):329–392, ???? 1982. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0079661182900027.

Valiente:2022:DSO

[VFCC^{+}22] S. Valiente, B. Fernández-Castro, R. Campanero, A. Marrero-Díaz, A. Rodríguez-Santana, M. D. Gelado-Cabellero, M. Nieto-Cid, A. Delgado-Huertas, J. Arístegui, and X. A. Álvarez-Salgado. Dissolved and suspended organic matter dynamics in the Cape Verde Frontal Zone (NW Africa). Progress in Oceanography, 201(??):??, February 2022. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S007966112100210X.

Vecchione:2015:ENA

[VFS^{+}15] Michael Vecchione, Tone Falkenhaug, Tracey Sutton, April Cook, Astthor Gislason, Hege Øverbø Hansen, Mikko Heino, Peter I. Miller, Uwe Piatkowski, Filipe Porteiro, Henrik Søiland, and Odd Aksel Bergstad. The effect of the North Atlantic subpolar front as a boundary in pelagic biogeography decreases with increasing depth and organism size. Progress in Oceanography, 138 (part A)(??):105–115, November 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S007966111500186X.

VanEngeland:2019:COO

[VGJ^{+}19] Tom Van Engeland, Olav Rune Godø, Espen Johnsen, Gerard C. A. Duineveld, and Dick van Oevelen. Cabled ocean observatory data reveal food supply mechanisms to a coldwater coral reef. Progress in Oceanography, 172(??):51–64, March 2019. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661118301800.

Vargas:2006:FMV

[VGLCS06] J. M. Vargas, J. García-Lafuente, J. Candela, and A. J. Sánchez. Fortnightly and monthly variability of the exchange through the Strait of Gibraltar. Progress in Oceanography, 70(2–4):466–485, August 2006. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661106000814.

Vallina:2023:SDP

[VGM^{+}23] Sergio M. Vallina, Charlie Gaborit, Celia Marrase, Josep M. Gasol, Nixon Bahamon, Michael J. Follows, Guillaume Le Gland, and Pedro Cermeño. Seasonal dynamics of phytoplankton community assembly at the Blanes Bay Microbial Observatory (BBMO), NW Mediterranean Sea. Progress in Oceanography, 219(??):??, December 2023. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661123001684.

Vihma:2009:EGS

[VH09a] Timo Vihma and Jari Haapala. Erratum to “Geophysics of sea ice in the Baltic Sea: a review” [Prog. Oceanogr. 80 (2009) 129–148]. Progress in Oceanography, 82(3):224, September 2009. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661109000512. See [VH09b].

Vihma:2009:GSI

[VH09b] Timo Vihma and Jari Haapala. Geophysics of sea ice in the Baltic Sea: a review. Progress in Oceanography, 80(3–4):129– 148, March 2009. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661109000123. See erra-

tum [VH09a].

vanHaren:2020:DSF

[vHCY^{+}20] Hans van Haren, Wu-Cheng Chi, Chu-Fang Yang, Yiing-Jang Yang, and Sen Jan. Deep sea floor observations of typhoon driven enhanced ocean turbulence. Progress in Oceanography, 184(??):??, May 2020. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661120300549.

Vuorinen:2003:TFM

[VHK03] I. Vuorinen, J. Hänninen, and G. Kornilovs. Transfer-function modelling between environmental variation and mesozooplankton in the Baltic Sea. Progress in Oceanography, 59(4):339–356, December 2003. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661103001757. See erra-

tum [VHK04].

Vuorinen:2004:ETF

[VHK04] I. Vuorinen, J. Hänninen, and G. Kornilovs. Erratum to: Transfer-function modelling between environmental variation and mesozooplankton in the Baltic Sea [Progress in Oceanography 59 (2003) 339–356]. Progress in Oceanography, 61(1):101– 102, April 2004. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661104000217. See [VHK03].

vanHaren:2014:HRT

[vHMDL14] Hans van Haren, Furu Mienis, Gerard C. A. Duineveld, and Marc S. S. Lavaleye. High-resolution temperature observations of a trapped nonlinear diurnal tide influencing cold-water corals on the logachev mounds. Progress in Oceanography, 125(??): 16–25, June 2014. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S007966111400069X.

VanOostende:2012:PCD

[VHV^{+}12] Nicolas Van Oostende, Jérôme Harlay, Bart Vanelslander, Lei Chou, Wim Vyverman, and Koen Sabbe. Phytoplankton community dynamics during late spring coccolithophore blooms at the continental margin of the Celtic Sea (North East Atlantic, 2006–2008). Progress in Oceanography, 104(??):1–16, October 2012. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661112000493.

vanHaren:2022:IWB

[vHVAT22] Hans van Haren, Gunnar Voet, Matthew H. Alford, and Daniel J. Torres. Internal wave breaking near the foot of a steep East-Pacific continental slope. Progress in Oceanography, 205 (??):??, July 2022. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661122000787.

Vidhya:2022:OMZ

[VJJ^{+}22] V. Vidhya, R. Jyothibabu, L. Jagadeesan, C. Rashid, K. T. Alok, N. Arunpandi, and R. Thirumurugan. Oxygen minimum zone copepods in the Arabian Sea and the Bay of Bengal: Their adaptations and status. Progress in Oceanography, 206(??):??, August 2022. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661122001008.

Vangriesheim:1990:NBP

[VK90] Annick Vangriesheim and Alexis Khripounoff. Near-bottom particle concentration and flux: Temporal variations observed with sediment traps and nepholometer on the Meriadzek Terrace, Bay of Biscay. Progress in Oceanography, 24(??):103– 116, ???? 1990. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/007966119090023U.

Venkatesan:1992:VLT

[VK92] M. I. Venkatesan and I. R. Kaplan. Vertical and lateral transport of organic carbon and the carbon budget in Santa Monica Basin, California. Progress in Oceanography, 30(1–4):291– 312, ???? 1992. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/007966119290016S.

Veit-Kohler:2018:OTC

[VKDS^{+}18] Gritta Veit-Köhler, Stephan Durst, Jan Schuckenbrock, Freija Hauquier, Laura Durán Suja, Boris Dorschel, Ann Vanreusel, and Pedro Martínez Arbizu. Oceanographic and topographic conditions structure benthic meiofauna communities in the Weddell Sea, Bransfield Strait and Drake Passage (Antarctic). Progress in Oceanography, 162(??):240–256, March 2018. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S007966111630204X.

Veit-Kohler:2013:CNS

[VKGP^{+}13] Gritta Veit-Köhler, Katja Guilini, Ilka Peeken, Petra Quillfeldt, and Christoph Mayr. Carbon and nitrogen stable isotope signatures of deep-sea meiofauna follow oceanographical gradients across the Southern Ocean. Progress in Oceanography, 110(??):69–79, March 2013. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661113000025.

Vijayan:2023:VAA

[VKJ^{+}23] Apsara P Vijayan, Siby Kurian, Duphrin Joseph, Megha Dixon, Harshada Kankonkar, Rakhee Khandeparker, Supriya Karapurkar, and Anand Methar. Variation of amino acid in the sinking particulates in the northeastern Arabian Sea and the northern Bay of Bengal. Progress in Oceanography, 219(??):??, December 2023. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661123002112.

Velaoras:2015:RIT

[VKT15] Dimitris Velaoras, George Krokos, and Alexander Theocharis. Recurrent intrusions of transitional waters of Eastern Mediterranean origin in the Cretan Sea as a tracer of Aegean Sea dense water formation events. Progress in Oceanography, 135(??):113–124, June 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661115000701.

Valle-Levinson:2014:TVM

[VLCCP14] Arnoldo Valle-Levinson, Leonardo Castro, Mario Cáceres, and Oscar Pizarro. Twilight vertical migrations of zooplankton in a Chilean fjord. Progress in Oceanography, 129(??):114–124, December 2014. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661114000342.

Valdes:2007:DSB

[VLUC^{+}07] Luis Valdés, A. López-Urrutia, J. Cabal, M. Alvarez-Ossorio, A. Bode, A. Miranda, M. Cabanas, I. Huskin, R. Anadón, F. Alvarez-Marqués, M. Llope, and N. Rodríguez. A decade of sampling in the Bay of Biscay: What are the zooplankton time series telling us? Progress in Oceanography, 74(2–3):98–114, August 2007. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661107000596.

Venkataramana:2024:HSO

[VMA^{+}24] V. Venkataramana, R. K. Mishra, N. Anilkumar, Melena A. Soares, P. Sabu, and A. Sarkar. Hydrography of the Southern Ocean fronts — the zooplankton link: a detailed study in the Indian Ocean sector. Progress in Oceanography, 221(??):??, February 2024. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661124000107.

VanAudenhaege:2022:LTM

[VMB^{+}22a] Loïc Van Audenhaege, Marjolaine Matabos, Anik Brind’Amour, Jonathan Drugmand, Agathe Laës-Huon, Pierre-Marie Sarradin, and Jozée Sarrazin. Long-term monitoring reveals unprecedented stability of a vent mussel assemblage on the Mid-Atlantic Ridge. Progress in Oceanography, 204(??):??, June 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661122000520.

Vedenin:2022:UBZ

[VMB^{+}22b] A. A. Vedenin, A. N. Mironov, B. A. Bluhm, M. Käß, R. Degen, S. V. Galkin, and A. V. Gebruk. Uniform bathymetric zonation of marine benthos on a Pan-Arctic scale. Progress in Oceanography, 202(??):??, March 2022. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S007966112200026X.

Villamana:2019:RMC

[VMC^{+}19] Marina Villamaña, Emilio Marañón, Pedro Cermeño, Marta Estrada, Bieito Fernández-Castro, Francisco G. Figueiras, Mikel Latasa, Jose Luis Otero-Ferrer, Beatriz Reguera, and Beatriz Mouriño-Carballido. The role of mixing in controlling resource availability and phytoplankton community composition. Progress in Oceanography, 178(??):??, November 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661118303653.

Vestfals:2021:MDP

[VMH^{+}21] Cathleen D. Vestfals, Franz J. Mueter, Katherine S. Hedstrom, Benjamin J. Laurel, Colleen M. Petrik, Janet T. Duffy-Anderson, and Seth L. Danielson. Modeling the dispersal of polar cod (Boreogadus saida) and saffron cod (Eleginus gracilis) early life stages in the Pacific Arctic using a biophysical transport model. Progress in Oceanography, 196(??):??, August 2021. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661121000586.

Vichi:2008:POM

[VMN08] M. Vichi, S. Masina, and F. Nencioli. A process-oriented model study of equatorial Pacific phytoplankton: the role of iron supply and tropical instability waves. Progress in Oceanography, 78(2):147–162, August 2008. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661108000888.

Vellojin:2023:PEC

[VMV^{+}23] Jurleys P. Vellojin, Jorge I. Mardones, Valentina Vargas, Pablo P. Leal, Andrea Corredor-Acosta, and José L. Iriarte. Potential effects of climate change on the growth response of the toxic dinoflagellate Karenia selliformis from Patagonian waters of Chile. Progress in Oceanography, 211(??):??, February 2023. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661122002154.

Vonnahme:2023:IWF

[VNH^{+}23] Tobias Reiner Vonnahme, Aga Nowak, Mark James Hopwood, Lorenz Meire, Dorte H. Søgaard, Diana Krawczyk, Kjersti Kalhagen, and Thomas Juul-Pedersen. Impact of winter freshwater from tidewater glaciers on fjords in Svalbard and Greenland; a review. Progress in Oceanography, 219(??):??, December 2023. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661123001878.

VanCamp:1991:UBC

[VNMS91] L. Van Camp, L. Nykjaer, E. Mittelstaedt, and P. Schlittenhardt. Upwelling and boundary circulation off Northwest Africa as depicted by infrared and visible satellite observations. Progress in Oceanography, 26(4):357–402, ???? 1991. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/007966119190012B.

Valdes:2008:YSA

[VOG^{+}08] Jorge Valdés, Luc Ortlieb, Dimitri Gutierrez, Luis Marinovic, Gabriel Vargas, and Abdel Sifeddine. 250 years of sardine and anchovy scale deposition record in Mejillones Bay, northern Chile. Progress in Oceanography, 79(2–4):198–207, October 2008. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661108001730.

Vitorino:2002:WDNa

[VOJD02a] J. Vitorino, A. Oliveira, J. M. Jouanneau, and T. Drago. Winter dynamics on the northern Portuguese shelf. Part 1: physical processes. Progress in Oceanography, 52(??):129–153, ???? 2002. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661102000034.

Vitorino:2002:WDNb

[VOJD02b] J. Vitorino, A. Oliveira, J. M. Jouanneau, and T. Drago. Winter dynamics on the northern Portuguese shelf. Part 2: bottom boundary layers and sediment dispersal. Progress in Oceanography, 52(??):155–170, ???? 2002. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661102000046.

Velegrakis:1999:IWR

[VOT^{+}99] A. F. Velegrakis, E. Oikonomou, A. Theocharis, M. B. Collins, H. Kontoyiannis, V. Papadopoulos, G. Voulgaris, T. Wells, and E. Balopoulos. Internal waves revealed by Synthetic Aperture Radar (SAR) imagery in the vicinity of the eastern Cretan Arc Straits (Eastern Mediterranean). Progress in Oceanography, 44(4):553–572, December 1999. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661199000464.

Voss:2012:RPB

[VPH^{+}12] Rüdiger Voss, Myron A. Peck, Hans-Harald Hinrichsen, Catriona Clemmesen, Hannes Baumann, Daniel Stepputtis, Matthias Bernreuther, Jörn O. Schmidt, Axel Temming, and Fritz W. Köster. Recruitment processes in Baltic sprat — a re-evaluation of GLOBEC Germany hypotheses. Progress in Oceanography, 107(??):61–79, December 2012. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661112000602.

Valdivia:2019:DEM

[VPM^{+}19] Nelson Valdivia, Luis M. Pardo, Erasmo C. Macaya, Pirjo Huovinen, and Iván Gómez. Different ecological mechanisms lead to similar grazer controls on the functioning of periphyton Antarctic and sub-Antarctic communities. Progress in Oceanography, 174(??):7–16, May 2019. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S007966111730335X.

vanPraet:1990:RTD

[vPRT90] M. van Praët, A. L. Rice, and M. H. Thurston. Reproduction in two deep-sea anemones (Actiniaria); Phelliactis hertwigi and P. robusta. Progress in Oceanography, 24(??):207– 222, ???? 1990. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/007966119090031V.

Veitch:2009:BLC

[VPS09] Jennifer Veitch, Pierrick Penven, and Frank Shillington. The Benguela: a laboratory for comparative modeling studies. Progress in Oceanography, 83(1–4):296–302, December 2009. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661109000913.

Varnavas:2001:BPS

[VPW01] S. P. Varnavas, D. Panagiotaras, and G. A. Wolff. Biogeochemical processes at the sediment–water interface in a Northeastern Atlantic abyssal locality (porcupine abyssal plain). Progress in Oceanography, 50(??):223–243, ???? 2001. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661101000556.

Vezzulli:2003:CSG

[VR03] L. Vezzulli and P. C. Reid. The CPR survey (1948–1997): a gridded database browser of plankton abundance in the North Sea. Progress in Oceanography, 58(2–4):327–336, August 2003. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661103001368.

vanRuth:2010:IMP

[vRGW10] Paul D. van Ruth, George G. Ganf, and Tim M. Ward. The influence of mixing on primary productivity: a unique application of classical critical depth theory. Progress in Oceanography, 85(3–4):224–235, June 2010. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661110000339.

Valsala:2021:OSS

[VSA^{+}21] Vinu Valsala, M. G. Sreeush, M. Anju, Pentakota Sreenivas, Yogesh K. Tiwari, Kunal Chakraborty, and S. Sijikumar. An observing system simulation experiment for Indian Ocean surface pCO_{2} measurements. Progress in Oceanography, 194(??): ??, June 2021. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661121000574.

Vangriesheim:2001:TVN

[VSC01] A. Vangriesheim, B. Springer, and P. Crassous. Temporal variability of near-bottom particle resuspension and dynamics at the Porcupine Abyssal Plain, Northeast Atlantic. Progress in Oceanography, 50(??):123–145, ???? 2001. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661101000519.

Valdes:2021:OMZ

[VSGC21] Jorge Valdés, Abdel Sifeddine, Marcos Guiñez, and Alexis Castillo. Oxygen minimum zone variability during the last 700 years in a coastal upwelling area of the Humboldt system (Mejillones, 23^{o}S, Chile). a new approach from geochemical signature. Progress in Oceanography, 193(??):??, April– May 2021. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661121000100.

Venello:2021:ZFG

[VSGD21] Theresa A. Venello, Akash R. Sastri, Moira D. Galbraith, and John F. Dower. Zooplankton functional group responses to environmental drivers off the west coast of Vancouver Island, Canada. Progress in Oceanography, 190(??):??, January 2021. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661120302172.

Vilas:2014:SVM

[VSPP14] Luis González Vilas, Evangelos Spyrakos, Jesus M. Torres Palenzuela, and Yolanda Pazos. Support Vector Machinebased method for predicting Pseudo-nitzschia spp. blooms in coastal waters (Galician rias, NW Spain). Progress in Oceanography, 124(??):66–77, May 2014. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661114000299.

Vincent:2019:ISW

[VTGC19] Charles L. Vincent, Jim Thomson, Hans C. Graber, and Clarence O. Collins. Impact of swell on the wind-sea and resulting modulation of stress. Progress in Oceanography, 178(??):??, November 2019. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661118302209.

Varalakshmi:2021:TCP

[VVV21] P. Varalakshmi, N. Vasumathi, and R. Venkatesan. Tropical cyclone prediction based on multi-model fusion across Indian coastal region. Progress in Oceanography, 193(??):??, April–May 2021. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661121000446.

Vallina:2014:MFA

[VWDF14] S. M. Vallina, B. A. Ward, S. Dutkiewicz, and M. J. Follows. Maximal feeding with active prey-switching: a kill-the-winner functional response and its effect on global diversity and biogeography. Progress in Oceanography, 120(??):93–109, January 2014. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661113001468.

vanWeering:2002:RST

[vWdSBdH02] T. C. E. van Weering, H. C. de Stigter, W. Boer, and H. de Haas. Recent sediment transport and accumulation on the NW Iberian margin. Progress in Oceanography, 52(??):349– 371, ???? 2002. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661102000150.

vanWeering:1998:RSS

[vWHdS^{+}98] Tj. C. E. van Weering, I. R. Hall, H. C. de Stigter, I. N. McCave, and L. Thomsen. Recent sediments, sediment accumulation and carbon burial at Goban Spur, N.W. European Continental Margin (47–50^{o}N). Progress in Oceanography, 42(1–4):5–35, December 1998. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661198000263.

vanWeering:2002:BPD

[vWM02a] T. C. E. van Weering and I. N. McCave. Benthic processes and dynamics at the NW Iberian margin: an introduction. Progress in Oceanography, 52(??):123–128, ???? 2002. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661102000022.

vanWeering:2002:DRW

[vWM02b] Tj. C. E. van Weering and I. N. McCave. Dedication to Roland Wollast. Progress in Oceanography, 52(??):121–122, ???? 2002. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661102000253.

vanWeering:1998:OME

[vWMH98] Tjeerd C. E. van Weering, I. Nicholas McCave, and Ian R. Hall. Ocean Margin Exchange (OMEX I) benthic processes study. Progress in Oceanography, 42(1–4):1–4, December 1998. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661198000251.

Vargas-Yanez:2017:UTS

[VYGMM^{+}17] M. Vargas-Yáñez, M. C. García-Martínez, F. Moya, R. Balbín, J. L. López-Jurado, M. Serra, P. Zunino, J. Pascual, and J. Salat. Updating temperature and salinity mean values and trends in the Western Mediterranean: the RADMED project. Progress in Oceanography, 157(??):27–46, September 2017. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661117300848.

Wakelin:2020:CNB

[WAH^{+}20] Sarah L. Wakelin, Yuri Artioli, Jason T. Holt, Momme Butenschön, and Jeremy Blackford. Controls on near-bed oxygen concentration on the Northwest European continental shelf under a potential future climate scenario. Progress in Oceanography, 187(??):??, August–September 2020. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661120301397.

Wainwright:2021:ERS

[Wai21] Thomas C. Wainwright. Ephemeral relationships in salmon forecasting: a cautionary tale. Progress in Oceanography, 193(??):??, April–May 2021. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661121000124.

Walsh:1983:DSE

[Wal83] John J. Walsh. Death in the sea: Enigmatic phytoplankton losses. Progress in Oceanography, 12(1):1–86, ???? 1983. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/007966118390006X.

Warren:1973:EP

[War73] Bruce A. Warren. Editorial preface. Progress in Oceanography, 6(??):vii, ???? 1973. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0079661173900025.

Warren:2006:FLT

[War06] Bruce A. Warren. The first law of thermodynamics in a salty ocean. Progress in Oceanography, 70(2–4):149–167, August 2006. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661106000589.

Wassmann:2006:SFC

[Was06] Paul Wassmann. Structure and function of contemporary food webs on Arctic shelves: an introduction. Progress in Oceanography, 71(2–4):123–128, October 2006. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661106001194.

Wassmann:2011:AME

[Was11] Paul Wassmann. Arctic marine ecosystems in an era of rapid climate change. Progress in Oceanography, 90(1–4):1–17, July 2011. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661111000176.

Wassmann:2015:OPC

[Was15] Paul Wassmann. Overarching perspectives of contemporary and future ecosystems in the Arctic Ocean. Progress in Oceanography, 139(??):1–12, December 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661115001767.

Wiebe:2003:HNT

[WB03] Peter H. Wiebe and Mark C. Benfield. From the Hensen net toward four-dimensional biological oceanography. Progress in Oceanography, 56(1):7–136, January 2003. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661102001404.

Watts:2022:ISI

[WBA^{+}22] Jennifer Watts, Thomas G. Bell, Karen Anderson, Brian J. Butterworth, Scott Miller, Brent Else, and Jamie Shutler. Impact of sea ice on air–sea CO_{2} exchange — a critical review of polar eddy covariance studies. Progress in Oceanography, 201(??):??, February 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661122000039.

Willebrand:2001:CCT

[WBB^{+}01] Jürgen Willebrand, Bernard Barnier, Claus Böning, Christian Dieterich, Peter D. Killworth, Christian Le Provost, Yanli Jia, Jean-Marc Molines, and Adrian L. New. Circulation characteristics in three eddy-permitting models of the North Atlantic. Progress in Oceanography, 48(??):123–161, ???? 2001. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661101000039.

Wiebe:2022:DPC

[WBC^{+}22] P. H. Wiebe, M. F. Baumgartner, N. J. Copley, G. L. Lawson, C. Davis, R. Ji, and C. H. Greene. Does predation control the diapausing stock of Calanus finmarchicus in the Gulf of Maine? Progress in Oceanography, 206(??):??, August 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661122001203.

Wood:2015:DEC

[WBD^{+}15] Kevin R. Wood, Nicholas A. Bond, Seth L. Danielson, James E. Overland, Sigrid A. Salo, Phyllis J. Stabeno, and Jonathan Whitefield. A decade of environmental change in the Pacific Arctic region. Progress in Oceanography, 136(??):12–31, August 2015. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661115001020.

Wilson:2021:PMC

[WBF^{+}21] Samantha M. Wilson, Thomas W. Buehrens, Jennifer L. Fisher, Kyle L. Wilson, and Jonathan W. Moore. Phenological mismatch, carryover effects, and marine survival in a wild steelhead trout Oncorhynchus mykiss population. Progress in Oceanography, 193(??):??, April–May 2021. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661121000239.

Wilton:2015:MTC

[WBH15] David J. Wilton, Grant R. Bigg, and Edward Hanna. Modelling twentieth century global ocean circulation and iceberg flux at 48^{o}n: implications for west Greenland iceberg discharge. Progress in Oceanography, 138 (part A)(??):194–210, November 2015. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S007966111500155X.

Williams:2015:LIS

[WC15] William J. Williams and Eddy C. Carmack. The ‘interior’ shelves of the Arctic Ocean: Physical oceanographic setting, climatology and effects of sea-ice retreat on cross-shelf exchange. Progress in Oceanography, 139(??):24–41, December 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661115001603.

Wang:2020:DNS

[WCB20a] Zhengui Wang, Fei Chai, and Damian Brady. Development of a new sediment flux model — application in Chesapeake Bay. Progress in Oceanography, 185(??):??, June 2020. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661120300719.

Wassmann:2020:TUP

[WCB^{+}20b] P. Wassmann, E. C. Carmack, B. A. Bluhm, C. M. Duarte, J. Berge, K. Brown, J. M. Grebmeier, J. Holding, K. Kosobokova, R. Kwok, P. Matrai, S. Agusti, M. Babin, U. Bhatt, H. Eicken, I. Polyakov, S. Rysgaard, and H. P. Huntington. Towards a unifying pan-Arctic perspective: a conceptual modelling toolkit. Progress in Oceanography, 189(??):??, November–December 2020. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661120301920.

Wang:2020:MWP

[WCC^{+}20] Bo Wang, Min Chen, Feng Chen, Renming Jia, Xiaopeng Li, Minfang Zheng, and Yusheng Qiu. Meteoric water promotes phytoplankton carbon fixation and iron uptake off the eastern tip of the Antarctic Peninsula (eAP). Progress in Oceanography, 185(??):??, June 2020. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661120300860.

White:2005:SCD

[WCN^{+}05] Warren B. White, Daniel R. Cayan, Peter P. Niiler, John Moisan, Gary Lagerloef, Fabrice Bonjean, and David Legler. The seasonal cycle of diabatic heat storage in the Pacific Ocean. Progress in Oceanography, 64(1):1–29, January 2005. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S007966110500011X.

Wei:2023:CCD

[WCS^{+}23] Yuqiu Wei, Zhengguo Cui, Yongqiang Shi, Xiujuan Shan, Bingzhang Chen, Keming Qu, Quandong Xin, Tao Jiang, and Jufa Chen. Contrasting currents drive geographic variability in the biomass of Pacific saury (Cololabis saira), zooplankton, and phytoplankton in the northwestern Pacific. Progress in Oceanography, 217(??):??, September–October 2023. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661123001428.

Wang:2021:IMS

[WCX^{+}21] Tao Wang, Fei Chai, Xiaogang Xing, Jue Ning, Wensheng Jiang, and Stephen C. Riser. Influence of multi-scale dynamics on the vertical nitrate distribution around the Kuroshio Extension: an investigation based on BGC-Argo and satellite data. Progress in Oceanography, 193(??):??, April–May 2021. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661121000331.

Walsh:1994:CCC

[WD94] John J. Walsh and Dwight A. Dieterle. CO_{2} cycling in the coastal ocean. I — a numerical analysis of the southeastern Bering Sea with applications to the Chukchi Sea and the northern Gulf of Mexico. Progress in Oceanography, 34(4):335– 392, ???? 1994. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661194900191.

Walsh:2011:TCF

[WDC^{+}11] John J. Walsh, Dwight A. Dieterle, F. Robert Chen, Jason M. Lenes, Wieslaw Maslowski, John J. Cassano, Terry E. Whitledge, Dean Stockwell, Mikhail Flint, Irina N. Sukhanova, and John Christensen. Trophic cascades and future harmful algal blooms within ice-free Arctic Seas north of Bering Strait: a simulation analysis. Progress in Oceanography, 91(3):312–343, November 2011. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661111000164.

Witbaard:2001:ROM

[WDK^{+}01] R. Witbaard, G. C. A. Duineveld, A. Kok, J. van der Weele, and E. M. Berghuis. The response of Oneirophanta mutabilis (Holothuroidea) to the seasonal deposition of phytopigments at the Porcupine Abyssal Plain in the Northeast Atlantic. Progress in Oceanography, 50(??):423–441, ???? 2001. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661101000647.

Wilkerson:2002:HNC

[WDMC02] F. P. Wilkerson, R. C. Dugdale, A. Marchi, and C. A. Collins. Hydrography, nutrients and chlorophyll during El Niño and La Niña 1997–99 winters in the Gulf of the Farallones, California. Progress in Oceanography, 54(1–4):293– 310, July 2002. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661102000551.

Walsh:1989:NEC

[WDMMK89] John J. Walsh, Dwight A. Dieterle, Mark B. Meyers, and Frank E. Müller-Karger. Nitrogen exchange at the continental margin: a numerical study of the Gulf of Mexico. Progress in Oceanography, 23(4):245–301, ???? 1989. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0079661189900025.

Webster:1969:RDD

[Web69] Ferris Webster. On the representativeness of direct deep-sea current measurements. Progress in Oceanography, 5(??):3– 15, ???? 1969. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661169900238.

Wensu:1988:YBW

[Wen88] Wang Wensu. Yangtze brackish water plume — circulation and diffusion. Progress in Oceanography, 21(3–4):373–385, ???? 1988. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661188900158.

Wang:2006:EVE

[WF06] Chunzai Wang and Paul C. Fiedler. ENSO variability and the eastern tropical Pacific: a review. Progress in Oceanography, 69(2–4):239–266, May 2006. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661106000334. See erratum [WF07].

Wang:2007:EEV

[WF07] Chunzai Wang and Paul C. Fiedler. Erratum to “ENSO variability and the eastern tropical Pacific: a review” [Prog. Oceanogr. 69 (2006) 239–266]. Progress in Oceanography, 72(4):384, March 2007. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661106001583.

See [WF06].

Warrick:2017:CRP

[WF17] Jonathan A. Warrick and Katherine L. Farnsworth. Coastal river plumes: Collisions and coalescence. Progress in Oceanography, 151(??):245–260, February 2017. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661116300532.

White:2013:NFG

[WFBN^{+}13] Angelicque E. White, Rachel A. Foster, Claudia R. Benitez-Nelson, Pere Masqué, Elisabet Verdeny, Brian N. Popp, Karen E. Arthur, and Fredrick G. Prahl. Nitrogen fixation in the Gulf of California and the Eastern tropical North Pacific. Progress in Oceanography, 109(??):1–17, February 2013. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661112001115.

Wanless:2007:BLK

[WFD^{+}07] S. Wanless, M. Frederiksen, F. Daunt, B. E. Scott, and M. P. Harris. Black-legged kittiwakes as indicators of environmental change in the North Sea: Evidence from longterm studies. Progress in Oceanography, 72(1):30–38, January 2007. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661106001017.

Waga:2022:PPP

[WFH^{+}22] Hisatomo Waga, Amane Fujiwara, Toru Hirawake, Koji Suzuki, Kazuhiro Yoshida, Hiroto Abe, and Daiki Nomura. Primary productivity and phytoplankton community structure in surface waters of the western subarctic Pacific and the Bering Sea during summer with reference to bloom stages. Progress in Oceanography, 201(??):??, February 2022. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661121002214.

Weijerman:2015:HMC

[WFJ^{+}15] Mariska Weijerman, Elizabeth A. Fulton, Annette B. G. Janssen, Jan J. Kuiper, Rik Leemans, Barbara J. Robson, Ingrid A. van de Leemput, and Wolf M. Mooij. How models can support ecosystem-based management of coral reefs. Progress in Oceanography, 138(??):559–570, November 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661114002274.

Whitney:2007:PDO

[WFR07] Frank A. Whitney, Howard J. Freeland, and Marie Robert. Persistently declining oxygen levels in the interior waters of the eastern subarctic Pacific. Progress in Oceanography, 75(2):179–199, October 2007. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661107001516.

Williamson:2015:CSS

[WFS^{+}15] Robert I. Williamson, John G. Field, Frank A. Shillington, Astrid Jarre, and Anet Potgieter. Combining shipboard in situ data with satellite data to estimate daily primary production in a coastal upwelling system: a data mining approach. Progress in Oceanography, 138 (part A)(??):58–76, November 2015. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661115001949.

Wunsch:1982:TGC

[WG82] Carl Wunsch and Barbara Grant. Towards the general circulation of the North Atlantic Ocean. Progress in Oceanography, 11(1):1–59, ???? 1982. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0079661182900106.

Watling:2013:PBD

[WGCS13] Les Watling, John Guinotte, Malcolm R. Clark, and Craig R. Smith. A proposed biogeography of the deep ocean floor. Progress in Oceanography, 111(??):91–112, April 2013. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661112001693.

Wishner:2008:VZD

[WGG^{+}08] Karen F. Wishner, Celia Gelfman, Marcia M. Gowing, Dawn M. Outram, Mary Rapien, and Rebecca L. Williams. Vertical zonation and distributions of calanoid copepods through the lower oxycline of the Arabian Sea oxygen minimum zone. Progress in Oceanography, 78(2):163–191, August 2008. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S007966110800044X.

Wang:2024:SIT

[WGM^{+}24] Shuya Wang, Xinyu Guo, Akihiko Morimoto, Anzhou Cao, Eisuke Tsutsumi, Yasumasa Miyazawa, and Sergey M. Varlamov. Semidiurnal internal tides in a shelf sea south of Japan: Characteristics, energetics, and temporal variations. Progress in Oceanography, 222(??):??, March 2024. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661124000351.

Wang:2019:SVN

[WGZZ19] Yucheng Wang, Xinyu Guo, Liang Zhao, and Jing Zhang. Seasonal variations in nutrients and biogenic particles in the upper and lower layers of East China Sea Shelf and their export to adjacent seas. Progress in Oceanography, 176(??):??, September 2019. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661118301034.

Wroblewski:1989:UIM

[WH89] J. S. Wroblewski and Eileen E. Hofmann. U.S. interdisciplinary modeling studies of coastal-offshore exchange processes: Past and future. Progress in Oceanography, 23(2):65– 99, ???? 1989. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661189900116.

Warner:1994:SCP

[WH94] A. J. Warner and G. C. Hays. Sampling by the continuous plankton recorder survey. Progress in Oceanography, 34(2–3):237–256, ???? 1994. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0079661194900116.

Wang:2020:AFC

[WH20] Bo-Shian Wang and Tung-Yuan Ho. Aerosol Fe cycling in the surface water of the Northwestern Pacific Ocean. Progress in Oceanography, 183(??):??, April 2020. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S007966112030029X.

Wiggert:2005:MDB

[WHBK05] J. D. Wiggert, R. R. Hood, K. Banse, and J. C. Kindle. Monsoon-driven biogeochemical processes in the Arabian Sea. Progress in Oceanography, 65(2–4):176–213, May 2005. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661105000467.

Wigham:2003:LTC

[WHBW03] Benjamin D. Wigham, Ian R. Hudson, David S. M. Billett, and George A. Wolff. Is long-term change in the abyssal Northeast Atlantic driven by qualitative changes in export flux? Evidence from selective feeding in deep-sea holothurians. Progress in Oceanography, 59(4):409–441, December 2003. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661103001794.

Wheeler:1993:NPS

[Whe93] Patricia A. Wheeler. New production in the subarctic Pacific Ocean: Net changes in nitrate concentrations, rates of nitrate assimilation and accumulation of particulate nitrogen. Progress in Oceanography, 32(??):137–161, ???? 1993. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0079661193900112.

Wheatcroft:2006:TSM

[Whe06] Robert A. Wheatcroft. Time-series measurements of macrobenthos abundance and sediment bioturbation intensity on a flood-dominated shelf. Progress in Oceanography, 71(1): 88–122, October 2006. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661106000796.

Wiebe:2016:IWG

[WHG^{+}16] Peter H. Wiebe, Roger Harris, Astthor Gislason, Piotr Margonski, Hein Rune Skjoldal, Mark Benfield, Steve Hay, Todd O’Brien, and Luis Valdés. The ICES Working Group on Zooplankton Ecology: Accomplishments of the first 25 years.

Progress in Oceanography, 141(??):179–201, February 2016. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661115002554.

White:1994:VBO

[Whi94] Brian N. White. Vicariance biogeography of the open-ocean Pacific. Progress in Oceanography, 34(2–3):257–284, ???? 1994. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0079661194900124.

White:1995:DGO

[Whi95] Warren B. White. Design of a global observing system for gyre-scale upper ocean temperature variability. Progress in Oceanography, 36(3):169–217, ???? 1995. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0079661195000178.

Wilderbuer:2002:FRR

[WHI^{+}02] T. K. Wilderbuer, A. B. Hollowed, W. J. Ingraham, P. D. Spencer, M. E. Conners, N. A. Bond, and G. E. Walters. Flatfish recruitment response to decadal climatic variability and ocean conditions in the eastern Bering Sea. Progress in Oceanography, 55(1–2):235–247, October 2002. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661102000812.

Wunsch:1997:DLP

[WHIH97] Carl Wunsch, Dale B. Haidvogel, Mohamed Iskandarani, and R. Hughes. Dynamics of the long-period tides. Progress in Oceanography, 40(??):81–108, ???? 1997. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661197000244.

Winans:2023:STV

[WHK23] Amanda K. Winans, BethElLee Herrmann, and Julie E. Keister. Spatio-temporal variation in zooplankton community composition in the southern Salish Sea: Changes during the 2015–2016 Pacific marine heatwave. Progress in Oceanography, 214(??):??, June 2023. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661123000654.

Wang:2017:VFP

[WHS17] Zeliang Wang, James Hamilton, and Jie Su. Variations in freshwater pathways from the Arctic Ocean into the North Atlantic Ocean. Progress in Oceanography, 155(??):54–73, June 2017. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661117300320.

Wold:2023:AIZ

[WHS^{+}23] Anette Wold, Haakon Hop, Camilla Svensen, Janne E. Søreide, Karen M. Assmann, Mateusz Ormanczyk, and Slawomir Kwasniewski. Atlantification influences zooplankton communities seasonally in the northern Barents Sea and Arctic Ocean. Progress in Oceanography, 219(??):??, December 2023. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661123001763.

Watanabe:1986:AYY

[WHT86] Tomowo Watanabe, Kimio Hanawa, and Yoshiaki Toba. Analysis of year-to-year variation of water temperature along the coast of the Japan Sea. Progress in Oceanography, 17(3–4):337– 357, ???? 1986. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661186900534.

Wiles:1965:PCP

[Wil65] William W. Wiles. Pleistocene changes in the pore concentration of a planktonic foraminiferal species from the Pacific Ocean. Progress in Oceanography, 4(??):153–160, September 2, 1965. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661165900480.

Williamson:1987:ILO

[Wil87] D. I. Williamson. Incongruous larvae and the origin of some invertebrate life-histories. Progress in Oceanography, 19(2):87– 116, ???? 1987. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/007966118790005X.

Wiseman:1965:CMC

[Wis65] John D. H. Wiseman. Calcium and magnesium carbonate in some Indian Ocean sediments. Progress in Oceanography, 3(??):373–383, ???? 1965. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0079661165900327.

Wong:1992:RCF

[WJE^{+}92] Kai M. Wong, Terrence A. Jokela, Rodney J. Eagle, James L. Brunk, and Victor E. Noshkin. Radionuclide concentrations, fluxes, and residence times at Santa Monica and San Pedro Basins. Progress in Oceanography, 30(1–4):353–391, ???? 1992. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/007966119290019V.

Woodworth-Jefcoats:2015:TTE

[WJPHB15] Phoebe A. Woodworth-Jefcoats, Jeffrey J. Polovina, Evan A. Howell, and Julia L. Blanchard. Two takes on the ecosystem impacts of climate change and fishing: Comparing a size-based and a species-based ecosystem model in the central North Pacific. Progress in Oceanography, 138(??):533–545, November 2015. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661115000646.

Wassmann:2015:CDA

[WKS^{+}15] P. Wassmann, K. N. Kosobokova, D. Slagstad, K. F. Drinkwater, R. R. Hopcroft, S. E. Moore, I. Ellingsen, R. J. Nelson, E. Carmack, E. Popova, and J. Berge. The contiguous domains of Arctic Ocean advection: Trails of life and death. Progress in Oceanography, 139(??):42–65, December 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661115001548.

Wu:2016:SDP

[WL16] Zhengchao Wu and Qian P. Li. Spatial distributions of polyunsaturated aldehydes and their biogeochemical implications in the Pearl River Estuary and the adjacent northern South China Sea. Progress in Oceanography, 147(??):1–9, September 2016. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661116300027.

Wang:2023:BEN

[WLCG23] Shuya Wang, Qun Li, Anzhou Cao, and Xinyu Guo. Burst events of near-inertial waves in the Beaufort Sea. Progress in Oceanography, 212(??):??, March 2023. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661123000290.

Wang:2015:VSS

[WLD^{+}15] Zeliang Wang, Youyu Lu, Frederic Dupont, John W. Loder, Charles Hannah, and Daniel G. Wright. Variability of sea surface height and circulation in the North Atlantic: Forcing mechanisms and linkages. Progress in Oceanography, 132(??):273–286, March 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661113002231.

Weimerskirch:2010:FMG

[WLKM10] Henri Weimerskirch, Matthieu Le Corre, Emilie Tew Kai, and Francis Marsac. Foraging movements of great frigatebirds from Aldabra Island: Relationship with environmental variables and interactions with fisheries. Progress in Oceanography, 86(1–2):204–213, July 2010. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661110000376.

Willett:2006:ETI

[WLL06] Cynthia S. Willett, Robert R. Leben, and Miguel F. Lavín. Eddies and tropical instability waves in the eastern tropical Pacific: a review. Progress in Oceanography, 69(2–4):218– 238, May 2006. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661106000322.

Wang:2023:PDC

[WLL^{+}23] Peixuan Wang, Edward A. Laws, Xin Liu, Jixin Chen, Yanping Zhong, Chao Xu, and Bangqin Huang. Proportion of dead cells in phytoplankton modulates community structure. Progress in Oceanography, 219(??):??, December 2023. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661123001854.

Wilczak:2007:UOT

[WLM07] James M. Wilczak, Robert R. Leben, and David S. Mc-Collum. Upper-ocean thermal structure and heat content off the US West Coast during the 1997–1998 El Niño event based on AXBT and satellite altimetry data. Progress in Oceanography, 74(1):48–70, July 2007. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661107000389.

Weller:2013:TVD

[WLM^{+}13] D. I. Weller, C. S. Law, A. Marriner, S. D. Nodder, F. H. Chang, J. A. Stephens, S. W. Wilhelm, P. W. Boyd, and P. J. H. Sutton. Temporal variation of dissolved methane in a subtropical mesoscale eddy during a phytoplankton bloom in the southwest Pacific Ocean. Progress in Oceanography, 116(??):193–206, September 2013. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661113001146.

Wilkin:2022:DAM

[WLM^{+}22] John Wilkin, Julia Levin, Andrew Moore, Hernan Arango, Alexander López, and Elias Hunter. A data-assimilative model reanalysis of the U.S. Mid Atlantic Bight and Gulf of Maine: Configuration and comparison to observations and global ocean models. Progress in Oceanography, 209(??):??, December 2022. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661122001781.

Wu:2021:IND

[WLP^{+}21] Jing-Ying Wu, Siou-Yan Lin, Shao-Hung Peng, Jia-Jang Hung, Chen-Tung Arthur Chen, and Li-Lian Liu. Isotopic niche differentiation in benthic consumers from shallow-water hydrothermal vents and nearby non-vent rocky reefs in northeastern Taiwan. Progress in Oceanography, 195(??):??, July 2021. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661121000835.

Waite:2013:STV

[WM13] Jason N. Waite and Franz J. Mueter. Spatial and temporal variability of chlorophyll-a concentrations in the coastal Gulf of Alaska, 1998–2011, using cloud-free reconstructions of SeaWiFS and MODIS-Aqua data. Progress in Oceanography, 116(??):179–192, September 2013. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661113001122.

Werdell:2018:OAC

[WMB^{+}18] P. Jeremy Werdell, Lachlan I. W. McKinna, Emmanuel Boss, Steven G. Ackleson, Susanne E. Craig, Watson W. Gregg, Zhongping Lee, Stéphane Maritorena, Collin S. Roesler, Cécile S. Rousseaux, Dariusz Stramski, James M. Sullivan, Michael S. Twardowski, Maria Tzortziou, and Xiaodong Zhang. An overview of approaches and challenges for retrieving marine inherent optical properties from ocean color remote sensing. Progress in Oceanography, 160(??):186–212, January 2018. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661117302999.

Wakefield:2021:SDH

[WMB^{+}21] Ewan D. Wakefield, David L. Miller, Sarah L. Bond, Fabrice le Bouard, Paloma C. Carvalho, Paulo Catry, Ben J. Dilley, David A. Fifield, Carina Gjerdrum, Jacob González-Solís, Holly Hogan, Vladimir Laptikhovsky, Benjamin Merkel, Julie A. O. Miller, Peter I. Miller, Simon J. Pinder, Tânia Pipa, Peter M. Ryan, Laura A. Thompson, Paul M. Thompson, and Jason Matthiopoulos. The summer distribution, habitat associations and abundance of seabirds in the sub-polar frontal zone of the Northwest Atlantic. Progress in Oceanography, 198(??):??, November 2021. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661121001427.

Walsh:1989:CNC

[WMC^{+}89] J. J. Walsh, C. P. McRoy, L. K. Coachman, J. J. Goering, J. J. Nihoul, T. E. Whitledge, T. H. Blackburn, P. L. Parker, C. D. Wirick, P. G. Shuert, J. M. Grebmeier, A. M. Springer, R. D. Tripp, D. A. Hansell, S. Djenidi, E. Deleersnijder, K. Henriksen, B. A. Lund, P. Andersen, F. E. Müller-Karger, and K. Dean. Carbon and nitrogen cycling within the Bering/Chukchi Seas: Source regions for organic matter effecting AOU demands of the Arctic Ocean. Progress in Oceanography, 22(4):277–359, ???? 1989. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0079661189900062.

Ward:2008:SPC

[WMWR08] Peter Ward, Michael P. Meredith, Mick J. Whitehouse, and Peter Rothery. The summertime plankton community at South Georgia (Southern Ocean): Comparing the historical (1926/1927) and modern (post 1995) records. Progress in Oceanography, 78(3):241–256, September 2008. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661108001067.

Watanabe:2021:CNC

[WNNI21] Yutaka W. Watanabe, Yuto Nakano, Jun Nishioka, and Masanori Ito. Change of nitrogen chemical tracer N_{2} * between the Okhotsk Sea and the subarctic North Pacific. Progress in Oceanography, 199(??):??, December 2021. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661121001774.

Warren:1985:SPR

[WO85] Bruce A. Warren and W. Brechner Owens. Some preliminary results concerning deep northern-boundary currents in the North Pacific. Progress in Oceanography, 14(??):537–551, ???? 1985. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661185900278.

Wang:2015:PFD

[WO15] Muyin Wang and James E. Overland. Projected future duration of the sea-ice-free season in the Alaskan Arctic. Progress in Oceanography, 136(??):50–59, August 2015. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661115000038.

Woods:2005:LEM

[Woo05] J. D. Woods. The Lagrangian Ensemble metamodel for simulating plankton ecosystems. Progress in Oceanography, 67(1–2):84–159, October 2005. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661105000893.

Woodgate:2018:IPI

[Woo18] Rebecca A. Woodgate. Increases in the Pacific inflow to the Arctic from 1990 to 2015, and insights into seasonal trends and driving mechanisms from year-round Bering Strait mooring data. Progress in Oceanography, 160(??):124–154, January 2018. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661117302215.

Wang:2014:FAP

[WOW^{+}14] M. Wang, R. O’Rorke, A. M. Waite, L. E. Beckley, P. Thompson, and A. G. Jeffs. Fatty acid profiles of phyllosoma larvae of western rock lobster (Panulirus cygnus) in cyclonic and anticyclonic eddies of the Leeuwin Current off Western Australia. Progress in Oceanography, 122(??):153–162, March 2014. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661114000123.

Waldron:1991:STV

[WP91] H. N. Waldron and T. A. Probyn. Short-term variability during an anchor station study in the southern Benguela upwelling system: Nitrogen supply to the euphotic zone during a quiscent phase in the upwelling cycle. Progress in Oceanography, 28(1–2):153–166, ???? 1991. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/007966119190024G.

Westgaard:2024:TEI

[WPA^{+}24] Jon-Ivar Westgaard, Kim Præbel, Per Arneberg, Brian P. Ulaski, Randi Ingvaldsen, Owen S. Wangensteen, and Torild Johansen. Towards eDNA informed biodiversity studies — comparing water derived molecular taxa with traditional survey methods. Progress in Oceanography, 222(??): ??, March 2024. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661124000363.

Woods:2005:SPV

[WPB05] John Woods, Angelo Perilli, and Wolfgang Barkmann. Stability and predictability of a virtual plankton ecosystem created with an individual-based model. Progress in Oceanography, 67(1–2):43–83, October 2005. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661105000881.

Wichard:2008:IDC

[WPB^{+}08] Thomas Wichard, Serge A. Poulet, Anne-Laure Boulesteix, Jean Baptiste Ledoux, Benoit Lebreton, Julie Marchetti, and Georg Pohnert. Influence of diatoms on copepod reproduction. II. Uncorrelated effects of diatom-derived α, β, γ, δunsaturated aldehydes and polyunsaturated fatty acids on Calanus helgolandicus in the field. Progress in Oceanography, 77(1):30–44, April 2008. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661108000451.

Wall-Palmer:2018:BGD

[WPBG^{+}18] Deborah Wall-Palmer, Alice K. Burridge, Erica Goetze, Frank R. Stokvis, Arie W. Janssen, Lisette Mekkes, María Moreno-Alcántara, Nina Bednaršek, Tom Schiøtte, Martin Vinther Sørensen, Christopher W. Smart, and Katja T. C. A. Peijnenburg. Biogeography and genetic diversity of the atlantid heteropods. Progress in Oceanography, 160(??):1–25, January 2018. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661117300757.

Woillez:2010:SMS

[WPH^{+}10] Mathieu Woillez, Pierre Petitgas, Martin Huret, Caroline Struski, and Fabien Léger. Statistical monitoring of spatial patterns of environmental indices for integrated ecosystem assessment: Application to the Bay of Biscay pelagic zone. Progress in Oceanography, 87(1–4):83–93, October 2010. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661110001199.

Wolhowe:2014:BPC

[WPW^{+}14] Matthew D. Wolhowe, Fredrick G. Prahl, Angelicque E. White, Brian N. Popp, and Anaid Rosas-Navarro. A biomarker perspective on coccolithophorid growth and export in a stratified sea. Progress in Oceanography, 122(??):65–76, March 2014. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661113002292.

Wilson:2008:GDS

[WQ08] Cara Wilson and Xuemei Qiu. Global distribution of summer chlorophyll blooms in the oligotrophic gyres. Progress in Oceanography, 78(2):107–134, August 2008. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661108001079.

Willson:2000:CMF

[WR00] Hester R. Willson and N. W. Rees. Classification of mesoscale features in the Brazil–Falkland Current confluence zone. Progress in Oceanography, 45(3–4):415–426, April 2000. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661100000112.

Wiebe:2003:TMV

[WR03] P. H. Wiebe and J. L. Reid. A tribute to Martin V. Angel and Robert L. Smith for their years as Co-Editors of Progress in Oceanography. Progress in Oceanography, 56(1):3– 5, January 2003. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S007966110300003X.

Wassmann:2006:FWC

[WRH^{+}06] Paul Wassmann, Marit Reigstad, Tore Haug, Bert Rudels, Michael L. Carroll, Haakon Hop, Geir Wing Gabrielsen, Stig Falk-Petersen, Stanislav G. Denisenko, Elena Arashkevich, Dag Slagstad, and Olga Pavlova. Food webs and carbon flux in the Barents Sea. Progress in Oceanography, 71(2–4):232–287, October 2006. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661106001315.

Williams:1992:ISC

[WRS^{+}92] P. M. Williams, K. J. Robertson, A. Soutar, S. M. Griffin, and E. R. M. Druffel. Isotopic signatures (^{14}C, ^{13}C, 15N) as tracers of sources and cycling of soluble and particulate organic matter in the Santa Monica Basin, California. Progress in Oceanography, 30(1–4):253–290, ???? 1992. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/007966119290015R.

Wei:2021:SCP

[WSC^{+}21] Yuqiu Wei, Jun Sun, Zhuo Chen, Zhaoyi Zhang, Guicheng Zhang, and Xiaofang Liu. Significant contribution of picoplankton size fraction to biogenic silica standing stocks in the Western Pacific Ocean. Progress in Oceanography, 192(??): ??, March 2021. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661121000069.

Welschmeyer:1993:PPS

[WSG^{+}93] Nicholas A. Welschmeyer, Suzanne Strom, Ralf Goericke, Giacomo DiTullio, Marcia Belvin, and Willis Petersen. Primary production in the subarctic Pacific Ocean: Project SUPER. Progress in Oceanography, 32(??):101–135, ???? 1993. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/007966119390010B.

Wilson:2015:SLD

[WSH15] Robert J. Wilson, Douglas C. Speirs, and Michael R. Heath. On the surprising lack of differences between two congeneric calanoid copepod species, Calanus finmarchicus and C. helgolandicus. Progress in Oceanography, 134(??):413–431, May 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661114002183.

Wihsgott:2022:OVM

[WSH^{+}22] Juliane U. Wihsgott, Jonathan Sharples, Joanne E. Hopkins, E. Malcolm S. Woodward, Tom Hull, Naomi Greenwood, and David B. Sivyer. Observations of vertical mixing in autumn and its effect on the autumn phytoplankton bloom. Progress in Oceanography, 177(??):??, October 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661118302544.

Wang:2020:ETI

[WSL20] Yuan Wang, Jinyu Sheng, and Youyu Lu. Examining tidal impacts on seasonal circulation and hydrography variability over the eastern Canadian shelf using a coupled circulation-ice regional model. Progress in Oceanography, 189(??):??, November–December 2020. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661120301865.

Wenzel:2001:ACG

[WSO01] Manfred Wenzel, Jens Schröter, and Dirk Olbers. The annual cycle of the global ocean circulation as determined by 4D VAR data assimilation. Progress in Oceanography, 48(1):73– 119, ???? 2001. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661100000483.

Ward:2013:WBM

[WSO^{+}13] Ben A. Ward, Markus Schartau, Andreas Oschlies, Adrian P. Martin, Michael J. Follows, and Thomas R. Anderson. When is a biogeochemical model too complex? Objective model reduction and selection for North Atlantic time-series sites.

Progress in Oceanography, 116(??):49–65, September 2013. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661113000839.

Watson:2015:ERM

[WSS15] James R. Watson, Charles A. Stock, and Jorge L. Sarmiento. Exploring the role of movement in determining the global distribution of marine biomass using a coupled hydrodynamic — size-based ecosystem model. Progress in Oceanography, 138(??):521–532, November 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661114001463.

Wurl:2016:BLP

[WST^{+}16] Oliver Wurl, Christian Stolle, Chu Van Thuoc, Pham The Thu, and Xavier Mari. Biofilm-like properties of the sea surface and predicted effects on air–sea CO_{2} exchange. Progress in Oceanography, 144(??):15–24, May 2016. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661115300616.

Wang:2021:SSI

[WST^{+}21] Guizhi Wang, Shengyao Sun, Ehui Tan, Liwen Chen, Lifang Wang, Tao Huang, Kuanbo Zhou, Weifang Chen, and Xianghui Guo. A strong summer intrusion of the Kuroshio and residence time in the northern South China Sea revealed by radium isotopes. Progress in Oceanography, 197(??):??, September– October 2021. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661121001063.

Wu:2012:CEC

[WTH12] Yongsheng Wu, Charles Tang, and Charles Hannah. The circulation of eastern Canadian seas. Progress in Oceanography, 106(??):28–48, November 2012. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661112000687.

Ward:2014:MSO

[WTT14] Peter Ward, Geraint A. Tarling, and Sally E. Thorpe. Mesozooplankton in the Southern Ocean: Spatial and temporal patterns from Discovery Investigations. Progress in Oceanography, 120(??):305–319, January 2014. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661113002036.

Wu:2013:IMP

[Wu13] Chau-Ron Wu. Interannual modulation of the Pacific Decadal Oscillation (PDO) on the low-latitude western North Pacific. Progress in Oceanography, 110(??):49–58, March 2013. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661112001735.

Wunsch:2024:TAO

[Wun24] Carl Wunsch. A time-average ocean: Thermal wind and flow spirals. Progress in Oceanography, 221(??):??, February 2024. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661124000120.

Wust:1964:MDS

[Wüs64] Georg Wüst. The major deep-sea expeditions and research vessels 1873–1960: a contribution to the history of oceanography. Progress in Oceanography, 2(??):1, 3–6, IN1–IN4, 7– 10, IN6–IN7, 11–14, IN9–IN12, 15–18, IN13–IN14, 19–52, ???? 1964. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/0079661164900023.

Whitney:2002:INN

[WW02] F. A. Whitney and D. W. Welch. Impact of the 1997–1998 El Niño and 1999 La Niña on nutrient supply in the Gulf of Alaska. Progress in Oceanography, 54(1–4):405–421, July 2002. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661102000617.

Walsh:2022:IED

[WWL^{+}22] J. J. Walsh, R. H. Weisberg, J. M. Lenes, F. R. Chen, D. A. Dieterle, L. Zheng, K. L. Carder, G. A. Vargo, J. A. Havens, E. Peebles, D. J. Hollander, R. He, C. A. Heil, B. Mahmoudi, and J. H. Landsberg. Isotopic evidence for dead fish maintenance of Florida red tides, with implications for coastal fisheries over both source regions of the West Florida shelf and within downstream waters of the South Atlantic Bight. Progress in Oceanography, 80(1–2):51–73, ???? 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661108002085.

Wu:2024:ASC

[WWL^{+}24] Zelun Wu, Hongjie Wang, Enhui Liao, Chuanmin Hu, Kelsea Edwing, Xiao-Hai Yan, and Wei-Jun Cai. Air-sea CO_{2} flux in the Gulf of Mexico from observations and multiple machinelearning data products. Progress in Oceanography, 223(??):??, April–May 2024. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661124000508.

Whitworth:1999:DWB

[WWN^{+}99] T. Whitworth III, B. A. Warren, W. D. Nowlin, Jr., S. B. Rutz, R. D. Pillsbury, and M. I. Moore. On the deep western-boundary current in the Southwest Pacific Basin. Progress in Oceanography, 43(1):1–54, January 1999. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661199000051.

Wu:2007:DNA

[WWSJ07] Peili Wu, Richard Wood, Peter Stott, and Gareth S. Jones. Deep North Atlantic freshening simulated in a coupled climate model. Progress in Oceanography, 73(3–4):370–383, May 2007. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661107000730.

Wang:2023:PCC

[WWW^{+}23] Chaofeng Wang, Xiaoyu Wang, Yuanyuan Wei, Guijun Guo, Haibo Li, Aiyong Wan, and Wuchang Zhang. Pelagic ciliate (Ciliophora) communities in the Southern Ocean: Bioindicator to water mass, habitat suitability classification and potential response to global warming. Progress in Oceanography, 216(??):??, August 2023. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661123001246.

Wei:2019:MWS

[WWZ19] Yanzhou Wei, Hongna Wang, and Rong-Hua Zhang. Mesoscale wind stress-SST coupled perturbations in the Kuroshio Extension. Progress in Oceanography, 172(??):108–123, March 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661117302471.

Wong:2007:VNC

[WXH07] C. S. Wong, Liusen Xie, and William W. Hsieh. Variations in nutrients, carbon and other hydrographic parameters related to the 1976/77 and 1988/89 regime shifts in the sub-arctic Northeast Pacific. Progress in Oceanography, 75(2):326–342, October 2007. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661107001541.

Wyrtki:2006:P

[Wyr06] Klaus Wyrtki. Preface. Progress in Oceanography, 69(2–4):91– 93, May 2006. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661106000401.

Wassmann:2000:VFF

[WYT00] Paul Wassmann, Jan Erik Ypma, and Anastassios Tselepides. Vertical flux of faecal pellets and microplankton on the shelf of the oligotrophic Cretan Sea (NE Mediterranean Sea). Progress in Oceanography, 46(2–4):241–258, August 2000. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661100000215.

Wooster:2004:RSN

[WZ04] Warren S. Wooster and Chang Ik Zhang. Regime shifts in the North Pacific: early indications of the 1976–1977 event. Progress in Oceanography, 60(2–4):183–200, February 2004. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661104000278.

Weydmann-Zwolicka:2021:MSD

[WZBK^{+}21] Agata Weydmann-Zwolicka, Piotr Balazy, Piotr Kuklinski, Janne E. Søreide, Weronika Patuła, and Marta Ronowicz. Meroplankton seasonal dynamics in the high Arctic fjord: Comparison of different sampling methods. Progress in Oceanography, 190(??):??, January 2021. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661120302196.

Wu:2020:ETK

[WZC20] Renhao Wu, Han Zhang, and Dake Chen. Effect of Typhoon Kalmaegi (2014) on northern South China Sea explored using multi-platform satellite and buoy observations data. Progress in Oceanography, 180(??):??, January 2020. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661119303982.

Wang:2016:SRD

[WZFW16] Jianing Wang, Liang Zhao, Renfu Fan, and Hao Wei. Scaling relationships for diffusive boundary layer thickness and diffusive flux based on in situ measurements in coastal seas. Progress in Oceanography, 144(??):1–14, May 2016. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S007966111600029X.

Xiu:2014:CBP

[XC14] Peng Xiu and Fei Chai. Connections between physical, optical and biogeochemical processes in the Pacific Ocean. Progress in Oceanography, 122(??):30–53, March 2014. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661113002279.

Xue:2016:SSC

[XCH^{+}16] Liang Xue, Wei-Jun Cai, Xinping Hu, Christopher Sabine, Stacy Jones, Adrienne J. Sutton, Li-Qing Jiang, and Janet J.

Reimer. Sea surface carbon dioxide at the Georgia time series site (2006–2007): Air–sea flux and controlling processes. Progress in Oceanography, 140(??):14–26, January 2016. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661115001986.

Xing:1995:ATD

[XD95] Jiuxing Xing and Alan M. Davies. Application of three dimensional turbulence energy models to the determination of tidal mixing and currents in a shallow sea. Progress in Oceanography, 35(2):153–205, ???? 1995. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0079661195000074.

Xing:1996:PII

[XD96] Jiuxing Xing and Alan M. Davies. Processes influencing the internal tide, its higher harmonics, and tidally induced mixing on the malin-hebrides shelf. Progress in Oceanography, 38(2):155– 204, ???? 1996. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661197000013.

Xie:2023:RSC

[XDG^{+}23] Jieshuo Xie, Hui Du, Yankun Gong, Jianwei Niu, Yinghui He, Zhiwu Chen, Guangping Liu, Le Liu, Lindan Zhang, and Shuqun Cai. The role of seasonal circulation in the variability of dynamic parameters of internal solitary waves in the Sulu Sea. Progress in Oceanography, 217(??):??, September– October 2023. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S007966112300143X.

Xu:2020:ODE

[XHC^{+}20] Jiexin Xu, Yinghui He, Zhiwu Chen, Haigang Zhan, Yuqi Wu, Jieshuo Xie, Xiaodong Shang, Dezhi Ning, Wendong Fang, and Shuqun Cai. Observations of different effects of an anti-cyclonic eddy on internal solitary waves in the South China Sea. Progress in Oceanography, 188(??):??, October 2020. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661120301610.

Xie:2020:DMC

[XHW^{+}20] Zhang-Xian Xie, Yan-Bin He, Ming-Hua Wang, Shu-Feng Zhang, Ling-Fen Kong, Lin Lin, Si-Qi Liu, and Da-Zhi Wang. Dissecting microbial community structure and metabolic activities at an oceanic deep chlorophyll maximum layer by size-fractionated metaproteomics. Progress in Oceanography, 188(??):??, October 2020. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661120301786.

Xu:2020:IZM

[XLL^{+}20] Zhimeng Xu, Yifan Li, Yanhong Lu, Yingdong Li, Zhongwei Yuan, Minhan Dai, and Hongbin Liu. Impacts of the zhemin coastal current on the biogeographic pattern of microbial eukaryotic communities. Progress in Oceanography, 183(??): ??, April 2020. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661120300483.

Xie:2020:SSP

[XLX^{+}20] Yuyuan Xie, Lizhen Lin, Wupeng Xiao, Xiaolong Yu, Wenlu Lan, and Bangqin Huang. Striking seasonal pattern of primary production in the river-dominated ocean margin of the northern South China Sea (NSCS-RiOMar) revealed by new field and remotely sensed data. Progress in Oceanography, 189(??):??, November–December 2020. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661120302068.

Xu:2017:IFS

[XNT^{+}17] Yi Xu, Karen Nieto, Steven L. H. Teo, Sam McClatchie, and John Holmes. Influence of fronts on the spatial distribution of albacore tuna (Thunnus alalunga) in the Northeast Pacific over the past 30 years (1982–2011). Progress in Oceanography, 150(??):72–78, January 2017. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661115000737.

Xu:2015:DSV

[XRC^{+}15] Yi Xu, Kenneth A. Rose, Fei Chai, Francisco P. Chavez, and Patricia Ayón. Does spatial variation in environmental conditions affect recruitment? A study using a 3-D model of Peruvian anchovy. Progress in Oceanography, 138(??):417–430, November 2015. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661115000567.

Xiao:2018:RNE

[XWL^{+}18] Wupeng Xiao, Lei Wang, Edward Laws, Yuyuan Xie, Jixin Chen, Xin Liu, Bingzhang Chen, and Bangqin Huang. Realized niches explain spatial gradients in seasonal abundance of phytoplankton groups in the South China Sea. Progress in Oceanography, 162(??):223–239, March 2018. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661117302318.

Xie:2022:ENB

[XWL^{+}22] Xiaohui Xie, Yan Wang, Xiaohui Liu, Jun Wang, Dongfeng Xu, Tongya Liu, Jinlin Ji, Dongsheng Zhang, Chunsheng Wang, and Dake Chen. Enhanced near-bottom circulation and mixing driven by the surface eddies over abyssal seamounts. Progress in Oceanography, 208(??):??, November 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661122001550.

Xu:2021:SIV

[XWW^{+}21] Yi Xu, Ying Wu, Huiwu Wang, Zhenqiu Zhang, Jian Li, and Jing Zhang. Seasonal and interannual variabilities of chlorophyll across the eastern equatorial Indian Ocean and Bay of Bengal. Progress in Oceanography, 198(??):??, November 2021. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661121001464.

Xu:2020:EFW

[XY20] Yue Xu and Xiping Yu. Enhanced formulation of wind energy input into waves in developing sea. Progress in Oceanography, 186(??):??, July 2020. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661120301154.

Xu:2021:EAW

[XY21] Yue Xu and Xiping Yu. Enhanced atmospheric wave boundary layer model for evaluation of wind stress over waters of finite depth. Progress in Oceanography, 198(??):??, November 2021. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S007966112100149X.

Xiao:2023:EMC

[XYGJ23] Xilin Xiao, Youhei Yamashita, Michael Gonsior, and Nianzhi Jiao. The efficiency of the microbial carbon pump as seen from the relationship between apparent oxygen utilization and fluorescent dissolved organic matter. Progress in Oceanography, 210(??):??, January 2023. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661122001884.

Xia:2022:MCD

[XYK^{+}22] Shujuan Xia, Takashi Yamakawa, Mari Kuroki, Toshihiro Horiguchi, Keita Kodama, Hiroaki Shiraishi, and Makoto Shimizu. Multidecadal changes in the demersal community structure in an urban bay: Shift from a bottom- to a topheavy pyramid. Progress in Oceanography, 203(??):??, April– May 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661122000283.

Xing:2022:AFH

[XYL^{+}22] Qinwang Xing, Haiqing Yu, Yang Liu, Jianchao Li, Yongjun Tian, Andrew Bakun, Chang Cao, Hao Tian, and Wenjia Li. Application of a fish habitat model considering mesoscale oceanographic features in evaluating climatic impact on distribution and abundance of Pacific saury (Cololabis saira). Progress in Oceanography, 201(??):??, February 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661122000052.

Xing:2023:SWT

[XYWY23] Qinwang Xing, Haiqing Yu, Hui Wang, and Huaming Yu. A sliding-window-threshold algorithm for identifying global mesoscale ocean fronts from satellite observations. Progress in Oceanography, 216(??):??, August 2023. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661123001155.

Yasuda:2014:DTP

[YAI^{+}14] Tomoki Yasuda, Yoshihiro Asahara, Ryo Ichikawa, Takeshi Nakatsuka, Hideki Minami, and Seiya Nagao. Distribution and transport processes of lithogenic material from the Amur River revealed by the Sr and Nd isotope ratios of sediments from the Sea of Okhotsk. Progress in Oceanography, 126(??):155–167, August 2014. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661114000639.

Yatsu:2008:EDR

[YAK^{+}08] A. Yatsu, K. Y. Aydin, J. R. King, G. A. McFarlane, S. Chiba, K. Tadokoro, M. Kaeriyama, and Y. Watanabe. Elucidating dynamic responses of North Pacific fish populations to climatic forcing: Influence of life-history strategy. Progress in Oceanography, 77(2–3):252–268, May 2008. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661108000608.

Yoon:2013:RMC

[YAK13] Seokjin Yoon, Hiroya Abe, and Michio J. Kishi. Responses of Manila clam growth and its food sources to global warming in a subarctic lagoon in Japan. Progress in Oceanography, 119(??):48–58, December 2013. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661113000864.

YaoXiang:1988:TSD

[Yao88] Huang YaoXiang. Temperature and salinity distributions in the South China Sea and adjacent waters. Progress in Oceanography, 21(3–4):493–501, ???? 1988. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0079661188900237.

Yashayaev:2007:CFC

[Yas07a] Igor Yashayaev. Changing freshwater content: Insights from the subpolar North Atlantic and new oceanographic challenges. Progress in Oceanography, 73(3–4):203–209, May 2007. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661107000833.

Yashayaev:2007:EHC

[Yas07b] Igor Yashayaev. Erratum: Hydrographic changes in the Labrador Sea, 1960–2005 [Prog. Oceanogr. 73 (2007) 242– 276]. Progress in Oceanography, 75(4):857–859, December 2007. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S007966110700184X. See [Yas07c].

Yashayaev:2007:HCL

[Yas07c] Igor Yashayaev. Hydrographic changes in the Labrador Sea, 1960–2005. Progress in Oceanography, 73(3–4):242–276, May 2007. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661107000857. See erra-

tum [Yas07b].

Yoo:2008:SIE

[YBPS08] Sinjae Yoo, Harold P. Batchelder, William T. Peterson, and William J. Sydeman. Seasonal, interannual and event scale variation in North Pacific ecosystems. Progress in Oceanography, 77(2–3):155–181, May 2008. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661108000529.

Yanez:2001:CVP

[YBS^{+}01] E. Yáñez, M. A. Barbieri, C. Silva, K. Nieto, and F. Espíndola. Climate variability and pelagic fisheries in northern Chile. Progress in Oceanography, 49(??):581–596, ???? 2001. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661101000428.

Yannicelli:2012:DPM

[YCP^{+}12] Beatriz Yannicelli, Leonardo Castro, Carolina Parada, Wolfgang Schneider, Francois Colas, and David Donoso. Distribution of Pleuroncodes monodon larvae over the continental shelf of south-central Chile: Field and modeling evidence for partial local retention and transport. Progress in Oceanography, 92–95(??):206–227, January 2012. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661111000590.

Yamada:2021:RBO

[YFK21] Sylvia Behrens Yamada, Jennifer L. Fisher, and P. Michael Kosro. Relationship between ocean ecosystem indicators and year class strength of the invasive European green crab (Carcinus maenas). Progress in Oceanography, 196(??):??, August 2021. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661121001051.

Yoshie:2005:ECA

[YFY05] Naoki Yoshie, Masahiko Fujii, and Yasuhiro Yamanaka. Ecosystem changes after the SEEDS iron fertilization in the western North Pacific simulated by a one-dimensional ecosystem model. Progress in Oceanography, 64(2–4):283–306, February 2005. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661105000352.

Yoneda:2022:BPD

[YFY^{+}22] Michio Yoneda, Tatsunori Fujita, Masayuki Yamamoto, Kazuaki Tadokoro, Yuji Okazaki, Masahiro Nakamura, Masanori Takahashi, Naoaki Kono, Tadashi Matsubara, Katsuyuki Abo, Guo Xinyu, and Naoki Yoshie. Bottom-up processes drive reproductive success of Japanese anchovy in an oligotrophic sea: a case study in the central Seto Inland Sea, Japan. Progress in Oceanography, 206(??):??, August 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661122001197.

Yang:2021:SDC

[YGC^{+}21] Wei Yang, Xianghui Guo, Zhimian Cao, Yi Xu, Lifang Wang, Liguo Guo, Tao Huang, Yan Li, Yanping Xu, Jianping Gan, and Minhan Dai. Seasonal dynamics of the carbonate system under complex circulation schemes on a large continental shelf: the northern South China Sea. Progress in Oceanography, 197(??):??, September–October 2021. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661121001166.

Young:2010:DPS

[YGL^{+}10] Jock W. Young, Michaela A. Guest, Matt Lansdell, Charles F. Phleger, and Peter D. Nichols. Discrimination of prey species of juvenile swordfish Xiphias gladius (Linnaeus, 1758) using signature fatty acid analyses. Progress in Oceanography, 86(1–2):139–151, July 2010. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661110000625.

Yunda-Guarin:2023:TEE

[YGMR^{+}23] Gustavo Yunda-Guarin, Loïc N. Michel, Virginie Roy, Noémie Friscourt, Michel Gosselin, Christian Nozais, and Philippe Archambault. Trophic ecology of epibenthic communities exposed to different sea-ice concentrations across the Canadian Arctic Ocean. Progress in Oceanography, 217(??):??, September– October 2023. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661123001489.

Yebra:2004:EUF

[YHLA^{+}04] L. Yebra, S. Hernández-León, C. Almeida, P. Bécognée, and J. M. Rodríguez. The effect of upwelling filaments and islandinduced eddies on indices of feeding, respiration and growth in copepods. Progress in Oceanography, 62(2–4):151–169, August 2004. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661104000977.

Yebra:2018:ZPC

[YHM^{+}18] Lidia Yebra, Inma Herrera, Jesús M. Mercado, Dolores Cortés, Francisco Gómez-Jakobsen, Aitor Alonso, Antonio Sánchez, Soluna Salles, and Nerea Valcárcel-Pérez. Zooplankton production and carbon export flux in the western Alboran Sea Gyre (SW Mediterranean). Progress in Oceanography, 167(??):64– 77, October 2018. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661117302070.

Yanez:2022:CSP

[YHRT22] Sonia Yáñez, Pamela Hidalgo, Paula Ruz, and Kam W. Tang. Copepod secondary production in the sea: Errors due to uneven molting and growth patterns and incidence of carcasses. Progress in Oceanography, 165(??):257–267, ???? 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661117302422.

Yang:2022:TDS

[YHZ^{+}22] Yunchao Yang, Xiaodong Huang, Chun Zhou, Zhiwei Zhang, Wei Zhao, and Jiwei Tian. Three-dimensional structures of internal solitary waves in the northern south China sea revealed by mooring array observations. Progress in Oceanography, 209(??):??, December 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661122001665.

Yin:1988:SWT

[Yin88] Fuh Yin. Sea water temperature variation east of Taiwan. Progress in Oceanography, 21(3–4):457–467, ???? 1988. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0079661188900213.

Yamada:2004:SIV

[YIY^{+}04] Keiko Yamada, Joji Ishizaka, Sinjae Yoo, Hyun cheol Kim, and Sanae Chiba. Seasonal and interannual variability of sea surface chlorophyll-a concentration in the Japan/East Sea (JES). Progress in Oceanography, 61(2–4):193–211, May 2004. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S007966110400062X.

Yaochu:1988:CKCa

[YJ88] Yuan Yaochu and Su Jilan. The calculation of Kuroshio Current structure in the East China Sea — early summer 1986. Progress in Oceanography, 21(3–4):343–361, ???? 1988. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0079661188900134.

Yaochu:1986:DMS

[YJS86] Yuan Yaochu, Su Jilan, and Xia Sangyun. A diagnostic model of summer circulation on the northwest shelf of the East China Sea. Progress in Oceanography, 17(3–4):163–176, ???? 1986. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/007966118690042X.

oYaochu:1988:CKCb

[YJW88] Yuan Yaochu, Su Jilan, and Zhou Weidong. Calculation of the Kuroshio Current South of Japan in May–June 1986. Progress in Oceanography, 21(3–4):503–514, ???? 1988. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/0079661188900249.

Yamazaki:2024:RCW

[YKH^{+}24] Kaihe Yamazaki, Katsuro Katsumata, Daisuke Hirano, Daiki Nomura, Hiroko Sasaki, Hiroto Murase, and Shigeru Aoki. Revisiting circulation and water masses over the East Antarctic margin (80–150^{o}E). Progress in Oceanography, 225(??): ??, July 2024. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661124000910.

Yasuda:2023:VMP

[YKNO23] Tohya Yasuda, Junji Kinoshita, Yohei Niino, and Junichi Okuyama. Vertical migration patterns linked to body and environmental temperatures in chub mackerel. Progress in Oceanography, 213(??):??, April–May 2023. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661123000605.

Yamaguchi:2012:SSI

[YKS^{+}12] Hisashi Yamaguchi, Hyun-Cheol Kim, Young Baek Son, Sang Woo Kim, Kazumaro Okamura, Yoko Kiyomoto, and Joji Ishizaka. Seasonal and summer interannual variations of SeaWiFS chlorophyll-a in the Yellow Sea and East China Sea. Progress in Oceanography, 105(??):22–29, October 2012. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661112000377.

Yabe:2021:AST

[YKWF21] Itsuka Yabe, Yusuke Kawaguchi, Taku Wagawa, and Shinzou Fujio. Anatomical study of Tsushima Warm Current system: Determination of principal pathways and its variation. Progress in Oceanography, 194(??):??, June 2021. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S007966112100077X.

Yang:2019:TIM

[YLL19] Eun Jin Yang, Youngju Lee, and SangHoon Lee. Trophic interactions of micro- and mesozooplankton in the Amundsen Sea polynya and adjacent sea ice zone during austral late summer. Progress in Oceanography, 174(??):117–130, May 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661118301265.

Yuan:2014:SKI

[YLY^{+}14] Yaochu Yuan, Guanghong Liao, Chenghao Yang, Zenghong Liu, Hong Chen, and Zhang-Gui Wang. Summer Kuroshio Intrusion through the Luzon Strait confirmed from observations and a diagnostic model in summer 2009. Progress in Oceanography, 121(??):44–59, February 2014. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661113001894.

Yamaguchi:2017:LVA

[YMA^{+}17] A. Yamaguchi, K. Matsuno, Y. Abe, D. Arima, and I. Imai. Latitudinal variations in the abundance, biomass, taxonomic composition and estimated production of epipelagic mesozooplankton along the 155^{o}E longitude in the western North Pacific during spring. Progress in Oceanography, 150(??):13–19, January 2017. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661115000713.

Yamashiro:1988:MTC

[YMI88] Toru Yamashiro, Akio Maeda, and Hiroshi Ichikawa. M_{2} tidal currents near the continental shelf margin in the East China Sea. Progress in Oceanography, 21(3–4):431–440, ???? 1988. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/0079661188900195.

Yun:2004:STV

[YMK^{+}04] Jae-Yul Yun, Lorenz Magaard, Kuh Kim, Chang-Woong Shin, Cheolsoo Kim, and Sang-Kyung Byun. Spatial and temporal variability of the North Korean cold water leading to the near-bottom cold water intrusion in Korea Strait. Progress in Oceanography, 60(1):99–131, January 2004. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661104000047.

Yoshida:2003:NED

[YN03a] J. Yoshida and H. Nagashima. Numerical experiments on double-diffusive intrusions in the ocean and their relation to laboratory experiments. Progress in Oceanography, 56(3– 4):549–557, March 2003. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661103000302.

Yoshida:2003:NES

[YN03b] Jiro Yoshida and Hideki Nagashima. Numerical experiments on salt-finger convection. Progress in Oceanography, 56(3–4):435– 459, March 2003. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661103000326.

Yeo:2020:SSV

[YN20] Da-Eun Yeo and SungHyun Nam. Seasonal and spatial variations of air-sea heat exchange in the seas around the Korean Peninsula: Based on the observations and reanalysis products from 2011 to 2016. Progress in Oceanography, 181(??):??, February 2020. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661119304197.

Yanagimoto:2002:ICB

[YNM^{+}02] T. Yanagimoto, A. Nishimura, K. Mito, Y. Takao, and N. J. Williamson. Interannual changes of biological properties of walleye pollock Theragra chalcogramma in the central Bering Sea. Progress in Oceanography, 55(1–2):195–208, October 2002. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661102000782.

Yamamae:2023:VCZ

[YNMY23] Kunito Yamamae, Yasuhide Nakamura, Kohei Matsuno, and Atsushi Yamaguchi. Vertical changes in zooplankton abundance, biomass, and community structure at seven stations down to 3000m in neighboring waters of Japan during the summer: Insights from ZooScan imaging analysis. Progress in Oceanography, 219(??):??, December 2023. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661123001982.

Yan:2022:WMC

[YNTS22] Dong Yan, Jun Nishioka, Takenobu Toyota, and Koji Suzuki. Winter microalgal communities of the southern Sea of Okhotsk: a comparison of sea ice, coastal, and basinal seawater. Progress in Oceanography, 204(??):??, June 2022. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661122000672.

Yoshida:1980:CUR

[Yos80] Kozo Yoshida. The coastal undercurrent — a role of longshore scales in coastal upwelling dynamics. Progress in Oceanography, 9(2):83–131, ???? 1980. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0079661180900014.

Yanez:2010:AER

[YPGE^{+}10] Eleuterio Yáñez, Francisco Plaza, Juan Carlos Gutiérrez-Estrada, Nibaldo Rodríguez, M. A. Barbieri, Inmaculada Pulido-Calvo, and Cinthya Bórquez. Anchovy (Engraulis ringens) and sardine (Sardinops sagax) abundance forecast off northern Chile: a multivariate ecosystemic neural network approach. Progress in Oceanography, 87(1–4):242–250, October 2010. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661110001278.

Yeh:2010:ACS

[YPM^{+}10] Sang-Wook Yeh, Young-Gyu Park, HongSik Min, Cheol-Ho Kim, and Jae-Hak Lee. Analysis of characteristics in the sea surface temperature variability in the East/Japan Sea. Progress in Oceanography, 85(3–4):213–223, June 2010. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661110000327.

Yebra:2022:STV

[YPVP^{+}22] Lidia Yebra, Marta Puerto, Nerea Valcárcel-Pérez, Sébastien Putzeys, Francisco Gómez-Jakobsen, Candela García-Gómez, and Jesús M. Mercado. Spatio-temporal variability of the zooplankton community in the SW Mediterranean 1992– 2020: Linkages with environmental drivers. Progress in Oceanography, 203(??):??, April–May 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S007966112200043X.

You:2008:MTD

[YRKC08] Yuzhu You, Joachim Ribbe, Johannes Karstensen, and John Church. Matthias tomczak: a distinguished regional oceanographer and water-mass expert. Progress in Oceanography, 77(4):269–272, June 2008. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S007966110800061X.

Yashayaev:2015:RAW

[YS15] Igor Yashayaev and Dan Seidov. The role of the Atlantic Water in multidecadal ocean variability in the Nordic and Barents Seas. Progress in Oceanography, 132(??):68–127, March 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661114001980.

Yashayaev:2015:NCV

[YSD15] Igor Yashayaev, Dan Seidov, and Entcho Demirov. A new collective view of oceanography of the Arctic and North Atlantic basins. Progress in Oceanography, 132(??):1–21, March 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661114002225.

Yumruktepe:2020:CCE

[YSN20] Veli Çağlar Yumruktepe, Barış Salihoğlu, and Susanne Neuer. Controls on carbon export in the subtropical North Atlantic. Progress in Oceanography, 187(??):??, August–September 2020. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661120301191.

Yoda:2014:FSS

[YSS14] Ken Yoda, Kozue Shiomi, and Katsufumi Sato. Foraging spots of streaked shearwaters in relation to ocean surface currents as identified using their drift movements. Progress in Oceanography, 122(??):54–64, March 2014. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661113002309.

Yang:2019:LSP

[YSY^{+}19] Shixiong Yang, Bing Song, Siyuan Ye, Edward A. Laws, Lei He, Jie Li, Jinxia Chen, Guangming Zhao, Jingtao Zhao, Xi Mei, and Hermann Behling. Large-scale pollen distribution in marine surface sediments from the Bohai Sea, China: Insights into pollen provenance, transport, deposition, and coastalshelf paleoenvironment. Progress in Oceanography, 178(??):??, November 2019. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661119300576.

Yamada:2006:ICC

[YT06] Namiha Yamada and Eiichiro Tanoue. The inventory and chemical characterization of dissolved proteins in oceanic waters. Progress in Oceanography, 69(1):1–18, April 2006. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S007966110500162X.

Yamashita:2021:WDA

[YTB^{+}21] Youhei Yamashita, Tetsu Tosaka, Rise Bamba, Ryuichi Kamezaki, Shuji Goto, Jun Nishioka, Ichiro Yasuda, Toru Hirawake, Joji Oida, Hajime Obata, and Hiroshi Ogawa. Widespread distribution of allochthonous fluorescent dissolved organic matter in the intermediate water of the North Pacific. Progress in Oceanography, 191(??):??, February 2021. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661120302457.

Yoo:2019:SNK

[YTL^{+}19] Hyunsu Yoo, Hayato Tanaka, Wonchoel Lee, Simone Nunes Brandão, and Ivana Karanovic. Six new Krithe from the Kuril– Kamchatka Trench, with the first insight into phylogeography of deep-sea ostracods. Progress in Oceanography, 176(??):??, September 2019. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661119301326.

Yasuda:2000:HBR

[YTNK00] Ichiro Yasuda, Tomoki Tozuka, Masayuki Noto, and Shinya Kouketsu. Heat balance and regime shifts of the mixed layer in the Kuroshio Extension. Progress in Oceanography, 47(2–4):257–278, October 2000. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661100000380.

Yu:2023:CSS

[Yu23] Lisan Yu. Connecting subtropical salinity maxima to tropical salinity minima: Synchronization between ocean dynamics and the water cycle. Progress in Oceanography, 219(??):??, December 2023. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S007966112300215X.

Yuxiang:1988:NMT

[Yux88] Tang Yuxiang. Numerical modelling of the tide-induced residual current in the East China Sea. Progress in Oceanography, 21(3–4):417–429, ???? 1988. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0079661188900183.

Yao:2022:MHC

[YW22] Yulong Yao and Chunzai Wang. Marine heatwaves and coldspells in global coral reef zones. Progress in Oceanography, 209(??):??, December 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/S0079661122001793.

Yoon:2015:PHC

[YWUK15] Seokjin Yoon, Eiji Watanabe, Hiromichi Ueno, and Michio J. Kishi. Potential habitat for chum salmon (Oncorhynchus keta) in the Western Arctic based on a bioenergetics model coupled with a three-dimensional lower trophic ecosystem model. Progress in Oceanography, 131(??):146–158, February 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661114002195.

Yang:2018:OIK

[YYC^{+}18] Dezhou Yang, Baoshu Yin, Fei Chai, Xingru Feng, Huijie Xue, Guandong Gao, and Fei Yu. The onshore intrusion of Kuroshio subsurface water from February to July and a mechanism for the intrusion variation. Progress in Oceanography, 167(??):97–115, October 2018. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661118300235.

Yuan:2017:ALP

[YYhT^{+}17] Yaochu Yuan, Chenghao Yang, Yu heng Tseng, Xiao-Hua Zhu, Huiqun Wang, and Hong Chen. Analysis of longer period variation of the Kuroshio Current intrusion into the Luzon Strait using rectified wavelet power spectra. Progress in Oceanography, 156(??):61–77, August 2017. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661116301987.

Yasui:1988:WTP

[YYK88] Masashi Yasui, Makoto Yamano, and Masataka Kinoshita. Warped temperature profiles observed in the sediment of geothermally active areas in the Japan Sea — a hypothesis of double diffusive sinking. Progress in Oceanography, 21(3–4):255–267, ???? 1988. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/0079661188900067.

Yoon:2012:URS

[YYK^{+}12] Jinhee Yoon, Sang-Wook Yeh, Young-Ho Kim, Jong-Seong Kug, and Hong Sik Min. Understanding the responses of sea surface temperature to the two different types of El Niño in the western North Pacific. Progress in Oceanography, 105(??):81– 89, October 2012. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661112000407.

Yagi:2014:RET

[YYT^{+}14] Masahiro Yagi, Ichiro Yasuda, Takahiro Tanaka, Yuki Tanaka, Kazuya Ono, Kay I. Ohshima, and Katsurou Katsumata. Reevaluation of turbulent mixing vertical structure in the Bussol’ Strait and its impact on water masses in the Okhotsk Sea and the North Pacific. Progress in Oceanography, 126(??):121–134, August 2014. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661114000718.

Yin:2023:PIO

[YZX^{+}23] Jie Yin, Chongliang Zhang, Ying Xue, Yupeng Ji, Binduo Xu, and Yiping Ren. Potential impacts of ocean warming on the trophic control of a threatened marine ecosystem. Progress in Oceanography, 211(??):??, February 2023. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661123000198.

Zhang:2023:CSE

[ZAC^{+}23] Weifeng (Gordon) Zhang, Philip Alatalo, Taylor Crockford, Andrew J. Hirzel, Meredith G. Meyer, Hilde Oliver, Emily Peacock, Christian M. Petitpas, Zoe Sandwith, Walker O. Smith, Heidi M. Sosik, Rachel H. R. Stanley, Bethany L. F. Stevens, Jefferson T. Turner, and Dennis J. McGillicuddy. Cross-shelf exchange associated with a shelf-water streamer at the Mid-Atlantic Bight shelf edge. Progress in Oceanography, 210(??): ??, January 2023. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661122001902.

Zavatarelli:1999:MSM

[Zav99] Marco Zavatarelli. Mediterranean Sea multiscale variability and environmental management issues: a scientific perspective. Progress in Oceanography, 44(1–3):401–409, August 1999. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661199000348.

Zou:2023:DOC

[ZBLF23] Sijia Zou, Amy S. Bower, M. Susan Lozier, and Heather H. Furey. Deep ocean circulation in the subpolar North Atlantic observed by acoustically-tracked floats. Progress in Oceanography, 211(??):??, February 2023. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661123000186.

Ziegler:2023:WSB

[ZBRJ23] Amanda F. Ziegler, Bodil A. Bluhm, Paul E. Renaud, and Lis L. Jørgensen. Weak seasonality in benthic food web structure within an Arctic inflow shelf region. Progress in Oceanography, 217(??):??, September–October 2023. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661123001520.

Zhang:2022:AMP

[ZBY^{+}22] Miming Zhang, Dennis Booge, Jinpei Yan, Suqing Xu, Chen Liang, Yanfang Wu, Bo Yang, Jianjun Wang, Jun Zhao, Dong Li, Jianming Pan, and Keyhong Park. Abundant microzooplankton possibly cause ultrahigh seawater dimethylsulfide during Southern Ocean algal blooms. Progress in Oceanography, 202(??):??, March 2022. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661122000064.

Zhang:2021:EPI

[ZCA21] Weifeng (Gordon) Zhang, Zhen Cheng, and Andrew D. Ashton. Exploring the potential for internal tides to reshape the continental shelf edge seafloor. Progress in Oceanography, 195(??): ??, July 2021. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661121000628.

Zahariev:2008:PHF

[ZCD08] Konstantin Zahariev, James R. Christian, and Kenneth L. Denman. Preindustrial, historical, and fertilization simulations using a global ocean carbon model with new parameterizations of iron limitation, calcification, and N_{2} fixation. Progress in Oceanography, 77(1):56–82, April 2008. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661108000104.

Zhou:2017:IHC

[ZCH^{+}17] Feng Zhou, Fei Chai, Daji Huang, Huijie Xue, Jianfang Chen, Peng Xiu, Jiliang Xuan, Jia Li, Dingyong Zeng, Xiaobo Ni, and Kui Wang. Investigation of hypoxia off the Changjiang Estuary using a coupled model of ROMS-CoSiNE. Progress in Oceanography, 159(??):237–254, December 2017. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661117301210.

Ziegler:2020:IDR

[ZCLS20] A. F. Ziegler, M. Cape, Ø. Lundesgaard, and C. R. Smith. Intense deposition and rapid processing of seafloor phytodetritus in a glaciomarine fjord, Andvord Bay (Antarctica). Progress in Oceanography, 187(??):??, August–September 2020. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S007966112030152X.

Zenteno:2019:UMB

[ZCV^{+}19] L. Zenteno, L. Cárdenas, N. Valdivia, I. Gómez, J. Höfer, I. Garrido, and L. M. Pardo. Unraveling the multiple bottomup supplies of an Antarctic nearshore benthic community. Progress in Oceanography, 174(??):55–63, May 2019. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661118300909.

Zheng:2017:UAS

[ZD17] Guangming Zheng and Paul M. DiGiacomo. Uncertainties and applications of satellite-derived coastal water quality products. Progress in Oceanography, 159(??):45–72, December 2017. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661116302592.

Zhao:2021:RDD

[ZDG^{+}21] Huade Zhao, Minhan Dai, Jianping Gan, Xiaozheng Zhao, Zhongming Lu, Linlin Liang, Zhiqiang Liu, Jianzhong Su, and Zhimian Cao. River-dominated pCO_{2} dynamics in the northern South China Sea during summer: a modeling study. Progress in Oceanography, 190(??):??, January 2021. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661120301944.

Zhou:2020:IPB

[ZDM^{+}20] Kuanbo Zhou, Minhan Dai, Kanchan Maiti, Weifang Chen, Junhui Chen, Qingquan Hong, Yifan Ma, Peng Xiu, Lei Wang, and Yuyuan Xie. Impact of physical and biogeochemical forcing on particle export in the South China Sea. Progress in Oceanography, 187(??):??, August–September 2020. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661120301427.

Zenk:2008:TFC

[Zen08] Walter Zenk. Temperature fluctuations and current shear in Antarctic Bottom Water at the Vema Sill. Progress in Oceanography, 77(4):276–284, June 2008. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661108000633.

Zezina:1990:WDS

[Zez90] O. N. Zezina. What do the simple morphological features of deep-sea animals mean? Progress in Oceanography, 24 (??):253–255, ???? 1990. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic). URL http://www.

sciencedirect.com/science/article/pii/007966119090035Z.

Zuniga:2009:PFD

[ZFSV^{+}09] Diana Zúñiga, M. Mar Flexas, Anna Sanchez-Vidal, Johan Coenjaerts, Antoni Calafat, Gabriel Jordà, Jordi García-Orellana, Joan Puigdefàbregas, Miquel Canals, Manuel Espino, Francesc Sardà, and Joan B. Company. Particle fluxes dynamics in Blanes submarine canyon (Northwestern Mediterranean). Progress in Oceanography, 82(4):239–251, September 2009. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661109000548.

Zagami:2020:SCV

[ZGB^{+}20] Giacomo Zagami, Antonia Granata, Cinzia Brugnano, Roberta Minutoli, Vincenzo Bonanzinga, and Letterio Guglielmo. Spring copepod vertical zonation pattern and diel migration in the open Ligurian Sea (north-western Mediterranean). Progress in Oceanography, 183(??):??, April 2020. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661120300367.

Zhang:2019:TES

[ZGZ19] Jing Zhang, Xinyu Guo, and Liang Zhao. Tracing external sources of nutrients in the East China Sea and evaluating their contributions to primary production. Progress in Oceanography, 176(??):??, September 2019. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661118301046.

Zonneveld:2001:GDO

[ZHBW01] Karin A. F. Zonneveld, Ramses P. Hoek, Henk Brinkhuis, and Helmut Willems. Geographical distributions of organic-walled dinoflagellate cysts in surficial sediments of the Benguela upwelling region and their relationship to upper ocean conditions. Progress in Oceanography, 48(1):25–72, ???? 2001.

CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661100000471.

Zhang:2020:LPE

[ZHD^{+}20] Zhaoru Zhang, Eileen E. Hofmann, Michael S. Dinniman, Christian Reiss, Walker O. Smith, and Meng Zhou. Linkage of the physical environments in the northern Antarctic Peninsula region to the southern annular mode and the implications for the phytoplankton production. Progress in Oceanography, 188(??):??, October 2020. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661120301555.

Zheng:2024:ESW

[ZHF^{+}24] Youchang Zheng, Jingwen Huang, Yang Feng, Huijie Xue, Xiaohui Xie, Hanqin Tian, Yuanzhi Yao, Lin Luo, Xianghui Guo, and Yonglin Liu. The effects of seasonal wind regimes on the evolution of hypoxia in Chesapeake Bay: Results from a terrestrial–estuarine–ocean biogeochemical modeling system. Progress in Oceanography, 222(??):??, March 2024. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661124000132.

Zapata-Hernandez:2014:BFW

[ZHSMM14] Germán Zapata-Hernández, Javier Sellanes, Christoph Mayr, and Práxedes Muñoz. Benthic food web structure in the Comau fjord, Chile (≈ 42^{o}S): Preliminary assessment including a site with chemosynthetic activity. Progress in Oceanography, 129(??):149–158, December 2014. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661114000317.

Zhuang:2021:IAC

[ZJZ^{+}21] Yanpei Zhuang, Haiyan Jin, Yang Zhang, Hongliang Li, Tianzhen Zhang, Yangjie Li, Youcheng Bai, Jian Ren, and Jianfang Chen. Incursion of Alaska Coastal Water as a mechanism promoting small phytoplankton in the western Arctic Ocean. Progress in Oceanography, 197(??):??, September– October 2021. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661121001245.

Zheng:2006:RVE

[ZK06] Jie Zheng and Gordon H. Kruse. Recruitment variation of eastern Bering Sea crabs: Climate-forcing or top-down effects? Progress in Oceanography, 68(2–4):184–204, February 2006. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S007966110600005X.

Zhou:2016:IDO

[ZKK^{+}16] J. Zhou, M. Kotovitch, H. Kaartokallio, S. Moreau, J.- L. Tison, G. Kattner, G. Dieckmann, D. N. Thomas, and B. Delille. The impact of dissolved organic carbon and bacterial respiration on pCO_{2} in experimental sea ice. Progress in Oceanography, 141(??):153–167, February 2016. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661115002517.

Zhang:1988:PCS

[ZKT88] Ruo-Chao Zhang, Hiroshi Kawamura, and Yoshiaki Toba. Propagation characteristics of swell incoming from the deep ocean into the seas East of China. Progress in Oceanography, 21(3–4):441–455, ???? 1988. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/0079661188900201.

Zhang:2001:SAM

[ZL01] Chang Ik Zhang and Jae Bong Lee. Stock assessment and management implications of horse mackerel (Trachurus japonicus) in Korean waters, based on the relationships between recruitment and the ocean environment. Progress in Oceanography, 49(??):513–537, ???? 2001. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661101000386.

Zhu:2024:OHC

[ZL24] Qing Zhu and Zhongping Lee. Offshore high chlorophyll eddies: Separating upwelling from nearshore migration from ocean color remote sensing. Progress in Oceanography, 223(??):??, April–May 2024. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661124000387.

Zhao:2015:SST

[ZLC^{+}15] Jingtao Zhao, Jun Li, Feng Cai, Helong Wei, Bangqi Hu, Yanguang Dou, Libo Wang, Rong Xiang, Hongwei Cheng, Liang Dong, and Chuanlun L. Zhang. Sea surface temperature variation during the last deglaciation in the southern Okinawa Trough: Modulation of high latitude teleconnections and the Kuroshio Current. Progress in Oceanography, 138 (part A)(??):238–248, November 2015. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661115001470.

Zarubin:2017:DCM

[ZLG17a] Margarita Zarubin, Yoav Lindemann, and Amatzia Genin. The dispersion-confinement mechanism: Phytoplankton dynamics and the spring bloom in a deeply-mixing subtropical sea. Progress in Oceanography, 155(??):13–27, June 2017. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661117301659.

Zarubin:2017:WDC

[ZLG17b] Margarita Zarubin, Yoav Lindemann, and Amatzia Genin. Withdrawn: the dispersion-confinement mechanism: Phytoplankton dynamics and the spring bloom in a deeply-mixing subtropical sea. Progress in Oceanography, 155(??):74, June 2017. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661117301519.

Zhang:2000:CRS

[ZLKO00] Chang Ik Zhang, Jae Bong Lee, Suam Kim, and Jai-Ho Oh. Climatic regime shifts and their impacts on marine ecosystem and fisheries resources in Korean waters. Progress in Oceanography, 47(2–4):171–190, October 2000. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661100000355.

Zhang:2007:NGE

[ZLR^{+}07] J. Zhang, S. M. Liu, J. L. Ren, Y. Wu, and G. L. Zhang. Nutrient gradients from the eutrophic Changjiang (Yangtze River) Estuary to the oligotrophic Kuroshio waters and re-evaluation of budgets for the East China Sea Shelf. Progress in Oceanography, 74(4):449–478, September 2007. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661107000961.

Zhou:2024:CML

[ZLRVB24a] Xing Zhou, Luisa Lopera, Adela Roa-Varón, and Annalisa Bracco. Corrigendum to “Modeling the larval dispersal and connectivity of Red Snapper (Lutjanus campechanus) in the Northern Gulf of Mexico” [Prog. Oceanogr. 224 (2024) 103265]. Progress in Oceanography, 225(??):??, July 2024. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661124000909.

Zhou:2024:MLD

[ZLRVB24b] Xing Zhou, Luisa Lopera, Adela Roa-Varón, and Annalisa Bracco. Modeling the larval dispersal and connectivity of red snapper (Lutjanus campechanus) in the Northern Gulf of Mexico. Progress in Oceanography, 224(??):??, June 2024. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661124000715.

Zhang:2004:VAF

[ZLS^{+}04] Chang Ik Zhang, Jae Bong Lee, Young Il Seo, Sang Cheol Yoon, and Suam Kim. Variations in the abundance of fisheries resources and ecosystem structure in the Japan/ East Sea. Progress in Oceanography, 61(2–4):245–265, May 2004. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661104000655.

Zhong:2020:PCP

[ZLX^{+}20] Yanping Zhong, Xin Liu, Wupeng Xiao, Edward A. Laws, Jixin Chen, Lei Wang, Siguang Liu, Fan Zhang, and Bangqin Huang. Phytoplankton community patterns in the Taiwan Strait match the characteristics of their realized niches. Progress in Oceanography, 186(??):??, July 2020. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661120301051.

Zou:2017:OMP

[ZLZ^{+}17] Sijia Zou, Susan Lozier, Walter Zenk, Amy Bower, and William Johns. Observed and modeled pathways of the Iceland Scotland Overflow Water in the eastern North Atlantic. Progress in Oceanography, 159(??):211–222, December 2017. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661117301404.

Zeppilli:2011:MVM

[ZMCD11] Daniela Zeppilli, Marianna Mea, Cinzia Corinaldesi, and Roberto Danovaro. Mud volcanoes in the Mediterranean Sea are hot spots of exclusive meiobenthic species. Progress in Oceanography, 91(3):260–272, November 2011. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661111000024.

Zheng:2023:MDO

[ZMW^{+}23] Yiling Zheng, Wentao Ma, Yuntao Wang, Zhongfang Liu, and Peng Xiu. Modeling dissolved organic carbon exchange across major straits and shelf breaks in the South China Sea. Progress in Oceanography, 210(??):??, January 2023. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661122001872.

Zhang:2024:MOT

[ZNR^{+}24] J. Zhang, Z. T. Ni, J. L. Ren, F. Yu, X. Y. Diao, Y. Wang, S. J. Zhang, H. Su, S. L. Cong, Z. J. Lu, S. Jiang, J. Ou, Y. Chen, Q. Wang, Z. B. Zhang, J. T. Ai, C. B. Wang, and Z. D. Tao. Modular ocean trace elements sampling for the international GEOTRACES studies – evidence from analysis of dissolved Fe and Pb. Progress in Oceanography, 221(??):??, February 2024. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661124000181.

Zoccarato:2016:WMD

[ZPC^{+}16] Luca Zoccarato, Alberto Pallavicini, Federica Cerino, Serena Fonda Umani, and Mauro Celussi. Water mass dynamics shape Ross Sea protist communities in mesopelagic and bathypelagic layers. Progress in Oceanography, 149(??):16–26, December 2016. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661116301203.

Zhang:2020:ADS

[ZPY^{+}20] Miming Zhang, Ki-Tae Park, Jinpei Yan, Keyhong Park, Yanfang Wu, Eunho Jang, Wei Gao, Guobin Tan, Jianjun Wang, and Liqi Chen. Atmospheric dimethyl sulfide and its significant influence on the sea-to-air flux calculation over the Southern Ocean. Progress in Oceanography, 186(??):??, July 2020. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661120301312.

Zhang:2023:ESI

[ZQWP23] Hua Zhang, Yu-Kun Qian, Xixi Wen, and Shiqiu Peng. Energetics of semidiurnal internal tides near Madagascar in the Southwest Indian Ocean. Progress in Oceanography, 211(??): ??, February 2023. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661123000150.

Zubkov:2000:PCS

[ZSBL00] Mikhail V. Zubkov, Michael A. Sleigh, Peter H. Burkill, and Raymond J. G. Leakey. Picoplankton community structure on the Atlantic Meridional Transect: a comparison between seasons. Progress in Oceanography, 45(3–4):369–386, April 2000. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661100000082.

Zhang:2024:SEE

[ZSH^{+}24] Linlin Zhang, Weiqi Song, Yuchao Hui, Zhenxiao Wang, and Dunxin Hu. Subsurface eddies east of the Philippines: Geographic characteristics, vertical structures, volume and thermohaline transport. Progress in Oceanography, 222(??):??, March 2024. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S007966112400034X.

Zenk:2005:PVA

[ZSI^{+}05] Walter Zenk, Gerold Siedler, Akio Ishida, Jürgen Holfort, Yuji Kashino, Yoshifumi Kuroda, Toru Miyama, and Thomas J. Müller. Pathways and variability of the Antarctic intermediate water in the western equatorial Pacific Ocean. Progress in Oceanography, 67(1–2):245–281, October 2005. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661105001096.

Zhu:2022:OML

[ZSW^{+}22] Yaohua Zhu, Junchuan Sun, Yonggang Wang, Shujiang Li, Tengfei Xu, Zexun Wei, and Tangdong Qu. Overview of the multi-layer circulation in the South China Sea. Progress in Oceanography, 175(??):171–182, ???? 2022. CODEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (electronic).

URL http://www.sciencedirect.com/science/article/
pii/S0079661118301083.

Zhou:2022:IKS

[ZSY^{+}22] Peng Zhou, Xiuxian Song, Yongquan Yuan, Wentao Wang, Lianbao Chi, Xihua Cao, and Zhiming Yu. Intrusion of the Kuroshio subsurface water in the southern East China Sea and its variation in 2014 and 2015 traced by dissolved inorganic iodine species. Progress in Oceanography, 165(??):287– 298, ???? 2022. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661117302902.

Zhang:2015:ULP

[ZWM^{+}15] Qingtian Zhang, Richard M. Warwick, Caroline L. McNeill, Claire E. Widdicombe, Aaron Sheehan, and Stephen Widdicombe. An unusually large phytoplankton spring bloom drives rapid changes in benthic diversity and ecosystem function. Progress in Oceanography, 137(??):533–545, September 2015. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661115000968.

Zimmermann:2023:GAD

[ZWP23] Mark Zimmermann, Rebecca A. Woodgate, and Megan M. Prescott. Gateway to the Arctic: Defining the eastern channel of the Bering Strait. Progress in Oceanography, 215(??): ??, July 2023. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661123000952.

Zhao:2024:EHT

[ZYN^{+}24] Ruixiang Zhao, Fei Yu, Feng Nan, Xiao-Hua Zhu, and Hua Zheng. Eddy heat transport across the Luzon Strait derived from current- and pressure-recording inverted echo sounder observations. Progress in Oceanography, 224(??):??, June 2024. CODEN POCNA8. ISSN 0079-6611 (print), 1873-

4472 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0079661124000405.

Zhang:2024:ECE

[ZZM^{+}24] Mengli Zhang, Chunling Zhang, Kefeng Mao, Xing Liu, and Zhizu Wang. Experimental construction of eddy real-time structure based on gradient-dependent OI in the Kuroshio– Oyashio confluence region. Progress in Oceanography, 224(??): ??, June 2024. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661124000685.

Zhao:2018:ITN

[ZZPL18] Ruixiang Zhao, Xiao-Hua Zhu, Jae-Hun Park, and Qiang Li. Internal tides in the northwestern South China Sea observed by pressure-recording inverted echo sounders. Progress in Oceanography, 168(??):112–122, November 2018. CO-DEN POCNA8. ISSN 0079-6611 (print), 1873-4472 (elec-

tronic). URL http://www.sciencedirect.com/science/
article/pii/S0079661117303580.

Zhu:2006:BPO

[ZZWL06] Zhuo Yi Zhu, Jing Zhang, Ying Wu, and Jing Lin. Bulk particulate organic carbon in the East China Sea: Tidal influence and bottom transport. Progress in Oceanography, 69(1):37– 60, April 2006. CODEN POCNA8. ISSN 0079-6611 (print),

1873-4472 (electronic). URL http://www.sciencedirect.
com/science/article/pii/S0079661106000164.